A Bibliography of Publications in the book series Boston Studies in the Philosophy of Science and Boston Studies in the

Philosophy and History of Science

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
WWW URL: http://www.math.utah.edu/~beebe/

30 March 2022 Version 1.03

Title word cross-reference

2 [Lip17]. ^{∘} [Onn91i]. _{2} [Cha12d, Cha12e, HD95, SVD95]. β [Fra09b]. G [PP15a, PP15b, PP16]. E = mc^{2} [Jan03b]. N [Bri96, Nor03, PP03, Gre83]. N_{2} [PLH95]. Q [Gan15]. S [Cos87a, Mun15].

&HPS [Ric12].

* [Nee11j].

-Dimensional [Bri96]. -Dualities [Mun15]. -Matrix [Cos87a]. -Stein [Nor03].

Ṡaṅkarā [Ano06-30].

/On [Ano06i].

1 [JNR^{+}07b, Sch08a, UM86a]. 100 [FB17]. 119 [Wer01]. 120MN [Sun18]. 12th [Gre75]. 14 [Ras00]. 1690s [LP01]. 16th [Cha14, Hel08, Vil08, Bro08, Duh91l]. 16th-Century [Hel08]. 17-Kev [Fra99a]. 1779 [Bou02a]. 17th [Bla94, Fre86i, Van08, GHQ96b]. 1800s [Wan18c]. 1850s [Bog15b]. 1870s [Gov15]. 1890s [Kat06b]. 18th [Bla94, Fre86j, GP03, Sil11]. 1913 [Jan07]. 1920s [Che97, Vil02e]. 1925 [Hen98]. 1927e [CS79t]. 1930s [Kar10d]. 1932c [CS79u]. 1933b [CS79-36]. 1936b [CS79m]. 1938/1942 [Ano08o]. 1938/42 [Ano08-32]. 1941a [CS79p]. 1942d [CS79k]. 1945 [CS79-30]. 1945e [CS79-30]. 1947b [CS79-35]. 1950a [CS79l]. 1952a [CS79s]. 1953c [CS79-44]. 1954c [CS79-40]. 1955b [CS79-34]. 1956a [CS79-47]. 1956e [CS79b]. 1956f [CS79-41]. 1957a [CS79-27]. 1958b [CS79d]. 1959b [CS79w]. 1959c [CS79r]. 1960s [NG15b, Wen14]. 1961 [War63]. 1961/1962 [War63]. 1961e [CS79o]. 1962 [War63]. 1963a [CS79-33]. 1963c [CS79h]. 1963d [CS79v]. 1963h [CS79-31]. 1964/1966 [CW67]. 1965c [CS79x]. 1965d [CS79-28]. 1966 [CW69a, CW69b]. 1966/1968 [CW69a, CW69b]. 1966a [CS79-37]. 1967a [CS79g]. 1968 [CW69a, CW69b]. 1968e [CS79e]. 1968f [CS79-39]. 1968g [CS79z]. 1969b [CS79c]. 1969c [CS79-29]. 1969e [CS79n]. 1969f [CS79-45]. 1970b [CS79j]. 1970s [Cer14, Fu93a, Wen14]. 1971a [CS79i]. 1971b [CS79-46]. 1971c [CS79-38]. 1971d [CS79y]. 1971e [CS79-42]. 1972a [GP79]. 1972d [CS79-43]. 1972f [CS79f]. 1973 [MS75b]. 1974a [CS79-43]. 1977 [GCC81b]. 1977b [CS79q]. 1980s [Fu93a]. 1984 [HFD88]. 1985 [CW85]. 19th [Cha14, Hor93].

2 [JRN^{+}07b, Sch08b, UM86c]. 2.IV.1948 [Coh94l]. 2000s [Bog15b, Gov15]. 2008 [dF14a]. 2009 [Bar14b]. 2016 [PP18]. 20th [Bar02b, CT09, Wis71, FS96, Lop92]. 20th-Century [Wis71]. 21st [Cho14, Ful02]. 28.IX.1882 [Coh94l]. 2WW [Pap14].

3 [RSS^{+}07a]. 30-Year [Nik15]. 30th [CS79x]. 35 [Mir14]. 35-26-7 [Mir14]. 368 [Rit05].

4 [Pos95f, RSS^{+}07b]. 4-Vectors [Wal07b]. 4.19 [Onn91i].

50th [Nia96]. 5B [Koh95]. 5th [Yil05, HFD88].

60th [BM89].

7 [Mir14]. 70th [CS79-34].

80th [CS79h].

90th [BAG17].

Âge [Bea75].

A. [CS79b, Øhr97, Tob96]. A.D. [Har10]. AA [TO91]. Abaci [Cai14]. Abandoned [Lau80a]. Abdera [Vam09d]. Abduction [Des00]. Ability [Dam96k]. Abner [Aga87c, CHS97a, CHS97b].

Abraham [Ren07c]. Absolute [Bar76a, Bec91a, Bec91o, Bru76, Duh76a, Duh76e, Eul76, Föp07, Fre86q, Fri07a, Gas76a, Grü73l, Hor16, Max76c, Max76b, Nee11b, Neu76b, New76a, Pos95a, Rus76b, dB87s]. Absolutes [Fun86]. Absolutism [OP82]. Abstract [Dam96h, FA76h, LE97, Zin73d]. Abstracting [Cle02]. Abstraction [Dam96a, Dam96b]. Abstractions [GS09]. Absurdum} [Far03e]. Abuse [Pro01]. Abyss [Ken17]. Academia [Mei17b]. Academic [Abb14, Bel15, Her14a, Mün14, Pat76a, RT01, SDG15b]. Academy [Bro08, Seg03, Wen14]. Acausality [Bel96]. Accelerator [SS74]. Accept [And84c]. Acceptability [WLC18]. Acceptable [Bar16, DF93]. Acceptance [Coh95, DF93, Fet81a, Hop74, Mic71, Zim83]. Accepts [Coh02o]. Accident [DS18]. accidente [Ela11d]. Accidents [Ela11d]. Accommodation [Pal03, Ung91b]. Accomplishing [Har99]. Accomplishment [Wal81n]. Accomplishments [HTG14]. Accorded [Feh99b]. According [Bai76a, Bai76b, Duh76a, Duh76c, Duh76f, Duh02l, Gas76b, Klo79, Pal12, Pal13, Ras94a, vW94]. Account [Buc95, Bun12c, Dil96, Dil06f, Fab84, Fis15, Imo16, PP15a, PP15b, SM96, Wen96, KV11]. Accountability [WG11, Pes07]. Accounts [Ban80, Sch12a]. Accumulation [Ams75h]. accurate [Onn91n]. Achievement [Mor63]. Achievements [Kap80b, Sol12a, Tob96]. Achieving [Tri12]. Acquaintance [Koh98f]. Acquiescence [Zam92l]. Acquired [Ges74b, Gru74a]. Acquisition [KK07, Mar78e, Mar78d]. Acragas [Vam09e]. Across [Elk10]. Acted [Lag97j]. Action [Ada01, Bou02d, Cha12f, Dam96c, Enç05, HOO88c, Jaw90, Kow16k, Len82, Lev16, Mar78a, PR97, Sau07, Tur86j, Tur86k, Tur86l, War79a, War79g, War79m, War79r, War86, Waw82, WW03]. Action-at-a-Distance [WW03]. Action-Guiding [Kow16k]. Actions [Mar79c, Rot87d]. Active [AB12, Cha12a, Fin80a]. Activism [Bar01b, AM01b]. Activities [DPW86b, Ols82]. Activity [Kap80-40, McC02, Ton73a, Ton81c]. Actor [Sam92]. Actors [Che10]. Acts [Sta07a]. Actual [Bec91b, Bec91s]. Actuality [Kuz87e].

Acute [Ach67, Ges74-27]. Ad [Aga75s, Far03e, Sas03h]. Adaptation [Akb11, DD12, GR11b, Has11, MC80i, Mun76, Rai11]. Adaptational [Sta79]. Adaptations [Wan18a]. Adaption [Kum11]. Adaptive [Gro82a, Nik15]. Addis [Hem76]. Additional [Gro09a]. Additive [Kał82]. Additivity [Wei01b]. Address [Bad92]. Addresses [Kap80j]. Adequacy [Gil09a, Sim11b]. Adequate [Bro86, Hol83, Tur86h]. Adjustable [HL17]. Adler [BIT01d]. Administration [Cla10, Kum92]. Admissible [Bun12r]. Adolf [CL83]. Adoption [Chu11, GR11b, Has11, Kum11]. Adult [Dvo91, Sta91c]. Adults [Aga76b]. Advaita [ZC88h, Ano06a]. Advaita-Vedānta [Ano06a]. Advanced [Bar14a, OJ14]. Advancement [FS96, SC74b]. Advances [Pes11]. Advantages [And03b]. Adventurer [De 17]. Advocacy [Mar91]. Advocate [RB09]. Advocating [Mai01]. Aesthetic [CK97, Fax97, Koh97, Mar97b, McA97, McC02, RB97, Tru95, Wol76-29]. Aesthetics [Bab02b, Fin80d, GF97, Par00, Tau97a, Wan17, Wet87]. Aether [CS79-29]. Aethers [Mcg74]. Aetia [And90]. Affair [Ela11f, Fin09e]. Affairs [Fin09b]. Affection [Buc89]. Affine [Car07]. Affinities [Mil11a]. Affinity [Wei01b]. Affordances [Har15b]. Africa [Coh17a]. After [CS79v, EF18, Gou94b, HH15, STNW12, Whi88, Boc96, Ger15a, Ges74j, Vil02c]. Aftermath [Bla95e, Rot86b, Sch18]. Afterword [Are89a, Fre86c, Hee02, Kuz87a]. Again [All99a, Bal90, Bue11, Joh11, PP03, Res17b, Sup03, Sza03]. Against [Bab97, Duh91k, Gav96, Haa99, Lau89, Mac90, Pop84, Rad76, Rey18, Rop99, Sch91b, Sep87, She82, Str84, Toa15, Tuc76, Yeh17]. Agassi [Bai17, BAG17, Cha17, Don17, Fin17, Fol17, JL95a, JL95b, Lao17, Lip17, Shi87a, Win17a]. Age [Aga81-36, Gab01, Lao95, Lek15c, Wan17, Ágo11, Fre86f, KB92b, Pog94]. Agency [Bus03, O’C02, Pic12]. Agenda [Kar10d]. Ages [CS79-45, MS75b, Bea75, Duh91n, Duh91p, Duh91o, Wic91]. Aggregation [Sim77a, Tur86f]. Aggregativity [Wim86]. Agnosticism [Aga75a, Aga76a]. Agraphia [Ges74c]. Agrarian [Hei08f]. Agreement [FC16, McL71]. Agricultural [Fin91, Hei08a, Hei08e, KS83, PHL95].

Agriculture [Cul11, Fin91, GC95, WRL95]. Agronomists [Sch94b]. Ahistorical [Ung94]. Ahjār [Haq94f, Haq94a, Haq94b, Haq94e]. Aid [Aga81g]. Aids [Dam96e]. Aim [Car08, D’A89, RH14]. Aiming [D’E97]. Aims [Hod74b, Kra95, Wuß91]. Air [CS79b, Don98]. Airy [Bur74a]. Akademie [Bie74]. Aken [SPB18]. Al [Fre91, Haq94f, Iru01, Ras94a, Ras00, She75, Haq94a, Haq94b, Haq94e]. Al- [Fre91]. al-Ahjār [Haq94f, Haq94a, Haq94b, Haq94e]. Al-Haytham [Ras94a]. Al-Jassās [She75]. Al-Sijzī [Ras00]. Alan [FB17]. Albert [Ber74b, Duh91c, Duh91j, Sch03, Zam92a, Zam92d, Zam92b, Zem09, vD17]. Albertus [Wal81h, Zam92h]. Alchemical [Kno89]. Alchemist [Haq94f]. Alchemy [Dob92]. Alcmeon [And90]. Alcott [Kon88]. Aldo [MP16]. Alexander [Bie74, Rit01]. Alexandre [Kap80a]. Alexandrian [Duh91a]. Alexandrovich [Kap80a]. Alexia [Ges74f]. Alexias [Ges74a]. Alfarabi [Mah75]. Algebra [Pec96, Ras75, Ras94b, Ras94c, Ras74, Scr74a]. Algebraic [Cif05]. Algebras [BT96]. Algébre [Ras74, Ras75]. Algebrization [Phi94]. Algorithms [Par17a, Sha93]. Alhazen [Smi90]. Alienation [Wro88]. alla [BAK11]. Allan [Byn89]. Allegory [Kno89]. alley [SVD95]. Allgemeine [Hof74]. Allies [Row06]. Allometric [Nik15]. Alluvial [Hie95b]. Almagest [Ber14, Yav17]. Almost [Ren07c, Sim77n, Voi96]. Almost-Everywhere [Sim77n]. Almqvist [CS79g]. Alois [BIT01a]. Alone [Zam92m]. Alpha [Lip86, Stu86]. Alpha-Decay [Lip86, Stu86]. Already [Kle91, Bou17]. Alteration [Sha84a]. Alternative [Alt87, AZ87, Ano08-42, Ano09l, Fre86d, Grü73p, Grü73u, Jos17, MC80a, Pan83a, Sac83, Tow71, Urb89]. Alternatives [Jar86h, Pan83b]. Althusser [Bal89a]. Altitude [Cue01]. Altruism [Cla12, Con96]. Altruistic [Cla12, Kle12]. Always [Loc97, Pro14, Rus09]. am [Mar79w]. Ambassador [Gün11]. Ambiguities [Gun97, Mue95]. Ambiguity [Coh02m, Gra95]. Ambivalence [Mer76, Sig94]. Amendment [Čap71e]. America [GPSR01, Jon91, Pop88, Sha84d, Wan17, Whi88, You95]. American [SC74b, WC91, Bar01b, Eas84, Hor88, Lau09, Sla89, Ste92, Wal95]. Amiss [HSS09a, HSS09b].

Amnesia [Ges74x]. Among [CS79-34, Cro99, Har08a, Mar95d, McA97, Dou87, Ols82]. Ampère [D’A01b, Onn91k]. Amsterdam [Pop88]. Analogical [Hil94]. Analogies [KL17, Nor13]. Analogs [Gro82h]. Analogy [Ano06b, Ano06c, GC06, Ito11, Nor81, Sca90b]. analyse [Ras74]. Analyses [Nik15]. Analysis [Ack92, AM63, Ano08-37, Bin96, Blo76, Buc84, CS79b, Coh94a, Com84, Cur80, D’A01k, Dev15, Dun89, Ela11h, Gon03, Hag16, HGR^{+}89, Hec01, HR74b, HR74a, HR74c, HR74d, HR74g, HR74f, HR74e, HR76, Jos17, Kat11f, Koe80, Koe86, Lin93d, Mar80, Mar90, Mat89, May74, May86, MC80c, Mil15, OP97, Rad12b, Rap81c, Ras74, Ras94a, Ras94f, Ras94e, Rat70, Rod15, Roq11, Sha84k, Tok09, Ton73c, Ton81f, BJ95]. Analyst [Ned90a]. Analytic [Fre86h, Kan84, Liu93, Mar67, Rad82]. Analytical [Bit96a, BV97, Coh04, De 81, Hac94b, HR74b, Rap81a, Zel95, Seg03]. Analyticity [Dav97, Grü05b, Win71]. Analytics [Hof74, Wal92e, Wal92c]. Analytik [Hof74]. analytique [BV97]. Anarchism [Far03d, Hoo91, Mar91, Mus78]. Anarchist [Die91]. Anarchistic [Sup91]. Anarchy [Ams92i]. Anatomical [Ges74t]. Anatomy [Ges69a, Ges74b, Hin94, Lam12, Lam13]. Anaxagoras [Vam09a]. Anaximander [Vam09b]. Anaximenes [Vam09c]. Ancient [Ana90a, Har10, HR76, Kno04b, Kno04c, Ras95, Sza84, Ung04a, Wan18c, Bai76c, Cha09c, Cha09l, Chu11, Dam96e, DL05, Fil05, Kno04a, Lia96a, LY96, Nie74, SM90]. Ancylostoma [Bel81b]. and/or [Zel95]. Andrea [Fie14, Wai14]. Andrews [Hoo99c]. Anecdotes [Bol95b]. Anglican [Ols87]. Anhui [Bin96]. Aniline [RT01]. Animal [Fen12, GWT00i, Gre74j, Gre95, Per87, Reg13, Tha86a]. Animals [Ges74k, Gre74o, Mar78k]. Animate [Hon09b]. Animism [Han88]. Anisotropy [Grü73a]. Ankara [KC95]. Anneliese [Wal81a]. Anniversary [CS79x, Kap80-38, Nia96]. Annotated [Amr87a, Nai86]. Annotations [Sch01]. Anomalies [Ams75e, Wim92]. Anomalous [Hen98]. Anouk [Ly08]. Answer [Rud67, Ung94]. Answered [Ror86]. Antagonisms [Sta91b]. Antecedent [Kam93]. Anterior [For01].

Anthropocentrism [Gal86]. Anthropological [Ber94, Gru74a, Hol83, Lee74, Wol89]. Anthropologists [Jar86a, Jar86d]. Anthropology [Ano08o, Ano08-45, Ano08-40, Ano08-32, Arm18, Col74, Fel89, GWT00a, Jar74, Jar86b, Jar86c, Jar86h, Jar86m, Jar86n, Jar86o, MAM69, PS01b, Sch08d]. Anthropomorphism [Grü73m]. Anti [Bal94, Ber96b, Coh88a, CHR96, Din96, Fol96, Kau74, Lee84, Nol03k, TB96, Von91]. Anti-Individualist [Ber96b]. Anti-Liberal [Von91]. Anti-Rationalist [Nol03k]. Anti-Realism [Coh88a, CHR96, Din96, Fol96, TB96]. Anti-Realist [Tul16]. Anti-Reductionism [Kau74]. Anti-Sociobiology [Lee84]. Anti-Whiggism [Bal94]. Antibody [HH89]. Anticipate [Stö11]. Anticumulative [PM82]. Antidemocratic [Aga95a]. Antidote [Rad76]. Antiessentialist [Gin03b]. Antifeudal [Fre86e]. Antilogikē [Fey89]. Antiquité [Fre04]. Antiquity [CS79-40, Fre86b, Fre04, Kno04d, Nen18, Glu91]. Antirealism [Aga97b, Kre96, Ren96, Sha96]. Antirealist [Sch96b]. Antirealists [Fra96a, Fra99k]. Antireductionism [Băe15]. Antoine [Sil11]. Antonelli [MP16]. Antonism [Bai76c]. Any [SF94, Sta97, Pan91i]. Anyone [Lag96]. Anything [Bun12b, Gou94a]. Apart [Eng01]. Aphasia [Ges74g, Ges74e, Ges74s, Ges74z, Wer69]. Aphasias [Ges74t]. Aphasic [Ges74q]. Apocalypse [McK92]. Apollonius [Ras00]. Apologetics [Aga95a]. Apologia [Pit89]. Apparatus [Bry98, Goo77f, GR89]. Appeal [Meg74, Mey85a]. Appearance [Con99, Fra99a, Goo77j, Han99i, Onn91j]. Appearances [Alm89, Gau99]. Appendices [Mac97a]. Appendix [Bun10b, Bun10a]. Applicability [Gin15, Gin18]. Application [CN11b, CN11a, Gai86, Gie71, Gre99, Imo16, Kör86, Lag97e, Pfa96, Rad12e, Ras95, SdV92, dV11]. Application-Oriented [SdV92]. Applications [Lin93a, MPC91, Nee11g]. Applied [Hin76, Joh11, Mor11, Ng17, Yav17]. Applying [Dem89, Kap80b, Seg09]. Appraisal [CFG96, Grü73p, McM76a, Pan83f, Pan91k]. Appreciation [Ber76c]. Apprenticeship [Val10a].

Approach [ÁNS91, AM83, Asl18, BBD18, COG90, Cat93, Che11, Chu95, Chu00, CG90, D’A95, D’A92, Gon96, Gon03, Goo77a, Gru80, Han17, Hel00g, Hon16, LMR95, Maj06, Man15a, Mil15, Nag84b, Nor13, dJAO14, Pan96, Pat82, Pol74b, Pos95p, Rim97, SM74, Ste79, Szt79a, Ton73d, Uye97, Wat78b, Zim99, GC95]. Approaches [Abr11, Ano06c, Bra11, Bur87b, Cur80, Dud86, Lon12, MC80a]. Approaching [Eck17]. Appropriate [Car83, GW97]. Appropriated [Wal92e, Wal92c]. Appropriation [GP03, Pat11a, Gar16]. Appropriations [Wat01b]. Approximation [Lag97f]. Approximations [Say05]. Approximative [Kra95]. Apraxia [Ges74c]. Apraxias [Ges74d]. April [Hen98]. Apriori [Mis00]. Apriorism [Fin80j, Noé92]. Aprioristic [Mey85b]. APS [LL14]. APS/CPS [LL14]. Aquinas [Syl75]. arabe [Ras74]. Arabic [Fre00, Ras74, Ras94c, Too04]. Araki [Yan98]. Arbitrary [Gro82k, Lag97d, Lag97e, Lag97f, Lag97l, Lag97t]. Archaeoastronomy [Pap94]. Archaeological [Bel92a, Emb92c, HK92, Nol03a, Sab92]. Archaeologists [Emb92b]. Archaeology [Dre92, Dun92, Gel16, Sal92c, Sal92b, Wat92a, Wat92b, Wyl92, Wyl15]. Archimedes [Duh91b, Dun96]. Architect [Jar86z]. Architectonic [Aga03k]. Architecture [Nye01, Sol12b]. Archive [Lek15g]. Area [Ges74o, Hin95]. Argentina [Mon01]. Argon [AG16]. Argument [Aga81v, Čap71x, Cla12, Dav95, Eul76, Fin80d, Fin80f, Gav96, Goo67, Goo86, HH94, Kom99, PS15, Pet69a, RL97, Rös95, Ste18]. Argumentation [Rad03]. Arguments [Aga75e, Duc96, Ier90, Mar83b, Pol71, Pop84, Seg00, vZ11]. arid [WRL95]. Aristides [Obe08]. Aristocratic [Mor80]. Aristole [Nee06]. Aristoteles [Bec04]. Aristotelian [Bec91c, Bec91f, Bec91m, Fer98, Kra00, Lai08, Mar12, Mar13, RM18, Wol69, vL16a]. Aristotelianism [Alt00]. Aristotelis [Gat18]. Aristotle [Aga97a, Ant90, Bec91d, Bec04, Cle02, Duh76f, Duh91b, Ela11c, Fey78b, Gas76b, Gre74a, Gre74l, Gre76a, Jam76, LMR95, Ruf69, Sas03l, Uga18, Wal92c, Wal92e, Wed86]. Arithmetic [CPM16, Dam96e, Dam96i, Gan15, Gar03, Mat86, McC02, Neg81, Nic18, Ras94c]. Arithmetica [Chr15]. Arithmetical [Dam96e, Pan16]. Arithmetics [Lam14]. Arles [Mar02]. Arnold [Wes95]. Aromatic [Kar10g]. Array [Fra96b]. Array-Based [Fra96b]. Arrow [HS16, MC80a, MC80h]. Arrowed [Pos95b]. Arrowless [dB87q]. Arsenal [Val10e]. Art [Aga03n, Aga03p, Aga03-32, Bab97, Bab99c, Bab99d, Bab02b, Cha10, Che97, Cif05, Dav10, Fax97, Fey95a, Fin80h, FH10b, GSW95b, Hei97, Hoo99m, Hym10, Jan89h, Jie96, Kel94, Lan94, Nyí97, Sie63, Ung04a, VN12, War79c, War79b, Wes90a, Wes95, Wet87]. Arte [Bus74]. Articles [Kap80j]. Articulating [Mor09]. Articulation [Fra01, Kau71, Kau76]. Artifact [Nic12]. Artifacts [GC94, Kel94, Koh98a, Ols05, Par09a, Pos18, Sta94b, Şte15b]. Artificial [Aga03a, Gav14b, Ger95, Lef01b, Mar86d, NG01g, SPB18, Tho96, dS14]. Artisans [NG01a]. Artist [Val10a]. Artistic [Bar02b, DD12, Gai94]. Artists [NG01a]. Arts [Fey67, Jar86l, Jar95, Kiv94, RB97, Hel08]. Ascent [Tur86d, Tur86g]. Ascertain [Grü73e, Mer97]. Ascetic [Owe99]. Asia [Ágo11, Bre11, BE10b, GR11b]. Asian [Jhu07]. Ask [Hål07]. Aspect [Fra96b, Mar79t, Mur75, Sha74c, Ste07, Pos95t]. Aspects [BN17, BJ89, Čap71-31, Čap91g, Gre74r, Hul76b, Kör86, Nik15, Pos95k, RT01, Sch76d, Sie63, Sta91a, Toe81, Twe00, Zaj87]. Assembly [Kel09]. Assessing [WvB07, WPJ95]. Assessment [Car83, Kle12, Kow16g, Kow16h, Med96, Mic83, Mis95a, Rap83, Set76, SF83, Sko76b, Zim83, Zol89f, tH09, BBT95]. Assessments [Kow16a, Kow16c, Kow16e, Kow16j]. Assimilation [Ela11a, Ela11g]. associated [BdOU^{+}95]. Association [Ges74f, SC74a, SC74b, Tur12, RL95, BC71, CHMV76]. Assumption [Tra15]. Assumptions [Bel12j, Gra00, Kno16]. Assurance [Aga75a, Aga76a]. Astral [Bro10]. Astro [Coh90]. Astrology [Lan00, Zam92d, Zam92e, Zam92h]. Astronomer [Gen91, She90]. Astronomiae [Zam92k, Zam92n, Zam92h]. Astronomical [Ano07a, Car15b, LE97, Mus91]. Astronomy [Bru95, Fin09k, Gol91, Kal90, Lan91, Rai10, RSS^{+}07a, Ros76, Swe90, Ung91b, Vla90].

Asymmetries [Ges74m]. Asymmetry [Ano08l, Grü73b, Hon93]. Asymptotic [Fre00]. Asymptotics [Pos95c]. Atkinson [And03b, Stö03]. Atlantis [Che94]. Atom [Cha09a, Fre86f, MC80v, Whe74]. Atomic [CS79y, CS79-43, CS79-45, Duh02a, Kap80q]. Atomism [Čap71-37, Čap91l, Cay99, Cha09c, Cha09b, Cha09d, Cha09e, Cha09g, Cha09f, Cha09h, Cha09n, Hie70, Len89, Roc01, Str76]. Atomistic [Čap71r, Val10f]. Atomists [Duh02l]. Atoms [Cha09a, Cha12d]. Attaching [Pag96]. Attack [Str76]. Attain [Tac17]. Attempt [Bel15, Buc95, Cam03, Ghi93, Hon03, Mey91h, Phi94, Sas03e, Ung94]. Attempts [Duh02c]. Attention [Gro82i, Max76a]. Attest [Grü89b]. Attitude [Hon96a]. Attitudes [Kim11, Mar11b, Pap14, Rec00, Ton81k, Wan18c]. Attraction [Cha12c, Lag97j]. Attributes [Sch76e]. Attunement [Sca90b]. Auctoritas [Zam92a]. Auditory [Čap71-36]. Aufbau [Goo77k, PS01a]. August [Chu01, Heg91, Jan07]. Auld [Koh98f]. Aura [Smi94]. Ausdehnungslehre [Lew96, Beu96]. Austral [Syl97a]. Australia [Stu97]. Austria [Smi97, Stu97]. Austrian [Aic15, Bar74, Hal91e, LM97, Syl97a, Ueb91b]. Authenticity [KT76, Wol95a]. Author [Vil02g]. Authority [Cah05, Orv89]. Authors [Wol76z]. autobiographic [HOO88c]. Autobiographical [Hal97]. Autobiography [Wol95b, vD17]. Autocratic [Yal96]. Autoimmunity [RB91]. Automata [Sim77y]. Automaton [Hon09b]. Autonomization [Böh83]. Autonomous [Aya76, But93a, Ned90a, Syl75, RR15]. Autonomy [Aga81a, Aga03-28, CM89, Gin07, Lom15]. Autopoiesis [MV80a, MV80f, MV80g, MV80k]. Autour [Pet92a]. aux [Ras75]. Auxiliary [HR74i, Ram01]. Avait [Fre04]. Avait-Il [Fre04]. Averting [Gor07]. Avogadro [Cha09k]. Avoid [Fra99e]. Avoidance [Gro82j]. Awakening [Kar10c]. Away [Boš79]. Axiology [Gon15b]. Axiom [Mar74b, Sim77f]. Axiomatic [Cat93, Cor07, Maj06, MC80t]. Axiomatics [Kno04d, Øhr97]. Axiomatization [Kar90, Sim77b, Sim77g]. Axiomatizations [Bre74b]. Axiomatizing [Gra83h, Leb95]. Axioms [Bud93, FT16, Sim77c, vH77e]. Azarya [GC97].

B [Bun10b, Duh91k, Eas84, Hil94, Smi92, Str76, GP79, Onn91c]. B. [Kit74]. B.C. [Duh91b, Har10, Vam09a, Vam09b, Vam09c, Vam09d, Vam09e, Vam09h, Vam09l, Vam09m, Vam09n, Vam09o]. Babbage [Seg03]. Babylonia [Cla10]. Babylonian [Dam96e, Pro14]. Babylonians [Krö12, Vog04]. Babyloniern [Vog04]. Bach [Ihd79a]. Bachelard [Čap71-31]. Bacillus [AA86, Ros86]. Back [Ano69a, Ano70a, Ano71a, Ano71b, Ano73a, Ano73c, Ano73d, Ano73b, Ano74c, Ano74b, Ano74d, Ano74e, Ano74a, Ano75a, Ano75c, Ano75b, Ano76c, Ano76e, Ano76g, Ano76f, Ano76b, Ano76d, Ano76a, Ano77a, Ano77b, Ano77c, Ano78a, Ano78b, Ano78c, Ano79b, Ano79c, Ano79e, Ano79d, Ano79f, Ano79a, Ano80a, Ano80c, Ano80d, Ano80e, Ano80b, Ano80f, Ano80g, Ano81f, Ano81c, Ano81d, Ano81g, Ano81b, Ano81a, Ano81e, Ano82b, Ano82c, Ano82a, Ano83e, Ano83f, Ano83g, Ano83h, Ano83a, Ano83c, Ano83d, Ano83b, Ano84b, Ano84a, Ano84d, Ano84e, Ano84g, Ano84f, Ano84c, Ano85a, Ano85b, Ano85c, Ano86b, Ano86c, Ano86f, Ano86g, Ano86k, Ano86a, Ano86d, Ano86e, Ano86i, Ano86j, Ano86h, Ano87e, Ano87b, Ano87c, Ano87d, Ano87f, Ano87g, Ano87h]. Back [Ano87a, Ano88a, Ano88b, Ano88c, Ano89e, Ano89f, Ano89g, Ano89h, Ano89i, Ano89d, Ano89a, Ano89b, Ano89c, Ano90a, Ano90d, Ano90b, Ano90c, Ano91b, Ano91c, Ano91k, Ano91h, Ano91j, Ano91i, Ano91e, Ano91d, Ano91g, Ano91f, Ano91a, Ano91-42, Ano92k, Ano91-41, Ano92c, Ano92j, Ano92i, Ano92g, Ano92f, Ano92h, Ano92e, Ano92d, Ano92a, Ano92b, Ano93d, Ano93a, Ano93b, Ano93e, Ano93c, Ano94g, Ano94h, Ano94i, Ano94e, Ano94k, Ano94j, Ano94d, Ano94c, Ano94f, Ano94a, Ano94b, Ano95k, Ano95l, Ano95g, Ano95h, Ano95a, Ano95e, Ano95b, Ano95m, Ano95f, Ano95c, Ano95i, Ano95d, Ano95j, Ano96c, Ano96f, Ano96h, Ano96b, Ano96a, Ano96e, Ano96d, Ano96g, Ano97b, Ano97d, Ano97f, Ano97a, Ano97c, Ano97e, Ano97g, Ano98a, Ano98c, Ano98b]. Back [Ano99b, Ano99a, Ano99f, Ano99e, Ano99d, Ano99c, Ano99g, Ano00e, Ano00b, Ano00a, Ano00d, Ano00f, Ano00c, Ano01g, Ano01e, Ano01a, Ano01b, Ano01d, Ano01f, Ano01h, Ano01c, Ano01i, Ano02d, Ano02c, Ano02e, Ano02f, Ano02a, Ano02b, Ano03h, Ano03e, Ano03b, Ano03a, Ano03g, Ano03i, Ano03f, Ano03c, Ano03d, Ano04a, Ano04b, Ano05a, Ano05b, Ano06e, Ano06f, Ano06g, Ano06h, Ano06d, Ano07b, Ano07c, Ano08h, Ano08a, Ano08b, Ano08c, Ano08e, Ano08f, Ano08g, Ano08d, Ano09a, Ano09d, Ano09e, Ano09f, Ano09g, Ano09c, Ano09b, Ano09h, Ano09i, Ano10d, Ano10e, Ano10b, Ano10a, Ano10c, Ano10f, Ano11b, Ano11h, Ano11e, Ano11c, Ano11d, Ano11g, Ano11f, Ano12a, Ano12b, Ano12c, Ano12d, Ano12e, Ano12f, Ano12g, Ano12h, Ano12i, Ano12j]. Back [Ano11a, Ano13a, Ano13b, Ano13c, Ano14d, Ano14b, Ano14a, Ano14c, Ano15a, Ano15d, Ano15e, Ano15f, Ano15b, Ano15c, Ano16a, Ano17a, Ano17b, Ano17d, Ano17c, Ano18b, Ano18c, Ano18a, Bla95b, Gea76, Ger15a, JR07, Mic18, Res17b, Sos95, Sup03, Sza03, Wei07]. Back-Reference [Gea76]. Background [Aga13l, Aga13p, And99, Buc90, Den99c, Den99d, FA76e, FA76d, FA76j, FA76k, Mar86e, Nee11g, PQQSe96, Sas03h, Sta91a, Wal92c]. Backlash [Gov15]. Backwards [Fit71]. Bacon [Aga13b, Aga13a, Aga13c, Aga13e, Aga13j, Aga13o, Aga13q, Bec91e, Dei17, Gig12, Gig13]. Baconian [Coh02q, Gig12, Gig13]. Baconianism [Coh02c]. Bacteria [Sar91, RL95]. Bad [Fin09f, Fin09g, Pie12]. Baer [Pog94]. Baglivi [And18]. Baker [Kit74]. Balance [Haq94b, Kow16b, RM18, SVD95]. Ball [Kap80y]. Baltas [Obe08]. Bancroft [Duh02t]. Bane [HL17]. Banner [NG03]. Bao [Lih93a]. Bar [Ara15, Wol89, Kas76b, UM92]. Bar-Hillel [Kas76b, UM92]. Barberousse [Ly08]. Barbò [Bur83]. Barcelona [RR15]. Bargain [Leb95, Ano08o]. Baroda [RH92]. Barreda [Gut95]. Base [Aga84, Aga03k, Mar94c]. Based [Bro96a, Duh02b, Duw17, Fra96b, Lag97f, Tau98, MPC91, Kow16m]. Baseline [Hon16]. Bases [Goo77d]. Basic [Kap80c, Krü91, Lin93d, Sch76d, SG11]. Basics [Sos95]. Basis [CS79i, Dor96, Fin80i, Ger15a, Hüb78b, Mac12a, Pan91k, Sim77v, Ton73c, Zah89]. Batches [Ita92]. Battista [Gat18, Omo18a]. Bay [Fei74, Fei86, Har08b]. Bayes [Bar17]. Bayesian [Ano09j, Dou03, FH10a, Gar03, Grü05a, How03, Riv96, Say05, Urb89]. Bayesianism [Coh02c, Hin71, Jur08, TG88]. BB [DO91]. Be [And18, Bun10g, Coh16, Flo15, Gar12b, Lag96, Le 08, Mar84p, Ror86, Sal18, War91, Wet09, Zwi08, BR18b, Cao07, Coh02d, Coh02e, Coh02p, Fra99c, Gin03b, Grü73d, Kap80q, Koh98f, Kra97, Nic80a, Nor98a, Pol92]. Beard [Ore00]. Bearing [Grü73c, Grü73r]. Beautiful [GM14, Hol97]. Beauty [Smi94]. Became [AG16, Bal14, DS18, Gig12, Gig13]. Because [Jan89j]. Become [Aga76b, Bun10g, FF07, Gou89, Kle91, Sza04]. Becomes [Bar99]. Becoming [Čap71a, Čap71-41, Grü73n, ML99, Pau97, Tha86a, Tha86b]. Becquerel [dAM11]. Beddoes [Lev07]. Bedeutung [Hof74]. Bedroom [Mar02]. Beeckman [dB12, dB13]. Been [Pap89, Wet17]. Bees [Hei17]. Before [Alt14b, Fre86n, HH15, LR08, dR15b, Bar01b, Vog91]. Began [Hac94a]. Beginning [Coo84, Vel14, Wol76k, Wol76-38]. Beginnings [Kat06a, MC80j, Pat92, Ras94b, Ret74, vE75]. Behavior [Fis96, Kit74, Lon12, Pos95i, Sza79b, Tha86a, Tur12]. Behavior-Shaper [Kit74]. Behavioral [Gro82b, PR12a, PR12b, Web12]. Behaviour [DD12]. Behaviouralist [Nic89a]. Beijing [CHR96]. Being [Bol17, Byr02, Coh02b, FC87, Kuz87c, Kuz87h, Kuz87k, Mar95c, dF14b, Bun12u]. Beispiel [Hof74]. Belief [Ano09-42, Ano09-44, Cam87, Coh95, Enç05, Kau96, Mar79m, Mus09, Nol99, Nol03c, Nol03i, Pan96, PR17, San90, Wri80]. Belief-Sentences [Pan96]. Beliefs [Ano09p, Grü95b, Mar78k, Pol16]. Believe [Arm88, BE88, Han67, Too10]. Believes [Coh02o]. Bell [Per97, DS12, Koç05, Pos95t]. Below [Har15a, Onn91i]. Ben [Gol91, Ezr14, Her14b, Her14d]. Ben-David [Ezr14, Her14b, Her14d]. Bench [NG09]. Beneath [Sha90]. Benedetti [Duh91k, Omo18a]. Benefit [Mic71]. benefits [GC95]. Benjamin [CW84b, Kap80-40, Smi94]. Benzene [Kar10a]. Berechnung [Vog04]. Bergmann [vHBM01]. Bergson [BB71, Can16, Čap71e, Čap71a, Čap71b, Čap71c, Čap71d, Čap71h, Čap71n, Čap71q, Čap71v, Čap71-31, Čap71-32, Čap71-33, Čap71-36, Čap71-40, Čap91a, Čap91j]. Bergsonian [Čap71i, Čap91r]. Bergsonism [Čap91k]. Berkeley [Bec91f]. Berlin [Bou02a, Cha18, Wen14, HH91, Hof98, Kóh90, MP13, Mil11a, Pec11, Pul11, Res11, Sch95b, Spr91a, Spr91b]. Bern [Sch95b]. Bernal [CS79-41]. Bernard [Sto75, Web08]. Bernd [Bro95]. Bernoulli [Bos74]. Bernstein [Tau91e]. Berthelot [Duh02w]. Bertrand [Gar95b]. Besso [Jan07]. Betrayed [Wei04]. Better [Fol18, Mai12, McB94]. Between [Aga81b, Aga13j, Ágo11, Akb11, Ams75a, Ams92b, Ara12, Arm18, Bea75, BZ96, Bou02i, BH17a, Bre11, But93f, Čap71j, Cek92b, Cha15, Cro99, DR15a, Die14, Ela11j, Elk86, Fel18, Fox07, Fre86l, Fre86r, Gäh89, Gol79, Grü89b, Ham91, Hoo91, Jah94, Jan89a, Kar10a, Kat76, KJ14, KT01, Lam14, Lef01a, Lev07, MM91, Mal12, Mal13, Mas15a, Mas15b, Mor76, Mor11, Mun89, Näp01, Ols87, Pap15, Pro95, Rad82, Ras94c, RSS^{+}07a, Sch07a, Sch89, Sim11a, Sle99, Soy05, Spr91b, Sun18, Tan01, Tym15, War79p, Wen14, Wol69, Wüt16, Živ79, Aga75c, Aga75d, Ano07n, Bal02, Bel12g, BB00, Ber13c, Bit96d, Bru76, Bus97b, CS79h, Coh94a, Cul11, Dam96g, Dor97, Fre99, Gou90, GWT00i, GG03, Gre74m, GR11b, Hal71, Hie94, Hül96, Ima01, Jar86p, Lei03]. between [Li18, Lim94, Mic74, Nik74, Øhr97, Pul06, Sch96e, Sch99d, Shi98, Tab03, Vam09g, War76]. Beuler [CS79q]. Bewandtnis [Kis02]. Beweis [Str94]. Beyond [Aga75b, ACE17, BS15, Cha12b, Con99, FH10b, Gau99, Gin06, Gou83, Mac12a, Mun91a, Pła95, Tur86a, dR15b]. BhāSya’ [Ano06a]. Bhavan [RH92]. Bias [Art15, IdMM16, Kow16b]. Biases [Wim80]. Bible [McC02, Pop74, Pop86b]. Biblical [Fin09c]. Bibliographic [Bas14]. Bibliography [Amr87a, Coh94l, Mac97b, Nai86, Nor98b]. Bibliometrics [MC14]. Biblionomia [Zam92b]. Bicentenary [Set76]. Biennial [BC71, CHMV76, SC74a]. Bifurcations [Pet03a]. Big [Ash07a, Ash07b, Voi96, vD14]. Bild [D’A01m]. Bild-Conception [D’A01m]. Bilingual [Aga76b]. Bilingualism [Cif14]. Billy [SPB18]. Binary [Gro82b]. Bio [Bec86, Coh94l]. Bio-Bibliography [Coh94l]. Bio-Social [Bec86]. Biocentric [Koh98b]. Biochemical [Gro82l, Zal86]. Biochemist [Sar92c]. Biodiversity [SG11]. Bioethics [Gro95]. Biographical [Für92, Kuh09, Cai14]. Biographies [Kra15b]. Biography [FM09a, Kra15b, Nye15, Sha88, Söd07]. Biological [Atl86, Bal15, BdOU^{+}95, Bro91b, Bur86, Čap69, Čap71e, Čap71g, Che11, Con96, Deb09, Gon01, Goo01, Gro82c, Gro86, KK07, Len87, Mac01, Mey85c, Mey91a, Mor09, Mun76, Pat76b, PHL95, Pra76, RL95, SA82, Web12, ZRKC03e, BJ95, GC95, SVD95, WRL95]. Biologie [Sch99c, Sch99c]. Biologism [Moc91]. Biologist [dR17]. Biologists [Gri15]. Biologization [Urs96]. Biology [Aya76, BMP09b, Bar94, Bee91, BH17a, Bur86, CC90, Cro92a, Cro92b, Daf96, Dei17, Fin84, Gay09, GS09, Goo69, Gre74a, Gre74d, Gre74c, Gre74b, Gre74t, Gre76a, GM76, HN16, Heg91, Hod92, HH94, Hun17, Jah94, Kam93, Kau71, Kau76, Kay01, KL17, Łas82, Lov15b, Lov15a, Mac17, Mac97c, Man71, MS11, MC96, May69, Men76, MO82, Moc91, Nag84a, NNEH11, OS09, PR12a, PR12b, Pog94, Sar96, Sar97, Sch74, Sch76a, Sha96, Tri12, Wag15, Wim15, GWT00e, Sch99c]. Biomechanical [Nik15]. Biomedical [Coh17a, IdMM16, Sch80a]. Biomimetic [BBD18]. Biomolecular [MPC91]. Bion [Gün11]. Biophysics [Mir14]. Biopower [Kay01]. Biosciences [Hoh83]. Biosocial [Coh17a]. Biosynthesis [Lel00a]. Biotechnology [Cap15]. Biotypology [Gom15a]. Bipolar [Wal95]. Bird [HH08]. Birds [Lev76, Mac01]. Birth [Bar94, Che95, Coh88b, Kal90, Pet87, ZC88a]. Birthday [BAG17, BM89, CS79h, CS79q, CS79-30, CS79-34, DPW86a]. Black [Web11, Dav74b]. Black-Boxing [Web11]. Blackstone [Swe94]. Blame [Zim86]. Blessing [Gou89, Hei94b]. Blindness [Zwi09]. Block [Bou17, Saw17]. Blocks [Wis71]. Blood [Toe81]. Bloodletting [Aga03b]. Blumenbach [Ber94]. Bob [Sil95]. Bockemühl [Amr87b]. Bodhisattva [ZC88g]. Bodies [Ano08q, Ano08-47, Bun10d, Čap71-28, Duh02r, Hel99, Lag97d, Lag97g, Lag97l, Lag97j, Lag97t, Lek15a]. Body [Bu96, Bun10k, Fre15, Ito11, Mar78n, Max76a, Wat63, Wol76-36]. Boer [Dol86, Lee86]. Boerhaave [Lõh01]. Bogus [Bun10h]. Bohemia [Nád96b]. Bohm [BS96, Bel96, Čap71-33, Cus93, DH96, Dic96, Fre11, GG96a, Gre74d, Pos95t]. Bohmian [AV96, Ber17a, Ber96a, Bub96, Col96, Cus94, CFG96, Dau96, DGZ96, Fin96, Har96, Ken96, Pag96, Squ96]. Bohr [Alb92, Bel92b, BF94, Che94, CS79h, CS79l, CS79-31, CS79-30, CS79-34, CS79-36, Cus94, D’A01c, Dai96, Fav94, FF94, Fol94, Fol96, Hoo94, HH94, Jon86, Kai94, Kan96, Kat11a, Kat11d, Kat11e, Kra15a, Kri94, Mac94c, Mur94, Rös94, Rös95, Shi83, Shi86]. Bolivia [Iru01]. Boltzmann [Bla95g, Bla95i, D’A01m, Kle74, Koc91, Sch00]. Bolzano [Kas11]. Bomb [Bar01b]. Bond [Wei01b]. Bonding [Ban80, Sch91a]. Book [Bla92b, Bla95g, Che05b, Duh02t, Duh02w, Duh02x, Gou06b, Gua96a, Haq94f, Wol17, Wol76-39, Zam92j, Sai04a]. Books [Ach80, Aga03-34, Hol95b]. Boon [HL17]. Bootstrap [Aga75r]. Bootstrapping [Grü05a]. Bootstraps [NMN02]. Borderline [Gäh89]. Borders [BS15]. Bordet [Zal86]. Boris [FM09a, FM09c]. Born [CS79h]. Borromean [Ara97]. Bosch [dV74]. Boscovich [Hei15, Whi99]. Boson [DH96, Kal96]. Boston [CW69a, CW69b, War63, CW67, CW69a, CW85]. Botanical [San92a]. Botanizes [Das15]. Botany [Pin01]. Both [Aga03-32, Rai11, Tac17, Wor89]. Bottle [Bar89]. Bottles [Gol87]. Bottom [UM86b]. Bound [Dra99]. Boundaries [Mil15, Sch08d]. Boundary [Ano08i, EP17, Mor09]. Bounds [Ano08j]. Bourbaki [Zad96]. Bourgeois [Fre86g, Spr91a, Vrc79]. Boutroux [BB71]. Boxing [Web11]. Boyle [Aga13d, Aga13q, Cha09b, Chr14]. Bradwardine [Cel08, Syl08]. Brain [AB12, Bun10c, Bun10m, ÉA99, Ges69a, Ges74h, Ges74m, Ges74r, Ges74u, Ges74w, Gro82g, Gro82m, TB96, Tho01, Wol18]. Brains [Sha93]. Branch [MC80y]. Branched [McC95a]. Brand [Mar79c]. Brandeis [Sch03]. Brandt [NPS17]. Brazil [Dan92, DFL11, DS01, GPSR01, Pat92, VdOM11]. Brazilian [Gom15a, KV11, Lop92]. Bread [Rus09, Rya94]. Break [Sch11]. Breaking [Jan89c, Wal07b]. Breakthrough [Kar10g]. Breast [Seg15]. Breeding [Hei08f, HD95]. Brentano [Bla95c, Pot97]. Brésil [Dan92, Pat92]. Breslau [Ges74e]. Bressan [Bre74c, Van74a]. Bridge [Ela11j, Gäh89, Gol79, Kah91]. Bridges [SV18]. Bridging [Bun12a, CCG93]. Bridgman [Čap71-31, Grü73w]. Brief [Ano08k, Goo95, Kat11a, Wen96]. Briefly [Bun93d]. Bright [Ken17]. Bringing [Bid11].

Britain

[Ara15, FG92, Hei15, Jon91, Lub15, McG87, SR87]. British [Kum92, San92a]. Broad [Zem09]. Broglie [Čap71b, Čap91a, DH96]. Broken [Wan17]. Bronson [Kon88]. Brooklyn [Kah91]. Brooks [BC94]. Brought [Sta94a, Onn91j]. Brouwer [Čap71-40]. Brown [Rit01]. Bruno [Höf76]. Brutes [Gre74m]. Buchdahl [Com84]. Buddhism [ZC88f]. Buddhist [Baa92, Eck98]. Buffer [Yeh17]. Buffon [Hoq18]. Build [Gro82g]. Building [Gom15a, Wis71]. Bulgarian [Gin03a]. Bunge [AC82, Wet82]. Buonamici [Hel08]. Burdach [Pog94]. Bureaucracy [Mić79, Tad79]. Burgerlijk [RV74]. Burning [Ras95]. Burtt [Vil02g, Vil02b, Vil02a, Vil02c, Vil02m, Woo02]. Business [Jar86e]. Butcher [Kit74]. Butts [BM89, Wol89]. Bypassed [Ror86]. Byproduct [DD12]. Byzantine [DN94].

C [Jan89a, NA07, Rub88, Har97b, Onn91d]. C. [GP79, Tis74]. ca [FG92, Vam09a, Vam09b, Vam09c, Vam09d, Vam09e, Vam09h, Vam09l, Vam09m, Vam09n, Vam09o]. Cahn [Rhe05]. Cai [Pin96]. Cailletet [Pap15]. Calculating [Dam96e, Vog04]. Calculation [Hac91]. Calculational [Gou89]. Calculative [DD86]. Calculatores [Wal81b]. Calculators [Syl91]. Calculi [AT15, Ser81]. Calculus [Bec91b, Coh74a, Phi94, Rei96, Sig91, Van67]. Calendrical [Chu11]. Calibration [Fra99b, Van83]. California [Bla95c]. Call [Aga03x, Bla95d, Cha12f]. Called [D’A92, McC74, Sza79a, Kis02]. Calmly [Byn89]. Calories [Hei08a, Hei08f]. Cambridge [Flo17b, Nav15, Sar92c]. Came [Krü88]. Can [Aga75b, Aga76b, Arm88, BE88, Bar15, Bun12b, Bun12c, Byn89, Byr83, Caw89, CLR^{+}92, Coh76b, Coh02d, Coh02e, Enç05, Fal86, Fra99c, Gav96, Gou94a, Grü73d, Grü73e, Hal71, Mei17b, Mic18, Nic80a, OJ14, Sal18, Sha84t, Tac17, Tis03, Wet09, Zwi08, Kle91, Le 08]. Canada [Duc92]. Cancer [Hoh83, Seg15]. cane [BdOU^{+}95]. Cannizzaro [Cha09k]. Cannon [Bu96]. Cannot [Coh16, Mer97]. Canon [Che10, Hei94b]. Canon-Maker [Hei94b]. Canonic [Aga86]. Canonical [ZC88b]. Canons [Mar89, Pan83d].

Can’t [Caw89, Set95]. Cantica [Sto75]. Caoutchouc [Hei08b, Hei08f]. Capacities [Aun71, Fis71, McM71]. Capek [Shi87e]. Capital [Mac97f, Mac97h, Mac97i, Now79a]. Capitalism [Sas17]. Cardan [Duh91h]. Career [Jar86e, Kar10f, ZM11]. Careers [Hei08c, Hei08f, Kra92b]. Carl [Die14, Ges74e, Mil11b]. Carnap [BC71, BIT01b, Dav97, Grü73h, MC80t, Rec11, Tso15]. Carnot [Nee11a, Nee11b]. Carpenter [Hil94, Sac83]. Carriers [Hol95a]. Carry [Voi96]. Cartesian [Gre95, Mor89, Rou04, Čap71-35, Don86, Lec86, She89, Zol89b]. Cartography [Var92]. Carus [Pog94]. Casati [EF18]. Case [Aga87c, Ágo11, Ano05c, Are89b, Bro95, Cap15, Cha12b, Cin89, Cor00, De 74, Fet81i, Fin80q, FC16, Fu93a, Fu93b, Gab01, Gru81, Hon16, Kal90, Kin16, Kos94, Kra92a, Lev07, Lim94, Mir14, NB12, NB13, Nic80e, Nol08, OP82, Omo18a, Par09a, Pie16, Pit87, Rad12e, Rol16, Rus80, Sab92, SS16b, Sau03, Sha16, Sil89, Sim11b, Škv88, Tri12, Wal84, Wol95l, Wüt16, Zah78, Zam92c, dV11, vD14, BLA95a, Wan18a]. Case-Studies [Cha12b, Sil89]. Case-Study [Rus80, Zam92c]. Cases [KL16]. Cassini [Coh90]. Cassirer [Hei11]. Catalan [NG03]. Catalogue [Hoq18]. Catch [Wol76-32, Wol76-30, Wol76-31, Wol76-28, Wol95d, Wol95f, Wol95g, Wol95j, Wol95h, Wol95i, Wol95e, Wol95m, Wol95l]. Catching [Pet03b]. Categorial [Ant90, Bar76c]. Categorical [Flo15, Mas15a]. Categoricity [Pol16]. Categories [Hag16, Han99e, Koh98b, Mac12b, Mar84b, Mar84a, Tam86, Wit86]. Categorization [Gur03]. Category [AM83, LMR95, Law96, Möl01]. Category-Mistake [Möl01]. Catholic [Bre11, Ful02]. Cauchy [Ana90b]. Causal [Ano08l, Ano09j, Ano09k, Bun93a, Bun93b, Coh80, Cus96, Dem05, Dra88, Eel10a, Fet81i, Gje07, Grü73f, Grü89b, Hau10, Hef87, Mar79l, MC80b, MC80g, MC80i, MC80j, MC80w, MC80t, MC80u, Mit76b, Nor12, Pâs15, PV90, Pes11, Psi07, RS69, Sim77d, Sim77p, Sim77u, Toa15, WvB07]. Causality [Aga81-37, Ber74a, CR81, Čap91p, CS79k, CS79-43, CS79-46, Coh80, Eel10a, GG90, Gou83, Hau10, Hav69, Mad69, MC80q, MC80z, Müü01, Reh99, Sak98, Tor83, Wal81j, Wat98, dB87q, Sta69]. Causally [Ber17a]. Causation [Ano09l, Ber17a, Den99a, Den05, Fay97a, Fet81h, Fit71, Hen06, Kra82a, Mar12, Mar13, Poe99, Spo09, Tur86b]. Cause [Bou02c, Bra87, BEW96, Den95, Mor63, Sim77e, Tur86g, Tur86h, Tur86l, dB00]. Causes [Ano09q, Bin96, Bou02d, Bro86, Enç05, Gre74e, Kat06g, Nol03c, Tuo76, Wal81c]. Cautious [Sil89]. Cavendish [McC95b]. celebi [Kil03]. Celebration [CT98]. Celestial [CS79c, Gra91, Kra91, Pao84]. Celestiall [Gal02c]. Cell [Fag12, SA12, Tri12]. Cells [Gro82h]. Cellular [Ger15a, HH89]. Centenary [Kap80d, vH77a, Sar92a]. Center [Das15, Duh91j, Moo67, Nen18, Sch95b]. Centers [NG15a]. Central [Haq94e, Ret74, Svo15]. Centre [BS03, Kil03, Nav15, Pal03, TP90]. Centres [Lei03]. Centrobaric [Nen18]. Centuries [Bla94, Cha14, Tab03, dB87r, dB87s, Bla04, McK75, Ras75, SDG15b, Wel15, Wil18]. Century [Bar02b, Cho14, Cin89, CT09, Fre86i, Fre86j, Har10, Hel08, Hoo04, Jhu07, Mau86, Ric91, Vil08, Wis71, D’A01l, Ara15, Are89j, Arm18, BR18b, Bro08, CALP03, Car15b, Chr14, Cif05, Coh90, Coo11, Cro92b, D’A81, D’A00, D’A01n, D’A01o, Das91, Duh91l, Duh02i, Duh02p, Duh02o, Duh02q, EP17, FEE18, FS96, Fra91, Ful02, GP03, Goo01, Gre75, Gre91b, GHQ96b, Hal90b, Hec01, Hey86, Hor93, Irz15, Jam92, Jol18, Jon91, Kah91, Kel10, Kle01a, Koy76b, Kra74, KF86, Kra92a, Kuz87n, Lef01a, Lek15a, Lek15f, LvHV92, Lop92, Lub15, Men76, Mül91b, NB12, NB13, NRS92, PT04, Pat03a, Pau17, Pes91, Pop86a, Rai10, Rau91, Roc01, Sal92a, Sas03k, Sas03j, Sch75, Sil11, Svo15, Tur86l, Van08, Wal81d, Wal81b, Wal81e, Wal81q]. Cerebellar [Gro82h]. Cerebral [Ges74n]. Certainty [Ano06i, Cha18, Jan89d, Jan89i, Mor89, Tha86a]. Cesalpino [Hel08]. Ceteris [Ano09-44, Llo15]. CG [BT96]. Chain [Pap14]. Chair [Ace92]. Challenge [Ano07s, Ano08-30, But93c, Hör91b, JW14, Par09c, Rös94, Sch09, Sch18, SG11, Sta09, dF14b]. Challenged [Des03]. Challenges [Har01, Jhu07, Rav18, Rod15, Wia79a]. Challenging [Büt08, Gov15]. Chance [Ano09m, Das91, Gay13, Smi82, Spo10]. Chances [Bar10, Das91, Pie95]. Change [Ari76, Bal08, Bar08, Bar01a, Bro86, Bur87a, But93b, Čap71f, CN11b, Emb92c, Est96, Fre11, GGN89, Ger15a, Hoo04, Hör89, Jar86w, Kor71, Lov15b, Lov15a, ML99, Nic08, NG01g, Obe75, Onn91e, Onn91g, Phi90b, Rus76c, Sch11, SM90, Sha80, Sha84b, Sha84f, Sha84p, Sha01, Sie71, SSHH08, Tou74b, Tua16, Wil12, Yin16, Onn91c, Onn91d, Onn91j, Ly08, Obe08]. Changes [Gua96a, GHQ96a, Nee11a, PM82, PP87, Sta94a, War96]. Changing [Ano09n, Bog15b, Car01, D’A01b, Feh99b, HB76, Lec94, Mar11b, Mar01, Nai95, Pal15, Rip11, Sab92, Sch06a, Svo15]. Chaos [Bla95j]. Chapter [Bus74, Pet76, Wal95, Wol76t, Wol76u, Wol76v, Wol76l, Wol76m, Wol76n, Wol76o, Wol76p, Wol76q, Wol76r, Wol76s, PP15b]. Chapters [Jie96]. Character [Ano09o, Coh02l, Die83a, Fin80l, Jar86v, Jar86o, Mur75, Nov74, Ric07, Sco80, Sha80, Sha84b]. Characteristics [Kał82, Lia96a, vL16f]. Characterization [Gay09, Hil89]. Characterizing [STNW12]. Characters [Gre74f, Gre74g, Gre74h]. Charcot [Sav15]. Charge [CS79l]. Chariot [Jin96]. Chariots [Kim11]. Charity [Fin17]. Charles [Ano05c, Gru80]. Charting [Fre11, Wes94]. Chase [Aga03-36]. Chemical [Abb94, ÁNS91, Bal15, BR15, Ban80, BV01, Bhu06, Cha09f, Cha12g, Duh02h, Duh02b, Duh02c, Duh02d, Duh02e, Duh02m, Duh02p, Duh02o, Ear06, Fra12, Fra13, Ham07, Har15b, Het15, Hol95a, Lev90, Lom15, Mah15, Nee06, Roc01, SH91, Syn82, TBC91, Von84a, Wil01, vSB91, Fre91]. Chemist [NG01b]. Chemistry [Ace92, BSM06b, Bar91b, BV09, Cha09h, Cha15, Duh02n, Ear15, Fra01, Gav96, Gro01, Hel00c, Hen06, Hoh83, Hun06, Jol18, Kar10c, Kar10b, Kar10d, Kle01a, KS83, Llo15, Mah15, Man15a, Nye91, Nye01, Pas01, Pau17, Pec94, Ram01, RS06, Sce93, Sce01, Sce06, SM15b, Sch06b, Sch15, SB15, Van06, Vem06, Vih01a]. Chemists [Cha12d, Jon91]. Chernyak [Tau91f]. Chess [Whi01]. Chez [You74]. Chiara [Sta83a]. Chicago [Vil02d]. Chiefdoms [Dre92]. Children [Ges74l, Mar76b]. Chimeras [Wil91]. China [BZ96, Bré14, BE10a, BE14, Clu05, Dai89, FQ18, GHQ96b, Har01, Hsü92, Jam92, Jam10, Kan96, Lia96a, LMZ18, LY96, Nia96, Pin96, Sch88, Wan18a, Wei96, Wen96, Xiu96, Yan93, YCMY18, Yus96a, Yus96b]. Chinese [Bra74b, Bu96, Che10, CW18, Chu11, Coo11, DC96, DL05, Har10, Hor93, Lue18, Nia96, Rav18, Wan18c, Wen96, Yuz96, ZC96]. Choice [Aga81c, Bou03, Bre09, Bun95b, Das88, DR15a, Els88, Nor13, Tve86]. Choices [Sim77v]. Chomsky [Pia82]. Choosing [But93f, Har08a]. Christiaan [Vil04]. Christian [BCK07, Gre07, Pop88]. Christoph [Sas03i]. Chroniclers [Aga08a]. Chronometry [Jan85a, Jan85b, Jan85d, dB87t]. Church [Par17a, San18, Web83]. Ciencia [RH14]. Circle [Ano06-29, Ber13b, Ber13c, Dil06k, Gel89, Coh94g, Fis99, Hal91a, Köh91a, Mil11a, Rut91a, Rut91b, Sta91a, Sta91b, Ueb91b, Ueb95]. Circles [Car89a, Dun96, KT93]. Circuits [Ara15, Onn91l]. Circular [Vil08]. Circularity [Bec91g, Car89a]. Circulating [NG01b]. Circulation [Toe81]. circumstance [Kis02]. Citizenship [Iru01]. City [Ihd79h]. Civil [BN17, Fre86d, Fre86h, Fre86s, Hes07, JJ07, Yeh17]. Civilization [Sab92, Sko76b]. Civilizations [NNW74]. Clagett [CS79n]. Claims [Mül94, Tau98]. Clairvaux [Sto75]. Clarification [Men70]. Clarity [She89]. Clarke [Sch01]. Clash [Kuz87g]. Class [Hoc79, Hor16, JN79, Kuv79, Ser81]. Classes [Gro82h, Kau74, Mas15a, Zin73a]. Classic [Fin87, Sab92]. Classical [Aga86, Aga13p, Ana97, BHN98, BB00, Bun10d, Čap71e, Čap71g, Cla89, CŻ82, DP92, FRSV18a, FM09b, Gra95, Gro82d, Grü73m, Hee74, Hel00a, Hoo74, How94, Hun06, JM06, Kuz87f, Mac12d, Mac12c, Mac12i, Mey76, Mou89, Pos95c, Ras95, Ren07a, RSS^{+}07a, RSS^{+}07b, RS07b, RS07a, Roc11, Sch07a, Sim77g, Sto91, Tem95, Thi11, Wet87, FRSV18b, Ras92]. Classical/Quantum [Mac12d]. Classics [Chr04a]. Classification [Gro82a, Pra76, Uye97]. Classificatory [Kri90]. classique [Ras92]. Classroom [Pur93]. Claude [Die14]. Clause [Llo15]. Clausiu [Nee11e]. Clavius [Sas03i]. Clazomenae [Vam09a]. Clebsch [Tob96]. Cleopatra [Hoo99b]. Clericalism [Fin17]. Clerk [Duh15c]. Climate [Bun12d, Kow16b, Kow16g, Kow16j, Kow16l, LW11, Tua16]. Clinical [FA76a, Sch83d, Wei14]. Clock [Duh76e, Grü73o, McR80]. Clockmaker [BAK11]. Clocks [Has11]. Clos [Fra12, Fra13]. Close [Eng01]. Closed [GL03, Mor74, Syn82]. Closest [Mar95d]. Closet [Smi95a]. Closing [Ano06-29, Dil06k]. Closure [Dum92]. Cloud [Sco80]. Clubs [Lub15]. Clue [PW17]. CNRS [Bel15]. Co [Čap91h, CS79-41, Koh98e, Pat76b, Wim92]. Co-Dependency [Koh98e]. Co-opted [Wim92]. Co-Ordination [Pat76b]. Co-Presence [Čap91h]. Coarse [Grü73m]. Coarse-Grained [Grü73m]. Code [Gro82g]. Codes [Gro82n]. Coding [Gro82a]. Coffer [Gün11]. Cogito [Coh86b]. Cognition [Bun10e, Chu00, CS86, Dam96c, Gel76, Gro82m, MV80a, MV80m, Ton97, Tym95]. Cognitive [And09b, Blo86, Fen12, Fer85d, Gen03, Gig03, Gin03b, Gro82g, Gru81, Gur03, HH97, Hil89, MS11, MC96, MV80b, MV80c, RDS^{+}18, Roc99, Sch02b, Ska92, Ste88, Ton73a, Tve86, WP96]. Cohen [BC94, Coh02j, Edm95, GSW95a, GSW95c, GSW95b, Mue95, Sil95]. Coherence [Ano09-48, JS03, Leh00, Spo09]. Coherentist [Ano09-42]. Cohomology [Pos95o]. Coincidences [Pos95s]. Cold [Chr14, Gav09, Gav14b, Gav14a, KJ14, Pap14, Wat01a, dS14, Ada01]. Colegio [Rui95b]. coli [Fos91]. Colin [Dou03, Nii03]. Collaboration [Lec14]. Collaborations [Cho14]. Collaborative [Wyl15]. Collaborators [Zam92a]. Collapse [GS82, Sab92]. Collapsing [Whe74]. Colleagues [DJ89]. Collection [Che10, Spo09, Lan90]. Collections [Cla10]. Collective [Tur86b]. Collectives [Gie86]. Collegio [Wal81c]. Colloquium [CW69a, CW69b, CW85, MS75b, War63, CW67, UM86a, UM86c, UM88, UM92]. Colloquy [Byr02]. Colocation [Ric15]. Colombia [Obr92]. Colonial [AB14, Kri92, San92a, Kum11]. Colonisation [Dor92]. Colonization [Dor92]. Colophon [Vam09o]. Color [Ano09o, Böh87, CS79u, Ges74f, Gög87, Kre91, Mil87]. Color-Naming [Ges74f]. Colored [McG96]. Colorimetry [Tur96]. Colouring [NG01c]. Columbia [Vil02e]. Combination [Duh02m, San15]. Combinations [Duh02r]. combinatoire [Ras74]. Combinatorial [Ras74, Ras94e]. Combined [Ste79]. Combining [Coo90]. Come [Gig03, Wil00, Was91]. Comets [Gen91]. Coming [Lao95]. Commandments [Ber95b]. Commensurability [Boy01, Buc94a, Kle71, Ner89, Ner01]. Commensurable [Pan83d]. Commensurate [Nye91]. Comment [AA86, Aga87a, Aga87b, Bar92, Cam03, Com84, Dol86, Dud86, EF86, Ezr86, Fel92, Gai86, Hey86, Irz03, KP86, Lip86, Mal92, Mar86c, Mar92, Mau80, McM74b, Min67, Pit92, Réd03, Rot86a, Sac83, Sca03, Sch86e, Sch86d, Shi87a, Shi87e, Shi87b, Shi87c, Shi87d, Shi87g, Shi87f, Stö03, UM86b, Wol03, Yan98, Zem86]. Commentaries [Che10, Kel10]. Commentary [Ber14, GHQ96a, Haq94h, Hee84, HH08, Jan84a, JRN^{+}07a, JRN^{+}07b, Le 08, Ly08, MM94, Nee11c, Obe08, Ras00, Sch84, Wal92e, Web08, vH77a]. Commenting [Bu96]. Comments [Ach67, Aga69a, And03b, Bal03, Bre74a, Bre74c, Bur83, Čap71u, Coh63, CS79e, Di 84, Dou03, Fey78b, Fri76, Gil03, Goo63, Hüb78b, Jan84b, Mar67, May69, Mil67, Nii03, Pap03, Pei03, Pos74, Put67b, Qui63, Sal71, Sch75, Shi63, Shi76, Sol84, Sol63, Sta69, Sta74b, Sta74a, Sta83a, Sza03, Tau91b, Tau91c, Tau91d, Tau91e, Tau91a, Tau91f, Tau91g, Tho67, Tis67, Tis74, War67b, Zal86, Zuc69]. Commerce [Car11]. Commercial [IdMM16]. Commercialization [Bid11, Wis11]. Commissures [Ges74j]. Commit [Coh02a]. Commitment [Han95, Kle71, Lev94, Mol18]. Commitments [Mac74, Rav91, Sta74a]. Common [Flo17b, Pro00, RB97, SM90, Sim15, Unt01, Wat96, You95]. Common-Sense [Wat96]. Commonsense [Aga03q]. Communication [Aga81-32, Can16, Den99c, Hil07a, Hil07b, Hil07c, Koe84, Krü91, Lam97, Mar84c, Mić79, Mor80, Nei15, Sau07, Ton81d]. Communication-Oriented [Krü91]. communications [Onn91f]. Communicative [McC94]. Communis [Lõh01]. Communism [Hor88]. Communist [Yuz96]. Communitarianism [Koc09, Ste09b]. Community [GSW95c, Har02, Hei91, How15, LvHV92, Nic89b, PB07, Sch95c, Sig94, Sim94, Wes94, Wol76-33]. Comparability [Ost76, Sch76c, Tow71]. Comparative [DP92, Fu93b, Gli87a, GWT00b, Nak95]. Comparativity [Fu93b]. Compared [Čap71v]. Comparison [Böh83, Dun96, Grü73w, GHQ96b, Han99e, HR74j, SSHH08, Wat78a, Xiu96]. Compass [Yav17]. Compatibilism [Gar12b]. Compatibilist [Abr11]. Compatibility [Čap71x, Pos95d]. Competing [Ada11, Ano08m, Hal71, Hop74, Lev71]. Competition [Gro82c, Gro82e, Gro86, WPJ95, Bou02a]. Complementarity [Bel92b, CS79d, CS79o, CS79-44, Cus94, D’A01c, Fol94, Jon86, Kat11d, Kat11e, Kat11f, Kat11g, Kau74, Li96, Rös95, TB96]. Complemetarity [Bit96b]. Completely [Čap71-35]. Completeness [Ana90b, Bal18, Res79]. Completion [Boc96, Swi96]. Complex [BBD18, HCP18, Kap80e, Kap86a, Mar84j, Sim77q, SC14, Wer69, Zin73g]. Complexities [Ada11]. Complexity [Bed09a, Bör93, Bro74, Fin80c, McG87, Rus09, Skl76, Wim74, Wim76]. Complications [Gou89]. Component [Kau96]. Composition [AT15, Coh02p, Duh76d, Gou06b, Pal12, Pal13]. Compossibility [Grü73b]. Compounding [Gal02c]. Compounds [Kar10g]. Comprehension [Kno94, Nic16]. Comprehensive [BAC14, Dvo91, Zim99]. Comprehensively [Chi17]. Compressibility [Knu07]. Compressible [Lag97b, Lag97k]. Comprising [Pan91h]. Compromise [Boy95]. Compromising [Oya09]. Compton [Stu76]. Computability [CPM16, Dor00]. Computation [Ayd05, Coo17, Duw17, Mac17]. Computational [Hum11, Mac86a]. Computer [Bun10c, CL17, Cro69, Fil17, Ihd79b, LA86, Len86, Mah11, Ric17, Sch86a, Win86]. Computers [Duw17, Mit86b, MH86, Sim77-27]. Computing [KJ14, Win17b]. Comte [Sch02a, Tur86a, Mey91f]. Con [Syl97b]. Concealed [Bou02h]. Conceivability [Uga18]. Conceiving [Fen12]. Concept [And90, Bec91s, Bou02a, Bou02f, Bou02i, Buc67, Bun10g, Čap87a, Che11, Chr07, Coh76a, CS79-45, Coh02k, Coh02u, Dam96d, Dam96h, Dam96g, Dov84, Ech95, Ela11j, EN96, Fon96, Fre86i, Fre86j, Fre86k, Fre86l, Fri90b, Gro09b, Grü84, Han07, Han99a, Hec01, Hoc79, How94, Kol74, KC95, Liz96, Mey91q, Mis82b, Mit76c, Mot82, Mül94, Nag98, Ned90b, Ner01, Pat03b, Pei03, Pet79, Por87, Rap81b, Ren96, Sal74, Sch98, Sch06a, Sha84c, Sha84j, Sta74d, Sta86b, Ste15a, Str82, Str74, Ton73a, Ton81a, Tuo82, Tym15, Wal81d, War79f, WW03, Wim71, Wro88, You74, ZM11, ZRKC03c, ZRKC03d, Żyt82, dB87q, dF81]. Conception [Akr09, Ano08n, Bec91k, Bie95, Bun93a, Cas96, CS79e, Coh02q, Coh04, D’A01m, Gre84, Hon09b, Klo79, Lau74, Maj98, Nag84a, Per94, PT99, Ska87, Ste03, Val10f, Vla90, Wes90a]. Conceptions [Abb14, AB12, Bec91c, Dro13, Edg15, EF18, Fre86d, Grü73w, Hül96, Kra82a, Mil79, Pia82]. Concepts [Ano09p, Ano09n, AN15, Bal90, Bal03, BMP09b, Bar16, Ber83d, Bog15b, Bre74a, Bri15, Bro74, Bun12m, Čap76, Car01, CS79j, Gav90, Han07, Kóh90, Krü91, Kün95, Len63, LY96, Mac12b, Mar84b, May76, Met89, NG09, Nie74, Pos95l, Sar08, Sha84e, Sha84n, Spo09, Sta94a, Ste09a, Str76, Sup74, Ton73b, Ton73f, Ton81g, Ton81j]. Conceptual [Are89a, Are89e, Are89g, Are89i, Art15, Bar01a, Bec91r, Bel12g, Ben06, Buc84, Bur87a, But93b, CN11b, CF11, Coh02g, Don98, Ela11j, GF97, Gol89, Jar86i, Kai94, Lim94, LF93, Lov15b, Lov15a, MC80c, Min67, Sar91, Tou67, Tou71, Tou86a, Com84]. Conceptualizing [Sla89]. Concerning [Ber74a, Hei07b, Jim95, Kal90, Koç05, Mar79s, Mie07c, Mis74, Pan91i, Pol95, Ste97, Tau91h, Ton73f, ZRKC03a]. Concerns [Aga95c, Wat01a]. Conciliation [Jan89g]. Concluded [Wol76d, Wol76h]. Concluding [Čap71h, Fey91, Kar10i]. Conclusion [Aga13e, Bel12a, Bit96c, Bun12e, Duh02f, Duh15a, FA76b, Han99b, MC80d, Mur12b, Sas03b, Vil02f, Zam92f, Zin73b, Zol89a, vL16b]. Conclusions [D’A01a, Hei08d, Mac97g, Mar67, MV80d, dB87k]. Concordia [Ste12]. Concreta [Goo77b]. Concrete [Čap71n, Kos76a, Kos76b, Tha86g]. Concreteness [Fin80b]. Concreton [Sie84]. Condemnation [Zam92k]. Condemnations [Zam92i]. Condensed [Duh02g, Hoy74]. Condillac [CS79f]. Condition [Fey78b, Wat76]. Conditional [Ano09-45, Zin73c]. Conditionals [Sta76b, Von76]. Conditioning [Rot87a]. Conditions [Ano05n, Ano09-44, Kam93, Len87, Mar84c, Nie82, Rog71, Sch76c, WG11, MA83, Rat18]. Condorcet [You74]. Conducted [BC94, Wet17]. Conducting [Led18, Onn91l]. Conduction [Duh15b, Ges74g]. conductive [Onn91h]. conductivity [TO91]. conductor [Onn91k]. conductors [Onn91j]. Conference [FKR99, HFD88, Sch96f, CS79-38, CHR96, Gua96c, PQQSe96, Stu95]. Conferences [Mis95c]. Confidence [Kid17]. Configuration [Gon15b]. Configurations [Lec94]. Confirm [Tri12]. Confirmation [Dil06j, Erw71, Fet81b, FH10a, Grü05a, Lin93a, Lin93b, Max76a, Roz71, Say05]. Confirmatory [Sim77r]. Conflation [Kow16d]. Conflict [CS79h, Kre79, WW90]. Conflicting [Gil88, War79b]. Conflicts [Wan18a]. Confrontation [Rad82, Shi95]. Confronting [Jer15]. Confusion [Aga75c, Aga75d, Cha12c]. Confusional [Ges74-27]. Congruence [Grü73w, Mas71]. Conic [Ras95, Ras00, Sai04a]. Conjointly [Che05b]. Conjunction [Duc89]. Connect [Kar17]. Connecting [Het15]. Connection [Bun93a, Car07, Duh02t, Duh02w, Duh02x, New15, Nol03d, Nol03c, Nol03m, RS69, Tso15]. Connectionism [Hou93, Ska92]. Connections [Den99g, Men70, Nee11d, Nol03l]. Connexion [Fox15]. Connoisseurship [Hei07a]. Conscious [Pro94]. Consciousness [Bun10e, Čap91n, Mey16, Pag96, Sch89, Sch10, Tha84a, Tha84b, Tha86g, Tru94]. Conscription [Mei17b]. Consensus [Aga89, Can16, Gro82e]. Consequences [Are89a, Cas96, Gro82l, Hal91c, Lee86, Mic98, Pan83a, Per91]. Consequent [LC07]. Conservation [Ela11b, Ela11e, Nee11i, Ord96]. Conservationists [Bar01b]. Conservatism [Jan89e]. Conservative [Luc88, McA97, Pos93]. Conservativity [Str18]. Consider [Le 08, Rus80, Zwi08]. Consideration [Dil06l, D’A01b, Jam79, Met89, Sas03f]. Considerations [CS79-46, Ein01, Fin80c, Has97, Hol83, Lor07a, Pis15, She87, Sch99c]. Consilience [But93b, For10, Har89a]. Consistency [CS79-39, Grü95b, Neg81, Zam92n]. Consistent [Ken96]. Consolations [Aga03c]. Constancies [Gro82f]. Constancy [War79q]. Constant [Bek86, Sos74]. Constants [Grü05b, Lag97f]. Constituent [RS07c]. Constitution [Gin02, Gin06, Gol83, Mar86h, War96, Win83a]. Constitutionalism [Tac17]. Constitutionally [MA83]. Constitutions [RT01]. Constrained [Lag97l]. Constraint [Bri15]. Constraints [Coh76d, Coh86c, Fra01, Nic80a, Tul03, Wie12]. Construal [BM88, Hem88]. Construct [Neu89]. Constructibility [Nai15]. Constructing [BS03, Gar95a, Mar84f]. Construction [Ano08-29, Bro91a, D’A01d, Dar80, Fet81c, Ham07, Hül96, Kae90b, Kap80f, Kno04b, Koh97, Lep80, Löw91, Mos00, OS09, Rud86, Sch74, Ste09c, Szt79b, War79a, Wie12, Yav17]. Constructional [Goo77c]. Constructions [HR74i]. Constructive [Fis95, Gau95, Jan85e, Joh96b, Mil15, Neg81, Nic89a, WP96]. Constructivism [Fra99g, Jan97]. Constructivist [Kom99]. Constructs [Len82]. Consultationes [Zam92i]. Consumer [Pap14]. Contact [Cha09i, Cha09j]. Contagion [Hod99]. Container [Čap71f]. Contemporaries [Mul98, Zam92a, Zam92h]. Contemporary [Aga95a, Alt14a, Ber74a, Ber70, BG09, Čap71-27, CW74a, Coh76c, Deb09, FF94, Fen12, Grü63, Grü86, Han15, Hul76a, KPC90, Kai94, KT76, Lau09, MG11, Mai16, Man81, Now79b, Phi90a, Roc11, Sha84s, VD92]. Content [Čap71i, Che05b, Fey78b, Fin80g, Fin80s, Fin80r, Mar78k, Pan91d, RH92, RT89, Wat78a, Wal92c]. Content-Comparison [Wat78a]. Content-Indeterminacy [Pan91d]. Contents [Ano09-43]. Contest [Pin10]. Context [Aga00a, Ams75b, Băe15, Bun93c, CN11b, CN11a, Cha14, Chm82, Fel18, Fet81e, Fis95, Gin02, Gin06, Haq94b, Hat80, Hee97, Hie94, Ito11, Kno94, Koe93, Mor80, MS75b, Nor11, Pes91, Pro01, RH92, RT01, Ren07d, San92a, Sie95, Wer01, Wun01, dV11]. Contexts [AN15, BP14b, Fu93a, Pal15, Rot87b, VdOM11]. Contextual [O’C02]. Contextuality [D’E00, Har96]. Contextualizing [VdOM11]. Contextually [Pan16]. Continent [Kim11]. Contingency [Nev98]. Contingent [Bra74a, San81]. Continua [Grü73t]. Continued [Duh91e, Duh91p, Onn91h]. Continuing [Jar09, Nev02]. Continuity [Aga81d, Bai76a, Čap71o, Čap74, Čap91g, Čap91q, Cin89, Fre11, Koc91, Lev16, Nol08, PM82, Sch07a, Vot11]. Continuous [Čap71-41, Car07, Fra09b, Gro82b]. Continuum [Art89, Jim95, Van86]. Contour [Gro82f]. Contra [Lar99, Rut91b]. Contradict [Str84]. Contradiction [Pet03b]. Contradictions [Sup79]. Contrast [Čap71j, Gro82b]. Contrasting [KL16]. Contrastive [Yli07]. Contre [Sar91]. Contribute [San14]. Contributed [Sch01, Van74d, Woo91]. Contribution [Avg90, CT91, CS79-31, Duh91d, Duh91e, Fle91a, Ges74z, Hee86a, Hül96, Jaw90, Kar90, Mat95, Mer70, Ton73f, Ton81f, SVD95]. Contributions [BdOU^{+}95, BVH^{+}83d, BVH^{+}83e, BVH^{+}83f, Cho67, Cho86, Hat17, NNEH11, Pin96, Rat70, Rui95b, Sil11]. Control [Bal15, Čap71j, Gro82m, Web11]. Controlled [Kap80-32]. Controlling [MA83]. Controversial [Ray82, Tri12]. Controversies [Coh04, Hel00h]. Controversy [Aga03r, Ber74a, Ber74d, Bit96d, BJ89, Dem89, Fin09a, Gou90, Jan89b, Kar10a, Kay76, Miś82a, Pro92, Rud86, Sar91, Sch86e, Wim80, Vas92]. Conundrum [Art89]. Convenience [Mcm76b]. Convention [FH10b, Hei00]. Conventionalism [Aga75g, Ams75c, Cus93, Sie82, Str84, Zol89b]. Conventionalist [Hoo74]. Conventionalities [Las01]. Conventionality [Bow76, Jan83]. Conventions [Ano08-38]. Convergent [Pol74b]. Conversation [BC94, Fey91, Gol12, Jan89f, Tym15]. Conversational [Kas76a]. Conversationalist [Coh02d]. Conversion [Don98]. Convex [CL93]. Conveying [Coo11]. Convincing [Sil89]. Cooperation [Cer14, Die14, Gem90, Sch96e, Vog91, Wen14, ZRKC03d]. Cooperative [Pos95i]. Coordinates [Omo18b]. Cope [Hun06]. Copenhagen [CS79-38, Bit96d, Cus93, Fje02, Hei88, Pos95e, Pos95r, Unt01]. Copernican [Bar01a, Bec91h, Fin09a, Fin09d, Fin09k]. Copernicus [Fin09b, Koy76a, Lev83d, Obe75, Wri80, ZRKC03b]. Coping [Ano05d]. Copy [Eco92]. Core [Ano06x, Dil06d, Hin95, RA01]. Corpus [Dal89].

Corpuscles [BR15]. Corpuscular [Čap71y]. Corpuses [Bro10]. Corrected [CS79-30]. Correlated [Enz00]. Correlation [Čap71k, Koç05, Sim77u]. Correlations [Cos87a, dB87n]. Correspondence [Ams75h, Bro93, D’A01c, D’A01d, DF93, FK93, Har08b, Nol91, Now82, Ost92, Pet69a, Pos93, Sce93]. Corresponds [Sau93]. Corroborate [Coh76b]. Corroboration [Fet81b, Fet81h, Nii89, Wat78a, Zah89]. Cortical [Ges74l]. Cosmic [Lay74, dB87f]. Cosmological [Dai16, DR16, Fre91, Nak95]. Cosmology [Ber74c, De 74, Ell84, Gra91, LY96, Mos96, Mos00, Nen18, Sch07a, Smo03, Sta74c, Ung91b, Val96, You76, dB87k, Ell03]. Cost [Mic71]. Cost-Benefit [Mic71]. Costa [Hay03]. Could [Lag96, Nor98a, Van74d]. Count [Cha12d]. Counter [Ams75f, Jan89a]. Counter-Enlightenment [Jan89a]. Counter-Reformation [Ams75f]. Counterexample [Hin76]. Counterfactual [Ran89, SS03, Sim77e]. Counterpart [Fer85a]. Counterproposition [Are89c]. Counting [Hac94a]. Country [Bar14a, Mar01, RH14, San14]. Courage [Saa17]. Course [Ano05o, Tel08, Val10c]. Court [Seg91]. Courts [Aga08a]. Covariance [Sau03, Sch82, dB87b, dB87c]. Covariant [Nor07d]. Covering [Add76, Hem76, Rec11]. Covering-Law [Add76, Hem76]. Cows [Voi96]. CPS [LL14]. CPT [Cos87a, dB87o, dB00]. Crafted [Gün11]. Craigian [Lin93c]. Cranes [Ros81]. Create [Fax97]. Creating [Bar91b, Gil91, PB07]. Creation [Gol69a, Hac91, Koh98a, Pet92a, Rus13, Pet92a]. Creationism [BZ96]. Creative [Ano05e, Ano05g, Cha09f, Cla02, Gru80, Jan89j, Kap80-42, Kle01a, RH14, RB97]. Creativity [And84a, Ano05f, Gru81, GB05, Jan89j, Jar86t, Wal95, War80]. Creator [Coh94d]. Credit [Koe95]. Crescas [Jam76]. Crete [Das15]. Crick [Sch76b]. Crime [Škv88]. Crise [Fre04]. Crises [Bun12c, McA97]. Crisis [Ams92j, Fle86b, Fle86a, Fre04, Hoo94, Kno89, Lar13, Sch95c, Tha84c, TB15, VK83, Wol76-37]. Criteria [Aga75e, Aga03j, Ams92a, Bre09, Kal90, Pal01, RA78a]. Criterion [Bun12e, D’A01d, Grü95b, Pos76b, Ton81b]. Critical [Aga95b, Aga03-33, Alb84, And84b, And09a, Bab99c, Bab99b, Ban17, Bas14, Chi17, Coh94b, Fel92, Fey84b, Fin09b, Gay09, Grü73w, Grü73v, Han91, Haq94a, Hüb78b, Jan89a, Jan89i, JL95a, JL95b, Mac97f, Mar76a, dAM11, Mat17, Met89, Mic18, Mil09, Ng17, Nol03b, Pog16, Qui82, Ric17, Sal84, Seg17, She09, Spi99, Sta09, Str84, Swe90, Swi99, Top84, Vic92, War79s, Wat01b, Wol95f, Zem09, BBT95, Duh15c]. Criticism [And84a, Ber76b, Čap71-27, Čap71-31, Fin80c, Fin95, Hüb76, Ito98, Jar86l, Koe71, Plo76, Pop74, Ste09b, Wor78a]. Criticisms [Bur84, Imo15, Lom15, Wat76]. Criticist [And84c]. Criticizing [Ela11c]. Critics [Cur89a, Lak71b, Sch83a]. Critique [Aga03d, Bec91f, Bun92, But93l, Čap91r, Col96, Duh02a, Ezr14, Fey76, Fin80e, Fin80j, Fin80n, Fin80p, Fin80o, Gen71, Gie76, Grü73g, Grü73h, Grü73q, Haa99, Jan89f, Joh96b, Kau74, Kop99, KM99, Kri94, Kuz87c, Lar99, Mar86f, Nol03k, Pia82, Rad12b, Rop83, Rut91c, Sch11, Sch99b, Sko76b, Ste15c, Tau99, Zim99, Zol89b, Zum96]. Critiques [Fin09c]. Croce [Čap71n]. crop [HD95]. cropping [SVD95]. Crossing [Sch08d, SH91]. Crossroad [Sas17]. Crossroads [Ham18, Stu95]. Crucial [CM89, Pet92b, Zah78]. Cruciaux [Pet92b]. Crucible [Vem06]. crucis [Ham18]. Crude [Duh02h]. Crum [Rit01]. Crustacea [GWT00h]. Cryogenic [vD14, Onn91o]. Cryogenics [Die14]. cryostat [Onn91f]. Crystal [Hal95a]. Crystallographic [Sch96d]. Crystals [Hol95a, PP81, Sør17]. Cuba [Bas14, Alo14, Alt14b, AB14, BAC14, BRW14a, Bar14b, Cas14, Goo01, Han14, MC14, MA14, dJAO14, RH14, SC14, Vel14, Wen14, dF14b]. Cuban [Alt14a, AAF^{+}14, Bar14a, Cer14, Gon01, HTG14, Lec14, LL14, Mir14, Tam14]. Cuban/US [Tam14]. Cults [SS74]. Cultural [Aga81e, AM63, Ano06-27, Ano06-33, Bel15, Cah05, CN11b, Dam96b, DD12, Die83a, Elk76, Fel89, Fin14, Gav14b, Gli87b, Gol69a, Jan89f, Kim11, Lek95, Mar78h, Mar86e, MAM69, Moo90, MS75b, Oru95, Vro90, War79b, Woo91, ZC88d, ZC88i, Ano06y]. Culture [Aga03s, Bab99c, Bre07, Cos79, Elk76, Elz95, Gai94, Gen91, Gol79, Jan89b, Klo79, Lav88, Liv92, McG87, Sca90a, Sie01, Spr91a, Urs96, Živ79, Bur87b, GSW95a]. Cultures [Joh17, McK75, Tan01]. Cumulative [Web08, d’E08]. Cuneiform [Bro10]. Cunning [Wol76x]. Curating [Rhe15]. Curiosity [Gün11, See70]. Current [Alo14, Dah90, Duh15b, Hüb78b, Max76a, Sch79, Var92, Vod15, Onn91i, Onn91k]. currents [Onn91l, TO91]. Curricula [LP01]. Curvature [LC07]. Curve [SH91]. Curves [Eve74, Ste09a]. Cut [AT15]. Cut-Like [AT15]. Cuts [Nik74]. Cybernetic [Bec86]. Cycle [NG09, Nee11a]. Cyclic [Pos95b]. Cyclotron [Pos95k].

D [Bal89a, CS79-41, Coh90, Duh02t, Onn91e].

D.D.R [Bie74]. D.D.R. [Bie74]. d’Acclimatation [Osb92]. Dahlem [Wag15, dR15b]. Daily [Bré10, Tho14]. D’Alembert [De 01]. Dalla [Sta83a]. Dalton [Cha09f]. Dance [Pic12]. Dancing [Rip92]. Danger [Kow16d, Pet03b]. Dangers [Rop86]. Danish [Hes07, Jac07, Kra15a]. Dans [Fre04, Gün92, Ras74]. Danto [Kel94]. Dao [Rav18]. Dappled [Het15]. D’Arcy [Gou76]. Daring [All99a]. Dark [Sal01]. Darkness [CS79n, Wic91]. Darwin [Ano05c, Fle91b, GH01, Gre91a, Gre74i, Gru80, Hod90, Koh86, Koh97, Lef03, Lim94, Pan94, PB07, Sar91, Sch86d, Sch94b, dF11a]. Darwinian [Cas96, Nor13, Rus80]. Darwinism [Cou96, DS01, GPSR01, Gli01, Gon01, GWT00a, Gut95, Iru01, Pin01, PS01b, RA01, RT89]. Data [Bun12a, Coh94j, Coh02a, Fin80i, Jos17, Mei17a, Pap94, SG11, Zim86]. Dated [Jan07]. Dating [Wal90]. David [And03b, Ezr14, Fre11, Her14b, Her14d, IG05a, Stö03]. Davidic [McC02]. Davidson [Fis96, Mar94a, Szu00]. Davy [Abb94]. Davydovitch [Kap80t]. DAW [Wen14]. Dawn [Cha09l, Raf15, dB87b]. Day [Duh02i, Jan85d, Mod15]. De-Ruse [Życ95a]. De-Stroyed [Syl97b]. Dead [Čap91r]. Deaf [Zam92c]. Dealing [HCP18, Kel10]. Dear [Wol76w]. Debate [Dil06b, Fol94, Fol96, Gur03, Pie12, Pro17b, Ren96, Sch83a, Vil02b, Whi01]. Debates [Bog15b, Mac94c, Zam92i]. Debating [Row06]. Debt [Cha09h, Smi90]. Débuts [Pat92]. Decade [Sar92c]. Decades [RH92, Rei96, SN87].

Decay [Fra09b, Lip86, Stu86]. December [dF14a]. Deception [Jan89i]. Decide [Hal71]. Decision [Bör93, Bou03, Cou96, Gje00, Gro82c, Gro82e, Gro86, Kau96, Myr97, Pet89, Sim77k, Sim77l, Sun18, Wei10, Zol89c]. Decision-Making [Gje00]. Decisions [Rot87a]. Decisive [Hod90]. Decomposing [Wyl92]. Deconnection [Ges74n]. Deconstruction [Hon94, Kat11b, Kat11d]. Découpage [Gra83a]. Dedekind [Nik74]. Dedicated [CS79h, CS79-30, CS79-34]. Deduced [Lag97g, Lag97h]. Deducible [Mey85d]. Deduction [AT15, Cel93, Dov84, Mey85v, Mey91b]. Deductions [vZ11]. Deductive [BM88, Bun93d, Hem88, Mus89a, Tuo76, Sza04]. Deductive-Nomological [Bun93d]. Deduktiven [Sza04]. Deep [dR17]. Deeper [Att13, Bla92a, Čap71l, Chu91]. Defects [Ges74f, GL03, Grü09]. Defence [Bon87, Fey78b, FL11, Haa99, Prz95, vdW04]. Defended [Coh02d]. Defender [Gin03b]. Defending [Add76, Fin09b, Imo15]. Defense [Bid11, Hoo74, Kiv94, Łod82, Rus76b]. Definability [Sim77n]. Definable [Sim77f]. Defining [Alt00]. Definite [Duh02k]. Definition [Cap81a, Cin89, CS79t, CS79w, Goo77c, Mar84e, Neu76a, Sim77p, Str82]. Definitions [Čap91h, Coh02p, Hin76, May76, Nee11h]. Definitively [Suc82]. Deflationary [BG18, PP15a, PP15b]. Deflationism [Str18]. Degenerating [Per89]. Degeneration [Grü89a, MH01]. Degree [Chr94a, Coh02k]. Degrees [Čap71k, Čap71m]. Deity [McC02]. Delaying [GHQ96b]. Delbrück [Bee91]. Deleuze [Yin16]. Deliberative [Kow16h]. Delineation [Bry98]. Demands [Mey91q]. Demarcation [Ams75i, Aya96, Bar84, FA76b, Grü89a, Grü05b, Lau83, Lia96c, Tok09, Nor91]. Demarcations [Hie94]. Demise [Lau83, Nor07d]. Democracy [Lia96b, RR15, Swa17, Tac17]. Democratic [WC91]. Democritean [Cha09g]. Democritus [Vam09d]. Demonstration [Gal92a, Wal92a]. Demonstrations [vZ11]. Denials [Čap71s]. Denotation [Mar79k, Ton81l]. Density [Ana90b]. Deontological [D’A95]. Department [Edm95]. Dependence [Ano09j, Koh98e]. Dependency [Chr11, Koh98e]. Dependent [Fay97b]. Depending [Bur74a]. Depth [Hoo94, Imo15, Kow16h, Kra95, SF94]. Derivability [Ned12]. Derivations [Sol12b, Tos81]. Derrida [McK92]. Derridean [Kat11b]. Descartes [Kuz87b, Bec91i, Bor99, Bro86, CS79g, Coh86b, Duh91e, Dun89, Fra86, Gro09b, Hüb76, Man12, Man13, Mar12, Mar13, Mat01, Mey91g, Sas03c, Sas03d, She04, Tau97b, dB12, dB13, CS79g]. Descent [Hal16]. Describe [Caw89]. Describing [D’E97]. Description [ÁNS91, DP92, Hon94, Le 08, Mar79m, Zwi08]. Descriptive [Are89d, Kat06e, Sce06]. Design [Coh74a, Kro92, Nic12, Sei92, SPB18, Spr91b, Vin92]. Designation [Mar79p]. Desire [Enç05, Kau96]. d’Espagnat [Web08]. Despair [Arm18, Bur80]. Despite [Kow16i]. Destiny [Que95, Rog81]. Destruction [ZC88h]. Desvío [Iru01]. Details [KS95]. Detected [Pia90]. Detective [Aga03y]. Detectors [Gro82a]. Determinacy [Hyd06]. Determinants [Blo86]. Determination [Bec86, Mar74a, Wie12]. determinations [DO91]. Determinism [Bun12g, Čap71x, Cla89, Das91, Kir76, Mil79, Qui83, Roc11, Rop83, Tem95]. Deterministic [Čap71-38, Fet10d, Sch63]. Dethroning [Hal84, Jan84b]. Deus [Ide00]. Develop [OJ14]. Developed [Duh02g]. Developing [Che05b]. Development [Ada11, Ams92c, Ano05o, Arm18, Bar14a, BAC14, Bar14b, Bin96, Bun12u, Čap76, Čap91b, CLR^{+}92, CM89, Dam96d, Dam96e, DP92, Ech95, EN96, Fal86, Fra80, Gav90, Ges74h, Gil88, GWT00h, Gro82a, Gro82m, GHQ96b, Hel00f, HH97, Hoo99c, Hör89, How76, HH15, Jah94, Kap80d, Kap80-39, Kat11a, KB92a, KB92b, Kuh87, Lau80b, Lov15b, Mer70, Miś82a, Mur12a, New15, Nia96, Nic17, Nij15, Nik15, Pap14, PJM92, RA79, Ras94c, Rav18, Ree14, Roq11, San14, Sch98, Sch11, Sha84a, Som76a, Syn82, Tru82, TP90, Wak15b, Wal81e, War79d, Yan93, Yuz96]. Developmental [Art15, Bri15, Fle91b, Ger15a, Ges74i, GWT00g, Lov15a, Nij15, OS09, Wim15]. Developmental-Physiological [Nij15]. Developments [Eag83, Gav89, Hoh83]. développement [CLR^{+}92]. Device [ÉA99].

Devices [Kri90]. Devil [Gen91]. devo [Han15, Str15, Gil09a, Art15, Ger15b, Gri15, Raf15, Win15]. Devonian [Rud86, Sch86e]. Dewar [Gav09, Row14]. Dewey [Rat88, Vil02i]. d’Expansion [Pye92]. Dia [OP91a]. Dia-magnetism [OP91a]. diachronic [Vam09g]. Diagnoses [Coh02c]. Diagnosis [FA76g]. Diagnostic [LA86]. Diagrammatic [vL16c]. Diagrams [PS01a, SH91, vL16a, vL16e]. Dialectic [Kra00, Mar79a, Tha86a, Tha86b]. Dialectical [CT91, Coh86d, Fin80b, Har84a, Hör91a, Mar84d, McM84, Now82, O’C99, Pet69b, Zel95]. Dialectics [Kos76a, Kos76b, Kos84, Law96, Sab84, Tru94, ZC88c]. Dialogism [Byr02]. Dialogue [Ágo11, Ara12, Bar83, Bel92b, Bur83, Fey79, Fin80g, Fin80s, Koh86, Kou89, Nev02, Wüt16]. Dialogues [Oru95]. Diane [Aga75n]. Dichotic [Ges74j]. Dichotomies [PB07]. Dichotomy [Ede88]. Dickel [Chu01]. Dictatorship [Kre79]. Dictionaries* [Bel81a]. Dictum [Ide00]. Did [Ber14, Jan07, Rös95, Stö11, Wil00, Sza04, Tur96, Was91, ZC96]. Didactic [Ber14]. Didactical [Vol14]. Didactics [Seg09]. Diderot [War79d]. Dielectrics [Duh15f]. Dies [Kar17]. Diet [Tho14]. Diffeo [Pos95f]. Diffeo-4 [Pos95f]. Differ [Krü88]. Difference [Fre99, Jar86p, Loc76a, Mor76, Sim94, Wol89, Onn91i]. Differences [Cor00, Dam96k, HN16, Lih93b]. Different [Aga95d, Čap71k, Čap71m, Mar95a, Mor09, Nor98a, Sim15, Vog91]. Differential [Gro82k, Lag97a, Roq11]. Differentiation [Bie76, Lec90, Mün14]. Differing [Rap81b]. Differs [McC95a]. Difficult [Moc91, Pul11, Pes07]. Difficulties [Akr09, CŻ82, Ghi93, Goh81, Kow16h, OP82, Sch95a, Sol12a]. Difficulty [Aga13f, Mar11b]. Diffusion [TP90]. Dig [Nol03a]. Digital [Bun12n]. Dilemma [Car74, Kat76, Koh98b, Sch96b, Str76, Vro90]. Dilemmas [Now79b]. Dilthey [Are89d, Kóh90]. Dimension [Bal15, CT98, Dor96, Koh98k, Kro89, McC94]. Dimensional [Bri96, D’A01k, DG15, Gab76, Pec94, Wal07b, dB87d]. Dimensionality [Grü73j, Lay92]. Dimensionally [Des03]. Dimensions [D’A01o, Góm15b, Max97, Roz71, Vol14, WW90].

Dioni [Lam12]. Dionis [Lam13]. Dionysus [Lar99]. Diophantus [Chr15, Hea04]. Dios [dF14a]. Dipole [Buc98]. Direct [Ano09q]. Directed [Fos91, Sar91]. Direction [Fay97a, HR74g, Mei97, Rot87e, Teg97]. Directionem [Sas03h]. Directions [Sim15]. Dirk [CSW74, Rie74]. Dirt [Hal95a]. Disagreement [Räz16]. Disappearance [CW18, Fra99a, Onn91e, Onn91h]. Disappears [Onn91g]. Disarray [Ren07a]. Disastrous [Tru82]. Discarded [JR07]. Discarding [PB07, Tul16]. Discharge [Hie95a]. Disciplinary [Bar91b, Bri15, Cin89, Gab01, Mün14]. Discipline [Aic15, BSM06b, Gom15a, Har99, Hie07, Pap15, Rau91, SM15b, Str83, Woo91, WC91]. Disciplines [Abb14, Bra91, BCK07, Chr94b, GL91, How79, Lie76, Rai11, Ros76, Sta07b, WP94, ZC88e]. Disconnexion [Ges74k]. Discontents [Nol03b]. Discontinuity [Aga81d, Cin89]. Discordant [Fra99i, Ste12]. Discors [Ste12]. Discourse [Are89c, Cif05, Coh02e, Das76, Fek99, Gav94, HCP18, Nor91, Pos76a, Ton73c]. Discourses [Bee89, Rou89]. Discoveries [Fin80q, Grm81, Nic80d, Rog81, Sta94b]. Discovering [Lek15b, Myl03]. Discovery [Ach80, Ada11, Aga80, Aga13c, Ams75b, AG16, Ban80, Bel81b, Bry98, Bur80, Cap81b, Chm82, Cra93, Cur80, D’A81, D’A01i, De 81, Fel89, Fow94, Fra99d, Gig03, Gio81, Goh81, GCC81b, Hat80, Ina05, Kat06d, Koe80, Koe86, Koe93, Lau80a, Lug89, Mac80, Mor80, Ner89, Nic80c, Nic80e, Nic80f, Obe75, Ord96, Pei03, Per81, Per87, Rus80, Sch80a, Sco80, Şen05, Sha74a, Sig91, Sim77i, Sim77m, Sim77t, Som81, Toe81, Tou74a, Van74d, VdOM11, War80, ZC96, vF04, Mit04, Wal92c]. Discrepancy [AG16]. Discrete [AB17, Ela11h, Fra09b]. Discrimination [Ban95, Gro82i]. Discursive [Byr02]. Discussing [Jan89c]. Discussion [FTM^{+}63, Fre86l, Grü73o, Grü73v, MQK^{+}63, MP16, Nic80b, Omo18b, Rut91b, SRI^{+}92, Sim77g, Sim77h, Sim77y]. Diseases [Boc96]. Disk [Vis03]. Disorders [Ges74b, Ges74l, Ges74q]. Dispelling [PW17]. Dispensability [MV80e]. Displacement [Duh15b, Nee11l]. Displaying [Hei97]. Disposal [SF94]. Dispositio [Cif05]. Disposition [Ano09-46]. Dispositional [Fet71, Fet81c, Hem76]. Dispositions [Dwy96]. Dispute [Din96, Mur94]. Disputes [Ams92b, Ara15]. Disquotation [Mar79i]. Disruptive [Nic08]. Disseminate [Kar10e]. Disseminator [Yus96a]. Dissent [Swe94]. Dissolution [Kre96]. Distance [CHS97b, PR97, Wei01a, WW03, PO91]. Distended [Čap71w]. Distinction [Elk86, GM89b, Irz15, Mas15a, Soy05, Syl89, Tra15]. Distinctions [Gre74m]. Distinguishing [Gru69b, Par09a]. Distributed [Ste99]. Distribution [Vol14]. Distrust [Rap82]. disturbance [TO91]. Disturbances [Ges74-27]. Divergence [Koh86, Sch86d]. Divergences [Mil11a]. Divergent [Pol74b, Pul11]. Diversity [Aga69b, Aga75t, Gra07a, MV80f, Mey85e, Sca03]. Divide [Mac12d, Mai16]. Divided [Mit76a, Rou12, Rou13]. Divides [Can12]. Divine [Lel00b, Zam92g]. Division [Lim94]. Do [Bar83, Bar97, De 16, Grü89b, Kam11, Pet76, SV18, Str15, Swa09, Aga81i, AV96, Dam96k, Gig03, Mod15, VdOM11, Tis03]. Doctores [Wal81k]. Doctrinal [Haq94b]. Doctrine [Aga13g, Aga13k, Bec04, Buc94b, Das91, Duh91c, Mar82, Rie99, Sam76a]. Doctrines [Duh15f]. Documentalist [De 17]. Documentary [Bla95g]. Documents [Bla92a, BE10a, Für92, Mac92]. Does [And18, BR18b, Bun12f, Bun12g, Gro82g, HR74j, Jam10, LL92, Pit92, Pro14, Rop86, Sch83e, Sim77i, Skl06, Str84, Bon00, Cha09m, GG96b, Sch99c, Stö99]. Dogma [Čap71-35]. Dogmas [Gre83]. Dogmatis [Par17b]. Dogmatism [Mey85f]. Domain [Ano05f, Fu93b, Sha84n]. Domains [Sha84o]. Domestic [RA14]. Domination [MA83, Ros76]. Domingo [Wal81f]. Dominican [Zam92e]. Dominion [Hie95a]. Donald [Mar94a]. Done [Flo15]. Don’t [Bun17, Han67]. Dorado [Bol95b]. Dörffel [Hel74]. Dorpat [LP01]. Double [Bun12i, Ege99, Lih93b, Ohm98, RS07a, Wul96]. Doubts [D’A01e, Zam92l]. Down [Rus09, Sie84]. Downward [Hen06]. Dr. [CS79x, Lev07]. Dragons [Bai17]. Drake [Lan90, LS90, Mac90]. Drama [Sam92]. Draughts [KT01]. Dream [Tau97b]. Dreams [Gly83]. Dressed [Ree14]. Drift [Fre09]. Drive [Saw17]. Drosophila [Wim92]. Drug [Ada11]. DSS [D’I03]. Dualism [Bun10c, Don86, Mor83]. Dualistic [Mar89]. Dualities [Mun15]. Duality [Joh96a, Sim11a]. Dubislav [Kas11, Thi11, vZ11]. Duhem [Grü63, Grü86, Lau09, Nee11c, Qui74, Ueb97, Wal81p, Wor78a]. Duhemian [Mon09]. Dumb [Zam92c]. Dumbleton [Syl91]. Dungeons [Bai17]. Duodenal [Bel81b]. Durant [Dor92]. Duration [Čap71o, Čap71n, Čap71l, Čap71p, Čap71u, Čap71s, Čap71w, Čap71-34, Čap71-40, Loc76b, MC80w]. Durationless [Čap71-41]. Durée [Čap91n]. During [Cla14, Dor92, Duh91n, Kar10d, Koe80, Koe86, Sil11, Vic92, Vla94, Ano08-32, Bar01a, Bin96, Boc96, Gav90, Row96]. Durkheim [Lim94, Mac86b, Meš89, Tur86c, Tur86l]. Dutch [Maa17, Sne94, Vis92]. Dutton [Kel94]. Dyeing [NG01a, NG01f]. Dyers [NG01b]. Dyes [RT01]. Dyestuffs [NG01d, NG01e, NG01c]. Dynamic [Ano09-45, Čap71o, Knu07, Nic12, Sim77a]. Dynamical [AB12, Bur74b, GG96a, GW97, Roq11]. Dynamics [Ano09-44, Bun93i, Car07, Dou99, Duh02s, Fre86p, HH97, Ito98, Kos84, Lag97a, Lag97d, Lag97f, Lag97r, Mer70, Ned90a, Poi07, Rap97, Rey18, Ste79, Stö03]. Dynamism [Kuz87c]. Dynamist [Heu96]. Dynasties [Bin96]. Dynasty [Yal96]. Dystopia [Gor07]. Dzierzon [Chu01].

E. [Grü73p, Vil02g]. Each [Rös95]. Earliest [FQ18, Hei88]. Early [Ara15, BD05, Bau17, Büt08, BRS18, CW18, Chr14, Dam96i, Dra94, Edg15, Gru80, Hie70, HR74e, Hor94, Jol18, Kap80f, Kle01a, Kno04d, Knu95, LP01, Nel67, Nor07d, Omo18b, Rai10, Ree87b, Rei96, Rit01, Rub83, Rus76b, Top84, Tru82, Vil02b, Wel15, vD14, vL16d, Clu05, Fow04, Jam10, Min10, Nav15, Scu98, Van08, Zam92e]. Early-Modern [Büt08]. Earth [NNEH11, TB15, Aga03f, Bod18, Fri74, Ham07, Hoo99k, Nen18, Zum96]. East [Jhu07, SC74a, Cul11, MLNZ18, NNW74, Woo91]. Eastern [Kim11, Nak95]. EBI [Per07]. Ebola [Coh17a]. Eccleciastes [Fis95]. Echoes [Roc11]. Eclecticism [Pat11a]. Eclipse [Gün11]. Eclipses [RS03]. Ecocide [Got94]. EcoEvoEthics [Bla13]. Ecological [Bir09, Jer15, Koh98b, Koh98i, Ree87a, GC95]. Ecologically [Ber13b]. Ecology [Ber13a, Ber13c, Cal13, Dro13, Hör91b, Koh98e, NNEH11, Pâs15, Tau98]. Econometric [GG90]. Economic [Bau88, Bol95a, Bun12c, Bun12u, FM09c, Hag16, Hes09, Kow16a, Kow16g, Kow16j, Rad89a, San14, Spa03]. Economics [Aga81f, Aga09, Dil06l, Bre88, CC90, Fol17, Gem90, Gon03, Kos76c, Kow16b, Kow16l, Kow16m, Mac97d, Mac97f, Mäk96, MC96, Nad96a, Sch94a, Sel03, Wet09]. Economist [Bol17]. Economy [Gir01, Hal82, Hal92, Hei08a, Jie96, Kos76d, Nem91b, Osb92, Sch94a, SN11, Sch94b]. Ecosystems [OS09]. Ecstasy [Žva79]. ed [CS79n]. Eddington [Bek86, Grü73w, Sha86a, Sta86a, Sta17]. Edgar [Byn89]. Edge [Gin06]. Edges [Woo02]. Edification [Jan89f]. Edinburgh [Ara94, Nol03d, Nol03c]. Editing [vL16e]. Edition [CS79-30, Haq94a, Scr74a, vH77a, BV97]. Editions [Akb11]. Editor [Ros74a, Sar92b, Tau91b, Tau91c, Tau91d, Tau91e, Tau91f, Tau91g]. Editorial [Tau91a]. Editors [vH77a]. Edmund [Mat01]. Education [BS15, Che14, CW18, Dam96f, Dam96k, Dvo91, Gon01, Kap80-42, Kar10c, Kay76, Mar76c, Mat95, Pal76, Sas03c, Sie01, Swa09, Swa17, War79b, Wel15]. Educational [Bar74, Win17a]. Educationalist [Sta91c]. Edwards [Lim94, Sel74]. Eells [Fet10a]. Effect [BEW96, Hen98, Hor88, Kat06d, Yal96, Onn91j]. Effective [Han15, Mar84c]. Effectiveness [Kap80g, Sar11]. Effects [Bal94, Die14, Hum11, Tap91, Tru82]. Efficient [Bou02d, dB00, PHL95]. Efforts [Fra91, Hel00h, Kar10e, Wet95]. Egalitarianism [Gel84]. Egg [Wal84]. Egypt [Kno04a]. Egyptian [Dam96e]. Eighteenth [Rai10, Bla04, BR18b, Duh02p, Duh02o, Hec01, Jol18, Kra92a, Lef01a, Pat03a, Wel15, Wil18]. Eighteenth-Century [Hec01, Jol18, Kra92a]. Eighty [Wol95d]. Einstein [Alb92, Ano07d, Bar07, Ber74b, Bie87, Can16, Cao07, Čap71a, CS79h, Cos87a, D’A01d, D’A01e, DD11, Elz95, Fin76, Fol96, Fox15, Gia71, GH01, Goe70, Grü73g, Grü73p, Har97a, Hol70, Hol92, Hol95b, Hol07, Ita92, JNR^{+}07b, JRN^{+}07b, JR07, Jan07, Kan87, Kox95, Kuz87d, Kuz87m, Mey92, Mie07c, Mur94, Nag71, Nor07d, Nor07e, Pao84, Pat86, PW17, Pul06, Räz16, Ree87b, RS07a, Ren07d, Row06, Sau07, Sch00, Sch95b, Sch03, Sta86a, Sta15, Tan92, Web93, Wie92, Zuo96, dB87n, vD17]. Einsteinian [MC80i]. Elaboration [Kat06f, Maj98]. Elastic [Lag97b, Lag97k]. Elea [Vam09l]. Eleatic [Kuz87g]. Electric [Hie95a, Duh15c, Onn91d, Onn91i]. Electrical [FG92, Mey85g, Onn91e, Onn91g, Onn91i, Onn91h]. Electricity [Ito11, Per87]. Electrodynamics [Ano07n, Bor07, Buc90, CS79l, CS79-34, CS79-47, D’A81, D’A01b, D’A01h]. Electrolysis [Cha12c]. Electromagnetic [Ano07f, CS79-36, D’A01j, D’A01k, D’A01i, Duh15d, Duh15i, JM06, Knu95, Sie95]. Electromagnetism [Fri07b]. electromotive [Onn91l]. Electron [JM06, Nol08, Poi07]. Electronic [MPC91]. Electronics [Kap80o]. Electrons [Cha09i, Lea96]. Electrostatic [TBC91]. Electrostatics [Duh15e, Duh15h, Duh15j, Sto91]. Electrotherapeutics [Ito11]. Element [AG16, Fre86i, Fre86j, Fre86k, Fre86m]. Elementarity [Sac67, Wal67]. Elementary [Alb96, CG90, Fin03, Hin95, Sto97]. Elements [BR15, BV01, Ber95a, Cai14, Duh76d, Kow16a, Van74d, Sai04a]. Eleventh [Ras75]. Elie [GWT00e]. Eliminability [Sim77s]. Eliminate [Fek99]. Elimination [Cor76a, Mey76]. Eliminative [Hou93]. elle [CLR^{+}92]. Ellery [Fet10a]. Ellis [Aga13f]. Ellsworth [O’D02]. Elsasser [Kau74]. Elusive [Čap91c, Gav09, Mor11, Tau97a]. Emancipation [Gua96b, Jam76]. Embeddedness [Byr02]. Embedding [Şte15b]. Embodied [Dol03]. Embodiments [Cam87, MV80g]. Embodying [Coo17]. Embrace [Kai07]. Embryological [GWT00b, GWT00c]. Embryonic [Fre15, Gil91, GWT00h]. Embryos [GF97, Str15]. Emerge [Dam96k]. Emergence [Bro93, Bun10f, Cha09h, Chr14, Duc92, FRSV18a, FRSV18b, Gom15a, Jar86c, Kri92, Lek15e, Man15a, Man15b, Mou89, Pap15, Pas01, Rei96, Ren07a, Rip11, Sel03, Duc92]. Emergent [Con99]. Emerging [GF97, Wat01a, WØ97]. Émigré [Wol95k]. Emmanuel [Hay03]. Emotion [Fin80d, Mai16]. Emotional [Cla12]. Empedocles [Vam09e]. Empire [Rou12, Rou13, Ágo11, Ele92, GG03, Gün92, Osb92, Pat11a, Sas92]. Empires [PJM92, Pet92b]. Empirical [Ano06j, Ano09-42, Bar17, Chi81a, Chi83, Coh76c, D’E97, Dil06a, Ech96, Gai86, Grü95b, Hoo94, How79, Jeb17, Joh17, Jol18, Kai96, Kat06b, Kör86, Led18, MC80p, Pan91h, Pat82, PC69, Put69, Put86b, Sch18, She04, Sie82, Sta83a, SL82, Ton73c, Vil02h, Wor78a, Zah89, Zin73d, Zol89h, Zwi09]. Empiricism [Ams75d, Ams75e, Are89e, Avg90, Bec91n, Ber00, Bod18, Bue11, Cha18, Coh76c, Dem89, Dil06b, Dil06c, Fei74, Fei86, Gre83, Grü73i, Grü73j, Ham18, Joh96b, Ler95, Mar84p, McM74a, McM86, MP13, Nea00, Nic89a, PV90, Pot97, Rat18, Rut91c, Sha00, War79j, Wil18, Wol18, ZRKC03j, Cha18]. Empiricisms [BR18b]. Empiricist [And18, BR18b, Duh02o, Koc97, Sie71]. Empiricists [Mar78d, Rös95]. Empirio [Čap71-31]. Empirio-Criticism [Čap71-31]. Empyrean [Duh76b]. Encounter [Bal89a]. Encouraging [BAG17]. Encryption [Bau17]. Encyclopedias [Bré10]. Encyclopédie [Roe01]. Encyclopedist [Sta91c]. End [Aga03e, Bro08, Čap91d, Guy07, Hac17, Kim11, Lel00n, Mar94c, McB94, Pet76, Que97, Red93, Tur86d, Wol76j, Goo01]. Endeavor [Qua84, Sch84]. Ended [Aga03k, Eşa15]. Endogenously [AB12]. Endowed [Zam92m]. Ends [Guy07]. Enemies [Row06]. Energetic [Hel00b, Hel00c, PP81]. Energeticists [Del95]. Energetics [Hel00f, Hel00g]. Energizes [Ito11]. Energy [Bor07, Bun12m, Duh15i, Gav89, Gol03, Hei07b, Hel00d, Kap80h, Kap80m, Knu95, Nee11i, Ord96, Pet87]. Enforced [Ano07r]. Engaged [Har01]. Engagements [HN16]. Engaging [LL14]. Engels [Wei84]. Engine [Kro92, Seg03]. Engineer [Asl18, Dib74, DS18, Kle91, Val10a, Val10b, Val10g, Yin18]. Engineering [Arm18, Asl18, Bri18, CW18, DS18, FG92, HCP18, Kap80b, Kro92, Lay92, Li18, LvHV92, LMZ18, Lue18, Mic18, Mir14, MZ18, MLNZ18, Moo18, SV18, SPB18, Vin92, WLC18, Wan18b, Yin18, vdP15]. Engineers [FQ18, Mod15, Vin92]. England [Fre86n, Fre86s, Hal90b, Pop88, Ric91]. English [Jon91, Kau14, Mil87, Sch08a, Sch08b, Scr74a]. English-Speaking [Jon91]. Enhancement [Gro82f]. Enhancing [HD95, PLH95]. Enigma [Bit90, FR08, Wal81f, Zam92h]. Enjeux [Pet92b]. Enjoyment [Gar16]. Enlightenment [Abb94, Dvo91, GM89a, Kit90, Pos76a, Tur86a, Wil01, Zol89b, Ear94, Jan89a, Lev07, O’C99, Pat15, Rat18, Vla94]. Enough [Wol76-40]. Enquiry [Ric17]. Enriques [Gra06]. enseignement [CLR^{+}92]. Ensembles [Pos95p]. Entailment [Zin73f, Zin73h, Zin73l]. Entanglement [Ara97, CHS97b, JS03, Kow16i, Myr97, Sta97]. Entdeckung [Mit04]. Entering [dB87d]. Enterprise [Aga03-27, Dev15, Mar76a, Pan91i, Pan91h, San92a, UM92]. Entertainment [Lau87]. Entities [Aga69a, Ara12, Čap71-33, Dor00, GC10, Gre71, Mar78h, Max76a, Sal71, ZC88d]. Entoptische [Bur87c]. entre [Bea75]. Entrench [SG11]. Entrenchment [LW11, Wim15]. Entropy [CS79w, Grü73m, Nee11e]. Enumeration [Mah75]. Enumerative [Ano09r]. Envelope [Mic18]. Environment [Bar15, Mat89, Ste94, Wal07a]. Environmental [Gen83, Lar13, Nor13, Reg13]. Environmentalism [Wat01a]. Environmentalist [Att13]. Ephesus [Vam09h]. Epicurus [Bai76a, Cha09m, Kuz87d]. Epilepsy [Ges74p, Lek15b, Lek15c, Lek15e, Lek15g]. Epileptic [Lek15e, Lek15g]. Epileptics [Lek15b, Lek15g]. Epilogue [Bec91j, Ber13b, Nol03e, Pan91a, Pos95g, SM91, Vam09f, Wal92b]. Epiphenomenalism [Wat95a]. Episode [Elz95]. Epistemic [Ano09-45, Aub17, Ber13b, Ber13c, Far03j, Kar17, Kow16l, Pan83b, Pan83f, Pan91k, Von76, Zem88, ZC88e, vDP17a, vD17, vDP17b]. Epistemological [Aga95c, Bor99, Cas96, CW74a, CS79e, CS79h, Col96, Eag83, Gin06, Goo67, Goo86, Kat14, Mey85v, Mey91c, Mil15, Mot82, Pan91f, Qui00a, Ric15, Shi76, Str86, Wet95, Zol89i, vH77a, vH77c]. Epistemologist [Fin09f, Kal87, Löw86]. Epistemology [Aga81g, Aga87c, And99, Ant87, Avg90, Bab99a, Bab99d, Bar99, Bro96a, Čap71-31, Čap91b, CS79i, CS79-31, CW83a, Coh02t, Dam96c, Elz95, Fag12, Fle86c, Gje00, GG96c, Han99a, Han99c, Hee86a, Imo16, Irz95, Kai07, Kri90, Lee74, Lee84, Löw86, Pan83b, Pan91i, Rau91, Ree87a, Ren95, Sag87, Shi87h, SN87, Ska87, Som81, Sup91, Twe00, War79e, War79m, War86, War87, Wra15, Zol89g, GSW95b]. Epistle [Jar86d]. Epoch [Sch11]. Epoch-Making [Sch11]. Epochal [Sch11]. Epochal-Break-Thesis [Sch11]. époque [Ras92]. EPR [BF94, Pos95t, RL97]. Equal [Pie95]. Equality [Hoo88a, Neu76a, Vrc79]. Equation [RS07a, Onn91a]. Equations [Duh15i, JR07, Lag97a, Mel03, Ras94g, Roq11, Sau07, Vog04]. Equilibria [Kuo07, PP81]. Equilibrium [Lag97b, Lag97c, Lag97e, Lag97g, Lag97i, Nee11j, Nee11l]. Equipollent [Har08a]. Equity [Gel89]. Equivalence [Gly71, Ren90, Sig91, dB87t]. Equivalency [Chr11]. Equivalent [Duh02h]. Era [Hug88, Obe75, Ada01, Ano08o]. Erazim [CT98]. Ergebnisse [Bie74]. Erice [GCC81b]. Erich [Kar10a, Kar10c, Kar10b, Kar10d]. erkenntnistheoretikers [Str76]. Erleben [Ned90b]. Ernest [Kap80r, Wal81g]. Ernst [Sta92a, Baa92, Ber70, Bla92a, Bla92b, BIT01c, Blü70, Čap69, Coh70, CS70, Hal82, Hal92, Hei11, Hof92, Pec96, San92b]. Errata [Haq94c]. Erratum [GR12, GR13a, Mur12c, PP15b]. Error [Aga13a, BM09, Bro91a, Cam03, Hon09a, Hon09b, Mil09, Sch09, Wol03]. Errors [Fra09b, Ges74y, Zan74]. Erudition [Fin80e]. Erwin [NRS92, dR01]. Esau [Hac17]. Escape [Ams92d, Lay92]. Escherichia [Fos91]. Eschinardi [EF18]. Essay [Bon08, Bou02a, CS79-30, Coh80, GWT00d, Gou06a, Hel14, Hoo94, Lia96b, Mac97f, Mic97, Nic80c, dR17, HOO88c]. Essays [AC82, Aga08a, Aga08b, ARS15, Bab02c, BAG17, But93d, CSW74, CW74b, CFW76, CS79-34, CL83, Coh02j, DJ89, D’A00, Dam96b, DPW86a, Duh02m, FK93, GSW95c, GSW95b, GC94, Gre74t, HOO88c, JRN^{+}07b, JL95a, JL95b, Kas76b, KS95, Kra82c, LM97, LS90, MC95, MC96, MP79, Mil09, Mun91a, NRS92, Pan91f, PP87, Spo09, Wal81q, Wia79b, Aga08c, D’A01l, GSW95a, GC97, Fra99c]. Essence [Bou02g, Han99f, Han99h, Mey85h]. Essences [Ano09p, Tha86d]. Essential [Bre74a, DR15a, Fre86g, Han99d, Now82, Sup74]. Essentialism [Han99g, Pan91e, She95b]. Essentiality [Mar79j]. Establishing [Car15b]. Establishment [Hel00e, Höf76]. Establishments [FS96]. Esteem [CS79q]. Estimation [SVD95, Sei92]. Estonian [Vih01b]. état [Onn91r]. etc [Onn91c, Onn91d, Onn91e, Onn91g, Onn91i, Onn91h]. Eternal [Čap91e, Fre91, Len87, Sam76a]. Eternity [Goo77g, Llo86]. Ether [Ein07a, Fre09]. Ether-Drift [Fre09]. Ethic [Gel76]. Ethical [Coh02b, De 95, Gre74j, Jar86q, Koe84, Koh98b, Kou11, Moo18, Sch86a]. Ethics [Ber13a, Ber13c, Cap15, Coh74b, D’A76, EPL69, Gen83, Gon15a, Gon15b, Gro95, Köh91b, Kov06, Kow16b, Lue18, Mit15, Reg13, Rut91a, She09, Tau98, vM89, GSW95b]. Ethnographic [Jar86r]. Ethology [Lij99]. Ethos [Ste15c]. Etiological [Che11]. Etiquette [Bai17]. Eucharist [Ela11t, Syl75]. Euclid [Bec04, Ros74b, Sai04a, Wei04]. Euclidean [Gra07b, Mat86, Mit76d, dB87i]. Euclides [Bağ05]. Eudoxic [Bec04]. Eudoxos [Bec04]. Eudoxus [Nik74]. Eugenics [Gon01, LG01, Ste92, Ano08o, Ano08-45, Ano08-40, Ano08-32, Sch08d]. Euklid [Bec04]. Euler [Elk74]. Eurasian [Kim11]. Europae [Lõh01]. Europe [Ágo11, Bre11, Coo11, GG03, GR11b, NG01c, PT04, PB94, SCD03b, SDG15b, Svo15]. European [Bed09b, Cif14, Dah90, DP92, Ele92, Gai94, Ihs92, NG01b, Pat15, PJM92, Ras92, Rhe15, Van15]. européenne [Ras92]. Evaluating [Bun12h, Kow16m]. Evaluation [Ano08q, Ano08p, Avg89, Bru95, Coh02c, Fin80a, Grü63, Grü86, Kow16c, Rot87c, Tap91, Ton81c, Ton81d, Zol89c, BJ95, Čap71c]. Evaluative [Res17a]. Evangelista [Duh91j]. Evaporation [Buc95]. Eve [Arb92]. Even [Piz81]. Event [Sch83f]. Events [BB71, Čap71-33, Čap84b, Čap91i, Eel10b, Feu88, Grü89b, Hee84, Mar79c, Nak98, Rot87d]. Ever [vM89]. Everett [AM01b, Ezr01, McC95a]. Every [Mer97]. Everybody [Kam11]. Everyday [Bit96g, Fin14, Koh98g]. Everything [Joh11, RS07c, Nor98a]. Everywhere [Sim77n, Wei11]. Evidence [Aga75k, Aga75j, Aga75u, Ano08-41, Blu16, Bon08, Cha12e, Ele92, Fra99c, Gru74a, Har99, Hem84, How03, Pro14, Şen05, Sim77r, Ste12, TG88, Wyl92]. Evidential [Mus78]. Evil [Kek17, Koh95, Koh98k]. Evo [Han15, Str15, Art15, Ger15b, Gil09a, Gri15, Raf15, Win15]. Evo-devo [Han15, Str15, Art15, Ger15b, Gri15, Raf15, Win15]. Evolution [Abr11, Ams75a, Art15, Atl86, Ban80, Bed09a, Ber13a, Ber13c, Bro74, Cha93b, CS79k, CS79j, CS79m, CS79-47, Coh94k, Dam96b, Dav74b, Dav86, Den99e, Dud86, Duh02i, Eşa15, Fet10b, Gay13, GM17, Ges74h, Goo92, Gre74k, Gre86a, Gut95, Heu96, Kau92, Lay74, Lel00c, Lev86, Lew76, Lov15b, MG11, Mey85i, Nag98, New15, NMN02, Nij15, OS09, Raf15, Rip11, Roc01, Rus13, Skl74, Slo76, Smo03, Ste15a, Ste09c, VN12, Wak15b, ZRKC03e]. Evolutionary [Ber13b, BZ96, Bro92, Cro92a, Cue01, Ech95, GS09, GWT00e, GWT00c, Gre74s, Heg91, Hun17, Kal87, Kri90, Lov15a, Pia82, Rus09, Sar92b, Sch12b, Sha96, Sko76a, Slo83, Sob10, Som81, Wag15, War12, Wes91, Wim15, Sar92a]. Evolutionary-biology [GWT00e]. Evolutionism [Goo01, Mun89, Pel01, Pia95]. Evolutionist [Mon01]. Evolvability [Bri15, Wag15]. Evolving [BMP09b, Fre11, Gru80, Mar15, Wil12]. Ex [Wal81i, Wal76]. Exact [Ano08-49, Pye92, Scr74b, SG92]. Exactes [Pye92, SG92]. Exactness [Kir82]. exaltation [For07]. Examination [Cof83, Grü69a, Hod90, Mah15, Nee11f, Whi99, Bar09]. Example [Bla94, Coh90, DS12, HR74j, Hof74, Mat86, Ras95, SR07, Şen05, Sch08a, Sch08b]. Examples [Hsü92, Lev71, Min10, Sla89]. Exceeds [Sim77x]. Excellence [Abb14]. Excentric [TP90]. Except [KS95]. Exception [Bar14a, Man12, Man13]. Excerpt [Lor07b, Wey07]. Excerpts [LC07, Mie07a, Poi07]. Excess [Don86]. Excessive [Bal94]. Exchanges [Lim94, Lop92]. Excited [MPC91]. Exercise [McM71, Par00]. Exercitationes [Gat18]. Exhibitions [Hei97]. Exist [Bar97]. Existence [Ber96a, Bun12i, Dor00, GWT00i, Har86, Hor00, Kno04b, KSW97, Kuz87e, Len87, Mar79d, Mey85h, Ols87, PP18, Sam92, Sei97, Uga18, Urc97]. Existential [Ihd79b, Mar67, Zin73e]. Existentialists [Ihd79h]. Exogeneity [GG90]. Exorcism [Cla14]. Expanding [And03a, Ihd99a]. Expansion [CS79b, DP92, Ele92, FRSV18a, FRSV18b, PJM92, Pye92, Ras92]. Expectation [Gro82h]. Expectations [RR15]. Expected [Mic71, Mon09]. Experience [Aga75n, Ams75a, And76, Ano06j, Bro91a, But93k, Dai16, Den95, Dil06a, Ihd79c, Ihd79k, Koh98d, Koh98g, Koh98i, Mar84l, Mey85b, Mus09, O’D02, PV90, Pro00, Pro94, Sto75, SA12, Str94, Wol76-29, Wol76-32]. Experienced [Die16]. Experiences [PQQSe96]. Experiential [Hal91f]. Experiment [Bra07, Bro95, Cha09b, D’A98, Fag12, Fin91, Fra99c, Fra99h, Fre09, Gal03, Guc81, Hie95a, Hon03, Hun10, Jan03b, Kap80i, Kap80j, Kap86b, Lel00h, LS90, McC95b, Mel03, Mos89, Pat82, Sau16, Sch10, Spr91b, Ste92, Sup03, Tru82, Ung91a, WP96, ZC88h, van91]. Experimental [Ara12, Bry98, Cam03, Cha09i, Cha09j, Chr14, CHS97a, Coh02a, Cre99, Cre02, Dei17, Dud86, DS12, Fra09b, Gon03, Hee97, HSS09a, Kap80d, Mal12, Mal13, Men76, Mül91b, Pro01, Rhe09, San08, San14, Sel03, Sol08a, Sol12b, Spr91b, Wan18b, War67b]. Experimentation [Gil03, Hon98, Irz03, Kap80-29, Kle03, Rad12a, Rad12d]. Experimenter [Fra99e]. Experiments [AAF^{+}14, Ano08q, Ano08p, Atk03, Bre07, Bry98, Buc90, CM89, CS79b, D’A01j, Enz00, Gol91, Hof09, Hol97, Kap80-35, Onn91e, Onn91h, Onn91j, Rat18, Réd03, Rey18, Sca03, Sch96a, Sei92, Sey96, Ste09a, Stö03, Sza03, Zah78, DO91, Onn91m, Onn91f, OB91a, Onn91b, OB91b, Onn91c, Onn91d, Onn91g, Onn91i, Onn91k, Onn91l, TO91]. Experimentum [Ham18]. Expert [Coo90, DD86]. Expertise [Fra99f, Kow16k, Rei11]. Experts [Ara15]. Explain [Bun12b, Dwy96, Fal86]. Explaining [Aga81q, Duc89, Gie97, Kuo07]. Explains [Web12]. Explanandum [Yli07]. Explanans} [Kra97]. Explanation [Ams92e, Ana90a, Ano09-47, Băe15, Bec91c, Bel92a, Ber17a, Bir07, Bri15, Bun93c, Bur74b, Car86, Chu95, Das88, Dil06i, DF11b, Dra88, Els88, Fay07, Fet81i, Fet81j, Fet81h, Fis96, Fis15, Gje07, Gre71, Gre74k, Gre86a, Gro76, Grü73b, Hål07, Hem76, Hun06, JS03, Kau71, Kau76, KK07, Kle74, Man89, MC80w, Mey85r, Mey85-27, Mey91d, Mey91e, Mey91l, Mey91r, MO82, Möl01, Mor09, Nic80d, Nor12, Now79a, PV90, PY07, PP18, Psi07, Rub89, Sal71, Sal92c, Sch07d, Tay92, Toa15, Ton73e, Ton73f, Tuo76, Tur86i, Wal07a, War79o, Wat92a, Wig74, Wri96, Sal71]. Explanations [Ano09-47, Băe15, Cau80, Cou96, Enç05, Fet74, Han07, Hop74, NNEH11, Pâs15, Per95, Rös91, Sto97, Tur86f, WvB07]. Explanatory [Bal18, Col74, Jef71, Put67a, Put86a, Sal18, WvB07]. Explication [Grü84]. Explicit [AT15]. Exploitation [Bar15, DR15a, VN12]. Exploitative [Gou83]. Exploiting [Web11]. Exploration [Ano08n, Att13, DR15a]. Explorations [FB17, Jar86e]. Exploratory [MZ18]. Explore [Sza84]. Exploring [Duw17, Vla94]. Exponentiation [Heu96]. Exposition [Grü73v]. Expositions [dS14]. Exposure [Whi01]. Expression [Amr89, Onn91a, Tru74]. Expressions [But93a]. Expressive [CK97, Nic18]. Expressivism [Buc94b]. extend [dBJD95]. Extended [ÁNS91, Len83]. Extendible [Uga18]. Extending [Mat89]. Extensa [Van86]. Extension [Čap71k, Gra07b, Grü73t, Höf76, Mor76, Ste97, Wat01b]. Extensionless [Čap71-35]. Extensions [Duh02s]. Extensive [Čap71a]. Extent [Gat09, Stö99]. Extérieure [SG92]. External [Elk86, Gäh89, Irz15, Rim97, Ytu95]. Externalism [Aga81h]. Extra [Som74]. Extra-Logical [Som74]. Extralogical [Goo77d]. Extramolecular [TBC91]. Extraordinary [Ano05d]. Extrapolating [Pos95g]. Extraterrestrial [Lam97]. Extravagance [SA82]. Extremely [Hon09b]. Extremes [Aga89]. Extremist [Hor88]. Extrusion [Fin09i]. Eye [Cro02]. Eyes [Aga13d, Alt02, Bab02c, Hor94].

F [Gre94, Kit74, Pul11]. F. [Nol91, Rub88]. Fables [Car10]. Fabri [Ela11g, Ela11e, Ela11h, Ela11i, Ela11f, Ela11j, EF18]. Face [Grü05a]. Faces [Han99m, Moc91]. Facets [Ano05g]. Facilitated [Kir15]. Facing [Lom15]. Facsimile [Ano07d]. Fact [CS86, Elg84, Fin80f, Hoo99c, Kow16d, Kow16i]. Fact/Value [Kow16d, Kow16i]. Factlike [dB87g]. Factor [Blo76]. Factories [Ret74]. Factors [Art15, GHQ96b, Hel00d, Kap80c, Mam94, McA97, Mur75]. Factory [NG01f]. Facts [Ams75c, Ams75d, Bit96b, Buc01, De 95, Hoq18, Mic83, Nor13, Wor91, Zaj87, vH77b, vH77d]. Factual [McL71]. Factuals [Bun93b]. Faculties [Eng01]. Faculty [DCS15, Hel08, Mot15]. Faded [Mün14]. Fail [Hun06, Ste09a, Wor89, ZC96]. Failed [Čap71-42, Cha09a, Maj06]. Failures [Bro11]. Faith [Aga81i, Aga03e, Fin80g, Hoo99c, Kea02]. Faithful [Rya94]. Fall [Abr07a, Ano08q, Ano08p, Ano08-37, Ano08-47, Lei04, Pal04a, SS74, Sta74e, Woo91]. Fallacies [Coh02a]. Fallacy [Čap71f]. Fallibilism [KZ09]. Fallibility [Fet81g, Lev83c, Wir95]. Fallible [And84c, Sch83e]. Falling [Ano08q]. Fallout [Wat01a]. Falls [Rus09]. False [Gat09, Mar97b, Mar95c]. Falsifiability [Dav86, Grü63, Grü73v, Grü76, Grü86, Pan91b]. Falsification [Apt74, Hüb78a, Kou90, Mus78]. Falsificationism [RA78a, Rad89a]. Falsifications [And91]. Falsifying [Suc82]. Falsity [Grü73e]. Familiar [Pos95l]. Family [BAK11, Nor03, ZC88d]. Far [BE10a]. Faraday [Ste09a, Cra93, Duh15f, Hie95a]. Farben [Bur87c]. Farbenlehre [Mil87]. Fascism [Hor88, Yeh17]. Fascist [SS91]. Fashion [Lag97l, Ned90a]. Fatalism [Eil95]. Fate [Cha09n, Jaw90, Sch96g, Syl08]. Fateful [Nor07e]. Faustian [Ano08o]. Favor [Pos95g]. Favour [Aga75u, Cla12]. Fears [Wat01a]. Feasibility [WLC18]. Feature [Glo76, Gro82a]. Feature-Placing [Glo76]. Features [Fle86f]. Fechner [Hei94a]. Fedoroff [Gro01]. Feedback [Fab84, GG96c, Man71, Wim71]. Felix [CH14, Hel14]. Fellowship [Fie14]. Feltre [Zam92e]. Feminism [Gel16, Got94]. Feminist [Aga95c, AV16a, AV16b, Blu16, CG16, Hag16, HN16, Har16, IdMM16, Rol16, Tua16]. Feng [Mat17]. Ferdinand [Chu01]. Fermentation [Böh83]. Fertility [McM76a, BBT95]. Festschrift [RDS03]. Feuer [HOO88b, HOO88c, Škv88]. Feuerbach [Caw94, Mac94b]. Fevers [Boc96]. Few [Mar83a]. Feyerabend [And91, Bur84, Del91, Die91, Far03f, Far03d, Far03a, Far03g, Far03e, Far03b, Far03c, Far03k, Lau89, Mar91, Max91, Mun91a, Mun91b, Nae91, Sch91b, Sup91, Tow71, Wor91]. Feyerabendian [Chu91]. Feynman [Sta97]. Fiction [Aga03-29, Čap91f, Fri10, Hoo99c, Jan03a, Mos89, Zhe96]. Fictional [Dav10, GC10, Ram01]. Fictionalism [Blo76, Dum15, Tor82]. Fictions [KL17]. Fictitious [Rat18]. Fidelity [War79q]. Field [Ano07e, Chr14, Cin89, CS79e, CS79l, CS79-36, D’A01e, JR07, Kap80x, Mot15, RSS^{+}07a, RS07a, Sam96, Sha84n, dB87c, Onn91j, WPJ95]. Fields [Bun10d, CS79-33, Kal96, Kap80-35, Mcg74, Val96, TO91]. Fieldwork [Jar86o]. Fifth [DPW86a, Fra93, Fra99j, Kel10]. Fifty [CS79v]. Figurative [GC10]. Figure [Bri15, Kae90a, Kap80-30, Phi90a]. Figures [HR74b, HR74a]. FilMat [BS16, PP18]. filosofien [CS79g]. Filth [Hof09]. Final [Bla95e, Bou02d, CLR^{+}92, dB00]. Finality [Nag84a]. Finalization [Böh83, BVH83a, Sch83a, Sch83b, Sch83c]. Finalized [KS83]. Finding [Sch95a]. Fine [Bek86, Jar95, Pos95s]. Fine-Structure [Bek86]. Finetti [Hin71]. Finite [Grü73d, Kau74, Koy76a]. Finiteness [CPM16]. Fire [Hoo99l]. Fired [Gro82h]. First [Ace92, Chr94a, CS79m, Duh02c, Duh15e, Eas84, EN96, FKR99, Hal91a, Hat95, Heg91, Hel00e, Hil07a, Hil07b, Lel00d, Mar97a, Mar78e, MS75b, Neg81, Phi94, RH92, Sas03e, Sau07, Sta07a, Tos81, Wol95b, Arm18, Cla10, CS79-38, Lek15f]. First-Order [Tos81]. Fisher [Pro92, Hod92, Sei92]. Fisk [Aun71]. Fission [Hor88]. Fit [Cam87, GG96b]. Fitness [COG90]. Fitting [Pos95h]. Five [Gie12, CW85]. Fix [Ros86]. Fixation [FA76h, BJ95, BdOU^{+}95, GC95, HD95, PHL95, PLH95, RL95, SVD95, WRL95, WPJ95]. Fixing [KN01, dBJD95]. Fizeau [Fre09]. Flat [Bel12g, Hon09b]. Fleck [CS86, Gie86, Hee86a, Löw86, Mar86e, Mou86, Sch86b, Sch86c, Tou86b, Wit86, Wol86, Zal86]. Fleeting [Nee97]. Flicker [CS79n]. Fling [Coh99]. Flocks [Ros81]. Flow [Den99b, Eck17, Grü73n, Ism17]. Fluid [Böh83, Eck17]. Fluide [Onn91r]. Fluids [Cob74, Gav90, Lag97b, Lag97c, Lag97k, Lag97m]. Flux [Aga67, Aga75l, Aga75o, Aga75p, Mei97, Pos95k]. Focus [Kas76b]. Following [Pis15, Sma92]. Fonction [You74]. Fondements [Fre04]. Food [Hei08a]. Foods [Tho14]. Footnote [Ng17]. Forbidden [Flo15]. Force [Ano07n, Bou02a, Bou02b, Bou02f, Bou02g, Bou02i, Chr07, CS79v, Fin80d, Fra93, Fra99j, Hey86, Poe99, Pop86a, Sta07c, Wat01b, Onn91l]. Forces [Bec91a, Lag97d, Lag97e, Lag97j, Mcg74, Tur86b, Von84a, Wal81c, Wil00, vdA91]. Forcing [NPS17, Ven16]. Forecast [Zim83]. Foreign [SG92]. Foreknowledges [Gal92b]. Foreknowns [Gal92b]. Forerunner [Alt87, Fey84b, Yus96a, Yus96b]. Forgetting [May95]. Forging [Sun18]. Forgot [Koh98f]. Forgotten [Pro10, Ueb91b]. Form [Bra87, Bré14, Chu91, Dou87, Fin80g, GF97, Gou76, Gre74l, Har10, HGR^{+}89, Kar17, Len87, Mar79d, NB12, NB13, RT89, Sar97, Sch76d, Ano06-28, Bur87b, HH91, Tha84a]. Formal [AM63, Cas00, Chr11, Gra83b, Leb82, Mor83, Pol16, Pos74, Sta74b, Stö99, Str82, Str18, Van74b]. Formalised [Coh02e]. Formalism [Ano07s, Hoo91, Seg95]. Formalisms [Gra83g]. Formality [Han99d]. Formalization [Car15a, Pis15, Zin73f]. Formation [Aic15, Cla10, Gav90, Gil09a, Gro82c, Han07, Len63, Mar01, Ner01, Rau91, Tuo82, Woo91, WC91, Gro86]. Forms [Ban95, Goo92, Gra83e, Kir76, Mey91e, Pat82, Top84, War79m, War86, Wim86, Win83b]. Formula [Fin87, Lag97d, Lag97e, Lag97g, Lag97h, Lag97i, Las01]. Formulae [Cha09f]. Formulas [Duh02h, Duh02g, Har89d, Har89c, Rit01, RG89]. Formulation [Fu93a, Sam96, Van95]. Formulations [Hil96]. Forschung [Bie74]. Fortifications [Val10c]. Fortran [Gal91]. Fortune [Pos95g]. Fortunes [Bab02a]. Foster [Tau91b]. Foucault [Nol03a]. Fouillé [Čap71x]. Found [Har95, JR07]. Foundation [Abr81, Car15b, Coh86b, D’A01f, DGZ96, Hel00b, Jan85e, Okr89, Pra93, RS07c]. Foundations [Acc90, Ano09-42, Aub17, BC02, Bye74, CS79o, CS79-27, CC90, CCG93, D’A01e, Dop01, Fef93, Fer85b, Fin80h, Fre74b, Fre86p, Fre91, Fre04, Gal02c, Gei67, Goo77e, Gro95, Gro09c, Grü73r, Grü74b, Han99j, Hil07a, Hil07b, Hil07c, Ion15, JN79, Kas76b, Krö12, Maj06, MC80f, Mie07a, Pul06, Qui82, RDS03, Sau07, Sch96e, SC74b, Ste94, Vil02g, Vil02k, Vil02m, Vra79, Zin73g, Zwi09]. Founders [Sar92b, Vam09g, Sar92a]. Founding [Ano08-40, Nia96, Sch03, Yuz96]. Four [Kóh90, Kra82a, Och96, San92b, Sch18, Seg00, Wal07b, dB87d, Wol95d]. Four-Dimensional [Wal07b, dB87d]. Fourier [Gar95a]. Fourteenth [Koy76b]. Fourth [Har10, Mat86]. Fraassen [Bre74c, Joh96b]. Fractal [Dou99]. Fractions [Bur74a, Fow04, Kno04a]. Fragment [Bra07]. Fragmentation [ÁNS91, Čap91m]. Frame [Duh76a, Neu76b]. Framework [Alb95, Ano07e, AT15, Coh04, Fet81f, Fol94, Mir15, Pop87, Ska87, Tul16]. Frameworks [Ela11j]. Francais [Duc92]. Francaise [Pat92, Pet92a]. France [BG09, Cif05, Det07, Dor92, Duc92, FG92, Pat87, Sal92a]. Francesco [EF18, Ros81]. Francis [Aga13q, Dei17]. Franco [Cha18]. Frank [BIT01g]. Franklin [Kap80-40]. Free [Abr07a, Ano08q, Ano08p, Ben81, Bla95j, Bun10e, Dum15, Gar12b, Han91, Koe91, Kom99, Lag97j, Leb82, Leb95, Lei04, Mun91b, Pal04a, Pro17b, Qio96, Voi96]. Freedom [Bec86, DJ89, Ezr14, Fit71, MA87, Sel74, Wat95a, WG11]. Freeing [Grm81]. Freeman [Jar86f]. Freezing [Fin14]. Frege [Pat90, Boc15, Cur89a, Mar79p]. French [Osb92, Bal03, Bea75, BG09, Cer14, CLR^{+}92, CS79f, Dal89, Dan92, Dor92, Duc92, Duh91d, Duh91e, Fre04, Gre75, Gün92, Hül96, Lam14, Loc97, Par09b, Pat92, Pet92a, Pet92b, Pye92, Rai10, Ras74, Ras75, Ras92, Rou12, Rou13, Sal92a, SG92, You74]. Frequencies [Fet10e]. Frequency [CG90, Van83]. Fresh [Ano09-46, Pap14]. Fresher [Pap14]. Freud [Are89f, GH01, Laz77, Rub83, Sav15]. Freudian [Grü09, Sim11a, TP90]. Friction [Hel99]. Friedmann [Kap80a]. Friedrich [Ber94, Jah91, BIT01d, Wei84, Whi99]. Friendly [Hin95, San15]. Friends [DJ89]. Friendship [Ber86, CS79q, Gre86b]. Fries [Gre94, Jah91, Pul11]. Friesian [Pul11]. Fritz [Hei91]. Front [Ano63, Ano67, Ano69b, Ano69c, Ano70b, Ano71c, Ano71d, Ano71e, Ano71f, Ano71k, Ano71g, Ano71h, Ano71i, Ano71j, Ano73e, Ano73g, Ano73h, Ano73i, Ano73j, Ano73k, Ano73f, Ano74u, Ano74-33, Ano74-34, Ano74-35, Ano74-36, Ano74n, Ano74p, Ano74q, Ano74r, Ano74s, Ano74t, Ano74f, Ano74g, Ano74h, Ano74i, Ano74j, Ano74k, Ano74l, Ano74m, Ano74o, Ano74v, Ano74w, Ano74x, Ano74y, Ano74z, Ano74-27, Ano74-28, Ano74-29, Ano74-30, Ano74-31, Ano74-32, Ano75d, Ano75e, Ano75f, Ano75g, Ano75i, Ano75h, Ano76s, Ano76p, Ano76q, Ano76r, Ano76-33, Ano76-34, Ano76-35, Ano76-36, Ano76-37, Ano76-38, Ano76-39, Ano76-40, Ano76h, Ano76t, Ano76u, Ano76v, Ano76w, Ano76x, Ano76y, Ano76z, Ano76-27, Ano76-28, Ano76-29, Ano76-30, Ano76-31, Ano76-32, Ano76-41, Ano76-45, Ano76-42]. Front [Ano76-43, Ano76-44, Ano76i, Ano76j, Ano76k, Ano76l, Ano76m, Ano76n, Ano76o, Ano77d, Ano77e, Ano77f, Ano77g, Ano77h, Ano77i, Ano77j, Ano77k, Ano77l, Ano77m, Ano77n, Ano77o, Ano78d, Ano78e, Ano78f, Ano78g, Ano78h, Ano78i, Ano78j, Ano78k, Ano78l, Ano78m, Ano78n, Ano79g, Ano79h, Ano79i, Ano79n, Ano79o, Ano79p, Ano79q, Ano79r, Ano79s, Ano79u, Ano79t, Ano79v, Ano79w, Ano79x, Ano79y, Ano79j, Ano79k, Ano79l, Ano79m, Ano80h, Ano80i, Ano80j, Ano80n, Ano80o, Ano80p, Ano80q, Ano80r, Ano80s, Ano80t, Ano80u, Ano80v, Ano80w, Ano80x, Ano80y, Ano80k, Ano80l, Ano80m, Ano80z, Ano80-27, Ano80-28, Ano80-29, Ano80-30, Ano81v, Ano81o, Ano81p, Ano81q, Ano81r, Ano81s, Ano81t, Ano81w, Ano81x, Ano81y, Ano81z, Ano81k, Ano81l, Ano81m]. Front [Ano81n, Ano81h, Ano81i, Ano81j, Ano81u, Ano82e, Ano82f, Ano82g, Ano82h, Ano82i, Ano82j, Ano82k, Ano82l, Ano82m, Ano82d, Ano83m, Ano83n, Ano83o, Ano83p, Ano83q, Ano83r, Ano83s, Ano83t, Ano83u, Ano83v, Ano83w, Ano83x, Ano83y, Ano83z, Ano83i, Ano83k, Ano83l, Ano83j, Ano84l, Ano84m, Ano84n, Ano84o, Ano84p, Ano84h, Ano84i, Ano84j, Ano84k, Ano84u, Ano84v, Ano84w, Ano84x, Ano84y, Ano84-27, Ano84z, Ano84q, Ano84r, Ano84s, Ano84t, Ano85d, Ano85e, Ano85f, Ano85g, Ano86o, Ano86y, Ano86z, Ano86-27, Ano86-28, Ano86-29, Ano86-30, Ano86-46, Ano86-47, Ano86-48, Ano86-49, Ano86-42, Ano86-43, Ano86-44, Ano86-45, Ano86p, Ano86q, Ano86r, Ano86s, Ano86t, Ano86u, Ano86v, Ano86w, Ano86-31, Ano86-32, Ano86-33, Ano86-34, Ano86-35, Ano86-36, Ano86-37]. Front [Ano86-38, Ano86-39, Ano86-40, Ano86-41, Ano86-50, Ano86-51, Ano86l, Ano86m, Ano86n, Ano86x, Ano87z, Ano87-27, Ano87-28, Ano87-29, Ano87m, Ano87n, Ano87o, Ano87p, Ano87q, Ano87r, Ano87s, Ano87t, Ano87u, Ano87v, Ano87w, Ano87x, Ano87y, Ano87-30, Ano87-31, Ano87-32, Ano87i, Ano87j, Ano87k, Ano87l, Ano88d, Ano88e, Ano88f, Ano89w, Ano89x, Ano89y, Ano89z, Ano89-27, Ano89-28, Ano89-29, Ano89s, Ano89t, Ano89u, Ano89v, Ano89j, Ano89k, Ano89l, Ano89m, Ano89n, Ano89o, Ano89p, Ano89q, Ano89r, Ano90e, Ano90l, Ano90f, Ano90g, Ano90h, Ano90i, Ano90j, Ano90k, Ano91m, Ano91-28, Ano91-29, Ano91l, Ano91n, Ano91o, Ano91p, Ano91q, Ano91-40, Ano91-31, Ano91-32, Ano91-33, Ano91-35, Ano91-36, Ano91-37, Ano91-38, Ano91-39, Ano91-34, Ano91x, Ano91r, Ano91s, Ano91t]. Front [Ano91u, Ano91v, Ano91w, Ano91-30, Ano91y, Ano91z, Ano91-27, Ano91-43, Ano92l, Ano92m, Ano92n, Ano92o, Ano92p, Ano92q, Ano92r, Ano92s, Ano92t, Ano92-37, Ano92-38, Ano92-39, Ano92-40, Ano91-44, Ano91-45, Ano91-46, Ano91-47, Ano91-48, Ano92u, Ano92v, Ano92w, Ano92x, Ano92y, Ano92-34, Ano92-35, Ano92-36, Ano92-33, Ano92-31, Ano92-30, Ano92-32, Ano92-29, Ano92z, Ano92-27, Ano92-28, Ano93m, Ano93f, Ano93g, Ano93h, Ano93i, Ano93j, Ano93k, Ano93n, Ano93o, Ano93p, Ano93q, Ano93l, Ano94r, Ano94s, Ano94t, Ano94p, Ano94x, Ano94u, Ano94w, Ano94v, Ano94o, Ano94n, Ano94q, Ano94l, Ano94m, Ano95-32, Ano95-33, Ano95y, Ano95z, Ano95n, Ano95t, Ano95o, Ano95p, Ano95q, Ano95-34, Ano95-35, Ano95-36, Ano95-37, Ano95-38, Ano95u, Ano95v, Ano95w, Ano95x, Ano95r, Ano95-27]. Front [Ano95s, Ano95-28, Ano95-29, Ano95-30, Ano95-31, Ano96k, Ano96n, Ano96t, Ano96j, Ano96i, Ano96m, Ano96l, Ano96o, Ano96p, Ano96q, Ano96r, Ano96s, Ano97i, Ano97n, Ano97p, Ano97h, Ano97j, Ano97k, Ano97l, Ano97m, Ano97o, Ano97q, Ano98d, Ano98h, Ano98e, Ano98f, Ano98g, Ano99i, Ano99h, Ano99m, Ano99l, Ano99k, Ano99j, Ano99n, Ano99o, Ano99p, Ano99q, Ano99r, Ano00k, Ano00l, Ano00m, Ano00n, Ano00o, Ano00p, Ano00h, Ano00g, Ano00j, Ano00q, Ano00i, Ano01-28, Ano01q, Ano01r, Ano01s, Ano01t, Ano01u, Ano01j, Ano01k, Ano01l, Ano01m, Ano01n, Ano01p, Ano01v, Ano01w, Ano01x, Ano01y, Ano01z, Ano01-27, Ano01-29, Ano01o, Ano01-30, Ano01-31, Ano01-32, Ano01-33, Ano01-34, Ano02s, Ano02o, Ano02p, Ano02q, Ano02r, Ano02t, Ano02u, Ano02g, Ano02h]. Front [Ano02i, Ano02j, Ano02k, Ano02l, Ano02m, Ano02n, Ano03-37, Ano03t, Ano03u, Ano03v, Ano03w, Ano03x, Ano03y, Ano03z, Ano03-27, Ano03-28, Ano03-29, Ano03m, Ano03n, Ano03o, Ano03p, Ano03q, Ano03j, Ano03k, Ano03l, Ano03-31, Ano03-32, Ano03-33, Ano03-34, Ano03-35, Ano03-36, Ano03-38, Ano03-39, Ano03-40, Ano03-30, Ano03r, Ano03s, Ano04c, Ano04d, Ano04e, Ano04f, Ano04g, Ano04h, Ano04i, Ano04j, Ano05h, Ano05i, Ano05j, Ano06m, Ano06n, Ano06o, Ano06p, Ano06q, Ano06r, Ano06s, Ano06t, Ano06u, Ano06l, Ano06v, Ano06w, Ano06k, Ano07j, Ano07i, Ano07k, Ano07l, Ano07m, Ano08-28, Ano08r, Ano08s, Ano08t, Ano08u, Ano08v, Ano08w, Ano08y, Ano08z, Ano08-27, Ano08x, Ano09s, Ano09v, Ano09w, Ano09x, Ano09y, Ano09z, Ano09-27, Ano09-28, Ano09-29, Ano09-30]. Front [Ano09-31, Ano09-32, Ano09-33, Ano09-34, Ano09-35, Ano09-36, Ano09-37, Ano09u, Ano09t, Ano09-38, Ano09-39, Ano09-40, Ano09-41, Ano10o, Ano10p, Ano10q, Ano10r, Ano10s, Ano10t, Ano10h, Ano10i, Ano10j, Ano10g, Ano10k, Ano10l, Ano10m, Ano10n, Ano10u, Ano10v, Ano10w, Ano10x, Ano11p, Ano11q, Ano11r, Ano11s, Ano11t, Ano11u, Ano11v, Ano11-40, Ano11-33, Ano11w, Ano11x, Ano11y, Ano11z, Ano11-27, Ano11-28, Ano11-29, Ano11-30, Ano11-31, Ano11-32, Ano11-35, Ano11-36, Ano11-37, Ano11-38, Ano11-39, Ano11-34, Ano12k, Ano12-37, Ano12l, Ano12m, Ano12n, Ano12o, Ano12p, Ano12q, Ano12r, Ano12s, Ano12t, Ano12u, Ano12v, Ano12w, Ano12x, Ano12y, Ano12z, Ano12-27, Ano12-28, Ano12-29, Ano12-30, Ano12-31, Ano12-32, Ano12-33, Ano12-34, Ano12-35, Ano12-36, Ano11i, Ano11j, Ano11k, Ano11l, Ano11m]. Front [Ano11n, Ano11o, Ano13d, Ano13e, Ano13f, Ano13g, Ano13h, Ano13i, Ano13j, Ano13k, Ano14u, Ano14i, Ano14j, Ano14k, Ano14l, Ano14m, Ano14n, Ano14e, Ano14f, Ano14g, Ano14h, Ano14o, Ano14p, Ano14q, Ano14r, Ano14s, Ano14t, Ano14v, Ano14w, Ano14x, Ano14y, Ano14z, Ano14-27, Ano15g, Ano15-45, Ano15-46, Ano15-47, Ano15h, Ano15i, Ano15j, Ano15k, Ano15l, Ano15m, Ano15n, Ano15s, Ano15t, Ano15u, Ano15v, Ano15w, Ano15o, Ano15p, Ano15q, Ano15r, Ano15-28, Ano15x, Ano15y, Ano15z, Ano15-27, Ano15-29, Ano15-30, Ano15-31, Ano15-32, Ano15-33, Ano15-34, Ano15-35, Ano15-36, Ano15-37, Ano15-38, Ano15-39, Ano15-40, Ano15-41, Ano15-42, Ano15-43, Ano15-44, Ano16b, Ano16c, Ano16d, Ano16e, Ano16f, Ano16g, Ano16l, Ano16m, Ano16n, Ano16o, Ano16h, Ano16i, Ano16j]. Front [Ano16k, Ano16p, Ano16q, Ano16r, Ano16s, Ano16t, Ano17e, Ano17f, Ano17g, Ano17h, Ano17i, Ano17q, Ano17r, Ano17s, Ano17t, Ano17u, Ano17v, Ano17w, Ano17x, Ano17y, Ano17j, Ano17k, Ano17l, Ano17m, Ano17n, Ano17o, Ano17p, Ano18h, Ano18i, Ano18j, Ano18k, Ano18l, Ano18m, Ano18n, Ano18o, Ano18p, Ano18q, Ano18r, Ano18d, Ano18e, Ano18f, Ano18g, Ano18s, Ano18t, Ano18u]. Frontiers [Duw17]. Frontline [MA14]. Frozen [Tho14]. Fruitful [Vog91]. Frustrated [Nak74, RR15]. FSGO [MPC91]. FSGO-based [MPC91]. Function [Aga13g, Bec76, BM88, Bur74a, Che11, Fab84, Ges74l, Hem88, Lau96, MV80b, MV80c, Pan91d, Pro95, Pro09, Sar97, You74]. Functional [Chr11, Gro82k, Men70, NNEH11, Per95, Str15]. Functionalism [Ayd05, Blo76, Jar86h, Man71, Pro95, Szt79a]. Functionalist [Bun12b]. Functions [Aga03g, Ano09-45, Ges69a, Hal84, Jan84b, Koç05, Web12, Wri76]. Fundamental [Ano06x, Čap71-39, Dil06d, Duh15f, Grü09, How97, Kow16e, Pal01, PQQSe96, Phi90b, Sal74]. Fundamentalism [Pop92, Yeh17]. Fundamentals [Oru95]. Further [Ano05n, Car89b, DO91, Jah94, Mil09, Onn91m, OB91a, OB91b, Onn91c, Onn91d, Onn91e, Onn91g, Onn91i, Onn91h, Onn91j, Onn91k, Onn91l, Rip11, TO91, Vin92, Zam92i, Coh02a, Onn91f, Onn91b]. Fusion [Hor88, Man15b]. Future [Ano08-45, Ano08-35, BMP09b, Bou17, Dil06g, Emb92a, For01, Gel87, Gie71, GC95, Grü73b, HN16, Jan89l, Kap80l, Kap80n, Kow16a, Lan94, McB94, Mey85m, Nic80c, Res83, Sel03, Sha74b, Wil12]. Fuzhou [CW18]. Fuzzy [AM83].

G [Sal71, Onn91g]. G. [Coh90, Duh91k, Jan89a, MP16]. Gabino [Gut95]. Gadfly [Cro69]. Gain [Cha09a, Wet95]. Galilean [Ela11f, Cha93a, Fin09c, Fin09d, Giu04, Góm12, Góm13, Lei04, PT04, Vil04]. Galilei [Duh91f, She90]. Galileian [Gat18]. Galileo [Lan90, Aga81q, And91, Ano08-42, Bec91k, Bra95, BDR04, But93k, Cha93a, Coh04, De 16, Duh91f, Ela11a, EF18, FR08, Fin80b, Fin80d, Fin80e, Fin80f, Fin80g, Fin80h, Fin80i, Fin80l, Fin80s, Fin80r, Fin09b, Fin09e, Fin09f, Fin09g, Fin09i, Fin09j, Fin09h, Fin09k, Fin09m, Fin17, Hal90a, Hem84, Ito98, LS90, Mar11a, Pal04a, Pit87, Pit89, Ren90, Sch08a, Sch08b, She90, She04, Val10a, Val10b, Val10c, Val10g, Wal76, Wal81a, Wal81g, Wal81h, Wal81j, Wal81k, Wal81i, Wal81l, Wal81m, Wal81p, Wal81q, Wal90, Wal92c, Wal92d, Wal92e, Wal92f, Wal92g]. Gall [Pog94]. Galvani [Per87]. galvanic [Onn91j]. Gambling [Res89]. Games [Lew76]. Gamow [Lip86, Stu86]. Ganot [Akb11]. Gap [CCG93, Rop86]. Gaps [Cam87]. Garden [Cha09m]. Gas [Mer70]. Gases [Gav90, Hie95a, Knu07, Onn91a]. Gassendi [Pal04a]. Gates [Ion15]. Gauge [Hoo04]. Gauss [Bot94]. Gavroglu [ARS15]. GDR [Krö91]. Gedanken [Bel03]. Geht [Vaj99]. Gender [HK92]. Gene [Fon96]. Genealogy [Nol03f, Nol03i, O’C02, Smi94]. General [BP14a, Bre09, Bun10g, CG16, CS79-43, Coh02s, D’A01d, D’A01e, Fle86e, FA76b, Goo77f, Grü73k, Grü73l, Grü73o, Grü73u, HR74d, Hof74, Jar17, Lag97d, Lag97e, Lag97f, Lag97g, Lag97h, Lag97i, LC07, Mar78b, Mar84e, Mar84f, MV80b, Pap74a, Sac67, Şak11, Sau03, SH91, Yil05, Zin73f, Zin73h, dB87j, vL16f, JNR^{+}07b, JRN^{+}07b, JNR^{+}07a, RSS^{+}07a, RSS^{+}07b, RS07c, Ren07d]. Générale [Onn91r]. Generalities [Wim92]. Generality [Has97]. Generalization [Dra88, Nic76]. Generalizations [Eel10a, Hau10]. Generalized [Gin18, Kau74, Lap95]. Generate [Buc95]. Generation [Men76]. Generations [Tur86m]. Generic [Goo92]. Genesis [Bel92b, Čap87a, CS79p, Fre86f, Gai94, Hie70, JNR^{+}07b, JRN^{+}07b, JNR^{+}07a, Kuz87f, Mar86e, Mar84g, Mül91b, RSS^{+}07a, RSS^{+}07b, Rog81, Tor99, Vil02h, ZRKC03c, ZRKC03d, Koh95]. Genetic [Cap15, Dam96c, Ech95, MO82, Nor12, Vil02i, War79e]. Genetics [Dar80, Dav86, Fal86, GM17, Gro01, Gua96c, MG11, PQQSe96, Pia90, Rus76a, Sar92b, Sar92a]. Genidentity [Pad11]. Genius [Aga81j, Per86, Pol74a]. Genome [For01, Tur12]. Gentlemen [Rud86]. Gentlemen-Specialists [Rud86]. Gentzen [Neg81]. Genuine [Bun10h]. Genuinely [Aga76b]. Genus [Gre74l]. Geographic [ZC96]. Geographical [Bra91]. Géographie [Dor92]. geography [SCD03b, Dor92]. Geological [CALP03, SF94]. Geology [Fre91, Ham91, Mot15, Şen05, Yan93]. Geometrical [Duh76d, Heu96, HR74c, HR74d, HR74f, HR76, LC07, Mil15]. Géométrie [Sas03a]. Geometrisation [Lüt98]. Geometrodynamics [Grü73k, Grü74b, Mis74, Sta74e]. Geometry [Ana97, Bot94, Bri96, Cof83, CS79n, Fre74b, Gra07b, Gra06, Grü73g, Grü73h, Grü73i, Kar90, Kli96, Kno04b, Koç05, MA15, Men70, Mit76d, Nai15, Pfa96, Pul06, Rob03, Sig94, Web87, Wey07, dB87h, dB87i, vH77d, vH77e, Grü69a]. Georg [Dah90, Fri90b, Fri90a, Hel74, KPC90, Ned90b, Ned90a, Sca90a, Vro90, WW90]. George [GP79]. Georges [Die14]. Geostatic [Fin09l]. Geostaticians [Duh91c]. Gerard [And84b, Zam92e]. Gerd [Pap03, Pei03]. Gerlach [Sau16]. German [Bec04, Bie74, Buc91, Che97, D’A01h, Det07, Fin91, Gre91a, Hof74, Hül96, Kat14, Kis02, Mit04, Mül91b, Mün14, Nol03l, Pes91, Sch86c, Sch99c, Sch99d, Ste15c, Str84, Sza04, Vaj99, Van15, Vog04, Vog91, Wen14, Wol17, Woo91, WC91]. Germany [Chr94b, Fra91, Gre91b, Jon91, Kar10f, Moc91, Pog94, Pye87, Rau91, Row96, Row06, SS91, Tho14]. Germs [Tul16]. Gerson [Gol91]. Gerstmann [Ges74i]. Gestalt [Chr11]. Gesture [Tha84a]. Gets [Sta92b]. Gettier [PR17]. Getting [O’D02]. Geworden [Sza04]. GHZ [Ara97]. Gibbs [Del95, Nee11g]. Gifford [Hoo99c]. Gigerenzer [Pap03, Pei03]. Gilbert [Jam76, Tau91c, ZRKC03g]. Giora [Cam03, Wol03]. Giordano [Höf76]. Giovan [Gat18]. Giovanni [Omo18a]. Girard [Liv92, McK92]. Giuseppe [Lai08]. Given [Hoo99c]. Giving [Rad79]. Glance [BE10a]. Glass [Har95, Nic89b]. Gleichungen [Vog04]. Global [BBD18, Ber96a, Dol16, Fel15, Kap80m, Kap80n, Pos95q]. Globalization [Ren15]. Globule [Sch09]. Glory [Uus01, Web87]. Gloss [Mue95]. Glycolysis [Oes91]. Gnomon [Lih93b]. Gnomons [Lih93b].

Go [Aga75b, Ned90a]. Goal [Fab84, Zol89i]. Goals [Sha84a, Bie74]. God [Ano08-43, Bab02c, Kea02, Mar95d, Ols87, Ric91, Wil07, Zam92m]. Gödel [PP15b, DD11, Ken17, PP15a, PP16, Web83]. Gods [Whi88, Scu98]. Goethe [Alt87, Amr87a, AZW87, AZ87, Bar87, Böh87, Bra87, Bur87c, Ham91, Heg87, Kuh87, Len87, Mar87, MA87, Mil87, Por87, Sep87, Wet87, Zaj87, vW87]. Goethean [Amr87b]. Gogh [Alt02, Bab02c, Mar02]. Going [Fre09, HSS09a, HSS09b, Sta09]. Golden [Aga81-36, Wim92]. Goldstein [Ges74s]. Gombrich [Hor94]. Gone [Räz16]. Gong [Fey78a]. Good [Can16, Kap80k, Koh94, Koh95, Koh98h, Mar84p, Pap89, Rad79, War91]. Goodman [Lin93b]. Goose [Aga03-36]. Got [Gav09]. Göttingen [Bit96d, Sig94]. Governance [Bog15b]. Government [Byr83]. Gradual [Jam76]. grain [WRL95]. Grained [Grü73m]. Grammar [But93c, Har97b, Hül96, Sau02, Sim99]. Grammarian [Mar79w]. Grammatical [Ano06-28, Bar76c, Sch76d, Wun76]. Grammaticalness [Coh02k]. Granada [Arb92]. Grand [McB94]. Grandeur [Mün14]. Graphic [BV01, Ton81d]. Graphical [Ano08-29, Rit01]. Grasp [Fer85d]. Grassmann [Beu96, Bra96, Bri96, BT96, Dor96, Fra96b, GG96b, Hes96, Heu96, Hül96, Law96, Lew96, Nád96b, Pec96, Pfa96, Row96, Sch96c, Sch96d, Sch96e, Sch96g, Sch96h, Swi96, Tob96, Tur96, Sch96f]. Grassmannian [Kli96]. Gravesande [Rey18]. Gravitation [Abr07b, Abr07d, Abr07c, Ano07a, Ano07f, Ano07e, Ano07o, Ano07r, Ano07s, Ano07h, Bel12j, CS79-29, CS79-37, Ein07b, Fin80q, Goe70, Lor07a, Mey85j, Mie07c, Nor07c, Nor07b, Nor07d, Ren07a, RS07b, Ren07c, SM07, Val96, Wal07b, Zen07, CS79-28, RSS^{+}07a, RSS^{+}07b]. Gravitational [Ano07n, EFN03, Mie07b, Nor07a, RS07a]. Gravities [Des03]. Gravity [Dob92, Duh91j, Hat80, JRN^{+}07a, Ric12, Sta07c]. Great [ML99, ZC96, SR87]. Greater [Hoo88a, Web87]. Greece [Kno04a, Nie74, Pap14, Pap94]. Greek [Cha09c, Chr94a, Chr04a, Fow04, GP03, Kal90, Kit90, Kno04d, Mac97h, Mac97i, Net05, Nic90, Pat03a, Pat11a, Scu98, SM90, Too04, Ung94, Ung04c, Vla94, Was91]. Greeks [Cha09l]. Green [BLA95a, Nae91, Hie95b]. Greeno [Sal71]. Grelling [Chr11, Pec11]. Grene [DPW86a, DPW86b, DPW86c]. Grice [Coh02r]. Grossman [Kuh09]. Grossmann [FM09b, FM09c]. Grossmann-Thesis [FM09b]. Ground [Jan89c, Mor11, SM90]. Grounding [Das88, Pog16]. Grounds [Hal91c]. Groundwork [Sha90]. Group [Bir09, DD12, Rud86, MP13, Mil11a, Pec11, Pul11, Res11, Wim92]. Grow [Cha09m]. Growth [Bun12u, Fin80q, Gie71, Hee79, Koe71, Lev83c, Lev83d, Pan83e, Pan83f, PM82, SM15b, Sne89]. Grue [Hüb78a]. Grünbaum [Qui83, CL83]. Grundgesetze [Boc15]. Guesses [Boy09]. Guessing [Coh02f]. Guide [Bas14, Fre00]. Guideline [Kow16a]. Guiding [Kow16k]. Guido [Duh91k]. Gunpowder [Cul11, Ágo11]. Günther [Heu96, Pec96, Sch96c, Sch96f]. Gustav [Bor07, Hei94a].

H [Cha12a, CS79x, Duh02x, Scr74a, Wei06, Cha12d, Cha12e, Onn91i, Onn91h]. H. [Bri96, Cro92a, NA07, Sch06a, Tob96]. Habermas [Dav99, Fre99, Pen99, Rad12b, Rad12c, Rad12d, Roc99, Str99, Tau99]. Habermasian [Hod99]. Halakah [Twe00]. Haldane [Sar92c, Smi92, Wim92]. Half [D’A01n]. Halfmann [Jan84b]. Hall [Onn91j]. Hall-effect [Onn91j]. Haller [Jac97, LM97, Mar97a, Nyí97, Pot97, Smi97, Wol97]. Hamann [But93c]. Hand [Kiv01]. Hander [Ges74c]. Handled [Kap80c]. Handles [FH10a]. Handmaiden [Syl75]. Hands [Hoo99j]. Hangover [Žva79]. Hans [BCK07, Gre07, Hei08c, Hei11, Ihd79j, Kam11, NG09, Pad11, Zem09]. Hanson [CW67]. Happened [Vil02m, vM89]. Happy [Wol03]. Harakiri [Sag87]. Hard [RA01]. Hardware [Bun10c, Gen71]. Hardware/Software [Bun10c]. Harmful [Bal94]. Harmony [Ano08-44, Bec91h, Hoo99d, Toa15]. Harré [Suá08]. Harriot [Ano08-30, Ano08-34, Ano08-42, Ano08-47, Ano08-46, Sch08a, Sch08b]. Harris [Mar76d]. Harvey [Vil02i, BC94]. Hats [Pet89]. Hatton [CS79b]. Hauser [Wes95]. Havana [dMPC14]. Haven [Mar95b]. Hayek [vHBM01]. Haytham [Ras94a]. Hayyān [Haq94f]. Hazardousness [Tym15]. Hazel [Hie95b]. Health [PT99, Roc11, Wat01a]. Healthily [Byn89]. Healthy [Buc01, Man12, Man13]. Hearing [Sav15]. Heart [Fra00, Byn89]. Heat [Duh02v, Nee11m, Val10f, Van74d, DO91]. Heats [Kox95]. Heavenly [Ros81]. Heavens [Gen91, Wal92g, Zam92m]. Heavily [Wyl92]. Heaviness [FEE18]. Hebrew [Fre00, Kat14]. Heelan [Bab02c, Cre02, Cro02, Mar02, Nev02]. Hegel [Bri98, Com84, Bie95, Bub84, Buc84, Buc94b, Bur87d, Čap84a, CW84a, Dam96c, Di 84, Dov84, Eas84, Fer98, Gau84, Gre84, Han99e, Ley84, Mey91h, Mey91f, Mey91g, Pao84, She82, Sol84, Von84a, Von84b, Wol76k]. Hegelian [Fin84, Kos84]. Heidegger [Bab02b, Bam99, Fre86a, Gia99, Ihd79d, Jan89k, Rou16, Sch99d, Tam02]. Heideggers [Sch99d]. Heike [GG91, SR15, vD14]. Heinrich [BHN98, Bry98, Buc90, Buc91, Buc95, Hen98, Hof98, Lüt98, Lüt06, Maj98, Mul98, Nor98b, Sch98]. Heinz [FK93]. Heisenberg [Bel92b, Gou90, Hen98, Shi83, Shi86, Wüt16]. Heisst [Sch99c, Kis02]. Heister [Tra15]. Hélène [Fre88, Ker88]. Helium [Kap80v, Kap80-34, Onn91e, Onn91g, Onn91j, DO91, Onn91m, Onn91f, OB91a, Onn91b, OB91b, Onn91c, Onn91d, Onn91i, Onn91h, Onn91k, Onn91l, Onn91q, TO91]. helium-cryostat [Onn91f]. Hellenic [Ano06y, Pat15]. Hellenistic [Cla14]. Hells [Gen91]. Helmholtz [Bar87, Čap71t, Edg15, Hei98, Hör95, Hyd06, Ler95, Ord96, Pul06, Sch98, Tur96]. Helmont [BR15]. Hemoglobin [Hol95a]. Hempel [Lin93c, Mil11b, Rec11]. Heng [Zum96]. Hengfang [Yus96a]. Henkin [Bal18]. Henri [dAM11]. Henry [McC95b]. Henryk [FM09c, Kuh09]. Henseleit [NG09]. Her [DPW86a]. Heraclitean [Kuz87g]. Heraclitus [Lam82, Vam09h]. Herbart [Are89g]. Herbert [Hil94, SPB18]. Here [Seg15]. Heredity [Ano08o, Ano08-45, Ano08-40, Ano08-32, Sch08d]. Heresies [Mar79q]. Heretical [Ano07g, Her74]. Hering [Kre91]. Heritability [Dan76]. Heritage [Vod15]. Herman [Lõh01]. Hermann [Fin91, Hül96, Pec96, Sch98, Sch96c, Sch96d, Sch96e, Sch96f, Sch96h]. Hermeneutic [Bab02c, Bab02d, Ege99, ÉA99, Gin02, Hee99, Kat11c, Koc97, Sch99c]. Hermeneutical [Fre88, Ker88, Lij99]. Hermeneutics [Aga03e, CS02, Cre99, Dra99, Feh99a, FKR99, Føl99, Fre99, Gin97, Hee97, Ihd99a, Ihd99b, Jan89i, Jan02, Kam99, Kat11d, Mar86g, Sch02a, See02, Tou02, Wol95g, Zen99]. Hermeneutischen [Sch99c]. Hermeneutisches [Ape99]. Hermes [Lar99]. Hero [Cao07, Nae91]. Heron [FR08, Sch08c]. Herrings [Wyl92]. Herschel [Aga81-29]. Hertz [vH77a, BHN98, Bry98, Buc90, Buc91, Buc95, Buc98, D’A98, D’A01f, D’A01j, D’A01i, Don98, Gra98, Hei98, Hen98, Hof98, Hon98, Jan02, Koc02, Lüt98, Lüt06, Maj98, Mul98, Nor98b, Sau98, Sch98]. Hertz/Moritz [vH77a]. Hertzian [Buc98]. Herve [Le 08]. Hesitations [Bru76]. Hess [Fra80]. Hesse [Mor80]. Hessen [FM09a, FM09b, FM09c]. Heterochrony [Han15]. Heterogeneity [Čap71l, Čap71p]. Heterogeneous [Kuz87h]. heterotrophic [RL95]. Heuristic [Bağ05, Pol74b, Sch12b, Sim77k, War67a, War79h]. Heuristics [Alb84, Bro93, FK93, Gig03, Kam93, Koe93, Mic18, Mus78, Mus89a, Nic87, Pos93, Sca03, SPB18]. Hiżian [Mar79q]. Hidden [Ara12, Bun95a, Čap91k, Rös94, Sie63]. Hidden-Variable [Sie63]. Hiebert [NRS92]. Hierarchical [De 74]. Hierarchies [Wak15b]. Hierarchy [Vin92]. High [Cue01, Gav89, Har08b, Kap80o]. High-Altitude [Cue01]. High-Temperature [Har08b]. Higher [Coo17, Ges69a, Ges74l, SA12, Woo00]. Highly [Hoy74]. Hilary [Grü69a]. Hilbert [Abr81, Cor07, Hal95b, Maj98, Maj06, Pra93, RS07c, Sau07, Sig94, Sta15, Tor95, dC95]. Hillel [Kas76b, UM92]. Him [CS79-30]. Hindsight [Sim77r]. Hindu [Coh88b, ZC88d, ZC88e, ZC88f]. Hippasus [vF04]. Hippocrates [And90]. Hire [Gün11]. histoire [CLR^{+}92]. Historian [Bla92b, Che14, SR07, Vil02g]. Historians [Can12, FF07, Gol12, Hac94b, Stu07]. Historical [Aga08b, Ara79, BV94, BN17, Bra74b, But93d, CSW74, CW74b, CS79c, Coh86a, Coh94b, Coh02q, Coh04, D’A01k, Dam96d, Dea12, Gol83, GHQ96a, GR11b, Ham91, Hel00f, HR74c, Hon16, Ihd83, Jie96, Kam11, Kat11f, KP86, MO82, Moo18, Nee11g, Nic80d, dJAO14, PQQSe96, PJM92, Ren95, Rös91, Ros09, SS16b, Sha90, She87, Sta94b, Sto79, Tou74b, Tou86b, War79m, War86, War96, Yuz96, tH09, Duh15c]. Historicism [Sza79a]. Historicity [Rös94]. Historicized [War87]. Historico [Gay09, ZRKC03i, Zol89i]. Histories [Aga75c, Fu93b, Ken96, Ste94]. Historiogenesis [Dam96g]. Historiographic [Aga08c]. Historiographical [D’A92, Gav14a, Kno94, Ott94, Sch09, Zam92c]. Historiography [Aga08d, Aga08e, Bec91o, Bir15, DN94, FM09b, GCN94, Hal91e, Kin16, Mac94a, Mam94, NG15a, Wuß91]. History [Aga81d, Aga08a, Aga08d, All01, Ams92b, Ano06z, Ano08k, Apt74, Ara12, ARS15, Ara79, Are89b, Ash07a, Ash07b, Bam99, BRW14a, Bau17, Bel12b, Bel81a, Bel90, Ber95a, Ber17b, BP14b, BS15, Bla95g, Bla94, BE10b, BM89, BRS18, Cao07, Cha12b, CLR^{+}92, Che05a, Che05b, Che14, Che97, Chr94a, Chr04a, Chr14, CS79-35, CS79y, CS79-41, CW84b, CW84c, CL17, Cue01, D’A00, DC96, Dam96i, Dar07, Dra88, Fin80j, FS96, Fox07, Fre88, Fu93b, Gäh89, Gav94, GSW95c, GR07, Gav14b, Gav14a, Ges74e, Ges74s, Gie12, Gin07, Gol69b, Gou94a, GWT00g, GG96b, Grm81, GB05, Grü73c, Grü73r, GHQ96b, Hal71, Hal91b, Ham91, Har99, Hol95c, HOO88c, Hoq18, HFD88, IG05c, Irz15, Jan85b, Jan85e, Joh11, Jur08, Kap80p, Kat11d, Ker88, Kno04d, Knu95]. History [Kox95, Kra07, KV11, Krö91, Lac94, Lak71a, LS90, Li18, Lia96c, LF93, Lin82, Lop92, Lug89, Mai01, Mai12, MC95, Mat95, MS12b, May95, McR80, Mcm76b, Mil12, Mit89, Moc91, Nak74, Nic90, Nou90, Nye07, Nye15, Omo18a, Ott94, OP97, Pal76, Pau17, Pec94, Pel01, Per95, Per94, Pie12, Pol92, Pol95, Rad12e, RH92, RC95, Ren15, Ric15, Rim97, Rit01, Rot87h, Rot87j, Rou15, Rui95b, San92a, Sar96, Sar08, Sch94a, SR16, Sha84t, Sha86b, Sim15, Ste15a, Str95, Syl08, Sza84, Tis03, Tou76, Tym15, UM86a, UM86c, UM88, UM92, Ung04a, Ung04c, Vih01b, Vil02i, Vil02m, Wal95, War76, War79p, Wen96, Wes95, Wet17, Wis07, Wol95k, Wüt16, Ytu95, ZRKC03e, AM01b, D’A01l, For07, Gar12a, Gar13, NG01c, RG07, CS79-41]. Hittorf [Hie95a]. Hoax [Bab02d]. Hobbes [Fre86e, Fre86k, Lei04, Wat76]. Hoc [Aga75s].

Hodes [Hod74a]. Hoff [Duh02x]. Höfler [BIT01a]. Holism [Gib00, LW11]. Holland [Pop88]. Holocaust [Got94, Sch86c]. Holonomic [Tul03]. Homeostasis [Bu96]. Homo [Dju79]. Homology [Wag15]. Homoplasy [Wak15a]. Hon [Cam03, Wol03]. Honest [And03b]. Honor [AC82, ARS15, Bab02c, CSW74, DJ89, Fet10a, GSW95c, HOO88c, KS95, LM97, MC95, MC96, NRS92, RDS03, GSW95a, GSW95b]. Honoré [Ela11j, EF18]. Honour [CL83, FK93, LS90, GC97]. Hooke [DS11b, Gal02c, Hun10]. Hooker [Tis74]. Hope [Nem91a]. Horror [Par17b]. Horse [Kim11]. Hospital [Lek15e, Lek15g]. Hotel [SC74a]. Houses [Mit76a]. Howson [Dou03, Nii03]. Hua [Yus96a]. Hübner [Hüb78a]. Hubris [Kid17]. Hückel [Kar10a, Kar10c, Kar10b, Kar10d, Kar10e, Kar10f, Kar10g]. Hugo [Bar14b]. Hugues [MC95, MC96]. Hui [Dun96, Ron96]. Human [Abr11, Ano05n, Ano08o, Ano08-45, Ano08-40, Ano08-32, Bri98, Bus03, Cho76, Chu00, CT98, Dan76, DD12, DPW86a, Fek99, For01, Gal91, Gar12b, Gel89, Ges74m, Ges74n, Ges74u, Gro82n, Hun83, Ihd79c, Ihd79j, Ism16, Jan89b, Kae90b, Koh98c, Kul01, Lee84, Man12, Man13, Mar89, Mey91k, Mic98, Näp01, O’C02, Pan91c, Pet03a, Pia90, Pro01, Qua84, Ros98, San97a, Saw98, Sch08d, Sch84, Sim11a, Ste17, Tha86b, Wat76, Wat95a, Web11, Wim12a]. Humanism [Kri03, Vil02l, ZRKC03f]. Humanist [Hel08]. Humanistic [ZRKC03k, GSW95b]. Humanities [Ber95b, Fel15, JL95b, KT01, LMZ18, vDP17a, vDP17b]. Humanity [Aga03-31, Ng17, Rod17]. Humanizing [War79c]. Humans [Lel00e]. Humboldt [Bie74, Det07]. Hume [Dem05, Mun89, Ror86, War79f]. Humean [Ano08-31, Ano09m]. Humeanism [Pil97]. HumeanProjection [Ano09m]. Humility [Kid17]. Humphry [Abb94]. Hun [Lih93a]. Hunches [Was91]. Husserl [BIT01e, Car15a, Hin97, Mat01, May02, Mic97, Tau97b]. Husserlian [CS02]. Huygens [Vil04]. Hybrid [Kiv94, Kra92b]. Hydraulic [Sun18]. Hydrodynamics [Lag97n]. Hydrogen [Row14, Sch91a, Onn91p]. Hydrostatics [Lag97o, Mal12, Mal13]. Hypercomplex [Sch96c]. Hypernumbers [SM74]. Hypersurfaces [Rob03]. Hypertext [Rad03]. Hyperuniverse [FT16]. Hypostasis [Chr15]. Hypostatization [EE17]. Hypotheses [Aga75s, But93f, Coh69, Goo69, Grü05a, Kar90, Kra82b, Lau80b, May69, Mor89, Ram01, Rog71, Tri12]. Hypothesi [Put67a, Put86a]. Hypothesis [Aga75u, Coh02d, Duc89, Dum92, Fal01, Fra80, Gat09, Grü73e, MC80y, Pat76a, Pet89, Sch09, Sim77x, Vil02h, Zam92l, TO91]. Hypothetical [Flo15]. Hypotheticism [Kra82b]. Hypotyposes [Koc89]. Hysteria [Rub83, Sav15].

I. [BC94, GP79]. IAM [Kow16m]. IBE [Per07]. Iberian [GPSR01]. ibn [Haq94f, Fre91, Ras94a]. ICBM [KJ14]. Ice [RA14]. Ice-Making [RA14]. Iconoclasm [Hor94]. Icons [All99b]. Idea [Abb94, Aga13q, Ano06-31, CS79k, Gre75, HR74c, Hin97, Jan89k, Kit90, Kra74, Meš89, Mil88, Sam76b, Smi94, Vra79, Yli07]. Ideal [Ams92h, Ams92j, Ber84, But93a, Čap71w, Eck17, Kir82, Lel00d, Owe99]. Idealisation [Mus95]. Idealism [Ste15c]. Idealists [Vil02i]. Idealization [Eck17, Fis15, PT99, Sta82]. Idealizational [Bra95, Han99f, Łas82, Szy82]. Idealizations [GS09, Jeb17, Say05, Tel08, Toa15]. Idealized [Gen03]. Ideals [Ams92c, Ams92e, Sha84e]. Ideas [Ber17b, BCK07, BC02, Cho67, Cho86, CS79y, Cue01, D’A00, Feu88, Fre11, Fre91, Gig03, Her74, Kum11, Mar86e, Men70, Nád96b, Øhr97, Ree87a, Rot86b, Šuš79, Var92, Woo02, D’A01l]. Idée [Gre75]. Identical [Mar95c]. Identifiability [Sim77d, Sim77j]. Identification [Hol95a, Mey91i]. Identifying [Wol76z]. Identity [Aye76, Bri15, Bun93e, CdF95, Eel10b, Fey67, Hah91, How97, Kat76, Mar78c, Mar78f, Mar78h, Mey91i, Omo18a, Par00, Pat15, Stu97, War79f]. Ideological [Aga81k, Bre88, Hal84, Jan84b, Kap80-28, Rav91]. Ideologies [Hor88, Rot87g]. Ideology [Die83b, Hod92, Kim96, Koe91, Koe93, Kuv79, Mar78a, McG87, Mon01, Pan83c, Sal84, SS91, Tru74]. If [Mar95c, Piz81, Voi96]. Ignorance [Hoo94]. Ignore [Aga75u]. Ihre [Bec04]. II [Ras00, AM63, Ano08-32, Bab99a, CS79d, FS94, Gro82j, JL95b, Lag97i, Mar78e, MC96, MC80e, MH86, Nol03d, Roz71, Sai04a, Sha84i, Syl91, Vog91].

III [Kow16c, Onn91c, OP91a, Syl91]. Ill [Sim77w]. Ill-Structured [Sim77w]. Illicit [Mar79v]. Illness [Von84b]. Illusion [Bau88, Čap91d, Hor94, Mar86c, McB94]. Illusions [And76, Tve86]. Illustration [War67b]. Ilya [Näp01]. Image [Bab97, Lüt06, LE97, Sch98, Syl97a]. Imageless [Bro91a]. Imagery [Čap71j, Ols87]. Images [ARG12, Glu91, GC97]. Imagination [Bar92, Bor99, Cla02, Coh17b, Hal82, Hal92, Mey85w, Saa17, War92]. Imitation [Cif05, Onn91k]. Immanuel [Web87]. Immature [Met89]. Immediacy [Čap71-29]. Immediate [Arm76, Kap80n]. Immodest [War79i]. Immortality [Čap71q]. Immunological [Löw91]. Immunology [CT91, HGR^{+}89]. Impact [AP00b, Fre74b, Hüb76, Kan87, Kno04c, Sie63, Vas92, BJ95]. Impasse [Lao17]. Imperatives [Flo15]. Imperfect [Aga75f]. Imperial [BE10a, BE14, Jam10]. Impetus [Ela11d, Ela11b, Ela11i, Ela11j, Ela11k, Ela11l, Ela11m, EF18, FEE18, Sar08, Wal81a]. Implications [Čap71-34, Cho14, Coh83, Dai16, De 81, Ear06, Ech95, Grü84, Jam79, Koç05, Kow16h, Len86, Lev71, Mar74b, Men70, Mey85v, Pos95i, Res97, Sch15, Sol12a, Soy05, Swa09]. Implicit [Bre74a, Sup74, Wal92j]. Import [Aga81k, Hol83, Ihd79b, Mol18]. Importance [Bro91b, Hof74, Onn91n]. Imposing [Kom99]. Imposition [Fel89]. Impossibility [Duh91g, Pan91i, Ste97]. Improve [Pie95]. Improvement [Fey67, BdOU^{+}95]. Improves [Wor78b]. Improving [Irz03, BBT95]. Impure [Kop99]. Imre [Ara94, Bar76b, CFW76, Dai89, GGN89, Gre76b, Wor76]. In-Between [Pap15, Bal02]. In-De [Syl97b]. In-Depth [Kow16h]. Inadequacy [Čap71r]. Inclined [Ano08n, FR08]. Inclined-Plane [Ano08n]. Including [Ano07o, Hod92]. Inclusive [Coh86a, KP86, Leb95]. Incommensurabilities [SSHH08]. Incommensurability [Bab02a, Bal89b, Bar01a, Bir08, Bro01, Bur84, Dev01, Dop01, Dra99, Far03f, Fow94, Gie08, HHS01, KN01, Lac01, Łod82, Mar15, Odd89a, Per91, San08, Sha01, Sie01, Sol08a, SL84, vF04]. Incommensurable [Pan83d, Tel08].

Incommensurate [Nye91]. Incompleteness [Čap71s, Qui96]. Incompressible [Lag97c, Lag97m]. Inconsistencies [Pia82]. Inconsistency [DF93, Tra15]. Incontinence [Mit86a]. Incorporating [Bro92]. Incorporeality [Lin91]. Incorruptibility [Gra91]. Increase [Fey78b]. Increased [Len83]. Indefinite [Woo00]. Independence [Arb92, Čap71-27, San15, Tel76, Wet17, Hin95]. Independence-Friendly [San15, Hin95]. Independent [Die14, Hug88, Şen05, Wet09]. Indeterminacy [Ano08-48, BB71, Pag00, Pan91d, Par00, Que97, Seg00]. Indeterminate [Chr94a]. Indeterminism [Bac90, Pos95a, Stö11, Web93]. Indeterministic [Fet10d, MC80q]. Index [Ano07p, Ano07q]. Indexical [Cla96, Tho76]. India [Kri92, Kum92, Kum11, San92a]. Indian [Ano06c, Ano06z, Ano06-27, Ano06-28, GC06, Rai10, Vas92]. Indicative [Sta76b, Tha84a]. Indicator [TBC91]. Indirect [Ano09q, Coh02e]. Indiscernibles [Sau03]. Indispensability [PS15]. Indispensable [PS15]. Indistinguishability [Bac90]. Individual [Fre86f, Fre86g, Gra83i, Mar79h, Sto00, Tur86c, Zim86]. Individualism [Cur89b, O’H09, Tuo82]. Individualist [Att13, Ber96b]. Individualistic [Cou96, Mit00]. Individuality [Phi90b, Kra11]. Individualizing [Tym95]. Individuals [CD93, Rot87b, Sob91, War79l]. Individuation [Coh02h, How97, Nak98]. Individuations [War79f]. Indivisible [Čap71l]. Induced [Ges74p]. Induction [Ano09r, Arm95, Bec91p, But93n, Fei71, Fet81d, Gar03, Gil91, Gil09b, Lug89, Nol08, Pop84, Pos93, Riv96, Sal18, Tok09, Wor89]. Inductions [But93b]. Inductive [Aga13h, Ano09k, Bun12f, Coh02i, Con81, Hal90b, Lin93c, Mil95, Per81, Pol71, Sch97b]. Inductivism [Grü76, Grü78, RA78a]. Industrial [Pes07, RT01, RR15, dV11, NG01c, KB92b]. Industry [FQ18, NG03, Şak11]. Inequalities [Per97, DS12, Koç05]. Inequality [Vrc79]. Inertia [Fre74a, Mey85o]. Inertial [Bec91a, Nor07a]. Inevitability [Bel96, Bel03]. Inexhaustibility [Ela11b]. Infamous [Sch09]. Inference [AT15, BA09, Kur98b, Mil95, Oid98, Riv96, TG88]. Inferences [Lap95]. Inferential [Wim12a].

Infinite [Gas76b, Kno94, Koy76b, Loc76a, Toa15]. Infinitely [Uga18]. Infinitesimal [Bar10, Mil15, Phi94, Wey07]. Infinitude [Grü73d]. Infinity [Bai76a, Bec91b, Hor16]. Inflection [Gal02a]. Influence [Aga13b, Aic15, Bri96, Čap71j, Cek92a, CW18, CS79f, Ges69b, Jah91, Jon91, Kos94, Lam82, Lew96, Lip17, NA07, Pat92, Pec96, San92b, Sch86b, Sch96d, Sch94b, She75, Ste09c, Wes90a, Wet95, PO91]. Informal [Hul76b, Mor80, Pol16]. Informatics [Bun12j]. Information [Cac82, Cér86, Dal89, Gra83c, GG96c, HGR^{+}89, Har89d, Har89b, Haw67, KF86, Lev86, Lon93, Lub82a, Mar86d, Mit86b, Mit86a, MH86, Nei15, Rap86, Rop86, RG89, Spa03, Ton81c, Win86, Xiu96, Zim86, dB87g, Str86]. Informational [Mat89, Ols82, Ton81e]. Informative [Bec91e, Bec91i]. Informativity [Bec91k]. Ingarden [Čap71u]. Ingenii [Sas03h]. Inherent [Ela11l, Goh81, Han99h, Han99g]. Initiation [Are89g]. Initiatives [Hel00h]. Initio [MPC91, Sch91a]. Innate [Cho67, Cho86, Kri90]. Innateness [KK07, Put67a, Put86a]. Inner [Aga03h]. Innocent [Cro02]. Innovation [BRS18, Coh02g]. Innovations [Coh83, Löw14, Rod15]. Inostensible [Ina05]. Inquiring [Tör76]. Inquiry [Aga81n, Ant90, Bun10j, Cre99, DG15, Ful02, Koe78, Mis74, Sch18, Wol76-32]. Inquisition [Bab02d]. Insanity [Byn89]. Inseparability [Bel03]. Insider [Alt14a]. Insight [Ano05g, Čap71j, Gög87]. Insights [Aga09, See74]. Inspirations [Cha10]. Instances [Elg10]. Instantaneous [Čap71-30]. Instants [Čap71-41, Čap91f]. Institute [Ano08o, Ano08-45, Ano08-40, Ano08-32, Sch08d, Bie74, Kap80f, Kap80z, Kap80-38, WC91]. Institutes [RH92, Hei08f, Mou92]. Institution [Aga75j, Her14b]. Institutional [Bre88, Hei94b, MS11]. Institutionalisation [İhs05]. Institutionalization [Ams92f, Rus79, Sil11]. Institutions [Her14b, Kai07, Krö91, Rai11, Rod15, Wet17]. Instruction [Swa09]. Instrument [Aub17]. Instrumental [Fra99f, Gro82d, Jur08, MS11, Och96, Par09a, RS06]. Instrumentalism [Gie76, Mus91, Nie82, Ska83].

Instrumentation

[Gou94a, Ihd79e, Ihd79f, Ihd79g]. Instruments [Bar74, Lev90, RDS^{+}18, Val10d, Zam92m]. Insufficiencies [Sup79]. Insulated [Coh02p]. Integral [Bun12u, Lel00m, Shi87h]. Integrals [Pos95h, Pos95m]. Integrated [Kow16a, Kow16g, Kow16j, SR16, Wun76]. Integrating [Kou11, MS12b, Mit76a]. Integration [SRI^{+}92, Ste17, Tho14, Wak15b, Zim92]. Integrity [Jar86q]. Intellectual [Aga03g, Aga03j, Čap71j, Cha17, Ezr01, GP03, Gra07a, How15, Jar95, Mar86e, Mur75, Omo18a, Sch79]. Intellectuals [Kan87]. Intelligence [Aga03a, Ger95, Lam97, Mar86d, Ste17, Tho96, Win17b]. Intelligibility [De 01, Hoo94]. Intension [Wei82]. Intensional [Mar63]. Intensionality [Mar79i, Per18]. Intentional [Ano09-43, Web12, Wil12]. Intentionality [Kal87, Nad96a]. Intentions [Kau96]. Inter [Gie86, Koe71, Möl01]. Inter-Level [Möl01]. Inter-Theoretic [Koe71]. Inter-War [Gie86]. Interact [Lag97l]. Interaction [Ger15b, Kno04d, Kra92b, Mar78l, Pan91c, Pro95, Rot87d, War79g]. Interactions [Coh94a, HJLP06, MM91, Res11, Sto91, Tam14, vSB91, dBJD95]. Interactive [Coh02p, Pan83e, Pan91g]. Interchange [Win17a]. Interchangeability [Bac90]. Interchanges [GG03]. Interdisciplinarity [Mün14, Ren95]. Interdisciplinary [Gem90, Gol12, Mül91b]. Interest [Hoc79, Pen99, Vil02m]. Interesting [Vih01a]. Interests [BD88, Kul01, Len88]. Interfaciology [Ros99]. Interim [Sas03f]. Intermediate [Čap71t]. Intermolecular [vdA91]. Internal [Art15, Bur86, Elk86, Gäh89, Gau95, Irz15, Koh86, Nei15, Nii96, Pal01, Sol12b, Ytu95]. International [CHR96, Ezr01, FKR99, HFD88, Lop92, MS75b, Ree14, Vod15]. Internationalisation [CALP03]. Interpersonal [Aga03z]. Interplay [Elz95, Ols87]. Interpretability [Str18]. Interpretation [BV94, BS96, BEW96, BD74, Bun93f, Car74, Coh02a, Cus93, D’A01g, Eas84, Fin96, Fje02, Fow04, Fre99, GW97, Gre07, Hoo74, Jar86m, Kal96, Lam95, Mac12c, MM91, Mar02, McC95a, MC80g, Mey85k, Mit00, Odd89a, Ray82, Sch08a, SM89, Sha84h, Sha84i, Shi96, Sim77u, Sol84, Tou86b, Tur86f, Unt01, Von76, Zab95, Zal86, Żyt82, dB84]. Interpretations [Bel12c, Bub96, Kat11g, MC80n, Mel03, Sch63, Sha84d]. Interpretative [Len82]. Interpreted [Pal04a]. Interpreting [Mac12d, Mac12e, Sab92]. Interpretive [Koc02]. Interregnum [Tur86e]. Interrogationes [Zam92g]. Intersections [Fax97, Kat11b, Kat11c, Kat11d, NRS92]. Intervals [Bos76, Neu76a]. Interview [Bar14b, IG05a, dF14a]. Interwar [De 17]. Intimate [BM89]. Intra [Bar08, Ly08]. Intra-Theoretical [Bar08, Ly08]. Intractable [You76]. Introduction [Aga81l, AP00a, AM01a, AV16b, Ams92g, Ano05k, Ano05l, Ano06c, Ano08-33, Ano08-34, BSM06a, BRW14b, BMP09a, Bec91l, Bel93, Ber13a, Ber83a, BP14a, Bla95f, BIT01f, BR18a, BVH^{+}83b, Bör93, Bou02e, BH17a, Chr04b, CS79-43, D’A01f, DP92, Den99c, DB15a, Dil06e, DS01, Duh15g, DV92, Dur83, Ela11m, Elk76, Far03h, FSU97a, Fer85c, Flo17a, Fra99g, Fre86o, FA76c, Gal02b, GM89a, GWT00f, GCC81a, GR11a, Gut95, Haq94d, Hei08e, Her14c, HSS09b, Hoo99e, JNR^{+}07b, JNR^{+}07a, Kae90a, Kat06c, Kow16f, KB92a, Krü91, Kuz87i, Lek15d, LC17a, Mac97e, Mar78b, Mar84h, MV80h, MV80i, MS12a, MC80e, MM94, Mit86b, Mit76e, MS75a, Net04, Nol03g, Pal04b, Pan83c, Pat11b, Phi90a, Ras94d, Ree87a, RS07b, Sai04b, Sar92b, Sas03g, SS16a, SM15a, Ses04, Shi87i, SDG15a, SAR15, Sol08b, Sol12a, Sta74c, Tau91h, Tau97c, Tha86c, Tho14]. Introduction [TRP15, Ung04b, Vam09i, Vam09j, Vil02j, Wal92h, Was04, vDP17b, vL16g, SR15]. Introductory [BVH^{+}83c, Gou06a, Hel14, IG05b, Nic80c, Pos95j, Zim92, dB87e]. Intuition [Ano06-30, Bor99, Čap71i, Car15a, Gio81, Pul06, Sch80a, Tha86d]. Intuitionistic [Leb82, Ser81]. Intuitive [MC80f, Tos81]. Invalidities [Fin80c]. Invariance [Bro93, Cos87a, Fra96b, FK93, Pan91b, Pan91h, Pos93, Sch82, dB87o]. Invention [Bec91p, Cif05, Cor76b, Gut80, Hac86, NRS92, Ree14]. Inventions [Ara15]. Inventors [Die14, Hug88]. Inverse [Ass11]. Investigating [Llo15]. Investigation [Are89h, Fra09b, Ton81a]. Investigations [CS79u, PO91, Sha84l, Tha84b]. Involvement [Fin80a]. IPCC [Kow16a, Kow16c, Kow16h]. IQ [Dan76, Kay76].

Iron [FQ18]. Irony [MA87, Lan94]. Irrational [CS79t, Jar86a, Lag96, Mey91j, Sch80a]. Irrationalism [Aga03i, Sle99]. Irreducibility [HH94]. Irreducible [Pol76]. Irrelevancy [Ren96]. Irreversibility [Čap71u, CS79q, CS79-43, Wat98, dB87f, dB87g, dB87p]. Irritant [Cha17]. Irwin [Sch10]. Isaac [Coh74a, DS11b, Pal12, Pal13, Wes92]. Islamic [Ebr11, vE75]. Island [Bee89, Zim02, Pop88]. Isolation [Ges74o]. Isomers [Duh02j]. Israel [UM86a, UM86c, UM88]. Israeli [Han17, Mei17b]. Isserles [Lan91]. Issue [AP00b, Coh88a, Gre74p, Rat18, You69, You76]. Issues [Ano08-49, Ber74d, DN94, Eşa15, Gav14b, Gav14a, GC97, Gli87b, Hun17, Pet92b, Res79, Sar91, Syl97a]. Istanbul [Bre11]. Italian [Bri96, Chi81b, Gom15a, Gov15, LPV15]. Italy [FG92, Ree87b]. IUHPS [HFD88]. IV [Mor80, CLR^{+}92, Onn91d]. Ivan [Kap80s]. IX [Onn91j].

J [CSW74, Cro92a, KS95, Nol91, Whi99, Onn91k]. J. [CS79-41, Duh02x, Duh15c, Eas84, Gre94, Pul11, Scr74a, Smi92, Str94, Str76]. Jābir [Haq94b, Haq94f]. Jack [Car89b]. Jacob [Rey18, Hac17]. Jacques [CS79r, Sch01]. Jail [Lay92]. Jakob [Jah91]. Jakobson [Kar17]. James [Čap71v, Gav09, NMN02, Row14, Sal71, Vil02i]. Jammy [Rus09]. Japan [Has11]. Japanese [Ito11, Kan87, NC98, San92b, Sas92]. Jassās [She75]. Jean [Sil11]. Jena [Bre07, Jah94]. Jerome [Duh91h]. Jerusalem [BIT01e, Kat14]. Jesuit [EF18, FEE18, Gat18, Rai10, Sas03c, Van08]. Jevonian [Sch94a]. Jewish [Lev83d, Pop88]. Jews [Whi88]. Jheronimus [dV74]. Jiao [Hor93]. Jiaoji [Wan18a]. Joan [SPB18]. Jobs [Saw98]. Jochen [Amr87b]. Joe [Cha17]. Johan [Maa17]. Johann [Ber94]. Johannes [Bus74, Kra91]. John [Aga81-29, DJ89, Duh76c, Hon09b, Jaf02, Pal12, Pal13, Phi94, Rat88, RDS03, Scr74a, Syl91, Vil02i, Zah89]. Joint [HFD88]. Jonas [Ihd79j]. Jonathan [Sel74]. Jordanus [Duh91k, Duh91n, Duh91l, Duh91p]. Joseph [Bal89a, BAG17, Das15, Don17, Her14b, Her14d, JL95a, JL95b, Kra92a, Lip17, NNW74, Win17a]. Jost [Jan84b]. Joule [CS79b, CS79s]. Journal [DCS15, Kap80d]. Journalist [Gal15]. Journals [Kap80d]. Journey [Nen18]. Judaism [Lav88, Lev95]. Judeocide [May95]. Judgements [Bro96a]. Judging [ML99, Sim77o]. Judgment [But93l, Coh02a, Dra88, Fin80b, Fin95, Tve86, War80]. Judgments [Scr74b, SF94]. Julio [SR15]. Juncture [Ger15b, Vam09k]. Jury [Coh02o]. Just [Bun12q, Rös95, Sta07c]. Justice [Boy09, Ezr01]. Justification [Ams75b, BV94, Bur86, Chm82, Fet81d, Fra99d, Fre86q, Gin06, Gre99, Hal91f, Leh00, Per84, Ste09b, Tok09, Van83, VdOM11, Wri80]. Justified [PR17]. Justifying [Rad79]. Justus [Fin91, Heu96, Sch96d]. Juxtaposition [Čap71w].

Kairetic [Fis15]. Kaiser [Ano08o, Ano08-45, Ano08-40, Ano08-32, Sch08d]. Kaiserwilhelm [Hei08f]. Kaiserwilhelm-Institutes [Hei08f]. Kala [RH92]. Kaleidoscope [UM86a]. Kamerlingh [GG91, SR15, vD14]. Kant [Bur74b, Are89g, Buc89, Bur87d, But93c, But93e, But93h, But93l, Cap81a, CK97, Elk74, Fer85b, Fer85e, Fer85f, Fer85h, Fri07b, Gre94, Hyd06, Jac97, Lef01a, Mar94b, MC80g, Mey91g, Mun89, NA07, Pau00, Pla94a, Pla94b, Ste03, Van06, Wat01b, Web87, Wun01, Zel74, ZC88c, vW94]. Kantian [And99, Edg15, Fer85a, Fun86, Hes07, Mas11]. Kantianism [Che94, Rau91]. Kantians [Cro99]. Kaplan [IG05a]. Karl [Dei17, Grü76, Klo79, Now79a, Sch07a, Zem09]. Kassel [Mor80]. Kaufmann [CH14, Hel14, Kau14]. Kazimierz [Coh94d]. Keeping [BAK11, KT93]. Keeps [Beu91]. Keill [Pal12, Pal13]. Kelly [O’D02]. Kelvin [Sha88]. Kepler [Kra91, Lin91, Lug89, She82]. Kev [Fra99a]. Key [AV16b]. Keynes [Coh02u]. Keystone [Ren90]. Khrushchev [Ada01]. Kim [Mar79c]. Kind [GS82, Har89a]. Kinds [Aga95d, Dil06h, Bhu06, Buc94a, Car86, CD93, Dut11, Fel89, FC16, Mor71, Sch67, Sie82, Ste07, Tor82]. Kinematical [dB87r]. Kinematics [Van93]. Kinetic [Cha09o]. King [Sim94]. Kingdoms [NG01d]. Kinships [Grü89b]. Kitāb [Haq94a, Haq94b, Haq94e, Haq94f].

Klappholz [Fol17]. Klaus [Hei08c]. Klein [Bal03, Gil03, Irz03, KS95]. Knocking [Ion15]. Knot [JR07]. Know [Ágo11, Bar83, De 16, Fle86g, Jan07, Ren96, San97a, Vin92]. Know-How [Ágo11]. Knowing [Are89j, Cau95, Koh95, Kün95, Życ95b]. Knowledge [Aga75f, Aga81-32, Aga95d, All99a, Ams75e, Ams75h, Ams92c, Ano08-42, Bab99b, Bam99, Bar89, Ben15, Bon08, Bré10, Bré14, BE10a, Bun10h, Bun12k, Byr02, Čap69, Čap71e, Čap71g, Car11, Cha09a, Cof74, CS79-42, Coh95, Coh02j, Coo86, D’A92, Dar07, De 95, DPW86a, Dop01, Ezr86, Ezr14, Fel18, Fen12, Fer85b, Fer85a, Fer85g, Fet81e, Fet81h, Gie71, Gje07, Gre75, Grü95b, GR11b, Hab99, HK92, Har10, Hee79, Hin89, Hör91a, Ina05, Jan89m, Jar86x, Kah91, KC95, Kuç95, Kul01, Lek95, Lel00b, Lel00f, Lel00o, Lev83c, Lon12, Mar83a, MC80f, Men86, MA87, Nol03f, Nol03j, Nol03k, Ols05, OK00, Pan83e, Pan83f, Pan91c, Pan91i, Pan91j, Pen99, Pet03a, Pit89, PC69, Que95, RH14, Ren15, RDS^{+}18, Res79, Rud86, Sad89, San97a, San90, Sch99a, Sch99b, Sch97b, Sha84m]. Knowledge [Sha84l, Sha84t, ST69, Sne89, Swi99, Szu00, Tym95, Val10e, Vin92, Wat78b, Wel99, Wir95, Wit86, Wol95e, Wol95m, Zal86, ZC88b, Zin73g, Zin73j, Życ95a, dS11a, dF11a, GG91, HFD88]. Koch [AA86, Ros86]. Koen [SPB18]. Koestler [Fin17]. Kohák [Cah98, CT98]. Kohn [Woo06]. Konrad [Kal87]. Kosok [Sol84]. Kostas [ARS15]. Krebs [NG09]. Kreis [Wol97]. Kripke [Mar79u]. Kronecker [Mar95b]. Kuhn [Bir15, Bur84, Dev15, DB15b, Far03i, Gol12, Hat89, HH15, Mar15, Mar95a, Mar97b, Mas15b, Pol95, Ric15, San97b, Sch99d, Sha15, Shi76, Tor99, Tso15, Wra15]. Kuhns [Sch99d]. Kulish [Luc88]. Kulturnation“ [Hül96]. Kurt [Chr11, Ges74s, Hei11, NG09, Pad11, Pec11]. Kuwaki [Tan92].

L [Fet81f, SC74b, MW76, Onn91l]. L. [Coh02j, Rub88]. Labor [Dam96d, Hei17]. Laboratorium [dV11]. Laboratory [Fra96a, Fra99k, NG01e, Van15, vD14, Kle01b, Onn91o, vD15]. Labour [Lim94]. Labyrinth [Van74c]. Lack [Pie12]. Ladening [Wyl92]. Ladenness [Rap86]. Lady [And91]. Lag [Aga81e, Pat76a]. Lagrange [Bou02f, Bou02i]. Lai [Hor93]. Laid [Wal84]. Laird [Hem76]. Lakatos [Ara94, Avg89, Bar76b, Ber76c, CFW76, Dai89, Far03d, GGN89, Gra95, Gre76b, Koc79, Kuh71, Wor76]. Lakatosian [Aga76c, Met89]. Lamarck [Bar94, Sar91]. Lambert [Led18]. Land [Elz95]. Landau [Kap80t]. Landen [Phi94]. Landscapes [SDG15b]. Lane [FG92]. Lange [Edg15]. Langevin [Kap80-30]. Language [Ant87, Bar74, Bie76, Bur87d, Cho63, Cho76, CW83b, Coh02j, Coh02p, Coh02r, Den99j, Ege99, Fek99, Fet81f, Ges74h, Ges74p, Ges74r, Ges74w, Gra83c, Hal95a, Kar17, Kas76b, KK07, Laz77, Lel00d, Lel00i, Len63, Lev90, Mac12d, Mar78e, Mar78d, Mar86f, Mar86g, Mar79l, Mar79f, Mar79o, MA83, MA87, Mon96, OK00, Pau17, Pia82, Put76, Sch89, Sha84a, Sha84k, Sha84q, ST69, Ste00, Sto00, Str94, Tha84b, Tha84d, Ton73b, Ton81f, Ung15, Van67, ZC88h]. Language-Induced [Ges74p]. Languages [Gra83g, Hin76, Lam95, Mar63, Mar78e, Mor83]. Lanka [DP92]. Lansing [SC74a]. Laplace [Čap71h]. Laplacian [Čap91d]. Large [Bro91b, Sha86a, Tur86f]. Large-Scale [Tur86f]. Larmor [Pos95k]. Larvae [Raf15, Str15]. Last [McC95b, Mur12e]. Late [Car15b, Mur75, BD05, Bré10, BE10a, BE14, Glu91, Hel08, Mar12, Mar13, McG87, Pau17, Rai10, Saa17]. Later [Bla95g, Gie12, Jan89e, Jan89k, Mül94, Bla95i]. Latin [Fre00, Hül96, Wal92e]. Lattice [Hee79]. Lavoisier [BR15, Hoo99l, Lev90]. Law [Add76, Ano07r, Bel12j, Ber74a, Buc67, Bun12q, Caw89, Cel08, Dil06i, Duh02k, Duh02t, Ela11n, Fre74a, Gem90, Han99a, Han99f, Han99h, Han99g, Han99i, Hel00e, Hel00i, Hem76, KT76, Kra74, Lei04, Mit76b, Nee11f, Rec11, Res97, Suc74, TG88, ZRKC03c, Bor07, CS79-28]. Lawlike [dB87g]. Lawlikeness [Ano09r, Fet76]. Laws [Ano06j, Ano09k, Ano09-44, Bra95, Bun93g, Car01, CS79p, Cop76, Dil96, Dil06a, Dil06f, Duh91q, Fra99h, Hon93, Len87, MC80i, MC80h, SM96, Szy82, Tem82, Wes76, Yos98, ZRKC03i]. Lay [CS79q, Feh99b]. Layers [Che10]. Lazarus [Kóh90]. Lead [ZC96, Onn91h]. Leadership [Kap80w]. Leads [Mar97b, Pie12]. Learn [CLR^{+}92, Mic18, Pet76, Sha84t, Tis03, Tur96]. Learnability [Dem89]. Learnable [Gro82h]. Learned [Ant16, Cha12d, Cho14, MP16]. Learning [Aga75n, BM09, BE14, Bro89, Cha14, Cla14, Dav10, Gro82d, Gro82i, Gro82k, Gro82m, Kow16h, Mar79f, MS75b, Mur75, SCD03a, SP10, SA12, Swa09, Wel15, Wer01, SCD03b]. Least [Bou02d, Wel15]. Leblanc [MC95, MC96]. Lechalas [MC80i]. Lectures [Cor95, CCG93, GCC81b, Lai08, Hoo99c]. Left [Ges74c, Ges74m, Hor88]. Left-Hander [Ges74c]. Left-Right [Ges74m]. Left-Wing [Hor88]. Legacy [BCK07, FB17, FS94, Her14b, Her14d, Kai94, Lau09, Van06, Zol89g]. Legal [She75]. Legend [FL09]. Legends [Wet17]. Legitimate [Rod15]. Legitimation [Aga81m, Bod18, VK83, Wri80]. legume [BBT95, HD95, PLH95]. legumes [Tho95, WRL95]. Leibniz [Bec91m, Ana90b, Ano08-35, Ano08-36, Art89, Bou02g, But93f, Čap71x, Duc89, Fre86l, Fre86m, Fre86p, Fut08, Hec01, Kno94, Kno05, Lef01a, MC80j, Mug81, Pau00, San81, Sle99, Som76a, Sot03, Wat01b]. Leibnizian [Rey18]. Leiden [Onn91o, vD15]. Leipzig [Bla95d]. Lemniscate [Bos74]. Lens [Pis15, Tua16]. Lensing [EFN03]. Léon [CS79a]. Leonardo [Cai14, Dib74, Duh91i, Duh91o, Ret74, See74, Sha74c, Sti74]. Lesson [Bai17, Kam93, Krü90, Pâs15, Rös91]. Lessons [Cho14, Lem82, Mei17a, Now01, Ray00, Sto97]. Let [Hac15a]. Letter [Set82, Von91, Wol76i]. Letting [Wei07]. Leucippus [Bai76b]. Lev [Kap80t]. Level [Ada11, GP79, Möl01, Vin92]. Levels [Bun10f, Das76, Gre74c, Gro76]. Lever [RM18]. Levi [Gol91, MS83]. Levialdi [Fie14, Wai14]. Levinas [Got94]. Levine [Aga87a, Shi87b]. Lewin [Hei11, Pad11]. Lewis [HOO88b, HOO88c, Škv88]. Leyva [dF14a]. Li [Hor93, Yus96b]. Liber [Cai14]. Liberal [Ber84, Von91]. Liberalism [Kra92a]. Libraries [Min10]. Lichtenberg [Jan89a]. Liebig [Fin91]. Liebmann [Edg15]. Lieu [FM09a]. Life [Ano05l, Bab99d, Bit96g, Bla95g, Blü70, Coh94m, Cre02, D’A01e, Dou87, Ele94, Fin84, Fin14, FG92, GP03, Gen91, GWT00g, Har84b, Hol97, Hon09b, Ito11, Jan89m, Kam99, Kel09, Mül94, NMN02, Pol76, Put86c, Qui83, Rhe11, Rus09, Sch86c, Sch96h, Sha88, Ste15a, Šuš79, Tho96, Tym95, VD92, WW90, Wer01, Wig74, Win83b, Zam92m, Zem86, Bla95i, Bré10]. Life-History [GWT00g]. Life-Long [D’A01e]. Light [Bab97, CS79-47, D’A81, D’A01k, Duh15d, Góm12, Góm13, Heg87, Irz15, Lin91, Mac17, Pop86b, Sai04a, Sch96a, Syn82]. Lightness [FEE18]. Lightning [Kap80y]. Like [Aga91, AT15, Bou02b]. Likeness [Fet76]. Limit [Čap71w, Pan83b, Str74]. Limitation [Tho10]. Limitations [Čap71z, Čap71y, Čap71-36, Han17, Lev71, Pol95, Tem95, BLA95a]. Limits [Are89h, BM88, Bon08, Čap71-27, Das88, De 01, DD86, Hem88, Hod79, Jar86h, Jar86m, Mac17, Net05, Nor98a, Pis15, Pit89, Res79, Sha84p, Ste88, TG88, Vot11, Zim83]. Lindborg [CS79g]. Linde [Die14]. Lindemann [Fox15]. Line [KJ14, Mas15b, UM86b]. Lineale [Beu96]. Linear [Bus97b]. Linearity [Aga03d, Ela11l, Kos84]. Lines [Uga18, Fre00]. Linguistic [Fre99, Has97, Mar84i, Zol89d]. Linguistics [Chr94b, Lie76, Put67a, Put86a, Ras74]. linguistique [Ras74]. Linguosynthesis [Lel00g]. Link [Aga13j, Cek92b]. Linkage [Grü89b, Wim92]. Links [Ber13c]. Linus [Kar10a, Kar10g, Nye01]. Liquefaction [Kap80v, Pap15, Row14, Onn91q, Onn91p]. Liquid [Kap80-34, Onn91e, Onn91g, Onn91j, DO91, Onn91m, Onn91f, OB91a, Onn91b, OB91b, Onn91c, Onn91d, Onn91i, Onn91h, Onn91k, Onn91l, OP91b, PO91, TO91]. Liquids [Gav90, Knu07, Onn91a]. Lisbon [Car15b, DCS15, Mot15]. List [HOO88b]. Listening [Ges74j]. Lists [Pro10]. Literal [Wol76-40]. Literary [Cla10, Sla89, Zam92j]. Literature [Amr89, Fil05, Gir92, Jan89h, Orv89]. Little [Mil95]. Liu [Dun96, Ron96]. Live [Aga81-31]. Lives [Nye07]. Living [Čap91r, Gre74f, Gre74g, Gre74h, Hah91, Hon09b, Lek15a, MV80a, MV80j]. Local [Dic96, Fel15, Fu93a, Gol03, Kri94, Squ96, Wal92f]. Locality [Bit90, Fay94, SS03]. Locating [RH92]. Location [Čap71f, Čap71-27, Čap91m, Har86, SG11]. Locke [Bec91n, Wol18]. Lockean [Den99h]. Logarithmic [Res97]. Logic [Aga81n, Aga89, AD94, Ano06z, Ano06-28, Bec91e, Bec91i, Bec91q, Bel95, Bre74b, Bun12d, Bun12f, But93n, Čap71-28, Chi81a, CW83b, Coh94e, Coh94m, Coh02e, Coh02i, Con81, Cur80, Dam96c, Dou03, Dum15, Fin69, Fin80e, Fin80h, Fin80k, Fin80n, Fin80p, Fin17, Fra91, Fra96b, Gau84, Gau95, Gio81, Gol83, GG96b, Guc81, Gut80, Hal95b, Han99c, Hee74, HR74b, HR76, Hin95, Ism17, Köh91b, LMR95, Lau80a, Leb82, Leb95, Lep95, Lin93a, LPV15, Man81, MC95, Mar74b, Mar79f, Mic97, Mir15, Mit76f, Mit76h, Mur12f, Nar09, Nic80f, Nii03, Øhr97, OK00, Pec96, Pis15, Put69, Put86b, Sab84, San15, Sch80a, She75, Sim77i, Sim77k, Sim77l, Sol84, Som76a, Sta76a, Tab03, Tam86, Thi11, Van67, Van95, Wal92j, WZ76, Wat63, WØ97, You76]. Logic [Zin73e, Zin73g, Zin73i, Zin73p, Zin73w, Wal92c]. Logicae [Syl91]. Logical [Bar84, BG18, Bör93, Cat93, Chi83, CW74a, Coh76c, CS79t, Coh02r, Fin80f, Fin80l, Gau84, Grü05b, Has97, How76, JX96, Jim95, Lin93d, Mar84f, Mar79d, MP13, Pan84, Sie71, Som74, Som76b, Sos74, Sot03, Sta83a, Tis63, Toe81, Ton81j, Tou71, Wal92d, Wal92e, Zin73f, Zin73g, Zin73h, Zin73j, Zin83b, Rös95]. Logicality [McM74b]. Logician [Fin80i]. Logicism [Seg95]. Logico [ZM11, Rad03]. Logico-Philosophical [ZM11]. Logico-Philosophicus [Rad03]. Logics [AM63, Gab76, Mar63, Pog16, Pro17a, Van74c]. Logistic [Fow04]. Logos [War79e]. Loma [Wol76-33, Wol95a]. Lomonosov [Kap80u]. London [CS79n, CS79-41, Gou90, Hei91, Lek15b, Pop88]. Lonely [Kar17]. Long [D’A01e, Grü95b, Hon09a, Mar90, NG01g]. Long-Standing [NG01g]. Look [Ano09-46, Bla92a, Fle86g, Ohm98, Wal95, dB87k]. Looked [Rad89a]. Looking [BE10b, Ger15a, GF97, Har95, Nic89b]. Looking-Glass [Har95]. Looks [Cao95]. Lora [NG15b]. Lora-Tamayo [NG15b]. Lord [Kap80-37, Sha88]. Lorentz [Cos87a, Gia71, Nor07d]. Lorentz-and-CPT [Cos87a]. Lorenz [Kal87]. Loss [Sch98, Wei01a]. Lost [Sch10, Too04, dB87b]. Lotka [KL16]. Louis [Bal89a, Čap71b, Čap91a, Pap15]. Love [Ant16, Ber86, MP16]. Lovejoy [Čap71p]. Lover [All99a]. Low [DG15, Pap15, SR15, BBO91, Onn91o, Onn91c, Onn91d, Onn91e, Onn91j, Onn91n, OP91a, WO91]. Low-Dimensional [DG15]. Low-Temperature [Pap15, SR15]. Lower [GWT00h, BE10a]. lowest [Onn91m]. lowland [BLA95a]. Löwy [Tau91d]. Loyalty [Fra99f]. Ltd. [Lan94]. Lucretius [Bai76a]. Ludwig [Bla95g, Bla95i]. Ludwik [CS86, Gie86, Löw86, Sch86b, Sch86c, Tou86b, Wit86, Wol86]. Lukács [Wes90a]. l’Université [Pet92a]. Lwów [Coh94d, Coh94f, Mar86e, Sch86b, Sch86c]. LWS [Coh94d, Coh94g].

M [CS79n, Lip17]. Mīmāmsā [Ano06i]. Mach [Are89h, Baa92, Bar07, Ber70, Bla92a, Bla92b, BIT01c, Blü70, Bun92, Čap69, Čap71t, Cek92a, Cek92b, Coh70, CS70, Goe70, Hal82, Hal92, Hie70, Hof92, Hol70, Hol92, Ima01, Ita92, Kre91, Men70, Mer70, Mey92, Noé92, Ost92, Pla92, Rat70, Ren07d, San92b, See70, Sta92a, Tan92, Vis92, Wie92]. Machida [Yan98]. Machine [Erw71, Gün11, Hee86b, Ihd79c, Sne89, Ste17, Web11]. Machinery [Haa99, Win17b]. Machines [BV11, Dre86, Mac86a, Man12, Man13, MV80j, RA14, Sha93, Val10d]. Machist [BIT01b, Vis01]. Macro [Gin97, Sob10]. Macro-Hermeneutics [Gin97]. Macro-Probabilities [Sob10]. Macroevolution [Hun17]. Macromolecular [Wie12]. Macroscopic [Čap71j, GP79]. Macrosociology [Lec90]. Maddy [Imo15]. Madison [CS79n]. Madness [Mus76]. Madrid [Bro08]. Magic [Müü01, dB87h]. magnet [Onn91k]. Magnetic [Kap80-35, PO91, Ste09a, Onn91j, TO91, Mir14]. Magnetism [Kap80x, OP91a]. magnetization [OP91b, WO91]. Magnetochemistry [dAM07]. Magnitude [Cos87b, Pas76]. Magnitudes [Ber83b]. Magnus [Wal81h, Zam92h]. Maier [Wal81a]. Maimonidean [Goo00]. Maimonides [Aga00a, Alt00, CL00, Fre00, Ide00, Kra00, Lan00, Ras00, Ste00, Twe00]. Main [Ela11o, Kra95]. Mainstream [NG15a]. Major [Aga13f]. Majority [Mir15]. Make [Bar15, Kil03, Kow16j, RH14, Rop86, Ste09a, Stö99, Too10, ZC96].

Make-Believe [Too10]. Maker [Hei94b]. Makes [Aga81-36, Arm95, How94, Mah11]. Making [Boy09, Fre09, Gje00, Hem84, Hes07, Mai12, McG96, Mey85p, RA14, Sch11, Sim77k, Str18, Sun18, Tas17]. Man [Ano06-31, Ano08-40, Ges74j, Ges74k, Gra83b, KT76, Kos76a, Pec11, Sam92, Slo83, You95, Zim02]. Management [Rav18, SVD95, Vra79, PLH95]. Managing [Sup79]. Manchu [Jam10]. Mandate [Pos95f]. Mandatory [Mei17b]. Mandonnet [Zam92k, Zam92l]. Mang [Tho10]. Manifestation [JX96]. Manifesto [Gel16]. Manifold [LC07]. Manifolds [Kli96]. Manipulation [BBT95, Cap15, Psi07]. Mannheim [Nol03l]. Manuel [NG15b]. Manufacture [Gro09c]. manure [BLA95a]. Manuscripts [Ano08-34, Har10, Kno05, Rai10, Wal90]. Many [Gou89, Hil96, Pop84]. Many-Particle [Gou89]. Mao [Sch88]. Mapmaking [Bre11]. Mapping [BMP09b, HSS09b, Wim92]. Maps [Gro82n]. Marcelin [Duh02w]. Marci [CS79u]. Marconi [Ara15]. Marcus [CS79u, Mar79j]. Margalit [Lev83a]. Marginalization [Tho10]. Maria [IG05a]. Marine [Raf15]. Mario [AC82, Set82, Wet82]. Marjorie [DPW86a, DPW86b, DPW86c]. Mark [Fay97b]. Market [Bro11, Sch94a]. Marketplace [Bid11]. Marr [Ber96b]. Marriage [Mcm76b, Ric15]. Mars [Lug89]. Martin [KS95]. Marx [GC94, Gou94b, Han99e, Jan89o, JN79, Kan79, Klo79, Lef03, Nol03l, Now79a, Pan94, Sch88, Zel74]. Marxian [Lei83, Mar86h, McB94]. Marxism [All01, GSW95a, Gou94b, Kur88, Meg74, Miś82a, Rui95a, Szt79a, Tha84c, Zol89b]. Marxist [Bun12l, FM09b, Lau74, Wia79a]. Masked [Aga95a]. Mass [Bel12b, Bel12e, Lag97j, Nor07a]. Masses [Duh02h]. Master [Giu04, Pro14]. Masters [Hel08]. Mastery [Pet76]. Material [Hun10, Kni07, Rad12d, Rat18, Rog71, Wim12a]. Materialism [Bun10i, Bun12g, Bun12s, Coh86d, Coh86a, Dam96d, Dvo91, Hei94a, Hou93, KP86, Mar78f, Mar78i, Mar78m, Mar86h, Max91, Pet69b, Sha90, Sto79, War79s, Wol18]. Materialist [Ano09o, Dav74a, Qui82, War79d]. Materialistic [Tho01]. Materiality [Sch96a]. Materials [BV11, CS86]. math [Sza04]. Mathematic [Mar95b, Sas03l]. Mathematical [Abr81, Ano08-37, Bré14, Bro10, Cap81a, Cas95c, Cas95a, Cel08, Cif14, Dam96k, Das91, Dis06, Fin09i, Fol18, Fre00, Grü73t, Hoo99h, Imo15, Ion15, Jie96, Kol74, Mar79e, Men70, Mis74, Pat92, Pin11, Ras00, Sab84, Sal18, Sas03c, Sas03d, Sas03h, Sas03i, Sha84e, Sør17, Syl91, Tob96, Too04, Tor82, Van08, Vol14, Xiu96, dS11a]. Mathematically [Čap71-41]. Mathematician [Ano08-46, MS17, Omo18a, Sch96f]. Mathematicized [Hon96a]. Mathematics [Alt14b, Ano07h, Aub17, BS16, Bré10, Čap71-32, Che10, CK97, Chr94a, Chr04a, Cif14, Cle02, CS02, CCG93, D’A01b, Dam96f, Dam96i, Dam96k, Dav74a, DD11, Ech96, Fil05, Fit76, Fow04, Fre04, Gai86, Gar95a, Gau95, GSW95c, Gio81, Gon96, Goo95, Gua96a, HJLP06, Hen01, Her74, Hor93, Jam92, Kar90, Kel10, Kno94, Kno04d, Kno04c, Kno16, Kör86, Krö12, Kuu01, Lam82, LC17a, LC17b, LL92, Llo04, LPV15, Lub82b, Mal12, Mal13, Mar76a, Mar80, MC95, Mey85r, Mol18, Net05, NRS92, Ott94, OP97, Per18, Per94, PP18, Pit92, Pra93, Ras94c, Ras95, RSS^{+}07b, Ric91, Ron96, SV15, Sas03e, Sch96e, Sch96h, Sha16, SS91, Sim11b, Sør17, Str95, Syl08, Sza84, Ung94, Ung04a, Ung04c, Wor76, Wuß91, Yav17, dB12, dB13, Bro08, Gar95b]. Mathematik [Sza04]. Mathématique [Pat92]. Mathématiques [Fre04]. Mathematization [Bla94, Bla04, But93k, KL17]. Mathesis [Sas03k, Sas03j]. Matric [Mul74]. Matrix [Bun12o, Cos87a]. Matter [Ano63, Ano67, Ano69a, Ano69b, Ano69c, Ano70a, Ano70b, Ano71a, Ano71b, Ano71c, Ano71d, Ano71e, Ano71f, Ano71k, Ano71g, Ano71h, Ano71i, Ano71j, Ano73a, Ano73c, Ano73d, Ano73b, Ano73e, Ano73g, Ano73h, Ano73i, Ano73j, Ano73k, Ano73f, Ano74c, Ano74b, Ano74d, Ano74e, Ano74a, Ano74u, Ano74-33, Ano74-34, Ano74-35, Ano74-36, Ano74n, Ano74p, Ano74q, Ano74r, Ano74s, Ano74t, Ano74f, Ano74g, Ano74h, Ano74i, Ano74j, Ano74k, Ano74l, Ano74m, Ano74o, Ano74v, Ano74w, Ano74x, Ano74y, Ano74z, Ano74-27, Ano74-28, Ano74-29, Ano74-30, Ano74-31, Ano74-32, Ano75a, Ano75c, Ano75b, Ano75d, Ano75e, Ano75f, Ano75g, Ano75i, Ano75h, Ano76c, Ano76e, Ano76g, Ano76f, Ano76b, Ano76d, Ano76a, Ano76s, Ano76p, Ano76q, Ano76r, Ano76-33, Ano76-34]. Matter [Ano76-35, Ano76-36, Ano76-37, Ano76-38, Ano76-39, Ano76-40, Ano76h, Ano76t, Ano76u, Ano76v, Ano76w, Ano76x, Ano76y, Ano76z, Ano76-27, Ano76-28, Ano76-29, Ano76-30, Ano76-31, Ano76-32, Ano76-41, Ano76-45, Ano76-42, Ano76-43, Ano76-44, Ano76i, Ano76j, Ano76k, Ano76l, Ano76m, Ano76n, Ano76o, Ano77a, Ano77b, Ano77c, Ano77d, Ano77e, Ano77f, Ano77g, Ano77h, Ano77i, Ano77j, Ano77k, Ano77l, Ano77m, Ano77n, Ano77o, Ano78a, Ano78b, Ano78c, Ano78d, Ano78e, Ano78f, Ano78g, Ano78h, Ano78i, Ano78j, Ano78k, Ano78l, Ano78m, Ano78n, Ano79b, Ano79c, Ano79e, Ano79d, Ano79f, Ano79a, Ano79g, Ano79h, Ano79i, Ano79n, Ano79o, Ano79p, Ano79q, Ano79r, Ano79s, Ano79u, Ano79t, Ano79v, Ano79w, Ano79x, Ano79y, Ano79j, Ano79k, Ano79l, Ano79m, Ano80a]. Matter [Ano80c, Ano80d, Ano80e, Ano80b, Ano80f, Ano80g, Ano80h, Ano80i, Ano80j, Ano80n, Ano80o, Ano80p, Ano80q, Ano80r, Ano80s, Ano80t, Ano80u, Ano80v, Ano80w, Ano80x, Ano80y, Ano80k, Ano80l, Ano80m, Ano80z, Ano80-27, Ano80-28, Ano80-29, Ano80-30, Ano81f, Ano81c, Ano81d, Ano81g, Ano81b, Ano81a, Ano81e, Ano81v, Ano81o, Ano81p, Ano81q, Ano81r, Ano81s, Ano81t, Ano81w, Ano81x, Ano81y, Ano81z, Ano81k, Ano81l, Ano81m, Ano81n, Ano81h, Ano81i, Ano81j, Ano81u, Ano82b, Ano82c, Ano82a, Ano82e, Ano82f, Ano82g, Ano82h, Ano82i, Ano82j, Ano82k, Ano82l, Ano82m, Ano82d, Ano83e, Ano83f, Ano83g, Ano83h, Ano83a, Ano83c, Ano83d, Ano83b, Ano83m, Ano83n, Ano83o, Ano83p, Ano83q, Ano83r, Ano83s, Ano83t, Ano83u, Ano83v, Ano83w, Ano83x]. Matter [Ano83y, Ano83z, Ano83i, Ano83k, Ano83l, Ano83j, Ano84b, Ano84a, Ano84d, Ano84e, Ano84g, Ano84f, Ano84c, Ano84l, Ano84m, Ano84n, Ano84o, Ano84p, Ano84h, Ano84i, Ano84j, Ano84k, Ano84u, Ano84v, Ano84w, Ano84x, Ano84y, Ano84-27, Ano84z, Ano84q, Ano84r, Ano84s, Ano84t, Ano85a, Ano85b, Ano85c, Ano85d, Ano85e, Ano85f, Ano85g, Ano86b, Ano86c, Ano86f, Ano86g, Ano86k, Ano86a, Ano86d, Ano86e, Ano86i, Ano86j, Ano86h, Ano86o, Ano86y, Ano86z, Ano86-27, Ano86-28, Ano86-29, Ano86-30, Ano86-46, Ano86-47, Ano86-48, Ano86-49, Ano86-42, Ano86-43, Ano86-44, Ano86-45, Ano86p, Ano86q, Ano86r, Ano86s, Ano86t, Ano86u, Ano86v, Ano86w, Ano86-31, Ano86-32, Ano86-33, Ano86-34, Ano86-35, Ano86-36, Ano86-37, Ano86-38, Ano86-39, Ano86-40, Ano86-41, Ano86-50, Ano86-51]. Matter [Ano86l, Ano86m, Ano86n, Ano86x, Ano87e, Ano87b, Ano87c, Ano87d, Ano87f, Ano87g, Ano87h, Ano87a, Ano87z, Ano87-27, Ano87-28, Ano87-29, Ano87m, Ano87n, Ano87o, Ano87p, Ano87q, Ano87r, Ano87s, Ano87t, Ano87u, Ano87v, Ano87w, Ano87x, Ano87y, Ano87-30, Ano87-31, Ano87-32, Ano87i, Ano87j, Ano87k, Ano87l, Ano88a, Ano88b, Ano88c, Ano88d, Ano88e, Ano88f, Ano89e, Ano89f, Ano89g, Ano89h, Ano89i, Ano89d, Ano89a, Ano89b, Ano89c, Ano89w, Ano89x, Ano89y, Ano89z, Ano89-27, Ano89-28, Ano89-29, Ano89s, Ano89t, Ano89u, Ano89v, Ano89j, Ano89k, Ano89l, Ano89m, Ano89n, Ano89o, Ano89p, Ano89q, Ano89r, Ano90a, Ano90d, Ano90b, Ano90c, Ano90e, Ano90l, Ano90f, Ano90g, Ano90h, Ano90i, Ano90j, Ano90k, Ano91b, Ano91c, Ano91k, Ano91h, Ano91j]. Matter [Ano91i, Ano91e, Ano91d, Ano91g, Ano91f, Ano91a, Ano91m, Ano91-28, Ano91-29, Ano91l, Ano91n, Ano91o, Ano91p, Ano91q, Ano91-40, Ano91-31, Ano91-32, Ano91-33, Ano91-35, Ano91-36, Ano91-37, Ano91-38, Ano91-39, Ano91-34, Ano91x, Ano91r, Ano91s, Ano91t, Ano91u, Ano91v, Ano91w, Ano91-30, Ano91y, Ano91z, Ano91-27, Ano91-42, Ano92k, Ano91-41, Ano92c, Ano92j, Ano92i, Ano92g, Ano92f, Ano92h, Ano92e, Ano92d, Ano92a, Ano92b, Ano91-43, Ano92l, Ano92m, Ano92n, Ano92o, Ano92p, Ano92q, Ano92r, Ano92s, Ano92t, Ano92-37, Ano92-38, Ano92-39, Ano92-40, Ano91-44, Ano91-45, Ano91-46, Ano91-47, Ano91-48, Ano92u, Ano92v, Ano92w, Ano92x, Ano92y, Ano92-34, Ano92-35, Ano92-36, Ano92-33, Ano92-31, Ano92-30, Ano92-32, Ano92-29, Ano92z, Ano92-27, Ano92-28, Ano93d, Ano93a, Ano93b, Ano93e]. Matter [Ano93c, Ano93m, Ano93f, Ano93g, Ano93h, Ano93i, Ano93j, Ano93k, Ano93n, Ano93o, Ano93p, Ano93q, Ano93l, Ano94g, Ano94h, Ano94i, Ano94e, Ano94k, Ano94j, Ano94d, Ano94c, Ano94f, Ano94a, Ano94b, Ano94r, Ano94s, Ano94t, Ano94p, Ano94x, Ano94u, Ano94w, Ano94v, Ano94o, Ano94n, Ano94q, Ano94l, Ano94m, Ano95k, Ano95l, Ano95g, Ano95h, Ano95a, Ano95e, Ano95b, Ano95m, Ano95f, Ano95c, Ano95i, Ano95d, Ano95j, Ano95-32, Ano95-33, Ano95y, Ano95z, Ano95n, Ano95t, Ano95o, Ano95p, Ano95q, Ano95-34, Ano95-35, Ano95-36, Ano95-37, Ano95-38, Ano95u, Ano95v, Ano95w, Ano95x, Ano95r, Ano95-27, Ano95s, Ano95-28, Ano95-29, Ano95-30, Ano95-31, Ano96c, Ano96f, Ano96h, Ano96b, Ano96a, Ano96e, Ano96d, Ano96g, Ano96k, Ano96n, Ano96t, Ano96j, Ano96i]. Matter [Ano96m, Ano96l, Ano96o, Ano96p, Ano96q, Ano96r, Ano96s, Ano97b, Ano97d, Ano97f, Ano97a, Ano97c, Ano97e, Ano97g, Ano97i, Ano97n, Ano97p, Ano97h, Ano97j, Ano97k, Ano97l, Ano97m, Ano97o, Ano97q, Ano98a, Ano98c, Ano98b, Ano98d, Ano98h, Ano98e, Ano98f, Ano98g, Ano99b, Ano99a, Ano99f, Ano99e, Ano99d, Ano99c, Ano99g, Ano99i, Ano99h, Ano99m, Ano99l, Ano99k, Ano99j, Ano99n, Ano99o, Ano99p, Ano99q, Ano99r, Ano00e, Ano00b, Ano00a, Ano00d, Ano00f, Ano00c, Ano00k, Ano00l, Ano00m, Ano00n, Ano00o, Ano00p, Ano00h, Ano00g, Ano00j, Ano00q, Ano00i, Ano01g, Ano01e, Ano01a, Ano01b, Ano01d, Ano01f, Ano01h, Ano01c, Ano01i, Ano01-28, Ano01q, Ano01r, Ano01s, Ano01t, Ano01u, Ano01j, Ano01k, Ano01l, Ano01m, Ano01n, Ano01p]. Matter [Ano01v, Ano01w, Ano01x, Ano01y, Ano01z, Ano01-27, Ano01-29, Ano01o, Ano01-30, Ano01-31, Ano01-32, Ano01-33, Ano01-34, Ano02d, Ano02c, Ano02e, Ano02f, Ano02a, Ano02b, Ano02s, Ano02o, Ano02p, Ano02q, Ano02r, Ano02t, Ano02u, Ano02g, Ano02h, Ano02i, Ano02j, Ano02k, Ano02l, Ano02m, Ano02n, Ano03h, Ano03e, Ano03b, Ano03a, Ano03g, Ano03i, Ano03f, Ano03c, Ano03d, Ano03-37, Ano03t, Ano03u, Ano03v, Ano03w, Ano03x, Ano03y, Ano03z, Ano03-27, Ano03-28, Ano03-29, Ano03m, Ano03n, Ano03o, Ano03p, Ano03q, Ano03j, Ano03k, Ano03l, Ano03-31, Ano03-32, Ano03-33, Ano03-34, Ano03-35, Ano03-36, Ano03-38, Ano03-39, Ano03-40, Ano03-30, Ano03r, Ano03s, Ano04a, Ano04b, Ano04c, Ano04d, Ano04e, Ano04f, Ano04g, Ano04h, Ano04i, Ano04j, Ano05a, Ano05b, Ano05h]. Matter [Ano05i, Ano05j, Ano06e, Ano06f, Ano06g, Ano06h, Ano06d, Ano06m, Ano06n, Ano06o, Ano06p, Ano06q, Ano06r, Ano06s, Ano06t, Ano06u, Ano06l, Ano06v, Ano06w, Ano06k, Ano07b, Ano07c, Ano07f, Ano07j, Ano07i, Ano07k, Ano07l, Ano07m, Ano08h, Ano08a, Ano08b, Ano08c, Ano08e, Ano08f, Ano08g, Ano08d, Ano08-28, Ano08r, Ano08s, Ano08t, Ano08u, Ano08v, Ano08w, Ano08y, Ano08z, Ano08-27, Ano08x, Ano09a, Ano09d, Ano09e, Ano09f, Ano09g, Ano09c, Ano09b, Ano09h, Ano09i, Ano09s, Ano09v, Ano09w, Ano09x, Ano09y, Ano09z, Ano09-27, Ano09-28, Ano09-29, Ano09-30, Ano09-31, Ano09-32, Ano09-33, Ano09-34, Ano09-35, Ano09-36, Ano09-37, Ano09u, Ano09t, Ano09-38, Ano09-39, Ano09-40, Ano09-41, Ano10d, Ano10e, Ano10b, Ano10a, Ano10c, Ano10f, Ano10o, Ano10p, Ano10q]. Matter [Ano10r, Ano10s, Ano10t, Ano10h, Ano10i, Ano10j, Ano10g, Ano10k, Ano10l, Ano10m, Ano10n, Ano10u, Ano10v, Ano10w, Ano10x, Ano11b, Ano11h, Ano11e, Ano11c, Ano11d, Ano11g, Ano11f, Ano11p, Ano11q, Ano11r, Ano11s, Ano11t, Ano11u, Ano11v, Ano11-40, Ano11-33, Ano11w, Ano11x, Ano11y, Ano11z, Ano11-27, Ano11-28, Ano11-29, Ano11-30, Ano11-31, Ano11-32, Ano11-35, Ano11-36, Ano11-37, Ano11-38, Ano11-39, Ano11-34, Ano12a, Ano12b, Ano12c, Ano12d, Ano12e, Ano12f, Ano12g, Ano12h, Ano12i, Ano12j, Ano12k, Ano12-37, Ano12l, Ano12m, Ano12n, Ano12o, Ano12p, Ano12q, Ano12r, Ano12s, Ano12t, Ano12u, Ano12v, Ano12w, Ano12x, Ano12y, Ano12z, Ano12-27, Ano12-28, Ano12-29, Ano12-30, Ano12-31, Ano12-32, Ano12-33, Ano12-34, Ano12-35, Ano12-36, Ano11a, Ano13a, Ano13b, Ano13c]. Matter [Ano11i, Ano11j, Ano11k, Ano11l, Ano11m, Ano11n, Ano11o, Ano13d, Ano13e, Ano13f, Ano13g, Ano13h, Ano13i, Ano13j, Ano13k, Ano14d, Ano14b, Ano14a, Ano14c, Ano14u, Ano14i, Ano14j, Ano14k, Ano14l, Ano14m, Ano14n, Ano14e, Ano14f, Ano14g, Ano14h, Ano14o, Ano14p, Ano14q, Ano14r, Ano14s, Ano14t, Ano14v, Ano14w, Ano14x, Ano14y, Ano14z, Ano14-27, Ano15a, Ano15d, Ano15e, Ano15f, Ano15b, Ano15c, Ano15g, Ano15-45, Ano15-46, Ano15-47, Ano15h, Ano15i, Ano15j, Ano15k, Ano15l, Ano15m, Ano15n, Ano15s, Ano15t, Ano15u, Ano15v, Ano15w, Ano15o, Ano15p, Ano15q, Ano15r, Ano15-28, Ano15x, Ano15y, Ano15z, Ano15-27, Ano15-29, Ano15-30, Ano15-31, Ano15-32, Ano15-33, Ano15-34, Ano15-35, Ano15-36, Ano15-37, Ano15-38, Ano15-39, Ano15-40, Ano15-41, Ano15-42]. Matter [Ano15-43, Ano15-44, Ano16a, Ano16b, Ano16c, Ano16d, Ano16e, Ano16f, Ano16g, Ano16l, Ano16m, Ano16n, Ano16o, Ano16h, Ano16i, Ano16j, Ano16k, Ano16p, Ano16q, Ano16r, Ano16s, Ano16t, Ano17a, Ano17b, Ano17d, Ano17c, Ano17e, Ano17f, Ano17g, Ano17h, Ano17i, Ano17q, Ano17r, Ano17s, Ano17t, Ano17u, Ano17v, Ano17w, Ano17x, Ano17y, Ano17j, Ano17k, Ano17l, Ano17m, Ano17n, Ano17o, Ano17p, Ano18b, Ano18c, Ano18a, Ano18h, Ano18i, Ano18j, Ano18k, Ano18l, Ano18m, Ano18n, Ano18o, Ano18p, Ano18q, Ano18r, Ano18d, Ano18e, Ano18f, Ano18g, Ano18s, Ano18t, Ano18u, Bai76b, Bel12g, Bun10d, Bun10g, Bun10m, Bun10j, Bun10p, Bun12m, Bur74b, CS79v, Fin69, Glo76, Gou90, Gre74l, Grü74b, Jam10, Jer15, Kle09, Knu07, Lek15a, Loc76a, Mey85l]. Matter [Mie07a, Mis82b, Mor76, NB12, NB13, Oya09, Pal12, Pal13, Sch06a, SM07, Syn82, Tho96, War79g]. Matters [FA76a, HHS01, Mar84p]. Matthias [Bro95, Jah91]. Mature [Met89]. Maturity [Del91]. Maull [Sch80b]. Maupertuis [Bod18]. Max [Aga81o, Bee91, CS79h, CS79w, Ren07c]. Maximal [Ste97]. Maximizers [Mic74]. Maxims [Kas76a]. Maxwell [D’A81, D’A01k, Duh15c, Duh15e, Duh15h, Duh15i, Duh15j, Har95, Por94, Sie95, Sta17, Tur96]. May [Gat09, Bar14b]. Maya [Sab92]. Maybe [Aga03-29]. Mayerhöfer [vHBM01]. McClintock [Gro01]. Mead [Jar86f]. Mean [And18, Bar83, BR18b, Bon00, Sim77x, Sch99c]. Meanest [Gal02c]. Meaning [Bal90, Bec91n, Byr02, Čap71l, Čap71-29, Cel81, Chr15, Coh88b, Dam96j, Das76, Den99d, Den99e, Den99f, Den99h, Den99i, Dre86, Fit76, Kan79, Llo15, Mar95a, Mar84d, Mar84e, Mar84f, Mar84g, Mar84i, Mar84j, Mar84k, Mar84m, Mar84n, Mic97, Ner89, Odd89a, Pat90, Put86c, Rot87j, Sha84f, Var92, Wes76, Wun76, Zam92g, Zan74, Zem86, ZC88a]. Meaningful [Kis99]. Meanings [Boy01, Hor00, Nag98]. Means [Fek99, Fin14, Pos76a, Wie12, Onn91a, Onn91k]. Meant [Qui74]. Measurability [CS79-36, Dor00]. Measure [Han99m, Skl06, Sky83]. Measurement [Bel03, Ber83d, Ber83e, Ber83c, Ber83f, Ber83i, Ber83j, Bye74, Duw17, Fin93, Gru69b, Jan85e, KF86, MC80k, Mit76i, Nag84b, Sac67, Ste97, GG91]. Measurements [AV96, AV97, CS79l, Onn91n]. Measures [CL93]. Measuring [CS79x, Hag16, Sta83b, vH77c]. Meat [Ree14]. Mécanique [BV97]. Mechanic [Sta74a]. Mechanica [Lai08, Vil08]. Mechanical [Bou02i, Čap71-42, Cha09b, CHS97a, CHS97b, Duh91j, Gab04, Gar12a, Gar13, Has11, Hel00g, Kle74, Koç05, LvHV92, Lüt06, MPC91, Mit76i, Nag84b, She95a, Stö11, Van08]. Mechanicis [Gat18]. Mechanics [ACE17, Ano07n, Ano07g, AB17, Bal14, BS96, BC81, BB00, Ber17a, Bie82, Bit96e, Böh83, BV97, Bou02i, Bro08, BEW96, BD74, Bub96, Bun92, Büt08, Car00, CL93, CS79x, CS79-39, Col96, Cre02, Cus93, CFG96, CŻ82, D’A01m, D’A01g, Duh02b, Duh02c, DGZ96, FRSV18a, FRSV18b, Fel18, Fin96, Fje02, For10, Fre11, Gat18, GP79, Grü73m, Grü73u, Har96, Hel08, Hel00b, Hen98, Hol96, Hoo74, Ion15, JM06, Kan00, Ken96, Koc02, LR08, Lai08, Le 08, Lec86, Lüt98, Mac74, Mac12e, Mac12f, Mac97f, Mac97h, McC95a, Min07, Mit97, Mit00, Nor98a, Nor07a, Omo18b, Omo18a, Pag96, Pao84, Pos95d, Pos95p, Rad12e, RSS^{+}07a, RM18, Ren96, Rös91, Rou04, Sar08, Sch94a, Shi96, Sim77c, Sim77g, Skl76, Skl06, Squ96, Tem95, Unt01, Van08, Vil04]. Mechanics [Wal81e, Wal81n, Wal92f, Zab95, ZRKC03b, Zwi08, dCB11, dB84, dB87b, dB87c, dB87s, vL16a, vL16d, Sch08a, Sch08b, Lüt06, Mul98, Sch08c]. Mechanik [D’A01f]. Mechanism [Aga13i, Ano06-28, Gre95, Gro09a, Grü73b, Lam12, Lam13, Mey85u, Web83]. Mechanisms [Fos91, Sco80]. Mechanistic [Băe15, Duw17, Fre86f, Gro09b, Gro09c, Noé92, Ols87, Pâs15]. Mechanization [Fra12, Fra13, GR12, GR13a, GR13b]. Mechanizing [Góm12, Góm13]. Media [Car07, Wei11]. Mediation [Tru74]. Mediator [Han17]. Mediators [Fey95b]. medica [And18]. Medical [BE10a, Bro11, CF11, Coh02c, DP92, Fle86f, Gro95, Hac94a, Ito11, Kum11, Löw14, PV90, Tho10, WP94]. Medicine [And18, BE14, Bu96, Cla14, Coo11, Dol86, Ebr11, Fre86b, Gom15a, Hoh83, Hsü92, Izq74, LA86, Lee86, Mai12]. Medieval [CS79n, Duh91a, Fre00, Gra91, Har10, MS75b, Mur75, Ung91a, Wal81n, Wal81o, Wal81q]. Medium [Ano08p, HH91]. Medreses [İhs05]. Medusae [GWT00c]. Meeting [BC71, CHMV76, SC74a, Spr91a, dS14]. Meets [Gur03, Jur08, Ric12, SPB18]. Mehmed [Kil03]. Melquíades [dF14a]. Member [Mer97]. Memo [Jan07]. Memoriam [Wai14]. Memorization [Cha14]. Memory [Bog15a, BC71, CFW76, CW84b, Gro82b, Gro82f, Gro82n, Jed90, Kap80r, Kap80s, Kas76b, May95, CW67]. Men [CS79y, Gre74m, Mac86a, Mar95d]. Mendel [Fal01]. Mendeleev [TB15, Whe74]. Mendelism [Gon01]. Mendelsohn [AM01b, Ezr01]. Mendes [Hay03]. Meno [Eva95, Llo04, Sim77h]. Mensurandi [Bus74]. Mensuration [Lac94]. Mental [Čap71-33, Gen03, Mar78g, Mar84k, Pro09, Put86c, RDS^{+}18, Zem86]. Mentality [Mon01]. Mercury [Onn91g, Onn91d, Onn91e, Onn91i]. Mereological [Har15b]. Merit [Pos76b]. Merleau [Coh86b, Tam02]. Merleau-Ponty [Tam02]. Merry [Ned90a]. Merry-Go-Round [Ned90a]. Meshwork [Var91]. Mesmer [Wol89]. Mesmerism [Wol89]. Meson [CS79e, CS79x]. Mesopotamian [Pro10]. Message [Aga03-29]. Meta [AV16a, Ber13a, Gro95, Nol08, Sch99b]. Meta-Critique [Sch99b]. Meta-Induction [Nol08]. Meta-paradigm [Ber13a]. Meta-Philosophical [AV16a]. Metaarchaeology [Emb92a, Emb92b]. Metachemistry [Nor06]. Metacritical [Str84]. Metal [vSB91]. Metalaws [NT82]. Metallurgical [Lia96a]. Metals [Onn91e, Onn91g, Onn91c, Onn91d, Onn91i]. Metamathematics [Tos81]. Metamorphoses [EF86, Man86]. Metamorphosis [Ide00, Por87]. Metanarrative [Far03k]. Metaphor [Coh02n, Mac86a, Mah15, Mar94c, McR80, Mic98]. Metaphorical [Pet76, Zen99]. Metaphysical [Bun12m, BC02, Cas95c, Cur89b, Elk74, Fek99, FA76d, Góm12, Góm13, Len89, Par00, Sha84e, Sim99, Vil02g, Vil02k, Vil02m, Wri96]. Metaphysicist [Hal90a]. Metaphysics [Aga75c, Aga75h, Aga75m, Aga81b, Akr09, Alb96, Ams75a, AD94, Bel12d, Bot94, Bou02h, Bou02i, BH17a, But93g, But93h, CW84b, CHS97a, Dev01, Dil96, Dil06f, Don17, EL11, Fet10c, Fut08, Gab01, Gre74d, Jan89n, JL95a, Koh98a, Köh91a, Kra11, Nor06, Sha74b, Syl97b, Vil02h, War67a, War79h, War79j, Wes90b, Wis69]. Metapolitics [Chi17]. Metapontum [vF04]. Metasciences [Hal84, Jan84b]. Metazoa [Raf15]. Metchnikoff [GWT00e]. Météores [Mar12, Mar13]. Meteorology [Mar12, Mar13]. Meter [Gün92]. Method [Alb84, Amr87b, Ams92b, Ano08-41, BV94, Bel12j, Bit08, But93l, CS79-35, CS79z, CS79-39, CW83b, CH14, Cor07, Dun89, Fay97b, Fey79, Fin80h, Fin80s, Fre86h, Gen91, Gru81, Hil94, HR74b, HR74d, HR74e, Jan85c, Jan89m, Kap80v, Kau14, Kno94, Koe80, Koe86, Kos94, Krü90, Lag97f, Lau80b, Lau89, Lih93b, Lin93c, Mar80, McL71, Mus76, Nol03h, Pal01, Per81, Per86, Pol74b, Sav86, Sha84s, She95b, Ver09, Vil02i, War79l, WP94, Wyl92, ZRKC03g]. Methode [Swi96]. Methodical [Jan97]. Methodological [Aga95c, Ber83e, But93h, Car08, CN11b, CW74b, Duc96, Far03j, Fin80r, Fol17, FA76e, Gem90, Hat83, Jar86i, JN79, Kle71, Lih93b, Mar84p, Mot82, Now79b, Pan83d, Pan91b, Pan91e, Pan91f, Per89, Rap81c, SA82, Sch15, Ton73f, Wet95]. Methodologies [Hal71, Pan83a]. Methodology [Aga81b, Aga81p, Bol95a, But93g, CW83a, CW84b, Coh94h, Coh94m, Ela11o, Est96, Fin80b, Fin80l, Fri76, Gav89, Gou89, Hat83, How79, Koc79, Lam97, MP79, Meš89, Nic87, Pan83c, Pan83f, Pan91g, Pan91f, Rad81, SR16, Sma89a, Tur86l, Ung15, Urb89, Vla94, Wal76, Wal92d, Wat89, Wet17, Wia79b, Wor78b]. Methods [Bro91b, Chu11, CC90, D’A98, Ech96, Hea04, Hod74b, Kas76b, Kri03, MPC91, MA15, Mar91, Mül91b, Pat11a, Rei11, Sch12a, Sch18, SF94, Sim77m, Tuc76, Wis69, Wuß91, ZRKC03f]. Mètre“ [Gün92]. Metrical [Grü73t]. Metrizations [Grü73u]. Metrological [Gün92, Pro10]. Métrologique [Gün92]. Metrology [dB87t]. Metzger [Fre88, Ker88]. Mexican [LG01, RC95]. Mexico [Ace92, Bel95, Gut95, Izq74, Sal92a]. Mexique [Sal92a]. Meyer [BAK11]. Meyerson [Lau09]. Michael [Cra93, Dei17, Hol95c, Rhe05]. Michel [Lar99]. Michelson [Mel03, Pos95f]. Michigan [SC74a]. Micro [Gin97]. Micro- [Gin97]. Microbiology [Sch86c]. Microcausality [dB87q]. Microchronos [Čap71m]. Microcosmos [Čap71m]. Microeconomic [Lei83]. Microelectronic [Bor86]. Microelectronics [Byr86]. Microevolution [Hun17]. microflora [BBT95]. Microphysical [BB71]. Microphysics [D’A01g]. Mid [Fra91]. Mid-Nineteenth [Fra91]. Middle [Hoo91, Kuv79, Mor11, MS75b, Bea75, Duh91n, Duh91p, Duh91o, Wic91]. Mie [Bor07, SM07]. Migration [Kul01]. Migrations [Pia90]. Milestone [Nea00]. Miletus [Vam09b, Vam09c, Vam09n]. Military [Ágo11, Dib74, FS96, Mac01, Şak11]. Mill [Tur86g]. Millenarianism [Mau83, Pop86a, Pop88]. Millennia [Hac17, dB87i]. Millennium [Boc96, Cla10]. Millikan [Wei96]. Milne [Grü73p, Lim94]. Milne-Edwards [Lim94]. Mimesis [FH10b]. Mind [Ayd05, Bog15a, Bun10k, Bun10j, Bun10l, Čap71-35, GSW95b, Gro82m, Gua96b, Hei94a, Llo86, Lon93, Mar78n, Max76a, Ree87a, Sak98, Saw98, Set82, Sta94c, TB96, Tho01, VD92, Vil02a, Wat63]. Minding [Bun10m]. Minding [Oya09]. Minds [Dre86, Hel99, Kom99, Kur98a, Mar78i, Sha93]. Minería [Rui95b]. Ming [Bin96, Bré10]. Minh [Tho10]. Minimal [Aga03j, Mac12a]. Mining [Mei17a]. Minkowski [Ano07s, Jan03b]. Minute [SG11]. Miraculous [For10]. Mirroring [dF11a]. Mirrors [Ras95]. Miscellaneous [Boc96]. Miscellanies [Har10]. Misdiagnosis [Grü09]. Misery [Uus01]. Mises [Tok09]. Misplaced [Hel00h, Sie84]. Missed [PW17, Pul11]. Missing [Aga13j, Par09a, Ren90, Sau07]. Mission [Pet92a, Pet92a]. Missionaries [Hei88]. Mistake [Möl01]. Mistakes [Kis99]. Misunderstandings [CS79-27]. Misunderstood [Rös95]. Mitchell [Sch12a]. Mixt [Duh02p, Duh02o, Duh02q]. Mixts [Duh02l]. Mixture [Duh02m]. Mixtures [Car74, Duh02r, PO91]. Moa [Han99k]. Mobility [Gin88]. Modal [AM63, Bel93, Ber81, Bre74a, Bre74b, Bub96, Mar63, Pan16, Ser81, Sup74, WØ97, Zin73k]. Modalities [Ano05m, Mar63]. Modality [Bra11, Bre74c, Mar79t, Ray00, Van74a]. Mode [Rad89b]. Model [Add76, Ber17a, Bun93d, CPM16, Dav95, Dil06i, Fin80i, GS82, Gil09a, Gom15a, Gou83, Gro82i, Hem76, Hen98, Kau96, KL16, Kri92, LW11, Man71, Pan83e, Pol95, Rec11, Sca90b, Sch76b, Sim11a, Sør17, Syn82, TB96, Van15, War79i, Wat63, Waw82]. Modeling [Eşa15, Gil09a, Har08b, Jeb17, Les17, Wie12]. Modelling [Gie08]. Models [Bra11, Čap71y, Čap71-36, Čap71-42, CF11, Car10, DP92, Duc89, Dut11, Ele92, Est96, Fer98, Fra01, Für69, GG90, GC10, Gen03, Gro82b, Gru74a, HL17, Hun06, Hun10, JM06, Jos17, Kat06g, Kow16d, Kow16k, Lep80, Pal15, Pet76, PP81, Put67a, Put86a, RDS^{+}18, Sad89, Sal92b, SM90, SH91, Sim77m, Sne89, Sza79b, Ton73d, Too10, War79b, War79j, War79k, War79r]. Modern [Aga13q, AZ87, Ams92h, Ams92j, Ano06x, Ano08-46, Avg90, BHN98, Bec91o, Bra74b, Büt08, Čap71c, Cha09h, Chr94b, Chr14, Clu05, Dei17, Dil96, Dil06d, Dil06f, Dil06g, Dis06, Dra94, Duc92, Enz00, Erp91, Gai94, Gel84, Gou92, Gre74a, Gre76a, Haa99, Hat17, HR74e, HR76, Jer15, JW14, Kno04c, KM99, Lam82, Lel00h, Lel00i, Lel00j, Ler95, Lev83d, LY96, Mar11a, Mau83, Min10, Mit83, Mit76g, Nel67, Nia96, Omo18b, Raf15, Rap81d, Rös91, SM90, Sha84g, Sha15, Sie86, Str76, Tor99, Uus01, Vil02g, Vil02b, Vil02k, WW90, Whi88, Yus96a, Yus96b, Zam92c, Zel74, Zil03, Živ79, vL16d, vW87, Ras92, Kit90, Wan18a]. moderne [Ras92]. Modernes [Duc92]. Modernism [Lav88]. Modernity [Fri90a, Lec90, Lel00n, Lel00l, Lel00m, Que97, Von99]. Modernization [Fin14]. Moderns [Dou87]. Modes [Amr89, Hon98, Mar79m, Mor09, Pie16, War79r, Kle01b]. Modification [Cha09f, Gru74a]. Modified [Aga75g, She95b]. Modular [Wat63]. Möglichkeit [Mit04]. Molecular [Atl86, Dav86, Fra01, Hen01, Kay01, Mon96, Nye01, Sar96, Sch74, SA82, Sto91, Vol06, vdA91, Onn91k]. Molecularization [Sar97]. Molecules [Bro91b, Cha09j, Hef06, MM91, PS01a, PO91]. Moletti [Lai08]. Moment [Ley84, Sch94a, Sha15]. Momentous [Ung15]. Momentum [Bor07]. Monadic [Thi11]. Monads [Art89]. Monarchy [Yal96]. Money [Gou92, Orl92, Wel15]. Monism [Cao98, War79d]. Monist [Lon12]. Monistic [Add74, Mar89]. Monocycles [Kle74]. Montague [Mar79w]. Months [Bla95e]. Moody [Ito98, Wal81g]. Moon [Ren96, De 16]. Moral [Aga03k, Ano05f, Ben15, Byn89, Cah98, Cal13, Con96, Das76, Fra00, Gir01, Kle12, Koe09, Qui83, SR07, Ste09b, Tru94, Wan17, Zim86, For07]. Morality [Cah98, Grü95a, Koh98j, Nol03i, Qio96, Smi95b, Wol95j]. Morals [Smi95b]. Morgan [Mar11b]. Moritz [vH77a, Pul06, Sta91b]. Morphogenetic [Art15, Dum92, Ger15a]. Morphology [Bra87, Ham91]. Morphotypes [Len87]. Moses [Lan91]. Mossotti [Duh15f]. Most [Pog16, Wel15]. Motibus [Syl08]. Motion [Ano08n, Ano08p, Ano08-30, Ano08-34, Ano08-37, Ano08-42, Ano08-47, Ari76, Bec91a, Bel12h, Bla94, Bla04, Coh04, Duh91g, Ela11e, Ela11p, Ela11q, Ela11s, Föp07, Fri07a, Giu04, Hec01, Lag97d, Lag97h, Lag97l, Lag97k, Lag97m, Lag97j, Lag97p, Liz96, Mil15, Nee11k, New76a, Pal04a, PT04, Ren90, Rus76c, She89, Tel76, Vil08, Wal81d, Wal81p, Wal92f, Sch08a, Sch08b]. Motions [Gal02c, CS79n]. Motivated [Ung94]. Motivation [Cla12]. Motivations [Fra86]. Motive [War79l]. Motor [Gro82m, Gro82n]. Mountain [Wol89]. Mountainous [De 16]. Move [Fle91b]. Moved [Lag97d]. Movement [Bre88, Lai08, Rus79, Wal07b]. Movers [Zam92c]. Movies [Jar86e]. Moving [Wak15a]. Moyen [Bea75]. MS [Wal92a, Wal92i, Wal92j]. Much [Bar16, Mül94]. Muddle [War79i]. Muller [Cro92a, Wim92]. Multi [Ada11, Cho14]. Multi-Level [Ada11]. Multi-organizational [Cho14]. Multicultural [Lac01, Mat95]. Multidimensional [Whi01]. Multifaceted [Han17]. Multimerit [Res17a]. Multimodal [Ste12]. Multiple [Bun12j, Lau80b, Ned12, Sau16, Wie12]. Multiplicity [Čap71-40]. Multirealizability [Pro95]. Multivariate [Pia90]. Multiverse [Ven16]. Multiversism [Bar16]. Mundane [Kat06a]. Muris [Bus74]. Museum [Hei94b, Hei97, Mot15]. Museums [Lop92]. Music [Chr07, CL17, Kiv94]. Musical [BJ89, Ihd79a, Kiv94]. Musicology [McC02]. Musil [Ima01]. Must [Kra97, Wel15]. Mutagenesis [Sar91]. Mutation [Ech95, Fos91]. Mutual [Kai07, PO91]. My [Goo95, Lec14, Lip17]. Mysteries [Cho76, Llo04]. Mystery [DS11b, Ela11t, Rai10, Sta97]. Mystique [Ric17]. Myth [Alb95, Grm81, Jan89d, KF86, Mus91, Pop87, Swa09, War79b, Wei95, Win86]. Mythological [Kno89]. Myzostomum [GWT00g].

N [NRS92, Øhr97, RL95, Tho95, HD95, SVD95]. N. [CS79l]. Nôm [Tho10]. Nachlaß [Sch96g]. Naive [Aga97b, Leg03]. Names [CD93, Coh02h, Haq94c, Haq94e, Haq94f]. Namiki [Yan98]. Naming [Ges74f, Ges74y]. Nanoscale [MS11]. Nanotechnologies [Cas14]. Napoleonic [Det07]. Narrative [Feu88, Kno89, Res11]. Narratives [Dav10, Ric15]. Narratology [Rai10]. Nation [Maa17]. National [FS96, Kri92, Ano08o, Kar10f, Lek15e, Lek15g]. Nationalism [Arb92, Dav74b, Jar86j, Yeh17]. Natur [Swi96]. Natur-Wissenschaftliche [Swi96]. Natura [Ide00]. Natural [Dil06h, AT15, Atk03, Bhu06, Bon08, BH17a, BH17b, Bri98, BC94, Bro86, Che97, Cla96, CW74b, Coh94a, Coh94b, Coh02p, Coh02r, Cue01, Den99d, Den99e, Det07, DPW86a, Dum96, Dut11, Ege99, Ela11n, Føl99, Fre86i, Fre86r, Gal03, GR12, GR13a, GR13b, GSW95c, Gre07, Gre74n, Guy07, Ham91, Har89a, HH91, Hel08, Hin76, Hoo99f, Hoq18, Jar86p, Koh98k, Kra82c, LR08, Lam95, Lec94, Lef01b, Les17, Lim94, Lop92, Mar89, Mar11b, Mor83, NG01d, NG01e, Nou90, NRS92, Obr92, Ols87, Pel01, Pla94a, Pla94b, Qio96, Rad12a, Rad12c, Réd03, Rou12, Rou13, San92a, Sch67, Sha74b, Shi90, Sim11a, Tou02, Vam09g, Wes91, Wim12a, ZRKC03a, ZRKC03h, vW94, NG01c]. Natural-Language [Coh02p]. Naturalis [Syl91]. Naturalised [Gje00, Lel00f]. Naturalism [Bun10n, Goo00, Imo15, Jan89i, Kid17, Nol99, Nol03h, Sha15]. Naturalist [Arm88, BE88]. Naturalistic [Aga87c, Gra00, Sag87, SN87, Ska87]. Naturalists [Bar01b]. Naturalized [Ant87, Bir08, Bro96a, Den99f]. Naturalizing [Ant00]. Naturalness [SV15]. Nature [Aga75h, Ano08-40, Aya96, Bec91a, Bec91d, Bec91g, Bec91k, Bec91m, Bec91s, Ber74d, BS83, BH17b, Boy09, Bri98, Buc84, Cah98, Čap84a, Čap91c, Cau95, Chu95, Coh86e, CT98, Com84, Com02, Dan76, De 17, De 81, Di 84, DPW86a, Eck98, Els88, Fer98, Fey95a, Fin80b, Fre99, Gal91, Gar12b, Gir01, Gra98, Gre75, Gre74n, Gre74t, Gre95, GL91, Guy07, Heg87, Hei94a, Hon93, Hoo99d, Hoo99m, Kap80y, Koc02, Koh98c, Koh98d, Koh98j, Kon88, Lai08, Lee84, LS90, Max76b, MC80b, MC80h, MA83, MA87, Mic98, Neu89, Nev98, Nic12, NG01d, Ols87, Ols98, Pan94, Pan83d, Pet03b, Pul01, Roe01, Ros98, Rot87k, Scu98, SM96, Siv84, Sta94c, Von84a, Wal81j, Wal84, War79l, vW84, Wil07, dS11a, Ell03, Mar78h, Gre94]. Nature-Domination [MA83]. Natures [Aun71, Fis71, Haq94c, Haq94e, Haq94f, McM71]. Naturphilosophie [Fri07b, Heu96, WP94, Hah91]. Naturwissenschaften [Vaj99]. Natuurkundig [dV11]. Navier [Kah91]. Navya [Ano06b]. NEC [dB87t]. Necessarily [Fre99]. Necessary [Aga81-35, But93i, San81, Sch76e, Onn91i]. Necessity [Ano09m, Gre74n, Mad69, Neu76b, Spo10, Tho76]. Need [Kow16g, LL92, Mil95, Pit92, Ung04c]. Needs [Aga97a, Kau96, Suá08]. Negation [Čap71-30, Shr97]. Negativa [Par09c]. Negative [Čap71-31, Man71, Sma89b, Sta79]. Negentropy [Cac82, dB87g]. Negentropy-Information [dB87g]. Neglect [Hon09a]. Nelson [CW84b]. Nemore [Duh91n]. Neo [Cro99, Edg15, Fun86, Mas11, Pat15, Rau91, RT89, Tru95, Zol89b]. Neo-Darwinism [RT89]. Neo-Hellenic [Pat15]. Neo-Kantian [Edg15, Fun86, Mas11]. Neo-Kantianism [Rau91]. Neo-Kantians [Cro99]. Neo-Pragmatism [Tru95]. Neocortical [Ges74j]. Neohumanist [Sch96f]. Neoliberal [Bid11]. Nernst [Bar91a]. Nerve [Rou89]. Nesting [Fra96b]. Netherlands [LvHV92, Sie86]. Nets [Ano09j, Sne89]. Network [Mou92, NG01b, SP10]. Networks [Ada01, Gro82f, Gro82d, Gro82k]. Neumann [Bit97, Hon09b, Pos95t]. Neural [Gro82a, Gro82f, Gro82d, Gro82h, Gro82i, Gro82j, Gro82k, Gro82m, War12]. Neurath [Hal95a, Dvo91, Fle91a, Hal91b, Hal91c, Hal91d, Köh91b, Mül91a, Nem91a, Neu91, Rut91b, Rut91c, Rut91d, Smi97, Sta91c, Sta91b, Stö01, Ueb91a, Ueb91b, Ueb97, Zol89e, Zol89g]. Neurobiology [Kle12]. Neurocomputational [Chu91]. Neurological [Cam87, Gen71, Lek15e]. Neurology [Lek15c, Pog94]. Neurophysiology [MV80m]. Neuroscience [Blu16, Hym10, Sch02b, Web12]. Neurosciences [Sie84]. Neurosynthesis [Lel00k]. Neutral [Din96, Par01]. Neutrino [Fra99a]. Never [Gav09]. Newcomb [Voi96]. Newport [Pop88]. Newstein [Sta07c]. Newton [Aga81k, Ass11, Bec91b, Bec91f, Bec91g, Bec91o, Bec91p, Bec91q, Bec91r, Bel12j, Ber76b, Bou02i, But93m, Cha09n, Coh69, Coh74a, CS79u, CS79-28, CS79-29, DS11b, Fin80q, Fre86f, Fre86g, Fre86l, Fre86q, Fre86r, Gab01, Gal02c, Gal02d, Grü73w, Ham18, Hat80, Hes09, Lef01a, LP01, Okr89, Pal12, Pal13, Pao84, Pop92, Rey18, Rus76b, Sch18, Sch01, Sep87, She82, Vil02b, Wat01b, Wes92]. Newtonian [Bec91o, Bel12b, Bun92, Duh02p, Fer98, Grü73u, Sim77c, Swe94, dB87s]. NFCPS [Pan91h]. Niche [OS09, Ste09c]. Niches [Bir09]. Nicole [CS79n]. Nicolo [Duh91l]. Niels [Che94, CS79-31, CS79-30, CS79-34, CS79-36, D’A01c, Dai96, Fav94, FF94, HH94, Jon86, Kai94, Kan96, Kat11d]. Nietzsche [And99, Bab97, Bab99c, Bab99a, Bab99b, Bab99d, Bam99, Bor99, Cay99, Coh99, Con99, Cro99, Dav99, Fis99, Gau99, Haa99, Hab99, Hee99, Kop99, KM99, Lar99, ML99, Nol99, Nol03i, O’C99, Pen99, PT99, Poe99, Reh99, Roc99, Sch99a, Sim99, Sle99, Spi99, Str99, Swi99, Tau99, Van99, Von99, Whi99, Zim99]. Nietzschean [Dou99, Hod99, Mic98, Sch99b]. Nightmare [Tau97b]. Nine [Jie96]. Nineteen [Wol95d]. Nineteen-Eighty-Four [Wol95d]. nineteenth [D’A01l, Are89j, CALP03, Car15b, D’A81, D’A00, D’A01n, D’A01o, EP17, Fra91, Goo01, Gre91b, Hal90b, Jam92, Jon91, Kah91, Kle01a, Kra74, Kuz87n, Lek15a, LvHV92, Lub15, Mül91b, Pau17, Pes91, Rau91, Ric91, Roc01, Sal92a, SDG15b, Tur86l, Wel15]. Nineteenth- [Kuz87n, Lub15]. Nineteenth-Century [D’A01o, EP17, Gre91b, Hal90b, Kle01a, Mül91b, Pau17, Rau91, Tur86l]. nitrogen [BJ95, BdOU^{+}95, GC95, PHL95, PO91, WRL95, WPJ95, dBJD95]. nitrogen-fixing [dBJD95]. No [AZ87, Cha09h, Fra96a, KS95, Sim77v, Suá08, Zim02, Fra99k]. Noûs [Wed86]. Nobel [Elz95]. Nobler [Gal02c]. Nobody [Kle91]. nodulated [dBJD95]. Nollet [Sil11]. Nominalism [Ant16, Mar79e, MP16, Wal81g]. Nominalistic [Lam95, Pan96]. Nomological [Bun93d]. Non [Cob74, CG90, Des00, Fay94, Ges74q, Gou83, Gra07b, HN16, Hei94a, Hoo74, How79, Kos84, Kra11, Lie76, MC80l, Nag84b, Oes91, Ord96, Pan91d, SR15, Sau03, SS03, Zim83, Zin83a, dBJD95]. Non-Acceptance [Zim83]. Non-Aphasic [Ges74q]. Non-Conventionalist [Hoo74]. Non-Discovery [Ord96]. Non-Empirical [How79]. Non-Exploitative [Gou83]. Non-feminist [HN16]. Non-individuality [Kra11]. Non-introduction [SR15]. Non-Linearity [Kos84]. Non-Locality [Fay94, SS03]. non-nodulated [dBJD95]. Non-Observability [Des00]. Non-Phosphorylating [Oes91]. Non-Physical [MC80l]. Non-Quantum [Nag84b]. Non-Reductive [Hei94a, Sau03]. Non-Semiotic [Lie76]. Non-Separability [Fay94]. Non-Traditional [Zin83a]. Non-Viscous [Cob74]. Nonclassical [Pos95c]. Nonhuman [Fen12]. Nonidentity [Kuz87f]. Nonlocal [ACE17]. Nonlocality [Har97a]. Nonnatural [Den99f]. Nonpure [AP97]. Nonrational [Ber95a]. Nonreductive [Mar78i]. Nonself [RB91]. Nonsense [Bal08, Obe08]. Nonseparability [Web93]. Nontraditional [Akr09]. Norbert [Hon96b]. Nordström [Nor07d]. Norm [Bre11]. Normal [Jer15, Sim77x]. Normalcy [Mod15]. Normality [Lau96]. Normative [Dop01, Fre99, Sce06, Sch83c]. Norms [Car83, Fis96, Nol03h]. Norway [Fin14]. Norwood [CW67]. Nose [Hoo99b]. Not-True [Mar95c]. Note [Ano06-32, BVH^{+}83c, Čap91h, Haq94e, Hod74a, MA15, MP16, Sim77n, Sta74d, Sta86b, Tul03, Voi96, Wol76y]. Notebook [JNR^{+}07b, JRN^{+}07b, Ano07d, JRN^{+}07a, JNR^{+}07a, JR07]. Notes [Ano08-38, Ano08-47, Bla92b, Fol17, Haq94h, Hyd06, JRN^{+}07a, Kuh71, Len83, Len86, Mac86b, Nic89a, Sch07d, Sha84h, Sha84i, Tos81, Tym15, Wal92e, Wol95b, vH77a]. Nothing [Aga81i, Kuz87j, Wol95c]. Notion [BG18, Čap91p, Car07, Cat95, Coh94k, Duh02p, Duh02o, Duh02q, Har99, Jar86k, LC07, Mad69, Mar15, Ols82, Sco95, Wil00, Yin16, You74]. Notions [Boo12, KK07, Rou16]. Noumena [Cam87]. Nouvelle [Gre75]. Nova [Duc89]. Novel [Aga03y, Buc95, Man15a, Pos76b]. novelle [BV97].

Novelties [Čap91g, Hac15b]. Novelty [Aga75i, Čap71s, Ste09c]. November [Räz16]. Now-Ambiguities [Gun97]. Nowness [Fit74]. Nuclear [CS79-32, SF94, Stu95, Wan17]. Null [Eve74, Mar79h, Rob03, Shi90]. Number [Coh90, Dam96g, EN96, Ras94e, Sch96c, Tos81, Voi96, War79l]. Number-Free [Voi96]. Numbering [vH77c]. Numbers [Caw89, CS79t, Cox74, Ela11n, EN96, Sha86a]. Numerical [Hac94a, Ras94f, Ras94g]. Numerological [Coh90]. Nyāya [Ano06b, ZC88c]. Nyāya-Sūtra [ZC88c]. nya [CS79g]. Nyíri [Jan89e].

O [Cha12a, Wei06, Wol03, Cha12d, Cha12e]. Obedience [JJ07]. Obituary [Die91]. Object [Büt08, Fra09a, Góm12, Góm13, Jos17, Kuç95, Mur12d, War96]. Objectification [Ihd79l]. Objectifications [Ano09k]. Objection [Fin09c]. Objections [Čap71p, Mey91o, Str84]. Objectivation [Mar86g]. Objective [Den99c, Den99g, Fer85b, Fer85d, Fer85g, Lee74, Lel00o, Mar84a, Mar84l, Mar84m, Moo90, RA78a, Şen05, Tur86h, Urb78]. Objectivism [Pos78]. Objectivities [Csi15]. Objectivity [Aga74b, Aga03-30, Bit08, BS16, Cre99, Dav99, Die83b, Edg15, Fet10d, Gal15, Hac15a, Har15a, Hem83, Hem86, How15, Jar74, Jar86n, Jar86l, Kor71, Kow16i, Maa17, Mar97b, Mei97, Mus96, PRT15, Que00, Rad12b, Sav15, Seg15, TRP15, Zab95, Lid89]. Objects [AM63, Bal03, Bra11, Bun10a, EL11, Hil96, Hop74, Hoy74, Lek95, McG96, Pan91j, Par00, Qui76, SG11, Vol06, Wie12, Woo00, Zin73d]. Observability [Aga00b, AP00b, Des00, Enz00, Hei00]. Observable [Bus97a, Cor00]. Observables [Car00, Mer97]. Observation [D’E00, Emb92c, Far03f, Fle86e, Gal03, Geo00, Jar86q, Mac71, MAM69, Mos00, Sha84c, Tow71, Vla90]. Observational [Izq74, Sup91, Zan74]. Observationality [Ost76]. Observations [Bra74b, Haq94g, Led18, Şak11, Sto75, Wis71]. Observatories [Aub17]. Observatory [Car15b]. Observe [Bon00, Fal00, Byn89]. Observer [Kan00, Rim97]. Observing [De 16, Fay00]. Obstacles [Oru95]. Obstructing [BN17]. obtained [Onn91m]. Occasion [BAG17, CS79q, CS79-30, CS79-34, DPW86a]. Occasional [Aga81y]. Occidental [Vas92]. Occult [Fel92, Har10, Vic92]. Occupational [Gin88]. Ocean [Mac01]. Ockham [Ore00, Res17b, Syl75]. October [SC74a, CS79-30]. Odyssey [Ihd79a]. Oedipal [Tha84c]. Oersted [CS79j]. off [Was91]. Old [Cha09k, Fin76, Gol87, Joh11, Lor07b, Nov74, Pro17b, Pat03b, Pro14]. Olds [SC74a]. om [CS79g]. Omitting [Cas81]. Omnia [Cah05]. On-Line [KJ14]. Once [Bal90, Čap91e, SL82, VN12]. Oncological [Mar90]. One [AD94, Bar08, Ber76c, Bla95g, Bun12j, Coh76b, CHS97a, Coh02p, Gol12, Heg91, Hon09a, Hor94, Kae90a, Lih93b, Ly08, Mar95c, Mil79, Nee15, Nor91, Nov74, Pop84, Sim77x, Sta09, Wal95, Wet17, GHQ96b]. One-Sided [Hon09a, Mil79]. Oneness [Mey91k]. Oneself [Bra97]. Ongoing [Wei14]. Online [Ben15]. Only [BE10a, Yeh17]. Onnes [GG91, SR15, vD14]. Ontic [EL11, FL11, Mac74, Sta74a]. Ontogenesis [Dam96g]. Ontogenetic [Fle91b]. Ontogeny [Fin87, Oya92]. Ontological [Han99j, Ihd83, Kal96, Kle71, Lom15, Mar86a, MC80m, Mol18, Nol08, Per91, San97b, SM89, Str86, War96]. Ontologies [BR15, Wol69]. Ontology [ACE17, Apo82, Bit96h, Cal13, Chr11, Den99g, Fet81k, Grü73k, Hee97, Imo16, McM71, Mur12g, Pan91f, Pos18, Rub89, Sma89a, Teg97, Ton81i, Ung15, War79o, Wel99, Wol18, Zin73i]. Opacity [Rec00]. Open [Aga03e, Aga03k, Bed09b, Boy09, Eşa15, GL03, Mor74, Nar09, She95b]. Open-Ended [Aga03k, Eşa15]. Opening [Sam86]. Openness [BAG17, Gol89]. Opera [Cah05]. Operation [Aga75r, McC74]. Operational [Sie63, Żyt82]. Operations [Grü73d]. Opinion [Coo90, Wal92i]. Opinions [Aga13m]. Oppenheim [ZM11]. Opportunities [Pul11]. Opportunity [Ano08-32]. Opposition [Bel15]. opted [Wim92]. Optic [Fu93b]. Optics [Buc84, Com84, Smi90, Ung91a, Wal81n, dB87r]. Optimality [NPS17]. Optimism [Aga94, Arm18, Cla08]. Optimistic [Nol08]. Optimization [Dup92]. Optimizing [Fet10b, Pos95l]. Optimum [D’A89]. Optional [Eks84]. Options [Kow16g]. Oral [Fil05]. Orbit [DS11b, Lug89]. Orbital [Fis15, Mal03, vSB91]. Order [Ams92i, Ano06-30, Bra91, Den95, Goo77h, Gou90, Kau92, Max76b, Pul01, Tos81, Woo00, Yin16, ZM11]. Ordering [Sim77d]. Orderings [Bus97b]. Orders [Pul01, Vog91]. Ordinal [Ano09-45]. Ordinary [Ano05o, Onn91h]. Ordination [Pat76b]. Oresme [CS79n]. Organ [Mar94b]. Organic [BD88, Čap84a, Di 84, Kar10d, Kle09, Len88, Ols87, Bur87b]. Organicism [Bon87, Bur87d, Dou87, Fin87, Gel87, McG87, Pet87, Wet87]. Organisation [Bel15, Näp01]. Organism [GWT00i, Gre94, Hil94, Tau91i, Var91, Wag15]. Organisms [Gil09a, Sob91, Web11, dBJD95]. Organisms/Modeling [Gil09a]. Organization [Ano06-30, DL05, Ges74r, Gro82n, Her14a, Kap80c, Kap80z, Kau92, Kel09, McK75, Rus09, Şak11, Sim77q, Syn82, Tad79, Wim74, Wim76]. Organizational [Gro76, HH91, Kat14, O’T88, PP81, Cho14]. Organized [Gre94, HFD88]. Organs [Edg15]. Oriental [Vas92]. Orientalism [EP17]. Orientation [Fab84, Kow16j, Sch83b]. Orientations [Rhe11]. Oriente [PM14]. Oriented [Jos17, Krü91, SdV92]. Origin [Aga13a, Bou02d, Cor07, GWT00d, GL91, Har84b, HR74d, Jah94, Kel09, KS83, Liv92, Orl92, Rot86a, Sch00, Sch98, Syl08, Tha84b, VD92, vH77e]. Original [Eco92, Tha84a]. Origins [Aga81s, Bro92, Duh91r, DV92, Gir92, GWT00c, Gro09b, Kat14, Kau92, Mac97f, MS11, Mau83, Mit83, Pop92, Por94, RM18, Sup91, Tau91h, Tau91i, Tha84c, Tur96, VD92, Zam92c, ZRKC03g, Zil03]. Orphans [Koh98a]. Ørsted [Das07b, JJ07, BCK07, Can07, Chr07, Gre07, Ham07, Knu07, Kra15a, dAM07, NA07]. Orthodox [Mac12f]. Orthodoxy [Kat76]. Oscillations [Gro82c, Lag97t, Gro86]. Osianderian [Boy95]. Other [Bar08, Can07, Gen03, JJ07, Kur98a, Ly08, Mül94, Pie95, Ros76, Rös95, Rou16, Sau03, Sim77m, Gre74o, Mar84b, Ols82]. Otherwise [MV80j]. Otto [Dvo91, Fle91a, Hal91b, Hal91d, Nem91a, Neu91, Rut91d, Sta91c, Sta91b, Stö01, Ueb91a, Ueb91b, Zol89g]. Ottoman [Ágo11, BD05, Ele92, Gün92, Gün11, İhs05, Kil05, Pat11a, Şak11]. Ottomans [Ihs92]. Ought [Flo15, Rus80]. Our [Bek86, Bro10, Duh02i, Fer85d, Hoo94, Kel94, Pan91j, Put86c, Sch83a, Sch95a, Wia79a, Zem86, Urs96, Wol76x, Wol76-37]. Outline [Čap71-32, Hal97, Pan96, Ton81c, Ton81d, Waw82]. Outlook [CS79s, Löw86, Mey85m]. Outlooks [Nak95]. Outrunning [Hus17]. Outsider [dB87k]. Outsmart [Mar92]. Outsolutions [Ein01]. Outstanding [Kap80-39]. Overall [dB87k]. Overcome [Ghi93]. Overcoming [Sim11a]. Overdetermination [Wim12b]. Overlooked [Mat17]. Overseas [Mou92]. Overview [DFL11, Ebr11, Kat11e, Mac12g, Pog16, Sas92, dB87l, dB87m]. Own [Bek86, Can07]. Oxford [Bog15b, Fox15, Syl91, Wel15]. Oxygen [Kap80-27, Pap15, OP91b, PO91].

P [CS79b, Onn91m]. P. [vW94]. Pacific [Mac00, Mac01]. Page [Sau07]. Pain [Ric02]. Painting [LE97, Sha74c]. Palacios [SR15]. Paleobiologist [dR17]. Paleontological [Tas17]. Paleontology [Hus17]. Palestinian [Han17]. Pambour [Kro92]. Pan [Ste92]. Pancritical [RB09]. Panel [Grü73o, Grü73v]. Panmathematism [Čap71z]. Panteleimon [Luc88]. Pantheon [Jaf02]. Paolo [EF18]. Paper [Bre74a, Bur83, Hen98, HH08, Ito98, Kle01a, Kox95, Nic80b, Nye01, Ram01, RT01, Rhe05, Sce01, Zuc69]. Papers [CS79a, D’A01c, Ges74w, GWT00e, Sch96f, Wol76z, WC91, GG91]. Pappus [HR74a, HR74g, HR74f, HR74j]. Para [Fin17, OP91a]. Para- [OP91a]. Para-Clericalism [Fin17]. Parables [Car10]. Paradigm [Bal08, Hac17, Han15, Mar86g, Mar86h, Näp01, Nou90, Obe08, Sha84j, Spa03, Woo06, dB87d, dB87n, Ber13a]. Paradigms [Ams75f, Ano06-28, BJ89, Far03i, Kou89, Lei83, Mar86f, Tör76, Web11]. Paradise [Lev83d]. Paradox [Bec91k, FH10a, Grü73t, Han95, Hin89, IdMM16, Lev76, Lin93b, Mey91c, Sky83, Ste18, Tra15, dB87n, Nag71, Sim77h, Web93, Zuo96]. Paradoxes [Ack92, Bud93, FA76c, FA76f, Kuz87k, Lin93b, RB91, Rus76d]. Paradoxical [Ges74s]. Paraíso [Iru01]. Parallel [Bun12n, Das07a, Gro82a, Gro82b, Nye07, Ste99]. Parallelen [Sch99d]. Parallelism [Jon86, LC07, MC80c, MC80p, Wol69, Vam09g].

Parallels [Gor07, HN16, Sch99d]. Paralogism [Wun01]. Paralysed [Lek15e, Lek15g]. paramagnetism [PO91]. Parameters [HL17]. Paranoia [FA76c, FA76h, FA76f, FA76g]. Paraphrase [Mar76d]. Paribus [Ano09-44, Llo15]. Paris [Bie87, Gün11]. Parisienses [Wal81k]. Parity [dB87o]. Parmenides [Cor76a, Vam09l]. Parodoxes [Ber76a]. Parsimony [Mol18, SA82]. Parsons [Jaw90]. Part [Bus74, GHQ96a, Sha84i, Tha86c, Waw82, CLR^{+}92, GHQ96b, MC95, MC96, Sha84h]. Partial [DF93, Grü63, Grü86, Lep95, Odd89a]. Participant [Jar86q, MAM69]. Participation [KT76, Pes07]. Particle [AV96, CG90, Fre86g, Gou89, Joh96a, Kai94, Sto97, Wal67]. Particles [Čap84b, Čap91i, Coh02r, Hee84, vSB91]. Particular [MV80c, Rad89b]. Partisanship [Tru94]. Partners [Rip92]. Partnership [LL14]. Parts [Den05, GWT00i, Kau71, Kau76, Syl91]. Party [Yuz96]. Passage [Ism17]. Passion [CHS97b, Hod99, PR97, War79a]. Passion-at-a-Distance [CHS97b]. Passive [Tur16]. Past [Ano08-35, Bar91b, Bro89, Čap71q, Čap71v, Čap91c, Emb92a, For01, Grü73b, Jin96, LM97, Liz96, Mar86c, Roc11]. Pasteur [Mou92]. pasture [Tho95]. Patchwork [Bre11]. Patent [Ara15]. Paternalism [Par01]. Path [Gro82k, Kar10c, Sch95b]. Pathogenic [Bel81b]. Pathology [Hac94a]. Paths [Sta97, Mat01]. Pathways [Bel12g, Ger15a, RS07a]. Patriarchal [Mou92]. Patrick [Bab02c, Cro02, Mar02, Nev02, Sza03]. Patrizi [Ros81]. Patronage [Coh90]. Pattern [Gro82a, Gro82c, Gro82h, Gro82k, Gro86]. Patterns [EPL69, GP03]. Paul [Pap15, Die91, Kap80-30, Mun91a, Nae91, ZM11, vH77a]. Pauling [Kar10a, Kar10g, Nye01]. Paulo [Pet92a]. Pavlov [Kap80s]. PDP [Chu95]. Peace [Ano05n, Bar15, Rat88]. Peacemaking [Han17]. Peano [Seg95]. Pearson [Ska83]. Peculiar [Mas71, Slo83]. Peculiarities [Mäk96]. Pedagogical [Dam96h]. Pedagogy [Goh81, Sas95]. Peirce [Fra96b, Ska83, Ska87]. Pell [Scr74a]. Pendulum [Büt08]. Penrose [Ste18]. People [Coh02a, Gre74o, Kap80k, Yuz96]. Peraqim [Fra00]. Perceive [Ano06-30].

Perception [Ano06i, Arm76, Cho63, Coh70, Efr69, Fax97, Fle86e, Gre74r, Gro82m, Hee86b, Hef87, Ihd99b, Jar86w, Mar78e, Mor63, Sau02, Ska92, War79m, War86, Yil74, vH77b]. Perceptions [Del91, Lue18, Von99]. Perceptual [Ihd99b, Mus09, Ohm98, Szu00, Var92]. Perennial [Dil06b]. Pérez [Bar14b]. Perfect [Han17, Koç05]. Performance [Cre99, Gro82n]. Performed [Grü73d]. Perils [Kow16e]. Period [Cla14, Gie86, Moc91, Pos95h, Pos95m, Min10]. Periodic [BV01, Sce01, Van74d]. Periodicals [BD05]. Periodics [Hef06]. Periods [Lia96c]. Peripatetics [Duh02l]. Peripheral [Ano07h, Nav15]. Peripheries [NG15a]. Periphery [BS03, Das15, Lei03, Nak95, Sch95b, Van15]. Permanent [Aga81y, Nee97, Onn91k]. Permits [MA83]. Permitted [All99a]. permutations [Bur87b]. Perpetual [Duh91g]. Perplexed [Fre00, Gel95]. Persecution [Fun86]. Persistence [Sta17, Onn91l]. Persistent [Čap71j]. Persisting [Roc11]. Person [Ano05e, Kle91, Mar97a]. Personal [Alo14, Alt14b, Ber74b, Feu88, Gin88, Mis95b, Res11, Sco80, Van83, Wol95k]. Personalities [Hac86]. Personality [Čap91j, Kar17]. Persons [Mar78i, Mar78j, Mar78o]. Perspectivalism [Hee99]. Perspective [Alo14, BRS18, Cha09c, Che14, CS79c, Dau96, EFN03, FEE18, Gie08, Gin88, Han14, Hel14, Krö12, MC14, Mas15a, Mas11, Mil87, Nij15, Nik15, Pos95q, Roc99, Sal92a, Sha74a, Str18, Str86, War79q]. Perspectives [BP14b, Bla92a, BH17b, Coh94b, FSU97b, Gil88, Gon15a, Ism17, KB92b, LMZ18, Lov15b, Moo18, PRT15, Pap94, Rav18, Sab92, Sar96, Sau16, Svo15]. Perspectivism [Sle99]. Perspectivist [Wel99]. Persuasion [BJ89, Fin80d]. Peru [Cue01]. Pessimism [Aga94]. Pessimist [Cla08]. Petrovich [Kap80s]. peut [CLR^{+}92]. peut-elle [CLR^{+}92]. Phantasy [Wol76w]. Pharmacology [Fra09a]. Phase [BIT01b, CS79m, Duh02t, MC80t, MC80u]. Phenomena [Bro96b, Cam87, Duc89, Gav09, Gra83a, Hac91, Hem84, Kat06g, Mey85c, Mey85g, Mey91a, Pat03b, Ray82]. Phenomenalism [BIT01c, Cek92b, Kre91]. Phenomenalist [MC80g]. Phenomenological [Com02, dB87p]. Phenomenology [Are89i, Bar02b, Coh86d, Dol16, Emb92c, Fje02, Haa99, Hin97, Ihd79e, Ihd79f, Ihd79g, Koh98e, Mar86b, Nev02, O’C02, Tha86f, Ung15, Wol76-28, Wol95h, ZRKC03a, ZRKC03h]. Phenomenon [Bou02c, Sch83f, Sep87, Zim83, dB87f]. Phenotypic [Nij15, Ste15a, Ste09c]. Philipp [BIT01g]. Philips [dV11]. Philology [Are89e, Chr94b]. Philopon [Duh76c]. Philosopher [BHN98, Cao95, Cha09a, CS70, Fol96, Hoo99g, Mar84p, Mil11b, Nol91, Rut91d, Vil02g, Wor76]. Philosophers [Bun17, Bur80, But93j, Can12, Emb92b, Gol12, Hac94b, Mac90, Res89, Rus80, Vil02e]. Philosophiae [Syl91]. Philosophic [Mey91p]. Philosophical [Aga13l, AV16a, Ams75a, Ano06-30, Ber83f, BIT01h, Bor86, BH17b, Bun10j, Bun12o, But93d, Čap87a, Dam96h, Dev15, Dis06, Ear06, Erp91, FB17, GC06, Gro95, Grü73p, Grü73q, Grü73r, Grü73s, Hat17, Hon16, Hoo94, Jam79, Jia96, Kat11f, Kör86, Len86, Lov15b, Löw86, Mac92, Mac12g, Mar94a, MC80n, Men76, Mis74, Mit76g, Nag84a, Nag98, Oru95, Pet69a, Pet79, Ras00, Rau91, Rös94, Ros09, Sal92b, Sch15, SC74b, Sha84g, She95b, Shi94, Sol12a, Sta91b, Str83, Von84b, Wal81o, Wil18, Yin18, Zan74, ZM11, Zol89g, Zuo96, GC97]. Philosophico [Wil01]. Philosophicus [Rad03]. Philosophies [Aga75d, AV16a, AV16b, Bun12h, Bun12p, CT98, Far03e, Hul76a, Lar13, Pul11]. Philosophizing [But93m, HH91]. Philosophy [Abb94, Aga81t, Aga81-29, Aga81-30, AC82, Aga94, Aga95a, Aga03l, Aga03u, Aga13c, Alm89, Ams75g, Ana90a, AD94, And09b, Ano06c, Ano08k, Ara12, Ara79, Asl18, Aug67, Bab97, Bab99a, Bab99d, Bab02c, BSM06b, Bal89a, Bar02a, Bar74, Bec91d, Bec91g, Bec91m, Bec91s, BV09, Ber84, Bit96e, Bla95g, BIT01c, Blu16, BS16, BG09, Bre09, Bri18, Bro86, BM89, Buc84, BC71, Bun10o, Bun12l, Cam03, CG16, Čap71-32, Cas95c, Cek92a, Cha09b, Cha09d, Cha09a, Cha12b, Che95, Che97, Chi81b, Chr14, Chu00, Cle02, Cof83, CT09, CW69a, Coh69, CW69b, Coh70, CHMV76, Coh80, CL83, CW85, Coh86b, CHR96, Com84, Com02, Cor95, CCG93, Cro99, DC96, De 01, Dea12, Deb09, Det07, Di 84, Dil06b, Dur83, DR83, Ear15, Eas84, Eck98, Fag12, FL09]. Philosophy [FF94, Fer98, Fey84a, Fra12, Fra13, Fre86a, Fre86b, Fre86e, Fre86i, Fre86j, Fre86k, Fre86m, Fre86r, Gab04, Gal00, Gar16, Gar12a, GR13a, GR13b, Gar13, GSW95c, Gay09, Gie86, Gie71, Gie12, Gin03a, Goo95, Gra83e, Gra98, Gre94, Gre74i, Gre74t, GM76, Gre83, Gro09c, Grü69a, Grü73c, Grü73g, Grü73h, Grü73v, Gur03, Hah91, Hal84, Hal91e, HN16, Han99a, Har16, Har84b, HCP18, Hee99, Hei98, Hel08, HJLP06, Hod92, Hol83, Hon03, HOO88c, Hör91b, HFD88, Ihd79d, IG05c, Irz15, Jan84b, Jan89e, Jan89k, Jan89l, Jar86c, Kan79, Kan84, Kap80-28, Kat11c, Kay76, Kea02, Kir82, Kno04d, Kno16, Koc97, Koe74, Koe09, Koh98g, Kop99, Kos76c, Kos76d, Kos94, Kra82c, KV11, Kro89, Krü88, LR08, Lao95, Lau09, Laz77, Lef01a, LM97, Li18, LF93]. Philosophy [LPV15, Lub82b, Mah75, Mai12, Mai16, Mal12, Mal13, MC95, MC96, MP79, Mar76c, Mat95, MS12b, Max76a, McM78, Mcm76b, Med96, MC80z, Mei17a, Mes90, Mic18, MP13, Mil12, Mis95c, Mit86b, MH86, MLNZ18, Mit76a, Mit89, Mos96, ML99, Mun91a, Mun96, MS75b, Mur12a, Mur12f, NC98, Nar09, Nel67, Nic80c, Nic89b, Nic90, NA07, NNEH11, NRS92, Okr89, OP82, Oru95, OP97, Pal76, Pan91e, Par09b, PST15, Pǎr15b, Pat90, Pes91, Pet69b, PP18, Pie12, PR12a, PR12b, Pul11, Put76, Rad12c, Rad12e, Rad76, Rad82, RC95, Rap81a, Rap81e, Rat88, Rav91, Ric15, Rou12, Rou13, Sag87, Sar96, SS16b, Sce06, SM15b, SC74a, SV18, Sch75, Sch86b, Sch86c, SR16, Sch06b, SM90, Sha74a, Sha84c, Sha84k, Sha84l, Sha16, She82, She95a]. Philosophy [Shi83, Shi86, Shi98, Smi97, STNW12, Soy05, Sto00, Str83, Str84, Syl97a, Tor95, Tua16, UM86a, UM86c, UM88, UM92, Ung15, Vam09g, Van06, Van08, Vih01a, Vih01b, Vil02a, Vil02c, Vis92, Von84a, Wan18c, Wan18b, War63, War76, War79l, War79p, Wes95, Wol86, Wüt16, Yil74, Yin18, YCMY18, Zaj87, Zim83, Bla95i, CW67, GSW95a, GC97, HH91, GR12]. Phobosophies [Bun12p]. phosphorescence [BBO91]. Phosphorylating [Oes91]. Photosynthetic [MPC91]. phototrophic [RL95]. Phrenology [Jac07]. Phyla [Fre15]. Phylogenetics [SA82]. Phylogeny [Fin87, Fre15, Oya92]. Physica [Duc89, Ela11r]. Physical [Ana97, Bal90, Bar91b, Bec91k, Bel12g, Bel12e, Bel12f, Bie82, Bon00, Bre81, Bun12m, Čap71a, Čap71-33, Čap71-34, CW84c, Coh86e, Cos87b, D’A01m, D’A01e, De 01, Die16, Duh91r, Duh02i, Duh02n, Ein01, Eks84, Fra09a, Gav94, Gol91, Het15, Hoo99h, Hyd06, Ism16, JX96, Jon86, Kap80f, Kap80z, Kap80-29, Kno89, Kuu01, Lan67, Len89, MM91, MC80l, MC80o, Mis74, Näp01, Nia96, NRS92, Pas01, Pat76b, Pau97, Pol03, Pos74, PS01b, Que97, Qui76, Sch95c, See74, Sim77b, Sta74b, Van74b, Vil02g, Vil02k, ZRKC03c, Zin73l, AAF^{+}14, Ein01]. Physicalism [Ueb95]. Physicist [BHN98, Bee91, Buc91, CS70, Kap80-30, Omo18a, SR07, Stu07, Web08, d’E08]. Physicist-Historians [Stu07]. Physicists [Bit96d, Hei91, Jhu07, Lec14, LL14]. Physico [Kol74, dB12, dB13, Kap80-38]. Physico-Mathematics [dB12]. Physics [Aga75c, Aga75r, Alo14, Alt14a, Alt14b, AB14, Ano07o, Bab02d, BAC14, BRW14a, Bar14b, Bec91f, Bec91r, BD88, Ber74a, Ber70, Bit90, BH17a, Bra11, Bra96, Bra95, Bre74b, Bru95, Bun12g, Bun12n, Bus03, Cao07, Čap71c, Čap71-30, Cek92b, Cer14, Cha15, Che97, CD93, Cin89, Cla89, CT09, Coh70, CW74a, CS79c, CS79z, CS79-46, CL83, CW84c, CC90, CCG93, D’A98, D’A00, D’A01f, D’A01n, D’A01o, Daf96, Dis06, Dra94, Edm95, Fin09i, Für69, Gal91, Gar95a, Gau95, Gav89, GSW95c, Gia99, Gou89, Grü73w, HTG14, Han14, Har16, Hat17, Hec01, HJLP06, Her74, Hil07a, Hil07b, Hil07c, How97, HH94, JS03, Jam79, Jan85c, Jan03a, JW14, Kai94, Kap80d, Kap80h, Kap80-33, Kar10b, Kat06a, KJ14, Köh91b, Kra91, Krü90, Kuu01, Len88, LL92, Lor07b]. Physics [Mac12d, Mac12c, Mac12i, Mar82, MC95, MC14, MC80i, MC80n, Mir14, Mit76g, NT82, Nav15, New15, Nia96, Nor98a, Nye91, dJAO14, Pat92, Pau17, dMPC14, PM14, Pet76, Pit92, Por94, Pos95h, Pos95m, Pro17a, Qua84, Réd03, Red93, Rei96, RDS03, Ren07a, RSS^{+}07a, RSS^{+}07b, RS07b, RS07c, RS07a, Rom69, Rös91, SR15, San14, Sau07, Sau93, Sch01, Sch84, Sch95c, Sco80, Shi98, Sil11, Sim11b, SC14, Sto97, Str76, Stu95, Syl75, Syl91, Teg97, Tem82, Tis63, Ung91b, Vem06, Wal67, Wan17, War79e, Wig74, Wil01, ZRKC03i, Zin83b, Zol89i, Zwi09, dF14b, dF81, vW69, D’A01l, Hof98, vD15]. Physiological [Gro82l, Hol95a, Nij15]. Physiology [Cue01, Edg15]. Physique [Pat92, Sch01]. Piaget [Čap71d, Čap71-37, Čap87a, Čap91l, Dam96c, Mar76b, May02]. Pictorial [Čap71-42, Mül91a]. Picture [Hei98, Hin94, Maj98, Que97]. Picture-Conception [Maj98]. Pictures [Hel00g, Mit89, Vol06, War79n]. Piecemeal [SdV92]. Pierre [Nee11c, Bla04, Pal04a, Wal81p]. Piezoelectric [Kat06d, Kat06g]. Piezoelectricity [Kat06a]. Piles [Cha12c]. Pilot [BEW96, Val96]. Pilot-Wave [BEW96, Val96]. Pincock [Imo16]. Pioneer [Ano08-46, Gav09, Ren07c]. Pisa [Cai14, Hel08]. Pisan [Ito98, Wal90]. Pitfall [EE17]. pitfalls [dBJD95]. Pitting [She82]. Pitton [Das15]. Plaass [vW94]. Place [Car89b, Cro92b, Di 84, Duh76b, Duh76c, Fet10a, Jar95, Joh11, MS11, Nar09, Ols82, Sta94c, Str15, Wat92b, Wet82]. Placebo [Grü84]. Places [Spr91a]. Placing [Glo76]. Plague [Mit76a]. Planck [CS79w, Koc91, Pla92, Sch00]. Plane [Ano08n, FR08, MA15]. Planets [Coh90]. Planetts [Gal02c]. Planned [Nem91b]. Planning [Kap80-31, Sto75]. Plans [Fre15, Gro82n]. Plant [Hei08f, PLH95]. plants [dBJD95]. Plasma [Kap80-32]. Plastic [Bun10m]. Plasticity [Nij15, Ste15a, Ste09c, War12]. Plate [Lau80b]. platinum [Onn91c]. Plato [Bec91s, Čap71-28, Cop76, Duh76d, Glu91, Kar90, Kou90, Nai95, Ore00, PR17, Ros98, San90, Vla90]. Platonic [Bec91c]. Platonism [Sig91]. Plausibility [Sim77o]. Plausible [Aga75e]. Plaza [SC74a]. Plea [Che05b, Grm81, Mar84p, Sil89]. Plenum [Des76]. Plücker [Hie95a]. Pluralism [Aga03v, Cha12f, Cha12e, Far03a, Far03j, Gra07a, Kin16, Mun15, Sch15]. Pluralisms [Wyl15]. Pluralist [Lon12]. Pluralistic [Pat82]. Plurality [Wyl15]. Plus [dB87t]. Plutarchian [Söd07]. Plutonium [Whi01]. Pneumatics [Val10f]. Podolsky [Cos87a, Har97a, Nag71, Web93, Zuo96, dB87n, Alb92]. Poe [Byn89]. Poem [Ele94]. Poems [Hie95b, Shi94]. Poetic [Hal82, Hal92, Scu98]. Poetics [Car89b, Kea02].

Poetry [Bon87, Wol95j]. Poincaré [Čap71t, Gia71, Grü73w, SL84]. Point [Ano05g, Dah90, Del91, Dis06, Gau84, Has97, Rad89a, vdW04]. Pointing [Jin96]. Points [Bus97b, Lag97j, Wei76]. Poisonous [Sch12a]. Poland [Sre87]. Polanyi [Cat95, Dei17, Hol95c, Mis95c, Sco95]. Polarity [Bur87c]. Polemics [Pla92]. Poles [Cha12c]. Policies [Her76]. Policy [BC94, CLR^{+}92, Kow16g, Kow16k, Kow16h, Kow16j, Mar01, MZ18, SG92, Yuz96]. Polikarov [GC97]. Polish [Gie86, Kra82c, Mis95c, Wia79b, Wol86]. Politeia [Win83a]. Political [Aga86, Bar01b, Bel15, BN17, CSW74, Hol83, Jan89n, Kre79, KT01, Lee86, Mit83, Nar09, O’T88, Omo18b, Omo18a, Osb92, Pal15, Pap74b, Pet89, Sar11, Sch94a, SN11, Sch94b, Sta91b, Syl89, Zol89c]. Politicization [Wis11]. Politicized [Ree87b]. Politics [Aga03m, Aga03o, Aga03-30, All01, And84a, And84d, Bam99, BJ89, Bur86, Car11, Don17, Elz95, Ezr14, Fin14, Hei08e, Jan89g, Jan89k, Jan89l, Kre79, Mal92, Str99, Sun18, Wei95, Win86, GSW95a]. Politicus [Dju79]. Politique [SG92, CLR^{+}92]. Pollen [COG90]. Pollution [Zim86]. Polyhedral [Cox74, Lih93a]. Polymath [BIT01a]. Polytechnic [Car15b]. Ponty [Coh86b, Tam02]. Popper [Far03j, Aga75n, Aga09, Akr09, And84c, Bar74, Bar09, Cat95, Cur89a, Dei17, Fer85b, Fer85e, Fer85f, Grü76, Grü78, Grü89a, Grü09, Han99k, Irz95, Jar86p, Jar09, Koc09, Koe74, Koe09, KZ09, Lao17, Lin93d, Mis95c, Nar09, Pap89, PC09, Pos95t, Rad82, Ros09, SM89, Ste09b, Str84, Tok09, Ver09, Wet09, Wor78b, Wor89, Zem09, Zol89e, tH09]. Popperian [Bol17, Fer85a, Fey78a, Koe95, O’H09, Par09c, Rad76, Rad81, Seg09, Swa09, Wat78b, Zab95]. Popperianism [Hüb78b]. Popular [Bra91, Elz95, Gen91, GH01]. Popularising [Gra06, Sch01]. Popularization [Cas14]. Population [GM17, MG11, Sim77x]. Populations [Tou71]. Pornography [Jar86i]. Portrait [CW85, Kap80p]. Portrayal [Bar09]. Portuguese [CALP03]. Position [AV96, Ges74s, Omo18a, Sta91a]. Positioning [GR07, RG07]. Positions [Krö91, Oya09]. Positive [Aga75k, Aga75j]. Positively [Cas81]. Positivism [And84b, Coh99, Gut95, Mac86b, Mar86g, Sil89, Dan92]. Positivisme [Dan92]. Positivist [Prz95]. Positivistic [Men70, Sha84h, Sha84i]. Possibilia [Mar79j]. Possibilities [Mey91l, Moo90]. Possibility [Bar10, Bre81, Har16, Sch02a, Tur86h, Mit04]. Possible [Coh02g, Dun92, Fey67, Hec01, Kea02, Lin82, McB94, Pet03a, Ste94]. Post [DN94, Dum92, Sha84h, Sha84i, Wag15, FK93, Pap14]. Post-2WW [Pap14]. Post-Byzantine [DN94]. Post-Dahlem [Wag15]. Post-Positivistic [Sha84h, Sha84i]. Post-Structuralism [Dum92]. Postcolonial [Moo18]. Posterior [Wal92e, Wal92c]. Posterity [Aga13d]. Postface [Gra83d]. postmodern [For07]. Postmodernism [Bab97, Fax97, Fra99g]. Postprocessual [Sal92b]. Postscript [AZ87, FA76i, Hab99, Pen99, Sch07c, Sie63, Str99]. Postselection [Per97]. Postulate [Min07]. Postulates [Gro82l, Wes76]. Postwar [Kar10h, Whi01]. Potential [BLA95a, Kow16h, Mie07b, Mil15, SM96, TBC91, dBJD95, Onn91i]. Potentiality [Bec91d, CHS97b, Mey91s]. pour [Pet92a]. Pouvoir [Sal92a]. Poverty [Bun95b, Grü95a, Jar83, Jar86u, May95]. Power [Bal18, Bar15, Bel81b, Ezr86, Ezr01, Gal02e, Jef71, Kap80o, Kle01a, Len83, Men86, Nic18, Nol03f, Ols87, Pai95, Pan91k, Ret74, Sal92a, Ste99, WvB07]. Powers [Mcg74]. pp [CS79g, CS79n, CS79-41]. Practical [Ano08-46, Ber13b, Ber86, Ber74d, Fel18, Gon15a, HCP18, Hon96a, Koh98h, Lam14, Mac86b, Mar84n, Pea86, Rie99, Sta82, Tay92, UM86b]. Practice [Aga84, Aga95b, Aga03-33, Ara12, Bea75, Bré10, Car11, DG15, Dra94, Ede88, Fin91, Fin80r, Gou94a, GC94, HR74a, Jar86y, Kap80i, Kap80j, Kap86b, Kek91, Kos94, LPV15, Mar74a, McC94, MH86, San18, Sch08c, STNW12, Sta74d, Sta86b, Sta09, Tor00, Vod15, Wei14, Wim92, Wyl15, GSW95a]. Practices [Boo12, BCK07, Bre11, Elk10, Kom99, Moo18, Sim15, Sol08a]. Praeceptor [Lõh01]. Pragmatic [But93g, Far03f, Fay07, KT93, SM96, Sig91, Tow71].

Pragmatical [Wun76]. Pragmatics [Bun93h, But93d, Kou90, Lie76, Mar79n, Mar79o, Soy05]. Pragmatism [Ear15, Kow16i, Mar83b, Nev02, Ska83, Tru95, Vil02d]. Pragmatist [Kow16j, Lon12]. Prague [Hof92]. Praise [Pos93]. Pratique [Bea75]. praxi [And18]. Praxical [Mar86d]. Praxis [Coo84, Hoh83, Ihd79i, Kos76e, May02, Pet69b, Ric02, Sto75, Tou76, War79c, War79e, MP79]. Praxis-Related [Hoh83]. Pre [Bec04, FRSV18a, FRSV18b, Gar12a, Gar13, MC80i, RR15, İhs05]. Pre-Classical [FRSV18a, FRSV18b]. Pre-Einsteinian [MC80i]. Pre-Eudoxic [Bec04]. Pre-history [Gar12a, Gar13]. Pre-Ottoman [İhs05]. Pre-war [RR15]. Preceding [Lag97g, Lag97h, Onn91f]. Preclassical [Fel18, Sch08a, Sch08b]. Precursors [Ito98]. Predicament [Ano08-31]. Predicate [Mas71, Zin73r]. Predicates [Ano09-46, Par83]. Predication [Glo76, Som76b]. Predicative [Nic16]. Predicativity [Boc15]. Predictability [Grü73b]. Prediction [Bru95, Gon96, Grü73b, Gün11, Joh17, Ram01, Sim77r]. Predictions [Pos76b]. Preemption [Eel10b]. Preface [Bee80, Mac97e, Mey85n, Ren07b]. Prefatory [Haq94e, Wol76y]. Preference [Ton81k]. Preferences [McA97]. Preferring [Rad79]. Prehistory [Eis03]. Prejudice [Aga13a, Aga13g, Aga13k, Han91, Nor07e]. Prejudices [Aga13m, Aga13n]. Preliminary [Aga08a, Bin96, Bla95h, Dam96i, DL05, Haq94g, Lia96a, Nee11h, Omo18b, Sch75, Wet95, DO91, Sch99c]. Prelude [Wal81q]. Premises [Mar67, PS15]. Preparation [Hel00i]. Prerequisites [Ton81a]. Prescription [Zol89c]. Prescriptions [Fol17]. Presence [Čap91h, Ela11r, Koh98d, MV80k]. Present [Aga81t, Ano08-35, Bou17, Col74, DP92, Ein07b, Fon96, For01, Jan85d, LM97, Liz96, MC80p, Mey16, Mor09, Zal86]. Presentation [Can07]. Presented [BM89, DPW86a, Lag97i]. Presentism [Coh16]. presently [dBJD95]. Preservation [Coo86]. Presocratic [Vam09g]. Presocratics [Vam09j, Vam09g]. Press [CS79n, Sun18]. Pressure [Car11]. Presuppose [Mil95]. Presuppositional [Her76]. Presuppositions [And79, Hod79]. Pretty [Sta07c]. Preunderstanding [Kel94]. Prevalent [Kow16k]. Prevented [Kap80q]. Preyer [Fle91b]. Priestley [Kra92a]. Priestly [McC02]. Prigogine [GP79, Näp01]. Primacy [Ela11m, Fin80l, Koh98h, Zol89i]. Primary [GWT00c]. Primate [Fen12]. Prime [Glo76]. Primitive [Bel12h, Gou92]. Primitives [Sim77f]. Primordial [Liz96, Rou16]. Prince [Wie95]. Principal [Nee11g]. Principia [Hes09, LP01]. Principium [War79f]. Principle [And09a, Bar07, Bel12c, Bel92b, Bou02d, Byc16, Dil06i, Hal91c, Jan84a, Knu95, Koh86, Lat83, Mey85o, Mie07b, NPS17, Nee11i, PS01a, Pfa84, Sch86d, Ueb97, Goe70, Nor07c, Nor07b, PS01a]. Principles [Ano06x, Ano09-48, Dam96i, Dil96, Dil06d, Dil06f, Gro82m, Har15b, Kap80-42, Lag97o, Lag97n, Lag97r, Lag97s, NT82, Nor98a, Per86, Sau98, Gar95b, Lüt06, Mul98, Nee11c]. Print [Csi15]. Printing [Cif14, Cul11, NG01a, NG01f]. Prinzipien [D’A01f]. Prior [Kat11g, Wan18c, Yuz96, Øhr97]. Priorean [Nee97]. Priori [Ano09-46, Nie82, Pan16, Per86]. Priority [Dev01, Ihd83, Zol89i]. Prism [CS79u, UM86c, Vod15]. Privacy [Sak98, Sch86a]. Private [Bro91a, Val10c]. Pro [Pit89]. Probabilistic [Eel10a, Fet10c, Hau10, Lec94, Mit97, Ray82, San15, Sch07d]. Probabilities [Bun12r, Fet71, Hin71, Sob10, Wei10]. Probability [Acc90, CL93, Cla89, Coh02a, Coh02q, Das91, Dil06j, Fet10d, Fet10a, Gei67, How76, Leg03, Nii03, San15, SM89, TG88, Tur86b, Tur86l, Van83, Van93]. Probable [Hac15b]. Problem [Alb84, Ams75i, And78, Ano07s, Ape99, Aya96, Bec91p, Ber17b, Bre09, Buc89, Bun10k, Bur87c, But93b, Can16, CS79t, CS79-39, Coh02m, Cop76, Den99h, Duh76e, Ein07b, Ein07c, Ela11d, Ela11r, Fax97, Fer85f, Fer85h, Fle86c, Fu93b, Gil09b, Goo77h, Gra83e, Gre74c, HR74h, Jan85d, Jar86q, Jar86r, Jim95, Jon86, Kal87, Kar10a, Kek17, Koe95, Kos94, Kre96, Kuz87f, KZ09, Lau83, Lea96, Lia96c, Mar69, Mar86h, MV80l, McC74, MC80q, Mis95c, Mit86a, Myr97, Näp01, Nic76, Nol03j, Now79a, Pan91g, Par83, Pat76a, Pfa96, Pol74b, Rut91a, Sal84, Sch97a, Sch86a, SF83, Sim77t, Sim77z, Sim11b, Skl74, Sol12a, Sta79, Syn82, Teg97, Tho96, Ton81e, Ton81h, Tur86a, Tur86l, Uye97, Voi96, War12, Wat78a, Wor78a, Wor89, ZRKC03i, Zol89f, dC95]. Problem-Solving [Gil09b, Pol74b]. Problems [Aga75h, Aga81c, Aga81u, Ams75a, And79, Ano08-39, Ass11, BBD18, BC81, Ber83d, Ber83e, Ber83f, BP14b, Bör93, Chi83, Cho76, Chr94a, D’A01n, Elk74, Erp91, Fer85e, Fle86d, Gem90, Ges74t, Gin18, Gio81, Goo77a, Gra06, Grü73s, HCP18, How97, Jar86i, Jar86v, Kap80e, Kap80m, Kap80n, Kap80-34, Kap80-33, Kap86a, Koe84, Kos76a, Kum92, Lag97a, Lag97f, Lag97q, Mar78j, Mar84f, MV80m, MS12b, Min10, Mit76g, OJ14, Pan83e, Pap94, Pet76, Ray82, Şak11, Sal74, Sch76a, SRI^{+}92, Sim77w, Sta83a, Sta07b, Str83, Str95, Swe90, Tha86e, Tok09, Ton73f, Ton81a, Ton81f, Ton81i, Vla94, Vol14, Was91, Wim71, ZRKC03j, Zim92, Zuo96, De 16, Kap80f, Kap80z]. Procedures [Suc82, Ton73d, Ton73f, Żyt82]. Proceedings [BC71, CW67, CW69a, CHMV76, SC74b, HFD88, SC74a, CW69b, FKR99, MS75b, War63]. Process [Aic15, Ano05g, Čap91j, Cha93b, CS79x, Fet10b, Kno05, Mit76i, Rot87i, Sei97, Sti74]. Processes [BE10b, Fin03, Von84a]. Processing [LE97, Ste99]. Proclus [AD94]. Product [Bra96, Fey95a, Heu96]. Production [Cah05, Kap80-35, Mar86f, Mar86h, TP90, Onn91o, PHL95]. Productive [Cre99, Mon80]. Productivity [Her14a, BBT95]. Products [BT96, Zad96]. Professional [But93j, DS18, DPW86b, Kar10f, Kov06, LvHV92, Löw86]. Professionalization [Ams92f]. Professor [Kap80-36, Sta83a, Bai17, Bur83, CS79q, Ito98, Sac83, dF14a]. Profit [San92a]. Profound [Jon91]. Program [Cus96, KJ14, Maj06, Pra93, Som76a, Stu76, Wet09, Mac01]. Programmatic [Krö91]. Programme [FT16, Mou89, Nol03c, Tor95, Urb78]. Programmed [Coh02a]. Programmes [Fei71, Gav89, Gra95, Kou89, Nic89b, Urb89, Wat89, Wor78a, Wor78b]. Programming [Atl86]. Programs [Ams75f, D’A76, Koc79, Per94, PP81]. Progress [Aga81-33, Aga03j, Aga03n, And79, Ban15, Bun12o, Cif14, Ger15a, Hil89, Kai96, Lef03, McM84, Mey85p, Mey85r, Mon01, Per84, RA78a, RA78b, Rad81, Rot87e, Sal01, Sch83b, Sch76c, Sha74a, Tym95, Váz96, Wes95, ZRKC03d, dF81]. Progressive [BT96]. Prohibition [Tho10].

Project [Aga81w, Bit96f, Cai14, Coh02l, For01, Mir14, Mun15, Nee97, Per89, Rai10, Sun18]. Projectile [Ano08n, Ano08-29, Ano08-30, Ano08-47, Ela11s, Pal04a]. Projective [Bri96]. Projects [Aga81p]. Prolegomena [Jan89i]. Proliferation [Pol03]. Proliferitis [Ach67]. Prolog [Sne89]. Prologue [Aga75l, Web83]. Prometheus [Gel95]. Prometteur [Pet92b]. Promise [Bun12l, RSS^{+}07b, Ueb91a, Urb78]. Promising [Pet92b]. Promotion [Lub15]. Proof [Bal18, BA09, BS16, Cas95a, Cel81, Kno04b, Mar86a, Neg81, Sol12b, You69, Wal92c, Abr81]. Proofs [Ana89, Sal18, Sau07, Tos81]. Propaganda [But93k]. Propagation [Cob74, Hon98]. Propensities [Fet10e, SM89]. Propensity [Eel10b, Wat83]. Proper [Coh02h, Wet95]. Properties [Bry98, Den05, Duh91j, Duh15f, Fre86g, Lag97g, Lag97h, Nee11j, Nee11k, vdA91]. Proportion [Bec04]. Proportional [Mat86]. Proportionenlehre [Bec04]. Proportionibus [Syl08]. Proportions [Duh02k, Nik74]. Proposal [Pos18]. Proposals [Och96, War79i]. Proposition [BC81, Ras00]. Propositional [Ano09-43, Fin80m, Gab76, Lep95, Mar78k]. Propositions [SPB18, Van95]. Prospective [Kau96]. Prospects [HN16, Man15b, Mar78i, MS12b, Rap83, Sch76a, BdOU^{+}95]. Protected [Rip11]. Protection [De 17]. Protective [AV97]. Protestant [Bre11]. Proto [Čap71-33, Rot86b]. Proto-Ideas [Rot86b]. Proto-Mental [Čap71-33]. Protolinguistic [Mar79r]. Protophysics [Jan84a, Jan85e, Mou82, Pfa84]. Protoscience [Tuo87]. Prototype [Bro96a]. Prototype-Based [Bro96a]. Prove [TV95]. Provide [Ano09-47, Gav96]. Providence [Zam92g]. Provides [Gje07]. providing [Tho95]. Province [Bin96]. Proving [DG15]. Prudence [Pil97]. PSA [BC71, CHMV76, SC74a]. Pseudo [Lai08]. Pseudo-Aristotelian [Lai08]. Pseudorationalism [Zol89d]. Pseudoscience [Tuo87]. Psycho [Jon86]. Psycho-Physical [Jon86]. Psychoanalysis [Are89f, CL83, Grü09, Jan89h, Mar87, Rui95a, Sab84, Sim11a, Tha84c]. Psychoanalytic [Eag83, Ric02]. Psychogenesis [Čap71-37, Čap91l]. Psychological [BN17, Cla12, DR16, FA76j, Gen71, Gro82l, Spr91b, Toe81, Wet95]. Psychologies [Are89j]. Psychologism [Cur89a, Wet95]. Psychology [Are89a, Are89c, Are89b, Are89d, Are89e, Are89f, Are89g, Are89h, Are89i, Bun12b, Cek92b, Erp91, Fin80n, Fle91b, Fra00, Gre84, Gru81, GB05, Lev87, Mai16, MC96, McR80, MC80r, Mit76a, Mül91b, Pos95n, Sas95, Sch12b, Sim77t, Sma92, War12, War79a, dR17]. Psychopathology [FA76b]. Psychophysical [Mar78l, MC80c, MC80p]. Psychophysics [Are89e, Kre91]. Ptolemy [Smi90]. Public [Aga03f, Aga03-27, Aga03t, BC94, Ezr01, Feh99b, HB76, Kap80-30, Mić79, Pes07, Wat01a, Wei11]. Publications [Aga81f, Alt14a, DPW86c]. Publicness [Pag00]. Published [HOO88b]. Publishing [Wol76z]. Punishment [Gro82j, Kle12]. pupils [Giu04]. Pure [Mor11, Onn91e, Onn91g, Onn91c, Onn91d, Onn91i]. Purely [Ano08-40, But93a, Wey07]. Purity [MA15]. Purpose [Ano05l, Ber14]. Pursuing [Aga74a, Aga81r, Mat01]. Pursuit [Fra99d, Lau87, Mar83a, San92a]. Put [War79q]. Putnam [Cof83, Grü69a]. Putting [Gol87]. Puzzle [Bec91o, Eva95, Ueb91a]. Puzzling [Kae90a, Pap74a, Phi90a]. Pyramids [Bun12k]. Pythagoras [Vam09m]. Pythagoric [Gar03].

Qi [Har01, Lih93a]. Qi-Shou [Lih93a]. Qing [Jam10, Bin96, Yal96]. Qingdao [Gua96c, PQQSe96]. Qinghua [Hon96b]. QM [Din96]. Qohelet [Fis95]. Quadratic [Vog04]. Quadratischen [Vog04]. Qualification [Goo77b]. Qualitative [GM89b, Hoo99l, Roq11, Smi89, Syl89]. Qualitative-Quantitative [GM89b, Syl89]. Qualities [CS79n, Haq94b]. Quality [Abb14, Caw89, Goo77l, Gra83f, Kow16l]. Quanta [Pos95k]. Quantification [Ber83g, Das91, GM89a, Man89, SF83]. Quantificational [Thi11]. Quantifiers [Lap95, Zin73o, Zin73u, Zin83a]. Quantisation [D’A01g]. Quantitative [De 81, GM89b, Gro82j, Mac12b, Mey85q, Smi89, Syl89]. Quantitie [Pos74, Sta74b]. Quantities [CS79-36, Kał82, Van74b]. Quantity [Caw89, Gra83f, Nee11m]. Quantization [CS79-33]. Quantized [MC80z]. Quantum [Acc90, ACE17, Alb96, AP97, Ana97, ÁNS91, Bal14, Bar91a, BS96, BC81, BB00, Ber17a, Bie82, Bit96e, BEW96, BD74, Bun10p, Bun12g, Bus97a, Car89b, Car74, Car00, CL93, Cat93, Che95, CdF95, CS79l, CS79m, CS79o, CS79-27, CS79v, CS79x, CS79-34, CS79-39, CS79-37, CS79-38, Coh80, CHS97a, CHS97b, Cre02, Cus93, CFG96, Cus96, DGZ96, Duw17, Enz00, Fin93, Fje02, For10, Fre11, GP79, Ghi93, Guc81, Hee74, Hei91, Hol96, Ion15, JS03, Kan00, Kar10c, Kar10b, Kar10d, Kat11a, Koç05, Le 08, Mac80, Mac12d, Mac74, Mac12e, Mac12f, MPC91, Mar74b, Mau96, McC95a, MC80n, MC80z, Mit76f, Mit76h, Mit76i, Mit97, Mit00, Nag84b, Nav15, Pag96, Par15a, Pat03b, Pau00, Pos95d, Pos95o, Pos95p, Pos95q, PW17, Pro00, Rad12e, Ray82, Ren96]. Quantum [Ric12, Rös91, Sam96, Sch63, SB15, SS03, Shi96, Sie63, Sta74a, Sta76a, Sta97, Sta94c, Ste97, Stö11, Unt01, Van74c, Wüt16, Zab95, Zwi08, dCB11, dB84, dB87b, dB87c, dB87a]. Quantum- [Wüt16]. Quantum-Mechanical [Mit76i]. Quarks [Fal00]. Quasi [Bus97b, Gol03, Sør17, Spr91b, dCB11]. Quasi-crystals [Sør17]. Quasi-Experimental [Spr91b]. Quasi-Linear [Bus97b]. Quasi-Local [Gol03]. Quasi-truth [dCB11]. Québec [MC95, MC96]. Queries [Nii96]. Quest [Pap74b, Rou95, Sha84m, She82, She87]. Question [Bab02b, CS79-36, Dav99, Dup92, Elg84, Gia99, Oli16, Pes07]. Questions [Aga75m, Čap71-39, CdF95, CS79-39, Coh02u, Dur83, GWT00d, Lor07b, OK00, Pap74a, Pol95, Ros09, Wal92e]. Quetelet [Tur86i, Por94]. Qui [Loc97]. Quine [Ber00, Čap71-28, Dav97, Geo00, Gib00, Grü63, Grü86, Leh00, Lev87, Mur12a, Mur12f, OK00, Ore00, Qui00b, Ray00, Sto00, Szu00, Yal05]. Quinean [Pan91d]. Quining [Mis00]. Quoque [Aga81v]. Quotation [Tho76]. Quotations [Bla95h].

R&D [Mar01]. R. [Pro92, Rub88, Sei92]. Rabbi [Lan91, Wie95]. Racial [Ban95]. Radical [All01, Ams75d, Ams75e, Ant00, Kra92a, Lev71, Mar78m, Pan83b, Roe01, Sha01, Vil02e, Whi88]. Radically [Gau95]. Radicals [Rut15]. Radio [Bry98]. Radioactive [Wat01a]. Radioactivity [dAM11]. Radnitzky [And84b]. Rahn [Scr74a]. Railway [Wan18a]. Rain [Par09a]. Raising [Kiv01]. Ramírez [Rui95b]. Ramsey [Maj98, Sim77s]. Ramsification [KN01]. Random [Gal00, Ges74u, Goo74]. Randomness [Cla89, Cof74, Goo74, Kyb74, Sal74]. Range [Pol95]. Rapport [Lid89]. Rapports [Bea75]. Rapture [O’D02]. Rare [TB15]. Rare-Earth [TB15]. Rarefied [Hie95a]. Rate [Onn91g]. Rates [Tur86i]. Rather [Čap91o]. Ratio [Moo90, Par09c]. Rational [Alb84, Ara94, Arm95, Ber95a, Bre09, Bun95b, Das88, Das07b, Els88, Izq74, Joh17, Kau71, Kau76, Koe76, Koh94, Lak71a, Lao95, McM78, Mey85p, Mey85r, Nic80d, Nol03m, PP87, She04, Sim77l, Sza79b, War74]. Rational-Choice [Das88, Els88]. Rationalism [Aga95b, Aga03-33, Alb84, And84b, And09a, Ban17, CS79-40, Dem89, Fey84b, Fra86, Han88, Jan89a, JL95a, JL95b, Kuz87l, Lan67, Mil09, Ng17, Qui82, RB09, Sal84, Seg17, She09, Str84, Top84, Wil18, Zem09, Zol89b]. Rationalist [Far03e, Nol03k]. Rationalists [Mar78d, Mat17]. Rationality [Aga80, Aga81i, Aga81v, Ams82, Ams92b, And84d, Ano08-41, Bou03, Boy09, Bro96a, Cat95, DJ89, D’A95, D’A76, DD86, Far03b, Far03k, Fet81g, Fin80o, Fin09d, Gar16, Góm15b, Gon03, Grü76, Jar76, Jar86t, Jar86s, Kas76a, Koe93, Kou11, Lat83, Lau87, Lev94, McM74b, Mey91m, Mey91n, Mic74, Mis95b, Mis95c, Mit15, Mot82, Nic80f, Nic87, Per87, Pit87, Pła95, RA78b, Rad81, Rou15, SdV92, Sco95, Sha74a, She87, Sie01, ST69, Sko76a, Tam01, Tou74a, Tsi90, Tur86j, Vil02b, Wat87, Wet09, Zel74, Życ95b]. Rationalize [Enç05]. Rationalizing [Aga03o]. Rationally [Nic80a]. Raumbuch [Car15a]. Ravens [FH10a]. Rays [Kox95]. Razor [Ore00]. Re [Bar09, Čap71c, Fen12, Gar16, Hod90, Whi99, Wel99, Ina05, Ray00]. Re-appropriation [Gar16]. Re-Conceiving [Fen12]. Re-evaluation [Čap71c]. Re-Examination [Hod90, Whi99, Bar09]. Reaching [MA14]. Reaction [Duh91k, Kap80-32, Kar10g, Nee06, Rhe05, Zal86]. Reactions [Fel92, Lev83d, Rut15, Vic92, Wat92a]. Reactive [AB12]. Reactivity [SH91, TBC91, vSB91]. Reader [Wes95]. Readers [Gar95a]. Reading [Cha14, Che10, Gar95a, Gav94, Ges74b, Mar79p, O’C99, Rit05, Tam02]. Readings [Har08a, Rad03]. Real [Bun10p, Ela11r, Jar86r, Kle91, Mar95b, Mey85d, Mey91m, Mey91n, Nee97, Par09a, Per94]. Realism [Aga97b, AP00b, Alb84, Bah96, BS16, Bra74a, Buc89, Bun95a, Bun12g, Bur84, Cas00, Cha12a, Cha12e, Coh88a, CHR96, Con99, Cor00, Cus93, D’E00, Dai96, Dil06b, Dil06c, Din96, Enz00, EL11, Fav94, Fer85g, Fey94, Fol94, Fol96, FL11, Gav96, Gly71, Hil96, Jar83, Jar86u, JX96, Joh96a, Kan00, Kre96, Kri94, Lau09, Mac94c, Mac12h, Mäk96, Mas11, Mas15b, Mun15, Mus95, Mus96, Nic89a, Nie82, Nii96, Pin11, Que00, Rad12d, Rad12f, Ren96, Rut15, Sch00, Sha96, Suá08, TB96, Tis74, Tor00, Tur86k, Ven16, Vot11, WP96, War79i, Wat96, Wri96, Zem88]. Realism/Antirealism [Kre96]. Realist [Bye74, Coh02b, Sch96b, Smi95a, Tul16, War91]. Realist/Anti [Tul16]. Realist/Anti-Realist [Tul16]. Realistic [Goo77e, Lin82, Sam96, Shi96]. Realists [Woo02]. Realities [Näp01]. Reality [AP00b, Bau88, Bec91k, Bon00, Bro01, Bun12l, Čap71-34, Čap91m, Cha12a, D’E97, Dor97, Eck17, EE17, Eul76, Fle86b, Gas76a, Gas76b, Gau99, Gol79, Gre74c, Hol70, Jaf02, JX96, Kuu01, Le 08, Li96, Mar84a, MC80s, Mey85s, Mey85t, Pau97, San97a, SM96, Sha95, Sob10, Wal67, Wul96, Zwi08]. Realization [BN17, MV80a, Rad12d]. Really [Aga81-35, Qui74, Sta92a]. Realm [Ano08-32]. Reappraisal [AZW87, Chm82, Gie76, Sar92a]. Reason [Aga75b, Aga03k, And09a, Ano09-47, Ber86, But93c, FC87, Fey76, Fin80g, Gar16, Han99a, Kek91, Koh98h, Kop99, KM99, Kuz87n, Mai16, Mey85i, Mey91k, Mun91a, Nol03h, Nol03k, Rap82, Sha84m, Sha84l, Tau99, Tay92, Tha86b, Tha86e, Wei95, Wol76x, And91]. Reason/Emotion [Mai16]. Reasonable [Hoo91]. Reasoning [Fin80a, Fin80f, Fin80h, Fin80k, Fin80n, Fin80l, Fin80m, Fin95, Fin09b, Hac94b, Hac15b, Har97a, Har89a, Ihd99b, Mac86b, Mon80, Pea86, Pie16, PP87, SS03, UM86b, Wal76, Wal81i, vL16c]. Reasons [Ano09-46, Ano09-47, Coh02b, Fin09j, Har08a, Ita92, Rad79, Sha01]. Reassessing [Pos95r]. Reassessment [Aga08d, Ana89, Boc15]. Rebellion [Wol76-34]. Rebuttals [Wat92a]. Recapitulates [Fin87]. Recapitulating [Oya92].

Reception [Bie87, Fre09, Gli87b, Gli87a, GPSR01, Gli01, Hsü92, Kar10e, Mul98, Nád96b, PT04, Pat87, Ree87b, Row96, SR87, Sre87, Tob96, Ueb91a, VG87]. Receptions [GH01]. Receptors [SA12]. Recettes [Loc97]. Recipes [Loc97]. Reciprocal [Lim94]. Reciprocals [dB00]. Reciprocity [Gou83]. Recoding [Gro82a]. Recollection [Alt14b]. Recollections [CS79-38, Gre76b, Kap80-37]. Recombination [Cha93b]. Recommencements [Ras75]. Reconciled [D’A98]. Reconciling [Att13]. Reconsidered [Gal86, Mey91n, Sha88]. Reconsidering [Tso15]. Reconstruction [Fer85h, Gra95, JN79, Pel95, Sot03]. Reconstructions [Dea12, Koe76, Lak71a, McM78, Sha86b]. Recording [SG11]. Recovered [dB87c]. Recovery [Wol76-27]. Rectilinear [Ela11e, Vil08]. Recurrence [Čap91e, Sam76a]. Recycling [Fra99f]. Red [Wyl92]. Redefinition [McC74]. Rediscovering [Ier90, Ueb91b]. Rediscovery [Dah90]. Reducibility [Gre74p]. Reducing [Har89c]. Reductio [Far03e]. Reduction [CŻ82, Elg84, GG96a, GW97, Het15, Hul76b, Krü89, Neg81, Nic76, Pos95l, Ran89, Rus76a, Sar91, Sce93, Sch97a, War79o]. Reductionism [Ano08-36, Ara94, Ber86, Bun93j, Cao98, Cha15, Kau74, Krü89, Mac12h, Mar82, Sch76a, Sch76b, Băe15]. Reductionist [JS03]. Reductionistic [Fab84, Wim80]. Reductions [Skl76]. Reductive [Buc89, Hei94a, Sau03]. Redundancy [Har97b]. Redux [Rus82]. Reed [Hoo99i, Shi87c]. Réelle [Čap91n]. Reexamined [Nag71]. Referee [Aga81x]. Reference [Aye76, Boy01, Buc84, Com84, Duh76a, Gea76, Kor71, KN01, LMR95, Mar79n, Mar79s, Mur12d, Neu76b, Rad12f, Sha84m, Zam92j]. Referent [AP97]. Referential [Rad12d]. Referentiality [Aga00b]. Refined [Ano06x, Dil06d]. Reflection [AV16a, ML99, UM88]. Reflections [Ams82, Bor86, Buc98, CS79-42, Coh17a, Coh17b, Cul11, De 16, Del91, Den99i, Emb92b, Gie71, Gol89, Hie07, Hon16, Kay01, Koc79, Koh98h, MZ18, Moc91, Ols98, San08, Shi83, Shi86, Sim94, Siv84, Ung04a, Yin18, Zem88, Bea75]. Reflective [Vil02i]. Réflexions [Bea75]. Reflexive [Zol89g]. Reflexivity [Zol89d]. Reform [Bar74, Fra91, JX96, Kow16h, NG15b]. Reformation [Ams75f, Obe75, Ols98]. Reformer [Sta91c]. Reforming [Sas03e]. Reforms [Aic15, Vod15]. Reformulation [Gre94]. Refrigerated [Ree14]. Refrigeration [dS14]. Refutability [Ros09]. Refutation [Wis69]. Refutations [Ana89]. Refute [Bun12g]. Refuted [Coh16]. Regard [Buc01]. Regarding [Mar78j, Nei15, Res79]. Regeneration [Mar11b]. Regenerative [Mai12]. Regiomontanus [Swe90]. Region [Ges74m, BE10a]. Regional [Syl97a]. Regress [Fra99e]. Regressive [BT96, Zad96]. Regulae [Sas03h]. Regularities [Den99c]. Regularity [Suc74]. Regulate [Byr83, Rod15]. Regulation [Cac82, Sta79]. Regulative [McC02]. Reichenbach [Čap71d, Čap91b, Grü73h, Hei11, IG05a, Kam11, MC80u, Pad11, Stö11]. Reification [Wei01b]. Reified [Tad79]. Reinforcement [Gro82i]. Reinhard [Gon03, Sca03]. Reinterpretation [Čap71c]. Reinventing [Wag15]. Rejection [Čap71-35, Ela11s, Fet81a, Ita92]. Rejoinder [MS83]. Related [Duh02m, Hoh83, HHS01, MPC91]. Relating [Lie76, Wes91]. Relation [Art89, BM89, Cha15, CS79u, Cro99, Fra96b, Gen71, Gra07b, Lev16, Ley84, Mar78n, Mes90, Nik74, Øhr97, Rou15, Sim77p, Syl08, War76, War79p, Wat63, dC95]. Relational [Bai76c, Bru76, Gan15, JX96, Plo76, Smi95b]. Relationalism [Bel12h, Sau03]. Relations [Aga03t, Bus97b, Čap71-31, Dem05, Fer85b, Gou83, Hef87, Ihd79c, McK75, Mug81, Pul11, Smi95b, Vog91, Zin73m, Pop88]. Relationship [Bea75, Dam96i, Dam96g, Die14, Dor97, Fre86r, HB76, KB92b, Kuh87, Len63, Löw86, Näp01, Sch89, Shi98, Sim99]. Relationships [BB00, Lam14, Li18, Mar84j]. Relative [CM89, Föp07, Fri07a, Mal03, Pap74b, Pos95a, Str18]. Relativism [Bar16, Far03g, Mas15b, Mey85a, Mey85u, Mey85t, Noé92, Rad12f, Rad76, RA78a, San97b, Sie01, Tsi90, Yal05, Jar86s]. Relativistic [Čap71-30, CŻ82, Gra95, JM06, KT93, MC80t, MC80u, Mey85v, Mey85w, Nor07a, RL97, RDS03, Ren07c, Sch07a, dB87b, dB87c]. Relativists [But93j]. Relativity [Ano07g, Bel12c, Bel12i, Bos76, Bou17, Bro93, Bud93, CR81, Čap91r, Cha93a, Coh16, D’A01d, D’A01e, Dol16, Ein07a, Ein07c, Eis03, Gia71, Gli87b, Grü73c, Grü73k, Grü73l, Grü73q, Grü73o, Grü73r, Grü73u, Guc87, Har97b, Hav69, Ita92, Jan84a, Mal03, MC80v, Mie07b, Min07, Pap74a, Pas76, Pat87, Pfa84, Pos95d, Pye87, Ree87b, Row06, Sac67, SR87, Sch97a, SS03, Sim11b, Sre87, Sta83b, Sta94a, Tor83, VG87, Wie92, Yil05, dB87j, Ano07e, Ano07r, Gli87a, JNR^{+}07b, JRN^{+}07b, JNR^{+}07a, Mey85v, Nor07c, Nor07b, RSS^{+}07a, RSS^{+}07b, RS07c, Ren07d, Sta69]. Relevance [Coh02s, Fet81j, Gie71, Jar09, Mar76c, Nik15, Rop86, Sal71]. Relevant [Zuo96]. Reliability [Ned12, Wim12b]. Religion [Aga75b, Aga81o, Alb95, Bal02, Bar02a, Boš79, Fel18, Fin09h, Fri74, Gen91, Jar86e, Kur88, Lao95, Mah75, Rou95, Set95, Val99, Wol76-35, GSW95b]. Religious [Mit83, Sta17, Vil02l]. Relocating [ARS15]. Remark [Gan15]. Remarks [Apo82, Čap71h, Coh02q, Coh02r, Dam96h, Dam96i, DL05, Gar12a, Gar13, IG05b, Jef71, Jim95, Kap80-38, Kar10i, Ken96, Kri03, Łod82, Lon93, Mie07c, Mor18, Mul74, Onn91p, Pos95j, Pra93, Ros98, Sam86, Sch96g, Sha84n, Str84, Str74, Wes95, dB87e, Onn91f]. Remedied [Wet09]. Remedy [Bro11]. Remembering [May95, Sha95]. Remembrances [Ber74b]. Reminders [Mar79s]. Reminiscences [CS79e, CS79-32, Edm95]. Renaissance [Mar13, Aga81-34, Cif14, Gai94, Kno89, Kuz87l, Mar12, Nen18, Omo18a, Vic92]. René [Duh91e]. Renewal [Mün14]. Renewing [Ara12]. Reopening [Str94]. Reorganization [Gro01]. Reorientation [Are89b, CN11b]. Reorientations [Rhe09, Rhe11]. Repeated [Şen05]. Repercussions [Pel01]. Repetition [Ano05g, GM17]. Repetitions [Ras75]. Replace [Mei17b]. Replies [Aga75n, Bre74c, Lak71b, Wor78a]. Reply [Čap87b, Grü69a, Grü73v, Hüb78a, MS83, Sac83, Sch83a, Sch80b]. Report [Mar76d, Wet95, Zim92]. Reports [Coh94i, Ges74u]. Represent [Vol06]. Representation [Ano08-39, AB17, Bit96b, Cas00, Cha10, Cla96, Clu05, Dam96a, Dam96b, Dam96j, FH10b, Fri10, Hen01, Hun10, Kiv94, Mül91a, Pos74, Smi89, Sta74b, TJ11, Van74b, Vol06, War79k, War79m, War79n, War79q, War86, Kle01b, Hin71]. Representational [Kom99]. Representations [BV01, EN96, Gen03, GC94, PS01a]. Representing [MW76]. Repression [NG15b]. Reprogramming [Pos95p]. Republic [Ano08-45, Dor92, NG01b, WC91, Yuz96, Kar90, San90]. Republican [BD05, Yan93]. République [Dor92]. Repudiated [Grü73l]. Repudiation [Lan00]. Requirements [Hav69, Sta69, Str15, Wal92a]. Rereading [Fin17]. Rerum [Ste12]. Res [Van86]. Rescuing [Nol03k]. Research [Aga81p, Aga81w, Ams75f, Ams92f, Ano08-45, BJ95, Bel15, Bid11, Bie74, Böh83, BG09, Bro95, Buc91, CF11, D’A76, Deb09, DCS15, Fag12, Fei71, Flo15, FG92, Gav89, Ger15b, Hei08a, Hei08e, Hei08f, Hod74b, HSS09a, IdMM16, Kap80z, Kap80x, Kar10d, Koc79, Kou89, Lub15, Mai12, Maj06, MS11, Mar90, Mon80, Mor09, Mou89, Nic89b, Ols82, Pap15, Per94, Pes91, Rad81, Rol16, SdV92, Sch12a, Sti94, Stu76, Tam14, Urb78, Urb89, Vel14, Wal95, Wat89, Wei14, Wim80, Wor78a, Wor78b, Hoh83]. Researches [Čap71-37, Čap87a, Čap91l, OP91a, Uye97, WO91, Zam92n, OP91b]. Reserve [Hei08b]. Residual [Kox95]. Resistance [Chu11, Det07, Eck17, Onn91e, Onn91g, Onn91c, Onn91d, Onn91i, Onn91h, Onn91j]. Resolution [Coo90, Fra99i, Grü73t, HR74h]. Resolving [Pan91k, Sch96b]. Resonance [Mir14]. Resonances [Flo17b]. Respect [Bun17, Her74, Sie82]. Respective [Hee74]. Response [Alb92, BF94, Cus94, Hem76, Koh98j, Lev83a, Lev83b, Wol76-37]. Responses [Qui00b]. Responsibility [Aga03f, Bar01b, BN17, Ben15, Len83, Zim86]. Responsible [Bec86]. Rest [Sch07b]. Restoration [Ols98]. Restraint [Saa17]. Restricted [Coh86a, KP86]. Result [DS12, Har89d, Tha86f]. Resulting [Tho10]. Results [Fra99i, Nei15, Par09a, Sol12b, Ton81c, Bie74]. Retarded [dB87q]. Rethinking [PC09, PY07, SSHH08]. Retinal [Gro82h]. Retracted [Kox95]. Retreat [Sch88]. Retrodictability [Grü73b]. Retrograde [Ges74x]. Retrospect [Wat89]. Return [BC02, Cas95b, Sch88]. Returns [Res97].

Réussissent [Loc97]. Revelation [Ano06-28, Cra93, ZC88h]. Revelations [Xiu96]. Revenge [Maa17]. Reverberating [Gro82f]. Reversibility [Fay97a, GM17, dB87g]. Reversible [Nee11a]. Review [CS79g, CS79n, CS79-41, Ges74v, Goo95, Gre74j, HH89, Kat11a, Kil05]. Reviews [Bla92b]. Revising [Aga81x, Nar09, Tra15]. Revision [Fer85e, dAM11]. Revisions [Mur12e]. Revisited [Ayd05, BVH83a, Bun93d, Coh86e, Dil06c, Fow94, FA76f, GL03, NG15a, Sey96, Vis03, Wro88]. Revisiting [Fol17, RDS03]. Revolution [Aga76c, Arb92, Bec91r, Bel15, Bor86, Bra95, Cha12g, Coh76a, Coh94c, Cos79, Duh02p, Duh02o, Fre86n, FM09c, Grl79, Koe80, Koe86, LR08, Len86, Lev90, Nel67, Obe75, Ols98, Pet79, Pos95n, Pye87, SB15, Spa03, Win86, Žva79, Bar01a, Fin09d]. Revolutionary [McA97, Rus79, Tor99]. Revolutions [Aga81y, Ams75h, Ams75g, Bar02b, Ber74d, DB15b, Gie97, Koc91, Min67, Pan83c, Ric15, Sha84r, Tou67, Tou86a, Tri08]. Rewrite [Ung04c]. Rheticus [Ros74a]. Rhetoric [Bod18, Fin80p, Jan89k, Seg15]. Rhetorical [Fay07, Fin80d, Fin80g, Fin80h, Kop99]. Rhizobia [BBT95, WPJ95]. Rhode [Pop88]. Rhythm [Rot87f]. Rhythms [Čap71k]. rice [BLA95a, BdOU^{+}95, dBJD95]. Richer [Bur83]. Riddle [Aga13e, Aga13o, Ren07a]. Riemann [Bot94, Grü73w]. Riemannian [LC07, Mit76d]. Right [Fin09j, Fra99c, Fre09, Gav09, Ges74m, Hor88, Hor94, Jos17, Kam11, San97a, Sch95a]. Right-Wing [Hor88]. Rightness [Koh98c]. Rights [Byr86, Gel89, Hun83, Koh98c, Reg13, You95]. Rigidly [Vis03]. Rigor [Gio81]. Rise [Aga13e, Bar14b, Fra93, Fra99j, Fre86s, Kay01, Maa17, She95a, Sie86, SS74, Smi97, Sta74e]. Rises [Ren96]. Risk [Góm15b, Rod15]. Ritter [dAM07, Str94]. Rival [But93f, Cor00]. Rivalry [Rus13]. River [BE10a]. Rivetti [Bur83]. Road [Ano07a, Kat06e, Kir15, Mac94b, Rap81d, Sig07]. Robb [MC80t]. Robert [Aga13q, BM89, Cha09b, Chr14, DS11b, Edm95, GSW95a, GSW95c, GSW95b, Hod74a, Hun10, Mue95, Pec96, Sch96e]. Roberval [Duh91d]. Robinson [Gan15, Vil02i]. Robotics [Pfa96]. Robust [Yal05]. Robustness [ARG12, Boo12, DS12, Nic12, Pic12, STNW12, Sol12b, Ste12, Tri12, Wie12, Wim12a, Wim12b]. Rock [Ihd79a]. Rodney [Sim94]. Rods [Sta83b]. Roebling [Kah91]. Roger [Whi99]. Rohault [Sch01]. Rojas [Bar14b]. Roland [Imo15]. Role [Bar76c, Boo12, Bre74a, Bur74b, CALP03, Cha14, Cho14, CM89, Cro92a, D’A01b, D’A01k, Dam96e, Est96, Feh99b, Für69, Gil91, Gon15b, HL17, HR74i, Imo15, Kap80d, Kap80-39, Kro92, Lat83, Lep80, Lia96b, Löw14, Lub15, Mar15, Mar97b, Sce06, Scr74b, Sup74, Tho95, TP90, Vla90, Zam92j]. Roles [Hee74, LC17b]. Rolf [CS79g]. Roman [Kar17]. Romania [Pǎr15b]. Romanian [PST15]. Romano [Wal81c]. Romantic [BCK07, Bra07, Jah94, Luc88, Ric15, Str94, Win86, WP94, Pog94]. Romanticism [Abb94, Det07, Jac07, Lev07, PB94, Sne94]. Rooster [Wal84]. Root [Tau91e]. Roots [Aga75h, Bur74a, FM09c, Gel84, Hes09, Kam11, Mur12d, Pop86b, YCMY18, ZRKC03l, ZRKC03m]. Rorty [Jan89f]. Rosen [Nag71, Web93, Zuo96, Alb92, Cos87a, Har97a, dB87n]. Rosenfeld [CS79a]. Rosenstiel [Hei08c]. Rosettes [Bun12k]. Rotating [Vis03]. Rotation [Mal03]. Rotational [Lag97p]. Round [CLR^{+}92, Ned90a]. Rousseau [Fre86r]. Route [Kou90, Rod17, Row14]. Roy [SL84]. Royal [Aga81s, Aga13j]. Royce [Čap71u]. Rubbish [Aga03g]. Rudolf [BC71, LM97]. Rudolph [Maa17]. Rui [Hor93]. Rule [Ach80, Cel08, Mir15, Okr89, Pis15, Sma92, Şte15b]. Rule-Books [Ach80]. Rule-Following [Pis15]. Rules [Ams92e, Bun12q, But93e, But93m, Far03i, Fet81a, Flo15, Gro82c, Hat83, Hüb76, Mar92, Mic71, Tra15, War79q, Wei10, Gro86]. Rumford [Sch96a]. Run [Grü95b]. Ruptures [SSHH08]. Ruse [Życ95a]. Russell [CW67, Art89, Čap71-40, Čap91k, Grü73w, MC80w, Pel95]. Russella [Gar95b]. Russia [VG87, Vod15]. Russian [GG03]. Ruth [Mar79j]. Rutherford [Hon98, Kap80p, Kap80r, Kap80-37, Kap80-41].

S [GSW95a, GSW95c, GSW95b, HOO88b, Mar97b, Mue95, Škv88, Smi92, Sau02]. S. [Sch12a]. S.J. [Bab02c]. Sūtra [ZC88c]. Sînâ [Fre91]. Saccheri [Bağ05]. Sacrobosco [Ros74a]. Safe [Mar95b]. Sagal [Aga87b, Shi87d]. Sagnac [Pos95f]. Sahasri [Ano06-32]. Salamander [Gri15]. Salons [Spr91a]. Sam [PB07]. Same [Wet09]. Samos [Vam09m]. Samuel [Hel74, Sch01]. Sanskrit [Fil05, Kel10, Min10]. Santiago [Rui95b]. Sao [Pet92a, Pet92a]. ‘Śariraka [Ano06a]. ‘Śariraka-BhāSya’ [Ano06a]. Sarkar [Tau91b, Tau91a]. Satisfaction [Mar79s]. Satisfactory [Băe15]. Satisficers [Mic74]. Savants [Dol03]. Save [Byc16]. Saving [Mus89b, Sep87]. Savoir [Gre75]. Saw [Pur93]. Saxony [Duh91c, Duh91j]. Say [Coh02o, Sta92b]. Says [HR74j]. Scalar [Nor07d]. Scale [Hun17, Tur86f]. Scaled [Ren07c]. Scales [Ber83h, Ber83i, Ber83k, Les17]. Scandal [Fre86a]. Scarcity [Mac97f]. Scattering [Dau96]. Scent [Hof09]. Scepticism [Mus89b]. Schaffner [Mau80]. Scheffler [Lev83b]. Scheler [Tor99]. Schelling [Hah91, Mey91o, Nol91]. Schemata [But93e, Koe89]. Scheme [Höf76, Jar95]. Schemes [Ano09k, Čap71-38]. Schleiden [Jah91]. Schlick [Avg90, Pul06, Rut91b, Sta91b, vH77a]. Schmoller [Kóh90]. Scholar [Ber84, Sch96f]. Scholarly [Cla10, EP17]. Scholars [Pat03a, Pat11a]. Scholarship [Fin80e, Har01, Räz16]. Scholastic [Ela11i]. School [Car15b, Cha18, Coh94d, Duh02o, Hof98, Hül96, Kap80-29, Sta91c, Tob96, Vil02d, Ara94, CW18, Coh94f, Duh91l, Duh91p, Duh02p, Ges74e, Giu04, Pul11, Zam92e]. Schools [Aga81-27, Pro14]. Schoolteachers [Bar74]. Schopenhauer [Hör95, Meš89, Mes90]. Schröder [Pec96]. Schrödinger [Bit96d, Bit96e, Bit96f, D’A01g, Shi83, Shi86, Woo06, dR01]. Schroeder [Tra15]. Schwarzschild [Sch07a]. Schweber [PB07, SR07]. Schweberian [Gor07]. SCI [Wer01]. Science [Ada01, Aga67, Aga69b, Aga75d, Aga75k, Aga75m, Aga75o, Aga75p, Aga75t, Aga81a, Aga81d, Aga81e, Aga81g, Aga81j, Aga81l, Aga81m, Aga81t, Aga81y, Aga81-30, Aga81z, Aga81-34, Aga95a, Aga03c, Aga03k, Aga03n, Aga03l, Aga03m, Aga03p, Aga03x, Aga03r, Aga03q, Aga03s, Aga03y, Aga03-29, Aga03z, Aga03-27, Aga03-30, Aga03t, Aga03v, Aga03u, Aga03w, Aga03-28, Aga03-32, Aga08a, Aga08d, Aga08e, Aga13p, Aga13q, Alb78, Alb95, All01, AM01b, Alt87, AV16a, AV16b, AZ87, Amr89, Ams75a, Ams75e, Ams75h, Ams75g, Ams82, Ams92b, Ams92c, Ams92e, Ams92h, Ams92j, And84a, And84d, And09b, Ano06x, Ano08-46, Ara12, ARS15, Ara79, Arb92, Are89h, Atk03, Aya76, Aya96, Bab97, Bab99c, Bab99a, Bab99d, Bab02c, BD05, Bal02, Bal89a, Bar15, Bar87, Bar74, Bel81a, Bel90, Ber74c, Ber95a, Ber84]. Science [BS15, Bid11, Bir15, Bla92b, BIT01c, Bla94, Bla04, Blo86, BS83, Böh87, BMM17, BCK07, Bre07, BG09, Bre09, Bre74a, BE10b, BM89, Bru07, BC71, Bun12r, Bur87a, But93b, But93h, But93i, But93a, CG16, Can16, Can12, Cao95, CALP03, Car11, CN11b, CN11a, Cau76, Cau80, Cha10, Cha09l, Cha09d, Cha09a, Cha93b, Cha12f, Cha18, CLR^{+}92, Che05a, Che05b, Che14, Che97, Chi81b, Cho14, CM89, Chu00, CW69a, CW69b, Coh70, Coh74b, Coh76d, CHMV76, CS79-35, CS79-41, Coh80, CW84b, CW85, Coh86c, Coh94m, CHR96, Coh17b, Col74, Coo84, CC90, CCG93, Cro99, D’A89, D’A95, DC96, Dan92, Dav10, De 17, De 01, DR15a, De 81, DP92, Dea12, DN94, Dil96, Dil06b, Dil06d, Dil06f, Dil06g, DCS15, Dov84, Dra99, Duh02n, Dvo91, Eas84, EPL69, Ege99, Elk76]. Science [Elk86, Elz95, Eng01, FL09, Fax97, Feh99b, FKR99, Fel18, Fet10a, Fey84a, Fin91, Fin80j, Fin80k, Fin80o, Fin80r, Fin95, Fin09h, Fit76, Fle86a, Fle86d, Føl99, FS96, Fox07, Fra99c, Fra99g, Fre86l, Fre88, FM09b, Fre99, FH10b, Fu93b, Fun86, Gäh89, Gai94, Gau99, GSW95a, GCN94, GSW95b, GR07, Gie71, Gie12, Gin97, Gin03a, Gin03b, Gin07, Gol12, Góm12, Góm13, Gou76, Gov15, Gra06, Gre91b, Gre07, Gre99, Gre83, GB05, GHQ96a, GHQ96b, GR11b, Gur03, Hal84, Hal71, Hal82, Hal91b, Hal92, Hal95a, Han99a, Haq94b, Har99, Har02, HGR^{+}89, Har84b, HH91, Hee99, Heg87, Hei98, Hei97, Hem83, Hem86, Her14b, HH97, Hes07, Hie94, HR74e, Hod79, HB76, Hol95c, Hoo99c, Hör89, HFD88, Hum11, Ihd83, Ihs92, İhs05, IG05c]. Science [Irz15, Jam10, Jan85d, Jan84b, Jan89h, Jan03a, Jar76, Jar86b, Jar86k, JL95a, Jar17, Joh11, Jol18, Jur08, Kai96, Kal90, Kap80b, Kap80c, Kap80d, Kap80l, Kap80w, Kap80u, Kap80-31, Kap80-39, Kap80-43, Kar17, Ker88, Kil05, Kim96, Kir82, Kit90, Kno89, Koc97, Koe71, Koe74, Koe84, Koe91, Koe95, Kóh90, Kov06, Kra07, Kra74, Kra95, Kra92a, Kra92b, KV11, Kri92, Krö91, Kro89, KB92a, KB92b, KS83, Krü88, Krü91, Kum92, Kur98b, KT01, Kuz87f, Kuz87n, Lac01, Lak71a, Lau74, Lau09, Lay92, Led18, Lef96, Lef01a, Lel00h, Lel00i, Lel00j, Lel00n, Lel00l, Lel00m, LS90, Lev83d, Lia96b, Lia96c, LF93, Lin82, LMZ18, Löw86, Mac00, Mac01, Mah75, Mah11, Mai01, Mai12, Maj06, Mal92, Mam94, Mar86b, Mar11a, MC95, MC96, Mar76c, Mar84p, Mat95, Mat89]. Science [Mau86, MS12b, Max76a, May02, McL71, McM74b, McM76a, McM78, McM84, Mcm76b, MC80q, MC80z, Mei17a, Men74, Met89, Mey76, Mey85f, Mey91q, Mey91p, Mey91r, Mil12, Min67, Miś82a, Mis95c, Mit76a, Mit89, Mor89, Mou92, Mue95, Mun91b, Mun96, MS75b, Mus89b, Müü01, NC98, Nak74, Nel67, Nem91b, Nic76, Nic80c, Nic90, Nie82, NG01e, NG15a, Nii89, Nol99, Nol03k, Nor11, Nor81, NNEH11, NRS92, Nye07, Nye15, O’C99, Ohm98, Osb92, Ott94, Owe99, PRT15, Pal01, PT04, Pal76, Pan91b, Pan91e, Pan91f, Pan91h, PST15, Pat90, PJM92, Pic12, PM82, Pin96, PP87, Pla94a, Pla94b, Pol92, Pol74a, Pol95, Pop74, Pop86b, Pop86a, Prz95, Pul01, Pul11, Put76, Qio96, Rad12d, Rad76, RA78b, RA79, Rad81, Rad89b, RC95, Rap97, Ras92]. Science [Rav91, Réd03, Ren15, Res79, Res83, Ric15, Rie74, Rip92, Rip11, Rou95, Rou15, Saa17, Sal92a, Sar11, Sas92, Say05, Sce06, Sch83b, SC74a, Sch74, Sch02a, Sch11, Sch75, Sch86c, SR16, Sch99d, Scr74b, SC74b, Seg91, Şen05, Set95, SM90, Sha74a, Sha74b, Sha84a, Sha84c, Sha84d, Sha84g, Sha84h, Sha84i, Sha84l, Sha84s, Sha01, Sha86b, She82, Shi07, Sie86, Sie01, Sim15, Sim77m, Ska92, Skl74, STNW12, Soy05, Sta92b, Ste88, Ste94, Ste00, Str94, Suc82, Sup74, Syl75, Syl89, Tau97a, Tau98, Ton73b, Ton73a, Ton81f, Tor99, Tou67, Tou86a, Tou86b, TRP15, Tuo87, Tur12, Tur86a, Tur86l, Twe00, Tym15, UM86a, UM86c, UM88, UM92, Uus01, Val99, VK83, Van15, Vas92, VdOM11, Vih01a, Vih01b, Vil02b, Vil02k, Vil02m, Von84a]. Science [Wal81p, Wal81o, Wal92g, Wal92i, Wal95, WP96, War63, War67a, War74, War76, War79h, War79p, Was91, Wat92b, Waw82, Web08, Wei01a, Wei11, Wen96, Wer01, Wet17, Wet87, Whi01, WG11, Wil12, Wis71, Wis07, Wis11, Woo91, WC91, Wor76, Wra15, Wri96, Wüt16, Yal96, Yil74, Ytu95, Yus96a, Yus96b, Zaj87, Zam92c, Zen99, ZRKC03a, ZRKC03h, ZRKC03k, Zil03, ZRKC03l, ZRKC03m, Zim99, Zol89i, Życ95b, Żyt82, d’E08, dV11, vD14, vW87, vW94, Bur87b, CW67, For07, GC97, Mit04, PB94, Ras74, RG07, SCD03b, Sza04, Vaj99, Zam92h, Bar15, CLR^{+}92, FKR99, Pet92b, Sal92a, SC74b, CS79-41, Grü73v, Vil02g]. Science-Based [Tau98]. Sciences [Aga74b, Aga75c, Amr87a, AZW87, Dil06l, Ano08-49, Bab99a, Bab99b, Ben06, BR18b, BH17a, BH17b, BC94, Bub84, But93g, Chi81a, CW74b, CW83a, CW84a, CW84c, Coh94a, Coh94b, Coh94c, CL00, CH14, DFL11, Duc92, Ezr86, Feh99a, Fey67, Fle91a, Fu93a, Gal03, Gav94, GSW95c, Gig03, GM89b, Goh81, Gou83, Gra83b, Gra83e, GN86, Hac94b, Hal90b, Har15a, Hee97, Hol97, Jar86j, Jar86p, JL95b, Jar95, Kol74, Kos84, Kow16j, Kra15a, Kra82c, Lat83, Lec94, LC17b, LS90, Ley84, Lim94, Lon12, Löw86, Mah75, Mar89, Mar97b, MP79, McK75, Men86, Mey91d, Mor71, Mül94, Nai86, NNW74, Nou90, Now01, Nye91, O’C02, Obr92, Pet92b, Pye92, Rad12a, Rad12c, Rhe11, Roc11, Rol16, RB97, Sch80a, SG92, Sie82, SPB18, Sim11a, SDG15b, Spr91a, Sza79a, Tok09]. Sciences [Tou02, Tuo82, Uye97, Vam09g, Vic92, Wal76, Wal92f, Wen14, Wes91, Wia79b, Wic91, Wim12a, vDP17a, vDP17b, CLR^{+}92, GSW95a, Kle01b, DCS15, Mot15, Seg03, Kau14]. Scientific [Aga75h, Aga81c, Aga81f, Aga81n, Aga81o, Aga81u, Aga81-27, Aga81-33, Aga81-36, AC82, Aga84, AP00b, ARG12, Alt14a, Ams92f, Ano06-29, Ant90, Avg89, Ban15, Bar14a, Bec91r, Bel12j, BM88, Ber74d, BP14b, Bid11, Bin96, BB11, Boo12, Bou03, Bra74a, Buc84, Bun12o, Bur80, Bur84, But93f, But93a, But93k, But93l, Cas00, Chr94b, CSW74, Coh76a, CS79f, CS79j, CS79i, CS79s, Coh94c, Coh02b, Com84, De 81, DP92, DB15b, Dil06i, Dil06k, Dop01, DF11b, Dun92, Dut11, Ele92, Elk74, Est96, Far03b, Fer85a, Fet81h, Fey76, Fin80h, Fin80q, Fin80s, Fis15, Fle86e, Flo15, FM09c, Fri10, Ful02, GGN89, Gav89, GSW95c, Gav14b, Gel87, Gie71, Gie97, GH01, Góm12, Góm13, Gou94a, Gra83g, Gre75, Grm81, GCC81b, Gru80, Gru81, Grü73e, Grü76, Grü09, GL91, Hac94b]. Scientific [Han99a, Har84a, HH91, Hem88, Her14a, Hod74b, Hoo04, Hop74, Hör91a, HFD88, Izq74, Jan89m, Jar86o, Kap80e, Kap80g, Kap80n, Kap80-40, Kap80-41, Kap86a, Kar10c, Kid17, Kil03, Koc91, Koc79, Koe78, Koe80, Koe86, Kor71, Kos94, Kou11, Kow16e, Kow16k, Kra15b, Kra92a, Kum92, LR08, Lam97, Lov15b, Mäk96, Mar91, Mey91l, Min10, Mod15, Mon80, Ner89, Nic80a, Nic80c, Nic80e, Nic80f, Nic08, Nic89b, Nie74, Nol03c, Pal01, Pan83c, Pan83e, Pan83f, Pan91k, Pǎr15b, Pat03a, Pat11a, Pat87, Per81, Per84, Per94, Per86, Pes11, Pes91, PJM92, Pia95, Pin10, PP87, Pul11, Que00, RA78a, Rai11, Res79, Rie99, Rog81, Rud86, Rus80, Sab84, Sad89, Sav86, Sch83b, Sch00, Sch94b, Sha80, Sha84b, Sha84f, Sha84p, Sha84r, Sha84o, She87, Sie63, Sie71, Sim77i, Sim77s]. Scientific [Sim77t, Sma92, Smi97, Söd07, SSHH08, Sol12a, Som81, Sta74d, Sta86b, Sta94b, Sta17, TJ11, Ton73d, Ton73e, Ton73f, Ton81c, Tor00, Tou74a, Tou74b, Tri08, UM92, Urb89, Uye97, Váz96, Ver09, Vla94, Vog91, War79a, War79k, War80, War96, Wat78b, Wat87, Wat89, Wat95b, Wat96, Wen14, Wes94, Wil18, Wis71, Woo91, WC91, WP94, Wor78b, Wri80, Wyl92, Yos98, Yuz96, Zah89, Zem88, Zhe96, ZRKC03d, ZRKC03g, Zin73g, Zin73j, GC97, Hül96, Tor00]. Scientification [BVK83]. Scientifique [Gre75, CLR^{+}92]. Scientist [Baa92, Cha09a, EF86, Gal15, Kap80-39, Löw14, Man86, McA97, Sav86, Sch12a, Sch96f]. Scientists [Aga81-28, Aga95c, Bar01b, Bun17, FF07, FS94, Kra15b, Mer76, Mod15, RH14, Sne94, Wei07]. Scion [Kat14]. Scope [Els88, Sha84p, Zem09]. Scotists [Wal81l]. Scotland [Mal12, Mal13, Wil01]. Scribal [Pro14]. Scriven [Mar79l]. Scrutiny [Ano08-45]. Sea [McM74a, McM86, ZC96]. Seafloor [Fra80]. Search [Alb78, Ber13a, DR15a, FT16, Gin07, Hol70, Kau71, Kau76, Mor11, Pat15, RS07a, Sha84l, Sør17, Tur86l, Wal92g]. Searchlight [tH09]. Seat [BE10a]. Second [Bus74, CS79-30, D’A01n, D’A01g, Duh15h, Hel00i, Hil07c, Lel00o, Mar78e, Nee11f, Wun01, Ela11f, dC95]. Second-Quantisation [D’A01g]. Secret [dV74]. Section [Ges74j, Lag97i, SC74b]. Sections [Ras95, Ras00, Sai04a]. Sects [O’T88]. Secular [Ihd79h, Kek17]. See [Fle86g, Pur93]. See-Saw [Pur93]. Seeing [De 16, O’D02, Sav15]. Seeks [Mey91r]. Seem [Oes91]. Seen [Bla92b, DCS15]. Seine [Hof74]. Seldom [Oes91]. Select [Haq94a]. Selected [CS79a, Coh94m, Coh02j, Ges74w, GG91, HFD88]. Selection [Ams92a, Bir07, Bir09, Cha93b, Cla96, DD12, Dum96, Fab84, Gre74q, Kau92, Mac94a, Mar11b, Res17a, Shi90, Sob91, VN12, Wim80, HD95]. Selectionism [Swa09]. Selections [Coh94k]. Self [CT91, Coo86, Gil91, Gro82n, Jan89a, Jan89i, Kau92, Kel09, Kor71, Löw91, Mar74a, Mar79n, MA87, Näp01, Pan91i, RB91, Rus09, SM91, Set82, Sta79, Sup79, Tau91h, Tau91i, Tym95, Vra79, ZC88b, Pol71]. Self-Assembly [Kel09]. Self-Critical [Jan89a]. Self-Deception [Jan89i]. Self-Determination [Mar74a]. Self-Individualizing [Tym95]. Self-Knowledge [Coo86, MA87, Pan91i]. Self-Management [Vra79]. Self-Managing [Sup79]. Self-Organisation [Näp01]. Self-Organization [Gro82n, Kau92, Kel09, Rus09]. Self-Preservation [Coo86]. Self-Reference [Kor71, Mar79n]. Self-Regulation [Sta79]. Self-Supporting [Pol71]. Selfish [Rop99]. Selfless [Var91]. Sellars [Rot87k]. Selten [Gon03, Sca03]. Selves [Var91]. Semantic [Ant16, Buc67, Kno16, Ton81a, Ton81c, Ton81d, Ton81g, Ton81h, Zem88, ZC88h, MP16]. Semantical [But93e]. Semantics [Bel93, Ben81, Ber81, Cof83, Coh02m, Coh02n, Fet81k, Glo76, Mur12g, Øhr97, Ser81, Soy05, Ton81a, Ton81f, Ton81j, Ton81k, Ton81i]. semi [WRL95]. semi-arid [WRL95]. Semiclassical [Bro91b]. Semiconductor [San14, Vel14]. Semigroups [Mul74]. Seminar [Lip17, KC95]. Semiotic [Lie76]. Sen [Wet09]. Sensation [Bar87]. Sensationalism [Aga75q]. Sensations [Rat70]. Sense [Ber14, Cah98, Edg15, Flo17b, Mar79p, Mot82, Rop86, Stö99, Str18, Tha86a, Ton81b, Ton81l, Unt01, Wat96, You95]. Senses [Aga13n]. Sensibility [Sch02b]. Sensory [Gro82n, VN12]. Sensory-Motor [Gro82n]. Sensualists [Sma99]. Sentence [Coh02p, PP15a, PP15b, PP16]. Sentences [Mat89, Mor18, Pan96, Zin73c, Zin73k, Zin73o, Zin73n]. Sentential [Sim99, Zin73p]. Sentience [Mar78g]. Separability [Bun95a, Fay94, How97]. Separation [Aug82]. September [MS75b]. Sequent [Neg81]. Sequential [Ana90b, Wil91]. Serial [Gro82b]. Series [Bur74a, Onn91a]. Seriously [Kam93].

Sermon [CS79q]. Sermones [Sto75]. Serres [Lar99]. Serve [Kow16j]. Sesquicentennial [Sch96f]. Set [Bar16, Bla95j, Hof09, Mar74b, Mer97, Żyt82]. Set-Up [Hof09]. SETI [Lam97]. Sets [Bus97b, CL93, Hal16, Hüb78b, Mas15a]. Setting [Aga81l, Eve74, Wal81o]. Settings [Now01, Rot87g]. Seventeenth [Hoo04, Rai10, Bla04, Chr14, Coh90, Coo11, Das91, Duh02i, Duh02q, FEE18, Hey86, Jam92, Men76, NRS92, PT04, Pop86a, Sas03j, Wil18]. Seventeenth-Century [Hoo04, Coh90, Coo11, Hey86, PT04, Pop86a]. Seventh [Stu95]. Seventy [DPW86a]. Seventy-Fifth [DPW86a]. Several [Xiu96]. Sewall [Pro92, Cro92b]. Sexual [VN12]. Shadow [Hol07, Roc11, dB87b]. Shall [Bla95j]. Shanghai [Sun18]. Shanlan [Yus96b]. Shape [Bod18, Goo77i, Lay92]. Shaped [BE10b, Bro10, Che10]. Shaper [Kit74]. Shaping [Sør17]. Share [RB97]. Shared [Ano08-42]. Sharpened [Woo02]. Sheds [Irz15]. Shelley [Bon87]. Shells [GL03]. Shemonah [Fra00]. Shift [Bor99]. Shifting [Car01]. Shimony [Aga87c, Čap87b, CHS97a, CHS97b]. Shipping [CW18]. Ships [Coh74a]. Shocking [vdW04]. Shockwaves [See70]. Shopkeepers [Nic89a]. Short [BRW14b, Bas14, Coh94l, Gro82b, Gro82f, Mar90]. Shortcomings [Lin93c]. Shou [Lih93a]. Should [Aga75u, Coh02o, Coh02p, FF07, Gar12b, Joh11, Koh98f, Men74, Ror86, Bur80]. Show [AV96, TV95, Fey78a]. Shrinking [Ges74x]. Shui [Mat17]. Sibling [Rus13]. Siblinghood [Aga03-31, Ng17, Rod17]. Side [Cas00, Die14, Gre74p, JJ07, Rus09, Sal01]. Sided [Hon09a, Mil79]. Siècle [Gre75, Sal92a]. Siècles [Ras75]. Sight [Red93]. Sigma [CL93]. Sigma-Convex [CL93]. Signal [Lub82a]. Signaling [Ger15a]. Significance [Bub84, Čap71-36, Čap71-37, Čap87a, Čap91l, Chu00, Grü73u, HR74c, HR74d, HR74e, Pan84, Pet69a, Pol16, Smi89, Wal90, tH09, vH77e]. Signs [Mar84g, Zin73q]. Sijzī [Ras00]. Silence [Fin93]. Silsbee [TO91]. similarity [Onn91p]. Simme [Ned90b]. Simmel [Kae90a, Phi90a, Phi90b, Dah90, Fri90b, Fri90a, Jaw90, Jed90, KPC90, Kóh90, Lec90, Mes90, Moo90, Ned90a, Sca90a, Sha90, Vro90, WW90, Wes90a, Wes90b]. Simon [Duh91m, SPB18]. Simple [Čap71f, Čap71-27, Čap91m, Duh02u, GS82, Mir15]. Simpler [Lag97i]. Simplicio [Pit89]. Simplicity [Kol74, Str82]. Simply [Zam92m]. Simulation [HL17, Waw82]. Simulations [Fil17]. Simultaneity [Grü73o, Jan83]. Simultaneously [Gin03b]. Since [NRS92, Que97, Tam14, GHQ96b, Piz81]. Single [Bun12i, Fet81i, Pos95p]. Sinicization [Fu93a]. Sinification [Tho10]. Sir [Abb94, Aga81-29]. Sites [NG01f]. Situations [Fet10d, Res17a, Rot87a]. Sive [Ide00]. Six [Duh15i]. Sixteenth [Cif05, NB12, NB13, Sas03k, Sch75, Wal81d, Wal81b, Wal81e, Wal81q]. Sixteenth-Century [Cif05, NB12, NB13, Sch75, Wal81q]. Sixtieth [CS79q, CS79-30]. Size [Goo77i]. Skeptic [ZC88g]. Skepticism [Bro96a, Fis95, Mey85f, Pop86a, Van99]. Sketch [Aga81p, Kuh09, Leb82]. Sketched [Mar78o]. Sketching [Wil12]. Skew [Eve74]. Skids [Ros74b]. Skills [Dol03, NG01b]. Skinner [Kit74]. Slant [van91]. Slave [Hei17]. Sleepwalkers [Aga81-28]. Slice [Jan03b]. Slit [Guc81]. Slow [Bow76, FG92, Grü73o]. Slow-Transport [Bow76]. Slowly [dB87c]. Small [Lag97t, Mar01]. Smith [Fre86r]. Smithsonian [Mac01]. Snakes [Sch12a]. Snapshot [Hon16]. Sneed [Bal89a]. So-Called [D’A92, Sza79a]. Sober [Tau91a]. SOC [Wer01]. Social [Aga74b, Aga75j, Aga81l, Aga84, Aga95c, Aga13p, Dil06l, Aub17, Bal15, Bar01b, Bec86, BD88, Bie76, BS83, Bro91a, BC94, Bub84, Cau80, CW74b, CW83a, Coh83, Coh94a, Coh94b, Coh94c, CH14, Cou96, De 95, EE17, Fag12, Fle91a, Fre86e, Fre86j, Fre86r, FM09c, Fu93a, Gal03, Gel84, GM89b, Gou83, GC94, Gro09b, Gro09c, Hac94a, Hes09, Hil94, Hol83, HOO88c, Iru01, Jan89j, Jar86c, Jar86e, Jar86k, Jar86j, Jar86p, Jar86v, Jar86w, Jar86o, JL95b, Kau14, Kle12, Koe74, Kóh90, Kow16j, Kra92a, Lat83, Lec90, Lec94, Lee86, Lee74, Len88, Lim94, MP79, Mar86h, MAM69, McC94, Mil79, Mit76a, Mor71, Mur75, Nie74, Nol03c, Nol03m, O’C02, Omo18a, OJ14, Ott94, Par17a, Pop74, Por94, RH92, Rol16, Rop86, Sca03, Sch83b, Sch02b, Sta91a, Ste94, Sto00, Sza79a]. Social [Tok09, Tuo82, Tur12, Tur86a, Tur86l, Vog91, Vrc79, Wer01, Wia79a, Wia79b, Wit86, Wra15, Zil03, ZRKC03l, AM01b, For07, GSW95a]. Socialism [Boš79, Sup79, Vrc79]. Socialist [Ano08o, Kar10f]. Socialization [Jan89j]. socialized [HFD88]. Socially [Rod15]. Société [Osb92]. Societies [Lub15, Tha86b, Wet95]. Society [Add74, Aga81z, Aga03e, Bed09b, Ber13b, Boy09, Bun10l, Cap15, Dam96f, Eng01, FKR99, Fre86d, Fre86h, Fre86s, Fri90b, Gil09a, HH91, Hes07, Jar86y, Jer15, Jie96, KT76, Koe91, Kow16j, Kre79, Lia96b, Mun91b, Nar09, Nia96, Sch86a, She95b, VD92, Whi01, Win83a, Aga81s, Aga13j, BIT01h, LG01, Pel01]. Socio [Dam96d, Len86, Omo18b, Omo18a, Pap74b]. Socio-Historical [Dam96d]. Socio-Philosophical [Len86]. Socio-Political [Omo18b, Omo18a, Pap74b]. Sociobiological [Życ95a]. Sociobiology [Lee84]. Sociocultural [Mam94]. Sociological [Ara94, BN17, BJ89, FA76k, Jar83, Jar86u, Neu91, Sha86b, Str95, ZRKC03i, ZRKC03m, Zol89i]. Sociologism [Aga81-30, Pos78]. Sociology [Aga81z, Ban17, Dah90, Elk86, Ezr14, Hac94a, Jan89o, Jar17, KPC90, Koe95, Mac86b, Mes90, Ned90b, Now79b, Phi90a, Rie74, Sza79b, Szt79b, UM86a, UM86c, UM88, UM92, Wes90a, Wol76-28, Wol95c, Wol95j, Wol95i, Wol95e, Wol95m, Wra15, Zol89h, Zol89i]. Sociosynthesis [Lel00j]. Socrates [Bar09, DD86, Nai95]. Socratic [Söd07]. Soemmerring [Mar94b]. Software [Bun10c, SV18]. soil [BBT95, PLH95]. Sokal [Bab02d]. Solar [Bra74b]. Solid [Čap71-28, MA15, Sol12b, OP91b]. Solidity [Sol12a]. Solids [vdA91]. Solomon [CS79r]. Solution [And84c, Den99h, Ela11r, Hea04, Lag97a, Lag97q]. Solutions [Ano08-42, Lin93b]. Solvay [Stu95]. Solve [OJ14, Wor89]. Solvent [Tap91]. Solving [Gil09b, Pol74b, Sim77t, Sim77z]. Some [AV16b, Ana90a, Apo82, Are89a, Art15, Ber74d, Blo86, Bur84, But93k, Cau95, Chi83, Chu91, Cla08, CS79e, CS79-42, Coh94b, Coh94j, Coh02q, Coh02r, Coh02s, Dai16, Den99i, Dro13, Duh02s, Dur83, Ear06, Edm95, Fle86f, Fri76, Für92, Gav89, Gil03, Gio81, Gre74m, Gre76b, Gro82l, Her74, Hil96, Hüb78b, Ier90, Jim95, Kap80-42, Kir82, Kör86, Lem82, Ley84, Łod82, Mäk96, Mar79q, Mar79r, Mar79s, Max76a, Mil79, Mis74, Moc91, Mor18, NB12, NB13, Nol03b, Nol03l, Pal01, Pap74a, Pol95, Pop86b, Res79, Ros09, San08, Sch75, Sch07d, She87, Sim15, Sla89, Str84, Str74, Str95, Sup79, Ton81i, Tos81, Tuc76, Ung04a, Was91, Wim71, Zal86, Kap80x, Van08, Pro14, Sta83a]. Sophisticationism [Per89]. Sort [MA83]. Soto [Wal81f]. Soul [Bri98, Mar94b, Sca90b, Fra00]. Sound [BA09]. Sounding [GR03]. Source [Ele92, HH89, JNR^{+}07b, Ret74, PHL95]. Sources [Ams92j, BP14b, BE10b, Buc67, Che14, Duh91a, Duh91r, Fle91b, Pro10, Sch08b, Ton81h, Vis01, Wal81q]. South [BE10a, Jin96]. South-Pointing [Jin96]. Sovereignty [Fel18]. Soviet [Ada01, Vod15, Vog91]. Space [AAF^{+}14, Ano08k, Ano08j, Ano08-43, Ano08-44, Ano08-50, Aug67, Bai76a, Bar91b, Bec91o, Bec91q, Bel12g, Bot94, Čap71-30, Čap76, Čap91o, Car07, Cor76b, Des76, Duh76d, Duh76f, EFN03, Eks84, Eul76, Eve74, Fin69, Fre86q, Fut08, GL03, Grü73i, Grü73j, Grü73l, Grü73s, Grü73t, Grü73v, Grü74b, How97, Jan03b, Koy76b, LY96, Loc76a, Mau96, Max76c, Mit76j, New76a, Pal12, Pal13, Pul06, Rus76b, Sam76b, Sch76e, Sch10, Sha84q, Sta94a, Ste03, Van86, Vol06, dB87s, dB87t]. Spaces [Rip11]. Spaceship [Aga03f]. Spacetime [Bel12c, GS82, Har97b, Tor83, WW03, dB87d, dB87h]. Spain [Bro08, GPSR01, Guc87, MH01, NB12, NB13, RR15, SR15]. Span [Čap71m]. Spanish [GPSR01, Mun96, NG15b, Pel01, PS01b]. Spatial [DL05, Grü73w, Mey85x, Mey91e, Pul06]. Speaker [Den99i]. Speaking [Jon91, Pat11a, Ezr01]. Special [Bel12i, Bou17, Bro93, Buc84, Bud93, Coh16, Com84, Grü73c, Grü73o, Grü73r, Har97b, LMR95, Max76a, Pat82, PP16, SR87, Sch97a, SS03, Sta83b, Ano07e, Ano07r]. Specialists [Rud86]. Species [Beu91, GWT00d, Gre74l, May76, SG11]. Specific [Fle86f, Kox95, Van74d, DO91]. Specification [LC07, Rad12e]. Specificity [Gin03b]. Spectator [Sam92]. Spectrum [Fra09b]. Speculation [Hac91, Mos00]. Speculations [Gen91, Tau91h]. Speculative [Sim99]. Speculum [Zam92j, Zam92k, Zam92h, Zam92n]. Speech [Ges74m, Ges74o, Ges74q, Rya94]. Speeches [Gua96c]. Spencer [Hil94]. Spheres [Ros81]. Spherical [Zum96]. Spicer [Car89b]. Spinoza [Sav86, Fra86, GN86, Ide00, Jon86, Kuz87m, Lec86, Llo86, Mar86a, Mat86, Mau86, Nai86, Pat86, Pop86b, Sie86, War79a, War79l]. Spirit [Bel15, Hei88]. Spiritual [Gre07, Kni07, For07]. Split [Ges74u, Hac86]. Spontaneity [Rot86a]. Spontaneous [Men76]. Spooks [PW17]. Spreading [Fra80]. Spreads [dB87a]. Springs [dB87a]. Spuren [Bec04]. Spurious [Sim77u]. Sri [DP92]. St [Hoo99c]. St. [Duh76a, Syl75]. Stabilities [SSHH08]. Stability [Aga75l, Nee11l, Per86]. Stabilization [Ned12]. Stable [Ano09-47]. Stachel [RDS03]. Stage [Čap71t]. Stages [Kap80x, Rap81e, War96]. Stahl [Hoo99l]. Stallo [Eas84, Str76]. Stance [Bit96a, Ful02, Web12]. Stances [Fin09k]. Standard [Aga75c, Aga75d, Bun12c, CPM16, Skl06]. Standardization [Şak11]. Standards [Aga81-31, Pap74b, Ste09b]. Standing [NG01g]. Standpoint [Nor07b]. Stanley [Sac83]. Star [Whe74]. Stars [RS03, Zam92c, Zam92m]. Started [Ano05c]. State [Aga81t, Aga09, AV97, Alo14, AP97, Ara97, BC81, Bus97a, Ein07b, FS94, Maa17, Mey91s, Pat03b, Rad89b, Ung04a, Onn91a, Onn91h]. Statements [And84c, Lin93d]. States [Ano09-45, CL93, Ges74-27, MPC91, Mar78g, Obr92, Meg74]. Statics [Bun93i, Duh91a, Duh91d, Duh91e, Duh91r, Duh91n, Duh91q, Duh91p, Duh91o, Duh02s, Lag97e, Lag97q, Lag97s]. Stationary [Kuz87c]. Stations [Fin91, Sch86c]. Statistical [Cla89, CS79w, CS79-43, CC90, Fet74, Fet81j, Gre71, Grü73m, Kra74, Mac97f, Mac97h, Mit00, Mül91a, Oid98, Por94, Riv96, Rog71, Sal71, Sei92, Sim77v, Skl76, Skl06, dB87q, COG90]. Statistics [CS79d, Con81, Coo90, CC90, CG90, Gre74q, Koe89, Krü90, Pia90, Por94, Sar91]. Status [Col74, De 01, Din96, Eag83, Fre88, Ker88, Łas82, Lat83, MC80m, MC80p, Pan16, Rot87h, Sim77j, War96]. Staudt [Fre74b]. Steam [Kro92]. Steel [FQ18]. Steffens [Ham07]. Stein [Nor03]. Stem [Fag12]. Step [Car15b, Car89b, NG15a]. Stephen [Min67]. Steps [Coh02s]. Stereochemistry [Duh02j].

Stereotypes [Koe89]. Stern [Sau16]. Stevin [Duh91m]. Still [Cao07, Han15, Hoo99a, Krö12, Stö99]. Stillman [Lan90, LS90]. Stimulus [Bro91a]. Stirrings [Wol18]. Stirrups [Kim11]. Stockholm [CS79g]. Stoic [Ier90, Sam76a, Sam76b, Sam76c, She75]. Stone [Cha09a, Hoo99g]. Stones [Ham91, Haq94f]. Stood [Hoo99a]. Stop [Ant16, MP16]. Stopped [Gol12]. Storage [Aga81-32]. Storey [Sha90]. Stories [Sil89]. Story [Byn89, Cha09k, Fow94, GM14, Now01, Oes91, Ord96, Sta07c, Wol17]. STR [Grü73p]. Straight [Uga18]. Straightedge [Yav17]. Strange [Bal89a, Whi88]. Strangeness [Har97b]. Strasbourg [Las15, Rit05]. Strategic [Rod15]. Strategies [Ada11, Ela11g, Fu93a, Mar94a, Pan83d, Pye92, Tou74b, Wie12, Wim80, Pye92]. Strategy [RS07a, Slo76, Slo83, Szt79b]. straw [RL95]. Strawson [Čap71-36]. Stream [Čap91n]. Strength [Bar84, Pes11]. Strengths [Gra07a, Zol89h]. Strictly [Mac01]. Striderna [CS79g]. Strife [CS79-44]. String [Mun15]. Striptease [FL09]. Strong [Kap80-35, Nol03d, Nol03c]. Stroyed [Syl97b]. structible [Syl97b]. Structural [Ban80, Bel12j, Ber13c, Bue11, Cau80, EL11, Fra01, FL11, Mas11, Mun15, Pin11, Tem82, Vot11]. Structuralism [BB11, DF11b, Dum92, Pan83b, Sch79, TJ11]. Structuralist [Bra11]. Structure [Ano09-43, Art15, Ban80, Bec91r, Bek86, BC81, Ben06, Bie82, Bie76, But93h, Čap76, Car08, CF11, DB15b, Dis06, Fin80f, Fin80m, Gie71, Gin15, Goo77j, GHQ96a, GHQ96b, Har84a, Har89b, Has97, Hen01, Hon98, HH15, Jar86x, Kah18, Kat11d, KS83, Krö12, Mod15, Mon80, Mug81, Nie74, Oid98, Pal01, Pan83e, Pan83f, Pol76, Pos95s, Pro95, Qua84, RA79, Roz71, Sch84, SH91, Sha84r, Sha15, Sim77w, Smo03, Sol12a, Tis63, Tor83, Wul96]. Structured [Sim77w]. Structures [Are89j, CL93, DF93, Gäh89, Kah18, Kou89, Lei83, RT01, Xiu96, Zin73r]. Structuring [Gra83h]. Struggle [BZ96, GWT00i, Han17, Kap80-28, PT99]. Struik [CSW74, Rie74]. Stubbe [Hei08c]. Stubble [Ore00]. Student [Kap80-36, Lue18]. Students [Gov15, Pro14, Wol76z]. Studied [Eck17]. Studien [Bec04]. Studies [Aga81z, Aug67, BS16, BG09, BM89, Cha12b, Chi81b, CW69a, CW74a, Coh80, CHS97a, CHS97b, DC96, Dun96, Fen12, FS96, Gin03a, GWT00b, GWT00c, Gro82m, GR11b, Han15, Hat89, Hod90, Hon16, IG05c, Jan89b, Jin96, Kil05, Kin16, KV11, Krü91, Kul01, LPV15, MC95, MC96, MC80z, Mil87, Mun96, NC98, NB12, NB13, Nic80e, Nic90, PRT15, PST15, PB07, dMPC14, PJM92, PP18, Pie16, Rad12e, RC95, SS16b, Sch91a, Sha16, Sil89, Tur12, UM86a, UM86c, UM88, UM92, VdOM11, Vih01b, Wen96, Woo91, WC91, ZRKC03k, Ueb91b, Bec04]. Study [Ano05c, Ano08-40, Ant87, Bağ05, BAC14, Bru07, COG90, Cho14, Cho76, Fel15, Fis95, FC16, FA76g, Fri90a, Fu93a, Fu93b, Ges74z, Gru81, Han99a, Hod92, Hoh83, Hon98, Jar86i, Kat06a, Kin16, Kos76a, Kos94, Kum92, Kum11, Lia96a, Lih93b, Lim94, Mou82, Omo18a, Pan83f, Pop86b, Rus80, Sab92, Tri12, Vil02g, Wol76-33, Zah78, Zam92c, BLA95a, Duh15c]. studying [WPJ95]. Stuff [Nee97]. Stumpf [Ima01]. Style [Aga13h, Blo86, Bol95a, Fey78a, Jan89k, Jan89l, Mar76b, Mou86]. Styles [Gie86, Gil03, Hac94b, Irz03, Kle03]. Subject [Gol69a, Han99l, Nak74, Rot87i, Vih01a, Zin73r]. Subjective [Bun12r, Cas00, Moo90, Pan91j, SF94, Lee74]. Subjectivism [Fin80o]. Subjects [Chr94b]. Sublanguage [GR89, HGR^{+}89, RG89]. Sublimatory [Sti74]. Subsequent [DP92]. Substance [Bol95a, Bou02g, Han99l, Haq94b, Mit76c, Nee15, Pan84, Sei97, WZ76]. Substances [Duh02u, Nee06, Von84a]. Substitute [RH14]. Substitution [Duh02d, Pel95]. Succeed [Loc97]. Succeeded [Cha09a]. Success [Aga81-29, Aga81-27, Men74, Nii89, Wri96]. Successfully [dF14b]. Succession [Čap71x, Čap71-38, Jon86, Loc76b]. Successive [Neu76a]. Such [Dou87, dBJD95, Dou87]. Sudden [Nov74, Onn91g, Onn91h]. Suddenly [Hor94]. Sufficient [And09a]. Sugar [RH14, BdOU^{+}95]. Suggested [Chi83, Sta83a]. Suggestions [Bru07]. Suicide [Mac86b]. Suicides [Hac94a]. Summa [Syl91]. Summary [BVH^{+}83d, BVH^{+}83e, BVH^{+}83f, Kar10i, Mac97g].

Summation [Bur74a]. Summer [WC91]. Summit [Ren07c]. Sun [Hoo99a, Lem82]. Super [Cao07, Onn91h]. super-conductive [Onn91h]. Super-Hero [Cao07]. Superconductivity [Bro95, CM89, Har08b, JW14, KJ14, MA14]. Superfluidity [Bal14]. Superfluity [Čap71-38]. Supernatural [Bri98]. Superposed [Ohm98]. Superposition [Ela11s]. Superstructure [Mar94c, Sta79]. Superstructures [Gal02c]. Supervenience [Ano09m, Bun93j]. Supervention [Ano06j, Dil06a]. Suppes [Bre74a, Bre74c, Mar79l, Sza03, Van74a]. Supplement [McK92]. Supplementary [Zuc69]. Support [Mus78, SL82, Wor78a]. Supporting [Pol71]. Supposed [Jar83, Jar86u]. Suppositio [Mar79k]. Suppositione [Wal76, Wal81i]. supra [Onn91j, Onn91k, Onn91l, TO91]. supra-conducting [Onn91l]. supra-conductivity [TO91]. supra-conductor [Onn91k]. supra-conductors [Onn91j]. Surface [RS06, SA12]. Surfaces [vSB91]. Surgery [Lam12, Lam13]. Surrender [Wol76-29, Wol76-36, Wol76-32, Wol76-33, Wol76-30, Wol76-28, Wol76-37, Wol76-35, Wol76-34, Wol95b, Wol95d, Wol95f, Wol95g, Wol95j, Wol95h, Wol95i, Wol95e, Wol95m, Wol95l, Wol76-31]. Surrender-and-Catch [Wol95d, Wol95f, Wol95g, Wol95j, Wol95h, Wol95i, Wol95e, Wol95m, Wol95l]. Surrendered [Lev83d]. Surrogate [Lel00d, Lel00o]. Surrogates [Mar79v]. Surrounding [Rai10]. Survey [Beu96, D’A01o, Jia96, Jin96, Mac01]. Surveys [Kum92]. Survival [Mar90, Mac00]. Survive [Men74]. susceptibility [PO91]. Sustainability [HCP18]. Sustainable [Ber13b, WRL95, PHL95, Tho95]. Switzerland [Myl03]. Syllogisms [LMR95]. symbiotic [dBJD95]. Symbol [Fin09h, Tho96]. Symbolic [Che97, Jar86m]. Symbolism [Gög87, Smi82]. Symbols [Mar84o]. Symmetric [ACE17, Ber17a, Sim77x]. Symmetries [Sau03, Sto97, Van93]. Symmetry [Byc16, Ear06, Fis15, Hon93, MC80x, MC80y, Rom69, Tam01, Wan17, dB84]. Symposium [SN87, Sta74c]. Symptom [Wer69]. Symptomatology [Van99]. Synchrony [Bow76].

Syncretic [Tha84d]. Syndrome [Ges74i, Ges74n]. Syndromes [Ges74k, Ges74u]. Synonymity [Ton81e]. Syntactic [Ayd05]. Syntax [Bar76c, Fet81k, Som76b]. Syntheses [Tur16]. Synthesis [BSM06b, Dun89, Gal02d, HR74g, OP97, Pol03, Ram01, Ras94a, Tau97a]. Synthetic [Fre86h, Heg91, Kan84, KL17, Liu93, Per86]. System [Aga81x, Aic15, Bar14a, BV01, Cér86, Chu01, Elk76, Eve74, Fre86m, FA76h, Goo77e, Goo77k, Hal91b, Hoq18, Hüb78b, Kra15a, Lag97d, Lag97e, Lag97g, Lag97j, Lag97t, Mac94a, Mey85y, Mün14, Nee11j, Nee11k, NG01f, NNEH11, Ton97, Uye97, Van74d, Von84a]. Systematic [Bel15, Fre86k, Hul76a, Tis74]. Systematization [Duh91q]. Systemics [Bun12s]. Systems [ÁNS91, AM83, Bel12g, Bel12e, Bel12f, DD86, Enz00, Ger15b, Gil09a, Gru80, Hef06, Hon09b, Kas76b, Lef01b, Les17, Mac17, MPC91, Mar76d, MC80t, Mey91p, Nic16, Pan91g, Pan91h, Pos95p, Roq11, Sch96c, Sim77a, Sim77f, Sim77q, Sot03, SC14, Szt79a, Tör76, Tou71, Xiu96, ZC88e, SVD95, Tho95]. Systems-Approach [Szt79a].

’t [Duh02x, Lip17]. T. [Mar97b]. Table [CLR^{+}92, Sce01]. Tables [Pro10]. Tablets [Cla10]. Tachyons [CR81, Fit71]. Tacit [Jan89m, Życ95b]. Tactics [But93k]. Tait [Saa17]. Taiwan [Fu93a, LF93]. Take [Pro17b]. Taken [Mac94b]. Taking [Kam93]. Tale [Byn89, Pos95s, Rit05]. Tales [Pur93]. Talk [Bur87a, Hac15a, Mar79l]. Talking [Bra97, Gol12, Man81, Mar76b]. Tamayo [NG15b]. Taming [Rad89b]. Tao [Wul96]. Target [Wak15a]. Tarski [Hin76, Nar09, Wol17]. Tarski-Type [Hin76]. Tartaglia [Duh91l]. Tartu [LP01]. Tasks [Wuß91]. Tauber [Tau91f]. Taught [Gri15]. Taxonomy [Fu93b, Gre74l, Haw67, Yos98]. Teacher [Blü70, Sch96h]. Teachers [dJAO14]. Teaching [AB14, Ano06-30, BP14b, BE14, Cha14, Cif14, Cla14, Dam96i, Goh81, Lam14, Mot15, Sti94]. Technē [Win83a]. Technical [Čap71j, Hal84, Jan84b, Kap80-38, Pos18, RH92, Win83a]. Technicalities [SG11]. Technics [Ihd79i, Qui96, Wan18c]. Technique [Clu05, War79r]. Techniques [Kno04a, Kum11, Pia90, RS06, Ver09, WPJ95].

Techno [AAF^{+}14, Spa03]. Techno-physical [AAF^{+}14]. Technoaxiology [Car83]. Technocracy [Aga81-33]. Technological [Ada11, Ágo11, Cér86, Coh83, Gav14b, Góm15b, Jar86v, Kro89, KB92a, KB92b, Len83, Mar83b, Nak95, NG01g, Ols82, Que00, Rap81f, Rod15, Rop83, Ros86]. Technologies [Ben15, Cul11, Fin14, Kim11, Lev86, Nei15, Win83b]. Technology [Aga75d, Aga75k, Aga03w, Aga03-32, Ágo11, Bal15, BVK83, BRS18, Byr83, Car83, Cér86, DC96, Die83a, Dur83, DR83, FS96, Gon15a, Gon15b, GHQ96a, GHQ96b, Haa99, Hol83, Hun83, Ihd79b, Ihd79c, Ihd79d, Ihd79j, Ihd79k, Ihd83, Jan89c, Jar86x, Jar86g, Jer15, Kap80-43, KF86, Kra92b, KB92b, Lia96a, LvHV92, LMZ18, Mau83, Med96, MA83, Mic83, Mit83, Mit86b, Mit86a, MH86, Mit15, Nor11, OJ14, Orv89, PRT15, Pan94, Par09b, Rap81a, Rap81b, Rap81d, Rap81e, Rap83, Ric17, Rip92, Rop86, Sac83, Sar11, Sch83f, Set76, She95a, SF83, Sko76b, Sla89, Spa03, Sta92b, Str83, Sun18, Tym15, Wan18c, Wan18a, War79b, Wen96, Yal96, YCMY18, Zim83, Zim86, vdP15, BLA95a]. Technoscience [Bun12t, Pes07, SN11]. Technosynthesis [Lel00p]. Tectonic [Lau80b]. Teleological [Car86, Duc96, May74, May86]. Teleology [Bec76, Bie95, But93l, Gre74b, Grü73b, Rus82, Stö99, Tur86k]. Teleonomic [May74, May86]. Teleonomy [MV80e]. Tell [Byn89, Gou94a, Now01, SV18]. Tell-Tale [Byn89]. Telling [Elg10, Sil89]. Telos [War79r]. Temper [But93a]. Temperature [Har08b, Nee11b, Pap15, SR15, Onn91m]. Temperatures [Onn91e, BBO91, Onn91o, Onn91c, Onn91d, Onn91j, Onn91n, OP91a, WO91]. Templates [KL17]. Tempo [WØ97]. Temporal [Ano08l, Bos76, Čap71k, Čap71m, Dai16, Ges74m, Grü73w, Grü73n, Ism17, Øhr97, Wat98]. Temporality [Lau96, Qua84, Rou16, Sch84]. Temporalization [Ber81]. Temptation [Hoq18]. Ten [Shi94]. Tenable [Mor83]. Tendency [Mey85b]. Tennant [Tra15]. Tense [Dol16, Gab76, Has97, Mar79t, Urc97]. Tenses [Has97]. Tension [BS96, DR15a, Gol89, Sto79, Ung91b, Wei14, WG11]. Tentative [Bah96]. Tenth [Har10, dC95]. Term [Gro82b, Gro82f, LMR95, Mar90]. Terminological [Kle71]. Terminology [HR74a]. Terms [Ano09o, Dil06f, Grü05b, Pap90, Sim77f, Sim77j, Sla89, Ton81c, Wat63, Zin73v, Zin73s, Dil96]. Terrestrial [CS79c, Fin09i]. Terror [Grl79]. Test [And84c, War67b, Yil05]. Testability [Sha00, Sim77s]. Testimony [Coh76b]. Testing [Aga75r, Chu01, Pat82, Szy82]. Tests [DS12, Led18, Rog71, Sim77v, Sim77x]. Teutsche [Scr74a]. Text [Che05a, Che05b, Clu05, Kno05, Net05, Sie95, vL16a, Haq94a, Mac97h, Mac97i]. Textbook [Pet76]. Textbooks [Gom15a, Rai11, Van08]. Textiles [NG01f, NG01c]. Texts [Ano06-30, Bro10, Cha14, Che10, DL05, Ela11o, FM09c, Gro09a, Har89c, Nol03a, Pin10, See02, Str94, Tho10]. Textual [Haq94h, Har10, vL16h]. Thales [Vam09n]. Theft [And03b]. Their [Aga81-27, AP00b, Ano08q, Ano08p, BAK11, Bro11, Bry98, Čap76, Gem90, Grü73r, Grü89b, Hee74, Hon98, Hüb76, Les17, Lub15, MM91, Min10, RH14, Rip11, Rog81, Rot86b, Sha84o, Wim80, Zam92c, Aga75h, Aic15, CS79u, Tur86i, Wol76z]. Theistic [Grü95a]. Them [Ghi93, Kau76, Kow16l, Kau71]. Thematic [Grü89b]. Thematization [Tha86g]. Theme [Haq94e, Pet92b, Pet92b]. Themes [Chu91, Mit86b, Mon09, Ton81h]. Theodor [Hei94a]. Theologian [Fin09g, Wes92]. Theologians [Gre91a, Gre91b]. Theological [Ols98]. Theology [Hoo99f, Hor16, MS75b, NRS92, Ols87, Vil02h, vE75, Zam92h]. Theon [Ber14]. Theorem [Bit97, Koç05, Nee11b, Pol16, Ren90, Ste97, Hin71, Nee11e]. Theorems [Har97a, Web83]. Theoretic [AM83, Cha93b, Dav95, Koe71, LMR95]. Theoretical [Aga69a, Ano08-40, AN15, Bar08, Bit96f, CM89, Coh02m, D’A00, D’A01f, D’A01i, D’A01n, Dor00, Fel18, Fin80c, Für69, GC10, GM17, Gly71, Gon15a, Gre71, Grü05a, Han99a, Hoh83, Hon98, Kai96, Kap80d, Kat06f, Kuu01, Łas82, Ly08, Max76a, Pap90, Pat92, Sar92b, Sim77j, Tap91, Tok09, Vra79, Wim15, Win71, D’A01l, Sal71]. Theoria [CS02]. Théorie [Bea75, Onn91r]. Theories [Abr07d, Add74, Aga69a, Ams75c, Ams75d, Ams75h, Ano06-29, Att13, Avg89, Bab99b, BMP09b, BM88, Ber83d, Bie82, Bun12a, Bur86, Cau76, Chi83, Cor00, D’A01e, D’A01o, Dil96, Dil06f, Dil06k, Duh02i, Duh02v, Dum96, Ein01, Fet81c, Fun86, GGN89, Gav90, Gav96, Gen71, GW97, Giu04, GR03, Goo69, Gre91b, Gre74s, Grü63, Grü73o, Grü86, Har08a, Hem88, Het15, Hoo04, Hoo99h, Hoo99l, Jar83, Jar86u, Jar86o, Kat06g, Kle71, Kro92, Kuh87, Lin93b, Maj98, Mar86g, May69, McA97, Mis95a, Mis74, Ned12, Nor07d, Odd89b, Ost76, Pag96, Pal04a, Pat76b, Plo76, Pol03, Pol16, Ron96, Rub83, Sch76c, Sha84o, Sie63, Sim77b, Sim77o, Sim77s, SM07, Ste79, Str76, Tho01, Tow71, Wal81a, Zah89, Zin73t, Duh15c, Sta83a]. Théorique [Pat92]. Theorist [Hal82, Hal92, Sav86, vD17]. Theorists [Heg91]. Theory [Abr07b, Abr07c, Abr81, Acc90, Aga75s, Aga84, Aga86, Aga95b, Aga03-33, And84b, Ano07a, Ano07f, Ano07e, Ano07g, Ano07o, Ano07h, Ano09-45, Ant90, Ayd05, Bab99c, Bab99b, Bai76c, Bar91a, Bar74, Bar76c, Bea75, Bec91d, Bel12i, Ber83j, Ber83k, Ber96b, Ber94, BZ96, Böh87, Bou17, Bro95, Bro92, Bru76, Bru95, Buc84, Bun95b, Bun12c, Bur83, Bur87d, BRS18, But93i, Čap69, Čap71e, Čap71g, Čap71r, Car08, CF11, Cat93, Cel81, Cel93, Cha09o, Che95, CdF95, Cho67, Cho86, Chr07, CM89, CS79m, CS79o, CS79-27, CS79v, CS79y, CS79u, CS79x, CS79-37, CS79-38, CS79-43, Coh80, Coh02s, CH14, Com84, Coo84, Coo90, Cou96, CFG96, Cus96, D’A98, D’A01m, D’A01k, D’A01g, DF93, Dam96d, Dar80, Dau96, DR15a, De 81, Dem89, Dic96, Dil06i, Dra94, Duh76d, Duh91r]. Theory [Duh02a, Duh15d, Eag83, Eck17, Ede88, Ein07a, Ela11a, Elg84, Far03f, Fay07, Fer85b, Fer85a, Fet81i, Fin80r, FH10a, Fle91a, Fos91, Fre86q, Fre99, Gei67, GG96a, Gia71, Gie86, Gly83, Goe70, Gög87, Gol69b, Gra07b, Gro82j, Gro82n, Grü73c, Grü73f, Grü73l, Grü73q, Grü73r, Grü73u, Grü89a, Hål07, Hab99, Hav69, Heg87, Hei98, Hin94, Hor16, Hüb78b, Hul76b, Hun10, Ita92, Jan85d, Jar86y, JN79, Jaw90, Kap80i, Kap80j, Kap86b, Kar10d, Kar10e, Kar10g, Kat11a, Kat06e, Kat06f, Kau96, Knu07, Koe78, Koh97, Kra92a, Kro92, Krü90, Lau80b, Law96, Led18, Lel00m, Lep80, Lew96, Lew76, Lip86, Mac80, Mal03, Mar78c, Mar78e, Mar78o, MH01, Mar74a, Mar84d, Mar84f, Mar11b, May02, McM74b, McM76a, MC80g, MC80i, MC80j, MC80t, MC80u]. Theory [MC80z, Mey85t, Mic97, Mie07a, Mie07c, MH86, Mit97, Mon09, Mor71, Mül91a, Nag84b, Nar09, Nee06, Nee11g, Nen18, Nic76, Nik74, Nor07b, Nou90, Nov74, Pan91g, Pan91k, Pat90, Pel95, Pen99, Pet89, Pia95, Pla94a, Pla94b, Pro00, Rad79, Rap86, Ras94e, Ren90, RSS^{+}07a, RS07c, Ren07c, Rie99, Roc01, Rop99, Roq11, Roz71, Rus09, Sac67, Sai04a, Sam96, Say05, Sca90a, Sch74, Sch83d, Sch97a, Sch08c, Sch63, Sch96c, Sch76d, Sch95c, Sch99d, Sha84t, She75, Shi90, Sie95, Sie71, Sim77y, Sim77z, Sim11b, Sne89, Sob10, SSHH08, Spi99, Sre87, Sta69, Sta94c, Stu97, Stu86, Sup03, Swi99, Sza03, Szt79b, Tho76, Tos81, Tow71, TP90, Val96, VN12, VG87, Wat83, Wei82, Wes91, Wia79a, Wim92, Wol95f, WP94, Wri80, Wun76, Wyl92, Xiu96, Yan98]. Theory [Yil05, Zaj87, Zhe96, Zin73f, Zin73g, Zin73h, Zin73j, Zin73u, Zin73v, Zin83a, Zum96, Życ95a, dB87c, tH09, vW94, Mey85v, Kau14]. Theoryhood [Wat95b]. Therapy [CF11, Sch83d, Seg17]. There [Aga95d, Ams92a, Ano09j, Coh02b, Eck98, Fra96a, Fra99k, Fre04, Grü73n, Har95, Hen06, Hil96, Par17a, Ros86, Seg15, Sta97, Wol76-38, Wol76j, Chi81a, Elk86, Suc82]. Thermochemistry [Duh02w]. thermodynamical [Onn91p]. Thermodynamics [Cha09o, CS79p, Duh02b, Hel00a, Mac97f, Mac97h, Mac97i, Mou89, Nee11f, Skl76, Tru82, Nee11c]. Thermoelectricity [Kos94]. Thermonuclear [Kap80-32]. Thermoscope [Val10f]. Theses [Cas95c, Mac94b, Shi76]. Thesis [Bra74a, FM09b, Gen71, Grü63, Grü86, Łod82, Mar15, San18, Sch11, Smi97, Ueb97, Ver09, Web83, Par17a]. Thick [Kle09]. Thing [Bit96g, Mey91q, Pap89]. Things [EE17, Fek99, Gre74f, Gre74g, Gre74h, Hah91, Haq94c, Haq94e, Haq94f, Jar95, Mey85p, Nee97, Oes91, Pie95, Sch07b]. Thinker [Blü70]. Thinkers [San92b]. Thinking [Bar02a, Bra97, Dam96b, Dam96e, Den99j, Fle86f, Hac94b, Har84a, Hoo99j, Hoo99i, Jar86y, Jar86z, Lec94, Mic18, Rad89b, Rat18, Réd03, Sab84, Sim77-27, Vil02i, Zel95, ZC88i]. Thinly [Aga95a]. Third [Coh02t, Duh15j, Gil88, Hey86, Kau96, Okr89, Pec11, Pop86a, Ren07d, Sol08a, Tou76, Dor92]. Thirdness [SM96]. Thirteenth [McK75]. Thirty [Gie12]. Thirty-Five [Gie12]. Thisworldly [Lev95]. Thomas [Lev07, Syl75, Ano08-47, Ano08-46, Duh76a, Gol12, Mar95a, Ric15, Sch08a, Sch08b, Shi76, Syl08]. Thomists [Wal81m]. Thompson [Gou76]. Thorbecke [Maa17]. Thorndike [Zam92n]. Those [Fer85e, Wel15]. Thought [Atk03, Bel12f, Bro91a, Čap91j, CS79f, Coh88b, Cra93, Fre86g, Gie86, GC06, Gra83b, Gru80, Hag16, Hey86, Hoo94, Len87, Mar78e, Mar84b, Neu91, Pop86a, Réd03, Rös94, Rot87k, Sas03d, Sas03i, Sch98, Sch06a, Sch94b, Scu98, Sey96, Stö03, Vam09g, Vla94, Wal81q, War80, Wis71, Zen99, ZC88e]. Thought-Collectives [Gie86]. Thought-Styles [Gie86]. Thoughts [Ana90a, Art15, Cau95, Coh02a, Goo74, Kir82, Pan94, Vin92, Win17a, Wis11]. Thread [Lek15g]. Three [Aga81-34, Ams92d, Aug67, Cas95c, Cur80, Dor97, Gre74r, Grü73j, Han99m, Has97, Hod90, Irz95, Ita92, KL16, Mir15, Rad03, Sha16, Tor82, War96, WG11, dB87r, dB87s, Sey96]. Three-Dimensionality [Grü73j]. Three-Point [Has97]. Three-Valued [Mir15]. Three-Way [WG11]. threshold [Onn91j]. Thrice [Rit05]. Thrice-Told [Rit05]. Throughout [BP14b, Wet17]. Tides [Hoo04, Lew96, Pit87]. Tier [Pos95p]. Tightrope [Far03b, Far03k, Sun18]. Tightrope-Walking [Far03b, Far03k]. Till [Räz16]. Time [Aga81-37, ACE17, Ano08k, Ano08j, Ano08-43, Ano08-44, Ano08-50, AB17, Ari76, Aug76, Aug67, Aug82, Bai76c, Bar76a, Bec91q, Bek86, Bel12d, Ber17a, BAK11, BH17a, BH17b, BMM17, Bru76, Bus97b, Čap71r, Čap76, Čap87a, Čap91g, Čap91o, Car07, Coh80, Coh94k, Coh94m, Cor76a, Cos87b, DP92, Den99b, Die16, DR16, Dor97, DD11, Dro13, EFN03, Fay97a, Fay97b, FSU97b, Fit74, Fut08, Gas76a, GL03, Goo77g, Grü73a, Grü73d, Grü73f, Grü73s, Grü73t, Grü73u, Grü73v, Grü73n, Grü74b, HS16, How97, Hun17, Hus17, Ism16, Jan84a, Jan85e, Jan03b, Jia96, KT93, KK01, KSW97, Kuh87, Lea96, Les17, LY96, Mar79h, Mau96, Max97, Max76b, MC80a, MC80b, MC80f, MC80g, MC80i, MC80j, MC80h, MC80k, MC80l, MC80m, MC80o, MC80r, MC80t, MC80u, MC80s, MC80x]. Time [MC80y, MC80z, Mei97, Mey76, Mey85z, Mic97, Mic98, Mit76j, Nee97, Neu76a, New76b, Pad11, Pal12, Pal13, Pfa84, Plo76, Pos95b, Pro17a, Pro94, Rot87f, Rot87j, Rou16, Rus76c, Sam76c, Sch76e, Sei97, Sha84q, Shr97, Smo03, Sta94a, Ste03, Tas17, Teg97, Tel76, Tur16, Van86, Von84a, Wia79a, Wol76x, Yin16, dB84, dB87s, dB87t, dR17, Bel12g, Ras92]. Timekeeping [And03a]. Times [AB14, Bar97, KL16]. Timing [Nic17]. tin [Onn91h]. tirer [CLR^{+}92]. Tissues [GWT00c]. Today [AC82, Aga03i, Kap80-42, Mar79a, O’H09, Pes07, Sch88, Wol76k, Wol76-32, Hac17]. Todays [dB87t]. Together [Beu91, Das07a]. Toil [And03b]. Token [CS79q]. Told [Rit05]. Too [Wol76-40]. Tool [Aub17, Bar17, Hac94b, LA86, LC17a, LC17b, Pos95m, Sce01, Sør17, Wim15]. Tools [Joh06, Kle01a, Kle01b, Lef96, Nye01, Ram01, RT01, Yos98]. Topic [Mat17]. Topics [AV16b, GM76, Mis74, Sim77m]. Topology [DG15, Goo77l, Pad11]. Torricelli [Duh91j]. Total [Grü63, Grü86]. Totality [Kos76b, Kos76e]. Touchstone [Grü76]. Toujours [Loc97]. Toulmin [Jar76, McM74b, Min67]. Tournefort [Das15]. Tower [van91]. Traces [Bec04]. Tracing [LC17b]. Tracking [Ano05o, Sie84, Wed86]. Tract [Fre00]. Tractatus [Gra98, Rad03]. Tractional [Mil15]. Trade [Obr92, Ree14]. Trading [Win15]. Tradition [Ano06y, Ano06-27, Ano06-28, Ano06-31, Ano06-34, Ano06-33, Bur74b, Che97, Ela11i, Fil05, HR74f, Hül96, Nol03b, Pǎr15b, Pat92, Sha84g, Sim99, Spr91b, Zam92j]. Traditional [Bu96, Car74, D’A92, Kow16d, Pan83b, Pan91i, Zin83a]. Traditions [Aga08a, Fre00, ZC88i]. Trained [FQ18]. Training [Aga03-28, Buc91, Cas14, Jon91, dJAO14]. Traité [Akb11, Sch01]. Traits [Nor12, Ste15a, Web12]. Trajectories [Ano08-29, Ber96a, Eel10b]. Transcend [Lao17]. Transcendental [Alb84, Ano06-30, Hin89, Koh98g, Mar83b, vZ11]. Transcription [Ano07d]. Transfer [Kah91, Wan18a]. Transfers [Nak95]. Transformation [Ada01, CN11b, Cul11, Gün92, Ihd79k, Pes91, RDS^{+}18, Wol95l]. Transformations [CS79-45, Gab01, GR89, Kno89, Mar79r, vL16d].

Transforming [SG11]. Transgressions [Ano08i]. Transition [Hah91, Lev16, Lop92, Nov74, Wei01a, dB87g]. Translated [BV97]. Translating [Caw94]. Translation [Akb11, Ant00, Gro01, Haq94h, Har99, Kau14, Pan91d, Seg00, Too04, Wal92e, Wol17]. Translations [Lev71, Web11]. Translator [MM94]. Transmission [Bré14, Cul11, Fre00, GR11b, vL16f, vL16h]. Transport [Bow76, Grü73o, Ree14]. Transvaluation [Reh99, Tau95]. Travelling [Dol03, GG03, Rai11]. Travelogue [Kil03]. Travels [CALP03, Pat03a, SCD03a, SCD03b]. Treated [Lag97j]. Treatise [Boc96, Gal92a, Gal92b, Gat18, Vol14]. Treatises [BDR04, Wal92e, Wal92c]. Treatment [Fin84, Hel00c]. Tree [Art15, Kat14]. Trends [FL09, GCN94, Max76a, Ott94, Sie63]. Trial [Aga81q, Fin09m, Siv84, Wal84]. Trials [Sch83d]. Triangle [Qui00a, Ung15]. Tribal [Kel94]. tribute [AM01b]. Trigonometry [Yav17]. Troisième [Dor92]. tropical [Tho95]. tropics [WRL95]. Troubled [Ric15, Svo15]. Troubles [Ams75i]. Trouton [Jan03b]. True [Hin95, Koh98h, Kra97, Mar95c, MW76, Mor18, Oli16, PR17]. True-in-L [MW76]. Trumbull [Jaf02]. Trust [Kow16l, Rod15]. Truth [Alb78, And99, Bab99d, Bab02b, Bac99, Bam99, Ber00, BG18, But93i, Can12, Cha10, Coh02m, Con99, Dav97, Dev15, Fin80s, GC10, Gat09, Gre91b, Grü05b, Hin76, Hüb78a, KS95, Kra95, KSW97, Lev83c, Mar69, Mar79o, Mar79s, Mar79u, Mar79m, Mar79v, Mus96, Nar09, Nic16, Nor91, Now82, Pen99, PP15a, PP15b, PP18, Rad12b, Ric91, Ric07, Ric02, Rou95, Rut91b, San15, Sch99a, Sim99, Ska87, Tho76, WØ97, Wir95, Wol17, Zhe96, Zim99, dCB11]. Truth-Definitions [Hin76]. Truth-Theoretical [Coh02m]. Truths [Ano08m, Bun10b, Pan16, San81]. Trying [Wol76-27, dBJD95, Wol95m]. Tu [Aga81v, Lih93a]. Tunnel [Hoo99k]. Tunneling [Lea96]. Tunneling-Time [Lea96]. turca} [BAK11]. Turin [Seg03]. Turing [Par17a, CL17, FB17, Flo17b, Ken17, MS17, PR17, Pro17b, Res17b, San18, Ste17]. Turkey [İhs05]. Turkish [Akb11, IG05c]. Turn [Bla94, Bla04, Che97, Hor93, STNW12]. Turning [Dis06]. Turning-Point [Dis06]. Turns [Wet09]. Turrell [Sch10]. Tuscan [Seg91]. Twardowski [Coh94d]. Twelfth [Kel10, McK75, Ras75]. Twelve [Coh02u]. Twentieth [D’A00, SDG15b, D’A01l, Ara15, Cro92b, Irz15, Jhu07, KF86, Kuz87n, Lub15, Svo15]. Twentieth- [Kuz87n]. Twentieth-Century [Ara15, Cro92b, Irz15, Lub15]. Twenty [CW85, Arm18, Lek15f]. Twenty-First [Arm18, Lek15f]. Twenty-five [CW85]. Twilight [All99b, RSS^{+}07a, RSS^{+}07b, RS07b]. Twisting [PP03]. Two [Aga03-34, AV97, Ano09-48, Ber76c, Ber95b, Čap71-39, Čap74, Čap91q, Car86, Cha10, Che10, CHS97b, Cur89a, Die14, Fey67, Fin09b, Fra86, FC16, Fre00, Gab76, Gre74s, GHQ96a, Guc81, Hac17, Hal91f, Hat17, Hon98, Kir76, Kün95, Lac94, Lar13, Lei03, Lih93b, Lin93b, MC80z, Moc91, Mor71, Mul74, Pec94, Phi90a, Pie16, Pos95p, Rad12e, RH92, Rog81, Shi76, SN87, SS74, Sie82, Sta07a, Stu97, Tab03, Tha86c, Tur86m, Wat76, Wet09, Živ79, dB87i, Mit76a, Tan01, Wal76]. Two-By-Two [Mul74]. Two-Dimensional [Gab76, Pec94]. Two-Slit [Guc81]. Two-State [AV97]. Two-Tier [Pos95p]. Twofold [Ela11t]. Tyndall [Saa17]. Type [Ano06y, Ano06-27, Ano06-33, Dum96, Hin76, PP03, Sol08a, Vin92]. Type- [PP03]. Types [Čap74, Čap91q, Cas81, Coo17, Duh02e, Duh02g, Gav94, Gru69b, Hör89, ZC88i, Živ79]. Typologie [Pye92]. Typology [Hon09a, Pye92]. Tyrants [Fey95b].

U.S.S.R. [Kap80c]. Ubaldo [Duh91k]. Ubiquitous [Dav17]. Ukrainian [Luc88]. Ultimate [Cas96, Kan00, Sce01]. Ultra [dB87t]. Umbrellas [Voi96]. Umdeutung} [Wüt16]. Unattainable [Aga74a, Aga81r]. Unavoidable [Jer15]. Unbestimmten [Hof74]. Uncertainty [Góm15b, Mac17, Roz71]. Unconscious [Das07b]. Uncut [SM91]. Undecidable [Mor18]. Undecided [You69]. Underdetermination [Ano08-48, Ano08-49, Bon08, Cus93, Pie12]. Underdeveloped [Bar14a]. Underdevelopment [Gil88]. Underestimated [VN12].

Undergraduate [Wel15]. Underlying [vH77d]. Underpinning [Kar10d]. Underpinnings [Hon98, Koe09]. Understand [Ano09-42]. Understanding [Ban15, Boo12, Bur80, DV92, Fin80m, Fin80q, Fit74, Fol18, Ges74t, Gra83i, Gre74t, Hör91a, Kow16m, Kuz87n, MAM69, Nai95, Pos95q, VD92, Von84b, Wal92j, War79k, War79n, ZC88i, GSW95b, Ber17b]. Understood [Wal92i]. Undertakings [Sim15]. Undisputed [Zam92d]. Undying [Hoo99l]. Uneasy [Ric15]. Unexpected [Nov74]. Unfathomed [Bar89]. Unification [Ban15, Bun93e, Mac12i, Sig94]. Unificationism [Sch07d]. Unified [Ano07o, Cau76, D’A01e, GR03, SM74]. Uniformity [Mil95]. Union [Bed09b]. Unique [Eks84, Ell03]. Uniqueness [Ber96a]. Unit [McM76a]. Unitary [Mur75]. United [Meg74, Obr92]. Units [D’A01o, Dal89, RG89, Sob91, Wim80]. Unity [Aga69b, Aga75t, Ano08-50, Ant90, But93g, Čap71-40, Chu91, Daf96, Dvo91, GR03, Hei94a, Hör91a, Kap80-43, Nem91b, Nic76, Odd89b, Sch74, Sha84s, Skl74, Sti94, Ver09, vW84, Zol89i, vW69]. Universal [Bel12j, Čap71n, Coh02l, Gro82a, Hol96, Mey85-27, Pos95m, dB87s, Sas03l]. Universali [Sas03k, Sas03j]. Universalism [Tru94]. Universality [Dav17, Zin73w]. Universalization [Bel92a]. Universals [Ant16, Arm88, BE88, Fey95b, Kri90, MP16]. Universe [And03a, Bou17, MC80z, Nor91, Sch07b, Ton73c, Duh76b]. Universes [Bun12n]. Universities [Gov15, OJ14, Sch75, SDG15b, Svo15, Vod15]. University [Aic15, Car15b, Cho14, Chr94b, CS79n, Elk10, Kat14, Kra15a, Lub15, NG15b, Pal15, PM14, Rhe15, RR15, Tan01, BIT01h, Bog15b, DCS15, Hon96b, Hoo99c, Jah94, Kóh90, Las15, LP01, Mot15, dMPC14, Pet92a, Sch03, Sta91a, Vil02e, Wel15, vD15]. Unjust [Bun12q]. Unknown [Hin71]. Unlearned [Gro82h]. Unlocked [Dav95]. Unnatural [Hon93, Jar86g]. Unobservability [AP00b]. Unobservable [AP00b, Fay00, Wil00]. Unphilosophical [CS79-46, Fey91]. Unpredictability [Res83]. Unpredictable [Web11]. Unpublished [BDR04].

Unreality [Čap71-41]. Unrolling [Lek15g]. Unsafe [SF94]. Unspecified [Hof74]. Unsteady [Ede88]. Untying [JR07]. Upadeśa [Ano06-32]. Upbringing [Kap80-42]. Update [Sal92c]. Upon [Bur74a, Lag97j, Den99i, PO91, Tru82]. Uppsala [CS79g, CS79g]. Urea [NG09]. Ursula [Bal03, Gil03, Irz03]. Uruguay [Gli01]. Uruk [Cla14]. USA [Res11]. Usable [Bar91b]. usage [BLA95a]. Use [Bra95, Bra74b, EPL69, Kno05, Kra00, Lag97i, Saa17, Tho10, Wit86, Hal71, Onn91p, Wal92c]. Used [Wet17]. Usefulness [Bre74b, Čap71y]. Uses [Gro01, McG87, Min10, TG88, War79b]. Ushenko [Čap71p]. USSR [VG87]. Utilitarianism [Sma89b]. Utility [Ber86, Kle12, Mic71, Mon09, WG11]. Utopia [EF86, Gel87, Gol79, Man86]. Utopian [Jar86z]. Utopias [Nem91a]. Utrecht [Vis92].

V [Oli16, Car10, Onn91e]. V. [Rub88]. Vacuum [Kuz87j, Par15a, PP03, Onn91p]. Vagaries [War79j]. Vague [Par83]. Vagueness [Ton81m]. Valency [Duh02g]. Validation [Aga81u, Aga03-35, LW11, Rey18, Sch83d]. Validity [Bit97, Pol95]. Valuation [Bun12e, Hem83, Hem86]. Value [Dra88, Elg84, Hag16, Jie96, Kom99, Kow16d, Kow16i, Lac94, Owe99, Par01, Rut91a, Scr74b, Zim83, Onn91j]. Value-Free [Kom99]. Value-Imposing [Kom99]. Value-Judgment [Dra88]. Value-Neutral [Par01]. Valued [Mir15]. Values [Blu16, DS18, Far03b, Far03i, Far03j, Gon15a, Gon15b, IdMM16, Jer15, Mer97, Mic83, Moo90, Nei15, Nor13, Pap74b, Rol16, Rot87c, Tau95, ZC88f, vdP15]. Vanishing [Hat80]. Varela [Tau91f]. Variable [Les17, Sie63]. Variables [Bun95a, Rob03, Sim77a]. Variance [Odd89a, Pan83d, Pan91f, Pan91h]. Variation [Kir15, Lag97f, Ste09c]. Variations [Coh74a, Mel03]. Varieties [BMM17, Ges74y, GR03, Koh98i, Per91]. Variety [Fra09b]. Varignon [Bla04]. Various [AT15, Lag97q, Lag97r, Lag97s]. Vaughn [SPB18]. VC [Coh94g]. Vector [Rei96]. Vectors [AV97, Wal07b]. Vedānta [ZC88h, Ano06a]. Vehicle [Čap71f, Gra83c]. Vehicles [Lev86]. Velikovsky [Fri76, Pat76a]. Velocitatum [Syl08].

Velocity [CS79-47, D’A81]. Venetian [Val10e]. Venom [Sch12a]. Verein [Sta92a]. Verifiability [Ell84, Rad12f]. Verification [You69]. Verificationism [Mon96]. Verisimilitude [And78, Nii89, Nii17]. Versa [Mei17a]. Versified [Bré14]. Version [Aga86, Bar08, Gan15, Llo86, Ly08]. Versions [Rap81b, Stu97]. Verstehen [Sch99c, McC74]. Versus [Abb94, Aga69a, AD94, Ano09-43, Ara94, BA09, CG16, Coh88a, Coh02c, Fin80g, Fin09h, Ger95, GG96a, Grü76, Kai96, Lei83, Mic71, Noé92, Pat76a, Pos95a, Pos95e, Rus13, Sar91, Söd07, SL84, Ver09, Zol89e, Atl86, Fol96, Wat96]. Very [Aga13q, Lag97t, BBO91, Onn91o, Onn91c, Onn91d, Onn91e, Onn91n, OP91a, WO91]. vessels [Onn91p]. Veszprém [HFD88]. VI [Onn91g]. Via [And84b, Cha09f]. Viability [Gen83]. Vice [Mei17a]. Vices [EP17]. Vicissitudes [Boc96]. Victorian [McG87, But93a, Kid17, Lek15b, Saa17]. Vienna [BIT01h, Bla95b, BIT01c, Cas95b, Coh94g, Fis99, Hal91a, Köh91a, Mil11a, Rut91a, Rut91b, Sta91a, Sta91b, Ueb91b, Ueb95]. Vietnamese [Vol14]. View [Alt14a, Ano05g, Ano09o, Bou17, Čap84a, Dah90, Del91, Des76, Duw17, Fin09l, Fre86f, Gir92, Hon09a, JX96, Kim11, MG11, McC02, Mon96, Nak95, Noé92, Pec94, Prz95, Rad89a, Sas95, Seg15, Tri08, Wat87, Wei76, Zal86, vdW04]. View1 [Gau84]. Viewed [Rim97]. Viewpoint [Duc96, Rad81, Som81, d’E08, vH77c, Web08]. Views [Aga81-34, And99, Aug76, Čap71v, CS79-29, Coh02r, D’A98, Dor97, Far03c, Fer85f, Fre88, Hie70, Ker88, Nol03k, Pat03b, Sam76c, VD92, Wei11, Woo91, WC91]. Vignettes [Lac94]. VII [Onn91i]. VIII [Onn91h]. Vinci [Duh91i, Duh91o, Ret74, See74, Sti74]. Vindicated [Cha09e]. Vindication [Fil17, Wei84]. Violated [Nic80a]. Violations [dB87o]. Violence [Jar86i]. Virgin [SS74]. Virtual [Jaf02, Joh06]. Virtue [De 17, How15, Jaf02, Pau17, Wan17]. Virtues [Kar17, Saa17, Sta17, War79q, vDP17a, vD17, vDP17b]. Virtuous [Ber13c, vD17]. Virus [Eil95]. Vis [Bro95]. Vis-a-Vis [Bro95]. Viscous [Cob74]. Visible [Bal14, Hem84]. Vision [Ber96b, Bra91, CK97, Gre95, Hes96, Ihd79l, KT76, Kre91, McB94, ZC88d]. Visionary [Sch96f].

Visions [DCS15]. Visual [Elk10, GF97, Grü73i, Hef87, O’D02, PS01a]. Visualizing [RS06]. Vital [Hei08b]. Viva [RH14]. Vocation [Cap81b, DS18]. Voice [Bau17]. Void [Bai76b, Duh76c, Duh76f, Ela11q, Gas76b]. Voigt [Kat06f]. Volitionism [Kam11]. Volterra [KL16]. Volume [BRW14b, Sta09, UM88, UM92, Coh80, CHS97a, CHS97b, JNR^{+}07b, JRN^{+}07b, JL95a, JL95b, MC80z, RSS^{+}07a, RSS^{+}07b, Sch08a, Sch08b, UM86a, UM86c]. Volumes [Ano07p, Ano07q, JNR^{+}07a, MC80e]. Voluntary [Lai08]. Von-Humboldt-Forschung [Bie74]. Voreudoxische [Bec04]. Vorüberlegungen [Sch99c]. Voyage [ZC96]. vs [BIT01e, Bol95a, Cus93, Dil06b, Dil06c, Din96, Grü73b, Grü78, Hon98, Jos17, Koc97, Mar78d, Men76, Mis95c, Pos78, Rad79, Ren96, Ska83, Stu97, Uga18, War67b, Wie92, Wim92].

W [Nol91, Str94]. Waerden [Dor96]. Waitz [Pau17]. Wake [Sim94]. Walking [Far03b, Far03k, Mas15b, Sun18]. Wall [Woo91]. Walter [Smi94, vZ11]. Walther [Bar91a]. Wang [Hor93]. Waning [Ric17]. Want [Tis03]. Wants [Kam11]. War [Ano08-32, FS94, Hat95, Hei08a, Hei08b, Kap80q, KJ14, Rat88, Vog91, RR15, Wat01a, Ada01, Dol86, Gie86, Lee86, Mac00]. Warranted [Lef03]. Wars [Gar12b, Las15]. Warsaw [Coh94d, Coh94f]. Wartofsky [GC94, Nic80b, Shi87f]. Was [Aga81-35, Bou02i, Eck17, Gab04, Hal90a, Jol18, Nor07e, Ros86, Sta92a, Val10g, Lau80a, Fre04, Kis02, Sch99c]. Wassermann [Zal86]. Waste [Hof09, Jer15, SF94]. Watch [BAK11]. Water [Bou02b, Cha12g, Cha12e, Rav18, Wei06]. Watkins [DJ89, Hüb78a, Mun89, Wor89, Zah89]. Watson [Sch76b]. Watts [CS79-41]. Wave [BEW96, Cob74, D’A01g, Joh96a, Val96]. Wavefunctions [ACE17]. Waves [Bry98, D’A01j, D’A01i, Don98]. Way [Ano09-42, Fle86f, Hoo91, JR07, Lag97i, Ren07d, Wer01, WG11, Wil07]. Ways [Hal91f, Mac97f, Rai11, Wor78b]. Weak [MM91]. Weaknesses [Zol89h]. Wealth [Bun12u]. Weapons [Bal15, O’T88]. Webb [Shi87g]. Weber [Aga81o, D’A81, D’A01b, Pau17, Tur86l].

Wechselwirkung [Mül94]. We’d [Voi96]. Weighing [BRS18]. Weight [Ano08q, Coh02u]. Weighted [Vol14]. Weights [Coo90, Gro82k]. Weimar [Ano08-45]. Weir [Sch12a]. Weird [Bun10p]. Weismann [Chu01, Heg91]. Weizsäcker [OP82, Zuc69]. Welcome [Bad92]. Welfare [Ano05n]. Well [Bre11, Bun12u]. Well-being [Bun12u]. Werner [Hen98]. Wernicke [Ges69b, Ges74e, Ges74z]. West [Coh17a, Cul11, GHQ96b, MLNZ18, NNW74, Tho14, Xiu96]. Western [Ano06c, Ano06-33, Bre11, FQ18, GC06, GG03, Hsü92, Jam92, Vam09g]. Western-Trained [FQ18]. Weyl [Maj98, Sch06a, Sig94]. WG [Kow16c]. Whence [Var92]. Where [Gig03, GG96b, Wil00]. Whewell [But93i, But93n, But93m, Coh76a, Lug89]. Which [AV96, GW97, Lag97l, Mar95c, Smi95b, McC95a, Onn91g, Onn91j, Wet17, Wor78b, Cah98, Koh98j]. Whiggism [Bal94]. Whitehead [Čap71q, Čap71-33, Čap91h, Grü73q, Grü73w]. Whiteness [Wes87]. Whither [Qui76]. Who [Aga97a, Gav09, Ito98, Kle91, Sta92b, Wei04, Woo02, Zim86]. Whole [Ano05e, Byn89, Gre94, Kle91, SA12]. Wholly [Han91]. Whose [Cah98, Mil11b, Koh98j]. Wide [Tur12]. Widening [Čap91p]. Wie [Sza04]. Wiener [Hon96b, Wol97]. Wiksell [CS79g]. Wild [Aga03x, Aga03-36]. Wilder [Duh02t]. Wilfrid [Rot87k]. Wilhelm [Ano08o, Ano08-45, Ano08-40, Ano08-32, Sch08d, Mor80]. Will [Bac99, Bun10e, Cao07, Gar12b, Meš89, Nem91a, Pro17b, San14, Sel74, Ste99]. Willem [Rey18]. William [Coh76a, Fle91b, Hil94, Syl75, ZRKC03g]. Williamson [Tau91g]. Wimsattian [Krö12]. Window [Bog15a]. Wine [Gol87]. Wing [Hor88]. Wings [Voi96]. Wire [Don98]. Wireless [Ara15]. Wisconsin [CS79n]. Wisdom [Bu96, Fis95, Ols87]. Wissenschaft [Sza04, Mit04]. Wissenschaften [Bie74]. Wissenschaftliche [Swi96]. Wissenschaftsauffassung [Sch99d]. Wissenschaftsgeschichte [Ape99]. Witch [Hie95b]. Witch-Hazel [Hie95b]. Withering [Boš79]. Within [Bam99, McC02, Nar09, Pan91i, AZ87, Sas95]. Without [BM09, Das91, Gor07, Mar78f, Swa09, Aga03l, Čap71f, Car89a, Car00, Cor00, Hat83, Koç05, Krü89, LvHV92, Onn91l]. Witness [Cer14]. Wittgenstein [Sma92, Tsi90, Aga81-35, Bar74, Cas95a, Ger95, Gib00, Gra98, Hin94, Hin97, Jac97, Jan89e, Jan89o, Jan02, Kou90, Laz77, Nol03m, Nyí97, Pis15, Ruf69, Ueb95, Vis01]. Wittgensteinian [Gon96, Jan89n, Wol69]. Wittgensteinians [Rad82]. Wöhler [Pau17]. Wolniewicz [Ruf69]. Woman [Gar16, Hol07]. Women [Gov15, Hag16, Vro90, dF14b]. Wonders [Kil03]. Wood [Hie95b]. Word [Mül94, Mur12d, Pai95]. Words [Che94, Rad03, Sch95a]. Work [Abr81, Amr87b, Ano08-42, Bun12f, Cai14, Can07, Cha09b, Coh94m, Duh02x, EP17, Fey95a, Fre86r, Ges69b, Goo95, Gru80, Har02, Hod90, Kap80f, Kap80g, Kap80-41, Kat06b, Knu07, Lip17, Mac97i, Mai12, MC80u, Nee11m, Row96, Sch07a, Sch10, Sch96h, Sha88, She90, Skl06, Sto75, Sch08a, Sch08b]. Worker [Byr86, Hei17]. Working [Das07a, Fef93, Jan89m, Lau80b]. Works [Bar92, Hoo99m, Löw86, Rhe05, Sch96d, Too04, War92, Zam92d, KV11]. Workshop [NG01f]. World [Aga03h, Ams92d, AM83, Čap71a, Čap71h, Čap71-34, Coh86e, Coh02t, Dis06, Dou99, Far03c, Fer85d, Fin09l, FS94, Fre86f, Fre91, Gil88, GPSR01, Gro09b, Hat95, Het15, Hil96, Höf76, Irz95, Kae90b, Kap80u, Koh98b, Kos76a, Koy76a, Lan67, Las15, Lom15, Mar86h, Mar01, Mau96, Mit89, Moo67, ML99, Nai95, Noé92, Ohm98, Pol92, Que97, Raf15, Rap81f, Rip11, Sch98, SA12, Tou76, Vrc79, Wal95, Woo91, WC91, ZC88b, Ano08-32, Vog91]. World-Science [Pol92]. World-Views [Far03c]. Worlds [Hec01, Kom99, Mar95a, Ram01]. Worldview [Bri18, Bun10o, Ghi93, Pau97]. Worried [Gar12b]. Worries [Mar83a]. Worrying [Ant16, MP16]. Worst [Don86]. Worum [Vaj99]. Wright [Cro92b, Hod92, Pro92]. Write [Pro14]. Writer [Škv88]. Writing [Ano06-34, Ges74-27, Jhu07, Las01, Rhe05, Wol95l]. Writings [Coh94m, HOO88b, Köh91b, vH77a]. Written [Fil05, Jam10, PS01a, Pol92]. Wrong [Fin09j, Fre09, Nol03d, Wet09].

Xenophanes [Fey84b, Vam09o]. XIe [Ras75]. XIIe [Gre75, Ras75]. xiii [CS79n, PO91]. XIXe [Sal92a]. Xun [Hor93]. xvi [CS79g]. xxiv [CS79-41]. XXth [Cin89].

Yang [Koh98k]. Yangzi [BE10a]. Year [Nik15]. Years [CS79v, CW85, DB15b, Ger15a, Gie12, Hod90, Jin96, Kar10h]. Yehoshua [Kas76b]. Yehuda [Wie95]. Yehuda-the-Prince [Wie95]. Yes [Sim77v]. Yes-No [Sim77v]. Yield [Gat09]. Yin [Koh98k]. Yirmisekiz [Kil03]. Yongjing [Chu11]. Young [Fin76, NA07, She82, Vil02e, Kah91]. Youth [Kap80-42]. Yuan [Lih93a]. Yuanpei [Pin96]. Yugoslav [MP79, Sup79]. Yugoslavia [Cek92a]. Yukawa [CS79x]. Yunnan [Hsü92].

Zahar [Gra95]. Zawirski [Coh94m, Coh94l]. Zeeman [Hen98]. Zeno [Ber76a, Grü73t, Rus76d, Sky83]. Zhang [Chu11, Zum96]. Zheng [ZC96]. Ziele [Bie74]. Zilsel [Kri03]. Zira [Gün92]. Zira“ [Gün92]. Zone [Win15]. Zoologique [Osb92]. zt [Fet81e]. Zupi [Gat18]. zur [Hof74, Vog04]. Zurich [JNR^{+}07b, JRN^{+}07b, Ano07d, JRN^{+}07a, JNR^{+}07a, JR07]. Zwirn [Le 08]. zwischen [Sch99d]. Zygmunt [Coh94m, Coh94l]. Zymes [Tul16].

References

Abir-Am:1986:KBC

[AA86]  Pnina Abir-Am.  Koch’s bacillus:

a comment. In Ullmann-Margalit [UM86c], chapter 12, pages 161–164. ISBN 90-277-2160-2, 90-277-2161-0 (paperback), 94-009-4566-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 95

                                                        Q175. URL https://link.springer.
                                                        com/chapter/10.1007/978-94-009-

4566-1_12. See [Ros86].

Altshuler:2014:CTP

[AAF^{+}14] José Altshuler, Ocatvio Calzadilla Amaya, Federico Falcon, Juan E. Fuentes, Jorge Lodos, and Elena Vigil Santos. Cuban techno-physical experiments in space. In The History of Physics in Cuba [BRW14a], chapter 14, pages 295–300. ISBN 94-017-8040-4, 94-017-8041-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC5.53; QC6.9; QC9 C9H57 2014; Q124.6-127.2. URL https:/

/link.springer.com/chapter/10.
1007/978-94-017-8041-4_14.

Abrahamsen:2012:REA

[AB12] Adele Abrahamsen and William Bechtel. From reactive to endogenously active dynamical conceptions of the brain. In Philosophy of Behavioral Biology [PR12b], chapter 16, pages 329–366. ISBN 94-007-1950-7, 94-007-1951-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QH331 .P4679 2012; Q175 .B657 v.282.

URL https://link.springer.com/
chapter/10.1007/978-94-007-1951-
4_16.

Altshuler:2014:TPC

[AB14] José Altshuler and Angelo Baracca.

The teaching of physics in Cuba from colonial times to 1959. In The History of Physics in Cuba [BRW14a], chapter 4, pages 57–106. ISBN 94-017-8040-4, 94-017-8041-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC5.53; QC6.9; QC9 C9H57 2014; Q124.6-127.2. URL https:/

/link.springer.com/chapter/10.
1007/978-94-017-8041-4_4.

Ardourel:2017:RTD

[AB17] Vincent Ardourel and Anouk Barberousse. The representation of time in discrete mechanics. In Time of Nature and the Nature of Time: Philosophical Perspectives of Time in Natural Sciences [BH17b], chapter 9, pages 173–208. ISBN 3-319-53723-7 (hardcover), 3-319-53725-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC173.59.S65 T56

                                              2017. URL https://link.springer.
                                              com/chapter/10.1007/978-3-319-
                                              53725-2_9.

Abbri:1994:RVE

[Abb94]  Ferdinando Abbri.  Romanticism

versus enlightenment: Sir Humphry Davy’s idea of chemical philosophy. In Poggi and Bossi [PB94], chapter 3, pages 31–45. ISBN 0-7923-2336-X, 94-017-2921-2. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 152. URL https://link.

                                              springer.com/chapter/10.1007/
                                              978-94-017-2921-5_3.

Abbott:2014:EIC

[Abb14] Andrew Abbott. The excellence of IT: Conceptions of quality in academic disciplines. In The Institution of Science and the Science of Institutions: The Legacy of Joseph Ben-David [Her14b], chapter 8, pages 147–165. ISBN 94-007-7406-0 (hardcover), 94-007-7407-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 v.302; Q125 .I58 2014. URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-007-7407-0_8.

Abrusci:1981:PTF

[Abr81] Vito Michele Abrusci. ‘Proof’, ‘theory’, and ‘foundations’ in Hilbert’s mathematical work from 1885 to 1900. In Italian Studies in the Philosophy of Science [Chi81b], chapter 24, pages 453–491. ISBN 90-277-1073-2, 94-009-8937-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

                                              ???? URL https://link.springer.
                                              com/chapter/10.1007/978-94-009-
                                              8937-5_24.

Abraham:2007:FF

[Abr07a] Max Abraham. The free fall. In Renn et al. [RSS^{+}07a], pages 341–345. ISBN 1-4020-3999-9. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 no. 250 v. 3; QC173.6 .G469

2007. URL https://link.springer.
com/chapter/10.1007/978-1-4020-

4000-9_24. Originally published in Rendiconti del R. Instituto Lombardo di scienze e lettere. German translation published as “Der freie Fall”, Physikalische Zeitschrift 13, 310–311 (1912).

Abraham:2007:NTG

[Abr07b] Max Abraham. A new theory of gravitation. In Renn et al. [RSS^{+}07a], pages 347–362. ISBN 1-4020-3999-9. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 no. 250 v. 3; QC173.6 .G469 2007. URL https:

//link.springer.com/chapter/10.
1007/978-1-4020-4000-9_25. Lecture

presented on 19 October 1912 to the Società italiana per il progresso delle scienze. German translation published as “Eine neue Gravitationstheorie”, Archiv der Mathematik und Physik, Third Series 20, 193–209 (1913).

Abraham:2007:TG

[Abr07c] Max Abraham. On the theory of gravitation. In Renn et al. [RSS^{+}07a], pages 331–339. ISBN 1-4020-3999-9. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 no. 250 v. 3; QC173.6 .G469 2007. URL https:

//link.springer.com/chapter/
10.1007/978-1-4020-4000-9_23.

Originally published in Rendiconti della R. Accademia dei Lincei. German translation published as “Zur Theorie der Gravitation”, Physikalische Zeitschrift 13, 1–4, 176 (1912).

Abraham:2007:RTG

[Abr07d] Max Abraham. Recent theories of gravitation. In Renn et al. [RSS^{+}07a], pages 363–410. ISBN 1-4020-3999-9. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 no. 250 v. 3; QC173.6 .G469 2007.

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-1-4020-4000-

9_26. Originally published as “Neuere Gravitationstheorien”, Jahrbuch der Relativität und Elektronik 11, 470–520 (1915).

Abrantes:2011:HEC

[Abr11] Paulo C. Abrantes. Human evolution: Compatibilist approaches. In Brazilian Studies in Philosophy and History of Science: An account of recent works [KV11], chapter 12, pages 171–183. ISBN 90-481-9421-0 (hardcover), 90-481-9422-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175.3 .B73 2011. URL https:/

                                              /link.springer.com/chapter/10.
                                              1007/978-90-481-9422-3_12.

Agassi:1982:SPT

[AC82] Joseph Agassi and Robert S. Cohen. Scientific Philosophy Today: Essays in Honor of Mario Bunge, volume 67 of Boston Studies in the Philosophy of Science. D. Reidel, Dordrecht, The Netherlands; Boston, MA, USA; Lancaster, UK; Tokyo, Japan, 1982. ISBN 90-277-1262-X, 90-277-1263-8, 94-009-8462-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). x + 513 pp. LCCN Q174 .B67 vol. 67.

                                              URL https://link.springer.com/
                                              book/10.1007/978-94-009-8462-2.

Accardi:1990:QPF

[Acc90] Luigi Accardi. Quantum probability and the foundations of quantum theory. In Statistics in Science: The Foundations of Statistical Methods in Biology, Physics and Economics [CC90], chapter 8, pages 119–147. ISBN 94-010-6765-1, 94-009-0619-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QA276-280.

URL https://link.springer.com/
chapter/10.1007/978-94-009-0619-
8_8.

Aceves:1992:FCC

[Ace92] Patricia Aceves. The first chair of chemistry in Mexico (1796–1810). In Science and Empires: Historical Studies about Scientific Development and European Expansion [PJM92], chapter 15, pages 137–146. ISBN 0-7923-1518-9 (hardcover), 94-010-5145-3, 94-011-2594-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 136. URL https://link.

springer.com/chapter/10.1007/
978-94-011-2594-9_15.

Aharonov:2017:BWT

[ACE17] Yakir Aharonov, Eliahu Cohen, and Avshalom C. Elitzur. Beyond wavefunctions: A time–symmetric nonlocal ontology for quantum mechanics. In Encouraging Openness: Essays for Joseph Agassi on the Occasion of His 90th Birthday [BAG17], chapter 19, pages 235–239. ISBN 3-319-57668-2 (hardcover), 3-319-57669-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN B833.E53 2017.

URL https://link.springer.com/
chapter/10.1007/978-3-319-57669-
5_19.

Achinstein:1967:CAP

[Ach67] Peter Achinstein. Comments: Acute proliferitis. In Cohen and Wartofsky [CW67], pages 416–424. ISBN 90-277-0013-3, 94-010-3508-3. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175 .B731 v.3. URL https://link.

                                              springer.com/chapter/10.1007/
                                              978-94-010-3508-8_23.

Achinstein:1980:DRB

[Ach80] Peter Achinstein. Discovery and rulebooks. In Scientific Discovery, Logic, and Rationality [Nic80f], chapter 3, pages 117–137. ISBN 90-277-1070-8, 94-009-8986-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

                                              ???? URL https://link.springer.
                                              com/chapter/10.1007/978-94-009-
                                              8986-3_3.

Ackerman:1992:AP

[Ack92] Felicia Ackerman. Analysis and its paradoxes. In The Scientific Enterprise: The Bar-Hillel Colloquium: Studies in History, Philosophy, and Sociology of Science, Volume 4 [UM92], chapter 11, pages 169–178. ISBN 94-010-5190-9, 94-011-2688-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .S354 1992. URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-011-2688-5_11.

Anapolitanos:1994:OPP

[AD94] Dionysios A. Anapolitanos and Apostolos K. Demis. The one in the philosophy of Proclus: Logic versus metaphysics. In Trends in the Historiography of Science [GCN94], chapter 13, pages 169–175. ISBN 90-481-4264-4, 94-017-3596-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

                                              ???? URL https://link.springer.
                                              com/chapter/10.1007/978-94-017-
                                              3596-4_13.

Adams:2001:NAK

[Ada01] Mark B. Adams. Networks in action: The Khrushchev Era, the Cold War, and the transformation of Soviet science. In Allen and MacLeod [AM01b], chapter 16, pages 255–276. ISBN 1-4020-0495-8 (hardcover), 94-017-2956-5. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 228; Q175.55. URL https:

//link.springer.com/chapter/10.
1007/978-94-017-2956-7_16.

Adam:2011:MLC

[Ada11] Matthias Adam. Multi-level complexities in technological development: Competing strategies for drug discovery. In Science in the Context of Application [CN11a], chapter 5, pages 67–83. ISBN 90-481-9050-9, 90-481-9051-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q180.55.M4 S385 2011. URL https:

//link.springer.com/chapter/10.
1007/978-90-481-9051-5_5.

Addis:1974:MTS

[Add74] Laird Addis. Monistic theories of society. In Schaffner and Cohen [SC74a], pages 209–216. ISBN 90-277-0408-2, 90-277-0409-0, 94-010-2140-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175 .B7312.

URL https://link.springer.com/
chapter/10.1007/978-94-010-2140-
1_14.

Addis:1976:DCL

[Add76]  Laird Addis.  On defending the

covering-law “model”. In Cohen et al. [CHMV76], pages 361–368. ISBN 90-277-0647-6, 90-277-0648-4 (paperback), 94-010-1449-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic).

LCCN ????  URL https://link.
springer.com/chapter/10.1007/
978-94-010-1449-6_17.

Arabatzis:2016:DDH

[AG16] Theodore Arabatzis and Kostas Gavroglu. From discrepancy to discovery: How argon became an element. In Sauer and Scholl [SS16b], chapter 10, pages 203–222. ISBN 3-319-30227-2, 3-319-30229-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q126.8 .P45 2016. URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-3-319-30229-4_10.

Agassi:1967:SF

[Aga67] Joseph Agassi. Science in flux. In Cohen and Wartofsky [CW67], pages 293–323. ISBN 90-277-0013-3, 94-010-3508-3. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175 .B731 v.3.

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-010-3508-
                                              8_16.

Agassi:1969:CTE

[Aga69a] Joseph Agassi. Comments: Theoretical entities versus theories. In Cohen and Wartofsky [CW69a], pages 457–459. ISBN 94-010-3383-8, 94-010-3263-7, 94-010-3381-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

                                              ???? URL https://link.springer.
                                              com/chapter/10.1007/978-94-010-
                                              3381-7_16.

Agassi:1969:UDS

[Aga69b] Joseph Agassi. Unity and diversity in science. In Cohen and Wartofsky [CW69b], pages 463–522. ISBN 94-010-3380-3, 94-010-3378-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-010-3378-7_17.

Agassi:1974:PU

[Aga74a] Joseph Agassi. On pursuing the unattainable. In Seeger and Cohen [SC74b], pages 431–444. ISBN 90-277-0390-6, 90-277-0376-0 (paperback), 94-010-2126-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 11. URL https:

//link.springer.com/chapter/10.
1007/978-94-010-2126-5_25.

Agassi:1974:OSS

[Aga74b] Judith Buber Agassi. Objectivity in the social sciences. In Seeger and Cohen [SC74b], pages 305–316. ISBN 90-277-0390-6, 90-277-0376-0 (paperback), 94-010-2126-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 11. URL https:

//link.springer.com/chapter/10.
1007/978-94-010-2126-5_16.

Agassi:1975:AA

[Aga75a] Joseph Agassi. Assurance and agnosticism. In Science in Flux [Aga75p], chapter 20, pages 515–523. ISBN 90-277-0612-3, 94-010-1810-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-010-1810-4_20.

Agassi:1975:CRG

[Aga75b] Joseph Agassi. Can religion go beyond reason? In Science in Flux [Aga75p], chapter 19, pages 469–514. ISBN 90-277-0612-3, 94-010-1810-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-010-1810-4_19.

Agassi:1975:CBP

[Aga75c] Joseph Agassi. The confusion between physics and metaphysics in the standard histories of sciences. In Science in Flux [Aga75p], chapter 11, pages 270–281. ISBN 90-277-0612-3, 94-010-1810-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-010-
1810-4_11.

Agassi:1975:CBS

[Aga75d] Joseph Agassi. The confusion between science and technology in the standard philosophies of science. In Science in Flux [Aga75p], chapter 12, pages 282–305. ISBN 90-277-0612-3, 94-010-1810-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-010-
1810-4_12.

Agassi:1975:CPA

[Aga75e] Joseph Agassi. Criteria for plausible arguments. In Science in Flux [Aga75p], chapter 16, pages 353–364. ISBN 90-277-0612-3, 94-010-1810-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-010-1810-4_16.

Agassi:1975:IK

[Aga75f] Joseph Agassi. Imperfect knowledge. In Science in Flux [Aga75p], chapter 15, pages 338–352. ISBN 90-277-0612-3, 94-010-1810-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-010-
1810-4_15.

Agassi:1975:MC

[Aga75g] Joseph Agassi. Modified conventionalism. In Science in Flux [Aga75p], chapter 17, pages 365–403. ISBN 90-277-0612-3, 94-010-1810-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-010-1810-4_17.

Agassi:1975:NSP

[Aga75h] Joseph Agassi. The nature of scientific problems and their roots in metaphysics. In Science in Flux [Aga75p], chapter 9, pages 208–239. ISBN 90-277-0612-3, 94-010-1810-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-010-1810-4_9.

Agassi:1975:N

[Aga75i] Joseph Agassi. On novelty. In Science in Flux [Aga75p], chapter 3, pages 51– 80. ISBN 90-277-0612-3, 94-010-1810-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-010-1810-
4_3.

Agassi:1975:PESb

[Aga75j] Joseph Agassi. Positive evidence as a social institution. In Science in Flux [Aga75p], chapter 14, pages 322–337. ISBN 90-277-0612-3, 94-010-1810-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-010-1810-
4_14.

Agassi:1975:PESa

[Aga75k] Joseph Agassi. Positive evidence in science and technology. In Science in Flux [Aga75p], chapter 13, pages 306–321. ISBN 90-277-0612-3, 94-010-1810-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

                                              ???? URL https://link.springer.
                                              com/chapter/10.1007/978-94-010-
                                              1810-4_13.

Agassi:1975:PSF

[Aga75l] Joseph Agassi. A prologue: On stability and flux. In Science in Flux [Aga75p], chapter 1, pages 1–8. ISBN 90-277-0612-3, 94-010-1810-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-010-1810-4_1.

Agassi:1975:QSM

[Aga75m] Joseph Agassi. Questions of science and metaphysics. In Science in Flux [Aga75p], chapter 10, pages 240–269. ISBN 90-277-0612-3, 94-010-1810-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-010-1810-
                                              4_10.

Agassi:1975:RDP

[Aga75n] Joseph Agassi. Replies to Diane: Popper on learning from experience. In Science in Flux [Aga75p], chapter 4, pages 81–91. ISBN 90-277-0612-3, 94-010-1810-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

                                              ???? URL https://link.springer.
                                              com/chapter/10.1007/978-94-010-
                                              1810-4_4.

Agassi:1975:SFa

[Aga75o] Joseph Agassi. Science in flux. In Science in Flux [Aga75p], chapter 2, pages 9–50. ISBN 90-277-0612-3, 94-010-1810-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

                                              ???? URL https://link.springer.
                                              com/chapter/10.1007/978-94-010-
                                              1810-4_2.

Agassi:1975:SFb

[Aga75p] Joseph Agassi. Science in Flux, volume 28 of Boston Studies in the Philosophy of Science. D. Reidel, Dordrecht, The Netherlands; Boston, MA, USA; Lancaster, UK; Tokyo, Japan, 1975. ISBN 90-277-0612-3, 94-010-1810-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). xxvi + 559 pp. LCCN ???? URL

https://link.springer.com/book/
10.1007/978-94-010-1810-4.

Agassi:1975:S

[Aga75q] Joseph Agassi. Sensationalism. In Science in Flux [Aga75p], chapter 5, pages 92–126. ISBN 90-277-0612-3, 94-010-1810-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-010-
1810-4_5.

Agassi:1975:TBO

[Aga75r] Joseph Agassi. Testing as a bootstrap operation in physics. In Science in Flux [Aga75p], chapter 7, pages 155–188. ISBN 90-277-0612-3, 94-010-1810-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-010-1810-
4_7.

Agassi:1975:TTA

[Aga75s] Joseph Agassi. Towards a theory of ad hoc hypotheses. In Science in Flux [Aga75p], chapter 8, pages 189–207. ISBN 90-277-0612-3, 94-010-1810-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-010-1810-
4_8.

Agassi:1975:UDS

[Aga75t] Joseph Agassi. Unity and diversity in science. In Science in Flux [Aga75p], chapter 18, pages 404–468. ISBN 90-277-0612-3, 94-010-1810-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-010-1810-4_18.

Agassi:1975:WSW

[Aga75u] Joseph Agassi. When should we ignore evidence in favour of a hypothesis? In Science in Flux [Aga75p], chapter 6, pages 127–154. ISBN 90-277-0612-3, 94-010-1810-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

                                              ???? URL https://link.springer.
                                              com/chapter/10.1007/978-94-010-
                                              1810-4_6.

Agassi:1976:AA

[Aga76a] Joseph Agassi. Assurance and agnosticism. In Cohen et al. [CHMV76], pages 449–457. ISBN 90-277-0647-6, 90-277-0648-4 (paperback), 94-010-1449-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-010-1449-
                                              6_23.

Agassi:1976:CAB

[Aga76b] Joseph Agassi. Can adults become genuinely bilingual? In Language in Focus: Foundations, Methods and Systems: Essays in Memory of Yehoshua Bar-Hillel [Kas76b], chapter 20, pages 473–484. ISBN 90-277-0644-1, 90-277-0645-X, 94-010-1876-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN P101-120. URL https://link.

                                              springer.com/chapter/10.1007/
                                              978-94-010-1876-0_20.

Agassi:1976:LR

[Aga76c] Joseph Agassi. The Lakatosian revolution. In Essays in Memory of Imre Lakatos [CFW76], chapter 2, pages 9– 21. ISBN 90-277-0655-7, 94-010-1451-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-010-1451-
9_2.

Agassi:1980:RD

[Aga80] Joseph Agassi. The rationality of discovery. In Scientific Discovery, Logic, and Rationality [Nic80f], chapter 7, pages 185–199. ISBN 90-277-1070-8, 94-009-8986-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-009-
8986-3_7.

Agassi:1981:AS

[Aga81a] Joseph Agassi. The autonomy of science. In Science and Society: Studies in the Sociology of Science [Aga81z], chapter 6, pages 68–76. ISBN 94-011-6458-4, 94-011-6456-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-011-6456-6_6.

Agassi:1981:BMM

[Aga81b] Joseph Agassi. Between metaphysics and methodology. In Science and Society: Studies in the Sociology of Science [Aga81z], chapter 19, pages 253–261. ISBN 94-011-6458-4, 94-011-6456-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-011-6456-
6_19.

Agassi:1981:CSP

[Aga81c] Joseph Agassi. The choice of scientific problems. In Science and Society: Studies in the Sociology of Science [Aga81z], chapter 18, pages 239–252. ISBN 94-011-6458-4, 94-011-6456-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-011-6456-
                                              6_18.

Agassi:1981:CDH

[Aga81d] Joseph Agassi. Continuity and discontinuity in the history of science. In Science and Society: Studies in the Sociology of Science [Aga81z], chapter 22, pages 283–299. ISBN 94-011-6458-4, 94-011-6456-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-011-6456-6_22.

Agassi:1981:CLS

[Aga81e] Joseph Agassi. Cultural lag in science. In Science and Society: Studies in the Sociology of Science [Aga81z], chapter 10, pages 119–131. ISBN 94-011-6458-4, 94-011-6456-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-011-6456-6_10.

Agassi:1981:ESP

[Aga81f] Joseph Agassi. The economics of scientific publications. In Science and Society: Studies in the Sociology of Science [Aga81z], chapter 12, pages 142–155. ISBN 94-011-6458-4, 94-011-6456-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

                                              ???? URL https://link.springer.

com/chapter/10.1007/978-94-011-
6456-6_12.

Agassi:1981:EAS

[Aga81g] Joseph Agassi. Epistemology as an aid to science. In Science and Society: Studies in the Sociology of Science [Aga81z], chapter 4, pages 45–54. ISBN 94-011-6458-4, 94-011-6456-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-011-6456-6_4.

Agassi:1981:E

[Aga81h] Joseph Agassi. Externalism. In Science and Society: Studies in the Sociology of Science [Aga81z], chapter 5, pages 55– 67. ISBN 94-011-6458-4, 94-011-6456-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-011-6456-
6_5.

Agassi:1981:FND

[Aga81i] Joseph Agassi. Faith has nothing to do with rationality. In Science and Society: Studies in the Sociology of Science [Aga81z], chapter 31, pages 457–464. ISBN 94-011-6458-4, 94-011-6456-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-011-
6456-6_31.

Agassi:1981:GS

[Aga81j] Joseph Agassi. Genius in science.

In Science and Society: Studies in the Sociology of Science [Aga81z], chapter 15, pages 192–209. ISBN 94-011-6458-4, 94-011-6456-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-011-6456-6_15.

Agassi:1981:IIN

[Aga81k] Joseph Agassi. The ideological import of Newton. In Science and Society: Studies in the Sociology of Science [Aga81z], chapter 26, pages 372–387. ISBN 94-011-6458-4, 94-011-6456-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-011-6456-
                                              6_26.

Agassi:1981:ISS

[Aga81l] Joseph Agassi. Introduction: Science in its social setting. In Science and Society: Studies in the Sociology of Science [Aga81z], chapter 1, pages 1–17. ISBN 94-011-6458-4, 94-011-6456-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-011-6456-6_1.

Agassi:1981:LS

[Aga81m] Joseph Agassi. The legitimation of science. In Science and Society: Studies in the Sociology of Science [Aga81z], chapter 7, pages 77–84. ISBN 94-011-6458-4, 94-011-6456-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-011-6456-6_7.

Agassi:1981:LSI

[Aga81n] Joseph Agassi. The logic of scientific inquiry. In Science and Society: Studies in the Sociology of Science [Aga81z], chapter 17, pages 223–238. ISBN 94-011-6458-4, 94-011-6456-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-011-6456-6_17.

Agassi:1981:MWS

[Aga81o] Joseph Agassi. Max Weber’s scientific religion. In Science and Society: Studies in the Sociology of Science [Aga81z], chapter 29, pages 436–443. ISBN 94-011-6458-4, 94-011-6456-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-011-6456-6_29.

Agassi:1981:MRP

[Aga81p] Joseph Agassi. The methodology of research projects: A sketch. In Science and Society: Studies in the Sociology of Science [Aga81z], chapter 21, pages 273–282. ISBN 94-011-6458-4, 94-011-6456-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-011-
6456-6_21.

Agassi:1981:ETG

[Aga81q] Joseph Agassi. On explaining the trial of Galileo. In Science and Society: Studies in the Sociology of Science [Aga81z], chapter 24, pages 321–351. ISBN 94-011-6458-4, 94-011-6456-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-011-6456-
6_24.

Agassi:1981:PU

[Aga81r]  Joseph Agassi.  On pursuing the

unattainable. In Science and Society: Studies in the Sociology of Science [Aga81z], chapter 30, pages 444–456. ISBN 94-011-6458-4, 94-011-6456-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-011-6456-
                                              6_30.

Agassi:1981:ORS

[Aga81s] Joseph Agassi. The origins of the Royal Society. In Science and Society: Studies in the Sociology of Science [Aga81z], chapter 25, pages 352–371. ISBN 94-011-6458-4, 94-011-6456-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-011-6456-6_25.

Agassi:1981:PSP

[Aga81t] Joseph Agassi. The present state of the philosophy of science. In Science and Society: Studies in the Sociology of Science [Aga81z], chapter 2, pages 18– 32. ISBN 94-011-6458-4, 94-011-6456-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-011-6456-
                                              6_2.

Agassi:1981:PSV

[Aga81u] Joseph Agassi. The problems of scientific validation. In Grmek et al. [GCC81b], pages 103–114. ISBN 90-277-1123-2, 94-010-1284-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-010-1284-3_4.

Agassi:1981:RTQ

[Aga81v] Joseph Agassi. Rationality and the tu quoque argument. In Science and Society: Studies in the Sociology of Science [Aga81z], chapter 32, pages 465–476. ISBN 94-011-6458-4, 94-011-6456-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-011-
6456-6_32.

Agassi:1981:RP

[Aga81w] Joseph Agassi. Research project.

In Science and Society: Studies in the Sociology of Science [Aga81z], chapter 20, pages 262–272. ISBN 94-011-6458-4, 94-011-6456-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-011-6456-6_20.

Agassi:1981:RRS

[Aga81x] Joseph Agassi. Revising the referee system. In Science and Society: Studies in the Sociology of Science [Aga81z], chapter 13, pages 156–163. ISBN 94-011-6458-4, 94-011-6456-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-011-6456-6_13.

Agassi:1981:RSO

[Aga81y] Joseph Agassi. Revolutions in science, occasional or permanent? In Science and Society: Studies in the Sociology of Science [Aga81z], chapter 9, pages 104–118. ISBN 94-011-6458-4, 94-011-6456-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-011-
6456-6_9.

Agassi:1981:SSS

[Aga81z] Joseph Agassi. Science and Society:

Studies in the Sociology of Science, volume 65 of Boston Studies in the Philosophy of Science. D. Reidel, Dordrecht, The Netherlands; Boston, MA, USA; Lancaster, UK; Tokyo, Japan, 1981. ISBN 94-011-6458-4, 94-011-6456-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). 553 (est.) pp. LCCN ???? URL

https://link.springer.com/book/
10.1007/978-94-011-6456-6.

Agassi:1981:SST

[Aga81-27] Joseph Agassi. Scientific schools and their success. In Science and Society: Studies in the Sociology of Science [Aga81z], chapter 14, pages 164–191. ISBN 94-011-6458-4, 94-011-6456-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-011-6456-
6_14.

Agassi:1981:SS

[Aga81-28] Joseph Agassi. Scientists as sleepwalkers. In Science and Society: Studies in the Sociology of Science [Aga81z], chapter 16, pages 210–222. ISBN 94-011-6458-4, 94-011-6456-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-011-6456-6_16.

Agassi:1981:SJH

[Aga81-29] Joseph Agassi. Sir John Herschel’s philosophy of success. In Science and Society: Studies in the Sociology of Science [Aga81z], chapter 27, pages 388–420. ISBN 94-011-6458-4, 94-011-6456-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-011-
6456-6_27.

Agassi:1981:SPS

[Aga81-30] Joseph Agassi. Sociologism in philosophy of science. In Science and Society:

Studies in the Sociology of Science [Aga81z], chapter 8, pages 85–103. ISBN 94-011-6458-4, 94-011-6456-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-011-6456-
6_8.

Agassi:1981:SL

[Aga81-31] Joseph Agassi. Standards to live by.

In Science and Society: Studies in the Sociology of Science [Aga81z], chapter 34, pages 491–501. ISBN 94-011-6458-4, 94-011-6456-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-011-6456-6_34.

Agassi:1981:SCK

[Aga81-32] Joseph Agassi. Storage and communication of knowledge. In Science and Society: Studies in the Sociology of Science [Aga81z], chapter 11, pages 132–141. ISBN 94-011-6458-4, 94-011-6456-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-011-
6456-6_11.

Agassi:1981:TSP

[Aga81-33] Joseph Agassi. Technocracy and scientific progress. In Science and Society: Studies in the Sociology of Science [Aga81z], chapter 33, pages 477–490. ISBN 94-011-6458-4, 94-011-6456-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-011-6456-
6_33.

Agassi:1981:TVR

[Aga81-34] Joseph Agassi. Three views of the Renaissance of science. In Science and Society: Studies in the Sociology of Science [Aga81z], chapter 23, pages 300–320. ISBN 94-011-6458-4, 94-011-6456-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

                                              ???? URL https://link.springer.
                                              com/chapter/10.1007/978-94-011-
                                              6456-6_23.

Agassi:1981:WWR

[Aga81-35] Joseph Agassi. Was Wittgenstein really necessary? In Science and Society: Studies in the Sociology of Science [Aga81z], chapter 3, pages 33–44. ISBN 94-011-6458-4, 94-011-6456-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-011-6456-6_3.

Agassi:1981:WMS

[Aga81-36] Joseph Agassi. What makes for a scientific golden age? In Science and Society: Studies in the Sociology of Science [Aga81z], chapter 28, pages 421–435. ISBN 94-011-6458-4, 94-011-6456-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

                                              ???? URL https://link.springer.
                                              com/chapter/10.1007/978-94-011-
                                              6456-6_28.

Agazzi:1981:TC

[Aga81-37] Evandro Agazzi. Time and causality. In Italian Studies in the Philosophy of Science [Chi81b], chapter 17, pages 299–321. ISBN 90-277-1073-2, 94-009-8937-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

                                              ???? URL https://link.springer.
                                              com/chapter/10.1007/978-94-009-
                                              8937-5_17.

Agassi:1984:SBS

[Aga84] Joseph Agassi. The social base of scientific theory and practice. In Rationality in Science and Politics [And84d], chapter 2, pages 15–28. ISBN 90-277-1953-5, 94-009-6254-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-009-6254-
5_2.

Agassi:1986:TCV

[Aga86] Joseph Agassi. Towards a canonic version of classical political theory. In Spinoza and the Sciences [GN86], chapter 6, pages 153–170. ISBN 94-010-8511-0, 94-009-4514-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-009-4514-2_6.

Agassi:1987:CL

[Aga87a] Joseph Agassi. Comment on Levine.

In Shimony and Nails [SN87], pages 295–298. ISBN 94-010-8168-9, 94-009-3735-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-009-

3735-2_17. See [Lev87].

Agassi:1987:CS

[Aga87b] Joseph Agassi. Comment on Sagal.

In Shimony and Nails [SN87], pages 337–340. ISBN 94-010-8168-9, 94-009-3735-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-009-

3735-2_21. See [Sag87].

Agassi:1987:NEC

[Aga87c] Joseph Agassi. Naturalistic epistemology: The case of Abner Shimony. In Shimony and Nails [SN87], pages 341–351. ISBN 94-010-8168-9, 94-009-3735-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-009-

3735-2_22. See comment [Shi87a].

Agassi:1989:LCE

[Aga89] Joseph Agassi. The logic of consensus and of extremes. In Freedom and Rationality: Essays in Honor of John Watkins From his Colleagues and Friends [DJ89], chapter 1, pages 3–21. ISBN 94-010-7571-9, 94-009-2380-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-009-2380-
5_1.

Agassi:1991:YLI

[Aga91] Joseph Agassi. As you like it. In Beyond Reason: Essays on the Philosophy of Paul Feyerabend [Mun91a], chapter 17, pages 379–387. ISBN 94-010-5406-1, 94-011-3188-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-011-3188-9_17.

Agassi:1994:POP

[Aga94] Joseph Agassi. The philosophy of optimism and pessimism. In Artifacts, Representations and Social Practice: Essays for Marx Wartofsky [GC94], chapter 20, pages 349–359. ISBN 94-010-4390-6, 94-011-0902-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN B73 .A77 1994.

URL https://link.springer.com/
chapter/10.1007/978-94-011-0902-
4_20.

Agassi:1995:CPS

[Aga95a] Joseph Agassi. Contemporary philosophy of science as a thinly masked antidemocratic apologetics. In Gavroglu et al. [GSW95c], chapter 8, pages 153–170. ISBN 0-7923-2991-0 (set), 0-7923-2988-0, 94-017-2658-2. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 163.

URL https://link.springer.com/
chapter/10.1007/978-94-017-2658-
0_8.

Agassi:1995:TPC

[Aga95b] Joseph Agassi. The theory and practice of critical rationalism. In Misiek [Mis95c], pages 1–18. ISBN 94-010-4206-3, 94-011-0461-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-011-0461-6_1.

Agassi:1995:EMC

[Aga95c] Judith Buber Agassi. Epistemological and methodological concerns of feminist social scientists. In Critical Rationalism, the Social Sciences and the Humanities: Essays for Joseph Agassi. Volume II [JL95b], chapter 9, pages 153–165. ISBN 0-7923-2961-9, 90-481-4430-2, 94-017-0441-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN B29 .C75 1995.

URL https://link.springer.com/
chapter/10.1007/978-94-017-0441-
0_9.

Agazzi:1995:TDK

[Aga95d] Evandro Agazzi. Are there different kinds of knowledge? In The Concept of Knowledge: The Ankara Seminar [KC95], chapter 9, pages 103–118. ISBN 90-481-4495-7, 94-017-3263-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-017-3263-5_9.

Agassi:1997:WNA

[Aga97a] Joseph Agassi. Who needs Aristotle? In Ginev and Cohen [GC97], chapter 1, pages 1–11. ISBN 0-7923-4444-8 (hardcover), 94-011-5788-X. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 192. URL https:/

                                              /link.springer.com/chapter/10.
                                              1007/978-94-011-5788-9_1.

Agazzi:1997:NRN

[Aga97b] Evandro Agazzi. Naive realism and naive antirealism. In Ginev and Cohen [GC97], chapter 2, pages 13–26. ISBN 0-7923-4444-8 (hardcover), 94-011-5788-X. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 192. URL https://link.

                                              springer.com/chapter/10.1007/
                                              978-94-011-5788-9_2.

Agassi:2000:MC

[Aga00a] Joseph Agassi. Maimonides in context. In Maimonides and the Sciences [CL00], chapter 2, pages 9–24. ISBN 90-481-5348-4, 94-017-2128-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-017-2128-8_2.

Agazzi:2000:OR

[Aga00b] Evandro Agazzi. Observability and referentiality. In The Reality of the Unobservable: Observability, Unobservability and Their Impact on the Issue of Scientific Realism [AP00b], chapter 3, pages 45–57. ISBN 90-481-5458-8, 94-015-9391-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-015-
9391-5_3.

Agassi:2003:AI

[Aga03a] Joseph Agassi. Artificial intelligence.

In Science and Culture [Aga03s], chapter 25, pages 279–293. ISBN 90-481-6234-3, 94-017-2946-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174. URL https:

//link.springer.com/chapter/10.
1007/978-94-017-2946-8_25.

Agassi:2003:B

[Aga03b] Joseph Agassi. Bloodletting. In Science and Culture [Aga03s], chapter 16, pages 196–203. ISBN 90-481-6234-3, 94-017-2946-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

Q174. URL https://link.springer.
com/chapter/10.1007/978-94-017-
2946-8_16.

Agassi:2003:CS

[Aga03c] Joseph Agassi. The consolations of science. In Science and Culture [Aga03s], chapter 2, pages 7–23. ISBN 90-481-6234-3, 94-017-2946-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174. URL https:

//link.springer.com/chapter/10.
1007/978-94-017-2946-8_2.

Agassi:2003:CL

[Aga03d] Joseph Agassi. The critique of linearity. In Science and Culture [Aga03s], chapter 17, pages 204–212. ISBN 90-481-6234-3, 94-017-2946-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174. URL https:

//link.springer.com/chapter/10.
1007/978-94-017-2946-8_17.

Agassi:2003:FOS

[Aga03e] Joseph Agassi. Faith in the open society: The end of hermeneutics. In Science and Culture [Aga03s], chapter 8, pages 83–96. ISBN 90-481-6234-3, 94-017-2946-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic).

                                              LCCN Q174. URL https://link.
                                              springer.com/chapter/10.1007/
                                              978-94-017-2946-8_8.

Agassi:2003:PRS

[Aga03f] Joseph Agassi. For public responsibility for Spaceship Earth. In Science and Culture [Aga03s], chapter 35, pages 397–403. ISBN 90-481-6234-3, 94-017-2946-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

                                              Q174. URL https://link.springer.
                                              com/chapter/10.1007/978-94-017-
                                              2946-8_35.

Agassi:2003:FIR

[Aga03g] Joseph Agassi. The functions of intellectual rubbish. In Science and Culture [Aga03s], chapter 9, pages 97–112. ISBN 90-481-6234-3, 94-017-2946-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174.

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-017-2946-
                                              8_9.

Agassi:2003:IW

[Aga03h] Joseph Agassi. The inner world. In Science and Culture [Aga03s], chapter 28, pages 319–326. ISBN 90-481-6234-3, 94-017-2946-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic).

                                              LCCN Q174. URL https://link.
                                              springer.com/chapter/10.1007/
                                              978-94-017-2946-8_28.

Agassi:2003:IT

[Aga03i] Joseph Agassi. Irrationalism today.

In Science and Culture [Aga03s], chapter 14, pages 164–177. ISBN 90-481-6234-3, 94-017-2946-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174. URL https:

//link.springer.com/chapter/10.
1007/978-94-017-2946-8_14.

Agassi:2003:MCI

[Aga03j] Joseph Agassi. Minimal criteria for intellectual progress. In Science and Culture [Aga03s], chapter 15, pages 181–195. ISBN 90-481-6234-3, 94-017-2946-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

Q174. URL https://link.springer.
com/chapter/10.1007/978-94-017-
2946-8_15.

Agassi:2003:MBS

[Aga03k] Joseph Agassi. The moral base of science, or, the architectonic of openended reason. In Science and Culture [Aga03s], chapter 3, pages 24–36. ISBN 90-481-6234-3, 94-017-2946-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174. URL https:

//link.springer.com/chapter/10.
1007/978-94-017-2946-8_3.

Agassi:2003:PSb

[Aga03l] Joseph Agassi. Philosophy without science. In Science and Culture [Aga03s], chapter 26, pages 294–305. ISBN 90-481-6234-3, 94-017-2946-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174. URL https:

//link.springer.com/chapter/10.
1007/978-94-017-2946-8_26.

Agassi:2003:PSc

[Aga03m] Joseph Agassi. The politics of science. In Science and Culture [Aga03s], chapter 30, pages 344–358. ISBN 90-481-6234-3, 94-017-2946-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174. URL https:

//link.springer.com/chapter/10.
1007/978-94-017-2946-8_30.

Agassi:2003:PSA

[Aga03n] Joseph Agassi. Progress in science and in art. In Science and Culture [Aga03s], chapter 23, pages 260–269. ISBN 90-481-6234-3, 94-017-2946-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174. URL https:

//link.springer.com/chapter/10.
1007/978-94-017-2946-8_23.

Agassi:2003:RP

[Aga03o] Joseph Agassi. Rationalizing politics. In Science and Culture [Aga03s], chapter 33, pages 376–388. ISBN 90-481-6234-3, 94-017-2946-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174. URL https:

//link.springer.com/chapter/10.
1007/978-94-017-2946-8_33.

Agassi:2003:SA

[Aga03p] Joseph Agassi. Science and art. In Science and Culture [Aga03s], chapter 27, pages 306–318. ISBN 90-481-6234-3, 94-017-2946-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic).

LCCN Q174. URL https://link.
springer.com/chapter/10.1007/
978-94-017-2946-8_27.

Agassi:2003:SCb

[Aga03q] Joseph Agassi. Science, and commonsense. In Science and Culture [Aga03s], chapter 32, pages 372–375. ISBN 90-481-6234-3, 94-017-2946-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174. URL https:

//link.springer.com/chapter/10.
1007/978-94-017-2946-8_32.

Agassi:2003:SCa

[Aga03r] Joseph Agassi. Science and controversy. In Science and Culture [Aga03s], chapter 7, pages 72–79. ISBN 90-481-6234-3, 94-017-2946-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174. URL https:

//link.springer.com/chapter/10.
1007/978-94-017-2946-8_7.

Agassi:2003:SCc

[Aga03s] Joseph Agassi. Science and Culture, volume 231 of Boston Studies in the Philosophy of Science. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 2003. ISBN 90-481-6234-3, 94-017-2946-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). xxxii + 416 pp. LCCN Q174. URL

https://link.springer.com/book/
10.1007/978-94-017-2946-8.

Agassi:2003:SPR

[Aga03t] Joseph Agassi. Science and its public relations. In Science and Culture [Aga03s], chapter 13, pages 152–163. ISBN 90-481-6234-3, 94-017-2946-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174.

URL https://link.springer.com/
chapter/10.1007/978-94-017-2946-
8_13.

Agassi:2003:SPb

[Aga03u] Joseph Agassi. Science and philosophy. In Science and Culture [Aga03s], chapter 22, pages 251–259. ISBN 90-481-6234-3, 94-017-2946-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174. URL https:

//link.springer.com/chapter/10.
1007/978-94-017-2946-8_22.

Agassi:2003:SPa

[Aga03v] Joseph Agassi. Science and pluralism. In Science and Culture [Aga03s], chapter 11, pages 126–139. ISBN 90-481-6234-3, 94-017-2946-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174. URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-017-2946-8_11.

Agassi:2003:ST

[Aga03w] Joseph Agassi. Science and technology. In Science and Culture [Aga03s], chapter 21, pages 239–248. ISBN 90-481-6234-3, 94-017-2946-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174. URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-017-2946-8_21.

Agassi:2003:SCW

[Aga03x] Joseph Agassi. Science and the call of the wild. In Science and Culture [Aga03s], chapter 6, pages 57–71. ISBN 90-481-6234-3, 94-017-2946-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174. URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-017-2946-8_6.

Agassi:2003:SDN

[Aga03y] Joseph Agassi. Science and the detective novel. In Science and Culture [Aga03s], chapter 19, pages 221–231. ISBN 90-481-6234-3, 94-017-2946-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174.

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-017-2946-
                                              8_19.

Agassi:2003:SI

[Aga03z] Joseph Agassi. Science and the interpersonal. In Science and Culture [Aga03s], chapter 10, pages 113–125.

ISBN 90-481-6234-3, 94-017-2946-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174.

URL https://link.springer.com/
chapter/10.1007/978-94-017-2946-
8_10.

Agassi:2003:SPE

[Aga03-27] Joseph Agassi. Science as a public enterprise. In Science and Culture [Aga03s], chapter 31, pages 359–371. ISBN 90-481-6234-3, 94-017-2946-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174.

URL https://link.springer.com/
chapter/10.1007/978-94-017-2946-
8_31.

Agassi:2003:STA

[Aga03-28] Joseph Agassi. Science as training for autonomy. In Science and Culture [Aga03s], chapter 5, pages 48–56. ISBN 90-481-6234-3, 94-017-2946-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174. URL https:

//link.springer.com/chapter/10.
1007/978-94-017-2946-8_5.

Agassi:2003:SFM

[Aga03-29] Joseph Agassi. Science fiction: This message is for you. Maybe. In Science and Culture [Aga03s], chapter 1, pages 3–6. ISBN 90-481-6234-3, 94-017-2946-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174.

URL https://link.springer.com/
chapter/10.1007/978-94-017-2946-
8_1.

Agassi:2003:SPO

[Aga03-30] Joseph Agassi. Science, politics and objectivity. In Science and Culture [Aga03s], chapter 18, pages 213–220. ISBN 90-481-6234-3, 94-017-2946-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174.

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-017-2946-
                                              8_18.

Agassi:2003:SH

[Aga03-31] Joseph Agassi. The siblinghood of humanity. In Science and Culture [Aga03s], chapter 34, pages 389–396. ISBN 90-481-6234-3, 94-017-2946-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174.

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-017-2946-
                                              8_34.

Agassi:2003:TBS

[Aga03-32] Joseph Agassi. Technology as both science and art. In Science and Culture [Aga03s], chapter 24, pages 270–278. ISBN 90-481-6234-3, 94-017-2946-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174.

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-017-2946-
                                              8_24.

Agassi:2003:TPC

[Aga03-33] Joseph Agassi. The theory and practice of critical rationalism. In Science and Culture [Aga03s], chapter 4, pages 37– 47. ISBN 90-481-6234-3, 94-017-2946-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174.

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-017-2946-
                                              8_4.

Agassi:2003:TB

[Aga03-34] Joseph Agassi. The two books. In Science and Culture [Aga03s], chapter 20, pages 232–238. ISBN 90-481-6234-3, 94-017-2946-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic).

                                              LCCN Q174. URL https://link.

springer.com/chapter/10.1007/
978-94-017-2946-8_20.

Agassi:2003:V

[Aga03-35] Joseph Agassi. Validation. In Science and Culture [Aga03s], chapter 29, pages 329–343. ISBN 90-481-6234-3, 94-017-2946-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

Q174. URL https://link.springer.
com/chapter/10.1007/978-94-017-
2946-8_29.

Agassi:2003:WGC

[Aga03-36] Joseph Agassi. Wild goose chase.

In Science and Culture [Aga03s], chapter 12, pages 140–151. ISBN 90-481-6234-3, 94-017-2946-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174. URL https:

//link.springer.com/chapter/10.
1007/978-94-017-2946-8_12.

Agassi:2008:CCS

[Aga08a] Joseph Agassi. Chroniclers in the courts of science: Preliminary essays on the traditions and the history of science. In Science and its History: A Reassessment of the Historiography of Science [Aga08d], chapter 1, pages 1– 118. ISBN 1-4020-5631-1 (hardcover), 1-4020-5632-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174.8 .A33 2008. URL https:

//link.springer.com/chapter/10.
1007/978-1-4020-5632-1_1.

Agassi:2008:HEb

[Aga08b] Joseph Agassi. Historical essays. In Science and its History: A Reassessment of the Historiography of Science [Aga08d], chapter 4, pages 335–500. ISBN 1-4020-5631-1 (hardcover), 1-4020-5632-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174.8 .A33 2008. URL https:/

/link.springer.com/chapter/10.
1007/978-1-4020-5632-1_4.

Agassi:2008:HEa

[Aga08c] Joseph Agassi. Historiographic essays. In Science and its History: A Reassessment of the Historiography of Science [Aga08d], chapter 3, pages 243–334. ISBN 1-4020-5631-1 (hardcover), 1-4020-5632-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174.8 .A33 2008. URL https:/

/link.springer.com/chapter/10.
1007/978-1-4020-5632-1_3.

Agassi:2008:SHR

[Aga08d] Joseph Agassi. Science and its History: A Reassessment of the Historiography of Science, volume 253 of Boston Studies in the Philosophy of Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2008. ISBN 1-4020-5631-1 (hardcover), 1-4020-5632-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). xxi + 514 pp. LCCN Q174.8 .A33 2008. URL

https://link.springer.com/book/
10.1007/978-1-4020-5632-1.

Agassi:2008:THS

[Aga08e] Joseph Agassi. Towards an historiography of science. In Science and its History: A Reassessment of the Historiography of Science [Aga08d], chapter 2, pages 119–242. ISBN 1-4020-5631-1 (hardcover), 1-4020-5632-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174.8 .A33 2008.

URL https://link.springer.com/
chapter/10.1007/978-1-4020-5632-
1_2.

Agassi:2009:PIS

[Aga09] Joseph Agassi. Popper’s insights into the state of economics. In Rethinking Popper [PC09], chapter 26, pages 357–368. ISBN 1-4020-9337-3, 1-4020-9338-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN B1649.P64 R48 2009; Q174 .B67 2009.

URL https://link.springer.com/
chapter/10.1007/978-1-4020-9338-
8_26.

Agassi:2013:BOE

[Aga13a] Joseph Agassi. Bacon on the origin of error and prejudice. In The Very Idea of Modern Science: Francis Bacon and Robert Boyle [Aga13q], chapter 5, pages 49–56. ISBN 94-007-5350-0, 94-007-5351-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175 .A318 2013; Q174 .B67 v. 298.

URL https://link.springer.com/
chapter/10.1007/978-94-007-5351-
8_5.

Agassi:2013:BI

[Aga13b] Joseph Agassi. Bacon’s influence.

In The Very Idea of Modern Science: Francis Bacon and Robert Boyle [Aga13q], chapter 8, pages 109–120. ISBN 94-007-5350-0, 94-007-5351-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175 .A318 2013; Q174 .B67 v. 298.

URL https://link.springer.com/
chapter/10.1007/978-94-007-5351-
8_8.

Agassi:2013:BPD

[Aga13c] Joseph Agassi. Bacon’s philosophy of discovery. In The Very Idea of Modern Science: Francis Bacon and Robert Boyle [Aga13q], chapter 2, pages 15– 33. ISBN 94-007-5350-0, 94-007-5351-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175 .A318 2013; Q174 .B67 v. 298. URL https:

//link.springer.com/chapter/10.
1007/978-94-007-5351-8_2.

Agassi:2013:BEP

[Aga13d] Joseph Agassi. Boyle in the eyes of posterity. In The Very Idea of Modern Science: Francis Bacon and Robert Boyle [Aga13q], chapter 13, pages 167–178. ISBN 94-007-5350-0, 94-007-5351-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175 .A318 2013; Q174 .B67 v. 298.

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-007-5351-
                                              8_13.

Agassi:2013:CRR

[Aga13e] Joseph Agassi. Conclusion: The rise of the riddle of Bacon. In The Very Idea of Modern Science: Francis Bacon and Robert Boyle [Aga13q], chapter 9, pages 121–124. ISBN 94-007-5350-0, 94-007-5351-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175 .A318 2013; Q174 .B67 v.

                                              298. URL https://link.springer.
                                              com/chapter/10.1007/978-94-007-
                                              5351-8_9.

Agassi:2013:EMD

[Aga13f] Joseph Agassi. Ellis’ major difficulty. In The Very Idea of Modern Science: Francis Bacon and Robert Boyle [Aga13q], chapter 3, pages 35–38. ISBN 94-007-5350-0, 94-007-5351-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175 .A318 2013; Q174 .B67 v. 298. URL https:/

                                              /link.springer.com/chapter/10.
                                              1007/978-94-007-5351-8_3.

Agassi:2013:FDP

[Aga13g] Joseph Agassi. The function of the doctrine of prejudice. In The Very Idea of Modern Science: Francis Bacon and Robert Boyle [Aga13q], chapter 4, pages 39–47. ISBN 94-007-5350-0, 94-007-5351-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175 .A318 2013; Q174 .B67 v. 298.

URL https://link.springer.com/
chapter/10.1007/978-94-007-5351-
8_4.

Agassi:2013:IS

[Aga13h] Joseph Agassi. The inductive style.

In The Very Idea of Modern Science: Francis Bacon and Robert Boyle [Aga13q], chapter 14, pages 179–224. ISBN 94-007-5350-0, 94-007-5351-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175 .A318 2013; Q174 .B67 v. 298.

URL https://link.springer.com/
chapter/10.1007/978-94-007-5351-
8_14.

Agassi:2013:M

[Aga13i] Joseph Agassi. Mechanism. In The Very Idea of Modern Science: Francis Bacon and Robert Boyle [Aga13q], chapter 15, pages 225–248. ISBN 94-007-5350-0, 94-007-5351-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175 .A318 2013; Q174 .B67 v. 298. URL https:/

/link.springer.com/chapter/10.
1007/978-94-007-5351-8_15.

Agassi:2013:MLB

[Aga13j]  Joseph Agassi.  The missing link

between Bacon and the Royal Society. In The Very Idea of Modern Science: Francis Bacon and Robert Boyle [Aga13q], chapter 12, pages 157–166. ISBN 94-007-5350-0, 94-007-5351-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175 .A318 2013; Q174 .B67 v. 298.

URL https://link.springer.com/
chapter/10.1007/978-94-007-5351-
8_12.

Agassi:2013:NDP

[Aga13k] Joseph Agassi. The new doctrine of prejudice. In The Very Idea of Modern Science: Francis Bacon and Robert Boyle [Aga13q], chapter 16, pages 249–264. ISBN 94-007-5350-0, 94-007-5351-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175 .A318 2013; Q174 .B67 v. 298.

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-007-5351-
                                              8_16.

Agassi:2013:PB

[Aga13l] Joseph Agassi. Philosophical background. In The Very Idea of Modern Science: Francis Bacon and Robert Boyle [Aga13q], chapter 10, pages 131–137. ISBN 94-007-5350-0, 94-007-5351-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175 .A318 2013; Q174 .B67 v. 298.

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-007-5351-
                                              8_10.

Agassi:2013:PO

[Aga13m] Joseph Agassi. Prejudices of opinions. In The Very Idea of Modern Science: Francis Bacon and Robert Boyle [Aga13q], chapter 7, pages 81–108. ISBN 94-007-5350-0, 94-007-5351-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175 .A318 2013; Q174 .B67 v. 298.

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-007-5351-
                                              8_7.

Agassi:2013:PS

[Aga13n] Joseph Agassi. Prejudices of the senses. In The Very Idea of Modern Science: Francis Bacon and Robert Boyle [Aga13q], chapter 6, pages 57–80. ISBN 94-007-5350-0, 94-007-5351-9 (e-book).

ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175 .A318 2013; Q174 .B67 v. 298. URL https:/

/link.springer.com/chapter/10.
1007/978-94-007-5351-8_6.

Agassi:2013:RB

[Aga13o] Joseph Agassi. The riddle of Bacon.

In The Very Idea of Modern Science: Francis Bacon and Robert Boyle [Aga13q], chapter 1, pages 3–13. ISBN 94-007-5350-0, 94-007-5351-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175 .A318 2013; Q174 .B67 v. 298. URL https:/

/link.springer.com/chapter/10.
1007/978-94-007-5351-8_1.

Agassi:2013:SBC

[Aga13p] Joseph Agassi. The social background of classical science. In The Very Idea of Modern Science: Francis Bacon and Robert Boyle [Aga13q], chapter 11, pages 139–155. ISBN 94-007-5350-0, 94-007-5351-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175 .A318 2013; Q174 .B67 v.

298. URL https://link.springer.
com/chapter/10.1007/978-94-007-
5351-8_11.

Agassi:2013:VIM

[Aga13q] Joseph Agassi. The Very Idea of Modern Science: Francis Bacon and Robert Boyle, volume 298 of Boston Studies in the Philosophy and History of Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2013. ISBN 94-007-5350-0, 94-007-5351-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). xvii + 315 pp. LCCN Q175 .A318 2013; Q174 .B67 v. 298.

URL https://link.springer.com/
book/10.1007/978-94-007-5351-8.

Agoston:2011:OET

[Ágo11] Gábor Ágoston. The Ottoman Empire and the technological dialogue between Europe and Asia: The case of military technology and know-how in the Gunpowder Age. In Science between Europe and Asia: Historical Studies on the Transmission, Adoption and Adaptation of Knowledge [GR11b], chapter 3, pages 27–39. ISBN 90-481-9968-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q125 .S354 2010; Q175.52.A78 S35

                                              2010. URL https://link.springer.
                                              com/chapter/10.1007/978-90-481-
                                              9968-6_3.

Aichner:2015:RAU

[Aic15] Christof Aichner. The reforms of the Austrian university system 1848–1860 and their influence on the process of discipline formation. In Sciences in the Universities of Europe, Nineteenth and Twentieth Centuries: Academic Landscapes [SDG15b], chapter 17, pages 293–309. ISBN 94-017-9635-1, 94-017-9636-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q183.4.E85 2015. URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-017-9636-1_17.

Akbas:2011:BTA

[Akb11] Meltem Akbaş. Between translation and adaptation: Turkish editions of Ganot’s Traité. In Science between Europe and Asia: Historical Studies on the Transmission, Adoption and Adaptation of Knowledge [GR11b], chapter 12, pages 177–191. ISBN 90-481-9968-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q125 .S354 2010; Q175.52.A78 S35

                                              2010. URL https://link.springer.

com/chapter/10.1007/978-90-481-
9968-6_12.

Akrami:2009:DPN

[Akr09] Musa Akrami. The difficulties with Popper’s nontraditional conception of metaphysics. In Rethinking Popper [PC09], chapter 30, pages 397–416. ISBN 1-4020-9337-3, 1-4020-9338-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN B1649.P64 R48 2009; Q174 .B67 2009.

URL https://link.springer.com/
chapter/10.1007/978-1-4020-9338-
8_30.

Albert:1978:SST

[Alb78] Hans Albert. Science and the search for truth. In Progress and Rationality in Science [RA78b], chapter 9, pages 203–220. ISBN 90-277-0922-X, 94-009-9866-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-009-
9866-7_9.

Albert:1984:TRR

[Alb84] Hans Albert. Transcendental realism and rational heuristics: Critical rationalism and the problem of method. In Rationality in Science and Politics [And84d], chapter 3, pages 29–46. ISBN 90-277-1953-5, 94-009-6254-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-009-6254-5_3.

Albert:1992:BRE

[Alb92] David Z. Albert. Bohr’s response to Einstein, Podolsky, and Rosen. In The Scientific Enterprise: The Bar-Hillel Colloquium: Studies in History, Philosophy, and Sociology of Science, Volume 4 [UM92], chapter 17, pages 269–272. ISBN 94-010-5190-9, 94-011-2688-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174

.S354 1992. URL https://link.
springer.com/chapter/10.1007/
978-94-011-2688-5_17.

Albert:1995:RSM

[Alb95] Hans Albert. Religion, science, and the myth of the framework. In Critical Rationalism, Metaphysics and Science: Essays for Joseph Agassi. Volume I [JL95a], chapter 5, pages 41–58. ISBN 94-010-4211-X, 94-011-0471-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-011-0471-5_5.

Albert:1996:EQM

[Alb96] David Z. Albert. Elementary quantum metaphysics. In Bohmian Mechanics and Quantum Theory: an Appraisal [CFG96], chapter 19, pages 277–284. ISBN 90-481-4698-4, 94-015-8715-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC173.96-174.52. URL https://link.

springer.com/chapter/10.1007/
978-94-015-8715-0_19.

Allen:1999:ADL

[All99a] Barry Allen. All the daring of the lover of knowledge is permitted again. In Nietzsche, Epistemology, and Philosophy of Science: Nietzsche and the Sciences II [Bab99a], chapter 10, pages 123–140. ISBN 90-481-5234-8, 94-017-2428-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-017-2428-
9_10.

Allison:1999:TI

[All99b] David B. Allison. Twilight of the icons.

In Nietzsche, Epistemology, and Philosophy of Science: Nietzsche and the Sciences II [Bab99a], chapter 15, pages 179–185. ISBN 90-481-5234-8, 94-017-2428-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-017-
2428-9_15.

Allen:2001:RPM

[All01] Garland E. Allen. Radical politics and Marxism in the history of science. In Allen and MacLeod [AM01b], chapter 12, pages 185–201. ISBN 1-4020-0495-8 (hardcover), 94-017-2956-5. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 228; Q175.55. URL https:

//link.springer.com/chapter/10.
1007/978-94-017-2956-7_12.

Almasi:1989:PA

[Alm89] Miklós Almási. The Philosophy of Appearances, volume 106 of Boston Studies in the Philosophy of Science. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 1989. ISBN 90-277-2150-5. xviii + 294 pp. LCCN BD331 .A3713 1989; Q174 .B67 vol. 106.

Alonso:2014:CSP

[Alo14] Marcelo Alonso. The current state of physics in Cuba: A personal perspective. In The History of Physics in Cuba [BRW14a], chapter 28, pages 399–401. ISBN 94-017-8040-4, 94-017-8041-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC5.53; QC6.9; QC9 C9H57 2014; Q124.6-127.2. URL https://link.

                                              springer.com/chapter/10.1007/
                                              978-94-017-8041-4_28.

Altner:1987:GFA

[Alt87] Günter Altner. Goethe as a forerunner of alternative science. In Amrine et al. [AZW87], chapter 17, pages 341–350. ISBN 90-277-2265-X, 94-009-3761-X. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol.

97. US$79.00. URL https://link.

                                              springer.com/chapter/10.1007/
                                              978-94-009-3761-1_17.

Altmann:2000:DMA

[Alt00] Alexander Altmann. Defining Maimonides’ Aristotelianism. In Maimonides and the Sciences [CL00], chapter 1, pages 1–7. ISBN 90-481-5348-4, 94-017-2128-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-017-2128-8_1.

Altizer:2002:VGE

[Alt02] Thomas J. J. Altizer. Van Gogh’s eyes. In Hermeneutic Philosophy of Science, Van Gogh’s Eyes, and God: Essays in Honor of Patrick A. Heelan, S.J. [Bab02c], chapter 34, pages 393–402. ISBN 90-481-5926-1, 94-017-1767-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 225. URL https://link.

                                              springer.com/chapter/10.1007/
                                              978-94-017-1767-0_34.

Altshuler:2014:CCP

[Alt14a]  Ernesto Altshuler.  Contemporary

Cuban physics through scientific publications: An insider’s view. In The History of Physics in Cuba [BRW14a], chapter 34, pages 439–446. ISBN 94-017-8040-4, 94-017-8041-2 (e-book).

ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC5.53; QC6.9; QC9 C9H57 2014; Q124.6-127.2.

URL https://link.springer.com/
chapter/10.1007/978-94-017-8041-
4_34.

Altshuler:2014:MPC

[Alt14b]  José Altshuler.  Mathematics and

physics in Cuba before 1959: A personal recollection. In The History of Physics in Cuba [BRW14a], chapter 5, pages 107–113. ISBN 94-017-8040-4, 94-017-8041-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC5.53; QC6.9; QC9 C9H57 2014; Q124.6-127.2. URL https:/

/link.springer.com/chapter/10.
1007/978-94-017-8041-4_5.

Anderson:1963:MLI

[AM63]  Alan  Ross  Anderson  and

Omar Khayyam Moore. Modal logics II: Toward a formal analysis of cultural objects. In Wartofsky [War63], pages 117–143. ISBN 94-010-3265-3, 94-010-3263-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B678 1970. URL https:

//link.springer.com/chapter/10.
1007/978-94-010-3263-6_10.

Arbib:1983:CTA

[AM83] Michael A. Arbib and Ernest G. Manes.

A category–theoretic approach to systems in a fuzzy world. In Language, Logic and Method [CW83b], chapter 11, pages 199–224. ISBN 94-009-7704-2, 94-009-7702-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-009-
7702-0_11.

Allen:2001:I

[AM01a] Garland E. Allen and Roy MacLeod. Introduction. In Allen and MacLeod [AM01b], chapter 1, pages 1–20. ISBN 1-4020-0495-8 (hardcover), 94-017-2956-5. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 228; Q175.55. URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-017-2956-7_1.

Allen:2001:SHS

[AM01b] Garland E. Allen and Roy M. MacLeod, editors. Science, history, and social activism: a tribute to Everett Mendelsohn, volume 228 of Boston Studies in the Philosophy of Science. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 2001. ISBN 1-4020-0495-8 (hardcover), 94-017-2956-5. ISSN 0068-0346 (print), 2214-7942 (electronic). x + 388 pp. LCCN Q174 .B67 vol.

                                              228; Q175.55. URL http://www.loc.
                                              gov/catdir/enhancements/fy0821/
                                              2002023597-d.html;   http://
                                              www.loc.gov/catdir/toc/fy022/
                                              2002023597.html;   https://link.
                                              springer.com/book/10.1007/978-
                                              94-017-2956-7.

Amrine:1987:GSA

[Amr87a] Frederick Amrine. Goethe and the sciences: An annotated bibliography. In Amrine et al. [AZW87], chapter 20, pages 389–437. ISBN 90-277-2265-X, 94-009-3761-X. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 97. US$79.00. URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-009-3761-1_20.

Amrine:1987:GMW

[Amr87b] Frederick Amrine. Goethean method in the work of Jochen Bockemühl. In Amrine et al. [AZW87], chapter 15, pages 301–318. ISBN 90-277-2265-X, 94-009-3761-X. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 97. US$79.00. URL https:

//link.springer.com/chapter/10.
1007/978-94-009-3761-1_15.

Amrine:1989:LSM

[Amr89]  Frederick Amrine.  Literature and

Science as Modes of Expression, volume 115 of Boston Studies in the Philosophy of Science. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 1989. ISBN 94-010-7531-X, 94-009-2297-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). xxvi + 201 pp. LCCN ???? URL

https://link.springer.com/book/
10.1007/978-94-009-2297-6.

Amsterdamski:1975:BEM

[Ams75a] Stefan Amsterdamski. Between Experience and Metaphysics: Philosophical Problems of the Evolution of Science, volume 35 of Boston Studies in the Philosophy of Science. D. Reidel, Dordrecht, The Netherlands; Boston, MA, USA; Lancaster, UK; Tokyo, Japan, 1975. ISBN 90-277-0580-1, 94-010-1797-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). xviii + 198 pp. LCCN Q174 .A478 1975. URL

https://link.springer.com/book/
10.1007/978-94-010-1797-8.

Amsterdamski:1975:CDC

[Ams75b] Stefan Amsterdamski. The context of discovery and the context of justification. In Between Experience and Metaphysics: Philosophical Problems of the Evolution of Science [Ams75a], chapter 3, pages 47–68. ISBN 90-277-0580-1, 94-010-1797-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .A478 1975.

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-010-1797-
                                              8_3.

Amsterdamski:1975:FTC

[Ams75c] Stefan Amsterdamski. Facts and theories: Conventionalism. In Between Experience and Metaphysics: Philosophical Problems of the Evolution of Science [Ams75a], chapter 5, pages 88–114. ISBN 90-277-0580-1, 94-010-1797-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .A478 1975. URL https://link.

                                              springer.com/chapter/10.1007/
                                              978-94-010-1797-8_5.

Amsterdamski:1975:FTR

[Ams75d] Stefan Amsterdamski. Facts and theories: Radical empiricism. In Between Experience and Metaphysics: Philosophical Problems of the Evolution of Science [Ams75a], chapter 4, pages 69– 87. ISBN 90-277-0580-1, 94-010-1797-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .A478

                                              1975. URL https://link.springer.
                                              com/chapter/10.1007/978-94-010-
                                              1797-8_4.

Amsterdamski:1975:REA

[Ams75e] Stefan Amsterdamski. Radical empiricism and the anomalies in the knowledge of science. In Between Experience and Metaphysics: Philosophical Problems of the Evolution of Science [Ams75a], chapter 1, pages 1–23. ISBN 90-277-0580-1, 94-010-1797-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .A478 1975.

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-010-1797-
                                              8_1.

Amsterdamski:1975:RCR

[Ams75f] Stefan Amsterdamski. Reformation and counter-reformation: Paradigms and research programs. In Between Experience and Metaphysics: Philosophical Problems of the Evolution of Science [Ams75a], chapter 6, pages 115–143. ISBN 90-277-0580-1, 94-010-1797-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174

.A478 1975. URL https://link.
springer.com/chapter/10.1007/
978-94-010-1797-8_6.

Amsterdamski:1975:RSS

[Ams75g] Stefan Amsterdamski. Revolutions in science: Science and philosophy. In Between Experience and Metaphysics: Philosophical Problems of the Evolution of Science [Ams75a], chapter 8, pages 160–176. ISBN 90-277-0580-1, 94-010-1797-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .A478 1975. URL https://link.

springer.com/chapter/10.1007/
978-94-010-1797-8_8.

Amsterdamski:1975:RSA

[Ams75h] Stefan Amsterdamski. Revolutions in science: The accumulation of knowledge and the correspondence of theories. In Between Experience and Metaphysics: Philosophical Problems of the Evolution of Science [Ams75a], chapter 7, pages 144–159. ISBN 90-277-0580-1, 94-010-1797-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .A478 1975. URL https://link.

springer.com/chapter/10.1007/
978-94-010-1797-8_7.

Amsterdamski:1975:TPD

[Ams75i] Stefan Amsterdamski. Troubles with the problem of demarcation. In Between Experience and Metaphysics:

Philosophical Problems of the Evolution of Science [Ams75a], chapter 2, pages 24–46. ISBN 90-277-0580-1, 94-010-1797-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .A478 1975. URL https://link.

springer.com/chapter/10.1007/
978-94-010-1797-8_2.

Amsterdamski:1982:RSR

[Ams82]  Stefan Amsterdamski.  Reflections

on science and rationality. In Polish Essays in the Philosophy of the Natural Sciences [Kra82c], chapter 1, pages 3–17. ISBN 94-009-7705-0 (e-book), 90-277-1287-5. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-009-7705-
1_1.

Amsterdamski:1992:TSC

[Ams92a] Stefan Amsterdamski. Are there selection criteria? In Between History and Method: Disputes about the Rationality of Science [Ams92b], chapter 8, pages 122–152. ISBN 94-010-5199-2, 94-011-2706-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 145. URL https:/

/link.springer.com/chapter/10.
1007/978-94-011-2706-6_8.

Amsterdamski:1992:BHM

[Ams92b] Stefan Amsterdamski. Between History and Method: Disputes about the Rationality of Science, volume 145 of Boston Studies in the Philosophy of Science. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 1992. ISBN 94-010-5199-2, 94-011-2706-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). xi + 227 pp. LCCN Q174 .B67 vol. 145. URL

https://link.springer.com/book/
10.1007/978-94-011-2706-6.

Amsterdamski:1992:DKI

[Ams92c] Stefan Amsterdamski. The development of knowledge and the ideals of science. In Between History and Method: Disputes about the Rationality of Science [Ams92b], chapter 2, pages 9–26. ISBN 94-010-5199-2, 94-011-2706-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 145. URL https:/

/link.springer.com/chapter/10.
1007/978-94-011-2706-6_2.

Amsterdamski:1992:EWT

[Ams92d] Stefan Amsterdamski. Escape to world three. In Between History and Method: Disputes about the Rationality of Science [Ams92b], chapter 7, pages 101–121. ISBN 94-010-5199-2, 94-011-2706-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 145. URL https:/

/link.springer.com/chapter/10.
1007/978-94-011-2706-6_7.

Amsterdamski:1992:ISR

[Ams92e] Stefan Amsterdamski. Ideals of science and rules of explanation. In Between History and Method: Disputes about the Rationality of Science [Ams92b], chapter 3, pages 27–43. ISBN 94-010-5199-2, 94-011-2706-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 145.

URL https://link.springer.com/
chapter/10.1007/978-94-011-2706-
6_3.

Amsterdamski:1992:IPS

[Ams92f] Stefan Amsterdamski. The institutionalization and professionalization of scientific research. In Between History and Method: Disputes about the Rationality of Science [Ams92b], chapter 5, pages 65–78. ISBN 94-010-5199-2, 94-011-2706-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 145. URL https:/

/link.springer.com/chapter/10.
1007/978-94-011-2706-6_5.

Amsterdamski:1992:I

[Ams92g] Stefan Amsterdamski. Introduction. In Between History and Method: Disputes about the Rationality of Science [Ams92b], chapter 1, pages 1–8. ISBN 94-010-5199-2, 94-011-2706-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 145.

URL https://link.springer.com/
chapter/10.1007/978-94-011-2706-
6_1.

Amsterdamski:1992:MIS

[Ams92h] Stefan Amsterdamski. The modern ideal of science. In Between History and Method: Disputes about the Rationality of Science [Ams92b], chapter 4, pages 44–64. ISBN 94-010-5199-2, 94-011-2706-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 145. URL https:/

/link.springer.com/chapter/10.
1007/978-94-011-2706-6_4.

Amsterdamski:1992:OA

[Ams92i] Stefan Amsterdamski. Order and anarchy. In Between History and Method: Disputes about the Rationality of Science [Ams92b], chapter 9, pages 153–176. ISBN 94-010-5199-2, 94-011-2706-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 145. URL https:/

/link.springer.com/chapter/10.
1007/978-94-011-2706-6_9.

Amsterdamski:1992:SCM

[Ams92j] Stefan Amsterdamski. The sources of the crisis of the modern ideal of science. In Between History and Method: Disputes about the Rationality of Science [Ams92b], chapter 6, pages 79–100. ISBN 94-010-5199-2, 94-011-2706-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 145. URL https://link.

springer.com/chapter/10.1007/
978-94-011-2706-6_6.

Arabatzis:2015:CTC

[AN15] Theodore Arabatzis and Nancy J.

Nersessian. Concepts out of theoretical contexts. In Arabatzis et al. [ARS15], chapter 15, pages 225–238. ISBN 3-319-14552-5 (paperback), 3-319-14553-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q125 .R45

2015. URL https://link.springer.
com/chapter/10.1007/978-3-319-
14553-2_15.

Anapolitanos:1989:PRR

[Ana89] D. A. Anapolitanos. Proofs and refutations: A reassessment. In Imre Lakatos and Theories of Scientific Change [GGN89], chapter 24, pages 337–345. ISBN 94-010-7860-2, 94-009-3025-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-009-3025-
4_24.

Anagnostopoulos:1990:STE

[Ana90a] Georgios Anagnostopoulos. Some thoughts on explanation in ancient philosophy. In Greek Studies in the Philosophy and History of Science [Nic90], chapter 4, pages 61–80. ISBN 94-010-7403-8, 94-009-2015-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-009-2015-6_4.

Anapolitanos:1990:LDS

[Ana90b] D. A. Anapolitanos. Leibniz on density and sequential or Cauchy completeness. In Greek Studies in the Philosophy and History of Science [Nic90], chapter 23, pages 361–372. ISBN 94-010-7403-8, 94-009-2015-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

                                              ???? URL https://link.springer.
                                              com/chapter/10.1007/978-94-009-
                                              2015-6_23.

Anandan:1997:CQP

[Ana97] Jeeva S. Anandan. Classical and quantum physical geometry. In Cohen et al. [CHS97b], pages 31–52. ISBN 90-481-4809-X, 94-017-2732-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC173.96-174.52.

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-017-2732-
                                              7_3.

Anderson:1976:IE

[And76] Robert M. Anderson, Jr. The illusions of experience. In Cohen et al. [CHMV76], pages 549–561. ISBN 90-277-0647-6, 90-277-0648-4 (paperback), 94-010-1449-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic).

                                              LCCN ????  URL https://link.
                                              springer.com/chapter/10.1007/
                                              978-94-010-1449-6_31.

Andersson:1978:PV

[And78] Gunnar Andersson. The problem of verisimilitude. In Progress and Rationality in Science [RA78b], chapter 13, pages 291–310. ISBN 90-277-0922-X, 94-009-9866-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic).

LCCN ????  URL https://link.
springer.com/chapter/10.1007/
978-94-009-9866-7_13.

Andersson:1979:PPP

[And79] Gunnar Andersson. Presuppositions, problems, progress. In The Structure and Development of Science [RA79], chapter 1, pages 3–15. ISBN 90-277-0995-5, 94-009-9459-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .S778 1979.

URL https://link.springer.com/
chapter/10.1007/978-94-009-9459-
1_1.

Andersson:1984:CCS

[And84a] Gunnar Andersson. Creativity and criticism in science and politics. In Rationality in Science and Politics [And84d], chapter 1, pages 1–14. ISBN 90-277-1953-5, 94-009-6254-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-009-6254-5_1.

Andersson:1984:GRP

[And84b] Gunnar Andersson. Gerard Radnitzky:

From positivism, via critical theory, to critical rationalism. In Rationality in Science and Politics [And84d], chapter 17, pages 289–295. ISBN 90-277-1953-5, 94-009-6254-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-009-6254-5_17.

Andersson:1984:HAF

[And84c] Gunnar Andersson. How to accept fallible test statements? Popper’s criticist solution. In Rationality in Science and Politics [And84d], chapter 4, pages 47– 68. ISBN 90-277-1953-5, 94-009-6254-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-009-6254-
5_4.

Andersson:1984:RSP

[And84d] Gunnar Andersson. Rationality in Science and Politics, volume 79 of Boston Studies in the Philosophy of Science. D. Reidel, Dordrecht, The Netherlands; Boston, MA, USA; Lancaster, UK; Tokyo, Japan, 1984. ISBN 90-277-1953-5, 94-009-6254-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). xi + 307 pp. LCCN ????

URL https://link.springer.com/
book/10.1007/978-94-009-6254-5.

Andriopoulos:1990:AHC

[And90] D. Z. Andriopoulos. Alcmeon’s and Hippocrates’s concept of Aetia. In Greek Studies in the Philosophy and History of Science [Nic90], chapter 5, pages 81–90. ISBN 94-010-7403-8, 94-009-2015-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-009-
2015-6_5.

Andersson:1991:FFG

[And91] Gunnar Andersson. Feyerabend on falsifications, Galileo, and Lady Reason. In Beyond Reason: Essays on the Philosophy of Paul Feyerabend [Mun91a], chapter 12, pages 281–295. ISBN 94-010-5406-1, 94-011-3188-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-011-3188-
9_12.

Anderson:1999:NVT

[And99] R. Lanier Anderson. Nietzsche’s views on truth and the Kantian background of his epistemology. In Nietzsche, Epistemology, and Philosophy of Science: Nietzsche and the Sciences II [Bab99a], chapter 4, pages 47–59. ISBN 90-481-5234-8, 94-017-2428-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-017-2428-9_4.

Anderson:2003:TEU

[And03a] James L. Anderson. Timekeeping in an expanding universe. In Renn et al. [RDS03], chapter 11, pages 275–280. ISBN 1-4020-1284-5, 1-4020-1285-3 (paperback). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 234. URL https://link.

springer.com/chapter/10.1007/
978-94-010-0111-3_11.

Andler:2003:ATH

[And03b] Daniel Andler. The advantages of theft over honest toil: Comments on David Atkinson. In Descartes’s Mathematical Thought [Sas03d], chapter 15, pages 227–236. ISBN 90-481-6487-7, 94-017-1225-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/0-306-48123-
5_15.

Andersson:2009:CRP

[And09a] Gunnar Andersson. Critical rationalism and the principle of sufficient reason. In Rethinking Popper [PC09], chapter 2, pages 21–30. ISBN 1-4020-9337-3, 1-4020-9338-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN B1649.P64 R48 2009; Q174 .B67 2009.

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-1-4020-9338-
                                              8_2.

Andler:2009:PCS

[And09b] Daniel Andler. Philosophy of cognitive science. In Brenner and Gayon [BG09], chapter 12, pages 255–300. ISBN 1-4020-9367-5, 1-4020-9368-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .F746 2009.

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-1-4020-9368-
                                              5_12.

Andrault:2018:WDI

[And18] Raphaële Andrault. What does it mean to be an empiricist in medicine? Baglivi’s De praxi medica (1696). In What Does it Mean to be an Empiricist?: Empiricisms in Eighteenth Century Sciences [BR18b], chapter 9, pages 169–188. ISBN 3-319-69858-3 (hardcover), 3-319-69860-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN B816 .W43 2018.

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-3-319-69860-
                                              1_9.

Anonymous:1963:FM

[Ano63] Anonymous. Front matter. In Wartofsky [War63], pages i–viii. ISBN 94-010-3265-3, 94-010-3263-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B678 1970.

                                              URL https://link.springer.com/
                                              book/10.1007/978-94-010-3263-6.

Anonymous:1967:FM

[Ano67] Anonymous. Front matter. In Cohen and Wartofsky [CW67], pages I–XLIX. ISBN 90-277-0013-3, 94-010-3508-3. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175 .B731 v.3.

URL https://link.springer.com/
book/10.1007/978-94-010-3508-8.

Anonymous:1969:BM

[Ano69a] Anonymous. Back matter. In Cohen and Wartofsky [CW69a], pages 483–484. ISBN 94-010-3383-8, 94-010-3263-7, 94-010-3381-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL

https://link.springer.com/book/
10.1007/978-94-010-3381-7.

Anonymous:1969:FMa

[Ano69b] Anonymous. Front matter. In Cohen and Wartofsky [CW69b], pages I–VIII. ISBN 94-010-3380-3, 94-010-3378-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
book/10.1007/978-94-010-3378-7.

Anonymous:1969:FMb

[Ano69c] Anonymous. Front matter. In Cohen and Wartofsky [CW69a], pages I–VIII. ISBN 94-010-3383-8, 94-010-3263-7, 94-010-3381-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL

https://link.springer.com/book/
10.1007/978-94-010-3381-7.

Anonymous:1970:BM

[Ano70a] Anonymous. Back matter. In Ernst Mach: Physicist and Philosopher [CS70], pages 217–298. ISBN 90-481-8318-9, 90-277-0016-8, 94-017-1462-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
book/10.1007/978-94-017-1462-4.

Anonymous:1970:FM

[Ano70b] Anonymous. Front matter. In Ernst Mach: Physicist and Philosopher [CS70], pages N3–VIII. ISBN 90-481-8318-9, 90-277-0016-8, 94-017-1462-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
book/10.1007/978-94-017-1462-4.

Anonymous:1971:BMa

[Ano71a] Anonymous. Back matter. In Bergson and Modern Physics: A Reinterpretation and Re-evaluation [Čap71c], pages 335–417. ISBN 94-010-3098-7, 94-010-3096-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic).

LCCN ????  URL https://link.
springer.com/book/10.1007/978-
94-010-3096-0.

Anonymous:1971:BMb

[Ano71b] Anonymous. Back matter. In Buck and Cohen [BC71], page 616. ISBN 90-277-0309-4, 94-010-3142-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL

https://link.springer.com/book/
10.1007/978-94-010-3142-4.

Anonymous:1971:FMa

[Ano71c] Anonymous. Front matter. In Bergson and Modern Physics: A Reinterpretation and Re-evaluation [Čap71c], pages I–XV. ISBN 94-010-3098-7, 94-010-3096-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic).

LCCN ????  URL https://link.
springer.com/book/10.1007/978-
94-010-3096-0.

Anonymous:1971:FMb

[Ano71d] Anonymous. Front matter. In Bergson and Modern Physics: A Reinterpretation and Re-evaluation [Čap71c], page 1. ISBN 94-010-3098-7, 94-010-3096-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/book/10.1007/978-94-010-
3096-0.

Anonymous:1971:FMc

[Ano71e] Anonymous. Front matter. In Bergson and Modern Physics: A Reinterpretation and Re-evaluation [Čap71c], page 81. ISBN 94-010-3098-7, 94-010-3096-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/book/10.1007/978-94-010-
3096-0.

Anonymous:1971:FMd

[Ano71f] Anonymous. Front matter. In Bergson and Modern Physics: A Reinterpretation and Re-evaluation [Čap71c], page 187. ISBN 94-010-3098-7, 94-010-3096-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
book/10.1007/978-94-010-3096-0.

Anonymous:1971:FMf

[Ano71g] Anonymous. Front matter. In Buck and Cohen [BC71], page 1. ISBN 90-277-0309-4, 94-010-3142-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL

https://link.springer.com/book/
10.1007/978-94-010-3142-4.

Anonymous:1971:FMg

[Ano71h] Anonymous. Front matter. In Buck and Cohen [BC71], page 47. ISBN 90-277-0309-4, 94-010-3142-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL

https://link.springer.com/book/
10.1007/978-94-010-3142-4.

Anonymous:1971:FMh

[Ano71i] Anonymous. Front matter. In Buck and Cohen [BC71], page 89. ISBN 90-277-0309-4, 94-010-3142-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL

https://link.springer.com/book/
10.1007/978-94-010-3142-4.

Anonymous:1971:FMi

[Ano71j] Anonymous. Front matter. In Buck and Cohen [BC71], page 183. ISBN 90-277-0309-4, 94-010-3142-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL

https://link.springer.com/book/
10.1007/978-94-010-3142-4.

Anonymous:1971:FMe

[Ano71k] Anonymous. Front matter. In Buck and Cohen [BC71], pages N1–LXVI. ISBN 90-277-0309-4, 94-010-3142-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
book/10.1007/978-94-010-3142-4.

Anonymous:1973:BMa

[Ano73a] Anonymous. Back matter. In Foundations of the Logical Theory of Scientific Knowledge (Complex Logic) [Zin73g], pages 265–304. ISBN 90-277-0324-8, 94-010-2501-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL

https://link.springer.com/book/
10.1007/978-94-010-2501-0.

Anonymous:1973:BMd

[Ano73b] Anonymous. Back matter. In Scientific Procedures: A Contribution Concerning the Methodological Problems of Scientific Concepts and Scientific Explanation [Ton73f], pages 265–271. ISBN 90-277-0147-4 (hardcover), 90-277-0323-X (paperback), 94-010-2916-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 10; Q175 .T663s 1973. URL

https://link.springer.com/book/
10.1007/978-94-010-2916-2.

Anonymous:1973:BMb

[Ano73c] Anonymous. Back matter. In Philosophical Problems of Space and Time [Grü73s], pages 429–446. ISBN 90-277-0358-2, 94-010-2622-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL

https://link.springer.com/book/
10.1007/978-94-010-2622-2.

Anonymous:1973:BMc

[Ano73d] Anonymous. Back matter. In Philosophical Problems of Space and Time [Grü73s], pages 804–884. ISBN 90-277-0358-2, 94-010-2622-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL

https://link.springer.com/book/
10.1007/978-94-010-2622-2.

Anonymous:1973:FMa

[Ano73e] Anonymous. Front matter. In Foundations of the Logical Theory of Scientific Knowledge (Complex Logic) [Zin73g], pages I–XXII. ISBN 90-277-0324-8, 94-010-2501-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL

https://link.springer.com/book/
10.1007/978-94-010-2501-0.

Anonymous:1973:FMg

[Ano73f] Anonymous. Front matter. In Scientific Procedures: A Contribution Concerning the Methodological Problems of Scientific Concepts and Scientific Explanation [Ton73f], pages I–XIII. ISBN 90-277-0147-4 (hardcover), 90-277-0323-X (paperback), 94-010-2916-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 10; Q175 .T663s 1973. URL

                                              https://link.springer.com/book/
                                              10.1007/978-94-010-2916-2.

Anonymous:1973:FMb

[Ano73g] Anonymous. Front matter. In Philosophical Problems of Space and Time [Grü73s], pages I–XXIII. ISBN 90-277-0358-2, 94-010-2622-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL

                                              https://link.springer.com/book/
                                              10.1007/978-94-010-2622-2.

Anonymous:1973:FMc

[Ano73h] Anonymous. Front matter. In Philosophical Problems of Space and Time [Grü73s], page 1. ISBN 90-277-0358-2, 94-010-2622-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic).

                                              LCCN ????  URL https://link.
                                              springer.com/book/10.1007/978-
                                              94-010-2622-2.

Anonymous:1973:FMd

[Ano73i]  Anonymous.  Front matter.  In

Philosophical Problems of Space and Time [Grü73s], page 177. ISBN 90-277-0358-2, 94-010-2622-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL

                                              https://link.springer.com/book/
                                              10.1007/978-94-010-2622-2.

Anonymous:1973:FMe

[Ano73j]  Anonymous.  Front matter.  In

Philosophical Problems of Space and Time [Grü73s], page 339. ISBN 90-277-0358-2, 94-010-2622-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL

                                              https://link.springer.com/book/
                                              10.1007/978-94-010-2622-2.

Anonymous:1973:FMf

[Ano73k] Anonymous.  Front matter.  In

Philosophical Problems of Space and Time [Grü73s], page 447. ISBN 90-277-0358-2, 94-010-2622-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL

https://link.springer.com/book/
10.1007/978-94-010-2622-2.

Anonymous:1974:BMe

[Ano74a] Anonymous. Back matter. In Logical and Epistemological Studies in Contemporary Physics [CW74a], pages 397–466. ISBN 90-277-0391-4 (hardcover), 90-277-0377-9 (paperback), 94-010-2656-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 13; QC5.56.

URL https://link.springer.com/
book/10.1007/978-94-010-2656-7.

Anonymous:1974:BMb

[Ano74b] Anonymous. Back matter. In Methodological and Historical Essays in the Natural and Social Sciences [CW74b], pages 406–410. ISBN 90-277-0378-7, 94-010-2128-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 14. URL

https://link.springer.com/book/
10.1007/978-94-010-2128-9.

Anonymous:1974:BMa

[Ano74c] Anonymous. Back matter. In Selected Papers on Language and the Brain [Ges74w], pages 542–555. ISBN 90-277-0263-2, 94-010-2093-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL

https://link.springer.com/book/
10.1007/978-94-010-2093-0.

Anonymous:1974:BMc

[Ano74d] Anonymous. Back matter. In The Understanding of Nature: Essays in the Philosophy of Biology [Gre74t], pages 361–379. ISBN 90-277-0463-5, 94-010-2224-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 23; QH331. URL

                                              https://link.springer.com/book/
                                              10.1007/978-94-010-2224-8.

Anonymous:1974:BMd

[Ano74e] Anonymous. Back matter. In The Method of Analysis: Its Geometrical Origin and Its General Significance [HR74d], pages 118–149. ISBN 90-277-0543-7, 94-010-2296-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL

                                              https://link.springer.com/book/
                                              10.1007/978-94-010-2296-5.

Anonymous:1974:FMg

[Ano74f] Anonymous. Front matter. In Seeger and Cohen [SC74b], pages I–IX. ISBN 90-277-0390-6, 90-277-0376-0 (paperback), 94-010-2126-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 11. URL

                                              https://link.springer.com/book/
                                              10.1007/978-94-010-2126-5.

Anonymous:1974:FMh

[Ano74g] Anonymous. Front matter. In Seeger and Cohen [SC74b], page 1. ISBN 90-277-0390-6, 90-277-0376-0 (paperback), 94-010-2126-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 11. URL

                                              https://link.springer.com/book/
                                              10.1007/978-94-010-2126-5.

Anonymous:1974:FMi

[Ano74h] Anonymous. Front matter. In Seeger and Cohen [SC74b], page 117. ISBN 90-277-0390-6, 90-277-0376-0 (paperback), 94-010-2126-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 11. URL

https://link.springer.com/book/
10.1007/978-94-010-2126-5.

Anonymous:1974:FMj

[Ano74i] Anonymous. Front matter. In Seeger and Cohen [SC74b], page 153. ISBN 90-277-0390-6, 90-277-0376-0 (paperback), 94-010-2126-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 11. URL

https://link.springer.com/book/
10.1007/978-94-010-2126-5.

Anonymous:1974:FMk

[Ano74j] Anonymous. Front matter. In Seeger and Cohen [SC74b], page 177. ISBN 90-277-0390-6, 90-277-0376-0 (paperback), 94-010-2126-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 11. URL

https://link.springer.com/book/
10.1007/978-94-010-2126-5.

Anonymous:1974:FMl

[Ano74k] Anonymous. Front matter. In Seeger and Cohen [SC74b], page 303. ISBN 90-277-0390-6, 90-277-0376-0 (paperback), 94-010-2126-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 11. URL

https://link.springer.com/book/
10.1007/978-94-010-2126-5.

Anonymous:1974:FMm

[Ano74l] Anonymous. Front matter. In Seeger and Cohen [SC74b], page 363. ISBN 90-277-0390-6, 90-277-0376-0 (paperback), 94-010-2126-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 11. URL

https://link.springer.com/book/
10.1007/978-94-010-2126-5.

Anonymous:1974:FMn

[Ano74m] Anonymous. Front matter. In Seeger and Cohen [SC74b], page 495. ISBN 90-277-0390-6, 90-277-0376-0 (paperback), 94-010-2126-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 11. URL

                                              https://link.springer.com/book/
                                              10.1007/978-94-010-2126-5.

Anonymous:1974:FMae

[Ano74n] Anonymous. Front matter. In Logical and Epistemological Studies in Contemporary Physics [CW74a], pages I–VIII. ISBN 90-277-0391-4 (hardcover), 90-277-0377-9 (paperback), 94-010-2656-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 13; QC5.56.

                                              URL https://link.springer.com/
                                              book/10.1007/978-94-010-2656-7.

Anonymous:1974:FMo

[Ano74o] Anonymous. Front matter. In Methodological and Historical Essays in the Natural and Social Sciences [CW74b], pages I–VIII. ISBN 90-277-0378-7, 94-010-2128-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 14. URL

                                              https://link.springer.com/book/
                                              10.1007/978-94-010-2128-9.

Anonymous:1974:FMb

[Ano74p] Anonymous. Front matter. In For Dirk Struik: Scientific, Historical and Political Essays in Honor of Dirk J. Struik [CSW74], pages I–XXVII. ISBN 90-277-0379-5, 94-010-2115-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

                                              URL https://link.springer.com/
                                              book/10.1007/978-94-010-2115-9.

Anonymous:1974:FMc

[Ano74q] Anonymous. Front matter. In For Dirk Struik: Scientific, Historical and Political Essays in Honor of Dirk J. Struik [CSW74], page 1. ISBN 90-277-0379-5, 94-010-2115-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL

https://link.springer.com/book/
10.1007/978-94-010-2115-9.

Anonymous:1974:FMd

[Ano74r] Anonymous. Front matter. In For Dirk Struik: Scientific, Historical and Political Essays in Honor of Dirk J. Struik [CSW74], page 141. ISBN 90-277-0379-5, 94-010-2115-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL

https://link.springer.com/book/
10.1007/978-94-010-2115-9.

Anonymous:1974:FMe

[Ano74s] Anonymous. Front matter. In For Dirk Struik: Scientific, Historical and Political Essays in Honor of Dirk J. Struik [CSW74], page 293. ISBN 90-277-0379-5, 94-010-2115-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL

https://link.springer.com/book/
10.1007/978-94-010-2115-9.

Anonymous:1974:FMf

[Ano74t] Anonymous. Front matter. In For Dirk Struik: Scientific, Historical and Political Essays in Honor of Dirk J. Struik [CSW74], page 491. ISBN 90-277-0379-5, 94-010-2115-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL

https://link.springer.com/book/
10.1007/978-94-010-2115-9.

Anonymous:1974:FMa

[Ano74u] Anonymous. Front matter. In Selected Papers on Language and the Brain [Ges74w], pages I–XII. ISBN 90-277-0263-2, 94-010-2093-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL

                                              https://link.springer.com/book/
                                              10.1007/978-94-010-2093-0.

Anonymous:1974:FMp

[Ano74v] Anonymous. Front matter. In Schaffner and Cohen [SC74a], pages i–ix. ISBN 90-277-0408-2, 90-277-0409-0, 94-010-2140-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175 .B7312. URL https://link.

                                              springer.com/book/10.1007/978-
                                              94-010-2140-1.

Anonymous:1974:FMq

[Ano74w] Anonymous. Front matter. In Schaffner and Cohen [SC74a], page 1. ISBN 90-277-0408-2, 90-277-0409-0, 94-010-2140-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175 .B7312. URL https://link.

                                              springer.com/book/10.1007/978-
                                              94-010-2140-1.

Anonymous:1974:FMr

[Ano74x] Anonymous. Front matter. In Schaffner and Cohen [SC74a], page 55. ISBN 90-277-0408-2, 90-277-0409-0, 94-010-2140-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175 .B7312. URL https://link.

                                              springer.com/book/10.1007/978-
                                              94-010-2140-1.

Anonymous:1974:FMs

[Ano74y] Anonymous. Front matter. In Schaffner and Cohen [SC74a], page 99. ISBN 90-277-0408-2, 90-277-0409-0, 94-010-2140-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175 .B7312. URL https://link.

springer.com/book/10.1007/978-
94-010-2140-1.

Anonymous:1974:FMt

[Ano74z] Anonymous. Front matter. In Schaffner and Cohen [SC74a], page 151. ISBN 90-277-0408-2, 90-277-0409-0, 94-010-2140-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175 .B7312. URL https://link.

springer.com/book/10.1007/978-
94-010-2140-1.

Anonymous:1974:FMu

[Ano74-27] Anonymous. Front matter. In Schaffner and Cohen [SC74a], page 165. ISBN 90-277-0408-2, 90-277-0409-0, 94-010-2140-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175 .B7312. URL https://link.

springer.com/book/10.1007/978-
94-010-2140-1.

Anonymous:1974:FMv

[Ano74-28] Anonymous. Front matter. In Schaffner and Cohen [SC74a], page 217. ISBN 90-277-0408-2, 90-277-0409-0, 94-010-2140-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175 .B7312. URL https://link.

springer.com/book/10.1007/978-
94-010-2140-1.

Anonymous:1974:FMw

[Ano74-29] Anonymous. Front matter. In Schaffner and Cohen [SC74a], page 249. ISBN 90-277-0408-2, 90-277-0409-0, 94-010-2140-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175 .B7312. URL https://link.

springer.com/book/10.1007/978-
94-010-2140-1.

Anonymous:1974:FMx

[Ano74-30] Anonymous. Front matter. In Schaffner and Cohen [SC74a], page 283. ISBN 90-277-0408-2, 90-277-0409-0, 94-010-2140-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175 .B7312. URL https://link.

                                              springer.com/book/10.1007/978-
                                              94-010-2140-1.

Anonymous:1974:FMy

[Ano74-31] Anonymous. Front matter. In Schaffner and Cohen [SC74a], page 335. ISBN 90-277-0408-2, 90-277-0409-0, 94-010-2140-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175 .B7312. URL https://link.

                                              springer.com/book/10.1007/978-
                                              94-010-2140-1.

Anonymous:1974:FMz

[Ano74-32] Anonymous. Front matter. In Schaffner and Cohen [SC74a], page 365. ISBN 90-277-0408-2, 90-277-0409-0, 94-010-2140-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175 .B7312. URL https://link.

                                              springer.com/book/10.1007/978-
                                              94-010-2140-1.

Anonymous:1974:FMaa

[Ano74-33] Anonymous. Front matter. In Schaffner and Cohen [SC74a], page 385. ISBN 90-277-0408-2, 90-277-0409-0, 94-010-2140-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175 .B7312. URL https://link.

                                              springer.com/book/10.1007/978-
                                              94-010-2140-1.

Anonymous:1974:FMab

[Ano74-34] Anonymous. Front matter. In Schaffner and Cohen [SC74a], page 421. ISBN 90-277-0408-2, 90-277-0409-0, 94-010-2140-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175 .B7312. URL https://link.

springer.com/book/10.1007/978-
94-010-2140-1.

Anonymous:1974:FMac

[Ano74-35] Anonymous. Front matter. In The Understanding of Nature: Essays in the Philosophy of Biology [Gre74t], pages I–XII. ISBN 90-277-0463-5, 94-010-2224-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 23; QH331. URL

https://link.springer.com/book/
10.1007/978-94-010-2224-8.

Anonymous:1974:FMad

[Ano74-36] Anonymous. Front matter. In The Method of Analysis: Its Geometrical Origin and Its General Significance [HR74d], pages I–XVIII. ISBN 90-277-0543-7, 94-010-2296-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL

https://link.springer.com/book/
10.1007/978-94-010-2296-5.

Anonymous:1975:BMa

[Ano75a] Anonymous. Back matter. In Murdoch and Sylla [MS75b], pages 539–571. ISBN 90-277-0587-9, 94-010-1781-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN B721

.I584 1975.  URL https://link.
springer.com/book/10.1007/978-
94-010-1781-7.

Anonymous:1975:BMc

[Ano75b] Anonymous. Back matter. In Science in Flux [Aga75p], pages 525–535. ISBN 90-277-0612-3, 94-010-1810-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
book/10.1007/978-94-010-1810-4.

Anonymous:1975:BMb

[Ano75c] Anonymous. Back matter. In Between Experience and Metaphysics: Philosophical Problems of the Evolution of Science [Ams75a], pages 177–198. ISBN 90-277-0580-1, 94-010-1797-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .A478 1975.

                                              URL https://link.springer.com/
                                              book/10.1007/978-94-010-1797-8.

Anonymous:1975:FMa

[Ano75d] Anonymous. Front matter. In Murdoch and Sylla [MS75b], pages I–X. ISBN 90-277-0587-9, 94-010-1781-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN B721 .I584 1975.

                                              URL https://link.springer.com/
                                              book/10.1007/978-94-010-1781-7.

Anonymous:1975:FMb

[Ano75e] Anonymous. Front matter. In Murdoch and Sylla [MS75b], page 31. ISBN 90-277-0587-9, 94-010-1781-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN B721 .I584 1975.

                                              URL https://link.springer.com/
                                              book/10.1007/978-94-010-1781-7.

Anonymous:1975:FMc

[Ano75f] Anonymous. Front matter. In Murdoch and Sylla [MS75b], page 149. ISBN 90-277-0587-9, 94-010-1781-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN B721 .I584 1975.

                                              URL https://link.springer.com/
                                              book/10.1007/978-94-010-1781-7.

Anonymous:1975:FMd

[Ano75g] Anonymous. Front matter. In Murdoch and Sylla [MS75b], page 269. ISBN 90-277-0587-9, 94-010-1781-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN B721 .I584 1975.

URL https://link.springer.com/
book/10.1007/978-94-010-1781-7.

Anonymous:1975:FMf

[Ano75h] Anonymous. Front matter. In Science in Flux [Aga75p], pages II–XXVI. ISBN 90-277-0612-3, 94-010-1810-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
book/10.1007/978-94-010-1810-4.

Anonymous:1975:FMe

[Ano75i] Anonymous. Front matter. In Between Experience and Metaphysics: Philosophical Problems of the Evolution of Science [Ams75a], pages I–XVIII. ISBN 90-277-0580-1, 94-010-1797-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .A478 1975.

URL https://link.springer.com/
book/10.1007/978-94-010-1797-8.

Anonymous:1976:BMg

[Ano76a] Anonymous. Back matter. In Philosophical Problems of Modern Physics [Mit76g], pages 205–211. ISBN 90-277-0506-2, 94-010-9617-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC6 .M588 1976.

URL https://link.springer.com/
book/10.1007/978-94-010-9617-1.

Anonymous:1976:BMe

[Ano76b] Anonymous. Back matter. In Topics in the Philosophy of Biology [GM76], pages 441–454. ISBN 90-277-0596-8, 94-010-1829-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic).

LCCN ????  URL https://link.
springer.com/book/10.1007/978-
94-010-1829-6.

Anonymous:1976:BMa

[Ano76c] Anonymous. Back matter. In Essays in Memory of Imre Lakatos [CFW76], pages 755–768. ISBN 90-277-0655-7, 94-010-1451-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL

https://link.springer.com/book/
10.1007/978-94-010-1451-9.

Anonymous:1976:BMf

[Ano76d] Anonymous. Back matter. In Language in Focus: Foundations, Methods and Systems: Essays in Memory of Yehoshua Bar-Hillel [Kas76b], pages 675–685. ISBN 90-277-0644-1, 90-277-0645-X, 94-010-1876-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN P101-120. URL

https://link.springer.com/book/
10.1007/978-94-010-1876-0.

Anonymous:1976:BMb

[Ano76e] Anonymous. Back matter. In Paranoia: a Study in Diagnosis [FA76g], pages 108–218. ISBN 90-277-0705-7, 94-010-1506-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic).

LCCN ????  URL https://link.
springer.com/book/10.1007/978-
94-010-1506-6.

Anonymous:1976:BMd

[Ano76f] Anonymous. Back matter. In Surrender and Catch: Experience and Inquiry Today [Wol76-32], pages 394–410. ISBN 90-277-0765-0, 94-010-1526-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
book/10.1007/978-94-010-1526-4.

Anonymous:1976:BMc

[Ano76g] Anonymous. Back matter. In Dialectics of the Concrete: A Study on Problems of Man and World [Kos76a], pages 156– 164. ISBN 94-011-9766-0, 94-010-1520-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL

https://link.springer.com/book/
10.1007/978-94-010-1520-2.

Anonymous:1976:FMal

[Ano76h] Anonymous. Front matter. In Philosophical Problems of Modern Physics [Mit76g], pages III–X. ISBN 90-277-0506-2, 94-010-9617-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC6 .M588 1976. URL

https://link.springer.com/book/
10.1007/978-94-010-9617-1.

Anonymous:1976:FMt

[Ano76i] Anonymous. Front matter. In The Concepts of Space and Time: Their Structure and Their Development [Čap76], pages I–LVII. ISBN 90-277-0375-2, 94-010-1727-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN BD632. URL

https://link.springer.com/book/
10.1007/978-94-010-1727-5.

Anonymous:1976:FMu

[Ano76j] Anonymous. Front matter. In The Concepts of Space and Time: Their Structure and Their Development [Čap76], page 1. ISBN 90-277-0375-2, 94-010-1727-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic).

LCCN BD632. URL https://link.
springer.com/book/10.1007/978-
94-010-1727-5.

Anonymous:1976:FMv

[Ano76k] Anonymous. Front matter. In The Concepts of Space and Time: Their Structure and Their Development [Čap76], page 135. ISBN 90-277-0375-2, 94-010-1727-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic).

LCCN BD632. URL https://link.
springer.com/book/10.1007/978-
94-010-1727-5.

Anonymous:1976:FMw

[Ano76l] Anonymous. Front matter. In The Concepts of Space and Time: Their Structure and Their Development [Čap76], page 265. ISBN 90-277-0375-2, 94-010-1727-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic).

                                              LCCN BD632. URL https://link.
                                              springer.com/book/10.1007/978-
                                              94-010-1727-5.

Anonymous:1976:FMx

[Ano76m] Anonymous. Front matter. In Topics in the Philosophy of Biology [GM76], pages I–XIII. ISBN 90-277-0596-8, 94-010-1829-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic).

                                              LCCN ????  URL https://link.
                                              springer.com/book/10.1007/978-
                                              94-010-1829-6.

Anonymous:1976:FMy

[Ano76n] Anonymous. Front matter. In Topics in the Philosophy of Biology [GM76], page 1. ISBN 90-277-0596-8, 94-010-1829-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

                                              ???? URL https://link.springer.
                                              com/book/10.1007/978-94-010-
                                              1829-6.

Anonymous:1976:FMz

[Ano76o] Anonymous. Front matter. In Topics in the Philosophy of Biology [GM76], page 99. ISBN 90-277-0596-8, 94-010-1829-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

                                              ???? URL https://link.springer.
                                              com/book/10.1007/978-94-010-
                                              1829-6.

Anonymous:1976:FMaa

[Ano76p] Anonymous. Front matter. In Topics in the Philosophy of Biology [GM76], page 143. ISBN 90-277-0596-8, 94-010-1829-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
book/10.1007/978-94-010-1829-6.

Anonymous:1976:FMab

[Ano76q] Anonymous. Front matter. In Topics in the Philosophy of Biology [GM76], page 265. ISBN 90-277-0596-8, 94-010-1829-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
book/10.1007/978-94-010-1829-6.

Anonymous:1976:FMac

[Ano76r] Anonymous. Front matter. In Topics in the Philosophy of Biology [GM76], page 351. ISBN 90-277-0596-8, 94-010-1829-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
book/10.1007/978-94-010-1829-6.

Anonymous:1976:FMa

[Ano76s] Anonymous. Front matter. In Cohen et al. [CHMV76], pages I–XIII. ISBN 90-277-0647-6, 90-277-0648-4 (paperback), 94-010-1449-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL

https://link.springer.com/book/
10.1007/978-94-010-1449-6.

Anonymous:1976:FMb

[Ano76t] Anonymous. Front matter. In Cohen et al. [CHMV76], page 1. ISBN 90-277-0647-6, 90-277-0648-4 (paperback), 94-010-1449-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL

https://link.springer.com/book/
10.1007/978-94-010-1449-6.

Anonymous:1976:FMc

[Ano76u] Anonymous. Front matter. In Cohen et al. [CHMV76], page 77. ISBN 90-277-0647-6, 90-277-0648-4 (paperback), 94-010-1449-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL

https://link.springer.com/book/
10.1007/978-94-010-1449-6.

Anonymous:1976:FMd

[Ano76v] Anonymous. Front matter. In Cohen et al. [CHMV76], page 125. ISBN 90-277-0647-6, 90-277-0648-4 (paperback), 94-010-1449-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL

https://link.springer.com/book/
10.1007/978-94-010-1449-6.

Anonymous:1976:FMe

[Ano76w] Anonymous. Front matter. In Cohen et al. [CHMV76], page 189. ISBN 90-277-0647-6, 90-277-0648-4 (paperback), 94-010-1449-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL

https://link.springer.com/book/
10.1007/978-94-010-1449-6.

Anonymous:1976:FMf

[Ano76x] Anonymous. Front matter. In Cohen et al. [CHMV76], page 291. ISBN 90-277-0647-6, 90-277-0648-4 (paperback), 94-010-1449-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL

https://link.springer.com/book/
10.1007/978-94-010-1449-6.

Anonymous:1976:FMg

[Ano76y] Anonymous. Front matter. In Cohen et al. [CHMV76], page 323. ISBN 90-277-0647-6, 90-277-0648-4 (paperback), 94-010-1449-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL

https://link.springer.com/book/
10.1007/978-94-010-1449-6.

Anonymous:1976:FMh

[Ano76z] Anonymous. Front matter. In Cohen et al. [CHMV76], page 435. ISBN 90-277-0647-6, 90-277-0648-4 (paperback), 94-010-1449-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL

https://link.springer.com/book/
10.1007/978-94-010-1449-6.

Anonymous:1976:FMi

[Ano76-27] Anonymous. Front matter. In Cohen et al. [CHMV76], page 467. ISBN 90-277-0647-6, 90-277-0648-4 (paperback), 94-010-1449-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL

https://link.springer.com/book/
10.1007/978-94-010-1449-6.

Anonymous:1976:FMj

[Ano76-28] Anonymous. Front matter. In Cohen et al. [CHMV76], page 499. ISBN 90-277-0647-6, 90-277-0648-4 (paperback), 94-010-1449-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL

https://link.springer.com/book/
10.1007/978-94-010-1449-6.

Anonymous:1976:FMk

[Ano76-29] Anonymous. Front matter. In Cohen et al. [CHMV76], page 527. ISBN 90-277-0647-6, 90-277-0648-4 (paperback), 94-010-1449-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL

https://link.springer.com/book/
10.1007/978-94-010-1449-6.

Anonymous:1976:FMl

[Ano76-30] Anonymous. Front matter. In Cohen et al. [CHMV76], page 563. ISBN 90-277-0647-6, 90-277-0648-4 (paperback), 94-010-1449-3 (e-book).

ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL

https://link.springer.com/book/
10.1007/978-94-010-1449-6.

Anonymous:1976:FMm

[Ano76-31] Anonymous. Front matter. In Cohen et al. [CHMV76], page 611. ISBN 90-277-0647-6, 90-277-0648-4 (paperback), 94-010-1449-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL

https://link.springer.com/book/
10.1007/978-94-010-1449-6.

Anonymous:1976:FMn

[Ano76-32] Anonymous. Front matter. In Essays in Memory of Imre Lakatos [CFW76], pages I–XI. ISBN 90-277-0655-7, 94-010-1451-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic).

LCCN ????  URL https://link.
springer.com/book/10.1007/978-
94-010-1451-9.

Anonymous:1976:FMad

[Ano76-33] Anonymous. Front matter. In Language in Focus: Foundations, Methods and Systems: Essays in Memory of Yehoshua Bar-Hillel [Kas76b], pages I–XXVIII. ISBN 90-277-0644-1, 90-277-0645-X, 94-010-1876-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN P101-120. URL

https://link.springer.com/book/
10.1007/978-94-010-1876-0.

Anonymous:1976:FMae

[Ano76-34] Anonymous. Front matter. In Language in Focus: Foundations, Methods and Systems: Essays in Memory of Yehoshua Bar-Hillel [Kas76b], page 1. ISBN 90-277-0644-1, 90-277-0645-X, 94-010-1876-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN P101-120. URL

https://link.springer.com/book/
10.1007/978-94-010-1876-0.

Anonymous:1976:FMaf

[Ano76-35] Anonymous. Front matter. In Language in Focus: Foundations, Methods and Systems: Essays in Memory of Yehoshua Bar-Hillel [Kas76b], page 105. ISBN 90-277-0644-1, 90-277-0645-X, 94-010-1876-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN P101-120. URL

https://link.springer.com/book/
10.1007/978-94-010-1876-0.

Anonymous:1976:FMag

[Ano76-36] Anonymous. Front matter. In Language in Focus: Foundations, Methods and Systems: Essays in Memory of Yehoshua Bar-Hillel [Kas76b], page 177. ISBN 90-277-0644-1, 90-277-0645-X, 94-010-1876-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN P101-120. URL

https://link.springer.com/book/
10.1007/978-94-010-1876-0.

Anonymous:1976:FMah

[Ano76-37] Anonymous. Front matter. In Language in Focus: Foundations, Methods and Systems: Essays in Memory of Yehoshua Bar-Hillel [Kas76b], page 279. ISBN 90-277-0644-1, 90-277-0645-X, 94-010-1876-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN P101-120. URL

https://link.springer.com/book/
10.1007/978-94-010-1876-0.

Anonymous:1976:FMai

[Ano76-38] Anonymous. Front matter. In Language in Focus: Foundations, Methods and Systems: Essays in Memory of Yehoshua Bar-Hillel [Kas76b], page 405. ISBN 90-277-0644-1, 90-277-0645-X, 94-010-1876-6 (e-book).

ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN P101-120. URL

                                              https://link.springer.com/book/
                                              10.1007/978-94-010-1876-0.

Anonymous:1976:FMaj

[Ano76-39] Anonymous. Front matter. In Language in Focus: Foundations, Methods and Systems: Essays in Memory of Yehoshua Bar-Hillel [Kas76b], page 485. ISBN 90-277-0644-1, 90-277-0645-X, 94-010-1876-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN P101-120. URL

                                              https://link.springer.com/book/
                                              10.1007/978-94-010-1876-0.

Anonymous:1976:FMak

[Ano76-40] Anonymous. Front matter. In Language in Focus: Foundations, Methods and Systems: Essays in Memory of Yehoshua Bar-Hillel [Kas76b], page 585. ISBN 90-277-0644-1, 90-277-0645-X, 94-010-1876-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN P101-120. URL

                                              https://link.springer.com/book/
                                              10.1007/978-94-010-1876-0.

Anonymous:1976:FMo

[Ano76-41] Anonymous. Front matter. In Paranoia: a Study in Diagnosis [FA76g], pages I– XV. ISBN 90-277-0705-7, 94-010-1506-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL

                                              https://link.springer.com/book/
                                              10.1007/978-94-010-1506-6.

Anonymous:1976:FMq

[Ano76-42] Anonymous. Front matter. In Surrender and Catch: Experience and Inquiry Today [Wol76-32], pages i–xii. ISBN 90-277-0765-0, 94-010-1526-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
book/10.1007/978-94-010-1526-4.

Anonymous:1976:FMr

[Ano76-43] Anonymous. Front matter. In Surrender and Catch: Experience and Inquiry Today [Wol76-32], page 1. ISBN 90-277-0765-0, 94-010-1526-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
book/10.1007/978-94-010-1526-4.

Anonymous:1976:FMs

[Ano76-44] Anonymous. Front matter. In Surrender and Catch: Experience and Inquiry Today [Wol76-32], page 215. ISBN 90-277-0765-0, 94-010-1526-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
book/10.1007/978-94-010-1526-4.

Anonymous:1976:FMp

[Ano76-45] Anonymous. Front matter. In Dialectics of the Concrete: A Study on Problems of Man and World [Kos76a], pages I– VIII. ISBN 94-011-9766-0, 94-010-1520-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
book/10.1007/978-94-010-1520-2.

Anonymous:1977:BMa

[Ano77a] Anonymous. Back matter. In von Helmholtz [vH77a], pages 186–214. ISBN 90-277-0582-8, 94-010-1115-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
book/10.1007/978-94-010-1115-0.

Anonymous:1977:BMb

[Ano77b] Anonymous. Back matter. In The Structure of Appearance [Goo77j], pages 277–294. ISBN 90-277-0774-X, 94-010-1184-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic).

LCCN ????  URL https://link.
springer.com/book/10.1007/978-
94-010-1184-6.

Anonymous:1977:BMc

[Ano77c] Anonymous. Back matter. In Models of Discovery: and Other Topics in the Methods of Science [Sim77m], pages 441–464. ISBN 90-277-0970-X, 94-010-9521-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic).

LCCN ????  URL https://link.
springer.com/book/10.1007/978-
94-010-9521-1.

Anonymous:1977:FMa

[Ano77d] Anonymous. Front matter. In von Helmholtz [vH77a], pages I–XXXVII. ISBN 90-277-0582-8, 94-010-1115-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
book/10.1007/978-94-010-1115-0.

Anonymous:1977:FMb

[Ano77e] Anonymous. Front matter. In The Structure of Appearance [Goo77j], pages I–L. ISBN 90-277-0774-X, 94-010-1184-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
book/10.1007/978-94-010-1184-6.

Anonymous:1977:FMc

[Ano77f] Anonymous. Front matter. In The Structure of Appearance [Goo77j], page 1. ISBN 90-277-0774-X, 94-010-1184-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/book/10.1007/978-94-010-
1184-6.

Anonymous:1977:FMd

[Ano77g] Anonymous. Front matter. In The Structure of Appearance [Goo77j], page 91. ISBN 90-277-0774-X, 94-010-1184-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/book/10.1007/978-94-010-
1184-6.

Anonymous:1977:FMe

[Ano77h] Anonymous. Front matter. In The Structure of Appearance [Goo77j], page 191. ISBN 90-277-0774-X, 94-010-1184-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
book/10.1007/978-94-010-1184-6.

Anonymous:1977:FMf

[Ano77i] Anonymous. Front matter. In Models of Discovery: and Other Topics in the Methods of Science [Sim77m], pages i– xix. ISBN 90-277-0970-X, 94-010-9521-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL

https://link.springer.com/book/
10.1007/978-94-010-9521-1.

Anonymous:1977:FMg

[Ano77j] Anonymous. Front matter. In Models of Discovery: and Other Topics in the Methods of Science [Sim77m], pages 1–6. ISBN 90-277-0970-X, 94-010-9521-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
book/10.1007/978-94-010-9521-1.

Anonymous:1977:FMh

[Ano77k] Anonymous. Front matter. In Models of Discovery: and Other Topics in the Methods of Science [Sim77m], pages 47– 52. ISBN 90-277-0970-X, 94-010-9521-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
book/10.1007/978-94-010-9521-1.

Anonymous:1977:FMi

[Ano77l] Anonymous. Front matter. In Models of Discovery: and Other Topics in the Methods of Science [Sim77m], pages 133–136. ISBN 90-277-0970-X, 94-010-9521-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic).

LCCN ????  URL https://link.
springer.com/book/10.1007/978-
94-010-9521-1.

Anonymous:1977:FMj

[Ano77m] Anonymous. Front matter. In Models of Discovery: and Other Topics in the Methods of Science [Sim77m], pages 177–181. ISBN 90-277-0970-X, 94-010-9521-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic).

LCCN ????  URL https://link.
springer.com/book/10.1007/978-
94-010-9521-1.

Anonymous:1977:FMk

[Ano77n] Anonymous. Front matter. In Models of Discovery: and Other Topics in the Methods of Science [Sim77m], pages 263–267. ISBN 90-277-0970-X, 94-010-9521-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic).

LCCN ????  URL https://link.
springer.com/book/10.1007/978-
94-010-9521-1.

Anonymous:1977:FMl

[Ano77o] Anonymous. Front matter. In Models of Discovery: and Other Topics in the Methods of Science [Sim77m], pages 343–348. ISBN 90-277-0970-X, 94-010-9521-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic).

LCCN ????  URL https://link.

springer.com/book/10.1007/978-
94-010-9521-1.

Anonymous:1978:BMa

[Ano78a] Anonymous. Back matter. In Persons and Minds: The Prospects of Nonreductive Materialism [Mar78i], pages 263–310. ISBN 90-277-0863-0, 94-009-9801-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic).

LCCN ????  URL https://link.
springer.com/book/10.1007/978-
94-009-9801-8.

Anonymous:1978:BMb

[Ano78b] Anonymous. Back matter. In Progress and Rationality in Science [RA78b], pages 379–388. ISBN 90-277-0922-X, 94-009-9866-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic).

LCCN ????  URL https://link.
springer.com/book/10.1007/978-
94-009-9866-7.

Anonymous:1978:BMc

[Ano78c] Anonymous. Back matter. In Progress and Rationality in Science [RA78b], pages 397–427. ISBN 90-277-0922-X, 94-009-9866-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic).

LCCN ????  URL https://link.
springer.com/book/10.1007/978-
94-009-9866-7.

Anonymous:1978:FMa

[Ano78d] Anonymous. Front matter. In Persons and Minds: The Prospects of Nonreductive Materialism [Mar78i], pages I–XIII. ISBN 90-277-0863-0, 94-009-9801-5 (ebook). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
book/10.1007/978-94-009-9801-8.

Anonymous:1978:FMb

[Ano78e] Anonymous. Front matter. In Persons and Minds: The Prospects of Nonreductive Materialism [Mar78i], page 1. ISBN 90-277-0863-0, 94-009-9801-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

                                              URL https://link.springer.com/
                                              book/10.1007/978-94-009-9801-8.

Anonymous:1978:FMc

[Ano78f] Anonymous. Front matter. In Persons and Minds: The Prospects of Nonreductive Materialism [Mar78i], page 27. ISBN 90-277-0863-0, 94-009-9801-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

                                              URL https://link.springer.com/
                                              book/10.1007/978-94-009-9801-8.

Anonymous:1978:FMd

[Ano78g] Anonymous. Front matter. In Persons and Minds: The Prospects of Nonreductive Materialism [Mar78i], page 79. ISBN 90-277-0863-0, 94-009-9801-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

                                              URL https://link.springer.com/
                                              book/10.1007/978-94-009-9801-8.

Anonymous:1978:FMe

[Ano78h] Anonymous. Front matter. In Persons and Minds: The Prospects of Nonreductive Materialism [Mar78i], page 197. ISBN 90-277-0863-0, 94-009-9801-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

                                              URL https://link.springer.com/
                                              book/10.1007/978-94-009-9801-8.

Anonymous:1978:FMf

[Ano78i] Anonymous. Front matter. In Progress and Rationality in Science [RA78b], pages i–x. ISBN 90-277-0922-X, 94-009-9866-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/book/10.1007/978-94-009-
9866-7.

Anonymous:1978:FMg

[Ano78j] Anonymous. Front matter. In Progress and Rationality in Science [RA78b], page 1. ISBN 90-277-0922-X, 94-009-9866-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/book/10.1007/978-94-009-
9866-7.

Anonymous:1978:FMh

[Ano78k] Anonymous. Front matter. In Progress and Rationality in Science [RA78b], page 21. ISBN 90-277-0922-X, 94-009-9866-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/book/10.1007/978-94-009-
9866-7.

Anonymous:1978:FMi

[Ano78l] Anonymous. Front matter. In Progress and Rationality in Science [RA78b], page 115. ISBN 90-277-0922-X, 94-009-9866-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/book/10.1007/978-94-009-
9866-7.

Anonymous:1978:FMj

[Ano78m] Anonymous. Front matter. In Progress and Rationality in Science [RA78b], page 319. ISBN 90-277-0922-X, 94-009-9866-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/book/10.1007/978-94-009-
9866-7.

Anonymous:1978:FMk

[Ano78n] Anonymous. Front matter. In Progress and Rationality in Science [RA78b], page 385. ISBN 90-277-0922-X, 94-009-9866-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

                                              ???? URL https://link.springer.
                                              com/book/10.1007/978-94-009-
                                              9866-7.

Anonymous:1979:BMf

[Ano79a] Anonymous. Back matter. In Technics and Praxis [Ihd79i], pages 151–163. ISBN 90-277-0954-8, 94-009-9900-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

                                              URL https://link.springer.com/
                                              book/10.1007/978-94-009-9900-8.

Anonymous:1979:BMa

[Ano79b] Anonymous. Back matter. In Polish Essays in the Methodology of the Social Sciences [Wia79b], pages 249–260. ISBN 90-277-0956-4, 94-009-9353-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174

                                              .P655 1979. URL https://link.
                                              springer.com/book/10.1007/978-
                                              94-009-9353-2.

Anonymous:1979:BMb

[Ano79c] Anonymous. Back matter. In Praxis: Yugoslav Essays in the Philosophy and Methodology of the Social Sciences [MP79], pages 371–415. ISBN 90-277-0968-8, 94-009-9355-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL

                                              https://link.springer.com/book/
                                              10.1007/978-94-009-9355-6.

Anonymous:1979:BMd

[Ano79d] Anonymous. Back matter. In Pragmatics, Truth, and Language [Mar79o], pages 296–312. ISBN 90-277-0993-9, 94-009-9457-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic).

LCCN ????  URL https://link.
springer.com/book/10.1007/978-
94-009-9457-7.

Anonymous:1979:BMc

[Ano79e] Anonymous. Back matter. In Models:

Representation and the Scientific Understanding [War79k], pages 371–398. ISBN 90-277-0947-5, 94-009-9357-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 48; Q175.3. URL https://

link.springer.com/book/10.1007/
978-94-009-9357-0.

Anonymous:1979:BMe

[Ano79f] Anonymous. Back matter. In The Structure and Development of Science [RA79], pages 267–282. ISBN 90-277-0995-5, 94-009-9459-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .S778 1979.

URL https://link.springer.com/
book/10.1007/978-94-009-9459-1.

Anonymous:1979:FMa

[Ano79g] Anonymous. Front matter. In Cohen and Stachel [CS79a], pages I–XXXIV. ISBN 90-277-0651-4, 90-277-0652-2 (paperback), 94-009-9349-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 21 QC7.

URL https://link.springer.com/
book/10.1007/978-94-009-9349-5.

Anonymous:1979:FMb

[Ano79h] Anonymous. Front matter. In Cohen and Stachel [CS79a], page 1. ISBN 90-277-0651-4, 90-277-0652-2 (paperback), 94-009-9349-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 21 QC7. URL

https://link.springer.com/book/
10.1007/978-94-009-9349-5.

Anonymous:1979:FMc

[Ano79i] Anonymous. Front matter. In Cohen and Stachel [CS79a], page 355. ISBN 90-277-0651-4, 90-277-0652-2 (paperback), 94-009-9349-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 21 QC7.

                                              URL https://link.springer.com/
                                              book/10.1007/978-94-009-9349-5.

Anonymous:1979:FMp

[Ano79j] Anonymous. Front matter. In Technics and Praxis [Ihd79i], pages i–xxviii. ISBN 90-277-0954-8, 94-009-9900-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

                                              URL https://link.springer.com/
                                              book/10.1007/978-94-009-9900-8.

Anonymous:1979:FMq

[Ano79k] Anonymous. Front matter. In Technics and Praxis [Ihd79i], page 1. ISBN 90-277-0954-8, 94-009-9900-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL

                                              https://link.springer.com/book/
                                              10.1007/978-94-009-9900-8.

Anonymous:1979:FMr

[Ano79l] Anonymous. Front matter. In Technics and Praxis [Ihd79i], page 51. ISBN 90-277-0954-8, 94-009-9900-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

                                              URL https://link.springer.com/
                                              book/10.1007/978-94-009-9900-8.

Anonymous:1979:FMs

[Ano79m] Anonymous. Front matter. In Technics and Praxis [Ihd79i], page 101. ISBN 90-277-0954-8, 94-009-9900-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

                                              URL https://link.springer.com/
                                              book/10.1007/978-94-009-9900-8.

Anonymous:1979:FMd

[Ano79n] Anonymous. Front matter. In Polish Essays in the Methodology of the Social Sciences [Wia79b], pages i–xiii. ISBN 90-277-0956-4, 94-009-9353-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .P655 1979.

URL https://link.springer.com/
book/10.1007/978-94-009-9353-2.

Anonymous:1979:FMe

[Ano79o] Anonymous. Front matter. In Praxis:

Yugoslav Essays in the Philosophy and Methodology of the Social Sciences [MP79], pages i–xxxvii. ISBN 90-277-0968-8, 94-009-9355-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL

https://link.springer.com/book/
10.1007/978-94-009-9355-6.

Anonymous:1979:FMf

[Ano79p] Anonymous. Front matter. In Praxis:

Yugoslav Essays in the Philosophy and Methodology of the Social Sciences [MP79], page 1. ISBN 90-277-0968-8, 94-009-9355-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic).

LCCN ????  URL https://link.
springer.com/book/10.1007/978-
94-009-9355-6.

Anonymous:1979:FMg

[Ano79q] Anonymous. Front matter. In Praxis:

Yugoslav Essays in the Philosophy and Methodology of the Social Sciences [MP79], page 99. ISBN 90-277-0968-8, 94-009-9355-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic).

LCCN ????  URL https://link.
springer.com/book/10.1007/978-
94-009-9355-6.

Anonymous:1979:FMh

[Ano79r] Anonymous. Front matter. In Praxis: Yugoslav Essays in the Philosophy and Methodology of the Social Sciences [MP79], page 165. ISBN 90-277-0968-8, 94-009-9355-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic).

LCCN ????  URL https://link.
springer.com/book/10.1007/978-
94-009-9355-6.

Anonymous:1979:FMi

[Ano79s] Anonymous. Front matter. In Praxis: Yugoslav Essays in the Philosophy and Methodology of the Social Sciences [MP79], page 227. ISBN 90-277-0968-8, 94-009-9355-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic).

LCCN ????  URL https://link.
springer.com/book/10.1007/978-
94-009-9355-6.

Anonymous:1979:FMk

[Ano79t] Anonymous. Front matter. In Pragmatics, Truth, and Language [Mar79o], pages i–xv. ISBN 90-277-0993-9, 94-009-9457-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic).

LCCN ????  URL https://link.
springer.com/book/10.1007/978-
94-009-9457-7.

Anonymous:1979:FMj

[Ano79u] Anonymous. Front matter. In Models: Representation and the Scientific Understanding [War79k], pages i–xxvi. ISBN 90-277-0947-5, 94-009-9357-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 48; Q175.3. URL https://

link.springer.com/book/10.1007/
978-94-009-9357-0.

Anonymous:1979:FMl

[Ano79v] Anonymous. Front matter. In The Structure and Development of Science [RA79], pages i–ix. ISBN 90-277-0995-5, 94-009-9459-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .S778 1979. URL

https://link.springer.com/book/
10.1007/978-94-009-9459-1.

Anonymous:1979:FMm

[Ano79w] Anonymous. Front matter. In The Structure and Development of Science [RA79], page 1. ISBN 90-277-0995-5, 94-009-9459-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .S778 1979. URL

https://link.springer.com/book/
10.1007/978-94-009-9459-1.

Anonymous:1979:FMn

[Ano79x] Anonymous. Front matter. In The Structure and Development of Science [RA79], page 17. ISBN 90-277-0995-5, 94-009-9459-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .S778 1979. URL

https://link.springer.com/book/
10.1007/978-94-009-9459-1.

Anonymous:1979:FMo

[Ano79y] Anonymous. Front matter. In The Structure and Development of Science [RA79], page 149. ISBN 90-277-0995-5, 94-009-9459-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .S778 1979. URL

https://link.springer.com/book/
10.1007/978-94-009-9459-1.

Anonymous:1980:BMa

[Ano80a] Anonymous. Back matter. In Time, Causality, and the Quantum Theory: Studies in the Philosophy of Science. Volume 1: Essay on the Causal Theory of Time [Coh80], pages 286–307. ISBN 90-277-1074-0, 94-009-8935-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

                                              URL https://link.springer.com/
                                              book/10.1007/978-94-009-8935-1.

Anonymous:1980:BMe

[Ano80b] Anonymous. Back matter. In Time, Causality, and the Quantum Theory: Studies in the Philosophy of Science. Volume Two: Time in a Quantized Universe [MC80z], pages 262–294. ISBN 90-277-1076-7, 94-009-8988-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN BJ1-1725. URL

                                              https://link.springer.com/book/
                                              10.1007/978-94-009-8988-7.

Anonymous:1980:BMb

[Ano80c] Anonymous. Back matter. In Autopoiesis and Cognition: The Realization of the Living [MV80a], pages 57– 58. ISBN 90-277-1016-3, 94-009-8947-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

                                              URL https://link.springer.com/
                                              book/10.1007/978-94-009-8947-4.

Anonymous:1980:BMc

[Ano80d] Anonymous. Back matter. In Autopoiesis and Cognition: The Realization of the Living [MV80a], pages 124– 145. ISBN 90-277-1016-3, 94-009-8947-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL

                                              https://link.springer.com/book/
                                              10.1007/978-94-009-8947-4.

Anonymous:1980:BMd

[Ano80e] Anonymous. Back matter. In Scientific Discovery, Logic, and Rationality [Nic80f], pages 375–388. ISBN 90-277-1070-8, 94-009-8986-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL

                                              https://link.springer.com/book/
                                              10.1007/978-94-009-8986-3.

Anonymous:1980:BMf

[Ano80f] Anonymous. Back matter. In Scientific Discovery: Case Studies [Nic80e], pages 367–382. ISBN 90-277-1093-7, 94-009-9015-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 60. URL

https://link.springer.com/book/
10.1007/978-94-009-9015-9.

Anonymous:1980:BMg

[Ano80g] Anonymous. Back matter. In Galileo and the Art of Reasoning: Rhetorical Foundations of Logic and Scientific Method [Fin80h], pages 432–481. ISBN 90-277-1095-3, 94-009-9017-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL

https://link.springer.com/book/
10.1007/978-94-009-9017-3.

Anonymous:1980:FMa

[Ano80h] Anonymous. Front matter. In Time, Causality, and the Quantum Theory: Studies in the Philosophy of Science. Volume 1: Essay on the Causal Theory of Time [Coh80], pages i–xix. ISBN 90-277-1074-0, 94-009-8935-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
book/10.1007/978-94-009-8935-1.

Anonymous:1980:FMb

[Ano80i] Anonymous. Front matter. In Time, Causality, and the Quantum Theory: Studies in the Philosophy of Science. Volume 1: Essay on the Causal Theory of Time [Coh80], pages 37–41. ISBN 90-277-1074-0, 94-009-8935-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
book/10.1007/978-94-009-8935-1.

Anonymous:1980:FMc

[Ano80j] Anonymous. Front matter. In Time, Causality, and the Quantum Theory: Studies in the Philosophy of Science. Volume 1: Essay on the Causal Theory of Time [Coh80], page 161. ISBN 90-277-1074-0, 94-009-8935-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL

https://link.springer.com/book/
10.1007/978-94-009-8935-1.

Anonymous:1980:FMp

[Ano80k] Anonymous. Front matter. In Time, Causality, and the Quantum Theory: Studies in the Philosophy of Science. Volume Two: Time in a Quantized Universe [MC80z], pages i–x. ISBN 90-277-1076-7, 94-009-8988-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN BJ1-1725. URL

https://link.springer.com/book/
10.1007/978-94-009-8988-7.

Anonymous:1980:FMq

[Ano80l] Anonymous. Front matter. In Time, Causality, and the Quantum Theory: Studies in the Philosophy of Science. Volume Two: Time in a Quantized Universe [MC80z], page 1. ISBN 90-277-1076-7, 94-009-8988-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN BJ1-1725. URL

https://link.springer.com/book/
10.1007/978-94-009-8988-7.

Anonymous:1980:FMr

[Ano80m] Anonymous. Front matter. In Time, Causality, and the Quantum Theory: Studies in the Philosophy of Science. Volume Two: Time in a Quantized Universe [MC80z], page 209. ISBN 90-277-1076-7, 94-009-8988-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN BJ1-1725. URL

https://link.springer.com/book/
10.1007/978-94-009-8988-7.

Anonymous:1980:FMd

[Ano80n] Anonymous. Front matter. In Autopoiesis and Cognition: The Realization of the Living [MV80a], pages i–xxx. ISBN 90-277-1016-3, 94-009-8947-4 (ebook). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
book/10.1007/978-94-009-8947-4.

Anonymous:1980:FMe

[Ano80o] Anonymous.  Front matter.  In

Autopoiesis and Cognition: The Realization of the Living [MV80a], pages 1–3. ISBN 90-277-1016-3, 94-009-8947-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
book/10.1007/978-94-009-8947-4.

Anonymous:1980:FMf

[Ano80p] Anonymous. Front matter. In Autopoiesis and Cognition: The Realization of the Living [MV80a], pages 59– 61. ISBN 90-277-1016-3, 94-009-8947-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
book/10.1007/978-94-009-8947-4.

Anonymous:1980:FMg

[Ano80q] Anonymous. Front matter. In Experiment, Theory, Practice: Articles and Addresses [Kap80j], pages i–xxvi. ISBN 90-277-1062-7, 94-009-8977-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
book/10.1007/978-94-009-8977-1.

Anonymous:1980:FMh

[Ano80r] Anonymous. Front matter. In Experiment, Theory, Practice: Articles and Addresses [Kap80j], page 1. ISBN 90-277-1062-7, 94-009-8977-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL

https://link.springer.com/book/
10.1007/978-94-009-8977-1.

Anonymous:1980:FMi

[Ano80s] Anonymous. Front matter. In Experiment, Theory, Practice: Articles and Addresses [Kap80j], page 101. ISBN 90-277-1062-7, 94-009-8977-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
book/10.1007/978-94-009-8977-1.

Anonymous:1980:FMj

[Ano80t] Anonymous. Front matter. In Experiment, Theory, Practice: Articles and Addresses [Kap80j], page 133. ISBN 90-277-1062-7, 94-009-8977-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
book/10.1007/978-94-009-8977-1.

Anonymous:1980:FMk

[Ano80u] Anonymous. Front matter. In Experiment, Theory, Practice: Articles and Addresses [Kap80j], page 195. ISBN 90-277-1062-7, 94-009-8977-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
book/10.1007/978-94-009-8977-1.

Anonymous:1980:FMl

[Ano80v] Anonymous. Front matter. In Experiment, Theory, Practice: Articles and Addresses [Kap80j], page 227. ISBN 90-277-1062-7, 94-009-8977-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
book/10.1007/978-94-009-8977-1.

Anonymous:1980:FMm

[Ano80w] Anonymous. Front matter. In Experiment, Theory, Practice: Articles and Addresses [Kap80j], page 277. ISBN 90-277-1062-7, 94-009-8977-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
book/10.1007/978-94-009-8977-1.

Anonymous:1980:FMn

[Ano80x] Anonymous. Front matter. In Experiment, Theory, Practice: Articles and Addresses [Kap80j], page 335. ISBN 90-277-1062-7, 94-009-8977-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
book/10.1007/978-94-009-8977-1.

Anonymous:1980:FMo

[Ano80y] Anonymous.  Front matter.  In

Scientific Discovery, Logic, and Rationality [Nic80f], pages i–xii. ISBN 90-277-1070-8, 94-009-8986-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL

https://link.springer.com/book/
10.1007/978-94-009-8986-3.

Anonymous:1980:FMs

[Ano80z] Anonymous. Front matter. In Scientific Discovery: Case Studies [Nic80e], pages i–xxv. ISBN 90-277-1093-7, 94-009-9015-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 60. URL

https://link.springer.com/book/
10.1007/978-94-009-9015-9.

Anonymous:1980:FMt

[Ano80-27] Anonymous. Front matter. In Galileo and the Art of Reasoning: Rhetorical Foundations of Logic and Scientific Method [Fin80h], pages i–xx. ISBN 90-277-1095-3, 94-009-9017-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL

https://link.springer.com/book/
10.1007/978-94-009-9017-3.

Anonymous:1980:FMu

[Ano80-28] Anonymous. Front matter. In Galileo and the Art of Reasoning: Rhetorical Foundations of Logic and Scientific Method [Fin80h], page 1. ISBN 90-277-1095-3, 94-009-9017-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL

https://link.springer.com/book/
10.1007/978-94-009-9017-3.

Anonymous:1980:FMv

[Ano80-29] Anonymous. Front matter. In Galileo and the Art of Reasoning: Rhetorical Foundations of Logic and Scientific Method [Fin80h], page 143. ISBN 90-277-1095-3, 94-009-9017-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL

https://link.springer.com/book/
10.1007/978-94-009-9017-3.

Anonymous:1980:FMw

[Ano80-30] Anonymous. Front matter. In Galileo and the Art of Reasoning: Rhetorical Foundations of Logic and Scientific Method [Fin80h], page 309. ISBN 90-277-1095-3, 94-009-9017-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL

https://link.springer.com/book/
10.1007/978-94-009-9017-3.

Anonymous:1981:BMf

[Ano81a] Anonymous. Back matter. In Grmek et al. [GCC81b], pages 323–336. ISBN 90-277-1123-2, 94-010-1284-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL

https://link.springer.com/book/
10.1007/978-94-010-1284-3.

Anonymous:1981:BMe

[Ano81b] Anonymous. Back matter. In Italian Studies in the Philosophy of Science [Chi81b], pages 517–529. ISBN 90-277-1073-2, 94-009-8937-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL

https://link.springer.com/book/
10.1007/978-94-009-8937-5.

Anonymous:1981:BMb

[Ano81c] Anonymous. Back matter. In Prelude to Galileo: Essays on Medieval and Sixteenth-Century Sources of Galileo’s Thought [Wal81q], pages 349–372. ISBN 90-277-1215-8, 90-277-1216-6 (paperback), 94-009-8404-9. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 62. URL

https://link.springer.com/book/
10.1007/978-94-009-8404-2.

Anonymous:1981:BMc

[Ano81d] Anonymous. Back matter. In Analytical Philosophy of Technology [Rap81a], pages 187–202. ISBN 90-277-1222-0, 94-009-8412-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL

https://link.springer.com/book/
10.1007/978-94-009-8412-7.

Anonymous:1981:BMg

[Ano81e] Anonymous. Back matter. In Science and Society: Studies in the Sociology of Science [Aga81z], pages 502–535. ISBN 94-011-6458-4, 94-011-6456-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
book/10.1007/978-94-011-6456-6.

Anonymous:1981:BMa

[Ano81f] Anonymous. Back matter. In Problems of Semantics: A Contribution to the Analysis of the Language Science [Ton81f], pages 354–407. ISBN 90-277-0316-7, 94-009-8364-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL

                                              https://link.springer.com/book/
                                              10.1007/978-94-009-8364-9.

Anonymous:1981:BMd

[Ano81g] Anonymous. Back matter. In Scientific Knowledge: Causation, Explanation, and Corroboration [Fet81h], pages 297– 327. ISBN 90-277-1336-7, 94-009-8558-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL

                                              https://link.springer.com/book/
                                              10.1007/978-94-009-8558-2.

Anonymous:1981:FMp

[Ano81h] Anonymous. Front matter. In Grmek et al. [GCC81b], pages i–vii. ISBN 90-277-1123-2, 94-010-1284-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL

                                              https://link.springer.com/book/
                                              10.1007/978-94-010-1284-3.

Anonymous:1981:FMq

[Ano81i] Anonymous. Front matter. In Grmek et al. [GCC81b], page 7. ISBN 90-277-1123-2, 94-010-1284-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL

                                              https://link.springer.com/book/
                                              10.1007/978-94-010-1284-3.

Anonymous:1981:FMr

[Ano81j] Anonymous. Front matter. In Grmek et al. [GCC81b], page 227. ISBN 90-277-1123-2, 94-010-1284-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL

https://link.springer.com/book/
10.1007/978-94-010-1284-3.

Anonymous:1981:FMl

[Ano81k] Anonymous. Front matter. In Italian Studies in the Philosophy of Science [Chi81b], pages i–xi. ISBN 90-277-1073-2, 94-009-8937-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL

https://link.springer.com/book/
10.1007/978-94-009-8937-5.

Anonymous:1981:FMm

[Ano81l] Anonymous. Front matter. In Italian Studies in the Philosophy of Science [Chi81b], page 1. ISBN 90-277-1073-2, 94-009-8937-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic).

LCCN ????  URL https://link.
springer.com/book/10.1007/978-
94-009-8937-5.

Anonymous:1981:FMn

[Ano81m] Anonymous. Front matter. In Italian Studies in the Philosophy of Science [Chi81b], page 185. ISBN 90-277-1073-2, 94-009-8937-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL

https://link.springer.com/book/
10.1007/978-94-009-8937-5.

Anonymous:1981:FMo

[Ano81n] Anonymous. Front matter. In Italian Studies in the Philosophy of Science [Chi81b], page 361. ISBN 90-277-1073-2, 94-009-8937-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL

https://link.springer.com/book/
10.1007/978-94-009-8937-5.

Anonymous:1981:FMb

[Ano81o] Anonymous. Front matter. In Prelude to Galileo: Essays on Medieval and Sixteenth-Century Sources of Galileo’s Thought [Wal81q], pages i–xvi. ISBN 90-277-1215-8, 90-277-1216-6 (paperback), 94-009-8404-9. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 62.

URL https://link.springer.com/
book/10.1007/978-94-009-8404-2.

Anonymous:1981:FMc

[Ano81p] Anonymous. Front matter. In Prelude to Galileo: Essays on Medieval and Sixteenth-Century Sources of Galileo’s Thought [Wal81q], page 1. ISBN 90-277-1215-8, 90-277-1216-6 (paperback), 94-009-8404-9. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 62. URL

https://link.springer.com/book/
10.1007/978-94-009-8404-2.

Anonymous:1981:FMd

[Ano81q] Anonymous. Front matter. In Prelude to Galileo: Essays on Medieval and Sixteenth-Century Sources of Galileo’s Thought [Wal81q], page 49. ISBN 90-277-1215-8, 90-277-1216-6 (paperback), 94-009-8404-9. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 62. URL

https://link.springer.com/book/
10.1007/978-94-009-8404-2.

Anonymous:1981:FMe

[Ano81r] Anonymous. Front matter. In Prelude to Galileo: Essays on Medieval and Sixteenth-Century Sources of Galileo’s Thought [Wal81q], page 127. ISBN 90-277-1215-8, 90-277-1216-6 (paperback), 94-009-8404-9. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 62. URL

https://link.springer.com/book/
10.1007/978-94-009-8404-2.

Anonymous:1981:FMf

[Ano81s] Anonymous. Front matter. In Prelude to Galileo: Essays on Medieval and Sixteenth-Century Sources of Galileo’s Thought [Wal81q], page 301. ISBN 90-277-1215-8, 90-277-1216-6 (paperback), 94-009-8404-9. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 62. URL

https://link.springer.com/book/
10.1007/978-94-009-8404-2.

Anonymous:1981:FMg

[Ano81t] Anonymous. Front matter. In Analytical Philosophy of Technology [Rap81a], pages i–xiv. ISBN 90-277-1222-0, 94-009-8412-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL

https://link.springer.com/book/
10.1007/978-94-009-8412-7.

Anonymous:1981:FMs

[Ano81u] Anonymous. Front matter. In Science and Society: Studies in the Sociology of Science [Aga81z], pages i–xxi. ISBN 94-011-6458-4, 94-011-6456-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
book/10.1007/978-94-011-6456-6.

Anonymous:1981:FMa

[Ano81v] Anonymous. Front matter. In Problems of Semantics: A Contribution to the Analysis of the Language Science [Ton81f], pages i–xiv. ISBN 90-277-0316-7, 94-009-8364-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL

https://link.springer.com/book/
10.1007/978-94-009-8364-9.

Anonymous:1981:FMh

[Ano81w] Anonymous. Front matter. In Scientific Knowledge: Causation, Explanation, and Corroboration [Fet81h], pages i– xvi. ISBN 90-277-1336-7, 94-009-8558-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

                                              URL https://link.springer.com/
                                              book/10.1007/978-94-009-8558-2.

Anonymous:1981:FMi

[Ano81x] Anonymous. Front matter. In Scientific Knowledge: Causation, Explanation, and Corroboration [Fet81h], page 1. ISBN 90-277-1336-7, 94-009-8558-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

                                              URL https://link.springer.com/
                                              book/10.1007/978-94-009-8558-2.

Anonymous:1981:FMj

[Ano81y] Anonymous. Front matter. In Scientific Knowledge: Causation, Explanation, and Corroboration [Fet81h], page 75. ISBN 90-277-1336-7, 94-009-8558-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

                                              URL https://link.springer.com/
                                              book/10.1007/978-94-009-8558-2.

Anonymous:1981:FMk

[Ano81z] Anonymous. Front matter. In Scientific Knowledge: Causation, Explanation, and Corroboration [Fet81h], page 173. ISBN 90-277-1336-7, 94-009-8558-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

                                              URL https://link.springer.com/
                                              book/10.1007/978-94-009-8558-2.

Anonymous:1982:BMc

[Ano82a] Anonymous. Back matter. In Scientific Philosophy Today: Essays in Honor of Mario Bunge [AC82], pages 487–517. ISBN 90-277-1262-X, 90-277-1263-8, 94-009-8462-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 67. URL

https://link.springer.com/book/
10.1007/978-94-009-8462-2.

Anonymous:1982:BMa

[Ano82b] Anonymous. Back matter. In Polish Essays in the Philosophy of the Natural Sciences [Kra82c], pages 431– 491. ISBN 94-009-7705-0 (e-book), 90-277-1287-5. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
book/10.1007/978-94-009-7705-1.

Anonymous:1982:BMb

[Ano82c] Anonymous. Back matter. In Studies of Mind and Brain: Neural Principles of Learning, Perception, Development, Cognition, and Motor Control [Gro82m], pages 640–662. ISBN 94-009-7758-1 (e-book), 90-277-1360-X. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL

https://link.springer.com/book/
10.1007/978-94-009-7758-7.

Anonymous:1982:FMj

[Ano82d] Anonymous. Front matter. In Scientific Philosophy Today: Essays in Honor of Mario Bunge [AC82], pages i–x. ISBN 90-277-1262-X, 90-277-1263-8, 94-009-8462-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 67. URL

https://link.springer.com/book/
10.1007/978-94-009-8462-2.

Anonymous:1982:FMa

[Ano82e] Anonymous. Front matter. In Polish Essays in the Philosophy of the Natural Sciences [Kra82c], pages i–xxviii. ISBN 94-009-7705-0 (e-book), 90-277-1287-5. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL

                                              https://link.springer.com/book/
                                              10.1007/978-94-009-7705-1.

Anonymous:1982:FMb

[Ano82f] Anonymous. Front matter. In Polish Essays in the Philosophy of the Natural Sciences [Kra82c], page 1. ISBN 94-009-7705-0 (e-book), 90-277-1287-5. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL

                                              https://link.springer.com/book/
                                              10.1007/978-94-009-7705-1.

Anonymous:1982:FMc

[Ano82g] Anonymous. Front matter. In Polish Essays in the Philosophy of the Natural Sciences [Kra82c], page 167. ISBN 94-009-7705-0 (e-book), 90-277-1287-5. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL

                                              https://link.springer.com/book/
                                              10.1007/978-94-009-7705-1.

Anonymous:1982:FMd

[Ano82h] Anonymous. Front matter. In Polish Essays in the Philosophy of the Natural Sciences [Kra82c], page 209. ISBN 94-009-7705-0 (e-book), 90-277-1287-5. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL

                                              https://link.springer.com/book/
                                              10.1007/978-94-009-7705-1.

Anonymous:1982:FMe

[Ano82i] Anonymous. Front matter. In Polish Essays in the Philosophy of the Natural Sciences [Kra82c], page 237. ISBN 94-009-7705-0 (e-book), 90-277-1287-5. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL

                                              https://link.springer.com/book/
                                              10.1007/978-94-009-7705-1.

Anonymous:1982:FMf

[Ano82j] Anonymous. Front matter. In Polish Essays in the Philosophy of the Natural Sciences [Kra82c], page 275. ISBN 94-009-7705-0 (e-book), 90-277-1287-5. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL

https://link.springer.com/book/
10.1007/978-94-009-7705-1.

Anonymous:1982:FMg

[Ano82k] Anonymous. Front matter. In Polish Essays in the Philosophy of the Natural Sciences [Kra82c], page 333. ISBN 94-009-7705-0 (e-book), 90-277-1287-5. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL

https://link.springer.com/book/
10.1007/978-94-009-7705-1.

Anonymous:1982:FMh

[Ano82l] Anonymous. Front matter. In Polish Essays in the Philosophy of the Natural Sciences [Kra82c], page 375. ISBN 94-009-7705-0 (e-book), 90-277-1287-5. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL

https://link.springer.com/book/
10.1007/978-94-009-7705-1.

Anonymous:1982:FMi

[Ano82m] Anonymous. Front matter. In Studies of Mind and Brain: Neural Principles of Learning, Perception, Development, Cognition, and Motor Control [Gro82m], pages i–xvii. ISBN 94-009-7758-1 (e-book), 90-277-1360-X. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL

https://link.springer.com/book/
10.1007/978-94-009-7758-7.

Anonymous:1983:BMe

[Ano83a] Anonymous. Back matter. In Language, Logic and Method [CW83b], pages 459– 464. ISBN 94-009-7704-2, 94-009-7702-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL

https://link.springer.com/book/
10.1007/978-94-009-7702-0.

Anonymous:1983:BMh

[Ano83b] Anonymous. Back matter. In Epistemology, Methodology, and the Social Sciences [CW83a], pages 265–270. ISBN 90-481-8376-6, 94-017-1458-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
book/10.1007/978-94-017-1458-7.

Anonymous:1983:BMf

[Ano83c] Anonymous. Back matter. In Measurement: Its Concepts, Theories and Problems [Ber83d], pages 218–250. ISBN 94-009-7830-8, 94-009-7828-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
book/10.1007/978-94-009-7828-7.

Anonymous:1983:BMg

[Ano83d] Anonymous. Back matter. In The Structure and Growth of Scientific Knowledge: A Study in the Methodology of Epistemic Appraisal [Pan83f], pages 184–231. ISBN 90-481-8375-8, 94-015-7680-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 73. URL

https://link.springer.com/book/
10.1007/978-94-015-7680-2.

Anonymous:1983:BMa

[Ano83e] Anonymous. Back matter. In Formal Thought and the Sciences of Man [Gra83b], pages 194–210. ISBN 94-009-7039-0, 94-009-7037-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL

https://link.springer.com/book/
10.1007/978-94-009-7037-3.

Anonymous:1983:BMb

[Ano83f] Anonymous. Back matter. In Cohen and Laudan [CL83], pages 321–339. ISBN 94-009-7055-2, 94-009-7057-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
book/10.1007/978-94-009-7055-7.

Anonymous:1983:BMc

[Ano83g] Anonymous. Back matter. In Finalization in Science: The Social Orientation of Scientific Progress [Sch83b], pages 273–299. ISBN 94-009-7082-X, 94-009-7080-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL

https://link.springer.com/book/
10.1007/978-94-009-7080-9.

Anonymous:1983:BMd

[Ano83h] Anonymous. Back matter. In Philosophy and Technology [DR83], pages 337– 343. ISBN 94-009-7126-5, 94-009-7124-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL

https://link.springer.com/book/
10.1007/978-94-009-7124-0.

Anonymous:1983:FMo

[Ano83i] Anonymous. Front matter. In Language, Logic and Method [CW83b], pages i–viii. ISBN 94-009-7704-2, 94-009-7702-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic).

LCCN ????  URL https://link.
springer.com/book/10.1007/978-
94-009-7702-0.

Anonymous:1983:FMr

[Ano83j] Anonymous. Front matter. In Epistemology, Methodology, and the Social Sciences [CW83a], pages i–viii. ISBN 90-481-8376-6, 94-017-1458-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL

https://link.springer.com/book/
10.1007/978-94-017-1458-7.

Anonymous:1983:FMp

[Ano83k] Anonymous. Front matter. In Measurement: Its Concepts, Theories and Problems [Ber83d], pages i–xii. ISBN 94-009-7830-8, 94-009-7828-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
book/10.1007/978-94-009-7828-7.

Anonymous:1983:FMq

[Ano83l] Anonymous. Front matter. In The Structure and Growth of Scientific Knowledge: A Study in the Methodology of Epistemic Appraisal [Pan83f], pages i–xxii. ISBN 90-481-8375-8, 94-015-7680-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 73. URL

https://link.springer.com/book/
10.1007/978-94-015-7680-2.

Anonymous:1983:FMa

[Ano83m] Anonymous. Front matter. In Formal Thought and the Sciences of Man [Gra83b], pages i–xxviii. ISBN 94-009-7039-0, 94-009-7037-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL

https://link.springer.com/book/
10.1007/978-94-009-7037-3.

Anonymous:1983:FMb

[Ano83n] Anonymous. Front matter. In Cohen and Laudan [CL83], pages i–xviii. ISBN 94-009-7055-2, 94-009-7057-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
book/10.1007/978-94-009-7055-7.

Anonymous:1983:FMc

[Ano83o] Anonymous. Front matter. In Finalization in Science: The Social Orientation of Scientific Progress [Sch83b], pages i–xviii. ISBN 94-009-7082-X, 94-009-7080-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL

https://link.springer.com/book/
10.1007/978-94-009-7080-9.

Anonymous:1983:FMd

[Ano83p] Anonymous. Front matter. In Finalization in Science: The Social Orientation of Scientific Progress [Sch83b], page 1. ISBN 94-009-7082-X, 94-009-7080-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
book/10.1007/978-94-009-7080-9.

Anonymous:1983:FMe

[Ano83q] Anonymous. Front matter. In Finalization in Science: The Social Orientation of Scientific Progress [Sch83b], page 13. ISBN 94-009-7082-X, 94-009-7080-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
book/10.1007/978-94-009-7080-9.

Anonymous:1983:FMf

[Ano83r] Anonymous. Front matter. In Finalization in Science: The Social Orientation of Scientific Progress [Sch83b], page 127. ISBN 94-009-7082-X, 94-009-7080-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
book/10.1007/978-94-009-7080-9.

Anonymous:1983:FMg

[Ano83s] Anonymous. Front matter. In Finalization in Science: The Social Orientation of Scientific Progress [Sch83b], page 233. ISBN 94-009-7082-X, 94-009-7080-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

                                              URL https://link.springer.com/
                                              book/10.1007/978-94-009-7080-9.

Anonymous:1983:FMh

[Ano83t] Anonymous. Front matter. In Finalization in Science: The Social Orientation of Scientific Progress [Sch83b], page 271. ISBN 94-009-7082-X, 94-009-7080-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

                                              URL https://link.springer.com/
                                              book/10.1007/978-94-009-7080-9.

Anonymous:1983:FMi

[Ano83u] Anonymous. Front matter. In Philosophy and Technology [DR83], pages i– xiv. ISBN 94-009-7126-5, 94-009-7124-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

                                              URL https://link.springer.com/
                                              book/10.1007/978-94-009-7124-0.

Anonymous:1983:FMj

[Ano83v] Anonymous. Front matter. In Philosophy and Technology [DR83], page 15. ISBN 94-009-7126-5, 94-009-7124-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

                                              URL https://link.springer.com/
                                              book/10.1007/978-94-009-7124-0.

Anonymous:1983:FMk

[Ano83w] Anonymous. Front matter. In Philosophy and Technology [DR83], page 113. ISBN 94-009-7126-5, 94-009-7124-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

                                              URL https://link.springer.com/
                                              book/10.1007/978-94-009-7124-0.

Anonymous:1983:FMl

[Ano83x] Anonymous. Front matter. In Philosophy and Technology [DR83], page 185.

ISBN 94-009-7126-5, 94-009-7124-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
book/10.1007/978-94-009-7124-0.

Anonymous:1983:FMm

[Ano83y] Anonymous. Front matter. In Philosophy and Technology [DR83], page 233. ISBN 94-009-7126-5, 94-009-7124-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
book/10.1007/978-94-009-7124-0.

Anonymous:1983:FMn

[Ano83z] Anonymous. Front matter. In Philosophy and Technology [DR83], page 311. ISBN 94-009-7126-5, 94-009-7124-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
book/10.1007/978-94-009-7124-0.

Anonymous:1984:BMb

[Ano84a] Anonymous. Back matter. In Investigations into the Origin of Language and Consciousness [Tha84b], pages 199–214. ISBN 94-009-6238-X, 94-009-6236-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .T736 1984. URL https://link.

springer.com/book/10.1007/978-
94-009-6236-1.

Anonymous:1984:BMa

[Ano84b] Anonymous. Back matter. In Hegel and the Sciences [CW84a], pages 365–374. ISBN 94-009-6235-5, 94-009-6233-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
book/10.1007/978-94-009-6233-0.

Anonymous:1984:BMg

[Ano84c] Anonymous. Back matter. In Reason and the Search for Knowledge: Investigations in the Philosophy of Science [Sha84l], pages 419–441. ISBN 90-277-1641-2, 94-010-9731-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL

                                              https://link.springer.com/book/
                                              10.1007/978-94-010-9731-4.

Anonymous:1984:BMc

[Ano84d] Anonymous. Back matter. In Rationality in Science and Politics [And84d], pages 297–307. ISBN 90-277-1953-5, 94-009-6254-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic).

                                              LCCN ????  URL https://link.
                                              springer.com/book/10.1007/978-
                                              94-009-6254-5.

Anonymous:1984:BMd

[Ano84e] Anonymous. Back matter. In Dialectical Theory of Meaning [Mar84d], pages 396–399. ISBN 94-009-6258-4, 94-009-6256-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic).

                                              LCCN ????  URL https://link.
                                              springer.com/book/10.1007/978-
                                              94-009-6256-9.

Anonymous:1984:BMf

[Ano84f] Anonymous. Back matter. In Physical Sciences and History of Physics [CW84c], pages 255–259. ISBN 94-009-7180-X, 94-009-7178-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .P497 1984.

                                              URL https://link.springer.com/
                                              book/10.1007/978-94-009-7178-3.

Anonymous:1984:BMe

[Ano84g] Anonymous. Back matter. In Methodology, Metaphysics and the History of Science: In Memory of Benjamin Nelson [CW84b], pages 359–373. ISBN 94-009-6333-5, 94-009-6331-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL

https://link.springer.com/book/
10.1007/978-94-009-6331-3.

Anonymous:1984:FMf

[Ano84h] Anonymous. Front matter. In Investigations into the Origin of Language and Consciousness [Tha84b], pages i–xi. ISBN 94-009-6238-X, 94-009-6236-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174

.T736 1984. URL https://link.
springer.com/book/10.1007/978-
94-009-6236-1.

Anonymous:1984:FMg

[Ano84i] Anonymous. Front matter. In Investigations into the Origin of Language and Consciousness [Tha84b], page 1. ISBN 94-009-6238-X, 94-009-6236-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .T736 1984.

URL https://link.springer.com/
book/10.1007/978-94-009-6236-1.

Anonymous:1984:FMh

[Ano84j] Anonymous. Front matter. In Investigations into the Origin of Language and Consciousness [Tha84b], page 31. ISBN 94-009-6238-X, 94-009-6236-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .T736 1984.

URL https://link.springer.com/
book/10.1007/978-94-009-6236-1.

Anonymous:1984:FMi

[Ano84k] Anonymous. Front matter. In Investigations into the Origin of Language and Consciousness [Tha84b], page 143. ISBN 94-009-6238-X, 94-009-6236-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174

                                              .T736 1984. URL https://link.
                                              springer.com/book/10.1007/978-
                                              94-009-6236-1.

Anonymous:1984:FMa

[Ano84l] Anonymous. Front matter. In Hegel and the Sciences [CW84a], pages i–ix. ISBN 94-009-6235-5, 94-009-6233-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

                                              URL https://link.springer.com/
                                              book/10.1007/978-94-009-6233-0.

Anonymous:1984:FMb

[Ano84m] Anonymous. Front matter. In Hegel and the Sciences [CW84a], page 1. ISBN 94-009-6235-5, 94-009-6233-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

                                              URL https://link.springer.com/
                                              book/10.1007/978-94-009-6233-0.

Anonymous:1984:FMc

[Ano84n] Anonymous. Front matter. In Hegel and the Sciences [CW84a], page 11. ISBN 94-009-6235-5, 94-009-6233-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

                                              URL https://link.springer.com/
                                              book/10.1007/978-94-009-6233-0.

Anonymous:1984:FMd

[Ano84o] Anonymous. Front matter. In Hegel and the Sciences [CW84a], page 193. ISBN 94-009-6235-5, 94-009-6233-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

                                              URL https://link.springer.com/
                                              book/10.1007/978-94-009-6233-0.

Anonymous:1984:FMe

[Ano84p] Anonymous. Front matter. In Hegel and the Sciences [CW84a], page 301. ISBN 94-009-6235-5, 94-009-6233-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
book/10.1007/978-94-009-6233-0.

Anonymous:1984:FMq

[Ano84q] Anonymous. Front matter. In Reason and the Search for Knowledge: Investigations in the Philosophy of Science [Sha84l], pages i–xlv. ISBN 90-277-1641-2, 94-010-9731-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL

https://link.springer.com/book/
10.1007/978-94-010-9731-4.

Anonymous:1984:FMr

[Ano84r] Anonymous. Front matter. In Reason and the Search for Knowledge: Investigations in the Philosophy of Science [Sha84l], page 1. ISBN 90-277-1641-2, 94-010-9731-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL

https://link.springer.com/book/
10.1007/978-94-010-9731-4.

Anonymous:1984:FMs

[Ano84s] Anonymous. Front matter. In Reason and the Search for Knowledge: Investigations in the Philosophy of Science [Sha84l], page 121. ISBN 90-277-1641-2, 94-010-9731-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL

https://link.springer.com/book/
10.1007/978-94-010-9731-4.

Anonymous:1984:FMt

[Ano84t] Anonymous. Front matter. In Reason and the Search for Knowledge: Investigations in the Philosophy of Science [Sha84l], page 203. ISBN 90-277-1641-2, 94-010-9731-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL

https://link.springer.com/book/
10.1007/978-94-010-9731-4.

Anonymous:1984:FMj

[Ano84u] Anonymous. Front matter. In Rationality in Science and Politics [And84d], pages i–xi. ISBN 90-277-1953-5, 94-009-6254-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

                                              ???? URL https://link.springer.
                                              com/book/10.1007/978-94-009-
                                              6254-5.

Anonymous:1984:FMk

[Ano84v] Anonymous. Front matter. In Dialectical Theory of Meaning [Mar84d], pages i–xxviii. ISBN 94-009-6258-4, 94-009-6256-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic).

                                              LCCN ????  URL https://link.
                                              springer.com/book/10.1007/978-
                                              94-009-6256-9.

Anonymous:1984:FMl

[Ano84w] Anonymous. Front matter. In Dialectical Theory of Meaning [Mar84d], page 29. ISBN 94-009-6258-4, 94-009-6256-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

                                              ???? URL https://link.springer.
                                              com/book/10.1007/978-94-009-
                                              6256-9.

Anonymous:1984:FMm

[Ano84x] Anonymous. Front matter. In Dialectical Theory of Meaning [Mar84d], page 169. ISBN 94-009-6258-4, 94-009-6256-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

                                              URL https://link.springer.com/
                                              book/10.1007/978-94-009-6256-9.

Anonymous:1984:FMn

[Ano84y] Anonymous. Front matter. In Dialectical Theory of Meaning [Mar84d], page 329. ISBN 94-009-6258-4, 94-009-6256-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
book/10.1007/978-94-009-6256-9.

Anonymous:1984:FMp

[Ano84z] Anonymous. Front matter. In Physical Sciences and History of Physics [CW84c], pages i–ix. ISBN 94-009-7180-X, 94-009-7178-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .P497 1984. URL

https://link.springer.com/book/
10.1007/978-94-009-7178-3.

Anonymous:1984:FMo

[Ano84-27] Anonymous. Front matter. In Methodology, Metaphysics and the History of Science: In Memory of Benjamin Nelson [CW84b], pages i–x. ISBN 94-009-6333-5, 94-009-6331-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL

https://link.springer.com/book/
10.1007/978-94-009-6331-3.

Anonymous:1985:BMa

[Ano85a] Anonymous. Back matter. In Protophysics of Time: Constructive Foundation and History of Time Measurement [Jan85e], pages 214–241. ISBN 94-010-8794-6, 94-009-5189-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
book/10.1007/978-94-009-5189-1.

Anonymous:1985:BMb

[Ano85b] Anonymous. Back matter. In The Relativistic Deduction: Epistemological Implications of the Theory of Relativity [Mey85v], pages 252–268. ISBN 90-277-1699-4, 94-009-5211-2. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC173.585; Q174 .B67 vol. 83. URL

https://link.springer.com/book/
10.1007/978-94-009-5211-9. With a

review by Albert Einstein.

Anonymous:1985:BMc

[Ano85c] Anonymous. Back matter. In Foundations of Objective Knowledge: The Relations of Popper’s Theory of Knowledge to that of Kant [Fer85b], pages 252–272. ISBN 90-481-8402-9, 94-015-7704-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

                                              ???? URL https://link.springer.
                                              com/book/10.1007/978-94-015-
                                              7704-5.

Anonymous:1985:FMa

[Ano85d] Anonymous. Front matter. In Protophysics of Time: Constructive Foundation and History of Time Measurement [Jan85e], pages i–xxvi. ISBN 94-010-8794-6, 94-009-5189-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

                                              URL https://link.springer.com/
                                              book/10.1007/978-94-009-5189-1.

Anonymous:1985:FMb

[Ano85e] Anonymous. Front matter. In The Relativistic Deduction: Epistemological Implications of the Theory of Relativity [Mey85v], pages i–liii. ISBN 90-277-1699-4, 94-009-5211-2. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC173.585; Q174 .B67 vol. 83. URL

                                              https://link.springer.com/book/
                                              10.1007/978-94-009-5211-9. With a

review by Albert Einstein.

Anonymous:1985:FMc

[Ano85f] Anonymous. Front matter. In The Relativistic Deduction: Epistemological Implications of the Theory of Relativity [Mey85v], page 1. ISBN 90-277-1699-4, 94-009-5211-2. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC173.585; Q174 .B67 vol. 83. URL

                                              https://link.springer.com/book/

10.1007/978-94-009-5211-9. With a

review by Albert Einstein.

Anonymous:1985:FMd

[Ano85g] Anonymous. Front matter. In Foundations of Objective Knowledge: The Relations of Popper’s Theory of Knowledge to that of Kant [Fer85b], pages i–xvii. ISBN 90-481-8402-9, 94-015-7704-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/book/10.1007/978-94-015-
7704-5.

Anonymous:1986:BMf

[Ano86a] Anonymous. Back matter. In Phenomenology and Dialectical Materialism [Coh86d], pages 219–244. ISBN 94-010-8795-4, 94-009-5191-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL

https://link.springer.com/book/
10.1007/978-94-009-5191-4.

Anonymous:1986:BMa

[Ano86b] Anonymous. Back matter. In Cognition and Fact: Materials on Ludwik Fleck [CS86], pages 443–468. ISBN 94-010-8504-8, 94-009-4498-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL

https://link.springer.com/book/
10.1007/978-94-009-4498-5.

Anonymous:1986:BMb

[Ano86c] Anonymous. Back matter. In Atom and Individual in the Age of Newton: On the Genesis of the Mechanistic World View [Fre86f], pages 215–276. ISBN 94-010-8505-6, 94-009-4500-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .F748 1986.

URL https://link.springer.com/
book/10.1007/978-94-009-4500-5.

Anonymous:1986:BMg

[Ano86d] Anonymous. Back matter. In Cohen and Wartofsky [CW85], pages 318–327. ISBN 90-277-1971-3, 94-009-5345-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

                                              URL https://link.springer.com/
                                              book/10.1007/978-94-009-5345-1.

Anonymous:1986:BMh

[Ano86e] Anonymous. Back matter. In Cohen and Wartofsky [CW85], pages 375–381. ISBN 90-277-1971-3, 94-009-5345-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

                                              URL https://link.springer.com/
                                              book/10.1007/978-94-009-5345-1.

Anonymous:1986:BMc

[Ano86f] Anonymous. Back matter. In Philosophy and Technology II: Information Technology and Computers in Theory and Practice [MH86], pages 307–355. ISBN 94-010-8510-2, 94-009-4512-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

                                              URL https://link.springer.com/
                                              book/10.1007/978-94-009-4512-8.

Anonymous:1986:BMd

[Ano86g] Anonymous. Back matter. In Spinoza and the Sciences [GN86], pages 315– 339. ISBN 94-010-8511-0, 94-009-4514-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL

                                              https://link.springer.com/book/
                                              10.1007/978-94-009-4514-2.

Anonymous:1986:BMk

[Ano86h] Anonymous. Back matter. In The Search for a Methodology of Social Science: Durkheim, Weber, and the Nineteenth-Century Problem of Cause, Probability, and Action [Tur86l], pages 228–255. ISBN 90-481-8417-7, 94-017-3461-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/book/10.1007/978-94-017-
3461-5.

Anonymous:1986:BMi

[Ano86i] Anonymous. Back matter. In Thinking about Society: Theory and Practice [Jar86y], pages 476–523. ISBN 94-010-8894-2, 94-009-5424-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN HM545. URL

https://link.springer.com/book/
10.1007/978-94-009-5424-3.

Anonymous:1986:BMj

[Ano86j] Anonymous. Back matter. In Ullmann-Margalit [UM86a], pages 247–250. ISBN 90-277-2159-9 (paperback), 90-277-2158-0 (hardcover), 94-009-5496-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 94 Q175. URL https://

link.springer.com/book/10.1007/
978-94-009-5496-0.

Anonymous:1986:BMe

[Ano86k] Anonymous. Back matter. In Language and Production: A Critique of the Paradigms [Mar86f], pages 126–190. ISBN 94-010-8532-3, 94-009-4574-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
book/10.1007/978-94-009-4574-6.

Anonymous:1986:FMw

[Ano86l]  Anonymous.  Front matter.  In

Phenomenology and Dialectical Materialism [Coh86d], pages i–xxx. ISBN 94-010-8795-4, 94-009-5191-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL

                                              https://link.springer.com/book/
                                              10.1007/978-94-009-5191-4.

Anonymous:1986:FMx

[Ano86m] Anonymous.  Front matter.  In

Phenomenology and Dialectical Materialism [Coh86d], pages 1–2. ISBN 94-010-8795-4, 94-009-5191-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL

                                              https://link.springer.com/book/
                                              10.1007/978-94-009-5191-4.

Anonymous:1986:FMy

[Ano86n] Anonymous. Front matter. In Phenomenology and Dialectical Materialism [Coh86d], pages 131–132. ISBN 94-010-8795-4, 94-009-5191-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL

                                              https://link.springer.com/book/
                                              10.1007/978-94-009-5191-4.

Anonymous:1986:FMa

[Ano86o] Anonymous. Front matter. In Cognition and Fact: Materials on Ludwik Fleck [CS86], pages i–xxxiii. ISBN 94-010-8504-8, 94-009-4498-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL

                                              https://link.springer.com/book/
                                              10.1007/978-94-009-4498-5.

Anonymous:1986:FMb

[Ano86p] Anonymous. Front matter. In Cognition and Fact: Materials on Ludwik Fleck [CS86], page 1. ISBN 94-010-8504-8, 94-009-4498-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL

                                              https://link.springer.com/book/
                                              10.1007/978-94-009-4498-5.

Anonymous:1986:FMc

[Ano86q] Anonymous. Front matter. In Cognition and Fact: Materials on Ludwik Fleck [CS86], page 37. ISBN 94-010-8504-8, 94-009-4498-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL

https://link.springer.com/book/
10.1007/978-94-009-4498-5.

Anonymous:1986:FMd

[Ano86r] Anonymous. Front matter. In Cognition and Fact: Materials on Ludwik Fleck [CS86], page 159. ISBN 94-010-8504-8, 94-009-4498-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL

https://link.springer.com/book/
10.1007/978-94-009-4498-5.

Anonymous:1986:FMe

[Ano86s] Anonymous. Front matter. In Atom and Individual in the Age of Newton: On the Genesis of the Mechanistic World View [Fre86f], pages i–xv. ISBN 94-010-8505-6, 94-009-4500-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .F748 1986.

URL https://link.springer.com/
book/10.1007/978-94-009-4500-5.

Anonymous:1986:FMf

[Ano86t] Anonymous. Front matter. In Atom and Individual in the Age of Newton: On the Genesis of the Mechanistic World View [Fre86f], page 11. ISBN 94-010-8505-6, 94-009-4500-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .F748 1986.

URL https://link.springer.com/
book/10.1007/978-94-009-4500-5.

Anonymous:1986:FMg

[Ano86u] Anonymous. Front matter. In Atom and Individual in the Age of Newton: On the Genesis of the Mechanistic World View [Fre86f], pages 73–76. ISBN 94-010-8505-6, 94-009-4500-0 (e-book).

ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .F748 1986.

                                              URL https://link.springer.com/
                                              book/10.1007/978-94-009-4500-5.

Anonymous:1986:FMh

[Ano86v] Anonymous. Front matter. In Atom and Individual in the Age of Newton: On the Genesis of the Mechanistic World View [Fre86f], pages 105–107. ISBN 94-010-8505-6, 94-009-4500-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .F748 1986.

                                              URL https://link.springer.com/
                                              book/10.1007/978-94-009-4500-5.

Anonymous:1986:FMi

[Ano86w] Anonymous. Front matter. In Atom and Individual in the Age of Newton: On the Genesis of the Mechanistic World View [Fre86f], pages 169–172. ISBN 94-010-8505-6, 94-009-4500-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .F748 1986.

                                              URL https://link.springer.com/
                                              book/10.1007/978-94-009-4500-5.

Anonymous:1986:FMz

[Ano86x] Anonymous. Front matter. In Cohen and Wartofsky [CW85], pages i–viii. ISBN 90-277-1971-3, 94-009-5345-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

                                              URL https://link.springer.com/
                                              book/10.1007/978-94-009-5345-1.

Anonymous:1986:FMaa

[Ano86y] Anonymous. Front matter. In Cohen and Wartofsky [CW85], pages i–xvii. ISBN 90-277-1971-3, 94-009-5345-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

                                              URL https://link.springer.com/
                                              book/10.1007/978-94-009-5345-1.

Anonymous:1986:FMab

[Ano86z] Anonymous. Front matter. In Cohen and Wartofsky [CW85], page 1. ISBN 90-277-1971-3, 94-009-5345-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
book/10.1007/978-94-009-5345-1.

Anonymous:1986:FMac

[Ano86-27] Anonymous. Front matter. In Cohen and Wartofsky [CW85], page 105. ISBN 90-277-1971-3, 94-009-5345-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
book/10.1007/978-94-009-5345-1.

Anonymous:1986:FMad

[Ano86-28] Anonymous. Front matter. In Cohen and Wartofsky [CW85], page 199. ISBN 90-277-1971-3, 94-009-5345-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
book/10.1007/978-94-009-5345-1.

Anonymous:1986:FMae

[Ano86-29] Anonymous. Front matter. In Cohen and Wartofsky [CW85], page 293. ISBN 90-277-1971-3, 94-009-5345-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
book/10.1007/978-94-009-5345-1.

Anonymous:1986:FMaf

[Ano86-30] Anonymous. Front matter. In Cohen and Wartofsky [CW85], page 353. ISBN 90-277-1971-3, 94-009-5345-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
book/10.1007/978-94-009-5345-1.

Anonymous:1986:FMj

[Ano86-31] Anonymous. Front matter. In Philosophy and Technology II: Information Technology and Computers in Theory and Practice [MH86], pages i–xxii. ISBN 94-010-8510-2, 94-009-4512-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
book/10.1007/978-94-009-4512-8.

Anonymous:1986:FMk

[Ano86-32] Anonymous. Front matter. In Philosophy and Technology II: Information Technology and Computers in Theory and Practice [MH86], page 15. ISBN 94-010-8510-2, 94-009-4512-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
book/10.1007/978-94-009-4512-8.

Anonymous:1986:FMl

[Ano86-33] Anonymous. Front matter. In Philosophy and Technology II: Information Technology and Computers in Theory and Practice [MH86], page 83. ISBN 94-010-8510-2, 94-009-4512-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
book/10.1007/978-94-009-4512-8.

Anonymous:1986:FMm

[Ano86-34] Anonymous. Front matter. In Philosophy and Technology II: Information Technology and Computers in Theory and Practice [MH86], page 187. ISBN 94-010-8510-2, 94-009-4512-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
book/10.1007/978-94-009-4512-8.

Anonymous:1986:FMn

[Ano86-35] Anonymous. Front matter. In Spinoza and the Sciences [GN86], pages i–xix. ISBN 94-010-8511-0, 94-009-4514-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
book/10.1007/978-94-009-4514-2.

Anonymous:1986:FMo

[Ano86-36] Anonymous. Front matter. In Spinoza and the Sciences [GN86], page 1. ISBN 94-010-8511-0, 94-009-4514-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
book/10.1007/978-94-009-4514-2.

Anonymous:1986:FMp

[Ano86-37] Anonymous. Front matter. In Spinoza and the Sciences [GN86], page 93. ISBN 94-010-8511-0, 94-009-4514-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
book/10.1007/978-94-009-4514-2.

Anonymous:1986:FMq

[Ano86-38] Anonymous. Front matter. In Spinoza and the Sciences [GN86], page 151. ISBN 94-010-8511-0, 94-009-4514-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
book/10.1007/978-94-009-4514-2.

Anonymous:1986:FMr

[Ano86-39] Anonymous. Front matter. In Spinoza and the Sciences [GN86], page 189. ISBN 94-010-8511-0, 94-009-4514-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
book/10.1007/978-94-009-4514-2.

Anonymous:1986:FMs

[Ano86-40] Anonymous. Front matter. In Spinoza and the Sciences [GN86], page 235. ISBN 94-010-8511-0, 94-009-4514-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
book/10.1007/978-94-009-4514-2.

Anonymous:1986:FMt

[Ano86-41] Anonymous. Front matter. In Spinoza and the Sciences [GN86], page 303. ISBN 94-010-8511-0, 94-009-4514-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
book/10.1007/978-94-009-4514-2.

Anonymous:1986:FMak

[Ano86-42] Anonymous. Front matter. In The Search for a Methodology of Social Science: Durkheim, Weber, and the Nineteenth-Century Problem of Cause, Probability, and Action [Tur86l], pages i–xv. ISBN 90-481-8417-7, 94-017-3461-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
book/10.1007/978-94-017-3461-5.

Anonymous:1986:FMal

[Ano86-43] Anonymous. Front matter. In The Search for a Methodology of Social Science: Durkheim, Weber, and the Nineteenth-Century Problem of Cause, Probability, and Action [Tur86l], page 1. ISBN 90-481-8417-7, 94-017-3461-5 (ebook). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
book/10.1007/978-94-017-3461-5.

Anonymous:1986:FMam

[Ano86-44] Anonymous. Front matter. In The Search for a Methodology of Social Science: Durkheim, Weber, and the Nineteenth-Century Problem of Cause, Probability, and Action [Tur86l], page 105. ISBN 90-481-8417-7, 94-017-3461-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
book/10.1007/978-94-017-3461-5.

Anonymous:1986:FMan

[Ano86-45] Anonymous. Front matter. In The Search for a Methodology of Social Science: Durkheim, Weber, and the Nineteenth-Century Problem of Cause, Probability, and Action [Tur86l], page 161. ISBN 90-481-8417-7, 94-017-3461-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
book/10.1007/978-94-017-3461-5.

Anonymous:1986:FMag

[Ano86-46] Anonymous. Front matter. In Thinking about Society: Theory and Practice [Jar86y], pages i–xix. ISBN 94-010-8894-2, 94-009-5424-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN HM545. URL

https://link.springer.com/book/
10.1007/978-94-009-5424-3.

Anonymous:1986:FMah

[Ano86-47] Anonymous.  Front matter.  In

Thinking about Society: Theory and Practice [Jar86y], page 1. ISBN 94-010-8894-2, 94-009-5424-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN HM545. URL

https://link.springer.com/book/
10.1007/978-94-009-5424-3.

Anonymous:1986:FMai

[Ano86-48] Anonymous.  Front matter.  In

Thinking about Society: Theory and Practice [Jar86y], page 263. ISBN 94-010-8894-2, 94-009-5424-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN HM545. URL

                                              https://link.springer.com/book/
                                              10.1007/978-94-009-5424-3.

Anonymous:1986:FMaj

[Ano86-49] Anonymous.  Front matter.  In

Ullmann-Margalit [UM86a], pages i–xii. ISBN 90-277-2159-9 (paperback), 90-277-2158-0 (hardcover), 94-009-5496-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 94 Q175. URL

                                              https://link.springer.com/book/
                                              10.1007/978-94-009-5496-0.

Anonymous:1986:FMu

[Ano86-50] Anonymous. Front matter. In Ullmann-Margalit [UM86c], pages i–ix. ISBN 90-277-2160-2, 90-277-2161-0 (paperback), 94-009-4566-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 95 Q175.

                                              URL  http://www.springerlink.
                                              com/content/978-94-009-4566-1;
                                              https://link.springer.com/book/
                                              10.1007/978-94-009-4566-1.

Anonymous:1986:FMv

[Ano86-51] Anonymous. Front matter. In Language and Production: A Critique of the Paradigms [Mar86f], pages i–xv. ISBN 94-010-8532-3, 94-009-4574-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

                                              URL https://link.springer.com/
                                              book/10.1007/978-94-009-4574-6.

Anonymous:1987:BMh

[Ano87a] Anonymous. Back matter. In Reason and Being [FC87], pages 427–444. ISBN 90-277-2181-5, 94-009-4590-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 17.

URL https://link.springer.com/
book/10.1007/978-94-009-4590-6.

Anonymous:1987:BMb

[Ano87b] Anonymous. Back matter. In Amrine et al. [AZW87], pages 439–445. ISBN 90-277-2265-X, 94-009-3761-X. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol.

97. US$79.00. URL https://link.

springer.com/book/10.1007/978-
94-009-3761-1.

Anonymous:1987:BMc

[Ano87c] Anonymous. Back matter. In Pitt and Pera [PP87], pages 221–227. ISBN 90-277-2417-2, 94-009-3779-2. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 98. US$64.00.

URL https://link.springer.com/
book/10.1007/978-94-009-3779-6.

Anonymous:1987:BMd

[Ano87d] Anonymous. Back matter. In Time, The Physical Magnitude [Cos87b], pages 299–340. ISBN 94-010-8195-6, 94-009-3811-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic).

LCCN ????  URL https://link.
springer.com/book/10.1007/978-
94-009-3811-3.

Anonymous:1987:BMa

[Ano87e] Anonymous. Back matter. In Shimony and Nails [SN87], pages 379–388. ISBN 94-010-8168-9, 94-009-3735-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
book/10.1007/978-94-009-3735-2.

Anonymous:1987:BMe

[Ano87f] Anonymous. Back matter. In Time and Meaning in History [Rot87j], pages 207–220. ISBN 94-010-8212-X, 94-009-3845-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic).

LCCN ????  URL https://link.
springer.com/book/10.1007/978-
94-009-3845-8.

Anonymous:1987:BMf

[Ano87g] Anonymous. Back matter. In Glick [Gli87a], pages 401–416. ISBN 90-277-2498-9, 94-009-3875-6. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 103; QC173.5. US$79.00 (US). URL

https://link.springer.com/book/
10.1007/978-94-009-3875-5.

Anonymous:1987:BMg

[Ano87h] Anonymous. Back matter. In Burwick [Bur87b], pages 275–290. ISBN 90-277-2541-1, 94-009-3917-5. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 105. US$64.00(U.S.).

URL https://link.springer.com/
book/10.1007/978-94-009-3917-2.

Anonymous:1987:FMu

[Ano87i] Anonymous. Front matter. In Reason and Being [FC87], pages i–xx. ISBN 90-277-2181-5, 94-009-4590-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 17.

URL https://link.springer.com/
book/10.1007/978-94-009-4590-6.

Anonymous:1987:FMv

[Ano87j] Anonymous. Front matter. In Reason and Being [FC87], page 1. ISBN 90-277-2181-5, 94-009-4590-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 17.

URL https://link.springer.com/
book/10.1007/978-94-009-4590-6.

Anonymous:1987:FMw

[Ano87k] Anonymous. Front matter. In Reason and Being [FC87], page 29. ISBN 90-277-2181-5, 94-009-4590-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 17.

URL https://link.springer.com/
book/10.1007/978-94-009-4590-6.

Anonymous:1987:FMx

[Ano87l] Anonymous. Front matter. In Reason and Being [FC87], page 277. ISBN 90-277-2181-5, 94-009-4590-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 17.

URL https://link.springer.com/
book/10.1007/978-94-009-4590-6.

Anonymous:1987:FMe

[Ano87m] Anonymous. Front matter. In Amrine et al. [AZW87], pages i–xvii. ISBN 90-277-2265-X, 94-009-3761-X. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 97. US$79.00.

URL https://link.springer.com/
book/10.1007/978-94-009-3761-1.

Anonymous:1987:FMf

[Ano87n] Anonymous. Front matter. In Amrine et al. [AZW87], page 1. ISBN 90-277-2265-X, 94-009-3761-X. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 97. US$79.00. URL

https://link.springer.com/book/
10.1007/978-94-009-3761-1.

Anonymous:1987:FMg

[Ano87o] Anonymous. Front matter. In Amrine et al. [AZW87], page 113. ISBN 90-277-2265-X, 94-009-3761-X. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 97. US$79.00. URL

https://link.springer.com/book/
10.1007/978-94-009-3761-1.

Anonymous:1987:FMh

[Ano87p] Anonymous. Front matter. In Amrine et al. [AZW87], page 255. ISBN 90-277-2265-X, 94-009-3761-X. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 97. US$79.00. URL

                                              https://link.springer.com/book/
                                              10.1007/978-94-009-3761-1.

Anonymous:1987:FMi

[Ano87q] Anonymous. Front matter. In Pitt and Pera [PP87], pages i–xiii. ISBN 90-277-2417-2, 94-009-3779-2. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 98. US$64.00. URL

                                              https://link.springer.com/book/
                                              10.1007/978-94-009-3779-6.

Anonymous:1987:FMj

[Ano87r] Anonymous. Front matter. In Pitt and Pera [PP87], page 1. ISBN 90-277-2417-2, 94-009-3779-2. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 98. US$64.00. URL

                                              https://link.springer.com/book/
                                              10.1007/978-94-009-3779-6.

Anonymous:1987:FMk

[Ano87s] Anonymous. Front matter. In Pitt and Pera [PP87], page 133. ISBN 90-277-2417-2, 94-009-3779-2. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 98. US$64.00. URL

                                              https://link.springer.com/book/
                                              10.1007/978-94-009-3779-6.

Anonymous:1987:FMl

[Ano87t] Anonymous. Front matter. In Time, The Physical Magnitude [Cos87b], pages i–xxiii. ISBN 94-010-8195-6, 94-009-3811-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic).

                                              LCCN ????  URL https://link.
                                              springer.com/book/10.1007/978-
                                              94-009-3811-3.

Anonymous:1987:FMm

[Ano87u] Anonymous. Front matter. In Time, The Physical Magnitude [Cos87b], page 1. ISBN 94-010-8195-6, 94-009-3811-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/book/10.1007/978-94-009-
3811-3.

Anonymous:1987:FMn

[Ano87v] Anonymous. Front matter. In Time, The Physical Magnitude [Cos87b], page 11. ISBN 94-010-8195-6, 94-009-3811-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/book/10.1007/978-94-009-
3811-3.

Anonymous:1987:FMo

[Ano87w] Anonymous. Front matter. In Time, The Physical Magnitude [Cos87b], page 95. ISBN 94-010-8195-6, 94-009-3811-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/book/10.1007/978-94-009-
3811-3.

Anonymous:1987:FMp

[Ano87x] Anonymous. Front matter. In Time, The Physical Magnitude [Cos87b], page 165. ISBN 94-010-8195-6, 94-009-3811-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
book/10.1007/978-94-009-3811-3.

Anonymous:1987:FMq

[Ano87y] Anonymous. Front matter. In Time, The Physical Magnitude [Cos87b], page 285. ISBN 94-010-8195-6, 94-009-3811-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
book/10.1007/978-94-009-3811-3.

Anonymous:1987:FMa

[Ano87z] Anonymous. Front matter. In Shimony and Nails [SN87], pages i–vi. ISBN 94-010-8168-9, 94-009-3735-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

                                              URL https://link.springer.com/
                                              book/10.1007/978-94-009-3735-2.

Anonymous:1987:FMb

[Ano87-27] Anonymous. Front matter. In Shimony and Nails [SN87], page 15. ISBN 94-010-8168-9, 94-009-3735-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL

                                              https://link.springer.com/book/
                                              10.1007/978-94-009-3735-2.

Anonymous:1987:FMc

[Ano87-28] Anonymous. Front matter. In Shimony and Nails [SN87], page 163. ISBN 94-010-8168-9, 94-009-3735-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

                                              URL https://link.springer.com/
                                              book/10.1007/978-94-009-3735-2.

Anonymous:1987:FMd

[Ano87-29] Anonymous. Front matter. In Shimony and Nails [SN87], page 319. ISBN 94-010-8168-9, 94-009-3735-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

                                              URL https://link.springer.com/
                                              book/10.1007/978-94-009-3735-2.

Anonymous:1987:FMr

[Ano87-30] Anonymous. Front matter. In Time and Meaning in History [Rot87j], pages i–ix. ISBN 94-010-8212-X, 94-009-3845-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

                                              URL https://link.springer.com/
                                              book/10.1007/978-94-009-3845-8.

Anonymous:1987:FMs

[Ano87-31] Anonymous. Front matter. In Glick [Gli87a], pages i–vii. ISBN 90-277-2498-9, 94-009-3875-6. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 103; QC173.5. US$79.00

(US). URL https://link.springer.
com/book/10.1007/978-94-009-
3875-5.

Anonymous:1987:FMt

[Ano87-32] Anonymous. Front matter. In Burwick [Bur87b], pages i–xix. ISBN 90-277-2541-1, 94-009-3917-5. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 105. US$64.00(U.S.).

URL https://link.springer.com/
book/10.1007/978-94-009-3917-2.

Anonymous:1988:BMa

[Ano88a] Anonymous. Back matter. In The Birth of Meaning in Hindu Thought [Coh88b], pages 350–368. ISBN 94-010-7141-1, 94-009-1431-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN B131 .Z55 1988; Q174 .B67 vol. 102 B131. URL

https://link.springer.com/book/
10.1007/978-94-009-1431-5.

Anonymous:1988:BMb

[Ano88b] Anonymous. Back matter. In Philosophy, History and Social Action: Essays in Honor of Lewis Feuer with an autobiographic essay by Lewis Feuer [HOO88c], pages 475–496. ISBN 94-010-7793-2, 94-009-2873-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL

https://link.springer.com/book/
10.1007/978-94-009-2873-2.

Anonymous:1988:BMc

[Ano88c] Anonymous. Back matter. In Science in Reflection: The Israel Colloquium:

Studies in History, Philosophy, and Sociology of Science Volume 3 [UM88], pages 235–237. ISBN 90-277-2713-9, 94-009-2957-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic).

LCCN ????  URL https://link.
springer.com/book/10.1007/978-
94-009-2957-9.

Anonymous:1988:FMa

[Ano88d] Anonymous. Front matter. In The Birth of Meaning in Hindu Thought [Coh88b], pages i–xxii. ISBN 94-010-7141-1, 94-009-1431-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN B131 .Z55 1988; Q174 .B67 vol. 102 B131. URL

https://link.springer.com/book/
10.1007/978-94-009-1431-5.

Anonymous:1988:FMb

[Ano88e] Anonymous. Front matter. In Philosophy, History and Social Action: Essays in Honor of Lewis Feuer with an autobiographic essay by Lewis Feuer [HOO88c], pages i–ix. ISBN 94-010-7793-2, 94-009-2873-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL

https://link.springer.com/book/
10.1007/978-94-009-2873-2.

Anonymous:1988:FMc

[Ano88f] Anonymous. Front matter. In Science in Reflection: The Israel Colloquium: Studies in History, Philosophy, and Sociology of Science Volume 3 [UM88], pages i–ix. ISBN 90-277-2713-9, 94-009-2957-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/book/10.1007/978-94-009-
2957-9.

Anonymous:1989:BMg

[Ano89a] Anonymous. Back matter. In The Form of Information in Science: Analysis of an Immunology Sublanguage [HGR^{+}89], pages 217–593. ISBN 94-010-7777-0, 94-009-2837-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN B832.A-832.Z.

URL https://link.springer.com/
book/10.1007/978-94-009-2837-4.

Anonymous:1989:BMh

[Ano89b] Anonymous. Back matter. In Imre Lakatos and Theories of Scientific Change [GGN89], pages 457–465. ISBN 94-010-7860-2, 94-009-3025-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
book/10.1007/978-94-009-3025-4.

Anonymous:1989:BMi

[Ano89c] Anonymous. Back matter. In The Qualitative-Quantitative Distinction in the Social Sciences [GM89b], pages 207–236. ISBN 90-481-8460-6, 94-017-3444-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 112. URL

https://link.springer.com/book/
10.1007/978-94-017-3444-8.

Anonymous:1989:BMf

[Ano89d] Anonymous. Back matter. In Structures of Knowing: Psychologies of the Nineteenth Century [Are89j], pages 343–430. ISBN 94-010-7692-8, 94-009-2641-3 (ebook). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
book/10.1007/978-94-009-2641-7.

Anonymous:1989:BMa

[Ano89e] Anonymous. Back matter. In Style, Politics and the Future of Philosophy [Jan89l], pages 264–278. ISBN 94-010-7508-5, 94-009-2251-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL

https://link.springer.com/book/
10.1007/978-94-009-2251-8.

Anonymous:1989:BMb

[Ano89f] Anonymous. Back matter. In Literature and Science as Modes of Expression [Amr89], pages 189–201. ISBN 94-010-7531-X, 94-009-2297-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
book/10.1007/978-94-009-2297-6.

Anonymous:1989:BMc

[Ano89g] Anonymous. Back matter. In An Intimate Relation: Studies in the History and Philosophy of Science Presented to Robert E. Butts on his 60th Birthday [BM89], pages 513–523. ISBN 94-010-7546-8, 94-009-2327-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
book/10.1007/978-94-009-2327-0.

Anonymous:1989:BMd

[Ano89h] Anonymous. Back matter. In Freedom and Rationality: Essays in Honor of John Watkins From his Colleagues and Friends [DJ89], pages 369–388. ISBN 94-010-7571-9, 94-009-2380-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
book/10.1007/978-94-009-2380-5.

Anonymous:1989:BMe

[Ano89i] Anonymous. Back matter. In Reflexive Epistemology: The Philosophical Legacy of Otto Neurath [Zol89g], pages 185– 203. ISBN 94-010-7588-3, 94-009-2415-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL

https://link.springer.com/book/
10.1007/978-94-009-2415-4.

Anonymous:1989:FMl

[Ano89j] Anonymous. Front matter. In The Form of Information in Science: Analysis of an Immunology Sublanguage [HGR^{+}89], pages i–xvii. ISBN 94-010-7777-0, 94-009-2837-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN B832.A-832.Z.

URL https://link.springer.com/
book/10.1007/978-94-009-2837-4.

Anonymous:1989:FMm

[Ano89k] Anonymous. Front matter. In Imre Lakatos and Theories of Scientific Change [GGN89], pages I–XI. ISBN 94-010-7860-2, 94-009-3025-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL

https://link.springer.com/book/
10.1007/978-94-009-3025-4.

Anonymous:1989:FMn

[Ano89l] Anonymous. Front matter. In Imre Lakatos and Theories of Scientific Change [GGN89], page 1. ISBN 94-010-7860-2, 94-009-3025-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL

https://link.springer.com/book/
10.1007/978-94-009-3025-4.

Anonymous:1989:FMo

[Ano89m] Anonymous. Front matter. In Imre Lakatos and Theories of Scientific Change [GGN89], page 81. ISBN 94-010-7860-2, 94-009-3025-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL

https://link.springer.com/book/
10.1007/978-94-009-3025-4.

Anonymous:1989:FMp

[Ano89n] Anonymous. Front matter. In Imre Lakatos and Theories of Scientific Change [GGN89], page 155. ISBN 94-010-7860-2, 94-009-3025-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL

                                              https://link.springer.com/book/
                                              10.1007/978-94-009-3025-4.

Anonymous:1989:FMq

[Ano89o] Anonymous. Front matter. In Imre Lakatos and Theories of Scientific Change [GGN89], page 227. ISBN 94-010-7860-2, 94-009-3025-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL

                                              https://link.springer.com/book/
                                              10.1007/978-94-009-3025-4.

Anonymous:1989:FMr

[Ano89p] Anonymous. Front matter. In Imre Lakatos and Theories of Scientific Change [GGN89], page 335. ISBN 94-010-7860-2, 94-009-3025-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL

                                              https://link.springer.com/book/
                                              10.1007/978-94-009-3025-4.

Anonymous:1989:FMs

[Ano89q] Anonymous. Front matter. In Imre Lakatos and Theories of Scientific Change [GGN89], page 397. ISBN 94-010-7860-2, 94-009-3025-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL

                                              https://link.springer.com/book/
                                              10.1007/978-94-009-3025-4.

Anonymous:1989:FMt

[Ano89r] Anonymous. Front matter. In The Qualitative-Quantitative Distinction in the Social Sciences [GM89b], pages i–vii. ISBN 90-481-8460-6, 94-017-3444-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 112. URL

https://link.springer.com/book/
10.1007/978-94-017-3444-8.

Anonymous:1989:FMh

[Ano89s] Anonymous. Front matter. In Structures of Knowing: Psychologies of the Nineteenth Century [Are89j], pages i–x. ISBN 94-010-7692-8, 94-009-2641-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
book/10.1007/978-94-009-2641-7.

Anonymous:1989:FMi

[Ano89t] Anonymous. Front matter. In Structures of Knowing: Psychologies of the Nineteenth Century [Are89j], page 1. ISBN 94-010-7692-8, 94-009-2641-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
book/10.1007/978-94-009-2641-7.

Anonymous:1989:FMj

[Ano89u] Anonymous. Front matter. In Structures of Knowing: Psychologies of the Nineteenth Century [Are89j], page 57. ISBN 94-010-7692-8, 94-009-2641-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
book/10.1007/978-94-009-2641-7.

Anonymous:1989:FMk

[Ano89v] Anonymous. Front matter. In Structures of Knowing: Psychologies of the Nineteenth Century [Are89j], page 151. ISBN 94-010-7692-8, 94-009-2641-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
book/10.1007/978-94-009-2641-7.

Anonymous:1989:FMa

[Ano89w] Anonymous. Front matter. In Style, Politics and the Future of Philosophy [Jan89l], pages i–xvi. ISBN 94-010-7508-5, 94-009-2251-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL

                                              https://link.springer.com/book/
                                              10.1007/978-94-009-2251-8.

Anonymous:1989:FMb

[Ano89x] Anonymous. Front matter. In Literature and Science as Modes of Expression [Amr89], pages i–xxv. ISBN 94-010-7531-X, 94-009-2297-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

                                              URL https://link.springer.com/
                                              book/10.1007/978-94-009-2297-6.

Anonymous:1989:FMc

[Ano89y] Anonymous. Front matter. In An Intimate Relation: Studies in the History and Philosophy of Science Presented to Robert E. Butts on his 60th Birthday [BM89], pages i–xx. ISBN 94-010-7546-8, 94-009-2327-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

                                              URL https://link.springer.com/
                                              book/10.1007/978-94-009-2327-0.

Anonymous:1989:FMd

[Ano89z] Anonymous. Front matter. In Freedom and Rationality: Essays in Honor of John Watkins From his Colleagues and Friends [DJ89], pages i–xi. ISBN 94-010-7571-9, 94-009-2380-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL

                                              https://link.springer.com/book/
                                              10.1007/978-94-009-2380-5.

Anonymous:1989:FMe

[Ano89-27] Anonymous. Front matter. In Freedom and Rationality: Essays in Honor of John Watkins From his Colleagues and Friends [DJ89], page 1. ISBN 94-010-7571-9, 94-009-2380-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL

https://link.springer.com/book/
10.1007/978-94-009-2380-5.

Anonymous:1989:FMf

[Ano89-28] Anonymous. Front matter. In Freedom and Rationality: Essays in Honor of John Watkins From his Colleagues and Friends [DJ89], page 183. ISBN 94-010-7571-9, 94-009-2380-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL

https://link.springer.com/book/
10.1007/978-94-009-2380-5.

Anonymous:1989:FMg

[Ano89-29] Anonymous. Front matter. In Reflexive Epistemology: The Philosophical Legacy of Otto Neurath [Zol89g], pages i–xx. ISBN 94-010-7588-3, 94-009-2415-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
book/10.1007/978-94-009-2415-4.

Anonymous:1990:BMa

[Ano90a] Anonymous. Back matter. In Georg Simmel and Contemporary Sociology [KPC90], pages 385–388. ISBN 94-010-6691-4 (hardcover), 94-009-0459-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN HM401-1281.

URL https://link.springer.com/
book/10.1007/978-94-009-0459-0.

Anonymous:1990:BMc

[Ano90b] Anonymous. Back matter. In Levere and Shea [LS90], pages 339–355. ISBN 0-7923-0420-9, 94-009-1878-X. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 120.

URL https://link.springer.com/
book/10.1007/978-94-009-1878-8.

Anonymous:1990:BMd

[Ano90c] Anonymous. Back matter. In Greek Studies in the Philosophy and History of Science [Nic90], pages 415–439. ISBN 94-010-7403-8, 94-009-2015-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
book/10.1007/978-94-009-2015-6.

Anonymous:1990:BMb

[Ano90d] Anonymous. Back matter. In Statistics in Science: The Foundations of Statistical Methods in Biology, Physics and Economics [CC90], pages 183–184. ISBN 94-010-6765-1, 94-009-0619-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QA276-280.

URL https://link.springer.com/
book/10.1007/978-94-009-0619-8.

Anonymous:1990:FMa

[Ano90e] Anonymous. Front matter. In Georg Simmel and Contemporary Sociology [KPC90], pages i–vi. ISBN 94-010-6691-4 (hardcover), 94-009-0459-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN HM401-1281. URL

https://link.springer.com/book/
10.1007/978-94-009-0459-0.

Anonymous:1990:FMc

[Ano90f] Anonymous. Front matter. In Levere and Shea [LS90], pages i–xiv. ISBN 0-7923-0420-9, 94-009-1878-X. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 120.

URL https://link.springer.com/
book/10.1007/978-94-009-1878-8.

Anonymous:1990:FMd

[Ano90g] Anonymous. Front matter. In Levere and Shea [LS90], page 1. ISBN 0-7923-0420-9, 94-009-1878-X. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 120.

URL https://link.springer.com/
book/10.1007/978-94-009-1878-8.

Anonymous:1990:FMe

[Ano90h] Anonymous. Front matter. In Levere and Shea [LS90], page 145. ISBN 0-7923-0420-9, 94-009-1878-X. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 120.

URL https://link.springer.com/
book/10.1007/978-94-009-1878-8.

Anonymous:1990:FMf

[Ano90i] Anonymous. Front matter. In Levere and Shea [LS90], page 197. ISBN 0-7923-0420-9, 94-009-1878-X. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 120.

URL https://link.springer.com/
book/10.1007/978-94-009-1878-8.

Anonymous:1990:FMg

[Ano90j] Anonymous. Front matter. In Levere and Shea [LS90], page 319. ISBN 0-7923-0420-9, 94-009-1878-X. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 120.

URL https://link.springer.com/
book/10.1007/978-94-009-1878-8.

Anonymous:1990:FMh

[Ano90k] Anonymous. Front matter. In Greek Studies in the Philosophy and History of Science [Nic90], pages i–xvii. ISBN 94-010-7403-8, 94-009-2015-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
book/10.1007/978-94-009-2015-6.

Anonymous:1990:FMb

[Ano90l] Anonymous. Front matter. In Statistics in Science: The Foundations of Statistical Methods in Biology, Physics and Economics [CC90], pages i–ix. ISBN 94-010-6765-1, 94-009-0619-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QA276-280. URL

                                              https://link.springer.com/book/
                                              10.1007/978-94-009-0619-8.

Anonymous:1991:BMk

[Ano91a] Anonymous. Back matter. In The Origins of Statics: The Sources of Physical Theory [Duh91r], pages 449–596. ISBN 94-010-5658-7, 94-011-3730-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 123.

                                              URL http://d-nb.info/101431545X/
                                              34;   https://link.springer.com/
                                              book/10.1007/978-94-011-3730-0.

Anonymous:1991:BMa

[Ano91b] Anonymous. Back matter. In Gavroglu and Goudaroulis [GG91], pages 491– 575. ISBN 94-010-7433-X (hardcover), 94-009-2079-2. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 124. URL https://link.

                                              springer.com/book/10.1007/978-
                                              94-009-2079-8.

Anonymous:1991:BMb

[Ano91c] Anonymous. Back matter. In The New Aspects of Time: Its Continuity and Novelties [Čap91g], pages 345–353. ISBN 94-010-7455-0, 94-009-2123-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

                                              URL https://link.springer.com/
                                              book/10.1007/978-94-009-2123-8.

Anonymous:1991:BMh

[Ano91d] Anonymous. Back matter. In Unguru [Ung91b], pages 309–327. ISBN 94-010-5476-2, 94-011-3342-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC5.53; QC6.9.

URL https://link.springer.com/
book/10.1007/978-94-011-3342-5.

Anonymous:1991:BMg

[Ano91e] Anonymous. Back matter. In Newton’s Physics and the Conceptual Structure of the Scientific Revolution [Bec91r], pages 513–588. ISBN 0-7923-1054-3, 94-011-3276-3. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 127. URL https://link.

springer.com/book/10.1007/978-
94-011-3276-3.

Anonymous:1991:BMj

[Ano91f] Anonymous. Back matter. In Explanation in the Sciences [Mey91d], pages 545–623. ISBN 94-010-5511-4, 94-011-3414-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic).

LCCN ????  URL https://link.
springer.com/book/10.1007/978-
94-011-3414-9.

Anonymous:1991:BMi

[Ano91g] Anonymous. Back matter. In Organism and the Origins of Self [Tau91i], pages 375–384. ISBN 0-7923-1185-X, 94-011-3406-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QR180-189.5. URL

https://link.springer.com/book/
10.1007/978-94-011-3406-4.

Anonymous:1991:BMd

[Ano91h] Anonymous. Back matter. In Methodological Variance: Essays in Epistemological Ontology and the Methodology of Science [Pan91f], pages 363–431. ISBN 94-010-5400-2, 94-011-3174-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

                                              URL https://link.springer.com/
                                              book/10.1007/978-94-011-3174-2.

Anonymous:1991:BMf

[Ano91i] Anonymous. Back matter. In Beyond Reason: Essays on the Philosophy of Paul Feyerabend [Mun91a], pages 529– 538. ISBN 94-010-5406-1, 94-011-3188-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL

                                              https://link.springer.com/book/
                                              10.1007/978-94-011-3188-9.

Anonymous:1991:BMe

[Ano91j] Anonymous. Back matter. In Uebel [Ueb91b], pages 293–331. ISBN 0-7923-1276-7 (hardcover), 94-011-3182-1. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 133.

                                              URL https://link.springer.com/
                                              book/10.1007/978-94-011-3182-7.

Anonymous:1991:BMc

[Ano91k] Anonymous. Back matter. In Woodward and Cohen [WC91], pages 431– 471. ISBN 94-010-5395-2, 94-011-3164-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL

                                              https://link.springer.com/book/
                                              10.1007/978-94-011-3164-3.

Anonymous:1991:FMac

[Ano91l] Anonymous. Front matter. In The Origins of Statics: The Sources of Physical Theory [Duh91r], pages i–9. ISBN 94-010-5658-7, 94-011-3730-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 123.

                                              URL http://d-nb.info/101431545X/
                                              34;   https://link.springer.com/
                                              book/10.1007/978-94-011-3730-0.

Anonymous:1991:FMa

[Ano91m] Anonymous. Front matter. In Gavroglu and Goudaroulis [GG91], pages i–cxv.

ISBN 94-010-7433-X (hardcover), 94-009-2079-2. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 124. URL https://link.

springer.com/book/10.1007/978-
94-009-2079-8.

Anonymous:1991:FMb

[Ano91n] Anonymous. Front matter. In The New Aspects of Time: Its Continuity and Novelties [Čap91g], pages i–xx. ISBN 94-010-7455-0, 94-009-2123-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
book/10.1007/978-94-009-2123-8.

Anonymous:1991:FMc

[Ano91o] Anonymous. Front matter. In The New Aspects of Time: Its Continuity and Novelties [Čap91g], page 1. ISBN 94-010-7455-0, 94-009-2123-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
book/10.1007/978-94-009-2123-8.

Anonymous:1991:FMd

[Ano91p] Anonymous. Front matter. In The New Aspects of Time: Its Continuity and Novelties [Čap91g], page 101. ISBN 94-010-7455-0, 94-009-2123-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
book/10.1007/978-94-009-2123-8.

Anonymous:1991:FMe

[Ano91q] Anonymous. Front matter. In The New Aspects of Time: Its Continuity and Novelties [Čap91g], page 219. ISBN 94-010-7455-0, 94-009-2123-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
book/10.1007/978-94-009-2123-8.

Anonymous:1991:FMq

[Ano91r] Anonymous. Front matter. In Unguru [Ung91b], pages i–viii. ISBN 94-010-5476-2, 94-011-3342-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC5.53; QC6.9.

                                              URL https://link.springer.com/
                                              book/10.1007/978-94-011-3342-5.

Anonymous:1991:FMr

[Ano91s] Anonymous. Front matter. In Unguru [Ung91b], page 1. ISBN 94-010-5476-2, 94-011-3342-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC5.53; QC6.9. URL

                                              https://link.springer.com/book/
                                              10.1007/978-94-011-3342-5.

Anonymous:1991:FMs

[Ano91t] Anonymous. Front matter. In Unguru [Ung91b], page 45. ISBN 94-010-5476-2, 94-011-3342-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC5.53; QC6.9. URL

                                              https://link.springer.com/book/
                                              10.1007/978-94-011-3342-5.

Anonymous:1991:FMt

[Ano91u] Anonymous. Front matter. In Unguru [Ung91b], page 99. ISBN 94-010-5476-2, 94-011-3342-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC5.53; QC6.9. URL

                                              https://link.springer.com/book/
                                              10.1007/978-94-011-3342-5.

Anonymous:1991:FMu

[Ano91v] Anonymous. Front matter. In Unguru [Ung91b], page 183. ISBN 94-010-5476-2, 94-011-3342-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC5.53; QC6.9. URL

                                              https://link.springer.com/book/
                                              10.1007/978-94-011-3342-5.

Anonymous:1991:FMv

[Ano91w] Anonymous. Front matter. In Unguru [Ung91b], page 251. ISBN 94-010-5476-2, 94-011-3342-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC5.53; QC6.9. URL

https://link.springer.com/book/
10.1007/978-94-011-3342-5.

Anonymous:1991:FMp

[Ano91x] Anonymous. Front matter. In Newton’s Physics and the Conceptual Structure of the Scientific Revolution [Bec91r], pages i–xviii. ISBN 0-7923-1054-3, 94-011-3276-3. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 127. URL https://link.

springer.com/book/10.1007/978-
94-011-3276-3.

Anonymous:1991:FMx

[Ano91y] Anonymous. Front matter. In Explanation in the Sciences [Mey91d], pages i–6. ISBN 94-010-5511-4, 94-011-3414-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
book/10.1007/978-94-011-3414-9.

Anonymous:1991:FMy

[Ano91z] Anonymous. Front matter. In Explanation in the Sciences [Mey91d], page 7. ISBN 94-010-5511-4, 94-011-3414-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
book/10.1007/978-94-011-3414-9.

Anonymous:1991:FMz

[Ano91-27] Anonymous. Front matter. In Explanation in the Sciences [Mey91d], page 45. ISBN 94-010-5511-4, 94-011-3414-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
book/10.1007/978-94-011-3414-9.

Anonymous:1991:FMaa

[Ano91-28] Anonymous. Front matter. In Explanation in the Sciences [Mey91d], page 261. ISBN 94-010-5511-4, 94-011-3414-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
book/10.1007/978-94-011-3414-9.

Anonymous:1991:FMab

[Ano91-29] Anonymous. Front matter. In Explanation in the Sciences [Mey91d], page 393. ISBN 94-010-5511-4, 94-011-3414-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
book/10.1007/978-94-011-3414-9.

Anonymous:1991:FMw

[Ano91-30] Anonymous. Front matter. In Organism and the Origins of Self [Tau91i], pages i–xix. ISBN 0-7923-1185-X, 94-011-3406-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QR180-189.5. URL https://link.

springer.com/book/10.1007/978-
94-011-3406-4.

Anonymous:1991:FMg

[Ano91-31] Anonymous. Front matter. In Methodological Variance: Essays in Epistemological Ontology and the Methodology of Science [Pan91f], pages i–xxiv. ISBN 94-010-5400-2, 94-011-3174-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
book/10.1007/978-94-011-3174-2.

Anonymous:1991:FMh

[Ano91-32] Anonymous. Front matter. In Methodological Variance: Essays in Epistemological Ontology and the Methodology of Science [Pan91f], page 1. ISBN 94-010-5400-2, 94-011-3174-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL

https://link.springer.com/book/
10.1007/978-94-011-3174-2.

Anonymous:1991:FMi

[Ano91-33] Anonymous. Front matter. In Methodological Variance: Essays in Epistemological Ontology and the Methodology of Science [Pan91f], page 97. ISBN 94-010-5400-2, 94-011-3174-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL

https://link.springer.com/book/
10.1007/978-94-011-3174-2.

Anonymous:1991:FMo

[Ano91-34] Anonymous. Front matter. In Beyond Reason: Essays on the Philosophy of Paul Feyerabend [Mun91a], pages i–xxi. ISBN 94-010-5406-1, 94-011-3188-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
book/10.1007/978-94-011-3188-9.

Anonymous:1991:FMj

[Ano91-35] Anonymous. Front matter. In Uebel [Ueb91b], pages i–xii. ISBN 0-7923-1276-7 (hardcover), 94-011-3182-1. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 133.

URL https://link.springer.com/
book/10.1007/978-94-011-3182-7.

Anonymous:1991:FMk

[Ano91-36] Anonymous. Front matter. In Uebel [Ueb91b], page 1. ISBN 0-7923-1276-7 (hardcover), 94-011-3182-1. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 133. URL

https://link.springer.com/book/
10.1007/978-94-011-3182-7.

Anonymous:1991:FMl

[Ano91-37] Anonymous. Front matter. In Uebel [Ueb91b], page 23. ISBN 0-7923-1276-7 (hardcover), 94-011-3182-1. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 133. URL

                                              https://link.springer.com/book/
                                              10.1007/978-94-011-3182-7.

Anonymous:1991:FMm

[Ano91-38] Anonymous. Front matter. In Uebel [Ueb91b], page 79. ISBN 0-7923-1276-7 (hardcover), 94-011-3182-1. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 133. URL

                                              https://link.springer.com/book/
                                              10.1007/978-94-011-3182-7.

Anonymous:1991:FMn

[Ano91-39] Anonymous. Front matter. In Uebel [Ueb91b], page 253. ISBN 0-7923-1276-7 (hardcover), 94-011-3182-1. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 133.

                                              URL https://link.springer.com/
                                              book/10.1007/978-94-011-3182-7.

Anonymous:1991:FMf

[Ano91-40] Anonymous. Front matter. In Woodward and Cohen [WC91], pages i–xvi. ISBN 94-010-5395-2, 94-011-3164-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

                                              URL https://link.springer.com/
                                              book/10.1007/978-94-011-3164-3.

Anonymous:1992:BMc

[Ano91-41] Anonymous. Back matter. In The Invention of Physical Science: Intersections of Mathematics, Theology and Natural Philosophy Since the Seventeenth Century: Essays in Honor of Erwin N. Hiebert [NRS92], pages 261– 278. ISBN 94-010-5097-X, 94-011-2488-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
book/10.1007/978-94-011-2488-1.

Anonymous:1992:BMa

[Ano91-42] Anonymous. Back matter. In The Invention of Physical Science: Intersections of Mathematics, Theology and Natural Philosophy Since the Seventeenth Century: Essays in Honor of Erwin N. Hiebert [NRS92], pages 459– 460. ISBN 94-010-5097-X, 94-011-2488-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
book/10.1007/978-94-011-2488-1.

Anonymous:1992:FMa

[Ano91-43] Anonymous. Front matter. In The Invention of Physical Science: Intersections of Mathematics, Theology and Natural Philosophy Since the Seventeenth Century: Essays in Honor of Erwin N. Hiebert [NRS92], pages i–x. ISBN 94-010-5097-X, 94-011-2488-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
book/10.1007/978-94-011-2488-1.

Anonymous:1992:FMf

[Ano91-44] Anonymous. Front matter. In The Invention of Physical Science: Intersections of Mathematics, Theology and Natural Philosophy Since the Seventeenth Century: Essays in Honor of Erwin N. Hiebert [NRS92], pages i–xxxiv. ISBN 94-010-5097-X, 94-011-2488-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
book/10.1007/978-94-011-2488-1.

Anonymous:1992:FMg

[Ano91-45] Anonymous. Front matter. In The Invention of Physical Science: Intersections of Mathematics, Theology and Natural Philosophy Since the Seventeenth Century: Essays in Honor of Erwin N. Hiebert [NRS92], page 1. ISBN 94-010-5097-X, 94-011-2488-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
book/10.1007/978-94-011-2488-1.

Anonymous:1992:FMh

[Ano91-46] Anonymous. Front matter. In The Invention of Physical Science: Intersections of Mathematics, Theology and Natural Philosophy Since the Seventeenth Century: Essays in Honor of Erwin N. Hiebert [NRS92], page 79. ISBN 94-010-5097-X, 94-011-2488-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
book/10.1007/978-94-011-2488-1.

Anonymous:1992:FMi

[Ano91-47] Anonymous. Front matter. In The Invention of Physical Science: Intersections of Mathematics, Theology and Natural Philosophy Since the Seventeenth Century: Essays in Honor of Erwin N. Hiebert [NRS92], page 117. ISBN 94-010-5097-X, 94-011-2488-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
book/10.1007/978-94-011-2488-1.

Anonymous:1992:FMj

[Ano91-48] Anonymous. Front matter. In The Invention of Physical Science: Intersections of Mathematics, Theology and Natural Philosophy Since the Seventeenth Century: Essays in Honor of Erwin N. Hiebert [NRS92], page 203. ISBN 94-010-5097-X, 94-011-2488-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
book/10.1007/978-94-011-2488-1.

Anonymous:1992:BMl

[Ano92a] Anonymous. Back matter. In Understanding Origins: Contemporary Views on the Origin of Life, Mind and Society [VD92], pages 305–329. ISBN 90-481-4090-0, 94-015-8054-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
book/10.1007/978-94-015-8054-0.

Anonymous:1992:BMm

[Ano92b] Anonymous. Back matter. In The Speculum astronomiae and its enigma: astrology, theology, and science in Albertus Magnus and his contemporaries [Zam92h], pages 127–352. ISBN 0-7923-1380-1 (hardcover), 94-017-3467-4. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174

.B67 vol. 135. URL http://www.loc.
gov/catdir/enhancements/fy0826/
91024566-d.html;   http://www.
loc.gov/catdir/enhancements/
fy0826/91024566-t.html;   https:
//link.springer.com/book/10.
1007/978-94-017-3467-7.

Anonymous:1992:BMd

[Ano92c] Anonymous. Back matter. In Science and Empires: Historical Studies about Scientific Development and European Expansion [PJM92], pages 393–418. ISBN 0-7923-1518-9 (hardcover), 94-010-5145-3, 94-011-2594-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 136.

URL https://link.springer.com/
book/10.1007/978-94-011-2594-9.

Anonymous:1992:BMk

[Ano92d] Anonymous. Back matter. In Galileo’s logic of discovery and proof: the background, content, and use of his appropriated treatises on Aristotle’s Posterior analytics [Wal92c], pages 304–333. ISBN 0-7923-1577-4, 0-7923-1579-0 (set), 94-015-8040-5. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 137.

URL https://link.springer.com/
book/10.1007/978-94-015-8040-3.

Anonymous:1992:BMj

[Ano92e] Anonymous. Back matter. In Galileo’s Logical Treatises: A Translation, with Notes and Commentary, of His Appropriated Latin Questions on Aristotle’s Posterior Analytics [Wal92e], pages 216–251. ISBN 90-481-4116-8, 94-015-8036-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic).

LCCN ????  URL https://link.
springer.com/book/10.1007/978-
94-015-8036-6.

Anonymous:1992:BMh

[Ano92f] Anonymous. Back matter. In Sarkar [Sar92a], pages 295–306. ISBN 0-7923-1777-7 (hardcover), 94-011-2856-1. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67

vol. 142.  URL http://www.loc.
gov/catdir/enhancements/fy0823/
92012823-d.html;   http://www.
loc.gov/catdir/enhancements/
fy0823/92012823-t.html;   https:
//link.springer.com/book/10.
1007/978-94-011-2856-8.

Anonymous:1992:BMg

[Ano92g] Anonymous. Back matter. In Ernst Mach — A Deeper Look: Documents and New Perspectives [Bla92a], pages 431–463. ISBN 94-010-5230-1, 94-011-2771-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN C5.53; QC6.9. URL https://link.

springer.com/book/10.1007/978-
94-011-2771-4.

Anonymous:1992:BMi

[Ano92h] Anonymous. Back matter. In Technological Development and Science in the Industrial Age: New Perspectives on the Science–Technology Relationship [KB92b], pages 271–280. ISBN 90-481-4186-9, 94-015-8010-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL

https://link.springer.com/book/
10.1007/978-94-015-8010-6.

Anonymous:1992:BMf

[Ano92i] Anonymous. Back matter. In Between History and Method: Disputes about the Rationality of Science [Ams92b], pages 177–227. ISBN 94-010-5199-2, 94-011-2706-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 145. URL

https://link.springer.com/book/
10.1007/978-94-011-2706-6.

Anonymous:1992:BMe

[Ano92j] Anonymous. Back matter. In The Scientific Enterprise: The Bar-Hillel Colloquium: Studies in History, Philosophy, and Sociology of Science, Volume 4 [UM92], pages 295–302. ISBN 94-010-5190-9, 94-011-2688-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .S354 1992.

URL https://link.springer.com/
book/10.1007/978-94-011-2688-5.

Anonymous:1992:BMb

[Ano92k] Anonymous. Back matter. In Metaarchaeology: Reflections by Archaeologists and Philosophers [Emb92b], pages 289– 343. ISBN 94-010-4806-1, 94-011-1826-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL

https://link.springer.com/book/
10.1007/978-94-011-1826-2.

Anonymous:1992:FMaa

[Ano92l] Anonymous. Front matter. In Understanding Origins: Contemporary Views on the Origin of Life, Mind and Society [VD92], pages i–viii. ISBN 90-481-4090-0, 94-015-8054-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL

https://link.springer.com/book/
10.1007/978-94-015-8054-0.

Anonymous:1992:FMab

[Ano92m] Anonymous. Front matter. In Understanding Origins: Contemporary Views on the Origin of Life, Mind and Society [VD92], page 43. ISBN 90-481-4090-0, 94-015-8054-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL

https://link.springer.com/book/
10.1007/978-94-015-8054-0.

Anonymous:1992:FMac

[Ano92n] Anonymous. Front matter. In Understanding Origins: Contemporary Views on the Origin of Life, Mind and Society [VD92], page 111. ISBN 90-481-4090-0, 94-015-8054-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL

https://link.springer.com/book/
10.1007/978-94-015-8054-0.

Anonymous:1992:FMad

[Ano92o] Anonymous. Front matter. In Understanding Origins: Contemporary Views on the Origin of Life, Mind and Society [VD92], page 151. ISBN 90-481-4090-0, 94-015-8054-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL

https://link.springer.com/book/
10.1007/978-94-015-8054-0.

Anonymous:1992:FMae

[Ano92p] Anonymous. Front matter. In Understanding Origins: Contemporary Views on the Origin of Life, Mind and Society [VD92], page 233. ISBN 90-481-4090-0, 94-015-8054-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL

https://link.springer.com/book/
10.1007/978-94-015-8054-0.

Anonymous:1992:FMaf

[Ano92q] Anonymous. Front matter. In The Speculum astronomiae and its enigma: astrology, theology, and science in Albertus Magnus and his contemporaries [Zam92h], pages I–XVI. ISBN 0-7923-1380-1 (hardcover), 94-017-3467-4. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174

.B67 vol. 135. URL http://www.loc.
gov/catdir/enhancements/fy0826/
91024566-d.html;   http://www.
loc.gov/catdir/enhancements/
fy0826/91024566-t.html;   https:
//link.springer.com/book/10.
1007/978-94-017-3467-7.

Anonymous:1992:FMag

[Ano92r] Anonymous. Front matter. In The Speculum astronomiae and its enigma: astrology, theology, and science in Albertus Magnus and his contemporaries [Zam92h], page 1. ISBN 0-7923-1380-1 (hardcover), 94-017-3467-4. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174

.B67 vol. 135. URL http://www.loc.
gov/catdir/enhancements/fy0826/
91024566-d.html;   http://www.
loc.gov/catdir/enhancements/
fy0826/91024566-t.html;   https:
//link.springer.com/book/10.
1007/978-94-017-3467-7.

Anonymous:1992:FMah

[Ano92s] Anonymous. Front matter. In The Speculum astronomiae and its enigma: astrology, theology, and science in Albertus Magnus and his contemporaries [Zam92h], page 43. ISBN 0-7923-1380-1 (hardcover), 94-017-3467-4. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174

.B67 vol. 135. URL http://www.loc.
gov/catdir/enhancements/fy0826/
91024566-d.html;   http://www.
loc.gov/catdir/enhancements/
fy0826/91024566-t.html;   https:
//link.springer.com/book/10.
1007/978-94-017-3467-7.

Anonymous:1992:FMai

[Ano92t] Anonymous. Front matter. In The Speculum astronomiae and its enigma: astrology, theology, and science in Albertus Magnus and his contemporaries [Zam92h], page 103. ISBN 0-7923-1380-1 (hardcover), 94-017-3467-4. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174

.B67 vol. 135. URL http://www.loc.
gov/catdir/enhancements/fy0826/
91024566-d.html;   http://www.
loc.gov/catdir/enhancements/
fy0826/91024566-t.html;   https:
//link.springer.com/book/10.
1007/978-94-017-3467-7.

Anonymous:1992:FMk

[Ano92u] Anonymous. Front matter. In Science and Empires: Historical Studies about Scientific Development and European Expansion [PJM92], pages i–xiii. ISBN 0-7923-1518-9 (hardcover), 94-010-5145-3, 94-011-2594-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 136.

URL https://link.springer.com/
book/10.1007/978-94-011-2594-9.

Anonymous:1992:FMl

[Ano92v] Anonymous. Front matter. In Science and Empires: Historical Studies about Scientific Development and European Expansion [PJM92], page 1. ISBN 0-7923-1518-9 (hardcover), 94-010-5145-3, 94-011-2594-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 136.

URL https://link.springer.com/
book/10.1007/978-94-011-2594-9.

Anonymous:1992:FMm

[Ano92w] Anonymous. Front matter. In Science and Empires: Historical Studies about Scientific Development and European Expansion [PJM92], page 31. ISBN 0-7923-1518-9 (hardcover), 94-010-5145-3, 94-011-2594-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 136.

URL https://link.springer.com/
book/10.1007/978-94-011-2594-9.

Anonymous:1992:FMn

[Ano92x] Anonymous. Front matter. In Science and Empires: Historical Studies about Scientific Development and European Expansion [PJM92], page 209. ISBN 0-7923-1518-9 (hardcover), 94-010-5145-3, 94-011-2594-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 136.

URL https://link.springer.com/
book/10.1007/978-94-011-2594-9.

Anonymous:1992:FMo

[Ano92y] Anonymous. Front matter. In Science and Empires: Historical Studies about Scientific Development and European Expansion [PJM92], page 363. ISBN 0-7923-1518-9 (hardcover), 94-010-5145-3, 94-011-2594-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 136.

                                              URL https://link.springer.com/
                                              book/10.1007/978-94-011-2594-9.

Anonymous:1992:FMx

[Ano92z] Anonymous. Front matter. In Galileo’s logic of discovery and proof: the background, content, and use of his appropriated treatises on Aristotle’s Posterior analytics [Wal92c], pages i–xxiii. ISBN 0-7923-1577-4, 0-7923-1579-0 (set), 94-015-8040-5. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 137.

                                              URL https://link.springer.com/
                                              book/10.1007/978-94-015-8040-3.

Anonymous:1992:FMy

[Ano92-27] Anonymous. Front matter. In Galileo’s logic of discovery and proof: the background, content, and use of his appropriated treatises on Aristotle’s Posterior analytics [Wal92c], page 31. ISBN 0-7923-1577-4, 0-7923-1579-0 (set), 94-015-8040-5. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 137. URL

                                              https://link.springer.com/book/
                                              10.1007/978-94-015-8040-3.

Anonymous:1992:FMz

[Ano92-28] Anonymous. Front matter. In Galileo’s logic of discovery and proof: the background, content, and use of his appropriated treatises on Aristotle’s Posterior analytics [Wal92c], page 191. ISBN 0-7923-1577-4, 0-7923-1579-0 (set), 94-015-8040-5. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 137. URL

                                              https://link.springer.com/book/
                                              10.1007/978-94-015-8040-3.

Anonymous:1992:FMw

[Ano92-29] Anonymous. Front matter. In Galileo’s Logical Treatises: A Translation, with Notes and Commentary, of His Appropriated Latin Questions on Aristotle’s Posterior Analytics [Wal92e], pages i– xix. ISBN 90-481-4116-8, 94-015-8036-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
book/10.1007/978-94-015-8036-6.

Anonymous:1992:FMu

[Ano92-30] Anonymous. Front matter. In Sarkar [Sar92a], pages i–v. ISBN 0-7923-1777-7 (hardcover), 94-011-2856-1. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67

vol. 142.  URL http://www.loc.
gov/catdir/enhancements/fy0823/
92012823-d.html;   http://www.
loc.gov/catdir/enhancements/
fy0823/92012823-t.html;   https:
//link.springer.com/book/10.
1007/978-94-011-2856-8.

Anonymous:1992:FMt

[Ano92-31] Anonymous. Front matter. In Ernst Mach — A Deeper Look: Documents and New Perspectives [Bla92a], pages i– xix. ISBN 94-010-5230-1, 94-011-2771-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN C5.53; QC6.9.

URL https://link.springer.com/
book/10.1007/978-94-011-2771-4.

Anonymous:1992:FMv

[Ano92-32] Anonymous. Front matter. In Technological Development and Science in the Industrial Age: New Perspectives on the Science–Technology Relationship [KB92b], pages i–vii. ISBN 90-481-4186-9, 94-015-8010-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
book/10.1007/978-94-015-8010-6.

Anonymous:1992:FMs

[Ano92-33] Anonymous. Front matter. In Between History and Method: Disputes about the Rationality of Science [Ams92b], pages i–xi. ISBN 94-010-5199-2, 94-011-2706-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 145. URL

                                              https://link.springer.com/book/
                                              10.1007/978-94-011-2706-6.

Anonymous:1992:FMp

[Ano92-34] Anonymous. Front matter. In The Scientific Enterprise: The Bar-Hillel Colloquium: Studies in History, Philosophy, and Sociology of Science, Volume 4 [UM92], pages i–ix. ISBN 94-010-5190-9, 94-011-2688-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .S354 1992.

                                              URL https://link.springer.com/
                                              book/10.1007/978-94-011-2688-5.

Anonymous:1992:FMq

[Ano92-35] Anonymous. Front matter. In The Scientific Enterprise: The Bar-Hillel Colloquium: Studies in History, Philosophy, and Sociology of Science, Volume 4 [UM92], page 203. ISBN 94-010-5190-9, 94-011-2688-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .S354 1992.

                                              URL https://link.springer.com/
                                              book/10.1007/978-94-011-2688-5.

Anonymous:1992:FMr

[Ano92-36] Anonymous. Front matter. In The Scientific Enterprise: The Bar-Hillel Colloquium: Studies in History, Philosophy, and Sociology of Science, Volume 4 [UM92], page 261. ISBN 94-010-5190-9, 94-011-2688-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .S354 1992.

URL https://link.springer.com/
book/10.1007/978-94-011-2688-5.

Anonymous:1992:FMb

[Ano92-37] Anonymous. Front matter. In Metaarchaeology: Reflections by Archaeologists and Philosophers [Emb92b], pages i– vii. ISBN 94-010-4806-1, 94-011-1826-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
book/10.1007/978-94-011-1826-2.

Anonymous:1992:FMc

[Ano92-38] Anonymous. Front matter. In Metaarchaeology: Reflections by Archaeologists and Philosophers [Emb92b], page 1. ISBN 94-010-4806-1, 94-011-1826-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
book/10.1007/978-94-011-1826-2.

Anonymous:1992:FMd

[Ano92-39] Anonymous. Front matter. In Metaarchaeology: Reflections by Archaeologists and Philosophers [Emb92b], page 51. ISBN 94-010-4806-1, 94-011-1826-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
book/10.1007/978-94-011-1826-2.

Anonymous:1992:FMe

[Ano92-40] Anonymous. Front matter. In Metaarchaeology: Reflections by Archaeologists and Philosophers [Emb92b], page 141. ISBN 94-010-4806-1, 94-011-1826-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
book/10.1007/978-94-011-1826-2.

Anonymous:1993:BMb

[Ano93a] Anonymous. Back matter. In Bridging the Gap: Philosophy, Mathematics, and Physics: Lectures on the Foundations of Science [CCG93], pages 315–329. ISBN 94-010-5101-1, 94-011-2496-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
book/10.1007/978-94-011-2496-6.

Anonymous:1993:BMc

[Ano93b] Anonymous. Back matter. In Philosophy and Conceptual History of Science in Taiwan [LF93], pages 265–276. ISBN 94-010-5103-8, 94-011-2500-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q127.T28 .P455

1993. URL https://link.springer.
com/book/10.1007/978-94-011-
2500-0.

Anonymous:1993:BMe

[Ano93c] Anonymous. Back matter. In Correspondence, Invariance and Heuristics: Essays in Honour of Heinz Post [FK93], pages 343–363. ISBN 90-481-4229-6, 94-017-1185-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL

https://link.springer.com/book/
10.1007/978-94-017-1185-2.

Anonymous:1993:BMa

[Ano93d] Anonymous. Back matter. In The Context of Explanation [Bun93c], pages 131–164. ISBN 94-010-4760-X, 94-011-1735-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic).

LCCN ????  URL https://link.
springer.com/book/10.1007/978-
94-011-1735-7.

Anonymous:1993:BMd

[Ano93e] Anonymous. Back matter. In Historical Pragmatics: Philosophical Essays [But93d], pages 339–370. ISBN 90-481-4329-2, 94-015-8188-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL

https://link.springer.com/book/
10.1007/978-94-015-8188-2.

Anonymous:1993:FMb

[Ano93f] Anonymous. Front matter. In Bridging the Gap: Philosophy, Mathematics, and Physics: Lectures on the Foundations of Science [CCG93], pages i–xi. ISBN 94-010-5101-1, 94-011-2496-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
book/10.1007/978-94-011-2496-6.

Anonymous:1993:FMc

[Ano93g] Anonymous. Front matter. In Bridging the Gap: Philosophy, Mathematics, and Physics: Lectures on the Foundations of Science [CCG93], page 1. ISBN 94-010-5101-1, 94-011-2496-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
book/10.1007/978-94-011-2496-6.

Anonymous:1993:FMd

[Ano93h] Anonymous. Front matter. In Bridging the Gap: Philosophy, Mathematics, and Physics: Lectures on the Foundations of Science [CCG93], page 147. ISBN 94-010-5101-1, 94-011-2496-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
book/10.1007/978-94-011-2496-6.

Anonymous:1993:FMe

[Ano93i] Anonymous. Front matter. In Philosophy and Conceptual History of Science in Taiwan [LF93], pages i–xiv. ISBN 94-010-5103-8, 94-011-2500-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q127.T28 .P455

1993. URL https://link.springer.

                                              com/book/10.1007/978-94-011-
                                              2500-0.

Anonymous:1993:FMf

[Ano93j] Anonymous. Front matter. In Philosophy and Conceptual History of Science in Taiwan [LF93], page 1. ISBN 94-010-5103-8, 94-011-2500-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q127.T28 .P455 1993.

                                              URL https://link.springer.com/
                                              book/10.1007/978-94-011-2500-0.

Anonymous:1993:FMg

[Ano93k] Anonymous. Front matter. In Philosophy and Conceptual History of Science in Taiwan [LF93], page 121. ISBN 94-010-5103-8, 94-011-2500-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q127.T28 .P455

                                              1993. URL https://link.springer.
                                              com/book/10.1007/978-94-011-
                                              2500-0.

Anonymous:1993:FMl

[Ano93l] Anonymous. Front matter. In Correspondence, Invariance and Heuristics: Essays in Honour of Heinz Post [FK93], pages i–xxiii. ISBN 90-481-4229-6, 94-017-1185-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL

                                              https://link.springer.com/book/
                                              10.1007/978-94-017-1185-2.

Anonymous:1993:FMa

[Ano93m] Anonymous. Front matter. In The Context of Explanation [Bun93c], pages i– xii. ISBN 94-010-4760-X, 94-011-1735-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

                                              URL https://link.springer.com/
                                              book/10.1007/978-94-011-1735-7.

Anonymous:1993:FMh

[Ano93n] Anonymous. Front matter. In Historical Pragmatics: Philosophical Essays [But93d], pages i–xiii. ISBN 90-481-4329-2, 94-015-8188-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL

https://link.springer.com/book/
10.1007/978-94-015-8188-2.

Anonymous:1993:FMi

[Ano93o] Anonymous. Front matter. In Historical Pragmatics: Philosophical Essays [But93d], page 1. ISBN 90-481-4329-2, 94-015-8188-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL

https://link.springer.com/book/
10.1007/978-94-015-8188-2.

Anonymous:1993:FMj

[Ano93p] Anonymous. Front matter. In Historical Pragmatics: Philosophical Essays [But93d], page 65. ISBN 90-481-4329-2, 94-015-8188-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL

https://link.springer.com/book/
10.1007/978-94-015-8188-2.

Anonymous:1993:FMk

[Ano93q] Anonymous. Front matter. In Historical Pragmatics: Philosophical Essays [But93d], page 187. ISBN 90-481-4329-2, 94-015-8188-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL

https://link.springer.com/book/
10.1007/978-94-015-8188-2.

Anonymous:1994:BMj

[Ano94a] Anonymous. Back matter. In Cohen [Coh94b], pages 399–412. ISBN 0-7923-2223-1, 94-017-3391-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 150.

URL https://link.springer.com/
book/10.1007/978-94-017-3391-5.

Anonymous:1994:BMk

[Ano94b] Anonymous. Back matter. In Trends in the Historiography of Science [GCN94], pages 445–460. ISBN 90-481-4264-4, 94-017-3596-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL

https://link.springer.com/book/
10.1007/978-94-017-3596-4.

Anonymous:1994:BMh

[Ano94c] Anonymous. Back matter. In Poggi and Bossi [PB94], pages 237–256. ISBN 0-7923-2336-X, 94-017-2921-2. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67

vol. 152.  URL http://www.loc.
gov/catdir/enhancements/fy0821/
93001728-d.html;   http://www.
loc.gov/catdir/enhancements/
fy0821/93001728-t.html;   https:
//link.springer.com/book/10.
1007/978-94-017-2921-5.

Anonymous:1994:BMg

[Ano94d] Anonymous. Back matter. In Faye and Folse [FF94], pages 353–388. ISBN 0-7923-2378-5 (hardcover), 94-015-8106-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 153 1994;

QC16 1994. URL http://www.loc.
gov/catdir/enhancements/fy0823/
93024825-d.html;   http://www.
loc.gov/catdir/enhancements/
fy0823/93024825-t.html;   https:
//link.springer.com/book/10.
1007/978-94-015-8106-6.

Anonymous:1994:BMd

[Ano94e] Anonymous. Back matter. In Artifacts, Representations and Social Practice: Essays for Marx Wartofsky [GC94], pages 551–572. ISBN 94-010-4390-6, 94-011-0902-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN B73 .A77 1994.

URL https://link.springer.com/
book/10.1007/978-94-011-0902-4.

Anonymous:1994:BMi

[Ano94f] Anonymous. Back matter. In The Development of Arabic Mathematics: Between Arithmetic and Algebra [Ras94c], pages 332–382. ISBN 90-481-4338-1, 94-017-3274-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QA27.A67 R37 1994.

URL https://link.springer.com/
book/10.1007/978-94-017-3274-1.

Anonymous:1994:BMa

[Ano94g] Anonymous. Back matter. In Zygmunt Zawirski: His Life and Work: With Selected Writings on Time, Logic and the Methodology of Science [Coh94m], pages 79–80. ISBN 94-010-4329-9, 94-011-0768-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL

https://link.springer.com/book/
10.1007/978-94-011-0768-6.

Anonymous:1994:BMb

[Ano94h] Anonymous. Back matter. In Zygmunt Zawirski: His Life and Work: With Selected Writings on Time, Logic and the Methodology of Science [Coh94m], pages 369–374. ISBN 94-010-4329-9, 94-011-0768-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL

https://link.springer.com/book/
10.1007/978-94-011-0768-6.

Anonymous:1994:BMc

[Ano94i] Anonymous. Back matter. In Zygmunt Zawirski: His Life and Work:

With Selected Writings on Time, Logic and the Methodology of Science [Coh94m], pages 375–384. ISBN 94-010-4329-9, 94-011-0768-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL

                                              https://link.springer.com/book/
                                              10.1007/978-94-011-0768-6.

Anonymous:1994:BMf

[Ano94j] Anonymous. Back matter. In Names, Natures and Things: The Alchemist Jābir ibn Hayyān and his Kitāb al-Ahjār (Book of Stones) [Haq94f], pages 249–284. ISBN 0-7923-2587-7, 94-011-1898-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 158 QD25. URL

                                              https://link.springer.com/book/
                                              10.1007/978-94-011-1898-9.

Anonymous:1994:BMe

[Ano94k] Anonymous. Back matter. In Kant’s Theory of Natural Science [Pla94b], pages 333–367. ISBN 94-010-4492-9, 94-011-1126-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic).

                                              LCCN ????  URL https://link.
                                              springer.com/book/10.1007/978-
                                              94-011-1126-3.

Anonymous:1994:FMl

[Ano94l] Anonymous. Front matter. In Cohen [Coh94b], pages i–xxxvi. ISBN 0-7923-2223-1, 94-017-3391-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 150.

                                              URL https://link.springer.com/
                                              book/10.1007/978-94-017-3391-5.

Anonymous:1994:FMm

[Ano94m] Anonymous. Front matter. In Trends in the Historiography of Science [GCN94], pages i–xi. ISBN 90-481-4264-4, 94-017-3596-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/book/10.1007/978-94-017-
3596-4.

Anonymous:1994:FMj

[Ano94n] Anonymous. Front matter. In Poggi and Bossi [PB94], pages i–xv. ISBN 0-7923-2336-X, 94-017-2921-2. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 152.

URL http://www.loc.gov/catdir/
enhancements/fy0821/93001728-
d.html;   http://www.loc.gov/
catdir/enhancements/fy0821/
93001728-t.html;   https://link.
springer.com/book/10.1007/978-
94-017-2921-5.

Anonymous:1994:FMi

[Ano94o] Anonymous. Front matter. In Faye and Folse [FF94], pages i–xxvii. ISBN 0-7923-2378-5 (hardcover), 94-015-8106-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 153 1994;

QC16 1994. URL http://www.loc.
gov/catdir/enhancements/fy0823/
93024825-d.html;   http://www.
loc.gov/catdir/enhancements/
fy0823/93024825-t.html;   https:
//link.springer.com/book/10.
1007/978-94-015-8106-6.

Anonymous:1994:FMd

[Ano94p] Anonymous. Front matter. In Artifacts, Representations and Social Practice: Essays for Marx Wartofsky [GC94], pages i–xiv. ISBN 94-010-4390-6, 94-011-0902-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN B73 .A77 1994. URL

https://link.springer.com/book/
10.1007/978-94-011-0902-4.

Anonymous:1994:FMk

[Ano94q] Anonymous. Front matter. In The Development of Arabic Mathematics: Between Arithmetic and Algebra [Ras94c], pages i–x. ISBN 90-481-4338-1, 94-017-3274-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QA27.A67 R37 1994.

                                              URL https://link.springer.com/
                                              book/10.1007/978-94-017-3274-1.

Anonymous:1994:FMa

[Ano94r] Anonymous. Front matter. In Zygmunt Zawirski: His Life and Work: With Selected Writings on Time, Logic and the Methodology of Science [Coh94m], pages i–xv. ISBN 94-010-4329-9, 94-011-0768-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL

                                              https://link.springer.com/book/
                                              10.1007/978-94-011-0768-6.

Anonymous:1994:FMb

[Ano94s] Anonymous. Front matter. In Zygmunt Zawirski: His Life and Work: With Selected Writings on Time, Logic and the Methodology of Science [Coh94m], page 1. ISBN 94-010-4329-9, 94-011-0768-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

                                              ???? URL https://link.springer.
                                              com/book/10.1007/978-94-011-
                                              0768-6.

Anonymous:1994:FMc

[Ano94t] Anonymous. Front matter. In Zygmunt Zawirski: His Life and Work: With Selected Writings on Time, Logic and the Methodology of Science [Coh94m], page 81. ISBN 94-010-4329-9, 94-011-0768-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

                                              ???? URL https://link.springer.

com/book/10.1007/978-94-011-
0768-6.

Anonymous:1994:FMf

[Ano94u] Anonymous. Front matter. In Names, Natures and Things: The Alchemist Jābir ibn Hayyān and his Kitāb al-Ahjār (Book of Stones) [Haq94f], pages i–xx. ISBN 0-7923-2587-7, 94-011-1898-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 158 QD25. URL

https://link.springer.com/book/
10.1007/978-94-011-1898-9.

Anonymous:1994:FMh

[Ano94v] Anonymous. Front matter. In Names, Natures and Things: The Alchemist Jābir ibn Hayyān and his Kitāb al-Ahjār (Book of Stones) [Haq94f], page 109. ISBN 0-7923-2587-7, 94-011-1898-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 158 QD25. URL

https://link.springer.com/book/
10.1007/978-94-011-1898-9.

Anonymous:1994:FMg

[Ano94w] Anonymous. Front matter. In Names, Natures and Things: The Alchemist Jābir ibn Hayyān and his Kitāb al-Ahjār (Book of Stones) [Haq94f], pages N1–N2. ISBN 0-7923-2587-7, 94-011-1898-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 158 QD25. URL

https://link.springer.com/book/
10.1007/978-94-011-1898-9.

Anonymous:1994:FMe

[Ano94x] Anonymous. Front matter. In Kant’s Theory of Natural Science [Pla94b], pages i–xvi. ISBN 94-010-4492-9, 94-011-1126-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic).

LCCN ????  URL https://link.
springer.com/book/10.1007/978-
94-011-1126-3.

Anonymous:1995:BMe

[Ano95a] Anonymous. Back matter. In Misiek [Mis95c], pages 265–276. ISBN 94-010-4206-3, 94-011-0461-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL

https://link.springer.com/book/
10.1007/978-94-011-0461-6.

Anonymous:1995:BMg

[Ano95b] Anonymous. Back matter. In Critical Rationalism, Metaphysics and Science: Essays for Joseph Agassi. Volume I [JL95a], pages 239–266. ISBN 94-010-4211-X, 94-011-0471-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
book/10.1007/978-94-011-0471-5.

Anonymous:1995:BMj

[Ano95c] Anonymous. Back matter. In Critical Rationalism, the Social Sciences and the Humanities: Essays for Joseph Agassi. Volume II [JL95b], pages 235–288. ISBN 0-7923-2961-9, 90-481-4430-2, 94-017-0441-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN B29 .C75 1995.

URL https://link.springer.com/
book/10.1007/978-94-017-0441-0.

Anonymous:1995:BMl

[Ano95d] Anonymous. Back matter. In Gavroglu et al. [GSW95c], pages 385–388. ISBN 0-7923-2991-0 (set), 0-7923-2988-0, 94-017-2658-2. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 163. URL

http://catdir.loc.gov/catdir/
enhancements/fy0823/94022250-
d.html;   http://catdir.loc.

gov/catdir/enhancements/fy0823/
94022250-t.html;   https://link.
springer.com/book/10.1007/978-
94-017-2658-0.

Anonymous:1995:BMf

[Ano95e] Anonymous. Back matter. In Science, Mind and Art: Essays on science and the humanistic understanding in art, epistemology, religion and ethics In honor of Robert S. Cohen [GSW95b], pages 441–462. ISBN 0-7923-2990-2 (hardcover), 94-011-0469-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 165 BD418.3. URL

https://link.springer.com/book/
10.1007/978-94-011-0469-2.

Anonymous:1995:BMi

[Ano95f] Anonymous. Back matter. In Transformation in the Writing: A Case of Surrender-and-Catch [Wol95l], pages 205–223. ISBN 90-481-4478-7, 94-015-8412-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN B829.5.A-829.5.Z. URL

https://link.springer.com/book/
10.1007/978-94-015-8412-8.

Anonymous:1995:BMc

[Ano95g] Anonymous. Back matter. In No Truth Except in the Details: Essays in Honor of Martin J. Klein [KS95], pages 287– 291. ISBN 94-010-4097-4, 94-011-0217-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL

https://link.springer.com/book/
10.1007/978-94-011-0217-9.

Anonymous:1995:BMd

[Ano95h] Anonymous. Back matter. In No Truth Except in the Details: Essays in Honor of Martin J. Klein [KS95], pages 363– 382. ISBN 94-010-4097-4, 94-011-0217-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL

                                              https://link.springer.com/book/
                                              10.1007/978-94-011-0217-9.

Anonymous:1995:BMk

[Ano95i] Anonymous. Back matter. In Blackmore [Bla95g], pages 215–272. ISBN 0-7923-3231-8, 90-481-4492-2, 94-017-0489-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 168; QC16 .B64. URL

                                              https://link.springer.com/book/
                                              10.1007/978-94-017-0489-2.

Anonymous:1995:BMm

[Ano95j] Anonymous. Back matter. In The Concept of Knowledge: The Ankara Seminar [KC95], pages 231–246. ISBN 90-481-4495-7, 94-017-3263-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL

                                              https://link.springer.com/book/
                                              10.1007/978-94-017-3263-5.

Anonymous:1995:BMa

[Ano95k] Anonymous. Back matter. In Mexican Studies in the History and Philosophy of Science [RC95], pages 275–291. ISBN 94-010-6535-7 (hardcover), 94-009-0109-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

                                              ???? URL https://link.springer.
                                              com/book/10.1007/978-94-009-
                                              0109-4.

Anonymous:1995:BMb

[Ano95l] Anonymous. Back matter. In Québec Studies in the Philosophy of Science: Part I: Logic, Mathematics, Physics and History of Science Essays in Honor of Hugues Leblanc [MC95], pages 309– 331. ISBN 94-010-7204-3, 94-009-1575-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL

https://link.springer.com/book/
10.1007/978-94-009-1575-6.

Anonymous:1995:BMh

[Ano95m] Anonymous. Back matter. In Quantum Reprogramming: Ensembles and Single Systems: a Two-Tier Approach to Quantum Mechanics [Pos95p], pages 308–322. ISBN 0-7923-3565-1 (hardcover), 94-015-8410-9. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 181; Q174; Q174.12; QC174.12 P675

1995. URL http://books.google.
com/books?id=lK3aAAAAMAAJ;
http://catalog.hathitrust.org/
api/volumes/oclc/32469387.html;
 http://catdir.loc.gov/catdir/
enhancements/fy0823/95017463d.
html;   http://catdir.loc.gov/
catdir/enhancements/fy0823/
95017463t.html;   https://link.
springer.com/book/10.1007/978-
94-015-8410-4.

Anonymous:1995:FMe

[Ano95n] Anonymous. Front matter. In Misiek [Mis95c], pages i–xii. ISBN 94-010-4206-3, 94-011-0461-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL

https://link.springer.com/book/
10.1007/978-94-011-0461-6.

Anonymous:1995:FMg

[Ano95o] Anonymous. Front matter. In Critical Rationalism, Metaphysics and Science: Essays for Joseph Agassi. Volume I [JL95a], pages i–xxii. ISBN 94-010-4211-X, 94-011-0471-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL

https://link.springer.com/book/
10.1007/978-94-011-0471-5.

Anonymous:1995:FMh

[Ano95p] Anonymous. Front matter. In Critical Rationalism, Metaphysics and Science: Essays for Joseph Agassi. Volume I [JL95a], page 1. ISBN 94-010-4211-X, 94-011-0471-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL

                                              https://link.springer.com/book/
                                              10.1007/978-94-011-0471-5.

Anonymous:1995:FMi

[Ano95q] Anonymous. Front matter. In Critical Rationalism, Metaphysics and Science: Essays for Joseph Agassi. Volume I [JL95a], page 131. ISBN 94-010-4211-X, 94-011-0471-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL

                                              https://link.springer.com/book/
                                              10.1007/978-94-011-0471-5.

Anonymous:1995:FMs

[Ano95r] Anonymous. Front matter. In Critical Rationalism, the Social Sciences and the Humanities: Essays for Joseph Agassi. Volume II [JL95b], pages i–xvii. ISBN 0-7923-2961-9, 90-481-4430-2, 94-017-0441-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN B29 .C75 1995. URL

                                              https://link.springer.com/book/
                                              10.1007/978-94-017-0441-0.

Anonymous:1995:FMu

[Ano95s] Anonymous.  Front matter.  In

Gavroglu et al. [GSW95c], pages i–xxvii. ISBN 0-7923-2991-0 (set), 0-7923-2988-0, 94-017-2658-2. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 163. URL

                                              http://catdir.loc.gov/catdir/
                                              enhancements/fy0823/94022250-
                                              d.html;   http://catdir.loc.
                                              gov/catdir/enhancements/fy0823/

94022250-t.html;   https://link.
springer.com/book/10.1007/978-
94-017-2658-0.

Anonymous:1995:FMf

[Ano95t] Anonymous. Front matter. In Science, Mind and Art: Essays on science and the humanistic understanding in art, epistemology, religion and ethics In honor of Robert S. Cohen [GSW95b], pages i–ix. ISBN 0-7923-2990-2 (hardcover), 94-011-0469-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 165 BD418.3. URL

https://link.springer.com/book/
10.1007/978-94-011-0469-2.

Anonymous:1995:FMo

[Ano95u] Anonymous. Front matter. In Transformation in the Writing: A Case of Surrender-and-Catch [Wol95l], pages i–xxiii. ISBN 90-481-4478-7, 94-015-8412-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN B829.5.A-829.5.Z. URL

https://link.springer.com/book/
10.1007/978-94-015-8412-8.

Anonymous:1995:FMp

[Ano95v] Anonymous. Front matter. In Transformation in the Writing: A Case of Surrender-and-Catch [Wol95l], page 1. ISBN 90-481-4478-7, 94-015-8412-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN B829.5.A-829.5.Z. URL https://

link.springer.com/book/10.1007/
978-94-015-8412-8.

Anonymous:1995:FMq

[Ano95w] Anonymous. Front matter. In Transformation in the Writing: A Case of Surrender-and-Catch [Wol95l], page 53. ISBN 90-481-4478-7, 94-015-8412-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN B829.5.A-829.5.Z. URL https://

                                              link.springer.com/book/10.1007/
                                              978-94-015-8412-8.

Anonymous:1995:FMr

[Ano95x] Anonymous. Front matter. In Transformation in the Writing: A Case of Surrender-and-Catch [Wol95l], page 167. ISBN 90-481-4478-7, 94-015-8412-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN B829.5.A-829.5.Z. URL https://

                                              link.springer.com/book/10.1007/
                                              978-94-015-8412-8.

Anonymous:1995:FMc

[Ano95y] Anonymous. Front matter. In No Truth Except in the Details: Essays in Honor of Martin J. Klein [KS95], pages i–vi. ISBN 94-010-4097-4, 94-011-0217-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

                                              URL https://link.springer.com/
                                              book/10.1007/978-94-011-0217-9.

Anonymous:1995:FMd

[Ano95z] Anonymous. Front matter. In No Truth Except in the Details: Essays in Honor of Martin J. Klein [KS95], pages i–xxv. ISBN 94-010-4097-4, 94-011-0217-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

                                              URL https://link.springer.com/
                                              book/10.1007/978-94-011-0217-9.

Anonymous:1995:FMt

[Ano95-27] Anonymous. Front matter. In Blackmore [Bla95g], pages i–xvi. ISBN 0-7923-3231-8, 90-481-4492-2, 94-017-0489-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 168; QC16 .B64. URL

                                              https://link.springer.com/book/
                                              10.1007/978-94-017-0489-2.

Anonymous:1995:FMv

[Ano95-28] Anonymous. Front matter. In The Concept of Knowledge: The Ankara Seminar [KC95], pages i–lviii. ISBN 90-481-4495-7, 94-017-3263-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL

https://link.springer.com/book/
10.1007/978-94-017-3263-5.

Anonymous:1995:FMw

[Ano95-29] Anonymous. Front matter. In The Concept of Knowledge: The Ankara Seminar [KC95], page 1. ISBN 90-481-4495-7, 94-017-3263-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL

https://link.springer.com/book/
10.1007/978-94-017-3263-5.

Anonymous:1995:FMx

[Ano95-30] Anonymous. Front matter. In The Concept of Knowledge: The Ankara Seminar [KC95], page 81. ISBN 90-481-4495-7, 94-017-3263-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL

https://link.springer.com/book/
10.1007/978-94-017-3263-5.

Anonymous:1995:FMy

[Ano95-31] Anonymous. Front matter. In The Concept of Knowledge: The Ankara Seminar [KC95], page 165. ISBN 90-481-4495-7, 94-017-3263-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL

https://link.springer.com/book/
10.1007/978-94-017-3263-5.

Anonymous:1995:FMa

[Ano95-32] Anonymous. Front matter. In Mexican Studies in the History and Philosophy of Science [RC95], pages i–xii. ISBN 94-010-6535-7 (hardcover), 94-009-0109-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
book/10.1007/978-94-009-0109-4.

Anonymous:1995:FMb

[Ano95-33] Anonymous. Front matter. In Québec Studies in the Philosophy of Science: Part I: Logic, Mathematics, Physics and History of Science Essays in Honor of Hugues Leblanc [MC95], pages i–xi. ISBN 94-010-7204-3, 94-009-1575-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
book/10.1007/978-94-009-1575-6.

Anonymous:1995:FMj

[Ano95-34] Anonymous. Front matter. In Quantum Reprogramming: Ensembles and Single Systems: a Two-Tier Approach to Quantum Mechanics [Pos95p], pages i–xiv. ISBN 0-7923-3565-1 (hardcover), 94-015-8410-9. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 181; Q174; Q174.12; QC174.12 P675

1995. URL http://books.google.
com/books?id=lK3aAAAAMAAJ;
http://catalog.hathitrust.org/
api/volumes/oclc/32469387.html;
 http://catdir.loc.gov/catdir/
enhancements/fy0823/95017463d.
html;   http://catdir.loc.gov/
catdir/enhancements/fy0823/
95017463t.html;   https://link.
springer.com/book/10.1007/978-
94-015-8410-4.

Anonymous:1995:FMk

[Ano95-35] Anonymous. Front matter. In Quantum Reprogramming: Ensembles and Single Systems: a Two-Tier Approach to Quantum Mechanics [Pos95p], page 15. ISBN 0-7923-3565-1 (hardcover), 94-015-8410-9.

ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 181; Q174; Q174.12; QC174.12 P675

1995. URL http://books.google.
com/books?id=lK3aAAAAMAAJ;
http://catalog.hathitrust.org/
api/volumes/oclc/32469387.html;
 http://catdir.loc.gov/catdir/
enhancements/fy0823/95017463d.
html;   http://catdir.loc.gov/
catdir/enhancements/fy0823/
95017463t.html;   https://link.
springer.com/book/10.1007/978-
94-015-8410-4.

Anonymous:1995:FMl

[Ano95-36] Anonymous. Front matter. In Quantum Reprogramming: Ensembles and Single Systems: a Two-Tier Approach to Quantum Mechanics [Pos95p], page 84. ISBN 0-7923-3565-1 (hardcover), 94-015-8410-9. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 181; Q174; Q174.12; QC174.12 P675

1995. URL http://books.google.
com/books?id=lK3aAAAAMAAJ;
http://catalog.hathitrust.org/
api/volumes/oclc/32469387.html;
 http://catdir.loc.gov/catdir/
enhancements/fy0823/95017463d.
html;   http://catdir.loc.gov/
catdir/enhancements/fy0823/
95017463t.html;   https://link.
springer.com/book/10.1007/978-
94-015-8410-4.

Anonymous:1995:FMm

[Ano95-37] Anonymous. Front matter. In Quantum Reprogramming: Ensembles and Single Systems: a Two-Tier Approach to Quantum Mechanics [Pos95p], page 188. ISBN 0-7923-3565-1 (hardcover), 94-015-8410-9. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol.

181; Q174; Q174.12; QC174.12 P675

                                              1995. URL http://books.google.
                                              com/books?id=lK3aAAAAMAAJ;
                                              http://catalog.hathitrust.org/
                                              api/volumes/oclc/32469387.html;
                                               http://catdir.loc.gov/catdir/
                                              enhancements/fy0823/95017463d.
                                              html;   http://catdir.loc.gov/
                                              catdir/enhancements/fy0823/
                                              95017463t.html;   https://link.
                                              springer.com/book/10.1007/978-
                                              94-015-8410-4.

Anonymous:1995:FMn

[Ano95-38] Anonymous. Front matter. In Quantum Reprogramming: Ensembles and Single Systems: a Two-Tier Approach to Quantum Mechanics [Pos95p], page 239. ISBN 0-7923-3565-1 (hardcover), 94-015-8410-9. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 181; Q174; Q174.12; QC174.12 P675

                                              1995. URL http://books.google.
                                              com/books?id=lK3aAAAAMAAJ;
                                              http://catalog.hathitrust.org/
                                              api/volumes/oclc/32469387.html;
                                               http://catdir.loc.gov/catdir/
                                              enhancements/fy0823/95017463d.
                                              html;   http://catdir.loc.gov/
                                              catdir/enhancements/fy0823/
                                              95017463t.html;   https://link.
                                              springer.com/book/10.1007/978-
                                              94-015-8410-4.

Anonymous:1996:BMe

[Ano96a] Anonymous. Back matter. In Cohen et al. [CHR96], pages 481–495. ISBN 90-481-4493-0, 94-015-8638-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

                                              URL https://link.springer.com/
                                              book/10.1007/978-94-015-8638-2.

Anonymous:1996:BMd

[Ano96b] Anonymous. Back matter. In Abstraction and Representation: Essays on the Cultural Evolution of Thinking [Dam96b], pages 405–418. ISBN 0-7923-3816-2, 94-015-8624-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 175.

URL https://link.springer.com/
book/10.1007/978-94-015-8624-5.

Anonymous:1996:BMa

[Ano96c] Anonymous. Back matter. In Québec Studies in the Philosophy of Science: Part II: Biology, Psychology, Cognitive Science and Economics Essays in Honor of Hugues Leblanc [MC96], pages 295–308. ISBN 94-010-6537-3, 94-009-0113-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL

https://link.springer.com/book/
10.1007/978-94-009-0113-1.

Anonymous:1996:BMg

[Ano96d] Anonymous. Back matter. In Chinese Studies in the History and Philosophy of Science and Technology [DC96], pages 453–483. ISBN 90-481-4546-5, 94-015-8717-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 180; U390 .N19 1996.

URL https://link.springer.com/
book/10.1007/978-94-015-8717-4.

Anonymous:1996:BMf

[Ano96e] Anonymous. Back matter. In Bohmian Mechanics and Quantum Theory: an Appraisal [CFG96], pages 379– 408. ISBN 90-481-4698-4, 94-015-8715-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC173.96-174.52. URL https://link.

springer.com/book/10.1007/978-
94-015-8715-0.

Anonymous:1996:BMb

[Ano96f] Anonymous. Back matter. In Munévar [Mun96], pages 345–363. ISBN 94-010-6622-1, 94-009-0305-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL

                                              https://link.springer.com/book/
                                              10.1007/978-94-009-0305-0.

Anonymous:1996:BMh

[Ano96g] Anonymous. Back matter. In Schubring [Sch96f], pages 345–369. ISBN 90-481-4758-1, 94-015-8753-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QA8.9-10.3. URL

                                              https://link.springer.com/book/
                                              10.1007/978-94-015-8753-2.

Anonymous:1996:BMc

[Ano96h] Anonymous.  Back matter.  In

Schrödinger’s Philosophy of Quantum Mechanics [Bit96e], pages 267–292. ISBN 0-7923-4266-6, 94-009-1772-4. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 188;

                                              QC173.98. URL http://www.loc.
                                              gov/catdir/enhancements/fy0813/
                                              96042123-d.html;   http://www.
                                              loc.gov/catdir/enhancements/
                                              fy0813/96042123-t.html;   https:
                                              //link.springer.com/book/10.
                                              1007/978-94-009-1772-9.

Anonymous:1996:FMe

[Ano96i] Anonymous. Front matter. In Cohen et al. [CHR96], pages i–xxii. ISBN 90-481-4493-0, 94-015-8638-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL

                                              https://link.springer.com/book/
                                              10.1007/978-94-015-8638-2.

Anonymous:1996:FMd

[Ano96j] Anonymous. Front matter. In Abstraction and Representation: Essays on the Cultural Evolution of Thinking [Dam96b], pages i–xiv. ISBN 0-7923-3816-2, 94-015-8624-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 175.

URL https://link.springer.com/
book/10.1007/978-94-015-8624-5.

Anonymous:1996:FMa

[Ano96k] Anonymous. Front matter. In Québec Studies in the Philosophy of Science: Part II: Biology, Psychology, Cognitive Science and Economics Essays in Honor of Hugues Leblanc [MC96], pages i–xi. ISBN 94-010-6537-3, 94-009-0113-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
book/10.1007/978-94-009-0113-1.

Anonymous:1996:FMg

[Ano96l] Anonymous. Front matter. In Chinese Studies in the History and Philosophy of Science and Technology [DC96], pages i–xiii. ISBN 90-481-4546-5, 94-015-8717-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 180; U390 .N19 1996.

URL https://link.springer.com/
book/10.1007/978-94-015-8717-4.

Anonymous:1996:FMf

[Ano96m] Anonymous. Front matter. In Bohmian Mechanics and Quantum Theory: an Appraisal [CFG96], pages i–viii. ISBN 90-481-4698-4, 94-015-8715-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC173.96-174.52.

URL https://link.springer.com/
book/10.1007/978-94-015-8715-0.

Anonymous:1996:FMb

[Ano96n] Anonymous. Front matter. In Munévar [Mun96], pages i–xviii. ISBN 94-010-6622-1, 94-009-0305-7 (e-book).

ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL

                                              https://link.springer.com/book/
                                              10.1007/978-94-009-0305-0.

Anonymous:1996:FMh

[Ano96o] Anonymous.  Front matter.  In

Schubring [Sch96f], pages i–xxx. ISBN 90-481-4758-1, 94-015-8753-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QA8.9-10.3. URL

                                              https://link.springer.com/book/
                                              10.1007/978-94-015-8753-2.

Anonymous:1996:FMi

[Ano96p] Anonymous.  Front matter.  In

Schubring [Sch96f], pages 1–2. ISBN 90-481-4758-1, 94-015-8753-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QA8.9-10.3. URL

                                              https://link.springer.com/book/
                                              10.1007/978-94-015-8753-2.

Anonymous:1996:FMj

[Ano96q] Anonymous.  Front matter.  In

Schubring [Sch96f], pages 27–28. ISBN 90-481-4758-1, 94-015-8753-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QA8.9-10.3. URL

                                              https://link.springer.com/book/
                                              10.1007/978-94-015-8753-2.

Anonymous:1996:FMk

[Ano96r] Anonymous.  Front matter.  In

Schubring [Sch96f], page 115. ISBN 90-481-4758-1, 94-015-8753-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QA8.9-10.3. URL

                                              https://link.springer.com/book/
                                              10.1007/978-94-015-8753-2.

Anonymous:1996:FMl

[Ano96s] Anonymous.  Front matter.  In

Schubring [Sch96f], pages 229–230. ISBN 90-481-4758-1, 94-015-8753-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QA8.9-10.3.

URL https://link.springer.com/
book/10.1007/978-94-015-8753-2.

Anonymous:1996:FMc

[Ano96t] Anonymous.  Front matter.  In

Schrödinger’s Philosophy of Quantum Mechanics [Bit96e], pages I–XI. ISBN 0-7923-4266-6, 94-009-1772-4. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 188;

QC173.98. URL http://www.loc.
gov/catdir/enhancements/fy0813/
96042123-d.html;   http://www.
loc.gov/catdir/enhancements/
fy0813/96042123-t.html;   https:
//link.springer.com/book/10.
1007/978-94-009-1772-9.

Anonymous:1997:BMd

[Ano97a] Anonymous. Back matter. In Scarcity’s Ways: The Origins of Capital: A Critical Essay on Thermodynamics, Statistical Mechanics and Economics [Mac97f], pages 223–230. ISBN 90-481-4919-3, 94-015-8861-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN HB1-846.8. URL

https://link.springer.com/book/
10.1007/978-94-015-8861-4.

Anonymous:1997:BMa

[Ano97b] Anonymous. Back matter. In The Elusive Synthesis: Aesthetics and Science [Tau97a], pages 313–329. ISBN 0-7923-4763-3, 0-7923-3904-5, 94-009-1786-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
book/10.1007/978-94-009-1786-6.

Anonymous:1997:BMe

[Ano97c] Anonymous. Back matter. In Perspectives on Time [FSU97b], pages 459–472. ISBN 90-481-4774-3, 94-015-8875-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol.189. URL https://

link.springer.com/book/10.1007/
978-94-015-8875-1.

Anonymous:1997:BMb

[Ano97d] Anonymous. Back matter. In Austrian Philosophy Past and Present: Essays in Honor of Rudolf Haller [LM97], pages 251–274. ISBN 94-010-6412-1, 94-011-5720-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL

https://link.springer.com/book/
10.1007/978-94-011-5720-9.

Anonymous:1997:BMf

[Ano97e] Anonymous. Back matter. In Boissonnade and Vagliente [BV97], pages 575–594. ISBN 90-481-4779-4, 94-015-8903-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC5.53; QC6.9. URL https://link.

springer.com/book/10.1007/978-
94-015-8903-1.

Anonymous:1997:BMc

[Ano97f] Anonymous. Back matter. In Ginev and Cohen [GC97], pages 391–412. ISBN 0-7923-4444-8 (hardcover), 94-011-5788-X. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 192. URL https://link.

springer.com/book/10.1007/978-
94-011-5788-9.

Anonymous:1997:BMg

[Ano97g] Anonymous. Back matter. In Cohen et al. [CHS97b], pages 257–276. ISBN 90-481-4809-X, 94-017-2732-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC173.96-174.52.

URL https://link.springer.com/
book/10.1007/978-94-017-2732-7.

Anonymous:1997:FMd

[Ano97h] Anonymous.  Front matter.  In

Scarcity’s Ways: The Origins of Capital: A Critical Essay on Thermodynamics, Statistical Mechanics and Economics [Mac97f], pages i–ix. ISBN 90-481-4919-3, 94-015-8861-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN HB1-846.8. URL

https://link.springer.com/book/
10.1007/978-94-015-8861-4.

Anonymous:1997:FMa

[Ano97i] Anonymous. Front matter. In The Elusive Synthesis: Aesthetics and Science [Tau97a], pages I–XXIV. ISBN 0-7923-4763-3, 0-7923-3904-5, 94-009-1786-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
book/10.1007/978-94-009-1786-6.

Anonymous:1997:FMe

[Ano97j] Anonymous. Front matter. In Perspectives on Time [FSU97b], pages i–vii. ISBN 90-481-4774-3, 94-015-8875-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol.189. URL https://link.

springer.com/book/10.1007/978-
94-015-8875-1.

Anonymous:1997:FMf

[Ano97k] Anonymous. Front matter. In Perspectives on Time [FSU97b], page 59. ISBN 90-481-4774-3, 94-015-8875-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol.189.

URL https://link.springer.com/
book/10.1007/978-94-015-8875-1.

Anonymous:1997:FMg

[Ano97l] Anonymous. Front matter. In Perspectives on Time [FSU97b], page 201. ISBN 90-481-4774-3, 94-015-8875-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol.189. URL https://link.

springer.com/book/10.1007/978-
94-015-8875-1.

Anonymous:1997:FMh

[Ano97m] Anonymous. Front matter. In Perspectives on Time [FSU97b], page 299. ISBN 90-481-4774-3, 94-015-8875-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol.189. URL https://link.

springer.com/book/10.1007/978-
94-015-8875-1.

Anonymous:1997:FMb

[Ano97n] Anonymous. Front matter. In Austrian Philosophy Past and Present: Essays in Honor of Rudolf Haller [LM97], pages i– xiii. ISBN 94-010-6412-1, 94-011-5720-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
book/10.1007/978-94-011-5720-9.

Anonymous:1997:FMi

[Ano97o] Anonymous. Front matter. In Boissonnade and Vagliente [BV97], pages i–10. ISBN 90-481-4779-4, 94-015-8903-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC5.53; QC6.9. URL https://link.

springer.com/book/10.1007/978-
94-015-8903-1.

Anonymous:1997:FMc

[Ano97p] Anonymous. Front matter. In Ginev and Cohen [GC97], pages i–xvii. ISBN 0-7923-4444-8 (hardcover), 94-011-5788-X. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 192. URL https://link.

springer.com/book/10.1007/978-
94-011-5788-9.

Anonymous:1997:FMj

[Ano97q] Anonymous. Front matter. In Cohen et al. [CHS97b], pages i–xi. ISBN 90-481-4809-X, 94-017-2732-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC173.96-174.52.

URL https://link.springer.com/
book/10.1007/978-94-017-2732-7.

Anonymous:1998:BMa

[Ano98a] Anonymous. Back matter. In Japanese Studies in the Philosophy of Science [NC98], pages 189–202. ISBN 94-010-6176-9, 94-011-5175-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL

https://link.springer.com/book/
10.1007/978-94-011-5175-7.

Anonymous:1998:BMc

[Ano98b] Anonymous. Back matter. In Philosophies of Nature: The Human Dimension: In Celebration of Erazim Kohák [CT98], pages 315–334. ISBN 90-481-4859-6, 94-017-2614-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN BD581. URL

https://link.springer.com/book/
10.1007/978-94-017-2614-6.

Anonymous:1998:BMb

[Ano98c] Anonymous. Back matter. In Heinrich Hertz: Classical Physicist, Modern Philosopher [BHN98], pages 307–324. ISBN 90-481-4881-2, 94-015-8855-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
book/10.1007/978-94-015-8855-3.

Anonymous:1998:FMa

[Ano98d] Anonymous. Front matter. In Japanese Studies in the Philosophy of Science [NC98], pages i–xviii. ISBN 94-010-6176-9, 94-011-5175-X (e-book).

ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL

https://link.springer.com/book/
10.1007/978-94-011-5175-7.

Anonymous:1998:FMc

[Ano98e] Anonymous.  Front matter.  In

Philosophies of Nature: The Human Dimension: In Celebration of Erazim Kohák [CT98], pages i–xvii. ISBN 90-481-4859-6, 94-017-2614-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN BD581. URL

https://link.springer.com/book/
10.1007/978-94-017-2614-6.

Anonymous:1998:FMd

[Ano98f] Anonymous.  Front matter.  In

Philosophies of Nature: The Human Dimension: In Celebration of Erazim Kohák [CT98], page 1. ISBN 90-481-4859-6, 94-017-2614-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN BD581. URL

https://link.springer.com/book/
10.1007/978-94-017-2614-6.

Anonymous:1998:FMe

[Ano98g] Anonymous.  Front matter.  In

Philosophies of Nature: The Human Dimension: In Celebration of Erazim Kohák [CT98], page 207. ISBN 90-481-4859-6, 94-017-2614-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN BD581. URL

https://link.springer.com/book/
10.1007/978-94-017-2614-6.

Anonymous:1998:FMb

[Ano98h] Anonymous. Front matter. In Heinrich Hertz: Classical Physicist, Modern Philosopher [BHN98], pages i–xi. ISBN 90-481-4881-2, 94-015-8855-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
book/10.1007/978-94-015-8855-3.

Anonymous:1999:BMb

[Ano99a] Anonymous. Back matter. In The Natural Background of Meaning [Den99d], pages 238–252. ISBN 90-481-5126-0, 94-015-9084-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL

https://link.springer.com/book/
10.1007/978-94-015-9084-6.

Anonymous:1999:BMa

[Ano99b] Anonymous. Back matter. In Can that be Right?: Essays on Experiment, Evidence, and Science [Fra99c], pages 283–317. ISBN 94-010-6242-0, 94-011-5334-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic).

LCCN ????  URL https://link.
springer.com/book/10.1007/978-
94-011-5334-8.

Anonymous:1999:BMf

[Ano99c] Anonymous. Back matter. In Nietzsche, Theories of Knowledge, and Critical Theory: Nietzsche and the Sciences I [Bab99b], pages 315–348. ISBN 90-481-5233-X, 94-017-2430-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
book/10.1007/978-94-017-2430-2.

Anonymous:1999:BMe

[Ano99d] Anonymous. Back matter. In Nietzsche, Epistemology, and Philosophy of Science: Nietzsche and the Sciences II [Bab99a], pages 341–382. ISBN 90-481-5234-8, 94-017-2428-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL

https://link.springer.com/book/
10.1007/978-94-017-2428-9.

Anonymous:1999:BMd

[Ano99e] Anonymous. Back matter. In Hooykaas [Hoo99c], pages 373–462. ISBN 0-7923-5774-4, 94-015-9295-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 205.

                                              URL https://link.springer.com/
                                              book/10.1007/978-94-015-9295-6.

Anonymous:1999:BMc

[Ano99f] Anonymous. Back matter. In Fehér et al. [FKR99], pages 353–388. ISBN 90-481-5257-7, 94-015-9293-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL

                                              https://link.springer.com/book/
                                              10.1007/978-94-015-9293-2.

Anonymous:1999:BMg

[Ano99g] Anonymous. Back matter. In The Concept of Scientific Law in the Philosophy of Science and Epistemology: A Study of Theoretical Reason [Han99a], pages 201–230. ISBN 90-481-5275-5, 94-017-3265-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL

                                              https://link.springer.com/book/
                                              10.1007/978-94-017-3265-9.

Anonymous:1999:FMb

[Ano99h] Anonymous. Front matter. In The Natural Background of Meaning [Den99d], pages i–xi. ISBN 90-481-5126-0, 94-015-9084-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

                                              ???? URL https://link.springer.
                                              com/book/10.1007/978-94-015-
                                              9084-6.

Anonymous:1999:FMa

[Ano99i] Anonymous. Front matter. In Can that be Right?: Essays on Experiment, Evidence, and Science [Fra99c], pages i–ix. ISBN 94-010-6242-0, 94-011-5334-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
book/10.1007/978-94-011-5334-8.

Anonymous:1999:FMf

[Ano99j] Anonymous. Front matter. In Nietzsche, Theories of Knowledge, and Critical Theory: Nietzsche and the Sciences I [Bab99b], pages i–xx. ISBN 90-481-5233-X, 94-017-2430-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
book/10.1007/978-94-017-2430-2.

Anonymous:1999:FMe

[Ano99k] Anonymous. Front matter. In Nietzsche, Epistemology, and Philosophy of Science: Nietzsche and the Sciences II [Bab99a], pages i–xvii. ISBN 90-481-5234-8, 94-017-2428-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL

https://link.springer.com/book/
10.1007/978-94-017-2428-9.

Anonymous:1999:FMd

[Ano99l] Anonymous. Front matter. In Hooykaas [Hoo99c], pages i–xviii. ISBN 0-7923-5774-4, 94-015-9295-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 205.

URL https://link.springer.com/
book/10.1007/978-94-015-9295-6.

Anonymous:1999:FMc

[Ano99m] Anonymous. Front matter. In Fehér et al. [FKR99], pages i–xi. ISBN 90-481-5257-7, 94-015-9293-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL

https://link.springer.com/book/
10.1007/978-94-015-9293-2.

Anonymous:1999:FMg

[Ano99n] Anonymous. Front matter. In The Concept of Scientific Law in the Philosophy of Science and Epistemology: A Study of Theoretical Reason [Han99a], pages i–xiv. ISBN 90-481-5275-5, 94-017-3265-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL

                                              https://link.springer.com/book/
                                              10.1007/978-94-017-3265-9.

Anonymous:1999:FMh

[Ano99o] Anonymous. Front matter. In The Concept of Scientific Law in the Philosophy of Science and Epistemology: A Study of Theoretical Reason [Han99a], page 1. ISBN 90-481-5275-5, 94-017-3265-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

                                              URL https://link.springer.com/
                                              book/10.1007/978-94-017-3265-9.

Anonymous:1999:FMi

[Ano99p] Anonymous. Front matter. In The Concept of Scientific Law in the Philosophy of Science and Epistemology: A Study of Theoretical Reason [Han99a], page 31. ISBN 90-481-5275-5, 94-017-3265-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

                                              ???? URL https://link.springer.
                                              com/book/10.1007/978-94-017-
                                              3265-9.

Anonymous:1999:FMj

[Ano99q] Anonymous. Front matter. In The Concept of Scientific Law in the Philosophy of Science and Epistemology: A Study of Theoretical Reason [Han99a], page 61. ISBN 90-481-5275-5, 94-017-3265-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

                                              ???? URL https://link.springer.

com/book/10.1007/978-94-017-
3265-9.

Anonymous:1999:FMk

[Ano99r] Anonymous. Front matter. In The Concept of Scientific Law in the Philosophy of Science and Epistemology: A Study of Theoretical Reason [Han99a], page 155. ISBN 90-481-5275-5, 94-017-3265-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
book/10.1007/978-94-017-3265-9.

Anonymous:2000:BMc

[Ano00a] Anonymous. Back matter. In The Historical Development of Energetics [Hel00f], pages 405–420. ISBN 0-7923-5874-0, 94-011-4471-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 209.

URL https://link.springer.com/
book/10.1007/978-94-011-4471-1.

Anonymous:2000:BMb

[Ano00b] Anonymous. Back matter. In Knowledge, Language and Logic: Questions for Quine [OK00], pages 431–446. ISBN 1-4020-0253-X, 94-011-3933-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
book/10.1007/978-94-011-3933-5.

Anonymous:2000:BMf

[Ano00c] Anonymous. Back matter. In Maimonides and the Sciences [CL00], pages 245–256. ISBN 90-481-5348-4, 94-017-2128-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic).

LCCN ????  URL https://link.
springer.com/book/10.1007/978-
94-017-2128-8.

Anonymous:2000:BMd

[Ano00d] Anonymous. Back matter. In The Evolutionary-biology Papers of Elie Metchnikoff [GWT00e], pages 217–230. ISBN 0-7923-6067-2, 94-015-9381-7. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol.

212; QH366.2. URL http://www.loc.
gov/catdir/enhancements/fy0821/
99049680-d.html;   http://www.
loc.gov/catdir/enhancements/
fy0821/99049680-t.html;   https:
//link.springer.com/book/10.
1007/978-94-015-9381-6.

Anonymous:2000:BMa

[Ano00e] Anonymous. Back matter. In Science and Modernity: Toward an Integral Theory of Science [Lel00m], pages 281– 300. ISBN 1-4020-0247-5, 94-010-9036-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN BD581.

URL https://link.springer.com/
book/10.1007/978-94-010-9036-0.

Anonymous:2000:BMe

[Ano00f] Anonymous. Back matter. In The Reality of the Unobservable: Observability, Unobservability and Their Impact on the Issue of Scientific Realism [AP00b], pages 361–377. ISBN 90-481-5458-8, 94-015-9391-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL

https://link.springer.com/book/
10.1007/978-94-015-9391-5.

Anonymous:2000:FMh

[Ano00g] Anonymous. Front matter. In The Historical Development of Energetics [Hel00f], pages i–63. ISBN 0-7923-5874-0, 94-011-4471-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 209.

URL https://link.springer.com/
book/10.1007/978-94-011-4471-1.

Anonymous:2000:FMg

[Ano00h] Anonymous. Front matter. In Knowledge, Language and Logic: Questions for Quine [OK00], pages i–xi. ISBN 1-4020-0253-X, 94-011-3933-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL

https://link.springer.com/book/
10.1007/978-94-011-3933-5.

Anonymous:2000:FMk

[Ano00i] Anonymous. Front matter. In Maimonides and the Sciences [CL00], pages i–xv. ISBN 90-481-5348-4, 94-017-2128-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
book/10.1007/978-94-017-2128-8.

Anonymous:2000:FMi

[Ano00j] Anonymous. Front matter. In The Evolutionary-biology Papers of Elie Metchnikoff [GWT00e], pages i–x. ISBN 0-7923-6067-2, 94-015-9381-7. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol.

212; QH366.2. URL http://www.loc.
gov/catdir/enhancements/fy0821/
99049680-d.html;   http://www.
loc.gov/catdir/enhancements/
fy0821/99049680-t.html;   https:
//link.springer.com/book/10.
1007/978-94-015-9381-6.

Anonymous:2000:FMa

[Ano00k] Anonymous. Front matter. In Science and Modernity: Toward an Integral Theory of Science [Lel00m], pages i–xiii. ISBN 1-4020-0247-5, 94-010-9036-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN BD581.

URL https://link.springer.com/
book/10.1007/978-94-010-9036-0.

Anonymous:2000:FMb

[Ano00l] Anonymous. Front matter. In Science and Modernity: Toward an Integral Theory of Science [Lel00m], page 1. ISBN 1-4020-0247-5, 94-010-9036-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN BD581.

                                              URL https://link.springer.com/
                                              book/10.1007/978-94-010-9036-0.

Anonymous:2000:FMc

[Ano00m] Anonymous. Front matter. In Science and Modernity: Toward an Integral Theory of Science [Lel00m], page 47. ISBN 1-4020-0247-5, 94-010-9036-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN BD581.

                                              URL https://link.springer.com/
                                              book/10.1007/978-94-010-9036-0.

Anonymous:2000:FMd

[Ano00n] Anonymous. Front matter. In Science and Modernity: Toward an Integral Theory of Science [Lel00m], page 97. ISBN 1-4020-0247-5, 94-010-9036-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN BD581.

                                              URL https://link.springer.com/
                                              book/10.1007/978-94-010-9036-0.

Anonymous:2000:FMe

[Ano00o] Anonymous. Front matter. In Science and Modernity: Toward an Integral Theory of Science [Lel00m], page 171. ISBN 1-4020-0247-5, 94-010-9036-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN BD581.

                                              URL https://link.springer.com/
                                              book/10.1007/978-94-010-9036-0.

Anonymous:2000:FMf

[Ano00p] Anonymous. Front matter. In Science and Modernity: Toward an Integral Theory of Science [Lel00m], page 269. ISBN 1-4020-0247-5, 94-010-9036-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN BD581.

URL https://link.springer.com/
book/10.1007/978-94-010-9036-0.

Anonymous:2000:FMj

[Ano00q] Anonymous. Front matter. In The Reality of the Unobservable: Observability, Unobservability and Their Impact on the Issue of Scientific Realism [AP00b], pages i–vii. ISBN 90-481-5458-8, 94-015-9391-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL

https://link.springer.com/book/
10.1007/978-94-015-9391-5.

Anonymous:2001:BMc

[Ano01a] Anonymous. Back matter. In A history of the ideas of theoretical physics: essays on the nineteenth and twentieth century physics [D’A01l], pages 311– 381. ISBN 1-4020-0244-0, 94-010-9034-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC5.53; QC6.

Anonymous:2001:BMd

[Ano01b] Anonymous. Back matter. In Incommensurability and Related Matters [HHS01], pages 303–337. ISBN 90-481-5709-9, 94-015-9680-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL

https://link.springer.com/book/
10.1007/978-94-015-9680-0.

Anonymous:2001:BMh

[Ano01c] Anonymous. Back matter. In Colouring textiles: a history of natural dyestuffs in industrial Europe [NG01c], pages 205–246. ISBN 0-7923-7022-8, 94-017-1081-3. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN TP919

.N54 2001. URL http://www.loc.
gov/catdir/enhancements/fy0813/

2001033541-d.html;   http://www.
loc.gov/catdir/enhancements/
fy0813/2001033541-t.html;
https://link.springer.com/book/
10.1007/978-94-017-1081-7.

Anonymous:2001:BMe

[Ano01d] Anonymous. Back matter. In Blackmore et al. [BIT01c], pages 315–350. ISBN 90-481-5752-8, 94-015-9690-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
book/10.1007/978-94-015-9690-9.

Anonymous:2001:BMb

[Ano01e] Anonymous. Back matter. In Estonian Studies in the History and Philosophy of Science [Vih01b], pages 305–306. ISBN 94-010-3863-5, 94-010-0672-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
book/10.1007/978-94-010-0672-9.

Anonymous:2001:BMf

[Ano01f] Anonymous. Back matter. In Between Leibniz, Newton, and Kant: Philosophy and Science in the Eighteenth Century [Lef01a], pages 211–288. ISBN 90-481-5774-9, 94-015-9729-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
book/10.1007/978-94-015-9729-6.

Anonymous:2001:BMa

[Ano01g] Anonymous. Back matter. In The Reception of Darwinism in the Iberian World: Spain, Spanish America and Brazil [GPSR01], pages 263–283. ISBN 94-010-3885-6, 94-010-0602-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
book/10.1007/978-94-010-0602-6.

Anonymous:2001:BMg

[Ano01h] Anonymous. Back matter. In Klein [Kle01b], pages 253–264. ISBN 1-4020-0100-2, 94-015-9737-5. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol.

222; QD7. URL http://www.loc.
gov/catdir/enhancements/fy0814/
2001050224-d.html;   http://www.
loc.gov/catdir/enhancements/
fy0814/2001050224-t.html;
https://link.springer.com/book/
10.1007/978-94-015-9737-1.

Anonymous:2001:BMi

[Ano01i] Anonymous. Back matter. In Allen and MacLeod [AM01b], pages 369–399. ISBN 1-4020-0495-8 (hardcover), 94-017-2956-5. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 228; Q175.55. URL http://

www.loc.gov/catdir/enhancements/
fy0821/2002023597-d.html;
http://www.loc.gov/catdir/toc/
fy022/2002023597.html;   https:/
/link.springer.com/book/10.1007/
978-94-017-2956-7.

Anonymous:2001:FMg

[Ano01j] Anonymous. Front matter. In A history of the ideas of theoretical physics: essays on the nineteenth and twentieth century physics [D’A01l], pages i–xviii. ISBN 1-4020-0244-0, 94-010-9034-3 (ebook). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC5.53; QC6.

Anonymous:2001:FMh

[Ano01k] Anonymous. Front matter. In A history of the ideas of theoretical physics: essays on the nineteenth and twentieth century physics [D’A01l], pages 1– 6. ISBN 1-4020-0244-0, 94-010-9034-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC5.53; QC6.

Anonymous:2001:FMi

[Ano01l] Anonymous. Front matter. In A history of the ideas of theoretical physics: essays on the nineteenth and twentieth century physics [D’A01l], pages 107– 110. ISBN 1-4020-0244-0, 94-010-9034-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC5.53; QC6.

Anonymous:2001:FMj

[Ano01m] Anonymous. Front matter. In A history of the ideas of theoretical physics: essays on the nineteenth and twentieth century physics [D’A01l], pages 217– 222. ISBN 1-4020-0244-0, 94-010-9034-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC5.53; QC6.

Anonymous:2001:FMk

[Ano01n] Anonymous. Front matter. In Incommensurability and Related Matters [HHS01], pages i–xxxiv. ISBN 90-481-5709-9, 94-015-9680-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL

                                              https://link.springer.com/book/
                                              10.1007/978-94-015-9680-0.

Anonymous:2001:FMt

[Ano01o] Anonymous. Front matter. In Colouring textiles: a history of natural dyestuffs in industrial Europe [NG01c], pages i–xxv. ISBN 0-7923-7022-8, 94-017-1081-3. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN TP919

                                              .N54 2001. URL http://www.loc.
                                              gov/catdir/enhancements/fy0813/
                                              2001033541-d.html;   http://www.
                                              loc.gov/catdir/enhancements/
                                              fy0813/2001033541-t.html;
                                              https://link.springer.com/book/
                                              10.1007/978-94-017-1081-7.

Anonymous:2001:FMl

[Ano01p] Anonymous. Front matter. In Blackmore et al. [BIT01c], pages i–x. ISBN 90-481-5752-8, 94-015-9690-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL

https://link.springer.com/book/
10.1007/978-94-015-9690-9.

Anonymous:2001:FMb

[Ano01q] Anonymous. Front matter. In Estonian Studies in the History and Philosophy of Science [Vih01b], pages i–xiv. ISBN 94-010-3863-5, 94-010-0672-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
book/10.1007/978-94-010-0672-9.

Anonymous:2001:FMc

[Ano01r] Anonymous. Front matter. In Estonian Studies in the History and Philosophy of Science [Vih01b], page 1. ISBN 94-010-3863-5, 94-010-0672-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL

https://link.springer.com/book/
10.1007/978-94-010-0672-9.

Anonymous:2001:FMd

[Ano01s] Anonymous. Front matter. In Estonian Studies in the History and Philosophy of Science [Vih01b], page 89. ISBN 94-010-3863-5, 94-010-0672-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL

https://link.springer.com/book/
10.1007/978-94-010-0672-9.

Anonymous:2001:FMe

[Ano01t] Anonymous. Front matter. In Estonian Studies in the History and Philosophy of Science [Vih01b], page 201. ISBN 94-010-3863-5, 94-010-0672-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
book/10.1007/978-94-010-0672-9.

Anonymous:2001:FMf

[Ano01u] Anonymous. Front matter. In Estonian Studies in the History and Philosophy of Science [Vih01b], page 263. ISBN 94-010-3863-5, 94-010-0672-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
book/10.1007/978-94-010-0672-9.

Anonymous:2001:FMm

[Ano01v] Anonymous. Front matter. In Between Leibniz, Newton, and Kant: Philosophy and Science in the Eighteenth Century [Lef01a], pages I–XVI. ISBN 90-481-5774-9, 94-015-9729-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
book/10.1007/978-94-015-9729-6.

Anonymous:2001:FMn

[Ano01w] Anonymous. Front matter. In Between Leibniz, Newton, and Kant: Philosophy and Science in the Eighteenth Century [Lef01a], page 1. ISBN 90-481-5774-9, 94-015-9729-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL

https://link.springer.com/book/
10.1007/978-94-015-9729-6.

Anonymous:2001:FMo

[Ano01x] Anonymous. Front matter. In Between Leibniz, Newton, and Kant: Philosophy and Science in the Eighteenth Century [Lef01a], page 25. ISBN 90-481-5774-9, 94-015-9729-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL

https://link.springer.com/book/
10.1007/978-94-015-9729-6.

Anonymous:2001:FMp

[Ano01y] Anonymous. Front matter. In Between Leibniz, Newton, and Kant: Philosophy and Science in the Eighteenth Century [Lef01a], page 93. ISBN 90-481-5774-9, 94-015-9729-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL

https://link.springer.com/book/
10.1007/978-94-015-9729-6.

Anonymous:2001:FMq

[Ano01z] Anonymous. Front matter. In Between Leibniz, Newton, and Kant: Philosophy and Science in the Eighteenth Century [Lef01a], page 147. ISBN 90-481-5774-9, 94-015-9729-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
book/10.1007/978-94-015-9729-6.

Anonymous:2001:FMr

[Ano01-27] Anonymous. Front matter. In Between Leibniz, Newton, and Kant: Philosophy and Science in the Eighteenth Century [Lef01a], page 189. ISBN 90-481-5774-9, 94-015-9729-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
book/10.1007/978-94-015-9729-6.

Anonymous:2001:FMa

[Ano01-28] Anonymous. Front matter. In The Reception of Darwinism in the Iberian World: Spain, Spanish America and Brazil [GPSR01], pages i–xii. ISBN 94-010-3885-6, 94-010-0602-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL

https://link.springer.com/book/
10.1007/978-94-010-0602-6.

Anonymous:2001:FMs

[Ano01-29] Anonymous.  Front matter.  In

Klein [Kle01b], pages i–xv. ISBN 1-4020-0100-2, 94-015-9737-5. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol.

222; QD7. URL http://www.loc.
gov/catdir/enhancements/fy0814/
2001050224-d.html;   http://www.
loc.gov/catdir/enhancements/
fy0814/2001050224-t.html;
https://link.springer.com/book/
10.1007/978-94-015-9737-1.

Anonymous:2001:FMu

[Ano01-30] Anonymous. Front matter. In Allen and MacLeod [AM01b], pages i–x. ISBN 1-4020-0495-8 (hardcover), 94-017-2956-5. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 228; Q175.55. URL http://

www.loc.gov/catdir/enhancements/
fy0821/2002023597-d.html;
http://www.loc.gov/catdir/toc/
fy022/2002023597.html;   https:/
/link.springer.com/book/10.1007/
978-94-017-2956-7.

Anonymous:2001:FMv

[Ano01-31] Anonymous. Front matter. In Allen and MacLeod [AM01b], page 21. ISBN 1-4020-0495-8 (hardcover), 94-017-2956-5. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 228; Q175.55. URL http://

www.loc.gov/catdir/enhancements/
fy0821/2002023597-d.html;
http://www.loc.gov/catdir/toc/
fy022/2002023597.html;   https:/
/link.springer.com/book/10.1007/
978-94-017-2956-7.

Anonymous:2001:FMw

[Ano01-32] Anonymous. Front matter. In Allen and MacLeod [AM01b], page 103.

ISBN 1-4020-0495-8 (hardcover), 94-017-2956-5. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 228; Q175.55. URL http://

www.loc.gov/catdir/enhancements/
fy0821/2002023597-d.html;
http://www.loc.gov/catdir/toc/
fy022/2002023597.html;   https:/
/link.springer.com/book/10.1007/
978-94-017-2956-7.

Anonymous:2001:FMx

[Ano01-33] Anonymous. Front matter. In Allen and MacLeod [AM01b], page 215. ISBN 1-4020-0495-8 (hardcover), 94-017-2956-5. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 228; Q175.55. URL http://

www.loc.gov/catdir/enhancements/
fy0821/2002023597-d.html;
http://www.loc.gov/catdir/toc/
fy022/2002023597.html;   https:/
/link.springer.com/book/10.1007/
978-94-017-2956-7.

Anonymous:2001:FMy

[Ano01-34] Anonymous. Front matter. In Allen and MacLeod [AM01b], page 363. ISBN 1-4020-0495-8 (hardcover), 94-017-2956-5. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 228; Q175.55. URL http://

www.loc.gov/catdir/enhancements/
fy0821/2002023597-d.html;
http://www.loc.gov/catdir/toc/
fy022/2002023597.html;   https:/
/link.springer.com/book/10.1007/
978-94-017-2956-7.

Anonymous:2002:BMe

[Ano02a] Anonymous. Back matter. In Mixture and Chemical Combination: and Related Essays [Duh02m], pages 315–322. ISBN 90-481-5924-5, 94-017-2292-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

                                              URL https://link.springer.com/
                                              book/10.1007/978-94-017-2292-6.

Anonymous:2002:BMf

[Ano02b] Anonymous. Back matter. In What Was Mechanical About Mechanics: the Concept of Force Between Metaphysics and Mechanics from Newton to Lagrange [Bou02i], pages 241–280. ISBN 1-4020-0233-5, 94-017-3672-3. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 224; QC73.

                                              URL http://www.loc.gov/catdir/
                                              enhancements/fy0814/2001050645-
                                              b.html;   http://www.loc.gov/
                                              catdir/enhancements/fy0814/
                                              2001050645-t.html;   http://
                                              www.loc.gov/catdir/enhancements/
                                              fy0823/2001050645-d.html;
                                              https://link.springer.com/book/
                                              10.1007/978-94-017-3672-5.

Anonymous:2002:BMb

[Ano02c] Anonymous.  Back matter.  In

Hermeneutic Philosophy of Science, Van Gogh’s Eyes, and God: Essays in Honor of Patrick A. Heelan, S.J. [Bab02c], pages 461–495. ISBN 90-481-5926-1, 94-017-1767-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 225.

                                              URL https://link.springer.com/
                                              book/10.1007/978-94-017-1767-0.

Anonymous:2002:BMa

[Ano02d] Anonymous. Back matter. In E. A. Burtt, Historian and Philosopher: A Study of the Author of The Metaphysical Foundations of Modern Physical Science [Vil02g], pages 247–299. ISBN 90-481-5937-7, 94-017-1331-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL

                                              https://link.springer.com/book/
                                              10.1007/978-94-017-1331-3.

Anonymous:2002:BMc

[Ano02e] Anonymous. Back matter. In Knowledge and Language: Selected Essays of L. Jonathan Cohen [Coh02j], pages 323–324. ISBN 90-481-5955-5, 94-017-2020-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/book/10.1007/978-94-017-
2020-5.

Anonymous:2002:BMd

[Ano02f] Anonymous. Back matter. In Meanest Foundations and Nobler Superstructures: Hooke, Newton and the “Compounding of the Celestiall Motions of the Planetts” [Gal02c], pages 207–251. ISBN 90-481-6067-7, 94-017-2223-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QB351

.G35 2002.  URL https://link.
springer.com/book/10.1007/978-
94-017-2223-0.

Anonymous:2002:FMh

[Ano02g] Anonymous. Front matter. In Mixture and Chemical Combination: and Related Essays [Duh02m], pages i–2. ISBN 90-481-5924-5, 94-017-2292-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
book/10.1007/978-94-017-2292-6.

Anonymous:2002:FMi

[Ano02h] Anonymous. Front matter. In Mixture and Chemical Combination: and Related Essays [Duh02m], page 3. ISBN 90-481-5924-5, 94-017-2292-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
book/10.1007/978-94-017-2292-6.

Anonymous:2002:FMj

[Ano02i] Anonymous. Front matter. In Mixture and Chemical Combination: and Related Essays [Duh02m], page 29. ISBN 90-481-5924-5, 94-017-2292-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
book/10.1007/978-94-017-2292-6.

Anonymous:2002:FMk

[Ano02j] Anonymous. Front matter. In Mixture and Chemical Combination: and Related Essays [Duh02m], page 119. ISBN 90-481-5924-5, 94-017-2292-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
book/10.1007/978-94-017-2292-6.

Anonymous:2002:FMl

[Ano02k] Anonymous. Front matter. In What Was Mechanical About Mechanics: the Concept of Force Between Metaphysics and Mechanics from Newton to Lagrange [Bou02i], pages i–xvi. ISBN 1-4020-0233-5, 94-017-3672-3. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 224; QC73.

URL http://www.loc.gov/catdir/
enhancements/fy0814/2001050645-
b.html;   http://www.loc.gov/
catdir/enhancements/fy0814/
2001050645-t.html;   http://
www.loc.gov/catdir/enhancements/
fy0823/2001050645-d.html;
https://link.springer.com/book/
10.1007/978-94-017-3672-5.

Anonymous:2002:FMm

[Ano02l] Anonymous. Front matter. In What Was Mechanical About Mechanics: the Concept of Force Between Metaphysics and Mechanics from Newton to Lagrange [Bou02i], page 31. ISBN 1-4020-0233-5, 94-017-3672-3. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 224; QC73.

URL http://www.loc.gov/catdir/
enhancements/fy0814/2001050645-
b.html;   http://www.loc.gov/
catdir/enhancements/fy0814/
2001050645-t.html;   http://
www.loc.gov/catdir/enhancements/
fy0823/2001050645-d.html;
https://link.springer.com/book/
10.1007/978-94-017-3672-5.

Anonymous:2002:FMn

[Ano02m] Anonymous. Front matter. In What Was Mechanical About Mechanics: the Concept of Force Between Metaphysics and Mechanics from Newton to Lagrange [Bou02i], page 103. ISBN 1-4020-0233-5, 94-017-3672-3. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 224; QC73.

URL http://www.loc.gov/catdir/
enhancements/fy0814/2001050645-
b.html;   http://www.loc.gov/
catdir/enhancements/fy0814/
2001050645-t.html;   http://
www.loc.gov/catdir/enhancements/
fy0823/2001050645-d.html;
https://link.springer.com/book/
10.1007/978-94-017-3672-5.

Anonymous:2002:FMo

[Ano02n] Anonymous. Front matter. In What Was Mechanical About Mechanics: the Concept of Force Between Metaphysics and Mechanics from Newton to Lagrange [Bou02i], page 173. ISBN 1-4020-0233-5, 94-017-3672-3. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 224; QC73.

URL http://www.loc.gov/catdir/
enhancements/fy0814/2001050645-
b.html;   http://www.loc.gov/
catdir/enhancements/fy0814/

                                              2001050645-t.html;   http://
                                              www.loc.gov/catdir/enhancements/
                                              fy0823/2001050645-d.html;
                                              https://link.springer.com/book/
                                              10.1007/978-94-017-3672-5.

Anonymous:2002:FMb

[Ano02o] Anonymous.  Front matter.  In

Hermeneutic Philosophy of Science, Van Gogh’s Eyes, and God: Essays in Honor of Patrick A. Heelan, S.J. [Bab02c], pages i–xvii. ISBN 90-481-5926-1, 94-017-1767-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 225.

                                              URL https://link.springer.com/
                                              book/10.1007/978-94-017-1767-0.

Anonymous:2002:FMc

[Ano02p] Anonymous.  Front matter.  In

Hermeneutic Philosophy of Science, Van Gogh’s Eyes, and God: Essays in Honor of Patrick A. Heelan, S.J. [Bab02c], pages 19–23. ISBN 90-481-5926-1, 94-017-1767-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 225.

                                              URL https://link.springer.com/
                                              book/10.1007/978-94-017-1767-0.

Anonymous:2002:FMd

[Ano02q] Anonymous.  Front matter.  In

Hermeneutic Philosophy of Science, Van Gogh’s Eyes, and God: Essays in Honor of Patrick A. Heelan, S.J. [Bab02c], pages 231–232. ISBN 90-481-5926-1, 94-017-1767-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 225.

                                              URL https://link.springer.com/
                                              book/10.1007/978-94-017-1767-0.

Anonymous:2002:FMe

[Ano02r] Anonymous.  Front matter.  In

Hermeneutic Philosophy of Science, Van Gogh’s Eyes, and God: Essays in Honor of Patrick A. Heelan, S.J. [Bab02c], pages 335–338. ISBN 90-481-5926-1, 94-017-1767-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 225.

URL https://link.springer.com/
book/10.1007/978-94-017-1767-0.

Anonymous:2002:FMa

[Ano02s] Anonymous. Front matter. In E.

A. Burtt, Historian and Philosopher: A Study of the Author of The Metaphysical Foundations of Modern Physical Science [Vil02g], pages i–x. ISBN 90-481-5937-7, 94-017-1331-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
book/10.1007/978-94-017-1331-3.

Anonymous:2002:FMf

[Ano02t] Anonymous. Front matter. In Knowledge and Language: Selected Essays of L. Jonathan Cohen [Coh02j], pages i–xxviii. ISBN 90-481-5955-5, 94-017-2020-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/book/10.1007/978-94-017-
2020-5.

Anonymous:2002:FMg

[Ano02u] Anonymous. Front matter. In Meanest Foundations and Nobler Superstructures: Hooke, Newton and the “Compounding of the Celestiall Motions of the Planetts” [Gal02c], pages i–xiii. ISBN 90-481-6067-7, 94-017-2223-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QB351 .G35 2002.

URL https://link.springer.com/
book/10.1007/978-94-017-2223-0.

Anonymous:2003:BMd

[Ano03a] Anonymous. Back matter. In The Social Origins of Modern Science [Zil03], pages 225–267. ISBN 0-7923-6457-0, 94-011-4142-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol.

                                              200; Q175.5. URL http://www.loc.
                                              gov/catdir/enhancements/fy0813/
                                              00042657-d.html;   http://www.
                                              loc.gov/catdir/enhancements/
                                              fy0813/00042657-t.html;   https:
                                              //link.springer.com/book/10.
                                              1007/978-94-011-4142-0.  Edited

59 years after the author’s death by Diederick Raven, Wolfgang Krohn, and R. S. (Robert Sonné) Cohen.

Anonymous:2003:BMc

[Ano03b] Anonymous. Back matter. In Rescuing Reason: A Critique of Anti-Rationalist Views of Science and Knowledge [Nol03k], pages 543–571. ISBN 1-4020-1043-5, 94-010-0289-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 230; Q175.32.R45. URL https://link.

                                              springer.com/book/10.1007/978-
                                              94-010-0289-9.

Anonymous:2003:BMh

[Ano03c] Anonymous. Back matter. In Science and Culture [Aga03s], pages 404–416. ISBN 90-481-6234-3, 94-017-2946-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174.

                                              URL https://link.springer.com/
                                              book/10.1007/978-94-017-2946-8.

Anonymous:2003:BMi

[Ano03d] Anonymous. Back matter. In Simões et al. [SCD03b], pages 345–365. ISBN 1-4020-1259-4, 94-017-3584-0. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 v. 233. URL

https://link.springer.com/book/
10.1007/978-94-017-3584-1.

Anonymous:2003:BMb

[Ano03e] Anonymous. Back matter. In Renn et al. [RDS03], pages 641–649. ISBN 1-4020-1284-5, 1-4020-1285-3 (paperback). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol.

234. URL http://www.gbv.de/dms/
bowker/toc/9781402012846.pdf;
http://www.loc.gov/catdir/
enhancements/fy0813/2003046757-
d.html;   http://www.loc.gov/
catdir/enhancements/fy0813/
2003046757-t.html;   https://
link.springer.com/book/10.1007/
978-94-010-0111-3.

Anonymous:2003:BMg

[Ano03f] Anonymous. Back matter. In Feyerabend and Scientific Values: Tightrope-Walking Rationality [Far03b], pages 237–259. ISBN 90-481-6309-9, 94-017-1542-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .F377 2003. URL https://link.

springer.com/book/10.1007/978-
94-017-1542-3.

Anonymous:2003:BMe

[Ano03g] Anonymous. Back matter. In Bulgarian Studies in the Philosophy of Science [Gin03a], pages 207–221. ISBN 1-4020-1496-1, 94-017-0961-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
book/10.1007/978-94-017-0961-3.

Anonymous:2003:BMa

[Ano03h] Anonymous.  Back matter.  In

Descartes’s Mathematical Thought [Sas03d], pages 339–356. ISBN 90-481-6487-7, 94-017-1225-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL

https://link.springer.com/book/
10.1007/978-94-017-1225-5.

Anonymous:2003:BMf

[Ano03i]  Anonymous.  Back matter.  In

Descartes’s Mathematical Thought [Sas03d], pages 439–496. ISBN 90-481-6487-7, 94-017-1225-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL

https://link.springer.com/book/
10.1007/978-94-017-1225-5.

Anonymous:2003:FMl

[Ano03j] Anonymous. Front matter. In The Social Origins of Modern Science [Zil03], pages i–lx. ISBN 0-7923-6457-0, 94-011-4142-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 200;

Q175.5.  URL http://www.loc.
gov/catdir/enhancements/fy0813/
00042657-d.html;   http://www.
loc.gov/catdir/enhancements/
fy0813/00042657-t.html;   https:
//link.springer.com/book/10.
1007/978-94-011-4142-0.  Edited

59 years after the author’s death by Diederick Raven, Wolfgang Krohn, and R. S. (Robert Sonné) Cohen.

Anonymous:2003:FMm

[Ano03k] Anonymous. Front matter. In The Social Origins of Modern Science [Zil03], page 1. ISBN 0-7923-6457-0, 94-011-4142-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 200;

Q175.5.  URL http://www.loc.
gov/catdir/enhancements/fy0813/
00042657-d.html;   http://www.
loc.gov/catdir/enhancements/
fy0813/00042657-t.html;   https:
//link.springer.com/book/10.
1007/978-94-011-4142-0.  Edited

59 years after the author’s death by Diederick Raven, Wolfgang Krohn, and R. S. (Robert Sonné) Cohen.

Anonymous:2003:FMn

[Ano03l] Anonymous. Front matter. In The Social Origins of Modern Science [Zil03], page 169. ISBN 0-7923-6457-0, 94-011-4142-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 200;

Q175.5.  URL http://www.loc.
gov/catdir/enhancements/fy0813/
00042657-d.html;   http://www.
loc.gov/catdir/enhancements/
fy0813/00042657-t.html;   https:
//link.springer.com/book/10.
1007/978-94-011-4142-0.  Edited

59 years after the author’s death by Diederick Raven, Wolfgang Krohn, and R. S. (Robert Sonné) Cohen.

Anonymous:2003:FMg

[Ano03m] Anonymous.  Front matter.  In

Rescuing Reason: A Critique of Anti-Rationalist Views of Science and Knowledge [Nol03k], pages i–xi. ISBN 1-4020-1043-5, 94-010-0289-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 230; Q175.32.R45. URL

https://link.springer.com/book/
10.1007/978-94-010-0289-9.

Anonymous:2003:FMh

[Ano03n] Anonymous.  Front matter.  In

Rescuing Reason: A Critique of Anti-Rationalist Views of Science and Knowledge [Nol03k], pages 13–17. ISBN 1-4020-1043-5, 94-010-0289-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 230; Q175.32.R45. URL

https://link.springer.com/book/
10.1007/978-94-010-0289-9.

Anonymous:2003:FMi

[Ano03o] Anonymous.  Front matter.  In

Rescuing Reason: A Critique of Anti-Rationalist Views of Science and Knowledge [Nol03k], pages 173–178. ISBN 1-4020-1043-5, 94-010-0289-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 230; Q175.32.R45. URL

                                              https://link.springer.com/book/
                                              10.1007/978-94-010-0289-9.

Anonymous:2003:FMj

[Ano03p] Anonymous.  Front matter.  In

Rescuing Reason: A Critique of Anti-Rationalist Views of Science and Knowledge [Nol03k], pages 361–364. ISBN 1-4020-1043-5, 94-010-0289-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 230; Q175.32.R45. URL

                                              https://link.springer.com/book/
                                              10.1007/978-94-010-0289-9.

Anonymous:2003:FMk

[Ano03q] Anonymous.  Front matter.  In

Rescuing Reason: A Critique of Anti-Rationalist Views of Science and Knowledge [Nol03k], pages 461–464. ISBN 1-4020-1043-5, 94-010-0289-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 230; Q175.32.R45. URL

                                              https://link.springer.com/book/
                                              10.1007/978-94-010-0289-9.

Anonymous:2003:FMy

[Ano03r] Anonymous. Front matter. In Science and Culture [Aga03s], pages i–xxxii. ISBN 90-481-6234-3, 94-017-2946-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174.

                                              URL https://link.springer.com/
                                              book/10.1007/978-94-017-2946-8.

Anonymous:2003:FMz

[Ano03s] Anonymous. Front matter. In Science and Culture [Aga03s], page 1. ISBN 90-481-6234-3, 94-017-2946-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174. URL

https://link.springer.com/book/
10.1007/978-94-017-2946-8.

Anonymous:2003:FMaa

[Ano03t] Anonymous. Front matter. In Science and Culture [Aga03s], page 81. ISBN 90-481-6234-3, 94-017-2946-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174. URL

https://link.springer.com/book/
10.1007/978-94-017-2946-8.

Anonymous:2003:FMab

[Ano03u] Anonymous. Front matter. In Science and Culture [Aga03s], page 179. ISBN 90-481-6234-3, 94-017-2946-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174. URL

https://link.springer.com/book/
10.1007/978-94-017-2946-8.

Anonymous:2003:FMac

[Ano03v] Anonymous. Front matter. In Science and Culture [Aga03s], page 249. ISBN 90-481-6234-3, 94-017-2946-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174. URL

https://link.springer.com/book/
10.1007/978-94-017-2946-8.

Anonymous:2003:FMad

[Ano03w] Anonymous. Front matter. In Science and Culture [Aga03s], page 327. ISBN 90-481-6234-3, 94-017-2946-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174. URL

https://link.springer.com/book/
10.1007/978-94-017-2946-8.

Anonymous:2003:FMae

[Ano03x] Anonymous. Front matter. In Simões et al. [SCD03b], pages i–xiv. ISBN 1-4020-1259-4, 94-017-3584-0. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 v. 233. URL

https://link.springer.com/book/
10.1007/978-94-017-3584-1.

Anonymous:2003:FMb

[Ano03y] Anonymous. Front matter. In Renn et al. [RDS03], pages i–xxii. ISBN 1-4020-1284-5, 1-4020-1285-3 (paperback). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol.

234. URL http://www.gbv.de/dms/
bowker/toc/9781402012846.pdf;
 http://www.loc.gov/catdir/
enhancements/fy0813/2003046757-
d.html;   http://www.loc.gov/
catdir/enhancements/fy0813/
2003046757-t.html;   https://
link.springer.com/book/10.1007/
978-94-010-0111-3.

Anonymous:2003:FMc

[Ano03z] Anonymous. Front matter. In Renn et al. [RDS03], page 1. ISBN 1-4020-1284-5, 1-4020-1285-3 (paperback). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol.

234. URL http://www.gbv.de/dms/
bowker/toc/9781402012846.pdf;
 http://www.loc.gov/catdir/
enhancements/fy0813/2003046757-
d.html;   http://www.loc.gov/
catdir/enhancements/fy0813/
2003046757-t.html;   https://
link.springer.com/book/10.1007/
978-94-010-0111-3.

Anonymous:2003:FMd

[Ano03-27] Anonymous. Front matter. In Renn et al. [RDS03], page 191. ISBN 1-4020-1284-5, 1-4020-1285-3 (paperback).

ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol.

234. URL http://www.gbv.de/dms/
bowker/toc/9781402012846.pdf;
 http://www.loc.gov/catdir/
enhancements/fy0813/2003046757-
d.html;   http://www.loc.gov/
catdir/enhancements/fy0813/
2003046757-t.html;   https://
link.springer.com/book/10.1007/
978-94-010-0111-3.

Anonymous:2003:FMe

[Ano03-28] Anonymous. Front matter. In Renn et al. [RDS03], page 437. ISBN 1-4020-1284-5, 1-4020-1285-3 (paperback). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol.

234. URL http://www.gbv.de/dms/
bowker/toc/9781402012846.pdf;
 http://www.loc.gov/catdir/
enhancements/fy0813/2003046757-
d.html;   http://www.loc.gov/
catdir/enhancements/fy0813/
2003046757-t.html;   https://
link.springer.com/book/10.1007/
978-94-010-0111-3.

Anonymous:2003:FMf

[Ano03-29] Anonymous. Front matter. In Renn et al. [RDS03], page 543. ISBN 1-4020-1284-5, 1-4020-1285-3 (paperback). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol.

234. URL http://www.gbv.de/dms/
bowker/toc/9781402012846.pdf;
 http://www.loc.gov/catdir/
enhancements/fy0813/2003046757-
d.html;   http://www.loc.gov/
catdir/enhancements/fy0813/
2003046757-t.html;   https://
link.springer.com/book/10.1007/
978-94-010-0111-3.

Anonymous:2003:FMx

[Ano03-30] Anonymous. Front matter. In Feyerabend and Scientific Values: Tightrope-Walking Rationality [Far03b], pages i–x. ISBN 90-481-6309-9, 94-017-1542-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174

                                              .F377 2003. URL https://link.
                                              springer.com/book/10.1007/978-
                                              94-017-1542-3.

Anonymous:2003:FMo

[Ano03-31] Anonymous. Front matter. In Bulgarian Studies in the Philosophy of Science [Gin03a], pages i–xii. ISBN 1-4020-1496-1, 94-017-0961-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL

                                              https://link.springer.com/book/
                                              10.1007/978-94-017-0961-3.

Anonymous:2003:FMp

[Ano03-32] Anonymous. Front matter. In Bulgarian Studies in the Philosophy of Science [Gin03a], page 1. ISBN 1-4020-1496-1, 94-017-0961-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL

                                              https://link.springer.com/book/
                                              10.1007/978-94-017-0961-3.

Anonymous:2003:FMq

[Ano03-33] Anonymous. Front matter. In Bulgarian Studies in the Philosophy of Science [Gin03a], page 51. ISBN 1-4020-1496-1, 94-017-0961-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL

                                              https://link.springer.com/book/
                                              10.1007/978-94-017-0961-3.

Anonymous:2003:FMr

[Ano03-34] Anonymous. Front matter. In Bulgarian Studies in the Philosophy of Science [Gin03a], page 83. ISBN 1-4020-1496-1, 94-017-0961-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL

https://link.springer.com/book/
10.1007/978-94-017-0961-3.

Anonymous:2003:FMs

[Ano03-35] Anonymous. Front matter. In Bulgarian Studies in the Philosophy of Science [Gin03a], page 127. ISBN 1-4020-1496-1, 94-017-0961-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL

https://link.springer.com/book/
10.1007/978-94-017-0961-3.

Anonymous:2003:FMt

[Ano03-36] Anonymous. Front matter. In Bulgarian Studies in the Philosophy of Science [Gin03a], page 167. ISBN 1-4020-1496-1, 94-017-0961-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL

https://link.springer.com/book/
10.1007/978-94-017-0961-3.

Anonymous:2003:FMa

[Ano03-37] Anonymous.  Front matter.  In

Descartes’s Mathematical Thought [Sas03d], pages i–x. ISBN 90-481-6487-7, 94-017-1225-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL

https://link.springer.com/book/
10.1007/978-94-017-1225-5.

Anonymous:2003:FMu

[Ano03-38] Anonymous.  Front matter.  In

Descartes’s Mathematical Thought [Sas03d], pages i–xiv. ISBN 90-481-6487-7, 94-017-1225-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL

https://link.springer.com/book/
10.1007/978-94-017-1225-5.

Anonymous:2003:FMv

[Ano03-39] Anonymous.  Front matter.  In

Descartes’s Mathematical Thought [Sas03d], page 11. ISBN 90-481-6487-7, 94-017-1225-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic).

LCCN ????  URL https://link.
springer.com/book/10.1007/978-
94-017-1225-5.

Anonymous:2003:FMw

[Ano03-40] Anonymous.  Front matter.  In

Descartes’s Mathematical Thought [Sas03d], page 287. ISBN 90-481-6487-7, 94-017-1225-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL

https://link.springer.com/book/
10.1007/978-94-017-1225-5.

Anonymous:2004:BMa

[Ano04a] Anonymous. Back matter. In The Reception of the Galilean Science of Motion in Seventeenth-Century Europe [PT04], pages 261–287. ISBN 1-4020-2454-1, 1-4020-2455-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC133 .R43 2004.

URL https://link.springer.com/
book/10.1007/978-1-4020-2455-9.

Anonymous:2004:BMb

[Ano04b] Anonymous. Back matter. In Classics in the History of Greek Mathematics [Chr04a], pages 463–474. ISBN 1-4020-0081-2 (hardcover), 1-4020-2640-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QA22

.C596 2004. URL https://link.
springer.com/book/10.1007/978-1-
4020-2640-9.

Anonymous:2004:FMa

[Ano04c] Anonymous. Front matter. In The Reception of the Galilean Science of Motion in Seventeenth-Century Europe [PT04], pages i–ix. ISBN 1-4020-2454-1, 1-4020-2455-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC133 .R43 2004.

URL https://link.springer.com/
book/10.1007/978-1-4020-2455-9.

Anonymous:2004:FMb

[Ano04d] Anonymous. Front matter. In Classics in the History of Greek Mathematics [Chr04a], pages i–xi. ISBN 1-4020-0081-2 (hardcover), 1-4020-2640-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QA22 .C596 2004.

URL https://link.springer.com/
book/10.1007/978-1-4020-2640-9.

Anonymous:2004:FMc

[Ano04e] Anonymous. Front matter. In Classics in the History of Greek Mathematics [Chr04a], page 1. ISBN 1-4020-0081-2 (hardcover), 1-4020-2640-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QA22 .C596 2004.

URL https://link.springer.com/
book/10.1007/978-1-4020-2640-9.

Anonymous:2004:FMd

[Ano04f] Anonymous. Front matter. In Classics in the History of Greek Mathematics [Chr04a], page 111. ISBN 1-4020-0081-2 (hardcover), 1-4020-2640-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QA22 .C596 2004.

URL https://link.springer.com/
book/10.1007/978-1-4020-2640-9.

Anonymous:2004:FMe

[Ano04g] Anonymous. Front matter. In Classics in the History of Greek Mathematics [Chr04a], page 185. ISBN 1-4020-0081-2 (hardcover), 1-4020-2640-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QA22 .C596 2004.

URL https://link.springer.com/
book/10.1007/978-1-4020-2640-9.

Anonymous:2004:FMf

[Ano04h] Anonymous. Front matter. In Classics in the History of Greek Mathematics [Chr04a], page 255. ISBN 1-4020-0081-2 (hardcover), 1-4020-2640-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QA22 .C596 2004.

                                              URL https://link.springer.com/
                                              book/10.1007/978-1-4020-2640-9.

Anonymous:2004:FMg

[Ano04i] Anonymous. Front matter. In Classics in the History of Greek Mathematics [Chr04a], page 329. ISBN 1-4020-0081-2 (hardcover), 1-4020-2640-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QA22 .C596 2004.

                                              URL https://link.springer.com/
                                              book/10.1007/978-1-4020-2640-9.

Anonymous:2004:FMh

[Ano04j] Anonymous. Front matter. In Classics in the History of Greek Mathematics [Chr04a], page 381. ISBN 1-4020-0081-2 (hardcover), 1-4020-2640-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QA22 .C596 2004.

                                              URL https://link.springer.com/
                                              book/10.1007/978-1-4020-2640-9.

Anonymous:2005:BMa

[Ano05a] Anonymous.  Back matter.  In

History of Science, History of Text [Che05a], pages 247–266. ISBN 1-281-33589-4, 1-4020-2320-0 (hardcover), 1-4020-2321-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 238

                                              Q125. URL https://link.springer.
                                              com/book/10.1007/1-4020-2321-9.

Anonymous:2005:BMb

[Ano05b] Anonymous.  Back matter.  In

Creativity, Psychology and the History of Science [GB05], pages 475– 544. ISBN 1-4020-2455-X, 1-4020-3491-1 (hardcover), 1-4020-3509-8 (ebook). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN BF408 .G793

2005. URL https://link.springer.
com/book/10.1007/1-4020-3509-8.

Anonymous:2005:CSS

[Ano05c] Anonymous. The case study that started it all: Charles Darwin. In Creativity, Psychology and the History of Science [GB05], chapter 4, pages 105– 191. ISBN 1-4020-2455-X, 1-4020-3491-1 (hardcover), 1-4020-3509-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN BF408 .G793 2005.

URL https://link.springer.com/
chapter/10.1007/1-4020-3509-8_4.

Anonymous:2005:CE

[Ano05d] Anonymous. Coping with the extraordinary. In Creativity, Psychology and the History of Science [GB05], chapter 8, pages 365–422. ISBN 1-4020-2455-X, 1-4020-3491-1 (hardcover), 1-4020-3509-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN BF408 .G793 2005.

URL https://link.springer.com/
chapter/10.1007/1-4020-3509-8_8.

Anonymous:2005:CPW

[Ano05e] Anonymous. The creative person as a whole. In Creativity, Psychology and the History of Science [GB05], chapter 3, pages 35–104. ISBN 1-4020-2455-X, 1-4020-3491-1 (hardcover), 1-4020-3509-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN BF408 .G793 2005.

URL https://link.springer.com/
chapter/10.1007/1-4020-3509-8_3.

Anonymous:2005:CMD

[Ano05f] Anonymous. Creativity in the moral domain. In Creativity, Psychology and the History of Science [GB05], chapter 9, pages 423–449. ISBN 1-4020-2455-X, 1-4020-3491-1 (hardcover), 1-4020-3509-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN BF408 .G793 2005.

URL https://link.springer.com/
chapter/10.1007/1-4020-3509-8_9.

Anonymous:2005:FCP

[Ano05g] Anonymous. Facets of the creative process: Insight, point of view and repetition. In Creativity, Psychology and the History of Science [GB05], chapter 5, pages 193–230. ISBN 1-4020-2455-X, 1-4020-3491-1 (hardcover), 1-4020-3509-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN BF408 .G793 2005.

URL https://link.springer.com/
chapter/10.1007/1-4020-3509-8_5.

Anonymous:2005:FMa

[Ano05h] Anonymous. Front matter. In History of Science, History of Text [Che05a], pages i–xxvii. ISBN 1-281-33589-4, 1-4020-2320-0 (hardcover), 1-4020-2321-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 238 Q125. URL

https://link.springer.com/book/
10.1007/1-4020-2321-9.

Anonymous:2005:FMb

[Ano05i] Anonymous. Front matter. In Turkish Studies in the History and Philosophy of Science [IG05c], pages i–xvii. ISBN 1-4020-3332-X (hardcover), 1-4020-3333-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 244. URL

http://d-nb.info/1079906622/34;
 https://link.springer.com/book/
10.1007/1-4020-3333-8.

Anonymous:2005:FMc

[Ano05j]  Anonymous.  Front matter.  In

Creativity, Psychology and the History of Science [GB05], pages i– viii. ISBN 1-4020-2455-X, 1-4020-3491-1 (hardcover), 1-4020-3509-8 (ebook). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN BF408 .G793

2005. URL https://link.springer.
com/book/10.1007/1-4020-3509-8.

Anonymous:2005:I

[Ano05k] Anonymous. Introduction. In Creativity, Psychology and the History of Science [GB05], chapter 1, pages 1–17. ISBN 1-4020-2455-X, 1-4020-3491-1 (hardcover), 1-4020-3509-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN BF408 .G793 2005.

URL https://link.springer.com/
chapter/10.1007/1-4020-3509-8_1.

Anonymous:2005:ILP

[Ano05l] Anonymous. Introduction: A life with a purpose. In Creativity, Psychology and the History of Science [GB05], chapter 2, pages 19–33. ISBN 1-4020-2455-X, 1-4020-3491-1 (hardcover), 1-4020-3509-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN BF408 .G793 2005.

URL https://link.springer.com/
chapter/10.1007/1-4020-3509-8_2.

Anonymous:2005:M

[Ano05m] Anonymous. Modalities. In Creativity, Psychology and the History of Science [GB05], chapter 6, pages 231–291. ISBN 1-4020-2455-X, 1-4020-3491-1 (hardcover), 1-4020-3509-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN BF408 .G793 2005.

URL https://link.springer.com/
chapter/10.1007/1-4020-3509-8_6.

Anonymous:2005:PFC

[Ano05n] Anonymous. Peace and further conditions for human welfare. In Creativity, Psychology and the History of Science [GB05], chapter 10, pages 451–474. ISBN 1-4020-2455-X, 1-4020-3491-1 (hardcover), 1-4020-3509-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN BF408 .G793 2005.

                                              URL https://link.springer.com/
                                              chapter/10.1007/1-4020-3509-8_10.

Anonymous:2005:TOC

[Ano05o] Anonymous. Tracking the ordinary course of development. In Creativity, Psychology and the History of Science [GB05], chapter 7, pages 293–363. ISBN 1-4020-2455-X, 1-4020-3491-1 (hardcover), 1-4020-3509-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN BF408 .G793 2005.

                                              URL https://link.springer.com/
                                              chapter/10.1007/1-4020-3509-8_7.

Anonymous:2006:AVA

[Ano06a] Anonymous. Advaita-Vedānta: ‘Śariraka-BhāS.ya’. In Analogy in Indian and Western Philosophical Thought [GC06], chapter 10, pages 176–179. ISBN 1-4020-3339-7 (hardcover), 1-4020-3340-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN BD190 .Z52 2006; Q175.A1 B65 v. 243.

                                              URL https://link.springer.com/
                                              chapter/10.1007/1-4020-3340-0_10.

Anonymous:2006:ANN

[Ano06b] Anonymous. Analogy in Navya– Nyāya. In Analogy in Indian and Western Philosophical Thought [GC06], chapter 4, pages 56–109. ISBN 1-4020-3339-7 (hardcover), 1-4020-3340-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN BD190 .Z52 2006; Q175.A1 B65 v. 243.

URL https://link.springer.com/
chapter/10.1007/1-4020-3340-0_4.

Anonymous:2006:AWP

[Ano06c] Anonymous. Analogy in Western philosophy and Indian approaches to analogy: Introduction. In Analogy in Indian and Western Philosophical Thought [GC06], chapter 3, pages 45– 55. ISBN 1-4020-3339-7 (hardcover), 1-4020-3340-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN BD190 .Z52 2006; Q175.A1 B65 v. 243.

URL https://link.springer.com/
chapter/10.1007/1-4020-3340-0_3.

Anonymous:2006:BMf

[Ano06d] Anonymous. Back matter. In The Metaphysics of Science: An Account of Modern Science in Terms of Principles, Laws and Theories [Dil06f], pages 209–333. ISBN 1-4020-6327-X (paperback), 1-4020-3837-2 (hardcover), 1-4020-3838-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175 .D634195 2006. URL

https://link.springer.com/book/
10.1007/978-1-4020-3838-9.

Anonymous:2006:BMa

[Ano06e] Anonymous. Back matter. In Philosophy of Chemistry: Synthesis of a New Discipline [BSM06b], pages 363–374. ISBN 1-4020-2455-X, 1-4020-3256-0, 1-4020-3261-7. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 242

QD6. URL https://link.springer.
com/book/10.1007/1-4020-3261-7.

Anonymous:2006:BMb

[Ano06f] Anonymous. Back matter. In Analogy in Indian and Western Philosophical Thought [GC06], pages 228–285.

ISBN 1-4020-3339-7 (hardcover), 1-4020-3340-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN BD190 .Z52 2006; Q175.A1 B65 v.

243. URL https://link.springer.
com/book/10.1007/1-4020-3340-0.

Anonymous:2006:BMd

[Ano06g] Anonymous. Back matter. In The Beginnings of Piezoelectricity: A Study in Mundane Physics [Kat06a], pages 227–273. ISBN 1-4020-4669-3 (hardcover), 1-4020-4670-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC595.5 .K37 2006.

URL https://link.springer.com/
book/10.1007/978-1-4020-4670-4.

Anonymous:2006:BMe

[Ano06h] Anonymous. Back matter. In Hendricks et al. [HJLP06], pages 321–326. ISBN 1-4020-5194-8 (hardcover), 1-4020-5195-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175 .B69 v.251 2006; Q175 .I51377 2006.

URL http://www.loc.gov/catdir/
enhancements/fy0825/2007425756-
d.html;   http://www.loc.gov/
catdir/toc/fy0713/2007425756.
html;   https://link.springer.
com/book/10.1007/978-1-4020-
5195-1.

Anonymous:2006:MUR

[Ano06i] Anonymous. Mīmām .sā /on certainty of perception in Mīmām .sā. In Analogy in Indian and Western Philosophical Thought [GC06], chapter 9, pages 170– 175. ISBN 1-4020-3339-7 (hardcover), 1-4020-3340-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN BD190 .Z52 2006; Q175.A1 B65 v. 243.

URL https://link.springer.com/
chapter/10.1007/1-4020-3340-0_9.

Anonymous:2006:ELS

[Ano06j] Anonymous. Empirical laws: The supervention of experience. In The Metaphysics of Science: An Account of Modern Science in Terms of Principles, Laws and Theories [Dil06f], chapter 3, pages 73–94. ISBN 1-4020-6327-X (paperback), 1-4020-3837-2 (hardcover), 1-4020-3838-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175 .D634195

2006. URL https://link.springer.
com/chapter/10.1007/978-1-4020-
3838-9_3.

Anonymous:2006:FMm

[Ano06k] Anonymous. Front matter. In The Metaphysics of Science: An Account of Modern Science in Terms of Principles, Laws and Theories [Dil06f], pages i–x. ISBN 1-4020-6327-X (paperback), 1-4020-3837-2 (hardcover), 1-4020-3838-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175 .D634195 2006. URL

https://link.springer.com/book/
10.1007/978-1-4020-3838-9.

Anonymous:2006:FMj

[Ano06l] Anonymous. Front matter. In The Metaphysics of Science: An Account of Modern Science in Terms of Principles, Laws and Theories [Dil06f], pages i–xxiv. ISBN 1-4020-6327-X (paperback), 1-4020-3837-2 (hardcover), 1-4020-3838-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175 .D634195 2006. URL

https://link.springer.com/book/
10.1007/978-1-4020-3838-9.

Anonymous:2006:FMa

[Ano06m] Anonymous.  Front matter.  In

Philosophy of Chemistry: Synthesis of a New Discipline [BSM06b], pages i–vii. ISBN 1-4020-2455-X, 1-4020-3256-0, 1-4020-3261-7. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 242

                                              QD6. URL https://link.springer.
                                              com/book/10.1007/1-4020-3261-7.

Anonymous:2006:FMb

[Ano06n] Anonymous.  Front matter.  In

Philosophy of Chemistry: Synthesis of a New Discipline [BSM06b], page 1. ISBN 1-4020-2455-X, 1-4020-3256-0, 1-4020-3261-7. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 242

                                              QD6. URL https://link.springer.
                                              com/book/10.1007/1-4020-3261-7.

Anonymous:2006:FMc

[Ano06o] Anonymous.  Front matter.  In

Philosophy of Chemistry: Synthesis of a New Discipline [BSM06b], page 41. ISBN 1-4020-2455-X, 1-4020-3256-0, 1-4020-3261-7. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 242

                                              QD6. URL https://link.springer.
                                              com/book/10.1007/1-4020-3261-7.

Anonymous:2006:FMd

[Ano06p] Anonymous.  Front matter.  In

Philosophy of Chemistry: Synthesis of a New Discipline [BSM06b], page 93. ISBN 1-4020-2455-X, 1-4020-3256-0, 1-4020-3261-7. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 242

                                              QD6. URL https://link.springer.
                                              com/book/10.1007/1-4020-3261-7.

Anonymous:2006:FMe

[Ano06q] Anonymous.  Front matter.  In

Philosophy of Chemistry: Synthesis of a New Discipline [BSM06b], page 171. ISBN 1-4020-2455-X, 1-4020-3256-0, 1-4020-3261-7. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 242

QD6. URL https://link.springer.
com/book/10.1007/1-4020-3261-7.

Anonymous:2006:FMf

[Ano06r] Anonymous.  Front matter.  In

Philosophy of Chemistry: Synthesis of a New Discipline [BSM06b], page 205. ISBN 1-4020-2455-X, 1-4020-3256-0, 1-4020-3261-7. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 242

QD6. URL https://link.springer.
com/book/10.1007/1-4020-3261-7.

Anonymous:2006:FMg

[Ano06s] Anonymous.  Front matter.  In

Philosophy of Chemistry: Synthesis of a New Discipline [BSM06b], page 271. ISBN 1-4020-2455-X, 1-4020-3256-0, 1-4020-3261-7. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 242

QD6. URL https://link.springer.
com/book/10.1007/1-4020-3261-7.

Anonymous:2006:FMh

[Ano06t] Anonymous.  Front matter.  In

Philosophy of Chemistry: Synthesis of a New Discipline [BSM06b], page 325. ISBN 1-4020-2455-X, 1-4020-3256-0, 1-4020-3261-7. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 242

QD6. URL https://link.springer.
com/book/10.1007/1-4020-3261-7.

Anonymous:2006:FMi

[Ano06u] Anonymous. Front matter. In Analogy in Indian and Western Philosophical Thought [GC06], pages I–VIII. ISBN 1-4020-3339-7 (hardcover), 1-4020-3340-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN BD190 .Z52 2006; Q175.A1 B65 v.

243. URL https://link.springer.
com/book/10.1007/1-4020-3340-0.

Anonymous:2006:FMk

[Ano06v] Anonymous. Front matter. In The Beginnings of Piezoelectricity: A Study in Mundane Physics [Kat06a], pages I–XIII. ISBN 1-4020-4669-3 (hardcover), 1-4020-4670-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC595.5 .K37 2006.

URL https://link.springer.com/
book/10.1007/978-1-4020-4670-4.

Anonymous:2006:FMl

[Ano06w] Anonymous. Front matter. In Hendricks et al. [HJLP06], pages I–XIV. ISBN 1-4020-5194-8 (hardcover), 1-4020-5195-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175 .B69 v.251 2006; Q175

.I51377 2006. URL http://www.loc.
gov/catdir/enhancements/fy0825/
2007425756-d.html;   http://
www.loc.gov/catdir/toc/fy0713/
2007425756.html;   https://link.
springer.com/book/10.1007/978-1-
4020-5195-1.

Anonymous:2006:FRP

[Ano06x] Anonymous. Fundamental and refined principles: The core of modern science. In The Metaphysics of Science: An Account of Modern Science in Terms of Principles, Laws and Theories [Dil06f], chapter 2, pages 49–72. ISBN 1-4020-6327-X (paperback), 1-4020-3837-2 (hardcover), 1-4020-3838-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175 .D634195

2006. URL https://link.springer.
com/chapter/10.1007/978-1-4020-
3838-9_2.

Anonymous:2006:HTC

[Ano06y] Anonymous. The Hellenic type of cultural tradition. In Analogy in Indian and Western Philosophical Thought [GC06], chapter 14, pages 215–221. ISBN 1-4020-3339-7 (hardcover), 1-4020-3340-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN BD190 .Z52 2006; Q175.A1 B65 v. 243.

URL https://link.springer.com/
chapter/10.1007/1-4020-3340-0_14.

Anonymous:2006:HIL

[Ano06z] Anonymous. History of Indian logic. In Analogy in Indian and Western Philosophical Thought [GC06], chapter 5, pages 110–120. ISBN 1-4020-3339-7 (hardcover), 1-4020-3340-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN BD190 .Z52 2006; Q175.A1 B65 v. 243. URL https:

//link.springer.com/chapter/10.
1007/1-4020-3340-0_5.

Anonymous:2006:ITC

[Ano06-27] Anonymous. The Indian type of cultural tradition. In Analogy in Indian and Western Philosophical Thought [GC06], chapter 6, pages 121–142. ISBN 1-4020-3339-7 (hardcover), 1-4020-3340-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN BD190 .Z52 2006; Q175.A1 B65 v. 243.

URL https://link.springer.com/
chapter/10.1007/1-4020-3340-0_6.

Anonymous:2006:RMT

[Ano06-28] Anonymous. Revelation of the mechanism of tradition in a form of grammatical paradigms of Indian logic. In Analogy in Indian and Western Philosophical Thought [GC06], chapter 7, pages 143–164. ISBN 1-4020-3339-7 (hardcover), 1-4020-3340-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN BD190 .Z52 2006; Q175.A1 B65 v. 243.

URL https://link.springer.com/
chapter/10.1007/1-4020-3340-0_7.

Anonymous:2006:STC

[Ano06-29] Anonymous. Scientific theories: Closing the circle. In The Metaphysics of Science: An Account of Modern Science in Terms of Principles, Laws and Theories [Dil06f], chapter 4, pages 95–108. ISBN 1-4020-6327-X (paperback), 1-4020-3837-2 (hardcover), 1-4020-3838-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175 .D634195 2006. URL https://link.

springer.com/chapter/10.1007/
978-1-4020-3838-9_4.

Anonymous:2006:TMS

[Ano06-30] Anonymous. The teaching of Ṡaṅkarā on intuition and the organization of philosophical texts in order to perceive the transcendental. In Analogy in Indian and Western Philosophical Thought [GC06], chapter 8, pages 165– 169. ISBN 1-4020-3339-7 (hardcover), 1-4020-3340-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN BD190 .Z52 2006; Q175.A1 B65 v. 243.

URL https://link.springer.com/
chapter/10.1007/1-4020-3340-0_8.

Anonymous:2006:TIM

[Ano06-31] Anonymous. Tradition of the idea of man. In Analogy in Indian and Western Philosophical Thought [GC06], chapter 13, pages 204–214. ISBN 1-4020-3339-7 (hardcover), 1-4020-3340-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN BD190 .Z52 2006; Q175.A1 B65 v. 243.

URL https://link.springer.com/
chapter/10.1007/1-4020-3340-0_13.

Anonymous:2006:USN

[Ano06-32] Anonymous. Upadeśa–Sahasri (note 1). In Analogy in Indian and Western Philosophical Thought [GC06], chapter 11, pages 180–187. ISBN 1-4020-3339-7 (hardcover), 1-4020-3340-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN BD190 .Z52 2006; Q175.A1 B65 v. 243.

URL https://link.springer.com/
chapter/10.1007/1-4020-3340-0_11.

Anonymous:2006:WTC

[Ano06-33] Anonymous. The Western type of cultural tradition. In Analogy in Indian and Western Philosophical Thought [GC06], chapter 15, pages 222–227. ISBN 1-4020-3339-7 (hardcover), 1-4020-3340-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN BD190 .Z52 2006; Q175.A1 B65 v. 243.

URL https://link.springer.com/
chapter/10.1007/1-4020-3340-0_15.

Anonymous:2006:WT

[Ano06-34] Anonymous. Writing and tradition.

In Analogy in Indian and Western Philosophical Thought [GC06], chapter 12, pages 188–203. ISBN 1-4020-3339-7 (hardcover), 1-4020-3340-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN BD190 .Z52 2006; Q175.A1 B65 v. 243.

URL https://link.springer.com/
chapter/10.1007/1-4020-3340-0_12.

Anonymous:2007:ARN

[Ano07a] Anonymous. An astronomical road to a new theory of gravitation. In Renn et al. [RSS^{+}07a], pages 153–154. ISBN 1-4020-3999-9. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 no. 250 v. 3; QC173.6 .G469 2007.

Anonymous:2007:BMa

[Ano07b] Anonymous. Back matter. In Hans Christian Ørsted and The Romantic Legacy In Science: Ideas, Disciplines, Practices [BCK07], pages 433–442. ISBN 1-4020-2979-9 (hardcover), 1-4020-2987-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC515.O37 H36 2007. URL

https://link.springer.com/book/
10.1007/978-1-4020-2987-5.

Anonymous:2007:BMb

[Ano07c] Anonymous. Back matter. In Rethinking Explanation [PY07], pages 205–219. ISBN 1-4020-5580-3 (paperback), 1-4020-5581-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN BD237 .R48 2007; Q175.32.E97 R48 2007. URL

https://link.springer.com/book/
10.1007/978-1-4020-5581-2.

Anonymous:2007:EZN

[Ano07d] Anonymous. Einstein’s Zurich Notebook: Transcription and facsimile. In The Genesis of General Relativity: Volume 1. Einstein’s Zurich Notebook: Introduction and Source [JNR^{+}07b], pages 313–487. ISBN 1-4020-3999-9 (hardcover), 1-4020-4000-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 no. 250 v. 1; QC173.6 .G469 2007. URL https:

//link.springer.com/chapter/10.
1007/978-1-4020-4000-9_5.

Anonymous:2007:FTG

[Ano07e] Anonymous. A field theory of gravitation in the framework of Special Relativity. In Renn et al. [RSS^{+}07a], pages 411–412. ISBN 1-4020-3999-9. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 no. 250 v. 3; QC173.6 .G469 2007.

Anonymous:2007:ETM

[Ano07f] Anonymous. From an electromagnetic theory of matter to a new theory of gravitation. In Renn et al. [RSS^{+}07b], pages 621–622. ISBN 1-4020-3999-9 (hardcover), 1-4020-4000-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 no. 250 v. 4; QC173.6

.G469 2007. URL http://d-nb.info/
985765933/34;   http://www.loc.
gov/catdir/enhancements/fy0824/
2007440325-d.html;   http://www.
loc.gov/catdir/enhancements/
fy0824/2007440325-t.html;
https://link.springer.com/book/
10.1007/978-1-4020-4000-9.

Anonymous:2007:HMN

[Ano07g] Anonymous. From heretical mechanics to a new theory of relativity. In Renn et al. [RSS^{+}07a], pages 567–568. ISBN 1-4020-3999-9. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 no. 250 v. 3; QC173.6 .G469 2007.

Anonymous:2007:PMN

[Ano07h] Anonymous. From peripheral mathematics to a new theory of gravitation. In Renn et al. [RSS^{+}07b], pages 1039– 1040. ISBN 1-4020-3999-9 (hardcover), 1-4020-4000-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 no. 250 v. 4; QC173.6

.G469 2007. URL http://d-nb.info/
985765933/34;   http://www.loc.
gov/catdir/enhancements/fy0824/
2007440325-d.html;   http://www.
loc.gov/catdir/enhancements/
fy0824/2007440325-t.html;
https://link.springer.com/book/
10.1007/978-1-4020-4000-9.

Anonymous:2007:FMb

[Ano07i] Anonymous. Front matter. In Hans Christian Ørsted and The Romantic Legacy In Science: Ideas, Disciplines, Practices [BCK07], pages i–xix. ISBN 1-4020-2979-9 (hardcover), 1-4020-2987-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC515.O37 H36 2007. URL

                                              https://link.springer.com/book/
                                              10.1007/978-1-4020-2987-5.

Anonymous:2007:FMa

[Ano07j] Anonymous. Front matter. In Positioning the history of science [RG07], pages I–VII. ISBN 1-4020-5419-X (hardcover), 1-4020-5420-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q125 .P77 2007.

Anonymous:2007:FMd

[Ano07k] Anonymous. Front matter. In Rethinking Explanation [PY07], pages I–XIV. ISBN 1-4020-5580-3 (paperback), 1-4020-5581-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN BD237 .R48 2007; Q175.32.E97 R48 2007. URL

                                              https://link.springer.com/book/
                                              10.1007/978-1-4020-5581-2.

Anonymous:2007:FMe

[Ano07l] Anonymous. Front matter. In Rethinking Explanation [PY07], pages I–XIV. ISBN 1-4020-5580-3 (paperback), 1-4020-5581-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN BD237 .R48 2007; Q175.32.E97 R48 2007. URL

                                              https://link.springer.com/book/
                                              10.1007/978-1-4020-5581-2.

Anonymous:2007:FMf

[Ano07m] Anonymous. Front matter. In Rethinking Explanation [PY07], pages I–XIV. ISBN 1-4020-5580-3 (paperback), 1-4020-5581-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN BD237 .R48 2007; Q175.32.E97 R48 2007. URL

https://link.springer.com/book/
10.1007/978-1-4020-5581-2.

Anonymous:2007:GFB

[Ano07n] Anonymous. The gravitational force between mechanics and electrodynamics. In Renn et al. [RSS^{+}07a], pages 19–20. ISBN 1-4020-3999-9. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 no. 250 v. 3; QC173.6 .G469 2007.

Anonymous:2007:IGU

[Ano07o] Anonymous. Including gravitation in a unified theory of physics. In Renn et al. [RSS^{+}07b], pages 757–758. ISBN 1-4020-3999-9 (hardcover), 1-4020-4000-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 no. 250 v. 4; QC173.6

.G469 2007. URL http://d-nb.info/
985765933/34;   http://www.loc.
gov/catdir/enhancements/fy0824/
2007440325-d.html;   http://www.
loc.gov/catdir/enhancements/
fy0824/2007440325-t.html;
https://link.springer.com/book/
10.1007/978-1-4020-4000-9.

Anonymous:2007:IVa

[Ano07p] Anonymous. Index: Volumes 1 and 2.

In The Genesis of General Relativity: Volume 2. Einstein’s Zurich Notebook: Commentary and Essays [JRN^{+}07b], pages 927–938. ISBN 1-4020-3942-5. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 no. 250 v. 2; QC173.6 .G469 2007.

Anonymous:2007:IVb

[Ano07q] Anonymous. Index: Volumes 3 and 4. In Renn et al. [RSS^{+}07b], pages 1131– 1152. ISBN 1-4020-3999-9 (hardcover), 1-4020-4000-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 no. 250 v. 4; QC173.6

                                              .G469 2007. URL http://d-nb.info/
                                              985765933/34;   http://www.loc.
                                              gov/catdir/enhancements/fy0824/
                                              2007440325-d.html;   http://www.
                                              loc.gov/catdir/enhancements/
                                              fy0824/2007440325-t.html;
                                              https://link.springer.com/book/
                                              10.1007/978-1-4020-4000-9.

Anonymous:2007:NLG

[Ano07r] Anonymous. A new law of gravitation enforced by Special Relativity. In Renn et al. [RSS^{+}07a], pages 191–192. ISBN 1-4020-3999-9. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 no. 250 v. 3; QC173.6 .G469 2007.

Anonymous:2007:PGC

[Ano07s] Anonymous. The problem of gravitation as a challenge for the Minkowski formalism. In Renn et al. [RSS^{+}07a], pages 303–304. ISBN 1-4020-3999-9. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 no. 250 v. 3; QC173.6 .G469 2007.

Anonymous:2008:BMb

[Ano08a] Anonymous. Back matter. In Science and its History: A Reassessment of the Historiography of Science [Aga08d], pages 501–514. ISBN 1-4020-5631-1 (hardcover), 1-4020-5632-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174.8 .A33 2008.

                                              URL https://link.springer.com/
                                              book/10.1007/978-1-4020-5632-1.

Anonymous:2008:BMc

[Ano08b] Anonymous. Back matter. In Mechanics and Natural Philosophy Before the Scientific Revolution [LR08], pages 275–306. ISBN 1-4020-5966-3 (hardcover), 1-4020-5967-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QA802 .M43 2008. URL

https://link.springer.com/book/
10.1007/978-1-4020-5967-4.

Anonymous:2008:BMd

[Ano08c] Anonymous. Back matter. In Rethinking Scientific Change and Theory Comparison: Stabilities, Ruptures, Incommensurabilities? [SSHH08], pages 385–395. ISBN 1-4020-6274-5 (hardcover), 1-4020-6279-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175 .R397 2008.

URL https://link.springer.com/
book/10.1007/978-1-4020-6279-7.

Anonymous:2008:BMh

[Ano08d] Anonymous. Back matter. In Leibniz’s Metaphysics of Time and Space [Fut08], pages 195–205. ISBN 1-4020-8236-3 (hardcover), 1-4020-8237-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN B2599.S7 F88 2008.

URL https://link.springer.com/
book/10.1007/978-1-4020-8237-5.

Anonymous:2008:BMe

[Ano08e] Anonymous. Back matter. In Schmuhl [Sch08d], pages 419–467. ISBN 1-4020-6599-X (hardcover), 1-4020-6600-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN HQ755.5.G3 S37 2008. URL https:/

/link.springer.com/book/10.1007/
978-1-4020-6600-9.

Anonymous:2008:BMf

[Ano08f] Anonymous. Back matter. In Heim [Hei08f], pages 203–222. ISBN 1-4020-6717-8 (hardcover), 1-4020-6718-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN SB123.25.G3 H44913 2008. URL https://link.

springer.com/book/10.1007/978-1-
4020-6718-1.

Anonymous:2008:BMg

[Ano08g] Anonymous. Back matter. In Underdetermination: An Essay on Evidence and the Limits of Natural Knowledge [Bon08], pages 259–284. ISBN 1-4020-6898-0, 1-4020-6899-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175 .B72235 2008.

URL https://link.springer.com/
book/10.1007/978-1-4020-6899-7.

Anonymous:2008:BMa

[Ano08h] Anonymous. Back matter. In The English Galileo: Thomas Harriot’s work on motion as an example of preclassical mechanics. Volume 1: Interpretation [Sch08a], pages 243–388. ISBN 1-4020-5498-X (set: cased), 1-4020-5499-8 (set: e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC133 .S36 2008; Q174 .B67 v.268. URL

https://link.springer.com/book/
10.1007/978-1-4020-5499-0.

Anonymous:2008:BT

[Ano08i] Anonymous. Boundary transgressions. In Schmuhl [Sch08d], pages 409–417. ISBN 1-4020-6599-X (hardcover), 1-4020-6600-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN HQ755.5.G3 S37 2008. URL https:

//link.springer.com/chapter/10.
1007/978-1-4020-6600-9_6.

Anonymous:2008:BST

[Ano08j] Anonymous. The bounds of space and time. In Leibniz’s Metaphysics of Time and Space [Fut08], chapter 4, pages 81–103. ISBN 1-4020-8236-3 (hardcover), 1-4020-8237-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN B2599.S7 F88 2008.

URL https://link.springer.com/
chapter/10.1007/978-1-4020-8237-
5_4.

Anonymous:2008:BHP

[Ano08k] Anonymous. A brief history of the philosophy of space and time. In Leibniz’s Metaphysics of Time and Space [Fut08], chapter 1, pages 5–28. ISBN 1-4020-8236-3 (hardcover), 1-4020-8237-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN B2599.S7 F88 2008. URL https:/

/link.springer.com/chapter/10.
1007/978-1-4020-8237-5_1.

Anonymous:2008:CTA

[Ano08l] Anonymous. Causal and temporal asymmetry. In Leibniz’s Metaphysics of Time and Space [Fut08], chapter 5, pages 105–127. ISBN 1-4020-8236-3 (hardcover), 1-4020-8237-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN B2599.S7 F88 2008.

URL https://link.springer.com/
chapter/10.1007/978-1-4020-8237-
5_5.

Anonymous:2008:CT

[Ano08m] Anonymous. Competing truths. In Underdetermination: An Essay on Evidence and the Limits of Natural Knowledge [Bon08], chapter 4, pages 141–175. ISBN 1-4020-6898-0, 1-4020-6899-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175 .B72235 2008. URL https:/

/link.springer.com/chapter/10.
1007/978-1-4020-6899-7_4.

Anonymous:2008:EIP

[Ano08n] Anonymous. Exploration of the inclined-plane conception of projectile motion. In The English Galileo: Thomas Harriot’s work on motion as an example of preclassical mechanics. Volume 1: Interpretation [Sch08a], chapter 8, pages 175–227. ISBN 1-4020-5498-X (set: cased), 1-4020-5499-8 (set: e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC133 .S36 2008; Q174 .B67 v.268. URL https:

//link.springer.com/chapter/10.
1007/978-1-4020-5499-0_8.

Anonymous:2008:FBK

[Ano08o] Anonymous. The “Faustian Bargain”: The Kaiser Wilhelm Institute for Anthropology, Human Heredity and Eugenics in the National Socialist Era, 1933–1938/1942. In Schmuhl [Sch08d], pages 117–240. ISBN 1-4020-6599-X (hardcover), 1-4020-6600-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN HQ755.5.G3 S37

2008. URL https://link.springer.
com/chapter/10.1007/978-1-4020-
6600-9_4.

Anonymous:2008:FFEb

[Ano08p] Anonymous. Free fall experiments and their evaluation: Motion in a medium. In The English Galileo: Thomas Harriot’s work on motion as an example of preclassical mechanics. Volume 1: Interpretation [Sch08a], chapter 6, pages 133–151. ISBN 1-4020-5498-X (set: cased), 1-4020-5499-8 (set: e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC133 .S36 2008; Q174 .B67 v.268. URL https://link.

springer.com/chapter/10.1007/
978-1-4020-5499-0_6.

Anonymous:2008:FFEa

[Ano08q] Anonymous. Free fall experiments and their evaluation: The “weight” of falling bodies. In The English Galileo: Thomas Harriot’s work on motion as an example of preclassical mechanics. Volume 1: Interpretation [Sch08a], chapter 5, pages 97–131. ISBN 1-4020-5498-X (set: cased), 1-4020-5499-8 (set: e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC133 .S36 2008; Q174 .B67 v.268. URL https:

//link.springer.com/chapter/10.
1007/978-1-4020-5499-0_5.

Anonymous:2008:FMb

[Ano08r] Anonymous. Front matter. In Science and its History: A Reassessment of the Historiography of Science [Aga08d], pages I–XIX. ISBN 1-4020-5631-1 (hardcover), 1-4020-5632-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174.8 .A33 2008.

URL https://link.springer.com/
book/10.1007/978-1-4020-5632-1.

Anonymous:2008:FMc

[Ano08s] Anonymous. Front matter. In Mechanics and Natural Philosophy Before the Scientific Revolution [LR08], pages I–11. ISBN 1-4020-5966-3 (hardcover), 1-4020-5967-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QA802 .M43 2008. URL

https://link.springer.com/book/
10.1007/978-1-4020-5967-4.

Anonymous:2008:FMd

[Ano08t] Anonymous. Front matter. In Mechanics and Natural Philosophy Before the Scientific Revolution [LR08], page 12. ISBN 1-4020-5966-3 (hardcover), 1-4020-5967-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QA802 .M43 2008. URL

                                              https://link.springer.com/book/
                                              10.1007/978-1-4020-5967-4.

Anonymous:2008:FMe

[Ano08u] Anonymous. Front matter. In Mechanics and Natural Philosophy Before the Scientific Revolution [LR08], page 146. ISBN 1-4020-5966-3 (hardcover), 1-4020-5967-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QA802 .M43 2008. URL

                                              https://link.springer.com/book/
                                              10.1007/978-1-4020-5967-4.

Anonymous:2008:FMf

[Ano08v] Anonymous. Front matter. In Mechanics and Natural Philosophy Before the Scientific Revolution [LR08], page 221. ISBN 1-4020-5966-3 (hardcover), 1-4020-5967-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QA802 .M43 2008. URL

                                              https://link.springer.com/book/
                                              10.1007/978-1-4020-5967-4.

Anonymous:2008:FMg

[Ano08w] Anonymous. Front matter. In Rethinking Scientific Change and Theory Comparison: Stabilities, Ruptures, Incommensurabilities? [SSHH08], pages i–x. ISBN 1-4020-6274-5 (hardcover), 1-4020-6279-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175 .R397 2008. URL

                                              https://link.springer.com/book/
                                              10.1007/978-1-4020-6279-7.

Anonymous:2008:FMk

[Ano08x] Anonymous. Front matter. In Leibniz’s Metaphysics of Time and Space [Fut08], pages i–xiv. ISBN 1-4020-8236-3 (hardcover), 1-4020-8237-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN B2599.S7 F88 2008.

URL https://link.springer.com/
book/10.1007/978-1-4020-8237-5.

Anonymous:2008:FMh

[Ano08y] Anonymous.  Front matter.  In

Schmuhl [Sch08d], pages i–xiv. ISBN 1-4020-6599-X (hardcover), 1-4020-6600-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN HQ755.5.G3 S37 2008. URL https:/

/link.springer.com/book/10.1007/
978-1-4020-6600-9.

Anonymous:2008:FMi

[Ano08z] Anonymous. Front matter. In Heim [Hei08f], pages I–X. ISBN 1-4020-6717-8 (hardcover), 1-4020-6718-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN SB123.25.G3 H44913 2008. URL https://link.

springer.com/book/10.1007/978-1-
4020-6718-1.

Anonymous:2008:FMj

[Ano08-27] Anonymous. Front matter. In Underdetermination: An Essay on Evidence and the Limits of Natural Knowledge [Bon08], pages i–xii. ISBN 1-4020-6898-0, 1-4020-6899-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175 .B72235 2008.

URL https://link.springer.com/
book/10.1007/978-1-4020-6899-7.

Anonymous:2008:FMa

[Ano08-28] Anonymous. Front matter. In The English Galileo: Thomas Harriot’s work on motion as an example of preclassical mechanics. Volume 1: Interpretation [Sch08a], pages i–xxx. ISBN 1-4020-5498-X (set: cased), 1-4020-5499-8 (set: e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC133 .S36 2008; Q174 .B67 v.268. URL

https://link.springer.com/book/
10.1007/978-1-4020-5499-0.

Anonymous:2008:GCP

[Ano08-29] Anonymous. Graphical construction of projectile trajectories. In The English Galileo: Thomas Harriot’s work on motion as an example of preclassical mechanics. Volume 1: Interpretation [Sch08a], chapter 7, pages 153–174. ISBN 1-4020-5498-X (set: cased), 1-4020-5499-8 (set: e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC133 .S36 2008; Q174 .B67 v.268. URL https://link.

springer.com/chapter/10.1007/
978-1-4020-5499-0_7.

Anonymous:2008:HCP

[Ano08-30] Anonymous. Harriot and the challenge of projectile motion. In The English Galileo: Thomas Harriot’s work on motion as an example of preclassical mechanics. Volume 1: Interpretation [Sch08a], chapter 2, pages 25–37. ISBN 1-4020-5498-X (set: cased), 1-4020-5499-8 (set: e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC133 .S36 2008; Q174 .B67 v.268.

URL https://link.springer.com/
chapter/10.1007/978-1-4020-5499-
0_2.

Anonymous:2008:HP

[Ano08-31] Anonymous. A Humean predicament? In Underdetermination: An Essay on Evidence and the Limits of Natural Knowledge [Bon08], chapter 1, pages 1– 44. ISBN 1-4020-6898-0, 1-4020-6899-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175 .B72235 2008. URL https://link.

springer.com/chapter/10.1007/
978-1-4020-6899-7_1.

Anonymous:2008:ROK

[Ano08-32] Anonymous. In the realm of opportunity: The Kaiser Wilhelm Institute for Anthropology, Human Heredity and Eugenics during World War II, 1938/42–1945. In Schmuhl [Sch08d], pages 241–407. ISBN 1-4020-6599-X (hardcover), 1-4020-6600-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN HQ755.5.G3 S37

2008. URL https://link.springer.
com/chapter/10.1007/978-1-4020-
6600-9_5.

Anonymous:2008:I

[Ano08-33] Anonymous. Introduction. In Schmuhl [Sch08d], pages 1–6. ISBN 1-4020-6599-X (hardcover), 1-4020-6600-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN HQ755.5.G3 S37

2008. URL https://link.springer.
com/chapter/10.1007/978-1-4020-
6600-9_1.

Anonymous:2008:IHM

[Ano08-34] Anonymous. An introduction to Harriot’s manuscripts on motion. In The English Galileo: Thomas Harriot’s work on motion as an example of preclassical mechanics. Volume 1: Interpretation [Sch08a], chapter 3, pages 39–50. ISBN 1-4020-5498-X (set: cased), 1-4020-5499-8 (set: e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC133 .S36 2008; Q174 .B67 v.268.

URL https://link.springer.com/
chapter/10.1007/978-1-4020-5499-
0_3.

Anonymous:2008:LPP

[Ano08-35] Anonymous. Leibniz on past, present, and future. In Leibniz’s Metaphysics of Time and Space [Fut08], chapter 6, pages 129–142. ISBN 1-4020-8236-3 (hardcover), 1-4020-8237-1 (e-book).

ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN B2599.S7 F88 2008.

URL https://link.springer.com/
chapter/10.1007/978-1-4020-8237-
5_6.

Anonymous:2008:LR

[Ano08-36] Anonymous. Leibniz’s reductionism. In Leibniz’s Metaphysics of Time and Space [Fut08], chapter 2, pages 29–56. ISBN 1-4020-8236-3 (hardcover), 1-4020-8237-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN B2599.S7 F88 2008. URL https:/

/link.springer.com/chapter/10.
1007/978-1-4020-8237-5_2.

Anonymous:2008:MAM

[Ano08-37] Anonymous. Mathematical analysis of the motion of fall. In The English Galileo: Thomas Harriot’s work on motion as an example of preclassical mechanics. Volume 1: Interpretation [Sch08a], chapter 4, pages 53–95. ISBN 1-4020-5498-X (set: cased), 1-4020-5499-8 (set: e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC133 .S36 2008; Q174 .B67 v.268.

URL https://link.springer.com/
chapter/10.1007/978-1-4020-5499-
0_4.

Anonymous:2008:NC

[Ano08-38] Anonymous. Notes and conventions. In The English Galileo: Thomas Harriot’s work on motion as an example of preclassical mechanics. Volume 2: Sources [Sch08b], chapter 10, pages 391–393. ISBN 1-4020-5498-X (set: cased), 1-4020-5499-8 (set: e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC133 .S36 2008; Q174 .B67 v.268. URL https://link.

springer.com/chapter/10.1007/
978-1-4020-5499-0_10.

Anonymous:2008:PR

[Ano08-39] Anonymous. Problems of representation. In Underdetermination: An Essay on Evidence and the Limits of Natural Knowledge [Bon08], chapter 5, pages 177–205. ISBN 1-4020-6898-0, 1-4020-6899-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175 .B72235 2008. URL https:/

/link.springer.com/chapter/10.
1007/978-1-4020-6899-7_5.

Anonymous:2008:PTI

[Ano08-40] Anonymous. A “purely theoretical institute for the study of the nature of man”: The founding of the Kaiser Wilhelm Institute for Anthropology, Human Heredity, and Eugenics, 1920– 1927. In Schmuhl [Sch08d], pages 7–39. ISBN 1-4020-6599-X (hardcover), 1-4020-6600-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN HQ755.5.G3 S37 2008. URL https:

//link.springer.com/chapter/10.
1007/978-1-4020-6600-9_2.

Anonymous:2008:RME

[Ano08-41] Anonymous. Rationality, method, and evidence. In Underdetermination: An Essay on Evidence and the Limits of Natural Knowledge [Bon08], chapter 3, pages 89–139. ISBN 1-4020-6898-0, 1-4020-6899-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175 .B72235 2008. URL https:/

/link.springer.com/chapter/10.
1007/978-1-4020-6899-7_3.

Anonymous:2008:SKA

[Ano08-42] Anonymous. Shared knowledge and alternative solutions in Harriot’s and Galileo’s work on motion. In The English Galileo: Thomas Harriot’s work on motion as an example of preclassical mechanics. Volume 1: Interpretation [Sch08a], chapter 9, pages 231–241. ISBN 1-4020-5498-X (set: cased), 1-4020-5499-8 (set: e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC133 .S36 2008; Q174 .B67 v.268. URL https://link.

springer.com/chapter/10.1007/
978-1-4020-5499-0_9.

Anonymous:2008:STG

[Ano08-43] Anonymous. Space, time, and God. In Leibniz’s Metaphysics of Time and Space [Fut08], chapter 8, pages 171–194. ISBN 1-4020-8236-3 (hardcover), 1-4020-8237-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN B2599.S7 F88 2008.

URL https://link.springer.com/
chapter/10.1007/978-1-4020-8237-
5_8.

Anonymous:2008:STH

[Ano08-44] Anonymous. Space, time, and harmony. In Leibniz’s Metaphysics of Time and Space [Fut08], chapter 7, pages 143–170. ISBN 1-4020-8236-3 (hardcover), 1-4020-8237-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN B2599.S7 F88 2008.

URL https://link.springer.com/
chapter/10.1007/978-1-4020-8237-
5_7.

Anonymous:2008:HFU

[Ano08-45] Anonymous. “The human of the future under the scrutiny of research”: The Kaiser Wilhelm Institute for Anthropology, Human Heredity and Eugenics in the Weimar Republic, 1927–1933. In Schmuhl [Sch08d], pages 41–116. ISBN 1-4020-6599-X (hardcover), 1-4020-6600-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN HQ755.5.G3 S37 2008. URL https:

//link.springer.com/chapter/10.
1007/978-1-4020-6600-9_3.

Anonymous:2008:THP

[Ano08-46] Anonymous. Thomas Harriot: Practical mathematician and pioneer of modern science. In The English Galileo: Thomas Harriot’s work on motion as an example of preclassical mechanics. Volume 1: Interpretation [Sch08a], chapter 1, pages 15–23. ISBN 1-4020-5498-X (set: cased), 1-4020-5499-8 (set: e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC133 .S36 2008; Q174 .B67 v.268. URL https:

//link.springer.com/chapter/10.
1007/978-1-4020-5499-0_1.

Anonymous:2008:THN

[Ano08-47] Anonymous. Thomas Harriot’s notes on projectile motion and the fall of bodies. In The English Galileo: Thomas Harriot’s work on motion as an example of preclassical mechanics. Volume 2: Sources [Sch08b], chapter 11, pages 395–755. ISBN 1-4020-5498-X (set: cased), 1-4020-5499-8 (set: e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC133 .S36 2008; Q174 .B67 v.268. URL https://link.

springer.com/chapter/10.1007/
978-1-4020-5499-0_11.

Anonymous:2008:UI

[Ano08-48] Anonymous. Underdetermination and indeterminacy. In Underdetermination: An Essay on Evidence and the Limits of Natural Knowledge [Bon08], chapter 6, pages 207–258. ISBN 1-4020-6898-0, 1-4020-6899-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175 .B72235 2008. URL https:/

/link.springer.com/chapter/10.
1007/978-1-4020-6899-7_6.

Anonymous:2008:UIE

[Ano08-49] Anonymous. Underdetermination issues in the exact sciences. In Underdetermination: An Essay on Evidence and the Limits of Natural Knowledge [Bon08], chapter 2, pages 45–87. ISBN 1-4020-6898-0, 1-4020-6899-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175 .B72235

                                              2008. URL https://link.springer.
                                              com/chapter/10.1007/978-1-4020-
                                              6899-7_2.

Anonymous:2008:UTS

[Ano08-50] Anonymous. The unity of time and space. In Leibniz’s Metaphysics of Time and Space [Fut08], chapter 3, pages 57–80. ISBN 1-4020-8236-3 (hardcover), 1-4020-8237-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN B2599.S7 F88 2008.

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-1-4020-8237-
                                              5_3.

Anonymous:2009:BMa

[Ano09a] Anonymous. Back matter. In Causation, Coherence, and Concepts: A Collection of Essays [Spo09], pages 361–386. ISBN 1-4020-5473-4 (hardcover), 1-4020-5474-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN BD541; BD591 .S66 2009; Q175.3. URL https://link.

                                              springer.com/book/10.1007/978-1-
                                              4020-5474-7.

Anonymous:2009:BMg

[Ano09b] Anonymous. Back matter. In The Founders of Western Thought — The Presocratics: A diachronic parallelism between Presocratic Thought and Philosophy and the Natural Sciences [Vam09g], pages 255–289. ISBN 1-4020-9790-5 (hardcover), 1-4020-9791-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175 .V33413 2009. URL https://link.

springer.com/book/10.1007/978-1-
4020-9791-1.

Anonymous:2009:BMf

[Ano09c] Anonymous. Back matter. In Mapping the Future of Biology: Evolving Concepts and Theories [BMP09b], pages 155–173. ISBN 1-4020-9635-6 (hardcover), 1-4020-9636-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QH331 .M295 2009.

URL https://link.springer.com/
book/10.1007/978-1-4020-9636-5.

Anonymous:2009:BMb

[Ano09d] Anonymous. Back matter. In Going Amiss In Experimental Research [HSS09a], pages 273–277. ISBN 1-4020-8892-2 (hardcover), 1-4020-8893-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q172.5.E77

G647 2009.  URL https://link.
springer.com/book/10.1007/978-1-
4020-8893-3.

Anonymous:2009:BMc

[Ano09e] Anonymous. Back matter. In Rethinking Popper [PC09], pages 425–431. ISBN 1-4020-9337-3, 1-4020-9338-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN B1649.P64 R48 2009; Q174 .B67 2009. URL

https://link.springer.com/book/
10.1007/978-1-4020-9338-8.

Anonymous:2009:BMd

[Ano09f] Anonymous. Back matter. In Brenner and Gayon [BG09], pages 359–375. ISBN 1-4020-9367-5, 1-4020-9368-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174

.F746 2009. URL https://link.

springer.com/book/10.1007/978-1-
4020-9368-5.

Anonymous:2009:BMe

[Ano09g] Anonymous. Back matter. In Freudenthal and McLaughlin [FM09c], pages 257–273. ISBN 1-4020-9603-8, 1-4020-9604-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174.6 .S63 2009.

URL http://www.loc.gov/catdir/
enhancements/fy1312/2008942091-
d.html;   http://www.loc.gov/
catdir/enhancements/fy1312/
2008942091-t.html;   https://
link.springer.com/book/10.1007/
978-1-4020-9604-4.

Anonymous:2009:BMh

[Ano09h] Anonymous. Back matter. In The Scientist’s Atom and the Philosopher’s Stone: How Science Succeeded and Philosophy Failed to Gain Knowledge of Atoms [Cha09a], pages 269–287. ISBN 90-481-2361-5, 90-481-2362-3. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QD461 .C375 2009.

URL https://link.springer.com/
book/10.1007/978-90-481-2362-9.

Anonymous:2009:BMi

[Ano09i] Anonymous. Back matter. In Defending Copernicus and Galileo: Critical Reasoning in the Two Affairs [Fin09b], pages 315–350. ISBN 90-481-3200-2 (hardcover), 90-481-3201-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 2009; Q175 .B73 2009. URL

https://link.springer.com/book/
10.1007/978-90-481-3201-0.

Anonymous:2009:BNA

[Ano09j] Anonymous. Bayesian nets are all there is to causal dependence. In Causation, Coherence, and Concepts: A Collection of Essays [Spo09], chapter 4, pages 99–111. ISBN 1-4020-5473-4 (hardcover), 1-4020-5474-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN BD541; BD591 .S66 2009; Q175.3. URL https:/

/link.springer.com/chapter/10.
1007/978-1-4020-5474-7_4.

Anonymous:2009:CLO

[Ano09k] Anonymous. Causal laws are objectifications of inductive schemes. In Causation, Coherence, and Concepts: A Collection of Essays [Spo09], chapter 5, pages 113–134. ISBN 1-4020-5473-4 (hardcover), 1-4020-5474-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN BD541; BD591 .S66 2009; Q175.3. URL https:

//link.springer.com/chapter/10.
1007/978-1-4020-5474-7_5.

Anonymous:2009:CA

[Ano09l] Anonymous. Causation: An alternative. In Causation, Coherence, and Concepts: A Collection of Essays [Spo09], chapter 3, pages 75–97. ISBN 1-4020-5473-4 (hardcover), 1-4020-5474-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN BD541; BD591 .S66 2009; Q175.3.

URL https://link.springer.com/
chapter/10.1007/978-1-4020-5474-
7_3.

Anonymous:2009:CNH

[Ano09m] Anonymous. Chance and necessity: From Humean supervenience to HumeanProjection. In Causation, Coherence, and Concepts: A Collection of Essays [Spo09], chapter 8, pages 175–205. ISBN 1-4020-5473-4 (hardcover), 1-4020-5474-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN BD541; BD591 .S66 2009; Q175.3. URL https:/

                                              /link.springer.com/chapter/10.
                                              1007/978-1-4020-5474-7_8.

Anonymous:2009:CC

[Ano09n] Anonymous. Changing concepts. In Causation, Coherence, and Concepts: A Collection of Essays [Spo09], chapter 15, pages 329–333. ISBN 1-4020-5473-4 (hardcover), 1-4020-5474-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN BD541; BD591 .S66 2009; Q175.3. URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-1-4020-5474-7_15.

Anonymous:2009:CCT

[Ano09o] Anonymous. The character of color terms: A materialist view. In Causation, Coherence, and Concepts: A Collection of Essays [Spo09], chapter 13, pages 285–304. ISBN 1-4020-5473-4 (hardcover), 1-4020-5474-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN BD541; BD591 .S66 2009; Q175.3. URL https:/

                                              /link.springer.com/chapter/10.
                                              1007/978-1-4020-5474-7_13.

Anonymous:2009:CBA

[Ano09p] Anonymous. Concepts are beliefs about essences. In Causation, Coherence, and Concepts: A Collection of Essays [Spo09], chapter 14, pages 305–328. ISBN 1-4020-5473-4 (hardcover), 1-4020-5474-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN BD541; BD591 .S66 2009; Q175.3.

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-1-4020-5474-
                                              7_14.

Anonymous:2009:DIC

[Ano09q] Anonymous. Direct and indirect causes. In Causation, Coherence, and Concepts: A Collection of Essays [Spo09], chapter 2, pages 45–74. ISBN 1-4020-5473-4 (hardcover), 1-4020-5474-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN BD541; BD591 .S66 2009; Q175.3. URL https:

//link.springer.com/chapter/10.
1007/978-1-4020-5474-7_2.

Anonymous:2009:EIL

[Ano09r] Anonymous. Enumerative induction and lawlikeness. In Causation, Coherence, and Concepts: A Collection of Essays [Spo09], chapter 7, pages 155–173. ISBN 1-4020-5473-4 (hardcover), 1-4020-5474-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN BD541; BD591 .S66 2009; Q175.3. URL https://link.

springer.com/chapter/10.1007/
978-1-4020-5474-7_7.

Anonymous:2009:FMa

[Ano09s] Anonymous. Front matter. In Causation, Coherence, and Concepts: A Collection of Essays [Spo09], pages i–xvi. ISBN 1-4020-5473-4 (hardcover), 1-4020-5474-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN BD541; BD591 .S66 2009; Q175.3. URL

https://link.springer.com/book/
10.1007/978-1-4020-5474-7.

Anonymous:2009:FMs

[Ano09t] Anonymous. Front matter. In The Founders of Western Thought — The Presocratics: A diachronic parallelism between Presocratic Thought and Philosophy and the Natural Sciences [Vam09g], pages I–XVI. ISBN 1-4020-9790-5 (hardcover), 1-4020-9791-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175 .V33413 2009. URL https://link.

springer.com/book/10.1007/978-1-
4020-9791-1.

Anonymous:2009:FMr

[Ano09u] Anonymous. Front matter. In Mapping the Future of Biology: Evolving Concepts and Theories [BMP09b], pages I–XI. ISBN 1-4020-9635-6 (hardcover), 1-4020-9636-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QH331 .M295 2009. URL

                                              https://link.springer.com/book/
                                              10.1007/978-1-4020-9636-5.

Anonymous:2009:FMb

[Ano09v] Anonymous.  Front matter.  In

Going Amiss In Experimental Research [HSS09a], pages I–X. ISBN 1-4020-8892-2 (hardcover), 1-4020-8893-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q172.5.E77 G647 2009. URL https:/

                                              /link.springer.com/book/10.1007/
                                              978-1-4020-8893-3.

Anonymous:2009:FMc

[Ano09w] Anonymous. Front matter. In Going Amiss In Experimental Research [HSS09a], page 1. ISBN 1-4020-8892-2 (hardcover), 1-4020-8893-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q172.5.E77 G647

                                              2009. URL https://link.springer.
                                              com/book/10.1007/978-1-4020-
                                              8893-3.

Anonymous:2009:FMd

[Ano09x] Anonymous. Front matter. In Going Amiss In Experimental Research [HSS09a], page 9. ISBN 1-4020-8892-2 (hardcover), 1-4020-8893-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q172.5.E77 G647

                                              2009. URL https://link.springer.
                                              com/book/10.1007/978-1-4020-
                                              8893-3.

Anonymous:2009:FMe

[Ano09y] Anonymous. Front matter. In Going Amiss In Experimental Research [HSS09a], page 47. ISBN 1-4020-8892-2 (hardcover), 1-4020-8893-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q172.5.E77 G647

2009. URL https://link.springer.
com/book/10.1007/978-1-4020-
8893-3.

Anonymous:2009:FMf

[Ano09z] Anonymous. Front matter. In Going Amiss In Experimental Research [HSS09a], page 73. ISBN 1-4020-8892-2 (hardcover), 1-4020-8893-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q172.5.E77 G647

2009. URL https://link.springer.
com/book/10.1007/978-1-4020-
8893-3.

Anonymous:2009:FMg

[Ano09-27] Anonymous. Front matter. In Going Amiss In Experimental Research [HSS09a], page 159. ISBN 1-4020-8892-2 (hardcover), 1-4020-8893-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q172.5.E77 G647

2009. URL https://link.springer.
com/book/10.1007/978-1-4020-
8893-3.

Anonymous:2009:FMh

[Ano09-28] Anonymous. Front matter. In Going Amiss In Experimental Research [HSS09a], page 237. ISBN 1-4020-8892-2 (hardcover), 1-4020-8893-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q172.5.E77 G647

2009. URL https://link.springer.
com/book/10.1007/978-1-4020-
8893-3.

Anonymous:2009:FMi

[Ano09-29] Anonymous. Front matter. In Rethinking Popper [PC09], pages i–xiv. ISBN 1-4020-9337-3, 1-4020-9338-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN B1649.P64 R48 2009; Q174 .B67 2009. URL https:/

                                              /link.springer.com/book/10.1007/
                                              978-1-4020-9338-8.

Anonymous:2009:FMj

[Ano09-30] Anonymous. Front matter. In Brenner and Gayon [BG09], pages i–xxviii. ISBN 1-4020-9367-5, 1-4020-9368-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .F746 2009.

                                              URL https://link.springer.com/
                                              book/10.1007/978-1-4020-9368-5.

Anonymous:2009:FMk

[Ano09-31] Anonymous. Front matter. In Brenner and Gayon [BG09], page 23. ISBN 1-4020-9367-5, 1-4020-9368-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .F746 2009.

                                              URL https://link.springer.com/
                                              book/10.1007/978-1-4020-9368-5.

Anonymous:2009:FMl

[Ano09-32] Anonymous. Front matter. In Brenner and Gayon [BG09], page 71. ISBN 1-4020-9367-5, 1-4020-9368-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .F746 2009.

                                              URL https://link.springer.com/
                                              book/10.1007/978-1-4020-9368-5.

Anonymous:2009:FMm

[Ano09-33] Anonymous. Front matter. In Brenner and Gayon [BG09], page 113. ISBN 1-4020-9367-5, 1-4020-9368-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .F746 2009.

                                              URL https://link.springer.com/
                                              book/10.1007/978-1-4020-9368-5.

Anonymous:2009:FMn

[Ano09-34] Anonymous. Front matter. In Brenner and Gayon [BG09], page 187. ISBN 1-4020-9367-5, 1-4020-9368-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .F746 2009.

URL https://link.springer.com/
book/10.1007/978-1-4020-9368-5.

Anonymous:2009:FMo

[Ano09-35] Anonymous. Front matter. In Brenner and Gayon [BG09], page 224. ISBN 1-4020-9367-5, 1-4020-9368-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .F746 2009.

URL https://link.springer.com/
book/10.1007/978-1-4020-9368-5.

Anonymous:2009:FMp

[Ano09-36] Anonymous. Front matter. In Brenner and Gayon [BG09], page 301. ISBN 1-4020-9367-5, 1-4020-9368-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .F746 2009.

URL https://link.springer.com/
book/10.1007/978-1-4020-9368-5.

Anonymous:2009:FMq

[Ano09-37] Anonymous. Front matter. In Freudenthal and McLaughlin [FM09c], pages I–XI. ISBN 1-4020-9603-8, 1-4020-9604-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174.6

.S63 2009. URL http://www.loc.
gov/catdir/enhancements/fy1312/
2008942091-d.html;   http://www.
loc.gov/catdir/enhancements/
fy1312/2008942091-t.html;
https://link.springer.com/book/
10.1007/978-1-4020-9604-4.

Anonymous:2009:FMt

[Ano09-38] Anonymous. Front matter. In The Scientist’s Atom and the Philosopher’s Stone: How Science Succeeded and Philosophy Failed to Gain Knowledge of Atoms [Cha09a], pages I–XII. ISBN 90-481-2361-5, 90-481-2362-3. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QD461 .C375 2009.

URL https://link.springer.com/
book/10.1007/978-90-481-2362-9.

Anonymous:2009:FMu

[Ano09-39] Anonymous. Front matter. In Defending Copernicus and Galileo: Critical Reasoning in the Two Affairs [Fin09b], pages i–xliii. ISBN 90-481-3200-2 (hardcover), 90-481-3201-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 2009; Q175 .B73 2009. URL

https://link.springer.com/book/
10.1007/978-90-481-3201-0.

Anonymous:2009:FMv

[Ano09-40] Anonymous. Front matter. In Defending Copernicus and Galileo: Critical Reasoning in the Two Affairs [Fin09b], page 1. ISBN 90-481-3200-2 (hardcover), 90-481-3201-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 2009; Q175

.B73 2009.  URL https://link.
springer.com/book/10.1007/978-
90-481-3201-0.

Anonymous:2009:FMw

[Ano09-41] Anonymous. Front matter. In Defending Copernicus and Galileo: Critical Reasoning in the Two Affairs [Fin09b], page 136. ISBN 90-481-3200-2 (hardcover), 90-481-3201-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 2009; Q175 .B73 2009. URL

https://link.springer.com/book/
10.1007/978-90-481-3201-0.

Anonymous:2009:HUF

[Ano09-42] Anonymous. How to understand the foundations of empirical belief in a coherentist way. In Causation, Coherence, and Concepts: A Collection of Essays [Spo09], chapter 11, pages 251–263. ISBN 1-4020-5473-4 (hardcover), 1-4020-5474-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN BD541; BD591 .S66 2009; Q175.3.

URL https://link.springer.com/
chapter/10.1007/978-1-4020-5474-
7_11.

Anonymous:2009:IVP

[Ano09-43] Anonymous. The intentional versus the propositional structure of contents. In Causation, Coherence, and Concepts: A Collection of Essays [Spo09], chapter 16, pages 335–359. ISBN 1-4020-5473-4 (hardcover), 1-4020-5474-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN BD541; BD591 .S66 2009; Q175.3. URL https:

//link.springer.com/chapter/10.
1007/978-1-4020-5474-7_16.

Anonymous:2009:LCP

[Ano09-44] Anonymous. Laws, ceteris paribus conditions, and the dynamics of belief. In Causation, Coherence, and Concepts: A Collection of Essays [Spo09], chapter 6, pages 137–154. ISBN 1-4020-5473-4 (hardcover), 1-4020-5474-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN BD541; BD591 .S66 2009; Q175.3. URL https:

//link.springer.com/chapter/10.
1007/978-1-4020-5474-7_6.

Anonymous:2009:OCF

[Ano09-45] Anonymous. Ordinal conditional functions: A dynamic theory of epistemic states. In Causation, Coherence, and Concepts: A Collection of Essays [Spo09], chapter 1, pages 19–41. ISBN 1-4020-5473-4 (hardcover), 1-4020-5474-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN BD541; BD591 .S66 2009; Q175.3.

URL https://link.springer.com/
chapter/10.1007/978-1-4020-5474-
7_1.

Anonymous:2009:PRF

[Ano09-46] Anonymous. A priori reasons: A fresh look at disposition predicates. In Causation, Coherence, and Concepts: A Collection of Essays [Spo09], chapter 12, pages 267–283. ISBN 1-4020-5473-4 (hardcover), 1-4020-5474-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN BD541; BD591 .S66 2009; Q175.3. URL https:

//link.springer.com/chapter/10.
1007/978-1-4020-5474-7_12.

Anonymous:2009:REE

[Ano09-47] Anonymous. A reason for explanation: Explanations provide stable reasons. In Causation, Coherence, and Concepts: A Collection of Essays [Spo09], chapter 9, pages 209–232. ISBN 1-4020-5473-4 (hardcover), 1-4020-5474-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN BD541; BD591 .S66 2009; Q175.3. URL https:

//link.springer.com/chapter/10.
1007/978-1-4020-5474-7_9.

Anonymous:2009:TCP

[Ano09-48] Anonymous. Two coherence principles. In Causation, Coherence, and Concepts: A Collection of Essays [Spo09], chapter 10, pages 233–250. ISBN 1-4020-5473-4 (hardcover), 1-4020-5474-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN BD541; BD591 .S66 2009; Q175.3. URL https:

//link.springer.com/chapter/10.
1007/978-1-4020-5474-7_10.

Anonymous:2010:BMd

[Ano10a] Anonymous. Back matter. In Frigg and Hunter [FH10b], pages 263–265. ISBN 90-481-3850-7 (hardcover), 90-481-3851-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN BH301.M47 B49 2010; Q175 .B52 2010.

URL https://link.springer.com/
book/10.1007/978-90-481-3851-7.

Anonymous:2010:BMc

[Ano10b] Anonymous. Back matter. In Looking at it from Asia: the Processes that Shaped the Sources of History of Science [BE10b], pages 413–426. ISBN 90-481-3675-X (hardcover), 90-481-3676-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q127.A65 L66 2010. URL

https://link.springer.com/book/
10.1007/978-90-481-3676-6.

Anonymous:2010:BMe

[Ano10c] Anonymous. Back matter. In Galileo Engineer [Val10b], pages 213–320. ISBN 90-481-8644-7 (hardcover), 90-481-8645-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/book/10.1007/978-90-481-
8645-7.

Anonymous:2010:BMa

[Ano10d] Anonymous. Back matter. In Erich Hückel (1896–1980): From Physics to Quantum Chemistry [Kar10b], pages 179–200. ISBN 90-481-3559-1 (hardcover), 90-481-3560-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC174.12 .K37313 2010.

URL https://link.springer.com/
book/10.1007/978-90-481-3560-8.

Anonymous:2010:BMb

[Ano10e] Anonymous. Back matter. In The Place of Probability in Science: In Honor of Ellery Eells (1953–2006) [Fet10a], pages 353–368. ISBN 90-481-3614-8 (hardcover), 90-481-3615-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175 .P563 2010.

URL https://link.springer.com/
book/10.1007/978-90-481-3615-5.

Anonymous:2010:BMf

[Ano10f] Anonymous. Back matter. In Matter and Mind: a Philosophical Inquiry [Bun10j], pages 287–319. ISBN 90-481-9224-2 (hardcover), 90-481-9225-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN BD418.3 .B7664 2010; Q175 .B73 v.287. URL

https://link.springer.com/book/
10.1007/978-90-481-9225-0.

Anonymous:2010:FMj

[Ano10g] Anonymous. Front matter. In Frigg and Hunter [FH10b], pages i–xxx. ISBN 90-481-3850-7 (hardcover), 90-481-3851-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN BH301.M47 B49 2010; Q175 .B52 2010.

URL https://link.springer.com/
book/10.1007/978-90-481-3851-7.

Anonymous:2010:FMg

[Ano10h] Anonymous. Front matter. In Looking at it from Asia: the Processes that Shaped the Sources of History of Science [BE10b], pages i–xlvi. ISBN 90-481-3675-X (hardcover), 90-481-3676-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q127.A65 L66 2010. URL

https://link.springer.com/book/
10.1007/978-90-481-3676-6.

Anonymous:2010:FMh

[Ano10i] Anonymous. Front matter. In Looking at it from Asia: the Processes that Shaped the Sources of History of Science [BE10b], page 1. ISBN 90-481-3675-X (hardcover), 90-481-3676-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q127.A65 L66 2010. URL https://

link.springer.com/book/10.1007/
978-90-481-3676-6.

Anonymous:2010:FMi

[Ano10j] Anonymous. Front matter. In Looking at it from Asia: the Processes that Shaped the Sources of History of Science [BE10b], page 167. ISBN 90-481-3675-X (hardcover), 90-481-3676-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q127.A65 L66 2010. URL https://

link.springer.com/book/10.1007/
978-90-481-3676-6.

Anonymous:2010:FMk

[Ano10k] Anonymous. Front matter. In Galileo Engineer [Val10b], pages i–xxii. ISBN 90-481-8644-7 (hardcover), 90-481-8645-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/book/10.1007/978-90-481-
8645-7.

Anonymous:2010:FMl

[Ano10l] Anonymous. Front matter. In Galileo Engineer [Val10b], page 1. ISBN 90-481-8644-7 (hardcover), 90-481-8645-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
book/10.1007/978-90-481-8645-7.

Anonymous:2010:FMm

[Ano10m] Anonymous. Front matter. In Galileo Engineer [Val10b], page 115. ISBN 90-481-8644-7 (hardcover), 90-481-8645-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
book/10.1007/978-90-481-8645-7.

Anonymous:2010:FMn

[Ano10n] Anonymous. Front matter. In Galileo Engineer [Val10b], page 191. ISBN 90-481-8644-7 (hardcover), 90-481-8645-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
book/10.1007/978-90-481-8645-7.

Anonymous:2010:FMa

[Ano10o] Anonymous. Front matter. In Erich Hückel (1896–1980): From Physics to Quantum Chemistry [Kar10b], pages i–xvii. ISBN 90-481-3559-1 (hardcover), 90-481-3560-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC174.12 .K37313 2010. URL

https://link.springer.com/book/
10.1007/978-90-481-3560-8.

Anonymous:2010:FMb

[Ano10p] Anonymous. Front matter. In The Place of Probability in Science: In Honor of Ellery Eells (1953–2006) [Fet10a], pages i–xliii. ISBN 90-481-3614-8 (hardcover), 90-481-3615-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175 .P563 2010.

URL https://link.springer.com/
book/10.1007/978-90-481-3615-5.

Anonymous:2010:FMc

[Ano10q] Anonymous. Front matter. In The Place of Probability in Science: In Honor of Ellery Eells (1953–2006) [Fet10a], page 46. ISBN 90-481-3614-8 (hardcover), 90-481-3615-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175 .P563 2010.

URL https://link.springer.com/
book/10.1007/978-90-481-3615-5.

Anonymous:2010:FMd

[Ano10r] Anonymous. Front matter. In The Place of Probability in Science: In Honor of Ellery Eells (1953–2006) [Fet10a], page 100. ISBN 90-481-3614-8 (hardcover), 90-481-3615-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175 .P563 2010.

URL https://link.springer.com/
book/10.1007/978-90-481-3615-5.

Anonymous:2010:FMe

[Ano10s] Anonymous. Front matter. In The Place of Probability in Science: In Honor of Ellery Eells (1953–2006) [Fet10a], page 180. ISBN 90-481-3614-8 (hardcover), 90-481-3615-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175 .P563 2010.

URL https://link.springer.com/
book/10.1007/978-90-481-3615-5.

Anonymous:2010:FMf

[Ano10t] Anonymous. Front matter. In The Place of Probability in Science: In Honor of Ellery Eells (1953–2006) [Fet10a], page 246. ISBN 90-481-3614-8 (hardcover), 90-481-3615-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175 .P563 2010.

URL https://link.springer.com/
book/10.1007/978-90-481-3615-5.

Anonymous:2010:FMo

[Ano10u] Anonymous. Front matter. In Matter and Mind: a Philosophical Inquiry [Bun10j], pages i–xviii. ISBN 90-481-9224-2 (hardcover), 90-481-9225-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN BD418.3 .B7664 2010; Q175 .B73 v.287. URL

https://link.springer.com/book/
10.1007/978-90-481-9225-0.

Anonymous:2010:FMp

[Ano10v] Anonymous. Front matter. In Matter and Mind: a Philosophical Inquiry [Bun10j], page 1. ISBN 90-481-9224-2 (hardcover), 90-481-9225-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN BD418.3 .B7664 2010; Q175 .B73 v.287. URL

                                              https://link.springer.com/book/
                                              10.1007/978-90-481-9225-0.

Anonymous:2010:FMq

[Ano10w] Anonymous. Front matter. In Matter and Mind: a Philosophical Inquiry [Bun10j], page 141. ISBN 90-481-9224-2 (hardcover), 90-481-9225-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN BD418.3 .B7664 2010; Q175 .B73 v.287. URL

                                              https://link.springer.com/book/
                                              10.1007/978-90-481-9225-0.

Anonymous:2010:FMr

[Ano10x] Anonymous. Front matter. In Matter and Mind: a Philosophical Inquiry [Bun10j], page 265. ISBN 90-481-9224-2 (hardcover), 90-481-9225-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN BD418.3 .B7664 2010; Q175 .B73 v.287. URL

                                              https://link.springer.com/book/
                                              10.1007/978-90-481-9225-0.

Anonymous:2013:BMa

[Ano11a] Anonymous. Back matter. In The Berlin Group and the Philosophy of Logical Empiricism [MP13], pages 325–332. ISBN 94-007-5484-1 (hardcover), 94-007-5485-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175.32.A24 .M384 2011; Q175 .B47 2013. URL https:/

                                              /link.springer.com/book/10.1007/
                                              978-94-007-5485-0.

Anonymous:2011:BMa

[Ano11b] Anonymous. Back matter. In Science in the Context of Application [CN11a], pages 483–492. ISBN 90-481-9050-9, 90-481-9051-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q180.55.M4 S385 2011. URL http://

www.loc.gov/catdir/enhancements/
fy1114/2010934513-d.html;
http://www.loc.gov/catdir/toc/
fy11pdf03/2010934513.html;
https://link.springer.com/book/
10.1007/978-90-481-9051-5.

Anonymous:2011:BMd

[Ano11c] Anonymous. Back matter. In Science between Europe and Asia: Historical Studies on the Transmission, Adoption and Adaptation of Knowledge [GR11b], pages 271–279. ISBN 90-481-9968-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q125 .S354 2010; Q175.52.A78 S35 2010.

URL https://link.springer.com/
book/10.1007/978-90-481-9968-6.

Anonymous:2011:BMe

[Ano11d] Anonymous. Back matter. In Commentary on the Principles of Thermodynamics by Pierre Duhem [Nee11c], pages 211–215. ISBN 94-007-0310-4, 94-007-0311-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC311 .D79 2011. URL

https://link.springer.com/book/
10.1007/978-94-007-0311-7.

Anonymous:2011:BMc

[Ano11e] Anonymous. Back matter. In Bokulich and Bokulich [BB11], pages 177–182. ISBN 90-481-9596-9, 90-481-9597-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175.32.S78

S35 2011.  URL https://link.

springer.com/book/10.1007/978-
90-481-9597-8.

Anonymous:2011:BMg

[Ano11f] Anonymous. Back matter. In Niels Bohr’s Complementarity: Its Structure, History, and Intersections with Hermeneutics and Deconstruction [Kat11d], pages 153–176. ISBN 94-007-1747-4, 94-007-1748-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC174.17.C63 K38

2011. URL https://link.springer.
com/book/10.1007/978-94-007-
1748-0.

Anonymous:2011:BMf

[Ano11g] Anonymous. Back matter. In Honoré Fabri and the Concept of Impetus: a Bridge Between Conceptual Frameworks [Ela11j], pages 231–259. ISBN 94-007-1604-4 (hardcover), 94-007-1605-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175 .E43 2011 v.288; QA802

.E43 2011. URL http://www.loc.
gov/catdir/enhancements/fy1202/
2011928368-d.html;   http://www.
loc.gov/catdir/enhancements/
fy1202/2011928368-t.html;
https://link.springer.com/book/
10.1007/978-94-007-1605-6.

Anonymous:2011:BMb

[Ano11h] Anonymous. Back matter. In Brazilian Studies in Philosophy and History of Science: An account of recent works [KV11], pages 341–352. ISBN 90-481-9421-0 (hardcover), 90-481-9422-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175.3

.B73 2011.  URL https://link.
springer.com/book/10.1007/978-
90-481-9422-3.

Anonymous:2013:FMa

[Ano11i] Anonymous. Front matter. In The Berlin Group and the Philosophy of Logical Empiricism [MP13], pages i–x. ISBN 94-007-5484-1 (hardcover), 94-007-5485-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175.32.A24 .M384 2011; Q175

.B47 2013.  URL https://link.
springer.com/book/10.1007/978-
94-007-5485-0.

Anonymous:2013:FMb

[Ano11j] Anonymous. Front matter. In The Berlin Group and the Philosophy of Logical Empiricism [MP13], page 1. ISBN 94-007-5484-1 (hardcover), 94-007-5485-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175.32.A24 .M384 2011; Q175 .B47

2013. URL https://link.springer.
com/book/10.1007/978-94-007-
5485-0.

Anonymous:2013:FMc

[Ano11k] Anonymous. Front matter. In The Berlin Group and the Philosophy of Logical Empiricism [MP13], page 41. ISBN 94-007-5484-1 (hardcover), 94-007-5485-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175.32.A24 .M384 2011; Q175 .B47

2013. URL https://link.springer.
com/book/10.1007/978-94-007-
5485-0.

Anonymous:2013:FMd

[Ano11l] Anonymous. Front matter. In The Berlin Group and the Philosophy of Logical Empiricism [MP13], page 95. ISBN 94-007-5484-1 (hardcover), 94-007-5485-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175.32.A24 .M384 2011; Q175 .B47

2013. URL https://link.springer.

                                              com/book/10.1007/978-94-007-
                                              5485-0.

Anonymous:2013:FMe

[Ano11m] Anonymous. Front matter. In The Berlin Group and the Philosophy of Logical Empiricism [MP13], page 177. ISBN 94-007-5484-1 (hardcover), 94-007-5485-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175.32.A24 .M384 2011; Q175 .B47

                                              2013. URL https://link.springer.
                                              com/book/10.1007/978-94-007-
                                              5485-0.

Anonymous:2013:FMf

[Ano11n] Anonymous. Front matter. In The Berlin Group and the Philosophy of Logical Empiricism [MP13], page 229. ISBN 94-007-5484-1 (hardcover), 94-007-5485-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175.32.A24 .M384 2011; Q175 .B47

                                              2013. URL https://link.springer.
                                              com/book/10.1007/978-94-007-
                                              5485-0.

Anonymous:2013:FMg

[Ano11o] Anonymous. Front matter. In The Berlin Group and the Philosophy of Logical Empiricism [MP13], page 263. ISBN 94-007-5484-1 (hardcover), 94-007-5485-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175.32.A24 .M384 2011; Q175 .B47

                                              2013. URL https://link.springer.
                                              com/book/10.1007/978-94-007-
                                              5485-0.

Anonymous:2011:FMa

[Ano11p] Anonymous. Front matter. In Science in the Context of Application [CN11a], pages i–x. ISBN 90-481-9050-9, 90-481-9051-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q180.55.M4 S385 2011. URL http://

www.loc.gov/catdir/enhancements/
fy1114/2010934513-d.html;
http://www.loc.gov/catdir/toc/
fy11pdf03/2010934513.html;
https://link.springer.com/book/
10.1007/978-90-481-9051-5.

Anonymous:2011:FMb

[Ano11q] Anonymous. Front matter. In Science in the Context of Application [CN11a], page 9. ISBN 90-481-9050-9, 90-481-9051-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q180.55.M4 S385 2011. URL http://

www.loc.gov/catdir/enhancements/
fy1114/2010934513-d.html;
http://www.loc.gov/catdir/toc/
fy11pdf03/2010934513.html;
https://link.springer.com/book/
10.1007/978-90-481-9051-5.

Anonymous:2011:FMc

[Ano11r] Anonymous. Front matter. In Science in the Context of Application [CN11a], page 113. ISBN 90-481-9050-9, 90-481-9051-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q180.55.M4 S385 2011. URL http://

www.loc.gov/catdir/enhancements/
fy1114/2010934513-d.html;
http://www.loc.gov/catdir/toc/
fy11pdf03/2010934513.html;
https://link.springer.com/book/
10.1007/978-90-481-9051-5.

Anonymous:2011:FMd

[Ano11s] Anonymous. Front matter. In Science in the Context of Application [CN11a], page 195. ISBN 90-481-9050-9, 90-481-9051-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q180.55.M4 S385 2011. URL http://

www.loc.gov/catdir/enhancements/
fy1114/2010934513-d.html;

http://www.loc.gov/catdir/toc/
fy11pdf03/2010934513.html;
https://link.springer.com/book/
10.1007/978-90-481-9051-5.

Anonymous:2011:FMe

[Ano11t] Anonymous. Front matter. In Science in the Context of Application [CN11a], page 243. ISBN 90-481-9050-9, 90-481-9051-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q180.55.M4 S385 2011. URL http://

www.loc.gov/catdir/enhancements/
fy1114/2010934513-d.html;
http://www.loc.gov/catdir/toc/
fy11pdf03/2010934513.html;
https://link.springer.com/book/
10.1007/978-90-481-9051-5.

Anonymous:2011:FMf

[Ano11u] Anonymous. Front matter. In Science in the Context of Application [CN11a], page 349. ISBN 90-481-9050-9, 90-481-9051-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q180.55.M4 S385 2011. URL http://

www.loc.gov/catdir/enhancements/
fy1114/2010934513-d.html;
http://www.loc.gov/catdir/toc/
fy11pdf03/2010934513.html;
https://link.springer.com/book/
10.1007/978-90-481-9051-5.

Anonymous:2011:FMg

[Ano11v] Anonymous. Front matter. In Science in the Context of Application [CN11a], page 387. ISBN 90-481-9050-9, 90-481-9051-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q180.55.M4 S385 2011. URL http://

www.loc.gov/catdir/enhancements/
fy1114/2010934513-d.html;
http://www.loc.gov/catdir/toc/
fy11pdf03/2010934513.html;
https://link.springer.com/book/
10.1007/978-90-481-9051-5.

Anonymous:2011:FMj

[Ano11w] Anonymous. Front matter. In Science between Europe and Asia: Historical Studies on the Transmission, Adoption and Adaptation of Knowledge [GR11b], pages i–xiii. ISBN 90-481-9968-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q125 .S354 2010; Q175.52.A78 S35 2010.

URL https://link.springer.com/
book/10.1007/978-90-481-9968-6.

Anonymous:2011:FMk

[Ano11x] Anonymous. Front matter. In Science between Europe and Asia: Historical Studies on the Transmission, Adoption and Adaptation of Knowledge [GR11b], page 11. ISBN 90-481-9968-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q125 .S354 2010; Q175.52.A78 S35 2010.

URL https://link.springer.com/
book/10.1007/978-90-481-9968-6.

Anonymous:2011:FMl

[Ano11y] Anonymous. Front matter. In Science between Europe and Asia: Historical Studies on the Transmission, Adoption and Adaptation of Knowledge [GR11b], page 75. ISBN 90-481-9968-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q125 .S354 2010; Q175.52.A78 S35 2010.

URL https://link.springer.com/
book/10.1007/978-90-481-9968-6.

Anonymous:2011:FMm

[Ano11z] Anonymous. Front matter. In Science between Europe and Asia: Historical Studies on the Transmission, Adoption and Adaptation of Knowledge [GR11b], page 163. ISBN 90-481-9968-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q125 .S354 2010; Q175.52.A78 S35 2010.

                                              URL https://link.springer.com/
                                              book/10.1007/978-90-481-9968-6.

Anonymous:2011:FMn

[Ano11-27] Anonymous. Front matter. In Science between Europe and Asia: Historical Studies on the Transmission, Adoption and Adaptation of Knowledge [GR11b], page 207. ISBN 90-481-9968-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q125 .S354 2010; Q175.52.A78 S35 2010.

                                              URL https://link.springer.com/
                                              book/10.1007/978-90-481-9968-6.

Anonymous:2011:FMo

[Ano11-28] Anonymous.  Front matter.  In

Commentary on the Principles of Thermodynamics by Pierre Duhem [Nee11c], pages i–xv. ISBN 94-007-0310-4, 94-007-0311-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC311 .D79 2011. URL

                                              https://link.springer.com/book/
                                              10.1007/978-94-007-0311-7.

Anonymous:2011:FMp

[Ano11-29] Anonymous.  Front matter.  In

Commentary on the Principles of Thermodynamics by Pierre Duhem [Nee11c], pages 1–2. ISBN 94-007-0310-4, 94-007-0311-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC311 .D79 2011. URL

                                              https://link.springer.com/book/
                                              10.1007/978-94-007-0311-7.

Anonymous:2011:FMq

[Ano11-30] Anonymous.  Front matter.  In

Commentary on the Principles of Thermodynamics by Pierre Duhem [Nee11c], pages 35–36. ISBN 94-007-0310-4, 94-007-0311-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC311 .D79 2011. URL

https://link.springer.com/book/
10.1007/978-94-007-0311-7.

Anonymous:2011:FMr

[Ano11-31] Anonymous.  Front matter.  In

Commentary on the Principles of Thermodynamics by Pierre Duhem [Nee11c], page 83. ISBN 94-007-0310-4, 94-007-0311-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC311 .D79 2011. URL

https://link.springer.com/book/
10.1007/978-94-007-0311-7.

Anonymous:2011:FMs

[Ano11-32] Anonymous.  Front matter.  In

Commentary on the Principles of Thermodynamics by Pierre Duhem [Nee11c], page 141. ISBN 94-007-0310-4, 94-007-0311-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC311 .D79 2011. URL

https://link.springer.com/book/
10.1007/978-94-007-0311-7.

Anonymous:2011:FMi

[Ano11-33] Anonymous.  Front matter.  In

Bokulich and Bokulich [BB11], pages i–xvii. ISBN 90-481-9596-9, 90-481-9597-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175.32.S78 S35 2011. URL https:/

/link.springer.com/book/10.1007/
978-90-481-9597-8.

Anonymous:2011:FMy

[Ano11-34] Anonymous. Front matter. In Niels Bohr’s Complementarity: Its Structure, History, and Intersections with Hermeneutics and Deconstruction [Kat11d], pages i–xiv. ISBN 94-007-1747-4, 94-007-1748-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC174.17.C63 K38

2011. URL https://link.springer.

com/book/10.1007/978-94-007-
1748-0.

Anonymous:2011:FMt

[Ano11-35] Anonymous. Front matter. In Honoré Fabri and the Concept of Impetus: a Bridge Between Conceptual Frameworks [Ela11j], pages i–xxix. ISBN 94-007-1604-4 (hardcover), 94-007-1605-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175 .E43 2011 v.288; QA802

.E43 2011. URL http://www.loc.
gov/catdir/enhancements/fy1202/
2011928368-d.html;   http://www.
loc.gov/catdir/enhancements/
fy1202/2011928368-t.html;
https://link.springer.com/book/
10.1007/978-94-007-1605-6.

Anonymous:2011:FMu

[Ano11-36] Anonymous. Front matter. In Honoré Fabri and the Concept of Impetus: a Bridge Between Conceptual Frameworks [Ela11j], page 1. ISBN 94-007-1604-4 (hardcover), 94-007-1605-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175 .E43 2011 v.288; QA802 .E43 2011.

URL http://www.loc.gov/catdir/
enhancements/fy1202/2011928368-
d.html;   http://www.loc.gov/
catdir/enhancements/fy1202/
2011928368-t.html;   https://
link.springer.com/book/10.1007/
978-94-007-1605-6.

Anonymous:2011:FMv

[Ano11-37] Anonymous.  Front matter.  In

Honoré Fabri and the Concept of Impetus: a Bridge Between Conceptual Frameworks [Ela11j], page 61. ISBN 94-007-1604-4 (hardcover), 94-007-1605-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175 .E43 2011 v.288; QA802

.E43 2011. URL http://www.loc.
gov/catdir/enhancements/fy1202/
2011928368-d.html;   http://www.
loc.gov/catdir/enhancements/
fy1202/2011928368-t.html;
https://link.springer.com/book/
10.1007/978-94-007-1605-6.

Anonymous:2011:FMw

[Ano11-38] Anonymous.  Front matter.  In

Honoré Fabri and the Concept of Impetus: a Bridge Between Conceptual Frameworks [Ela11j], page 119. ISBN 94-007-1604-4 (hardcover), 94-007-1605-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175 .E43 2011 v.288; QA802

.E43 2011. URL http://www.loc.
gov/catdir/enhancements/fy1202/
2011928368-d.html;   http://www.
loc.gov/catdir/enhancements/
fy1202/2011928368-t.html;
https://link.springer.com/book/
10.1007/978-94-007-1605-6.

Anonymous:2011:FMx

[Ano11-39] Anonymous.  Front matter.  In

Honoré Fabri and the Concept of Impetus: a Bridge Between Conceptual Frameworks [Ela11j], page 189. ISBN 94-007-1604-4 (hardcover), 94-007-1605-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175 .E43 2011 v.288; QA802

.E43 2011. URL http://www.loc.
gov/catdir/enhancements/fy1202/
2011928368-d.html;   http://www.
loc.gov/catdir/enhancements/
fy1202/2011928368-t.html;
https://link.springer.com/book/
10.1007/978-94-007-1605-6.

Anonymous:2011:FMh

[Ano11-40] Anonymous. Front matter. In Brazilian Studies in Philosophy and History of Science: An account of recent works [KV11], pages i–xiii. ISBN 90-481-9421-0 (hardcover), 90-481-9422-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175.3 .B73 2011.

                                              URL https://link.springer.com/
                                              book/10.1007/978-90-481-9422-3.

Anonymous:2012:BMa

[Ano12a] Anonymous. Back matter. In Mauskopf and Schmaltz [MS12b], pages 241–249. ISBN 94-007-1744-X, 94-007-1745-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175 .I5134 2011. URL https://link.

                                              springer.com/book/10.1007/978-
                                              94-007-1745-9.

Anonymous:2012:BMb

[Ano12b] Anonymous. Back matter. In Physical Systems: Conceptual Pathways between Flat Space–time and Matter [Bel12g], pages 227–234. ISBN 94-007-2372-5, 94-007-2373-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL

                                              https://link.springer.com/book/
                                              10.1007/978-94-007-2373-3.

Anonymous:2012:BMc

[Ano12c] Anonymous. Back matter. In Philosophy of Behavioral Biology [PR12b], pages 303–338. ISBN 94-007-1950-7, 94-007-1951-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QH331 .P4679 2012; Q175 .B657 v.282.

                                              URL https://link.springer.com/
                                              book/10.1007/978-94-007-1951-4.

Anonymous:2012:BMd

[Ano12d] Anonymous. Back matter. In Philosophy of Behavioral Biology [PR12b], pages 367–373. ISBN 94-007-1950-7, 94-007-1951-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QH331 .P4679 2012; Q175 .B657 v.282.

URL https://link.springer.com/
book/10.1007/978-94-007-1951-4.

Anonymous:2012:BMe

[Ano12e] Anonymous. Back matter. In Interpreting Physics: Language and the Classical/Quantum Divide [Mac12d], pages 259–268. ISBN 94-007-2368-7, 94-007-2369-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC174.13 .M335 2012. URL

https://link.springer.com/book/
10.1007/978-94-007-2369-6.

Anonymous:2012:BMf

[Ano12f] Anonymous. Back matter. In The Development of Quine’s Philosophy [Mur12a], pages 247–283. ISBN 94-007-2423-3, 94-007-2424-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN B945.Q54 M87 2012.

URL https://link.springer.com/
book/10.1007/978-94-007-2424-2.

Anonymous:2012:BMg

[Ano12g] Anonymous. Back matter. In Characterizing the Robustness of Science: After the Practice Turn in Philosophy of Science [STNW12], pages 361–372. ISBN 94-007-2758-5, 94-007-2759-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175 .C4595 2012. URL https://link.

springer.com/book/10.1007/978-
94-007-2759-5.

Anonymous:2012:BMh

[Ano12h] Anonymous. Back matter. In Is Water H_{2}O?: Evidence, Realism and Pluralism [Cha12e], pages 303–316. ISBN 94-007-3931-1, 94-007-3932-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QD169.W3 C43 2012.

URL https://link.springer.com/
book/10.1007/978-94-007-3932-1.

Anonymous:2012:BMi

[Ano12i] Anonymous. Back matter. In The Material Realization of Science: From Habermas to Experimentation and Referential Realism [Rad12d], pages 155–205. ISBN 94-007-4106-5 (hardcover), 94-007-4107-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175 .R33 2012.

URL http://www.loc.gov/catdir/
enhancements/fy1313/2012935860-
d.html;   http://www.loc.gov/
catdir/enhancements/fy1313/
2012935860-t.html;   https://
link.springer.com/book/10.1007/
978-94-007-4107-2.

Anonymous:2012:BMj

[Ano12j] Anonymous. Back matter. In Evaluating Philosophies [Bun12h], pages 185–200. ISBN 94-007-4407-2, 94-007-4408-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 v. 295; Q175 .B8223 2012.

URL http://www.loc.gov/catdir/
enhancements/fy1506/2012941366-
d.html;   http://www.loc.gov/
catdir/enhancements/fy1506/
2012941366-t.html;   https://
link.springer.com/book/10.1007/
978-94-007-4408-0.

Anonymous:2012:FMa

[Ano12k] Anonymous. Front matter. In Mauskopf and Schmaltz [MS12b], pages i–x. ISBN 94-007-1744-X, 94-007-1745-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175 .I5134 2011.

URL https://link.springer.com/
book/10.1007/978-94-007-1745-9.

Anonymous:2012:FMb

[Ano12l] Anonymous. Front matter. In Mauskopf and Schmaltz [MS12b], page 11. ISBN 94-007-1744-X, 94-007-1745-8 (e-book).

ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175 .I5134 2011.

URL https://link.springer.com/
book/10.1007/978-94-007-1745-9.

Anonymous:2012:FMc

[Ano12m] Anonymous. Front matter. In Mauskopf and Schmaltz [MS12b], page 107. ISBN 94-007-1744-X, 94-007-1745-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175 .I5134 2011.

URL https://link.springer.com/
book/10.1007/978-94-007-1745-9.

Anonymous:2012:FMd

[Ano12n] Anonymous. Front matter. In Physical Systems: Conceptual Pathways between Flat Space–time and Matter [Bel12g], pages i–xiii. ISBN 94-007-2372-5, 94-007-2373-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL

https://link.springer.com/book/
10.1007/978-94-007-2373-3.

Anonymous:2012:FMe

[Ano12o] Anonymous. Front matter. In Philosophy of Behavioral Biology [PR12b], pages i–viii. ISBN 94-007-1950-7, 94-007-1951-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QH331 .P4679 2012; Q175 .B657 v.282.

URL https://link.springer.com/
book/10.1007/978-94-007-1951-4.

Anonymous:2012:FMf

[Ano12p] Anonymous. Front matter. In Philosophy of Behavioral Biology [PR12b], pages i–xviii. ISBN 94-007-1950-7, 94-007-1951-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QH331 .P4679 2012; Q175 .B657 v.282.

URL https://link.springer.com/
book/10.1007/978-94-007-1951-4.

See erratum [GR12].

Anonymous:2012:FMg

[Ano12q] Anonymous. Front matter. In Philosophy of Behavioral Biology [PR12b], page 1. ISBN 94-007-1950-7, 94-007-1951-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QH331 .P4679 2012; Q175 .B657 v.282.

                                              URL https://link.springer.com/
                                              book/10.1007/978-94-007-1951-4.

Anonymous:2012:FMh

[Ano12r] Anonymous. Front matter. In Philosophy of Behavioral Biology [PR12b], page 1. ISBN 94-007-1950-7, 94-007-1951-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QH331 .P4679 2012; Q175 .B657 v.282.

                                              URL https://link.springer.com/
                                              book/10.1007/978-94-007-1951-4.

Anonymous:2012:FMi

[Ano12s] Anonymous. Front matter. In Philosophy of Behavioral Biology [PR12b], page 41. ISBN 94-007-1950-7, 94-007-1951-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QH331 .P4679 2012; Q175 .B657 v.282.

                                              URL https://link.springer.com/
                                              book/10.1007/978-94-007-1951-4.

Anonymous:2012:FMj

[Ano12t] Anonymous. Front matter. In Philosophy of Behavioral Biology [PR12b], page 83. ISBN 94-007-1950-7, 94-007-1951-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QH331 .P4679 2012; Q175 .B657 v.282.

                                              URL https://link.springer.com/
                                              book/10.1007/978-94-007-1951-4.

Anonymous:2012:FMk

[Ano12u] Anonymous. Front matter. In Philosophy of Behavioral Biology [PR12b], page 97. ISBN 94-007-1950-7, 94-007-1951-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QH331 .P4679 2012; Q175 .B657 v.282.

URL https://link.springer.com/
book/10.1007/978-94-007-1951-4.

Anonymous:2012:FMl

[Ano12v] Anonymous. Front matter. In Philosophy of Behavioral Biology [PR12b], page 147. ISBN 94-007-1950-7, 94-007-1951-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QH331 .P4679 2012; Q175 .B657 v.282.

URL https://link.springer.com/
book/10.1007/978-94-007-1951-4.

Anonymous:2012:FMm

[Ano12w] Anonymous. Front matter. In Philosophy of Behavioral Biology [PR12b], page 189. ISBN 94-007-1950-7, 94-007-1951-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QH331 .P4679 2012; Q175 .B657 v.282.

URL https://link.springer.com/
book/10.1007/978-94-007-1951-4.

Anonymous:2012:FMn

[Ano12x] Anonymous. Front matter. In Philosophy of Behavioral Biology [PR12b], page 315. ISBN 94-007-1950-7, 94-007-1951-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QH331 .P4679 2012; Q175 .B657 v.282.

URL https://link.springer.com/
book/10.1007/978-94-007-1951-4.

Anonymous:2012:FMo

[Ano12y] Anonymous. Front matter. In Interpreting Physics: Language and the Classical/Quantum Divide [Mac12d], pages i–xiii. ISBN 94-007-2368-7, 94-007-2369-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC174.13 .M335 2012. URL

https://link.springer.com/book/
10.1007/978-94-007-2369-6.

Anonymous:2012:FMp

[Ano12z] Anonymous. Front matter. In Interpreting Physics: Language and the Classical/Quantum Divide [Mac12d], page 1. ISBN 94-007-2368-7, 94-007-2369-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC174.13 .M335 2012. URL https:/

                                              /link.springer.com/book/10.1007/
                                              978-94-007-2369-6.

Anonymous:2012:FMq

[Ano12-27] Anonymous. Front matter. In Interpreting Physics: Language and the Classical/Quantum Divide [Mac12d], pages 121–127. ISBN 94-007-2368-7, 94-007-2369-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC174.13 .M335 2012. URL

                                              https://link.springer.com/book/
                                              10.1007/978-94-007-2369-6.

Anonymous:2012:FMr

[Ano12-28] Anonymous. Front matter. In The Development of Quine’s Philosophy [Mur12a], pages i–xxv. ISBN 94-007-2423-3, 94-007-2424-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN B945.Q54 M87 2012.

                                              URL https://link.springer.com/
                                              book/10.1007/978-94-007-2424-2.

Anonymous:2012:FMs

[Ano12-29] Anonymous. Front matter. In Characterizing the Robustness of Science: After the Practice Turn in Philosophy of Science [STNW12], pages i–x. ISBN 94-007-2758-5, 94-007-2759-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175 .C4595 2012.

                                              URL https://link.springer.com/
                                              book/10.1007/978-94-007-2759-5.

Anonymous:2012:FMt

[Ano12-30] Anonymous. Front matter. In Is Water H_{2}O?: Evidence, Realism and Pluralism [Cha12e], pages i–xxi. ISBN 94-007-3931-1, 94-007-3932-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QD169.W3 C43 2012.

URL https://link.springer.com/
book/10.1007/978-94-007-3932-1.

Anonymous:2012:FMu

[Ano12-31] Anonymous. Front matter. In The Material Realization of Science: From Habermas to Experimentation and Referential Realism [Rad12d], pages i–xix. ISBN 94-007-4106-5 (hardcover), 94-007-4107-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175 .R33 2012.

URL http://www.loc.gov/catdir/
enhancements/fy1313/2012935860-
d.html;   http://www.loc.gov/
catdir/enhancements/fy1313/
2012935860-t.html;   https://
link.springer.com/book/10.1007/
978-94-007-4107-2.

Anonymous:2012:FMv

[Ano12-32] Anonymous. Front matter. In The Material Realization of Science: From Habermas to Experimentation and Referential Realism [Rad12d], page 1. ISBN 94-007-4106-5 (hardcover), 94-007-4107-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175 .R33 2012.

URL http://www.loc.gov/catdir/
enhancements/fy1313/2012935860-
d.html;   http://www.loc.gov/
catdir/enhancements/fy1313/
2012935860-t.html;   https://
link.springer.com/book/10.1007/
978-94-007-4107-2.

Anonymous:2012:FMw

[Ano12-33] Anonymous. Front matter. In The Material Realization of Science: From Habermas to Experimentation and Referential Realism [Rad12d], page 51. ISBN 94-007-4106-5 (hardcover), 94-007-4107-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175 .R33 2012.

URL http://www.loc.gov/catdir/
enhancements/fy1313/2012935860-
d.html;   http://www.loc.gov/
catdir/enhancements/fy1313/
2012935860-t.html;   https://
link.springer.com/book/10.1007/
978-94-007-4107-2.

Anonymous:2012:FMx

[Ano12-34] Anonymous. Front matter. In Evaluating Philosophies [Bun12h], pages i–xiv. ISBN 94-007-4407-2, 94-007-4408-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 v. 295; Q175 .B8223 2012.

URL http://www.loc.gov/catdir/
enhancements/fy1506/2012941366-
d.html;   http://www.loc.gov/
catdir/enhancements/fy1506/
2012941366-t.html;   https://
link.springer.com/book/10.1007/
978-94-007-4408-0.

Anonymous:2012:FMy

[Ano12-35] Anonymous. Front matter. In Evaluating Philosophies [Bun12h], page 1. ISBN 94-007-4407-2, 94-007-4408-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 v. 295; Q175 .B8223 2012.

URL http://www.loc.gov/catdir/
enhancements/fy1506/2012941366-
d.html;   http://www.loc.gov/
catdir/enhancements/fy1506/
2012941366-t.html;   https://

link.springer.com/book/10.1007/
978-94-007-4408-0.

Anonymous:2012:FMz

[Ano12-36] Anonymous. Front matter. In Evaluating Philosophies [Bun12h], page 49. ISBN 94-007-4407-2, 94-007-4408-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 v. 295; Q175 .B8223 2012.

URL http://www.loc.gov/catdir/
enhancements/fy1506/2012941366-
d.html;   http://www.loc.gov/
catdir/enhancements/fy1506/
2012941366-t.html;   https://
link.springer.com/book/10.1007/
978-94-007-4408-0.

Anonymous:2012:FMaa

[Ano12-37] Anonymous. Front matter. In Evaluating Philosophies [Bun12h], page 103. ISBN 94-007-4407-2, 94-007-4408-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 v. 295; Q175 .B8223 2012.

URL http://www.loc.gov/catdir/
enhancements/fy1506/2012941366-
d.html;   http://www.loc.gov/
catdir/enhancements/fy1506/
2012941366-t.html;   https://
link.springer.com/book/10.1007/
978-94-007-4408-0.

Anonymous:2013:BMb

[Ano13a] Anonymous. Back matter. In Bergandi [Ber13c], pages 159–179. ISBN 94-007-5066-8, 94-007-5067-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN GE42 .S77 2013; Q175 .B73 v. 296. URL https://

link.springer.com/book/10.1007/
978-94-007-5067-8.

Anonymous:2013:BMc

[Ano13b] Anonymous. Back matter. In The Very Idea of Modern Science: Francis Bacon and Robert Boyle [Aga13q], pages 265–315. ISBN 94-007-5350-0, 94-007-5351-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175 .A318 2013; Q174 .B67 v. 298.

                                              URL https://link.springer.com/
                                              book/10.1007/978-94-007-5351-8.

Anonymous:2013:BMd

[Ano13c] Anonymous. Back matter. In Garber and Roux [GR13b], pages 303–338. ISBN 94-007-4344-0 (hardcover), 94-007-4345-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC6 .M3616 2013; QC6 .M43 2013.

                                              URL https://link.springer.com/
                                              book/10.1007/978-94-007-4345-8.

Anonymous:2013:FMh

[Ano13d] Anonymous. Front matter. In Bergandi [Ber13c], pages i–xv. ISBN 94-007-5066-8, 94-007-5067-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN GE42 .S77 2013; Q175

                                              .B73 v. 296. URL https://link.
                                              springer.com/book/10.1007/978-
                                              94-007-5067-8.

Anonymous:2013:FMi

[Ano13e] Anonymous. Front matter. In The Very Idea of Modern Science: Francis Bacon and Robert Boyle [Aga13q], pages i–xvii. ISBN 94-007-5350-0, 94-007-5351-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175 .A318 2013; Q174 .B67 v. 298.

                                              URL https://link.springer.com/
                                              book/10.1007/978-94-007-5351-8.

Anonymous:2013:FMj

[Ano13f] Anonymous. Front matter. In The Very Idea of Modern Science: Francis Bacon and Robert Boyle [Aga13q], page 1. ISBN 94-007-5350-0, 94-007-5351-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175 .A318 2013; Q174 .B67 v. 298. URL

https://link.springer.com/book/
10.1007/978-94-007-5351-8.

Anonymous:2013:FMk

[Ano13g] Anonymous. Front matter. In The Very Idea of Modern Science: Francis Bacon and Robert Boyle [Aga13q], pages 125–130. ISBN 94-007-5350-0, 94-007-5351-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175 .A318 2013; Q174 .B67 v. 298.

URL https://link.springer.com/
book/10.1007/978-94-007-5351-8.

Anonymous:2013:FMl

[Ano13h] Anonymous. Front matter. In Garber and Roux [GR13b], pages i–xviii. ISBN 94-007-4344-0 (hardcover), 94-007-4345-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC6 .M3616 2013; QC6 .M43 2013.

URL https://link.springer.com/
book/10.1007/978-94-007-4345-8.

Anonymous:2013:FMm

[Ano13i] Anonymous. Front matter. In Garber and Roux [GR13b], page 1. ISBN 94-007-4344-0 (hardcover), 94-007-4345-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC6 .M3616 2013; QC6 .M43 2013. URL

https://link.springer.com/book/
10.1007/978-94-007-4345-8.

Anonymous:2013:FMn

[Ano13j] Anonymous. Front matter. In Garber and Roux [GR13b], page 97. ISBN 94-007-4344-0 (hardcover), 94-007-4345-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC6 .M3616 2013; QC6 .M43 2013. URL

https://link.springer.com/book/
10.1007/978-94-007-4345-8.

Anonymous:2013:FMo

[Ano13k] Anonymous. Front matter. In Garber and Roux [GR13b], page 189. ISBN 94-007-4344-0 (hardcover), 94-007-4345-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC6 .M3616 2013; QC6 .M43 2013. URL

                                              https://link.springer.com/book/
                                              10.1007/978-94-007-4345-8.

Anonymous:2014:BMc

[Ano14a] Anonymous. Back matter. In History of Artificial Cold, Scientific, Technological and Cultural Issues [Gav14b], pages 281–288. ISBN 94-007-7198-3 (hardcover), 94-007-7199-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC278 H57 2014; TP494 .H577 2014. URL

                                              https://link.springer.com/book/
                                              10.1007/978-94-007-7199-4.

Anonymous:2014:BMb

[Ano14b] Anonymous. Back matter. In Bernard and Proust [BP14b], pages 325–334. ISBN 94-007-5121-4 (hardcover), 94-007-5122-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q181 .S35 2014; Q181 S387 2014. URL

                                              https://link.springer.com/book/
                                              10.1007/978-94-007-5122-4.

Anonymous:2014:BMd

[Ano14c] Anonymous. Back matter. In The Institution of Science and the Science of Institutions: The Legacy of Joseph Ben-David [Her14b], pages 213–227. ISBN 94-007-7406-0 (hardcover), 94-007-7407-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 v.302; Q125 .I58 2014. URL

                                              https://link.springer.com/book/
                                              10.1007/978-94-007-7407-0.

Anonymous:2014:BMa

[Ano14d] Anonymous. Back matter. In Felix Kaufmann’s Theory and Method in the Social Sciences [CH14], pages 355–357. ISBN 3-319-02844-8 (hardcover), 3-319-02845-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN H61 .K313 2014. URL

https://link.springer.com/book/
10.1007/978-3-319-02845-3.

Anonymous:2014:FMh

[Ano14e] Anonymous. Front matter. In History of Artificial Cold, Scientific, Technological and Cultural Issues [Gav14b], pages i–xii. ISBN 94-007-7198-3 (hardcover), 94-007-7199-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC278 H57 2014; TP494 .H577 2014. URL

https://link.springer.com/book/
10.1007/978-94-007-7199-4.

Anonymous:2014:FMi

[Ano14f] Anonymous. Front matter. In History of Artificial Cold, Scientific, Technological and Cultural Issues [Gav14b], page 19. ISBN 94-007-7198-3 (hardcover), 94-007-7199-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC278 H57 2014; TP494 .H577 2014.

URL https://link.springer.com/
book/10.1007/978-94-007-7199-4.

Anonymous:2014:FMj

[Ano14g] Anonymous. Front matter. In History of Artificial Cold, Scientific, Technological and Cultural Issues [Gav14b], page 133. ISBN 94-007-7198-3 (hardcover), 94-007-7199-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC278 H57 2014; TP494 .H577 2014.

URL https://link.springer.com/
book/10.1007/978-94-007-7199-4.

Anonymous:2014:FMk

[Ano14h] Anonymous. Front matter. In History of Artificial Cold, Scientific, Technological and Cultural Issues [Gav14b], page 199. ISBN 94-007-7198-3 (hardcover), 94-007-7199-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC278 H57 2014; TP494 .H577 2014.

                                              URL https://link.springer.com/
                                              book/10.1007/978-94-007-7199-4.

Anonymous:2014:FMb

[Ano14i] Anonymous. Front matter. In Bernard and Proust [BP14b], pages i–ix. ISBN 94-007-5121-4 (hardcover), 94-007-5122-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q181 .S35 2014; Q181 S387 2014. URL

                                              https://link.springer.com/book/
                                              10.1007/978-94-007-5122-4.

Anonymous:2014:FMc

[Ano14j] Anonymous. Front matter. In Bernard and Proust [BP14b], page 17. ISBN 94-007-5121-4 (hardcover), 94-007-5122-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q181 .S35 2014; Q181 S387 2014. URL

                                              https://link.springer.com/book/
                                              10.1007/978-94-007-5122-4.

Anonymous:2014:FMd

[Ano14k] Anonymous. Front matter. In Bernard and Proust [BP14b], page 67. ISBN 94-007-5121-4 (hardcover), 94-007-5122-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q181 .S35 2014; Q181 S387 2014. URL

                                              https://link.springer.com/book/
                                              10.1007/978-94-007-5122-4.

Anonymous:2014:FMe

[Ano14l] Anonymous. Front matter. In Bernard and Proust [BP14b], page 123. ISBN 94-007-5121-4 (hardcover), 94-007-5122-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q181 .S35 2014; Q181 S387 2014. URL

https://link.springer.com/book/
10.1007/978-94-007-5122-4.

Anonymous:2014:FMf

[Ano14m] Anonymous. Front matter. In Bernard and Proust [BP14b], page 215. ISBN 94-007-5121-4 (hardcover), 94-007-5122-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q181 .S35 2014; Q181 S387 2014. URL

https://link.springer.com/book/
10.1007/978-94-007-5122-4.

Anonymous:2014:FMg

[Ano14n] Anonymous. Front matter. In Bernard and Proust [BP14b], page 303. ISBN 94-007-5121-4 (hardcover), 94-007-5122-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q181 .S35 2014; Q181 S387 2014. URL

https://link.springer.com/book/
10.1007/978-94-007-5122-4.

Anonymous:2014:FMl

[Ano14o] Anonymous. Front matter. In The Institution of Science and the Science of Institutions: The Legacy of Joseph Ben-David [Her14b], pages I–XII. ISBN 94-007-7406-0 (hardcover), 94-007-7407-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 v.302; Q125 .I58 2014. URL

https://link.springer.com/book/
10.1007/978-94-007-7407-0.

Anonymous:2014:FMm

[Ano14p] Anonymous. Front matter. In The Institution of Science and the Science of Institutions: The Legacy of Joseph Ben-David [Her14b], page 1. ISBN 94-007-7406-0 (hardcover), 94-007-7407-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 v.302; Q125 .I58 2014. URL

                                              https://link.springer.com/book/
                                              10.1007/978-94-007-7407-0.

Anonymous:2014:FMn

[Ano14q] Anonymous. Front matter. In The Institution of Science and the Science of Institutions: The Legacy of Joseph Ben-David [Her14b], page 37. ISBN 94-007-7406-0 (hardcover), 94-007-7407-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 v.302; Q125 .I58 2014. URL

                                              https://link.springer.com/book/
                                              10.1007/978-94-007-7407-0.

Anonymous:2014:FMo

[Ano14r] Anonymous. Front matter. In The Institution of Science and the Science of Institutions: The Legacy of Joseph Ben-David [Her14b], page 81. ISBN 94-007-7406-0 (hardcover), 94-007-7407-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 v.302; Q125 .I58 2014. URL

                                              https://link.springer.com/book/
                                              10.1007/978-94-007-7407-0.

Anonymous:2014:FMp

[Ano14s] Anonymous. Front matter. In The Institution of Science and the Science of Institutions: The Legacy of Joseph Ben-David [Her14b], page 145. ISBN 94-007-7406-0 (hardcover), 94-007-7407-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 v.302; Q125 .I58 2014. URL

                                              https://link.springer.com/book/
                                              10.1007/978-94-007-7407-0.

Anonymous:2014:FMq

[Ano14t] Anonymous. Front matter. In The Institution of Science and the Science of Institutions: The Legacy of Joseph Ben-David [Her14b], page 185. ISBN 94-007-7406-0 (hardcover), 94-007-7407-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 v.302; Q125 .I58 2014. URL

https://link.springer.com/book/
10.1007/978-94-007-7407-0.

Anonymous:2014:FMa

[Ano14u] Anonymous. Front matter. In Felix Kaufmann’s Theory and Method in the Social Sciences [CH14], pages i–x. ISBN 3-319-02844-8 (hardcover), 3-319-02845-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN H61 .K313 2014. URL

https://link.springer.com/book/
10.1007/978-3-319-02845-3.

Anonymous:2014:FMr

[Ano14v] Anonymous. Front matter. In The History of Physics in Cuba [BRW14a], pages i–xxvi. ISBN 94-017-8040-4, 94-017-8041-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC5.53; QC6.9; QC9 C9H57 2014; Q124.6-127.2. URL

https://link.springer.com/book/
10.1007/978-94-017-8041-4.

Anonymous:2014:FMs

[Ano14w] Anonymous. Front matter. In The History of Physics in Cuba [BRW14a], page 1. ISBN 94-017-8040-4, 94-017-8041-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC5.53; QC6.9; QC9 C9H57 2014; Q124.6-127.2. URL https://link.

springer.com/book/10.1007/978-
94-017-8041-4.

Anonymous:2014:FMt

[Ano14x] Anonymous. Front matter. In The History of Physics in Cuba [BRW14a], page 55. ISBN 94-017-8040-4, 94-017-8041-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC5.53; QC6.9; QC9 C9H57 2014; Q124.6-127.2. URL https://link.

springer.com/book/10.1007/978-
94-017-8041-4.

Anonymous:2014:FMu

[Ano14y] Anonymous. Front matter. In The History of Physics in Cuba [BRW14a], page 277. ISBN 94-017-8040-4, 94-017-8041-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC5.53; QC6.9; QC9 C9H57 2014; Q124.6-127.2.

URL https://link.springer.com/
book/10.1007/978-94-017-8041-4.

Anonymous:2014:FMv

[Ano14z] Anonymous. Front matter. In The History of Physics in Cuba [BRW14a], page 351. ISBN 94-017-8040-4, 94-017-8041-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC5.53; QC6.9; QC9 C9H57 2014; Q124.6-127.2.

URL https://link.springer.com/
book/10.1007/978-94-017-8041-4.

Anonymous:2014:FMw

[Ano14-27] Anonymous. Front matter. In The History of Physics in Cuba [BRW14a], page 421. ISBN 94-017-8040-4, 94-017-8041-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC5.53; QC6.9; QC9 C9H57 2014; Q124.6-127.2.

URL https://link.springer.com/
book/10.1007/978-94-017-8041-4.

Anonymous:2015:BMa

[Ano15a] Anonymous. Back matter. In The Neurological Emergence of Epilepsy: The National Hospital for the Paralysed and Epileptic (1870–1895) [Lek15e], pages 187–209. ISBN 3-319-06292-1, 3-319-06293-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic).

LCCN Q174 .B67 vol.305. URL

https://link.springer.com/book/
10.1007/978-3-319-06293-8.

Anonymous:2015:BMe

[Ano15b] Anonymous. Back matter. In Conceptual Change in Biology: Scientific and Philosophical Perspectives on Evolution and Development [Lov15b], pages 483– 490. ISBN 94-017-9411-1, 94-017-9412-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QH331 C66 2015; QH491; QH492 .C66 2014.

URL https://link.springer.com/
book/10.1007/978-94-017-9412-1.

Anonymous:2015:BMf

[Ano15c] Anonymous. Back matter. In Sciences in the Universities of Europe, Nineteenth and Twentieth Centuries: Academic Landscapes [SDG15b], pages 381–393. ISBN 94-017-9635-1, 94-017-9636-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q183.4.E85 2015. URL

https://link.springer.com/book/
10.1007/978-94-017-9636-1.

Anonymous:2015:BMb

[Ano15d] Anonymous. Back matter. In Kuhn’s Structure of Scientific Revolutions — 50 Years On [DB15b], pages 197–199. ISBN 3-319-13382-9, 3-319-13383-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175.3

.K84 2015.  URL https://link.
springer.com/book/10.1007/978-3-
319-13383-6.

Anonymous:2015:BMc

[Ano15e] Anonymous. Back matter. In The electric theories of J. Clerk Maxwell: a historical and critical study [Duh15c], pages 175–185. ISBN 3-319-18514-4 (hardcover), 3-319-18515-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC518 .D84 2015.

                                              URL https://link.springer.com/
                                              book/10.1007/978-3-319-18515-6.

English translation by Alan Aversa.

Anonymous:2015:BMd

[Ano15f] Anonymous. Back matter. In New Perspectives on Technology, Values, and Ethics: Theoretical and Practical [Gon15a], pages 197–219. ISBN 3-319-21869-7 (hardcover), 3-319-21870-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

                                              URL https://link.springer.com/
                                              book/10.1007/978-3-319-21870-0.

Anonymous:2015:FMa

[Ano15g] Anonymous. Front matter. In The Neurological Emergence of Epilepsy: The National Hospital for the Paralysed and Epileptic (1870–1895) [Lek15e], pages i–xiii. ISBN 3-319-06292-1, 3-319-06293-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol.305. URL

                                              https://link.springer.com/book/
                                              10.1007/978-3-319-06293-8.

Anonymous:2015:FMad

[Ano15h] Anonymous. Front matter. In Philosophy of Chemistry: Growth of a New Discipline [SM15b], pages i–xii. ISBN 94-017-9363-8 (hardcover), 94-017-9364-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QD6 .P455 2015; QD6 .S384 2014.

                                              URL https://link.springer.com/
                                              book/10.1007/978-94-017-9364-3.

Anonymous:2015:FMae

[Ano15i] Anonymous. Front matter. In Conceptual Change in Biology: Scientific and Philosophical Perspectives on Evolution and Development [Lov15b], pages i– xvii. ISBN 94-017-9411-1, 94-017-9412-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QH331 C66 2015; QH491; QH492 .C66 2014.

URL https://link.springer.com/
book/10.1007/978-94-017-9412-1.

Anonymous:2015:FMaf

[Ano15j] Anonymous. Front matter. In Conceptual Change in Biology: Scientific and Philosophical Perspectives on Evolution and Development [Lov15b], page 55. ISBN 94-017-9411-1, 94-017-9412-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QH331 C66 2015; QH491; QH492 .C66 2014.

URL https://link.springer.com/
book/10.1007/978-94-017-9412-1.

Anonymous:2015:FMag

[Ano15k] Anonymous. Front matter. In Conceptual Change in Biology: Scientific and Philosophical Perspectives on Evolution and Development [Lov15b], page 129. ISBN 94-017-9411-1, 94-017-9412-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QH331 C66 2015; QH491; QH492 .C66 2014.

URL https://link.springer.com/
book/10.1007/978-94-017-9412-1.

Anonymous:2015:FMah

[Ano15l] Anonymous. Front matter. In Conceptual Change in Biology: Scientific and Philosophical Perspectives on Evolution and Development [Lov15b], page 219. ISBN 94-017-9411-1, 94-017-9412-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QH331 C66 2015; QH491; QH492 .C66 2014.

URL https://link.springer.com/
book/10.1007/978-94-017-9412-1.

Anonymous:2015:FMai

[Ano15m] Anonymous. Front matter. In Conceptual Change in Biology: Scientific and Philosophical Perspectives on Evolution and Development [Lov15b], page 303. ISBN 94-017-9411-1, 94-017-9412-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QH331 C66 2015; QH491; QH492 .C66 2014.

                                              URL https://link.springer.com/
                                              book/10.1007/978-94-017-9412-1.

Anonymous:2015:FMaj

[Ano15n] Anonymous. Front matter. In Conceptual Change in Biology: Scientific and Philosophical Perspectives on Evolution and Development [Lov15b], page 403. ISBN 94-017-9411-1, 94-017-9412-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QH331 C66 2015; QH491; QH492 .C66 2014.

                                              URL https://link.springer.com/
                                              book/10.1007/978-94-017-9412-1.

Anonymous:2015:FMb

[Ano15o] Anonymous. Front matter. In From Logic to Practice: Italian Studies in the Philosophy of Mathematics [LPV15], pages i–xix. ISBN 3-319-10433-0, 3-319-10434-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QA8.4 .F76 2015; Q174 .B67 v.308.

                                              URL https://link.springer.com/
                                              book/10.1007/978-3-319-10434-8.

Anonymous:2015:FMc

[Ano15p] Anonymous. Front matter. In From Logic to Practice: Italian Studies in the Philosophy of Mathematics [LPV15], page 1. ISBN 3-319-10433-0, 3-319-10434-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QA8.4 .F76 2015; Q174 .B67 v.308.

                                              URL https://link.springer.com/
                                              book/10.1007/978-3-319-10434-8.

Anonymous:2015:FMd

[Ano15q] Anonymous. Front matter. In From Logic to Practice: Italian Studies in the Philosophy of Mathematics [LPV15], page 51. ISBN 3-319-10433-0, 3-319-10434-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QA8.4 .F76 2015; Q174 .B67 v.308.

URL https://link.springer.com/
book/10.1007/978-3-319-10434-8.

Anonymous:2015:FMe

[Ano15r] Anonymous. Front matter. In From Logic to Practice: Italian Studies in the Philosophy of Mathematics [LPV15], page 201. ISBN 3-319-10433-0, 3-319-10434-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QA8.4 .F76 2015; Q174 .B67 v.308.

URL https://link.springer.com/
book/10.1007/978-3-319-10434-8.

Anonymous:2015:FMak

[Ano15s] Anonymous. Front matter. In Sciences in the Universities of Europe, Nineteenth and Twentieth Centuries: Academic Landscapes [SDG15b], pages i–x. ISBN 94-017-9635-1, 94-017-9636-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q183.4.E85 2015. URL https://link.

springer.com/book/10.1007/978-
94-017-9636-1.

Anonymous:2015:FMal

[Ano15t] Anonymous. Front matter. In Sciences in the Universities of Europe, Nineteenth and Twentieth Centuries: Academic Landscapes [SDG15b], page 9. ISBN 94-017-9635-1, 94-017-9636-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q183.4.E85 2015. URL https://link.

springer.com/book/10.1007/978-
94-017-9636-1.

Anonymous:2015:FMam

[Ano15u] Anonymous. Front matter. In Sciences in the Universities of Europe, Nineteenth and Twentieth Centuries: Academic Landscapes [SDG15b], page 125. ISBN 94-017-9635-1, 94-017-9636-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q183.4.E85 2015. URL https://link.

                                              springer.com/book/10.1007/978-
                                              94-017-9636-1.

Anonymous:2015:FMan

[Ano15v] Anonymous. Front matter. In Sciences in the Universities of Europe, Nineteenth and Twentieth Centuries: Academic Landscapes [SDG15b], page 191. ISBN 94-017-9635-1, 94-017-9636-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q183.4.E85 2015. URL https://link.

                                              springer.com/book/10.1007/978-
                                              94-017-9636-1.

Anonymous:2015:FMao

[Ano15w] Anonymous. Front matter. In Sciences in the Universities of Europe, Nineteenth and Twentieth Centuries: Academic Landscapes [SDG15b], page 291. ISBN 94-017-9635-1, 94-017-9636-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q183.4.E85 2015. URL https://link.

                                              springer.com/book/10.1007/978-
                                              94-017-9636-1.

Anonymous:2015:FMg

[Ano15x] Anonymous. Front matter. In Objectivity in Science: New Perspectives from Science and Technology Studies [PRT15], pages i–vi. ISBN 3-319-14348-4 (hardcover), 3-319-14349-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175 O25 2015.

                                              URL https://link.springer.com/
                                              book/10.1007/978-3-319-14349-1.

Anonymous:2015:FMh

[Ano15y] Anonymous. Front matter. In Objectivity in Science: New Perspectives from Science and Technology Studies [PRT15], page 17. ISBN 3-319-14348-4 (hardcover), 3-319-14349-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175 O25 2015.

URL https://link.springer.com/
book/10.1007/978-3-319-14349-1.

Anonymous:2015:FMi

[Ano15z] Anonymous. Front matter. In Objectivity in Science: New Perspectives from Science and Technology Studies [PRT15], page 77. ISBN 3-319-14348-4 (hardcover), 3-319-14349-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175 O25 2015.

URL https://link.springer.com/
book/10.1007/978-3-319-14349-1.

Anonymous:2015:FMj

[Ano15-27] Anonymous. Front matter. In Objectivity in Science: New Perspectives from Science and Technology Studies [PRT15], page 171. ISBN 3-319-14348-4 (hardcover), 3-319-14349-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175 O25 2015.

URL https://link.springer.com/
book/10.1007/978-3-319-14349-1.

Anonymous:2015:FMf

[Ano15-28] Anonymous. Front matter. In Kuhn’s Structure of Scientific Revolutions — 50 Years On [DB15b], pages i–xi. ISBN 3-319-13382-9, 3-319-13383-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175.3 .K84 2015.

URL https://link.springer.com/
book/10.1007/978-3-319-13383-6.

Anonymous:2015:FMk

[Ano15-29] Anonymous. Front matter. In Arabatzis et al. [ARS15], pages i–vii. ISBN 3-319-14552-5 (paperback), 3-319-14553-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q125

.R45 2015. URL http://www.loc.
gov/catdir/enhancements/fy1516/
2015936523-d.html;   http://www.
loc.gov/catdir/enhancements/
fy1516/2015936523-t.html;
https://link.springer.com/book/
10.1007/978-3-319-14553-2.

Anonymous:2015:FMl

[Ano15-30] Anonymous. Front matter. In Arabatzis et al. [ARS15], page 7. ISBN 3-319-14552-5 (paperback), 3-319-14553-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q125

.R45 2015. URL http://www.loc.
gov/catdir/enhancements/fy1516/
2015936523-d.html;   http://www.
loc.gov/catdir/enhancements/
fy1516/2015936523-t.html;
https://link.springer.com/book/
10.1007/978-3-319-14553-2.

Anonymous:2015:FMm

[Ano15-31] Anonymous. Front matter. In Arabatzis et al. [ARS15], page 67. ISBN 3-319-14552-5 (paperback), 3-319-14553-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q125

.R45 2015. URL http://www.loc.
gov/catdir/enhancements/fy1516/
2015936523-d.html;   http://www.
loc.gov/catdir/enhancements/
fy1516/2015936523-t.html;
https://link.springer.com/book/
10.1007/978-3-319-14553-2.

Anonymous:2015:FMn

[Ano15-32] Anonymous. Front matter. In Arabatzis et al. [ARS15], page 175. ISBN 3-319-14552-5 (paperback), 3-319-14553-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q125

.R45 2015. URL http://www.loc.
gov/catdir/enhancements/fy1516/
2015936523-d.html;   http://www.
loc.gov/catdir/enhancements/
fy1516/2015936523-t.html;
https://link.springer.com/book/
10.1007/978-3-319-14553-2.

Anonymous:2015:FMo

[Ano15-33] Anonymous. Front matter. In Arabatzis et al. [ARS15], page 239. ISBN 3-319-14552-5 (paperback), 3-319-14553-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q125

.R45 2015. URL http://www.loc.
gov/catdir/enhancements/fy1516/
2015936523-d.html;   http://www.
loc.gov/catdir/enhancements/
fy1516/2015936523-t.html;
https://link.springer.com/book/
10.1007/978-3-319-14553-2.

Anonymous:2015:FMp

[Ano15-34] Anonymous. Front matter. In Arabatzis et al. [ARS15], page 313. ISBN 3-319-14552-5 (paperback), 3-319-14553-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q125

.R45 2015. URL http://www.loc.
gov/catdir/enhancements/fy1516/
2015936523-d.html;   http://www.
loc.gov/catdir/enhancements/
fy1516/2015936523-t.html;
https://link.springer.com/book/
10.1007/978-3-319-14553-2.

Anonymous:2015:FMq

[Ano15-35] Anonymous. Front matter. In Romanian Studies in Philosophy of Science [PST15], pages i–viii. ISBN 3-319-16654-9 (hardcover), 3-319-16655-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

                                              URL https://link.springer.com/
                                              book/10.1007/978-3-319-16655-1.

Anonymous:2015:FMr

[Ano15-36] Anonymous. Front matter. In Romanian Studies in Philosophy of Science [PST15], page 1. ISBN 3-319-16654-9 (hardcover), 3-319-16655-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL

                                              https://link.springer.com/book/
                                              10.1007/978-3-319-16655-1.

Anonymous:2015:FMs

[Ano15-37] Anonymous. Front matter. In Romanian Studies in Philosophy of Science [PST15], page 43. ISBN 3-319-16654-9 (hardcover), 3-319-16655-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL

                                              https://link.springer.com/book/
                                              10.1007/978-3-319-16655-1.

Anonymous:2015:FMt

[Ano15-38] Anonymous. Front matter. In Romanian Studies in Philosophy of Science [PST15], page 103. ISBN 3-319-16654-9 (hardcover), 3-319-16655-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL

                                              https://link.springer.com/book/
                                              10.1007/978-3-319-16655-1.

Anonymous:2015:FMu

[Ano15-39] Anonymous. Front matter. In Romanian Studies in Philosophy of Science [PST15], page 165. ISBN 3-319-16654-9 (hardcover), 3-319-16655-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL

                                              https://link.springer.com/book/
                                              10.1007/978-3-319-16655-1.

Anonymous:2015:FMv

[Ano15-40] Anonymous. Front matter. In Romanian Studies in Philosophy of Science [PST15], page 237. ISBN 3-319-16654-9 (hardcover), 3-319-16655-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL

https://link.springer.com/book/
10.1007/978-3-319-16655-1.

Anonymous:2015:FMw

[Ano15-41] Anonymous. Front matter. In The electric theories of J. Clerk Maxwell: a historical and critical study [Duh15c], pages i–xv. ISBN 3-319-18514-4 (hardcover), 3-319-18515-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC518 .D84 2015. URL

https://link.springer.com/book/
10.1007/978-3-319-18515-6. English

translation by Alan Aversa.

Anonymous:2015:FMx

[Ano15-42] Anonymous. Front matter. In The electric theories of J. Clerk Maxwell: a historical and critical study [Duh15c], page 11. ISBN 3-319-18514-4 (hardcover), 3-319-18515-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC518 .D84 2015. URL

https://link.springer.com/book/
10.1007/978-3-319-18515-6. English

translation by Alan Aversa.

Anonymous:2015:FMy

[Ano15-43] Anonymous. Front matter. In The electric theories of J. Clerk Maxwell: a historical and critical study [Duh15c], page 79. ISBN 3-319-18514-4 (hardcover), 3-319-18515-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC518 .D84 2015. URL

https://link.springer.com/book/
10.1007/978-3-319-18515-6. English

translation by Alan Aversa.

Anonymous:2015:FMz

[Ano15-44] Anonymous. Front matter. In New Perspectives on Technology, Values, and Ethics: Theoretical and Practical [Gon15a], pages i–xiv. ISBN 3-319-21869-7 (hardcover), 3-319-21870-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

                                              URL https://link.springer.com/
                                              book/10.1007/978-3-319-21870-0.

Anonymous:2015:FMaa

[Ano15-45] Anonymous. Front matter. In New Perspectives on Technology, Values, and Ethics: Theoretical and Practical [Gon15a], page 1. ISBN 3-319-21869-7 (hardcover), 3-319-21870-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL

                                              https://link.springer.com/book/
                                              10.1007/978-3-319-21870-0.

Anonymous:2015:FMab

[Ano15-46] Anonymous. Front matter. In New Perspectives on Technology, Values, and Ethics: Theoretical and Practical [Gon15a], page 61. ISBN 3-319-21869-7 (hardcover), 3-319-21870-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL

                                              https://link.springer.com/book/
                                              10.1007/978-3-319-21870-0.

Anonymous:2015:FMac

[Ano15-47] Anonymous. Front matter. In New Perspectives on Technology, Values, and Ethics: Theoretical and Practical [Gon15a], page 145. ISBN 3-319-21869-7 (hardcover), 3-319-21870-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

                                              URL https://link.springer.com/
                                              book/10.1007/978-3-319-21870-0.

Anonymous:2016:BM

[Ano16a] Anonymous. Back matter. In The Aristotelian Mechanics: Text and Diagrams [vL16a], pages 193–253.

ISBN 3-319-25923-7 (hardcover), 3-319-25925-3 (electronic). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QA805 .L44 2016. URL

https://link.springer.com/book/
10.1007/978-3-319-25925-3.

Anonymous:2016:FMa

[Ano16b] Anonymous. Front matter. In Cosmological and Psychological Time [DR16], pages i–xiv. ISBN 3-319-22589-8 (hardcover), 3-319-22590-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN BD638 C67 2016.

URL https://link.springer.com/
book/10.1007/978-3-319-22590-6.

Anonymous:2016:FMb

[Ano16c] Anonymous. Front matter. In Cosmological and Psychological Time [DR16], page 1. ISBN 3-319-22589-8 (hardcover), 3-319-22590-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN BD638 C67 2016.

URL https://link.springer.com/
book/10.1007/978-3-319-22590-6.

Anonymous:2016:FMc

[Ano16d] Anonymous. Front matter. In Cosmological and Psychological Time [DR16], page 73. ISBN 3-319-22589-8 (hardcover), 3-319-22590-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN BD638 C67 2016.

URL https://link.springer.com/
book/10.1007/978-3-319-22590-6.

Anonymous:2016:FMd

[Ano16e] Anonymous. Front matter. In Cosmological and Psychological Time [DR16], page 163. ISBN 3-319-22589-8 (hardcover), 3-319-22590-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN BD638 C67 2016.

URL https://link.springer.com/
book/10.1007/978-3-319-22590-6.

Anonymous:2016:FMe

[Ano16f] Anonymous. Front matter. In The Aristotelian Mechanics: Text and Diagrams [vL16a], pages i–ix. ISBN 3-319-25923-7 (hardcover), 3-319-25925-3 (electronic). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QA805 .L44 2016.

                                              URL https://link.springer.com/
                                              book/10.1007/978-3-319-25925-3.

Anonymous:2016:FMf

[Ano16g] Anonymous. Front matter. In Meta-Philosophical Reflection on Feminist Philosophies of Science [AV16a], pages i–vii. ISBN 3-319-26346-3 (hardcover), 3-319-26348-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN HQ1154 .M48 2016; HQ1190 M48 2016. URL

                                              https://link.springer.com/book/
                                              10.1007/978-3-319-26348-9.

Anonymous:2016:FMk

[Ano16h] Anonymous. Front matter. In Objectivity, Realism, and Proof: FilMat Studies in the Philosophy of Mathematics [BS16], pages i–xl. ISBN 3-319-31642-7, 3-319-31644-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QA9 .O25 2016.

                                              URL https://link.springer.com/
                                              book/10.1007/978-3-319-31644-4.

Anonymous:2016:FMl

[Ano16i] Anonymous. Front matter. In Objectivity, Realism, and Proof: FilMat Studies in the Philosophy of Mathematics [BS16], page 11. ISBN 3-319-31642-7, 3-319-31644-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QA9 .O25 2016.

                                              URL https://link.springer.com/
                                              book/10.1007/978-3-319-31644-4.

Anonymous:2016:FMm

[Ano16j] Anonymous. Front matter. In Objectivity, Realism, and Proof: FilMat Studies in the Philosophy of Mathematics [BS16], page 101. ISBN 3-319-31642-7, 3-319-31644-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QA9 .O25 2016.

URL https://link.springer.com/
book/10.1007/978-3-319-31644-4.

Anonymous:2016:FMn

[Ano16k] Anonymous. Front matter. In Objectivity, Realism, and Proof: FilMat Studies in the Philosophy of Mathematics [BS16], page 243. ISBN 3-319-31642-7, 3-319-31644-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QA9 .O25 2016.

URL https://link.springer.com/
book/10.1007/978-3-319-31644-4.

Anonymous:2016:FMg

[Ano16l] Anonymous. Front matter. In Sauer and Scholl [SS16b], pages i–viii. ISBN 3-319-30227-2, 3-319-30229-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q126.8 .P45 2016.

URL https://link.springer.com/
book/10.1007/978-3-319-30229-4.

Anonymous:2016:FMh

[Ano16m] Anonymous. Front matter. In Sauer and Scholl [SS16b], page 9. ISBN 3-319-30227-2, 3-319-30229-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q126.8 .P45 2016.

URL https://link.springer.com/
book/10.1007/978-3-319-30229-4.

Anonymous:2016:FMi

[Ano16n] Anonymous. Front matter. In Sauer and Scholl [SS16b], page 93. ISBN 3-319-30227-2, 3-319-30229-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q126.8 .P45 2016.

URL https://link.springer.com/
book/10.1007/978-3-319-30229-4.

Anonymous:2016:FMj

[Ano16o] Anonymous. Front matter. In Sauer and Scholl [SS16b], page 201. ISBN 3-319-30227-2, 3-319-30229-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q126.8 .P45 2016.

URL https://link.springer.com/
book/10.1007/978-3-319-30229-4.

Anonymous:2016:FMo

[Ano16p] Anonymous. Front matter. In A Pragmatist Orientation for the Social Sciences in Climate Policy: How to Make Integrated Economic Assessments Serve Society [Kow16j], pages i–xvii. ISBN 3-319-43279-6, 3-319-43281-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN GE170; GE220; HC79.E5; QC903 .K683 2016; QC981.8.C5 K69 2016. URL

https://link.springer.com/book/
10.1007/978-3-319-43281-6.

Anonymous:2016:FMp

[Ano16q] Anonymous. Front matter. In A Pragmatist Orientation for the Social Sciences in Climate Policy: How to Make Integrated Economic Assessments Serve Society [Kow16j], page 17. ISBN 3-319-43279-6, 3-319-43281-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN GE170; GE220; HC79.E5; QC903 .K683 2016; QC981.8.C5 K69 2016. URL

https://link.springer.com/book/
10.1007/978-3-319-43281-6.

Anonymous:2016:FMq

[Ano16r] Anonymous. Front matter. In A Pragmatist Orientation for the Social Sciences in Climate Policy: How to Make Integrated Economic Assessments Serve Society [Kow16j], page 79. ISBN 3-319-43279-6, 3-319-43281-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN GE170; GE220; HC79.E5; QC903 .K683 2016; QC981.8.C5 K69 2016. URL

https://link.springer.com/book/
10.1007/978-3-319-43281-6.

Anonymous:2016:FMr

[Ano16s] Anonymous. Front matter. In A Pragmatist Orientation for the Social Sciences in Climate Policy: How to Make Integrated Economic Assessments Serve Society [Kow16j], page 171. ISBN 3-319-43279-6, 3-319-43281-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN GE170; GE220; HC79.E5; QC903 .K683 2016; QC981.8.C5 K69 2016. URL

https://link.springer.com/book/
10.1007/978-3-319-43281-6.

Anonymous:2016:FMs

[Ano16t] Anonymous. Front matter. In A Pragmatist Orientation for the Social Sciences in Climate Policy: How to Make Integrated Economic Assessments Serve Society [Kow16j], page 273. ISBN 3-319-43279-6, 3-319-43281-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN GE170; GE220; HC79.E5; QC903 .K683 2016; QC981.8.C5 K69 2016. URL

https://link.springer.com/book/
10.1007/978-3-319-43281-6.

Anonymous:2017:BMa

[Ano17a] Anonymous. Back matter. In van Dongen and Paul [vDP17a], pages 195–198. ISBN 3-319-48892-9 (hardcover), 3-319-48893-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 v. 321; Q175.32.V57 E65 2017. URL https:/

                                              /link.springer.com/book/10.1007/
                                              978-3-319-48893-6.

Anonymous:2017:BMb

[Ano17b] Anonymous. Back matter. In Floyd and Bokulich [FB17], pages 359–361. ISBN 3-319-53278-2, 3-319-53280-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QA29.T8 P45 2017.

                                              URL https://link.springer.com/
                                              book/10.1007/978-3-319-53280-6.

Anonymous:2017:BMd

[Ano17c] Anonymous. Back matter. In Encouraging Openness: Essays for Joseph Agassi on the Occasion of His 90th Birthday [BAG17], pages 563–584. ISBN 3-319-57668-2 (hardcover), 3-319-57669-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN B833.E53 2017. URL https://link.

                                              springer.com/book/10.1007/978-3-
                                              319-57669-5.

Anonymous:2017:BMc

[Ano17d] Anonymous. Back matter. In Time of Nature and the Nature of Time: Philosophical Perspectives of Time in Natural Sciences [BH17b], pages 391– 403. ISBN 3-319-53723-7 (hardcover), 3-319-53725-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC173.59.S65 T56 2017. URL

                                              https://link.springer.com/book/
                                              10.1007/978-3-319-53725-2.

Anonymous:2017:FMa

[Ano17e] Anonymous. Front matter. In van Dongen and Paul [vDP17a], pages i–vi. ISBN 3-319-48892-9 (hardcover), 3-319-48893-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 v. 321; Q175.32.V57

                                              E65 2017.  URL https://link.

springer.com/book/10.1007/978-3-
319-48893-6.

Anonymous:2017:FMb

[Ano17f] Anonymous. Front matter. In Floyd and Bokulich [FB17], pages i–xvii. ISBN 3-319-53278-2, 3-319-53280-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QA29.T8 P45 2017.

URL https://link.springer.com/
book/10.1007/978-3-319-53280-6.

Anonymous:2017:FMc

[Ano17g] Anonymous. Front matter. In Floyd and Bokulich [FB17], page 37. ISBN 3-319-53278-2, 3-319-53280-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QA29.T8 P45 2017.

URL https://link.springer.com/
book/10.1007/978-3-319-53280-6.

Anonymous:2017:FMd

[Ano17h] Anonymous. Front matter. In Floyd and Bokulich [FB17], page 151. ISBN 3-319-53278-2, 3-319-53280-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QA29.T8 P45 2017.

URL https://link.springer.com/
book/10.1007/978-3-319-53280-6.

Anonymous:2017:FMe

[Ano17i] Anonymous. Front matter. In Floyd and Bokulich [FB17], page 263. ISBN 3-319-53278-2, 3-319-53280-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QA29.T8 P45 2017.

URL https://link.springer.com/
book/10.1007/978-3-319-53280-6.

Anonymous:2017:FMo

[Ano17j] Anonymous. Front matter. In Encouraging Openness: Essays for Joseph Agassi on the Occasion of His 90th Birthday [BAG17], pages i–xiv. ISBN 3-319-57668-2 (hardcover), 3-319-57669-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN B833.E53 2017. URL

https://link.springer.com/book/
10.1007/978-3-319-57669-5.

Anonymous:2017:FMp

[Ano17k] Anonymous. Front matter. In Encouraging Openness: Essays for Joseph Agassi on the Occasion of His 90th Birthday [BAG17], page 1. ISBN 3-319-57668-2 (hardcover), 3-319-57669-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN B833.E53 2017. URL https://link.

springer.com/book/10.1007/978-3-
319-57669-5.

Anonymous:2017:FMq

[Ano17l] Anonymous. Front matter. In Encouraging Openness: Essays for Joseph Agassi on the Occasion of His 90th Birthday [BAG17], page 69. ISBN 3-319-57668-2 (hardcover), 3-319-57669-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN B833.E53 2017. URL https://link.

springer.com/book/10.1007/978-3-
319-57669-5.

Anonymous:2017:FMr

[Ano17m] Anonymous. Front matter. In Encouraging Openness: Essays for Joseph Agassi on the Occasion of His 90th Birthday [BAG17], page 165. ISBN 3-319-57668-2 (hardcover), 3-319-57669-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN B833.E53 2017. URL https://link.

springer.com/book/10.1007/978-3-
319-57669-5.

Anonymous:2017:FMs

[Ano17n] Anonymous. Front matter. In Encouraging Openness: Essays for Joseph Agassi on the Occasion of His 90th Birthday [BAG17], page 241. ISBN 3-319-57668-2 (hardcover), 3-319-57669-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN B833.E53 2017. URL https://link.

springer.com/book/10.1007/978-3-
319-57669-5.

Anonymous:2017:FMt

[Ano17o] Anonymous. Front matter. In Encouraging Openness: Essays for Joseph Agassi on the Occasion of His 90th Birthday [BAG17], page 365. ISBN 3-319-57668-2 (hardcover), 3-319-57669-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN B833.E53 2017. URL https://link.

springer.com/book/10.1007/978-3-
319-57669-5.

Anonymous:2017:FMu

[Ano17p] Anonymous. Front matter. In Encouraging Openness: Essays for Joseph Agassi on the Occasion of His 90th Birthday [BAG17], page 491. ISBN 3-319-57668-2 (hardcover), 3-319-57669-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN B833.E53 2017. URL https://link.

springer.com/book/10.1007/978-3-
319-57669-5.

Anonymous:2017:FMf

[Ano17q] Anonymous. Front matter. In Time of Nature and the Nature of Time: Philosophical Perspectives of Time in Natural Sciences [BH17b], pages i–xiii. ISBN 3-319-53723-7 (hardcover), 3-319-53725-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC173.59.S65 T56 2017. URL

https://link.springer.com/book/
10.1007/978-3-319-53725-2.

Anonymous:2017:FMg

[Ano17r] Anonymous. Front matter. In Time of Nature and the Nature of Time:

Philosophical Perspectives of Time in Natural Sciences [BH17b], page 21. ISBN 3-319-53723-7 (hardcover), 3-319-53725-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC173.59.S65 T56 2017. URL

https://link.springer.com/book/
10.1007/978-3-319-53725-2.

Anonymous:2017:FMh

[Ano17s] Anonymous. Front matter. In Time of Nature and the Nature of Time: Philosophical Perspectives of Time in Natural Sciences [BH17b], page 87. ISBN 3-319-53723-7 (hardcover), 3-319-53725-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC173.59.S65 T56 2017. URL

https://link.springer.com/book/
10.1007/978-3-319-53725-2.

Anonymous:2017:FMi

[Ano17t] Anonymous. Front matter. In Time of Nature and the Nature of Time: Philosophical Perspectives of Time in Natural Sciences [BH17b], page 209. ISBN 3-319-53723-7 (hardcover), 3-319-53725-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC173.59.S65 T56 2017. URL

https://link.springer.com/book/
10.1007/978-3-319-53725-2.

Anonymous:2017:FMj

[Ano17u] Anonymous. Front matter. In Time of Nature and the Nature of Time: Philosophical Perspectives of Time in Natural Sciences [BH17b], page 273. ISBN 3-319-53723-7 (hardcover), 3-319-53725-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC173.59.S65 T56 2017. URL

https://link.springer.com/book/
10.1007/978-3-319-53725-2.

Anonymous:2017:FMk

[Ano17v] Anonymous. Front matter. In Mathematics as a Tool: Tracing New Roles of Mathematics in the Sciences [LC17b], pages i–x. ISBN 3-319-54468-3 (hardcover), 3-319-54469-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN GE45.M37; GE45.M38; QA8.4. URL

https://link.springer.com/book/
10.1007/978-3-319-54469-4.

Anonymous:2017:FMl

[Ano17w] Anonymous.  Front matter.  In

Mathematics as a Tool: Tracing New Roles of Mathematics in the Sciences [LC17b], page 21. ISBN 3-319-54468-3 (hardcover), 3-319-54469-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN GE45.M37; GE45.M38; QA8.4. URL

https://link.springer.com/book/
10.1007/978-3-319-54469-4.

Anonymous:2017:FMm

[Ano17x] Anonymous.  Front matter.  In

Mathematics as a Tool: Tracing New Roles of Mathematics in the Sciences [LC17b], page 91. ISBN 3-319-54468-3 (hardcover), 3-319-54469-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN GE45.M37; GE45.M38; QA8.4. URL

https://link.springer.com/book/
10.1007/978-3-319-54469-4.

Anonymous:2017:FMn

[Ano17y] Anonymous. Front matter. In Mathematics as a Tool: Tracing New Roles of Mathematics in the Sciences [LC17b], page 175. ISBN 3-319-54468-3 (hardcover), 3-319-54469-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN GE45.M37; GE45.M38; QA8.4. URL

                                              https://link.springer.com/book/
                                              10.1007/978-3-319-54469-4.

Anonymous:2018:BMc

[Ano18a] Anonymous. Back matter. In Feldhay et al. [FRSV18a], pages 325–327. ISBN 3-319-90343-8 (hardcover), 3-319-90345-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic).

                                              LCCN QA802. URL https://link.
                                              springer.com/book/10.1007/978-3-
                                              319-90345-3.

Anonymous:2018:BMa

[Ano18b] Anonymous. Back matter. In Mitcham et al. [MLNZ18], pages 317–333. ISBN 3-319-62448-2, 3-319-62450-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN TA15 .P45 2018.

                                              URL https://link.springer.com/
                                              book/10.1007/978-3-319-62450-1.

Anonymous:2018:BMb

[Ano18c] Anonymous. Back matter. In What Does it Mean to be an Empiricist?: Empiricisms in Eighteenth Century Sciences [BR18b], pages 265–297. ISBN 3-319-69858-3 (hardcover), 3-319-69860-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN B816 .W43 2018. URL https://link.

                                              springer.com/book/10.1007/978-3-
                                              319-69860-1.

Anonymous:2018:FMl

[Ano18d] Anonymous. Front matter. In Feldhay et al. [FRSV18a], pages i–xvii. ISBN 3-319-90343-8 (hardcover), 3-319-90345-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QA802.

                                              URL https://link.springer.com/
                                              book/10.1007/978-3-319-90345-3.

Anonymous:2018:FMm

[Ano18e] Anonymous. Front matter. In Feldhay et al. [FRSV18a], page 1. ISBN 3-319-90343-8 (hardcover), 3-319-90345-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QA802.

URL https://link.springer.com/
book/10.1007/978-3-319-90345-3.

Anonymous:2018:FMn

[Ano18f] Anonymous. Front matter. In Feldhay et al. [FRSV18a], page 79. ISBN 3-319-90343-8 (hardcover), 3-319-90345-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QA802.

URL https://link.springer.com/
book/10.1007/978-3-319-90345-3.

Anonymous:2018:FMo

[Ano18g] Anonymous. Front matter. In Feldhay et al. [FRSV18a], page 179. ISBN 3-319-90343-8 (hardcover), 3-319-90345-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QA802.

URL https://link.springer.com/
book/10.1007/978-3-319-90345-3.

Anonymous:2018:FMa

[Ano18h] Anonymous. Front matter. In Mitcham et al. [MLNZ18], pages i–xl. ISBN 3-319-62448-2, 3-319-62450-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN TA15 .P45 2018.

URL https://link.springer.com/
book/10.1007/978-3-319-62450-1.

Anonymous:2018:FMb

[Ano18i] Anonymous. Front matter. In Mitcham et al. [MLNZ18], page 1. ISBN 3-319-62448-2, 3-319-62450-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN TA15 .P45 2018.

URL https://link.springer.com/
book/10.1007/978-3-319-62450-1.

Anonymous:2018:FMc

[Ano18j] Anonymous. Front matter. In Mitcham et al. [MLNZ18], page 83. ISBN 3-319-62448-2, 3-319-62450-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN TA15 .P45 2018.

URL https://link.springer.com/
book/10.1007/978-3-319-62450-1.

Anonymous:2018:FMd

[Ano18k] Anonymous. Front matter. In Mitcham et al. [MLNZ18], page 157. ISBN 3-319-62448-2, 3-319-62450-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN TA15 .P45 2018.

URL https://link.springer.com/
book/10.1007/978-3-319-62450-1.

Anonymous:2018:FMe

[Ano18l] Anonymous. Front matter. In Mitcham et al. [MLNZ18], page 215. ISBN 3-319-62448-2, 3-319-62450-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN TA15 .P45 2018.

URL https://link.springer.com/
book/10.1007/978-3-319-62450-1.

Anonymous:2018:FMf

[Ano18m] Anonymous. Front matter. In Mitcham et al. [MLNZ18], page 281. ISBN 3-319-62448-2, 3-319-62450-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN TA15 .P45 2018.

URL https://link.springer.com/
book/10.1007/978-3-319-62450-1.

Anonymous:2018:FMg

[Ano18n] Anonymous. Front matter. In What Does it Mean to be an Empiricist?: Empiricisms in Eighteenth Century Sciences [BR18b], pages i–ix. ISBN 3-319-69858-3 (hardcover), 3-319-69860-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN B816

.W43 2018.  URL https://link.
springer.com/book/10.1007/978-3-
319-69860-1.

Anonymous:2018:FMh

[Ano18o] Anonymous. Front matter. In What Does it Mean to be an Empiricist?: Empiricisms in Eighteenth Century Sciences [BR18b], page 13. ISBN 3-319-69858-3 (hardcover), 3-319-69860-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN B816

.W43 2018.  URL https://link.
springer.com/book/10.1007/978-3-
319-69860-1.

Anonymous:2018:FMi

[Ano18p] Anonymous. Front matter. In What Does it Mean to be an Empiricist?: Empiricisms in Eighteenth Century Sciences [BR18b], page 45. ISBN 3-319-69858-3 (hardcover), 3-319-69860-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN B816

.W43 2018.  URL https://link.
springer.com/book/10.1007/978-3-
319-69860-1.

Anonymous:2018:FMj

[Ano18q] Anonymous. Front matter. In What Does it Mean to be an Empiricist?: Empiricisms in Eighteenth Century Sciences [BR18b], page 121. ISBN 3-319-69858-3 (hardcover), 3-319-69860-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN B816

.W43 2018.  URL https://link.
springer.com/book/10.1007/978-3-
319-69860-1.

Anonymous:2018:FMk

[Ano18r] Anonymous. Front matter. In What Does it Mean to be an Empiricist?: Empiricisms in Eighteenth Century Sciences [BR18b], page 201. ISBN 3-319-69858-3 (hardcover), 3-319-69860-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN B816

.W43 2018.  URL https://link.

                                              springer.com/book/10.1007/978-3-
                                              319-69860-1.

Anonymous:2018:FMp

[Ano18s] Anonymous. Front matter. In Piazza and Pulcini [PP18], pages i–xii. ISBN 3-319-93341-8, 3-319-93342-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QA8.4. URL

                                              https://link.springer.com/book/
                                              10.1007/978-3-319-93342-9.

Anonymous:2018:FMq

[Ano18t] Anonymous. Front matter. In Piazza and Pulcini [PP18], page 1. ISBN 3-319-93341-8, 3-319-93342-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QA8.4. URL

                                              https://link.springer.com/book/
                                              10.1007/978-3-319-93342-9.

Anonymous:2018:FMr

[Ano18u] Anonymous. Front matter. In Piazza and Pulcini [PP18], page 107. ISBN 3-319-93341-8, 3-319-93342-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QA8.4. URL

                                              https://link.springer.com/book/
                                              10.1007/978-3-319-93342-9.

Angyan:1992:CFA

[ÁNS91] János G. Ángyán and Gábor Náray-Szabó. Chemical fragmentation approach to the quantum chemical description of extended systems. In The Invention of Physical Science: Intersections of Mathematics, Theology and Natural Philosophy Since the Seventeenth Century: Essays in Honor of Erwin N. Hiebert [NRS92], chapter 1, pages 1–49. ISBN 94-010-5097-X, 94-011-2488-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

                                              ???? URL https://link.springer.
                                              com/chapter/10.1007/978-3-642-
                                              58183-0_1.

Antony:1987:NES

[Ant87] Louise M. Antony. Naturalized epistemology and the study of language. In Shimony and Nails [SN87], pages 235–257. ISBN 94-010-8168-9, 94-009-3735-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-009-
3735-2_14.

Anton:1990:USI

[Ant90] John P. Anton. The unity of scientific inquiry and categorial theory in Aristotle. In Greek Studies in the Philosophy and History of Science [Nic90], chapter 2, pages 29–43. ISBN 94-010-7403-8, 94-009-2015-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-009-2015-6_2.

Antony:2000:NRT

[Ant00] Louise Antony. Naturalizing radical translation. In Knowledge, Language and Logic: Questions for Quine [OK00], chapter 11, pages 141–150. ISBN 1-4020-0253-X, 94-011-3933-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-011-3933-
5_11.

Antonelli:2016:SNH

[Ant16] G. Aldo Antonelli. Semantic nominalism: How I learned to stop worrying and love universals. In Objectivity, Realism, and Proof: FilMat Studies in the Philosophy of Mathematics [BS16], chapter 2, pages 13–32. ISBN 3-319-31642-7, 3-319-31644-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QA9 .O25 2016.

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-3-319-31644-

4_2. See discussion [MP16].

Amann:1997:WRN

[AP97] Anton Amann and Hans Primas. What is the referent of a nonpure quantum state? In Cohen et al. [CHS97b], pages 9–29. ISBN 90-481-4809-X, 94-017-2732-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC173.96-174.52. URL https://link.

                                              springer.com/chapter/10.1007/
                                              978-94-017-2732-7_2.

Agazzi:2000:I

[AP00a] Evandro Agazzi and Massimo Pauri. Introduction. In The Reality of the Unobservable: Observability, Unobservability and Their Impact on the Issue of Scientific Realism [AP00b], chapter 1, pages 1–29. ISBN 90-481-5458-8, 94-015-9391-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

                                              ???? URL https://link.springer.
                                              com/chapter/10.1007/978-94-015-
                                              9391-5_1.

Agazzi:2000:RUO

[AP00b] Evandro Agazzi and Massimo Pauri. The Reality of the Unobservable: Observability, Unobservability and Their Impact on the Issue of Scientific Realism, volume 215 of Boston Studies in the Philosophy of Science. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 2000. ISBN 90-481-5458-8, 94-015-9391-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). vii + 377 pp. LCCN ???? URL

                                              https://link.springer.com/book/
                                              10.1007/978-94-015-9391-5.

Apel:1999:WHP

[Ape99]  Karl-Otto Apel.  Wissenschafts-

geschichte als hermeneutisches problem. In Fehér et al. [FKR99], pages 101–115. ISBN 90-481-5257-7, 94-015-9293-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-015-9293-
2_9.

Apostel:1982:SRO

[Apo82] Leo Apostel. Some remarks on ontology.

In Scientific Philosophy Today: Essays in Honor of Mario Bunge [AC82], chapter 1, pages 1–44. ISBN 90-277-1262-X, 90-277-1263-8, 94-009-8462-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 67.

URL https://link.springer.com/
chapter/10.1007/978-94-009-8462-
2_1.

Aptheker:1974:FH

[Apt74] Herbert Aptheker. Falsification in history. In For Dirk Struik: Scientific, Historical and Political Essays in Honor of Dirk J. Struik [CSW74], chapter 38, pages 493–499. ISBN 90-277-0379-5, 94-010-2115-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-010-2115-9_38.

Arandjelovic:1979:HSP

[Ara79] Jovan Arandjelović. Historical science and the philosophy of history. In Praxis: Yugoslav Essays in the Philosophy and Methodology of the Social Sciences [MP79], chapter 5, pages 89–97. ISBN 90-277-0968-8, 94-009-9355-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-009-9355-6_5.

Arabatzis:1994:RVS

[Ara94] Theodore Arabatzis. Rational versus sociological reductionism: Imre Lakatos and the Edinburgh School. In Trends in the Historiography of Science [GCN94], chapter 14, pages 177–192. ISBN 90-481-4264-4, 94-017-3596-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-017-3596-4_14.

Aravind:1997:BEG

[Ara97] P. K. Aravind. Borromean entanglement of the GHZ state. In Cohen et al. [CHS97b], pages 53–59. ISBN 90-481-4809-X, 94-017-2732-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC173.96-174.52.

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-017-2732-
                                              7_4.

Arabatzis:2012:HEE

[Ara12] Theodore Arabatzis. Hidden entities and experimental practice: Renewing the dialogue between history and philosophy of science. In Mauskopf and Schmaltz [MS12b], chapter 9, pages 125–139. ISBN 94-007-1744-X, 94-007-1745-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175 .I5134 2011. URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-007-1745-9_9.

Arapostathis:2015:WBE

[Ara15]  Stathis Arapostathis.  Wireless at

the bar: Experts, circuits and Marconi’s inventions in patent disputes in early Twentieth-Century Britain. In Arabatzis et al. [ARS15], chapter 23, pages 343–356. ISBN 3-319-14552-5 (paperback), 3-319-14553-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q125 .R45 2015.

URL https://link.springer.com/
chapter/10.1007/978-3-319-14553-
2_23.

Arboleda:1992:SNN

[Arb92] Luis Carlos Arboleda. Science and nationalism in New Granada on the eve of the revolution of independence. In Science and Empires: Historical Studies about Scientific Development and European Expansion [PJM92], chapter 26, pages 247–258. ISBN 0-7923-1518-9 (hardcover), 94-010-5145-3, 94-011-2594-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 136.

URL https://link.springer.com/
chapter/10.1007/978-94-011-2594-
9_26.

Arens:1989:ASC

[Are89a] Katherine Arens. Afterword: Some consequences of conceptual psychology. In Structures of Knowing: Psychologies of the Nineteenth Century [Are89j], chapter 9, pages 333–341. ISBN 94-010-7692-8, 94-009-2641-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-009-2641-7_9.

Arens:1989:CRH

[Are89b] Katherine Arens. The case for a reorientation in the history of psychology. In Structures of Knowing: Psychologies of the Nineteenth Century [Are89j], chapter 1, pages 3–33. ISBN 94-010-7692-8, 94-009-2641-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-009-2641-7_1.

Arens:1989:CPD

[Are89c] Katherine Arens. Counterproposition: Psychology as discourse. In Structures of Knowing: Psychologies of the Nineteenth Century [Are89j], chapter 2, pages 34–55. ISBN 94-010-7692-8, 94-009-2641-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

                                              ???? URL https://link.springer.
                                              com/chapter/10.1007/978-94-009-
                                              2641-7_2.

Arens:1989:DDP

[Are89d] Katherine Arens. Dilthey and descriptive psychology. In Structures of Knowing: Psychologies of the Nineteenth Century [Are89j], chapter 5, pages 153–171. ISBN 94-010-7692-8, 94-009-2641-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic).

                                              LCCN ????  URL https://link.
                                              springer.com/chapter/10.1007/
                                              978-94-009-2641-7_5.

Arens:1989:ECP

[Are89e] Katherine Arens. Empiricism and conceptual psychology: Psychophysics and philology. In Structures of Knowing: Psychologies of the Nineteenth Century [Are89j], chapter 4, pages 105–149. ISBN 94-010-7692-8, 94-009-2641-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-009-2641-
                                              7_4.

Arens:1989:FPP

[Are89f] Katherine Arens. Freud: The psychology of psychoanalysis. In Structures of Knowing: Psychologies of the Nineteenth Century [Are89j], chapter 8, pages 263–331. ISBN 94-010-7692-8, 94-009-2641-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic).

LCCN ????  URL https://link.
springer.com/chapter/10.1007/
978-94-009-2641-7_8.

Arens:1989:KHI

[Are89g] Katherine Arens. Kant and Herbart:

The initiation of conceptual psychology. In Structures of Knowing: Psychologies of the Nineteenth Century [Are89j], chapter 3, pages 59–104. ISBN 94-010-7692-8, 94-009-2641-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-009-2641-7_3.

Arens:1989:MPI

[Are89h] Katherine Arens. Mach’s psychology of investigation and the limits of science. In Structures of Knowing: Psychologies of the Nineteenth Century [Are89j], chapter 7, pages 216–262. ISBN 94-010-7692-8, 94-009-2641-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-009-2641-7_7.

Arens:1989:PCP

[Are89i] Katherine Arens. Phenomenology and conceptual psychology. In Structures of Knowing: Psychologies of the Nineteenth Century [Are89j], chapter 6, pages 172–215. ISBN 94-010-7692-8, 94-009-2641-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic).

LCCN ????  URL https://link.
springer.com/chapter/10.1007/
978-94-009-2641-7_6.

Arens:1989:SKP

[Are89j] Katherine Arens. Structures of Knowing: Psychologies of the Nineteenth Century, volume 113 of Boston Studies in the Philosophy of Science. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 1989. ISBN 94-010-7692-8, 94-009-2641-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). 444 (est.) pp. LCCN ???? URL

https://link.springer.com/book/
10.1007/978-94-009-2641-7.

Allamel-Raffin:2012:SIR

[ARG12] Catherine Allamel-Raffin and Jean-Luc Gangloff. Scientific images and robustness. In Characterizing the Robustness of Science: After the Practice Turn in Philosophy of Science [STNW12], chapter 7, pages 169–188. ISBN 94-007-2758-5, 94-007-2759-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175 .C4595 2012.

URL https://link.springer.com/
chapter/10.1007/978-94-007-2759-
5_7.

Aristotle:1976:TMC

[Ari76] Aristotle. On time, motion and change. In The Concepts of Space and Time: Their Structure and Their Development [Čap76], chapter 26, pages 147–157. ISBN 90-277-0375-2, 94-010-1727-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN BD632.

URL https://link.springer.com/
chapter/10.1007/978-94-010-1727-
5_26.

Armstrong:1976:IP

[Arm76] D. M. Armstrong. Immediate perception. In Essays in Memory of Imre Lakatos [CFW76], chapter 3, pages 23– 35. ISBN 90-277-0655-7, 94-010-1451-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/

chapter/10.1007/978-94-010-1451-
9_3.

Armstrong:1988:CNB

[Arm88] David M. Armstrong. Can a naturalist believe in universals? In Science in Reflection: The Israel Colloquium: Studies in History, Philosophy, and Sociology of Science Volume 3 [UM88], chapter 9, pages 103–115. ISBN 90-277-2713-9, 94-009-2957-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-009-2957-9_9.

Armstrong:1995:WMI

[Arm95] D. M. Armstrong. What makes induction rational? In Misiek [Mis95c], pages 45–54. ISBN 94-010-4206-3, 94-011-0461-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-011-
0461-6_3.

Armstrong:2018:BOD

[Arm18] Rita Armstrong. Between optimism and despair: Engineering, anthropology, and development in the Twenty-First Century. In Mitcham et al. [MLNZ18], chapter 18, pages 217–228. ISBN 3-319-62448-2, 3-319-62450-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN TA15 .P45 2018.

URL https://link.springer.com/
chapter/10.1007/978-3-319-62450-
1_18.

Arabatzis:2015:RHS

[ARS15] Theodore Arabatzis, Jürgen Renn, and Anna Simoẽs, editors. Relocating the History of Science: Essays in Honor of Kostas Gavroglu, volume 312 of Boston Studies in the Philosophy and History of Science. Springer International Publishing AG, Cham, Switzerland, 2015. ISBN 3-319-14552-5 (paperback), 3-319-14553-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). vii + 383 pp. LCCN Q125

.R45 2015. URL http://www.loc.
gov/catdir/enhancements/fy1516/
2015936523-d.html;   http://www.
loc.gov/catdir/enhancements/
fy1516/2015936523-t.html;
https://link.springer.com/book/
10.1007/978-3-319-14553-2.

Arthur:1989:RCR

[Art89] Richard T. W. Arthur. Russell’s conundrum: On the relation of Leibniz’s monads to the continuum. In An Intimate Relation: Studies in the History and Philosophy of Science Presented to Robert E. Butts on his 60th Birthday [BM89], chapter 9, pages 171–201. ISBN 94-010-7546-8, 94-009-2327-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-009-2327-
0_9.

Arthur:2015:IFE

[Art15] Wallace Arthur. Internal factors in evolution: The morphogenetic tree, developmental bias, and some thoughts on the conceptual structure of Evodevo. In Conceptual Change in Biology: Scientific and Philosophical Perspectives on Evolution and Development [Lov15b], chapter 16, pages 343–363. ISBN 94-017-9411-1, 94-017-9412-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QH331 C66 2015; QH491; QH492 .C66 2014.

URL https://link.springer.com/
chapter/10.1007/978-94-017-9412-
1_16.

Ashrafi:2007:BHa

[Ash07a] Babak Ashrafi. Big history? In Positioning the history of science [RG07], chapter 1, pages 1–5. ISBN 1-4020-5419-X (hardcover), 1-4020-5420-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q125 .P77 2007. URL https://link.

springer.com/chapter/10.1007/1-
4020-5420-3_1.

Ashrafi:2007:BHb

[Ash07b] Babak Ashrafi. Big history? In Positioning the history of science [RG07], chapter 2, pages 7–11. ISBN 1-4020-5419-X (hardcover), 1-4020-5420-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q125 .P77 2007. URL https://link.

springer.com/chapter/10.1007/1-
4020-5420-3_2.

Aslaksen:2018:EAP

[Asl18] Erik W. Aslaksen. An engineer’s approach to the philosophy of engineering. In Mitcham et al. [MLNZ18], chapter 8, pages 85–93. ISBN 3-319-62448-2, 3-319-62450-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN TA15 .P45 2018. URL https:

//link.springer.com/chapter/10.
1007/978-3-319-62450-1_8.

Assis:2011:NIP

[Ass11] A. K. T. Assis. Newton and inverse problems. In Brazilian Studies in Philosophy and History of Science: An account of recent works [KV11], chapter 3, pages 71–76. ISBN 90-481-9421-0 (hardcover), 90-481-9422-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175.3 .B73 2011.

URL https://link.springer.com/
chapter/10.1007/978-90-481-9422-
3_3.

Arndt:2015:CLI

[AT15] Michael Arndt and Laura Tesconi. A cut-like inference in a framework of explicit composition for various calculi of natural deduction. In From Logic to Practice: Italian Studies in the Philosophy of Mathematics [LPV15], chapter 9, pages 163–183. ISBN 3-319-10433-0, 3-319-10434-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QA8.4 .F76 2015; Q174 .B67 v.308. URL https://link.

                                              springer.com/chapter/10.1007/
                                              978-3-319-10434-8_9.

Atkinson:2003:ETE

[Atk03] David Atkinson. Experiments and thought experiments in natural science. In Descartes’s Mathematical Thought [Sas03d], chapter 14, pages 209–225. ISBN 90-481-6487-7, 94-017-1225-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

                                              URL https://link.springer.com/
                                              chapter/10.1007/0-306-48123-5_14.

Atlan:1986:MVB

[Atl86] Henri Atlan. Molecular versus biological evolution and programming. In Ullmann-Margalit [UM86a], chapter 13, pages 137–145. ISBN 90-277-2159-9 (paperback), 90-277-2158-0 (hardcover), 94-009-5496-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 94 Q175.

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-009-5496-
                                              0_13.

Attfield:2013:RID

[Att13] Robin Attfield. Reconciling individualist and deeper environmentalist theories? An exploration. In Bergandi [Ber13c], chapter 9, pages 127–139. ISBN 94-007-5066-8, 94-007-5067-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN GE42 .S77 2013; Q175 .B73 v. 296. URL https:

//link.springer.com/chapter/10.
1007/978-94-007-5067-8_9.

Aubin:2017:ESF

[Aub17] David Aubin. On the epistemic and social foundations of mathematics as tool and instrument in observatories, 1793–1846. In Mathematics as a Tool: Tracing New Roles of Mathematics in the Sciences [LC17b], chapter 10, pages 177–196. ISBN 3-319-54468-3 (hardcover), 3-319-54469-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN GE45.M37; GE45.M38; QA8.4. URL https:/

/link.springer.com/chapter/10.
1007/978-3-319-54469-4_10.

Augustynek:1967:TSP

[Aug67] Zdzisław Augustynek. Three studies in the philosophy of space and time. In Cohen and Wartofsky [CW67], pages 447–466. ISBN 90-277-0013-3, 94-010-3508-3. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175

.B731 v.3.  URL https://link.
springer.com/chapter/10.1007/
978-94-010-3508-8_27.

Augustine:1976:VT

[Aug76] St. Augustine. Views on time. In The Concepts of Space and Time: Their Structure and Their Development [Čap76], chapter 30, pages 179–183. ISBN 90-277-0375-2, 94-010-1727-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN BD632.

URL https://link.springer.com/
chapter/10.1007/978-94-010-1727-
5_30.

Augustynek:1982:TS

[Aug82] Zdzisław Augustynek. Time separation. In Polish Essays in the Philosophy of the Natural Sciences [Kra82c], chapter 19, pages 215–222. ISBN 94-009-7705-0 (e-book), 90-277-1287-5. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-009-7705-1_19.

Aune:1971:FCN

[Aun71] Bruce Aune. Fisk on capacities and natures. In Buck and Cohen [BC71], pages 83–87. ISBN 90-277-0309-4, 94-010-3142-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

                                              ???? URL https://link.springer.
                                              com/chapter/10.1007/978-94-010-
                                              3142-4_6.

Aharonov:1996:APM

[AV96] Yakir Aharonov and Lev Vaidman. About position measurements which do not show the Bohmian particle position. In Bohmian Mechanics and Quantum Theory: an Appraisal [CFG96], chapter 10, pages 141–154. ISBN 90-481-4698-4, 94-015-8715-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC173.96-174.52.

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-015-8715-
                                              0_10.

Aharonov:1997:PMT

[AV97] Yakir Aharonov and Lev Vaidman. Protective measurements of two-state vectors. In Cohen et al. [CHS97b], pages 1–8. ISBN 90-481-4809-X, 94-017-2732-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC173.96-174.52. URL https://link.

                                              springer.com/chapter/10.1007/
                                              978-94-017-2732-7_1.

Amoretti:2016:MPR

[AV16a] Maria Cristina Amoretti and Nicla Vassallo. Meta-Philosophical Reflection on Feminist Philosophies of Science, volume 317 of Boston Studies in the Philosophy and History of Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2016. ISBN 3-319-26346-3 (hardcover), 3-319-26348-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). vi + 222 pp. LCCN HQ1154 .M48 2016; HQ1190 M48 2016. URL

https://link.springer.com/book/
10.1007/978-3-319-26348-9.

Amoretti:2016:SKT

[AV16b] Maria Cristina Amoretti and Nicla Vassallo. Some key topics in feminist philosophies of science: An introduction. In Meta-Philosophical Reflection on Feminist Philosophies of Science [AV16a], chapter 1, pages 1–13. ISBN 3-319-26346-3 (hardcover), 3-319-26348-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN HQ1154 .M48 2016; HQ1190 M48 2016.

URL https://link.springer.com/
chapter/10.1007/978-3-319-26348-
9_1.

Avgelis:1989:LES

[Avg89] Nikolaos Avgelis. Lakatos on the evaluation of scientific theories. In Imre Lakatos and Theories of Scientific Change [GGN89], chapter 13, pages 157–167. ISBN 94-010-7860-2, 94-009-3025-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-009-
3025-4_13.

Avgelis:1990:SEC

[Avg90] Nikolaos Avgelis. Schlick’s epistemology and its contribution to modern empiricism. In Greek Studies in the Philosophy and History of Science [Nic90], chapter 21, pages 335–351. ISBN 94-010-7403-8, 94-009-2015-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-009-2015-
6_21.

Ayala:1976:BAS

[Aya76] Francisco J. Ayala. Biology as an autonomous science. In Topics in the Philosophy of Biology [GM76], chapter 14, pages 312–329. ISBN 90-277-0596-8, 94-010-1829-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-010-1829-6_14.

Ayala:1996:NSP

[Aya96] Francisco J. Ayala. The nature of science and the problem of demarcation. In Munévar [Mun96], chapter 6, pages 123–142. ISBN 94-010-6622-1, 94-009-0305-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-009-
0305-0_6.

Aydede:2005:CFS

[Ayd05] Murat Aydede. Computation and functionalism: Syntactic theory of mind revisited. In Turkish Studies in the History and Philosophy of Science [IG05c], chapter 13, pages 177–204. ISBN 1-4020-3332-X (hardcover), 1-4020-3333-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 244.

URL https://link.springer.com/
chapter/10.1007/1-4020-3333-8_13.

Ayer:1976:IR

[Aye76] Alfred J. Ayer. Identity and reference.

In Language in Focus: Foundations, Methods and Systems: Essays in Memory of Yehoshua Bar-Hillel [Kas76b], chapter 1, pages 3–24. ISBN 90-277-0644-1, 90-277-0645-X, 94-010-1876-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN P101-120.

URL https://link.springer.com/
chapter/10.1007/978-94-010-1876-
0_1.

Amrine:1987:PGS

[AZ87] Frederick Amrine and Francis J.

Zucker. Postscript Goethe’s science: An alternative to modern science or within it — or no alternative at all? In Amrine et al. [AZW87], chapter 19, pages 373–388. ISBN 90-277-2265-X, 94-009-3761-X. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 97. US$79.00. URL https:

//link.springer.com/chapter/10.
1007/978-94-009-3761-1_19.

Amrine:1987:GSR

[AZW87] Frederick Amrine, Francis J. Zucker, and Harvey Wheeler, editors. Goethe and the Sciences: a Reappraisal, volume 97 of Boston Studies in the Philosophy of Science. D. Reidel, Dordrecht, The Netherlands; Boston, MA, USA; Lancaster, UK; Tokyo, Japan, 1987. ISBN 90-277-2265-X, 94-009-3761-X. ISSN 0068-0346 (print), 2214-7942 (electronic). xvii + 442 pp. LCCN Q174 .B67 vol. 97. US$79.00.

URL https://link.springer.com/
book/10.1007/978-94-009-3761-1.

Bar-Am:2009:PVS

[BA09] Nimrod Bar-Am. Proof versus sound inference. In Rethinking Popper [PC09], chapter 5, pages 63–70. ISBN 1-4020-9337-3, 1-4020-9338-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN B1649.P64 R48 2009; Q174 .B67 2009. URL https:

//link.springer.com/chapter/10.
1007/978-1-4020-9338-8_5.

Baatz:1992:EMS

[Baa92] Ursula Baatz. Ernst Mach — the scientist as a Buddhist? In Ernst Mach — A Deeper Look: Documents and New Perspectives [Bla92a], chapter 9, pages 183–199. ISBN 94-010-5230-1, 94-011-2771-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN C5.53; QC6.9. URL https:

//link.springer.com/chapter/10.
1007/978-94-011-2771-4_9.

Babich:1997:APN

[Bab97] Babette E. Babich. Against postmodernism and the “new” philosophy of science: Nietzsche’s image of science in the light of art. In Ginev and Cohen [GC97], chapter 3, pages 27–45. ISBN 0-7923-4444-8 (hardcover), 94-011-5788-X. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 192. URL https://link.

springer.com/chapter/10.1007/
978-94-011-5788-9_3.

Babich:1999:NEP

[Bab99a] Babette E. Babich. Nietzsche, Epistemology, and Philosophy of Science: Nietzsche and the Sciences II, volume 204 of Boston Studies in the Philosophy of Science. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 1999. ISBN 90-481-5234-8, 94-017-2428-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). xvii + 382 pp. LCCN ????

URL https://link.springer.com/
book/10.1007/978-94-017-2428-9.

Babich:1999:NTK

[Bab99b] Babette E. Babich. Nietzsche, Theories of Knowledge, and Critical Theory: Nietzsche and the Sciences I, volume 203 of Boston Studies in the Philosophy of Science. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 1999. ISBN 90-481-5233-X, 94-017-2430-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). xx + 348 pp. LCCN ???? URL

https://link.springer.com/book/
10.1007/978-94-017-2430-2.

Babich:1999:NCT

[Bab99c] Babette E. Babich. Nietzsche’s critical theory: The culture of science as art. In Nietzsche, Theories of Knowledge, and Critical Theory: Nietzsche and the Sciences I [Bab99b], chapter 1, pages 1–26. ISBN 90-481-5233-X, 94-017-2430-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-017-
2430-2_1.

Babich:1999:TAL

[Bab99d] Babette E. Babich. Truth, art, and life:

Nietzsche, epistemology, philosophy of science. In Nietzsche, Epistemology, and Philosophy of Science: Nietzsche and the Sciences II [Bab99a], chapter 1, pages 1–24. ISBN 90-481-5234-8, 94-017-2428-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-017-
2428-9_1.

Babich:2002:FI

[Bab02a] Babette E. Babich. The fortunes of incommensurability. In Hermeneutic Philosophy of Science, Van Gogh’s Eyes, and God: Essays in Honor of Patrick A. Heelan, S.J. [Bab02c], chapter 1, pages 1–18. ISBN 90-481-5926-1, 94-017-1767-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 225. URL https:

//link.springer.com/chapter/10.
1007/978-94-017-1767-0_1.

Babich:2002:HTA

[Bab02b] Babette E. Babich. Heidegger’s truth of art and the question of aesthetics. In Hermeneutic Philosophy of Science, Van Gogh’s Eyes, and God: Essays in Honor of Patrick A. Heelan, S.J. [Bab02c], chapter 22, pages 265–278. ISBN 90-481-5926-1, 94-017-1767-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 225. URL https://link.

springer.com/chapter/10.1007/
978-94-017-1767-0_22.

Babich:2002:HPS

[Bab02c] Babette E. Babich. Hermeneutic Philosophy of Science, Van Gogh’s Eyes, and God: Essays in Honor of Patrick A. Heelan, S.J., volume 225 of Boston Studies in the Philosophy and History of Science. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 2002. ISBN 90-481-5926-1, 94-017-1767-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). xvii + 484 pp. LCCN Q174 .B67 vol. 225. URL

https://link.springer.com/book/
10.1007/978-94-017-1767-0.

Babich:2002:SHH

[Bab02d] Babette E. Babich. Sokal’s hermeneutic hoax: Physics and the new inquisition. In Hermeneutic Philosophy of Science, Van Gogh’s Eyes, and God: Essays in Honor of Patrick A. Heelan, S.J. [Bab02c], chapter 6, pages 67–78. ISBN 90-481-5926-1, 94-017-1767-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 225.

URL https://link.springer.com/
chapter/10.1007/978-94-017-1767-
0_6.

Bach:1990:III

[Bac90] Alexander Bach. Indistinguishability, interchangeability, and indeterminism. In Statistics in Science: The Foundations of Statistical Methods in Biology, Physics and Economics [CC90], chapter 9, pages 149–166. ISBN 94-010-6765-1, 94-009-0619-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QA276-280.

URL https://link.springer.com/
chapter/10.1007/978-94-009-0619-
8_9.

Bacso:1999:WT

[Bac99] Béla Bacsó. The will to truth. In Nietzsche, Epistemology, and Philosophy of Science: Nietzsche and the Sciences II [Bab99a], chapter 13, pages 161–167. ISBN 90-481-5234-8, 94-017-2428-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-017-2428-
9_13.

Baracca:2014:CSD

[BAC14] Angelo Baracca, Víctor Luis Fajer Avila, and Carlos Rodríguez Castellanos. A comprehensive study of the development of physics in Cuba from 1959. In The History of Physics in Cuba [BRW14a], chapter 6, pages 115–234. ISBN 94-017-8040-4, 94-017-8041-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC5.53; QC6.9; QC9 C9H57 2014; Q124.6-127.2. URL https://link.

                                              springer.com/chapter/10.1007/
                                              978-94-017-8041-4_6.

Badran:1992:WA

[Bad92] Adnan Badran. Welcome address. In Science and Empires: Historical Studies about Scientific Development and European Expansion [PJM92], chapter 2, pages 13–15. ISBN 0-7923-1518-9 (hardcover), 94-010-5145-3, 94-011-2594-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 136. URL https:/

                                              /link.springer.com/chapter/10.
                                              1007/978-94-011-2594-9_2.

Baetu:2015:WME

[Băe15] Tudor M. Băetu. When is a mechanistic explanation satisfactory? Reductionism and antireductionism in the context of mechanistic explanations. In Romanian Studies in Philosophy of Science [PST15], chapter 16, pages 255–268. ISBN 3-319-16654-9 (hardcover), 3-319-16655-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

                                              ???? URL https://link.springer.
                                              com/chapter/10.1007/978-3-319-
                                              16655-1_16.

Bagce:2005:SHS

[Bağ05] S. Bağçe. A study on the heuristic of Saccheri’s Euclides. In Turkish Studies in the History and Philosophy of Science [IG05c], chapter 10, pages 137–150. ISBN 1-4020-3332-X (hardcover), 1-4020-3333-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 244.

                                              URL https://link.springer.com/
                                              chapter/10.1007/1-4020-3333-8_10.

Bar-Am:2017:EOE

[BAG17] Nimrod Bar-Am and Stefano Gattei. Encouraging Openness: Essays for Joseph Agassi on the Occasion of His 90th Birthday, volume 325 of Boston Studies in the Philosophy and History of Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2017. ISBN 3-319-57668-2 (hardcover), 3-319-57669-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). xiv + 584 pp. LCCN B833.E53 2017. URL

https://link.springer.com/book/
10.1007/978-3-319-57669-5.

Bahm:1996:TR

[Bah96] Archie J. Bahm. Tentative realism. In Cohen et al. [CHR96], pages 97–101. ISBN 90-481-4493-0, 94-015-8638-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-015-8638-
2_7.

Bailey:1976:CIS

[Bai76a] C. Bailey. The continuity and infinity of space according to Epicurus and Lucretius. In The Concepts of Space and Time: Their Structure and Their Development [Čap76], chapter 6, pages 33–37. ISBN 90-277-0375-2, 94-010-1727-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic).

LCCN BD632. URL https://link.
springer.com/chapter/10.1007/
978-94-010-1727-5_6.

Bailey:1976:MVA

[Bai76b] C. Bailey. Matter and the void according to Leucippus. In The Concepts of Space and Time: Their Structure and Their Development [Čap76], chapter 2, pages 17–19. ISBN 90-277-0375-2, 94-010-1727-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic).

LCCN BD632. URL https://link.

                                              springer.com/chapter/10.1007/
                                              978-94-010-1727-5_2.

Bailey:1976:RTT

[Bai76c] C. Bailey. The relational theory of time in Ancient Antonism. In The Concepts of Space and Time: Their Structure and Their Development [Čap76], chapter 25, pages 143–145. ISBN 90-277-0375-2, 94-010-1727-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN BD632.

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-010-1727-
                                              5_25.

Bailey:2017:DDL

[Bai17] Joan Bailey. Dungeons and dragons: A lesson in etiquette from Professor Agassi. In Encouraging Openness: Essays for Joseph Agassi on the Occasion of His 90th Birthday [BAG17], chapter 41, pages 543–545. ISBN 3-319-57668-2 (hardcover), 3-319-57669-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN B833.E53

                                              2017. URL https://link.springer.
                                              com/chapter/10.1007/978-3-319-
                                              57669-5_41.

Bir:2011:CFM

[BAK11] Atilla Bir, Şinasi Acar, and Mustafa Kaçar. The clockmaker family Meyer and their watch keeping the alla turca time. In Science between Europe and Asia: Historical Studies on the Transmission, Adoption and Adaptation of Knowledge [GR11b], chapter 8, pages 125–136. ISBN 90-481-9968-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q125 .S354 2010; Q175.52.A78 S35 2010. URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-90-481-9968-6_8.

Baltas:1989:LAJ

[Bal89a] Aristides Baltas. Louis Althusser and Joseph D. Sneed: A strange encounter in philosophy of science? In Imre Lakatos and Theories of Scientific Change [GGN89], chapter 20, pages 269–286. ISBN 94-010-7860-2, 94-009-3025-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-009-
3025-4_20.

Balzer:1989:I

[Bal89b] W. Balzer. On incommensurability. In Imre Lakatos and Theories of Scientific Change [GGN89], chapter 21, pages 287–304. ISBN 94-010-7860-2, 94-009-3025-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-009-
3025-4_21.

Baltas:1990:OAM

[Bal90] Aristides Baltas. Once again on the meaning of physical concepts. In Greek Studies in the Philosophy and History of Science [Nic90], chapter 19, pages 293–313. ISBN 94-010-7403-8, 94-009-2015-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-009-
2015-6_19.

Baltas:1994:HEE

[Bal94] Aristides Baltas. On the harmful effects of excessive anti-Whiggism. In Trends in the Historiography of Science [GCN94], chapter 7, pages 107–119. ISBN 90-481-4264-4, 94-017-3596-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-017-3596-
                                              4_7.

Balestra:2002:BSR

[Bal02] Dominic Balestra. In-between science and religion. In Hermeneutic Philosophy of Science, Van Gogh’s Eyes, and God: Essays in Honor of Patrick A. Heelan, S.J. [Bab02c], chapter 32, pages 377–384. ISBN 90-481-5926-1, 94-017-1767-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 225. URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-017-1767-0_32.

Baltas:2003:FCO

[Bal03] Aristides Baltas. On French concepts and objects: Comments on Ursula Klein. In Descartes’s Mathematical Thought [Sas03d], chapter 11, pages 187–197. ISBN 90-481-6487-7, 94-017-1225-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

                                              ???? URL https://link.springer.
                                              com/chapter/10.1007/0-306-48123-

5_11. See [Kle03].

Baltas:2008:NPC

[Bal08]  Aristides Baltas.  Nonsense and

paradigm change. In Rethinking Scientific Change and Theory Comparison: Stabilities, Ruptures, Incommensurabilities? [SSHH08], chapter 4, pages 49–70. ISBN 1-4020-6274-5 (hardcover), 1-4020-6279-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175 .R397 2008. URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-1-4020-6279-7_4.  See

comment [Obe08].

Balibar:2014:SHQ

[Bal14] Sébastien Balibar. Superfluidity: How quantum mechanics became visible. In History of Artificial Cold, Scientific, Technological and Cultural Issues [Gav14b], chapter 6, pages 93–117. ISBN 94-007-7198-3 (hardcover), 94-007-7199-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC278 H57 2014; TP494 .H577 2014.

URL https://link.springer.com/
chapter/10.1007/978-94-007-7199-
4_6.

Balmer:2015:SDT

[Bal15] Brian Balmer. The social dimension of technology: The control of chemical and biological weapons. In New Perspectives on Technology, Values, and Ethics: Theoretical and Practical [Gon15a], chapter 9, pages 167–182. ISBN 3-319-21869-7 (hardcover), 3-319-21870-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-3-319-
21870-0_9.

Baldwin:2018:EPN

[Bal18] John Baldwin. The explanatory power of a new proof: Henkin’s completeness proof. In Piazza and Pulcini [PP18], pages 147–162. ISBN 3-319-93341-8, 3-319-93342-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic).

LCCN QA8.4. URL https://link.
springer.com/chapter/10.1007/
978-3-319-93342-9_9.

Bambach:1999:PKN

[Bam99] Charles Bambach. The politics of knowledge: Nietzsche within Heidegger’s history of truth. In Nietzsche, Theories of Knowledge, and Critical Theory: Nietzsche and the Sciences I [Bab99b], chapter 8, pages 103–117. ISBN 90-481-5233-X, 94-017-2430-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-017-2430-
                                              2_8.

Bantz:1980:SDE

[Ban80] David A. Bantz. The structure of discovery: Evolution of structural accounts of chemical bonding. In Scientific Discovery: Case Studies [Nic80e], chapter 16, pages 291–329. ISBN 90-277-1093-7, 94-009-9015-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 60.

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-009-9015-
                                              9_16.

Banton:1995:AFR

[Ban95] Michael Banton. All forms of racial discrimination. In Critical Rationalism, the Social Sciences and the Humanities: Essays for Joseph Agassi. Volume II [JL95b], chapter 6, pages 101–114. ISBN 0-7923-2961-9, 90-481-4430-2, 94-017-0441-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN B29 .C75 1995. URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-017-0441-0_6.

Bangu:2015:SPU

[Ban15] Sorin Bangu. Scientific progress, understanding and unification. In Romanian Studies in Philosophy of Science [PST15], chapter 15, pages 239–253. ISBN 3-319-16654-9 (hardcover), 3-319-16655-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

                                              ???? URL https://link.springer.
                                              com/chapter/10.1007/978-3-319-
                                              16655-1_15.

Banton:2017:CRS

[Ban17] Michael Banton. Critical rationalism in sociology. In Encouraging Openness:

Essays for Joseph Agassi on the Occasion of His 90th Birthday [BAG17], chapter 31, pages 389–398. ISBN 3-319-57668-2 (hardcover), 3-319-57669-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN B833.E53

2017. URL https://link.springer.
com/chapter/10.1007/978-3-319-
57669-5_31.

Bartley:1974:TLP

[Bar74] W. W. Bartley III. Theory of language and philosophy of science as instruments of educational reform: Wittgenstein and Popper as Austrian schoolteachers. In Methodological and Historical Essays in the Natural and Social Sciences [CW74b], chapter 16, pages 307–337. ISBN 90-277-0378-7, 94-010-2128-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 14. URL https:

//link.springer.com/chapter/10.
1007/978-94-010-2128-9_16.

Barrow:1976:AT

[Bar76a] I. Barrow. Absolute time. In The Concepts of Space and Time: Their Structure and Their Development [Čap76], chapter 35, pages 203–208. ISBN 90-277-0375-2, 94-010-1727-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN BD632.

URL https://link.springer.com/
chapter/10.1007/978-94-010-1727-
5_35.

Bartley:1976:IL

[Bar76b] W. W. Bartley III. On Imre Lakatos.

In Essays in Memory of Imre Lakatos [CFW76], chapter 4, pages 37–38. ISBN 90-277-0655-7, 94-010-1451-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-010-1451-9_4.

Bartsch:1976:RCS

[Bar76c] Renate Bartsch. The role of categorial syntax in grammatical theory. In Language in Focus: Foundations, Methods and Systems: Essays in Memory of Yehoshua Bar-Hillel [Kas76b], chapter 22, pages 503–539. ISBN 90-277-0644-1, 90-277-0645-X, 94-010-1876-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN P101-120. URL https://link.

                                              springer.com/chapter/10.1007/
                                              978-94-010-1876-0_22.

Barbo:1983:DHD

[Bar83] Francesca Rivetti Barbò. Dialogue: How do we know what others mean and why? In Language, Logic and Method [CW83b], chapter 18, pages 409–444. ISBN 94-009-7704-2, 94-009-7702-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-009-7702-

0_18. See comments [Bur83].

Bartley:1984:LSD

[Bar84] W. W. Bartley III. Logical strength and demarcation. In Rationality in Science and Politics [And84d], chapter 5, pages 69–93. ISBN 90-277-1953-5, 94-009-6254-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

                                              ???? URL https://link.springer.
                                              com/chapter/10.1007/978-94-009-
                                              6254-5_5.

Barnouw:1987:GHS

[Bar87]  Jeffrey Barnouw.  Goethe and

Helmholtz: Science and sensation. In Amrine et al. [AZW87], chapter 4, pages 45–82. ISBN 90-277-2265-X, 94-009-3761-X. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 97. US$79.00. URL https:

//link.springer.com/chapter/10.
1007/978-94-009-3761-1_4.

Bartley:1989:UKB

[Bar89]  W. W. Bartley III.  Unfathomed

knowledge in a bottle. In Freedom and Rationality: Essays in Honor of John Watkins From his Colleagues and Friends [DJ89], chapter 13, pages 207–223. ISBN 94-010-7571-9, 94-009-2380-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-009-
2380-5_13.

Barkan:1991:WNQ

[Bar91a] Diana Barkan. Walther Nernst and quantum theory. In Woodward and Cohen [WC91], pages 151–162. ISBN 94-010-5395-2, 94-011-3164-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-011-3164-3_14.

Barkan:1992:UPC

[Bar91b] Diana Kormos Barkan. A usable past:

Creating disciplinary space for physical chemistry. In The Invention of Physical Science: Intersections of Mathematics, Theology and Natural Philosophy Since the Seventeenth Century: Essays in Honor of Erwin N. Hiebert [NRS92], chapter 8, pages 175–202. ISBN 94-010-5097-X, 94-011-2488-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-011-2488-1_8.

Barasch:1992:WIC

[Bar92] Moshe Barasch. Works of the imagination: A comment. In The Scientific Enterprise: The Bar-Hillel Colloquium: Studies in History, Philosophy, and Sociology of Science, Volume 4 [UM92], chapter 6, pages 117–121. ISBN 94-010-5190-9, 94-011-2688-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .S354 1992.

URL https://link.springer.com/
chapter/10.1007/978-94-011-2688-
5_6.

Barsanti:1994:LBB

[Bar94] Giulio Barsanti. Lamarck and the birth of biology 1740–1810. In Poggi and Bossi [PB94], chapter 4, pages 47–74. ISBN 0-7923-2336-X, 94-017-2921-2. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 152.

URL https://link.springer.com/
chapter/10.1007/978-94-017-2921-
5_4.

Bartels:1997:DTE

[Bar97] Andreas Bartels. Do times exist? In Perspectives on Time [FSU97b], chapter 9, pages 203–213. ISBN 90-481-4774-3, 94-015-8875-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol.189.

URL https://link.springer.com/
chapter/10.1007/978-94-015-8875-
1_9.

Barton:1999:HEB

[Bar99] Justin Barton. How epistemology becomes what it is. In Nietzsche, Epistemology, and Philosophy of Science: Nietzsche and the Sciences II [Bab99a], chapter 11, pages 141–149. ISBN 90-481-5234-8, 94-017-2428-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-017-2428-9_11.

Barker:2001:ICC

[Bar01a]  Peter Barker.  Incommensurability

and conceptual change during the Copernican Revolution. In Incommensurability and Related Matters [HHS01], chapter 10, pages 241–273. ISBN 90-481-5709-9, 94-015-9680-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-015-9680-0_10.

Barrow:2001:NCA

[Bar01b] Mark V. Barrow, Jr. Naturalists as conservationists: American scientists, social responsibility, and political activism before the bomb. In Allen and MacLeod [AM01b], chapter 14, pages 217–233. ISBN 1-4020-0495-8 (hardcover), 94-017-2956-5. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 228; Q175.55.

URL https://link.springer.com/
chapter/10.1007/978-94-017-2956-
7_14.

Barden:2002:TPR

[Bar02a] Garrett Barden. Thinking the philosophy of religion. In Hermeneutic Philosophy of Science, Van Gogh’s Eyes, and God: Essays in Honor of Patrick A. Heelan, S.J. [Bab02c], chapter 33, pages 385–392. ISBN 90-481-5926-1, 94-017-1767-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 225.

URL https://link.springer.com/
chapter/10.1007/978-94-017-1767-
0_33.

Barrett:2002:PCA

[Bar02b] D. Cyril Barrett. Phenomenology and 20th century artistic revolutions. In Hermeneutic Philosophy of Science, Van Gogh’s Eyes, and God: Essays in Honor of Patrick A. Heelan, S.J. [Bab02c], chapter 23, pages 279–286. ISBN 90-481-5926-1, 94-017-1767-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 225. URL https://link.

springer.com/chapter/10.1007/
978-94-017-1767-0_23.

Barbour:2007:EMP

[Bar07] Julian B. Barbour. Einstein and Mach’s principle. In Renn et al. [RSS^{+}07a], pages 569–604. ISBN 1-4020-3999-9. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 no. 250 v. 3; QC173.6 .G469 2007. URL https:

//link.springer.com/chapter/10.
1007/978-1-4020-4000-9_32.

Barberousse:2008:OVO

[Bar08] Anouk Barberousse. From one version to the other: Intra-theoretical change. In Rethinking Scientific Change and Theory Comparison: Stabilities, Ruptures, Incommensurabilities? [SSHH08], chapter 6, pages 87–101. ISBN 1-4020-6274-5 (hardcover), 1-4020-6279-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175 .R397 2008. URL https://link.

springer.com/chapter/10.1007/
978-1-4020-6279-7_6.

Baruch:2009:REP

[Bar09]  Herzl Baruch.  Re-examination of

Popper’s portrayal of Socrates. In Rethinking Popper [PC09], chapter 23, pages 305–320. ISBN 1-4020-9337-3, 1-4020-9338-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN B1649.P64 R48 2009; Q174 .B67 2009.

URL https://link.springer.com/
chapter/10.1007/978-1-4020-9338-
8_23.

Barrett:2010:PIC

[Bar10] Martin Barrett. The possibility of infinitesimal chances. In The Place of Probability in Science: In Honor of Ellery Eells (1953–2006) [Fet10a], chapter 3, pages 65–79. ISBN 90-481-3614-8 (hardcover), 90-481-3615-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175 .P563

2010. URL https://link.springer.
com/chapter/10.1007/978-90-481-
3615-5_3.

Baracca:2014:CED

[Bar14a] Angelo Baracca. The Cuban “exception”: The development of an advanced scientific system in an underdeveloped country. In The History of Physics in Cuba [BRW14a], chapter 2, pages 9–50. ISBN 94-017-8040-4, 94-017-8041-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC5.53; QC6.9; QC9 C9H57 2014; Q124.6-127.2. URL https://link.

springer.com/chapter/10.1007/
978-94-017-8041-4_2.

Baracca:2014:RDP

[Bar14b] Angelo Baracca. The rise and development of physics in Cuba: An interview with Hugo Pérez Rojas in May 2009. In The History of Physics in Cuba [BRW14a], chapter 11, pages 279–284. ISBN 94-017-8040-4, 94-017-8041-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC5.53; QC6.9; QC9 C9H57 2014; Q124.6-127.2. URL https://link.

springer.com/chapter/10.1007/
978-94-017-8041-4_11.

Baracca:2015:CSM

[Bar15] Angelo Baracca. Can science make peace with the environment? Science, power, exploitation. In Arabatzis et al.

[ARS15], chapter 25, pages 367–383. ISBN 3-319-14552-5 (paperback), 3-319-14553-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q125 .R45 2015. URL https://link.

springer.com/chapter/10.1007/
978-3-319-14553-2_25.

Barton:2016:MCS

[Bar16] Neil Barton. Multiversism and concepts of set: How much relativism is acceptable? In Objectivity, Realism, and Proof: FilMat Studies in the Philosophy of Mathematics [BS16], chapter 11, pages 189–209. ISBN 3-319-31642-7, 3-319-31644-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QA9 .O25 2016. URL https:

//link.springer.com/chapter/10.
1007/978-3-319-31644-4_11.

Barberousse:2017:EBT

[Bar17] Anouk Barberousse. Empirical Bayes as a tool. In Mathematics as a Tool: Tracing New Roles of Mathematics in the Sciences [LC17b], chapter 9, pages 157– 173. ISBN 3-319-54468-3 (hardcover), 3-319-54469-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN GE45.M37; GE45.M38; QA8.4.

URL https://link.springer.com/
chapter/10.1007/978-3-319-54469-
4_9.

Basosi:2014:CSC

[Bas14] Duccio Basosi. Cuba: A short critical bibliographic guide. In The History of Physics in Cuba [BRW14a], chapter 3, pages 51–54. ISBN 94-017-8040-4, 94-017-8041-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC5.53; QC6.9; QC9 C9H57 2014; Q124.6-127.2. URL https://link.

springer.com/chapter/10.1007/
978-94-017-8041-4_3.

Baumol:1988:REI

[Bau88] William J. Baumol. On the reality of economic illusion. In Philosophy, History and Social Action: Essays in Honor of Lewis Feuer with an autobiographic essay by Lewis Feuer [HOO88c], chapter 3, pages 103–117. ISBN 94-010-7793-2, 94-009-2873-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-009-2873-
2_3.

Bauer:2017:EHV

[Bau17] Craig Bauer. The early history of voice encryption. In Floyd and Bokulich [FB17], chapter 7, pages 159–187. ISBN 3-319-53278-2, 3-319-53280-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QA29.T8 P45 2017.

URL https://link.springer.com/
chapter/10.1007/978-3-319-53280-
6_7.

Bergson:1971:IME

[BB71] Bergson and Boutroux. The indeterminacy of microphysical events: Bergson and Boutroux. In Bergson and Modern Physics: A Reinterpretation and Reevaluation [Čap71c], chapter 37, pages 284–291. ISBN 94-010-3098-7, 94-010-3096-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-010-
3096-0_37.

Beltrametti:2000:RBC

[BB00] Enrico Beltrametti and Slawomir Bugajski. On the relationships between classical and quantum mechanics. In The Reality of the Unobservable: Observability, Unobservability and Their Impact on the Issue of Scientific Realism [AP00b], chapter 27, pages 353–359. ISBN 90-481-5458-8, 94-015-9391-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-015-
9391-5_27.

Bokulich:2011:SS

[BB11] Alisa Bokulich and Peter Bokulich, editors. Scientific Structuralism, volume 281 of Boston Studies in the Philosophy of Science. Springer-Ver-lag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2011. ISBN 90-481-9596-9, 90-481-9597-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). xvii + 182 pp. LCCN Q175.32.S78 S35 2011. URL

https://link.springer.com/book/
10.1007/978-90-481-9597-8.

Barnes:2018:BAC

[BBD18] James L. Barnes, Susan K. Barnes, and Michael J. Dyrenfurth. A biomimetic approach to complex global problems. In Mitcham et al. [MLNZ18], chapter 9, pages 95–102. ISBN 3-319-62448-2, 3-319-62450-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN TA15 .P45 2018. URL https:

//link.springer.com/chapter/10.
1007/978-3-319-62450-1_9.

Becquerel:1991:PVL

[BBO91] Henri Becquerel, Jean Becquerel, and H. Kamerlingh Onnes. On phosphorescence at very low temperatures. In Gavroglu and Goudaroulis [GG91], chapter 13, pages 236–251. ISBN 94-010-7433-X (hardcover), 94-009-2079-2. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 124.

URL https://link.springer.com/
chapter/10.1007/978-94-009-2079-
8_13.

Brockwell:1995:MRM

[BBT95] John Brockwell, Peter J. Bottomley, and Janice E. Thies. Manipulation of Rhizobia microflora for improving legume productivity and soil fertility: A critical assessment. In No Truth Except in the Details: Essays in Honor of Martin J. Klein [KS95], chapter 7, pages 143–180. ISBN 94-010-4097-4, 94-011-0217-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-011-
0055-7_7.

Buck:1971:PMR

[BC71] Roger C. Buck and Robert S. Cohen, editors. PSA 1970: In Memory of Rudolf Carnap. Proceedings of the 1970 Biennial Meeting Philosophy of Science Association, volume 8 of Boston Studies in the Philosophy of Science. D. Reidel, Dordrecht, The Netherlands; Boston, MA, USA; Lancaster, UK; Tokyo, Japan, 1971. ISBN 90-277-0309-4, 94-010-3142-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
book/10.1007/978-94-010-3142-4.

Beltrametti:1981:PPS

[BC81] E. G. Beltrametti and G. Cassinelli.

Problems of the proposition–state structure of quantum mechanics. In Italian Studies in the Philosophy of Science [Chi81b], chapter 14, pages 215–235. ISBN 90-277-1073-2, 94-009-8937-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-009-
8937-5_14.

Brooks:1994:CHB

[BC94] Harvey Brooks and I. Bernard Cohen. A conversation with Harvey Brooks on the social sciences, the natural sciences, and public policy — conducted by I. Bernard Cohen. In Cohen [Coh94b], chapter 12, pages 363–398. ISBN 0-7923-2223-1, 94-017-3391-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 150.

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-017-3391-
                                              5_12.

Burtt:2002:RIM

[BC02] Burtt and R. G. Collingwood. A return to the ideas in the metaphysical foundations. In E. A. Burtt, Historian and Philosopher: A Study of the Author of The Metaphysical Foundations of Modern Physical Science [Vil02g], chapter 13, pages 215–232. ISBN 90-481-5937-7, 94-017-1331-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-017-1331-3_13.

Brain:2007:HCO

[BCK07] Robert Michael Brain, Robert S. Cohen, and Ole Knudsen. Hans Christian Ørsted and The Romantic Legacy In Science: Ideas, Disciplines, Practices, volume 241 of Boston Studies in the Philosophy of Science. Springer-Ver-lag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2007. ISBN 1-4020-2979-9 (hardcover), 1-4020-2987-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). xix + 442 pp. LCCN QC515.O37 H36 2007.

                                              URL https://link.springer.com/
                                              book/10.1007/978-1-4020-2987-5.

Bub:1974:IQM

[BD74] Jeffrey Bub and William Demopoulos. The interpretation of quantum mechanics. In Logical and Epistemological Studies in Contemporary Physics [CW74a], chapter 2, pages 92–122. ISBN 90-277-0391-4 (hardcover), 90-277-0377-9 (paperback), 94-010-2656-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 13; QC5.56. URL https:

//link.springer.com/chapter/10.
1007/978-94-010-2656-7_2.

Ben-David:1988:SIO

[BD88] Joseph Ben-David. Social interests and the organic physics of 1847. In Science in Reflection: The Israel Colloquium: Studies in History, Philosophy, and Sociology of Science Volume 3 [UM88], chapter 15, pages 193–200. ISBN 90-277-2713-9, 94-009-2957-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-009-2957-9_15.

Bahadir:2005:LOE

[BD05]  Osman  Recep  Bahadir  and

H. H. Günhan Danişman. Late Ottoman and early republican science periodicals. In Turkish Studies in the History and Philosophy of Science [IG05c], chapter 19, pages 285–308. ISBN 1-4020-3332-X (hardcover), 1-4020-3333-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 244.

URL https://link.springer.com/
chapter/10.1007/1-4020-3333-8_19.

Boddey:1995:BNF

[BdOU^{+}95] R. M. Boddey, O. C. de Oliveira, S. Urquiaga, V. M. Reis, F. L. de Olivares, V. L. D. Baldani, and J. Döbereiner. Biological nitrogen fixation associated with sugar cane and rice: Contributions and prospects for improvement. In No Truth Except in the Details: Essays in Honor of Martin J. Klein [KS95], chapter 9, pages 195–209. ISBN 94-010-4097-4, 94-011-0217-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-011-
0055-7_9.

Buttner:2004:GUT

[BDR04] Jochen Büttner, Peter Damerow, and Jürgen Renn. Galileo’s unpublished treatises. In The Reception of the Galilean Science of Motion in Seventeenth-Century Europe [PT04], chapter 6, pages 99–117. ISBN 1-4020-2454-1, 1-4020-2455-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC133 .R43 2004.

URL https://link.springer.com/
chapter/10.1007/978-1-4020-2455-
9_6.

Bar-Elli:1988:CNB

[BE88] Gilead Bar-Elli. Can a naturalist believe in universals? In Science in Reflection: The Israel Colloquium: Studies in History, Philosophy, and Sociology of Science Volume 3 [UM88], chapter 10, pages 117–122. ISBN 90-277-2713-9, 94-009-2957-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-009-2957-9_10.

Bretelle-Establet:2010:LYR

[BE10a]  Florence Bretelle-Establet.  Is the

Lower Yangzi River Region the only seat of medical knowledge in Late Imperial China? A glance at the Far South Region and at its medical documents. In Looking at it from Asia: the Processes that Shaped the Sources of History of Science [BE10b], chapter 11, pages 331–369. ISBN 90-481-3675-X (hardcover), 90-481-3676-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q127.A65 L66

2010. URL https://link.springer.
com/chapter/10.1007/978-90-481-
3676-6_11.

Bretelle-Establet:2010:LIA

[BE10b] Florence Bretelle-Establet. Looking at it from Asia: the Processes that Shaped the Sources of History of Science, volume 265 of Boston Studies in the Philosophy of Science. Springer-Ver-lag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2010. ISBN 90-481-3675-X (hardcover), 90-481-3676-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). xlvi + 426 pp. LCCN Q127.A65 L66 2010.

URL https://link.springer.com/
book/10.1007/978-90-481-3676-6.

Bretelle-Establet:2014:LTM

[BE14] Florence Bretelle-Establet. Learning and teaching medicine in Late Imperial China. In Bernard and Proust [BP14b], chapter 11, pages 273–301. ISBN 94-007-5121-4 (hardcover), 94-007-5122-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q181 .S35 2014; Q181 S387 2014. URL https:

//link.springer.com/chapter/10.
1007/978-94-007-5122-4_11.

Beaujouan:1975:RRE

[Bea75] Guy Beaujouan. Réflexions sur les rapports entre théorie et pratique au moyen Âge. (French) [Reflections on the relationship between theory and practice in the Middle Ages]. In Murdoch and Sylla [MS75b], pages 437–484. ISBN 90-277-0587-9, 94-010-1781-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN B721 .I584 1975. URL https://link.

                                              springer.com/chapter/10.1007/
                                              978-94-010-1781-7_12.

Beckner:1976:FT

[Bec76] Morton Beckner. Function and teleology. In Topics in the Philosophy of Biology [GM76], chapter 9, pages 197–212. ISBN 90-277-0596-8, 94-010-1829-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

                                              ???? URL https://link.springer.
                                              com/chapter/10.1007/978-94-010-
                                              1829-6_9.

Beck:1986:BSC

[Bec86] Heinrich Beck. Bio-social cybernetic determination — or responsible freedom? In Philosophy and Technology II: Information Technology and Computers in Theory and Practice [MH86], chapter 7, pages 85–95. ISBN 94-010-8510-2, 94-009-4512-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

                                              ???? URL https://link.springer.
                                              com/chapter/10.1007/978-94-009-
                                              4512-8_7.

Bechler:1991:AMN

[Bec91a] Zev Bechler. Absolute motion and the nature of inertial forces. In Newton’s Physics and the Conceptual Structure of the Scientific Revolution [Bec91r], chapter 12, pages 287–318. ISBN 0-7923-1054-3, 94-011-3276-3. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 127.

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-011-3276-
                                              3_12.

Bechler:1991:AIN

[Bec91b] Zev Bechler. Actual infinity and Newton’s calculus. In Newton’s Physics and the Conceptual Structure of the Scientific Revolution [Bec91r], chapter 9, pages 238–252. ISBN 0-7923-1054-3, 94-011-3276-3. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 127. URL https://link.

springer.com/chapter/10.1007/
978-94-011-3276-3_9.

Bechler:1991:APC

[Bec91c] Zev Bechler. Aristotelian and Platonic conceptions of explanation. In Newton’s Physics and the Conceptual Structure of the Scientific Revolution [Bec91r], chapter 1, pages 1–14. ISBN 0-7923-1054-3, 94-011-3276-3. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 127. URL https:/

/link.springer.com/chapter/10.
1007/978-94-011-3276-3_1.

Bechler:1991:APN

[Bec91d] Zev Bechler. Aristotle’s philosophy of nature and theory of potentiality. In Newton’s Physics and the Conceptual Structure of the Scientific Revolution [Bec91r], chapter 2, pages 15–44. ISBN 0-7923-1054-3, 94-011-3276-3. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 127.

URL https://link.springer.com/
chapter/10.1007/978-94-011-3276-
3_2.

Bechler:1991:BIL

[Bec91e] Zev Bechler. Bacon’s informative logic.

In Newton’s Physics and the Conceptual Structure of the Scientific Revolution [Bec91r], chapter 6, pages 94–104. ISBN 0-7923-1054-3, 94-011-3276-3. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 127.

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-011-3276-
                                              3_6.

Bechler:1991:BAC

[Bec91f] Zev Bechler. Berkeley’s Aristotelian critique of Newton’s physics. In Newton’s Physics and the Conceptual Structure of the Scientific Revolution [Bec91r], chapter 17, pages 476–505. ISBN 0-7923-1054-3, 94-011-3276-3. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 127.

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-011-3276-
                                              3_17.

Bechler:1991:CNP

[Bec91g] Zev Bechler. Circularity and Newton’s philosophy of nature. In Newton’s Physics and the Conceptual Structure of the Scientific Revolution [Bec91r], chapter 15, pages 398–436. ISBN 0-7923-1054-3, 94-011-3276-3. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 127.

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-011-3276-
                                              3_15.

Bechler:1991:CH

[Bec91h] Zev Bechler. The Copernican harmony. In Newton’s Physics and the Conceptual Structure of the Scientific Revolution [Bec91r], chapter 5, pages 75–93. ISBN 0-7923-1054-3, 94-011-3276-3. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 127.

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-011-3276-
                                              3_5.

Bechler:1991:DIL

[Bec91i] Zev Bechler. Descartes’ informative logic. In Newton’s Physics and the Conceptual Structure of the Scientific Revolution [Bec91r], chapter 8, pages 172–237. ISBN 0-7923-1054-3, 94-011-3276-3. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 127. URL https://link.

springer.com/chapter/10.1007/
978-94-011-3276-3_8.

Bechler:1991:E

[Bec91j] Zev Bechler. Epilogue. In Newton’s Physics and the Conceptual Structure of the Scientific Revolution [Bec91r], chapter 18, pages 506–512. ISBN 0-7923-1054-3, 94-011-3276-3. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 127.

URL https://link.springer.com/
chapter/10.1007/978-94-011-3276-
3_18.

Bechler:1991:IPG

[Bec91k] Zev Bechler. Informativity and paradox: Galileo’s conception of the nature of physical reality. In Newton’s Physics and the Conceptual Structure of the Scientific Revolution [Bec91r], chapter 7, pages 105–171. ISBN 0-7923-1054-3, 94-011-3276-3. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 127. URL https://link.

springer.com/chapter/10.1007/
978-94-011-3276-3_7.

Bechler:1991:I

[Bec91l] Zev Bechler. Introduction. In Newton’s Physics and the Conceptual Structure of the Scientific Revolution [Bec91r], chapter 4, pages 72–74. ISBN 0-7923-1054-3, 94-011-3276-3. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 127. URL https:/

/link.springer.com/chapter/10.
1007/978-94-011-3276-3_4.

Bechler:1991:LAP

[Bec91m] Zev Bechler. Leibniz’ Aristotelian philosophy of nature. In Newton’s Physics and the Conceptual Structure of the Scientific Revolution [Bec91r], chapter 16, pages 437–475. ISBN 0-7923-1054-3, 94-011-3276-3. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 127.

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-011-3276-
                                              3_16.

Bechler:1991:LME

[Bec91n] Zev Bechler. Locke and the meaning of “empiricism”. In Newton’s Physics and the Conceptual Structure of the Scientific Revolution [Bec91r], chapter 13, pages 319–341. ISBN 0-7923-1054-3, 94-011-3276-3. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 127. URL https://link.

                                              springer.com/chapter/10.1007/
                                              978-94-011-3276-3_13.

Bechler:1991:MNH

[Bec91o] Zev Bechler. Modern Newtonian historiography and the puzzle of Newton’s absolute space. In Newton’s Physics and the Conceptual Structure of the Scientific Revolution [Bec91r], chapter 11, pages 269–286. ISBN 0-7923-1054-3, 94-011-3276-3. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 127. URL https://link.

                                              springer.com/chapter/10.1007/
                                              978-94-011-3276-3_11.

Bechler:1991:NIP

[Bec91p] Zev Bechler. Newton’s invention of the problem of induction. In Newton’s Physics and the Conceptual Structure of the Scientific Revolution [Bec91r], chapter 14, pages 342–397. ISBN 0-7923-1054-3, 94-011-3276-3. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 127.

URL https://link.springer.com/
chapter/10.1007/978-94-011-3276-
3_14.

Bechler:1991:NLS

[Bec91q] Zev Bechler. Newton’s logic of space and time. In Newton’s Physics and the Conceptual Structure of the Scientific Revolution [Bec91r], chapter 10, pages 253–268. ISBN 0-7923-1054-3, 94-011-3276-3. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 127. URL https://link.

springer.com/chapter/10.1007/
978-94-011-3276-3_10.

Bechler:1991:NPC

[Bec91r] Zev Bechler. Newton’s Physics and the Conceptual Structure of the Scientific Revolution, volume 127 of Boston Studies in the Philosophy of Science. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 1991. ISBN 0-7923-1054-3, 94-011-3276-3. ISSN 0068-0346 (print), 2214-7942 (electronic). xviii + 588 pp. LCCN Q174 .B67 vol. 127.

URL https://link.springer.com/
book/10.1007/978-94-011-3276-3.

Bechler:1991:PCA

[Bec91s] Zev Bechler. Plato’s concept of the actual and his philosophy of nature. In Newton’s Physics and the Conceptual Structure of the Scientific Revolution [Bec91r], chapter 3, pages 45–71. ISBN 0-7923-1054-3, 94-011-3276-3. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 127.

URL https://link.springer.com/
chapter/10.1007/978-94-011-3276-
3_3.

Becker:2004:ESV

[Bec04] Oskar Becker. Eudoxos — Studien I. Eine Voreudoxische Proportionenlehre und Ihre Spuren bei Aristoteles und Euklid. (German) [Eudoxos — Studies I. A pre-Eudoxic proportion doctrine and its traces in Aristotle and Euclid]. In Classics in the History of Greek Mathematics [Chr04a], chapter 10, pages 191–209. ISBN 1-4020-0081-2 (hardcover), 1-4020-2640-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QA22 .C596 2004.

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-1-4020-2640-
                                              9_10.

Bedau:2009:EC

[Bed09a] Mark A. Bedau. The evolution of complexity. In Mapping the Future of Biology: Evolving Concepts and Theories [BMP09b], chapter 8, pages 111–130. ISBN 1-4020-9635-6 (hardcover), 1-4020-9636-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QH331 .M295 2009.

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-1-4020-9636-
                                              5_8.

Bednar:2009:OSE

[Bed09b] Miloslav Bednář. Open society and the European Union. In Rethinking Popper [PC09], chapter 18, pages 237–244. ISBN 1-4020-9337-3, 1-4020-9338-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN B1649.P64 R48 2009; Q174 .B67 2009.

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-1-4020-9338-
                                              8_18.

Beer:1980:P

[Bee80] Stafford Beer. Preface. In Autopoiesis and Cognition: The Realization of the Living [MV80a], chapter 7, pages 63– 72. ISBN 90-277-1016-3, 94-009-8947-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-009-8947-
4_7.

Beer:1989:DI

[Bee89] Gillian Beer. Discourses of the island.

In Literature and Science as Modes of Expression [Amr89], chapter 1, pages 1–27. ISBN 94-010-7531-X, 94-009-2297-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-009-
2297-6_1.

Beese:1991:MDP

[Bee91] Wolfgang Beese. Max Delbrück: A physicist in biology. In Woodward and Cohen [WC91], pages 415–422. ISBN 94-010-5395-2, 94-011-3164-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-011-3164-3_38.

Bekenstein:1986:FSC

[Bek86] Jacob D. Bekenstein. The fine-structure constant: From Eddington’s time to our own. In Ullmann-Margalit [UM86c], chapter 17, pages 209–224. ISBN 90-277-2160-2, 90-277-2161-0 (paperback), 94-009-4566-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 95 Q175.

URL https://link.springer.com/
chapter/10.1007/978-94-009-4566-
1_17.

Bellone:1981:HSH

[Bel81a] Enrico Bellone. The history of science as the history of dictionaries*. In Italian Studies in the Philosophy of Science [Chi81b], chapter 25, pages 493–507. ISBN 90-277-1073-2, 94-009-8937-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-009-8937-
                                              5_25.

Belloni:1981:DDA

[Bel81b] Luigi Belloni. The discovery of duodenal ancylostoma and of its pathogenic power. In Grmek et al. [GCC81b], pages 261–279. ISBN 90-277-1123-2, 94-010-1284-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic).

                                              LCCN ????  URL https://link.
                                              springer.com/chapter/10.1007/
                                              978-94-010-1284-3_13.

Bellone:1990:SHS

[Bel90] Enrico Bellone. Science and history of science. In Levere and Shea [LS90], chapter 13, pages 307–317. ISBN 0-7923-0420-9, 94-009-1878-X. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 120.

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-009-1878-
                                              8_13.

Bell:1992:UAE

[Bel92a] James A. Bell. Universalization in archaeological explanation. In Metaarchaeology: Reflections by Archaeologists and Philosophers [Emb92b], chapter 6, pages 143–163. ISBN 94-010-4806-1, 94-011-1826-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

                                              ???? URL https://link.springer.
                                              com/chapter/10.1007/978-94-011-
                                              1826-2_6.

Beller:1992:GBC

[Bel92b] Mara Beller. The genesis of Bohr’s complementarity principle and the Bohr–Heisenberg dialogue. In The Scientific Enterprise: The Bar-Hillel Colloquium: Studies in History, Philosophy, and Sociology of Science, Volume 4 [UM92], chapter 18, pages 273–293. ISBN 94-010-5190-9, 94-011-2688-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174

.S354 1992. URL https://link.
springer.com/chapter/10.1007/
978-94-011-2688-5_18.

Bellissima:1993:IMS

[Bel93] Fabio Bellissima. An introduction to modal semantics. In Bridging the Gap: Philosophy, Mathematics, and Physics: Lectures on the Foundations of Science [CCG93], chapter 2, pages 39–69. ISBN 94-010-5101-1, 94-011-2496-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-011-2496-6_2.

Beller:1995:LM

[Bel95] Walter Beller. Logic in Mexico. In Mexican Studies in the History and Philosophy of Science [RC95], chapter 1, pages 1–23. ISBN 94-010-6535-7 (hardcover), 94-009-0109-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-009-0109-4_1.

Beller:1996:BIA

[Bel96] Mara Beller. Bohm and the “inevitability” of acausality. In Bohmian Mechanics and Quantum Theory: an Appraisal [CFG96], chapter 15, pages 211–230. ISBN 90-481-4698-4, 94-015-8715-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC173.96-174.52. URL https://link.

springer.com/chapter/10.1007/
978-94-015-8715-0_15.

Beller:2003:IIG

[Bel03] Mara Beller. Inevitability, inseparability and Gedanken measurement. In Renn et al. [RDS03], chapter 22, pages 439–450. ISBN 1-4020-1284-5, 1-4020-1285-3 (paperback). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 234. URL https:/

                                              /link.springer.com/chapter/10.
                                              1007/978-94-010-0111-3_22.

Belkind:2012:C

[Bel12a] Ori Belkind. Conclusion. In Physical Systems: Conceptual Pathways between Flat Space–time and Matter [Bel12g], chapter 9, pages 217–225. ISBN 94-007-2372-5, 94-007-2373-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-007-2373-3_9.

Belkind:2012:HNM

[Bel12b] Ori Belkind. The history of Newtonian mass. In Physical Systems: Conceptual Pathways between Flat Space–time and Matter [Bel12g], chapter 5, pages 119–144. ISBN 94-007-2372-5, 94-007-2373-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

                                              ???? URL https://link.springer.
                                              com/chapter/10.1007/978-94-007-
                                              2373-3_5.

Belkind:2012:ISP

[Bel12c] Ori Belkind. Interpretations of spacetime and the principle of relativity. In Physical Systems: Conceptual Pathways between Flat Space–time and Matter [Bel12g], chapter 2, pages 31–58. ISBN 94-007-2372-5, 94-007-2373-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-007-2373-3_2.

Belkind:2012:MT

[Bel12d] Ori Belkind. The metaphysics of time. In Physical Systems: Conceptual Pathways between Flat Space–time and Matter [Bel12g], chapter 4, pages 93–117. ISBN 94-007-2372-5, 94-007-2373-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-007-
2373-3_4.

Belkind:2012:PSM

[Bel12e] Ori Belkind. Physical systems and mass. In Physical Systems: Conceptual Pathways between Flat Space–time and Matter [Bel12g], chapter 6, pages 145–162. ISBN 94-007-2372-5, 94-007-2373-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-007-
2373-3_6.

Belkind:2012:PSP

[Bel12f] Ori Belkind. Physical systems and physical thought. In Physical Systems: Conceptual Pathways between Flat Space–time and Matter [Bel12g], chapter 1, pages 1–30. ISBN 94-007-2372-5, 94-007-2373-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic).

LCCN ????  URL https://link.
springer.com/chapter/10.1007/
978-94-007-2373-3_1.

Belkind:2012:PSC

[Bel12g] Ori Belkind. Physical Systems: Conceptual Pathways between Flat Space–time and Matter, volume 264 of Boston Studies in the Philosophy of Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2012. ISBN 94-007-2372-5, 94-007-2373-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). xiii + 234 pp. LCCN ???? URL

https://link.springer.com/book/
10.1007/978-94-007-2373-3.

Belkind:2012:PMR

[Bel12h] Ori Belkind. Primitive motion relationalism. In Physical Systems: Conceptual Pathways between Flat Space–time and Matter [Bel12g], chapter 3, pages 59– 91. ISBN 94-007-2372-5, 94-007-2373-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-007-2373-
3_3.

Belkind:2012:STR

[Bel12i] Ori Belkind. The special theory of relativity. In Physical Systems: Conceptual Pathways between Flat Space–time and Matter [Bel12g], chapter 8, pages 191–215. ISBN 94-007-2372-5, 94-007-2373-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-007-
2373-3_8.

Belkind:2012:SAN

[Bel12j] Ori Belkind. Structural assumptions, Newton’s scientific method, and the universal law of gravitation. In Physical Systems: Conceptual Pathways between Flat Space–time and Matter [Bel12g], chapter 7, pages 163–190. ISBN 94-007-2372-5, 94-007-2373-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-007-2373-3_7.

Belot:2015:PCR

[Bel15] Robert Belot. The political and cultural revolution of the CNRS: An attempt at the systematic organisation of research in opposition to “the academic spirit”. In Sciences in the Universities of Europe, Nineteenth and Twentieth Centuries: Academic Landscapes [SDG15b], chapter 15, pages 245–263. ISBN 94-017-9635-1, 94-017-9636-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q183.4.E85

2015. URL https://link.springer.
com/chapter/10.1007/978-94-017-
9636-1_15.

Bencivenga:1981:FS

[Ben81] E. Bencivenga. Free semantics. In Italian Studies in the Philosophy of Science [Chi81b], chapter 3, pages 31–48. ISBN 90-277-1073-2, 94-009-8937-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-009-8937-5_3.

Benfey:2006:CSS

[Ben06] Otto Theodor Benfey. The conceptual structure of the sciences. In Philosophy of Chemistry: Synthesis of a New Discipline [BSM06b], chapter 5, pages 95–117. ISBN 1-4020-2455-X, 1-4020-3256-0, 1-4020-3261-7. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 242 QD6. URL https:

//link.springer.com/chapter/10.
1007/1-4020-3261-7_5.

Bengoetxea:2015:KMR

[Ben15] Juan Bautista Bengoetxea. Knowledge and moral responsibility for online technologies. In New Perspectives on Technology, Values, and Ethics: Theoretical and Practical [Gon15a], chapter 5, pages 89–103. ISBN 3-319-21869-7 (hardcover), 3-319-21870-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-3-319-21870-
                                              0_5.

Bergmann:1970:EMC

[Ber70] Peter G. Bergmann. Ernst Mach and contemporary physics. In Ernst Mach: Physicist and Philosopher [CS70], chapter 5, pages 69–78. ISBN 90-481-8318-9, 90-277-0016-8, 94-017-1462-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-017-1462-
                                              4_5.

Bergman:1974:CCL

[Ber74a] H. Bergman. The controversy concerning the law of causality in contemporary physics. In Logical and Epistemological Studies in Contemporary Physics [CW74a], chapter 16, pages 395–396. ISBN 90-277-0391-4 (hardcover), 90-277-0377-9 (paperback), 94-010-2656-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 13; QC5.56. URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-010-2656-7_16.

Bergman:1974:PRA

[Ber74b] H. Bergman. Personal remembrances of Albert Einstein. In Logical and Epistemological Studies in Contemporary Physics [CW74a], chapter 15, pages 388–394. ISBN 90-277-0391-4 (hardcover), 90-277-0377-9 (paperback), 94-010-2656-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 13; QC5.56.

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-010-2656-
                                              7_15.

Bergmann:1974:CS

[Ber74c] Peter G. Bergmann. Cosmology as a science. In Seeger and Cohen [SC74b], pages 181–188. ISBN 90-277-0390-6, 90-277-0376-0 (paperback), 94-010-2126-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 11. URL https:/

/link.springer.com/chapter/10.
1007/978-94-010-2126-5_10.

Berkson:1974:SPI

[Ber74d] William Berkson. Some practical issues in the recent controversy on the nature of scientific revolutions. In Methodological and Historical Essays in the Natural and Social Sciences [CW74b], chapter 10, pages 197–210. ISBN 90-277-0378-7, 94-010-2128-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 14.

URL https://link.springer.com/
chapter/10.1007/978-94-010-2128-
9_10.

Bergson:1976:ZP

[Ber76a] H. Bergson. On Zeno’s parodoxes.

In The Concepts of Space and Time: Their Structure and Their Development [Čap76], chapter 43, pages 245–250. ISBN 90-277-0375-2, 94-010-1727-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN BD632.

URL https://link.springer.com/
chapter/10.1007/978-94-010-1727-
5_43.

Berkeley:1976:CN

[Ber76b] G. Berkeley. Criticism of Newton.

In The Concepts of Space and Time: Their Structure and Their Development [Čap76], chapter 46, pages 267–271. ISBN 90-277-0375-2, 94-010-1727-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN BD632.

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-010-1727-
                                              5_46.

Berkson:1976:LOL

[Ber76c] William Berkson. Lakatos one and Lakatos two: An appreciation. In Essays in Memory of Imre Lakatos [CFW76], chapter 5, pages 39–54. ISBN 90-277-0655-7, 94-010-1451-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-010-1451-
                                              9_5.

Bernini:1981:TMS

[Ber81] Sergio Bernini. A temporalization of modal semantics. In Italian Studies in the Philosophy of Science [Chi81b], chapter 4, pages 49–58. ISBN 90-277-1073-2, 94-009-8937-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-009-8937-5_4.

Berka:1983:I

[Ber83a] Karel Berka. Introduction. In Measurement: Its Concepts, Theories and Problems [Ber83d], chapter 1, pages 1– 13. ISBN 94-009-7830-8, 94-009-7828-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-009-7828-
                                              7_1.

Berka:1983:Mb

[Ber83b] Karel Berka. Magnitudes. In Measurement: Its Concepts, Theories and Problems [Ber83d], chapter 3, pages 35–82. ISBN 94-009-7830-8, 94-009-7828-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-009-
7828-7_3.

Berka:1983:Ma

[Ber83c] Karel Berka. Measurement. In Measurement: Its Concepts, Theories and Problems [Ber83d], chapter 2, pages 14–34. ISBN 94-009-7830-8, 94-009-7828-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-009-
7828-7_2.

Berka:1983:MCT

[Ber83d]  Karel Berka.  Measurement: Its

Concepts, Theories and Problems, volume 72 of Boston Studies in the Philosophy of Science. D. Reidel, Dordrecht, The Netherlands; Boston, MA, USA; Lancaster, UK; Tokyo, Japan, 1983. ISBN 94-009-7830-8, 94-009-7828-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). 268 (est.) pp. LCCN ???? URL

https://link.springer.com/book/
10.1007/978-94-009-7828-7.

Berka:1983:MPM

[Ber83e] Karel Berka. Methodological problems of measurement. In Measurement: Its Concepts, Theories and Problems [Ber83d], chapter 8, pages 178–204. ISBN 94-009-7830-8, 94-009-7828-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-009-7828-
7_8.

Berka:1983:PPM

[Ber83f] Karel Berka. Philosophical problems of measurement. In Measurement: Its Concepts, Theories and Problems [Ber83d], chapter 9, pages 205–217. ISBN 94-009-7830-8, 94-009-7828-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-009-7828-
7_9.

Berka:1983:Q

[Ber83g] Karel Berka. Quantification. In Measurement: Its Concepts, Theories and Problems [Ber83d], chapter 5, pages 101–111. ISBN 94-009-7830-8, 94-009-7828-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-009-
7828-7_5.

Berka:1983:S

[Ber83h] Karel Berka. Scales. In Measurement: Its Concepts, Theories and Problems [Ber83d], chapter 4, pages 83–100. ISBN 94-009-7830-8, 94-009-7828-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-009-7828-
7_4.

Berka:1983:SM

[Ber83i] Karel Berka. Scales of measurement. In Language, Logic and Method [CW83b], chapter 1, pages 1–73. ISBN 94-009-7704-2, 94-009-7702-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-009-7702-0_1.

Berka:1983:TM

[Ber83j] Karel Berka. Theory of measurement. In Measurement: Its Concepts, Theories and Problems [Ber83d], chapter 6, pages 112–157. ISBN 94-009-7830-8, 94-009-7828-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-009-
7828-7_6.

Berka:1983:TS

[Ber83k] Karel Berka. Theory of scales. In Measurement: Its Concepts, Theories and Problems [Ber83d], chapter 7, pages 158–177. ISBN 94-009-7830-8, 94-009-7828-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-009-
7828-7_7.

Bertocci:1984:SLI

[Ber84] Peter Bertocci. The scholar, the liberal ideal, and the philosophy of science. In Hegel and the Sciences [CW84a], chapter 1, pages 3–10. ISBN 94-009-6235-5, 94-009-6233-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-009-6233-0_1.

Berger:1986:LFU

[Ber86] Fred R. Berger. Love, friendship, and utility: On practical reason and reductionism. In Cohen and Wartofsky [CW85], chapter 1, pages 3–21. ISBN 90-277-1971-3, 94-009-5345-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-009-5349-9_1.

Bertoletti:1994:ATJ

[Ber94] Stefano Fabbri Bertoletti. The anthropological theory of Johann Friedrich Blumenbach. In Poggi and Bossi [PB94], chapter 7, pages 103–125. ISBN 0-7923-2336-X, 94-017-2921-2. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 152.

URL https://link.springer.com/
chapter/10.1007/978-94-017-2921-
5_7.

Berka:1995:RNE

[Ber95a] Karel Berka. Rational and nonrational elements in the history of science. In Gavroglu et al. [GSW95c], chapter 16, pages 273–285. ISBN 0-7923-2991-0 (set), 0-7923-2988-0, 94-017-2658-2. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 163.

URL https://link.springer.com/
chapter/10.1007/978-94-017-2658-
0_16.

Berkson:1995:TCH

[Ber95b] William Berkson. Two commandments for the humanities. In Critical Rationalism, the Social Sciences and the Humanities: Essays for Joseph Agassi. Volume II [JL95b], chapter 13, pages 221–233. ISBN 0-7923-2961-9, 90-481-4430-2, 94-017-0441-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN B29 .C75 1995.

URL https://link.springer.com/
chapter/10.1007/978-94-017-0441-
0_13.

Berndl:1996:GEU

[Ber96a] Karin Berndl. Global existence and uniqueness of Bohmian trajectories. In Bohmian Mechanics and Quantum Theory: an Appraisal [CFG96], chapter 5, pages 77–86. ISBN 90-481-4698-4, 94-015-8715-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC173.96-174.52. URL https:

//link.springer.com/chapter/10.
1007/978-94-015-8715-0_5.

Bernier:1996:WMT

[Ber96b] Paul Bernier. Why Marr’s theory of vision is not anti-individualist. In Québec Studies in the Philosophy of Science: Part II: Biology, Psychology, Cognitive Science and Economics Essays in Honor of Hugues Leblanc [MC96], chapter 7, pages 99–111. ISBN 94-010-6537-3, 94-009-0113-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-009-0113-1_7.

Bergstrom:2000:QET

[Ber00] Lars Bergström. Quine, empiricism, and truth. In Knowledge, Language and Logic: Questions for Quine [OK00], chapter 6, pages 63–79. ISBN 1-4020-0253-X, 94-011-3933-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-011-3933-5_6.

Bergandi:2013:EEE

[Ber13a] Donato Bergandi. Ecology, evolution, ethics: In search of a meta-paradigm — an introduction. In The Structural Links between Ecology, Evolution and Ethics: The Virtuous Epistemic Circle [Ber13c], chapter 1, pages 1–28. ISBN 94-007-5066-8, 94-007-5067-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN GE42 .S77 2013; Q175 .B73 v. 296. URL https:/

/link.springer.com/chapter/10.
1007/978-94-007-5067-8_1.

Bergandi:2013:EEP

[Ber13b] Donato Bergandi. Epilogue: The epistemic and practical circle in an evolutionary, ecologically sustainable society. In The Structural Links between Ecology, Evolution and Ethics: The Virtuous Epistemic Circle [Ber13c], chapter 11, pages 151–158. ISBN 94-007-5066-8, 94-007-5067-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN GE42 .S77 2013; Q175 .B73 v. 296. URL https:/

                                              /link.springer.com/chapter/10.
                                              1007/978-94-007-5067-8_11.

Bergandi:2013:SLB

[Ber13c] Donato Bergandi, editor. The Structural Links between Ecology, Evolution and Ethics: The Virtuous Epistemic Circle, volume 296 of Boston Studies in the Philosophy and History of Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2013. ISBN 94-007-5066-8, 94-007-5067-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). xv + 179 pp. LCCN GE42 .S77 2013; Q175 .B73 v. 296. URL

                                              https://link.springer.com/book/
                                              10.1007/978-94-007-5067-8.

Bernard:2014:WSD

[Ber14] Alain Bernard. In what sense did Theon’s commentary on the Almagest have a didactic purpose? In Bernard and Proust [BP14b], chapter 5, pages 95–121. ISBN 94-007-5121-4 (hardcover), 94-007-5122-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q181 .S35 2014; Q181

                                              S387 2014.  URL https://link.
                                              springer.com/chapter/10.1007/
                                              978-94-007-5122-4_5.

Berkovitz:2017:TCE

[Ber17a] Joseph Berkovitz. On time, causation and explanation in the causally symmetric Bohmian model of quantum mechanics. In Time of Nature and the Nature of Time: Philosophical Perspectives of Time in Natural Sciences [BH17b], chapter 8, pages 139–172. ISBN 3-319-53723-7 (hardcover), 3-319-53725-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC173.59.S65 T56 2017. URL https:

//link.springer.com/chapter/10.
1007/978-3-319-53725-2_8.

Berkson:2017:WPU

[Ber17b] William Berkson. What is the problem?

Understanding the history of ideas. In Encouraging Openness: Essays for Joseph Agassi on the Occasion of His 90th Birthday [BAG17], chapter 16, pages 207–214. ISBN 3-319-57668-2 (hardcover), 3-319-57669-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN B833.E53 2017.

URL https://link.springer.com/
chapter/10.1007/978-3-319-57669-
5_16.

Beurton:1991:WKS

[Beu91] Peter Beurton. What keeps a species together? In Woodward and Cohen [WC91], pages 293–298. ISBN 94-010-5395-2, 94-011-3164-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-011-3164-3_26.

Beutelspacher:1996:SGL

[Beu96] Albrecht Beutelspacher. A survey of Grassmann’s Lineale Ausdehnungslehre. In Schubring [Sch96f], pages 3–6. ISBN 90-481-4758-1, 94-015-8753-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QA8.9-10.3.

URL https://link.springer.com/
chapter/10.1007/978-94-015-8753-
2_1.

Brown:1996:CEP

[BEW96] Harvey R. Brown, Andrew Elby, and Robert Weingard. Cause and effect in the pilot-wave interpretation of quantum mechanics. In Bohmian Mechanics and Quantum Theory: an Appraisal [CFG96], chapter 21, pages 309–319. ISBN 90-481-4698-4, 94-015-8715-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC173.96-174.52. URL https://link.

springer.com/chapter/10.1007/
978-94-015-8715-0_21.

Beller:1994:BRE

[BF94] Mara Beller and Arthur Fine. Bohr’s response to EPR. In Faye and Folse [FF94], chapter 1, pages 1–31. ISBN 0-7923-2378-5 (hardcover), 94-015-8106-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 153 1994; QC16 1994. URL https:

//link.springer.com/chapter/10.
1007/978-94-015-8106-6_1.

Brenner:2009:FSP

[BG09] Anastasios Brenner and Jean Gayon, editors. French Studies in the Philosophy of Science: Contemporary Research in France, volume 276 of Boston Studies in the Philosophy of Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2009. ISBN 1-4020-9367-5, 1-4020-9368-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). vi + 375 pp. LCCN Q174 .F746 2009. URL

https://link.springer.com/book/
10.1007/978-1-4020-9368-5.

Bonnay:2018:DTL

[BG18] D. Bonnay and H. Galinon. Deflationary truth is a logical notion. In Piazza and Pulcini [PP18], pages 71– 88. ISBN 3-319-93341-8, 3-319-93342-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QA8.4.

URL https://link.springer.com/
chapter/10.1007/978-3-319-93342-
9_5.

Bouton:2017:ITB

[BH17a] Christophe Bouton and Philippe Huneman. Introduction: Time between metaphysics and natural sciences: From physics to biology. In Time of Nature and the Nature of Time: Philosophical Perspectives of Time in Natural Sciences [BH17b], chapter 1, pages 1–20. ISBN 3-319-53723-7 (hardcover), 3-319-53725-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC173.59.S65 T56 2017. URL https:

//link.springer.com/chapter/10.
1007/978-3-319-53725-2_1.

Bouton:2017:TNN

[BH17b] Christophe Bouton and Philippe Huneman. Time of Nature and the Nature of Time: Philosophical Perspectives of Time in Natural Sciences, volume 326 of Boston Studies in the Philosophy and History of Science. Springer-Ver-lag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2017. ISBN 3-319-53723-7 (hardcover), 3-319-53725-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). xiii + 403 pp. LCCN QC173.59.S65 T56 2017.

URL https://link.springer.com/
book/10.1007/978-3-319-53725-2.

Baird:1998:HHC

[BHN98] Davis Baird, R. I. G. Hughes, and Alfred Nordmann. Heinrich Hertz: Classical Physicist, Modern Philosopher, volume 198 of Boston Studies in the Philosophy of Science. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 1998. ISBN 90-481-4881-2, 94-015-8855-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). xi + 324 pp. LCCN ???? URL

https://link.springer.com/book/
10.1007/978-94-015-8855-3.

Bhushan:2006:CKN

[Bhu06] Nalini Bhushan. Are chemical kinds natural kinds? In Philosophy of Chemistry: Synthesis of a New Discipline [BSM06b], chapter 17, pages 327–336. ISBN 1-4020-2455-X, 1-4020-3256-0, 1-4020-3261-7. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 242 QD6. URL https:/

                                              /link.springer.com/chapter/10.
                                              1007/1-4020-3261-7_17.

Biddle:2011:BMS

[Bid11] Justin Biddle. Bringing the marketplace into science: On the neoliberal defense of the commercialization of scientific research. In Science in the Context of Application [CN11a], chapter 15, pages 245–269. ISBN 90-481-9050-9, 90-481-9051-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q180.55.M4 S385 2011. URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-90-481-9051-5_15.

Biermann:1974:AHF

[Bie74] Kurt-R. Biermann. Die Alexander–Von-Humboldt-Forschung an der Akademie der Wissenschaften der D.D.R. — Ergebnisse und Ziele. (German) [The Alexander von Humboldt Research Institute in the D.D.R. — results and goals]. In For Dirk Struik: Scientific, Historical and Political Essays in Honor of Dirk J. Struik [CSW74], chapter 24, pages 295–305. ISBN 90-277-0379-5, 94-010-2115-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-010-2115-9_24.

Bierwisch:1976:SDL

[Bie76]  Manfred Bierwisch.  Social differ-

entiation of language structure. In Language in Focus: Foundations, Methods and Systems: Essays in Memory of Yehoshua Bar-Hillel [Kas76b], chapter 18, pages 407–456. ISBN 90-277-0644-1, 90-277-0645-X, 94-010-1876-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

P101-120.  URL https://link.
springer.com/chapter/10.1007/
978-94-010-1876-0_18.

Bielecki:1982:QMS

[Bie82] Marek Bielecki. Quantum mechanics and the structure of physical theories. In Polish Essays in the Philosophy of the Natural Sciences [Kra82c], chapter 27, pages 309–317. ISBN 94-009-7705-0 (e-book), 90-277-1287-5. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-009-7705-1_27.

Biezunski:1987:ERP

[Bie87] Michel Biezunski. Einstein’s reception in Paris in 1922. In Glick [Gli87a], chapter 5, pages 169–188. ISBN 90-277-2498-9, 94-009-3875-6. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 103; QC173.5. US$79.00

(US). URL https://link.springer.
com/chapter/10.1007/978-94-009-
3875-5_5.

Bienenstock:1995:HCT

[Bie95] Myriam Bienenstock. Hegel’s conception of teleology. In Science, Mind and Art: Essays on science and the humanistic understanding in art, epistemology, religion and ethics In honor of Robert S. Cohen [GSW95b], chapter 5, pages 55–70. ISBN 0-7923-2990-2 (hardcover), 94-011-0469-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol.

165 BD418.3. URL https://link.

                                              springer.com/chapter/10.1007/
                                              978-94-011-0469-2_5.

Binglun:1996:PAS

[Bin96] Zhang Binglun. A preliminary analysis of scientific development and its causes in Anhui Province during the Ming and Qing Dynasties. In Chinese Studies in the History and Philosophy of Science and Technology [DC96], chapter 27, pages 327–344. ISBN 90-481-4546-5, 94-015-8717-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 180; U390 .N19

                                              1996. URL https://link.springer.
                                              com/chapter/10.1007/978-94-015-
                                              8717-4_27.

Bird:2007:SE

[Bir07] Alexander Bird. Selection and explanation. In Rethinking Explanation [PY07], chapter 9, pages 131–136. ISBN 1-4020-5580-3 (paperback), 1-4020-5581-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN BD237 .R48 2007; Q175.32.E97 R48

                                              2007. URL https://link.springer.
                                              com/chapter/10.1007/978-1-4020-
                                              5581-2_9.

Bird:2008:IN

[Bir08] Alexander Bird. Incommensurability naturalized. In Rethinking Scientific Change and Theory Comparison: Stabilities, Ruptures, Incommensurabilities? [SSHH08], chapter 2, pages 21–39. ISBN 1-4020-6274-5 (hardcover), 1-4020-6279-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175 .R397 2008. URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-1-4020-6279-7_2.

Birner:2009:GSE

[Bir09] Jack Birner. From group selection to ecological niches. In Rethinking Popper [PC09], chapter 15, pages 185–202. ISBN 1-4020-9337-3, 1-4020-9338-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN B1649.P64 R48 2009; Q174 .B67 2009.

URL https://link.springer.com/
chapter/10.1007/978-1-4020-9338-
8_15.

Bird:2015:KHS

[Bir15] Alexander Bird. Kuhn and the historiography of science. In Kuhn’s Structure of Scientific Revolutions — 50 Years On [DB15b], chapter 3, pages 23–38. ISBN 3-319-13382-9, 3-319-13383-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175.3

.K84 2015.  URL https://link.
springer.com/chapter/10.1007/
978-3-319-13383-6_3.

Bitsakis:1990:LNE

[Bit90] Eftichios Bitsakis. Locality: A new enigma for physics. In Greek Studies in the Philosophy and History of Science [Nic90], chapter 20, pages 315–334. ISBN 94-010-7403-8, 94-009-2015-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-009-2015-
6_20.

Bitbol:1996:AS

[Bit96a] Michel Bitbol. The analytical stance.

In Schrödinger’s Philosophy of Quantum Mechanics [Bit96e], chapter 3, pages 79–90. ISBN 0-7923-4266-6, 94-009-1772-4. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 188; QC173.98. URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-009-1772-9_3.

Bitbol:1996:CRF

[Bit96b] Michel Bitbol. Complemetarity, representation and facts. In Schrödinger’s Philosophy of Quantum Mechanics [Bit96e], chapter 6, pages 211–261. ISBN 0-7923-4266-6, 94-009-1772-4. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 188; QC173.98. URL https://link.

                                              springer.com/chapter/10.1007/
                                              978-94-009-1772-9_6.

Bitbol:1996:C

[Bit96c]  Michel Bitbol.  Conclusion.  In

Schrödinger’s Philosophy of Quantum Mechanics [Bit96e], chapter 7, pages 263–265. ISBN 0-7923-4266-6, 94-009-1772-4. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 188; QC173.98. URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-009-1772-9_7.

Bitbol:1996:CBS

[Bit96d] Michel Bitbol. The controversy between Schrödinger and the Göttingen– Copenhagen physicists in the 1950’s. In Schrödinger’s Philosophy of Quantum Mechanics [Bit96e], chapter 1, pages 1– 33. ISBN 0-7923-4266-6, 94-009-1772-4. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 188; QC173.98. URL https://link.

                                              springer.com/chapter/10.1007/
                                              978-94-009-1772-9_1.

Bitbol:1996:SPQ

[Bit96e] Michel Bitbol. Schrödinger’s Philosophy of Quantum Mechanics, volume 188 of Boston Studies in the Philosophy of Science. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 1996. ISBN 0-7923-4266-6, 94-009-1772-4. ISSN 0068-0346 (print), 2214-7942 (electronic). xi + 285 pp. LCCN Q174 .B67 vol. 188;

QC173.98. URL http://www.loc.
gov/catdir/enhancements/fy0813/
96042123-d.html;   http://www.
loc.gov/catdir/enhancements/
fy0813/96042123-t.html;   https:
//link.springer.com/book/10.
1007/978-94-009-1772-9.

Bitbol:1996:STP

[Bit96f] Michel Bitbol. Schrödinger’s theoretical project. In Schrödinger’s Philosophy of Quantum Mechanics [Bit96e], chapter 2, pages 34–78. ISBN 0-7923-4266-6, 94-009-1772-4. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 188; QC173.98. URL https:

//link.springer.com/chapter/10.
1007/978-94-009-1772-9_2.

Bitbol:1996:TEL

[Bit96g] Michel Bitbol. The ‘thing’ of everyday life. In Schrödinger’s Philosophy of Quantum Mechanics [Bit96e], chapter 5, pages 159–210. ISBN 0-7923-4266-6, 94-009-1772-4. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 188;

QC173.98.  URL https://link.
springer.com/chapter/10.1007/
978-94-009-1772-9_5.

Bitbol:1996:TNO

[Bit96h] Michel Bitbol. Towards a new ontology.

In Schrödinger’s Philosophy of Quantum Mechanics [Bit96e], chapter 4, pages 91–158. ISBN 0-7923-4266-6, 94-009-1772-4. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 188; QC173.98. URL https:

//link.springer.com/chapter/10.
1007/978-94-009-1772-9_4.

Bitsakis:1997:VNT

[Bit97] Eftichios Bitsakis. On the validity of von Neumann’s theorem. In Ginev and Cohen [GC97], chapter 4, pages 47–62. ISBN 0-7923-4444-8 (hardcover), 94-011-5788-X. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 192. URL https://link.

                                              springer.com/chapter/10.1007/
                                              978-94-011-5788-9_4.

Blackmore:2001:AHP

[BIT01a] J. Blackmore, R. Itagaki, and S. Tanaka. Alois höfler — polymath. In Ernst Mach’s Vienna 1895–1930: Or Phenomenalism as Philosophy of Science [BIT01c], pages 237–276. ISBN 90-481-5752-8, 94-015-9690-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-015-9690-9_11.

Blackmore:2001:CMP

[BIT01b] J. Blackmore, R. Itagaki, and S. Tanaka. Carnap’s Machist “phase”. In Ernst Mach’s Vienna 1895–1930: Or Phenomenalism as Philosophy of Science [BIT01c], pages 159–186. ISBN 90-481-5752-8, 94-015-9690-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-015-9690-9_8.

Blackmore:2001:EMV

[BIT01c] J. Blackmore, R. Itagaki, and S. Tanaka, editors. Ernst Mach’s Vienna 1895–1930: Or Phenomenalism as Philosophy of Science, volume 218 of Boston Studies in the Philosophy of Science. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 2001. ISBN 90-481-5752-8, 94-015-9690-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
book/10.1007/978-94-015-9690-9.

Blackmore:2001:FA

[BIT01d] J. Blackmore, R. Itagaki, and S. Tanaka. Friedrich Adler. In Ernst Mach’s Vienna 1895–1930: Or Phenomenalism as Philosophy of Science [BIT01c], pages 29–59. ISBN 90-481-5752-8, 94-015-9690-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-015-9690-9_2.

Blackmore:2001:HVJ

[BIT01e] J. Blackmore, R. Itagaki, and S. Tanaka. Husserl vs. Jerusalem. In Ernst Mach’s Vienna 1895–1930: Or Phenomenalism as Philosophy of Science [BIT01c], pages 211–235. ISBN 90-481-5752-8, 94-015-9690-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-015-9690-9_10.

Blackmore:2001:I

[BIT01f] J. Blackmore, R. Itagaki, and S. Tanaka. Introduction. In Ernst Mach’s Vienna 1895–1930: Or Phenomenalism as Philosophy of Science [BIT01c], pages 1–27. ISBN 90-481-5752-8, 94-015-9690-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-015-9690-9_1.

Blackmore:2001:PF

[BIT01g] J. Blackmore, R. Itagaki, and S. Tanaka. Philipp Frank. In Ernst Mach’s Vienna 1895–1930:

Or Phenomenalism as Philosophy of Science [BIT01c], pages 61–83. ISBN 90-481-5752-8, 94-015-9690-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-015-9690-9_3.

Blackmore:2001:UVP

[BIT01h] J. Blackmore, R. Itagaki, and S. Tanaka. The University of Vienna Philosophical Society. In Ernst Mach’s Vienna 1895–1930: Or Phenomenalism as Philosophy of Science [BIT01c], pages 277–314. ISBN 90-481-5752-8, 94-015-9690-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-015-9690-9_12.

Bitbol:2008:MO

[Bit08] Michel Bitbol. Method and objectivity. In Rethinking Scientific Change and Theory Comparison: Stabilities, Ruptures, Incommensurabilities? [SSHH08], chapter 21, pages 291–296. ISBN 1-4020-6274-5 (hardcover), 1-4020-6279-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175 .R397 2008. URL https://link.

springer.com/chapter/10.1007/
978-1-4020-6279-7_21.

Brody:1989:PPP

[BJ89] Martin Brody and Allan Janik. Paradigms, politics and persuasion: Sociological aspects of musical controversy. In Style, Politics and the Future of Philosophy [Jan89l], chapter 15, pages 225–263. ISBN 94-010-7508-5, 94-009-2251-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic).

LCCN ????  URL https://link.
springer.com/chapter/10.1007/
978-94-009-2251-8_15.

Bantilan:1995:REI

[BJ95] M. C. S. Bantilan and C. Johansen.

Research evaluation and impact analysis of biological nitrogen fixation. In No Truth Except in the Details: Essays in Honor of Martin J. Klein [KS95], chapter 14, pages 279–286. ISBN 94-010-4097-4, 94-011-0217-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-011-0055-7_14.

Blackmore:1992:EMD

[Bla92a] John Blackmore. Ernst Mach — A Deeper Look: Documents and New Perspectives, volume 143 of Boston Studies in the Philosophy of Science. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 1992. ISBN 94-010-5230-1, 94-011-2771-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). xix + 463 pp. LCCN C5.53; QC6.9.

URL https://link.springer.com/
book/10.1007/978-94-011-2771-4.

Blackmore:1992:EMH

[Bla92b] John Blackmore. Ernst Mach as a historian of science as seen in book reviews and notes (1896 — 1932). In Ernst Mach — A Deeper Look: Documents and New Perspectives [Bla92a], chapter 3, pages 57–70. ISBN 94-010-5230-1, 94-011-2771-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN C5.53; QC6.9.

URL https://link.springer.com/
chapter/10.1007/978-94-011-2771-
4_3.

Blay:1994:HSH

[Bla94] Michel Blay. History of science and history of mathematization: The example the science of motion at the turn of the 17th and 18th centuries. In Trends in the Historiography of Science [GCN94], chapter 30, pages 405–420. ISBN 90-481-4264-4, 94-017-3596-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-017-3596-4_30.

Becker:1995:GMT

[BLA95a] M. Becker, J. K. Ladha, and M. Ali. Green manure technology: Potential, usage, and limitations. A case study for lowland rice. In No Truth Except in the Details: Essays in Honor of Martin J. Klein [KS95], chapter 8, pages 181–194. ISBN 94-010-4097-4, 94-011-0217-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-011-0055-
7_8.

Blackmore:1995:BV

[Bla95b] John Blackmore. Back to Vienna. In Ludwig Boltzmann: His Later Life and Philosophy, 1900–1906: Book One: a Documentary History [Bla95g], pages 87–110. ISBN 0-7923-3231-8, 90-481-4492-2, 94-017-0489-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 168; QC16 .B64. URL https://link.

springer.com/chapter/10.1007/
978-94-017-0489-2_4.

Blackmore:1995:BC

[Bla95c] John Blackmore. Brentano and California. In Ludwig Boltzmann: His Later Life and Philosophy, 1900–1906: Book One: a Documentary History [Bla95g], pages 145–170. ISBN 0-7923-3231-8, 90-481-4492-2, 94-017-0489-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol.

168; QC16 .B64. URL https://link.

springer.com/chapter/10.1007/
978-94-017-0489-2_6.

Blackmore:1995:CL

[Bla95d] John Blackmore. Call to Leipzig. In Ludwig Boltzmann: His Later Life and Philosophy, 1900–1906: Book One: a Documentary History [Bla95g], pages 55–86. ISBN 0-7923-3231-8, 90-481-4492-2, 94-017-0489-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 168; QC16 .B64. URL https://link.

springer.com/chapter/10.1007/
978-94-017-0489-2_3.

Blackmore:1995:FMA

[Bla95e] John Blackmore. Final months & aftermath. In Ludwig Boltzmann: His Later Life and Philosophy, 1900–1906: Book One: a Documentary History [Bla95g], pages 205–214. ISBN 0-7923-3231-8, 90-481-4492-2, 94-017-0489-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 168; QC16 .B64. URL https://link.

springer.com/chapter/10.1007/
978-94-017-0489-2_8.

Blackmore:1995:I

[Bla95f] John Blackmore. Introduction. In Ludwig Boltzmann: His Later Life and Philosophy, 1900–1906: Book One: a Documentary History [Bla95g], pages 1–42. ISBN 0-7923-3231-8, 90-481-4492-2, 94-017-0489-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 168; QC16 .B64. URL https://link.

springer.com/chapter/10.1007/
978-94-017-0489-2_1.

Blackmore:1995:LBHa

[Bla95g]  John Blackmore, editor.  Ludwig

Boltzmann: His Later Life and Philosophy, 1900–1906: Book One: a Documentary History, volume 168 of Boston Studies in the Philosophy and History of Science. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 1995. ISBN 0-7923-3231-8, 90-481-4492-2, 94-017-0489-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). xvi + 266 pp. LCCN Q174 .B67 vol. 168; QC16 .B64. URL

https://link.springer.com/book/
10.1007/978-94-017-0489-2.

Blackmore:1995:PQ

[Bla95h] John Blackmore. Preliminary quotations. In Ludwig Boltzmann: His Later Life and Philosophy, 1900–1906: Book One: a Documentary History [Bla95g], pages 43–54. ISBN 0-7923-3231-8, 90-481-4492-2, 94-017-0489-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 168; QC16 .B64. URL https://link.

springer.com/chapter/10.1007/
978-94-017-0489-2_2.

Blackmore:1995:LBHb

[Bla95i] John T. Blackmore, editor. Ludwig Boltzmann: his later life and philosophy, 1900–1906, volume 174 of Boston Studies in the Philosophy of Science. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 1995. ISBN 0-7923-3231-8 (volume 1), 0-7923-3464-7 (volume 2). xv + 315 pp. LCCN Q174 .B67 vol. 168;

QC16 .B64. URL http://www.loc.
gov/catdir/enhancements/fy0823/
94039753-d.html;   http://www.
loc.gov/catdir/enhancements/
fy0823/94039753-t.html.

Blais:1995:LCS

[Bla95j] Michel J. Blais. … And chaos shall set you free …. In Québec Studies in the Philosophy of Science: Part I: Logic, Mathematics, Physics and History of Science Essays in Honor of Hugues Leblanc [MC95], chapter 14, pages 243–258. ISBN 94-010-7204-3, 94-009-1575-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-009-
1575-6_14.

Blay:2004:MSM

[Bla04] Michel Blay. Mathematization of the science of motion at the turn of the Seventeenth and Eighteenth Centuries: Pierre Varignon. In The Reception of the Galilean Science of Motion in Seventeenth-Century Europe [PT04], chapter 12, pages 243–259. ISBN 1-4020-2454-1, 1-4020-2455-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC133 .R43 2004.

URL https://link.springer.com/
chapter/10.1007/978-1-4020-2455-
9_12.

Blandin:2013:TE

[Bla13]  Patrick Blandin.  Towards Eco-

EvoEthics. In Bergandi [Ber13c], chapter 6, pages 83–100. ISBN 94-007-5066-8, 94-007-5067-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN GE42 .S77 2013; Q175 .B73 v. 296. URL https:/

/link.springer.com/chapter/10.
1007/978-94-007-5067-8_6.

Block:1976:FFF

[Blo76] N. J. Block. Fictionalism, functionalism and factor analysis. In Cohen et al. [CHMV76], pages 127–141. ISBN 90-277-0647-6, 90-277-0648-4 (paperback), 94-010-1449-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-010-1449-6_6.

Bloor:1986:SDC

[Blo86] David Bloor. Some determinants of cognitive style in science. In Cognition and Fact: Materials on Ludwik Fleck [CS86], chapter 19, pages 387–397. ISBN 94-010-8504-8, 94-009-4498-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-009-4498-
                                              5_19.

Bluh:1970:EMH

[Blü70] Otto Blüh. Ernst Mach — his life as a teacher and thinker. In Ernst Mach: Physicist and Philosopher [CS70], chapter 1, pages 1–22. ISBN 90-481-8318-9, 90-277-0016-8, 94-017-1462-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-017-1462-4_1.

Bluhm:2016:VEF

[Blu16] Robyn Bluhm. Values and evidence in feminist philosophy and in neuroscience. In Meta-Philosophical Reflection on Feminist Philosophies of Science [AV16a], chapter 6, pages 91–111. ISBN 3-319-26346-3 (hardcover), 3-319-26348-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN HQ1154 .M48 2016; HQ1190 M48 2016. URL https:/

                                              /link.springer.com/chapter/10.
                                              1007/978-3-319-26348-9_6.

Ben-Menachem:1988:LDC

[BM88] Yemima Ben-Menachem. Limits of a deductive construal of the function of scientific theories. In Science in Reflection: The Israel Colloquium: Studies in History, Philosophy, and Sociology of Science Volume 3 [UM88], chapter 2, pages 17–22. ISBN 90-277-2713-9, 94-009-2957-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-009-
2957-9_2.

Brown:1989:IRS

[BM89] James Robert Brown and Jürgen Mittelstrass. An Intimate Relation: Studies in the History and Philosophy of Science Presented to Robert E. Butts on his 60th Birthday, volume 116 of Boston Studies in the Philosophy of Science. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 1989. ISBN 94-010-7546-8, 94-009-2327-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). xx + 523 pp. LCCN ????

URL https://link.springer.com/
book/10.1007/978-94-009-2327-0.

Braun:2009:LE

[BM09] Erez Braun and Shimon Marom. Learning without error. In Going Amiss In Experimental Research [HSS09a], chapter 4, pages 49–54. ISBN 1-4020-8892-2 (hardcover), 1-4020-8893-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q172.5.E77 G647

2009. URL https://link.springer.
com/chapter/10.1007/978-1-4020-
8893-3_4.

Braddon-Mitchell:2017:TVS

[BMM17] David Braddon-Mitchell and Kristie Miller. On time and the varieties of science. In Time of Nature and the Nature of Time: Philosophical Perspectives of Time in Natural Sciences [BH17b], chapter 5, pages 67–85. ISBN 3-319-53723-7 (hardcover), 3-319-53725-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC173.59.S65 T56 2017. URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-3-319-53725-2_5.

Barberousse:2009:I

[BMP09a] Anouk Barberousse, Michel Morange, and Thomas Pradeu. Introduction. In Mapping the Future of Biology: Evolving Concepts and Theories [BMP09b], chapter 1, pages 1–14. ISBN 1-4020-9635-6 (hardcover), 1-4020-9636-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QH331

                                              .M295 2009. URL https://link.
                                              springer.com/chapter/10.1007/
                                              978-1-4020-9636-5_1.

Barberousse:2009:MFB

[BMP09b] Anouk Barberousse, Michel Morange, and Thomas Pradeu. Mapping the Future of Biology: Evolving Concepts and Theories, volume 266 of Boston Studies in the Philosophy of Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2009. ISBN 1-4020-9635-6 (hardcover), 1-4020-9636-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). xi + 173 pp. LCCN QH331 .M295 2009.

                                              URL https://link.springer.com/
                                              book/10.1007/978-1-4020-9636-5.

Ben-Noon:2017:ORC

[BN17] Chemi Ben-Noon. Obstructing the realization of civil responsibility: Political, sociological, historical, and psychological aspects of civil responsibility. In Encouraging Openness: Essays for Joseph Agassi on the Occasion of His 90th Birthday [BAG17], chapter 28, pages 333–348. ISBN 3-319-57668-2 (hardcover), 3-319-57669-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN B833.E53 2017.

URL https://link.springer.com/
chapter/10.1007/978-3-319-57669-
5_28.

Bocong:1996:TFM

[Boc96] Li Bocong. The treatise on fevers and miscellaneous diseases: Vicissitudes during the millennium after its completion. In Chinese Studies in the History and Philosophy of Science and Technology [DC96], chapter 31, pages 419–426. ISBN 90-481-4546-5, 94-015-8717-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 180; U390 .N19 1996.

URL https://link.springer.com/
chapter/10.1007/978-94-015-8717-
4_31.

Boccuni:2015:FGR

[Boc15] Francesca Boccuni. Frege’s Grundgesetze and a reassessment of predicativity. In From Logic to Practice: Italian Studies in the Philosophy of Mathematics [LPV15], chapter 4, pages 53–70. ISBN 3-319-10433-0, 3-319-10434-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QA8.4 .F76 2015; Q174 .B67 v.308.

URL https://link.springer.com/
chapter/10.1007/978-3-319-10434-
8_4.

Bodenmann:2018:ERL

[Bod18] Siegfried Bodenmann. Empiricism as a rhetoric of legitimation: Maupertuis and the shape of the Earth. In What Does it Mean to be an Empiricist?: Empiricisms in Eighteenth Century Sciences [BR18b], chapter 6, pages 87– 119. ISBN 3-319-69858-3 (hardcover), 3-319-69860-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN B816 .W43 2018. URL https:

//link.springer.com/chapter/10.
1007/978-3-319-69860-1_6.

Bogdan:2015:MWM

[Bog15a] Radu J. Bogdan. Memory as window on the mind. In Romanian Studies in Philosophy of Science [PST15], chapter 3, pages 45–53. ISBN 3-319-16654-9 (hardcover), 3-319-16655-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-3-319-16655-1_3.

Boggs:2015:CCU

[Bog15b] Andrew M. Boggs. Changing concepts of ‘The University’ and Oxford’s governance debates, 1850s–2000s. In Sciences in the Universities of Europe, Nineteenth and Twentieth Centuries: Academic Landscapes [SDG15b], chapter 4, pages 49–68. ISBN 94-017-9635-1, 94-017-9636-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q183.4.E85 2015. URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-017-9636-1_4.

Bohme:1983:AFC

[Böh83] Gernot Böhme. Autonomization and finalization: a comparison of fermentation research and fluid mechanics. In Finalization in Science: The Social Orientation of Scientific Progress [Sch83b], chapter 4, pages 53–91. ISBN 94-009-7082-X, 94-009-7080-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-009-7080-9_4.

Bohme:1987:GTC

[Böh87] Gernot Böhme. Is Goethe’s theory of color science? In Amrine et al. [AZW87], chapter 9, pages 147–173. ISBN 90-277-2265-X, 94-009-3761-X. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol.

97. US$79.00. URL https://link.

springer.com/chapter/10.1007/
978-94-009-3761-1_9.

Boland:1995:SVS

[Bol95a] Lawrence A. Boland. Style vs. substance in economic methodology. In Critical Rationalism, the Social Sciences and the Humanities: Essays for Joseph Agassi. Volume II [JL95b], chapter 7, pages 115–128. ISBN 0-7923-2961-9, 90-481-4430-2, 94-017-0441-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN B29

.C75 1995.  URL https://link.
springer.com/chapter/10.1007/
978-94-017-0441-0_7.

Boltzmann:1995:DA

[Bol95b] Ludwig Boltzmann. El Dorado & anecdotes. In Blackmore [Bla95g], pages 171–204. ISBN 0-7923-3231-8, 90-481-4492-2, 94-017-0489-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 168; QC16 .B64. URL https://link.

springer.com/chapter/10.1007/
978-94-017-0489-2_7.

Boland:2017:BPE

[Bol17] Lawrence A. Boland. On being a Popperian economist. In Encouraging Openness: Essays for Joseph Agassi on the Occasion of His 90th Birthday [BAG17], chapter 34, pages 429–437. ISBN 3-319-57668-2 (hardcover), 3-319-57669-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN B833.E53 2017. URL https://link.

springer.com/chapter/10.1007/
978-3-319-57669-5_34.

Bone:1987:OSD

[Bon87] J. Drummond Bone. Organicism and Shelley’s a defence of poetry. In Burwick [Bur87b], chapter 7, pages 195–210. ISBN 90-277-2541-1, 94-009-3917-5. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol.

105. US$64.00(U.S.). URL https:

//link.springer.com/chapter/10.
1007/978-94-009-3917-2_7.

Boniolo:2000:WDI

[Bon00] Giovanni Boniolo. What does it mean to observe physical reality? In The Reality of the Unobservable: Observability, Unobservability and Their Impact on the Issue of Scientific Realism [AP00b], chapter 13, pages 177–190. ISBN 90-481-5458-8, 94-015-9391-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-015-9391-5_13.

Bonk:2008:UEE

[Bon08] Thomas Bonk. Underdetermination: An Essay on Evidence and the Limits of Natural Knowledge, volume 261 of Boston Studies in the Philosophy of Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2008. ISBN 1-4020-6898-0, 1-4020-6899-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). xii + 284 pp. LCCN Q175 .B72235 2008.

URL https://link.springer.com/
book/10.1007/978-1-4020-6899-7.

Boon:2012:USP

[Boo12] Mieke Boon. Understanding scientific practices: The role of robustness notions. In Characterizing the Robustness of Science: After the Practice Turn in Philosophy of Science [STNW12], chapter 12, pages 289–315. ISBN 94-007-2758-5, 94-007-2759-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175 .C4595 2012.

URL https://link.springer.com/
chapter/10.1007/978-94-007-2759-
5_12.

Borgmann:1986:PRM

[Bor86] Albert Borgmann. Philosophical reflections on the microelectronic revolution. In Philosophy and Technology II: Information Technology and Computers in Theory and Practice [MH86], chapter 13, pages 189–203. ISBN 94-010-8510-2, 94-009-4512-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-009-4512-8_13.

Borger:1993:CLD

[Bör93] Egon Börger. Complexity of logical decision problems: An introduction. In Bridging the Gap: Philosophy, Mathematics, and Physics: Lectures on the Foundations of Science [CCG93], chapter 3, pages 71–86. ISBN 94-010-5101-1, 94-011-2496-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-011-2496-6_3.

Borsche:1999:ESD

[Bor99] Tilman Borsche. The epistemological shift from Descartes to Nietzsche: Intuition and imagination. In Nietzsche, Theories of Knowledge, and Critical Theory: Nietzsche and the Sciences I [Bab99b], chapter 4, pages 51–66. ISBN 90-481-5233-X, 94-017-2430-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-017-2430-
2_4.

Born:2007:MEL

[Bor07] Max Born. The Momentum–Energy Law in the electrodynamics of Gustav Mie. In Renn et al. [RSS^{+}07b], pages 745–756. ISBN 1-4020-3999-9 (hardcover), 1-4020-4000-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 no. 250 v. 4; QC173.6 .G469 2007.

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-1-4020-4000-

9_39. Originally published as “Der Impuls–Energie-Satz in der Elektrodynamik von Gustav Mie”, Nachrichten von der Königlichen Gesellschaft der Wissenschaften zu Göttingen 1, 23–36 (1914).

Bos:1974:LB

[Bos74] H. J. M. Bos. The lemniscate of Bernoulli. In For Dirk Struik: Scientific, Historical and Political Essays in Honor of Dirk J. Struik [CSW74], chapter 1, pages 3–14. ISBN 90-277-0379-5, 94-010-2115-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-010-2115-9_1.

Boscovich:1976:RTI

[Bos76] R. J. Boscovich. On the relativity of temporal intervals. In The Concepts of Space and Time: Their Structure and Their Development [Čap76], chapter 38, pages 225–226. ISBN 90-277-0375-2, 94-010-1727-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic).

                                              LCCN BD632. URL https://link.
                                              springer.com/chapter/10.1007/
                                              978-94-010-1727-5_38.

Bosnjak:1979:WRS

[Boš79] Branko Bošnjak. The withering away of religion in socialism. In Praxis:

Yugoslav Essays in the Philosophy and Methodology of the Social Sciences [MP79], chapter 13, pages 209–215. ISBN 90-277-0968-8, 94-009-9355-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-009-9355-
6_13.

Bottazzini:1994:GMS

[Bot94] Umberto Bottazzini. Geometry and “metaphysics of space” in Gauss and Riemann. In Poggi and Bossi [PB94], chapter 2, pages 15–29. ISBN 0-7923-2336-X, 94-017-2921-2. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 152. URL https:/

/link.springer.com/chapter/10.
1007/978-94-017-2921-5_2.

Boudri:2002:CFB

[Bou02a] J. Christiaan Boudri. The concept of force in the 1779 Berlin Essay Competition. In What Was Mechanical About Mechanics: the Concept of Force Between Metaphysics and Mechanics from Newton to Lagrange [Bou02i], chapter 6, pages 174–206. ISBN 1-4020-0233-5, 94-017-3672-3. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 224; QC73. URL https:

//link.springer.com/chapter/10.
1007/978-94-017-3672-5_6.

Boudri:2002:FLW

[Bou02b] J. Christiaan Boudri. Force like water.

In What Was Mechanical About Mechanics: the Concept of Force Between Metaphysics and Mechanics from Newton to Lagrange [Bou02i], chapter 2, pages 32–69. ISBN 1-4020-0233-5, 94-017-3672-3. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 224; QC73. URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-017-3672-5_2.

Boudri:2002:CP

[Bou02c] J. Christiaan Boudri. From cause to phenomenon. In What Was Mechanical About Mechanics: the Concept of Force Between Metaphysics and Mechanics from Newton to Lagrange [Bou02i], chapter 4, pages 104–133. ISBN 1-4020-0233-5, 94-017-3672-3. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 224; QC73. URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-017-3672-5_4.

Boudri:2002:EFC

[Bou02d] J. Christiaan Boudri. From efficient to final causes: The origin of the principle of least action. In What Was Mechanical About Mechanics: the Concept of Force Between Metaphysics and Mechanics from Newton to Lagrange [Bou02i], chapter 5, pages 134–171. ISBN 1-4020-0233-5, 94-017-3672-3. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 224;

                                              QC73. URL https://link.springer.
                                              com/chapter/10.1007/978-94-017-
                                              3672-5_5.

Boudri:2002:I

[Bou02e] J. Christiaan Boudri. Introduction. In What Was Mechanical About Mechanics: the Concept of Force Between Metaphysics and Mechanics from Newton to Lagrange [Bou02i], chapter 1, pages 1–29. ISBN 1-4020-0233-5, 94-017-3672-3. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 224; QC73. URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-017-3672-5_1.

Boudri:2002:LCF

[Bou02f]  J. Christiaan Boudri.  Lagrange’s

concept of force. In What Was Mechanical About Mechanics: the Concept of Force Between Metaphysics and Mechanics from Newton to Lagrange [Bou02i], chapter 7, pages 207–227. ISBN 1-4020-0233-5, 94-017-3672-3. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 224;

QC73. URL https://link.springer.
com/chapter/10.1007/978-94-017-
3672-5_7.

Boudri:2002:LFE

[Bou02g] J. Christiaan Boudri. Leibniz: Force as the essence of substance. In What Was Mechanical About Mechanics: the Concept of Force Between Metaphysics and Mechanics from Newton to Lagrange [Bou02i], chapter 3, pages 70–102. ISBN 1-4020-0233-5, 94-017-3672-3. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 224;

QC73. URL https://link.springer.
com/chapter/10.1007/978-94-017-
3672-5_3.

Boudri:2002:MC

[Bou02h] J. Christiaan Boudri. Metaphysics concealed. In What Was Mechanical About Mechanics: the Concept of Force Between Metaphysics and Mechanics from Newton to Lagrange [Bou02i], chapter 8, pages 229–239. ISBN 1-4020-0233-5, 94-017-3672-3. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 224; QC73. URL https:

//link.springer.com/chapter/10.
1007/978-94-017-3672-5_8.

Boudri:2002:WWM

[Bou02i] Johan Christiaan Boudri. What Was Mechanical About Mechanics: the Concept of Force Between Metaphysics and Mechanics from Newton to Lagrange, volume 224 of Boston Studies in the Philosophy of Science. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 2002. ISBN 1-4020-0233-5, 94-017-3672-3. ISSN 0068-0346 (print), 2214-7942 (electronic). xvi + 276 pp. LCCN Q174 .B67 vol. 224; QC73.

URL http://www.loc.gov/catdir/
enhancements/fy0814/2001050645-
b.html;   http://www.loc.gov/
catdir/enhancements/fy0814/
2001050645-t.html;   http://
www.loc.gov/catdir/enhancements/
fy0823/2001050645-d.html;
https://link.springer.com/book/
10.1007/978-94-017-3672-5.

Bouzov:2003:SRD

[Bou03] Vihren Bouzov. Scientific rationality, decision and choice. In Bulgarian Studies in the Philosophy of Science [Gin03a], chapter 2, pages 17–29. ISBN 1-4020-1496-1, 94-017-0961-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-017-0961-3_2.

Bouton:2017:FAP

[Bou17] Christophe Bouton. Is the future already present? The special theory of relativity and the block universe view. In Time of Nature and the Nature of Time: Philosophical Perspectives of Time in Natural Sciences [BH17b], chapter 6, pages 89–121. ISBN 3-319-53723-7 (hardcover), 3-319-53725-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC173.59.S65 T56 2017. URL https:

//link.springer.com/chapter/10.
1007/978-3-319-53725-2_6.

Bowman:1976:CST

[Bow76] Peter A. Bowman. The conventionality of slow-transport synchrony. In Cohen et al. [CHMV76], pages 423–434. ISBN 90-277-0647-6, 90-277-0648-4 (paperback), 94-010-1449-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-010-1449-6_21.

Boyer:1995:OC

[Boy95] Alain Boyer. The Osianderian compromise. In Misiek [Mis95c], pages 55–63. ISBN 94-010-4206-3, 94-011-0461-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-011-0461-
6_4.

Boyd:2001:RCM

[Boy01]  Richard N. Boyd.  Reference,

(in)commensurability and meanings. In Incommensurability and Related Matters [HHS01], chapter 1, pages 1–63. ISBN 90-481-5709-9, 94-015-9680-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-015-9680-
0_1.

Boyer:2009:ORM

[Boy09] Alain Boyer. Open rationality: Making guesses about nature, society and justice. In Rethinking Popper [PC09], chapter 19, pages 245–255. ISBN 1-4020-9337-3, 1-4020-9338-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN B1649.P64 R48 2009; Q174 .B67 2009. URL https:

//link.springer.com/chapter/10.
1007/978-1-4020-9338-8_19.

Bernard:2014:GI

[BP14a] Alain Bernard and Christine Proust. General introduction. In Scientific Sources and Teaching Contexts Throughout History: Problems and Perspectives [BP14b], chapter 1, pages 1–15. ISBN 94-007-5121-4 (hardcover), 94-007-5122-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q181 .S35 2014; Q181 S387

2014. URL https://link.springer.
com/chapter/10.1007/978-94-007-
5122-4_1.

Bernard:2014:SST

[BP14b] Alain Bernard and Christine Proust, editors. Scientific Sources and Teaching Contexts Throughout History: Problems and Perspectives, volume 301 of Boston Studies in the Philosophy and History of Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2014. ISBN 94-007-5121-4 (hardcover), 94-007-5122-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). ix + 334 pp. LCCN Q181 .S35 2014; Q181 S387 2014. URL

https://link.springer.com/book/
10.1007/978-94-007-5122-4.

Banchetti-Robino:2015:CEC

[BR15] Marina Paola Banchetti-Robino. From corpuscles to elements: Chemical ontologies from Van Helmont to Lavoisier. In Philosophy of Chemistry: Growth of a New Discipline [SM15b], chapter 10, pages 141–154. ISBN 94-017-9363-8 (hardcover), 94-017-9364-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QD6 .P455 2015; QD6 .S384 2014. URL https://link.

springer.com/chapter/10.1007/
978-94-017-9364-3_10.

Bodenmann:2018:I

[BR18a] Siegfried Bodenmann and Anne-Lise Rey. Introduction. In What Does it Mean to be an Empiricist?: Empiricisms in Eighteenth Century Sciences [BR18b], chapter 1, pages 1–11. ISBN 3-319-69858-3 (hardcover), 3-319-69860-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN B816 .W43 2018. URL https://link.

springer.com/chapter/10.1007/
978-3-319-69860-1_1.

Bodenmann:2018:WDI

[BR18b] Siegfried Bodenmann and Anne-Lise Rey. What Does it Mean to be an Empiricist?: Empiricisms in Eighteenth Century Sciences, volume 331 of Boston Studies in the Philosophy and History of Science. Springer-Ver-lag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2018. ISBN 3-319-69858-3 (hardcover), 3-319-69860-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). ix + 297 pp. LCCN B816 .W43 2018. URL

https://link.springer.com/book/
10.1007/978-3-319-69860-1.

Bradie:1974:SRC

[Bra74a] Michael Bradie. Is scientific realism a contingent thesis? In Schaffner and Cohen [SC74a], pages 367–373. ISBN 90-277-0408-2, 90-277-0409-0, 94-010-2140-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175 .B7312. URL https:/

/link.springer.com/chapter/10.
1007/978-94-010-2140-1_25.

Bray:1974:MUH

[Bra74b] J. Roger Bray. The modern use of historical Chinese solar observations. In For Dirk Struik: Scientific, Historical and Political Essays in Honor of Dirk J. Struik [CSW74], chapter 12, pages 143–146. ISBN 90-277-0379-5, 94-010-2115-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

                                              ???? URL https://link.springer.
                                              com/chapter/10.1007/978-94-010-
                                              2115-9_12.

Brady:1987:FCG

[Bra87] Ronald H. Brady. Form and cause in Goethe’s morphology. In Amrine et al. [AZW87], chapter 14, pages 257–300. ISBN 90-277-2265-X, 94-009-3761-X. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol.

97. US$79.00. URL https://link.

                                              springer.com/chapter/10.1007/
                                              978-94-009-3761-1_14.

Brain:1991:GVP

[Bra91] Robert Brain. The geographical vision and the popular order of disciplines, 1848–1870. In Woodward and Cohen [WC91], pages 367–376. ISBN 94-010-5395-2, 94-011-3164-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-011-3164-3_34.

Bravo:1995:GRU

[Bra95] Elia Nathan Bravo. Galileo’s revolution: The use of idealizational laws in physics. In Mexican Studies in the History and Philosophy of Science [RC95], chapter 8, pages 109–128. ISBN 94-010-6535-7 (hardcover), 94-009-0109-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

                                              ???? URL https://link.springer.
                                              com/chapter/10.1007/978-94-009-
                                              0109-4_8.

Braunss:1996:GPP

[Bra96] Günter Braunss. The Grassmann product in physics. In Schubring [Sch96f], pages 297–301. ISBN 90-481-4758-1, 94-015-8753-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QA8.9-10.3. URL https:/

/link.springer.com/chapter/10.
1007/978-94-015-8753-2_25.

Brandl:1997:TTA

[Bra97] Johannes L. Brandl. Thinking and talking about oneself. In Austrian Philosophy Past and Present: Essays in Honor of Rudolf Haller [LM97], chapter 12, pages 177–187. ISBN 94-010-6412-1, 94-011-5720-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-011-5720-9_12.

Brain:2007:REF

[Bra07] Robert Michael Brain. The romantic experiment as fragment. In Hans Christian Ørsted and The Romantic Legacy In Science: Ideas, Disciplines, Practices [BCK07], chapter 11, pages 217–233. ISBN 1-4020-2979-9 (hardcover), 1-4020-2987-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC515.O37 H36 2007.

URL https://link.springer.com/
chapter/10.1007/978-1-4020-2987-
5_11.

Brading:2011:SAP

[Bra11] Katherine Brading. Structuralist approaches to physics: Objects, models and modality. In Bokulich and Bokulich [BB11], chapter 3, pages 43–65. ISBN 90-481-9596-9, 90-481-9597-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175.32.S78 S35

2011. URL https://link.springer.
com/chapter/10.1007/978-90-481-
9597-8_3.

Bressan:1974:CSP

[Bre74a] Aldo Bressan. Comments on Suppes’ paper: The essential but implicit role of modal concepts in science. In Schaffner and Cohen [SC74a], pages 315–321. ISBN 90-277-0408-2, 90-277-0409-0, 94-010-2140-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175 .B7312. URL https:/

                                              /link.springer.com/chapter/10.
                                              1007/978-94-010-2140-1_20.  See

[Sup74].

Bressan:1974:UML

[Bre74b] Aldo Bressan. On the usefulness of modal logic in axiomatizations of physics. In Schaffner and Cohen [SC74a], pages 285–303. ISBN 90-277-0408-2, 90-277-0409-0, 94-010-2140-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175 .B7312.

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-010-2140-
                                              1_18.

Bressan:1974:RVF

[Bre74c] Aldo Bressan. Replies to van Fraassen’s comments: Bressan and Suppes on modality. In Schaffner and Cohen [SC74a], pages 331–334. ISBN 90-277-0408-2, 90-277-0409-0, 94-010-2140-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175 .B7312.

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-010-2140-
                                              1_22.

Bressan:1981:PP

[Bre81] Aldo Bressan. On physical possibility. In Italian Studies in the Philosophy of Science [Chi81b], chapter 13, pages 197–214. ISBN 90-277-1073-2, 94-009-8937-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-009-
8937-5_13.

Breit:1988:IEI

[Bre88] William Breit. Institutional economics as an ideological movement. In Philosophy, History and Social Action: Essays in Honor of Lewis Feuer with an autobiographic essay by Lewis Feuer [HOO88c], chapter 4, pages 119–136. ISBN 94-010-7793-2, 94-009-2873-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-009-2873-
2_4.

Breidbach:2007:CSE

[Bre07] Olaf Breidbach. The culture of science and experiments in Jena around 1800. In Hans Christian Ørsted and The Romantic Legacy In Science: Ideas, Disciplines, Practices [BCK07], chapter 10, pages 177–216. ISBN 1-4020-2979-9 (hardcover), 1-4020-2987-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC515.O37 H36 2007. URL https:

//link.springer.com/chapter/10.
1007/978-1-4020-2987-5_10.

Brenner:2009:PGP

[Bre09] Anastasios Brenner. A problem in general philosophy of science: The rational criteria of choice. In Brenner and Gayon [BG09], chapter 3, pages 73–89. ISBN 1-4020-9367-5, 1-4020-9368-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .F746 2009.

URL https://link.springer.com/
chapter/10.1007/978-1-4020-9368-
5_3.

Breard:2010:KPM

[Bré10] Andrea Bréard. Knowledge and practice of mathematics in Late Ming daily life encyclopedias. In Looking at it from Asia: the Processes that Shaped the Sources of History of Science [BE10b], chapter 10, pages 305–329. ISBN 90-481-3675-X (hardcover), 90-481-3676-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q127.A65 L66

                                              2010. URL https://link.springer.
                                              com/chapter/10.1007/978-90-481-
                                              3676-6_10.

Brentjes:2011:PNM

[Bre11] Sonja Brentjes. Patchwork — the norm of mapmaking practices for Western Asia in Catholic and Protestant Europe as well as in Istanbul between 1550 and 1750? In Science between Europe and Asia: Historical Studies on the Transmission, Adoption and Adaptation of Knowledge [GR11b], chapter 6, pages 77–101. ISBN 90-481-9968-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q125 .S354 2010; Q175.52.A78 S35 2010. URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-90-481-9968-6_6.

Breard:2014:TMK

[Bré14] Andrea Bréard. On the transmission of mathematical knowledge in versified form in China. In Bernard and Proust [BP14b], chapter 7, pages 155–185. ISBN 94-007-5121-4 (hardcover), 94-007-5122-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q181 .S35 2014; Q181 S387 2014.

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-007-5122-
                                              4_7.

Brigaglia:1996:IHG

[Bri96] Aldo Brigaglia. The influence of H.

Grassmann on Italian projective Ndimensional geometry. In Schubring [Sch96f], pages 155–163. ISBN 90-481-4758-1, 94-015-8753-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QA8.9-10.3.

URL https://link.springer.com/
chapter/10.1007/978-94-015-8753-
2_13.

Brinkmann:1998:NSH

[Bri98] Klaus Brinkmann. The natural and the supernatural in human nature: Hegel on the soul. In Philosophies of Nature: The Human Dimension: In Celebration of Erazim Kohák [CT98], chapter 1, pages 3–18. ISBN 90-481-4859-6, 94-017-2614-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic).

LCCN BD581. URL https://link.
springer.com/chapter/10.1007/
978-94-017-2614-6_1.

Brigandt:2015:DCE

[Bri15] Ingo Brigandt. From developmental constraint to evolvability: How concepts figure in explanation and disciplinary identity. In Conceptual Change in Biology: Scientific and Philosophical Perspectives on Evolution and Development [Lov15b], chapter 14, pages 305–325. ISBN 94-017-9411-1, 94-017-9412-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QH331 C66 2015; QH491; QH492 .C66

2014. URL https://link.springer.
com/chapter/10.1007/978-94-017-
9412-1_14.

Bristol:2018:PEE

[Bri18] Terry Bristol. The philosophy of engineering and the engineering worldview. In Mitcham et al. [MLNZ18], chapter 10, pages 103–118. ISBN 3-319-62448-2, 3-319-62450-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN TA15 .P45 2018.

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-3-319-62450-
                                              1_10.

Bronowski:1974:NCE

[Bro74] J. Bronowski. New concepts in the evolution of complexity. In Seeger and Cohen [SC74b], pages 133–151. ISBN 90-277-0390-6, 90-277-0376-0 (paperback), 94-010-2126-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 11.

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-010-2126-
                                              5_7.

Broughton:1986:ACN

[Bro86] Janet Broughton. Adequate causes and natural change in Descartes’ philosophy. In Cohen and Wartofsky [CW85], chapter 6, pages 107–127. ISBN 90-277-1971-3, 94-009-5345-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-009-5349-9_6.

Brown:1989:LP

[Bro89] James Robert Brown. Learning from the past. In An Intimate Relation: Studies in the History and Philosophy of Science Presented to Robert E. Butts on his 60th Birthday [BM89], chapter 17, pages 343–367. ISBN 94-010-7546-8, 94-009-2327-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-009-2327-0_17.

Brock:1991:ITS

[Bro91a] Adrian Brock. Imageless thought or stimulus error? The social construction of private experience. In Woodward and Cohen [WC91], pages 97–106. ISBN 94-010-5395-2, 94-011-3164-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-011-3164-3_9.

Brooks:1992:SML

[Bro91b] Charles L. Brooks III. Semiclassical methods for large molecules of biological importance. In The Invention of Physical Science: Intersections of Mathematics, Theology and Natural Philosophy Since the Seventeenth Century: Essays in Honor of Erwin N. Hiebert [NRS92], chapter 2, pages 51– 64. ISBN 94-010-5097-X, 94-011-2488-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-3-642-58183-
0_2.

Brooks:1992:IOE

[Bro92] Daniel R. Brooks. Incorporating origins into evolutionary theory. In Understanding Origins: Contemporary Views on the Origin of Life, Mind and Society [VD92], chapter 10, pages 191–212. ISBN 90-481-4090-0, 94-015-8054-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-015-8054-
0_10.

Brown:1993:CIH

[Bro93] Harvey R. Brown. Correspondence, invariance and heuristics in the emergence of special relativity. In Correspondence, Invariance and Heuristics: Essays in Honour of Heinz Post [FK93], chapter 12, pages 227–260. ISBN 90-481-4229-6, 94-017-1185-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-017-1185-2_12.

Bromberg:1995:EVV

[Bro95] Joan Lisa Bromberg. Experiment vis-a-vis theory in superconductivity research. The case of Bernd Matthias. In Gavroglu et al. [GSW95c], chapter 1, pages 1–10. ISBN 0-7923-2991-0 (set), 0-7923-2988-0, 94-017-2658-2. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 163.

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-017-2658-
                                              0_1.

Broncano:1996:PBJ

[Bro96a] Fernando Broncano. Prototype-based judgements and skepticism about rationality in naturalized epistemology. In Munévar [Mun96], chapter 13, pages 259–273. ISBN 94-010-6622-1, 94-009-0305-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

                                              ???? URL https://link.springer.
                                              com/chapter/10.1007/978-94-009-
                                              0305-0_13.

Brown:1996:P

[Bro96b] James Robert Brown. Phenomena. In Cohen et al. [CHR96], pages 117–129. ISBN 90-481-4493-0, 94-015-8638-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-015-8638-
                                              2_9.

Brown:2001:IR

[Bro01] Harold I. Brown. Incommensurability and reality. In Incommensurability and Related Matters [HHS01], chapter 4, pages 123–142. ISBN 90-481-5709-9, 94-015-9680-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-015-
9680-0_4.

Brotons:2008:MSE

[Bro08] Victor Navarro Brotons. Mechanics in Spain at the end of the 16th Century and the Madrid Academy of Mathematics. In Mechanics and Natural Philosophy Before the Scientific Revolution [LR08], chapter 10, pages 239–258. ISBN 1-4020-5966-3 (hardcover), 1-4020-5967-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QA802 .M43 2008.

URL https://link.springer.com/
chapter/10.1007/978-1-4020-5967-
4_10.

Brown:2010:WSO

[Bro10] David Brown. What shaped our corpuses of astral and mathematical cuneiform texts? In Looking at it from Asia: the Processes that Shaped the Sources of History of Science [BE10b], chapter 9, pages 277–303. ISBN 90-481-3675-X (hardcover), 90-481-3676-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q127.A65 L66

2010. URL https://link.springer.
com/chapter/10.1007/978-90-481-
3676-6_9.

Brown:2011:MMF

[Bro11] James Robert Brown. Medical market failures and their remedy. In Science in the Context of Application [CN11a], chapter 16, pages 271–281. ISBN 90-481-9050-9, 90-481-9051-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q180.55.M4 S385

2011. URL https://link.springer.
com/chapter/10.1007/978-90-481-
9051-5_16.

Buttner:2018:EHW

[BRS18] Jochen Büttner, Jürgen Renn, and Matthias Schemmel. The early history of weighing technology from the perspective of a theory of innovation. In Feldhay et al. [FRSV18a], chapter 4, pages 81–109. ISBN 3-319-90343-8 (hardcover), 3-319-90345-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QA802.

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-3-319-90345-
                                              3_4.

Bruno:1976:HBA

[Bru76] G. Bruno. Hesitations between absolute and relational theory of time. In The Concepts of Space and Time: Their Structure and Their Development [Čap76], chapter 33, pages 189–193. ISBN 90-277-0375-2, 94-010-1727-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN BD632.

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-010-1727-
                                              5_33.

Brush:1995:PTE

[Bru95] Stephen G. Brush. Prediction and theory evaluation in physics and astronomy. In No Truth Except in the Details: Essays in Honor of Martin J. Klein [KS95], chapter 13, pages 299–318. ISBN 94-010-4097-4, 94-011-0217-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

                                              ???? URL https://link.springer.
                                              com/chapter/10.1007/978-94-011-
                                              0217-9_13.

Brush:2007:SSS

[Bru07] Stephen G. Brush. Suggestions for the study of science. In Positioning the history of science [RG07], chapter 3, pages 13–25. ISBN 1-4020-5419-X (hardcover), 1-4020-5420-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q125 .P77 2007.

URL https://link.springer.com/
chapter/10.1007/1-4020-5420-3_3.

Baracca:2014:HPC

[BRW14a] Angelo Baracca, Jürgen Renn, and Helge Wendt. The History of Physics in Cuba, volume 304 of Boston Studies in the Philosophy and History of Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2014. ISBN 94-017-8040-4, 94-017-8041-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). xxvi + 446 + 24 pp. LCCN QC5.53; QC6.9; QC9 C9H57 2014; Q124.6-127.2. URL

https://link.springer.com/book/
10.1007/978-94-017-8041-4.

Baracca:2014:SIV

[BRW14b] Angelo Baracca, Jürgen Renn, and Helge Wendt. A short introduction to this volume. In The History of Physics in Cuba [BRW14a], chapter 1, pages 3–7. ISBN 94-017-8040-4, 94-017-8041-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC5.53; QC6.9; QC9 C9H57 2014; Q124.6-127.2. URL https://link.

springer.com/chapter/10.1007/
978-94-017-8041-4_1.

Bryant:1998:HHE

[Bry98] John H. Bryant. Heinrich Hertz’s experiments and experimental apparatus: His discovery of radio waves and his delineation of their properties. In Heinrich Hertz: Classical Physicist, Modern Philosopher [BHN98], chapter 4, pages 39–58. ISBN 90-481-4881-2, 94-015-8855-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.

                                              com/chapter/10.1007/978-94-015-
                                              8855-3_4.

Bohme:1983:TSS

[BS83] Gernot Böhme and Wolf Schäfer. Towards a social science of nature. In Finalization in Science: The Social Orientation of Scientific Progress [Sch83b], chapter 12, pages 251–269. ISBN 94-009-7082-X, 94-009-7080-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-009-7080-
                                              9_12.

Baublitz:1996:TBI

[BS96] Millard Baublitz and Abner Shimony. Tension in Bohm’s interpretation of quantum mechanics. In Bohmian Mechanics and Quantum Theory: an Appraisal [CFG96], chapter 17, pages 251–264. ISBN 90-481-4698-4, 94-015-8715-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC173.96-174.52. URL https://link.

                                              springer.com/chapter/10.1007/
                                              978-94-015-8715-0_17.

Belmar:2003:CCP

[BS03] Antonio García Belmar and José Ramón Bertomeu Sánchez. Constructing the centre from the periphery. In Simões et al. [SCD03b], chapter 7, pages 143–188. ISBN 1-4020-1259-4, 94-017-3584-0. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174

                                              .B67 v. 233. URL https://link.
                                              springer.com/chapter/10.1007/
                                              978-94-017-3584-1_7.

Bertomeu-Sanchez:2015:BBH

[BS15] José Ramón Bertomeu-Sánchez. Beyond borders in the history of science education. In Arabatzis et al. [ARS15], chapter 11, pages 159–173. ISBN 3-319-14552-5 (paperback), 3-319-14553-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q125 .R45

2015. URL https://link.springer.
com/chapter/10.1007/978-3-319-
14553-2_11.

Boccuni:2016:ORP

[BS16] Francesca Boccuni and Andrea Sereni.

Objectivity, Realism, and Proof: FilMat Studies in the Philosophy of Mathematics, volume 318 of Boston Studies in the Philosophy and History of Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2016. ISBN 3-319-31642-7, 3-319-31644-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). xl + 344 + 22 pp. LCCN QA9 .O25 2016.

URL https://link.springer.com/
book/10.1007/978-3-319-31644-4.

Baird:2006:I

[BSM06a] Davis Baird, Eric Scerri, and Lee McIntyre. Introduction. In Philosophy of Chemistry: Synthesis of a New Discipline [BSM06b], chapter 1, pages 3–18. ISBN 1-4020-2455-X, 1-4020-3256-0, 1-4020-3261-7. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 242 QD6. URL https:

//link.springer.com/chapter/10.
1007/1-4020-3261-7_1.

Baird:2006:PCS

[BSM06b] Davis Baird, Eric Scerri, and Lee McIntyre. Philosophy of Chemistry: Synthesis of a New Discipline, volume 242 of Boston Studies in the Philosophy of Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2006. ISBN 1-4020-2455-X, 1-4020-3256-0, 1-4020-3261-7. ISSN 0068-0346 (print), 2214-7942 (electronic). vii + 362 pp. LCCN Q174 .B67 vol. 242

                                              QD6. URL https://link.springer.
                                              com/book/10.1007/1-4020-3261-7.

Brini:1996:GPR

[BT96] Andrea Brini and Antonio G. B. Teolis. Grassmann progressive and regressive products and CG–Algebras. In Schubring [Sch96f], pages 231–242. ISBN 90-481-4758-1, 94-015-8753-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QA8.9-10.3.

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-015-8753-
                                              2_19.

Bu:1996:HCT

[Bu96] Chen Bu. Homeostasis and Chinese traditional medicine: Commenting on Cannon’s the wisdom of the body. In Chinese Studies in the History and Philosophy of Science and Technology [DC96], chapter 11, pages 115–124. ISBN 90-481-4546-5, 94-015-8717-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 180; U390 .N19 1996.

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-015-8717-
                                              4_11.

Bubner:1984:HSS

[Bub84] Rüdiger Bubner. On Hegel’s significance for the social sciences. In Hegel and the Sciences [CW84a], chapter 10, pages 143–159. ISBN 94-009-6235-5, 94-009-6233-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

                                              ???? URL https://link.springer.
                                              com/chapter/10.1007/978-94-009-
                                              6233-0_10.

Bub:1996:MIB

[Bub96] Jeffrey Bub. Modal interpretations and Bohmian mechanics. In Bohmian Mechanics and Quantum Theory: an Appraisal [CFG96], chapter 23, pages 331–341. ISBN 90-481-4698-4, 94-015-8715-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC173.96-174.52. URL https://link.

springer.com/chapter/10.1007/
978-94-015-8715-0_23.

Buchdahl:1967:SSC

[Buc67] Gerd Buchdahl. Semantic sources of the concept of law. In Cohen and Wartofsky [CW67], pages 272–292. ISBN 90-277-0013-3, 94-010-3508-3. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175 .B731 v.3.

URL https://link.springer.com/
chapter/10.1007/978-94-010-3508-
8_15.

Buchdahl:1984:CAS

[Buc84] Gerd Buchdahl. Conceptual analysis and scientific theory in Hegel’s philosophy of nature (with special reference to Hegel’s optics). In Hegel and the Sciences [CW84a], chapter 2, pages 13– 36. ISBN 94-009-6235-5, 94-009-6233-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-009-6233-
0_2.

Buchdahl:1989:RRP

[Buc89] Gerd Buchdahl. Reductive realism and the problem of affection in Kant. In An Intimate Relation: Studies in the History and Philosophy of Science Presented to Robert E. Butts on his 60th Birthday [BM89], chapter 11, pages 221–241. ISBN 94-010-7546-8, 94-009-2327-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-009-
2327-0_11.

Buchwald:1990:BHH

[Buc90] Jed Z. Buchwald. The background to Heinrich Hertz’s experiments in electrodynamics. In Levere and Shea [LS90], chapter 12, pages 275–306. ISBN 0-7923-0420-9, 94-009-1878-X. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 120.

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-009-1878-
                                              8_12.

Buchwald:1992:TGR

[Buc91] Jed Z. Buchwald. The training of German research physicist Heinrich Hertz. In The Invention of Physical Science: Intersections of Mathematics, Theology and Natural Philosophy Since the Seventeenth Century: Essays in Honor of Erwin N. Hiebert [NRS92], chapter 6, pages 119–145. ISBN 94-010-5097-X, 94-011-2488-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-011-2488-1_6.

Buchwald:1994:KC

[Buc94a] Jed Z. Buchwald. Kinds and (in)commensurability. In Trends in the Historiography of Science [GCN94], chapter 4, pages 49–63. ISBN 90-481-4264-4, 94-017-3596-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-017-3596-4_4.

Buchwalter:1994:HDE

[Buc94b] Andrew Buchwalter. Hegel and the doctrine of expressivism. In Artifacts, Representations and Social Practice: Essays for Marx Wartofsky [GC94], chapter 10, pages 163–183. ISBN 94-010-4390-6, 94-011-0902-8 (e-book).

ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN B73 .A77 1994.

URL https://link.springer.com/
chapter/10.1007/978-94-011-0902-
4_10.

Buchwald:1995:HHA

[Buc95] Jed Z. Buchwald. Heinrich Hertz’s attempt to generate a novel account of evaporation. In No Truth Except in the Details: Essays in Honor of Martin J. Klein [KS95], chapter 7, pages 171–189. ISBN 94-010-4097-4, 94-011-0217-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-011-0217-
9_7.

Buchwald:1998:RHH

[Buc98] Jed Z. Buchwald. Reflections on Hertz and the Hertzian dipole. In Heinrich Hertz: Classical Physicist, Modern Philosopher [BHN98], chapter 15, pages 269–280. ISBN 90-481-4881-2, 94-015-8855-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic).

LCCN ????  URL https://link.
springer.com/chapter/10.1007/
978-94-015-8855-3_15.

Buck:2001:HRF

[Buc01] Peter Buck. A healthy regard for the facts. In Allen and MacLeod [AM01b], chapter 9, pages 137–151. ISBN 1-4020-0495-8 (hardcover), 94-017-2956-5. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 228; Q175.55. URL https:

//link.springer.com/chapter/10.
1007/978-94-017-2956-7_9.

Budinich:1993:APS

[Bud93] Paolo Budinich. Axioms and paradoxes in special relativity. In Bridging the Gap: Philosophy, Mathematics, and Physics: Lectures on the Foundations of Science [CCG93], chapter 8, pages 149–173. ISBN 94-010-5101-1, 94-011-2496-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-011-
2496-6_8.

Bueno:2011:SEA

[Bue11] Otávio Bueno. Structural empiricism, again. In Bokulich and Bokulich [BB11], chapter 5, pages 81–103. ISBN 90-481-9596-9, 90-481-9597-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175.32.S78 S35

2011. URL https://link.springer.
com/chapter/10.1007/978-90-481-
9597-8_5.

Bunge:1992:MCN

[Bun92] Mario Bunge. Mach’s critique of Newtonian mechanics. In Ernst Mach — A Deeper Look: Documents and New Perspectives [Bla92a], chapter 13, pages 243–261. ISBN 94-010-5230-1, 94-011-2771-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN C5.53; QC6.9. URL https:

//link.springer.com/chapter/10.
1007/978-94-011-2771-4_13.

Bunzl:1993:CCC

[Bun93a] Martin Bunzl. Causal conception and causal connection. In The Context of Explanation [Bun93c], chapter 2, pages 11–29. ISBN 94-010-4760-X, 94-011-1735-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-011-
1735-7_2.

Bunzl:1993:CF

[Bun93b] Martin Bunzl. Causal factuals. In The Context of Explanation [Bun93c], chapter 9, pages 111–130. ISBN 94-010-4760-X, 94-011-1735-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-011-1735-7_9.

Bunzl:1993:CE

[Bun93c] Martin Bunzl. The Context of Explanation, volume 149 of Boston Studies in the Philosophy of Science. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 1993. ISBN 94-010-4760-X, 94-011-1735-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). xii + 157 pp. LCCN ???? URL

https://link.springer.com/book/
10.1007/978-94-011-1735-7.

Bunzl:1993:DNM

[Bun93d] Martin Bunzl. The deductivenomological model briefly revisited. In The Context of Explanation [Bun93c], chapter 1, pages 1–10. ISBN 94-010-4760-X, 94-011-1735-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-011-1735-7_1.

Bunzl:1993:IU

[Bun93e] Martin Bunzl. Identity and unification. In The Context of Explanation [Bun93c], chapter 3, pages 30–41. ISBN 94-010-4760-X, 94-011-1735-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-011-1735-7_3.

Bunzl:1993:I

[Bun93f] Martin Bunzl. Interpretation. In The Context of Explanation [Bun93c], chapter 8, pages 98–110. ISBN 94-010-4760-X, 94-011-1735-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-011-1735-7_8.

Bunzl:1993:L

[Bun93g] Martin Bunzl. Laws. In The Context of Explanation [Bun93c], chapter 7, pages 85–97. ISBN 94-010-4760-X, 94-011-1735-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

                                              ???? URL https://link.springer.
                                              com/chapter/10.1007/978-94-011-
                                              1735-7_7.

Bunzl:1993:P

[Bun93h] Martin Bunzl. Pragmatics. In The Context of Explanation [Bun93c], chapter 4, pages 42–58. ISBN 94-010-4760-X, 94-011-1735-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

                                              ???? URL https://link.springer.
                                              com/chapter/10.1007/978-94-011-
                                              1735-7_4.

Bunzl:1993:SD

[Bun93i] Martin Bunzl. Statics and dynamics. In The Context of Explanation [Bun93c], chapter 5, pages 59–72. ISBN 94-010-4760-X, 94-011-1735-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-011-1735-7_5.

Bunzl:1993:SR

[Bun93j] Martin Bunzl. Supervenience and reductionism. In The Context of Explanation [Bun93c], chapter 6, pages 73–84. ISBN 94-010-4760-X, 94-011-1735-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-011-
1735-7_6.

Bunge:1995:HVS

[Bun95a] Mario Bunge. Hidden variables, separability, and realism. In Québec Studies in the Philosophy of Science: Part I: Logic, Mathematics, Physics and History of Science Essays in Honor of Hugues Leblanc [MC95], chapter 12, pages 217–227. ISBN 94-010-7204-3, 94-009-1575-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-009-
1575-6_12.

Bunge:1995:PRC

[Bun95b] Mario Bunge. The poverty of rational choice theory. In Critical Rationalism, Metaphysics and Science: Essays for Joseph Agassi. Volume I [JL95a], chapter 10, pages 149–168. ISBN 94-010-4211-X, 94-011-0471-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-011-0471-5_10.

Bunge:2010:AO

[Bun10a] Mario Bunge. Appendix a: Objects.

In Matter and Mind: a Philosophical Inquiry [Bun10j], chapter 14, pages 267–275. ISBN 90-481-9224-2 (hardcover), 90-481-9225-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN BD418.3 .B7664 2010; Q175 .B73 v.287. URL https:

//link.springer.com/chapter/10.
1007/978-90-481-9225-0_14.

Bunge:2010:ABT

[Bun10b] Mario Bunge. Appendix b: Truths. In Matter and Mind: a Philosophical Inquiry [Bun10j], chapter 15, pages 277–285. ISBN 90-481-9224-2 (hardcover), 90-481-9225-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN BD418.3 .B7664 2010; Q175 .B73 v.287. URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-90-481-9225-0_15.

Bunge:2010:BCH

[Bun10c] Mario Bunge. Brain and computer: The hardware/software dualism. In Matter and Mind: a Philosophical Inquiry [Bun10j], chapter 12, pages 227–237. ISBN 90-481-9224-2 (hardcover), 90-481-9225-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN BD418.3 .B7664 2010; Q175 .B73 v.287.

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-90-481-9225-
                                              0_12.

Bunge:2010:CMB

[Bun10d] Mario Bunge. Classical matter: Bodies and fields. In Matter and Mind: a Philosophical Inquiry [Bun10j], chapter 2, pages 23–40. ISBN 90-481-9224-2 (hardcover), 90-481-9225-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN BD418.3 .B7664 2010; Q175 .B73 v.287. URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-90-481-9225-0_2.

Bunge:2010:CCF

[Bun10e] Mario Bunge. Cognition, consciousness, and free will. In Matter and Mind: a Philosophical Inquiry [Bun10j], chapter 11, pages 201–225. ISBN 90-481-9224-2 (hardcover), 90-481-9225-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN BD418.3 .B7664 2010; Q175 .B73 v.287.

URL https://link.springer.com/
chapter/10.1007/978-90-481-9225-
0_11.

Bunge:2010:EL

[Bun10f] Mario Bunge. Emergence and levels. In Matter and Mind: a Philosophical Inquiry [Bun10j], chapter 5, pages 73–92. ISBN 90-481-9224-2 (hardcover), 90-481-9225-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN BD418.3 .B7664 2010; Q175 .B73 v.287.

URL https://link.springer.com/
chapter/10.1007/978-90-481-9225-
0_5.

Bunge:2010:GCM

[Bun10g] Mario Bunge. General concept of matter: To be is to become. In Matter and Mind: a Philosophical Inquiry [Bun10j], chapter 4, pages 61–72. ISBN 90-481-9224-2 (hardcover), 90-481-9225-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN BD418.3 .B7664 2010; Q175 .B73 v.287.

URL https://link.springer.com/
chapter/10.1007/978-90-481-9225-
0_4.

Bunge:2010:KGB

[Bun10h] Mario Bunge. Knowledge: Genuine and bogus. In Matter and Mind: a Philosophical Inquiry [Bun10j], chapter 13, pages 239–264. ISBN 90-481-9224-2 (hardcover), 90-481-9225-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN BD418.3 .B7664 2010; Q175 .B73 v.287. URL https:

//link.springer.com/chapter/10.
1007/978-90-481-9225-0_13.

Bunge:2010:M

[Bun10i] Mario Bunge. Materialism. In Matter and Mind: a Philosophical Inquiry [Bun10j], chapter 7, pages 121–140. ISBN 90-481-9224-2 (hardcover), 90-481-9225-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN BD418.3 .B7664 2010; Q175 .B73 v.287.

URL https://link.springer.com/
chapter/10.1007/978-90-481-9225-
0_7.

Bunge:2010:MMPb

[Bun10j] Mario Bunge. Matter and Mind: a Philosophical Inquiry, volume 287 of Boston Studies in the Philosophy of Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2010. ISBN 90-481-9224-2 (hardcover), 90-481-9225-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). xviii + 319 pp. LCCN BD418.3 .B7664 2010; Q175 .B73 v.287.

URL https://link.springer.com/
book/10.1007/978-90-481-9225-0.

Bunge:2010:MBP

[Bun10k] Mario Bunge. The mind–body problem. In Matter and Mind: a Philosophical Inquiry [Bun10j], chapter 8, pages 143–157. ISBN 90-481-9224-2 (hardcover), 90-481-9225-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN BD418.3 .B7664 2010; Q175 .B73 v.287. URL https://link.

springer.com/chapter/10.1007/
978-90-481-9225-0_8.

Bunge:2010:MS

[Bun10l] Mario Bunge. Mind and society. In Matter and Mind: a Philosophical Inquiry [Bun10j], chapter 10, pages 181–199. ISBN 90-481-9224-2 (hardcover), 90-481-9225-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN BD418.3 .B7664 2010; Q175 .B73 v.287. URL https://link.

springer.com/chapter/10.1007/
978-90-481-9225-0_10.

Bunge:2010:MMPa

[Bun10m] Mario Bunge. Minding matter: The plastic brain. In Matter and Mind: a Philosophical Inquiry [Bun10j], chapter 9, pages 159–180. ISBN 90-481-9224-2 (hardcover), 90-481-9225-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN BD418.3 .B7664 2010; Q175 .B73 v.287.

URL https://link.springer.com/
chapter/10.1007/978-90-481-9225-
0_9.

Bunge:2010:N

[Bun10n] Mario Bunge. Naturalism. In Matter and Mind: a Philosophical Inquiry [Bun10j], chapter 6, pages 93–119. ISBN 90-481-9224-2 (hardcover), 90-481-9225-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN BD418.3 .B7664 2010; Q175 .B73 v.287.

URL https://link.springer.com/
chapter/10.1007/978-90-481-9225-
0_6.

Bunge:2010:PW

[Bun10o] Mario Bunge. Philosophy as worldview.

In Matter and Mind: a Philosophical Inquiry [Bun10j], chapter 1, pages 3–22. ISBN 90-481-9224-2 (hardcover), 90-481-9225-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN BD418.3 .B7664 2010; Q175 .B73 v.287.

URL https://link.springer.com/
chapter/10.1007/978-90-481-9225-
0_1.

Bunge:2010:QMW

[Bun10p] Mario Bunge. Quantum matter: Weird but real. In Matter and Mind: a Philosophical Inquiry [Bun10j], chapter 3, pages 41–59. ISBN 90-481-9224-2 (hardcover), 90-481-9225-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN BD418.3 .B7664 2010; Q175 .B73 v.287. URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-90-481-9225-0_3.

Bunge:2012:BTD

[Bun12a] Mario Bunge. Bridging theories to data. In Evaluating Philosophies [Bun12h], chapter 13, pages 119–130. ISBN 94-007-4407-2, 94-007-4408-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 v. 295; Q175 .B8223 2012. URL https:/

                                              /link.springer.com/chapter/10.
                                              1007/978-94-007-4408-0_13.

Bunge:2012:CFP

[Bun12b] Mario Bunge. Can functionalist psychology explain anything? In Evaluating Philosophies [Bun12h], chapter 17, pages 159–165. ISBN 94-007-4407-2, 94-007-4408-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 v. 295; Q175 .B8223 2012. URL https:/

                                              /link.springer.com/chapter/10.
                                              1007/978-94-007-4408-0_17.

Bunge:2012:CSE

[Bun12c] Mario Bunge. Can standard economic theory account for crises? In Evaluating Philosophies [Bun12h], chapter 8, pages 77–82. ISBN 94-007-4407-2, 94-007-4408-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 v. 295; Q175 .B8223 2012.

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-007-4408-
                                              0_8.

Bunge:2012:CL

[Bun12d] Mario Bunge. Climate and logic. In Evaluating Philosophies [Bun12h], chapter 5, pages 57–59. ISBN 94-007-4407-2, 94-007-4408-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 v. 295; Q175 .B8223 2012. URL https:/

/link.springer.com/chapter/10.
1007/978-94-007-4408-0_5.

Bunge:2012:CVC

[Bun12e] Mario Bunge. Conclusion: Valuation criterion. In Evaluating Philosophies [Bun12h], chapter 20, pages 181–183. ISBN 94-007-4407-2, 94-007-4408-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 v. 295; Q175 .B8223 2012.

URL https://link.springer.com/
chapter/10.1007/978-94-007-4408-
0_20.

Bunge:2012:DIL

[Bun12f] Mario Bunge. Does inductive logic work? In Evaluating Philosophies [Bun12h], chapter 12, pages 115–118. ISBN 94-007-4407-2, 94-007-4408-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 v. 295; Q175 .B8223 2012.

URL https://link.springer.com/
chapter/10.1007/978-94-007-4408-
0_12.

Bunge:2012:DQP

[Bun12g] Mario Bunge. Does quantum physics refute realism, materialism and determinism? In Evaluating Philosophies [Bun12h], chapter 15, pages 139–149. ISBN 94-007-4407-2, 94-007-4408-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 v. 295; Q175 .B8223 2012.

URL https://link.springer.com/
chapter/10.1007/978-94-007-4408-
0_15.

Bunge:2012:EP

[Bun12h] Mario Bunge. Evaluating Philosophies, volume 295 of Boston Studies in the Philosophy and History of Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2012. ISBN 94-007-4407-2, 94-007-4408-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). xiv + 200 pp. LCCN Q174 .B67 v. 295; Q175

.B8223 2012. URL http://www.loc.
gov/catdir/enhancements/fy1506/
2012941366-d.html;   http://www.
loc.gov/catdir/enhancements/
fy1506/2012941366-t.html;
https://link.springer.com/book/
10.1007/978-94-007-4408-0.

Bunge:2012:ESD

[Bun12i] Mario Bunge. Existence: Single or double? In Evaluating Philosophies [Bun12h], chapter 19, pages 173–179. ISBN 94-007-4407-2, 94-007-4408-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 v. 295; Q175 .B8223 2012.

URL https://link.springer.com/
chapter/10.1007/978-94-007-4408-
0_19.

Bunge:2012:IOM

[Bun12j] Mario Bunge. Informatics: One or multiple? In Evaluating Philosophies [Bun12h], chapter 6, pages 61–64. ISBN 94-007-4407-2, 94-007-4408-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 v. 295; Q175 .B8223 2012. URL https:/

/link.springer.com/chapter/10.
1007/978-94-007-4408-0_6.

Bunge:2012:KPR

[Bun12k] Mario Bunge. Knowledge pyramids or rosettes? In Evaluating Philosophies [Bun12h], chapter 18, pages 167–171. ISBN 94-007-4407-2, 94-007-4408-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 v. 295; Q175 .B8223 2012.

URL https://link.springer.com/
chapter/10.1007/978-94-007-4408-
0_18.

Bunge:2012:MPP

[Bun12l] Mario Bunge. Marxist philosophy:

Promise and reality. In Evaluating Philosophies [Bun12h], chapter 9, pages 83–93. ISBN 94-007-4407-2, 94-007-4408-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 v. 295; Q175 .B8223 2012.

URL https://link.springer.com/
chapter/10.1007/978-94-007-4408-
0_9.

Bunge:2012:MEP

[Bun12m] Mario Bunge. Matter and energy:

Physical or metaphysical concepts? In Evaluating Philosophies [Bun12h], chapter 14, pages 131–138. ISBN 94-007-4407-2, 94-007-4408-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 v. 295; Q175 .B8223 2012. URL https:/

/link.springer.com/chapter/10.
1007/978-94-007-4408-0_14.

Bunge:2012:PUD

[Bun12n] Mario Bunge. Parallel universes?

Digital physics? In Evaluating Philosophies [Bun12h], chapter 16, pages 151–157. ISBN 94-007-4407-2, 94-007-4408-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 v. 295; Q175 .B8223 2012.

URL https://link.springer.com/
chapter/10.1007/978-94-007-4408-
0_16.

Bunge:2012:PMS

[Bun12o] Mario Bunge. The philosophical matrix of scientific progress. In Evaluating Philosophies [Bun12h], chapter 2, pages 15–34. ISBN 94-007-4407-2, 94-007-4408-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 v. 295; Q175 .B8223 2012.

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-007-4408-
                                              0_2.

Bunge:2012:PP

[Bun12p] Mario Bunge. Philosophies and phobosophies. In Evaluating Philosophies [Bun12h], chapter 1, pages 3–13. ISBN 94-007-4407-2, 94-007-4408-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 v. 295; Q175 .B8223 2012. URL https:/

                                              /link.springer.com/chapter/10.
                                              1007/978-94-007-4408-0_1.

Bunge:2012:RLJ

[Bun12q] Mario Bunge. Rules of law: Just and unjust. In Evaluating Philosophies [Bun12h], chapter 10, pages 95–102. ISBN 94-007-4407-2, 94-007-4408-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 v. 295; Q175 .B8223 2012.

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-007-4408-
                                              0_10.

Bunge:2012:SPA

[Bun12r] Mario Bunge. Subjective probabilities: Admissible in science? In Evaluating Philosophies [Bun12h], chapter 11, pages 105–113. ISBN 94-007-4407-2, 94-007-4408-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 v. 295; Q175 .B8223

                                              2012. URL https://link.springer.
                                              com/chapter/10.1007/978-94-007-
                                              4408-0_11.

Bunge:2012:SM

[Bun12s] Mario Bunge. Systemics and materialism. In Evaluating Philosophies [Bun12h], chapter 3, pages 35–47. ISBN 94-007-4407-2, 94-007-4408-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 v. 295; Q175 .B8223 2012. URL https:/

/link.springer.com/chapter/10.
1007/978-94-007-4408-0_3.

Bunge:2012:T

[Bun12t] Mario Bunge. Technoscience? In Evaluating Philosophies [Bun12h], chapter 4, pages 51–56. ISBN 94-007-4407-2, 94-007-4408-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 v. 295; Q175 .B8223 2012.

URL https://link.springer.com/
chapter/10.1007/978-94-007-4408-
0_4.

Bunge:2012:WWB

[Bun12u] Mario Bunge. Wealth and well-being, economic growth and integral development. In Evaluating Philosophies [Bun12h], chapter 7, pages 65–76. ISBN 94-007-4407-2, 94-007-4408-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 v. 295; Q175 .B8223 2012. URL https:/

/link.springer.com/chapter/10.
1007/978-94-007-4408-0_7.

Bunge:2017:WDS

[Bun17] Mario Bunge. Why don’t scientists respect philosophers? In Encouraging Openness: Essays for Joseph Agassi on the Occasion of His 90th Birthday [BAG17], chapter 1, pages 3–12. ISBN 3-319-57668-2 (hardcover), 3-319-57669-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN B833.E53 2017. URL https://link.

springer.com/chapter/10.1007/
978-3-319-57669-5_1.

Burgers:1974:SSF

[Bur74a] J. M. Burgers. Summation of series of fractions depending upon the roots of the Airy function. In For Dirk Struik: Scientific, Historical and Political Essays in Honor of Dirk J. Struik [CSW74], chapter 2, pages 15–19. ISBN 90-277-0379-5, 94-010-2115-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-010-2115-9_2.

Buroker:1974:KDT

[Bur74b] Jill Vance Buroker. Kant, the dynamical tradition, and the role of matter in explanation. In Schaffner and Cohen [SC74a], pages 153–164. ISBN 90-277-0408-2, 90-277-0409-0, 94-010-2140-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175 .B7312.

URL https://link.springer.com/
chapter/10.1007/978-94-010-2140-
1_11.

Burian:1980:WPS

[Bur80] Richard M. Burian. Why philosophers should not despair of understanding scientific discovery. In Scientific Discovery, Logic, and Rationality [Nic80f], chapter 14, pages 317–336. ISBN 90-277-1070-8, 94-009-8986-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-009-8986-3_14.

Burian:1983:TRT

[Bur83] Richard M. Burian. Towards a richer theory of dialogue: Comments on Professor Rivetti Barbò’s paper. In Language, Logic and Method [CW83b], chapter 19, pages 445–457. ISBN 94-009-7704-2, 94-009-7702-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-009-7702-0_19.  See

[Bar83].

Burian:1984:SRI

[Bur84] Richard M. Burian. Scientific realism and incommensurability: Some criticisms of Kuhn and Feyerabend. In Methodology, Metaphysics and the History of Science: In Memory of Benjamin Nelson [CW84b], chapter 1, pages 1–31. ISBN 94-009-6333-5, 94-009-6331-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-009-
6331-3_1.

Burian:1986:IPB

[Bur86] Richard M. Burian. The “internal politics” of biology and the justification of biological theories. In Cohen and Wartofsky [CW85], chapter 2, pages 23–45. ISBN 90-277-1971-3, 94-009-5345-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-009-
5349-9_2.

Burian:1987:HTA

[Bur87a] Richard M. Burian. How not to talk about conceptual change in science. In Pitt and Pera [PP87], chapter 1, pages 3–33. ISBN 90-277-2417-2, 94-009-3779-2. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 98. US$64.00. URL https:

//link.springer.com/chapter/10.
1007/978-94-009-3779-6_1.

Burwick:1987:AOF

[Bur87b] Frederick Burwick, editor. Approaches to organic form: permutations in science and culture, volume 105 of Boston Studies in the Philosophy of Science. D. Reidel, Dordrecht, The Netherlands; Boston, MA, USA; Lancaster, UK; Tokyo, Japan, 1987.

ISBN 90-277-2541-1, 94-009-3917-5. ISSN 0068-0346 (print), 2214-7942 (electronic). xix + 286 pp. LCCN Q174 .B67 vol. 105. US$64.00(U.S.). URL

https://link.springer.com/book/
10.1007/978-94-009-3917-2.

Burwick:1987:GEF

[Bur87c] Frederick Burwick. Goethe’s Entoptische Farben and the problem of polarity. In Amrine et al. [AZW87], chapter 3, pages 29–44. ISBN 90-277-2265-X, 94-009-3761-X. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 97. US$79.00. URL https:

//link.springer.com/chapter/10.
1007/978-94-009-3761-1_3.

Burwick:1987:KHO

[Bur87d] Frederick Burwick. Kant and Hegel: Organicism and language theory. In Approaches to organic form: permutations in science and culture [Bur87b], chapter 6, pages 153–193. ISBN 90-277-2541-1, 94-009-3917-5. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 105. US$64.00(U.S.).

URL https://link.springer.com/
chapter/10.1007/978-94-009-3917-
2_6.

Busard:1974:SPC

[Bus74] H. L. L. Busard. The second part of chapter 5 of the De Arte Mensurandi by Johannes De Muris. In For Dirk Struik: Scientific, Historical and Political Essays in Honor of Dirk J. Struik [CSW74], chapter 13, pages 147–167. ISBN 90-277-0379-5, 94-010-2115-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-010-2115-
9_13.

Busch:1997:QSO

[Bus97a] P. Busch. Is the quantum state (an)

observable? In Cohen et al. [CHS97b], pages 61–70. ISBN 90-481-4809-X, 94-017-2732-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC173.96-174.52. URL https:

//link.springer.com/chapter/10.
1007/978-94-017-2732-7_5.

Buszkowski:1997:RBS

[Bus97b] Wojciech Buszkowski. Relations between sets of time points and quasilinear orderings. In Perspectives on Time [FSU97b], chapter 13, pages 301–321. ISBN 90-481-4774-3, 94-015-8875-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol.189. URL https:/

/link.springer.com/chapter/10.
1007/978-94-015-8875-1_13.

Bushev:2003:HAP

[Bus03] Michael Bushev. On human agency in physics. In Bulgarian Studies in the Philosophy of Science [Gin03a], chapter 6, pages 69–82. ISBN 1-4020-1496-1, 94-017-0961-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-017-0961-3_6.

Butts:1993:PST

[But93a] Robert E. Butts. “A purely scientific temper”: Victorian expressions of the ideal of an autonomous science. In Historical Pragmatics: Philosophical Essays [But93d], chapter 13, pages 313–338. ISBN 90-481-4329-2, 94-015-8188-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-015-
8188-2_13.

Butts:1993:CIP

[But93b] Robert E. Butts. Consilience of inductions and the problem of conceptual change in science. In Historical Pragmatics: Philosophical Essays [But93d], chapter 11, pages 269–291. ISBN 90-481-4329-2, 94-015-8188-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-015-8188-2_11.

Butts:1993:GRH

[But93c] Robert E. Butts. The grammar of reason: Hamann’s challenge to Kant. In Historical Pragmatics: Philosophical Essays [But93d], chapter 7, pages 133–168. ISBN 90-481-4329-2, 94-015-8188-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

                                              ???? URL https://link.springer.
                                              com/chapter/10.1007/978-94-015-
                                              8188-2_7.

Butts:1993:HPP

[But93d] Robert E. Butts. Historical Pragmatics: Philosophical Essays, volume 155 of Boston Studies in the Philosophy of Science. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 1993. ISBN 90-481-4329-2, 94-015-8188-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). xiii + 370 pp. LCCN ????

                                              URL https://link.springer.com/
                                              book/10.1007/978-94-015-8188-2.

Butts:1993:KSS

[But93e] Robert E. Butts. Kant’s schemata as semantical rules. In Historical Pragmatics: Philosophical Essays [But93d], chapter 4, pages 67–78. ISBN 90-481-4329-2, 94-015-8188-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-015-8188-2_4.

Butts:1993:LCB

[But93f] Robert E. Butts. Leibniz on choosing between rival scientific hypotheses. In Historical Pragmatics: Philosophical Essays [But93d], chapter 2, pages 33– 52. ISBN 90-481-4329-2, 94-015-8188-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-015-8188-
2_2.

Butts:1993:MMP

[But93g] Robert E. Butts. Metaphysics, methodology and the pragmatic unity of the sciences. In Historical Pragmatics: Philosophical Essays [But93d], chapter 8, pages 171–185. ISBN 90-481-4329-2, 94-015-8188-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-015-8188-2_8.

Butts:1993:MSK

[But93h] Robert E. Butts. The methodological structure of Kant’s metaphysics of science. In Historical Pragmatics: Philosophical Essays [But93d], chapter 5, pages 79–113. ISBN 90-481-4329-2, 94-015-8188-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-015-8188-2_5.

Butts:1993:NTW

[But93i] Robert E. Butts. Necessary truth in Whewell’s theory of science. In Historical Pragmatics: Philosophical Essays [But93d], chapter 9, pages 189–233.

ISBN 90-481-4329-2, 94-015-8188-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-015-8188-
2_9.

Butts:1993:PPR

[But93j] Robert E. Butts. Philosophers as professional relativists. In Historical Pragmatics: Philosophical Essays [But93d], chapter 3, pages 55–63. ISBN 90-481-4329-2, 94-015-8188-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-015-8188-2_3.

Butts:1993:STG

[But93k] Robert E. Butts. Some tactics in Galileo’s propaganda for the mathematization of scientific experience. In Historical Pragmatics: Philosophical Essays [But93d], chapter 1, pages 3–31. ISBN 90-481-4329-2, 94-015-8188-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-015-8188-
2_1.

Butts:1993:TSM

[But93l] Robert E. Butts. Teleology and scientific method in Kant’s critique of judgment. In Historical Pragmatics: Philosophical Essays [But93d], chapter 6, pages 115–132. ISBN 90-481-4329-2, 94-015-8188-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-015-8188-2_6.

Butts:1993:WNR

[But93m] Robert E. Butts. Whewell on Newton’s rules of philosophizing. In Historical Pragmatics: Philosophical Essays [But93d], chapter 12, pages 293–312. ISBN 90-481-4329-2, 94-015-8188-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-015-8188-
2_12.

Butts:1993:WLI

[But93n] Robert E. Butts. Whewell’s logic of induction. In Historical Pragmatics: Philosophical Essays [But93d], chapter 10, pages 235–267. ISBN 90-481-4329-2, 94-015-8188-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-015-8188-2_10.

Buttner:2008:PCO

[Büt08] Jochen Büttner. The pendulum as a challenging object in early-modern mechanics. In Mechanics and Natural Philosophy Before the Scientific Revolution [LR08], chapter 9, pages 223–237. ISBN 1-4020-5966-3 (hardcover), 1-4020-5967-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QA802 .M43 2008. URL https:/

/link.springer.com/chapter/10.
1007/978-1-4020-5967-4_9.

Bashmakova:1994:JMH

[BV94] Izabella G. Bashmakova and Ioannis M.

Vandoulakis. On the justification of the method of historical interpretation. In Trends in the Historiography of Science [GCN94], chapter 19, pages 249–264. ISBN 90-481-4264-4, 94-017-3596-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-017-3596-
4_19.

Boissonnade:1997:AMT

[BV97] Auguste Boissonnade and Victor N. Vagliente, editors. Analytical Mechanics: Translated from the Mécanique analytique, novelle édition of 1811, volume 191 of Boston Studies in the Philosophy of Science. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 1997. ISBN 90-481-4779-4, 94-015-8903-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC5.53; QC6.9. URL

                                              https://link.springer.com/book/
                                              10.1007/978-94-015-8903-1.

Bensaude-Vincent:2001:GRP

[BV01] Bernadette Bensaude-Vincent. Graphic representations of the periodic system of chemical elements. In Klein [Kle01b], chapter 9, pages 133–161. ISBN 1-4020-0100-2, 94-015-9737-5. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 222; QD7. URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-015-9737-1_9.

Bensaude-Vincent:2009:PC

[BV09] Bernadette Bensaude-Vincent. Philosophy of chemistry. In Brenner and Gayon [BG09], chapter 7, pages 165–186. ISBN 1-4020-9367-5, 1-4020-9368-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .F746 2009.

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-1-4020-9368-
                                              5_7.

Bensaude-Vincent:2011:MM

[BV11] Bernadette Bensaude-Vincent. Materials as machines. In Science in the Context of Application [CN11a], chapter 7, pages 101–111. ISBN 90-481-9050-9, 90-481-9051-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q180.55.M4 S385

2011. URL https://link.springer.
com/chapter/10.1007/978-90-481-
9051-5_7.

Bohme:1983:FR

[BVH83a] Gernot Böhme, Wolfgang Van Den Daele, and Rainer Hohlfeld. Finalization revisited. In Finalization in Science: The Social Orientation of Scientific Progress [Sch83b], chapter 7, pages 131–172. ISBN 94-009-7082-X, 94-009-7080-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic).

LCCN ????  URL https://link.
springer.com/chapter/10.1007/
978-94-009-7080-9_7.

Bohme:1983:I

[BVH^{+}83b] Gernot Böhme, Wolfgang Van Den Daele, Rainer Hohlfeld, Wolfgang Krohn, and Wolf Schäfer. Introduction. In Finalization in Science: The Social Orientation of Scientific Progress [Sch83b], chapter 1, pages 3–11. ISBN 94-009-7082-X, 94-009-7080-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-009-7080-9_1.

Bohme:1983:IN

[BVH^{+}83c] Gernot Böhme, Wolfgang Van Den Daele, Rainer Hohlfeld, Wolfgang Krohn, and Wolf Schäfer. Introductory note. In Finalization in Science: The Social Orientation of Scientific Progress [Sch83b], chapter 13, pages 273–274. ISBN 94-009-7082-X, 94-009-7080-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-009-7080-
9_13.

Bohme:1983:SCa

[BVH^{+}83d] Gernot Böhme, Wolfgang Van Den Daele, Rainer Hohlfeld, Wolfgang Krohn, and Wolf Schäfer. Summary of contributions. In Finalization in Science: The Social Orientation of Scientific Progress [Sch83b], chapter 2, pages 15–16. ISBN 94-009-7082-X, 94-009-7080-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

                                              ???? URL https://link.springer.
                                              com/chapter/10.1007/978-94-009-
                                              7080-9_2.

Bohme:1983:SCb

[BVH^{+}83e] Gernot Böhme, Wolfgang Van Den Daele, Rainer Hohlfeld, Wolfgang Krohn, and Wolf Schäfer. Summary of contributions. In Finalization in Science: The Social Orientation of Scientific Progress [Sch83b], chapter 6, pages 129–130. ISBN 94-009-7082-X, 94-009-7080-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic).

                                              LCCN ????  URL https://link.
                                              springer.com/chapter/10.1007/
                                              978-94-009-7080-9_6.

Bohme:1983:SCc

[BVH^{+}83f] Gernot Böhme, Wolfgang Van Den Daele, Rainer Hohlfeld, Wolfgang Krohn, and Wolf Schäfer. Summary of contributions. In Finalization in Science: The Social Orientation of Scientific Progress [Sch83b], chapter 10, pages 235–236. ISBN 94-009-7082-X, 94-009-7080-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic).

                                              LCCN ????  URL https://link.
                                              springer.com/chapter/10.1007/
                                              978-94-009-7080-9_10.

Bohme:1983:ST

[BVK83] Gernot Böhme, Wolfgang Van Den Daele, and Wolfgang Krohn. The scientification of technology. In Finalization in Science: The Social Orientation of Scientific Progress [Sch83b], chapter 8, pages 173–205. ISBN 94-009-7082-X, 94-009-7080-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic).

LCCN ????  URL https://link.
springer.com/chapter/10.1007/
978-94-009-7080-9_8.

Bycroft:2016:HSS

[Byc16] Michael Bycroft. How to save the symmetry principle. In Sauer and Scholl [SS16b], chapter 2, pages 11–29. ISBN 3-319-30227-2, 3-319-30229-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q126.8 .P45 2016.

URL https://link.springer.com/
chapter/10.1007/978-3-319-30229-
4_2.

Byerly:1974:RFM

[Bye74] Henry C. Byerly. Realist foundations of measurement. In Schaffner and Cohen [SC74a], pages 375–384. ISBN 90-277-0408-2, 90-277-0409-0, 94-010-2140-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175 .B7312.

URL https://link.springer.com/
chapter/10.1007/978-94-010-2140-
1_26.

Bynum:1989:OHH

[Byn89] Paige Matthey Bynum. “Observe how healthily — how calmly I can tell you the whole story”: Moral insanity and Edgar Allan Poe’s ‘The Tell-Tale Heart’. In Literature and Science as Modes of Expression [Amr89], chapter 8, pages 141–152. ISBN 94-010-7531-X, 94-009-2297-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-009-2297-6_8.

Byrne:1983:CGR

[Byr83]  Edmund Byrne.  Can government

regulate technology? In Philosophy and Technology [DR83], chapter 2, pages 17–33. ISBN 94-009-7126-5, 94-009-7124-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

                                              ???? URL https://link.springer.
                                              com/chapter/10.1007/978-94-009-
                                              7124-0_2.

Byrne:1986:MWR

[Byr86] Edmund F. Byrne. Microelectronics and workers’ rights. In Philosophy and Technology II: Information Technology and Computers in Theory and Practice [MH86], chapter 14, pages 205–216. ISBN 94-010-8510-2, 94-009-4512-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-009-4512-
                                              8_14.

Byrnes:2002:DMD

[Byr02] Heidi Byrnes. The dialogism of meaning, the discursive embeddedness of knowledge, the colloquy of being. In Hermeneutic Philosophy of Science, Van Gogh’s Eyes, and God: Essays in Honor of Patrick A. Heelan, S.J. [Bab02c], chapter 36, pages 411–422. ISBN 90-481-5926-1, 94-017-1767-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 225. URL https://link.

                                              springer.com/chapter/10.1007/
                                              978-94-017-1767-0_36.

Binglun:1996:SBE

[BZ96] Zhang Binglun and Wang Zichun. The struggle between evolutionary theory and Creationism in China. In Chinese Studies in the History and Philosophy of Science and Technology [DC96], chapter 24, pages 289–302. ISBN 90-481-4546-5, 94-015-8717-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 180; U390 .N19 1996. URL https:

//link.springer.com/chapter/10.
1007/978-94-015-8717-4_24.

Cackowski:1982:IRN

[Cac82]  Zdzisław Cackowski.  Information,

regulation, negentropy. In Polish Essays in the Philosophy of the Natural Sciences [Kra82c], chapter 22, pages 249–264. ISBN 94-009-7705-0 (e-book), 90-277-1287-5. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-009-7705-
1_22.

Cahoone:1998:WNW

[Cah98] Lawrence Cahoone. Whose nature?

Which morality? On Kohák’s moral sense of nature. In Philosophies of Nature: The Human Dimension: In Celebration of Erazim Kohák [CT98], chapter 2, pages 19–33. ISBN 90-481-4859-6, 94-017-2614-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN BD581. URL https:

//link.springer.com/chapter/10.
1007/978-94-017-2614-6_2.

Cahn:2005:OOP

[Cah05] Michael Cahn. Opera Omnia: The production of cultural authority. In History of Science, History of Text [Che05a], chapter 3, pages 81–94. ISBN 1-281-33589-4, 1-4020-2320-0 (hardcover), 1-4020-2321-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 238 Q125.

URL https://link.springer.com/
chapter/10.1007/1-4020-2321-9_3.

Caianiello:2014:LPL

[Cai14] Eva Caianiello. Leonardo of Pisa and the Liber Abaci. Biographical elements and the project of the work. In Bernard and Proust [BP14b], chapter 9, pages 217–246. ISBN 94-007-5121-4 (hardcover), 94-007-5122-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q181 .S35 2014; Q181 S387 2014. URL https://link.

                                              springer.com/chapter/10.1007/
                                              978-94-007-5122-4_9.

Callicott:2013:EMO

[Cal13]  J. Baird Callicott.  Ecology and

moral ontology. In Bergandi [Ber13c], chapter 7, pages 101–116. ISBN 94-007-5066-8, 94-007-5067-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN GE42 .S77 2013; Q175 .B73 v. 296. URL https:/

                                              /link.springer.com/chapter/10.
                                              1007/978-94-007-5067-8_7.

Carneiro:2003:RTI

[CALP03] Ana Carneiro, Dores Areias, Vanda Leitão, and Luís Teixeira Pinto. The role of travels in the internationalisation of Nineteenth Century Portuguese geological science. In Simões et al. [SCD03b], chapter 11, pages 249–297. ISBN 1-4020-1259-4, 94-017-3584-0. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 v. 233.

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-017-3584-
                                              1_11.

Campbell:1987:NEB

[Cam87] Donald T. Campbell. Neurological embodiments of belief and the gaps in the fit of phenomena to noumena. In Shimony and Nails [SN87], pages 165–192. ISBN 94-010-8168-9, 94-009-3735-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-009-
3735-2_10.

Campaner:2003:APE

[Cam03]  Raffaella Campaner.  An attempt

at a philosophy of experimental error: a comment on Giora Hon. In Descartes’s Mathematical Thought [Sas03d], chapter 19, pages 285–293. ISBN 90-481-6487-7, 94-017-1225-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/0-306-48123-5_19.

See [Hon03].

Caneva:2007:OPH

[Can07] Kenneth L. Caneva. Ørsted’s presentation of others’ — and his own — work. In Hans Christian Ørsted and The Romantic Legacy In Science: Ideas, Disciplines, Practices [BCK07], chapter 15, pages 273–338. ISBN 1-4020-2979-9 (hardcover), 1-4020-2987-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC515.O37 H36 2007. URL https:

//link.springer.com/chapter/10.
1007/978-1-4020-2987-5_15.

Caneva:2012:WTD

[Can12] Kenneth L. Caneva. What in truth divides historians and philosophers of science? In Mauskopf and Schmaltz [MS12b], chapter 4, pages 49–57. ISBN 94-007-1744-X, 94-007-1745-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175 .I5134 2011.

URL https://link.springer.com/
chapter/10.1007/978-94-007-1745-
9_4.

Canales:2016:EBP

[Can16] Jimena Canales. Einstein’s Bergson problem: Communication, consensus and good science. In Cosmological and Psychological Time [DR16], chapter 4, pages 53–72. ISBN 3-319-22589-8 (hardcover), 3-319-22590-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN BD638 C67 2016.

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-3-319-22590-
                                              6_4.

Cao:1995:PLS

[Cao95] Tian Yu Cao. A philosopher looks at science. In Gavroglu et al. [GSW95c], chapter 9, pages 171–187. ISBN 0-7923-2991-0 (set), 0-7923-2988-0, 94-017-2658-2. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 163. URL https://link.

                                              springer.com/chapter/10.1007/
                                              978-94-017-2658-0_9.

Cao:1998:MTR

[Cao98] Tian Yu Cao. Monism, but not through reductionism. In Philosophies of Nature: The Human Dimension: In Celebration of Erazim Kohák [CT98], chapter 4, pages 39–51. ISBN 90-481-4859-6, 94-017-2614-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN BD581. URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-017-2614-6_4.

Cao:2007:WES

[Cao07] Tian Yu Cao. Will Einstein still be the super-hero of physics history in 2050? In Positioning the history of science [RG07], chapter 4, pages 27–32. ISBN 1-4020-5419-X (hardcover), 1-4020-5420-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q125 .P77 2007.

URL https://link.springer.com/
chapter/10.1007/1-4020-5420-3_4.

Capek:1969:EMB

[Čap69] Milič Čapek. Ernst Mach’s biological theory of knowledge. In Cohen and Wartofsky [CW69a], pages 400–420. ISBN 94-010-3383-8, 94-010-3263-7, 94-010-3381-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-010-
3381-7_13.

Capek:1971:BEP

[Čap71a] Milič Čapek. Bergson and Einstein.

The physical world as extensive becoming. In Bergson and Modern Physics: A Reinterpretation and Re-evaluation [Čap71c], chapter 33, pages 238–256. ISBN 94-010-3098-7, 94-010-3096-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-010-3096-
0_33.

Capek:1971:BLB

[Čap71b] Milič Čapek. Bergson and Louis de Broglie. In Bergson and Modern Physics: A Reinterpretation and Reevaluation [Čap71c], chapter 38, pages 292–301. ISBN 94-010-3098-7, 94-010-3096-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-010-
3096-0_38.

Capek:1971:BMP

[Čap71c] Milič Čapek. Bergson and Modern Physics: A Reinterpretation and Reevaluation, volume 7 of Boston Studies in the Philosophy of Science. D. Reidel, Dordrecht, The Netherlands; Boston, MA, USA; Lancaster, UK; Tokyo, Japan, 1971. ISBN 94-010-3098-7, 94-010-3096-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). xvi + 418 pp. LCCN ???? URL

https://link.springer.com/book/
10.1007/978-94-010-3096-0.

Capek:1971:BRP

[Čap71d] Milič Čapek. Bergson, Reichenbach and Piaget. In Bergson and Modern Physics: A Reinterpretation and Re-evaluation [Čap71c], chapter 8, pages 65–71. ISBN 94-010-3098-7, 94-010-3096-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-010-
3096-0_8.

Capek:1971:BAC

[Čap71e] Milič Čapek. Bergson’s amendment of the classical biological theory of knowledge. In Bergson and Modern Physics: A Reinterpretation and Re-evaluation [Čap71c], chapter 3, pages 30–39. ISBN 94-010-3098-7, 94-010-3096-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-010-3096-0_3.

Capek:1971:CVC

[Čap71f] Milič Čapek. Change without vehicle and container. Fallacy of simple location. In Bergson and Modern Physics: A Reinterpretation and Re-evaluation [Čap71c], chapter 35, pages 270–277. ISBN 94-010-3098-7, 94-010-3096-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-010-3096-
0_35.

Capek:1971:CBTa

[Čap71g] Milič Čapek. The classical biological theory of knowledge. In Bergson and Modern Physics: A Reinterpretation and Re-evaluation [Čap71c], chapter 1, pages 3–14. ISBN 94-010-3098-7, 94-010-3096-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-010-
3096-0_1.

Capek:1971:CRW

[Čap71h] Milič Čapek. Concluding remarks: The world of Laplace and the world of Bergson. In Bergson and Modern Physics: A Reinterpretation and Re-evaluation [Čap71c], chapter 41, pages 331–334. ISBN 94-010-3098-7, 94-010-3096-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-010-3096-
0_41.

Capek:1971:CBI

[Čap71i] Milič Čapek. Content of the Bergsonian intuition. In Bergson and Modern Physics: A Reinterpretation and Reevaluation [Čap71c], chapter 11, pages 89–91. ISBN 94-010-3098-7, 94-010-3096-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-010-
3096-0_11.

Capek:1971:CBTb

[Čap71j] Milič Čapek. The contrast between technical control and intellectual insight: The persistent influence of macroscopic imagery. In Bergson and Modern Physics: A Reinterpretation and Re-evaluation [Čap71c], chapter 5, pages 46–50. ISBN 94-010-3098-7, 94-010-3096-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-010-
3096-0_5.

Capek:1971:CDT

[Čap71k] Milič Čapek. The correlation of different temporal rhythms with different degrees of extension. In Bergson and Modern Physics: A Reinterpretation and Re-evaluation [Čap71c], chapter 30, pages 214–222. ISBN 94-010-3098-7, 94-010-3096-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-010-
3096-0_30.

Capek:1971:DMI

[Čap71l] Milič Čapek. The deeper meaning of the ‘indivisible heterogeneity’ of duration. In Bergson and Modern Physics: A Reinterpretation and Re-evaluation [Čap71c], chapter 17, pages 126–132. ISBN 94-010-3098-7, 94-010-3096-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-010-3096-
0_17.

Capek:1971:DDT

[Čap71m] Milič Čapek. Different degrees of temporal span. Microcosmos as microchronos. In Bergson and Modern Physics: A Reinterpretation and Reevaluation [Čap71c], chapter 27, pages 195–201. ISBN 94-010-3098-7, 94-010-3096-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-010-
3096-0_27.

Capek:1971:DCU

[Čap71n] Milič Čapek. Duration as concrete universal. Bergson and Croce. In Bergson and Modern Physics: A Reinterpretation and Re-evaluation [Čap71c], chapter 24, pages 170–175. ISBN 94-010-3098-7, 94-010-3096-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-010-3096-0_24.

Capek:1971:DCD

[Čap71o] Milič Čapek. The dynamic continuity of duration. In Bergson and Modern Physics: A Reinterpretation and Re-evaluation [Čap71c], chapter 12, pages 92–98. ISBN 94-010-3098-7, 94-010-3096-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-010-
3096-0_12.

Capek:1971:HDL

[Čap71p] Milič Čapek. The heterogeneity of duration: Lovejoy–Ushenko’s objections. In Bergson and Modern Physics: A Reinterpretation and Re-evaluation [Čap71c], chapter 16, pages 118–125. ISBN 94-010-3098-7, 94-010-3096-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-010-3096-
0_16.

Capek:1971:IPB

[Čap71q] Milič Čapek. Immortality of the past:

Bergson and Whitehead. In Bergson and Modern Physics: A Reinterpretation and Re-evaluation [Čap71c], chapter 21, pages 152–154. ISBN 94-010-3098-7, 94-010-3096-0 (e-book).

ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-010-3096-0_21.

Capek:1971:IAT

[Čap71r] Milič Čapek. The inadequacy of the atomistic theory of time. In Bergson and Modern Physics: A Reinterpretation and Re-evaluation [Čap71c], chapter 19, pages 142–146. ISBN 94-010-3098-7, 94-010-3096-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-010-3096-0_19.

Capek:1971:IDN

[Čap71s] Milič Čapek. The incompleteness of duration: Novelty and its denials. In Bergson and Modern Physics: A Reinterpretation and Re-evaluation [Čap71c], chapter 13, pages 99–105. ISBN 94-010-3098-7, 94-010-3096-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-010-3096-
                                              0_13.

Capek:1971:ISH

[Čap71t] Milič Čapek. The intermediate stage: Helmholtz, Mach and Poincaré. In Bergson and Modern Physics: A Reinterpretation and Re-evaluation [Čap71c], chapter 2, pages 15–29. ISBN 94-010-3098-7, 94-010-3096-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-010-3096-0_2.

Capek:1971:IDC

[Čap71u] Milič Čapek. The irreversibility of duration: The comments of Royce and Ingarden. In Bergson and Modern Physics: A Reinterpretation and Reevaluation [Čap71c], chapter 23, pages 164–169. ISBN 94-010-3098-7, 94-010-3096-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-010-
3096-0_23.

Capek:1971:JBV

[Čap71v] Milič Čapek. James’s and Bergson’s views of the past compared. In Bergson and Modern Physics: A Reinterpretation and Re-evaluation [Čap71c], chapter 22, pages 155–163. ISBN 94-010-3098-7, 94-010-3096-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-010-3096-0_22.

Capek:1971:JIL

[Čap71w] Milič Čapek. Juxtaposition as the ideal limit of distended duration. In Bergson and Modern Physics: A Reinterpretation and Re-evaluation [Čap71c], chapter 31, pages 223–225. ISBN 94-010-3098-7, 94-010-3096-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-010-3096-0_31.

Capek:1971:LFA

[Čap71x] Milič Čapek. The Leibniz–Fouillé argument for the compatibility of succession and determinism. In Bergson and Modern Physics: A Reinterpretation and Re-evaluation [Čap71c], chapter 15, pages 113–117. ISBN 94-010-3098-7, 94-010-3096-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-010-3096-0_15.

Capek:1971:LUC

[Čap71y] Milič Čapek. Limitations and usefulness of the corpuscular models. In Bergson and Modern Physics: A Reinterpretation and Re-evaluation [Čap71c], chapter 34, pages 257–269. ISBN 94-010-3098-7, 94-010-3096-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-010-3096-
                                              0_34.

Capek:1971:LP

[Čap71z] Milič Čapek. Limitations of panmathematism. In Bergson and Modern Physics: A Reinterpretation and Reevaluation [Čap71c], chapter 6, pages 51–57. ISBN 94-010-3098-7, 94-010-3096-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

                                              ???? URL https://link.springer.
                                              com/chapter/10.1007/978-94-010-
                                              3096-0_6.

Capek:1971:LCS

[Čap71-27] Milič Čapek. Limits of the criticism of simple location: Contemporary independence. In Bergson and Modern Physics: A Reinterpretation and Reevaluation [Čap71c], chapter 36, pages 278–283. ISBN 94-010-3098-7, 94-010-3096-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

                                              ???? URL https://link.springer.
                                              com/chapter/10.1007/978-94-010-
                                              3096-0_36.

Capek:1971:LSB

[Čap71-28] Milič Čapek. The logic of solid bodies from Plato to Quine. In Bergson and Modern Physics: A Reinterpretation and Re-evaluation [Čap71c], chapter 9, pages 72–80. ISBN 94-010-3098-7, 94-010-3096-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-010-
3096-0_9.

Capek:1971:MI

[Čap71-29] Milič Čapek. The meaning of immediacy. In Bergson and Modern Physics: A Reinterpretation and Re-evaluation [Čap71c], chapter 10, pages 83–88. ISBN 94-010-3098-7, 94-010-3096-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-010-3096-
0_10.

Capek:1971:NIS

[Čap71-30] Milič Čapek. The negation of instantaneous space in the relativistic physics. In Bergson and Modern Physics: A Reinterpretation and Re-evaluation [Čap71c], chapter 32, pages 226–237. ISBN 94-010-3098-7, 94-010-3096-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-010-3096-
0_32.

Capek:1971:NAB

[Čap71-31] Milič Čapek. Negative aspects of Bergson’s epistemology — its relations to Bachelard, Bridgman and empirio-criticism. In Bergson and Modern Physics: A Reinterpretation and Re-evaluation [Čap71c], chapter 7, pages 58–64. ISBN 94-010-3098-7, 94-010-3096-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.

                                              com/chapter/10.1007/978-94-010-
                                              3096-0_7.

Capek:1971:OBP

[Čap71-32] Milič Čapek. An outline of Bergson’s philosophy of mathematics. In Bergson and Modern Physics: A Reinterpretation and Re-evaluation [Čap71c], chapter 25, pages 176–186. ISBN 94-010-3098-7, 94-010-3096-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-010-3096-0_25.

Capek:1971:PEP

[Čap71-33] Milič Čapek. Physical events as protomental entities. Bergson, Whitehead and Bohm. In Bergson and Modern Physics: A Reinterpretation and Reevaluation [Čap71c], chapter 39, pages 302–312. ISBN 94-010-3098-7, 94-010-3096-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

                                              ???? URL https://link.springer.
                                              com/chapter/10.1007/978-94-010-
                                              3096-0_39.

Capek:1971:RDP

[Čap71-34] Milič Čapek. The reality of duration in the physical world and its implications. In Bergson and Modern Physics: A Reinterpretation and Re-evaluation [Čap71c], chapter 26, pages 189–194. ISBN 94-010-3098-7, 94-010-3096-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-010-3096-
                                              0_26.

Capek:1971:RCD

[Čap71-35] Milič Čapek. The rejection of the Cartesian dogma of the completely extensionless mind. In Bergson and Modern Physics: A Reinterpretation and Re-evaluation [Čap71c], chapter 29, pages 208–213. ISBN 94-010-3098-7, 94-010-3096-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-010-
3096-0_29.

Capek:1971:SLA

[Čap71-36] Milič Čapek. The significance and the limitations of auditory models. Bergson and Strawson. In Bergson and Modern Physics: A Reinterpretation and Reevaluation [Čap71c], chapter 40, pages 313–330. ISBN 94-010-3098-7, 94-010-3096-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-010-
3096-0_40.

Capek:1971:SPR

[Čap71-37] Milič Čapek. The significance of Piaget’s researches on the psychogenesis of atomism. In Buck and Cohen [BC71], pages 446–455. ISBN 90-277-0309-4, 94-010-3142-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-010-
3142-4_28.

Capek:1971:SSD

[Čap71-38] Milič Čapek. Superfluity of succession in the deterministic schemes. In Bergson and Modern Physics: A Reinterpretation and Re-evaluation [Čap71c], chapter 14, pages 106–112. ISBN 94-010-3098-7, 94-010-3096-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-010-3096-
0_14.

Capek:1971:TFQ

[Čap71-39] Milič Čapek. Two fundamental questions. In Bergson and Modern Physics: A Reinterpretation and Re-evaluation [Čap71c], chapter 28, pages 202–207. ISBN 94-010-3098-7, 94-010-3096-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-010-3096-
                                              0_28.

Capek:1971:UMD

[Čap71-40] Milič Čapek. The unity and multiplicity of duration: Bergson, Russell and Brouwer. In Bergson and Modern Physics: A Reinterpretation and Reevaluation [Čap71c], chapter 20, pages 147–151. ISBN 94-010-3098-7, 94-010-3096-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

                                              ???? URL https://link.springer.
                                              com/chapter/10.1007/978-94-010-
                                              3096-0_20.

Capek:1971:UDI

[Čap71-41] Milič Čapek. The unreality of durationless instants: Becoming not mathematically continuous. In Bergson and Modern Physics: A Reinterpretation and Re-evaluation [Čap71c], chapter 18, pages 133–141. ISBN 94-010-3098-7, 94-010-3096-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-010-3096-0_18.

Capek:1971:WMP

[Čap71-42] Milič Čapek. Why mechanical–pictorial models failed. In Bergson and Modern Physics: A Reinterpretation and Re-evaluation [Čap71c], chapter 4, pages 40–45. ISBN 94-010-3098-7, 94-010-3096-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-010-
3096-0_4.

Capek:1974:TTC

[Čap74] Milič Čapek. Two types of continuity.

In Logical and Epistemological Studies in Contemporary Physics [CW74a], chapter 13, pages 361–375. ISBN 90-277-0391-4 (hardcover), 90-277-0377-9 (paperback), 94-010-2656-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol.

13; QC5.56. URL https://link.
springer.com/chapter/10.1007/
978-94-010-2656-7_13.

Capek:1976:CST

[Čap76] Milič Čapek. The Concepts of Space and Time: Their Structure and Their Development, volume 22 of Boston Studies in the Philosophy of Science. D. Reidel, Dordrecht, The Netherlands; Boston, MA, USA; Lancaster, UK; Tokyo, Japan, 1976. ISBN 90-277-0375-2, 94-010-1727-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). lvii + 570 pp. LCCN BD632.

URL https://link.springer.com/
book/10.1007/978-94-010-1727-5.

Capozzi:1981:KMD

[Cap81a] Mirella Capozzi. Kant on mathematical definition. In Italian Studies in the Philosophy of Science [Chi81b], chapter 23, pages 423–452. ISBN 90-277-1073-2, 94-009-8937-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-009-8937-5_23.

Cappelletti:1981:DV

[Cap81b] Vincenzo Cappelletti. Discovery and vocation. In Grmek et al. [GCC81b], pages 211–226. ISBN 90-277-1123-2, 94-010-1284-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

                                              ???? URL https://link.springer.
                                              com/chapter/10.1007/978-94-010-
                                              1284-3_10.

Capek:1984:HOV

[Čap84a] Milič Čapek. Hegel and the organic view of nature. In Hegel and the Sciences [CW84a], chapter 8, pages 109–121. ISBN 94-009-6235-5, 94-009-6233-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

                                              ???? URL https://link.springer.
                                              com/chapter/10.1007/978-94-009-
                                              6233-0_8.

Capek:1984:PE

[Čap84b] Milič Čapek. Particles or events? In Physical Sciences and History of Physics [CW84c], chapter 1, pages 1– 28. ISBN 94-009-7180-X, 94-009-7178-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .P497

                                              1984. URL https://link.springer.
                                              com/chapter/10.1007/978-94-009-
                                              7178-3_1.

Capek:1987:PSP

[Čap87a] Milič Čapek. The philosophical significance of Piaget’s researches on the genesis of the concept of time. In Shimony and Nails [SN87], pages 91–111. ISBN 94-010-8168-9, 94-009-3735-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

                                              ???? URL https://link.springer.
                                              com/chapter/10.1007/978-94-009-

3735-2_5. See [Shi87e] and reply [Čap87b].

Capek:1987:RS

[Čap87b] Milič Čapek. Reply to Shimony. In Shimony and Nails [SN87], pages 115–118.

ISBN 94-010-8168-9, 94-009-3735-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-009-3735-

2_7. See [Čap87a, Shi87e].

Capek:1991:BLB

[Čap91a] Milič Čapek. Bergson and Louis de Broglie. In The New Aspects of Time: Its Continuity and Novelties [Čap91g], chapter 15, pages 286–295. ISBN 94-010-7455-0, 94-009-2123-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-009-2123-8_15.

Capek:1991:DRE

[Čap91b] Milič Čapek. The development of Reichenbach’s epistemology. In The New Aspects of Time: Its Continuity and Novelties [Čap91g], chapter 7, pages 103–128. ISBN 94-010-7455-0, 94-009-2123-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-009-
2123-8_7.

Capek:1991:ENP

[Čap91c] Milič Čapek. The elusive nature of the past. In The New Aspects of Time: Its Continuity and Novelties [Čap91g], chapter 2, pages 26–42. ISBN 94-010-7455-0, 94-009-2123-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-009-2123-8_2.

Capek:1991:ELI

[Čap91d] Milič Čapek. The end of the Laplacian illusion. In The New Aspects of Time: Its Continuity and Novelties [Čap91g], chapter 12, pages 221–264. ISBN 94-010-7455-0, 94-009-2123-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-009-2123-8_12.

Capek:1991:ERO

[Čap91e] Milič Čapek. Eternal recurrence — once more. In The New Aspects of Time: Its Continuity and Novelties [Čap91g], chapter 13, pages 265–277. ISBN 94-010-7455-0, 94-009-2123-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-009-2123-
                                              8_13.

Capek:1991:FI

[Čap91f] Milič Čapek. The fiction of instants. In The New Aspects of Time: Its Continuity and Novelties [Čap91g], chapter 3, pages 43–55. ISBN 94-010-7455-0, 94-009-2123-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

                                              ???? URL https://link.springer.
                                              com/chapter/10.1007/978-94-009-
                                              2123-8_3.

Capek:1991:NAT

[Čap91g] Milič Čapek. The New Aspects of Time: Its Continuity and Novelties, volume 125 of Boston Studies in the Philosophy of Science. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 1991. ISBN 94-010-7455-0, 94-009-2123-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). 380 (est.) pp. LCCN ????

                                              URL https://link.springer.com/
                                              book/10.1007/978-94-009-2123-8.

Capek:1991:NAW

[Čap91h] Milič Čapek. Note about Whitehead’s definitions of co-presence. In The New Aspects of Time: Its Continuity and Novelties [Čap91g], chapter 14, pages 278–285. ISBN 94-010-7455-0, 94-009-2123-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-009-
2123-8_14.

Capek:1991:PE

[Čap91i] Milič Čapek. Particles or events?

In The New Aspects of Time: Its Continuity and Novelties [Čap91g], chapter 11, pages 191–218. ISBN 94-010-7455-0, 94-009-2123-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-009-2123-8_11.

Capek:1991:PPB

[Čap91j] Milič Čapek. Process and personality in Bergson’s thought. In The New Aspects of Time: Its Continuity and Novelties [Čap91g], chapter 5, pages 71–88. ISBN 94-010-7455-0, 94-009-2123-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-009-2123-8_5.

Capek:1991:RHB

[Čap91k] Milič Čapek. Russell’s hidden Bergsonism. In The New Aspects of Time: Its Continuity and Novelties [Čap91g], chapter 6, pages 89–99. ISBN 94-010-7455-0, 94-009-2123-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-009-2123-8_6.

Capek:1991:SPR

[Čap91l] Milič Čapek. The significance of Piaget’s researches on the psychogenesis of atomism. In The New Aspects of Time: Its Continuity and Novelties [Čap91g], chapter 8, pages 129–138. ISBN 94-010-7455-0, 94-009-2123-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-009-2123-
                                              8_8.

Capek:1991:SLF

[Čap91m] Milič Čapek. Simple location and fragmentation of reality. In The New Aspects of Time: Its Continuity and Novelties [Čap91g], chapter 10, pages 167–190. ISBN 94-010-7455-0, 94-009-2123-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

                                              ???? URL https://link.springer.
                                              com/chapter/10.1007/978-94-009-
                                              2123-8_10.

Capek:1991:SCD

[Čap91n] Milič Čapek. Stream of consciousness and “durée réelle”. In The New Aspects of Time: Its Continuity and Novelties [Čap91g], chapter 1, pages 3–25. ISBN 94-010-7455-0, 94-009-2123-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-009-2123-8_1.

Capek:1991:TSR

[Čap91o] Milič Čapek. Time–space rather than space–time. In The New Aspects of Time: Its Continuity and Novelties [Čap91g], chapter 17, pages 324–343. ISBN 94-010-7455-0, 94-009-2123-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-009-2123-
8_17.

Capek:1991:TWN

[Čap91p] Milič Čapek. Toward a widening of the notion of causality. In The New Aspects of Time: Its Continuity and Novelties [Čap91g], chapter 9, pages 139–166. ISBN 94-010-7455-0, 94-009-2123-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-009-2123-
8_9.

Capek:1991:TTC

[Čap91q] Milič Čapek. Two types of continuity.

In The New Aspects of Time: Its Continuity and Novelties [Čap91g], chapter 4, pages 56–70. ISBN 94-010-7455-0, 94-009-2123-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-009-2123-8_4.

Capek:1991:WLW

[Čap91r] Milič Čapek. What is living and what is dead in the Bergsonian critique of relativity. In The New Aspects of Time: Its Continuity and Novelties [Čap91g], chapter 16, pages 296–323. ISBN 94-010-7455-0, 94-009-2123-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-009-2123-8_16.

Capella:2015:BES

[Cap15] Vicente Bellver Capella. Biotechnology, ethics, and society: The case of genetic manipulation. In New Perspectives on Technology, Values, and Ethics: Theoretical and Practical [Gon15a], chapter 7, pages 123–143. ISBN 3-319-21869-7 (hardcover), 3-319-21870-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-3-319-21870-
0_7.

Cartwright:1974:DTI

[Car74] Nancy Cartwright. A dilemma for the traditional interpretation of quantum mixtures. In Schaffner and Cohen [SC74a], pages 251–258. ISBN 90-277-0408-2, 90-277-0409-0, 94-010-2140-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175 .B7312.

URL https://link.springer.com/
chapter/10.1007/978-94-010-2140-
1_16.

Carpenter:1983:TAN

[Car83] Stanley R. Carpenter. Technoaxiology: Appropriate norms for technology assessment. In Philosophy and Technology [DR83], chapter 8, pages 115–136. ISBN 94-009-7126-5, 94-009-7124-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-009-7124-
0_8.

Cartwright:1986:TKT

[Car86] Nancy Cartwright. Two kinds of teleological explanation. In Cohen and Wartofsky [CW85], chapter 10, pages 201–210. ISBN 90-277-1971-3, 94-009-5345-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-009-
5349-9_10.

Carrier:1989:CC

[Car89a] Martin Carrier. Circles without circularity. In An Intimate Relation:

Studies in the History and Philosophy of Science Presented to Robert E. Butts on his 60th Birthday [BM89], chapter 20, pages 405–428. ISBN 94-010-7546-8, 94-009-2327-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-009-2327-0_20.

Carter:1989:PSF

[Car89b] Steven Carter. “A place to step further”: Jack Spicer’s quantum poetics. In Literature and Science as Modes of Expression [Amr89], chapter 11, pages 177–188. ISBN 94-010-7531-X, 94-009-2297-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-009-
2297-6_11.

Cartwright:2000:QMO

[Car00] Nancy Cartwright. Quantum mechanics without the observables. In The Reality of the Unobservable: Observability, Unobservability and Their Impact on the Issue of Scientific Realism [AP00b], chapter 18, pages 241–249. ISBN 90-481-5458-8, 94-015-9391-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-015-9391-5_18.

Carrier:2001:CLS

[Car01] Martin Carrier. Changing laws and shifting concepts. In Incommensurability and Related Matters [HHS01], chapter 2, pages 65–90. ISBN 90-481-5709-9, 94-015-9680-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-015-9680-0_2.

Cartan:2007:DCM

[Car07] Elie Cartan. The dynamics of continuous media and the notion of an affine connection on space–time. In Renn et al. [RSS^{+}07b], pages 1107–1129. ISBN 1-4020-3999-9 (hardcover), 1-4020-4000-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 no. 250 v. 4; QC173.6

.G469 2007. URL http://d-nb.info/
985765933/34;   http://www.loc.
gov/catdir/enhancements/fy0824/
2007440325-d.html;   http://www.
loc.gov/catdir/enhancements/
fy0824/2007440325-t.html;
https://link.springer.com/book/
10.1007/978-1-4020-4000-9. Orig-

inally published as “Sur les variétés à connexion affine et la théorie de la relativité généralisée”, Annales Scientifiques de l’École Normal Supérieure, 325–412 (1923).

Carrier:2008:ASM

[Car08] Martin Carrier. The aim and structure of methodological theory. In Rethinking Scientific Change and Theory Comparison: Stabilities, Ruptures, Incommensurabilities? [SSHH08], chapter 20, pages 273–290. ISBN 1-4020-6274-5 (hardcover), 1-4020-6279-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175 .R397

2008. URL https://link.springer.
com/chapter/10.1007/978-1-4020-
6279-7_20.

Cartwright:2010:MPV

[Car10] Nancy Cartwright. Models: Parables v fables. In Frigg and Hunter [FH10b], chapter 2, pages 19–31. ISBN 90-481-3850-7 (hardcover), 90-481-3851-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN BH301.M47 B49 2010; Q175 .B52 2010.

URL https://link.springer.com/
chapter/10.1007/978-90-481-3851-
7_2.

Carrier:2011:KPC

[Car11] Martin Carrier. Knowledge, politics, and commerce: Science under the pressure of practice. In Science in the Context of Application [CN11a], chapter 2, pages 11–30. ISBN 90-481-9050-9, 90-481-9051-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q180.55.M4 S385

2011. URL https://link.springer.
com/chapter/10.1007/978-90-481-
9051-5_2.

Caracciolo:2015:FIH

[Car15a] Edoardo Caracciolo. Formalization and intuition in Husserl’s Raumbuch. In From Logic to Practice: Italian Studies in the Philosophy of Mathematics [LPV15], chapter 3, pages 33–50. ISBN 3-319-10433-0, 3-319-10434-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QA8.4 .F76 2015; Q174 .B67 v.308. URL https://link.

springer.com/chapter/10.1007/
978-3-319-10434-8_3.

Carolino:2015:FLP

[Car15b] Luís Miguel Carolino. Foundation of the Lisbon Polytechnic School Astronomical Observatory in the late Nineteenth Century: A step towards establishing a university. In Sciences in the Universities of Europe, Nineteenth and Twentieth Centuries: Academic Landscapes [SDG15b], chapter 14, pages 227–244. ISBN 94-017-9635-1, 94-017-9636-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q183.4.E85 2015. URL https:

//link.springer.com/chapter/10.
1007/978-94-017-9636-1_14.

Casari:1981:POT

[Cas81] Ettore Casari. Positively omitting types. In Italian Studies in the Philosophy of Science [Chi81b], chapter 1, pages 3–11. ISBN 90-277-1073-2, 94-009-8937-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

                                              ???? URL https://link.springer.
                                              com/chapter/10.1007/978-94-009-
                                              8937-5_1.

Castaneda:1995:WMP

[Cas95a] Guillermo Zambrana Castañeda. Wittgenstein on mathematical proof. In Mexican Studies in the History and Philosophy of Science [RC95], chapter 15, pages 235–248. ISBN 94-010-6535-7 (hardcover), 94-009-0109-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-009-0109-
                                              4_15.

Castaneda:1995:RV

[Cas95b] Santiago Ramirez Castañeda. A return to Vienna. In Mexican Studies in the History and Philosophy of Science [RC95], chapter 17, pages 265–273. ISBN 94-010-6535-7 (hardcover), 94-009-0109-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

                                              ???? URL https://link.springer.
                                              com/chapter/10.1007/978-94-009-
                                              0109-4_17.

Castaneda:1995:TMT

[Cas95c] Santiago Ramirez Castañeda. Three metaphysical theses on mathematical philosophy. In Mexican Studies in the History and Philosophy of Science [RC95], chapter 13, pages 201–211. ISBN 94-010-6535-7 (hardcover), 94-009-0109-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-009-
0109-4_13.

Castrodeza:1996:UEC

[Cas96]  Carlos Castrodeza.  The ultimate

epistemological consequences of the Darwinian conception. In Munévar [Mun96], chapter 8, pages 153–178. ISBN 94-010-6622-1, 94-009-0305-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-009-0305-
0_8.

Casartelli:2000:FRS

[Cas00] Mario Casartelli. Formal representation and the subjective side of scientific realism. In The Reality of the Unobservable: Observability, Unobservability and Their Impact on the Issue of Scientific Realism [AP00b], chapter 8, pages 129–135. ISBN 90-481-5458-8, 94-015-9391-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-015-
9391-5_8.

Castellanos:2014:NCP

[Cas14] Carlos Rodríguez Castellanos. Nanotechnologies in Cuba: Popularization and training. In The History of Physics in Cuba [BRW14a], chapter 18, pages 323–328. ISBN 94-017-8040-4, 94-017-8041-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC5.53; QC6.9; QC9 C9H57 2014; Q124.6-127.2. URL https://link.

springer.com/chapter/10.1007/
978-94-017-8041-4_18.

Cattaneo:1993:LAA

[Cat93] Gianpiero Cattaneo. The ‘logical’ approach to axiomatic quantum theory.

In Bridging the Gap: Philosophy, Mathematics, and Physics: Lectures on the Foundations of Science [CCG93], chapter 12, pages 225–260. ISBN 94-010-5101-1, 94-011-2496-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-011-2496-6_12.

Cattani:1995:PPN

[Cat95] Adelino Cattani. Popper, Polanyi and the notion of rationality. In Misiek [Mis95c], pages 65–74. ISBN 94-010-4206-3, 94-011-0461-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-011-0461-6_5.

Causey:1976:UTU

[Cau76] Robert L. Causey. Unified theories and unified science. In Cohen et al. [CHMV76], pages 3–13. ISBN 90-277-0647-6, 90-277-0648-4 (paperback), 94-010-1449-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic).

                                              LCCN ????  URL https://link.
                                              springer.com/chapter/10.1007/
                                              978-94-010-1449-6_1.

Causey:1980:SES

[Cau80] Robert L. Causey. Structural explanations in social science. In Scientific Discovery, Logic, and Rationality [Nic80f], chapter 16, pages 355–373. ISBN 90-277-1070-8, 94-009-8986-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-009-8986-
                                              3_16.

Cauchy:1995:STN

[Cau95] Venant Cauchy. Some thoughts on the nature of knowing. In The Concept of Knowledge: The Ankara Seminar [KC95], chapter 5, pages 49–65. ISBN 90-481-4495-7, 94-017-3263-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-017-3263-5_5.

Caws:1989:LQQ

[Caw89] Peter Caws. The law of quality and quantity, or what numbers can and can’t describe. In The Qualitative-Quantitative Distinction in the Social Sciences [GM89b], chapter 2, pages 13– 28. ISBN 90-481-8460-6, 94-017-3444-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 112. URL https://link.

springer.com/chapter/10.1007/
978-94-017-3444-8_2.

Caws:1994:TF

[Caw94] Peter Caws. Translating Feuerbach. In Artifacts, Representations and Social Practice: Essays for Marx Wartofsky [GC94], chapter 11, pages 185–193. ISBN 94-010-4390-6, 94-011-0902-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN B73 .A77

1994. URL https://link.springer.
com/chapter/10.1007/978-94-011-
0902-4_11.

Caygill:1999:NA

[Cay99]  Howard Caygill.  Nietzsche and

atomism. In Nietzsche, Theories of Knowledge, and Critical Theory: Nietzsche and the Sciences I [Bab99b], chapter 2, pages 27–36. ISBN 90-481-5233-X, 94-017-2430-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-017-2430-2_2.

Cooke:1990:SSF

[CC90] Roger Cooke and Domenico Costantini. Statistics in Science: The Foundations of Statistical Methods in Biology, Physics and Economics, volume 122 of Boston Studies in the Philosophy of Science. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 1990. ISBN 94-010-6765-1, 94-009-0619-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). 200 (est.) pp. LCCN QA276-280. URL

https://link.springer.com/book/
10.1007/978-94-009-0619-8.

Corsi:1993:BGP

[CCG93] Giovanna Corsi, Maria Luisa Dalla Chiara, and Gian Carlo Ghirardi. Bridging the Gap: Philosophy, Mathematics, and Physics: Lectures on the Foundations of Science, volume 140 of Boston Studies in the Philosophy of Science. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 1993. ISBN 94-010-5101-1, 94-011-2496-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). xi + 329 pp. LCCN ????

URL https://link.springer.com/
book/10.1007/978-94-011-2496-6.

Chiara:1993:IKN

[CD93] M. L. Dalla Chiara and G. Toraldo Di Francia. Individuals, kinds and names in physics. In Bridging the Gap: Philosophy, Mathematics, and Physics: Lectures on the Foundations of Science [CCG93], chapter 13, pages 261–283. ISBN 94-010-5101-1, 94-011-2496-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-011-2496-
6_13.

Chiara:1995:IQQ

[CdF95] Maria L. Dalla Chiara and G. Toraldo di Francia. Identity questions from quantum theory. In Gavroglu et al. [GSW95c], chapter 3, pages 39–46. ISBN 0-7923-2991-0 (set), 0-7923-2988-0, 94-017-2658-2. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 163. URL https:/

/link.springer.com/chapter/10.
1007/978-94-017-2658-0_3.

Cekic:1992:IMP

[Cek92a] Miodrag Cekic. The influence of Mach’s philosophy in Yugoslavia 1900–1941. In Ernst Mach — A Deeper Look: Documents and New Perspectives [Bla92a], chapter 19, pages 379–402. ISBN 94-010-5230-1, 94-011-2771-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN C5.53; QC6.9.

URL https://link.springer.com/
chapter/10.1007/978-94-011-2771-
4_19.

Cekic:1992:MPL

[Cek92b] Miodrag Cekic. Mach’s phenomenalism as a link between physics and psychology. In Ernst Mach — A Deeper Look: Documents and New Perspectives [Bla92a], chapter 10, pages 201–214. ISBN 94-010-5230-1, 94-011-2771-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN C5.53;

QC6.9. URL https://link.springer.
com/chapter/10.1007/978-94-011-
2771-4_10.

Cellucci:1981:PTT

[Cel81] Carlo Cellucci. Proof theory and theory of meaning. In Italian Studies in the Philosophy of Science [Chi81b], chapter 2, pages 13–29. ISBN 90-277-1073-2, 94-009-8937-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-009-8937-5_2.

Cellucci:1993:TD

[Cel93] Carlo Cellucci. Theory of deduction. In Bridging the Gap: Philosophy, Mathematics, and Physics: Lectures on the Foundations of Science [CCG93], chapter 1, pages 3–37. ISBN 94-010-5101-1, 94-011-2496-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-011-2496-6_1.

Celeyrette:2008:BRM

[Cel08] Jean Celeyrette. Bradwardine’s rule: A mathematical law? In Mechanics and Natural Philosophy Before the Scientific Revolution [LR08], chapter 2, pages 51–66. ISBN 1-4020-5966-3 (hardcover), 1-4020-5967-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QA802 .M43 2008.

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-1-4020-5967-
                                              4_2.

Cerezuelle:1986:ITT

[Cér86] Daniel Cérézuelle. Information technology and the technological system. In Philosophy and Technology II: Information Technology and Computers in Theory and Practice [MH86], chapter 15, pages 217–225. ISBN 94-010-8510-2, 94-009-4512-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-009-4512-8_15.

Cernogora:2014:WFC

[Cer14] Jacqueline Cernogora. A witness to French–Cuban cooperation in physics in the 1970s. In The History of Physics in Cuba [BRW14a], chapter 24, pages 365–379. ISBN 94-017-8040-4, 94-017-8041-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC5.53; QC6.9; QC9 C9H57 2014; Q124.6-127.2. URL https://link.

springer.com/chapter/10.1007/
978-94-017-8041-4_24.

Carrier:2011:TTC

[CF11] Martin Carrier and Patrick Finzer.

Theory and therapy: On the conceptual structure of models in medical research. In Science in the Context of Application [CN11a], chapter 6, pages 85–99. ISBN 90-481-9050-9, 90-481-9051-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q180.55.M4 S385 2011. URL https:

//link.springer.com/chapter/10.
1007/978-90-481-9051-5_6.

Cushing:1996:BMQ

[CFG96] James T. Cushing, Arthur Fine, and Sheldon Goldstein. Bohmian Mechanics and Quantum Theory: an Appraisal, volume 184 of Boston Studies in the Philosophy of Science. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 1996. ISBN 90-481-4698-4, 94-015-8715-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). viii + 408 pp. LCCN QC173.96-174.52. URL

https://link.springer.com/book/
10.1007/978-94-015-8715-0.

Cohen:1976:EMI

[CFW76] R. S. Cohen, P. K. Feyerabend, and M. W. Wartofsky. Essays in Memory of Imre Lakatos, volume 39 of Boston Studies in the Philosophy of Science. D. Reidel, Dordrecht, The Netherlands; Boston, MA, USA; Lancaster, UK; Tokyo, Japan, 1976. ISBN 90-277-0655-7, 94-010-1451-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). xii + 768 pp. LCCN ???? URL

https://link.springer.com/book/
10.1007/978-94-010-1451-9.

Costantini:1990:NFA

[CG90] Domenico Costantini and Ubaldo Garibaldi. The non frequency approach to elementary particle statistics. In Statistics in Science: The Foundations of Statistical Methods in Biology, Physics and Economics [CC90], chapter 10, pages 167–181. ISBN 94-010-6765-1, 94-009-0619-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QA276-280.

URL https://link.springer.com/
chapter/10.1007/978-94-009-0619-
8_10.

Campaner:2016:FVG

[CG16] Raffaella Campaner and Maria Carla Galavotti. Feminist versus general philosophy of science. In Meta-Philosophical Reflection on Feminist Philosophies of Science [AV16a], chapter 12, pages 203–222. ISBN 3-319-26346-3 (hardcover), 3-319-26348-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN HQ1154 .M48 2016; HQ1190 M48 2016. URL https:

//link.springer.com/chapter/10.
1007/978-3-319-26348-9_12.

Cohen:2014:FKT

[CH14] Robert S. Cohen and Ingeborg K. Helling. Felix Kaufmann’s Theory and Method in the Social Sciences, volume 303 of Boston Studies in the Philosophy and History of Science. Springer-Ver-lag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2014. ISBN 3-319-02844-8 (hardcover), 3-319-02845-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). x + 357 pp. LCCN H61 .K313 2014. URL

https://link.springer.com/book/
10.1007/978-3-319-02845-3.

Chalmers:1993:GRG

[Cha93a] Alan Chalmers. Galilean relativity and Galileo’s relativity. In Correspondence, Invariance and Heuristics: Essays in Honour of Heinz Post [FK93], chapter 10, pages 189–205. ISBN 90-481-4229-6, 94-017-1185-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-017-1185-2_10.

Chang:1993:EST

[Cha93b] Fu Chang. The evolution of science:

Theoretic recombination as a selection process. In Philosophy and Conceptual History of Science in Taiwan [LF93], chapter 6, pages 73–90. ISBN 94-010-5103-8, 94-011-2500-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q127.T28 .P455

1993. URL https://link.springer.
com/chapter/10.1007/978-94-011-
2500-0_6.

Chalmers:2009:SAP

[Cha09a] A. F. (Alan Francis) Chalmers. The Scientist’s Atom and the Philosopher’s Stone: How Science Succeeded and Philosophy Failed to Gain Knowledge of Atoms, volume 279 of Boston Studies in the Philosophy of Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2009. ISBN 90-481-2361-5, 90-481-2362-3. ISSN 0068-0346 (print), 2214-7942 (electronic). xii + 287 pp. LCCN QD461 .C375 2009. URL

https://link.springer.com/book/
10.1007/978-90-481-2362-9.

Chalmers:2009:AEM

[Cha09b] Alan Chalmers. Atomism, experiment and the mechanical philosophy: The work of Robert Boyle. In The Scientist’s Atom and the Philosopher’s Stone: How Science Succeeded and Philosophy Failed to Gain Knowledge of Atoms [Cha09a], chapter 6, pages 97–122. ISBN 90-481-2361-5, 90-481-2362-3. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QD461 .C375 2009.

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-90-481-2362-
                                              9_6.

Chalmers:2009:AAG

[Cha09c] Alan Chalmers. Atomism in its Ancient Greek perspective. In The Scientist’s Atom and the Philosopher’s Stone: How Science Succeeded and Philosophy Failed to Gain Knowledge of Atoms [Cha09a], chapter 4, pages 59–74. ISBN 90-481-2361-5, 90-481-2362-3. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QD461 .C375 2009.

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-90-481-2362-
                                              9_4.

Chalmers:2009:ASP

[Cha09d] Alan Chalmers. Atomism: Science or philosophy? In The Scientist’s Atom and the Philosopher’s Stone: How Science Succeeded and Philosophy Failed to Gain Knowledge of Atoms [Cha09a], chapter 1, pages 1–17. ISBN 90-481-2361-5, 90-481-2362-3. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QD461 .C375 2009. URL https:/

                                              /link.springer.com/chapter/10.
                                              1007/978-90-481-2362-9_1.

Chalmers:2009:AV

[Cha09e] Alan Chalmers. Atomism vindicated? In The Scientist’s Atom and the Philosopher’s Stone: How Science Succeeded and Philosophy Failed to Gain Knowledge of Atoms [Cha09a], chapter 14, pages 261–268. ISBN 90-481-2361-5, 90-481-2362-3. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QD461 .C375 2009.

URL https://link.springer.com/
chapter/10.1007/978-90-481-2362-
9_14.

Chalmers:2009:DAC

[Cha09f] Alan Chalmers. Dalton’s atomism and its creative modification via chemical formulae. In The Scientist’s Atom and the Philosopher’s Stone: How Science Succeeded and Philosophy Failed to Gain Knowledge of Atoms [Cha09a], chapter 9, pages 173–198. ISBN 90-481-2361-5, 90-481-2362-3. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QD461 .C375 2009. URL https:/

/link.springer.com/chapter/10.
1007/978-90-481-2362-9_9.

Chalmers:2009:DA

[Cha09g] Alan Chalmers. Democritean atomism. In The Scientist’s Atom and the Philosopher’s Stone: How Science Succeeded and Philosophy Failed to Gain Knowledge of Atoms [Cha09a], chapter 2, pages 19–41. ISBN 90-481-2361-5, 90-481-2362-3. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QD461 .C375 2009. URL https:/

/link.springer.com/chapter/10.
1007/978-90-481-2362-9_2.

Chalmers:2009:EMC

[Cha09h] Alan Chalmers. The emergence of modern chemistry with no debt to atomism. In The Scientist’s Atom and the Philosopher’s Stone: How Science Succeeded and Philosophy Failed to Gain Knowledge of Atoms [Cha09a], chapter 8, pages 139–171. ISBN 90-481-2361-5, 90-481-2362-3. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QD461 .C375 2009. URL https:/

                                              /link.springer.com/chapter/10.
                                              1007/978-90-481-2362-9_8.

Chalmers:2009:ECE

[Cha09i] Alan Chalmers. Experimental contact with electrons. In The Scientist’s Atom and the Philosopher’s Stone: How Science Succeeded and Philosophy Failed to Gain Knowledge of Atoms [Cha09a], chapter 13, pages 247–260. ISBN 90-481-2361-5, 90-481-2362-3. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QD461 .C375 2009.

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-90-481-2362-
                                              9_13.

Chalmers:2009:ECM

[Cha09j] Alan Chalmers. Experimental contact with molecules. In The Scientist’s Atom and the Philosopher’s Stone: How Science Succeeded and Philosophy Failed to Gain Knowledge of Atoms [Cha09a], chapter 12, pages 233–246. ISBN 90-481-2361-5, 90-481-2362-3. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QD461 .C375 2009.

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-90-481-2362-
                                              9_12.

Chalmers:2009:ACO

[Cha09k] Alan Chalmers. From Avogadro to Cannizzaro: The old story. In The Scientist’s Atom and the Philosopher’s Stone: How Science Succeeded and Philosophy Failed to Gain Knowledge of Atoms [Cha09a], chapter 10, pages 199–213. ISBN 90-481-2361-5, 90-481-2362-3. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QD461

.C375 2009. URL https://link.
springer.com/chapter/10.1007/
978-90-481-2362-9_10.

Chalmers:2009:AGD

[Cha09l] Alan Chalmers. From the Ancient Greeks to the dawn of science. In The Scientist’s Atom and the Philosopher’s Stone: How Science Succeeded and Philosophy Failed to Gain Knowledge of Atoms [Cha09a], chapter 5, pages 75– 95. ISBN 90-481-2361-5, 90-481-2362-3. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QD461 .C375 2009.

URL https://link.springer.com/
chapter/10.1007/978-90-481-2362-
9_5.

Chalmers:2009:HDE

[Cha09m] Alan Chalmers. How does Epicurus’s garden grow? In The Scientist’s Atom and the Philosopher’s Stone: How Science Succeeded and Philosophy Failed to Gain Knowledge of Atoms [Cha09a], chapter 3, pages 43–58. ISBN 90-481-2361-5, 90-481-2362-3. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QD461 .C375 2009. URL https:/

/link.springer.com/chapter/10.
1007/978-90-481-2362-9_3.

Chalmers:2009:NAF

[Cha09n] Alan Chalmers. Newton’s atomism and its fate. In The Scientist’s Atom and the Philosopher’s Stone: How Science Succeeded and Philosophy Failed to Gain Knowledge of Atoms [Cha09a], chapter 7, pages 123–138. ISBN 90-481-2361-5, 90-481-2362-3. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QD461 .C375 2009. URL https:/

/link.springer.com/chapter/10.
1007/978-90-481-2362-9_7.

Chalmers:2009:TKT

[Cha09o] Alan Chalmers. Thermodynamics and the kinetic theory. In The Scientist’s Atom and the Philosopher’s Stone: How Science Succeeded and Philosophy Failed to Gain Knowledge of Atoms [Cha09a], chapter 11, pages 215–231. ISBN 90-481-2361-5, 90-481-2362-3. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QD461 .C375 2009.

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-90-481-2362-
                                              9_11.

Chakravartty:2010:TRS

[Cha10]  Anjan Chakravartty.  Truth and

representation in science: Two inspirations from art. In Frigg and Hunter [FH10b], chapter 3, pages 33–50. ISBN 90-481-3850-7 (hardcover), 90-481-3851-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN BH301.M47 B49 2010; Q175 .B52 2010.

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-90-481-3851-
                                              7_3.

Chang:2012:ARR

[Cha12a] Hasok Chang. Active realism and the reality of h 2 o. In Is Water H_{2}O?: Evidence, Realism and Pluralism [Cha12e], chapter 4, pages 203–251. ISBN 94-007-3931-1, 94-007-3932-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QD169.W3 C43

                                              2012. URL https://link.springer.
                                              com/chapter/10.1007/978-94-007-
                                              3932-1_4.

Chang:2012:BCS

[Cha12b] Hasok Chang. Beyond case-studies: History as philosophy. In Mauskopf and Schmaltz [MS12b], chapter 8, pages 109–124. ISBN 94-007-1744-X, 94-007-1745-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175 .I5134 2011. URL https:

//link.springer.com/chapter/10.
1007/978-94-007-1745-9_8.

Chang:2012:EPC

[Cha12c] Hasok Chang. Electrolysis: Piles of confusion and poles of attraction. In Is Water H_{2}O?: Evidence, Realism and Pluralism [Cha12e], chapter 2, pages 71–131. ISBN 94-007-3931-1, 94-007-3932-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QD169.W3 C43 2012. URL https:

//link.springer.com/chapter/10.
1007/978-94-007-3932-1_2.

Chang:2012:HHC

[Cha12d] Hasok Chang. HO or H_{2}O? How chemists learned to count atoms. In Is Water H_{2}O?: Evidence, Realism and Pluralism [Cha12e], chapter 3, pages 133–201. ISBN 94-007-3931-1, 94-007-3932-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QD169.W3 C43 2012. URL https:

//link.springer.com/chapter/10.
1007/978-94-007-3932-1_3.

Chang:2012:WHE

[Cha12e] Hasok Chang. Is Water H_{2}O?: Evidence, Realism and Pluralism, volume 293 of Boston Studies in the Philosophy and History of Science. Springer-Ver-lag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2012. ISBN 94-007-3931-1, 94-007-3932-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). xxi + 316 pp. LCCN QD169.W3 C43 2012. URL

https://link.springer.com/book/
10.1007/978-94-007-3932-1.

Chang:2012:PSC

[Cha12f] Hasok Chang. Pluralism in science: A call to action. In Is Water H_{2}O?: Evidence, Realism and Pluralism [Cha12e], chapter 5, pages 253–301. ISBN 94-007-3931-1, 94-007-3932-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QD169.W3 C43

                                              2012. URL https://link.springer.
                                              com/chapter/10.1007/978-94-007-
                                              3932-1_5.

Chang:2012:WCR

[Cha12g] Hasok Chang. Water and the chemical revolution. In Is Water H_{2}O?: Evidence, Realism and Pluralism [Cha12e], chapter 1, pages 1–70. ISBN 94-007-3931-1, 94-007-3932-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QD169.W3 C43

                                              2012. URL https://link.springer.
                                              com/chapter/10.1007/978-94-007-
                                              3932-1_1.

Chartier:2014:TCR

[Cha14] Anne-Marie Chartier. The teaching context and reading from the 16th to the 19th centuries: The role of the memorization of texts in learning. In Bernard and Proust [BP14b], chapter 2, pages 19–40. ISBN 94-007-5121-4 (hardcover), 94-007-5122-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q181 .S35 2014; Q181 S387 2014. URL https://link.

                                              springer.com/chapter/10.1007/
                                              978-94-007-5122-4_2.

Chang:2015:RRB

[Cha15] Hasok Chang. Reductionism and the relation between chemistry and physics. In Arabatzis et al. [ARS15], chapter 13, pages 193–209. ISBN 3-319-14552-5 (paperback), 3-319-14553-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q125 .R45 2015.

URL https://link.springer.com/
chapter/10.1007/978-3-319-14553-
2_13.

Champion:2017:JAI

[Cha17] Rafe Champion. Joe Agassi, intellectual irritant. In Encouraging Openness: Essays for Joseph Agassi on the Occasion of His 90th Birthday [BAG17], chapter 39, pages 511–528. ISBN 3-319-57668-2 (hardcover), 3-319-57669-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN B833.E53

2017. URL https://link.springer.
com/chapter/10.1007/978-3-319-
57669-5_39.

Charrak:2018:ECS

[Cha18] André Charrak. Empiricism and certainty in science: The Franco–Berlin School of Empiricism. In What Does it Mean to be an Empiricist?: Empiricisms in Eighteenth Century Sciences [BR18b], chapter 10, pages 189–200. ISBN 3-319-69858-3 (hardcover), 3-319-69860-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN B816 .W43 2018. URL https://link.

springer.com/chapter/10.1007/
978-3-319-69860-1_10.

Chevalley:1994:NBW

[Che94] Catherine Chevalley. Niels Bohr’s words and the Atlantis of Kantianism. In Faye and Folse [FF94], chapter 2, pages 33–55. ISBN 0-7923-2378-5 (hardcover), 94-015-8106-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 153 1994; QC16 1994. URL https:

//link.springer.com/chapter/10.
1007/978-94-015-8106-6_2.

Chevalley:1995:PBQ

[Che95] Catherine Chevalley. Philosophy and the birth of quantum theory. In Gavroglu et al. [GSW95c], chapter 2, pages 11–37. ISBN 0-7923-2991-0 (set), 0-7923-2988-0, 94-017-2658-2. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 163.

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-017-2658-
                                              0_2.

Chevalley:1997:PAG

[Che97] Catherine Chevalley. Physics as an art: The German tradition and the symbolic turn in philosophy, history of art and natural science in the 1920s. In The Elusive Synthesis: Aesthetics and Science [Tau97a], chapter 11, pages 227–249. ISBN 0-7923-4763-3, 0-7923-3904-5, 94-009-1786-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-009-1786-6_11.

Chemla:2005:HSH

[Che05a] Karine Chemla. History of Science, History of Text, volume 238 of Boston Studies in the Philosophy of Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2005. ISBN 1-281-33589-4, 1-4020-2320-0 (hardcover), 1-4020-2321-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). xxvii + 254 pp. LCCN Q174 .B67 vol. 238

                                              Q125. URL https://link.springer.
                                              com/book/10.1007/1-4020-2321-9.

Chemla:2005:WCB

[Che05b] Karine Chemla. What is the content of this book? A plea for developing history of science and history of text conjointly. In History of Science, History of Text [Che05a], chapter 10, pages 201–230. ISBN 1-281-33589-4, 1-4020-2320-0 (hardcover), 1-4020-2321-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 238

Q125. URL https://link.springer.
com/chapter/10.1007/1-4020-2321-
9_10.

Chemla:2010:CCM

[Che10] Karine Chemla. A Chinese canon in mathematics and its two layers of commentaries: Reading a collection of texts as shaped by actors. In Looking at it from Asia: the Processes that Shaped the Sources of History of Science [BE10b], chapter 6, pages 169–210. ISBN 90-481-3675-X (hardcover), 90-481-3676-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q127.A65 L66 2010. URL https:/

/link.springer.com/chapter/10.
1007/978-90-481-3676-6_6.

Chediak:2011:EAC

[Che11] Karla Chediak. The etiological approach to the concept of biological function. In Brazilian Studies in Philosophy and History of Science: An account of recent works [KV11], chapter 11, pages 161–169. ISBN 90-481-9421-0 (hardcover), 90-481-9422-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175.3 .B73

2011. URL https://link.springer.
com/chapter/10.1007/978-90-481-
9422-3_11.

Chemla:2014:SHS

[Che14] Karine Chemla. On the sources of the historian of science from the perspective of a history of education. In Bernard and Proust [BP14b], chapter 12, pages 305–324. ISBN 94-007-5121-4 (hardcover), 94-007-5122-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q181 .S35 2014; Q181 S387 2014. URL https://link.

                                              springer.com/chapter/10.1007/
                                              978-94-007-5122-4_12.

Chiara:1981:TLE

[Chi81a] Maria Luisa Dalla Chiara. Is there a logic of empirical sciences? In Italian Studies in the Philosophy of Science [Chi81b], chapter 12, pages 187–196. ISBN 90-277-1073-2, 94-009-8937-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-009-8937-
                                              5_12.

Chiara:1981:ISP

[Chi81b] Maria Luisa Dalla Chiara. Italian Studies in the Philosophy of Science, volume 47 of Boston Studies in the Philosophy of Science. D. Reidel, Dordrecht, The Netherlands; Boston, MA, USA; Lancaster, UK; Tokyo, Japan, 1981. ISBN 90-277-1073-2, 94-009-8937-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). xi + 529 pp. LCCN ???? URL

                                              https://link.springer.com/book/
                                              10.1007/978-94-009-8937-5.

Chiara:1983:SLP

[Chi83] Maria Luisa Dalla Chiara. Some logical problems suggested by empirical theories. In Language, Logic and Method [CW83b], chapter 2, pages 75–90. ISBN 94-009-7704-2, 94-009-7702-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-009-7702-0_2.

Chiariello:2017:CCM

[Chi17] Michael Chiariello. Comprehensively critical metapolitics. In Encouraging Openness: Essays for Joseph Agassi on the Occasion of His 90th Birthday [BAG17], chapter 5, pages 41–52. ISBN 3-319-57668-2 (hardcover), 3-319-57669-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN B833.E53 2017. URL https://link.

springer.com/chapter/10.1007/
978-3-319-57669-5_5.

Chmielecka:1982:CDC

[Chm82] Ewa Chmielecka. The context of discovery and the context of justification: A reappraisal. In Polish Essays in the Philosophy of the Natural Sciences [Kra82c], chapter 6, pages 63–74. ISBN 94-009-7705-0 (e-book), 90-277-1287-5. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-009-7705-1_6.

Cohen:1976:PPB

[CHMV76] R. S. Cohen, C. A. Hooker, A. C.

Michalos, and J. W. Van Evra, editors. PSA 1974: Proceedings of the 1974 Biennial Meeting Philosophy of Science Association, volume 32 of Boston Studies in the Philosophy of Science. D. Reidel, Dordrecht, The Netherlands; Boston, MA, USA; Lancaster, UK; Tokyo, Japan, 1976. ISBN 90-277-0647-6, 90-277-0648-4 (paperback), 94-010-1449-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL

https://link.springer.com/book/
10.1007/978-94-010-1449-6.

Chomsky:1963:PL

[Cho63] Noam Chomsky. Perception and language. In Wartofsky [War63], pages 199–205. ISBN 94-010-3265-3, 94-010-3263-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B678 1970. URL https://link.

                                              springer.com/chapter/10.1007/
                                              978-94-010-3263-6_16.

Chomsky:1967:RCT

[Cho67] Noam Chomsky. Recent contributions to the theory of innate ideas. In Cohen and Wartofsky [CW67], pages 81–90. ISBN 90-277-0013-3, 94-010-3508-3. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175 .B731 v.3.

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-010-3508-
                                              8_4.

Chomsky:1976:PMS

[Cho76] Noam Chomsky. Problems and mysteries in the study of human language. In Language in Focus: Foundations, Methods and Systems: Essays in Memory of Yehoshua Bar-Hillel [Kas76b], chapter 15, pages 281–357. ISBN 90-277-0644-1, 90-277-0645-X, 94-010-1876-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN P101-120. URL https://link.

                                              springer.com/chapter/10.1007/
                                              978-94-010-1876-0_15.

Chomsky:1986:RCT

[Cho86] Noam Chomsky. Recent contributions to the theory of innate ideas. In Cohen and Wartofsky [CW85], chapter 3, pages 31–40. ISBN 90-277-1971-3, 94-009-5345-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

                                              ???? URL https://link.springer.
                                              com/chapter/10.1007/978-94-009-
                                              5345-1_3.

Chompalov:2014:LLS

[Cho14] Ivan Chompalov. Lessons learned from the study of multi-organizational collaborations in science and implications for the role of the university in the 21st century. In The Institution of Science and the Science of Institutions: The Legacy of Joseph Ben-David [Her14b], chapter 9, pages 167–184. ISBN 94-007-7406-0 (hardcover), 94-007-7407-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 v.302; Q125 .I58 2014. URL https:

//link.springer.com/chapter/10.
1007/978-94-007-7407-0_9.

Christianidis:1994:HIP

[Chr94a] Jean Christianidis. On the history of indeterminate problems of the first degree in Greek mathematics. In Trends in the Historiography of Science [GCN94], chapter 18, pages 237–247. ISBN 90-481-4264-4, 94-017-3596-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-017-3596-
4_18.

Christmann:1994:LMP

[Chr94b] Hans Helmut Christmann. Linguistics and modern philology in Germany 1800–1840 as ‘scientific’ subjects and as university disciplines. In Poggi and Bossi [PB94], chapter 13, pages 203–214. ISBN 0-7923-2336-X, 94-017-2921-2. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 152. URL https://link.

springer.com/chapter/10.1007/
978-94-017-2921-5_13.

Cohen:1996:RAR

[CHR96] Robert S. Cohen, Risto Hilpinen, and Qiu Renzong, editors. Realism and Anti-Realism in the Philosophy of Science: Beijing International Conference, 1992, volume 169 of Boston Studies in the Philosophy of Science. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 1996. ISBN 90-481-4493-0, 94-015-8638-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

                                              URL https://link.springer.com/
                                              book/10.1007/978-94-015-8638-2.

Christianidis:2004:CHG

[Chr04a] Jean Christianidis. Classics in the History of Greek Mathematics, volume 240 of Boston Studies in the Philosophy of Science. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 2004. ISBN 1-4020-0081-2 (hardcover), 1-4020-2640-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). x + 461 pp. LCCN QA22 .C596 2004. URL

                                              https://link.springer.com/book/
                                              10.1007/978-1-4020-2640-9.

Christianidis:2004:I

[Chr04b] Jean Christianidis. Introduction. In Classics in the History of Greek Mathematics [Chr04a], chapter 18, pages 331–336. ISBN 1-4020-0081-2 (hardcover), 1-4020-2640-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QA22 .C596 2004.

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-1-4020-2640-
                                              9_18.

Christensen:2007:OCF

[Chr07] Dan Charly Christensen. Ørsted’s concept of force and theory of music. In Hans Christian Ørsted and The Romantic Legacy In Science: Ideas, Disciplines, Practices [BCK07], chapter 7, pages 115–133. ISBN 1-4020-2979-9 (hardcover), 1-4020-2987-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC515.O37 H36 2007. URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-1-4020-2987-5_7.

Chrudzimski:2013:GEF

[Chr11] Arkadiusz Chrudzimski. Gestalt, equivalency, and functional dependency: Kurt Grelling’s formal ontology. In The Berlin Group and the Philosophy of Logical Empiricism [MP13], chapter 12, pages 245–261. ISBN 94-007-5484-1 (hardcover), 94-007-5485-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175.32.A24 .M384 2011; Q175 .B47 2013. URL https:

//link.springer.com/chapter/10.
1007/978-94-007-5485-0_12.

Christopoulou:2014:EMH

[Chr14] Christiana Christopoulou. Early modern history of cold: Robert Boyle and the emergence of a new experimental field in Seventeenth Century experimental philosophy. In History of Artificial Cold, Scientific, Technological and Cultural Issues [Gav14b], chapter 2, pages 21–51. ISBN 94-007-7198-3 (hardcover), 94-007-7199-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC278 H57 2014; TP494 .H577 2014. URL https:/

/link.springer.com/chapter/10.
1007/978-94-007-7199-4_2.

Christianidis:2015:MHD

[Chr15] Jean Christianidis. The meaning of hypostasis in Diophantus’ Arithmetica. In Arabatzis et al. [ARS15], chapter 21, pages 315–327. ISBN 3-319-14552-5 (paperback), 3-319-14553-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q125 .R45 2015.

URL https://link.springer.com/
chapter/10.1007/978-3-319-14553-
2_21.

Cohen:1997:EMQ

[CHS97a] Robert S. Cohen, Michael Horne, and John Stachel, editors. Experimental Metaphysics: Quantum Mechanical Studies for Abner Shimony. Volume One, volume 193 of Boston Studies in the Philosophy of Science. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 1997. ISBN 0-7923-4452-9 (volume 1), 0-7923-4454-5 (set). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol.

193. URL http://www.lehmanns.ch/
shop/naturwissenschaften/422840-
9780792344520-experimental-
metaphysics;   http://www.loc.
gov/catdir/enhancements/fy0822/
97002399-d.html;   http://www.
loc.gov/catdir/enhancements/
fy0822/97002399-t.html.

Cohen:1997:PEP

[CHS97b] Robert S. Cohen, Michael Horne, and John Stachel, editors. Potentiality, Entanglement and Passion-at-a-Distance: Quantum Mechanical Studies for Abner Shimony. Volume Two, volume 194 of Boston Studies in the Philosophy of Science. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 1997. ISBN 90-481-4809-X, 94-017-2732-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC173.96-174.52.

URL https://link.springer.com/
book/10.1007/978-94-017-2732-7.

Churchland:1991:DUS

[Chu91] Paul M. Churchland. A deeper unity: Some Feyerabendian themes in neurocomputational form. In Beyond Reason: Essays on the Philosophy of Paul Feyerabend [Mun91a], chapter 1, pages 1–23. ISBN 94-010-5406-1, 94-011-3188-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.

com/chapter/10.1007/978-94-011-
3188-9_1.

Churchland:1995:NEP

[Chu95] Paul M. Churchland. On the nature of explanation: A PDP approach. In Misiek [Mis95c], pages 75–108. ISBN 94-010-4206-3, 94-011-0461-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-011-0461-6_6.

Churchland:2000:NAH

[Chu00] Paul M. Churchland. A new approach to human cognition and its significance for the philosophy of science. In The Reality of the Unobservable: Observability, Unobservability and Their Impact on the Issue of Scientific Realism [AP00b], chapter 4, pages 59–85. ISBN 90-481-5458-8, 94-015-9391-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-015-9391-5_4.

Churchill:2001:AWF

[Chu01] Frederick B. Churchill. August Weismann and Ferdinand Dickel: Testing the Dzierzon system. In Allen and MacLeod [AM01b], chapter 4, pages 61–76. ISBN 1-4020-0495-8 (hardcover), 94-017-2956-5. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 228; Q175.55. URL https:

//link.springer.com/chapter/10.
1007/978-94-017-2956-7_4.

Chu:2011:ARZ

[Chu11] Pingyi Chu. Adoption and resistance:

Zhang Yongjing and Ancient Chinese calendrical methods. In Science between Europe and Asia: Historical Studies on the Transmission, Adoption and Adaptation of Knowledge [GR11b], chapter 10, pages 151–161. ISBN 90-481-9968-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q125 .S354 2010; Q175.52.A78 S35

2010. URL https://link.springer.
com/chapter/10.1007/978-90-481-
9968-6_10.

Cifoletti:2005:AAD

[Cif05] Giovanna C. Cifoletti. The algebraic art of discourse algebraic dispositio, invention and imitation in Sixteenth-Century France. In History of Science, History of Text [Che05a], chapter 6, pages 123–135. ISBN 1-281-33589-4, 1-4020-2320-0 (hardcover), 1-4020-2321-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 238 Q125. URL https:/

/link.springer.com/chapter/10.
1007/1-4020-2321-9_6.

Cifoletti:2014:MPM

[Cif14] Giovanna C. Cifoletti. Mathematical progress or mathematical teaching? Bilingualism and printing in European Renaissance mathematics. In Bernard and Proust [BP14b], chapter 8, pages 187–214. ISBN 94-007-5121-4 (hardcover), 94-007-5122-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q181 .S35 2014; Q181

S387 2014.  URL https://link.
springer.com/chapter/10.1007/
978-94-007-5122-4_8.

Cini:1989:CDD

[Cin89] Marcello Cini. Continuity and discontinuity in the definition of a disciplinary field: The case of XXth century physics. In Imre Lakatos and Theories of Scientific Change [GGN89], chapter 7, pages 83–94. ISBN 94-010-7860-2, 94-009-3025-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-009-
3025-4_7.

Chernyak:1997:KAE

[CK97] Leon Chernyak and David Kazhdan.

Kant and the aesthetic–expressive vision of mathematics. In The Elusive Synthesis: Aesthetics and Science [Tau97a], chapter 10, pages 203–225. ISBN 0-7923-4763-3, 0-7923-3904-5, 94-009-1786-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-009-
1786-6_10.

Cohen:1983:PPP

[CL83] R. S. Cohen and L. Laudan, editors.

Physics, Philosophy and Psychoanalysis: Essays in Honour of Adolf Grünbaum, volume 76 of Boston Studies in the Philosophy of Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1983. ISBN 94-009-7055-2, 94-009-7057-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). xviii + 339 pp. LCCN ???? URL

https://link.springer.com/book/
10.1007/978-94-009-7055-7.

Cassinelli:1993:SCS

[CL93] Gianni Cassinelli and Pekka J. Lahti.

Sigma-convex structures of the sets of states and probability measures in quantum mechanics. In Bridging the Gap: Philosophy, Mathematics, and Physics: Lectures on the Foundations of Science [CCG93], chapter 11, pages 211–224. ISBN 94-010-5101-1, 94-011-2496-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-011-
2496-6_11.

Cohen:2000:MS

[CL00] Robert S. Cohen and Hillel Levine. Maimonides and the Sciences, volume 211 of Boston Studies in the Philosophy of Science. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 2000. ISBN 90-481-5348-4, 94-017-2128-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). xv + 256 pp. LCCN ????

                                              URL https://link.springer.com/
                                              book/10.1007/978-94-017-2128-8.

Copeland:2017:THC

[CL17] B. Jack Copeland and Jason Long. Turing and the history of computer music. In Floyd and Bokulich [FB17], chapter 8, pages 189–218. ISBN 3-319-53278-2, 3-319-53280-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QA29.T8 P45 2017.

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-3-319-53280-
                                              6_8.

Clark:1989:DPR

[Cla89] Peter Clark. Determinism, probability and randomness in classical statistical physics. In Imre Lakatos and Theories of Scientific Change [GGN89], chapter 8, pages 95–110. ISBN 94-010-7860-2, 94-009-3025-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-009-3025-4_8.

Clarke:1996:NSI

[Cla96] Murray Clarke. Natural selection and indexical representation. In Québec Studies in the Philosophy of Science: Part II: Biology, Psychology, Cognitive Science and Economics Essays in Honor of Hugues Leblanc [MC96], chapter 4, pages 53–62. ISBN 94-010-6537-3, 94-009-0113-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-009-0113-1_4.

Clarke:2002:CI

[Cla02] W. Norris Clarke. The creative imagination. In Hermeneutic Philosophy of Science, Van Gogh’s Eyes, and God: Essays in Honor of Patrick A. Heelan, S.J. [Bab02c], chapter 37, pages 423–428. ISBN 90-481-5926-1, 94-017-1767-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 225. URL https:/

/link.springer.com/chapter/10.
1007/978-94-017-1767-0_37.

Clarke:2008:SOP

[Cla08] Steve Clarke. Some optimism for the pessimist. In Rethinking Scientific Change and Theory Comparison: Stabilities, Ruptures, Incommensurabilities? [SSHH08], chapter 13, pages 203–206. ISBN 1-4020-6274-5 (hardcover), 1-4020-6279-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175 .R397 2008.

URL https://link.springer.com/
chapter/10.1007/978-1-4020-6279-
7_13.

Clancier:2010:FAC

[Cla10] Philippe Clancier. Formation and administration of the collections of literary and scholarly tablets in First Millennium Babylonia. In Looking at it from Asia: the Processes that Shaped the Sources of History of Science [BE10b], chapter 1, pages 3–35. ISBN 90-481-3675-X (hardcover), 90-481-3676-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q127.A65 L66 2010. URL https:/

                                              /link.springer.com/chapter/10.
                                              1007/978-90-481-3676-6_1.

Clavien:2012:AEM

[Cla12] Christine Clavien. Altruistic emotional motivation: An argument in favour of psychological altruism. In Philosophy of Behavioral Biology [PR12b], chapter 13, pages 275–296. ISBN 94-007-1950-7, 94-007-1951-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QH331 .P4679 2012; Q175 .B657 v.282. URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-007-1951-4_13.

Clancier:2014:TLM

[Cla14]  Philippe Clancier.  Teaching and

learning medicine and exorcism at Uruk during the Hellenistic period. In Bernard and Proust [BP14b], chapter 3, pages 41–66. ISBN 94-007-5121-4 (hardcover), 94-007-5122-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q181 .S35 2014; Q181 S387 2014. URL https://link.

                                              springer.com/chapter/10.1007/
                                              978-94-007-5122-4_3.

Cleary:2002:AAP

[Cle02] John J. Cleary. Abstracting Aristotle’s philosophy of mathematics. In Hermeneutic Philosophy of Science, Van Gogh’s Eyes, and God: Essays in Honor of Patrick A. Heelan, S.J. [Bab02c], chapter 13, pages 163–176. ISBN 90-481-5926-1, 94-017-1767-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 225. URL https://link.

                                              springer.com/chapter/10.1007/
                                              978-94-017-1767-0_13.

Chatelin:1992:PIF

[CLR^{+}92] Yvon Chatelin, José Leite Lopes, Abdur Rahman, Shigeru Nakayama, Juan-José Saldaña, Jean-Jacques Salomon, and José Israël Vargas. Part IV: Final round table: Science et développement: une politique scientifique peut-elle tirer un enseignement de l’histoire des sciences?. (French) [Science and development: Can science policy learn from the history of science?]. In Science and Empires: Historical Studies about Scientific Development and European Expansion [PJM92], chapter 35, pages 365–392. ISBN 0-7923-1518-9 (hardcover), 94-010-5145-3, 94-011-2594-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 136. URL https://link.

springer.com/chapter/10.1007/
978-94-011-2594-9_35.

Clunas:2005:TRT

[Clu05] Craig Clunas. Text, representation and technique in Early Modern China. In History of Science, History of Text [Che05a], chapter 5, pages 107–121. ISBN 1-281-33589-4, 1-4020-2320-0 (hardcover), 1-4020-2321-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 238

Q125. URL https://link.springer.
com/chapter/10.1007/1-4020-2321-
9_5.

Christides:1989:RAT

[CM89] T. M. Christides and M. Mikou. The relative autonomy of theoretical science and the role of crucial experiments in the development of superconductivity theory. In Imre Lakatos and Theories of Scientific Change [GGN89], chapter 12, pages 147–153. ISBN 94-010-7860-2, 94-009-3025-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

                                              ???? URL https://link.springer.
                                              com/chapter/10.1007/978-94-009-
                                              3025-4_12.

Carrier:2011:SCAb

[CN11a] Martin Carrier and Alfred Nordmann. Science in the Context of Application, volume 274 of Boston Studies in the Philosophy of Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2011. ISBN 90-481-9050-9, 90-481-9051-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). x + 492 pp. LCCN Q180.55.M4 S385 2011. URL http://

                                              www.loc.gov/catdir/enhancements/
                                              fy1114/2010934513-d.html;
                                              http://www.loc.gov/catdir/toc/
                                              fy11pdf03/2010934513.html;
                                              https://link.springer.com/book/
                                              10.1007/978-90-481-9051-5.

Carrier:2011:SCAa

[CN11b] Martin Carrier and Alfred Nordmann. Science in the context of application: Methodological change, conceptual transformation, cultural reorientation. In Science in the Context of Application [CN11a], chapter 1, pages 1–7. ISBN 90-481-9050-9, 90-481-9051-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q180.55.M4 S385

                                              2011. URL https://link.springer.
                                              com/chapter/10.1007/978-90-481-
                                              9051-5_1.

Coburn:1974:WPN

[Cob74] N. Coburn. Wave propagation in non-viscous fluids. In For Dirk Struik: Scientific, Historical and Political Essays in Honor of Dirk J. Struik [CSW74], chapter 3, pages 21–24. ISBN 90-277-0379-5, 94-010-2115-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-010-2115-9_3.

Coffa:1974:RK

[Cof74] J. Alberto Coffa. Randomness and knowledge. In Schaffner and Cohen [SC74a], pages 103–115. ISBN 90-277-0408-2, 90-277-0409-0, 94-010-2140-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175 .B7312.

URL https://link.springer.com/
chapter/10.1007/978-94-010-2140-
1_8.

Coffa:1983:GSE

[Cof83] Alberto Coffa. Geometry and semantics: An examination of Putnam’s philosophy of geometry. In Cohen and Laudan [CL83], chapter 1, pages 1–30. ISBN 94-009-7055-2, 94-009-7057-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-009-7055-
7_1.

Calinski:1990:SAS

[COG90] T. Caliński, E. Ottaviano, and M. Sari Gorla. A statistical approach to the study of pollen fitness. In Statistics in Science: The Foundations of Statistical Methods in Biology, Physics and Economics [CC90], chapter 6, pages 89–101. ISBN 94-010-6765-1, 94-009-0619-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QA276-280. URL https:/

/link.springer.com/chapter/10.
1007/978-94-009-0619-8_6.

Cohen:1963:C

[Coh63] Robert S. Cohen. Comments. In Wartofsky [War63], pages 196–198. ISBN 94-010-3265-3, 94-010-3263-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174

                                              .B678 1970. URL https://link.
                                              springer.com/chapter/10.1007/
                                              978-94-010-3263-6_15.

Cohen:1969:HNP

[Coh69] I. Bernard Cohen. Hypotheses in Newton’s philosophy. In Cohen and Wartofsky [CW69a], pages 304–326. ISBN 94-010-3383-8, 94-010-3263-7, 94-010-3381-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

                                              ???? URL https://link.springer.
                                              com/chapter/10.1007/978-94-010-
                                              3381-7_8.

Cohen:1970:EMPa

[Coh70] Robert S. Cohen. Ernst Mach: Physics, perception and the philosophy of science. In Ernst Mach: Physicist and Philosopher [CS70], chapter 8, pages 126–164. ISBN 90-481-8318-9, 90-277-0016-8, 94-017-1462-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-017-1462-4_8.

Cohen:1974:INC

[Coh74a] I. Bernard Cohen. Isaac Newton, the calculus of variations, and the design of ships. In For Dirk Struik: Scientific, Historical and Political Essays in Honor of Dirk J. Struik [CSW74], chapter 14, pages 169–187. ISBN 90-277-0379-5, 94-010-2115-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-010-2115-9_14.

Cohen:1974:ES

[Coh74b] Robert S. Cohen. Ethics and science. In For Dirk Struik: Scientific, Historical and Political Essays in Honor of Dirk J. Struik [CSW74], chapter 25, pages 307–323. ISBN 90-277-0379-5, 94-010-2115-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-010-
2115-9_25.

Cohen:1976:WWC

[Coh76a] I. Bernard Cohen. William Whewell and the concept of scientific revolution. In Essays in Memory of Imre Lakatos [CFW76], chapter 6, pages 55–63. ISBN 90-277-0655-7, 94-010-1451-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-010-1451-9_6.

Cohen:1976:HCO

[Coh76b] L. Jonathan Cohen. How can one testimony corroborate another? In Essays in Memory of Imre Lakatos [CFW76], chapter 7, pages 65–78. ISBN 90-277-0655-7, 94-010-1451-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-010-1451-
9_7.

Cohen:1976:HEC

[Coh76c] L. Jonathan Cohen. How empirical is contemporary logical empiricism? In Language in Focus: Foundations, Methods and Systems: Essays in Memory of Yehoshua Bar-Hillel [Kas76b], chapter 16, pages 359–376. ISBN 90-277-0644-1, 90-277-0645-X, 94-010-1876-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN P101-120. URL https://link.

springer.com/chapter/10.1007/
978-94-010-1876-0_16.

Cohen:1976:CS

[Coh76d] Robert S. Cohen. Constraints on science. In Essays in Memory of Imre Lakatos [CFW76], chapter 8, pages 79– 86. ISBN 90-277-0655-7, 94-010-1451-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-010-1451-
                                              9_8.

Cohen:1980:TCQ

[Coh80] Robert S. Cohen. Time, Causality, and the Quantum Theory: Studies in the Philosophy of Science. Volume 1: Essay on the Causal Theory of Time, volume 19-1 of Boston Studies in the Philosophy of Science. D. Reidel, Dordrecht, The Netherlands; Boston, MA, USA; Lancaster, UK; Tokyo, Japan, 1980. ISBN 90-277-1074-0, 94-009-8935-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). xx + 308 pp. LCCN ????

                                              URL https://link.springer.com/
                                              book/10.1007/978-94-009-8935-1.

Cohen:1983:SIR

[Coh83] Robert S. Cohen. Social implications of recent technological innovations. In Philosophy and Technology [DR83], chapter 3, pages 35–47. ISBN 94-009-7126-5, 94-009-7124-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-009-7124-0_3.

Cohen:1986:RIH

[Coh86a] G. A. Cohen. Restricted and inclusive historical materialism. In Ullmann-Margalit [UM86c], chapter 5, pages 57–83. ISBN 90-277-2160-2, 90-277-2161-0 (paperback), 94-009-4566-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 95 Q175. URL https:/

/link.springer.com/chapter/10.
1007/978-94-009-4566-1_5.

Cohen:1986:DMP

[Coh86b] Lesley Cohen. Descartes and Merleau– Ponty on the cogito as the foundation of philosophy. In Cohen and Wartofsky [CW85], chapter 15, pages 295–312. ISBN 90-277-1971-3, 94-009-5345-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-009-5349-
9_15.

Cohen:1986:CS

[Coh86c] Robert S. Cohen. Constraints on science. In Cohen and Wartofsky [CW85], chapter 13, pages 198–205. ISBN 90-277-1971-3, 94-009-5345-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-009-5345-1_13.

Cohen:1986:PDM

[Coh86d] Robert S. Cohen. Phenomenology and Dialectical Materialism, volume 49 of Boston Studies in the Philosophy of Science. D. Reidel, Dordrecht, The Netherlands; Boston, MA, USA; Lancaster, UK; Tokyo, Japan, 1986. ISBN 94-010-8795-4, 94-009-5191-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). 280 (est.) pp. LCCN ????

URL https://link.springer.com/
book/10.1007/978-94-009-5191-4.

Cohen:1986:NPW

[Coh86e] Solly G. Cohen. The nature of the physical world revisited. In Ullmann-Margalit [UM86c], chapter 15, pages 183–194. ISBN 90-277-2160-2, 90-277-2161-0 (paperback), 94-009-4566-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 95 Q175. URL https:/

/link.springer.com/chapter/10.
1007/978-94-009-4566-1_15.

Cohen:1988:RVA

[Coh88a] L. Jonathan Cohen. Realism versus anti-realism: What is the issue? In Science in Reflection: The Israel Colloquium: Studies in History, Philosophy, and Sociology of Science Volume 3 [UM88], chapter 7, pages 81–96. ISBN 90-277-2713-9, 94-009-2957-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-009-2957-9_7.

Cohen:1988:BMH

[Coh88b] Robert S. Cohen. The Birth of Meaning in Hindu Thought, volume 102 of Boston Studies in the Philosophy of Science. D. Reidel, Dordrecht, The Netherlands; Boston, MA, USA; Lancaster, UK; Tokyo, Japan, 1988. ISBN 94-010-7141-1, 94-009-1431-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). xxii + 368 pp. LCCN B131 .Z55 1988; Q174 .B67 vol. 102 B131.

URL https://link.springer.com/
book/10.1007/978-94-009-1431-5.

Cohen:1990:GDC

[Coh90] I. Bernard Cohen. G. D. Cassini and the number of the planets: An example of Seventeenth-Century astro– numerological patronage. In Levere and Shea [LS90], chapter 8, pages 199–205. ISBN 0-7923-0420-9, 94-009-1878-X. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 120.

URL https://link.springer.com/
chapter/10.1007/978-94-009-1878-
8_8.

Cohen:1994:AIB

[Coh94a] I. Bernard Cohen. An analysis of interactions between the natural sciences and the social sciences. In The Natural Sciences and the Social Sciences: Some Critical and Historical Perspectives [Coh94b], chapter 1, pages 1–99. ISBN 0-7923-2223-1, 94-017-3391-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 150.

URL https://link.springer.com/
chapter/10.1007/978-94-017-3391-
5_1.

Cohen:1994:NSS

[Coh94b] I. Bernard Cohen, editor. The Natural Sciences and the Social Sciences: Some Critical and Historical Perspectives, volume 150 of Boston Studies in the Philosophy of Science. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 1994. ISBN 0-7923-2223-1, 94-017-3391-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). xxxvi + 404 pp. LCCN Q174 .B67 vol. 150.

URL https://link.springer.com/
book/10.1007/978-94-017-3391-5.

Cohen:1994:SRS

[Coh94c] I. Bernard Cohen. The scientific revolution and the social sciences. In The Natural Sciences and the Social Sciences: Some Critical and Historical Perspectives [Coh94b], chapter 4, pages 153–203. ISBN 0-7923-2223-1, 94-017-3391-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 150. URL https:/

/link.springer.com/chapter/10.
1007/978-94-017-3391-5_4.

Cohen:1994:KTC

[Coh94d] Robert S. Cohen. Kazimierz Twardowski (1866–1938), the creator of the Lwów–Warsaw School (LWS). In Zygmunt Zawirski: His Life and Work: With Selected Writings on Time, Logic and the Methodology of Science [Coh94m], chapter 2, pages 8–14. ISBN 94-010-4329-9, 94-011-0768-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-011-0768-6_2.

Cohen:1994:L

[Coh94e] Robert S. Cohen. Logic. In Zygmunt Zawirski: His Life and Work: With Selected Writings on Time, Logic and the Methodology of Science [Coh94m], chapter 7, pages 112–168. ISBN 94-010-4329-9, 94-011-0768-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-011-0768-6_7.

Cohen:1994:LWS

[Coh94f] Robert S. Cohen. The Lwów–Warsaw School (1895–1939). In Zygmunt Zawirski: His Life and Work: With Selected Writings on Time, Logic and the Methodology of Science [Coh94m], chapter 3, pages 15–18. ISBN 94-010-4329-9, 94-011-0768-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-011-0768-6_3.

Cohen:1994:LVC

[Coh94g] Robert S. Cohen. The LWS and the Vienna Circle (VC). In Zygmunt Zawirski: His Life and Work: With Selected Writings on Time, Logic and the Methodology of Science [Coh94m], chapter 4, pages 19–32. ISBN 94-010-4329-9, 94-011-0768-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-011-0768-6_4.

Cohen:1994:M

[Coh94h] Robert S. Cohen. Methodology. In Zygmunt Zawirski: His Life and Work: With Selected Writings on Time, Logic and the Methodology of Science [Coh94m], chapter 8, pages 169–189. ISBN 94-010-4329-9, 94-011-0768-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-011-0768-
6_8.

Cohen:1994:R

[Coh94i] Robert S. Cohen. Reports. In Zygmunt Zawirski: His Life and Work: With Selected Writings on Time, Logic and the Methodology of Science [Coh94m], chapter 6, pages 83–111. ISBN 94-010-4329-9, 94-011-0768-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-011-0768-6_6.

Cohen:1994:SD

[Coh94j] Robert S. Cohen. Some data. In Zygmunt Zawirski: His Life and Work: With Selected Writings on Time, Logic and the Methodology of Science [Coh94m], chapter 1, pages 3–7. ISBN 94-010-4329-9, 94-011-0768-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-011-0768-6_1.

Cohen:1994:TSE

[Coh94k] Robert S. Cohen. Time: Selections from the evolution of the notion of time. In Zygmunt Zawirski: His Life and Work: With Selected Writings on Time, Logic and the Methodology of Science [Coh94m], chapter 9, pages 190–368. ISBN 94-010-4329-9, 94-011-0768-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-011-
0768-6_9.

Cohen:1994:ZZS

[Coh94l] Robert S. Cohen. Zygmunt Zawirski: A short bio-bibliography (28.IX.1882– 2.IV.1948). In Zygmunt Zawirski: His Life and Work: With Selected Writings on Time, Logic and the Methodology of Science [Coh94m], chapter 5, pages 33– 78. ISBN 94-010-4329-9, 94-011-0768-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-011-0768-
6_5.

Cohen:1994:ZZH

[Coh94m] Robert S. Cohen. Zygmunt Zawirski: His Life and Work: With Selected Writings on Time, Logic and the Methodology of Science, volume 157 of Boston Studies in the Philosophy of Science. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 1994. ISBN 94-010-4329-9, 94-011-0768-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). xv + 384 pp. LCCN ????

URL https://link.springer.com/
book/10.1007/978-94-011-0768-6.

Cohen:1995:BAK

[Coh95] L. Jonathan Cohen. Belief, acceptance and knowledge. In The Concept of Knowledge: The Ankara Seminar [KC95], chapter 2, pages 11–19. ISBN 90-481-4495-7, 94-017-3263-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-017-3263-5_2.

Cohen:1999:NFP

[Coh99] Jonathan Cohen. Nietzsche’s fling with positivism. In Nietzsche, Epistemology, and Philosophy of Science: Nietzsche and the Sciences II [Bab99a], chapter 8, pages 101–107. ISBN 90-481-5234-8, 94-017-2428-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-017-
2428-9_8.

Cohen:2002:PPC

[Coh02a] L. Jonathan Cohen. Are people programmed to commit fallacies? Further thoughts about the interpretation of experimental data on probability judgment. In Knowledge and Language: Selected Essays of L. Jonathan Cohen [Coh02j], chapter 14, pages 195–220. ISBN 90-481-5955-5, 94-017-2020-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-017-2020-
5_14.

Cohen:2002:TER

[Coh02b] L. Jonathan Cohen. Are there ethical reasons for being, or not being, a scientific realist? In Knowledge and Language: Selected Essays of L. Jonathan Cohen [Coh02j], chapter 20, pages 313–322. ISBN 90-481-5955-5, 94-017-2020-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-017-
2020-5_20.

Cohen:2002:BVB

[Coh02c] L. Jonathan Cohen. Bayesianism versus Baconianism in the evaluation of medical diagnoses. In Knowledge and Language: Selected Essays of L. Jonathan Cohen [Coh02j], chapter 13, pages 175–193. ISBN 90-481-5955-5, 94-017-2020-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

                                              ???? URL https://link.springer.
                                              com/chapter/10.1007/978-94-017-
                                              2020-5_13.

Cohen:2002:CCH

[Coh02d] L. Jonathan Cohen. Can the conversationalist hypothesis be defended? In Knowledge and Language: Selected Essays of L. Jonathan Cohen [Coh02j], chapter 6, pages 67–76. ISBN 90-481-5955-5, 94-017-2020-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-017-2020-5_6.

Cohen:2002:CLI

[Coh02e] L. Jonathan Cohen. Can the logic of indirect discourse be formalised? In Knowledge and Language: Selected Essays of L. Jonathan Cohen [Coh02j], chapter 4, pages 41–48. ISBN 90-481-5955-5, 94-017-2020-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-017-2020-5_4.

Cohen:2002:G

[Coh02f] L. Jonathan Cohen. Guessing. In Knowledge and Language: Selected Essays of L. Jonathan Cohen [Coh02j], chapter 12, pages 155–174. ISBN 90-481-5955-5, 94-017-2020-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-017-2020-5_12.

Cohen:2002:HCI

[Coh02g] L. Jonathan Cohen. How is conceptual innovation possible? In Knowledge and Language: Selected Essays of L. Jonathan Cohen [Coh02j], chapter 7, pages 77–93. ISBN 90-481-5955-5, 94-017-2020-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-017-
2020-5_7.

Cohen:2002:IPN

[Coh02h] L. Jonathan Cohen. The individuation of proper names. In Knowledge and Language: Selected Essays of L. Jonathan Cohen [Coh02j], chapter 10, pages 119–139. ISBN 90-481-5955-5, 94-017-2020-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-017-
2020-5_10.

Cohen:2002:IL

[Coh02i] L. Jonathan Cohen. Inductive logic 1945–1977. In Knowledge and Language: Selected Essays of L. Jonathan Cohen [Coh02j], chapter 15, pages 221–243. ISBN 90-481-5955-5, 94-017-2020-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-017-
2020-5_15.

Cohen:2002:KLS

[Coh02j] L. Jonathan Cohen. Knowledge and Language: Selected Essays of L. Jonathan Cohen, volume 227 of Boston Studies in the Philosophy of Science.

Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 2002. ISBN 90-481-5955-5, 94-017-2020-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). xxviii + 324 pp. LCCN ????

URL https://link.springer.com/
book/10.1007/978-94-017-2020-5.

Cohen:2002:CDG

[Coh02k] L. Jonathan Cohen. On a concept of degree of grammaticalness. In Knowledge and Language: Selected Essays of L. Jonathan Cohen [Coh02j], chapter 2, pages 15–26. ISBN 90-481-5955-5, 94-017-2020-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-017-
2020-5_2.

Cohen:2002:PUC

[Coh02l] L. Jonathan Cohen. On the project of a universal character. In Knowledge and Language: Selected Essays of L. Jonathan Cohen [Coh02j], chapter 1, pages 1–14. ISBN 90-481-5955-5, 94-017-2020-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-017-
2020-5_1.

Cohen:2002:PAA

[Coh02m] L. Jonathan Cohen. A problem about ambiguity in truth-theoretical semantics. In Knowledge and Language: Selected Essays of L. Jonathan Cohen [Coh02j], chapter 9, pages 113–118. ISBN 90-481-5955-5, 94-017-2020-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-017-2020-
5_9.

Cohen:2002:SM

[Coh02n] L. Jonathan Cohen. The semantics of metaphor. In Knowledge and Language: Selected Essays of L. Jonathan Cohen [Coh02j], chapter 3, pages 27–40. ISBN 90-481-5955-5, 94-017-2020-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-017-2020-5_3.

Cohen:2002:SJS

[Coh02o] L. Jonathan Cohen. Should a jury say what it believes or what it accepts? In Knowledge and Language: Selected Essays of L. Jonathan Cohen [Coh02j], chapter 19, pages 293–311. ISBN 90-481-5955-5, 94-017-2020-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-017-2020-5_19.

Cohen:2002:SNL

[Coh02p] L. Jonathan Cohen. Should naturallanguage definitions be insulated from, or interactive with, one another in sentence composition? In Knowledge and Language: Selected Essays of L. Jonathan Cohen [Coh02j], chapter 8, pages 95–112. ISBN 90-481-5955-5, 94-017-2020-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-017-
2020-5_8.

Cohen:2002:SHR

[Coh02q] L. Jonathan Cohen. Some historical remarks on the Baconian conception of probability. In Knowledge and Language: Selected Essays of L. Jonathan Cohen [Coh02j], chapter 16, pages 245–259. ISBN 90-481-5955-5, 94-017-2020-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

                                              ???? URL https://link.springer.
                                              com/chapter/10.1007/978-94-017-
                                              2020-5_16.

Cohen:2002:SRG

[Coh02r] L. Jonathan Cohen. Some remarks on Grice’s views about the logical particles of natural language. In Knowledge and Language: Selected Essays of L. Jonathan Cohen [Coh02j], chapter 5, pages 49–66. ISBN 90-481-5955-5, 94-017-2020-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

                                              ???? URL https://link.springer.
                                              com/chapter/10.1007/978-94-017-
                                              2020-5_5.

Cohen:2002:SST

[Coh02s] L. Jonathan Cohen. Some steps towards a general theory of relevance. In Knowledge and Language: Selected Essays of L. Jonathan Cohen [Coh02j], chapter 18, pages 279–292. ISBN 90-481-5955-5, 94-017-2020-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-017-2020-5_18.

Cohen:2002:TWE

[Coh02t] L. Jonathan Cohen. Third world epistemology. In Knowledge and Language: Selected Essays of L. Jonathan Cohen [Coh02j], chapter 11, pages 141–153. ISBN 90-481-5955-5, 94-017-2020-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-017-2020-
                                              5_11.

Cohen:2002:TQA

[Coh02u] L. Jonathan Cohen. Twelve questions about Keynes’s concept of weight. In Knowledge and Language: Selected Essays of L. Jonathan Cohen [Coh02j], chapter 17, pages 261–278. ISBN 90-481-5955-5, 94-017-2020-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-017-2020-5_17.

Cohen:2004:HAF

[Coh04] H. Floris Cohen. A historical–analytical framework for the controversies over Galileo’s conception of motion. In The Reception of the Galilean Science of Motion in Seventeenth-Century Europe [PT04], chapter 5, pages 83–97. ISBN 1-4020-2454-1, 1-4020-2455-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC133 .R43 2004.

URL https://link.springer.com/
chapter/10.1007/978-1-4020-2455-
9_5.

Cohen:2016:WPC

[Coh16] Yehiel Cohen. Why presentism cannot be refuted by special relativity. In Cosmological and Psychological Time [DR16], chapter 3, pages 41–51. ISBN 3-319-22589-8 (hardcover), 3-319-22590-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN BD638 C67 2016. URL https://link.

springer.com/chapter/10.1007/
978-3-319-22590-6_3.

Cohen:2017:EWA

[Coh17a] Daniel B. Cohen. Ebola in West Africa:

Biosocial and biomedical reflections. In Encouraging Openness: Essays for Joseph Agassi on the Occasion of His 90th Birthday [BAG17], chapter 13, pages 143–164. ISBN 3-319-57668-2 (hardcover), 3-319-57669-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN B833.E53 2017.

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-3-319-57669-
                                              5_13.

Cohen:2017:RSI

[Coh17b] Robert S. Cohen. Reflections on science and imagination. In Encouraging Openness: Essays for Joseph Agassi on the Occasion of His 90th Birthday [BAG17], chapter 17, pages 215–220. ISBN 3-319-57668-2 (hardcover), 3-319-57669-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN B833.E53 2017. URL https://link.

                                              springer.com/chapter/10.1007/
                                              978-3-319-57669-5_17.

Collins:1974:PSA

[Col74] Paul W. Collins. The present status of anthropology as an explanatory science. In Seeger and Cohen [SC74b], pages 337–348. ISBN 90-277-0390-6, 90-277-0376-0 (paperback), 94-010-2126-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 11. URL https://link.

                                              springer.com/chapter/10.1007/
                                              978-94-010-2126-5_19.

Collins:1996:ECB

[Col96]  Robin Collins.  An epistemologi-

cal critique of Bohmian mechanics. In Bohmian Mechanics and Quantum Theory: an Appraisal [CFG96], chapter 18, pages 265–276. ISBN 90-481-4698-4, 94-015-8715-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC173.96-174.52.

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-015-8715-
                                              0_18.

Compton:1984:CBC

[Com84] John J. Compton. A comment on Buchdahl’s ‘Conceptual Analysis and Scientific Theory In Hegel’s Philosophy of Nature (With Special Reference To Hegel’s Optics)’. In Hegel and the Sciences [CW84a], chapter 3, pages 37–40. ISBN 94-009-6235-5, 94-009-6233-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-009-6233-
0_3.

Compton:2002:TPP

[Com02] John J. Compton. Toward a phenomenological philosophy of nature. In Hermeneutic Philosophy of Science, Van Gogh’s Eyes, and God: Essays in Honor of Patrick A. Heelan, S.J. [Bab02c], chapter 16, pages 195–202. ISBN 90-481-5926-1, 94-017-1767-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 225. URL https://link.

springer.com/chapter/10.1007/
978-94-017-1767-0_16.

Constantini:1981:ILI

[Con81] Domenico Constantini. Inductive logic and inductive statistics. In Italian Studies in the Philosophy of Science [Chi81b], chapter 11, pages 169–183. ISBN 90-277-1073-2, 94-009-8937-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-009-8937-
5_11.

Conde:1996:BMA

[Con96] Camilo J. Cela Conde. Biological and moral altruism. In Munévar [Mun96], chapter 7, pages 143–151. ISBN 94-010-6622-1, 94-009-0305-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-009-0305-0_7.

Conway:1999:BTA

[Con99] Daniel Conway. Beyond truth and appearance: Nietzsche’s emergent realism. In Nietzsche, Epistemology, and Philosophy of Science: Nietzsche and the Sciences II [Bab99a], chapter 9, pages 109–122. ISBN 90-481-5234-8, 94-017-2428-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

                                              ???? URL https://link.springer.
                                              com/chapter/10.1007/978-94-017-
                                              2428-9_9.

Cook:1984:TPB

[Coo84] Daniel Cook. Theory and praxis and the beginning of science. In Hegel and the Sciences [CW84a], chapter 16, pages 283–286. ISBN 94-009-6235-5, 94-009-6233-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

                                              ???? URL https://link.springer.
                                              com/chapter/10.1007/978-94-009-
                                              6233-0_16.

Cook:1986:SKS

[Coo86] J. Thomas Cook. Self-knowledge as self-preservation? In Spinoza and the Sciences [GN86], chapter 8, pages 191–210. ISBN 94-010-8511-0, 94-009-4514-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

                                              ???? URL https://link.springer.
                                              com/chapter/10.1007/978-94-009-
                                              4514-2_8.

Cooke:1990:SER

[Coo90] Roger M. Cooke. Statistics in expert resolution: A theory of weights for combining expert opinion. In Statistics in Science: The Foundations of Statistical Methods in Biology, Physics and Economics [CC90], chapter 4, pages 41–72. ISBN 94-010-6765-1, 94-009-0619-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic).

LCCN QA276-280. URL https:/

/link.springer.com/chapter/10.
1007/978-94-009-0619-8_4.

Cook:2011:CCM

[Coo11] Harold J. Cook. Conveying Chinese medicine to Seventeenth-Century Europe. In Science between Europe and Asia: Historical Studies on the Transmission, Adoption and Adaptation of Knowledge [GR11b], chapter 14, pages 209–232. ISBN 90-481-9968-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q125 .S354 2010; Q175.52.A78 S35 2010. URL https:

//link.springer.com/chapter/10.
1007/978-90-481-9968-6_14.

Cooper:2017:ECH

[Coo17] S. Barry Cooper. Embodying computation at higher types. In Floyd and Bokulich [FB17], chapter 10, pages 233–262. ISBN 3-319-53278-2, 3-319-53280-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QA29.T8 P45 2017. URL https:/

/link.springer.com/chapter/10.
1007/978-3-319-53280-6_10.

Copi:1976:PPL

[Cop76] Irving M. Copi. A problem in Plato’s laws. In Language in Focus: Foundations, Methods and Systems: Essays in Memory of Yehoshua Bar-Hillel [Kas76b], chapter 26, pages 627–639. ISBN 90-277-0644-1, 90-277-0645-X, 94-010-1876-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN P101-120. URL https://link.

springer.com/chapter/10.1007/
978-94-010-1876-0_26.

Cornford:1976:ETP

[Cor76a] F. M. Cornford. The elimination of time by Parmenides. In The Concepts of Space and Time: Their Structure and Their Development [Čap76], chapter 24, pages 137–142. ISBN 90-277-0375-2, 94-010-1727-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic).

LCCN BD632. URL https://link.
springer.com/chapter/10.1007/
978-94-010-1727-5_24.

Cornford:1976:IS

[Cor76b] F. M. Cornford. The invention of space. In The Concepts of Space and Time: Their Structure and Their Development [Čap76], chapter 1, pages 3–16. ISBN 90-277-0375-2, 94-010-1727-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN BD632.

URL https://link.springer.com/
chapter/10.1007/978-94-010-1727-
5_1.

Correspondence:1995:PL

[Cor95] A. Correspondence. Philosophy lectures. In Blackmore [Bla95g], pages 111–144. ISBN 0-7923-3231-8, 90-481-4492-2, 94-017-0489-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 168; QC16 .B64. URL https://link.

springer.com/chapter/10.1007/
978-94-017-0489-2_5.

Cordero:2000:RCR

[Cor00] Alberto Cordero. Realism, and the case of rival theories without observable differences. In The Reality of the Unobservable: Observability, Unobservability and Their Impact on the Issue of Scientific Realism [AP00b], chapter 14, pages 191–206. ISBN 90-481-5458-8, 94-015-9391-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-015-9391-5_14.

Corry:2007:OHA

[Cor07] Leo Corry. The origin of Hilbert’s axiomatic method. In Renn et al. [RSS^{+}07b], pages 759–855. ISBN 1-4020-3999-9 (hardcover), 1-4020-4000-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 no. 250 v. 4; QC173.6 .G469 2007.

URL https://link.springer.com/
chapter/10.1007/978-1-4020-4000-
9_40.

Cosic:1979:CR

[Cos79] Dobrica Cosic. Culture and revolution.

In Praxis: Yugoslav Essays in the Philosophy and Methodology of the Social Sciences [MP79], chapter 14, pages 217–225. ISBN 90-277-0968-8, 94-009-9355-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-009-
9355-6_14.

CostadeBeauregard:1987:MLC

[Cos87a] Olivier Costa de Beauregard. S-matrix, Lorentz-and-CPT invariance, and Einstein–Podolsky–Rosen correlations. In Time, The Physical Magnitude [Cos87b], chapter 19, pages 265–284. ISBN 94-010-8195-6, 94-009-3811-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-009-3811-
3_19.

CostaDeBeauregard:1987:TPM

[Cos87b] Olivier Costa De Beauregard. Time, The Physical Magnitude, volume 99 of Boston Studies in the Philosophy of Science. D. Reidel, Dordrecht, The Netherlands; Boston, MA, USA; Lancaster, UK; Tokyo, Japan, 1987. ISBN 94-010-8195-6, 94-009-3811-X (e-book).

ISSN 0068-0346 (print), 2214-7942 (electronic). xxiii + 340 pp. LCCN ????

URL https://link.springer.com/
book/10.1007/978-94-009-3811-3.

Couture:1996:DTI

[Cou96] Jocelyne Couture. Decision theory, individualistic explanations and social Darwinism. In Québec Studies in the Philosophy of Science: Part II: Biology, Psychology, Cognitive Science and Economics Essays in Honor of Hugues Leblanc [MC96], chapter 14, pages 229–246. ISBN 94-010-6537-3, 94-009-0113-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-009-
0113-1_14.

Coxeter:1974:PN

[Cox74] H. S. M. Coxeter. Polyhedral numbers. In For Dirk Struik: Scientific, Historical and Political Essays in Honor of Dirk J. Struik [CSW74], chapter 4, pages 25–35. ISBN 90-277-0379-5, 94-010-2115-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-010-
2115-9_4.

Carrara:2016:CFS

[CPM16] Massimiliano Carrara, Matteo Plebani, and Enrico Martino. Computability, finiteness and the standard model of arithmetic. In Objectivity, Realism, and Proof: FilMat Studies in the Philosophy of Mathematics [BS16], chapter 16, pages 311–318. ISBN 3-319-31642-7, 3-319-31644-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QA9 .O25 2016. URL https:

//link.springer.com/chapter/10.
1007/978-3-319-31644-4_16.

Caldirola:1981:CTR

[CR81] P. Caldirola and E. Recami. Causality and tachyons in relativity. In Italian Studies in the Philosophy of Science [Chi81b], chapter 16, pages 249–298. ISBN 90-277-1073-2, 94-009-8937-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-009-8937-
5_16.

Crawford:1993:MFT

[Cra93] Elspeth Crawford. Michael Faraday’s thought: Discovery or revelation? In Correspondence, Invariance and Heuristics: Essays in Honour of Heinz Post [FK93], chapter 5, pages 105–124. ISBN 90-481-4229-6, 94-017-1185-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-017-1185-
2_5.

Crease:1999:POH

[Cre99] Robert P. Crease. Productive objectivity: The hermeneutics of performance in experimental inquiry. In Fehér et al. [FKR99], pages 25–34. ISBN 90-481-5257-7, 94-015-9293-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-015-9293-2_3.

Crease:2002:ELH

[Cre02] Robert P. Crease. Experimental life:

Heelan on quantum mechanics. In Hermeneutic Philosophy of Science, Van Gogh’s Eyes, and God: Essays in Honor of Patrick A. Heelan, S.J. [Bab02c], chapter 3, pages 31–41. ISBN 90-481-5926-1, 94-017-1767-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 225.

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-017-1767-
                                              0_3.

Crosson:1969:CG

[Cro69] F. J. Crosson. The computer as gadfly. In Cohen and Wartofsky [CW69b], pages 226–240. ISBN 94-010-3380-3, 94-010-3378-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic).

                                              LCCN ????  URL https://link.
                                              springer.com/chapter/10.1007/
                                              978-94-010-3378-7_8.

Crow:1992:HJM

[Cro92a] James F. Crow. H. J. Muller’s role in evolutionary biology. In Sarkar [Sar92a], chapter 5, pages 83–105. ISBN 0-7923-1777-7 (hardcover), 94-011-2856-1. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 142. URL https://link.

                                              springer.com/chapter/10.1007/
                                              978-94-011-2856-8_5.

Crow:1992:SWP

[Cro92b] James F. Crow. Sewall Wright’s place in Twentieth-Century biology. In Sarkar [Sar92a], chapter 7, pages 167– 200. ISBN 0-7923-1777-7 (hardcover), 94-011-2856-1. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 142. URL https://link.

                                              springer.com/chapter/10.1007/
                                              978-94-011-2856-8_7.

Crowell:1999:NAN

[Cro99] Steven Galt Crowell. Nietzsche among the neo-Kantians; or, the relation between science and philosophy. In Nietzsche, Theories of Knowledge, and Critical Theory: Nietzsche and the Sciences I [Bab99b], chapter 6, pages 77–86. ISBN 90-481-5233-X, 94-017-2430-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-017-
2430-2_6.

Crowell:2002:PHI

[Cro02] Steven Crowell. Patrick Heelan’s innocent eye. In Hermeneutic Philosophy of Science, Van Gogh’s Eyes, and God: Essays in Honor of Patrick A. Heelan, S.J. [Bab02c], chapter 20, pages 239–250. ISBN 90-481-5926-1, 94-017-1767-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 225. URL https:/

/link.springer.com/chapter/10.
1007/978-94-017-1767-0_20.

Cohen:1970:EMPb

[CS70] Robert S. Cohen and Raymond J.

Seeger. Ernst Mach: Physicist and Philosopher, volume 6 of Boston Studies in the Philosophy of Science. D. Reidel, Dordrecht, The Netherlands; Boston, MA, USA; Lancaster, UK; Tokyo, Japan, 1970. ISBN 90-481-8318-9, 90-277-0016-8, 94-017-1462-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). viii + 298 pp. LCCN ????

URL https://link.springer.com/
book/10.1007/978-94-017-1462-4.

Cohen:1979:SPL

[CS79a] R. S. (Robert Sonné) Cohen and John J. Stachel, editors. Selected Papers of Léon Rosenfeld, volume 21 of Boston Studies in the Philosophy of Science. D. Reidel, Dordrecht, The Netherlands; Boston, MA, USA; Lancaster, UK; Tokyo, Japan, 1979. ISBN 90-277-0651-4, 90-277-0652-2 (paperback), 94-009-9349-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). xxxiv + 929 pp. LCCN Q174 .B67 vol. 21 QC7. URL https:/

                                              /link.springer.com/book/10.1007/
                                              978-94-009-9349-5.

Cohen:1979:AJE

[CS79b] Robert S. Cohen and John J. Stachel. An analysis of Joule’s experiments on the expansion of air [with A. P. Hatton] [1956e]. In Cohen and Stachel [CS79a], chapter 13, pages 123–133. ISBN 90-277-0651-4, 90-277-0652-2 (paperback), 94-009-9349-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 21

                                              QC7. URL https://link.springer.
                                              com/chapter/10.1007/978-94-009-
                                              9349-5_13.

Cohen:1979:CTP

[CS79c] Robert S. Cohen and John J. Stachel. Celestial and terrestrial physics in historical perspective [1969b]. In Cohen and Stachel [CS79a], chapter 25, pages 346–353. ISBN 90-277-0651-4, 90-277-0652-2 (paperback), 94-009-9349-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 21 QC7. URL https://link.

                                              springer.com/chapter/10.1007/
                                              978-94-009-9349-5_25.

Cohen:1979:CSI

[CS79d] Robert S. Cohen and John J. Stachel. Complementarity and statistics I and II [1958b]. In Cohen and Stachel [CS79a], chapter 32, pages 484–494. ISBN 90-277-0651-4, 90-277-0652-2 (paperback), 94-009-9349-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 21

                                              QC7. URL https://link.springer.
                                              com/chapter/10.1007/978-94-009-
                                              9349-5_32.

Cohen:1979:CMF

[CS79e] Robert S. Cohen and John J. Stachel. The conception of the meson field some reminiscences and epistemological comments [1968e]. In Cohen and Stachel [CS79a], chapter 23, pages 327–334. ISBN 90-277-0651-4, 90-277-0652-2 (paperback), 94-009-9349-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 21

QC7. URL https://link.springer.
com/chapter/10.1007/978-94-009-
9349-5_23.

Cohen:1979:CIF

[CS79f] Robert S. Cohen and John J. Stachel.

Condillac’s influence on French scientific thought [1972f]. In Cohen and Stachel [CS79a], chapter 45, pages 655– 665. ISBN 90-277-0651-4, 90-277-0652-2 (paperback), 94-009-9349-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 21

QC7. URL https://link.springer.
com/chapter/10.1007/978-94-009-
9349-5_45.

Cohen:1979:DUR

[CS79g] Robert S. Cohen and John J. Stachel.

Descartes at Uppsala (review of Rolf Lindborg’s Descartes i Uppsala. Striderna om “nya filosofien” 1633–1689. [Almqvist & Wiksell, Stockholm, 1965. xvi + 367 pp.]) [1967a]. In Cohen and Stachel [CS79a], chapter 8, pages 55– 57. ISBN 90-277-0651-4, 90-277-0652-2 (paperback), 94-009-9349-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 21

QC7. URL https://link.springer.
com/chapter/10.1007/978-94-009-
9349-5_8.

Cohen:1979:ECB

[CS79h] Robert S. Cohen and John J. Stachel.

The epistemological conflict between Einstein and Bohr (dedicated to Max Born on his 80th birthday) [1963c]. In Cohen and Stachel [CS79a], chapter 35, pages 517–521. ISBN 90-277-0651-4, 90-277-0652-2 (paperback), 94-009-9349-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 21 QC7. URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-009-9349-5_35.

Cohen:1979:ESB

[CS79i] Robert S. Cohen and John J. Stachel. Epistemology on a scientific basis [1971a]. In Cohen and Stachel [CS79a], chapter 44, pages 643–654. ISBN 90-277-0651-4, 90-277-0652-2 (paperback), 94-009-9349-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 21 QC7.

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-009-9349-
                                              5_44.

Cohen:1979:EOS

[CS79j] Robert S. Cohen and John J. Stachel. The evolution of Oersted’s scientific concepts [1970b]. In Cohen and Stachel [CS79a], chapter 15, pages 178–192. ISBN 90-277-0651-4, 90-277-0652-2 (paperback), 94-009-9349-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 21

                                              QC7. URL https://link.springer.
                                              com/chapter/10.1007/978-94-009-
                                              9349-5_15.

Cohen:1979:EIC

[CS79k] Robert S. Cohen and John J. Stachel. The evolution of the idea of causality [1942d]. In Cohen and Stachel [CS79a], chapter 30, pages 446–464. ISBN 90-277-0651-4, 90-277-0652-2 (paperback), 94-009-9349-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 21

                                              QC7. URL https://link.springer.
                                              com/chapter/10.1007/978-94-009-
                                              9349-5_30.

Cohen:1979:FCM

[CS79l] Robert S. Cohen and John J. Stachel.

Field and charge measurements in quantum electrodynamics [with N. Bohr] [1950a]. In Cohen and Stachel [CS79a], chapter 27, pages 401–412. ISBN 90-277-0651-4, 90-277-0652-2 (paperback), 94-009-9349-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 21

QC7. URL https://link.springer.
com/chapter/10.1007/978-94-009-
9349-5_27.

Cohen:1979:FPE

[CS79m] Robert S. Cohen and John J. Stachel.

The first phase in the evolution of the quantum theory [1936b]. In Cohen and Stachel [CS79a], chapter 16, pages 193– 234. ISBN 90-277-0651-4, 90-277-0652-2 (paperback), 94-009-9349-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 21

QC7. URL https://link.springer.
com/chapter/10.1007/978-94-009-
9349-5_16.

Cohen:1979:FDR

[CS79n] Robert S. Cohen and John J. Stachel.

Flicker in the darkness (review of Nicole Oresme and the Medieval Geometry of Qualities and Motions (ed. M. Clagett)[University of Wisconsin Press, Madison and London 1968. xiii + 713 pp.]) [1969e]. In Cohen and Stachel [CS79a], chapter 6, pages 45– 48. ISBN 90-277-0651-4, 90-277-0652-2 (paperback), 94-009-9349-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 21

QC7. URL https://link.springer.
com/chapter/10.1007/978-94-009-
9349-5_6.

Cohen:1979:FQT

[CS79o] Robert S. Cohen and John J. Stachel. Foundations of quantum theory and complementarity [1961e]. In Cohen and Stachel [CS79a], chapter 34, pages 503– 516. ISBN 90-277-0651-4, 90-277-0652-2 (paperback), 94-009-9349-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 21

                                              QC7. URL https://link.springer.
                                              com/chapter/10.1007/978-94-009-
                                              9349-5_34.

Cohen:1979:GLT

[CS79p] Robert S. Cohen and John J. Stachel. The genesis of the laws of thermodynamics [1941a]. In Cohen and Stachel [CS79a], chapter 11, pages 99–111. ISBN 90-277-0651-4, 90-277-0652-2 (paperback), 94-009-9349-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 21

                                              QC7. URL https://link.springer.
                                              com/chapter/10.1007/978-94-009-
                                              9349-5_11.

Cohen:1979:ILS

[CS79q] Robert S. Cohen and John J. Stachel. Irreversibility — a lay sermon (on the occasion of Professor K. Beuler’s sixtieth birthday, as a token of esteem and friendship) [1977b]. In Cohen and Stachel [CS79a], chapter 47, pages 681– 685. ISBN 90-277-0651-4, 90-277-0652-2 (paperback), 94-009-9349-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 21

                                              QC7. URL https://link.springer.
                                              com/chapter/10.1007/978-94-009-
                                              9349-5_47.

Cohen:1979:JS

[CS79r] Robert S. Cohen and John J. Stachel. Jacques Solomon [1959c]. In Cohen and Stachel [CS79a], chapter 20, pages 297–301. ISBN 90-277-0651-4, 90-277-0652-2 (paperback), 94-009-9349-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 21 QC7. URL https://link.

springer.com/chapter/10.1007/
978-94-009-9349-5_20.

Cohen:1979:JSO

[CS79s] Robert S. Cohen and John J. Stachel.

Joule’s scientific outlook [1952a]. In Cohen and Stachel [CS79a], chapter 12, pages 112–122. ISBN 90-277-0651-4, 90-277-0652-2 (paperback), 94-009-9349-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 21 QC7. URL https:

//link.springer.com/chapter/10.
1007/978-94-009-9349-5_12.

Cohen:1979:LPD

[CS79t] Robert S. Cohen and John J. Stachel.

The logical problem of the definition of irrational numbers [1927e]. In Cohen and Stachel [CS79a], chapter 3, pages 16–28. ISBN 90-277-0651-4, 90-277-0652-2 (paperback), 94-009-9349-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 21 QC7. URL https://link.

springer.com/chapter/10.1007/
978-94-009-9349-5_3.

Cohen:1979:MMI

[CS79u] Robert S. Cohen and John J. Stachel.

Marcus Marci’s investigations of the prism and their relation to Newton’s theory of color [1932c]. In Cohen and Stachel [CS79a], chapter 7, pages 49– 54. ISBN 90-277-0651-4, 90-277-0652-2 (paperback), 94-009-9349-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 21

QC7. URL https://link.springer.
com/chapter/10.1007/978-94-009-
9349-5_7.

Cohen:1979:MFA

[CS79v] Robert S. Cohen and John J. Stachel. Matter and force after fifty years of quantum theory [1963d]. In Cohen and Stachel [CS79a], chapter 18, pages 247– 265. ISBN 90-277-0651-4, 90-277-0652-2 (paperback), 94-009-9349-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 21

                                              QC7. URL https://link.springer.
                                              com/chapter/10.1007/978-94-009-
                                              9349-5_18.

Cohen:1979:MPS

[CS79w] Robert S. Cohen and John J. Stachel. Max Planck and the statistical definition of entropy [1959b]. In Cohen and Stachel [CS79a], chapter 17, pages 235– 246. ISBN 90-277-0651-4, 90-277-0652-2 (paperback), 94-009-9349-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 21

                                              QC7. URL https://link.springer.
                                              com/chapter/10.1007/978-94-009-
                                              9349-5_17.

Cohen:1979:MPQ

[CS79x] Robert S. Cohen and John J. Stachel. The measuring process in quantum mechanics (on the 30th anniversary of the meson theory by Dr. H. Yukawa, 1965) [1965c]. In Cohen and Stachel [CS79a], chapter 37, pages 536–546. ISBN 90-277-0651-4, 90-277-0652-2 (paperback), 94-009-9349-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 21

                                              QC7. URL https://link.springer.
                                              com/chapter/10.1007/978-94-009-
                                              9349-5_37.

Cohen:1979:MIH

[CS79y] Robert S. Cohen and John J. Stachel. Men and ideas in the history of atomic theory [1971d]. In Cohen and Stachel [CS79a], chapter 19, pages 266–296. ISBN 90-277-0651-4, 90-277-0652-2 (paperback), 94-009-9349-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 21

QC7. URL https://link.springer.
com/chapter/10.1007/978-94-009-
9349-5_19.

Cohen:1979:MP

[CS79z] Robert S. Cohen and John J. Stachel.

The method of physics [1968g]. In Cohen and Stachel [CS79a], chapter 42, pages 614–636. ISBN 90-277-0651-4, 90-277-0652-2 (paperback), 94-009-9349-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 21 QC7. URL https:

//link.springer.com/chapter/10.
1007/978-94-009-9349-5_42.

Cohen:1979:MAF

[CS79-27] Robert S. Cohen and John J. Stachel.

Misunderstandings about the foundations of quantum theory [1957a]. In Cohen and Stachel [CS79a], chapter 33, pages 495–502. ISBN 90-277-0651-4, 90-277-0652-2 (paperback), 94-009-9349-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 21 QC7. URL https:

//link.springer.com/chapter/10.
1007/978-94-009-9349-5_33.

Cohen:1979:NLG

[CS79-28] Robert S. Cohen and John J. Stachel.

Newton and the Law of Gravitation [1965d]. In Cohen and Stachel [CS79a], chapter 9, pages 58–87. ISBN 90-277-0651-4, 90-277-0652-2 (paperback), 94-009-9349-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 21

QC7. URL https://link.springer.
com/chapter/10.1007/978-94-009-
9349-5_9.

Cohen:1979:NVA

[CS79-29] Robert S. Cohen and John J. Stachel. Newton’s views on aether and gravitation [1969c]. In Cohen and Stachel [CS79a], chapter 10, pages 88–98. ISBN 90-277-0651-4, 90-277-0652-2 (paperback), 94-009-9349-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 21

                                              QC7. URL https://link.springer.
                                              com/chapter/10.1007/978-94-009-
                                              9349-5_10.

Cohen:1979:NBE

[CS79-30] Robert S. Cohen and John J. Stachel. Niels Bohr: An essay dedicated to him on the occasion of his sixtieth birthday, October 7, 1945 [second edition corrected] 1961 [1945e]. In Cohen and Stachel [CS79a], chapter 22, pages 313–326. ISBN 90-277-0651-4, 90-277-0652-2 (paperback), 94-009-9349-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 21 QC7. URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-009-9349-5_22.

Cohen:1979:NBC

[CS79-31] Robert S. Cohen and John J. Stachel. Niels Bohr’s contribution to epistemology [1963h]. In Cohen and Stachel [CS79a], chapter 36, pages 522–535. ISBN 90-277-0651-4, 90-277-0652-2 (paperback), 94-009-9349-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 21

                                              QC7. URL https://link.springer.
                                              com/chapter/10.1007/978-94-009-
                                              9349-5_36.

Cohen:1979:NR

[CS79-32] Robert S. Cohen and John J. Stachel. Nuclear reminiscences. In Cohen and Stachel [CS79a], chapter 24, pages 335– 345. ISBN 90-277-0651-4, 90-277-0652-2 (paperback), 94-009-9349-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 21

QC7. URL https://link.springer.
com/chapter/10.1007/978-94-009-
9349-5_24.

Cohen:1979:QF

[CS79-33] Robert S. Cohen and John J. Stachel.

On quantization of fields [1963a]. In Cohen and Stachel [CS79a], chapter 29, pages 442–445. ISBN 90-277-0651-4, 90-277-0652-2 (paperback), 94-009-9349-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 21 QC7. URL https:

//link.springer.com/chapter/10.
1007/978-94-009-9349-5_29.

Cohen:1979:QEA

[CS79-34] Robert S. Cohen and John J. Stachel.

On quantum electrodynamics [among essays dedicated to Niels Bohr on the occasion of his 70th birthday] [1955b]. In Cohen and Stachel [CS79a], chapter 28, pages 413–441. ISBN 90-277-0651-4, 90-277-0652-2 (paperback), 94-009-9349-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 21 QC7.

URL https://link.springer.com/
chapter/10.1007/978-94-009-9349-
5_28.

Cohen:1979:MHS

[CS79-35] Robert S. Cohen and John J. Stachel.

On the method of history of science [1947b]. In Cohen and Stachel [CS79a], chapter 1, pages 3–6. ISBN 90-277-0651-4, 90-277-0652-2 (paperback), 94-009-9349-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 21 QC7.

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-009-9349-
                                              5_1.

Cohen:1979:QME

[CS79-36] Robert S. Cohen and John J. Stachel. On the question of the measurability of electromagnetic field quantities [with Niels Bohr] [1933b]. In Cohen and Stachel [CS79a], chapter 26, pages 357– 400. ISBN 90-277-0651-4, 90-277-0652-2 (paperback), 94-009-9349-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 21

                                              QC7. URL https://link.springer.
                                              com/chapter/10.1007/978-94-009-
                                              9349-5_26.

Cohen:1979:QTG

[CS79-37] Robert S. Cohen and John J. Stachel. Quantum theory and gravitation [1966a]. In Cohen and Stachel [CS79a], chapter 40, pages 599–608. ISBN 90-277-0651-4, 90-277-0652-2 (paperback), 94-009-9349-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 21 QC7.

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-009-9349-
                                              5_40.

Cohen:1979:QTR

[CS79-38] Robert S. Cohen and John J. Stachel. Quantum theory in 1929: Recollections from the First Copenhagen Conference [1971c]. In Cohen and Stachel [CS79a], chapter 21, pages 302–312. ISBN 90-277-0651-4, 90-277-0652-2 (paperback), 94-009-9349-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 21

                                              QC7. URL https://link.springer.
                                              com/chapter/10.1007/978-94-009-
                                              9349-5_21.

Cohen:1979:QMC

[CS79-39] Robert S. Cohen and John J. Stachel.

Questions of method in the consistency problem of quantum mechanics [1968f]. In Cohen and Stachel [CS79a], chapter 41, pages 609–613. ISBN 90-277-0651-4, 90-277-0652-2 (paperback), 94-009-9349-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 21 QC7.

URL https://link.springer.com/
chapter/10.1007/978-94-009-9349-
5_41.

Cohen:1979:RA

[CS79-40] Robert S. Cohen and John J. Stachel.

Rationalism in antiquity [1954c]. In Cohen and Stachel [CS79a], chapter 4, pages 29–31. ISBN 90-277-0651-4, 90-277-0652-2 (paperback), 94-009-9349-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 21 QC7. URL https://link.

springer.com/chapter/10.1007/
978-94-009-9349-5_4.

Cohen:1979:SHR

[CS79-41] Robert S. Cohen and John J. Stachel.

Science in history (review of J. D. Bernal’s Science in History [Watts & Co., London, 1954. xxiv + 967 pp.]) [1956f]. In Cohen and Stachel [CS79a], chapter 2, pages 7–15. ISBN 90-277-0651-4, 90-277-0652-2 (paperback), 94-009-9349-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 21 QC7.

URL https://link.springer.com/
chapter/10.1007/978-94-009-9349-
5_2.

Cohen:1979:SRK

[CS79-42] Robert S. Cohen and John J. Stachel.

Some reflections on knowledge [1971e].

In Cohen and Stachel [CS79a], chapter 43, pages 637–642. ISBN 90-277-0651-4, 90-277-0652-2 (paperback), 94-009-9349-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 21 QC7. URL https:

//link.springer.com/chapter/10.
1007/978-94-009-9349-5_43.

Cohen:1979:SCA

[CS79-43] Robert S. Cohen and John J. Stachel. Statistical causality in atomic theory (a general introduction to irreversibility) [1972d and 1974a]. In Cohen and Stachel [CS79a], chapter 38, pages 547– 570. ISBN 90-277-0651-4, 90-277-0652-2 (paperback), 94-009-9349-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 21

QC7. URL https://link.springer.
com/chapter/10.1007/978-94-009-
9349-5_38.

Cohen:1979:SAC

[CS79-44] Robert S. Cohen and John J. Stachel. Strife about complementarity [1953c]. In Cohen and Stachel [CS79a], chapter 31, pages 465–483. ISBN 90-277-0651-4, 90-277-0652-2 (paperback), 94-009-9349-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 21 QC7. URL https:

//link.springer.com/chapter/10.
1007/978-94-009-9349-5_31.

Cohen:1979:TAC

[CS79-45] Robert S. Cohen and John J. Stachel. The transformations of the atomic concept through the ages [1969f]. In Cohen and Stachel [CS79a], chapter 5, pages 32–44. ISBN 90-277-0651-4, 90-277-0652-2 (paperback), 94-009-9349-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 21 QC7. URL https://link.

springer.com/chapter/10.1007/
978-94-009-9349-5_5.

Cohen:1979:UCC

[CS79-46] Robert S. Cohen and John J. Stachel.

Unphilosophical considerations on causality in physics [1971b]. In Cohen and Stachel [CS79a], chapter 46, pages 666–680. ISBN 90-277-0651-4, 90-277-0652-2 (paperback), 94-009-9349-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 21 QC7. URL https://link.

springer.com/chapter/10.1007/
978-94-009-9349-5_46.

Cohen:1979:VLE

[CS79-47] Robert S. Cohen and John J. Stachel.

The velocity of light and the evolution of electrodynamics [1956a]. In Cohen and Stachel [CS79a], chapter 14, pages 134–177. ISBN 90-277-0651-4, 90-277-0652-2 (paperback), 94-009-9349-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 21 QC7. URL https://link.

springer.com/chapter/10.1007/
978-94-009-9349-5_14.

Cohen:1986:CFM

[CS86] Robert S. Cohen and Thomas Schnelle.

Cognition and Fact: Materials on Ludwik Fleck, volume 87 of Boston Studies in the Philosophy of Science. D. Reidel, Dordrecht, The Netherlands; Boston, MA, USA; Lancaster, UK; Tokyo, Japan, 1986. ISBN 94-010-8504-8, 94-009-4498-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). xxxiii + 468 pp. LCCN ???? URL

https://link.springer.com/book/
10.1007/978-94-009-4498-5.

Cobb-Stevens:2002:HHM

[CS02]  Richard Cobb-Stevens.  Husserlian

hermeneutics: Mathematics and Theoria. In Hermeneutic Philosophy of Science, Van Gogh’s Eyes, and God: Essays in Honor of Patrick A. Heelan, S.J. [Bab02c], chapter 12, pages 153–161. ISBN 90-481-5926-1, 94-017-1767-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 225. URL https:/

/link.springer.com/chapter/10.
1007/978-94-017-1767-0_12.

Csiszar:2015:OP

[Csi15] Alex Csiszar. Objectivities in print. In Objectivity in Science: New Perspectives from Science and Technology Studies [PRT15], chapter 8, pages 145– 169. ISBN 3-319-14348-4 (hardcover), 3-319-14349-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175 O25 2015. URL https:

//link.springer.com/chapter/10.
1007/978-3-319-14349-1_8.

Cohen:1974:DSS

[CSW74] Robert S. Cohen, John J. Stachel, and Marx W. Wartofsky. For Dirk Struik: Scientific, Historical and Political Essays in Honor of Dirk J. Struik, volume 15 of Boston Studies in the Philosophy of Science. D. Reidel, Dordrecht, The Netherlands; Boston, MA, USA; Lancaster, UK; Tokyo, Japan, 1974. ISBN 90-277-0379-5, 94-010-2115-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). xxvii + 657 pp. LCCN ???? URL

https://link.springer.com/book/
10.1007/978-94-010-2115-9.

Chernyak:1991:DSI

[CT91] Leon Chernyak and Alfred I. Tauber. The dialectical self: Immunology’s contribution. In Organism and the Origins of Self [Tau91i], chapter 6, pages 109–156. ISBN 0-7923-1185-X, 94-011-3406-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QR180-189.5. URL https:/

/link.springer.com/chapter/10.
1007/978-94-011-3406-4_6.

Cohen:1998:PNH

[CT98] Robert S. Cohen and Alfred I. Tauber.

Philosophies of Nature: The Human Dimension: In Celebration of Erazim Kohák, volume 195 of Boston Studies in the Philosophy of Science. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 1998. ISBN 90-481-4859-6, 94-017-2614-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). xvii + 334 pp. LCCN BD581. URL

https://link.springer.com/book/
10.1007/978-94-017-2614-6.

Cohen-Tannoudji:2009:PCP

[CT09] Gilles Cohen-Tannoudji. Philosophy and 20th century physics. In Brenner and Gayon [BG09], chapter 5, pages 115–140. ISBN 1-4020-9367-5, 1-4020-9368-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .F746 2009. URL https://link.

springer.com/chapter/10.1007/
978-1-4020-9368-5_5.

Cueto:2001:NHH

[Cue01] Marcos Cueto. Natural history, highaltitude physiology and evolutionary ideas in Peru. In The Reception of Darwinism in the Iberian World: Spain, Spanish America and Brazil [GPSR01], chapter 5, pages 83–93. ISBN 94-010-3885-6, 94-010-0602-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-010-0602-
6_5.

Cullen:2011:RTT

[Cul11] Christopher Cullen. Reflections on the transmission and transformation of technologies: Agriculture, printing and gunpowder between East and West. In Science between Europe and Asia: Historical Studies on the Transmission, Adoption and Adaptation of Knowledge [GR11b], chapter 2, pages 13–26. ISBN 90-481-9968-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q125 .S354 2010; Q175.52.A78 S35 2010. URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-90-481-9968-6_2.

Curd:1980:LDA

[Cur80] Martin V. Curd. The logic of discovery: An analysis of three approaches. In Scientific Discovery, Logic, and Rationality [Nic80f], chapter 8, pages 201–219. ISBN 90-277-1070-8, 94-009-8986-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

                                              ???? URL https://link.springer.
                                              com/chapter/10.1007/978-94-009-
                                              8986-3_8.

Currie:1989:FPT

[Cur89a] G. Currie. Frege and Popper: Two critics of psychologism. In Imre Lakatos and Theories of Scientific Change [GGN89], chapter 30, pages 413–430. ISBN 94-010-7860-2, 94-009-3025-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-009-3025-
                                              4_30.

Currie:1989:MI

[Cur89b] Gregory Currie. Metaphysical individualism. In Freedom and Rationality: Essays in Honor of John Watkins From his Colleagues and Friends [DJ89], chapter 4, pages 47–65. ISBN 94-010-7571-9, 94-009-2380-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-009-2380-5_4.

Cushing:1993:UCR

[Cus93] James T. Cushing. Underdetermination, conventionalism and realism: The Copenhagen vs. the Bohm interpretation of quantum mechanics. In Correspondence, Invariance and Heuristics: Essays in Honour of Heinz Post [FK93], chapter 13, pages 261–278. ISBN 90-481-4229-6, 94-017-1185-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-017-
1185-2_13.

Cushing:1994:BRB

[Cus94] James T. Cushing. A Bohmian response to Bohr’s complementarity. In Faye and Folse [FF94], chapter 3, pages 57–75. ISBN 0-7923-2378-5 (hardcover), 94-015-8106-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 153 1994; QC16 1994.

URL https://link.springer.com/
chapter/10.1007/978-94-015-8106-
6_3.

Cushing:1996:CQT

[Cus96] James T. Cushing. The causal quantum theory program. In Bohmian Mechanics and Quantum Theory: an Appraisal [CFG96], chapter 1, pages 1–19. ISBN 90-481-4698-4, 94-015-8715-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC173.96-174.52.

URL https://link.springer.com/
chapter/10.1007/978-94-015-8715-
0_1.

Cohen:1967:PBC

[CW67] Robert Sonné Cohen and Marx W. Wartofsky, editors. Proceedings of the Boston colloquium for the philosophy of science 1964/1966, in memory of Norwood Russell Hanson, volume 3 of Boston Studies in the Philosophy of Science. D. Reidel, Dordrecht, The Netherlands; Boston, MA, USA; Lancaster, UK; Tokyo, Japan, 1967. ISBN 90-277-0013-3, 94-010-3508-3. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175 .B731 v.3.

URL https://link.springer.com/
book/10.1007/978-94-010-3508-8.

Cohen:1969:BSP

[CW69a] Robert S. Cohen and Marx W. Wartofsky, editors. Boston Studies in the Philosophy of Science: Proceedings of the Boston Colloquium for the Philosophy of Science 1966/1968, volume 5 of Boston Studies in the Philosophy of Science. D. Reidel, Dordrecht, The Netherlands; Boston, MA, USA; Lancaster, UK; Tokyo, Japan, 1969. ISBN 94-010-3383-8, 94-010-3263-7, 94-010-3381-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL

https://link.springer.com/book/
10.1007/978-94-010-3381-7.

Cohen:1969:PBC

[CW69b] Robert S. Cohen and Marx W. Wartofsky, editors. Proceedings of the Boston Colloquium for the Philosophy of Science 1966/1968, volume 4 of Boston Studies in the Philosophy of Science. D. Reidel, Dordrecht, The Netherlands; Boston, MA, USA; Lancaster, UK; Tokyo, Japan, 1969. ISBN 94-010-3380-3, 94-010-3378-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL

https://link.springer.com/book/
10.1007/978-94-010-3378-7.

Cohen:1974:LES

[CW74a] Robert S. Cohen and Marx W. Wartofsky. Logical and Epistemological Studies in Contemporary Physics, volume 13 of Boston Studies in the Philosophy of Science. D. Reidel, Dordrecht, The Netherlands; Boston, MA, USA; Lancaster, UK; Tokyo, Japan, 1974. ISBN 90-277-0391-4 (hardcover), 90-277-0377-9 (paperback), 94-010-2656-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). viii + 462 pp. LCCN Q174 .B67 vol. 13; QC5.56.

URL https://link.springer.com/
book/10.1007/978-94-010-2656-7.

Cohen:1974:MHE

[CW74b] Robert S. Cohen and Marx W. Wartofsky. Methodological and Historical Essays in the Natural and Social Sciences, volume 14 of Boston Studies in the Philosophy of Science. D. Reidel, Dordrecht, The Netherlands; Boston, MA, USA; Lancaster, UK; Tokyo, Japan, 1974. ISBN 90-277-0378-7, 94-010-2128-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). viii + 405 pp. LCCN Q174 .B67 vol. 14. URL

https://link.springer.com/book/
10.1007/978-94-010-2128-9.

Cohen:1983:EMS

[CW83a] Robert S. Cohen and Marx W. Wartofsky. Epistemology, Methodology, and the Social Sciences, volume 71 of Boston Studies in the Philosophy of Science. D. Reidel, Dordrecht, The Netherlands; Boston, MA, USA; Lancaster, UK; Tokyo, Japan, 1983. ISBN 90-481-8376-6, 94-017-1458-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). viii + 270 pp. LCCN ????

                                              URL https://link.springer.com/
                                              book/10.1007/978-94-017-1458-7.

Cohen:1983:LLM

[CW83b] Robert S. Cohen and Marx W. Wartofsky. Language, Logic and Method, volume 31 of Boston Studies in the Philosophy of Science. D. Reidel, Dordrecht, The Netherlands; Boston, MA, USA; Lancaster, UK; Tokyo, Japan, 1983. ISBN 94-009-7704-2, 94-009-7702-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). 484 (est.) pp. LCCN ???? URL

                                              https://link.springer.com/book/
                                              10.1007/978-94-009-7702-0.

Cohen:1984:HS

[CW84a] Robert S. Cohen and Marx W. Wartofsky. Hegel and the Sciences, volume 64 of Boston Studies in the Philosophy of Science. D. Reidel, Dordrecht, The Netherlands; Boston, MA, USA; Lancaster, UK; Tokyo, Japan, 1984. ISBN 94-009-6235-5, 94-009-6233-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). ix + 374 pp. LCCN ????

                                              URL https://link.springer.com/
                                              book/10.1007/978-94-009-6233-0.

Cohen:1984:MMH

[CW84b] Robert S. Cohen and Marx W. Wartofsky. Methodology, Metaphysics and the History of Science: In Memory of Benjamin Nelson, volume 84 of Boston Studies in the Philosophy of Science. D. Reidel, Dordrecht, The Netherlands; Boston, MA, USA; Lancaster, UK; Tokyo, Japan, 1984. ISBN 94-009-6333-5, 94-009-6331-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). x + 373 pp. LCCN ????

                                              URL https://link.springer.com/
                                              book/10.1007/978-94-009-6331-3.

Cohen:1984:PSH

[CW84c] Robert Sonné Cohen and Mark W.

Wartofsky. Physical Sciences and History of Physics, volume 82 of Boston Studies in the Philosophy of Science. D. Reidel, Dordrecht, The Netherlands; Boston, MA, USA; Lancaster, UK; Tokyo, Japan, 1984. ISBN 94-009-7180-X, 94-009-7178-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). x + 260 pp. LCCN Q174 .P497 1984. URL

https://link.springer.com/book/
10.1007/978-94-009-7178-3.

Cohen:1985:PTF

[CW85] Robert S. Cohen and Marx W. Wartofsky, editors. A Portrait of Twenty-five Years: Boston Colloquium for the Philosophy of Science 1960–1985, volume 89 of Boston Studies in the Philosophy of Science. D. Reidel, Dordrecht, The Netherlands; Boston, MA, USA; Lancaster, UK; Tokyo, Japan, 1985. ISBN 90-277-1971-3, 94-009-5345-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
book/10.1007/978-94-009-5345-1.

Chen:2018:ECE

[CW18] Jia Chen and Jian Wang. Early Chinese engineering education: Influence and disappearance of the Fuzhou Shipping School. In Mitcham et al. [MLNZ18], chapter 14, pages 159–171. ISBN 3-319-62448-2, 3-319-62450-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN TA15 .P45 2018.

URL https://link.springer.com/
chapter/10.1007/978-3-319-62450-
1_14.

Czarnocka:1982:DRC

[CŻ82] Małgorzata Czarnocka and Jan M.

Żytkow. Difficulties with the reduction of classical to relativistic mechanics. In Polish Essays in the Philosophy of the Natural Sciences [Kra82c], chapter 28, pages 319–332. ISBN 94-009-7705-0 (ebook), 90-277-1287-5. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-009-
7705-1_28.

DAmour:1976:RPR

[D’A76] Gene D’Amour. Research programs, rationality, and ethics. In Essays in Memory of Imre Lakatos [CFW76], chapter 9, pages 87–98. ISBN 90-277-0655-7, 94-010-1451-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-010-1451-9_9.

DAgostino:1981:WMD

[D’A81] Salvo D’Agostino. Weber and Maxwell on the discovery of the velocity of light in Nineteenth Century electrodynamics. In Grmek et al. [GCC81b], pages 281–293. ISBN 90-277-1123-2, 94-010-1284-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-010-
1284-3_14.

DAgostino:1989:OAS

[D’A89] Fred D’Agostino. The ‘optimum’ aim for science. In Freedom and Rationality: Essays in Honor of John Watkins From his Colleagues and Friends [DJ89], chapter 15, pages 247–256. ISBN 94-010-7571-9, 94-009-2380-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-009-2380-5_15.

DAmbrosio:1992:NHA

[D’A92] Ubiratán D’Ambrosio. For a new historiographical approach of the so-called “traditional knowledge”. In Science and Empires: Historical Studies about Scientific Development and European Expansion [PJM92], chapter 3, pages 15–17. ISBN 0-7923-1518-9 (hardcover), 94-010-5145-3, 94-011-2594-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 136.

URL https://link.springer.com/
chapter/10.1007/978-94-011-2594-
9_3.

DAgostino:1995:DAR

[D’A95] Fred D’Agostino. A deontological approach to the rationality of science. In Misiek [Mis95c], pages 19–43. ISBN 94-010-4206-3, 94-011-0461-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-011-0461-6_2.

DAgostino:1998:HVM

[D’A98] Salvo D’Agostino. Hertz’s views on the methods of physics: Experiment and theory reconciled? In Heinrich Hertz: Classical Physicist, Modern Philosopher [BHN98], chapter 7, pages 89–102. ISBN 90-481-4881-2, 94-015-8855-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-015-8855-
3_7.

DAgostino:2000:HIT

[D’A00] Salvo D’Agostino. A History of the Ideas of Theoretical Physics: Essays on the Nineteenth and Twentieth Century Physics, volume 213 of Boston Studies in the Philosophy of Science. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 2000. ISBN 0-7923-6094-X, 94-010-9034-3. ISSN 0068-0346 (print), 2214-7942 (electronic). xviii + 381 pp. LCCN Q174 .B67 no. 213. URL

https://link.springer.com/book/
10.1007/978-94-010-9034-6.

DAgostino:2001:C

[D’A01a] Salvo D’Agostino. Conclusions. In A history of the ideas of theoretical physics: essays on the nineteenth and twentieth century physics [D’A01l], chapter 14, pages 303–310. ISBN 1-4020-0244-0, 94-010-9034-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC5.53; QC6.

URL https://link.springer.com/
chapter/10.1007/978-94-010-9034-
6_14.

DAgostino:2001:CCR

[D’A01b] Salvo D’Agostino. A consideration on the changing role of mathematics in Ampère’s and Weber’s electrodynamics. In A history of the ideas of theoretical physics: essays on the nineteenth and twentieth century physics [D’A01l], chapter 1, pages 7–37. ISBN 1-4020-0244-0, 94-010-9034-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC5.53; QC6.

URL https://link.springer.com/
chapter/10.1007/978-94-010-9034-
6_1.

DAgostino:2001:CCN

[D’A01c] Salvo D’Agostino. Correspondence and complementarity in Niels Bohr’s papers: 1925–1927. In A history of the ideas of theoretical physics: essays on the nineteenth and twentieth century physics [D’A01l], chapter 12, pages 253–271. ISBN 1-4020-0244-0, 94-010-9034-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC5.53; QC6. URL https://link.

springer.com/chapter/10.1007/
978-94-010-9034-6_12.

DAgostino:2001:ECC

[D’A01d] Salvo D’Agostino. Einstein’s correspondence criterion and the construction of general relativity. In A history of the ideas of theoretical physics: essays on the nineteenth and twentieth century physics [D’A01l], chapter 10, pages 223–237. ISBN 1-4020-0244-0, 94-010-9034-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC5.53; QC6. URL https://link.

springer.com/chapter/10.1007/
978-94-010-9034-6_10.

DAgostino:2001:ELL

[D’A01e] Salvo D’Agostino. Einstein’s life-long doubts on the physical foundations of the general relativity and unified field theories. In A history of the ideas of theoretical physics: essays on the nineteenth and twentieth century physics [D’A01l], chapter 11, pages 239–252. ISBN 1-4020-0244-0, 94-010-9034-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC5.53;

QC6. URL https://link.springer.
com/chapter/10.1007/978-94-010-
9034-6_11.

DAgostino:2001:FTP

[D’A01f] Salvo D’Agostino. A foundation for theoretical physics in Hertz’s introduction to Die Prinzipien der Mechanik. In A history of the ideas of theoretical physics: essays on the nineteenth and twentieth century physics [D’A01l], chapter 8, pages 187–199. ISBN 1-4020-0244-0, 94-010-9034-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC5.53; QC6.

URL https://link.springer.com/

                                              chapter/10.1007/978-94-010-9034-
                                              6_8.

DAgostino:2001:WMS

[D’A01g] Salvo D’Agostino. From the 1926 wave mechanics to a second-quantisation theory: Schrödinger’s new interpretation of wave mechanics and microphysics in the 1950’s. In A history of the ideas of theoretical physics: essays on the nineteenth and twentieth century physics [D’A01l], chapter 13, pages 273–302. ISBN 1-4020-0244-0, 94-010-9034-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC5.53;

                                              QC6. URL https://link.springer.
                                              com/chapter/10.1007/978-94-010-
                                              9034-6_13.

DAgostino:2001:GE

[D’A01h] Salvo D’Agostino. German electrodynamics in the 1870’s. In A history of the ideas of theoretical physics: essays on the nineteenth and twentieth century physics [D’A01l], chapter 5, pages 111–134. ISBN 1-4020-0244-0, 94-010-9034-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC5.53; QC6. URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-010-9034-6_5.

DAgostino:2001:HTD

[D’A01i] Salvo D’Agostino. Hertz’s 1884 theoretical discovery of electromagnetic waves. In A history of the ideas of theoretical physics: essays on the nineteenth and twentieth century physics [D’A01l], chapter 7, pages 167–185. ISBN 1-4020-0244-0, 94-010-9034-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC5.53;

                                              QC6. URL https://link.springer.
                                              com/chapter/10.1007/978-94-010-
                                              9034-6_7.

DAgostino:2001:HEE

[D’A01j] Salvo D’Agostino. Hertz’s experiments on electromagnetic waves. In A history of the ideas of theoretical physics: essays on the nineteenth and twentieth century physics [D’A01l], chapter 6, pages 135–166. ISBN 1-4020-0244-0, 94-010-9034-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC5.53; QC6. URL https:

//link.springer.com/chapter/10.
1007/978-94-010-9034-6_6.

DAgostino:2001:HRD

[D’A01k] Salvo D’Agostino. A historical role for dimensional analysis in Maxwell’s electromagnetic theory of light. In A history of the ideas of theoretical physics: essays on the nineteenth and twentieth century physics [D’A01l], chapter 3, pages 45–75. ISBN 1-4020-0244-0, 94-010-9034-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC5.53; QC6.

URL https://link.springer.com/
chapter/10.1007/978-94-010-9034-
6_3.

DAgostino:2001:HIT

[D’A01l] Salvo D’Agostino. A history of the ideas of theoretical physics: essays on the nineteenth and twentieth century physics, volume 213 of Boston Studies in the Philosophy of Science. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 2001. ISBN 1-4020-0244-0, 94-010-9034-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). xviii + 381 pp. LCCN QC5.53; QC6.

DAgostino:2001:BMH

[D’A01m] Salvo D’Agostino. On Boltzmann’s mechanics and his Bild-conception of physical theory. In A history of the ideas of theoretical physics: essays on the nineteenth and twentieth century physics [D’A01l], chapter 9, pages 201–216. ISBN 1-4020-0244-0, 94-010-9034-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC5.53; QC6. URL https://link.

springer.com/chapter/10.1007/
978-94-010-9034-6_9.

DAgostino:2001:PTP

[D’A01n] Salvo D’Agostino. Problems of theoretical physics in the second half of Nineteenth Century. In A history of the ideas of theoretical physics: essays on the nineteenth and twentieth century physics [D’A01l], chapter 4, pages 77–106. ISBN 1-4020-0244-0, 94-010-9034-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC5.53; QC6. URL https:

//link.springer.com/chapter/10.
1007/978-94-010-9034-6_4.

DAgostino:2001:STU

[D’A01o] Salvo D’Agostino. A survey of theories of units and dimensions in Nineteenth-Century physics. In A history of the ideas of theoretical physics: essays on the nineteenth and twentieth century physics [D’A01l], chapter 2, pages 39– 43. ISBN 1-4020-0244-0, 94-010-9034-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC5.53;

QC6. URL https://link.springer.
com/chapter/10.1007/978-94-010-
9034-6_2.

Dafu:1996:UBP

[Daf96] Ding Dafu. On the unity of biology and physics. In Chinese Studies in the History and Philosophy of Science and Technology [DC96], chapter 10, pages 103–113. ISBN 90-481-4546-5, 94-015-8717-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic).

LCCN Q174 .B67 vol. 180; U390 .N19

1996. URL https://link.springer.
com/chapter/10.1007/978-94-015-
8717-4_10.

Dahme:1990:CRG

[Dah90] Heinz-Jürgen Dahme. On the current rediscovery of Georg Simmel’s sociology — a European point of view. In Georg Simmel and Contemporary Sociology [KPC90], chapter 3, pages 13–37. ISBN 94-010-6691-4 (hardcover), 94-009-0459-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN HM401-1281. URL https://link.

springer.com/chapter/10.1007/
978-94-009-0459-0_3.

Dainian:1989:ILC

[Dai89] Fan Dainian. Imre Lakatos in China. In Imre Lakatos and Theories of Scientific Change [GGN89], chapter 5, pages 59– 67. ISBN 94-010-7860-2, 94-009-3025-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-009-3025-
4_5.

Dainian:1996:NBR

[Dai96] Fan Dainian. Niels Bohr and realism. In Cohen et al. [CHR96], pages 279–287. ISBN 90-481-4493-0, 94-015-8638-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-015-8638-
2_19.

Dainton:2016:SCI

[Dai16] Barry Dainton. Some cosmological implications of temporal experience. In Cosmological and Psychological Time [DR16], chapter 5, pages 75–105. ISBN 3-319-22589-8 (hardcover), 3-319-22590-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN BD638 C67 2016. URL https://link.

springer.com/chapter/10.1007/
978-3-319-22590-6_5.

Daladier:1989:IUF

[Dal89] Anne Daladier. Information units in a French corpus. In The Form of Information in Science: Analysis of an Immunology Sublanguage [HGR^{+}89], chapter 7, pages 170–191. ISBN 94-010-7777-0, 94-009-2837-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN B832.A-832.Z.

URL https://link.springer.com/
chapter/10.1007/978-94-009-2837-
4_7.

Damerow:1996:AR

[Dam96a] Peter Damerow. Abstraction and representation. In Abstraction and Representation: Essays on the Cultural Evolution of Thinking [Dam96b], chapter 10, pages 371–381. ISBN 0-7923-3816-2, 94-015-8624-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 175.

URL https://link.springer.com/
chapter/10.1007/978-94-015-8624-
5_10.

Damerow:1996:ARE

[Dam96b] Peter Damerow. Abstraction and Representation: Essays on the Cultural Evolution of Thinking, volume 175 of Boston Studies in the Philosophy of Science. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 1996. ISBN 0-7923-3816-2, 94-015-8624-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). xiv + 414 pp. LCCN Q174 .B67 vol. 175. URL

https://link.springer.com/book/
10.1007/978-94-015-8624-5.

Damerow:1996:ACP

[Dam96c] Peter Damerow. Action and cognition in Piaget’s genetic epistemology and in Hegel’s logic. In Abstraction and Representation: Essays on the Cultural Evolution of Thinking [Dam96b], chapter 1, pages 1–27. ISBN 0-7923-3816-2, 94-015-8624-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 175. URL https:

//link.springer.com/chapter/10.
1007/978-94-015-8624-5_1.

Damerow:1996:CLH

[Dam96d] Peter Damerow. The concept of labor in historical materialism and the theory of socio-historical development. In Abstraction and Representation: Essays on the Cultural Evolution of Thinking [Dam96b], chapter 11, pages 383–394. ISBN 0-7923-3816-2, 94-015-8624-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 175. URL https://link.

springer.com/chapter/10.1007/
978-94-015-8624-5_11.

Damerow:1996:DAT

[Dam96e] Peter Damerow. The development of arithmetical thinking: On the role of calculating aids in Ancient Egyptian and Babylonian arithmetic. In Abstraction and Representation: Essays on the Cultural Evolution of Thinking [Dam96b], chapter 7, pages 173–273. ISBN 0-7923-3816-2, 94-015-8624-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 175. URL https://link.

springer.com/chapter/10.1007/
978-94-015-8624-5_7.

Damerow:1996:MES

[Dam96f] Peter Damerow. Mathematics education and society. In Abstraction and Representation: Essays on the Cultural Evolution of Thinking [Dam96b], chapter 5, pages 111–148. ISBN 0-7923-3816-2, 94-015-8624-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 175.

URL https://link.springer.com/
chapter/10.1007/978-94-015-8624-
5_5.

Damerow:1996:RBO

[Dam96g] Peter Damerow. On the relationship between ontogenesis and historiogenesis of the number concept. In Abstraction and Representation: Essays on the Cultural Evolution of Thinking [Dam96b], chapter 9, pages 299–370. ISBN 0-7923-3816-2, 94-015-8624-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 175.

URL https://link.springer.com/
chapter/10.1007/978-94-015-8624-
5_9.

Damerow:1996:PPR

[Dam96h] Peter Damerow. Philosophical and pedagogical remarks on the concept “abstract”. In Abstraction and Representation: Essays on the Cultural Evolution of Thinking [Dam96b], chapter 3, pages 71–86. ISBN 0-7923-3816-2, 94-015-8624-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 175. URL https:

//link.springer.com/chapter/10.
1007/978-94-015-8624-5_3.

Damerow:1996:PRR

[Dam96i] Peter Damerow. Preliminary remarks on the relationship of the principles of teaching arithmetic to the early history of mathematics. In Abstraction and Representation: Essays on the Cultural Evolution of Thinking [Dam96b], chapter 6, pages 149–171. ISBN 0-7923-3816-2, 94-015-8624-1 (e-book).

ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 175.

URL https://link.springer.com/
chapter/10.1007/978-94-015-8624-
5_6.

Damerow:1996:RM

[Dam96j] Peter Damerow. Representation and meaning. In Abstraction and Representation: Essays on the Cultural Evolution of Thinking [Dam96b], chapter 2, pages 29–70. ISBN 0-7923-3816-2, 94-015-8624-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 175. URL https:

//link.springer.com/chapter/10.
1007/978-94-015-8624-5_2.

Damerow:1996:WMA

[Dam96k] Peter Damerow. What is mathematical ability and how do ability differences emerge in mathematics education? In Abstraction and Representation: Essays on the Cultural Evolution of Thinking [Dam96b], chapter 4, pages 87–110. ISBN 0-7923-3816-2, 94-015-8624-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 175. URL https://link.

springer.com/chapter/10.1007/
978-94-015-8624-5_4.

Martins:2007:ORM

[dAM07] Roberto de Andrade Martins. Ørsted, Ritter, and magnetochemistry. In Hans Christian Ørsted and The Romantic Legacy In Science: Ideas, Disciplines, Practices [BCK07], chapter 16, pages 339–385. ISBN 1-4020-2979-9 (hardcover), 1-4020-2987-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC515.O37 H36 2007.

URL https://link.springer.com/
chapter/10.1007/978-1-4020-2987-
5_16.

Martins:2011:HBR

[dAM11] Roberto de Andrade Martins. Henri Becquerel and radioactivity: A critical revision. In Brazilian Studies in Philosophy and History of Science: An account of recent works [KV11], chapter 6, pages 107–117. ISBN 90-481-9421-0 (hardcover), 90-481-9422-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175.3 .B73

                                              2011. URL https://link.springer.
                                              com/chapter/10.1007/978-90-481-
                                              9422-3_6.

Daniels:1976:IHH

[Dan76]  Norman Daniels.  IQ, heritability,

and human nature. In Cohen et al. [CHMV76], pages 143–180. ISBN 90-277-0647-6, 90-277-0648-4 (paperback), 94-010-1449-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic).

                                              LCCN ????  URL https://link.
                                              springer.com/chapter/10.1007/
                                              978-94-010-1449-6_7.

Dantes:1992:PSA

[Dan92] Maria Amelia M. Dantes. Le positivisme et la science au brésil. (French) [Positivism and science in Brazil]. In Science and Empires: Historical Studies about Scientific Development and European Expansion [PJM92], chapter 18, pages 165–172. ISBN 0-7923-1518-9 (hardcover), 94-010-5145-3, 94-011-2594-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 136.

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-011-2594-
                                              9_18.

Darden:1980:TCG

[Dar80] Lindley Darden. Theory construction in genetics. In Scientific Discovery: Case Studies [Nic80e], chapter 9, pages 151–170. ISBN 90-277-1093-7, 94-009-9015-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 60. URL https:/

/link.springer.com/chapter/10.
1007/978-94-009-9015-9_9.

Darrigol:2007:HK

[Dar07] Olivier Darrigol. For a history of knowledge. In Positioning the history of science [RG07], chapter 5, pages 33–34. ISBN 1-4020-5419-X (hardcover), 1-4020-5420-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q125 .P77 2007.

URL https://link.springer.com/
chapter/10.1007/1-4020-5420-3_5.

Dascal:1976:LMM

[Das76] Marcelo Dascal. Levels of meaning and moral discourse. In Language in Focus: Foundations, Methods and Systems: Essays in Memory of Yehoshua Bar-Hillel [Kas76b], chapter 25, pages 587–625. ISBN 90-277-0644-1, 90-277-0645-X, 94-010-1876-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN P101-120. URL https://link.

springer.com/chapter/10.1007/
978-94-010-1876-0_25.

Dascal:1988:RCE

[Das88] Marcelo Dascal. Rational-choice explanation: The limits to grounding. In Science in Reflection: The Israel Colloquium: Studies in History, Philosophy, and Sociology of Science Volume 3 [UM88], chapter 6, pages 67–79. ISBN 90-277-2713-9, 94-009-2957-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-009-2957-9_6.

Daston:1992:DCC

[Das91] Lorraine J. Daston. The doctrine of chances without chance: Determinism, mathematical probability, and quantification in the Seventeenth Century. In The Invention of Physical Science: Intersections of Mathematics, Theology and Natural Philosophy Since the Seventeenth Century: Essays in Honor of Erwin N. Hiebert [NRS92], chapter 2, pages 27–50. ISBN 94-010-5097-X, 94-011-2488-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

                                              ???? URL https://link.springer.
                                              com/chapter/10.1007/978-94-011-
                                              2488-1_2.

Daston:2007:WPW

[Das07a] Lorraine Daston. Working in parallel, working together. In Positioning the history of science [RG07], chapter 6, pages 35–38. ISBN 1-4020-5419-X (hardcover), 1-4020-5420-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q125 .P77 2007.

                                              URL https://link.springer.com/
                                              chapter/10.1007/1-4020-5420-3_6.

Daston:2007:ORU

[Das07b] Lorraine J. Daston. Ørsted and the rational unconscious. In Hans Christian Ørsted and The Romantic Legacy In Science: Ideas, Disciplines, Practices [BCK07], chapter 12, pages 235–246. ISBN 1-4020-2979-9 (hardcover), 1-4020-2987-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC515.O37 H36 2007. URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-1-4020-2987-5_12.

Daston:2015:CPJ

[Das15] Lorraine Daston. At the center and the periphery: Joseph Pitton de Tournefort botanizes in Crete. In Arabatzis et al.

[ARS15], chapter 7, pages 85–98. ISBN 3-319-14552-5 (paperback), 3-319-14553-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q125 .R45 2015. URL https://link.

springer.com/chapter/10.1007/
978-3-319-14553-2_7.

Daumer:1996:STB

[Dau96] Martin Daumer. Scattering theory from a Bohmian perspective. In Bohmian Mechanics and Quantum Theory: an Appraisal [CFG96], chapter 6, pages 87–97. ISBN 90-481-4698-4, 94-015-8715-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC173.96-174.52. URL https://link.

springer.com/chapter/10.1007/
978-94-015-8715-0_6.

Davis:1974:MM

[Dav74a] Chandler Davis. Materialist mathematics. In For Dirk Struik: Scientific, Historical and Political Essays in Honor of Dirk J. Struik [CSW74], chapter 5, pages 37–66. ISBN 90-277-0379-5, 94-010-2115-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-010-
2115-9_5.

Davis:1974:EBN

[Dav74b] Horace B. Davis. The evolution of Black nationalism (1971). In For Dirk Struik: Scientific, Historical and Political Essays in Honor of Dirk J. Struik [CSW74], chapter 39, pages 501–529. ISBN 90-277-0379-5, 94-010-2115-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-010-2115-
9_39.

Davis:1986:MGF

[Dav86] Bernard D. Davis. Molecular genetics and the falsifiability of evolution. In Ullmann-Margalit [UM86a], chapter 9, pages 95–110. ISBN 90-277-2159-9 (paperback), 90-277-2158-0 (hardcover), 94-009-5496-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 94 Q175.

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-009-5496-
                                              0_9.

Davies:1995:MTA

[Dav95] David Davies. The model–theoretic argument unlocked. In Québec Studies in the Philosophy of Science: Part I: Logic, Mathematics, Physics and History of Science Essays in Honor of Hugues Leblanc [MC95], chapter 16, pages 275–286. ISBN 94-010-7204-3, 94-009-1575-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

                                              ???? URL https://link.springer.
                                              com/chapter/10.1007/978-94-009-
                                              1575-6_16.

David:1997:ACQ

[Dav97] Marian David. Analyticity, Carnap, Quine, and truth. In Austrian Philosophy Past and Present: Essays in Honor of Rudolf Haller [LM97], chapter 14, pages 203–219. ISBN 94-010-6412-1, 94-011-5720-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-011-5720-9_14.

Davey:1999:NHQ

[Dav99] Nicholas Davey. Nietzsche, Habermas, and the question of objectivity. In Nietzsche, Theories of Knowledge, and Critical Theory: Nietzsche and the Sciences I [Bab99b], chapter 23, pages 295–306. ISBN 90-481-5233-X, 94-017-2430-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-017-
2430-2_23.

Davies:2010:LTF

[Dav10] David Davies. Learning through fictional narratives in art and science. In Frigg and Hunter [FH10b], chapter 4, pages 51–69. ISBN 90-481-3850-7 (hardcover), 90-481-3851-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN BH301.M47 B49 2010; Q175 .B52 2010. URL https:

//link.springer.com/chapter/10.
1007/978-90-481-3851-7_4.

Davis:2017:UU

[Dav17] Martin Davis. Universality is ubiquitous. In Floyd and Bokulich [FB17], chapter 6, pages 153–158. ISBN 3-319-53278-2, 3-319-53280-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QA29.T8 P45 2017.

URL https://link.springer.com/
chapter/10.1007/978-3-319-53280-
6_6.

deBeauregard:1984:TSI

[dB84] O. Costa de Beauregard. Time symmetry and interpretation of quantum mechanics. In Physical Sciences and History of Physics [CW84c], chapter 3, pages 35–56. ISBN 94-009-7180-X, 94-009-7178-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .P497 1984. URL https:

//link.springer.com/chapter/10.
1007/978-94-009-7178-3_3.

deBeauregard:1987:QSS

[dB87a] Olivier Costa de Beauregard. 1900– 1925: The quantum springs out, and spreads. In Time, The Physical Magnitude [Cos87b], chapter 14, pages 185–195. ISBN 94-010-8195-6, 94-009-3811-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-009-
3811-3_14.

deBeauregard:1987:DQM

[dB87b] Olivier Costa de Beauregard. 1925– 1927: The dawn of quantum mechanics with a shadow: Relativistic covariance lost. In Time, The Physical Magnitude [Cos87b], chapter 15, pages 196–218. ISBN 94-010-8195-6, 94-009-3811-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-009-3811-
3_15.

deBeauregard:1987:QMQ

[dB87c] Olivier Costa de Beauregard. 1927– 1949: From quantum mechanics to quantum field theory: Relativistic covariance slowly recovered. In Time, The Physical Magnitude [Cos87b], chapter 16, pages 219–232. ISBN 94-010-8195-6, 94-009-3811-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-009-3811-3_16.

deBeauregard:1987:EFD

[dB87d] Olivier Costa de Beauregard. Entering the four-dimensional spacetime paradigm. In Time, The Physical Magnitude [Cos87b], chapter 6, pages 53–66. ISBN 94-010-8195-6, 94-009-3811-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-009-
3811-3_6.

deBeauregard:1987:IR

[dB87e] Olivier Costa de Beauregard. Introductory remarks. In Time, The Physical Magnitude [Cos87b], chapter 1, pages 3–9. ISBN 94-010-8195-6, 94-009-3811-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-009-3811-
3_1.

deBeauregard:1987:ICP

[dB87f] Olivier Costa de Beauregard. Irreversibility as a cosmic phenomenon. In Time, The Physical Magnitude [Cos87b], chapter 11, pages 142–149. ISBN 94-010-8195-6, 94-009-3811-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-009-3811-
3_11.

deBeauregard:1987:LRF

[dB87g] Olivier Costa de Beauregard. Lawlike reversibility and factlike irreversibility in the negentropy-information transition. In Time, The Physical Magnitude [Cos87b], chapter 12, pages 150–163. ISBN 94-010-8195-6, 94-009-3811-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-009-3811-
3_12.

deBeauregard:1987:MSG

[dB87h] Olivier Costa de Beauregard. The magic of spacetime geometry. In Time, The Physical Magnitude [Cos87b], chapter 7, pages 67–94. ISBN 94-010-8195-6, 94-009-3811-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-009-3811-3_7.

deBeauregard:1987:MTT

[dB87i] Olivier Costa de Beauregard. The more than two millennia of Euclidean geometry. In Time, The Physical Magnitude [Cos87b], chapter 2, pages 13–16. ISBN 94-010-8195-6, 94-009-3811-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-009-3811-3_2.

deBeauregard:1987:GR

[dB87j] Olivier Costa de Beauregard. On general relativity. In Time, The Physical Magnitude [Cos87b], chapter 20, pages 287–294. ISBN 94-010-8195-6, 94-009-3811-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

                                              ???? URL https://link.springer.
                                              com/chapter/10.1007/978-94-009-
                                              3811-3_20.

deBeauregard:1987:OLC

[dB87k] Olivier Costa de Beauregard. An outsider’s look at cosmology, and overall conclusions. In Time, The Physical Magnitude [Cos87b], chapter 21, pages 295–298. ISBN 94-010-8195-6, 94-009-3811-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

                                              ???? URL https://link.springer.
                                              com/chapter/10.1007/978-94-009-
                                              3811-3_21.

deBeauregard:1987:Oa

[dB87l] Olivier Costa de Beauregard. Overview. In Time, The Physical Magnitude [Cos87b], chapter 8, pages 97–118. ISBN 94-010-8195-6, 94-009-3811-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

                                              URL https://link.springer.com/

chapter/10.1007/978-94-009-3811-
3_8.

deBeauregard:1987:Ob

[dB87m] Olivier Costa de Beauregard. Overview.

In Time, The Physical Magnitude [Cos87b], chapter 13, pages 167–184. ISBN 94-010-8195-6, 94-009-3811-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-009-3811-
3_13.

deBeauregard:1987:PPE

[dB87n] Olivier Costa de Beauregard. Paradox and paradigm: The Einstein– Podolsky–Rosen correlations. In Time, The Physical Magnitude [Cos87b], chapter 18, pages 244–264. ISBN 94-010-8195-6, 94-009-3811-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-009-3811-3_18.

deBeauregard:1987:PVC

[dB87o] Olivier Costa de Beauregard. Parity violations and CPT invariance. In Time, The Physical Magnitude [Cos87b], chapter 17, pages 233–243. ISBN 94-010-8195-6, 94-009-3811-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-009-3811-3_17.

deBeauregard:1987:PI

[dB87p] Olivier Costa de Beauregard. Phenomenological irreversibility. In Time, The Physical Magnitude [Cos87b], chapter 9, pages 119–128. ISBN 94-010-8195-6, 94-009-3811-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-009-3811-3_9.

deBeauregard:1987:RCS

[dB87q] Olivier Costa de Beauregard. Retarded causality as a statistical concept. Arrowless microcausality. In Time, The Physical Magnitude [Cos87b], chapter 10, pages 129–141. ISBN 94-010-8195-6, 94-009-3811-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-009-3811-3_10.

deBeauregard:1987:TCK

[dB87r] Olivier Costa de Beauregard. Three centuries of kinematical optics. In Time, The Physical Magnitude [Cos87b], chapter 4, pages 28–41. ISBN 94-010-8195-6, 94-009-3811-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-009-3811-3_4.

deBeauregard:1987:TCN

[dB87s] Olivier Costa de Beauregard. The three centuries of Newtonian mechanics: Universal time and absolute space. In Time, The Physical Magnitude [Cos87b], chapter 3, pages 17–27. ISBN 94-010-8195-6, 94-009-3811-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-009-3811-3_3.

deBeauregard:1987:TNP

[dB87t] Olivier Costa de Beauregard. Todays NEC plus ultra of metrology and chronometry: ‘equivalence’ of space and time. In Time, The Physical Magnitude [Cos87b], chapter 5, pages 42–52. ISBN 94-010-8195-6, 94-009-3811-X (e-book).

ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-009-3811-3_5.

deBeauregard:2000:EFC

[dB00] Olivier Costa de Beauregard. Efficient and final cause as CPT reciprocals. In The Reality of the Unobservable: Observability, Unobservability and Their Impact on the Issue of Scientific Realism [AP00b], chapter 21, pages 283–291. ISBN 90-481-5458-8, 94-015-9391-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-015-
9391-5_21.

deBuzon:2012:BDP

[dB12]  Frédéric de Buzon.  Beeckman,

Descartes and physico-mathematics. In Philosophy of Behavioral Biology [PR12b], chapter 6, pages 143–158. ISBN 94-007-1950-7, 94-007-1951-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QH331 .P4679 2012; Q175 .B657 v.282.

URL https://link.springer.com/
chapter/10.1007/978-94-007-4345-
8_6.

deBuzon:2013:BDP

[dB13]  Frédéric de Buzon.  Beeckman,

Descartes and physico–mathematics. In Garber and Roux [GR13b], chapter 6, pages 143–158. CODEN ???? ISBN 94-007-4344-0 (hardcover), 94-007-4345-9 (e-book). ISSN ???? LCCN QC6 .M3616 2013; QC6 .M43 2013.

URL https://link.springer.com/
chapter/10.1007/978-94-007-4345-
8_6.

Devlin:2015:I

[DB15a] William J. Devlin and Alisa Bokulich. Introduction. In Kuhn’s Structure of Scientific Revolutions — 50 Years On [DB15b], chapter 1, pages 1–9. ISBN 3-319-13382-9, 3-319-13383-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175.3 .K84 2015.

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-3-319-13383-
                                              6_1.

Devlin:2015:KSS

[DB15b] William J. Devlin and Alisa Bokulich. Kuhn’s Structure of Scientific Revolutions — 50 Years On, volume 311 of Boston Studies in the Philosophy and History of Science. Springer-Ver-lag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2015. ISBN 3-319-13382-9, 3-319-13383-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). xi + 199 pp. LCCN Q175.3 .K84 2015. URL

                                              https://link.springer.com/book/
                                              10.1007/978-3-319-13383-6.

deBruijn:1995:PPT

[dBJD95] F. J. de Bruijn, Y. Jing, and F. B. Dazzo. Potential and pitfalls of trying to extend symbiotic interactions of nitrogen-fixing organisms to presently non-nodulated plants, such as rice. In No Truth Except in the Details: Essays in Honor of Martin J. Klein [KS95], chapter 11, pages 225–240. ISBN 94-010-4097-4, 94-011-0217-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-011-0055-7_11.

deCastro:1995:RHS

[dC95] Max Fernandez de Castro. On the relation of Hilbert’s Second and Tenth Problem. In Mexican Studies in the History and Philosophy of Science [RC95], chapter 12, pages 187–199. ISBN 94-010-6535-7 (hardcover), 94-009-0109-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-009-
0109-4_12.

Dainian:1996:CSH

[DC96] Fan Dainian and Robert S. Cohen.

Chinese Studies in the History and Philosophy of Science and Technology, volume 179 of Boston Studies in the Philosophy of Science. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 1996. ISBN 90-481-4546-5, 94-015-8717-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). xiii + 483 pp. LCCN Q174 .B67 vol. 180; U390 .N19 1996. URL

https://link.springer.com/book/
10.1007/978-94-015-8717-4.

daCosta:2011:QTQ

[dCB11] Newton C. A. da Costa and Otávio Bueno. Quasi-truth and quantum mechanics. In Brazilian Studies in Philosophy and History of Science: An account of recent works [KV11], chapter 23, pages 301–312. ISBN 90-481-9421-0 (hardcover), 90-481-9422-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175.3 .B73

2011. URL https://link.springer.
com/chapter/10.1007/978-90-481-
9422-3_23.

Diogo:2015:VSR

[DCS15] Maria Paula Diogo, Ana Carneiro, and Ana Simões. Visions of science: Research at the Faculty of Sciences of the University of Lisbon seen through its journal. In Sciences in the Universities of Europe, Nineteenth and Twentieth Centuries: Academic Landscapes [SDG15b], chapter 16, pages 265–289. ISBN 94-017-9635-1, 94-017-9636-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q183.4.E85

                                              2015. URL https://link.springer.
                                              com/chapter/10.1007/978-94-017-
                                              9636-1_16.

Dreyfus:1986:SES

[DD86] Hubert L. Dreyfus and Stuart E. Dreyfus. From Socrates to expert systems: The limits of calculative rationality. In Philosophy and Technology II: Information Technology and Computers in Theory and Practice [MH86], chapter 9, pages 111–130. ISBN 94-010-8510-2, 94-009-4512-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

                                              ???? URL https://link.springer.
                                              com/chapter/10.1007/978-94-009-
                                              4512-8_9.

Doria:2011:EGM

[DD11] Francisco Antonio Doria and Manuel Doria. Einstein, Gödel, and the mathematics of time. In Brazilian Studies in Philosophy and History of Science: An account of recent works [KV11], chapter 20, pages 269–279. ISBN 90-481-9421-0 (hardcover), 90-481-9422-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175.3 .B73

                                              2011. URL https://link.springer.
                                              com/chapter/10.1007/978-90-481-
                                              9422-3_20.

DeSmedt:2012:HAB

[DD12] Johan De Smedt and Helen De Cruz. Human artistic behaviour: Adaptation, byproduct, or cultural group selection? In Philosophy of Behavioral Biology [PR12b], chapter 8, pages 167–187. ISBN 94-007-1950-7, 94-007-1951-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QH331 .P4679 2012; Q175 .B657 v.282.

URL https://link.springer.com/
chapter/10.1007/978-94-007-1951-
4_8.

DeVaucouleurs:1974:CHC

[De 74] G. De Vaucouleurs. The case for a hierarchical cosmology. In Seeger and Cohen [SC74b], pages 227–256. ISBN 90-277-0390-6, 90-277-0376-0 (paperback), 94-010-2126-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 11.

URL https://link.springer.com/
chapter/10.1007/978-94-010-2126-
5_14.

DeSollaPrice:1981:AQT

[De 81] Derek De Solla Price. The analytical (quantitative) theory of science and its implications for the nature of scientific discovery. In Grmek et al. [GCC81b], pages 179–189. ISBN 90-277-1123-2, 94-010-1284-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-010-
1284-3_8.

DeGeorge:1995:EKS

[De 95] Richard T. De George. Ethical knowledge and social facts. In The Concept of Knowledge: The Ankara Seminar [KC95], chapter 10, pages 119–126. ISBN 90-481-4495-7, 94-017-3263-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-017-3263-
5_10.

DeGandt:2001:LIS

[De 01] François De Gandt. The limits of intelligibility: The status of physical science in D’Alembert’s philosophy. In Between Leibniz, Newton, and Kant: Philosophy and Science in the Eighteenth Century [Lef01a], chapter 3, pages 47–60. ISBN 90-481-5774-9, 94-015-9729-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-015-9729-6_3.

DeAngelis:2016:HDW

[De 16] Simone De Angelis. “So how do we know that the Moon is mountainous?” Problems of seeing in Galileo’s reflections on observing the Moon. In Sauer and Scholl [SS16b], chapter 11, pages 223–250. ISBN 3-319-30227-2, 3-319-30229-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q126.8 .P45 2016. URL https:

//link.springer.com/chapter/10.
1007/978-3-319-30229-4_11.

DeBont:2017:ADS

[De 17] Raf De Bont. The adventurer and the documentalist: Science and virtue in interwar nature protection. In van Dongen and Paul [vDP17a], chapter 9, pages 129–147. ISBN 3-319-48892-9 (hardcover), 3-319-48893-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 v. 321; Q175.32.V57 E65 2017. URL https:

//link.springer.com/chapter/10.
1007/978-3-319-48893-6_9.

DEspagnat:1997:ADE

[D’E97] Bernard D’Espagnat. Aiming at describing empirical reality. In Cohen et al. [CHS97b], pages 71–87. ISBN 90-481-4809-X, 94-017-2732-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC173.96-174.52.

URL https://link.springer.com/
chapter/10.1007/978-94-017-2732-
7_6.

DEspagnat:2000:OCR

[D’E00] Bernard D’Espagnat. Observation, contextuality, and realism. In The Reality of the Unobservable: Observability, Unobservability and Their Impact on the Issue of Scientific Realism [AP00b], chapter 19, pages 251–255. ISBN 90-481-5458-8, 94-015-9391-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-015-9391-5_19.

dEspagnat:2008:SCP

[d’E08] Bernard d’Espagnat. Is science cumulative? A physicist viewpoint. In Rethinking Scientific Change and Theory Comparison: Stabilities, Ruptures, Incommensurabilities? [SSHH08], chapter 10, pages 145–151. ISBN 1-4020-6274-5 (hardcover), 1-4020-6279-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175 .R397 2008. URL https://

link.springer.com/chapter/10.
1007/978-1-4020-6279-7_10.  See

commentary [Web08].

Dear:2012:PSH

[Dea12] Peter Dear. Philosophy of science and its historical reconstructions. In Mauskopf and Schmaltz [MS12b], chapter 6, pages 67–82. ISBN 94-007-1744-X, 94-007-1745-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175 .I5134 2011.

URL https://link.springer.com/
chapter/10.1007/978-94-007-1745-
9_6.

Debru:2009:PCB

[Deb09] Claude Debru. Philosophy and contemporary biological research. In Brenner and Gayon [BG09], chapter 10, pages 213–224. ISBN 1-4020-9367-5, 1-4020-9368-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .F746 2009. URL https:

//link.springer.com/chapter/10.
1007/978-1-4020-9368-5_10.

Deichmann:2017:FBK

[Dei17] Ute Deichmann. Francis Bacon, Karl Popper, Michael Polanyi, and modern experimental biology. In Encouraging Openness: Essays for Joseph Agassi on the Occasion of His 90th Birthday [BAG17], chapter 12, pages 129–142. ISBN 3-319-57668-2 (hardcover), 3-319-57669-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN B833.E53 2017. URL https://link.

springer.com/chapter/10.1007/
978-3-319-57669-5_12.

Deloria:1991:PMR

[Del91] Vine Deloria. Perceptions and maturity: Reflections on Feyerabend’s point of view. In Beyond Reason: Essays on the Philosophy of Paul Feyerabend [Mun91a], chapter 18, pages 389–401. ISBN 94-010-5406-1, 94-011-3188-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-011-3188-
9_18.

Deltete:1995:GE

[Del95] Robert J. Deltete. Gibbs and the energeticists. In No Truth Except in the Details: Essays in Honor of Martin J. Klein [KS95], chapter 6, pages 135–169. ISBN 94-010-4097-4, 94-011-0217-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-011-0217-
9_6.

Demopoulos:1989:ALT

[Dem89] William Demopoulos. On applying learnability theory to the rationalism– empiricism controversy. In An Intimate Relation: Studies in the History and Philosophy of Science Presented to Robert E. Butts on his 60th Birthday [BM89], chapter 21, pages 429–440. ISBN 94-010-7546-8, 94-009-2327-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-009-2327-
0_21.

Demirli:2005:CRH

[Dem05] Sun Demirli. Causal relations in Hume. In Turkish Studies in the History and Philosophy of Science [IG05c], chapter 15, pages 217–230. ISBN 1-4020-3332-X (hardcover), 1-4020-3333-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 244. URL https:/

/link.springer.com/chapter/10.
1007/1-4020-3333-8_15.

Denkel:1995:EOC

[Den95] Arda Denkel. Experience, order and cause. In The Concept of Knowledge: The Ankara Seminar [KC95], chapter 4, pages 31–47. ISBN 90-481-4495-7, 94-017-3263-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-017-
3263-5_4.

Denkel:1999:C

[Den99a] Arda Denkel. Causation. In The Natural Background of Meaning [Den99d], chapter 8, pages 194–212. ISBN 90-481-5126-0, 94-015-9084-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-015-9084-6_8.

Denkel:1999:FT

[Den99b] Arda Denkel. The flow of time. In The Natural Background of Meaning [Den99d], chapter 9, pages 213–237. ISBN 90-481-5126-0, 94-015-9084-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-015-9084-
6_9.

Denkel:1999:IRO

[Den99c] Arda Denkel. Introduction: Regularities and the objective background of communication. In The Natural Background of Meaning [Den99d], chapter 1, pages 1–34. ISBN 90-481-5126-0, 94-015-9084-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-015-
9084-6_1.

Denkel:1999:NBM

[Den99d] Arda Denkel. The Natural Background of Meaning, volume 197 of Boston Studies in the Philosophy of Science. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 1999. ISBN 90-481-5126-0, 94-015-9084-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). xi + 252 pp. LCCN ????

URL https://link.springer.com/
book/10.1007/978-94-015-9084-6.

Denkel:1999:NME

[Den99e] Arda Denkel. Natural meaning and evolution. In The Natural Background of Meaning [Den99d], chapter 4, pages 86–116. ISBN 90-481-5126-0, 94-015-9084-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-015-
9084-6_4.

Denkel:1999:NMN

[Den99f] Arda Denkel. Nonnatural meaning naturalized. In The Natural Background of Meaning [Den99d], chapter 6, pages 135–166. ISBN 90-481-5126-0, 94-015-9084-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-015-
9084-6_6.

Denkel:1999:OOC

[Den99g] Arda Denkel. The ontology of objective connections. In The Natural Background of Meaning [Den99d], chapter 7, pages 167–193. ISBN 90-481-5126-0, 94-015-9084-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-015-
9084-6_7.

Denkel:1999:PML

[Den99h] Arda Denkel. The problem of meaning and its Lockean solution. In The Natural Background of Meaning [Den99d], chapter 2, pages 35–61. ISBN 90-481-5126-0, 94-015-9084-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-015-9084-6_2.

Denkel:1999:SRU

[Den99i] Arda Denkel. Some reflections upon the speaker’s meaning. In The Natural Background of Meaning [Den99d], chapter 5, pages 117–134. ISBN 90-481-5126-0, 94-015-9084-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-015-9084-6_5.

Denkel:1999:TL

[Den99j] Arda Denkel. Thinking and language. In The Natural Background of Meaning [Den99d], chapter 3, pages 62–85. ISBN 90-481-5126-0, 94-015-9084-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-015-9084-6_3.

Denkel:2005:CPP

[Den05] Arda Denkel. Causation, parts and properties. In Turkish Studies in the History and Philosophy of Science [IG05c], chapter 14, pages 207–216. ISBN 1-4020-3332-X (hardcover), 1-4020-3333-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 244. URL https:/

/link.springer.com/chapter/10.
1007/1-4020-3333-8_14.

Descartes:1976:VSP

[Des76] R. Descartes. View of space as plenum. In The Concepts of Space and Time: Their Structure and Their Development [Čap76], chapter 14, pages 73–84. ISBN 90-277-0375-2, 94-010-1727-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN BD632.

URL https://link.springer.com/
chapter/10.1007/978-94-010-1727-
5_14.

Descles:2000:ANO

[Des00] Jean-Pierre Desclés. Abduction and non-observability. In The Reality of the Unobservable: Observability, Unobservability and Their Impact on the Issue of Scientific Realism [AP00b], chapter 5, pages 87–112. ISBN 90-481-5458-8, 94-015-9391-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-015-
9391-5_5.

Deser:2003:DCG

[Des03] S. Deser. Dimensionally challenged gravities. In Renn et al. [RDS03], chapter 19, pages 397–401. ISBN 1-4020-1284-5, 1-4020-1285-3 (paperback). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 234. URL https://link.

springer.com/chapter/10.1007/
978-94-010-0111-3_19.

Dettelbach:2007:RRH

[Det07] Michael Dettelbach. Romanticism and resistance: Humboldt and “German” natural philosophy in Napoleonic France. In Hans Christian Ørsted and The Romantic Legacy In Science: Ideas, Disciplines, Practices [BCK07], chapter 13, pages 247–258. ISBN 1-4020-2979-9 (hardcover), 1-4020-2987-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC515.O37 H36 2007. URL https:

//link.springer.com/chapter/10.
1007/978-1-4020-2987-5_13.

Devitt:2001:IPM

[Dev01] Michael Devitt. Incommensurability and the priority of metaphysics. In Incommensurability and Related Matters [HHS01], chapter 5, pages 143–157. ISBN 90-481-5709-9, 94-015-9680-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-015-9680-
0_5.

Devlin:2015:ATK

[Dev15] William J. Devlin. An analysis of truth in Kuhn’s philosophical enterprise. In Kuhn’s Structure of Scientific Revolutions — 50 Years On [DB15b], chapter 11, pages 153–166. ISBN 3-319-13382-9, 3-319-13383-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175.3 .K84 2015.

URL https://link.springer.com/
chapter/10.1007/978-3-319-13383-
6_11.

diFrancia:1981:CPP

[dF81] Giuliano Toraldo di Francia. The concept of progress in physics. In Italian Studies in the Philosophy of Science [Chi81b], chapter 18, pages 323–339. ISBN 90-277-1073-2, 94-009-8937-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-009-8937-
5_18.

DaCosta:1993:TAT

[DF93] Newton C. A. Da Costa and Steven French. Towards an acceptable theory of acceptance: Partial structures inconsistency and correspondence. In Correspondence, Invariance and Heuristics: Essays in Honour of Heinz Post [FK93], chapter 7, pages 137–158. ISBN 90-481-4229-6, 94-017-1185-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-017-1185-
2_7.

deFreitas:2011:DKM

[dF11a] Renan Springer de Freitas. On Darwin, knowledge and mirroring. In Brazilian Studies in Philosophy and History of Science: An account of recent works [KV11], chapter 14, pages 201–209. ISBN 90-481-9421-0 (hardcover), 90-481-9422-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175.3 .B73 2011. URL https:/

/link.springer.com/chapter/10.
1007/978-90-481-9422-3_14.

Dorato:2011:SES

[DF11b] Mauro Dorato and Laura Felline.

Scientific explanation and scientific structuralism. In Bokulich and Bokulich [BB11], chapter 9, pages 161–176. ISBN 90-481-9596-9, 90-481-9597-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175.32.S78 S35 2011. URL https:

//link.springer.com/chapter/10.
1007/978-90-481-9597-8_9.

deFuentes:2014:IPM

[dF14a] Olimpia Arias de Fuentes. An interview with Professor Melquíades de Dios Leyva, December 2008. In The History of Physics in Cuba [BRW14a], chapter 12, pages 285–288. ISBN 94-017-8040-4, 94-017-8041-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC5.53; QC6.9; QC9 C9H57 2014; Q124.6-127.2.

URL https://link.springer.com/
chapter/10.1007/978-94-017-8041-
4_12.

deFuentes:2014:PWC

[dF14b] Olimpia Arias de Fuentes. Physics and women: A challenge being successfully met in Cuba. In The History of Physics in Cuba [BRW14a], chapter 20, pages 339–349. ISBN 94-017-8040-4, 94-017-8041-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC5.53; QC6.9; QC9 C9H57 2014; Q124.6-127.2. URL https://link.

springer.com/chapter/10.1007/
978-94-017-8041-4_20.

Dantes:2011:SBO

[DFL11] Maria Amélia Mascarenhas Dantes, Silvia Figueirôa, and Maria Margaret Lopes. Sciences in Brazil: An overview from 1870–1920. In Brazilian Studies in Philosophy and History of Science: An account of recent works [KV11], chapter 5, pages 95–105. ISBN 90-481-9421-0 (hardcover), 90-481-9422-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175.3 .B73

2011. URL https://link.springer.
com/chapter/10.1007/978-90-481-
9422-3_5.

DeToffoli:2015:IPP

[DG15] Silvia De Toffoli and Valeria Giardino. An inquiry into the practice of proving in low-dimensional topology. In From Logic to Practice: Italian Studies in the Philosophy of Mathematics [LPV15], chapter 15, pages 315–336. ISBN 3-319-10433-0, 3-319-10434-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QA8.4 .F76 2015; Q174 .B67 v.308. URL https://link.

springer.com/chapter/10.1007/
978-3-319-10434-8_15.

Durr:1996:BMF

[DGZ96] Detlef Dürr, Sheldon Goldstein, and Nino Zanghì. Bohmian mechanics as the foundation of quantum mechanics. In Bohmian Mechanics and Quantum Theory: an Appraisal [CFG96], chapter 2, pages 21–44. ISBN 90-481-4698-4, 94-015-8715-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC173.96-174.52. URL https:

//link.springer.com/chapter/10.
1007/978-94-015-8715-0_2.

Dewdney:1996:BBB

[DH96] Chris Dewdney and George Horton. De Broglie, Bohm and the boson. In Bohmian Mechanics and Quantum Theory: an Appraisal [CFG96], chapter 12, pages 169–189. ISBN 90-481-4698-4, 94-015-8715-9 (e-book).

ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC173.96-174.52.

URL https://link.springer.com/
chapter/10.1007/978-94-015-8715-
0_12.

DiGiovanni:1984:MCP

[Di 84] George Di Giovanni. More comments on the place of the organic in Hegel’s philosophy of nature. In Hegel and the Sciences [CW84a], chapter 7, pages 101–107. ISBN 94-009-6235-5, 94-009-6233-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-009-
6233-0_7.

DInverno:2003:D

[D’I03] Ray A. D’Inverno. DSS 2+2. In Renn et al. [RDS03], chapter 14, pages 317–347. ISBN 1-4020-1284-5, 1-4020-1285-3 (paperback). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 234. URL https:/

/link.springer.com/chapter/10.
1007/978-94-010-0111-3_14.

Dibner:1974:LME

[Dib74] Bern Dibner. Leonardo as military engineer. In Seeger and Cohen [SC74b], pages 55–95. ISBN 90-277-0390-6, 90-277-0376-0 (paperback), 94-010-2126-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 11. URL https://link.

springer.com/chapter/10.1007/
978-94-010-2126-5_4.

Dickson:1996:BTL

[Dic96] Michael Dickson. Is the Bohm theory local? In Bohmian Mechanics and Quantum Theory: an Appraisal [CFG96], chapter 22, pages 321–330. ISBN 90-481-4698-4, 94-015-8715-9 (e-book).

ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC173.96-174.52.

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-015-8715-
                                              0_22.

Diemer:1983:CCT

[Die83a] Alwin Diemer. The cultural character of technology. In Philosophy and Technology [DR83], chapter 21, pages 313–318. ISBN 94-009-7126-5, 94-009-7124-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

                                              ???? URL https://link.springer.
                                              com/chapter/10.1007/978-94-009-
                                              7124-0_21.

Diesing:1983:IO

[Die83b] Paul Diesing. Ideology and objectivity. In Epistemology, Methodology, and the Social Sciences [CW83a], chapter 1, pages 1–17. ISBN 90-481-8376-6, 94-017-1458-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

                                              ???? URL https://link.springer.
                                              com/chapter/10.1007/978-94-017-
                                              1458-7_1.

Diederich:1991:OAP

[Die91] Werner Diederich. Obituary on the “anarchist” Paul Feyerabend. In Beyond Reason: Essays on the Philosophy of Paul Feyerabend [Mun91a], chapter 9, pages 213–224. ISBN 94-010-5406-1, 94-011-3188-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

                                              ???? URL https://link.springer.
                                              com/chapter/10.1007/978-94-011-
                                              3188-9_9.

Dienel:2014:CLH

[Die14] Hans-Liudger Dienel. Carl Linde and his relationship with Georges Claude: The cooperation between two independent inventors in cryogenics and its side effects. In History of Artificial Cold, Scientific, Technological and Cultural Issues [Gav14b], chapter 9, pages 171–188. ISBN 94-007-7198-3 (hardcover), 94-007-7199-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC278 H57 2014; TP494 .H577 2014. URL https:/

/link.springer.com/chapter/10.
1007/978-94-007-7199-4_9.

Dieks:2016:PTE

[Die16] Dennis Dieks. Physical time and experienced time. In Cosmological and Psychological Time [DR16], chapter 1, pages 3–20. ISBN 3-319-22589-8 (hardcover), 3-319-22590-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN BD638 C67 2016.

URL https://link.springer.com/
chapter/10.1007/978-3-319-22590-
6_1.

Dilworth:1996:MSA

[Dil96]  Craig Dilworth.  The Metaphysics

of Science: An Account of Modern Science in terms of Principles, Laws and Theories, volume 173 of Boston Studies in the Philosophy of Science. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 1996. ISBN 94-015-8623-3, 94-015-8621-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). x + 246 pp. LCCN ????

URL https://link.springer.com/
book/10.1007/978-94-015-8621-4.

Dilworth:2006:ELS

[Dil06a] Craig Dilworth. Empirical laws: The supervention of experience. In The Metaphysics of Science: An Account of Modern Science in Terms of Principles, Laws and Theories [Dil06f], chapter 4, pages 73–94. ISBN 1-4020-6327-X (paperback), 1-4020-3837-2 (hardcover), 1-4020-3838-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175 .D634195

                                              2006. URL https://link.springer.
                                              com/chapter/10.1007/978-94-015-
                                              8621-4_4.

Dilworth:2006:EVRa

[Dil06b] Craig Dilworth. Empiricism vs. realism — the perennial debate in the philosophy of science. In The Metaphysics of Science: An Account of Modern Science in Terms of Principles, Laws and Theories [Dil06f], chapter 1 (or 2), pages 9–48. ISBN 1-4020-6327-X (paperback), 1-4020-3837-2 (hardcover), 1-4020-3838-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175 .D634195 2006.

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-1-4020-3838-
                                              9_1;   https://link.springer.
                                              com/chapter/10.1007/978-94-015-
                                              8621-4_2.

Dilworth:2006:EVRb

[Dil06c] Craig Dilworth. Empiricism vs. realism revisited. In The Metaphysics of Science: An Account of Modern Science in Terms of Principles, Laws and Theories [Dil06f], chapter 9 (or 10), pages 183–192. ISBN 1-4020-6327-X (paperback), 1-4020-3837-2 (hardcover), 1-4020-3838-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175 .D634195 2006.

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-1-4020-3838-
                                              9_9;   https://link.springer.
                                              com/chapter/10.1007/978-94-015-
                                              8621-4_10.

Dilworth:2006:FRP

[Dil06d] Craig Dilworth. Fundamental and refined principles: The core of modern science. In The Metaphysics of Science:

An Account of Modern Science in Terms of Principles, Laws and Theories [Dil06f], chapter 3, pages 49–72. ISBN 1-4020-6327-X (paperback), 1-4020-3837-2 (hardcover), 1-4020-3838-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175 .D634195 2006. URL https://link.

springer.com/chapter/10.1007/
978-94-015-8621-4_3.

Dilworth:2006:I

[Dil06e] Craig Dilworth. Introduction. In The Metaphysics of Science: An Account of Modern Science in Terms of Principles, Laws and Theories [Dil06f], chapter 1, pages 1–8. ISBN 1-4020-6327-X (paperback), 1-4020-3837-2 (hardcover), 1-4020-3838-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175 .D634195

2006. URL https://link.springer.
com/chapter/10.1007/978-94-015-
8621-4_1.

Dilworth:2006:MSA

[Dil06f] Craig Dilworth. The Metaphysics of Science: An Account of Modern Science in Terms of Principles, Laws and Theories, volume 173 of Boston Studies in the Philosophy of Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., second edition, 2006. ISBN 1-4020-6327-X (paperback), 1-4020-3837-2 (hardcover), 1-4020-3838-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). xiv + 333 pp. LCCN Q175 .D634195 2006.

URL https://link.springer.com/
book/10.1007/978-1-4020-3838-9.

Dilworth:2006:MSF

[Dil06g] Craig Dilworth. Modern science and the future. In The Metaphysics of Science: An Account of Modern Science in Terms of Principles, Laws and Theories [Dil06f], chapter 10 (or 11), pages 193–208. ISBN 1-4020-6327-X (paperback), 1-4020-3837-2 (hardcover), 1-4020-3838-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175 .D634195 2006.

URL https://link.springer.com/
chapter/10.1007/978-1-4020-3838-
9_10;   https://link.springer.
com/chapter/10.1007/978-94-015-
8621-4_11.

Anonymous:2006:NK

[Dil06h] Craig Dilworth. Natural kinds. In The Metaphysics of Science: An Account of Modern Science in Terms of Principles, Laws and Theories [Dil06f], chapter 7 (or 8), pages 148–169. ISBN 1-4020-6327-X (paperback), 1-4020-3837-2 (hardcover), 1-4020-3838-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175 .D634195 2006. URL https://link.

springer.com/chapter/10.1007/
978-1-4020-3838-9_7;   https:/
/link.springer.com/chapter/10.
1007/978-94-015-8621-4_8.

Dilworth:2006:PTL

[Dil06i] Craig Dilworth. The principle–theory– law model of scientific explanation. In The Metaphysics of Science: An Account of Modern Science in Terms of Principles, Laws and Theories [Dil06f], chapter 5 (or 6), pages 109–129. ISBN 1-4020-6327-X (paperback), 1-4020-3837-2 (hardcover), 1-4020-3838-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175 .D634195 2006. URL https://link.

springer.com/chapter/10.1007/
978-1-4020-3838-9_5;   https:/
/link.springer.com/chapter/10.
1007/978-94-015-8621-4_6.

Dilworth:2006:PC

[Dil06j]  Craig Dilworth.  Probability and

confirmation. In The Metaphysics of Science: An Account of Modern Science in Terms of Principles, Laws and Theories [Dil06f], chapter 8 (or 9), pages 170–182. ISBN 1-4020-6327-X (paperback), 1-4020-3837-2 (hardcover), 1-4020-3838-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175 .D634195 2006.

URL https://link.springer.com/
chapter/10.1007/978-1-4020-3838-
9_8;   https://link.springer.
com/chapter/10.1007/978-94-015-
8621-4_9.

Dilworth:2006:STC

[Dil06k] Craig Dilworth. Scientific theories:

Closing the circle. In The Metaphysics of Science: An Account of Modern Science in Terms of Principles, Laws and Theories [Dil06f], chapter 5, pages 95– 108. ISBN 1-4020-6327-X (paperback), 1-4020-3837-2 (hardcover), 1-4020-3838-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175 .D634195 2006. URL https:

//link.springer.com/chapter/10.
1007/978-94-015-8621-4_5.

Anonymous:2006:SSC

[Dil06l] Craig Dilworth. The social sciences:

A consideration of economics. In The Metaphysics of Science: An Account of Modern Science in Terms of Principles, Laws and Theories [Dil06f], chapter 6 (or 7), pages 130–147. ISBN 1-4020-6327-X (paperback), 1-4020-3837-2 (hardcover), 1-4020-3838-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175 .D634195 2006. URL https://link.

springer.com/chapter/10.1007/
978-1-4020-3838-9_6;   https:/

                                              /link.springer.com/chapter/10.
                                              1007/978-94-015-8621-4_7.

Dingguo:1996:NSQ

[Din96] Hong Dingguo. On the neutral status of QM in the dispute of realism vs. anti-realism. In Cohen et al. [CHR96], pages 307–316. ISBN 90-481-4493-0, 94-015-8638-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

                                              ???? URL https://link.springer.
                                              com/chapter/10.1007/978-94-015-
                                              8638-2_22.

Disalle:2006:MSW

[Dis06] Robert Disalle. Mathematical structure, “world structure,” and the philosophical turning-point in modern physics. In Hendricks et al. [HJLP06], pages 207–230. ISBN 1-4020-5194-8 (hardcover), 1-4020-5195-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175 .B69 v.251 2006; Q175 .I51377 2006. URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-1-4020-5195-1_7.

DAgostino:1989:FRE

[DJ89] Fred D’Agostino and I. C. Jarvie. Freedom and Rationality: Essays in Honor of John Watkins From his Colleagues and Friends, volume 117 of Boston Studies in the Philosophy of Science. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 1989. ISBN 94-010-7571-9, 94-009-2380-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). xii + 388 pp. LCCN ????

                                              URL https://link.springer.com/
                                              book/10.1007/978-94-009-2380-5.

Ortega:2014:TPT

[dJAO14] Diego de Jesús Alamino Ortega. The training of physics teachers in Cuba: A historical approach. In The History of Physics in Cuba [BRW14a], chapter 9, pages 261–268. ISBN 94-017-8040-4, 94-017-8041-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC5.53; QC6.9; QC9 C9H57 2014; Q124.6-127.2. URL https:/

/link.springer.com/chapter/10.
1007/978-94-017-8041-4_9.

Djuric:1979:HP

[Dju79] Mihailo Djurić. Homo politicus. In Praxis: Yugoslav Essays in the Philosophy and Methodology of the Social Sciences [MP79], chapter 6, pages 101–119. ISBN 90-277-0968-8, 94-009-9355-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-009-
9355-6_6.

Dorofeeva-Lichtmann:2005:SOA

[DL05] Vera Dorofeeva-Lichtmann. Spatial organization of Ancient Chinese texts (preliminary remarks). In History of Science, History of Text [Che05a], chapter 1, pages 3–47. ISBN 1-281-33589-4, 1-4020-2320-0 (hardcover), 1-4020-2321-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 238 Q125.

URL https://link.springer.com/
chapter/10.1007/1-4020-2321-9_1.

Pereira:2014:PSU

[dMPC14] Osvaldo de Melo Pereira and María Sánchez Colina. Physics studies at the University of Havana. In The History of Physics in Cuba [BRW14a], chapter 19, pages 329–338. ISBN 94-017-8040-4, 94-017-8041-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC5.53; QC6.9; QC9 C9H57 2014; Q124.6-127.2.

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-017-8041-
                                              4_19.

Dialetis:1994:IHP

[DN94] Dimitrios Dialetis and Efthymios Nicolaidis. Issues in the historiography of post-Byzantine science. In Trends in the Historiography of Science [GCN94], chapter 8, pages 121–127. ISBN 90-481-4264-4, 94-017-3596-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-017-3596-4_8.

Dana:1991:FEL

[DO91] L. I. Dana and H. Kamerlingh Onnes. Further experiments with liquid helium. BB. Preliminary determinations of the specific heat of liquid helium. In Gavroglu and Goudaroulis [GG91], chapter 12, pages 227–235. ISBN 94-010-7433-X (hardcover), 94-009-2079-2. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 124.

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-009-2079-
                                              8_12.

Dobbs:1992:GA

[Dob92] B. J. T. Dobbs. Gravity and alchemy. In The Scientific Enterprise: The Bar-Hillel Colloquium: Studies in History, Philosophy, and Sociology of Science, Volume 4 [UM92], chapter 13, pages 205–222. ISBN 94-010-5190-9, 94-011-2688-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .S354 1992. URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-011-2688-5_13.

Dolev:1986:MBW

[Dol86] Eran Dolev. Medicine and the Boer War: a comment. In Ullmann-Margalit [UM86c], chapter 10, pages 141–145. ISBN 90-277-2160-2, 90-277-2161-0 (paperback), 94-009-4566-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 95

Q175. URL https://link.springer.
com/chapter/10.1007/978-94-009-

4566-1_10. See [Lee86].

Dolan:2003:EST

[Dol03] Brian Dolan. Embodied skills and travelling savants. In Simões et al. [SCD03b], chapter 6, pages 115–141. ISBN 1-4020-1259-4, 94-017-3584-0. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 v. 233.

URL https://link.springer.com/
chapter/10.1007/978-94-017-3584-
1_6.

Dolev:2016:RGT

[Dol16] Yuval Dolev. Relativity, global tense and phenomenology. In Cosmological and Psychological Time [DR16], chapter 2, pages 21–40. ISBN 3-319-22589-8 (hardcover), 3-319-22590-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN BD638 C67 2016.

URL https://link.springer.com/
chapter/10.1007/978-3-319-22590-
6_2.

Donagan:1986:WEC

[Don86] Alan Donagan. The worst excess of Cartesian dualism. In Cohen and Wartofsky [CW85], chapter 16, pages 313–325. ISBN 90-277-1971-3, 94-009-5345-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-009-
5349-9_16.

Doncel:1998:HCC

[Don98] Manuel G. Doncel. On Hertz’s conceptual conversion: From wire waves to air waves. In Heinrich Hertz: Classical Physicist, Modern Philosopher [BHN98], chapter 6, pages 73–87. ISBN 90-481-4881-2, 94-015-8855-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-015-8855-3_6.

Donohue:2017:JAM

[Don17]  Christopher R. Donohue.  Joseph

Agassi from metaphysics to politics. In Encouraging Openness: Essays for Joseph Agassi on the Occasion of His 90th Birthday [BAG17], chapter 3, pages 19–26. ISBN 3-319-57668-2 (hardcover), 3-319-57669-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN B833.E53 2017.

URL https://link.springer.com/
chapter/10.1007/978-3-319-57669-
5_3.

Doppelt:2001:INF

[Dop01] Gerald Doppelt. Incommensurability and the normative foundations of scientific knowledge. In Incommensurability and Related Matters [HHS01], chapter 6, pages 159–179. ISBN 90-481-5709-9, 94-015-9680-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-015-9680-0_6.

Dory:1992:GCF

[Dor92] Daniel Dory. Géographie et colonisation en France durant la Troisième République (1870–1940). (French) [Geography and colonization in France during the Third Republic (1870– 1940)]. In Science and Empires: Historical Studies about Scientific Development and European Expansion [PJM92], chapter 32, pages 323–329.

ISBN 0-7923-1518-9 (hardcover), 94-010-5145-3, 94-011-2594-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 136.

URL https://link.springer.com/
chapter/10.1007/978-94-011-2594-
9_32.

Dorier:1996:BDG

[Dor96] Jean-Luc Dorier. Basis and dimension — from Grassmann to van der Waerden. In Schubring [Sch96f], pages 175–196. ISBN 90-481-4758-1, 94-015-8753-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QA8.9-10.3. URL https://link.

springer.com/chapter/10.1007/
978-94-015-8753-2_15.

Dorato:1997:TVR

[Dor97] Mauro Dorato. Three views on the relationship between time and reality. In Perspectives on Time [FSU97b], chapter 2, pages 61–92. ISBN 90-481-4774-3, 94-015-8875-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol.189.

URL https://link.springer.com/
chapter/10.1007/978-94-015-8875-
1_2.

Dorato:2000:MCE

[Dor00] Mauro Dorato. Measurability, computability and the existence of theoretical entities. In The Reality of the Unobservable: Observability, Unobservability and Their Impact on the Issue of Scientific Realism [AP00b], chapter 15, pages 207–217. ISBN 90-481-5458-8, 94-015-9391-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-015-9391-5_15.

Douglass:1987:SLS

[Dou87] Paul Douglass. “Such as the life is, such is the form”: Organicism among the moderns. In Burwick [Bur87b], chapter 9, pages 253–273. ISBN 90-277-2541-1, 94-009-3917-5. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 105. US$64.00(U.S.).

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-009-3917-
                                              2_9.

Douglas:1999:FDN

[Dou99] Peter Douglas. The fractal dynamics of a Nietzschean world. In Nietzsche, Epistemology, and Philosophy of Science: Nietzsche and the Sciences II [Bab99a], chapter 25, pages 323–330. ISBN 90-481-5234-8, 94-017-2428-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-017-2428-
                                              9_25.

Douven:2003:BLC

[Dou03] Igor Douven. On Bayesian logic: Comments on Colin Howson. In Descartes’s Mathematical Thought [Sas03d], chapter 22, pages 321–326. ISBN 90-481-6487-7, 94-017-1225-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

                                              URL https://link.springer.com/
                                              chapter/10.1007/0-306-48123-5_22.

Dove:1984:HDC

[Dov84] Kenley Royce Dove. Hegel’s ‘deduction of the concept of science’. In Hegel and the Sciences [CW84a], chapter 15, pages 271–281. ISBN 94-009-6235-5, 94-009-6233-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

                                              ???? URL https://link.springer.

com/chapter/10.1007/978-94-009-
6233-0_15.

DeZoysa:1992:MES

[DP92] Arjuna De Zoysa and C. D.

Palitharatna. Models of European scientific expansion: A comparative description of “classical” medical science at the time of introduction of European medical science to Sri Lanka, and subsequent development to present. In Science and Empires: Historical Studies about Scientific Development and European Expansion [PJM92], chapter 13, pages 111–120. ISBN 0-7923-1518-9 (hardcover), 94-010-5145-3, 94-011-2594-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 136.

URL https://link.springer.com/
chapter/10.1007/978-94-011-2594-
9_13.

Donagan:1986:HNN

[DPW86a] Alan Donagan, Anthony N. Perovich, Jr., and Michael V. Wedin. Human Nature and Natural Knowledge: Essays Presented to Marjorie Grene on the Occasion of Her Seventy-Fifth Birthday, volume 89 of Boston Studies in the Philosophy of Science. D. Reidel, Dordrecht, The Netherlands; Boston, MA, USA; Lancaster, UK; Tokyo, Japan, 1986. ISBN 94-010-8859-4, 94-009-5349-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). 404 (est.) pp. LCCN ???? URL

https://link.springer.com/book/
10.1007/978-94-009-5349-9.

Donagan:1986:PAM

[DPW86b] Alan Donagan, Anthony N. Perovich, Jr., and Michael V. Wedin. The professional activities of Marjorie Grene. In Cohen and Wartofsky [CW85], chapter 20, pages 369–370. ISBN 90-277-1971-3, 94-009-5345-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-009-5349-9_20.

Donagan:1986:PMG

[DPW86c] Alan Donagan, Anthony N. Perovich, Jr., and Michael V. Wedin. The publications of Marjorie Grene. In Cohen and Wartofsky [CW85], chapter 21, pages 371–374. ISBN 90-277-1971-3, 94-009-5345-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

                                              ???? URL https://link.springer.
                                              com/chapter/10.1007/978-94-009-
                                              5349-9_21.

Durbin:1983:PT

[DR83] Paul T. Durbin and Friedrich Rapp. Philosophy and Technology, volume 80 of Boston Studies in the Philosophy of Science. D. Reidel, Dordrecht, The Netherlands; Boston, MA, USA; Lancaster, UK; Tokyo, Japan, 1983. ISBN 94-009-7126-5, 94-009-7124-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). xiv + 344 pp. LCCN ????

                                              URL https://link.springer.com/
                                              book/10.1007/978-94-009-7124-0.

deRegt:2001:ES

[dR01] Henk W. de Regt. Erwin Schrödinger. In Blackmore et al. [BIT01c], pages 85–104. ISBN 90-481-5752-8, 94-015-9690-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

                                              ???? URL https://link.springer.
                                              com/chapter/10.1007/978-94-015-
                                              9690-9_4.

DeLanghe:2015:TCT

[DR15a] Rogier De Langhe and Peter Rubbens. From theory choice to theory search: The essential tension between exploration and exploitation in science. In Kuhn’s Structure of Scientific Revolutions — 50 Years On [DB15b], chapter 8, pages 105–114. ISBN 3-319-13382-9, 3-319-13383-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175.3 .K84 2015.

URL https://link.springer.com/
chapter/10.1007/978-3-319-13383-
6_8.

deRicqles:2015:DBB

[dR15b] Armand J. de Ricqlès. Dahlem 1981:

Before and beyond. In Conceptual Change in Biology: Scientific and Philosophical Perspectives on Evolution and Development [Lov15b], chapter 12, pages 259–269. ISBN 94-017-9411-1, 94-017-9412-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QH331 C66 2015; QH491; QH492 .C66

2014. URL https://link.springer.
com/chapter/10.1007/978-94-017-
9412-1_12.

Dolev:2016:CPT

[DR16] Yuval Dolev and Michael Roubach.

Cosmological and Psychological Time, volume 285 of Boston Studies in the Philosophy and History of Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2016. ISBN 3-319-22589-8 (hardcover), 3-319-22590-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). xiv + 218 pp. LCCN BD638 C67 2016. URL https://link.

springer.com/book/10.1007/978-3-
319-22590-6.

deRicqles:2017:BTD

[dR17] Armand J. de Ricqlès. The biologist’s time and deep time: Essay on the psychology of the paleobiologist. In Time of Nature and the Nature of Time: Philosophical Perspectives of Time in Natural Sciences [BH17b], chapter 11, pages 237–251. ISBN 3-319-53723-7 (hardcover), 3-319-53725-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC173.59.S65 T56

                                              2017. URL https://link.springer.
                                              com/chapter/10.1007/978-3-319-
                                              53725-2_11.

Dray:1988:GVJ

[Dra88] William H. Dray. Generalization, valuejudgment and causal explanation in history. In Philosophy, History and Social Action: Essays in Honor of Lewis Feuer with an autobiographic essay by Lewis Feuer [HOO88c], chapter 5, pages 137–155. ISBN 94-010-7793-2, 94-009-2873-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

                                              ???? URL https://link.springer.
                                              com/chapter/10.1007/978-94-009-
                                              2873-2_5.

Drake:1994:TPE

[Dra94] Stillman Drake. Theory and practice in early modern physics. In Trends in the Historiography of Science [GCN94], chapter 2, pages 15–30. ISBN 90-481-4264-4, 94-017-3596-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-017-3596-4_2.

Drago:1999:IBH

[Dra99] Antonino Drago. Incommensurability as a bound of hermeneutics in science. In Fehér et al. [FKR99], pages 135–155. ISBN 90-481-5257-7, 94-015-9293-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-015-9293-
                                              2_12.

Dretske:1986:MMM

[Dre86] Fred Dretske. Minds, machines and meaning. In Philosophy and Technology II: Information Technology and Computers in Theory and Practice [MH86], chapter 8, pages 97–109. ISBN 94-010-8510-2, 94-009-4512-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-009-4512-8_8.

Drennan:1992:WAC

[Dre92] Robert D. Drennan. What is the archaeology of chiefdoms about? In Metaarchaeology: Reflections by Archaeologists and Philosophers [Emb92b], chapter 2, pages 53–74. ISBN 94-010-4806-1, 94-011-1826-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-011-1826-
2_2.

Drouin:2013:SCT

[Dro13] Jean-Marc Drouin. Some conceptions of time in ecology. In Bergandi [Ber13c], chapter 4, pages 49–62. ISBN 94-007-5066-8, 94-007-5067-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN GE42 .S77 2013; Q175 .B73 v. 296. URL https:/

/link.springer.com/chapter/10.
1007/978-94-007-5067-8_4.

Domingues:2001:IDB

[DS01] Heloisa Maria Bertol Domingues and Magali Romero Sá. The introduction of Darwinism in Brazil. In The Reception of Darwinism in the Iberian World: Spain, Spanish America and Brazil [GPSR01], chapter 4, pages 65–81. ISBN 94-010-3885-6, 94-010-0602-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-010-0602-
6_4.

daSilva:2011:NMK

[dS11a] Jairo José da Silva. On the nature of mathematical knowledge. In Brazilian Studies in Philosophy and History of Science: An account of recent works [KV11], chapter 10, pages 151–160. ISBN 90-481-9421-0 (hardcover), 90-481-9422-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175.3 .B73 2011. URL https:/

/link.springer.com/chapter/10.
1007/978-90-481-9422-3_10.

Dias:2011:INR

[DS11b] Penha Maria Cardoso Dias and Teresinha J. Stuchi. Isaac Newton, Robert Hooke and the mystery of the orbit. In Brazilian Studies in Philosophy and History of Science: An account of recent works [KV11], chapter 4, pages 77–94. ISBN 90-481-9421-0 (hardcover), 90-481-9422-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175.3 .B73 2011.

URL https://link.springer.com/
chapter/10.1007/978-90-481-9422-
3_4.

Dufour:2012:RER

[DS12] Catherine Dufour and Léna Soler. Robustness of an experimental result: The example of the tests of Bell’s Inequalities. In Characterizing the Robustness of Science: After the Practice Turn in Philosophy of Science [STNW12], chapter 6, pages 147–167. ISBN 94-007-2758-5, 94-007-2759-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175 .C4595 2012.

URL https://link.springer.com/

chapter/10.1007/978-94-007-2759-
5_6.

deSyon:2014:MAC

[dS14] Guillaume de Syon. Meeting artificial cold: Expositions and refrigeration, 1896–1937. In History of Artificial Cold, Scientific, Technological and Cultural Issues [Gav14b], chapter 10, pages 189–198. ISBN 94-007-7198-3 (hardcover), 94-007-7199-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC278 H57 2014; TP494 .H577 2014. URL https:/

/link.springer.com/chapter/10.
1007/978-94-007-7199-4_10.

Didier:2018:BEA

[DS18] Christelle Didier and Patrick Simonnin.

“I became an engineer by accident!”: Engineering, vocation, and professional values. In Mitcham et al. [MLNZ18], chapter 19, pages 229–236. ISBN 3-319-62448-2, 3-319-62450-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN TA15 .P45 2018.

URL https://link.springer.com/
chapter/10.1007/978-3-319-62450-
1_19.

Duchesneau:1989:LHP

[Duc89] François Duchesneau. Leibniz’s ‘hypothesis physica nova’: A conjunction of models for explaining phenomena. In An Intimate Relation: Studies in the History and Philosophy of Science Presented to Robert E. Butts on his 60th Birthday [BM89], chapter 8, pages 153–170. ISBN 94-010-7546-8, 94-009-2327-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-009-
2327-0_8.

Duchesne:1992:FES

[Duc92] Raymond Duchesne. La France et l’Émergence des sciences modernes au Canada français (1900–1940). (French) [France and the emergence of modern sciences in French Canada (1900–1940)]. In Science and Empires: Historical Studies about Scientific Development and European Expansion [PJM92], chapter 33, pages 331–338. ISBN 0-7923-1518-9 (hardcover), 94-010-5145-3, 94-011-2594-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 136.

URL https://link.springer.com/
chapter/10.1007/978-94-011-2594-
9_33.

Duchesneau:1996:TAM

[Duc96]  François Duchesneau.  Teleological

arguments from a methodological viewpoint. In Québec Studies in the Philosophy of Science: Part II: Biology, Psychology, Cognitive Science and Economics Essays in Honor of Hugues Leblanc [MC96], chapter 1, pages 1–12. ISBN 94-010-6537-3, 94-009-0113-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-009-0113-
1_1.

Dudai:1986:EAE

[Dud86] Yadin Dudai. On experimental approaches and evolution: A comment. In Ullmann-Margalit [UM86a], chapter 10, pages 111–115. ISBN 90-277-2159-9 (paperback), 90-277-2158-0 (hardcover), 94-009-5496-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 94 Q175.

URL https://link.springer.com/
chapter/10.1007/978-94-009-5496-
0_10.

Duhem:1976:AFR

[Duh76a] P. Duhem. Absolute frame of reference according to St. Thomas. In The Concepts of Space and Time: Their Structure and Their Development [Čap76], chapter 8, pages 41–42. ISBN 90-277-0375-2, 94-010-1727-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN BD632.

URL https://link.springer.com/
chapter/10.1007/978-94-010-1727-
5_8.

Duhem:1976:EPU

[Duh76b] P. Duhem. The Empyrean as the place of the Universe. In The Concepts of Space and Time: Their Structure and Their Development [Čap76], chapter 9, pages 43–45. ISBN 90-277-0375-2, 94-010-1727-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic).

LCCN BD632. URL https://link.
springer.com/chapter/10.1007/
978-94-010-1727-5_9.

Duhem:1976:PVA

[Duh76c] P. Duhem. Place and the void according to John Philopon. In The Concepts of Space and Time: Their Structure and Their Development [Čap76], chapter 7, pages 39–40. ISBN 90-277-0375-2, 94-010-1727-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic).

LCCN BD632. URL https://link.
springer.com/chapter/10.1007/
978-94-010-1727-5_7.

Duhem:1976:PTS

[Duh76d] P. Duhem. Plato’s theory of space and the geometrical composition of the elements. In The Concepts of Space and Time: Their Structure and Their Development [Čap76], chapter 3, pages 21–25. ISBN 90-277-0375-2, 94-010-1727-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic).

                                              LCCN BD632. URL https://link.
                                              springer.com/chapter/10.1007/
                                              978-94-010-1727-5_3.

Duhem:1976:PAC

[Duh76e] P. Duhem. The problem of the absolute clock. In The Concepts of Space and Time: Their Structure and Their Development [Čap76], chapter 31, pages 185–186. ISBN 90-277-0375-2, 94-010-1727-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic).

                                              LCCN BD632. URL https://link.
                                              springer.com/chapter/10.1007/
                                              978-94-010-1727-5_31.

Duhem:1976:SVA

[Duh76f] P. Duhem. Space and the void according to Aristotle. In The Concepts of Space and Time: Their Structure and Their Development [Čap76], chapter 4, pages 27–29. ISBN 90-277-0375-2, 94-010-1727-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic).

                                              LCCN BD632. URL https://link.
                                              springer.com/chapter/10.1007/
                                              978-94-010-1727-5_4.

Duhem:1991:ASM

[Duh91a] Pierre Duhem. The Alexandrian sources of medieval statics. In The Origins of Statics: The Sources of Physical Theory [Duh91r], chapter 5, pages 47–74. ISBN 94-010-5658-7, 94-011-3730-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 123.

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-011-3730-
                                              0_5.

Duhem:1991:AAB

[Duh91b] Pierre Duhem. Aristotle and Archimedes (384—322 and 287—212 B.C.). In The Origins of Statics: The Sources of Physical Theory [Duh91r], chapter 1, pages 11–15. ISBN 94-010-5658-7, 94-011-3730-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 123.

URL https://link.springer.com/
chapter/10.1007/978-94-011-3730-
0_1.

Duhem:1991:DAS

[Duh91c] Pierre Duhem. The doctrine of Albert of Saxony and the geostaticians. In The Origins of Statics: The Sources of Physical Theory [Duh91r], chapter 16, pages 357–379. ISBN 94-010-5658-7, 94-011-3730-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 123. URL https:

//link.springer.com/chapter/10.
1007/978-94-011-3730-0_16.

Duhem:1991:FCSa

[Duh91d] Pierre Duhem. The French contribution to statics — Roberval. In The Origins of Statics: The Sources of Physical Theory [Duh91r], chapter 13, pages 201–225. ISBN 94-010-5658-7, 94-011-3730-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 123. URL https:/

/link.springer.com/chapter/10.
1007/978-94-011-3730-0_13.

Duhem:1991:FCSb

[Duh91e] Pierre Duhem. The French contribution to statics (continued) René Descartes (1596–1650). In The Origins of Statics: The Sources of Physical Theory [Duh91r], chapter 14, pages 226–243. ISBN 94-010-5658-7, 94-011-3730-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 123. URL https://link.

springer.com/chapter/10.1007/
978-94-011-3730-0_14.

Duhem:1991:GG

[Duh91f] Pierre Duhem. Galileo Galilei (1564– 1642). In The Origins of Statics: The Sources of Physical Theory [Duh91r], chapter 11, pages 166–183. ISBN 94-010-5658-7, 94-011-3730-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 123.

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-011-3730-
                                              0_11.

Duhem:1991:IPM

[Duh91g] Pierre Duhem. The impossibility of perpetual motion. In The Origins of Statics: The Sources of Physical Theory [Duh91r], chapter 4, pages 41–46. ISBN 94-010-5658-7, 94-011-3730-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 123.

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-011-3730-
                                              0_4.

Duhem:1991:JC

[Duh91h] Pierre Duhem. Jerome Cardan (1501– 1576). In The Origins of Statics: The Sources of Physical Theory [Duh91r], chapter 3, pages 30–40. ISBN 94-010-5658-7, 94-011-3730-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 123.

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-011-3730-
                                              0_3.

Duhem:1991:LV

[Duh91i] Pierre Duhem. Leonardo da Vinci (1452–1519). In The Origins of Statics: The Sources of Physical Theory [Duh91r], chapter 2, pages 16–29. ISBN 94-010-5658-7, 94-011-3730-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 123.

URL https://link.springer.com/
chapter/10.1007/978-94-011-3730-
0_2.

Duhem:1991:MPC

[Duh91j] Pierre Duhem. The mechanical properties of the center of gravity from Albert of Saxony to Evangelista Torricelli. In The Origins of Statics: The Sources of Physical Theory [Duh91r], chapter 15, pages 261–356. ISBN 94-010-5658-7, 94-011-3730-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 123. URL https:

//link.springer.com/chapter/10.
1007/978-94-011-3730-0_15.

Duhem:1991:RAJ

[Duh91k] Pierre Duhem. The reaction against Jordanus Guido Ubaldo — G. B. Benedetti. In The Origins of Statics: The Sources of Physical Theory [Duh91r], chapter 10, pages 148–165. ISBN 94-010-5658-7, 94-011-3730-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 123. URL https://link.

springer.com/chapter/10.1007/
978-94-011-3730-0_10.

Duhem:1991:SJC

[Duh91l] Pierre Duhem. The School of Jordanus in the 16th Century Nicolo Tartaglia. In The Origins of Statics: The Sources of Physical Theory [Duh91r], chapter 9, pages 138–147. ISBN 94-010-5658-7, 94-011-3730-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 123. URL https:

//link.springer.com/chapter/10.
1007/978-94-011-3730-0_9.

Duhem:1991:SS

[Duh91m] Pierre Duhem. Simon Stevin (1548– 1620). In The Origins of Statics: The Sources of Physical Theory [Duh91r], chapter 12, pages 184–200. ISBN 94-010-5658-7, 94-011-3730-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 123.

URL https://link.springer.com/
chapter/10.1007/978-94-011-3730-
0_12.

Duhem:1991:SDM

[Duh91n] Pierre Duhem. Statics during the Middle Ages Jordanus de Nemore. In The Origins of Statics: The Sources of Physical Theory [Duh91r], chapter 6, pages 75–91. ISBN 94-010-5658-7, 94-011-3730-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 123. URL https:/

/link.springer.com/chapter/10.
1007/978-94-011-3730-0_6.

Duhem:1991:SMAb

[Duh91o] Pierre Duhem. The statics of the Middle Ages and Leonardo da Vinci. In The Origins of Statics: The Sources of Physical Theory [Duh91r], chapter 8, pages 114–137. ISBN 94-010-5658-7, 94-011-3730-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 123. URL https:

//link.springer.com/chapter/10.
1007/978-94-011-3730-0_8.

Duhem:1991:SMAa

[Duh91p] Pierre Duhem. The statics of the Middle Ages (continued) the School of Jordanus. In The Origins of Statics: The Sources of Physical Theory [Duh91r], chapter 7, pages 92–113. ISBN 94-010-5658-7, 94-011-3730-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 123. URL https://link.

springer.com/chapter/10.1007/
978-94-011-3730-0_7.

Duhem:1991:SLS

[Duh91q] Pierre Duhem. The systematization of the laws of statics. In The Origins of Statics: The Sources of Physical Theory [Duh91r], chapter 17, pages 380–448. ISBN 94-010-5658-7, 94-011-3730-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 123. URL https://link.

springer.com/chapter/10.1007/
978-94-011-3730-0_17.

Duhem:1991:OSS

[Duh91r] Pierre Maurice Marie Duhem. The Origins of Statics: The Sources of Physical Theory, volume 123 of Boston Studies in the Philosophy of Science. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 1991. ISBN 94-010-5658-7, 94-011-3730-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). xxxv + 593 pp. LCCN Q174 .B67 vol. 123. URL

http://d-nb.info/101431545X/34;
 https://link.springer.com/book/
10.1007/978-94-011-3730-0.

Duhem:2002:ATC

[Duh02a] Pierre Duhem. The atomic theory:

Critique of this theory. In Mixture and Chemical Combination: and Related Essays [Duh02m], chapter 12, pages 85–94. ISBN 90-481-5924-5, 94-017-2292-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-017-
2292-6_12.

Duhem:2002:CMB

[Duh02b] Pierre Duhem. Chemical mechanics based on thermodynamics. In Mixture and Chemical Combination: and Related Essays [Duh02m], chapter 14, pages 107–113. ISBN 90-481-5924-5, 94-017-2292-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

                                              ???? URL https://link.springer.
                                              com/chapter/10.1007/978-94-017-
                                              2292-6_14.

Duhem:2002:CMF

[Duh02c] Pierre Duhem. Chemical mechanics: First attempts. In Mixture and Chemical Combination: and Related Essays [Duh02m], chapter 13, pages 95–106. ISBN 90-481-5924-5, 94-017-2292-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-017-2292-
                                              6_13.

Duhem:2002:CS

[Duh02d] Pierre Duhem. Chemical substitution. In Mixture and Chemical Combination: and Related Essays [Duh02m], chapter 8, pages 57–61. ISBN 90-481-5924-5, 94-017-2292-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-017-2292-6_8.

Duhem:2002:CT

[Duh02e] Pierre Duhem. Chemical types. In Mixture and Chemical Combination: and Related Essays [Duh02m], chapter 9, pages 62–67. ISBN 90-481-5924-5, 94-017-2292-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

                                              ???? URL https://link.springer.
                                              com/chapter/10.1007/978-94-017-
                                              2292-6_9.

Duhem:2002:C

[Duh02f] Pierre Duhem. Conclusion. In Mixture and Chemical Combination: and Related Essays [Duh02m], chapter 15, pages 114–118. ISBN 90-481-5924-5, 94-017-2292-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-017-
2292-6_15.

Duhem:2002:CTV

[Duh02g] Pierre Duhem. Condensed types, valency and developed formulas. In Mixture and Chemical Combination: and Related Essays [Duh02m], chapter 10, pages 68–77. ISBN 90-481-5924-5, 94-017-2292-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-017-2292-6_10.

Duhem:2002:CCF

[Duh02h] Pierre Duhem. Crude chemical formulas and equivalent masses. In Mixture and Chemical Combination: and Related Essays [Duh02m], chapter 7, pages 43– 56. ISBN 90-481-5924-5, 94-017-2292-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-017-2292-
6_7.

Duhem:2002:EPT

[Duh02i]  Pierre Duhem.  The evolution of

physical theories from the Seventeenth Century to our day. In Mixture and Chemical Combination: and Related Essays [Duh02m], chapter 17, pages 193–213. ISBN 90-481-5924-5, 94-017-2292-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-017-
2292-6_17.

Duhem:2002:IS

[Duh02j] Pierre Duhem. Isomers and stereochemistry. In Mixture and Chemical Combination: and Related Essays [Duh02m], chapter 11, pages 78–84. ISBN 90-481-5924-5, 94-017-2292-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-017-2292-
                                              6_11.

Duhem:2002:LDP

[Duh02k] Pierre Duhem. The law of definite proportions. In Mixture and Chemical Combination: and Related Essays [Duh02m], chapter 6, pages 34–42. ISBN 90-481-5924-5, 94-017-2292-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-017-2292-
                                              6_6.

Duhem:2002:MAA

[Duh02l] Pierre Duhem. Mixts according to the atomists and according to the peripatetics. In Mixture and Chemical Combination: and Related Essays [Duh02m], chapter 1, pages 5–8. ISBN 90-481-5924-5, 94-017-2292-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-017-2292-6_1.

Duhem:2002:MCC

[Duh02m] Pierre Duhem. Mixture and Chemical Combination: and Related Essays, volume 223 of Boston Studies in the Philosophy of Science. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 2002. ISBN 90-481-5924-5, 94-017-2292-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). xxx + 322 pp. LCCN ???? URL

https://link.springer.com/book/
10.1007/978-94-017-2292-6.

Duhem:2002:NSP

[Duh02n] Pierre Duhem. A new science: Physical chemistry. In Mixture and Chemical Combination: and Related Essays [Duh02m], chapter 20, pages 253–276. ISBN 90-481-5924-5, 94-017-2292-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-017-2292-
6_20.

Duhem:2002:NMEb

[Duh02o] Pierre Duhem. The notion of a mixt in the Eighteenth Century up to the chemical revolution: The empiricist school. In Mixture and Chemical Combination: and Related Essays [Duh02m], chapter 4, pages 22–28. ISBN 90-481-5924-5, 94-017-2292-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-017-2292-
6_4.

Duhem:2002:NMEa

[Duh02p] Pierre Duhem. The notion of a mixt in the Eighteenth Century up to the chemical revolution: The Newtonian School. In Mixture and Chemical Combination: and Related Essays [Duh02m], chapter 3, pages 17–21. ISBN 90-481-5924-5, 94-017-2292-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-017-2292-
6_3.

Duhem:2002:NMS

[Duh02q] Pierre Duhem. The notion of a mixt in the Seventeenth Century. In Mixture and Chemical Combination: and Related Essays [Duh02m], chapter 2, pages 9–16. ISBN 90-481-5924-5, 94-017-2292-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

                                              ???? URL https://link.springer.
                                              com/chapter/10.1007/978-94-017-
                                              2292-6_2.

Duhem:2002:BMC

[Duh02r] Pierre Duhem. On bodies and mixtures or combinations. In Mixture and Chemical Combination: and Related Essays [Duh02m], chapter 23, pages 311–313. ISBN 90-481-5924-5, 94-017-2292-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-017-2292-
                                              6_23.

Duhem:2002:SRE

[Duh02s] Pierre Duhem. On some recent extensions of statics and dynamics. In Mixture and Chemical Combination: and Related Essays [Duh02m], chapter 22, pages 291–309. ISBN 90-481-5924-5, 94-017-2292-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-017-2292-6_22.

Duhem:2002:PLC

[Duh02t] Pierre Duhem. The phase law, in connection with a recent book of Wilder D. Bancroft. In Mixture and Chemical Combination: and Related Essays [Duh02m], chapter 19, pages 235–251. ISBN 90-481-5924-5, 94-017-2292-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-017-2292-
6_19.

Duhem:2002:SS

[Duh02u] Pierre Duhem. Simple substances.

In Mixture and Chemical Combination: and Related Essays [Duh02m], chapter 5, pages 31–33. ISBN 90-481-5924-5, 94-017-2292-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-017-2292-6_5.

Duhem:2002:TH

[Duh02v] Pierre Duhem. Theories of heat.

In Mixture and Chemical Combination: and Related Essays [Duh02m], chapter 16, pages 121–191. ISBN 90-481-5924-5, 94-017-2292-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-017-2292-6_16.

Duhem:2002:TCR

[Duh02w] Pierre Duhem. Thermochemistry:

In connection with a recent book of Marcelin Berthelot. In Mixture and Chemical Combination: and Related Essays [Duh02m], chapter 18, pages 215–234. ISBN 90-481-5924-5, 94-017-2292-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-017-
2292-6_18.

Duhem:2002:WJH

[Duh02x] Pierre Duhem. The work of J. H. Van ’t Hoff, in connection with a recent book. In Mixture and Chemical Combination: and Related Essays [Duh02m], chapter 21, pages 277–290. ISBN 90-481-5924-5, 94-017-2292-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-017-2292-6_21.

Duhem:2015:C

[Duh15a] Pierre Maurice Marie Duhem. Conclusion. In The electric theories of J. Clerk Maxwell: a historical and critical study [Duh15c], chapter 9, pages 171–173. ISBN 3-319-18514-4 (hardcover), 3-319-18515-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC518 .D84 2015. URL https:/

/link.springer.com/chapter/10.
1007/978-3-319-18515-6_9. English

translation by Alan Aversa.

Duhem:2015:CCD

[Duh15b] Pierre Maurice Marie Duhem. Conduction current and displacement current. In The electric theories of J. Clerk Maxwell: a historical and critical study [Duh15c], chapter 6, pages 81–99. ISBN 3-319-18514-4 (hardcover), 3-319-18515-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC518 .D84 2015. URL https:/

/link.springer.com/chapter/10.
1007/978-3-319-18515-6_6. English

translation by Alan Aversa.

Duhem:2015:ETJ

[Duh15c] Pierre Maurice Marie Duhem. The electric theories of J. Clerk Maxwell: a historical and critical study, volume 314 of Boston Studies in the Philosophy and History of Science. Springer-Ver-lag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2015. ISBN 3-319-18514-4 (hardcover), 3-319-18515-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). xv + 185 pp. LCCN QC518 .D84 2015. URL

https://link.springer.com/book/
10.1007/978-3-319-18515-6. English

translation by Alan Aversa.

Duhem:2015:ETL

[Duh15d] Pierre Maurice Marie Duhem. The electromagnetic theory of light. In The electric theories of J. Clerk Maxwell: a historical and critical study [Duh15c], chapter 8, pages 141–169. ISBN 3-319-18514-4 (hardcover), 3-319-18515-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC518 .D84

2015. URL https://link.springer.
com/chapter/10.1007/978-3-319-

18515-6_8. English translation by Alan Aversa.

Duhem:2015:FEM

[Duh15e] Pierre Maurice Marie Duhem. The first electrostatics of Maxwell. In The electric theories of J. Clerk Maxwell: a historical and critical study [Duh15c], chapter 3, pages 35–42. ISBN 3-319-18514-4 (hardcover), 3-319-18515-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC518 .D84

2015. URL https://link.springer.
com/chapter/10.1007/978-3-319-

18515-6_3. English translation by Alan Aversa.

Duhem:2015:FPD

[Duh15f] Pierre Maurice Marie Duhem. The fundamental properties of dielectrics. The doctrines of Faraday and Mossotti. In The electric theories of J. Clerk Maxwell: a historical and critical study [Duh15c], chapter 2, pages 13–33. ISBN 3-319-18514-4 (hardcover), 3-319-18515-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC518 .D84 2015. URL https:/

/link.springer.com/chapter/10.
1007/978-3-319-18515-6_2. English

translation by Alan Aversa.

Duhem:2015:I

[Duh15g] Pierre Maurice Marie Duhem. Introduction. In The electric theories of J. Clerk Maxwell: a historical and critical study [Duh15c], chapter 1, pages 1–9. ISBN 3-319-18514-4 (hardcover), 3-319-18515-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC518 .D84 2015. URL https:/

                                              /link.springer.com/chapter/10.
                                              1007/978-3-319-18515-6_1. English

translation by Alan Aversa.

Duhem:2015:SEMa

[Duh15h] Pierre Maurice Marie Duhem. The second electrostatics of Maxwell. In The electric theories of J. Clerk Maxwell: a historical and critical study [Duh15c], chapter 4, pages 43–63. ISBN 3-319-18514-4 (hardcover), 3-319-18515-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC518 .D84

                                              2015. URL https://link.springer.
                                              com/chapter/10.1007/978-3-319-

18515-6_4. English translation by Alan Aversa.

Duhem:2015:SEMb

[Duh15i] Pierre Maurice Marie Duhem. The six equations of Maxwell and electromagnetic energy. In The electric theories of J. Clerk Maxwell: a historical and critical study [Duh15c], chapter 7, pages 101–140. ISBN 3-319-18514-4 (hardcover), 3-319-18515-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC518 .D84 2015.

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-3-319-18515-

6_7. English translation by Alan Aversa.

Duhem:2015:TEM

[Duh15j] Pierre Maurice Marie Duhem. The third electrostatics of Maxwell. In The electric theories of J. Clerk Maxwell: a historical and critical study [Duh15c], chapter 5, pages 65–78. ISBN 3-319-18514-4 (hardcover), 3-319-18515-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC518 .D84

2015. URL https://link.springer.
com/chapter/10.1007/978-3-319-

18515-6_5. English translation by Alan Aversa.

Dumouchel:1992:MHC

[Dum92] Paul Dumouchel. A morphogenetic hypothesis on the closure of poststructuralism. In Understanding Origins: Contemporary Views on the Origin of Life, Mind and Society [VD92], chapter 4, pages 77–90. ISBN 90-481-4090-0, 94-015-8054-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-015-8054-0_4.

Dumouchel:1996:NSS

[Dum96] Paul Dumouchel. Natural selection and selection type theories. In Québec Studies in the Philosophy of Science: Part II: Biology, Psychology, Cognitive Science and Economics Essays in Honor of Hugues Leblanc [MC96], chapter 2, pages 13–24. ISBN 94-010-6537-3, 94-009-0113-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-009-
0113-1_2.

Dumitru:2015:FLF

[Dum15] Mircea Dumitru. A free logic for fictionalism. In Romanian Studies in Philosophy of Science [PST15], chapter 10, pages 149–163. ISBN 3-319-16654-9 (hardcover), 3-319-16655-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-3-319-16655-
                                              1_10.

Duncan:1989:DMA

[Dun89] Howard Duncan. Descartes and the method of analysis and synthesis. In An Intimate Relation: Studies in the History and Philosophy of Science Presented to Robert E. Butts on his 60th Birthday [BM89], chapter 4, pages 65–80. ISBN 94-010-7546-8, 94-009-2327-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

                                              ???? URL https://link.springer.
                                              com/chapter/10.1007/978-94-009-
                                              2327-0_4.

Dunnell:1992:SAP

[Dun92] Robert C. Dunnell. Is a scientific archaeology possible? In Metaarchaeology: Reflections by Archaeologists and Philosophers [Emb92b], chapter 3, pages 75–97. ISBN 94-010-4806-1, 94-011-1826-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

                                              ???? URL https://link.springer.
                                              com/chapter/10.1007/978-94-011-
                                              1826-2_3.

Dun:1996:CAL

[Dun96] Liu Dun. A comparison of Archimedes’ and Liu Hui’s studies of circles. In Chinese Studies in the History and Philosophy of Science and Technology [DC96], chapter 23, pages 279–287. ISBN 90-481-4546-5, 94-015-8717-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 180; U390 .N19 1996.

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-015-8717-
                                              4_23.

Dupre:1992:OQ

[Dup92] John Dupré. Optimization in question.

In Understanding Origins: Contemporary Views on the Origin of Life, Mind and Society [VD92], chapter 9, pages 183–190. ISBN 90-481-4090-0, 94-015-8054-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-015-
8054-0_9.

Durbin:1983:ISQ

[Dur83] Paul T. Durbin. Introduction: Some questions for philosophy of technology. In Philosophy and Technology [DR83], chapter 1, pages 1–14. ISBN 94-009-7126-5, 94-009-7124-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-009-7124-0_1.

Dutra:2011:NKS

[Dut11] Luiz Henrique Dutra. Natural kinds as scientific models. In Brazilian Studies in Philosophy and History of Science: An account of recent works [KV11], chapter 9, pages 141–150. ISBN 90-481-9421-0 (hardcover), 90-481-9422-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175.3 .B73

2011. URL https://link.springer.
com/chapter/10.1007/978-90-481-
9422-3_9.

Duwell:2017:EFC

[Duw17] Armond Duwell. Exploring the frontiers of computation: Measurement based quantum computers and the mechanistic view of computation. In Floyd and Bokulich [FB17], chapter 9, pages 219–232. ISBN 3-319-53278-2, 3-319-53280-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QA29.T8 P45 2017. URL https:/

                                              /link.springer.com/chapter/10.
                                              1007/978-3-319-53280-6_9.

deVries:1974:SJB

[dV74] Theun de Vries. The secret of Jheronimus Bosch. In For Dirk Struik: Scientific, Historical and Political Essays in Honor of Dirk J. Struik [CSW74], chapter 40, pages 531–546. ISBN 90-277-0379-5, 94-010-2115-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-010-2115-
                                              9_40.

Dupuy:1992:UOI

[DV92] Jean-Pierre Dupuy and Francisco J. Varela. Understanding origins: An introduction. In Understanding Origins: Contemporary Views on the Origin of Life, Mind and Society [VD92], chapter 1, pages 1–25. ISBN 90-481-4090-0, 94-015-8054-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-015-8054-0_1.

deVries:2011:SCI

[dV11] Marc J. de Vries. Science in the context of industrial application: The case of the Philips Natuurkundig Laboratorium. In Science in the Context of Application [CN11a], chapter 4, pages 47–66. ISBN 90-481-9050-9, 90-481-9051-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q180.55.M4 S385 2011. URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-90-481-9051-5_4.

Dvorak:1991:ONA

[Dvo91] Johann Dvorak. Otto Neurath and adult education: Unity of science, materialism and comprehensive enlightenment. In Uebel [Ueb91b], chapter 20, pages 265–274. ISBN 0-7923-1276-7 (hardcover), 94-011-3182-1. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 133. URL https:

//link.springer.com/chapter/10.
1007/978-94-011-3182-7_20.

Dwyer:1996:DE

[Dwy96] Susan Dwyer. Dispositions to explain.

In Québec Studies in the Philosophy of Science: Part II: Biology, Psychology, Cognitive Science and Economics Essays in Honor of Hugues Leblanc [MC96], chapter 15, pages 247–263. ISBN 94-010-6537-3, 94-009-0113-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-009-0113-
1_15.

Erdi:1999:BHD

[ÉA99] Péter Érdi and Ildikó Aradi. The brain as a hermeneutic device. In Fehér et al. [FKR99], pages 183–198. ISBN 90-481-5257-7, 94-015-9293-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-015-9293-2_15.

Eagle:1983:ESR

[Eag83] Morris N. Eagle. The epistemological status of recent developments in psychoanalytic theory. In Cohen and Laudan [CL83], chapter 2, pages 31–55. ISBN 94-009-7055-2, 94-009-7057-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-009-7055-
7_2.

Earle:1994:E

[Ear94] William James Earle. Is the Enlightenment over? In Artifacts, Representations and Social Practice: Essays for Marx Wartofsky [GC94], chapter 12, pages 195–203. ISBN 94-010-4390-6, 94-011-0902-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN B73 .A77 1994.

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-011-0902-
                                              4_12.

Earley:2006:SPI

[Ear06] Joseph E. Earley. Some philosophical implications of chemical symmetry. In Philosophy of Chemistry: Synthesis of a New Discipline [BSM06b], chapter 11, pages 207–220. ISBN 1-4020-2455-X, 1-4020-3256-0, 1-4020-3261-7. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 242 QD6.

                                              URL https://link.springer.com/
                                              chapter/10.1007/1-4020-3261-7_11.

Earley:2015:PPC

[Ear15] Joseph E. Earley, Sr. Pragmatism and the philosophy of chemistry. In Philosophy of Chemistry: Growth of a New Discipline [SM15b], chapter 6, pages 73–89. ISBN 94-017-9363-8 (hardcover), 94-017-9364-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QD6 .P455 2015; QD6 .S384 2014. URL https://link.

                                              springer.com/chapter/10.1007/
                                              978-94-017-9364-3_6.

Easton:1984:FAI

[Eas84] L. D. Easton. The first American interpretation of Hegel in J. B. Stallo’s philosophy of science. In Hegel and the Sciences [CW84a], chapter 17, pages 287–299. ISBN 94-009-6235-5, 94-009-6233-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-009-
6233-0_17.

Ebrahimnejad:2011:WII

[Ebr11]  Hormoz Ebrahimnejad.  What is

‘Islamic’ in Islamic medicine? An overview. In Science between Europe and Asia: Historical Studies on the Transmission, Adoption and Adaptation of Knowledge [GR11b], chapter 17, pages 259–270. ISBN 90-481-9968-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q125 .S354 2010; Q175.52.A78 S35 2010.

URL https://link.springer.com/
chapter/10.1007/978-90-481-9968-
6_17.

Echeverria:1995:GMD

[Ech95] Ana Barahona Echeverria. Genetic mutation: The development of the concept and its evolutionary implications. In Mexican Studies in the History and Philosophy of Science [RC95], chapter 7, pages 89–107. ISBN 94-010-6535-7 (hardcover), 94-009-0109-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-009-0109-4_7.

Echeverria:1996:EMM

[Ech96] Javier Echeverría. Empirical methods in mathematics. In Munévar [Mun96], chapter 2, pages 19–55. ISBN 94-010-6622-1, 94-009-0305-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-009-0305-0_2.

Eckel:1998:TBP

[Eck98] Malcolm David Eckel. Is there a Buddhist philosophy of nature? In Philosophies of Nature: The Human Dimension: In Celebration of Erazim Kohák [CT98], chapter 5, pages 53–69. ISBN 90-481-4859-6, 94-017-2614-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN BD581.

URL https://link.springer.com/
chapter/10.1007/978-94-017-2614-
6_5.

Eckert:2017:ARI

[Eck17] Michael Eckert. Approaching reality by idealization: How fluid resistance was studied by ideal flow theory. In Mathematics as a Tool: Tracing New Roles of Mathematics in the Sciences [LC17b], chapter 11, pages 197–211. ISBN 3-319-54468-3 (hardcover), 3-319-54469-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN GE45.M37; GE45.M38; QA8.4.

URL https://link.springer.com/
chapter/10.1007/978-3-319-54469-
4_11.

Eco:1992:OC

[Eco92] Umberto Eco. The original and the copy. In Understanding Origins: Contemporary Views on the Origin of Life, Mind and Society [VD92], chapter 15, pages 273–303. ISBN 90-481-4090-0, 94-015-8054-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-015-
8054-0_15.

Edel:1988:TPU

[Ede88] Abraham Edel. Theory and practice: An unsteady dichotomy? In Philosophy, History and Social Action: Essays in Honor of Lewis Feuer with an autobiographic essay by Lewis Feuer [HOO88c], chapter 6, pages 157–172. ISBN 94-010-7793-2, 94-009-2873-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-009-2873-
2_6.

Edgar:2015:PSO

[Edg15] Scott Edgar. The physiology of the sense organs and early neo-Kantian conceptions of objectivity: Helmholtz, Lange, Liebmann. In Objectivity in Science: New Perspectives from Science and Technology Studies [PRT15], chapter 6, pages 101–122. ISBN 3-319-14348-4 (hardcover), 3-319-14349-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175 O25

2015. URL https://link.springer.
com/chapter/10.1007/978-3-319-
14349-1_6.

Edmonds:1995:SRR

[Edm95] Dean S. Edmonds. Some reminiscences of Robert Cohen’s physics department. In Gavroglu et al. [GSW95c], chapter 4, pages 47–72. ISBN 0-7923-2991-0 (set), 0-7923-2988-0, 94-017-2658-2. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 163.

URL https://link.springer.com/
chapter/10.1007/978-94-017-2658-
0_4.

Eidlin:2017:PHR

[EE17] Fred Eidlin and Barry Eidlin. The pitfall of hypostatization and the reality of social things. In Encouraging Openness: Essays for Joseph Agassi on the Occasion of His 90th Birthday [BAG17], chapter 33, pages 417–427. ISBN 3-319-57668-2 (hardcover), 3-319-57669-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN B833.E53 2017. URL https://link.

springer.com/chapter/10.1007/
978-3-319-57669-5_33.

Eells:2010:PCC

[Eel10a] Ellery Eells. Probabilistic causality and causal generalizations. In The Place of Probability in Science: In Honor of Ellery Eells (1953–2006) [Fet10a], chapter 1, pages 3–44. ISBN 90-481-3614-8 (hardcover), 90-481-3615-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175 .P563

                                              2010. URL https://link.springer.
                                              com/chapter/10.1007/978-90-481-
                                              3615-5_1.

Eells:2010:PTP

[Eel10b] Ellery Eells. Propensity trajectories, preemption, and the identity of events. In The Place of Probability in Science: In Honor of Ellery Eells (1953–2006) [Fet10a], chapter 8, pages 181–199. ISBN 90-481-3614-8 (hardcover), 90-481-3615-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175 .P563 2010. URL https://link.

                                              springer.com/chapter/10.1007/
                                              978-90-481-3615-5_8.

Eliav-Feldon:1986:MSU

[EF86] Miriam Eliav-Feldon. Metamorphoses of the scientist in utopia: A comment. In Ullmann-Margalit [UM86c], chapter 2, pages 17–20. ISBN 90-277-2160-2, 90-277-2161-0 (paperback), 94-009-4566-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 95 Q175.

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-009-4566-
                                              1_2.

Elazar:2018:JCI

[EF18] Michael Elazar and Rivka Feldhay. Jesuit conceptions of impetus after Galileo: Honoré Fabri, Paolo Casati, and Francesco Eschinardi. In Feldhay et al. [FRSV18a], chapter 10, pages 285–323. ISBN 3-319-90343-8 (hardcover), 3-319-90345-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QA802. URL https:

//link.springer.com/chapter/10.
1007/978-3-319-90345-3_10.

Ehlers:2003:GLS

[EFN03] Jürgen Ehlers, Simonetta Frittelli, and Ezra T. Newman. Gravitational lensing from a space–time perspective. In Renn et al. [RDS03], chapter 12, pages 281–304. ISBN 1-4020-1284-5, 1-4020-1285-3 (paperback). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 234. URL https:/

/link.springer.com/chapter/10.
1007/978-94-010-0111-3_12.

Efron:1969:WP

[Efr69] Robert Efron. What is perception? In Cohen and Wartofsky [CW69b], pages 137–173. ISBN 94-010-3380-3, 94-010-3378-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-010-
3378-7_4.

Eger:1999:LDH

[Ege99] Martin Eger. Language and the double hermeneutic in natural science. In Fehér et al. [FKR99], pages 265–280. ISBN 90-481-5257-7, 94-015-9293-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-015-9293-
2_21.

Eilstein:1995:VF

[Eil95] Helena Eilstein. The virus of fatalism.

In Science, Mind and Art: Essays on science and the humanistic understanding in art, epistemology, religion and ethics In honor of Robert S. Cohen [GSW95b], chapter 6, pages 71–88. ISBN 0-7923-2990-2 (hardcover), 94-011-0469-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 165 BD418.3.

URL https://link.springer.com/
chapter/10.1007/978-94-011-0469-
2_6.

Eintalu:2001:OPT

[Ein01] Jüri Eintalu. Outsolutions in physical theories. Physical considerations. In Estonian Studies in the History and Philosophy of Science [Vih01b], chapter 16, pages 215–230. ISBN 94-010-3863-5, 94-010-0672-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-010-0672-9_16.

Einstein:2007:ETR

[Ein07a] Albert Einstein. Ether and the theory of relativity. In Renn et al. [RSS^{+}07a], chapter 34, pages 613–619. ISBN 1-4020-3999-9. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 no. 250 v. 3; QC173.6 .G469 2007.

URL https://link.springer.com/
chapter/10.1007/978-1-4020-4000-

9_34. Originally published as “Äther und Relativitätstheorie at Springer, Berlin on the basis of an address delivered on 5 May 1920 at the University of Leyden.

Einstein:2007:PSP

[Ein07b] Albert Einstein. On the present state of the problem of gravitation. In Renn et al. [RSS^{+}07a], chapter 31, pages 543–566. ISBN 1-4020-3999-9. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 no. 250 v. 3; QC173.6 .G469 2007. URL https:

//link.springer.com/chapter/10.
1007/978-1-4020-4000-9_31. Origi-

nally published as “Zum gegenwärtigen Stande des Gravitationsproblems”, Physikalische Zeitschrift 14 1249–1262 (1913).

Einstein:2007:RP

[Ein07c] Albert Einstein. On the relativity problem. In Renn et al. [RSS^{+}07a], chapter 33, pages 605–612. ISBN 1-4020-3999-9. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 no. 250 v. 3; QC173.6 .G469

2007. URL https://link.springer.
com/chapter/10.1007/978-1-4020-

4000-9_33. Originally published as “Zum Relativitätsproblem”, Scientia 15, 337–348 (1914).

Eisenstaedt:2003:PR

[Eis03] Jean Eisenstaedt. The prehistory of relativity. In Renn et al. [RDS03], chapter 1, pages 3–12. ISBN 1-4020-1284-5, 1-4020-1285-3 (paperback). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 234.

URL https://link.springer.com/
chapter/10.1007/978-94-010-0111-
3_1.

Ekstein:1984:PSU

[Eks84] Hans Ekstein. Is physical space unique or optional? In Physical Sciences and History of Physics [CW84c], chapter 4, pages 57–74. ISBN 94-009-7180-X, 94-009-7178-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .P497 1984. URL https:

//link.springer.com/chapter/10.
1007/978-94-009-7178-3_4.

Esfeld:2011:OSR

[EL11] Michael Esfeld and Vincent Lam. Ontic structural realism as a metaphysics of objects. In Bokulich and Bokulich [BB11], chapter 8, pages 143–159. ISBN 90-481-9596-9, 90-481-9597-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175.32.S78 S35

2011. URL https://link.springer.
com/chapter/10.1007/978-90-481-
9597-8_8.

Elazar:2011:AGT

[Ela11a] Michael Elazar. The assimilation of Galileo’s theory. In Honoré Fabri and the Concept of Impetus: a Bridge Between Conceptual Frameworks [Ela11j], chapter 10, pages 97–105. ISBN 94-007-1604-4 (hardcover), 94-007-1605-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175 .E43 2011 v.288; QA802 .E43 2011.

URL https://link.springer.com/
chapter/10.1007/978-94-007-1605-
6_10.

Elazar:2011:CII

[Ela11b] Michael Elazar. The conservation and inexhaustibility of impetus. In Honoré Fabri and the Concept of Impetus: a Bridge Between Conceptual Frameworks [Ela11j], chapter 13, pages 127–133. ISBN 94-007-1604-4 (hardcover), 94-007-1605-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175 .E43 2011 v.288; QA802 .E43 2011. URL https:

//link.springer.com/chapter/10.
1007/978-94-007-1605-6_13.

Elazar:2011:CA

[Ela11c] Michael Elazar. Criticizing Aristotle. In Honoré Fabri and the Concept of Impetus: a Bridge Between Conceptual Frameworks [Ela11j], chapter 7, pages 67–78. ISBN 94-007-1604-4 (hardcover), 94-007-1605-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175 .E43 2011 v.288; QA802 .E43 2011. URL https:

//link.springer.com/chapter/10.
1007/978-94-007-1605-6_7.

Elazar:2011:AIA

[Ela11d] Michael Elazar. De accidente: Impetus and the accidents problem. In Honoré Fabri and the Concept of Impetus: a Bridge Between Conceptual Frameworks [Ela11j], chapter 18, pages 197–210. ISBN 94-007-1604-4 (hardcover), 94-007-1605-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175 .E43 2011 v.288; QA802 .E43 2011. URL https:

//link.springer.com/chapter/10.
1007/978-94-007-1605-6_18.

Elazar:2011:FCR

[Ela11e] Michael Elazar. Fabri and conservation of rectilinear motion. In Honoré Fabri and the Concept of Impetus: a Bridge Between Conceptual Frameworks [Ela11j], chapter 12, pages 121–125. ISBN 94-007-1604-4 (hardcover), 94-007-1605-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175 .E43 2011 v.288; QA802 .E43

2011. URL https://link.springer.
com/chapter/10.1007/978-94-007-
1605-6_12.

Elazar:2011:FSG

[Ela11f] Michael Elazar. Fabri and the “Second Galilean Affair”. In Honoré Fabri and the Concept of Impetus: a Bridge Between Conceptual Frameworks [Ela11j], chapter 6, pages 63–66. ISBN 94-007-1604-4 (hardcover), 94-007-1605-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175 .E43 2011 v.288; QA802 .E43 2011.

URL https://link.springer.com/
chapter/10.1007/978-94-007-1605-
6_6.

Elazar:2011:FAS

[Ela11g] Michael Elazar. Fabri’s assimilation strategies. In Honoré Fabri and the Concept of Impetus: a Bridge Between Conceptual Frameworks [Ela11j], chapter 11, pages 107–117. ISBN 94-007-1604-4 (hardcover), 94-007-1605-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175 .E43 2011 v.288; QA802 .E43 2011.

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-007-1605-
                                              6_11.

Elazar:2011:FDA

[Ela11h] Michael Elazar. Fabri’s discrete analysis. In Honoré Fabri and the Concept of Impetus: a Bridge Between Conceptual Frameworks [Ela11j], chapter 9, pages 87–96. ISBN 94-007-1604-4 (hardcover), 94-007-1605-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175 .E43 2011 v.288; QA802 .E43 2011. URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-007-1605-6_9.

Elazar:2011:FIS

[Ela11i] Michael Elazar. Fabri’s impetus and the scholastic tradition. In Honoré Fabri and the Concept of Impetus: a Bridge Between Conceptual Frameworks [Ela11j], chapter 5, pages 47–60. ISBN 94-007-1604-4 (hardcover), 94-007-1605-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175 .E43 2011 v.288; QA802 .E43

                                              2011. URL https://link.springer.
                                              com/chapter/10.1007/978-94-007-
                                              1605-6_5.

Elazar:2011:HFC

[Ela11j] Michael Elazar. Honoré Fabri and the Concept of Impetus: a Bridge Between Conceptual Frameworks, volume 288 of Boston Studies in the Philosophy of Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2011. ISBN 94-007-1604-4 (hardcover), 94-007-1605-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). xxix + 259 pp. LCCN Q175 .E43 2011 v.288; QA802

.E43 2011. URL http://www.loc.
gov/catdir/enhancements/fy1202/
2011928368-d.html;   http://www.
loc.gov/catdir/enhancements/
fy1202/2011928368-t.html;
https://link.springer.com/book/
10.1007/978-94-007-1605-6.

Elazar:2011:I

[Ela11k] Michael Elazar. Impetus. In Honoré Fabri and the Concept of Impetus: a Bridge Between Conceptual Frameworks [Ela11j], chapter 4, pages 31–45. ISBN 94-007-1604-4 (hardcover), 94-007-1605-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175 .E43 2011 v.288; QA802 .E43

2011. URL https://link.springer.
com/chapter/10.1007/978-94-007-
1605-6_4.

Elazar:2011:ILI

[Ela11l] Michael Elazar. The inherent linearity of impetus. In Honoré Fabri and the Concept of Impetus: a Bridge Between Conceptual Frameworks [Ela11j], chapter 15, pages 157–168. ISBN 94-007-1604-4 (hardcover), 94-007-1605-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175 .E43 2011 v.288; QA802 .E43 2011.

URL https://link.springer.com/
chapter/10.1007/978-94-007-1605-
6_15.

Elazar:2011:IPI

[Ela11m] Michael Elazar. Introduction: The primacy of impetus. In Honoré Fabri and the Concept of Impetus: a Bridge Between Conceptual Frameworks [Ela11j], chapter 1, pages 3–7. ISBN 94-007-1604-4 (hardcover), 94-007-1605-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175 .E43 2011 v.288; QA802 .E43 2011.

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-007-1605-
                                              6_1.

Elazar:2011:LNN

[Ela11n] Michael Elazar. The law of natural numbers. In Honoré Fabri and the Concept of Impetus: a Bridge Between Conceptual Frameworks [Ela11j], chapter 8, pages 79–86. ISBN 94-007-1604-4 (hardcover), 94-007-1605-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175 .E43 2011 v.288; QA802 .E43 2011. URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-007-1605-6_8.

Elazar:2011:MMT

[Ela11o] Michael Elazar. Methodology and main texts. In Honoré Fabri and the Concept of Impetus: a Bridge Between Conceptual Frameworks [Ela11j], chapter 2, pages 9–16. ISBN 94-007-1604-4 (hardcover), 94-007-1605-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175 .E43 2011 v.288; QA802 .E43 2011. URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-007-1605-6_2.

Elazar:2011:M

[Ela11p] Michael Elazar. Motion. In Honoré Fabri and the Concept of Impetus: a Bridge Between Conceptual Frameworks [Ela11j], chapter 3, pages 17–29. ISBN 94-007-1604-4 (hardcover), 94-007-1605-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175 .E43 2011 v.288; QA802 .E43

2011. URL https://link.springer.
com/chapter/10.1007/978-94-007-
1605-6_3.

Elazar:2011:MV

[Ela11q] Michael Elazar. Motion in the void.

In Honoré Fabri and the Concept of Impetus: a Bridge Between Conceptual Frameworks [Ela11j], chapter 14, pages 135–155. ISBN 94-007-1604-4 (hardcover), 94-007-1605-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175 .E43 2011 v.288; QA802 .E43 2011. URL https:

//link.springer.com/chapter/10.
1007/978-94-007-1605-6_14.

Elazar:2011:PSR

[Ela11r] Michael Elazar. Physica: Solution of the real presence problem. In Honoré Fabri and the Concept of Impetus: a Bridge Between Conceptual Frameworks [Ela11j], chapter 19, pages 211–229. ISBN 94-007-1604-4 (hardcover), 94-007-1605-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175 .E43 2011 v.288; QA802 .E43 2011. URL https:

//link.springer.com/chapter/10.
1007/978-94-007-1605-6_19.

Elazar:2011:PMR

[Ela11s] Michael Elazar. Projectile motion and the rejection of superposition. In Honoré Fabri and the Concept of Impetus: a Bridge Between Conceptual Frameworks [Ela11j], chapter 16, pages 169–187. ISBN 94-007-1604-4 (hardcover), 94-007-1605-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175 .E43 2011 v.288; QA802 .E43 2011. URL https:

//link.springer.com/chapter/10.
1007/978-94-007-1605-6_16.

Elazar:2011:TME

[Ela11t] Michael Elazar. The twofold mystery of the Eucharist. In Honoré Fabri and the Concept of Impetus: a Bridge Between Conceptual Frameworks [Ela11j], chapter 17, pages 191–195. ISBN 94-007-1604-4 (hardcover), 94-007-1605-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175 .E43 2011 v.288; QA802 .E43 2011.

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-007-1605-
                                              6_17.

Elena:1992:MES

[Ele92] Alberto Elena. Models of European scientific expansion: The Ottoman Empire as a source of evidence. In Science and Empires: Historical Studies about Scientific Development and European Expansion [PJM92], chapter 27, pages 259–267. ISBN 0-7923-1518-9 (hardcover), 94-010-5145-3, 94-011-2594-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 136. URL https://link.

                                              springer.com/chapter/10.1007/
                                              978-94-011-2594-9_27.

Elevitch:1994:LP

[Ele94] Bernard Elevitch. Life is not a poem? In Artifacts, Representations and Social Practice: Essays for Marx Wartofsky [GC94], chapter 21, pages 361–364. ISBN 94-010-4390-6, 94-011-0902-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN B73 .A77

                                              1994. URL https://link.springer.
                                              com/chapter/10.1007/978-94-011-
                                              0902-4_21.

Elgin:1984:TRQ

[Elg84] Catherine Z. Elgin. Theory reduction: A question of fact or a question of value? In Physical Sciences and History of Physics [CW84c], chapter 5, pages 75–92. ISBN 94-009-7180-X, 94-009-7178-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .P497 1984. URL https:

//link.springer.com/chapter/10.
1007/978-94-009-7178-3_5.

Elgin:2010:TI

[Elg10] Catherine Z. Elgin. Telling instances.

In Frigg and Hunter [FH10b], chapter 1, pages 1–17. ISBN 90-481-3850-7 (hardcover), 90-481-3851-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN BH301.M47 B49 2010; Q175 .B52 2010. URL https:

//link.springer.com/chapter/10.
1007/978-90-481-3851-7_1.

Elkana:1974:SMP

[Elk74] Yehuda Elkana. Scientific and metaphysical problems: Euler and Kant. In Methodological and Historical Essays in the Natural and Social Sciences [CW74b], chapter 15, pages 277–305. ISBN 90-277-0378-7, 94-010-2128-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 14. URL https://link.

springer.com/chapter/10.1007/
978-94-010-2128-9_15.

Elkana:1976:ICC

[Elk76] Yehuda Elkana. Introduction: Culture, cultural system and science. In Essays in Memory of Imre Lakatos [CFW76], chapter 10, pages 99–107. ISBN 90-277-0655-7, 94-010-1451-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-010-1451-9_10.

Elkana:1986:TDB

[Elk86] Yehuda Elkana. Is there a distinction between external and internal sociology of science. In Cognition and Fact: Materials on Ludwik Fleck [CS86], chapter 16, pages 309–316. ISBN 94-010-8504-8, 94-009-4498-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-009-4498-5_16.

Elkins:2010:VPA

[Elk10] James Elkins. Visual practices across the university. In Frigg and Hunter [FH10b], chapter 8, pages 169–192. ISBN 90-481-3850-7 (hardcover), 90-481-3851-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN BH301.M47 B49 2010; Q175 .B52 2010.

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-90-481-3851-
                                              7_8.

Ellis:1984:CV

[Ell84] George F. R. Ellis. Cosmology and verifiability. In Physical Sciences and History of Physics [CW84c], chapter 6, pages 93–113. ISBN 94-009-7180-X, 94-009-7178-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .P497 1984. URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-009-7178-3_6.

Ellis:2003:UNC

[Ell03] George F. R. Ellis. The unique nature of cosmology. In Renn et al. [RDS03], chapter 9, pages 193–220. ISBN 1-4020-1284-5, 1-4020-1285-3 (paperback). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 234.

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-010-0111-
                                              3_9.

Elster:1988:NSR

[Els88] Jon Elster. The nature and scope of rational-choice explanation. In Science in Reflection: The Israel Colloquium: Studies in History, Philosophy, and Sociology of Science Volume 3 [UM88], chapter 5, pages 51–65. ISBN 90-277-2713-9, 94-009-2957-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-009-2957-9_5.

Elzinga:1995:ELN

[Elz95] Aant Elzinga. Einstein in the Land of Nobel: An episode in the interplay of science, politics, epistemology and popular culture. In Gavroglu et al. [GSW95c], chapter 5, pages 73–103. ISBN 0-7923-2991-0 (set), 0-7923-2988-0, 94-017-2658-2. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 163. URL https:/

/link.springer.com/chapter/10.
1007/978-94-017-2658-0_5.

Embree:1992:FPM

[Emb92a] Lester Embree. The future and past of metaarchaeology. In Metaarchaeology: Reflections by Archaeologists and Philosophers [Emb92b], chapter 1, pages 3–50. ISBN 94-010-4806-1, 94-011-1826-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-011-
1826-2_1.

Embree:1992:MRA

[Emb92b] Lester Embree. Metaarchaeology:

Reflections by Archaeologists and Philosophers, volume 147 of Boston Studies in the Philosophy of Science. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 1992. ISBN 94-010-4806-1, 94-011-1826-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). vii + 343 pp. LCCN ????

                                              URL https://link.springer.com/
                                              book/10.1007/978-94-011-1826-2.

Embree:1992:PCA

[Emb92c] Lester Embree. Phenomenology of a change in archaeological observation. In Metaarchaeology: Reflections by Archaeologists and Philosophers [Emb92b], chapter 7, pages 165–193. ISBN 94-010-4806-1, 94-011-1826-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-011-1826-
                                              2_7.

Englund:1996:FRN

[EN96] Robert K. Englund and Hans J. Nissen. The first representations of numbers and the development of the number concept. In Abstraction and Representation: Essays on the Cultural Evolution of Thinking [Dam96b], chapter 8, pages 275–297. ISBN 0-7923-3816-2, 94-015-8624-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 175.

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-015-8624-
                                              5_8.

Enc:2005:HCC

[Enç05] Berent Enç. How causes can rationalize: Belief–desire explanations of action. In Turkish Studies in the History and Philosophy of Science [IG05c], chapter 16, pages 231–247. ISBN 1-4020-3332-X (hardcover), 1-4020-3333-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 244.

                                              URL https://link.springer.com/
                                              chapter/10.1007/1-4020-3333-8_16.

Engelbrecht:2001:SSF

[Eng01] Jüri Engelbrecht. Science and society — faculties close or apart? In Estonian Studies in the History and Philosophy of Science [Vih01b], chapter 6, pages 77–88. ISBN 94-010-3863-5, 94-010-0672-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-010-
0672-9_6.

Enz:2000:ORM

[Enz00] Charles Paul Enz. Observability and realism in modern experiments with correlated quantum systems. In The Reality of the Unobservable: Observability, Unobservability and Their Impact on the Issue of Scientific Realism [AP00b], chapter 22, pages 293–310. ISBN 90-481-5458-8, 94-015-9391-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-015-9391-
5_22.

Engberts:2017:SVB

[EP17] Christiaan Engberts and Herman Paul.

Scholarly vices: Boundary work in Nineteenth-Century orientalism. In van Dongen and Paul [vDP17a], chapter 6, pages 79–90. ISBN 3-319-48892-9 (hardcover), 3-319-48893-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 v. 321; Q175.32.V57 E65 2017. URL https:

//link.springer.com/chapter/10.
1007/978-3-319-48893-6_6.

Edel:1969:PUS

[EPL69] Abraham Edel, Ruth Anna Putnam, and John Ladd. Patterns of use of science in ethics. In Cohen and Wartofsky [CW69b], pages 350–391. ISBN 94-010-3380-3, 94-010-3378-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-010-3378-7_13.

Erpenbeck:1991:PPM

[Erp91] John Erpenbeck. Philosophical problems of modern psychology. In Woodward and Cohen [WC91], pages 209–222. ISBN 94-010-5395-2, 94-011-3164-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

                                              ???? URL https://link.springer.
                                              com/chapter/10.1007/978-94-011-
                                              3164-3_19.

Erwin:1971:CM

[Erw71]  Edward Erwin.  The confirmation

machine. In Buck and Cohen [BC71], pages 306–321. ISBN 90-277-0309-4, 94-010-3142-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

                                              ???? URL https://link.springer.
                                              com/chapter/10.1007/978-94-010-
                                              3142-4_21.

Esanu:2015:IMO

[Eşa15] Andreea Eşanu. Issues in modeling open-ended evolution. In Romanian Studies in Philosophy of Science [PST15], chapter 6, pages 87–101. ISBN 3-319-16654-9 (hardcover), 3-319-16655-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

                                              ???? URL https://link.springer.
                                              com/chapter/10.1007/978-3-319-
                                              16655-1_6.

Estany:1996:RMM

[Est96] Anna Estany. The role of methodology in models of scientific change. In Munévar [Mun96], chapter 14, pages 275–297. ISBN 94-010-6622-1, 94-009-0305-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

                                              ???? URL https://link.springer.
                                              com/chapter/10.1007/978-94-009-
                                              0305-0_14.

Euler:1976:ARA

[Eul76] L. Euler. Argument for the reality of absolute space. In The Concepts of Space and Time: Their Structure and Their Development [Čap76], chapter 20, pages 113–119. ISBN 90-277-0375-2, 94-010-1727-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic).

LCCN BD632. URL https://link.
springer.com/chapter/10.1007/
978-94-010-1727-5_20.

Evans:1995:MP

[Eva95] J. David G. Evans. Meno’s puzzle.

In The Concept of Knowledge: The Ankara Seminar [KC95], chapter 6, pages 67–80. ISBN 90-481-4495-7, 94-017-3263-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-017-
3263-5_6.

Eves:1974:SCS

[Eve74] Howard Eves. Skew curves setting up a null system in space. In For Dirk Struik: Scientific, Historical and Political Essays in Honor of Dirk J. Struik [CSW74], chapter 6, pages 67–74. ISBN 90-277-0379-5, 94-010-2115-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-010-2115-9_6.

Ezrahi:1986:KPS

[Ezr86] Yaron Ezrahi. Knowledge and power in the sciences: A comment. In Ullmann-Margalit [UM86a], chapter 19, pages 241–245. ISBN 90-277-2159-9 (paperback), 90-277-2158-0 (hardcover), 94-009-5496-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 94 Q175.

URL https://link.springer.com/
chapter/10.1007/978-94-009-5496-
0_19.

Ezrahi:2001:SJP

[Ezr01] Yaron Ezrahi. “Speaking justice to power”: Everett Mendelsohn as an international public intellectual. In Allen and MacLeod [AM01b], chapter 21, pages 365–368. ISBN 1-4020-0495-8 (hardcover), 94-017-2956-5. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 228; Q175.55.

URL https://link.springer.com/
chapter/10.1007/978-94-017-2956-
7_21.

Ezrahi:2014:BDC

[Ezr14] Yaron Ezrahi. Ben-David’s critique of the sociology of knowledge and his politics of freedom. In The Institution of Science and the Science of Institutions: The Legacy of Joseph Ben-David [Her14b], chapter 3, pages 39–45. ISBN 94-007-7406-0 (hardcover), 94-007-7407-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 v.302; Q125 .I58 2014.

URL https://link.springer.com/
chapter/10.1007/978-94-007-7407-
0_3.

Fried:1976:CM

[FA76a] Yehuda Fried and Joseph Agassi. Clinical matters. In Paranoia: a Study in Diagnosis [FA76g], chapter 8, pages 75– 98. ISBN 90-277-0705-7, 94-010-1506-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-010-1506-
6_8.

Fried:1976:CTG

[FA76b] Yehuda Fried and Joseph Agassi. Conclusion towards a general demarcation of psychopathology. In Paranoia: a Study in Diagnosis [FA76g], chapter 9, pages 99–106. ISBN 90-277-0705-7, 94-010-1506-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-010-
1506-6_9.

Fried:1976:IPP

[FA76c] Yehuda Fried and Joseph Agassi. Introduction: The paradoxes of paranoia. In Paranoia: a Study in Diagnosis [FA76g], chapter 1, pages 1–8. ISBN 90-277-0705-7, 94-010-1506-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-010-1506-6_1.

Fried:1976:MBb

[FA76d] Yehuda Fried and Joseph Agassi. Metaphysical background. In Paranoia: a Study in Diagnosis [FA76g], chapter 5, pages 40–51. ISBN 90-277-0705-7, 94-010-1506-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-010-
1506-6_5.

Fried:1976:MBa

[FA76e] Yehuda Fried and Joseph Agassi.

Methodological background. In Paranoia: a Study in Diagnosis [FA76g], chapter 4, pages 29–39. ISBN 90-277-0705-7, 94-010-1506-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-010-1506-6_4.

Fried:1976:PPR

[FA76f] Yehuda Fried and Joseph Agassi.

The paradoxes of paranoia revisited.

In Paranoia: a Study in Diagnosis [FA76g], chapter 6, pages 52–60. ISBN 90-277-0705-7, 94-010-1506-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-010-1506-6_6.

Fried:1976:PSD

[FA76g] Yehuda Fried and Joseph Agassi. Paranoia: a Study in Diagnosis, volume 50 of Boston Studies in the Philosophy of Science. D. Reidel, Dordrecht, The Netherlands; Boston, MA, USA; Lancaster, UK; Tokyo, Japan, 1976. ISBN 90-277-0705-7, 94-010-1506-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). xv + 218 pp. LCCN ????

URL https://link.springer.com/
book/10.1007/978-94-010-1506-6.

Fried:1976:PFA

[FA76h] Yehuda Fried and Joseph Agassi. Paranoia as a fixation of an abstract system. In Paranoia: a Study in Diagnosis [FA76g], chapter 7, pages 61–74. ISBN 90-277-0705-7, 94-010-1506-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-010-1506-6_7.

Fried:1976:P

[FA76i] Yehuda Fried and Joseph Agassi. Postscript. In Paranoia: a Study in Diagnosis [FA76g], chapter 10, page 107. ISBN 90-277-0705-7, 94-010-1506-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-010-1506-
6_10.

Fried:1976:PB

[FA76j] Yehuda Fried and Joseph Agassi. Psychological background. In Paranoia: a Study in Diagnosis [FA76g], chapter 2, pages 9–19. ISBN 90-277-0705-7, 94-010-1506-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-010-
1506-6_2.

Fried:1976:SB

[FA76k] Yehuda Fried and Joseph Agassi. Sociological background. In Paranoia: a Study in Diagnosis [FA76g], chapter 3, pages 20–28. ISBN 90-277-0705-7, 94-010-1506-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-010-
1506-6_3.

Faber:1984:FSF

[Fab84] Roger J. Faber. Feedback, selection, and function: A reductionistic account of goal–orientation. In Methodology, Metaphysics and the History of Science: In Memory of Benjamin Nelson [CW84b], chapter 3, pages 43–135. ISBN 94-009-6333-5, 94-009-6331-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-009-6331-
3_3.

Fagan:2012:SES

[Fag12] Melinda Bonnie Fagan. Social epistemology of stem cell research: Philosophy and experiment. In Mauskopf and Schmaltz [MS12b], chapter 13, pages 221–239. ISBN 94-007-1744-X, 94-007-1745-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175 .I5134 2011. URL https:

//link.springer.com/chapter/10.
1007/978-94-007-1745-9_13.

Falk:1986:CGE

[Fal86] Raphael Falk. Can genetics explain development? In Ullmann-Margalit [UM86c], chapter 13, pages 165–180. ISBN 90-277-2160-2, 90-277-2161-0 (paperback), 94-009-4566-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 95

                                              Q175. URL https://link.springer.
                                              com/chapter/10.1007/978-94-009-
                                              4566-1_13.

Falkenburg:2000:HOQ

[Fal00] Brigitte Falkenburg. How to observe quarks. In The Reality of the Unobservable: Observability, Unobservability and Their Impact on the Issue of Scientific Realism [AP00b], chapter 25, pages 329–341. ISBN 90-481-5458-8, 94-015-9391-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

                                              ???? URL https://link.springer.
                                              com/chapter/10.1007/978-94-015-
                                              9391-5_25.

Falk:2001:MH

[Fal01] Raphael Falk. Mendel’s hypothesis. In Allen and MacLeod [AM01b], chapter 5, pages 77–86. ISBN 1-4020-0495-8 (hardcover), 94-017-2956-5. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 228; Q175.55.

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-017-2956-
                                              7_5.

Farrell:2003:FP

[Far03a] Robert P. Farrell. Feyerabend and pluralism. In Feyerabend and Scientific Values: Tightrope-Walking Rationality [Far03b], chapter 6, pages 133–146. ISBN 90-481-6309-9, 94-017-1542-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174

.F377 2003. URL https://link.
springer.com/chapter/10.1007/
978-94-017-1542-3_6.

Farrell:2003:FSV

[Far03b] Robert P. Farrell. Feyerabend and Scientific Values: Tightrope-Walking Rationality, volume 235 of Boston Studies in the Philosophy of Science. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 2003. ISBN 90-481-6309-9, 94-017-1542-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). x + 259 pp. LCCN Q174 .F377 2003.

URL https://link.springer.com/
book/10.1007/978-94-017-1542-3.

Farrell:2003:FWV

[Far03c] Robert P. Farrell. Feyerabend and world-views. In Feyerabend and Scientific Values: Tightrope-Walking Rationality [Far03b], chapter 10, pages 214–236. ISBN 90-481-6309-9, 94-017-1542-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .F377 2003. URL https://link.

springer.com/chapter/10.1007/
978-94-017-1542-3_10.

Farrell:2003:FLA

[Far03d] Robert P. Farrell. Feyerabend, Lakatos and anarchism. In Feyerabend and Scientific Values: Tightrope-Walking Rationality [Far03b], chapter 3, pages 44–70. ISBN 90-481-6309-9, 94-017-1542-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .F377 2003. URL https://link.

springer.com/chapter/10.1007/
978-94-017-1542-3_3.

Farrell:2003:FRA

[Far03e] Robert P. Farrell. Feyerabend’s Reductio Ad Absurdum of ‘rationalist’ philosophies. In Feyerabend and Scientific Values: Tightrope-Walking Rationality [Far03b], chapter 2, pages 5–43. ISBN 90-481-6309-9, 94-017-1542-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .F377 2003. URL https://link.

springer.com/chapter/10.1007/
978-94-017-1542-3_2.

Farrell:2003:FIP

[Far03f] Robert P. Farrell. Feyerabend’s incommensurability and the pragmatic theory of observation. In Feyerabend and Scientific Values: Tightrope-Walking Rationality [Far03b], chapter 4, pages 71–99. ISBN 90-481-6309-9, 94-017-1542-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .F377 2003. URL https://link.

springer.com/chapter/10.1007/
978-94-017-1542-3_4.

Farrell:2003:FR

[Far03g] Robert P. Farrell. Feyerabend’s relativism. In Feyerabend and Scientific Values: Tightrope-Walking Rationality [Far03b], chapter 5, pages 100–132. ISBN 90-481-6309-9, 94-017-1542-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174

.F377 2003. URL https://link.
springer.com/chapter/10.1007/
978-94-017-1542-3_5.

Farrell:2003:I

[Far03h]  Robert P. Farrell.  Introduction.

In Feyerabend and Scientific Values: Tightrope-Walking Rationality [Far03b], chapter 1, pages 1–4. ISBN 90-481-6309-9, 94-017-1542-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .F377 2003.

URL https://link.springer.com/
chapter/10.1007/978-94-017-1542-
3_1.

Farrell:2003:KPR

[Far03i] Robert P. Farrell. Kuhn on paradigms, rules and values. In Feyerabend and Scientific Values: Tightrope-Walking Rationality [Far03b], chapter 8, pages 163–187. ISBN 90-481-6309-9, 94-017-1542-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .F377 2003. URL https://link.

springer.com/chapter/10.1007/
978-94-017-1542-3_8.

Farrell:2003:PMP

[Far03j] Robert P. Farrell. Popper, methodological pluralism and epistemic values. In Feyerabend and Scientific Values: Tightrope-Walking Rationality [Far03b], chapter 7, pages 147–162. ISBN 90-481-6309-9, 94-017-1542-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174

.F377 2003. URL https://link.
springer.com/chapter/10.1007/
978-94-017-1542-3_7.

Farrell:2003:TWR

[Far03k] Robert P. Farrell. Tightrope-walking rationality: Feyerabend’s metanarrative. In Feyerabend and Scientific Values: Tightrope-Walking Rationality [Far03b], chapter 9, pages 188–213. ISBN 90-481-6309-9, 94-017-1542-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174

.F377 2003. URL https://link.
springer.com/chapter/10.1007/
978-94-017-1542-3_9.

Favrholdt:1994:NBR

[Fav94] David Favrholdt. Niels Bohr and realism. In Faye and Folse [FF94], chapter 4, pages 77–96. ISBN 0-7923-2378-5 (hardcover), 94-015-8106-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 153 1994; QC16 1994. URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-015-8106-6_4.

Faxon:1997:IAS

[Fax97] Alicia Craig Faxon. Intersections of art and science to create aesthetic perception: The problem of postmodernism. In The Elusive Synthesis: Aesthetics and Science [Tau97a], chapter 12, pages 251–266. ISBN 0-7923-4763-3, 0-7923-3904-5, 94-009-1786-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-009-1786-6_12.

Faye:1994:NLN

[Fay94]  Jan Faye.  Non-locality or non-

separability? In Faye and Folse [FF94], chapter 5, pages 97–118. ISBN 0-7923-2378-5 (hardcover), 94-015-8106-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 153 1994; QC16 1994. URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-015-8106-6_5.

Faye:1997:CRD

[Fay97a] Jan Faye. Causation, reversibility and the direction of time. In Perspectives on Time [FSU97b], chapter 11, pages 237–266. ISBN 90-481-4774-3, 94-015-8875-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol.189. URL https:/

                                              /link.springer.com/chapter/10.
                                              1007/978-94-015-8875-1_11.

Faye:1997:MMT

[Fay97b] Jan Faye. Is the Mark method time dependent? In Perspectives on Time [FSU97b], chapter 10, pages 215–236. ISBN 90-481-4774-3, 94-015-8875-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol.189. URL https://link.

springer.com/chapter/10.1007/
978-94-015-8875-1_10.

Faye:2000:OU

[Fay00] Jan Faye. Observing the unobservable.

In The Reality of the Unobservable: Observability, Unobservability and Their Impact on the Issue of Scientific Realism [AP00b], chapter 12, pages 165–175. ISBN 90-481-5458-8, 94-015-9391-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-015-
9391-5_12.

Faye:2007:PRT

[Fay07] Jan Faye. The pragmatic–rhetorical theory of explanation. In Rethinking Explanation [PY07], chapter 4, pages 43–68. ISBN 1-4020-5580-3 (paperback), 1-4020-5581-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN BD237 .R48 2007; Q175.32.E97 R48 2007. URL https:

//link.springer.com/chapter/10.
1007/978-1-4020-5581-2_4.

Floyd:2017:PEL

[FB17] Juliet Floyd and Alisa Bokulich, editors. Philosophical Explorations of the Legacy of Alan Turing: Turing 100, volume 324 of Boston Studies in the Philosophy of Science. Springer-Ver-lag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2017. ISBN 3-319-53278-2, 3-319-53280-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). xvii + 361 pp. LCCN QA29.T8 P45 2017. URL

https://link.springer.com/book/
10.1007/978-3-319-53280-6.

Fawcett:1987:RB

[FC87] Carolyn R. Fawcett and Robert S. Cohen. Reason and Being, volume 17 of Boston Studies in the Philosophy of Science. D. Reidel, Dordrecht, The Netherlands; Boston, MA, USA; Lancaster, UK; Tokyo, Japan, 1987. ISBN 90-277-2181-5, 94-009-4590-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). xx + 441 pp. LCCN Q174 .B67 vol. 17. URL

                                              https://link.springer.com/book/
                                              10.1007/978-94-009-4590-6.

Franklin:2016:TKC

[FC16] Allan Franklin and Harry Collins. Two kinds of case study and a new agreement. In Sauer and Scholl [SS16b], chapter 6, pages 95–121. ISBN 3-319-30227-2, 3-319-30229-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q126.8 .P45 2016.

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-3-319-30229-
                                              4_6.

Feldhay:2018:HLI

[FEE18] Rivka Feldhay and Ayelet Even-Ezra. Heaviness, lightness and impetus in the Seventeenth Century: A Jesuit perspective. In Feldhay et al. [FRSV18a], chapter 9, pages 255–284. ISBN 3-319-90343-8 (hardcover), 3-319-90345-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QA802.

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-3-319-90345-
                                              3_9.

Feferman:1993:WF

[Fef93] Solomon Feferman. Working foundations — ’ 91. In Bridging the Gap: Philosophy, Mathematics, and Physics: Lectures on the Foundations of Science [CCG93], chapter 5, pages 99–124.

ISBN 94-010-5101-1, 94-011-2496-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-011-2496-
6_5.

Feher:1999:HS

[Feh99a] István M. Fehér. Hermeneutics and the sciences. In Fehér et al. [FKR99], pages 1–12. ISBN 90-481-5257-7, 94-015-9293-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-015-
9293-2_1.

Feher:1999:CRA

[Feh99b] Márta Fehér. The changing of the role accorded to the lay public in science. In Fehér et al. [FKR99], pages 247–258. ISBN 90-481-5257-7, 94-015-9293-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-015-9293-
2_19.

Feigl:1971:RPI

[Fei71] Herbert Feigl. Research programmes and induction. In Buck and Cohen [BC71], pages 147–150. ISBN 90-277-0309-4, 94-010-3142-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-010-3142-4_9.

Feigl:1974:EB

[Fei74] Herbert Feigl. Empiricism at bay? In Methodological and Historical Essays in the Natural and Social Sciences [CW74b], chapter 1, pages 1–20. ISBN 90-277-0378-7, 94-010-2128-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 14.

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-010-2128-
                                              9_1.

Feigl:1986:EB

[Fei86] Herbert Feigl. Empiricism at bay? In Cohen and Wartofsky [CW85], chapter 8, pages 101–120. ISBN 90-277-1971-3, 94-009-5345-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-009-5345-1_8.

Fekete:1999:EHD

[Fek99] László Fekete. To eliminate human discourse by means of the metaphysical language of things. In Fehér et al. [FKR99], pages 259–264. ISBN 90-481-5257-7, 94-015-9293-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-015-9293-2_20.

Feleppa:1989:CKI

[Fel89] Robert Feleppa. Cultural kinds: Imposition and discovery in anthropology. In The Qualitative-Quantitative Distinction in the Social Sciences [GM89b], chapter 8, pages 119–153. ISBN 90-481-8460-6, 94-017-3444-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 112.

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-017-3444-
                                              8_8.

Feldhay:1992:CRO

[Fel92] Rivka Feldhay. Critical reactions to the occult: A comment. In The Scientific Enterprise: The Bar-Hillel Colloquium: Studies in History, Philosophy, and Sociology of Science, Volume 4 [UM92], chapter 4, pages 93–99. ISBN 94-010-5190-9, 94-011-2688-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .S354 1992.

URL https://link.springer.com/
chapter/10.1007/978-94-011-2688-
5_4.

Feldhay:2015:GLS

[Fel15] Rivka Feldhay. The global and the local in the study of the humanities. In Arabatzis et al. [ARS15], chapter 17, pages 253–267. ISBN 3-319-14552-5 (paperback), 3-319-14553-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q125 .R45 2015.

URL https://link.springer.com/
chapter/10.1007/978-3-319-14553-
2_17.

Feldhay:2018:PMC

[Fel18] Rivka Feldhay. Preclassical mechanics in context: Practical and theoretical knowledge between sovereignty, religion, and science. In Feldhay et al. [FRSV18a], chapter 2, pages 29–53. ISBN 3-319-90343-8 (hardcover), 3-319-90345-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic).

LCCN QA802. URL https://link.
springer.com/chapter/10.1007/
978-3-319-90345-3_2.

Fenton:2012:RCN

[Fen12] Andrew Fenton. Re-conceiving nonhuman animal knowledge through contemporary primate cognitive studies. In Philosophy of Behavioral Biology [PR12b], chapter 6, pages 125–146. ISBN 94-007-1950-7, 94-007-1951-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QH331 .P4679 2012; Q175 .B657 v.282.

URL https://link.springer.com/
chapter/10.1007/978-94-007-1951-
4_6.

Fernandes:1985:KTS

[Fer85a] Sergio L. C. De Fernandes. The Kantian theory of scientific knowledge and its Popperian counterpart. In Foundations of Objective Knowledge: The Relations of Popper’s Theory of Knowledge to that of Kant [Fer85b], chapter 5, pages 70–123. ISBN 90-481-8402-9, 94-015-7704-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

                                              ???? URL https://link.springer.
                                              com/chapter/10.1007/978-94-015-
                                              7704-5_5.

Fernandes:1985:FOK

[Fer85b] Sergio L. De C. Fernandes. Foundations of Objective Knowledge: The Relations of Popper’s Theory of Knowledge to that of Kant, volume 86 of Boston Studies in the Philosophy of Science. D. Reidel, Dordrecht, The Netherlands; Boston, MA, USA; Lancaster, UK; Tokyo, Japan, 1985. ISBN 90-481-8402-9, 94-015-7704-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). xvii + 272 pp. LCCN ????

                                              URL https://link.springer.com/
                                              book/10.1007/978-94-015-7704-5.

Fernandes:1985:I

[Fer85c] Sergio L. De C. Fernandes. Introduction. In Foundations of Objective Knowledge: The Relations of Popper’s Theory of Knowledge to that of Kant [Fer85b], chapter 1, pages 1–6. ISBN 90-481-8402-9, 94-015-7704-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-015-7704-5_1.

Fernandes:1985:OCG

[Fer85d] Sergio L. De C. Fernandes. Our cognitive grasp of an objective world. In Foundations of Objective Knowledge:

The Relations of Popper’s Theory of Knowledge to that of Kant [Fer85b], chapter 6, pages 124–202. ISBN 90-481-8402-9, 94-015-7704-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-015-7704-5_6.

Fernandes:1985:PPR

[Fer85e] Sergio L. De C. Fernandes. Popper’s problems as a revision of those of Kant. In Foundations of Objective Knowledge: The Relations of Popper’s Theory of Knowledge to that of Kant [Fer85b], chapter 3, pages 18–35. ISBN 90-481-8402-9, 94-015-7704-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-015-7704-5_3.

Fernandes:1985:PVK

[Fer85f] Sergio L. De C. Fernandes. Popper’s views of Kant’s problem. In Foundations of Objective Knowledge: The Relations of Popper’s Theory of Knowledge to that of Kant [Fer85b], chapter 2, pages 7–17. ISBN 90-481-8402-9, 94-015-7704-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-015-
7704-5_2.

Fernandes:1985:ROK

[Fer85g] Sergio L. De C. Fernandes. Realism and objective knowledge. In Foundations of Objective Knowledge: The Relations of Popper’s Theory of Knowledge to that of Kant [Fer85b], chapter 7, pages 203–251. ISBN 90-481-8402-9, 94-015-7704-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.

                                              com/chapter/10.1007/978-94-015-
                                              7704-5_7.

Fernandes:1985:RKP

[Fer85h] Sergio L. De C. Fernandes. A reconstruction of Kant’s problem. In Foundations of Objective Knowledge: The Relations of Popper’s Theory of Knowledge to that of Kant [Fer85b], chapter 4, pages 36–69. ISBN 90-481-8402-9, 94-015-7704-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-015-7704-5_4.

Ferrarin:1998:ANM

[Fer98] Alfredo Ferrarin. Aristotelian and Newtonian models in Hegel’s philosophy of nature. In Philosophies of Nature: The Human Dimension: In Celebration of Erazim Kohák [CT98], chapter 6, pages 71–89. ISBN 90-481-4859-6, 94-017-2614-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic).

                                              LCCN BD581. URL https://link.
                                              springer.com/chapter/10.1007/
                                              978-94-017-2614-6_6.

Fetzer:1971:DP

[Fet71] James H. Fetzer. Dispositional probabilities. In Buck and Cohen [BC71], pages 473–482. ISBN 90-277-0309-4, 94-010-3142-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

                                              ???? URL https://link.springer.
                                              com/chapter/10.1007/978-94-010-
                                              3142-4_30.

Fetzer:1974:SE

[Fet74] James H. Fetzer. Statistical explanations. In Schaffner and Cohen [SC74a], pages 337–347. ISBN 90-277-0408-2, 90-277-0409-0, 94-010-2140-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175 .B7312.

URL https://link.springer.com/
chapter/10.1007/978-94-010-2140-
1_23.

Fetzer:1976:LL

[Fet76] James H. Fetzer. The likeness of lawlikeness. In Cohen et al. [CHMV76], pages 377–391. ISBN 90-277-0647-6, 90-277-0648-4 (paperback), 94-010-1449-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-010-
1449-6_19.

Fetzer:1981:ARR

[Fet81a] James H. Fetzer. Acceptance and rejection rules. In Scientific Knowledge: Causation, Explanation, and Corroboration [Fet81h], chapter 9, pages 231–263. ISBN 90-277-1336-7, 94-009-8558-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-009-
8558-2_9.

Fetzer:1981:CC

[Fet81b] James H. Fetzer. Confirmation and corroboration. In Scientific Knowledge: Causation, Explanation, and Corroboration [Fet81h], chapter 8, pages 202–230. ISBN 90-277-1336-7, 94-009-8558-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-009-
8558-2_8.

Fetzer:1981:DCT

[Fet81c] James H. Fetzer. The dispositional construction of theories. In Scientific Knowledge: Causation, Explanation, and Corroboration [Fet81h], chapter 6, pages 137–171. ISBN 90-277-1336-7, 94-009-8558-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-009-
8558-2_6.

Fetzer:1981:JI

[Fet81d] James H. Fetzer. The justification of induction. In Scientific Knowledge: Causation, Explanation, and Corroboration [Fet81h], chapter 7, pages 175–201. ISBN 90-277-1336-7, 94-009-8558-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-009-
8558-2_7.

Fetzer:1981:KCK

[Fet81e] James H. Fetzer. The knowledge context k zt. In Scientific Knowledge: Causation, Explanation, and Corroboration [Fet81h], chapter 1, pages 3–22. ISBN 90-277-1336-7, 94-009-8558-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-009-8558-
2_1.

Fetzer:1981:LF

[Fet81f]  James H. Fetzer.  The language

framework: L OR l ? In Scientific Knowledge: Causation, Explanation, and Corroboration [Fet81h], chapter 2, pages 23–45. ISBN 90-277-1336-7, 94-009-8558-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-009-
8558-2_2.

Fetzer:1981:RF

[Fet81g] James H. Fetzer. Rationality and fallibility. In Scientific Knowledge:

Causation, Explanation, and Corroboration [Fet81h], chapter 10, pages 264–296. ISBN 90-277-1336-7, 94-009-8558-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-009-
8558-2_10.

Fetzer:1981:SKC

[Fet81h] James H. Fetzer. Scientific Knowledge:

Causation, Explanation, and Corroboration, volume 69 of Boston Studies in the Philosophy of Science. D. Reidel, Dordrecht, The Netherlands; Boston, MA, USA; Lancaster, UK; Tokyo, Japan, 1981. ISBN 90-277-1336-7, 94-009-8558-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). xvi + 323 pp. LCCN ???? URL

https://link.springer.com/book/
10.1007/978-94-009-8558-2.

Fetzer:1981:SCT

[Fet81i] James H. Fetzer. A single case theory of causal explanation. In Scientific Knowledge: Causation, Explanation, and Corroboration [Fet81h], chapter 5, pages 104–136. ISBN 90-277-1336-7, 94-009-8558-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-009-
8558-2_5.

Fetzer:1981:SES

[Fet81j] James H. Fetzer. Statistical explanation and statistical relevance. In Scientific Knowledge: Causation, Explanation, and Corroboration [Fet81h], chapter 4, pages 77–103. ISBN 90-277-1336-7, 94-009-8558-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-009-
8558-2_4.

Fetzer:1981:SSO

[Fet81k] James H. Fetzer. Syntax, semantics, and ontology. In Scientific Knowledge: Causation, Explanation, and Corroboration [Fet81h], chapter 3, pages 46–73. ISBN 90-277-1336-7, 94-009-8558-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-009-8558-
                                              2_3.

Fetzer:2010:PPS

[Fet10a] J. H. Fetzer. The Place of Probability in Science: In Honor of Ellery Eells (1953–2006), volume 284 of Boston Studies in the Philosophy of Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2010. ISBN 90-481-3614-8 (hardcover), 90-481-3615-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). xliii + 368 pp. LCCN Q175 .P563 2010. URL

                                              https://link.springer.com/book/
                                              10.1007/978-90-481-3615-5.

Fetzer:2010:EOP

[Fet10b] James H. Fetzer. Is evolution an optimizing process? In The Place of Probability in Science: In Honor of Ellery Eells (1953–2006) [Fet10a], chapter 7, pages 163–177. ISBN 90-481-3614-8 (hardcover), 90-481-3615-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175 .P563

                                              2010. URL https://link.springer.
                                              com/chapter/10.1007/978-90-481-
                                              3615-5_7.

Fetzer:2010:PM

[Fet10c] James H. Fetzer. Probabilistic metaphysics. In The Place of Probability in Science: In Honor of Ellery Eells (1953–2006) [Fet10a], chapter 4, pages 81–98. ISBN 90-481-3614-8 (hardcover), 90-481-3615-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175 .P563 2010. URL https:

//link.springer.com/chapter/10.
1007/978-90-481-3615-5_4.

Fetzer:2010:POD

[Fet10d] James H. Fetzer. Probability and objectivity in deterministic and indeterministic situations. In The Place of Probability in Science: In Honor of Ellery Eells (1953–2006) [Fet10a], chapter 10, pages 229–244. ISBN 90-481-3614-8 (hardcover), 90-481-3615-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175 .P563

2010. URL https://link.springer.
com/chapter/10.1007/978-90-481-
3615-5_10.

Fetzer:2010:PF

[Fet10e] James H. Fetzer. Propensities and frequencies. In The Place of Probability in Science: In Honor of Ellery Eells (1953–2006) [Fet10a], chapter 14, pages 323–351. ISBN 90-481-3614-8 (hardcover), 90-481-3615-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175 .P563 2010.

URL https://link.springer.com/
chapter/10.1007/978-90-481-3615-
5_14.

Feuer:1988:NPE

[Feu88] Lewis S. Feuer. A narrative of personal events and ideas. In Philosophy, History and Social Action: Essays in Honor of Lewis Feuer with an autobiographic essay by Lewis Feuer [HOO88c], chapter 1, pages 1–85. ISBN 94-010-7793-2, 94-009-2873-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-009-2873-2_1.

Feyerabend:1967:ISA

[Fey67] Paul K. Feyerabend. On the improvement of the sciences and the arts, and the possible identity of the two. In Cohen and Wartofsky [CW67], pages 387–415. ISBN 90-277-0013-3, 94-010-3508-3. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175

                                              .B731 v.3.  URL https://link.
                                              springer.com/chapter/10.1007/
                                              978-94-010-3508-8_22.

Feyerabend:1976:CSR

[Fey76] P. K. Feyerabend. On the critique of scientific reason. In Essays in Memory of Imre Lakatos [CFW76], chapter 11, pages 109–143. ISBN 90-277-0655-7, 94-010-1451-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

                                              ???? URL https://link.springer.
                                              com/chapter/10.1007/978-94-010-
                                              1451-9_11.

Feyerabend:1978:GSP

[Fey78a] Paul Feyerabend. The Gong Show — Popperian style. In Progress and Rationality in Science [RA78b], chapter 17, pages 387–392. ISBN 90-277-0922-X, 94-009-9866-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic).

                                              LCCN ????  URL https://link.
                                              springer.com/chapter/10.1007/
                                              978-94-009-9866-7_17.

Feyerabend:1978:DAC

[Fey78b] Paul Feyerabend. In defence of Aristotle: Comments on the condition of content increase. In Progress and Rationality in Science [RA78b], chapter 7, pages 143–180. ISBN 90-277-0922-X, 94-009-9866-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic).

                                              LCCN ????  URL https://link.
                                              springer.com/chapter/10.1007/
                                              978-94-009-9866-7_7.

Feyerabend:1979:DM

[Fey79] Paul Feyerabend. Dialogue on method.

In The Structure and Development of Science [RA79], chapter 4, pages 63–131. ISBN 90-277-0995-5, 94-009-9459-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .S778 1979. URL https://link.

springer.com/chapter/10.1007/
978-94-009-9459-1_4.

Feyerabend:1984:PS

[Fey84a] Paul K. Feyerabend. Philosophy of science 2001. In Methodology, Metaphysics and the History of Science: In Memory of Benjamin Nelson [CW84b], chapter 4, pages 137–147. ISBN 94-009-6333-5, 94-009-6331-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-009-6331-3_4.

Feyerabend:1984:XFC

[Fey84b] Paul K. Feyerabend. Xenophanes: A forerunner of critical rationalism? In Rationality in Science and Politics [And84d], chapter 6, pages 95–109. ISBN 90-277-1953-5, 94-009-6254-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-009-6254-
5_6.

Feyerabend:1989:A

[Fey89] Paul Feyerabend. Antilogikē. In Freedom and Rationality: Essays in Honor of John Watkins From his Colleagues and Friends [DJ89], chapter 11, pages 185–189. ISBN 94-010-7571-9, 94-009-2380-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.

                                              com/chapter/10.1007/978-94-009-
                                              2380-5_11.

Feyerabend:1991:CUC

[Fey91] Paul K. Feyerabend. Concluding unphilosophical conversation. In Beyond Reason: Essays on the Philosophy of Paul Feyerabend [Mun91a], chapter 25, pages 487–527. ISBN 94-010-5406-1, 94-011-3188-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

                                              ???? URL https://link.springer.
                                              com/chapter/10.1007/978-94-011-
                                              3188-9_25.

Feyerabend:1994:R

[Fey94] Paul Feyerabend. Realism. In Artifacts, Representations and Social Practice: Essays for Marx Wartofsky [GC94], chapter 13, pages 205–222. ISBN 94-010-4390-6, 94-011-0902-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN B73 .A77 1994.

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-011-0902-
                                              4_13.

Feyerabend:1995:APN

[Fey95a] Paul Feyerabend. Art as a product of nature as a work of art. In Science, Mind and Art: Essays on science and the humanistic understanding in art, epistemology, religion and ethics In honor of Robert S. Cohen [GSW95b], chapter 1, pages 1–18. ISBN 0-7923-2990-2 (hardcover), 94-011-0469-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 165 BD418.3. URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-011-0469-2_1.

Feyerabend:1995:UTM

[Fey95b] Paul K. Feyerabend. Universals as tyrants and as mediators. In Critical Rationalism, Metaphysics and Science: Essays for Joseph Agassi. Volume I [JL95a], chapter 1, pages 3–14. ISBN 94-010-4211-X, 94-011-0471-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-011-0471-5_1.

Faye:1994:NBC

[FF94] Jan Faye and Henry J. Folse, editors. Niels Bohr and Contemporary Philosophy, volume 153 of Boston Studies in the Philosophy of Science. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 1994. ISBN 0-7923-2378-5 (hardcover), 94-015-8106-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). xxvii + 377 pp. LCCN Q174 .B67 vol. 153 1994;

QC16 1994. URL http://www.loc.
gov/catdir/enhancements/fy0823/
93024825-d.html;   http://www.
loc.gov/catdir/enhancements/
fy0823/93024825-t.html;   https:
//link.springer.com/book/10.
1007/978-94-015-8106-6.

Falk:2007:WSS

[FF07] Raphael Falk and Ruma Falk. Why should scientists become historians? In Positioning the history of science [RG07], chapter 8, pages 43–48. ISBN 1-4020-5419-X (hardcover), 1-4020-5420-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q125 .P77 2007. URL https://link.

springer.com/chapter/10.1007/1-
4020-5420-3_8.

Fox:1992:LSL

[FG92] Robert Fox and Anna Guagnini. Life in the slow lane: Research and electrical engineering in Britain, France, and Italy, ca. 1900. In Technological Development and Science in the Industrial Age: New Perspectives on the Science–Technology Relationship [KB92b], chapter 6, pages 133–153. ISBN 90-481-4186-9, 94-015-8010-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-015-8010-
6_6.

Fitelson:2010:HBC

[FH10a] Branden Fitelson and James Hawthorne. How Bayesian confirmation theory handles the paradox of the ravens. In The Place of Probability in Science: In Honor of Ellery Eells (1953–2006) [Fet10a], chapter 11, pages 247–275. ISBN 90-481-3614-8 (hardcover), 90-481-3615-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175 .P563 2010.

URL https://link.springer.com/
chapter/10.1007/978-90-481-3615-
5_11.

Frigg:2010:BMC

[FH10b] Roman Frigg and Matthew Hunter, editors. Beyond Mimesis and Convention: Representation in Art and Science, volume 262 of Boston Studies in the Philosophy of Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2010. ISBN 90-481-3850-7 (hardcover), 90-481-3851-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). xxx + 265 pp. LCCN BH301.M47 B49 2010; Q175 .B52 2010. URL

https://link.springer.com/book/
10.1007/978-90-481-3851-7.

Fieschi:2014:ALF

[Fie14] Roberto Fieschi. The Andrea Levialdi Fellowship. In The History of Physics in Cuba [BRW14a], chapter 23, pages 361–363. ISBN 94-017-8040-4, 94-017-8041-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC5.53; QC6.9; QC9 C9H57 2014; Q124.6-127.2. URL https://link.

springer.com/chapter/10.1007/
978-94-017-8041-4_23.

Filliozat:2005:ASM

[Fil05] Pierre-Sylvain Filliozat. Ancient Sanskrit mathematics: An oral tradition and a written literature. In History of Science, History of Text [Che05a], chapter 7, pages 137–157. ISBN 1-281-33589-4, 1-4020-2320-0 (hardcover), 1-4020-2321-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 238 Q125.

URL https://link.springer.com/
chapter/10.1007/1-4020-2321-9_7.

Fillion:2017:VCS

[Fil17] Nicolas Fillion. The vindication of computer simulations. In Mathematics as a Tool: Tracing New Roles of Mathematics in the Sciences [LC17b], chapter 8, pages 137–155. ISBN 3-319-54468-3 (hardcover), 3-319-54469-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN GE45.M37; GE45.M38; QA8.4. URL https:/

/link.springer.com/chapter/10.
1007/978-3-319-54469-4_8.

Finkelstein:1969:MSL

[Fin69]  David Finkelstein.  Matter, space

and logic. In Cohen and Wartofsky [CW69a], pages 199–215. ISBN 94-010-3383-8, 94-010-3263-7, 94-010-3381-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-010-3381-
7_4.

Fine:1976:YEO

[Fin76] Arthur Fine. The young Einstein and the old Einstein. In Essays in Memory of Imre Lakatos [CFW76], chapter 12, pages 145–159. ISBN 90-277-0655-7, 94-010-1451-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-010-1451-9_12.

Finocchiaro:1980:AIE

[Fin80a] Maurice A. Finocchiaro. Active involvement: The evaluation of reasoning. In Galileo and the Art of Reasoning: Rhetorical Foundations of Logic and Scientific Method [Fin80h], chapter 15, pages 332–342. ISBN 90-277-1095-3, 94-009-9017-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

                                              ???? URL https://link.springer.
                                              com/chapter/10.1007/978-94-009-
                                              9017-3_15.

Finocchiaro:1980:CJD

[Fin80b] Maurice A. Finocchiaro. Concreteness and judgment: The dialectical nature of Galileo’s methodology. In Galileo and the Art of Reasoning: Rhetorical Foundations of Logic and Scientific Method [Fin80h], chapter 6, pages 145–166. ISBN 90-277-1095-3, 94-009-9017-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

                                              ???? URL https://link.springer.
                                              com/chapter/10.1007/978-94-009-
                                              9017-3_6.

Finocchiaro:1980:CCI

[Fin80c] Maurice A. Finocchiaro. Criticism, complexity, and invalidities: Theoretical considerations. In Galileo and the Art of Reasoning: Rhetorical Foundations of Logic and Scientific Method [Fin80h], chapter 17, pages 413–431.

ISBN 90-277-1095-3, 94-009-9017-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-009-9017-
3_17.

Finocchiaro:1980:EAP

[Fin80d] Maurice A. Finocchiaro. Emotion, aesthetics, and persuasion: The rhetorical force of Galileo’s argument. In Galileo and the Art of Reasoning: Rhetorical Foundations of Logic and Scientific Method [Fin80h], chapter 3, pages 46– 66. ISBN 90-277-1095-3, 94-009-9017-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-009-9017-
3_3.

Finocchiaro:1980:ELL

[Fin80e] Maurice A. Finocchiaro. The erudition of logic and the logic of erudition: Critique of Galileo scholarship. In Galileo and the Art of Reasoning: Rhetorical Foundations of Logic and Scientific Method [Fin80h], chapter 10, pages 224–255. ISBN 90-277-1095-3, 94-009-9017-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-009-
9017-3_10.

Finocchiaro:1980:FRL

[Fin80f] Maurice A. Finocchiaro. Fact and reasoning: The logical structure of Galileo’s argument. In Galileo and the Art of Reasoning: Rhetorical Foundations of Logic and Scientific Method [Fin80h], chapter 2, pages 27–45. ISBN 90-277-1095-3, 94-009-9017-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-009-9017-3_2.

Finocchiaro:1980:FVR

[Fin80g] Maurice A. Finocchiaro. Faith versus reason: The rhetorical form and content of Galileo’s dialogue. In Galileo and the Art of Reasoning: Rhetorical Foundations of Logic and Scientific Method [Fin80h], chapter 1, pages 3–26. ISBN 90-277-1095-3, 94-009-9017-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-009-9017-3_1.

Finocchiaro:1980:GAR

[Fin80h] Maurice A. Finocchiaro. Galileo and the Art of Reasoning: Rhetorical Foundations of Logic and Scientific Method, volume 61 of Boston Studies in the Philosophy of Science. D. Reidel, Dordrecht, The Netherlands; Boston, MA, USA; Lancaster, UK; Tokyo, Japan, 1980. ISBN 90-277-1095-3, 94-009-9017-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). xx + 481 pp. LCCN ???? URL

                                              https://link.springer.com/book/
                                              10.1007/978-94-009-9017-3.

Finocchiaro:1980:GLM

[Fin80i] Maurice A. Finocchiaro. Galileo as a logician: A model and a data basis. In Galileo and the Art of Reasoning: Rhetorical Foundations of Logic and Scientific Method [Fin80h], chapter 16, pages 343–412. ISBN 90-277-1095-3, 94-009-9017-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

                                              ???? URL https://link.springer.
                                              com/chapter/10.1007/978-94-009-
                                              9017-3_16.

Finocchiaro:1980:HSS

[Fin80j] Maurice A. Finocchiaro. The history of science and the science of history: Critique of apriorism. In Galileo and the Art of Reasoning: Rhetorical Foundations of Logic and Scientific Method [Fin80h], chapter 9, pages 202–223. ISBN 90-277-1095-3, 94-009-9017-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-009-9017-
3_9.

Finocchiaro:1980:LSS

[Fin80k] Maurice A. Finocchiaro. The logic of science and the science of logic: Toward a science of reasoning. In Galileo and the Art of Reasoning: Rhetorical Foundations of Logic and Scientific Method [Fin80h], chapter 13, pages 293–307. ISBN 90-277-1095-3, 94-009-9017-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-009-9017-
3_13.

Finocchiaro:1980:PRL

[Fin80l] Maurice A. Finocchiaro. The primacy of reasoning: The logical character of Galileo’s methodology. In Galileo and the Art of Reasoning: Rhetorical Foundations of Logic and Scientific Method [Fin80h], chapter 7, pages 167–179. ISBN 90-277-1095-3, 94-009-9017-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-009-9017-
3_7.

Finocchiaro:1980:PSU

[Fin80m] Maurice A. Finocchiaro. Propositional structure: The understanding of reasoning. In Galileo and the Art of Reasoning: Rhetorical Foundations of Logic and Scientific Method [Fin80h], chapter 14, pages 311–331. ISBN 90-277-1095-3, 94-009-9017-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-009-9017-3_14.

Finocchiaro:1980:PLL

[Fin80n] Maurice A. Finocchiaro. The psychology of logic and the logic of psychology: Critique of the psychology of reasoning. In Galileo and the Art of Reasoning: Rhetorical Foundations of Logic and Scientific Method [Fin80h], chapter 11, pages 256–272. ISBN 90-277-1095-3, 94-009-9017-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-009-
9017-3_11.

Finocchiaro:1980:RSS

[Fin80o] Maurice A. Finocchiaro. The rationality of science and the science of rationality: Critique of subjectivism. In Galileo and the Art of Reasoning: Rhetorical Foundations of Logic and Scientific Method [Fin80h], chapter 8, pages 180–201. ISBN 90-277-1095-3, 94-009-9017-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-009-
9017-3_8.

Finocchiaro:1980:RLL

[Fin80p] Maurice A. Finocchiaro. The rhetoric of logic and the logic of rhetoric: Critique of the new rhetoric. In Galileo and the Art of Reasoning: Rhetorical Foundations of Logic and Scientific Method [Fin80h], chapter 12, pages 273–292. ISBN 90-277-1095-3, 94-009-9017-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-009-9017-
3_12.

Finocchiaro:1980:SDG

[Fin80q] Maurice A. Finocchiaro. Scientific discoveries as growth of understanding: The case of Newton’s gravitation. In Scientific Discovery, Logic, and Rationality [Nic80f], chapter 10, pages 235–255. ISBN 90-277-1070-8, 94-009-8986-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-009-
8986-3_10.

Finocchiaro:1980:TPM

[Fin80r] Maurice A. Finocchiaro. Theory and practice: The methodological content of Galileo’s science. In Galileo and the Art of Reasoning: Rhetorical Foundations of Logic and Scientific Method [Fin80h], chapter 5, pages 103–141. ISBN 90-277-1095-3, 94-009-9017-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-009-9017-
3_5.

Finocchiaro:1980:TMS

[Fin80s] Maurice A. Finocchiaro. Truth and method: The scientific content of Galileo’s dialogue. In Galileo and the Art of Reasoning: Rhetorical Foundations of Logic and Scientific Method [Fin80h], chapter 4, pages 67–102. ISBN 90-277-1095-3, 94-009-9017-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-009-9017-
3_4.

Findlay:1984:HTB

[Fin84] John N. Findlay. The Hegelian treatment of biology and life. In Hegel and the Sciences [CW84a], chapter 6, pages 87–100. ISBN 94-009-6235-5, 94-009-6233-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

                                              ???? URL https://link.springer.
                                              com/chapter/10.1007/978-94-009-
                                              6233-0_6.

Fink:1987:ORP

[Fin87] Karl J. Fink. Ontogeny recapitulates phylogeny: A classic formula of organicism. In Burwick [Bur87b], chapter 4, pages 87–112. ISBN 90-277-2541-1, 94-009-3917-5. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 105. US$64.00(U.S.).

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-009-3917-
                                              2_4.

Finlay:1991:SPG

[Fin91] Mark R. Finlay. Science and practice in German agriculture: Justus Von Liebig, Hermann Von Liebig, and the agricultural experiment stations. In Woodward and Cohen [WC91], pages 309–320. ISBN 94-010-5395-2, 94-011-3164-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

                                              ???? URL https://link.springer.
                                              com/chapter/10.1007/978-94-011-
                                              3164-3_28.

Fine:1993:MQS

[Fin93]  Arthur Fine.  Measurement and

quantum silence. In Correspondence, Invariance and Heuristics: Essays in Honour of Heinz Post [FK93], chapter 14, pages 279–294. ISBN 90-481-4229-6, 94-017-1185-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-017-1185-2_14.

Finocchiaro:1995:CRJ

[Fin95] Maurice A. Finocchiaro. Criticism, reasoning, and judgment in science. In Critical Rationalism, Metaphysics and Science: Essays for Joseph Agassi. Volume I [JL95a], chapter 11, pages 169–191. ISBN 94-010-4211-X, 94-011-0471-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-011-
0471-5_11.

Fine:1996:IBM

[Fin96] Arthur Fine. On the interpretation of Bohmian mechanics. In Bohmian Mechanics and Quantum Theory: an Appraisal [CFG96], chapter 16, pages 231–250. ISBN 90-481-4698-4, 94-015-8715-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC173.96-174.52. URL https://link.

springer.com/chapter/10.1007/
978-94-015-8715-0_16.

Finkelstein:2003:EP

[Fin03] David Ritz Finkelstein. Elementary processes. In Renn et al. [RDS03], chapter 24, pages 479–497. ISBN 1-4020-1284-5, 1-4020-1285-3 (paperback). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 234. URL https://link.

springer.com/chapter/10.1007/
978-94-010-0111-3_24.

Finocchiaro:2009:CC

[Fin09a] Maurice A. Finocchiaro. The Copernican controversy. In Defending Copernicus and Galileo: Critical Reasoning in the Two Affairs [Fin09b], chapter 2, pages 21–36. ISBN 90-481-3200-2 (hardcover), 90-481-3201-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 2009; Q175 .B73 2009. URL https:

//link.springer.com/chapter/10.
1007/978-90-481-3201-0_2.

Finocchiaro:2009:DCG

[Fin09b] Maurice A. Finocchiaro. Defending Copernicus and Galileo: Critical Reasoning in the Two Affairs, volume 280 of Boston Studies in the Philosophy of Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2009. ISBN 90-481-3200-2 (hardcover), 90-481-3201-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). xliii + 339 pp. LCCN Q174 .B67 2009; Q175 .B73 2009. URL

https://link.springer.com/book/
10.1007/978-90-481-3201-0.

Finocchiaro:2009:GCB

[Fin09c] Maurice A. Finocchiaro. Galilean critiques of the biblical objection. In Defending Copernicus and Galileo: Critical Reasoning in the Two Affairs [Fin09b], chapter 4, pages 65–96. ISBN 90-481-3200-2 (hardcover), 90-481-3201-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 2009; Q175 .B73 2009.

URL https://link.springer.com/
chapter/10.1007/978-90-481-3201-
0_4.

Finocchiaro:2009:GRC

[Fin09d] Maurice A. Finocchiaro. Galilean rationality in the Copernican Revolution. In Defending Copernicus and Galileo: Critical Reasoning in the Two Affairs [Fin09b], chapter 6, pages 121–134. ISBN 90-481-3200-2 (hardcover), 90-481-3201-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 2009; Q175 .B73 2009.

URL https://link.springer.com/
chapter/10.1007/978-90-481-3201-
0_6.

Finocchiaro:2009:GA

[Fin09e] Maurice A. Finocchiaro. The Galileo Affair, 1633–1992. In Defending Copernicus and Galileo: Critical Reasoning in the Two Affairs [Fin09b], chapter 8, pages 155–227. ISBN 90-481-3200-2 (hardcover), 90-481-3201-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 2009; Q175 .B73 2009. URL https://link.

springer.com/chapter/10.1007/
978-90-481-3201-0_8.

Finocchiaro:2009:GBE

[Fin09f] Maurice A. Finocchiaro. Galileo as a bad epistemologist? In Defending Copernicus and Galileo: Critical Reasoning in the Two Affairs [Fin09b], chapter 11, pages 277–290. ISBN 90-481-3200-2 (hardcover), 90-481-3201-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 2009; Q175 .B73 2009. URL https:

//link.springer.com/chapter/10.
1007/978-90-481-3201-0_11.

Finocchiaro:2009:GBT

[Fin09g] Maurice A. Finocchiaro. Galileo as a bad theologian? In Defending Copernicus and Galileo: Critical Reasoning in the Two Affairs [Fin09b], chapter 10, pages 251–276. ISBN 90-481-3200-2 (hardcover), 90-481-3201-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 2009; Q175 .B73 2009. URL https://link.

springer.com/chapter/10.1007/
978-90-481-3201-0_10.

Finocchiaro:2009:GSS

[Fin09h] Maurice A. Finocchiaro. Galileo as a symbol of science versus religion? In Defending Copernicus and Galileo: Critical Reasoning in the Two Affairs [Fin09b], chapter 12, pages 291–314. ISBN 90-481-3200-2 (hardcover), 90-481-3201-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 2009; Q175 .B73 2009.

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-90-481-3201-
                                              0_12.

Finocchiaro:2009:GMP

[Fin09i] Maurice A. Finocchiaro. Galileo on the mathematical physics of terrestrial extrusion. In Defending Copernicus and Galileo: Critical Reasoning in the Two Affairs [Fin09b], chapter 5, pages 97–120. ISBN 90-481-3200-2 (hardcover), 90-481-3201-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 2009; Q175 .B73 2009. URL https://link.

                                              springer.com/chapter/10.1007/
                                              978-90-481-3201-0_5.

Finocchiaro:2009:GRW

[Fin09j] Maurice A. Finocchiaro. Galileo right for the wrong reasons? In Defending Copernicus and Galileo: Critical Reasoning in the Two Affairs [Fin09b], chapter 9, pages 229–249. ISBN 90-481-3200-2 (hardcover), 90-481-3201-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 2009; Q175 .B73 2009. URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-90-481-3201-0_9.

Finocchiaro:2009:GST

[Fin09k] Maurice A. Finocchiaro. Galileo’s stances toward Copernican astronomy. In Defending Copernicus and Galileo:

Critical Reasoning in the Two Affairs [Fin09b], chapter 3, pages 37–64. ISBN 90-481-3200-2 (hardcover), 90-481-3201-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 2009; Q175 .B73 2009.

URL https://link.springer.com/
chapter/10.1007/978-90-481-3201-
0_3.

Finocchiaro:2009:GWV

[Fin09l] Maurice A. Finocchiaro. The geostatic world view. In Defending Copernicus and Galileo: Critical Reasoning in the Two Affairs [Fin09b], chapter 1, pages 3–19. ISBN 90-481-3200-2 (hardcover), 90-481-3201-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 2009; Q175 .B73

2009. URL https://link.springer.
com/chapter/10.1007/978-90-481-
3201-0_1.

Finocchiaro:2009:TG

[Fin09m] Maurice A. Finocchiaro. The trial of Galileo, 1613–1633. In Defending Copernicus and Galileo: Critical Reasoning in the Two Affairs [Fin09b], chapter 7, pages 137–153. ISBN 90-481-3200-2 (hardcover), 90-481-3201-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 2009; Q175 .B73 2009. URL https:

//link.springer.com/chapter/10.
1007/978-90-481-3201-0_7.

Finstad:2014:MMF

[Fin14] Terje Finstad. The means of modernization, freezing technologies and the cultural politics of everyday life, Norway 1940–1965. In History of Artificial Cold, Scientific, Technological and Cultural Issues [Gav14b], chapter 12, pages 231–249. ISBN 94-007-7198-3 (hardcover), 94-007-7199-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC278 H57 2014; TP494 .H577 2014. URL https:/

                                              /link.springer.com/chapter/10.
                                              1007/978-94-007-7199-4_12.

Finocchiaro:2017:CLP

[Fin17] Maurice A. Finocchiaro. Charity, logic, and para-clericalism: Rereading Agassi on Koestler on Galileo. In Encouraging Openness: Essays for Joseph Agassi on the Occasion of His 90th Birthday [BAG17], chapter 15, pages 191–206. ISBN 3-319-57668-2 (hardcover), 3-319-57669-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN B833.E53 2017. URL https://link.

                                              springer.com/chapter/10.1007/
                                              978-3-319-57669-5_15.

Fisk:1971:CN

[Fis71] Milton Fisk. Capacities and natures. In Buck and Cohen [BC71], pages 49–62. ISBN 90-277-0309-4, 94-010-3142-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-010-3142-
                                              4_4.

Fisch:1995:EQC

[Fis95] Menachem Fisch. Eccleciastes (Qohelet) in context — a study of wisdom as constructive skepticism. In Critical Rationalism, the Social Sciences and the Humanities: Essays for Joseph Agassi. Volume II [JL95b], chapter 10, pages 167–187. ISBN 0-7923-2961-9, 90-481-4430-2, 94-017-0441-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN B29 .C75 1995.

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-017-0441-
                                              0_10.

Fisette:1996:DNE

[Fis96] Denis Fisette. Davidson on norms and the explanation of behavior. In Québec Studies in the Philosophy of Science: Part II: Biology, Psychology, Cognitive Science and Economics Essays in Honor of Hugues Leblanc [MC96], chapter 9, pages 139–157. ISBN 94-010-6537-3, 94-009-0113-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-009-0113-1_9.

Fischer:1999:NVC

[Fis99] Kurt Rudolf Fischer. Nietzsche and the Vienna Circle. In Nietzsche, Theories of Knowledge, and Critical Theory: Nietzsche and the Sciences I [Bab99b], chapter 9, pages 119–128. ISBN 90-481-5233-X, 94-017-2430-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-017-2430-2_9.

Fisher:2015:OSI

[Fis15] Grant Fisher. Orbital symmetry, idealization, and the Kairetic account of scientific explanation. In Philosophy of Chemistry: Growth of a New Discipline [SM15b], chapter 13, pages 201–218. ISBN 94-017-9363-8 (hardcover), 94-017-9364-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QD6 .P455 2015; QD6 .S384 2014.

URL https://link.springer.com/
chapter/10.1007/978-94-017-9364-
3_13.

Fitzgerald:1971:TBC

[Fit71] Paul Fitzgerald. Tachyons, backwards causation, and freedom. In Buck and Cohen [BC71], pages 415–436. ISBN 90-277-0309-4, 94-010-3142-8 (e-book).

ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-010-3142-4_26.

Fitzgerald:1974:NUT

[Fit74] Paul Fitzgerald. Nowness and the understanding of time. In Schaffner and Cohen [SC74a], pages 259–281. ISBN 90-277-0408-2, 90-277-0409-0, 94-010-2140-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175 .B7312. URL https:/

                                              /link.springer.com/chapter/10.
                                              1007/978-94-010-2140-1_17.

Fitzgerald:1976:MSM

[Fit76] Paul Fitzgerald. Meaning in science and mathematics. In Cohen et al. [CHMV76], pages 235–269. ISBN 90-277-0647-6, 90-277-0648-4 (paperback), 94-010-1449-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic).

                                              LCCN ????  URL https://link.
                                              springer.com/chapter/10.1007/
                                              978-94-010-1449-6_11.

Fjelland:2002:CIQ

[Fje02] Ragnar Fjelland. The “Copenhagen interpretation” of quantum mechanics and phenomenology. In Hermeneutic Philosophy of Science, Van Gogh’s Eyes, and God: Essays in Honor of Patrick A. Heelan, S.J. [Bab02c], chapter 5, pages 53–65. ISBN 90-481-5926-1, 94-017-1767-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 225.

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-017-1767-
                                              0_5.

French:1993:CIH

[FK93] Steven French and Harmke Kamminga. Correspondence, Invariance and Heuristics: Essays in Honour of Heinz Post, volume 148 of Boston Studies in the Philosophy of Science. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 1993. ISBN 90-481-4229-6, 94-017-1185-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). xxiii + 363 pp. LCCN ????

URL https://link.springer.com/
book/10.1007/978-94-017-1185-2.

Feher:1999:HSP

[FKR99] Márta Fehér, Olga Kiss, and László Ropolyi, editors. Hermeneutics and Science: Proceedings of the First Conference of the International Society for Hermeneutics and Science, volume 206 of Boston Studies in the Philosophy of Science. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 1999. ISBN 90-481-5257-7, 94-015-9293-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
book/10.1007/978-94-015-9293-2.

Fagot-Largeault:2009:LPS

[FL09] Anne Fagot-Largeault. The legend of philosophy’s striptease (trends in philosophy of science). In Brenner and Gayon [BG09], chapter 1, pages 25–49. ISBN 1-4020-9367-5, 1-4020-9368-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174

.F746 2009. URL https://link.
springer.com/chapter/10.1007/
978-1-4020-9368-5_1.

French:2011:DOS

[FL11] Steven French and James Ladyman.

In defence of ontic structural realism. In Bokulich and Bokulich [BB11], chapter 2, pages 25–42. ISBN 90-481-9596-9, 90-481-9597-7 (e-book).

ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175.32.S78 S35

                                              2011. URL https://link.springer.
                                              com/chapter/10.1007/978-90-481-
                                              9597-8_2.

Fleck:1986:CS

[Fle86a] Ludwik Fleck. Crisis in science. In Cognition and Fact: Materials on Ludwik Fleck [CS86], chapter 8, pages 153–158. ISBN 94-010-8504-8, 94-009-4498-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-009-4498-
                                              5_8.

Fleck:1986:CR

[Fle86b] Ludwik Fleck. On the crisis of ‘reality’ [1929]. In Cognition and Fact: Materials on Ludwik Fleck [CS86], chapter 3, pages 47–57. ISBN 94-010-8504-8, 94-009-4498-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

                                              ???? URL https://link.springer.
                                              com/chapter/10.1007/978-94-009-
                                              4498-5_3.

Fleck:1986:PE

[Fle86c] Ludwik Fleck. The problem of epistemology [1936]. In Cognition and Fact: Materials on Ludwik Fleck [CS86], chapter 5, pages 79–112. ISBN 94-010-8504-8, 94-009-4498-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-009-4498-5_5.

Fleck:1986:PSS

[Fle86d] Ludwik Fleck. Problems of the science of science [1946]. In Cognition and Fact: Materials on Ludwik Fleck [CS86], chapter 6, pages 113–127. ISBN 94-010-8504-8, 94-009-4498-5 (e-book).

ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-009-4498-5_6.

Fleck:1986:SOP

[Fle86e] Ludwik Fleck. Scientific observation and perception in general [1935]. In Cognition and Fact: Materials on Ludwik Fleck [CS86], chapter 4, pages 59–78. ISBN 94-010-8504-8, 94-009-4498-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-009-
4498-5_4.

Fleck:1986:SSF

[Fle86f] Ludwik Fleck. Some specific features of the medical way of thinking [1927]. In Cognition and Fact: Materials on Ludwik Fleck [CS86], chapter 2, pages 39–46. ISBN 94-010-8504-8, 94-009-4498-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-009-
4498-5_2.

Fleck:1986:LSK

[Fle86g] Ludwik Fleck. To look, to see, to know [1947]. In Cognition and Fact: Materials on Ludwik Fleck [CS86], chapter 7, pages 129–151. ISBN 94-010-8504-8, 94-009-4498-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-009-4498-5_7.

Fleck:1991:ONC

[Fle91a] Lola Fleck. Otto Neurath’s contribution to the theory of the social sciences. In Uebel [Ueb91b], chapter 16, pages 203–208. ISBN 0-7923-1276-7 (hardcover), 94-011-3182-1. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 133. URL https:/

                                              /link.springer.com/chapter/10.
                                              1007/978-94-011-3182-7_16.

Flesher:1991:MDO

[Fle91b] Mary Mosher Flesher. Move over Darwin: The ontogenetic sources of William Preyer’s developmental psychology. In Woodward and Cohen [WC91], pages 117–128. ISBN 94-010-5395-2, 94-011-3164-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-011-3164-3_11.

Flonta:2015:WOD

[Flo15] Mircea Flonta. What ought to be done and what is forbidden: Rules of scientific research as categorical or hypothetical imperatives. In Romanian Studies in Philosophy of Science [PST15], chapter 2, pages 25–42. ISBN 3-319-16654-9 (hardcover), 3-319-16655-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

                                              ???? URL https://link.springer.
                                              com/chapter/10.1007/978-3-319-
                                              16655-1_2.

Floyd:2017:I

[Flo17a] Juliet Floyd. Introduction. In Floyd and Bokulich [FB17], chapter 1, pages 1–35. ISBN 3-319-53278-2, 3-319-53280-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QA29.T8 P45 2017. URL https:/

                                              /link.springer.com/chapter/10.
                                              1007/978-3-319-53280-6_1.

Floyd:2017:TCS

[Flo17b] Juliet Floyd. Turing on “common sense”: Cambridge resonances. In Floyd and Bokulich [FB17], chapter 5, pages 103–149. ISBN 3-319-53278-2, 3-319-53280-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QA29.T8 P45 2017. URL https:/

/link.springer.com/chapter/10.
1007/978-3-319-53280-6_5.

Freudenthal:2009:BHL

[FM09a] Gideon Freudenthal and Peter McLaughlin. Boris Hessen: In lieu of a biography. In The Social and Economic Roots of the Scientific Revolution: Texts by Boris Hessen and Henryk Grossmann [FM09c], chapter 7, pages 253–256. ISBN 1-4020-9603-8, 1-4020-9604-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174.6 .S63 2009. URL https:

//link.springer.com/chapter/10.
1007/978-1-4020-9604-4_7.

Freudenthal:2009:CMH

[FM09b] Gideon Freudenthal and Peter McLaughlin. Classical Marxist historiography of science: The Hessen– Grossmann-thesis. In The Social and Economic Roots of the Scientific Revolution: Texts by Boris Hessen and Henryk Grossmann [FM09c], chapter 1, pages 1–40. ISBN 1-4020-9603-8, 1-4020-9604-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174.6 .S63 2009. URL https://

link.springer.com/chapter/10.
1007/978-1-4020-9604-4_1.

Freudenthal:2009:SER

[FM09c] Gideon Freudenthal and Peter McLaughlin, editors. The Social and Economic Roots of the Scientific Revolution: Texts by Boris Hessen and Henryk Grossmann, volume 278 of Boston Studies in the Philosophy of Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2009. ISBN 1-4020-9603-8, 1-4020-9604-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). xi + 273 pp. LCCN Q174.6

                                              .S63 2009. URL http://www.loc.
                                              gov/catdir/enhancements/fy1312/
                                              2008942091-d.html;   http://www.
                                              loc.gov/catdir/enhancements/
                                              fy1312/2008942091-t.html;
                                              https://link.springer.com/book/
                                              10.1007/978-1-4020-9604-4.

Folse:1994:BFC

[Fol94] Henry J. Folse. Bohr’s framework of complementarity and the realism debate. In Faye and Folse [FF94], chapter 6, pages 119–139. ISBN 0-7923-2378-5 (hardcover), 94-015-8106-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 153 1994; QC16 1994. URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-015-8106-6_6.

Folse:1996:BED

[Fol96] Henry J. Folse. The Bohr–Einstein debate and the philosophers’ debate over realism versus anti-realism. In Cohen et al. [CHR96], pages 289–298. ISBN 90-481-4493-0, 94-015-8638-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-015-8638-
                                              2_20.

Follesdal:1999:HNS

[Føl99] Dagfinn Føllesdal. Hermeneutics and natural science. In Fehér et al. [FKR99], pages 293–298. ISBN 90-481-5257-7, 94-015-9293-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

                                              ???? URL https://link.springer.
                                              com/chapter/10.1007/978-94-015-
                                              9293-2_23.

Foldes:2017:NRK

[Fol17] Lucien Foldes. Notes on revisiting Klappholz and Agassi’s “methodological prescriptions in economics”. In Encouraging Openness: Essays for Joseph Agassi on the Occasion of His 90th Birthday [BAG17], chapter 36, pages 455–466. ISBN 3-319-57668-2 (hardcover), 3-319-57669-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN B833.E53 2017.

URL https://link.springer.com/
chapter/10.1007/978-3-319-57669-
5_36.

Folina:2018:TBU

[Fol18]  Janet Folina.  Towards a better

understanding of mathematical understanding. In Piazza and Pulcini [PP18], pages 121–146. ISBN 3-319-93341-8, 3-319-93342-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic).

LCCN QA8.4. URL https://link.
springer.com/chapter/10.1007/
978-3-319-93342-9_8.

Fontdevila:1996:PCG

[Fon96] Antonio Fontdevila. The present concept of the gene. In Munévar [Mun96], chapter 9, pages 179–199. ISBN 94-010-6622-1, 94-009-0305-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-009-0305-0_9.

Foppl:2007:ARM

[Föp07] August Föppl. On absolute and relative motion. In Renn et al. [RSS^{+}07a], pages 145–152. ISBN 1-4020-3999-9. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 no. 250 v. 3; QC173.6 .G469 2007. URL https:

//link.springer.com/chapter/10.
1007/978-1-4020-4000-9_15.

Fortun:2001:HGP

[For01] Michael A. Fortun. The human genome project: Past, present, and future anterior. In Allen and MacLeod [AM01b], chapter 20, pages 339–362. ISBN 1-4020-0495-8 (hardcover), 94-017-2956-5. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 228; Q175.55. URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-017-2956-7_20.

Forman:2007:SMS

[For07] Paul Forman. From the social to the moral to the spiritual: the postmodern exaltation of the history of science. In Positioning the history of science [RG07], chapter 9, pages 49–55. ISBN 1-4020-5419-X (hardcover), 1-4020-5420-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q125 .P77 2007.

                                              URL https://link.springer.com/
                                              chapter/10.1007/1-4020-5420-3_9.

Forster:2010:MCQ

[For10] Malcolm R. Forster. Miraculous consilience of quantum mechanics. In The Place of Probability in Science: In Honor of Ellery Eells (1953–2006) [Fet10a], chapter 9, pages 201–228. ISBN 90-481-3614-8 (hardcover), 90-481-3615-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175 .P563 2010. URL https://link.

                                              springer.com/chapter/10.1007/
                                              978-90-481-3615-5_9.

Foster:1991:DME

[Fos91] Patricia L. Foster. Directed mutation in Escherichia coli: Theory and mechanisms. In Organism and the Origins of Self [Tau91i], chapter 10, pages 213–234. ISBN 0-7923-1185-X, 94-011-3406-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QR180-189.5. URL https:/

/link.springer.com/chapter/10.
1007/978-94-011-3406-4_10.

Fowler:1994:SDI

[Fow94] David H. Fowler. The story of the discovery of incommensurability, revisited. In Trends in the Historiography of Science [GCN94], chapter 17, pages 221–235. ISBN 90-481-4264-4, 94-017-3596-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-017-
3596-4_17.

Fowler:2004:LFE

[Fow04] David H. Fowler. Logistic and fractions in Early Greek mathematics: A new interpretation. In Classics in the History of Greek Mathematics [Chr04a], chapter 20, pages 367–380. ISBN 1-4020-0081-2 (hardcover), 1-4020-2640-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QA22 .C596

2004. URL https://link.springer.
com/chapter/10.1007/978-1-4020-
2640-9_20.

FoxKeller:2007:BSH

[Fox07] Evelyn Fox Keller. Between science and history. In Positioning the history of science [RG07], chapter 10, pages 57–60. ISBN 1-4020-5419-X (hardcover), 1-4020-5420-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q125 .P77 2007.

URL https://link.springer.com/
chapter/10.1007/1-4020-5420-3_10.

Fox:2015:LEO

[Fox15] Robert Fox. Lindemann and Einstein:

The Oxford connexion. In Arabatzis et al. [ARS15], chapter 3, pages 23–31. ISBN 3-319-14552-5 (paperback), 3-319-14553-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q125 .R45 2015. URL https://link.

springer.com/chapter/10.1007/
978-3-319-14553-2_3.

Fang:2018:EWT

[FQ18] Yibing Fang and Wei Qian. The earliest Western-trained engineers in China’s iron and steel industry. In Mitcham et al. [MLNZ18], chapter 15, pages 173–186. ISBN 3-319-62448-2, 3-319-62450-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN TA15 .P45 2018. URL https:

//link.springer.com/chapter/10.
1007/978-3-319-62450-1_15.

Festa:2008:EIP

[FR08] Egidio Festa and Sophie Roux. The enigma of the inclined plane from Heron to Galileo. In Mechanics and Natural Philosophy Before the Scientific Revolution [LR08], chapter 8, pages 195–220. ISBN 1-4020-5966-3 (hardcover), 1-4020-5967-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QA802 .M43 2008.

URL https://link.springer.com/
chapter/10.1007/978-1-4020-5967-
4_8.

Frankel:1980:HDH

[Fra80] Hank Frankel. Hess’s development of his seafloor spreading hypothesis. In Scientific Discovery: Case Studies [Nic80e], chapter 18, pages 345–366. ISBN 90-277-1093-7, 94-009-9015-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 60. URL https://link.

springer.com/chapter/10.1007/
978-94-009-9015-9_18.

Frankfurt:1986:TMR

[Fra86] Harry G. Frankfurt. Two motivations for rationalism: Descartes and Spinoza. In Cohen and Wartofsky [CW85], chapter 3, pages 47–61. ISBN 90-277-1971-3, 94-009-5345-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-009-5349-9_3.

Frank:1991:REL

[Fra91] Hartwig Frank. Reform efforts of logic at Mid-Nineteenth Century in Germany. In Woodward and Cohen [WC91], pages 247–258. ISBN 94-010-5395-2, 94-011-3164-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-011-3164-3_22.

Franklin:1993:RFF

[Fra93] Allan Franklin. The rise of the ‘fifth force’. In Correspondence, Invariance and Heuristics: Essays in Honour of Heinz Post [FK93], chapter 4, pages 83–103. ISBN 90-481-4229-6, 94-017-1185-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-017-
1185-2_4.

Franklin:1996:TNA

[Fra96a] Allan Franklin. There are no antirealists in the laboratory. In Cohen et al. [CHR96], pages 131–148. ISBN 90-481-4493-0, 94-015-8638-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-015-8638-2_10.

Franksen:1996:IUN

[Fra96b] Ole Immanuel Franksen. Invariance under nesting — an aspect of arraybased logic with relation to Grassmann and Peirce. In Schubring [Sch96f], pages 303–335. ISBN 90-481-4758-1, 94-015-8753-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QA8.9-10.3. URL https:/

/link.springer.com/chapter/10.
1007/978-94-015-8753-2_26.

Franklin:1999:ADK

[Fra99a] Allan Franklin. The appearance and disappearance of the 17-kev neutrino. In Can that be Right?: Essays on Experiment, Evidence, and Science [Fra99c], chapter 3, pages 39–96. ISBN 94-010-6242-0, 94-011-5334-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-011-5334-8_3.

Franklin:1999:C

[Fra99b] Allan Franklin. Calibration. In Can that be Right?: Essays on Experiment, Evidence, and Science [Fra99c], chapter 9, pages 237–272. ISBN 94-010-6242-0, 94-011-5334-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-011-5334-8_9.

Franklin:1999:CRE

[Fra99c] Allan Franklin. Can that be Right?: Essays on Experiment, Evidence, and Science, volume 199 of Boston Studies in the Philosophy of Science. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 1999. ISBN 94-010-6242-0, 94-011-5334-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). ix + 318 pp. LCCN ???? URL

https://link.springer.com/book/
10.1007/978-94-011-5334-8.

Franklin:1999:DPJ

[Fra99d] Allan Franklin. Discovery, pursuit, and justification. In Can that be Right?: Essays on Experiment, Evidence, and Science [Fra99c], chapter 7, pages 163–182. ISBN 94-010-6242-0, 94-011-5334-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-011-
5334-8_7.

Franklin:1999:HAE

[Fra99e] Allan Franklin. How to avoid the experimenters’ regress. In Can that be Right?: Essays on Experiment, Evidence, and Science [Fra99c], chapter 2, pages 13–38. ISBN 94-010-6242-0, 94-011-5334-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-011-
5334-8_2.

Franklin:1999:ILR

[Fra99f] Allan Franklin. Instrumental loyalty and the recycling of expertise. In Can that be Right?: Essays on Experiment, Evidence, and Science [Fra99c], chapter 4, pages 97–131. ISBN 94-010-6242-0, 94-011-5334-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-011-5334-8_4.

Franklin:1999:ICP

[Fra99g] Allan Franklin. Introduction: Constructivism, postmodernism, and science. In Can that be Right?: Essays on Experiment, Evidence, and Science [Fra99c], chapter 1, pages 1–11. ISBN 94-010-6242-0, 94-011-5334-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-011-5334-8_1.

Franklin:1999:LE

[Fra99h] Allan Franklin. Laws and experiment. In Can that be Right?: Essays on Experiment, Evidence, and Science [Fra99c], chapter 10, pages 273–281. ISBN 94-010-6242-0, 94-011-5334-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-011-5334-
8_10.

Franklin:1999:RDR

[Fra99i] Allan Franklin. The resolution of discordant results. In Can that be Right?: Essays on Experiment, Evidence, and Science [Fra99c], chapter 8, pages 183–235. ISBN 94-010-6242-0, 94-011-5334-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-011-
5334-8_8.

Franklin:1999:RFF

[Fra99j] Allan Franklin. The rise of the “fifth force”. In Can that be Right?: Essays on Experiment, Evidence, and Science [Fra99c], chapter 5, pages 133–148. ISBN 94-010-6242-0, 94-011-5334-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-011-5334-
8_5.

Franklin:1999:TNA

[Fra99k] Allan Franklin. There are no antirealists in the laboratory. In Can that be Right?: Essays on Experiment, Evidence, and Science [Fra99c], chapter 6, pages 149–162. ISBN 94-010-6242-0, 94-011-5334-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-011-
5334-8_6.

Frank:2000:LAY

[Fra00] Daniel H. Frank. “… with all your heart and with all your soul… ”: The moral psychology of the Shemonah Peraqim. In Maimonides and the Sciences [CL00], chapter 3, pages 25–33. ISBN 90-481-5348-4, 94-017-2128-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-017-2128-8_3.

Francoeur:2001:MMA

[Fra01] Eric Francoeur. Molecular models and the articulation of structural constraints in chemistry. In Klein [Kle01b], chapter 7, pages 95–115. ISBN 1-4020-0100-2, 94-015-9737-5. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 222; QD7. URL https:

//link.springer.com/chapter/10.
1007/978-94-015-9737-1_7.

Francois:2009:PPO

[Fra09a] Dagognet François. Pharmacology as a physical object. In Brenner and Gayon [BG09], chapter 8, pages 189–200. ISBN 1-4020-9367-5, 1-4020-9368-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .F746 2009.

URL https://link.springer.com/
chapter/10.1007/978-1-4020-9368-
5_8.

Franklin:2009:SDC

[Fra09b] Allan Franklin. The spectrum of β decay: Continuous or discrete? A variety of errors in experimental investigation. In Going Amiss In Experimental Research [HSS09a], chapter 12, pages 211–235. ISBN 1-4020-8892-2 (hardcover), 1-4020-8893-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q172.5.E77 G647 2009.

URL https://link.springer.com/
chapter/10.1007/978-1-4020-8893-
3_12.

Franckowiak:2012:DCM

[Fra12] Rémi Franckowiak. Du clos and the mechanization of chemical philosophy. In Philosophy of Behavioral Biology [PR12b], chapter 12, pages 285–301. ISBN 94-007-1950-7, 94-007-1951-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QH331 .P4679 2012; Q175 .B657 v.282.

URL https://link.springer.com/
chapter/10.1007/978-94-007-4345-
8_12.

Franckowiak:2013:DCM

[Fra13] Rémi Franckowiak. Du Clos and the mechanization of chemical philosophy. In Garber and Roux [GR13b], chapter 12, pages 285–301. CODEN ???? ISBN 94-007-4344-0 (hardcover), 94-007-4345-9 (e-book). ISSN ???? LCCN QC6 .M3616 2013; QC6 .M43

2013. URL https://link.springer.
com/chapter/10.1007/978-94-007-
4345-8_12.

Frege:1974:LI

[Fre74a] Gottlob Frege. On the law of inertia. In Methodological and Historical Essays in the Natural and Social Sciences [CW74b], chapter 14, pages 257–276. ISBN 90-277-0378-7, 94-010-2128-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 14. URL https://link.

springer.com/chapter/10.1007/
978-94-010-2128-9_14.

Freudenthal:1974:ISF

[Fre74b] Hans Freudenthal. The impact of Von Staudt’s foundations of geometry. In For Dirk Struik: Scientific, Historical and Political Essays in Honor of Dirk J. Struik [CSW74], chapter 15, pages 189–200. ISBN 90-277-0379-5, 94-010-2115-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-010-
2115-9_15.

Frede:1986:HSP

[Fre86a] Dorothea Frede. Heidegger and the scandal of philosophy. In Cohen and Wartofsky [CW85], chapter 7, pages 129–151. ISBN 90-277-1971-3, 94-009-5345-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-009-
5349-9_7.

Frede:1986:PMA

[Fre86b]  Michael Frede.  Philosophy and

medicine in antiquity. In Cohen and Wartofsky [CW85], chapter 11, pages 211–232. ISBN 90-277-1971-3, 94-009-5345-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-009-
5349-9_11.

Freudenthal:1986:A

[Fre86c] Gideon Freudenthal. Afterword. In Atom and Individual in the Age of Newton: On the Genesis of the Mechanistic World View [Fre86f], chapter 16, pages 205–213. ISBN 94-010-8505-6, 94-009-4500-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic).

LCCN Q174 .F748 1986. URL https:

//link.springer.com/chapter/10.
1007/978-94-009-4500-5_16.

Freudenthal:1986:ACC

[Fre86d]  Gideon Freudenthal.  Alternative

conceptions of civil society. In Atom and Individual in the Age of Newton: On the Genesis of the Mechanistic World View [Fre86f], chapter 12, pages 153–160. ISBN 94-010-8505-6, 94-009-4500-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .F748 1986. URL https://link.

springer.com/chapter/10.1007/
978-94-009-4500-5_12.

Freudenthal:1986:ASP

[Fre86e] Gideon Freudenthal. The antifeudal social philosophy of Hobbes. In Atom and Individual in the Age of Newton: On the Genesis of the Mechanistic World View [Fre86f], chapter 10, pages 115–137. ISBN 94-010-8505-6, 94-009-4500-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .F748 1986. URL https://link.

springer.com/chapter/10.1007/
978-94-009-4500-5_10.

Freudenthal:1986:AIA

[Fre86f] Gideon Freudenthal. Atom and Individual in the Age of Newton: On the Genesis of the Mechanistic World View, volume 88 of Boston Studies in the Philosophy of Science. D. Reidel, Dordrecht, The Netherlands; Boston, MA, USA; Lancaster, UK; Tokyo, Japan, 1986. ISBN 94-010-8505-6, 94-009-4500-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). 288 (est.) pp. LCCN Q174 .F748 1986.

URL https://link.springer.com/
book/10.1007/978-94-009-4500-5.

Freudenthal:1986:BIE

[Fre86g] Gideon Freudenthal. The bourgeois individual and the essential properties of a particle in Newton’s thought. In Atom and Individual in the Age of Newton: On the Genesis of the Mechanistic World View [Fre86f], chapter 14, pages 173–188. ISBN 94-010-8505-6, 94-009-4500-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .F748 1986. URL https:

//link.springer.com/chapter/10.
1007/978-94-009-4500-5_14.

Freudenthal:1986:CSA

[Fre86h] Gideon Freudenthal. Civil society and the analytic–synthetic method. In Atom and Individual in the Age of Newton: On the Genesis of the Mechanistic World View [Fre86f], chapter 13, pages 161–168. ISBN 94-010-8505-6, 94-009-4500-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .F748 1986. URL https:

//link.springer.com/chapter/10.
1007/978-94-009-4500-5_13.

Freudenthal:1986:CECa

[Fre86i] Gideon Freudenthal. The concept of element in 17th century natural philosophy. In Atom and Individual in the Age of Newton: On the Genesis of the Mechanistic World View [Fre86f], chapter 5, pages 77–84. ISBN 94-010-8505-6, 94-009-4500-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .F748 1986.

URL https://link.springer.com/
chapter/10.1007/978-94-009-4500-
5_5.

Freudenthal:1986:CECb

[Fre86j] Gideon Freudenthal. The concept of element in 18th century social philosophy. In Atom and Individual in the Age of Newton: On the Genesis of the Mechanistic World View [Fre86f], chapter 7, pages 92–97. ISBN 94-010-8505-6, 94-009-4500-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .F748 1986.

URL https://link.springer.com/
chapter/10.1007/978-94-009-4500-
5_7.

Freudenthal:1986:CES

[Fre86k] Gideon Freudenthal. The concept of element in the systematic philosophy of Hobbes. In Atom and Individual in the Age of Newton: On the Genesis of the Mechanistic World View [Fre86f], chapter 6, pages 85–91. ISBN 94-010-8505-6, 94-009-4500-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .F748 1986.

URL https://link.springer.com/
chapter/10.1007/978-94-009-4500-
5_6.

Freudenthal:1986:DBL

[Fre86l] Gideon Freudenthal. The discussion between Leibniz and Newton on the concept of science. In Atom and Individual in the Age of Newton: On the Genesis of the Mechanistic World View [Fre86f], chapter 4, pages 44–71. ISBN 94-010-8505-6, 94-009-4500-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .F748 1986.

URL https://link.springer.com/
chapter/10.1007/978-94-009-4500-
5_4.

Freudenthal:1986:ESP

[Fre86m] Gideon Freudenthal. Element and system in the philosophy of Leibniz. In Atom and Individual in the Age of Newton: On the Genesis of the Mechanistic World View [Fre86f], chapter 15, pages 189–204. ISBN 94-010-8505-6, 94-009-4500-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .F748 1986. URL https:

//link.springer.com/chapter/10.
1007/978-94-009-4500-5_15.

Freudenthal:1986:EBR

[Fre86n] Gideon Freudenthal. England before the revolution. In Atom and Individual in the Age of Newton: On the Genesis of the Mechanistic World View [Fre86f], chapter 9, pages 108–114. ISBN 94-010-8505-6, 94-009-4500-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .F748 1986.

URL https://link.springer.com/
chapter/10.1007/978-94-009-4500-
5_9.

Freudenthal:1986:I

[Fre86o] Gideon Freudenthal. Introduction. In Atom and Individual in the Age of Newton: On the Genesis of the Mechanistic World View [Fre86f], chapter 1, pages 1–9. ISBN 94-010-8505-6, 94-009-4500-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .F748 1986. URL https://link.

springer.com/chapter/10.1007/
978-94-009-4500-5_1.

Freudenthal:1986:LFD

[Fre86p] Gideon Freudenthal. Leibniz’s foundations of dynamics. In Atom and Individual in the Age of Newton: On the Genesis of the Mechanistic World View [Fre86f], chapter 3, pages 29–43. ISBN 94-010-8505-6, 94-009-4500-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174

.F748 1986. URL https://link.
springer.com/chapter/10.1007/
978-94-009-4500-5_3.

Freudenthal:1986:NJT

[Fre86q] Gideon Freudenthal. Newton’s justification of the theory of absolute space. In Atom and Individual in the Age of Newton: On the Genesis of the Mechanistic World View [Fre86f], chapter 2, pages 13–28. ISBN 94-010-8505-6, 94-009-4500-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .F748 1986. URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-009-4500-5_2.

Freudenthal:1986:RBN

[Fre86r] Gideon Freudenthal. The relationship between natural and social philosophy in the work of Newton, Rousseau, and Smith. In Atom and Individual in the Age of Newton: On the Genesis of the Mechanistic World View [Fre86f], chapter 8, pages 98–103. ISBN 94-010-8505-6, 94-009-4500-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .F748 1986.

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-009-4500-
                                              5_8.

Freudenthal:1986:RCS

[Fre86s] Gideon Freudenthal. The rise of civil society in England. In Atom and Individual in the Age of Newton: On the Genesis of the Mechanistic World View [Fre86f], chapter 11, pages 138–152. ISBN 94-010-8505-6, 94-009-4500-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174

                                              .F748 1986. URL https://link.
                                              springer.com/chapter/10.1007/
                                              978-94-009-4500-5_11.

Freudenthal:1988:HSH

[Fre88] Gad Freudenthal. The hermeneutical status of the history of science: The views of Hélène Metzger. In Science in Reflection: The Israel Colloquium: Studies in History, Philosophy, and Sociology of Science Volume 3 [UM88], chapter 11, pages 123–144. ISBN 90-277-2713-9, 94-009-2957-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-009-2957-9_11.

Freudenthal:1991:ACF

[Fre91] Gad Freudenthal. (Al-)Chemical foundations for cosmological ideas: Ibn Sînâ on the geology of an eternal world. In Unguru [Ung91b], pages 47–73. ISBN 94-010-5476-2, 94-011-3342-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC5.53; QC6.9.

URL https://link.springer.com/
chapter/10.1007/978-94-011-3342-
5_3.

Freundlieb:1999:DBS

[Fre99]  Dieter Freundlieb.  The difference

between science and hermeneutics: Habermas’s theory of the necessarily normative nature of linguistic interpretation. In Fehér et al. [FKR99], pages 299–306. ISBN 90-481-5257-7, 94-015-9293-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-015-
9293-2_24.

Freudenthal:2000:MGP

[Fre00] Gad Freudenthal. Maimonides’ guide of the perplexed and the transmission of the mathematical tract “On Two Asymptotic Lines” in the Arabic, Latin and Hebrew medieval traditions. In Maimonides and the Sciences [CL00], chapter 4, pages 35–56. ISBN 90-481-5348-4, 94-017-2128-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-017-2128-8_4.

Freudenthal:2004:ACF

[Fre04] H. Freudenthal. Y avait-il une crise des fondements des mathématiques dans l’antiquité?. (French) [Was there a crisis of the foundations of mathematics in antiquity?]. In Classics in the History of Greek Mathematics [Chr04a], chapter 12, pages 233–241. ISBN 1-4020-0081-2 (hardcover), 1-4020-2640-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QA22 .C596 2004. URL https:/

/link.springer.com/chapter/10.
1007/978-1-4020-2640-9_12.

Frercks:2009:GRM

[Fre09] Jan Frercks. Going right and making it wrong: The reception of Fizeau’s ether-drift experiment of 1859. In Going Amiss In Experimental Research [HSS09a], chapter 11, pages 179–210. ISBN 1-4020-8892-2 (hardcover), 1-4020-8893-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q172.5.E77 G647 2009. URL https:

//link.springer.com/chapter/10.
1007/978-1-4020-8893-3_11.

Freire:2011:CCC

[Fre11] Olival Freire, Jr. Continuity and change: Charting David Bohm’s evolving ideas on quantum mechanics. In Brazilian Studies in Philosophy and History of Science: An account of recent works [KV11], chapter 22, pages 291–299. ISBN 90-481-9421-0 (hardcover), 90-481-9422-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175.3 .B73 2011.

URL https://link.springer.com/
chapter/10.1007/978-90-481-9422-
3_22.

Freeman:2015:PPE

[Fre15] Gary Freeman. Phyla, phylogeny, and embryonic body plans. In Conceptual Change in Biology: Scientific and Philosophical Perspectives on Evolution and Development [Lov15b], chapter 10, pages 221–241. ISBN 94-017-9411-1, 94-017-9412-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QH331 C66 2015; QH491; QH492 .C66

2014. URL https://link.springer.
com/chapter/10.1007/978-94-017-
9412-1_10.

Fritchman:1974:RE

[Fri74] Stephen H. Fritchman. A religion of Earth. In For Dirk Struik: Scientific, Historical and Political Essays in Honor of Dirk J. Struik [CSW74], chapter 26, pages 325–343. ISBN 90-277-0379-5, 94-010-2115-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-010-2115-9_26.

Friedlander:1976:SCV

[Fri76] M. W. Friedlander. Some comments on Velikovsky’s methodology. In Cohen et al. [CHMV76], pages 477–486. ISBN 90-277-0647-6, 90-277-0648-4 (paperback), 94-010-1449-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-010-1449-6_26.

Frisby:1990:GSS

[Fri90a] David P. Frisby. Georg Simmel and the study of modernity. In Georg Simmel and Contemporary Sociology [KPC90], chapter 5, pages 57–74. ISBN 94-010-6691-4 (hardcover), 94-009-0459-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN HM401-1281.

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-009-0459-
                                              0_5.

Frisby:1990:GSC

[Fri90b] David P. Frisby. Georg Simmel’s concept of society. In Georg Simmel and Contemporary Sociology [KPC90], chapter 4, pages 39–55. ISBN 94-010-6691-4 (hardcover), 94-009-0459-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN HM401-1281.

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-009-0459-
                                              0_4.

Friedlaender:2007:ARM

[Fri07a] Benedict Immanuel Friedlaender. Absolute or relative motion? In Renn et al. [RSS^{+}07a], pages 127–144. ISBN 1-4020-3999-9. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 no. 250 v. 3; QC173.6 .G469 2007.

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-1-4020-4000-
                                              9_14.

Friedman:2007:KNE

[Fri07b] Michael Friedman. Kant — Naturphilosophie — electromagnetism. In Hans Christian Ørsted and The Romantic Legacy In Science: Ideas, Disciplines, Practices [BCK07], chapter 8, pages 135–158. ISBN 1-4020-2979-9 (hardcover), 1-4020-2987-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC515.O37 H36

                                              2007. URL https://link.springer.
                                              com/chapter/10.1007/978-1-4020-
                                              2987-5_8.

Frigg:2010:FSR

[Fri10] Roman Frigg. Fiction and scientific representation. In Frigg and Hunter [FH10b], chapter 6, pages 97–138.

ISBN 90-481-3850-7 (hardcover), 90-481-3851-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN BH301.M47 B49 2010; Q175 .B52 2010.

URL https://link.springer.com/
chapter/10.1007/978-90-481-3851-
7_6.

Feldhay:2018:EEPa

[FRSV18a] Rivka Feldhay, Jürgen Renn, Matthias Schemmel, and Matteo Valleriani, editors. Emergence and Expansion of Pre-Classical Mechanics, volume 270 of Boston Studies in the Philosophy and History of Science. Springer-Ver-lag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2018. ISBN 3-319-90343-8 (hardcover), 3-319-90345-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). xvii + 327 pp. LCCN QA802. URL

https://link.springer.com/book/
10.1007/978-3-319-90345-3.

Feldhay:2018:EEPb

[FRSV18b] Rivka Feldhay, Jürgen Renn, Matthias Schemmel, and Matteo Valleriani, editors. Emergence and Expansion of Pre-classical Mechanics, volume 333 of Boston Studies in the Philosophy and History of Science. Springer-Ver-lag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2018. ISBN 3-319-90343-8, 3-319-90344-6, 3-319-90345-4 (ebook). xvii + 327 pp. LCCN QA802. URL

http://public.eblib.com/choice/
PublicFullRecord.aspx?p=5567632.

Fortun:1994:SSL

[FS94] Michael Fortun and Sylvan S. Schweber. Scientists and the state: The legacy of world war II. In Trends in the Historiography of Science [GCN94], chapter 24, pages 327–354. ISBN 90-481-4264-4, 94-017-3596-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-017-3596-4_24.

Forman:1996:NME

[FS96] Paul Forman and José M. (José Manuel) Sánchez Ron, editors. National Military Establishments and the Advancement of Science and Technology: Studies in 20th Century History, volume 180 of Boston Studies in the Philosophy of Science. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 1996. ISBN 0-7923-3541-4 (hardcover). xiv + 340 pp. LCCN Q174 .B67 vol. 180; U390 .N19 1996.

Faye:1997:I

[FSU97a] Jan Faye, Uwe Scheffler, and Max Urchs. Introduction. In Perspectives on Time [FSU97b], chapter 1, pages 1–58. ISBN 90-481-4774-3, 94-015-8875-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol.189. URL https://link.

                                              springer.com/chapter/10.1007/
                                              978-94-015-8875-1_1.

Faye:1997:PT

[FSU97b] Jan Faye, Uwe Scheffler, and Max Urchs. Perspectives on Time, volume 189 of Boston Studies in the Philosophy of Science. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 1997. ISBN 90-481-4774-3, 94-015-8875-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). vii + 472 pp. LCCN Q174 .B67 vol.189. URL

                                              https://link.springer.com/book/
                                              10.1007/978-94-015-8875-1.

Friedman:2016:SNA

[FT16] Sy-David Friedman and Claudio Ternullo. The search for new axioms in the hyperuniverse programme. In Objectivity, Realism, and Proof: FilMat Studies in the Philosophy of Mathematics [BS16], chapter 10, pages 165–188. ISBN 3-319-31642-7, 3-319-31644-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QA9

.O25 2016.  URL https://link.
springer.com/chapter/10.1007/
978-3-319-31644-4_10.

Frank:1963:D

[FTM^{+}63] Philipp G. Frank, Laszlo Tisza, Robert Maybury, Fritz Steckerl, William Wallace, Armand Siegel, and Abner Shimony. Discussion. In Wartofsky [War63], pages 72–76. ISBN 94-010-3265-3, 94-010-3263-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B678 1970.

URL https://link.springer.com/
chapter/10.1007/978-94-010-3263-
6_6.

Fu:1993:LCS

[Fu93a] Daiwie Fu. Local contexts, strategies and sinicization: A case study of the sinicization formulation in the social sciences of Taiwan (1970s–1980s). In Philosophy and Conceptual History of Science in Taiwan [LF93], chapter 13, pages 245–264. ISBN 94-010-5103-8, 94-011-2500-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q127.T28 .P455 1993. URL https:

//link.springer.com/chapter/10.
1007/978-94-011-2500-0_13.

Fu:1993:PDT

[Fu93b] Daiwie Fu. Problem domain, taxonomy, and comparativity in histories of science — with a case study in the comparative history of ‘optics’. In Philosophy and Conceptual History of Science in Taiwan [LF93], chapter 8, pages 123–147. ISBN 94-010-5103-8, 94-011-2500-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q127.T28 .P455 1993. URL https:

//link.springer.com/chapter/10.
1007/978-94-011-2500-0_8.

Fuller:2002:CST

[Ful02] Steve Fuller. A Catholic stance toward scientific inquiry for the 21st Century. In Hermeneutic Philosophy of Science, Van Gogh’s Eyes, and God: Essays in Honor of Patrick A. Heelan, S.J. [Bab02c], chapter 35, pages 403–410. ISBN 90-481-5926-1, 94-017-1767-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 225. URL https://link.

springer.com/chapter/10.1007/
978-94-017-1767-0_35.

Funkenstein:1986:PAK

[Fun86] Amos Funkenstein. The persecution of absolutes: On the Kantian and neo-Kantian theories of science. In Ullmann-Margalit [UM86a], chapter 5, pages 39–63. ISBN 90-277-2159-9 (paperback), 90-277-2158-0 (hardcover), 94-009-5496-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 94 Q175.

URL https://link.springer.com/
chapter/10.1007/978-94-009-5496-
0_5.

Furth:1969:RMT

[Für69] R. Fürth. The role of models in theoretical physics. In Cohen and Wartofsky [CW69a], pages 327–340. ISBN 94-010-3383-8, 94-010-3263-7, 94-010-3381-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-010-
3381-7_9.

Furst:1992:SBD

[Für92] P. Leo Fürst. Some biographical documents. In Ernst Mach — A Deeper Look: Documents and New Perspectives [Bla92a], chapter 1, pages 1–28. ISBN 94-010-5230-1, 94-011-2771-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN C5.53; QC6.9.

URL https://link.springer.com/
chapter/10.1007/978-94-011-2771-
4_1.

Futch:2008:LMT

[Fut08] Michael J. Futch. Leibniz’s Metaphysics of Time and Space, volume 258 of Boston Studies in the Philosophy of Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2008. ISBN 1-4020-8236-3 (hardcover), 1-4020-8237-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). x + 205 pp. LCCN B2599.S7 F88 2008. URL

https://link.springer.com/book/
10.1007/978-1-4020-8237-5.

Gabbay:1976:TDP

[Gab76] Dov M. Gabbay. Two-dimensional propositional tense logics. In Language in Focus: Foundations, Methods and Systems: Essays in Memory of Yehoshua Bar-Hillel [Kas76b], chapter 24, pages 569–583. ISBN 90-277-0644-1, 90-277-0645-X, 94-010-1876-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN P101-120.

URL https://link.springer.com/
chapter/10.1007/978-94-010-1876-
0_24.

Gabbey:2001:DTA

[Gab01] Alan Gabbey. Disciplinary transformations in the age of Newton: The case of metaphysics. In Between Leibniz, Newton, and Kant: Philosophy and Science in the Eighteenth Century [Lef01a], chapter 1, pages 3–23. ISBN 90-481-5774-9, 94-015-9729-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-015-9729-6_1.

Gabbey:2004:WWM

[Gab04] Alan Gabbey. What was “mechanical” about “the mechanical philosophy”? In The Reception of the Galilean Science of Motion in Seventeenth-Century Europe [PT04], chapter 2, pages 11–23. ISBN 1-4020-2454-1, 1-4020-2455-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC133 .R43

                                              2004. URL https://link.springer.
                                              com/chapter/10.1007/978-1-4020-
                                              2455-9_2.

Gahde:1989:BSB

[Gäh89] Ulrich Gähde. Bridge structures and the borderline between the internal and external history of science. In Imre Lakatos and Theories of Scientific Change [GGN89], chapter 17, pages 215–225. ISBN 94-010-7860-2, 94-009-3025-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

                                              ???? URL https://link.springer.
                                              com/chapter/10.1007/978-94-009-
                                              3025-4_17.

Gaifman:1986:EAM

[Gai86] Haim Gaifman. On the empirical application of mathematics: A comment. In Ullmann-Margalit [UM86a], chapter 2, pages 13–16. ISBN 90-277-2159-9 (paperback), 90-277-2158-0 (hardcover), 94-009-5496-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 94

Q175. URL https://link.springer.
com/chapter/10.1007/978-94-009-
5496-0_2.

Gaidenko:1994:ACR

[Gai94] Piama Gaidenko. The artistic culture of the Renaissance and the genesis of modern European science. In Trends in the Historiography of Science [GCN94], chapter 31, pages 421–432. ISBN 90-481-4264-4, 94-017-3596-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-017-3596-4_31.

Gale:1986:AR

[Gal86] George Gale. Anthropocentrism reconsidered. In Cohen and Wartofsky [CW85], chapter 12, pages 233–241. ISBN 90-277-1971-3, 94-009-5345-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-009-5349-
9_12.

Galison:1992:FPH

[Gal91] Peter Galison. Fortran, physics, and human nature. In The Invention of Physical Science: Intersections of Mathematics, Theology and Natural Philosophy Since the Seventeenth Century: Essays in Honor of Erwin N. Hiebert [NRS92], chapter 10, pages 225–260. ISBN 94-010-5097-X, 94-011-2488-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-011-
2488-1_10.

Galilei:1992:TD

[Gal92a] Galileo Galilei. Treatise on demonstration. In Galileo’s Logical Treatises: A Translation, with Notes and Commentary, of His Appropriated Latin Questions on Aristotle’s Posterior Analytics [Wal92e], chapter 3, pages 125–215. ISBN 90-481-4116-8, 94-015-8036-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

                                              ???? URL https://link.springer.
                                              com/chapter/10.1007/978-94-015-
                                              8036-6_3.

Galilei:1992:TFF

[Gal92b] Galileo Galilei. Treatise on foreknowledges and foreknowns. In Galileo’s Logical Treatises: A Translation, with Notes and Commentary, of His Appropriated Latin Questions on Aristotle’s Posterior Analytics [Wal92e], chapter 2, pages 85–123. ISBN 90-481-4116-8, 94-015-8036-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

                                              ???? URL https://link.springer.
                                              com/chapter/10.1007/978-94-015-
                                              8036-6_2.

Galison:2000:RP

[Gal00] Peter Galison. Random philosophy. In The Reality of the Unobservable: Observability, Unobservability and Their Impact on the Issue of Scientific Realism [AP00b], chapter 7, pages 123–128. ISBN 90-481-5458-8, 94-015-9391-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

                                              ???? URL https://link.springer.
                                              com/chapter/10.1007/978-94-015-
                                              9391-5_7.

Gal:2002:Ib

[Gal02a] Ofer Gal. Inflection. In Meanest Foundations and Nobler Superstructures: Hooke, Newton and the “Compounding of the Celestiall Motions of the Planetts” [Gal02c], chapter 2, pages 17–81. ISBN 90-481-6067-7, 94-017-2223-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QB351

.G35 2002.  URL https://link.
springer.com/chapter/10.1007/
978-94-017-2223-0_2.

Gal:2002:Ia

[Gal02b] Ofer Gal. Introduction. In Meanest Foundations and Nobler Superstructures: Hooke, Newton and the “Compounding of the Celestiall Motions of the Planetts” [Gal02c], chapter 1, pages 1–15. ISBN 90-481-6067-7, 94-017-2223-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QB351 .G35 2002. URL https:/

/link.springer.com/chapter/10.
1007/978-94-017-2223-0_1.

Gal:2002:MFN

[Gal02c] Ofer Gal. Meanest Foundations and Nobler Superstructures: Hooke, Newton and the “Compounding of the Celestiall Motions of the Planetts”, volume 229 of Boston Studies in the Philosophy of Science. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 2002. ISBN 90-481-6067-7, 94-017-2223-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). xiii + 239 pp. LCCN QB351 .G35 2002. URL

https://link.springer.com/book/
10.1007/978-94-017-2223-0.

Gal:2002:NS

[Gal02d] Ofer Gal. Newton’s synthesis. In Meanest Foundations and Nobler Superstructures: Hooke, Newton and the “Compounding of the Celestiall Motions of the Planetts” [Gal02c], chapter 4, pages 165–206. ISBN 90-481-6067-7, 94-017-2223-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QB351 .G35 2002. URL https:

//link.springer.com/chapter/10.
1007/978-94-017-2223-0_4.

Gal:2002:P

[Gal02e] Ofer Gal. Power. In Meanest Foundations and Nobler Superstructures: Hooke, Newton and the “Compounding of the Celestiall Motions of the Planetts” [Gal02c], chapter 3, pages 83–164. ISBN 90-481-6067-7, 94-017-2223-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QB351

.G35 2002.  URL https://link.
springer.com/chapter/10.1007/
978-94-017-2223-0_3.

Galavotti:2003:OEN

[Gal03] Maria Carla Galavotti. Observation and Experiment in the Natural and Social Sciences, volume 232 of Boston Studies in the Philosophy of Science. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 2003. ISBN 1-4020-1251-9 (hardcover), 90-481-6278-5 (paperback), 0-306-48123-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). x + 345 pp. LCCN Q175

.O24 2003. URL http://catdir.
loc.gov/catdir/enhancements/
fy0814/2003050110-d.html;
http://catdir.loc.gov/catdir/
enhancements/fy0814/2003050110-
t.html;   https://link.springer.
com/book/10.1007/0-306-48123-5.

Galison:2015:JSO

[Gal15] Peter Galison. The journalist, the scientist, and objectivity. In Objectivity in Science: New Perspectives from Science and Technology Studies [PRT15], chapter 4, pages 57–75. ISBN 3-319-14348-4 (hardcover), 3-319-14349-2 (e-book).

ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175 O25 2015.

URL https://link.springer.com/
chapter/10.1007/978-3-319-14349-
1_4.

Ganea:2015:RRV

[Gan15] Mihai Ganea. A remark on a relational version of Robinson’s arithmetic Q. In Romanian Studies in Philosophy of Science [PST15], chapter 8, pages 125–130. ISBN 3-319-16654-9 (hardcover), 3-319-16655-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-3-319-16655-1_8.

Garber:1995:RMC

[Gar95a] Elizabeth Garber. Reading mathematics, constructing physics: Fourier and his readers, 1822–1850. In No Truth Except in the Details: Essays in Honor of Martin J. Klein [KS95], chapter 2, pages 31–54. ISBN 94-010-4097-4, 94-011-0217-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-011-
0217-9_2.

Garciadiego:1995:PMB

[Gar95b] Alejandro R. Garciadiego. The Principles of Mathematics of Bertrand Russella. In Mexican Studies in the History and Philosophy of Science [RC95], chapter 14, pages 213–234. ISBN 94-010-6535-7 (hardcover), 94-009-0109-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-009-
0109-4_14.

Garbolino:2003:BIP

[Gar03] Paolo Garbolino. On Bayesian induction (and Pythagoric arithmetic). In Descartes’s Mathematical Thought [Sas03d], chapter 23, pages 327–331. ISBN 90-481-6487-7, 94-017-1225-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

                                              URL https://link.springer.com/
                                              chapter/10.1007/0-306-48123-5_23.

Garber:2012:RPH

[Gar12a] Daniel Garber. Remarks on the prehistory of the mechanical philosophy. In Philosophy of Behavioral Biology [PR12b], chapter 1, pages 3–26. ISBN 94-007-1950-7, 94-007-1951-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QH331 .P4679 2012; Q175 .B657 v.282. URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-007-4345-8_1.

Garvey:2012:FWC

[Gar12b] Brian Garvey. Free will, compatibilism, and the human nature wars: Should we be worried? In Philosophy of Behavioral Biology [PR12b], chapter 12, pages 255–274. ISBN 94-007-1950-7, 94-007-1951-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QH331 .P4679 2012; Q175 .B657

                                              v.282. URL https://link.springer.
                                              com/chapter/10.1007/978-94-007-
                                              1951-4_12.

Garber:2013:RPH

[Gar13] Daniel Garber. Remarks on the prehistory of the mechanical philosophy. In Garber and Roux [GR13b], chapter 1, pages 3–26. CODEN ???? ISBN 94-007-4344-0 (hardcover), 94-007-4345-9 (e-book). ISSN ???? LCCN QC6 .M3616 2013; QC6 .M43 2013.

URL https://link.springer.com/
chapter/10.1007/978-94-007-4345-
8_1.

Garavaso:2016:WRR

[Gar16] Pieranna Garavaso. The woman of reason: On the re-appropriation of rationality and the enjoyment of philosophy. In Meta-Philosophical Reflection on Feminist Philosophies of Science [AV16a], chapter 11, pages 185–202. ISBN 3-319-26346-3 (hardcover), 3-319-26348-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN HQ1154 .M48 2016; HQ1190 M48 2016.

URL https://link.springer.com/
chapter/10.1007/978-3-319-26348-
9_11.

Gassendi:1976:RAT

[Gas76a] P. Gassendi. Reality of absolute time.

In The Concepts of Space and Time: Their Structure and Their Development [Čap76], chapter 34, pages 195–201. ISBN 90-277-0375-2, 94-010-1727-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN BD632.

URL https://link.springer.com/
chapter/10.1007/978-94-010-1727-
5_34.

Gassendi:1976:RIV

[Gas76b] P. Gassendi. The reality of infinite void according to Aristotle. In The Concepts of Space and Time: Their Structure and Their Development [Čap76], chapter 17, pages 91–95. ISBN 90-277-0375-2, 94-010-1727-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN BD632.

URL https://link.springer.com/
chapter/10.1007/978-94-010-1727-
5_17.

Gattei:2009:WWE

[Gat09] Stefano Gattei. “Why, and to what extent, may a false hypothesis yield the truth?”. In Rethinking Popper [PC09], chapter 4, pages 47–61. ISBN 1-4020-9337-3, 1-4020-9338-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN B1649.P64 R48 2009; Q174 .B67 2009. URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-1-4020-9338-8_4.

Gatto:2018:TGM

[Gat18] Romano Gatto. A treatise on Galileian mechanics: Exercitationes in Mechanicis Aristotelis by the Jesuit Giovan Battista Zupi. In Feldhay et al. [FRSV18a], chapter 8, pages 215– 254. ISBN 3-319-90343-8 (hardcover), 3-319-90345-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic).

                                              LCCN QA802. URL https://link.
                                              springer.com/chapter/10.1007/
                                              978-3-319-90345-3_8.

Gauthier:1984:HLL

[Gau84] Yvon Gauthier. Hegel’s logic from a logical point of view1. In Hegel and the Sciences [CW84a], chapter 18, pages 303–310. ISBN 94-009-6235-5, 94-009-6233-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

                                              ???? URL https://link.springer.
                                              com/chapter/10.1007/978-94-009-
                                              6233-0_18.

Gauthier:1995:ILR

[Gau95] Yvon Gauthier. Internal logic. A radically constructive logic for mathematics and physics. In Québec Studies in the Philosophy of Science: Part I: Logic, Mathematics, Physics and History of Science Essays in Honor of Hugues Leblanc [MC95], chapter 8, pages 107–122. ISBN 94-010-7204-3, 94-009-1575-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-009-
1575-6_8.

Gaukroger:1999:BRN

[Gau99] Stephen Gaukroger. Beyond reality:

Nietzsche’s science of appearances. In Nietzsche, Theories of Knowledge, and Critical Theory: Nietzsche and the Sciences I [Bab99b], chapter 3, pages 37–49. ISBN 90-481-5233-X, 94-017-2430-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-017-
2430-2_3.

Gavroglu:1989:MSR

[Gav89] Kostas Gavroglu. The methodology of scientific research programmes and some developments in high energy physics. In Imre Lakatos and Theories of Scientific Change [GGN89], chapter 10, pages 123–133. ISBN 94-010-7860-2, 94-009-3025-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-009-3025-4_10.

Gavroglu:1990:GLF

[Gav90] Kostas Gavroglu. From gases and liquids to fluids: The formation of new concepts during the development of theories of liquids. In Greek Studies in the Philosophy and History of Science [Nic90], chapter 17, pages 251–277. ISBN 94-010-7403-8, 94-009-2015-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-009-2015-
6_17.

Gavroglu:1994:TDR

[Gav94] Kostas Gavroglu. Types of discourse and the reading of the history of the physical sciences. In Trends in the Historiography of Science [GCN94], chapter 5, pages 65–86. ISBN 90-481-4264-4, 94-017-3596-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-017-3596-4_5.

Gavroglu:1996:CTC

[Gav96] Kostas Gavroglu. Can theories of chemistry provide an argument against realism? In Cohen et al. [CHR96], pages 149–170. ISBN 90-481-4493-0, 94-015-8638-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

                                              ???? URL https://link.springer.
                                              com/chapter/10.1007/978-94-015-
                                              8638-2_11.

Gavroglu:2009:PWN

[Gav09] Kostas Gavroglu. A pioneer who never got it right: James Dewar and the elusive phenomena of cold. In Going Amiss In Experimental Research [HSS09a], chapter 9, pages 137–157. ISBN 1-4020-8892-2 (hardcover), 1-4020-8893-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q172.5.E77 G647 2009. URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-1-4020-8893-3_9.

Gavroglu:2014:HIH

[Gav14a] Kostas Gavroglu. Historiographical issues in the history of cold. In History of Artificial Cold, Scientific, Technological and Cultural Issues [Gav14b], chapter 1, pages 1–17. ISBN 94-007-7198-3 (hardcover), 94-007-7199-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC278 H57 2014; TP494 .H577 2014. URL https:/

/link.springer.com/chapter/10.
1007/978-94-007-7199-4_1.

Gavroglu:2014:HAC

[Gav14b] Kostas Gavroglu. History of Artificial Cold, Scientific, Technological and Cultural Issues, volume 299 of Boston Studies in the Philosophy and History of Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2014. ISBN 94-007-7198-3 (hardcover), 94-007-7199-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). xii + 288 pp. LCCN QC278 H57 2014; TP494 .H577 2014.

URL https://link.springer.com/
book/10.1007/978-94-007-7199-4.

Gayon:2009:PBH

[Gay09] Jean Gayon. Philosophy of biology: An historico–critical characterization. In Brenner and Gayon [BG09], chapter 9, pages 201–212. ISBN 1-4020-9367-5, 1-4020-9368-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .F746 2009. URL https:

//link.springer.com/chapter/10.
1007/978-1-4020-9368-5_9.

Gayon:2013:EC

[Gay13] Jean Gayon. Evolution and chance.

In Bergandi [Ber13c], chapter 3, pages 39–48. ISBN 94-007-5066-8, 94-007-5067-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN GE42 .S77 2013; Q175 .B73 v. 296.

URL https://link.springer.com/
chapter/10.1007/978-94-007-5067-
8_3.

Gruber:2005:CPH

[GB05] Howard E. Gruber and Katja Bödeker.

Creativity, Psychology and the History of Science, volume 245 of Boston Studies in the Philosophy of Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2005. ISBN 1-4020-2455-X, 1-4020-3491-1 (hardcover), 1-4020-3509-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). viii + 532 pp. LCCN BF408 .G793

2005. URL https://link.springer.
com/book/10.1007/1-4020-3509-8.

Gould:1994:ARS

[GC94] Carol C. Gould and Robert S. Cohen. Artifacts, Representations and Social Practice: Essays for Marx Wartofsky, volume 154 of Boston Studies in the Philosophy of Science. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 1994. ISBN 94-010-4390-6, 94-011-0902-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). xiv + 572 pp. LCCN B73 .A77 1994. URL

https://link.springer.com/book/
10.1007/978-94-011-0902-4.

Giller:1995:FBB

[GC95] K. E. Giller and G. Cadisch. Future benefits from biological nitrogen fixation: An ecological approach to agriculture. In No Truth Except in the Details: Essays in Honor of Martin J. Klein [KS95], chapter 13, pages 255–277. ISBN 94-010-4097-4, 94-011-0217-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-011-
0055-7_13.

Ginev:1997:IIP

[GC97] Dimitri Ginev and R. S. (Robert Sonné) Cohen, editors. Issues and images in the philosophy of science: scientific and philosophical essays in honour of Azarya Polikarov, volume 192 of Boston Studies in the Philosophy of Science. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 1997. ISBN 0-7923-4444-8 (hardcover), 94-011-5788-X. ISSN 0068-0346 (print), 2214-7942 (electronic). xvii + 403 pp. LCCN Q174 .B67 vol. 192. URL

https://link.springer.com/book/
10.1007/978-94-011-5788-9.

Gourko:2006:AIW

[GC06] Helena Gourko and Robert S. Cohen. Analogy in Indian and Western Philosophical Thought, volume 243 of Boston Studies in the Philosophy of Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2006. ISBN 1-4020-3339-7 (hardcover), 1-4020-3340-0 (ebook). ISSN 0068-0346 (print), 2214-7942 (electronic). viii + 273 pp. LCCN BD190 .Z52 2006; Q175.A1 B65 v.

243. URL https://link.springer.
com/book/10.1007/1-4020-3340-0.

Garcia-Carpintero:2010:FET

[GC10] Manuel García-Carpintero. Fictional entities, theoretical models and figurative truth. In Frigg and Hunter [FH10b], chapter 7, pages 139–168. ISBN 90-481-3850-7 (hardcover), 90-481-3851-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN BH301.M47 B49 2010; Q175 .B52 2010.

URL https://link.springer.com/
chapter/10.1007/978-90-481-3851-
7_7.

Grmek:1981:I

[GCC81a] Mirko Drazen Grmek, Robert S. Cohen, and Guido Cimino. Introduction. In Grmek et al. [GCC81b], pages 1–6. ISBN 90-277-1123-2, 94-010-1284-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-010-1284-
                                              3_1.

Grmek:1981:SDE

[GCC81b] Mirko Dražen Grmek, Robert S. Cohen, and Guido Cimino, editors. On Scientific Discovery: The Erice Lectures 1977, volume 34 of Boston Studies in the Philosophy of Science. D. Reidel, Dordrecht, The Netherlands; Boston, MA, USA; Lancaster, UK; Tokyo, Japan, 1981. ISBN 90-277-1123-2, 94-010-1284-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL

                                              https://link.springer.com/book/
                                              10.1007/978-94-010-1284-3.

Gavroglu:1994:THS

[GCN94] Kostas Gavroglu, Jean Christianidis, and Efthymios Nicolaidis. Trends in the Historiography of Science, volume 151 of Boston Studies in the Philosophy of Science. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 1994. ISBN 90-481-4264-4, 94-017-3596-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). xi + 460 pp. LCCN ????

                                              URL https://link.springer.com/
                                              book/10.1007/978-94-017-3596-4.

Geach:1976:BR

[Gea76] P. T. Geach. Back-reference. In Language in Focus: Foundations, Methods and Systems: Essays in Memory of Yehoshua Bar-Hillel [Kas76b], chapter 2, pages 25–39. ISBN 90-277-0644-1, 90-277-0645-X, 94-010-1876-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN P101-120.

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-010-1876-
                                              0_2.

Geiringer:1967:FPT

[Gei67] Hilda Geiringer. On the foundations of probability theory. In Cohen and Wartofsky [CW67], pages 199–227. ISBN 90-277-0013-3, 94-010-3508-3. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175 .B731 v.3.

URL https://link.springer.com/
chapter/10.1007/978-94-010-3508-
8_11.

Gellner:1976:EC

[Gel76] Ernest Gellner. An ethic of cognition.

In Essays in Memory of Imre Lakatos [CFW76], chapter 13, pages 161–177. ISBN 90-277-0655-7, 94-010-1451-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-010-1451-
9_13.

Gellner:1984:SRM

[Gel84] Ernest Gellner. The social roots of modern egalitarianism. In Rationality in Science and Politics [And84d], chapter 7, pages 111–130. ISBN 90-277-1953-5, 94-009-6254-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-009-6254-5_7.

Gelbart:1987:OFS

[Gel87] Nina Gelbart. Organicism and the future of scientific utopia. In Burwick [Bur87b], chapter 2, pages 49–70. ISBN 90-277-2541-1, 94-009-3917-5. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol.

105. US$64.00(U.S.). URL https:

//link.springer.com/chapter/10.
1007/978-94-009-3917-2_2.

Gellner:1989:HRN

[Gel89] Ernest Gellner. Human rights and the new circle of equity. In Freedom and Rationality: Essays in Honor of John Watkins From his Colleagues and Friends [DJ89], chapter 8, pages 125–140. ISBN 94-010-7571-9, 94-009-2380-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-009-
2380-5_8.

Gellner:1995:PP

[Gel95] Ernest Gellner. Prometheus perplexed. In Critical Rationalism, the Social Sciences and the Humanities: Essays for Joseph Agassi. Volume II [JL95b], chapter 1, pages 3–18. ISBN 0-7923-2961-9, 90-481-4430-2, 94-017-0441-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN B29

.C75 1995.  URL https://link.
springer.com/chapter/10.1007/
978-94-017-0441-0_1.

Geller:2016:MAF

[Gel16] Pamela L. Geller. This is not a manifesto: Archaeology and feminism. In Meta-Philosophical Reflection on Feminist Philosophies of Science [AV16a], chapter 9, pages 151–170. ISBN 3-319-26346-3 (hardcover), 3-319-26348-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN HQ1154 .M48 2016; HQ1190 M48 2016. URL https:

//link.springer.com/chapter/10.
1007/978-3-319-26348-9_9.

Gemtos:1990:LEM

[Gem90] Petros A. Gemtos. Law and economics: Methodological problems in their interdisciplinary cooperation. In Greek Studies in the Philosophy and History of Science [Nic90], chapter 16, pages 239–250. ISBN 94-010-7403-8, 94-009-2015-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-009-
2015-6_16.

Gendron:1971:RNP

[Gen71] Bernard Gendron. On the relation of neurological and psychological theories: A critique of the hardware thesis. In Buck and Cohen [BC71], pages 483–495. ISBN 90-277-0309-4, 94-010-3142-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-010-
3142-4_31.

Gendron:1983:VEE

[Gen83] Bernard Gendron. The viability of environmental ethics. In Philosophy and Technology [DR83], chapter 13, pages 187–194. ISBN 94-009-7126-5, 94-009-7124-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-009-
7124-0_13.

Genuth:1992:DHA

[Gen91] Sara Schechner Genuth. Devils’ hells and astronomers’ heavens: Religion, method, and popular culture in speculations about life on comets. In The Invention of Physical Science: Intersections of Mathematics, Theology and Natural Philosophy Since the Seventeenth Century: Essays in Honor of Erwin N. Hiebert [NRS92], chapter 1, pages 3–26. ISBN 94-010-5097-X, 94-011-2488-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-011-
2488-1_1.

Genova:2003:ICM

[Gen03] Dafina Genova. Idealized cognitive models and other mental representations. In Bulgarian Studies in the Philosophy of Science [Gin03a], chapter 9, pages 129–140. ISBN 1-4020-1496-1, 94-017-0961-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-017-0961-3_9.

George:2000:QO

[Geo00] Alexander George. Quine and observation. In Knowledge, Language and Logic: Questions for Quine [OK00], chapter 3, pages 21–45. ISBN 1-4020-0253-X, 94-011-3933-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-011-3933-5_3.

Gerrard:1995:WVA

[Ger95] Steve Gerrard. Wittgenstein versus artificial intelligence? In Science, Mind and Art: Essays on science and the humanistic understanding in art, epistemology, religion and ethics In honor of Robert S. Cohen [GSW95b], chapter 7, pages 89–98. ISBN 0-7923-2990-2 (hardcover), 94-011-0469-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 165 BD418.3. URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-011-0469-2_7.

Gerhart:2015:CBM

[Ger15a] John Gerhart. Cellular basis of morphogenetic change: Looking back after 30 years of progress on developmental signaling pathways. In Conceptual Change in Biology: Scientific and Philosophical Perspectives on Evolution and Development [Lov15b], chapter 8, pages 175–197. ISBN 94-017-9411-1, 94-017-9412-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QH331 C66 2015; QH491; QH492 .C66

2014. URL https://link.springer.
com/chapter/10.1007/978-94-017-
9412-1_8.

Gerson:2015:IRS

[Ger15b] Elihu M. Gerson. The interaction of research systems in the Evo-devo juncture. In Conceptual Change in Biology: Scientific and Philosophical Perspectives on Evolution and Development [Lov15b], chapter 20, pages 441–457. ISBN 94-017-9411-1, 94-017-9412-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QH331 C66 2015; QH491; QH492 .C66 2014.

URL https://link.springer.com/
chapter/10.1007/978-94-017-9412-
1_20.

Geschwind:1969:AHF

[Ges69a] Norman Geschwind. Anatomy and the higher functions of the brain. In Cohen and Wartofsky [CW69b], pages 98–136. ISBN 94-010-3380-3, 94-010-3378-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-010-3378-
7_3.

Geschwind:1969:WIW

[Ges69b] Norman Geschwind. The work and influence of Wernicke. In Cohen and Wartofsky [CW69b], pages 1–33. ISBN 94-010-3380-3, 94-010-3378-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-010-3378-7_1.

Geschwind:1974:Ab

[Ges74a] Norman Geschwind. The Alexias 1969. In Selected Papers on Language and the Brain [Ges74w], chapter 19, pages 382–430. ISBN 90-277-0263-2, 94-010-2093-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

                                              ???? URL https://link.springer.
                                              com/chapter/10.1007/978-94-010-
                                              2093-0_19.

Geschwind:1974:AAD

[Ges74b] Norman Geschwind. The anatomy of acquired disorders of reading. In Selected Papers on Language and the Brain [Ges74w], chapter 1, pages 1–17. ISBN 90-277-0263-2, 94-010-2093-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-010-2093-
                                              0_1.

Geschwind:1974:AAL

[Ges74c] Norman Geschwind. Apraxia and agraphia in a left-hander. In Selected Papers on Language and the Brain [Ges74w], chapter 26, pages 530–541. ISBN 90-277-0263-2, 94-010-2093-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-010-2093-
                                              0_26.

Geschwind:1974:Aa

[Ges74d] Norman Geschwind. The apraxias. In Selected Papers on Language and the Brain [Ges74w], chapter 14, pages 313–323. ISBN 90-277-0263-2, 94-010-2093-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

                                              ???? URL https://link.springer.
                                              com/chapter/10.1007/978-94-010-
                                              2093-0_14.

Geschwind:1974:CWB

[Ges74e] Norman Geschwind. Carl Wernicke, the Breslau School and the history of aphasia. In Selected Papers on Language and the Brain [Ges74w], chapter 4, pages 42–61. ISBN 90-277-0263-2, 94-010-2093-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-010-
2093-0_4.

Geschwind:1974:CND

[Ges74f] Norman Geschwind. Color-naming defects in association with alexia. In Selected Papers on Language and the Brain [Ges74w], chapter 9, pages 237–255. ISBN 90-277-0263-2, 94-010-2093-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-010-
2093-0_9.

Geschwind:1974:CA

[Ges74g] Norman Geschwind. Conduction aphasia. In Selected Papers on Language and the Brain [Ges74w], chapter 25, pages 509–529. ISBN 90-277-0263-2, 94-010-2093-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-010-
2093-0_25.

Geschwind:1974:DBE

[Ges74h] Norman Geschwind. The development of the brain and the evolution of language. In Selected Papers on Language and the Brain [Ges74w], chapter 7, pages 86–104. ISBN 90-277-0263-2, 94-010-2093-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.

                                              com/chapter/10.1007/978-94-010-
                                              2093-0_7.

Geschwind:1974:DGS

[Ges74i] Norman Geschwind. Developmental Gerstmann syndrome. In Selected Papers on Language and the Brain [Ges74w], chapter 18, pages 370–381. ISBN 90-277-0263-2, 94-010-2093-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-010-2093-
                                              0_18.

Geschwind:1974:DLM

[Ges74j] Norman Geschwind. Dichotic listening in man after section of neocortical commissures. In Selected Papers on Language and the Brain [Ges74w], chapter 15, pages 324–338. ISBN 90-277-0263-2, 94-010-2093-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-010-2093-0_15.

Geschwind:1974:DSA

[Ges74k] Norman Geschwind. Disconnexion syndromes in animals and man. In Selected Papers on Language and the Brain [Ges74w], chapter 8, pages 105–236. ISBN 90-277-0263-2, 94-010-2093-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-010-2093-
                                              0_8.

Geschwind:1974:DHC

[Ges74l] Norman Geschwind. Disorders of higher cortical function in children. In Selected Papers on Language and the Brain [Ges74w], chapter 22, pages 467–481. ISBN 90-277-0263-2, 94-010-2093-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-010-2093-
0_22.

Geschwind:1974:HBL

[Ges74m] Norman Geschwind. Human brain:

Left-right asymmetries in temporal speech region. In Selected Papers on Language and the Brain [Ges74w], chapter 17, pages 364–369. ISBN 90-277-0263-2, 94-010-2093-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-010-2093-0_17.

Geschwind:1974:HCD

[Ges74n] Norman Geschwind. A human cerebral deconnection syndrome. In Selected Papers on Language and the Brain [Ges74w], chapter 3, pages 22–41. ISBN 90-277-0263-2, 94-010-2093-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-010-2093-
0_3.

Geschwind:1974:ISA

[Ges74o] Norman Geschwind. Isolation of the speech area. In Selected Papers on Language and the Brain [Ges74w], chapter 16, pages 339–363. ISBN 90-277-0263-2, 94-010-2093-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-010-2093-0_16.

Geschwind:1974:LIE

[Ges74p] Norman Geschwind. Language-induced epilepsy. In Selected Papers on Language and the Brain [Ges74w], chapter 10, pages 256–267. ISBN 90-277-0263-2, 94-010-2093-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-010-2093-0_10.

Geschwind:1974:NAD

[Ges74q]  Norman Geschwind.  Non-aphasic

disorders of speech. In Selected Papers on Language and the Brain [Ges74w], chapter 6, pages 73–85. ISBN 90-277-0263-2, 94-010-2093-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-010-2093-0_6.

Geschwind:1974:OLB

[Ges74r] Norman Geschwind. The organization of language and the brain. In Selected Papers on Language and the Brain [Ges74w], chapter 21, pages 452–466. ISBN 90-277-0263-2, 94-010-2093-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-010-2093-
0_21.

Geschwind:1974:PPK

[Ges74s] Norman Geschwind. The paradoxical position of Kurt Goldstein in the history of aphasia 1964. In Selected Papers on Language and the Brain [Ges74w], chapter 5, pages 62–72. ISBN 90-277-0263-2, 94-010-2093-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-010-2093-
0_5.

Geschwind:1974:PAU

[Ges74t]  Norman Geschwind.  Problems in

the anatomical understanding of the aphasias. In Selected Papers on Language and the Brain [Ges74w], chapter 20, pages 431–451. ISBN 90-277-0263-2, 94-010-2093-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-010-2093-0_20.

Geschwind:1974:RRH

[Ges74u] Norman Geschwind. Random reports:

Human split–brain syndromes. In Selected Papers on Language and the Brain [Ges74w], chapter 2, pages 18– 21. ISBN 90-277-0263-2, 94-010-2093-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-010-2093-
0_2.

Geschwind:1974:R

[Ges74v] Norman Geschwind. A review 1972.

In Selected Papers on Language and the Brain [Ges74w], chapter 24, pages 498–508. ISBN 90-277-0263-2, 94-010-2093-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-010-
2093-0_24.

Geschwind:1974:SPL

[Ges74w] Norman Geschwind. Selected Papers on Language and the Brain, volume 16 of Boston Studies in the Philosophy of Science. D. Reidel, Dordrecht, The Netherlands; Boston, MA, USA; Lancaster, UK; Tokyo, Japan, 1974. ISBN 90-277-0263-2, 94-010-2093-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). xii + 555 pp. LCCN ????

URL https://link.springer.com/
book/10.1007/978-94-010-2093-0.

Geschwind:1974:SRA

[Ges74x] Norman Geschwind. Shrinking retrograde amnesia. In Selected Papers on Language and the Brain [Ges74w], chapter 13, pages 299–312. ISBN 90-277-0263-2, 94-010-2093-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-010-2093-0_13.

Geschwind:1974:VNE

[Ges74y] Norman Geschwind. The varieties of naming errors 1967. In Selected Papers on Language and the Brain [Ges74w], chapter 11, pages 268–283. ISBN 90-277-0263-2, 94-010-2093-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-010-2093-0_11.

Geschwind:1974:WCS

[Ges74z] Norman Geschwind. Wernicke’s contribution to the study of aphasia 1967. In Selected Papers on Language and the Brain [Ges74w], chapter 12, pages 284–298. ISBN 90-277-0263-2, 94-010-2093-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

                                              ???? URL https://link.springer.
                                              com/chapter/10.1007/978-94-010-
                                              2093-0_12.

Geschwind:1974:WDA

[Ges74-27] Norman Geschwind. Writing disturbances in acute confusional states. In Selected Papers on Language and the Brain [Ges74w], chapter 23, pages 482–497. ISBN 90-277-0263-2, 94-010-2093-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

                                              ???? URL https://link.springer.
                                              com/chapter/10.1007/978-94-010-
                                              2093-0_23.

Gilbert:1997:LEV

[GF97] Scott F. Gilbert and Marion Faber.

Looking at embryos: The visual and conceptual aesthetics of emerging form. In The Elusive Synthesis: Aesthetics and Science [Tau97a], chapter 6, pages 125–151. ISBN 0-7923-4763-3, 0-7923-3904-5, 94-009-1786-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-009-1786-6_6.

Galavotti:1990:CEE

[GG90] Maria Carla Galavotti and Guido Gambetta. Causality and exogeneity in econometric models. In Statistics in Science: The Foundations of Statistical Methods in Biology, Physics and Economics [CC90], chapter 3, pages 27–40. ISBN 94-010-6765-1, 94-009-0619-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QA276-280. URL https:/

/link.springer.com/chapter/10.
1007/978-94-009-0619-8_3.

Gavroglu:1991:TMK

[GG91] Kostas Gavroglu and Yorgos Goudaroulis, editors. Through measurement to knowledge: the selected papers of Heike Kamerlingh Onnes 1853–1926, volume 124 of Boston Studies in the Philosophy of Science. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 1991. ISBN 94-010-7433-X (hardcover), 94-009-2079-2. ISSN 0068-0346 (print), 2214-7942 (electronic). cxv + 570 pp. LCCN Q174 .B67 vol. 124. URL

https://link.springer.com/book/
10.1007/978-94-009-2079-8.

Ghirardi:1996:BTV

[GG96a] Giancarlo Ghirardi and Renata Grassi.

Bohm’s theory versus dynamical reduction. In Bohmian Mechanics and Quantum Theory: an Appraisal [CFG96], chapter 25, pages 353–377. ISBN 90-481-4698-4, 94-015-8715-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC173.96-174.52.

URL https://link.springer.com/
chapter/10.1007/978-94-015-8715-
0_25.

Grattan-Guinness:1996:WDG

[GG96b] Ivor Grattan-Guinness. Where does Grassmann fit in the history of logic? In Schubring [Sch96f], pages 211–216. ISBN 90-481-4758-1, 94-015-8753-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QA8.9-10.3.

URL https://link.springer.com/
chapter/10.1007/978-94-015-8753-
2_17.

Guantao:1996:IFE

[GG96c] Jin Guantao and Hua Guofan. Information and feedback in epistemology. In Chinese Studies in the History and Philosophy of Science and Technology [DC96], chapter 7, pages 61–75. ISBN 90-481-4546-5, 94-015-8717-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 180; U390 .N19 1996. URL https:

//link.springer.com/chapter/10.
1007/978-94-015-8717-4_7.

Gouzevitch:2003:TIB

[GG03] Irina Gouzevitch and Dmitri Gouzevitch. Travelling interchanges between the Russian Empire and Western Europe. In Simões et al. [SCD03b], chapter 9, pages 213–231. ISBN 1-4020-1259-4, 94-017-3584-0. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 v. 233. URL https:/

/link.springer.com/chapter/10.
1007/978-94-017-3584-1_9.

Gavroglu:1989:ILT

[GGN89] Kostas Gavroglu, Yorgos Goudaroulis, and Pantelis Nicolacopoulos. Imre Lakatos and Theories of Scientific Change, volume 111 of Boston Studies in the Philosophy of Science. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 1989. ISBN 94-010-7860-2, 94-009-3025-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). 484 (est.) pp. LCCN ???? URL

https://link.springer.com/book/
10.1007/978-94-009-3025-4.

Glick:2001:SPR

[GH01] Thomas F. Glick and Mark G. Henderson. The scientific and popular receptions of Darwin, Freud, and Einstein. In The Reception of Darwinism in the Iberian World: Spain, Spanish America and Brazil [GPSR01], chapter 14, pages 229–238. ISBN 94-010-3885-6, 94-010-0602-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-010-0602-6_14.

Ghirardi:1993:QWD

[Ghi93] Gian Carlo Ghirardi. The quantum worldview: Its difficulties and an attempt to overcome them. In Bridging the Gap: Philosophy, Mathematics, and Physics: Lectures on the Foundations of Science [CCG93], chapter 9, pages 175–197. ISBN 94-010-5101-1, 94-011-2496-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-011-
2496-6_9.

Guantao:1996:HCS

[GHQ96a] Jin Guantao, Fan Hongye, and Liu Qingfeng. Historical changes in the structure of science and technology (part two, a commentary). In Chinese Studies in the History and Philosophy of Science and Technology [DC96], chapter 14, pages 165–183. ISBN 90-481-4546-5, 94-015-8717-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 180; U390 .N19 1996. URL https:

//link.springer.com/chapter/10.
1007/978-94-015-8717-4_14.

Guantao:1996:SST

[GHQ96b] Jin Guantao, Fan Hongye, and Liu Qingfeng. The structure of science and technology in history: On the factors delaying the development of science and technology in China in comparison with the West since the 17th Century (Part One). In Chinese Studies in the History and Philosophy of Science and Technology [DC96], chapter 13, pages 137–164. ISBN 90-481-4546-5, 94-015-8717-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 180; U390 .N19 1996.

URL https://link.springer.com/
chapter/10.1007/978-94-015-8717-
4_13.

Giannoni:1971:ELP

[Gia71] Carlo Giannoni. Einstein and the Lorentz–Poincaré theory of relativity. In Buck and Cohen [BC71], pages 575–589. ISBN 90-277-0309-4, 94-010-3142-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-010-
3142-4_39.

Giannetto:1999:HQP

[Gia99] Enrico Giannetto. Heidegger and the question of physics. In Fehér et al.

[FKR99], pages 207–225. ISBN 90-481-5257-7, 94-015-9293-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-015-9293-2_17.

Gibson:2000:QWH

[Gib00] Roger F. Gibson. Quine, Wittgenstein, and holism. In Knowledge, Language and Logic: Questions for Quine [OK00], chapter 7, pages 81–93. ISBN 1-4020-0253-X, 94-011-3933-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-011-3933-5_7.

Giere:1971:SGA

[Gie71] Ronald N. Giere. The structure, growth and application of scientific knowledge: Reflections on relevance and the future of philosophy of science. In Buck and Cohen [BC71], pages 539–551. ISBN 90-277-0309-4, 94-010-3142-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-010-3142-4_36.

Giedymin:1976:ICR

[Gie76] Jerzy Giedymin. Instrumentalism and its critique: A reappraisal. In Essays in Memory of Imre Lakatos [CFW76], chapter 14, pages 179–207. ISBN 90-277-0655-7, 94-010-1451-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-010-1451-9_14.

Giedymin:1986:PPI

[Gie86] Jerzy Giedymin. Polish philosophy in the inter-War period and Ludwik Fleck’s theory of thought-styles and thought-collectives. In Cognition and Fact: Materials on Ludwik Fleck [CS86], chapter 10, pages 179–215. ISBN 94-010-8504-8, 94-009-4498-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-009-4498-
5_10.

Giere:1997:ESR

[Gie97] Ronald N. Giere. Explaining scientific revolutions. In Ginev and Cohen [GC97], chapter 5, pages 63–86. ISBN 0-7923-4444-8 (hardcover), 94-011-5788-X. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 192. URL https://link.

springer.com/chapter/10.1007/
978-94-011-5788-9_5.

Giere:2008:IMP

[Gie08] Ronald N. Giere. Incommensurability from a modelling perspective. In Rethinking Scientific Change and Theory Comparison: Stabilities, Ruptures, Incommensurabilities? [SSHH08], chapter 19, pages 265–269. ISBN 1-4020-6274-5 (hardcover), 1-4020-6279-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175 .R397

2008. URL https://link.springer.
com/chapter/10.1007/978-1-4020-
6279-7_19.

Giere:2012:HPS

[Gie12] Ronald N. Giere. History and philosophy of science: Thirty-five years later. In Mauskopf and Schmaltz [MS12b], chapter 5, pages 59–65. ISBN 94-007-1744-X, 94-007-1745-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175 .I5134 2011.

URL https://link.springer.com/
chapter/10.1007/978-94-007-1745-
9_5.

Gigerenzer:2003:WDN

[Gig03] Gerd Gigerenzer. Where do new ideas come from? A heuristics of discovery in the cognitive sciences. In Descartes’s Mathematical Thought [Sas03d], chapter 7, pages 99–139. ISBN 90-481-6487-7, 94-017-1225-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/0-306-48123-5_7.

Giglioni:2012:HBB

[Gig12] Guido Giglioni. How Bacon became Baconian. In Philosophy of Behavioral Biology [PR12b], chapter 2, pages 27– 54. ISBN 94-007-1950-7, 94-007-1951-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QH331 .P4679 2012; Q175 .B657 v.282. URL https:

//link.springer.com/chapter/10.
1007/978-94-007-4345-8_2.

Giglioni:2013:HBB

[Gig13] Guido Giglioni. How Bacon became Baconian. In Garber and Roux [GR13b], chapter 2, pages 27–54. CODEN ???? ISBN 94-007-4344-0 (hardcover), 94-007-4345-9 (e-book). ISSN ???? LCCN QC6 .M3616 2013; QC6 .M43

2013. URL https://link.springer.
com/chapter/10.1007/978-94-007-
4345-8_2.

Gilpin:1988:DUC

[Gil88]  Robert G. Gilpin.  Development

and underdevelopment: Conflicting perspectives on the third world. In Philosophy, History and Social Action: Essays in Honor of Lewis Feuer with an autobiographic essay by Lewis Feuer [HOO88c], chapter 7, pages 173–208. ISBN 94-010-7793-2, 94-009-2873-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-009-2873-
                                              2_7.

Gilbert:1991:REI

[Gil91] Scott F. Gilbert. The role of embryonic induction in creating self. In Organism and the Origins of Self [Tau91i], chapter 17, pages 341–360. ISBN 0-7923-1185-X, 94-011-3406-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QR180-189.5.

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-011-3406-

4_17. See comments [Tau91c].

Gillies:2003:SCS

[Gil03] Donald Gillies. Some comments on “styles of experimentation” by Ursula Klein. In Descartes’s Mathematical Thought [Sas03d], chapter 12, pages 199–201. ISBN 90-481-6487-7, 94-017-1225-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

                                              ???? URL https://link.springer.
                                              com/chapter/10.1007/0-306-48123-

5_12. See [Kle03].

Gilbert:2009:AMS

[Gil09a] Scott F. Gilbert. The adequacy of model systems for Evo–Devo: Modeling the formation of organisms/modeling the formation of society. In Mapping the Future of Biology: Evolving Concepts and Theories [BMP09b], chapter 5, pages 57–68. ISBN 1-4020-9635-6 (hardcover), 1-4020-9636-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QH331 .M295 2009.

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-1-4020-9636-
                                              5_5.

Gillies:2009:PSP

[Gil09b] Donald Gillies. Problem-solving and the problem of induction. In Rethinking Popper [PC09], chapter 9, pages 103–115. ISBN 1-4020-9337-3, 1-4020-9338-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN B1649.P64 R48 2009; Q174 .B67 2009.

URL https://link.springer.com/
chapter/10.1007/978-1-4020-9338-
8_9.

Ginzberg:1988:OMP

[Gin88] Eli Ginzberg. Occupational mobility:

A personal perspective. In Philosophy, History and Social Action: Essays in Honor of Lewis Feuer with an autobiographic essay by Lewis Feuer [HOO88c], chapter 8, pages 209–219. ISBN 94-010-7793-2, 94-009-2873-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-009-2873-
2_8.

Ginev:1997:MMH

[Gin97] Dimitri Ginev. Micro- and macrohermeneutics of science. In Ginev and Cohen [GC97], chapter 6, pages 87–93. ISBN 0-7923-4444-8 (hardcover), 94-011-5788-X. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 192. URL https://link.

springer.com/chapter/10.1007/
978-94-011-5788-9_6.

Ginev:2002:HCC

[Gin02] Dimitri Ginev. The hermeneutic context of constitution. In Hermeneutic Philosophy of Science, Van Gogh’s Eyes, and God: Essays in Honor of Patrick A. Heelan, S.J. [Bab02c], chapter 4, pages 43–52. ISBN 90-481-5926-1, 94-017-1767-2 (e-book).

ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 225.

URL https://link.springer.com/
chapter/10.1007/978-94-017-1767-
0_4.

Ginev:2003:BSP

[Gin03a] Dimitri Ginev. Bulgarian Studies in the Philosophy of Science, volume 236 of Boston Studies in the Philosophy of Science. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 2003. ISBN 1-4020-1496-1, 94-017-0961-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). xii + 209 pp. LCCN ????

URL https://link.springer.com/
book/10.1007/978-94-017-0961-3.

Ginev:2003:HSA

[Gin03b] Dimitri Ginev. How to be simultaneously an antiessentialist and a defender of science’s cognitive specificity. In Bulgarian Studies in the Philosophy of Science [Gin03a], chapter 13, pages 187–205. ISBN 1-4020-1496-1, 94-017-0961-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-017-
0961-3_13.

Ginev:2006:CCB

[Gin06]  Dimitri Ginev.  The Context of

Constitution: Beyond the Edge of Epistemological Justification, volume 247 of Boston Studies in the Philosophy of Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2006. ISBN 1-4020-4712-6 (hardcover), 1-4020-4713-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). ix + 261 pp. LCCN Q175.32.K45

G56 2006. URL http://www.loc.
gov/catdir/enhancements/fy0824/

2007464574-d.html;   http://
www.loc.gov/catdir/enhancements/
fy0824/2007464574-t.html.

Gingras:2007:SAH

[Gin07] Yves Gingras. The search for autonomy in history of science. In Positioning the history of science [RG07], chapter 11, pages 61–64. ISBN 1-4020-5419-X (hardcover), 1-4020-5420-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q125 .P77 2007.

URL https://link.springer.com/
chapter/10.1007/1-4020-5420-3_11.

Ginammi:2015:SA

[Gin15] Michele Ginammi. Structure and applicability. In From Logic to Practice: Italian Studies in the Philosophy of Mathematics [LPV15], chapter 11, pages 203–222. ISBN 3-319-10433-0, 3-319-10434-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QA8.4 .F76 2015; Q174 .B67 v.308.

URL https://link.springer.com/
chapter/10.1007/978-3-319-10434-
8_11.

Ginammi:2018:APG

[Gin18]  Michele Ginammi.  Applicability

problems generalized. In Piazza and Pulcini [PP18], pages 209–224. ISBN 3-319-93341-8, 3-319-93342-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QA8.4.

URL https://link.springer.com/
chapter/10.1007/978-3-319-93342-
9_12.

Giorello:1981:IRS

[Gio81] Giulio Giorello. Intuition and rigor:

Some problems of a ‘logic of discovery’ in mathematics. In Italian Studies in the Philosophy of Science [Chi81b], chapter 8, pages 113–135. ISBN 90-277-1073-2, 94-009-8937-7 (e-book).

ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-009-8937-5_8.

Girard:1992:OVL

[Gir92] René Girard. Origins: A view from the literature. In Understanding Origins: Contemporary Views on the Origin of Life, Mind and Society [VD92], chapter 2, pages 27–42. ISBN 90-481-4090-0, 94-015-8054-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-015-8054-0_2.

Giron:2001:MEN

[Gir01] Alvaro Girón. The moral economy of nature. In The Reception of Darwinism in the Iberian World: Spain, Spanish America and Brazil [GPSR01], chapter 12, pages 189–203. ISBN 94-010-3885-6, 94-010-0602-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-010-0602-6_12.

Giusti:2004:MHP

[Giu04] Enrico Giusti. A master and his pupils: Theories of motion in the Galilean School. In The Reception of the Galilean Science of Motion in Seventeenth-Century Europe [PT04], chapter 7, pages 119–135. ISBN 1-4020-2454-1, 1-4020-2455-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC133 .R43 2004.

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-1-4020-2455-
                                              9_7.

Gjelsvik:2000:EDM

[Gje00] Olav Gjelsvik. The epistemology of decision-making “naturalised”. In Knowledge, Language and Logic: Questions for Quine [OK00], chapter 9, pages 109–129. ISBN 1-4020-0253-X, 94-011-3933-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic).

LCCN ????  URL https://link.
springer.com/chapter/10.1007/
978-94-011-3933-5_9.

Gjelsvik:2007:CEP

[Gje07] Olav Gjelsvik. Causal explanation provides knowledge why. In Rethinking Explanation [PY07], chapter 5, pages 69–92. ISBN 1-4020-5580-3 (paperback), 1-4020-5581-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN BD237 .R48 2007; Q175.32.E97 R48 2007. URL https:

//link.springer.com/chapter/10.
1007/978-1-4020-5581-2_5.

Guntau:1991:ONS

[GL91] Martin Guntau and Hubert Laitko.

On the origin and nature of scientific disciplines. In Woodward and Cohen [WC91], pages 17–28. ISBN 94-010-5395-2, 94-011-3164-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-011-3164-3_2.

Gleiser:2003:STD

[GL03] Reinaldo J. Gleiser and Patricio S.

Letelier. Space–time defects: Open and closed shells revisited. In Renn et al. [RDS03], chapter 18, pages 383–395. ISBN 1-4020-1284-5, 1-4020-1285-3 (paperback). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 234. URL https://link.

springer.com/chapter/10.1007/
978-94-010-0111-3_18.

Glick:1987:CRR

[Gli87a] Thomas F. Glick, editor. The Comparative Reception of Relativity, volume 103 of Boston Studies in the Philosophy of Science. D. Reidel, Dordrecht, The Netherlands; Boston, MA, USA; Lancaster, UK; Tokyo, Japan, 1987. ISBN 90-277-2498-9, 94-009-3875-6. ISSN 0068-0346 (print), 2214-7942 (electronic). vii + 412 pp. LCCN Q174 .B67 vol. 103; QC173.5. US$79.00 (US).

                                              URL https://link.springer.com/
                                              book/10.1007/978-94-009-3875-5.

Glick:1987:CIR

[Gli87b] Thomas F. Glick. Cultural issues in the reception of relativity. In The Comparative Reception of Relativity [Gli87a], chapter 11, pages 381–400. ISBN 90-277-2498-9, 94-009-3875-6. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 103; QC173.5. US$79.00 (US). URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-009-3875-5_11.

Glick:2001:RDU

[Gli01] Thomas F. Glick. The reception of Darwinism in Uruguay. In The Reception of Darwinism in the Iberian World: Spain, Spanish America and Brazil [GPSR01], chapter 2, pages 29–52. ISBN 94-010-3885-6, 94-010-0602-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-010-0602-
                                              6_2.

Glouberman:1976:PMP

[Glo76]  M. Glouberman.  Prime matter,

predication, and the semantics of feature-placing. In Language in Focus: Foundations, Methods and Systems: Essays in Memory of Yehoshua Bar-Hillel [Kas76b], chapter 5, pages 75–104. ISBN 90-277-0644-1, 90-277-0645-X, 94-010-1876-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN P101-120. URL https://link.

springer.com/chapter/10.1007/
978-94-010-1876-0_5.

Glucker:1991:IPL

[Glu91] John Glucker. Images of Plato in Late Antiquity. In Unguru [Ung91b], pages 3–18. ISBN 94-010-5476-2, 94-011-3342-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC5.53; QC6.9. URL https://link.

springer.com/chapter/10.1007/
978-94-011-3342-5_1.

Glymour:1971:TRT

[Gly71] Clark Glymour. Theoretical realism and theoretical equivalence. In Buck and Cohen [BC71], pages 275–288. ISBN 90-277-0309-4, 94-010-3142-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-010-3142-
4_19.

Glymour:1983:TYD

[Gly83] Clark Glymour. The theory of your dreams. In Cohen and Laudan [CL83], chapter 3, pages 57–71. ISBN 94-009-7055-2, 94-009-7057-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-009-7055-7_3.

Grene:1976:TPB

[GM76] Marjorie Grene and Everett Mendelsohn. Topics in the Philosophy of Biology, volume 27 of Boston Studies in the Philosophy of Science. D. Reidel, Dordrecht, The Netherlands; Boston, MA, USA; Lancaster, UK; Tokyo, Japan, 1976. ISBN 90-277-0596-8, 94-010-1829-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). xiii + 454 pp. LCCN ???? URL

https://link.springer.com/book/
10.1007/978-94-010-1829-6.

Glassner:1989:IQE

[GM89a] Barry Glassner and Jonathan D. Moreno. Introduction: Quantification and enlightenment. In The Qualitative-Quantitative Distinction in the Social Sciences [GM89b], chapter 1, pages 1– 12. ISBN 90-481-8460-6, 94-017-3444-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 112. URL https://link.

springer.com/chapter/10.1007/
978-94-017-3444-8_1.

Glassner:1989:QQD

[GM89b] Barry Glassner and Jonathan D. Moreno. The Qualitative-Quantitative Distinction in the Social Sciences, volume 112 of Boston Studies in the Philosophy of Science. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 1989. ISBN 90-481-8460-6, 94-017-3444-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). vii + 231 pp. LCCN Q174 .B67 vol. 112.

URL https://link.springer.com/
book/10.1007/978-94-017-3444-8.

Garcia-Moliner:2014:BS

[GM14] Federico García-Moliner. A beautiful story. In The History of Physics in Cuba [BRW14a], chapter 27, pages 395–397. ISBN 94-017-8040-4, 94-017-8041-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC5.53; QC6.9; QC9 C9H57 2014; Q124.6-127.2. URL https://link.

springer.com/chapter/10.1007/
978-94-017-8041-4_27.

Gayon:2017:RRE

[GM17] Jean Gayon and Maël Montévil. Repetition and reversibility in evolution: Theoretical population genetics. In Time of Nature and the Nature of Time: Philosophical Perspectives of Time in Natural Sciences [BH17b], chapter 13, pages 275–314. ISBN 3-319-53723-7 (hardcover), 3-319-53725-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC173.59.S65 T56

2017. URL https://link.springer.
com/chapter/10.1007/978-3-319-
53725-2_13.

Grene:1986:SS

[GN86] Marjorie Grene and Debra Nails.

Spinoza and the Sciences, volume 91 of Boston Studies in the Philosophy of Science. D. Reidel, Dordrecht, The Netherlands; Boston, MA, USA; Lancaster, UK; Tokyo, Japan, 1986. ISBN 94-010-8511-0, 94-009-4514-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). xix + 339 pp. LCCN ????

URL https://link.springer.com/
book/10.1007/978-94-009-4514-2.

Goenner:1970:MPE

[Goe70] Hubert Goenner. Mach’s Principle and Einstein’s theory of gravitation. In Ernst Mach: Physicist and Philosopher [CS70], chapter 10, pages 200–215. ISBN 90-481-8318-9, 90-277-0016-8, 94-017-1462-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-017-
1462-4_10.

Gogelein:1987:TCS

[Gög87] Christoph Gögelein. The theory of color as the symbolism of insight. In Amrine et al. [AZW87], chapter 13, pages 247–254. ISBN 90-277-2265-X, 94-009-3761-X. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 97. US$79.00. URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-009-3761-1_13.

Gohau:1981:DIP

[Goh81] Gabriel Gohau. Difficulties inherent in a pedagogy of discovery in the teaching of the sciences. In Grmek et al. [GCC81b], pages 191–210. ISBN 90-277-1123-2, 94-010-1284-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-010-1284-3_9.

Goldmann:1969:SCC

[Gol69a] Lucien Goldmann. The subject of cultural creation. In Cohen and Wartofsky [CW69b], pages 241–260. ISBN 94-010-3380-3, 94-010-3378-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-010-3378-7_9.

Goldstein:1969:TH

[Gol69b] Leon J. Goldstein. Theory in history. In Cohen and Wartofsky [CW69b], pages 277–302. ISBN 94-010-3380-3, 94-010-3378-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

                                              ???? URL https://link.springer.
                                              com/chapter/10.1007/978-94-010-
                                              3378-7_11.

Golubovic:1979:CBB

[Gol79] Zagorka Golubović. Culture as a bridge between utopia and reality. In Praxis: Yugoslav Essays in the Philosophy and Methodology of the Social Sciences [MP79], chapter 10, pages 167–185.

ISBN 90-277-0968-8, 94-009-9355-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-009-9355-
6_10.

Goldstein:1983:TLH

[Gol83] Leon J. Goldstein. Toward a logic of historical constitution. In Epistemology, Methodology, and the Social Sciences [CW83a], chapter 2, pages 19–52. ISBN 90-481-8376-6, 94-017-1458-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-017-1458-7_2.

Goldberg:1987:PNW

[Gol87] Stanley Goldberg. Putting new wine in old bottles. In Glick [Gli87a], chapter 1, pages 1–26. ISBN 90-277-2498-9, 94-009-3875-6. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 103; QC173.5. US$79.00

(US). URL https://link.springer.
com/chapter/10.1007/978-94-009-
3875-5_1.

Goldstein:1989:RCO

[Gol89] Leon J. Goldstein. Reflections on conceptual openness and conceptual tension. In Freedom and Rationality: Essays in Honor of John Watkins From his Colleagues and Friends [DJ89], chapter 6, pages 87–110. ISBN 94-010-7571-9, 94-009-2380-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-009-2380-5_6.

Goldstein:1991:LBG

[Gol91]  Bernard R. Goldstein.  Levi Ben

Gerson: On astronomy and physical experiments. In Unguru [Ung91b], pages 75–82. ISBN 94-010-5476-2, 94-011-3342-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC5.53; QC6.9. URL https:

//link.springer.com/chapter/10.
1007/978-94-011-3342-5_4.

Goldberg:2003:QLE

[Gol03] Joshua N. Goldberg. Quasi-local energy. In Renn et al. [RDS03], chapter 17, pages 375–382. ISBN 1-4020-1284-5, 1-4020-1285-3 (paperback). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 234. URL https:

//link.springer.com/chapter/10.
1007/978-94-010-0111-3_17.

Golinski:2012:TKI

[Gol12] Jan Golinski. Thomas Kuhn and interdisciplinary conversation: Why historians and philosophers of science stopped talking to one another. In Mauskopf and Schmaltz [MS12b], chapter 2, pages 13–28. ISBN 94-007-1744-X, 94-007-1745-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175 .I5134 2011.

URL https://link.springer.com/
chapter/10.1007/978-94-007-1745-
9_2.

Gomez:2012:MSO

[Góm12] Susana Gómez. From a metaphysical to a scientific object: Mechanizing light in Galilean science. In Philosophy of Behavioral Biology [PR12b], chapter 8, pages 191–215. ISBN 94-007-1950-7, 94-007-1951-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QH331 .P4679 2012; Q175 .B657

v.282. URL https://link.springer.
com/chapter/10.1007/978-94-007-
4345-8_8.

Gomez:2013:MSO

[Góm13] Susana Gómez. From a metaphysical to a scientific object: Mechanizing light in Galilean science. In Garber and Roux [GR13b], chapter 8, pages 191–215. CODEN ???? ISBN 94-007-4344-0 (hardcover), 94-007-4345-9 (e-book). ISSN ???? LCCN QC6 .M3616 2013; QC6 .M43 2013. URL https://link.

springer.com/chapter/10.1007/
978-94-007-4345-8_8.

Gomes:2015:EBB

[Gom15a] Ana Carolina Vimieiro Gomes. The emergence of biotypology in Brazilian medicine: The Italian model, textbooks, and discipline building, 1930–1940. In Sciences in the Universities of Europe, Nineteenth and Twentieth Centuries: Academic Landscapes [SDG15b], chapter 21, pages 361–380. ISBN 94-017-9635-1, 94-017-9636-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q183.4.E85 2015. URL https://link.

springer.com/chapter/10.1007/
978-94-017-9636-1_21.

Gomez:2015:RUD

[Góm15b] Amparo Gómez. Risk, uncertainty, and the dimensions of technological rationality. In New Perspectives on Technology, Values, and Ethics: Theoretical and Practical [Gon15a], chapter 6, pages 105–121. ISBN 3-319-21869-7 (hardcover), 3-319-21870-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-3-319-21870-
0_6.

Gonzalez:1996:PMW

[Gon96] Wenceslao J. González. Prediction and mathematics: The Wittgensteinian approach. In Munévar [Mun96], chapter 15, pages 299–332. ISBN 94-010-6622-1, 94-009-0305-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-009-0305-0_15.

Gonzalez:2001:DEM

[Gon01] Armando García González. Darwinism, eugenics and Mendelism in Cuban biological education: 1900–1959. In The Reception of Darwinism in the Iberian World: Spain, Spanish America and Brazil [GPSR01], chapter 10, pages 153–169. ISBN 94-010-3885-6, 94-010-0602-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-010-
0602-6_10.

Gonzalez:2003:REE

[Gon03] Wenceslao J. Gonzalez. Rationality in experimental economics: An analysis of Reinhard Selten’s approach. In Descartes’s Mathematical Thought [Sas03d], chapter 5, pages 71–83. ISBN 90-481-6487-7, 94-017-1225-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/0-306-48123-5_5.

Gonzalez:2015:NPT

[Gon15a] Wenceslao J. Gonzalez. New Perspectives on Technology, Values, and Ethics: Theoretical and Practical, volume 315 of Boston Studies in the Philosophy and History of Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2015. ISBN 3-319-21869-7 (hardcover), 3-319-21870-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). xiv + 219 pp. LCCN ????

URL https://link.springer.com/
book/10.1007/978-3-319-21870-0.

Gonzalez:2015:RVC

[Gon15b] Wenceslao J. Gonzalez. On the role of values in the configuration of technology: From axiology to ethics. In New Perspectives on Technology, Values, and Ethics: Theoretical and Practical [Gon15a], chapter 1, pages 3–27. ISBN 3-319-21869-7 (hardcover), 3-319-21870-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-3-319-
21870-0_1.

Goodall:1963:C

[Goo63] Marcus Goodall. Comments. In Wartofsky [War63], pages 46–47. ISBN 94-010-3265-3, 94-010-3263-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B678 1970.

URL https://link.springer.com/
chapter/10.1007/978-94-010-3263-
6_3.

Goodman:1967:EA

[Goo67] Nelson Goodman. The epistemological argument. In Cohen and Wartofsky [CW67], pages 102–107. ISBN 90-277-0013-3, 94-010-3508-3. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175 .B731 v.3. URL https://link.

springer.com/chapter/10.1007/
978-94-010-3508-8_6.

Goodfield:1969:THB

[Goo69] June Goodfield. Theories and hypotheses in biology. In Cohen and Wartofsky [CW69a], pages 421–449. ISBN 94-010-3383-8, 94-010-3263-7, 94-010-3381-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/

                                              chapter/10.1007/978-94-010-3381-
                                              7_14.

Good:1974:RTA

[Goo74] Irving John Good. Random thoughts about randomness. In Schaffner and Cohen [SC74a], pages 117–135. ISBN 90-277-0408-2, 90-277-0409-0 (paperback). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175 .B7312.

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-010-2140-

1_9. Invited lecture in the symposium on the Concept of Randomness, dedicated to the memory of L. J. Savage, in the Biennial Meeting of the Philosophy of Science Association, Olds Plaza Hotel, Lansing, Michigan, October 27–29, 1972.

Goodman:1977:AP

[Goo77a] Nelson Goodman. Approach to the problems. In The Structure of Appearance [Goo77j], chapter 4, pages 93–109. ISBN 90-277-0774-X, 94-010-1184-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-010-1184-
                                              6_4.

Goodman:1977:CQ

[Goo77b] Nelson Goodman. Concreta and qualification. In The Structure of Appearance [Goo77j], chapter 7, pages 156–179. ISBN 90-277-0774-X, 94-010-1184-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-010-1184-
                                              6_7.

Goodman:1977:CD

[Goo77c] Nelson Goodman. Constructional definition. In The Structure of Appearance [Goo77j], chapter 1, pages 3–23. ISBN 90-277-0774-X, 94-010-1184-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-010-1184-6_1.

Goodman:1977:EB

[Goo77d] Nelson Goodman. Extralogical bases.

In The Structure of Appearance [Goo77j], chapter 3, pages 45–89. ISBN 90-277-0774-X, 94-010-1184-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-010-1184-6_3.

Goodman:1977:FRS

[Goo77e] Nelson Goodman. Foundations of a realistic system. In The Structure of Appearance [Goo77j], chapter 6, pages 135–155. ISBN 90-277-0774-X, 94-010-1184-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-010-
1184-6_6.

Goodman:1977:GA

[Goo77f] Nelson Goodman. The general apparatus. In The Structure of Appearance [Goo77j], chapter 2, pages 24–44. ISBN 90-277-0774-X, 94-010-1184-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-010-1184-6_2.

Goodman:1977:TE

[Goo77g] Nelson Goodman. Of time and eternity. In The Structure of Appearance [Goo77j], chapter 11, pages 258–276. ISBN 90-277-0774-X, 94-010-1184-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-010-1184-
                                              6_11.

Goodman:1977:PO

[Goo77h] Nelson Goodman. The problem of order. In The Structure of Appearance [Goo77j], chapter 9, pages 193–214. ISBN 90-277-0774-X, 94-010-1184-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-010-1184-
                                              6_9.

Goodman:1977:SS

[Goo77i] Nelson Goodman. Size and shape. In The Structure of Appearance [Goo77j], chapter 8, pages 180–189. ISBN 90-277-0774-X, 94-010-1184-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-010-1184-6_8.

Goodman:1977:SAb

[Goo77j] Nelson Goodman. The Structure of Appearance, volume 53 of Boston Studies in the Philosophy of Science. D. Reidel, Dordrecht, The Netherlands; Boston, MA, USA; Lancaster, UK; Tokyo, Japan, 1977. ISBN 90-277-0774-X, 94-010-1184-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). 335 (est.) pp. LCCN ???? URL

                                              https://link.springer.com/book/
                                              10.1007/978-94-010-1184-6.

Goodman:1977:SAa

[Goo77k] Nelson Goodman. The system of the Aufbau. In The Structure of Appearance [Goo77j], chapter 5, pages 110–134. ISBN 90-277-0774-X, 94-010-1184-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-010-1184-
6_5.

Goodman:1977:TQ

[Goo77l] Nelson Goodman. Topology of quality. In The Structure of Appearance [Goo77j], chapter 10, pages 215–257. ISBN 90-277-0774-X, 94-010-1184-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-010-1184-
6_10.

Goodman:1986:EA

[Goo86] Nelson Goodman. The epistemological argument. In Cohen and Wartofsky [CW85], chapter 5, pages 52–57. ISBN 90-277-1971-3, 94-009-5345-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-009-5345-1_5.

Goodwin:1992:EGF

[Goo92] Brian C. Goodwin. The evolution of generic forms. In Understanding Origins: Contemporary Views on the Origin of Life, Mind and Society [VD92], chapter 11, pages 213–226. ISBN 90-481-4090-0, 94-015-8054-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-015-8054-
0_11.

Good:1995:MPB

[Goo95] I. J. Good. The mathematics of philosophy: A brief review of my work. In Critical Rationalism, Metaphysics and Science: Essays for Joseph Agassi. Volume I [JL95a], chapter 13, pages 211–238. ISBN 94-010-4211-X, 94-011-0471-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

                                              ???? URL https://link.springer.
                                              com/chapter/10.1007/978-94-011-
                                              0471-5_13.

Goodman:2000:MN

[Goo00] L. E. Goodman. Maimonidean naturalism. In Maimonides and the Sciences [CL00], chapter 5, pages 57–85. ISBN 90-481-5348-4, 94-017-2128-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-017-2128-8_5.

Goodgall:2001:BEC

[Goo01] Pedro M. Pruna Goodgall. Biological evolutionism in Cuba at the end of the Nineteenth Century. In The Reception of Darwinism in the Iberian World: Spain, Spanish America and Brazil [GPSR01], chapter 3, pages 53–64. ISBN 94-010-3885-6, 94-010-0602-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-010-0602-
                                              6_3.

Gordin:2007:PAS

[Gor07] Michael D. Gordin. Without parallels?: Averting a Schweberian dystopia. In Positioning the history of science [RG07], chapter 12, pages 65–68. ISBN 1-4020-5419-X (hardcover), 1-4020-5420-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q125 .P77 2007. URL https://link.

                                              springer.com/chapter/10.1007/1-
                                              4020-5420-3_12.

Gottlieb:1994:LFH

[Got94] Roger S. Gottlieb. Levinas, feminism, Holocaust, ecocide. In Artifacts, Representations and Social Practice:

Essays for Marx Wartofsky [GC94], chapter 22, pages 365–376. ISBN 94-010-4390-6, 94-011-0902-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN B73 .A77 1994.

URL https://link.springer.com/
chapter/10.1007/978-94-011-0902-
4_22.

Gould:1976:DTS

[Gou76] Stephen Jay Gould. D’Arcy Thompson and the science of form. In Topics in the Philosophy of Biology [GM76], chapter 3, pages 66–97. ISBN 90-277-0596-8, 94-010-1829-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-010-1829-6_3.

Gould:1983:BCS

[Gou83] Carol C. Gould. Beyond causality in the social sciences: Reciprocity as a model of non-exploitative social relations. In Epistemology, Methodology, and the Social Sciences [CW83a], chapter 3, pages 53–88. ISBN 90-481-8376-6, 94-017-1458-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-017-
1458-7_3.

Goudaroulis:1989:MPP

[Gou89] Yorgos Goudaroulis. Many-particle physics: Calculational complications that become a blessing for methodology. In Imre Lakatos and Theories of Scientific Change [GGN89], chapter 11, pages 135–145. ISBN 94-010-7860-2, 94-009-3025-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-009-
3025-4_11.

Goudaroulis:1990:MOC

[Gou90] Yorgos Goudaroulis. A matter of order: A controversy between Heisenberg and London. In Greek Studies in the Philosophy and History of Science [Nic90], chapter 18, pages 279–291. ISBN 94-010-7403-8, 94-009-2015-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-009-2015-6_18.

Goux:1992:PMM

[Gou92] Jean-Joseph Goux. Primitive money, modern money. In Understanding Origins: Contemporary Views on the Origin of Life, Mind and Society [VD92], chapter 7, pages 145–149. ISBN 90-481-4090-0, 94-015-8054-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-015-8054-
                                              0_7.

Goudaroulis:1994:CHI

[Gou94a] Yorgos Goudaroulis. Can the history of instrumentation tell us anything about scientific practice? In Trends in the Historiography of Science [GCN94], chapter 12, pages 161–168. ISBN 90-481-4264-4, 94-017-3596-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-017-3596-4_12.

Gould:1994:MAM

[Gou94b] Carol C. Gould. Marx after Marxism. In Artifacts, Representations and Social Practice: Essays for Marx Wartofsky [GC94], chapter 23, pages 377–396. ISBN 94-010-4390-6, 94-011-0902-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN B73 .A77

1994. URL https://link.springer.
com/chapter/10.1007/978-94-011-
0902-4_23.

Gourko:2006:IE

[Gou06a] Helena Gourko. Introductory essay.

In Analogy in Indian and Western Philosophical Thought [GC06], chapter 1, pages 1–40. ISBN 1-4020-3339-7 (hardcover), 1-4020-3340-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN BD190 .Z52 2006; Q175.A1 B65 v. 243. URL https:

//link.springer.com/chapter/10.
1007/1-4020-3340-0_1.

Gourko:2006:CB

[Gou06b] Helena Gourko. On the composition of this book. In Analogy in Indian and Western Philosophical Thought [GC06], chapter 2, pages 41–44. ISBN 1-4020-3339-7 (hardcover), 1-4020-3340-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN BD190 .Z52 2006; Q175.A1 B65 v. 243.

URL https://link.springer.com/
chapter/10.1007/1-4020-3340-0_2.

Govoni:2015:CBW

[Gov15] Paola Govoni. Challenging the backlash: Women science students in Italian universities (1870s–2000s). In Sciences in the Universities of Europe, Nineteenth and Twentieth Centuries: Academic Landscapes [SDG15b], chapter 5, pages 69–88. ISBN 94-017-9635-1, 94-017-9636-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q183.4.E85 2015. URL https:

//link.springer.com/chapter/10.
1007/978-94-017-9636-1_5.

George:1979:MLQ

[GP79] C. George and I. Prigogine. The macroscopic level of quantum mechanics [1972a and b] (with C. George and I. Prigogine). In Cohen and Stachel [CS79a], chapter 39, pages 571–598. ISBN 90-277-0651-4, 90-277-0652-2 (paperback), 94-009-9349-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 21

QC7. URL https://link.springer.
com/chapter/10.1007/978-94-009-
9349-5_39.

Gavroglu:2003:PAG

[GP03] Kostas Gavroglu and Manolis Patiniotis. Patterns of appropriation in the Greek intellectual life of the 18th Century. In Renn et al. [RDS03], chapter 29, pages 569–591. ISBN 1-4020-1284-5, 1-4020-1285-3 (paperback). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 234. URL https://link.

springer.com/chapter/10.1007/
978-94-010-0111-3_29.

Glick:2001:RDI

[GPSR01] Thomas F. Glick, Miguel Angel Puig-Samper, and Rosaura Ruiz. The Reception of Darwinism in the Iberian World: Spain, Spanish America and Brazil, volume 221 of Boston Studies in the Philosophy of Science. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 2001. ISBN 94-010-3885-6, 94-010-0602-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). 282 (est.) pp. LCCN ???? URL

https://link.springer.com/book/
10.1007/978-94-010-0602-6.

Gottfried:1989:AST

[GR89] Michael Gottfried and Thomas Ryckman. The apparatus of sublanguage transformations. In The Form of Information in Science: Analysis of an Immunology Sublanguage [HGR^{+}89], chapter 5, pages 112–150. ISBN 94-010-7777-0, 94-009-2837-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN B832.A-832.Z.

URL https://link.springer.com/
chapter/10.1007/978-94-009-2837-
4_5.

Goldstein:2003:VUS

[GR03] Catherine Goldstein and Jim Ritter.

The varieties of unity: Sounding unified theories 1920–1930. In Renn et al. [RDS03], chapter 6, pages 93–149. ISBN 1-4020-1284-5, 1-4020-1285-3 (paperback). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 234. URL https://link.

springer.com/chapter/10.1007/
978-94-010-0111-3_6.

Gavroglu:2007:PHS

[GR07] Kostas Gavroglu and Jürgen Renn. Positioning the History of Science, volume 248 of Boston Studies in the Philosophy of Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2007. ISBN 1-4020-5195-6, 1-4020-5419-X (hardcover), 1-4020-5420-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). vii + 188 pp. LCCN Q125 .P77

2007. URL https://link.springer.
com/book/10.1007/1-4020-5420-3.

Gunergun:2011:I

[GR11a] Feza Günergun and Dhruv Raina.

Introduction. In Science between Europe and Asia: Historical Studies on the Transmission, Adoption and Adaptation of Knowledge [GR11b], chapter 1, pages 1–9. ISBN 90-481-9968-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q125 .S354 2010; Q175.52.A78 S35

2010. URL https://link.springer.

                                              com/chapter/10.1007/978-90-481-
                                              9968-6_1.

Gunergun:2011:SBE

[GR11b] Feza Günergun and Dhruv Raina. Science between Europe and Asia: Historical Studies on the Transmission, Adoption and Adaptation of Knowledge, volume 275 of Boston Studies in the Philosophy of Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2011. ISBN 90-481-9968-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). xiii + 279 pp. LCCN Q125 .S354 2010; Q175.52.A78 S35 2010. URL

                                              https://link.springer.com/book/
                                              10.1007/978-90-481-9968-6.

Garber:2012:EMN

[GR12] Daniel Garber and Sophie Roux. Erratum: The Mechanization of Natural Philosophy. In Philosophy of Behavioral Biology [PR12b], chapter 13, page E1. ISBN 94-007-1950-7, 94-007-1951-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QH331 .P4679 2012; Q175 .B657 v.282.

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-007-4345-

8_13. See [Ano12q].

Garber:2013:EMN

[GR13a] Daniel Garber and Sophie Roux. Erratum: The mechanization of natural philosophy. In The Mechanization of Natural Philosophy [GR13b], chapter 13, page E1. CODEN ???? ISBN 94-007-4344-0 (hardcover), 94-007-4345-9 (e-book). ISSN ???? LCCN QC6 .M3616 2013; QC6 .M43 2013.

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-007-4345-
                                              8_13.

Garber:2013:MNP

[GR13b] Daniel Garber and Sophie Roux, editors. The Mechanization of Natural Philosophy, volume 300 of Boston Studies in the Philosophy of Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2013. ISBN 94-007-4344-0 (hardcover), 94-007-4345-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC6 .M3616 2013; QC6 .M43 2013. URL https://link.

springer.com/book/10.1007/978-
94-007-4345-8.

Granger:1983:DP

[Gra83a] Gilles-Gaston Granger. The découpage of phenomena. In Formal Thought and the Sciences of Man [Gra83b], chapter 4, pages 48–84. ISBN 94-009-7039-0, 94-009-7037-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-009-7037-3_4.

Granger:1983:FTS

[Gra83b] Gilles-Gaston Granger. Formal Thought and the Sciences of Man, volume 75 of Boston Studies in the Philosophy of Science. D. Reidel, Dordrecht, The Netherlands; Boston, MA, USA; Lancaster, UK; Tokyo, Japan, 1983. ISBN 94-009-7039-0, 94-009-7037-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). 240 (est.) pp. LCCN ????

URL https://link.springer.com/
book/10.1007/978-94-009-7037-3.

Granger:1983:LVI

[Gra83c] Gilles-Gaston Granger. Language as a vehicle of information. In Formal Thought and the Sciences of Man [Gra83b], chapter 2, pages 13–30. ISBN 94-009-7039-0, 94-009-7037-4 (e-book).

ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-009-7037-3_2.

Granger:1983:P

[Gra83d] Gilles-Gaston Granger. Postface (1982). In Formal Thought and the Sciences of Man [Gra83b], chapter 8, pages 181–193. ISBN 94-009-7039-0, 94-009-7037-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

                                              ???? URL https://link.springer.
                                              com/chapter/10.1007/978-94-009-
                                              7037-3_8.

Granger:1983:PFP

[Gra83e] Gilles-Gaston Granger. The problem of forms and the philosophy of the sciences. In Formal Thought and the Sciences of Man [Gra83b], chapter 1, pages 1–12. ISBN 94-009-7039-0, 94-009-7037-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

                                              ???? URL https://link.springer.
                                              com/chapter/10.1007/978-94-009-
                                              7037-3_1.

Granger:1983:QQ

[Gra83f] Gilles-Gaston Granger. Quality and quantity. In Formal Thought and the Sciences of Man [Gra83b], chapter 5, pages 85–116. ISBN 94-009-7039-0, 94-009-7037-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

                                              ???? URL https://link.springer.
                                              com/chapter/10.1007/978-94-009-
                                              7037-3_5.

Granger:1983:SLF

[Gra83g] Gilles-Gaston Granger. Scientific languages and formalisms. In Formal Thought and the Sciences of Man [Gra83b], chapter 3, pages 31–47. ISBN 94-009-7039-0, 94-009-7037-4 (e-book).

ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-009-7037-3_3.

Granger:1983:SA

[Gra83h] Gilles-Gaston Granger. Structuring and axiomatizing. In Formal Thought and the Sciences of Man [Gra83b], chapter 6, pages 117–149. ISBN 94-009-7039-0, 94-009-7037-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-009-7037-3_6.

Granger:1983:UI

[Gra83i] Gilles-Gaston Granger. The understanding of the individual. In Formal Thought and the Sciences of Man [Gra83b], chapter 7, pages 150–180. ISBN 94-009-7039-0, 94-009-7037-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-009-7037-
3_7.

Grant:1991:CIM

[Gra91] Edward Grant. Celestial incorruptibility in medieval cosmology 1200–1687. In Unguru [Ung91b], pages 101–127. ISBN 94-010-5476-2, 94-011-3342-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC5.53; QC6.9.

URL https://link.springer.com/
chapter/10.1007/978-94-011-3342-
5_6.

Grabinska:1995:ALZ

[Gra95] Teresa Grabińska. The ambiguity of the Lakatos–Zahar reconstruction of classical and relativistic programmes. In Misiek [Mis95c], pages 109–117. ISBN 94-010-4206-3, 94-011-0461-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-011-0461-
                                              6_7.

Grasshoff:1998:HPN

[Gra98] Gerd Grasshoff. Hertz’s philosophy of nature in Wittgenstein’s Tractatus. In Heinrich Hertz: Classical Physicist, Modern Philosopher [BHN98], chapter 14, pages 243–268. ISBN 90-481-4881-2, 94-015-8855-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-015-8855-3_14.

Grayling:2000:NA

[Gra00] A. C. Grayling. Naturalistic assumptions. In Knowledge, Language and Logic: Questions for Quine [OK00], chapter 4, pages 47–56. ISBN 1-4020-0253-X, 94-011-3933-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-011-3933-5_4.

Gray:2006:EPS

[Gra06] Jeremy Gray. Enriques: Popularising science and the problems of geometry. In Hendricks et al. [HJLP06], pages 135–154. ISBN 1-4020-5194-8 (hardcover), 1-4020-5195-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175 .B69 v.251 2006; Q175 .I51377 2006. URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-1-4020-5195-1_4.

Graham:2007:ISP

[Gra07a] Loren Graham. The intellectual strengths of pluralism and diversity. In Positioning the history of science [RG07], chapter 13, pages 69–71. ISBN 1-4020-5419-X (hardcover), 1-4020-5420-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q125 .P77 2007. URL https://link.

springer.com/chapter/10.1007/1-
4020-5420-3_13.

Grassmann:2007:RNE

[Gra07b] Hermann Grassmann. On the relation of non–Euclidean geometry to extension theory. In Renn et al. [RSS^{+}07b], pages 1079–1080. ISBN 1-4020-3999-9 (hardcover), 1-4020-4000-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 no. 250 v. 4; QC173.6 .G469 2007.

URL https://link.springer.com/
chapter/10.1007/978-1-4020-4000-

9_47. Originally published as Appendix 1 (1877) to “A New Branch of Mathematics: The ‘Ausdehnungslehre’ of 1844 and other works” (Chicago: Open Court, 1995), 279–280.

Greeno:1971:TES

[Gre71] James G. Greeno. Theoretical entities in statistical explanation. In Buck and Cohen [BC71], pages 3–26. ISBN 90-277-0309-4, 94-010-3142-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-010-3142-4_1.

Grene:1974:AMB

[Gre74a] Marjorie Grene. Aristotle and modern biology. In The Understanding of Nature: Essays in the Philosophy of Biology [Gre74t], chapter 5, pages 74–107. ISBN 90-277-0463-5, 94-010-2224-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 23; QH331. URL https:

//link.springer.com/chapter/10.
1007/978-94-010-2224-8_5.

Grene:1974:BT

[Gre74b] Marjorie Grene. Biology and teleology. In The Understanding of Nature: Essays in the Philosophy of Biology [Gre74t], chapter 9, pages 172–179. ISBN 90-277-0463-5, 94-010-2224-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 23; QH331. URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-010-2224-8_9.

Grene:1974:BPL

[Gre74c] Marjorie Grene. Biology and the problem of levels of reality. In The Understanding of Nature: Essays in the Philosophy of Biology [Gre74t], chapter 3, pages 35–52. ISBN 90-277-0463-5, 94-010-2224-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 23; QH331.

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-010-2224-
                                              8_3.

Grene:1974:BMB

[Gre74d] Marjorie Grene. Bohm’s metaphysics and biology. In The Understanding of Nature: Essays in the Philosophy of Biology [Gre74t], chapter 10, pages 180–188. ISBN 90-277-0463-5, 94-010-2224-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 23; QH331. URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-010-2224-8_10.

Grene:1974:C

[Gre74e] Marjorie Grene. Causes. In The Understanding of Nature: Essays in the Philosophy of Biology [Gre74t], chapter 1, pages 1–12. ISBN 90-277-0463-5, 94-010-2224-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 23; QH331.

URL https://link.springer.com/
chapter/10.1007/978-94-010-2224-
8_1.

Grene:1974:CLTa

[Gre74f] Marjorie Grene. The characters of living things. In The Understanding of Nature: Essays in the Philosophy of Biology [Gre74t], chapter 16, pages 254–293. ISBN 90-277-0463-5, 94-010-2224-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 23; QH331. URL https:

//link.springer.com/chapter/10.
1007/978-94-010-2224-8_16.

Grene:1974:CLTb

[Gre74g] Marjorie Grene. The characters of living things. In The Understanding of Nature: Essays in the Philosophy of Biology [Gre74t], chapter 17, pages 294–319. ISBN 90-277-0463-5, 94-010-2224-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 23; QH331. URL https:

//link.springer.com/chapter/10.
1007/978-94-010-2224-8_17.

Grene:1974:CLTc

[Gre74h] Marjorie Grene. The characters of living things. In The Understanding of Nature: Essays in the Philosophy of Biology [Gre74t], chapter 18, pages 320–345. ISBN 90-277-0463-5, 94-010-2224-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 23; QH331. URL https:

//link.springer.com/chapter/10.
1007/978-94-010-2224-8_18.

Grene:1974:DP

[Gre74i] Marjorie Grene. Darwin and philosophy. In The Understanding of Nature: Essays in the Philosophy of Biology [Gre74t], chapter 11, pages 189–200.

ISBN 90-277-0463-5, 94-010-2224-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 23; QH331. URL https:

//link.springer.com/chapter/10.
1007/978-94-010-2224-8_11.

Grene:1974:EAR

[Gre74j] Marjorie Grene. The ethical animal: A review. In The Understanding of Nature: Essays in the Philosophy of Biology [Gre74t], chapter 12, pages 201–206. ISBN 90-277-0463-5, 94-010-2224-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 23; QH331. URL https:

//link.springer.com/chapter/10.
1007/978-94-010-2224-8_12.

Grene:1974:EE

[Gre74k] Marjorie Grene. Explanation and evolution. In The Understanding of Nature: Essays in the Philosophy of Biology [Gre74t], chapter 13, pages 207–227. ISBN 90-277-0463-5, 94-010-2224-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 23; QH331. URL https:

//link.springer.com/chapter/10.
1007/978-94-010-2224-8_13.

Grene:1974:GSM

[Gre74l] Marjorie Grene. Is genus to species as matter to form? Aristotle and taxonomy. In The Understanding of Nature: Essays in the Philosophy of Biology [Gre74t], chapter 6, pages 108–126. ISBN 90-277-0463-5, 94-010-2224-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 23; QH331. URL https:

//link.springer.com/chapter/10.
1007/978-94-010-2224-8_6.

Grene:1974:SDB

[Gre74m] Marjorie Grene. On some distinctions between men and brutes. In The Understanding of Nature: Essays in the Philosophy of Biology [Gre74t], chapter 15, pages 243–253. ISBN 90-277-0463-5, 94-010-2224-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol.

23; QH331. URL https://link.
springer.com/chapter/10.1007/
978-94-010-2224-8_15.

Grene:1974:NNN

[Gre74n] Marjorie Grene. On the nature of natural necessity. In The Understanding of Nature: Essays in the Philosophy of Biology [Gre74t], chapter 14, pages 228–242. ISBN 90-277-0463-5, 94-010-2224-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 23; QH331. URL https:

//link.springer.com/chapter/10.
1007/978-94-010-2224-8_14.

Grene:1974:POA

[Gre74o] Marjorie Grene. People and other animals. In The Understanding of Nature: Essays in the Philosophy of Biology [Gre74t], chapter 19, pages 346–360. ISBN 90-277-0463-5, 94-010-2224-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 23; QH331. URL https:

//link.springer.com/chapter/10.
1007/978-94-010-2224-8_19.

Grene:1974:RAS

[Gre74p] Marjorie Grene. Reducibility: Another side issue? In The Understanding of Nature: Essays in the Philosophy of Biology [Gre74t], chapter 4, pages 53– 73. ISBN 90-277-0463-5, 94-010-2224-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 23; QH331. URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-010-2224-8_4.

Grene:1974:SS

[Gre74q] Marjorie Grene. Statistics and selection. In The Understanding of Nature: Essays in the Philosophy of Biology [Gre74t], chapter 8, pages 154–171. ISBN 90-277-0463-5, 94-010-2224-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 23; QH331. URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-010-2224-8_8.

Grene:1974:TAP

[Gre74r] Marjorie Grene. Three aspects of perception. In The Understanding of Nature: Essays in the Philosophy of Biology [Gre74t], chapter 2, pages 13– 34. ISBN 90-277-0463-5, 94-010-2224-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 23; QH331. URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-010-2224-8_2.

Grene:1974:TET

[Gre74s] Marjorie Grene. Two evolutionary theories. In The Understanding of Nature: Essays in the Philosophy of Biology [Gre74t], chapter 7, pages 127–153. ISBN 90-277-0463-5, 94-010-2224-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 23; QH331. URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-010-2224-8_7.

Grene:1974:UNE

[Gre74t] Marjorie Grene. The Understanding of Nature: Essays in the Philosophy of Biology, volume 23 of Boston Studies in the Philosophy of Science.

D. Reidel, Dordrecht, The Netherlands; Boston, MA, USA; Lancaster, UK; Tokyo, Japan, 1974. ISBN 90-277-0463-5, 94-010-2224-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). 378 (est.) pp. LCCN Q174 .B67 vol. 23; QH331. URL

https://link.springer.com/book/
10.1007/978-94-010-2224-8.

Gregory:1975:NIN

[Gre75] Tullio Gregory. La nouvelle idée de nature et de savoir scientifique au XIIe siècle. (French) [The new idea of nature and scientific knowledge in the 12th Century]. In Murdoch and Sylla [MS75b], pages 193–218. ISBN 90-277-0587-9, 94-010-1781-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN B721 .I584 1975.

URL https://link.springer.com/
chapter/10.1007/978-94-010-1781-
7_7.

Grene:1976:AMB

[Gre76a] Marjorie Grene. Aristotle and modern biology. In Topics in the Philosophy of Biology [GM76], chapter 1, pages 3–36. ISBN 90-277-0596-8, 94-010-1829-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-010-1829-
6_1.

Grene:1976:ILS

[Gre76b] Marjorie Grene. Imre Lakatos: Some recollections. In Essays in Memory of Imre Lakatos [CFW76], chapter 15, pages 209–212. ISBN 90-277-0655-7, 94-010-1451-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-010-
1451-9_15.

Grene:1983:EPS

[Gre83] Marjorie Grene. Empiricism and the philosophy of science or, n dogmas of empiricism. In Epistemology, Methodology, and the Social Sciences [CW83a], chapter 4, pages 89–106. ISBN 90-481-8376-6, 94-017-1458-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-017-1458-7_4.

Greene:1984:HCP

[Gre84] Murray Greene. Hegel’s conception of psychology. In Hegel and the Sciences [CW84a], chapter 11, pages 161–191. ISBN 94-009-6235-5, 94-009-6233-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-009-6233-
                                              0_11.

Grene:1986:EE

[Gre86a] Marjorie Grene. Explanation and evolution. In Cohen and Wartofsky [CW85], chapter 12, pages 177–197. ISBN 90-277-1971-3, 94-009-5345-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-009-5345-
                                              1_12.

Grene:1986:F

[Gre86b] Marjorie Grene. In and on friendship. In Cohen and Wartofsky [CW85], chapter 19, pages 355–368. ISBN 90-277-1971-3, 94-009-5345-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-009-5349-9_19.

Gregory:1991:DGT

[Gre91a] Frederick Gregory. Darwin and the German theologians. In Woodward and Cohen [WC91], pages 269–278. ISBN 94-010-5395-2, 94-011-3164-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-011-3164-3_24.

Gregory:1992:TST

[Gre91b] Frederick Gregory. Theologians, science, and theories of truth in Nineteenth-Century Germany. In The Invention of Physical Science: Intersections of Mathematics, Theology and Natural Philosophy Since the Seventeenth Century: Essays in Honor of Erwin N. Hiebert [NRS92], chapter 4, pages 81–96. ISBN 94-010-5097-X, 94-011-2488-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-011-
2488-1_4.

Gregory:1994:NOW

[Gre94] Frederick Gregory. “Nature is an organized whole”: J. F. Fries’s reformulation of Kant’s philosophy of organism. In Poggi and Bossi [PB94], chapter 6, pages 91–101. ISBN 0-7923-2336-X, 94-017-2921-2. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 152. URL https://link.

springer.com/chapter/10.1007/
978-94-017-2921-5_6.

Grene:1995:AMC

[Gre95] Marjorie Grene. Animal mechanism and the cartesian vision of nature. In Gavroglu et al. [GSW95c], chapter 10, pages 189–204. ISBN 0-7923-2991-0 (set), 0-7923-2988-0, 94-017-2658-2. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 163.

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-017-2658-
                                              0_10.

Gremmen:1999:JAS

[Gre99] Bart Gremmen. The justification of the application of science. In Fehér et al. [FKR99], pages 315–323. ISBN 90-481-5257-7, 94-015-9293-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-015-9293-2_26.

Gregory:2007:HCO

[Gre07] Frederick Gregory. Hans Christian Ørsted’s spiritual interpretation of natural science. In Hans Christian Ørsted and The Romantic Legacy In Science: Ideas, Disciplines, Practices [BCK07], chapter 18, pages 399–416. ISBN 1-4020-2979-9 (hardcover), 1-4020-2987-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC515.O37 H36 2007. URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-1-4020-2987-5_18.

Griesemer:2015:WSB

[Gri15] James R. Griesemer. What salamander biologists have taught us about Evodevo. In Conceptual Change in Biology: Scientific and Philosophical Perspectives on Evolution and Development [Lov15b], chapter 13, pages 271–301. ISBN 94-017-9411-1, 94-017-9412-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QH331 C66 2015; QH491; QH492 .C66 2014.

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-017-9412-
                                              1_13.

Grlic:1979:RT

[Grl79] Danko Grlić. Revolution and terror.

In Praxis: Yugoslav Essays in the Philosophy and Methodology of the Social Sciences [MP79], chapter 8, pages 139–150. ISBN 90-277-0968-8, 94-009-9355-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-009-
9355-6_8.

Grmek:1981:PFH

[Grm81] Mirko D. Grmek. A plea for freeing the history of scientific discoveries from myth. In Grmek et al. [GCC81b], pages 9–42. ISBN 90-277-1123-2, 94-010-1284-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-010-
1284-3_2.

Grobstein:1976:OLE

[Gro76]  Clifford Grobstein.  Organizational

levels and explanation. In Topics in the Philosophy of Biology [GM76], chapter 6, pages 145–152. ISBN 90-277-0596-8, 94-010-1829-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-010-1829-6_6.

Grossberg:1982:APC

[Gro82a] Stephen Grossberg. Adaptive pattern classification and universal recoding I: Parallel development and coding of neural feature detectors. In Studies of Mind and Brain: Neural Principles of Learning, Perception, Development, Cognition, and Motor Control [Gro82m], chapter 12, pages 448–497. ISBN 94-009-7758-1 (e-book), 90-277-1360-X. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-009-7758-
7_12.

Grossberg:1982:BCS

[Gro82b] Stephen Grossberg. Behavioral contrast in short term memory: Serial binary memory models or parallel continuous memory models? In Studies of Mind and Brain: Neural Principles of Learning, Perception, Development, Cognition, and Motor Control [Gro82m], chapter 11, pages 425–447. ISBN 94-009-7758-1 (e-book), 90-277-1360-X. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-009-7758-
7_11.

Grossberg:1982:BCD

[Gro82c] Stephen Grossberg. Biological competition: Decision rules, pattern formation, and oscillations. In Studies of Mind and Brain: Neural Principles of Learning, Perception, Development, Cognition, and Motor Control [Gro82m], chapter 9, pages 379–398. ISBN 94-009-7758-1 (ebook), 90-277-1360-X. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-009-
7758-7_9.

Grossberg:1982:CIL

[Gro82d] Stephen Grossberg. Classical and instrumental learning by neural networks. In Studies of Mind and Brain: Neural Principles of Learning, Perception, Development, Cognition, and Motor Control [Gro82m], chapter 3, pages 65–156. ISBN 94-009-7758-1 (e-book), 90-277-1360-X. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-009-7758-
7_3.

Grossberg:1982:CDC

[Gro82e]  Stephen Grossberg.  Competition,

decision, and consensus. In Studies of Mind and Brain: Neural Principles of Learning, Perception, Development, Cognition, and Motor Control [Gro82m], chapter 10, pages 399–424. ISBN 94-009-7758-1 (e-book), 90-277-1360-X. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-009-7758-
7_10.

Grossberg:1982:CES

[Gro82f]  Stephen Grossberg.  Contour en-

hancement, short term memory, and constancies in reverberating neural networks. In Studies of Mind and Brain: Neural Principles of Learning, Perception, Development, Cognition, and Motor Control [Gro82m], chapter 8, pages 332–378. ISBN 94-009-7758-1 (ebook), 90-277-1360-X. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-009-
7758-7_8.

Grossberg:1982:HDB

[Gro82g] Stephen Grossberg. How does a brain build a cognitive code? In Studies of Mind and Brain: Neural Principles of Learning, Perception, Development, Cognition, and Motor Control [Gro82m], chapter 1, pages 1–52. ISBN 94-009-7758-1 (e-book), 90-277-1360-X. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-009-7758-7_1.

Grossberg:1982:NEC

[Gro82h] Stephen Grossberg. Neural expectation: Cerebellar and retinal analogs of cells fired by learnable or unlearned pattern classes. In Studies of Mind and Brain: Neural Principles of Learning, Perception, Development, Cognition, and Motor Control [Gro82m], chapter 7, pages 296–331. ISBN 94-009-7758-1 (ebook), 90-277-1360-X. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

                                              ???? URL https://link.springer.
                                              com/chapter/10.1007/978-94-009-
                                              7758-7_7.

Grossberg:1982:NMA

[Gro82i] Stephen Grossberg. A neural model of attention, reinforcement and discrimination learning. In Studies of Mind and Brain: Neural Principles of Learning, Perception, Development, Cognition, and Motor Control [Gro82m], chapter 6, pages 229–295. ISBN 94-009-7758-1 (ebook), 90-277-1360-X. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

                                              ???? URL https://link.springer.
                                              com/chapter/10.1007/978-94-009-
                                              7758-7_6.

Grossberg:1982:NTP

[Gro82j] Stephen Grossberg. A neural theory of punishment and avoidance, II: Quantitative theory. In Studies of Mind and Brain: Neural Principles of Learning, Perception, Development, Cognition, and Motor Control [Gro82m], chapter 5, pages 194–228. ISBN 94-009-7758-1 (ebook), 90-277-1360-X. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

                                              ???? URL https://link.springer.
                                              com/chapter/10.1007/978-94-009-
                                              7758-7_5.

Grossberg:1982:PLF

[Gro82k] Stephen Grossberg. Pattern learning by functional–differential neural networks with arbitrary path weights. In Studies of Mind and Brain: Neural Principles of Learning, Perception, Development, Cognition, and Motor Control [Gro82m], chapter 4, pages 157–193. ISBN 94-009-7758-1 (e-book), 90-277-1360-X. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-009-7758-
7_4.

Grossberg:1982:SPB

[Gro82l] Stephen Grossberg. Some physiological and biochemical consequences of psychological postulates. In Studies of Mind and Brain: Neural Principles of Learning, Perception, Development, Cognition, and Motor Control [Gro82m], chapter 2, pages 53–64. ISBN 94-009-7758-1 (e-book), 90-277-1360-X. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-009-7758-7_2.

Grossberg:1982:SMB

[Gro82m] Stephen Grossberg. Studies of Mind and Brain: Neural Principles of Learning, Perception, Development, Cognition, and Motor Control, volume 70 of Boston Studies in the Philosophy of Science. D. Reidel, Dordrecht, The Netherlands; Boston, MA, USA; Lancaster, UK; Tokyo, Japan, 1982. ISBN 94-009-7758-1 (e-book), 90-277-1360-X. ISSN 0068-0346 (print), 2214-7942 (electronic). 660 (est.) pp. LCCN ????

URL https://link.springer.com/
book/10.1007/978-94-009-7758-7.

Grossberg:1982:THM

[Gro82n] Stephen Grossberg. A theory of human memory: Self-organization and performance of sensory-motor codes, maps, and plans. In Studies of Mind and Brain: Neural Principles of Learning, Perception, Development, Cognition, and Motor Control [Gro82m], chapter 13, pages 498–639. ISBN 94-009-7758-1 (e-book), 90-277-1360-X. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-009-7758-7_13.

Grossberg:1986:BCD

[Gro86] Stephen Grossberg. Biological competition: Decision rules, pattern formation, and oscillations. In Cohen and Wartofsky [CW85], chapter 18, pages 257–276. ISBN 90-277-1971-3, 94-009-5345-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-009-5345-
                                              1_18.

Grodin:1995:MME

[Gro95] Michael A. Grodin, editor. Meta Medical Ethics: the Philosophical Foundations of Bioethics, volume 171 of Boston Studies in the Philosophy of Science. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 1995. ISBN 0-7923-3344-6. xiii + 205 pp. LCCN Q174 .B67 vol. 171; R724 .M47 1995.

Grosholz:2001:FTM

[Gro01] Emily Grosholz. Fedoroff’s translation of McClintock: The uses of chemistry in the reorganization of genetics. In Klein [Kle01b], chapter 12, pages 199–219. ISBN 1-4020-0100-2, 94-015-9737-5. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 222;

                                              QD7. URL https://link.springer.
                                              com/chapter/10.1007/978-94-015-
                                              9737-1_12.

Grossmann:2009:ATM

[Gro09a] Henryk Grossmann. Additional texts on mechanism. In Freudenthal and McLaughlin [FM09c], chapter 5, pages 231–237. ISBN 1-4020-9603-8, 1-4020-9604-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174.6 .S63 2009. URL https://

link.springer.com/chapter/10.
1007/978-1-4020-9604-4_5.

Grossmann:2009:DSO

[Gro09b] Henryk Grossmann. Descartes and the social origins of the mechanistic concept of the world. In Freudenthal and McLaughlin [FM09c], chapter 4, pages 157–229. ISBN 1-4020-9603-8, 1-4020-9604-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174.6 .S63 2009. URL https:

//link.springer.com/chapter/10.
1007/978-1-4020-9604-4_4.

Grossmann:2009:SFM

[Gro09c] Henryk Grossmann. The social foundations of the mechanistic philosophy and manufacture. In Freudenthal and McLaughlin [FM09c], chapter 3, pages 103–156. ISBN 1-4020-9603-8, 1-4020-9604-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174.6 .S63 2009. URL https:

//link.springer.com/chapter/10.
1007/978-1-4020-9604-4_3.

Grunbaum:1963:FTT

[Grü63] Adolf Grünbaum. The falsifiability of theories: Total or partial? A contemporary evaluation of the Duhem–Quine thesis. In Wartofsky [War63], pages 178–195. ISBN 94-010-3265-3, 94-010-3263-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B678 1970. URL https://link.

                                              springer.com/chapter/10.1007/
                                              978-94-010-3263-6_14.

Grunbaum:1969:RHP

[Grü69a] Adolf Grünbaum. Reply to Hilary Putnam’s ‘An Examination of Grünbaum’s Philosophy of Geometry’. In Cohen and Wartofsky [CW69a], pages 1–150. ISBN 94-010-3383-8, 94-010-3263-7, 94-010-3381-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

                                              ???? URL https://link.springer.
                                              com/chapter/10.1007/978-94-010-
                                              3381-7_1.

Grunstra:1969:DTM

[Gru69b] Bernard R. Grunstra. On distinguishing types of measurement. In Cohen and Wartofsky [CW69a], pages 253–303. ISBN 94-010-3383-8, 94-010-3263-7, 94-010-3381-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

                                              ???? URL https://link.springer.
                                              com/chapter/10.1007/978-94-010-
                                              3381-7_7.

Grunbaum:1973:AT

[Grü73a] Adolf Grünbaum. The anisotropy of time. In Philosophical Problems of Space and Time [Grü73s], chapter 8, pages 209–280. ISBN 90-277-0358-2, 94-010-2622-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

                                              ???? URL https://link.springer.
                                              com/chapter/10.1007/978-94-010-
                                              2622-2_8.

Grunbaum:1973:ARP

[Grü73b] Adolf Grünbaum. The asymmetry of retrodictability and predictability, the compossibility of explanation of the past and prediction of the future, and mechanism vs. teleology. In Philosophical Problems of Space and Time [Grü73s], chapter 9, pages 281–313.

ISBN 90-277-0358-2, 94-010-2622-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-010-2622-
2_9.

Grunbaum:1973:BPH

[Grü73c] Adolf Grünbaum. The bearing of philosophy on the history of the special theory of relativity. In Philosophical Problems of Space and Time [Grü73s], chapter 21, pages 709–727. ISBN 90-277-0358-2, 94-010-2622-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-010-2622-2_21.

Grunbaum:1973:CIO

[Grü73d] Adolf Grünbaum. Can an infinitude of operations be performed in a finite time? In Philosophical Problems of Space and Time [Grü73s], chapter 18, pages 630–645. ISBN 90-277-0358-2, 94-010-2622-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-010-
2622-2_18.

Grunbaum:1973:CWA

[Grü73e] Adolf Grünbaum. Can we ascertain the falsity of a scientific hypothesis? In Philosophical Problems of Space and Time [Grü73s], chapter 17, pages 569–629. ISBN 90-277-0358-2, 94-010-2622-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-010-
2622-2_17.

Grunbaum:1973:CTT

[Grü73f] Adolf Grünbaum. The causal theory of time. In Philosophical Problems of Space and Time [Grü73s], chapter 7, pages 179–208. ISBN 90-277-0358-2, 94-010-2622-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

                                              ???? URL https://link.springer.
                                              com/chapter/10.1007/978-94-010-
                                              2622-2_7.

Grunbaum:1973:CEP

[Grü73g] Adolf Grünbaum. Critique of Einstein’s philosophy of geometry. In Philosophical Problems of Space and Time [Grü73s], chapter 4, pages 106–151. ISBN 90-277-0358-2, 94-010-2622-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-010-2622-
                                              2_4.

Grunbaum:1973:CRC

[Grü73h] Adolf Grünbaum. Critique of Reichenbach’s and Carnap’s philosophy of geometry. In Philosophical Problems of Space and Time [Grü73s], chapter 3, pages 81–105. ISBN 90-277-0358-2, 94-010-2622-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

                                              ???? URL https://link.springer.
                                              com/chapter/10.1007/978-94-010-
                                              2622-2_3.

Grunbaum:1973:EGV

[Grü73i] Adolf Grünbaum. Empiricism and the geometry of visual space. In Philosophical Problems of Space and Time [Grü73s], chapter 5, pages 152–157. ISBN 90-277-0358-2, 94-010-2622-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-010-2622-
                                              2_5.

Grunbaum:1973:ETD

[Grü73j] Adolf Grünbaum. Empiricism and the three-dimensionality of space. In Philosophical Problems of Space and Time [Grü73s], chapter 11, pages 330–337. ISBN 90-277-0358-2, 94-010-2622-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-010-2622-
2_11.

Grunbaum:1973:GRG

[Grü73k] Adolf Grünbaum. General relativity, geometrodynamics and ontology. In Philosophical Problems of Space and Time [Grü73s], chapter 22, pages 728–803. ISBN 90-277-0358-2, 94-010-2622-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-010-
2622-2_22.

Grunbaum:1973:GTR

[Grü73l] Adolf Grünbaum. Has the general theory of relativity repudiated absolute space? In Philosophical Problems of Space and Time [Grü73s], chapter 14, pages 418–424. ISBN 90-277-0358-2, 94-010-2622-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-010-
2622-2_14.

Grunbaum:1973:CGE

[Grü73m] Adolf Grünbaum. Is the coarse-grained entropy of classical statistical mechanics an anthropomorphism? In Philosophical Problems of Space and Time [Grü73s], chapter 19, pages 646–665. ISBN 90-277-0358-2, 94-010-2622-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

                                              ???? URL https://link.springer.
                                              com/chapter/10.1007/978-94-010-
                                              2622-2_19.

Grunbaum:1973:TFT

[Grü73n] Adolf Grünbaum. Is there a “flow” of time or temporal “becoming”? In Philosophical Problems of Space and Time [Grü73s], chapter 10, pages 314–329. ISBN 90-277-0358-2, 94-010-2622-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

                                              ???? URL https://link.springer.
                                              com/chapter/10.1007/978-94-010-
                                              2622-2_10.

Grunbaum:1973:PDS

[Grü73o] Adolf Grünbaum. A panel discussion of simultaneity by slow clock transport in the special and general theories of relativity. In Philosophical Problems of Space and Time [Grü73s], chapter 20, pages 666–708. ISBN 90-277-0358-2, 94-010-2622-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

                                              ???? URL https://link.springer.
                                              com/chapter/10.1007/978-94-010-
                                              2622-2_20.

Grunbaum:1973:PAM

[Grü73p] Adolf Grünbaum. Philosophical appraisal of E. A. Milne’s alternative to Einstein’s STR. In Philosophical Problems of Space and Time [Grü73s], chapter 13, pages 410–417. ISBN 90-277-0358-2, 94-010-2622-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-010-2622-2_13.

Grunbaum:1973:PCW

[Grü73q] Adolf Grünbaum. Philosophical critique of Whitehead’s theory of relativity. In Philosophical Problems of Space and Time [Grü73s], chapter 15, pages 425–428. ISBN 90-277-0358-2, 94-010-2622-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-010-
2622-2_15.

Grunbaum:1973:PFS

[Grü73r]  Adolf Grünbaum.  Philosophical

foundations of the special theory of relativity, and their bearing on its history. In Philosophical Problems of Space and Time [Grü73s], chapter 12, pages 341–409. ISBN 90-277-0358-2, 94-010-2622-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-010-
2622-2_12.

Grunbaum:1973:PPS

[Grü73s] Adolf Grünbaum. Philosophical Problems of Space and Time, volume 12 of Boston Studies in the Philosophy of Science. D. Reidel, Dordrecht, The Netherlands; Boston, MA, USA; Lancaster, UK; Tokyo, Japan, second, enlarged edition, 1973. ISBN 90-277-0358-2, 94-010-2622-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). xxiv + 884 pp. LCCN ????

URL https://link.springer.com/
book/10.1007/978-94-010-2622-2.

Grunbaum:1973:RZM

[Grü73t] Adolf Grünbaum. The resolution of Zeno’s metrical paradox of extension for the mathematical continua of space and time. In Philosophical Problems of Space and Time [Grü73s], chapter 6, pages 158–176. ISBN 90-277-0358-2, 94-010-2622-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.

                                              com/chapter/10.1007/978-94-010-
                                              2622-2_6.

Grunbaum:1973:SAT

[Grü73u] Adolf Grünbaum. The significance of alternative time metrizations in Newtonian mechanics and in the general theory of relativity. In Philosophical Problems of Space and Time [Grü73s], chapter 2, pages 66–80. ISBN 90-277-0358-2, 94-010-2622-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-010-2622-2_2.

Grunbaum:1973:STF

[Grü73v] Adolf Grünbaum. Space, time and falsifiability critical exposition and reply to “A Panel Discussion of Grünbaum’s Philosophy of Science”. In Philosophical Problems of Space and Time [Grü73s], chapter 16, pages 449–568. ISBN 90-277-0358-2, 94-010-2622-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-010-2622-2_16.

Grunbaum:1973:STC

[Grü73w] Adolf Grünbaum. Spatial and temporal congruence in physics: A critical comparison of the conceptions of Newton, Riemann, Poincaré, Eddington, Bridgman, Russell, and Whitehead. In Philosophical Problems of Space and Time [Grü73s], chapter 1, pages 3–65. ISBN 90-277-0358-2, 94-010-2622-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-010-2622-
                                              2_1.

Gruber:1974:AMM

[Gru74a] Jacob W. Gruber. Acquired models and the modification of anthropological evidence. In Seeger and Cohen [SC74b], pages 325–336. ISBN 90-277-0390-6, 90-277-0376-0 (paperback), 94-010-2126-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 11. URL https:/

/link.springer.com/chapter/10.
1007/978-94-010-2126-5_18.

Grunbaum:1974:STM

[Grü74b] Adolf Grünbaum. Space, time and matter: The foundations of geometrodynamics. In Schaffner and Cohen [SC74a], pages 3–5. ISBN 90-277-0408-2, 90-277-0409-0, 94-010-2140-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175 .B7312.

URL https://link.springer.com/
chapter/10.1007/978-94-010-2140-
1_1.

Grunbaum:1976:FTS

[Grü76] Adolf Grünbaum. Is falsifiability the touchstone of scientific rationality? Karl Popper versus inductivism. In Essays in Memory of Imre Lakatos [CFW76], chapter 16, pages 213–252. ISBN 90-277-0655-7, 94-010-1451-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-010-1451-
9_16.

Grunbaum:1978:PVI

[Grü78] Adolf Grünbaum. Popper vs inductivism. In Progress and Rationality in Science [RA78b], chapter 6, pages 117–142. ISBN 90-277-0922-X, 94-009-9866-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

                                              ???? URL https://link.springer.
                                              com/chapter/10.1007/978-94-009-
                                              9866-7_6.

Gruber:1980:ESA

[Gru80] Howard E. Gruber. The evolving systems approach to creative scientific work: Charles Darwin’s early thought. In Scientific Discovery: Case Studies [Nic80e], chapter 7, pages 113–130. ISBN 90-277-1093-7, 94-009-9015-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 60. URL https://link.

                                              springer.com/chapter/10.1007/
                                              978-94-009-9015-9_7.

Gruber:1981:CPS

[Gru81] Howard E. Gruber. Cognitive psychology, scientific creativity, and the case study method. In Grmek et al. [GCC81b], pages 295–322. ISBN 90-277-1123-2, 94-010-1284-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-010-1284-3_15.

Grunbaum:1984:EIP

[Grü84] Adolf Grünbaum. Explication and implications of the placebo concept. In Rationality in Science and Politics [And84d], chapter 8, pages 131–158. ISBN 90-277-1953-5, 94-009-6254-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-009-6254-
                                              5_8.

Grunbaum:1986:FTT

[Grü86] Adolf Grünbaum. The falsifiability of theories: Total or partial? A contemporary evaluation of the Duhem–Quine thesis. In Cohen and Wartofsky [CW85], chapter 1, pages 1–18. ISBN 90-277-1971-3, 94-009-5345-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-009-5345-1_1.

Grunbaum:1989:DPT

[Grü89a] Adolf Grünbaum. The degeneration of Popper’s theory of demarcation. In Freedom and Rationality: Essays in Honor of John Watkins From his Colleagues and Friends [DJ89], chapter 9, pages 141–161. ISBN 94-010-7571-9, 94-009-2380-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-009-
2380-5_9.

Grunbaum:1989:WTK

[Grü89b] Adolf Grünbaum. Why thematic kinships between events do not attest their causal linkage. In An Intimate Relation: Studies in the History and Philosophy of Science Presented to Robert E. Butts on his 60th Birthday [BM89], chapter 24, pages 477–494. ISBN 94-010-7546-8, 94-009-2327-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-009-2327-
0_24.

Grunbaum:1995:PTM

[Grü95a] Adolf Grünbaum. The poverty of theistic morality. In Science, Mind and Art: Essays on science and the humanistic understanding in art, epistemology, religion and ethics In honor of Robert S. Cohen [GSW95b], chapter 13, pages 203–242. ISBN 0-7923-2990-2 (hardcover), 94-011-0469-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol.

165 BD418.3. URL https://link.

                                              springer.com/chapter/10.1007/
                                              978-94-011-0469-2_13.

Grunberg:1995:LRC

[Grü95b] Teo Grünberg. Long run consistency of beliefs as criterion of empirical knowledge. In The Concept of Knowledge: The Ankara Seminar [KC95], chapter 12, pages 149–163. ISBN 90-481-4495-7, 94-017-3263-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-017-3263-5_12.

Grunberg:2005:CTH

[Grü05a] David Grünberg. Confirmation of theoretical hypotheses: Bootstrapping with a Bayesian face. In Turkish Studies in the History and Philosophy of Science [IG05c], chapter 7, pages 87–101. ISBN 1-4020-3332-X (hardcover), 1-4020-3333-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 244.

                                              URL https://link.springer.com/
                                              chapter/10.1007/1-4020-3333-8_7.

Grunberg:2005:DLC

[Grü05b] Teo Grünberg. Demarcation of the logical constants and logical truth in terms of analyticity. In Turkish Studies in the History and Philosophy of Science [IG05c], chapter 3, pages 27–38. ISBN 1-4020-3332-X (hardcover), 1-4020-3333-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 244.

                                              URL https://link.springer.com/
                                              chapter/10.1007/1-4020-3333-8_3.

Grunbaum:2009:PFM

[Grü09] Adolf Grünbaum. Popper’s fundamental misdiagnosis of the scientific defects of Freudian psychoanalysis. In Rethinking Popper [PC09], chapter 10, pages 117–134. ISBN 1-4020-9337-3, 1-4020-9338-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN B1649.P64 R48 2009; Q174 .B67 2009.

URL https://link.springer.com/
chapter/10.1007/978-1-4020-9338-
8_10.

Garcia-Sucre:1982:KCS

[GS82] Máximo García-Sucre. A kind of collapse in a simple spacetime model. In Scientific Philosophy Today: Essays in Honor of Mario Bunge [AC82], chapter 2, pages 45–69. ISBN 90-277-1262-X, 90-277-1263-8, 94-009-8462-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 67. URL https://link.

springer.com/chapter/10.1007/
978-94-009-8462-2_2.

Godfrey-Smith:2009:AIE

[GS09] Peter Godfrey-Smith. Abstractions, idealizations, and evolutionary biology. In Mapping the Future of Biology: Evolving Concepts and Theories [BMP09b], chapter 4, pages 47–56. ISBN 1-4020-9635-6 (hardcover), 1-4020-9636-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QH331 .M295 2009. URL https:/

/link.springer.com/chapter/10.
1007/978-1-4020-9636-5_4.

Gavroglou:1995:SPS

[GSW95a] Kostas Gavroglou, John J. Stachel, and Marx W. Wartofsky, editors. Science, politics, and social practice: essays on Marxism and science, philosophy of culture and the social sciences: in honor of Robert S. Cohen, volume 164 of Boston Studies in the Philosophy of Science. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 1995. ISBN 0-7923-2989-9, 0-7923-2991-0 (series). ix + 395 pp. LCCN Q174 .B67 vol. 164; Q175

.S4225 1994. URL http://www.loc.
gov/catdir/enhancements/fy0823/
94003703-d.html;   http://www.
loc.gov/catdir/enhancements/
fy0823/94003703-t.html.

Gavroglu:1995:SMA

[GSW95b] Kostas Gavroglu, John Stachel, and Marx W. Wartofsky. Science, Mind and Art: Essays on science and the humanistic understanding in art, epistemology, religion and ethics In honor of Robert S. Cohen, volume 165 of Boston Studies in the Philosophy of Science. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 1995. ISBN 0-7923-2990-2 (hardcover), 94-011-0469-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). ix + 462 pp. LCCN Q174 .B67 vol. 165 BD418.3.

URL https://link.springer.com/
book/10.1007/978-94-011-0469-2.

Gavroglu:1995:PPS

[GSW95c] Kostas Gavroglu, John J. Stachel, and Marx W. Wartofsky, editors. Physics, Philosophy, and the Scientific Community: Essays in the Philosophy and History of the Natural Sciences and Mathematics in Honor of Robert S. Cohen, volume 163 of Boston Studies in the Philosophy of Science. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 1995. ISBN 0-7923-2991-0 (set), 0-7923-2988-0, 94-017-2658-2. ISSN 0068-0346 (print), 2214-7942 (electronic). xxvii + 383 pp. LCCN Q174 .B67 vol. 163. URL

http://catdir.loc.gov/catdir/
enhancements/fy0823/94022250-
d.html;   http://catdir.loc.

gov/catdir/enhancements/fy0823/
94022250-t.html;   https://link.
springer.com/book/10.1007/978-
94-017-2658-0.

Guangbi:1996:BCM

[Gua96a] Dong Guangbi. The book of changes and mathematics. In Chinese Studies in the History and Philosophy of Science and Technology [DC96], chapter 12, pages 125–135. ISBN 90-481-4546-5, 94-015-8717-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 180; U390 .N19

1996. URL https://link.springer.
com/chapter/10.1007/978-94-015-
8717-4_12.

Guangyuan:1996:EM

[Gua96b] Yu Guangyuan. On the emancipation of the mind. In Chinese Studies in the History and Philosophy of Science and Technology [DC96], chapter 1, pages 1–3. ISBN 90-481-4546-5, 94-015-8717-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 180; U390 .N19 1996.

URL https://link.springer.com/
chapter/10.1007/978-94-015-8717-
4_1.

Guangyuan:1996:SQG

[Gua96c] Yu Guangyuan. Speeches at the Qingdao Genetics Conference of 1956. In Chinese Studies in the History and Philosophy of Science and Technology [DC96], chapter 4, pages 27–40. ISBN 90-481-4546-5, 94-015-8717-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 180; U390 .N19 1996. URL https:

//link.springer.com/chapter/10.
1007/978-94-015-8717-4_4.

Guccione:1981:QLT

[Guc81] Salvatore Guccione. Quantum logic and the two-slit experiment. In Italian Studies in the Philosophy of Science [Chi81b], chapter 15, pages 237–247. ISBN 90-277-1073-2, 94-009-8937-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-009-8937-
                                              5_15.

Guck:1987:RS

[Guc87] Thomas F. Guck. Relativity in Spain. In Glick [Gli87a], chapter 7, pages 231–263. ISBN 90-277-2498-9, 94-009-3875-6. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 103; QC173.5. US$79.00

                                              (US). URL https://link.springer.
                                              com/chapter/10.1007/978-94-009-
                                              3875-5_7.

Gunergun:1992:DGZ

[Gün92] Feza Günergun. Du ,,zira“au ,,mètre“: Une transformation métrologique dans l’Empire Ottoman. (French) [from “zira” to “meter”: A metrological transformation in the Ottoman Empire]. In Science and Empires: Historical Studies about Scientific Development and European Expansion [PJM92], chapter 12, pages 103–110. ISBN 0-7923-1518-9 (hardcover), 94-010-5145-3, 94-011-2594-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 136.

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-011-2594-
                                              9_12.

Gundersen:1997:NA

[Gun97] Lars Gundersen. On now-ambiguities. In Perspectives on Time [FSU97b], chapter 3, pages 93–105. ISBN 90-481-4774-3, 94-015-8875-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol.189.

URL https://link.springer.com/
chapter/10.1007/978-94-015-8875-
1_3.

Gunergun:2011:OAC

[Gün11] Feza Günergun. The Ottoman Ambassador’s curiosity coffer: Eclipse prediction with De La Hire’s “machine” crafted by Bion of Paris. In Science between Europe and Asia: Historical Studies on the Transmission, Adoption and Adaptation of Knowledge [GR11b], chapter 7, pages 103–123. ISBN 90-481-9968-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q125 .S354 2010; Q175.52.A78 S35

2010. URL https://link.springer.
com/chapter/10.1007/978-90-481-
9968-6_7.

Gurova:2003:PSM

[Gur03] Lilia Gurova. Philosophy of science meets cognitive science: The categorization debate. In Bulgarian Studies in the Philosophy of Science [Gin03a], chapter 10, pages 141–162. ISBN 1-4020-1496-1, 94-017-0961-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-017-0961-3_10.

Gutting:1980:LI

[Gut80] Gary Gutting. The logic of invention.

In Scientific Discovery, Logic, and Rationality [Nic80f], chapter 9, pages 221–234. ISBN 90-277-1070-8, 94-009-8986-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-009-
8986-3_9.

Gutierrez:1995:GBI

[Gut95]  Rosaura Ruiz Gutierrez.  Gabino

Barreda and the introduction of Darwinism in Mexico: Positivism and evolution. In Mexican Studies in the History and Philosophy of Science [RC95], chapter 3, pages 29–48. ISBN 94-010-6535-7 (hardcover), 94-009-0109-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

                                              ???? URL https://link.springer.
                                              com/chapter/10.1007/978-94-009-
                                              0109-4_3.

Guyer:2007:NEE

[Guy07] Paul Guyer. Natural ends and the end of nature. In Hans Christian Ørsted and The Romantic Legacy In Science: Ideas, Disciplines, Practices [BCK07], chapter 5, pages 75–96. ISBN 1-4020-2979-9 (hardcover), 1-4020-2987-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC515.O37 H36 2007. URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-1-4020-2987-5_5.

Ghirardi:1997:IWA

[GW97] G. C. Ghirardi and T. Weber. An interpretation which is appropriate for dynamical reduction theories. In Cohen et al. [CHS97b], pages 89–104. ISBN 90-481-4809-X, 94-017-2732-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC173.96-174.52.

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-017-2732-
                                              7_7.

Gourko:2000:AD

[GWT00a] Helena Gourko, Donald I. Williamson, and Alfred I. Tauber. Anthropology and Darwinism (1875). In The Evolutionary-biology Papers of Elie Metchnikoff [GWT00e], chapter 4, pages 36–59. ISBN 0-7923-6067-2, 94-015-9381-7. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 212; QH366.2. URL https:

//link.springer.com/chapter/10.
1007/978-94-015-9381-6_4.

Gourko:2000:CES

[GWT00b] Helena Gourko, Donald I. Williamson, and Alfred I. Tauber. Comparative embryological studies (1881–1885). In The Evolutionary-biology Papers of Elie Metchnikoff [GWT00e], chapter 6, pages 92–143. ISBN 0-7923-6067-2, 94-015-9381-7. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 212; QH366.2. URL https:

//link.springer.com/chapter/10.
1007/978-94-015-9381-6_6.

Gourko:2000:ESM

[GWT00c] Helena Gourko, Donald I. Williamson, and Alfred I. Tauber. Embryological studies on Medusae: On the evolutionary origins of the primary tissues (1886). In The Evolutionary-biology Papers of Elie Metchnikoff [GWT00e], chapter 7, pages 144–206. ISBN 0-7923-6067-2, 94-015-9381-7. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 212; QH366.2. URL https:

//link.springer.com/chapter/10.
1007/978-94-015-9381-6_7.

Gourko:2000:EQA

[GWT00d] Helena Gourko, Donald I. Williamson, and Alfred I. Tauber. Essay on questions about the origin of species (1876). In The Evolutionary-biology Papers of Elie Metchnikoff [GWT00e], chapter 5, pages 60–91. ISBN 0-7923-6067-2, 94-015-9381-7. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 212; QH366.2. URL https:

//link.springer.com/chapter/10.
1007/978-94-015-9381-6_5.

Gourko:2000:EBP

[GWT00e] Helena Gourko, Donald I. Williamson, and Alfred I. Tauber. The Evolutionarybiology Papers of Elie Metchnikoff, volume 212 of Boston Studies in the Philosophy of Science. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 2000. ISBN 0-7923-6067-2, 94-015-9381-7. ISSN 0068-0346 (print), 2214-7942 (electronic). x + 221 pp. LCCN Q174 .B67 vol. 212;

                                              QH366.2.  URL http://www.loc.
                                              gov/catdir/enhancements/fy0821/
                                              99049680-d.html;   http://www.
                                              loc.gov/catdir/enhancements/
                                              fy0821/99049680-t.html;   https:
                                              //link.springer.com/book/10.
                                              1007/978-94-015-9381-6.

Gourko:2000:I

[GWT00f] Helena Gourko, Donald I. Williamson, and Alfred I. Tauber. Introduction. In The Evolutionary-biology Papers of Elie Metchnikoff [GWT00e], chapter 1, pages 1–21. ISBN 0-7923-6067-2, 94-015-9381-7. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 212; QH366.2. URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-015-9381-6_1.

Gourko:2000:DLH

[GWT00g] Helena Gourko, Donald I. Williamson, and Alfred I. Tauber. On the developmental life-history of Myzostomum (1866). In The Evolutionary-biology Papers of Elie Metchnikoff [GWT00e], chapter 2, pages 22–27. ISBN 0-7923-6067-2, 94-015-9381-7. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 212; QH366.2.

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-015-9381-
                                              6_2.

Gourko:2000:EDL

[GWT00h] Helena Gourko, Donald I. Williamson, and Alfred I. Tauber. On the embryonic development of lower crustacea (1866). In The Evolutionary-biology Papers of Elie Metchnikoff [GWT00e], chapter 3, pages 28–35. ISBN 0-7923-6067-2, 94-015-9381-7. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 212; QH366.2. URL https:

//link.springer.com/chapter/10.
1007/978-94-015-9381-6_3.

Gourko:2000:SEB

[GWT00i] Helena Gourko, Donald I. Williamson, and Alfred I. Tauber. The struggle for existence between parts of the animal organism (1892). In The Evolutionarybiology Papers of Elie Metchnikoff [GWT00e], chapter 8, pages 207–216. ISBN 0-7923-6067-2, 94-015-9381-7. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 212; QH366.2. URL https://link.

springer.com/chapter/10.1007/
978-94-015-9381-6_8.

Haase:1999:NCT

[Haa99] Ullrich Michael Haase. Nietzsche’s critique of technology: A defence of phenomenology against modern machinery. In Nietzsche, Epistemology, and Philosophy of Science: Nietzsche and the Sciences II [Bab99a], chapter 26, pages 331–339. ISBN 90-481-5234-8, 94-017-2428-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-017-2428-9_26.

Habermas:1999:NTK

[Hab99] Jürgen Habermas. On Nietzsche’s theory of knowledge: A postscript from 1968. In Nietzsche, Theories of Knowledge, and Critical Theory: Nietzsche and the Sciences I [Bab99b], chapter 15, pages 209–223. ISBN 90-481-5233-X, 94-017-2430-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-017-2430-2_15.

Hacking:1986:ISP

[Hac86] Ian Hacking. The invention of split personalities. In Cohen and Wartofsky [CW85], chapter 4, pages 63–85. ISBN 90-277-1971-3, 94-009-5345-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-009-5349-9_4.

Hacking:1991:SCC

[Hac91] Ian Hacking. Speculation, calculation and the creation of phenomena. In Beyond Reason: Essays on the Philosophy of Paul Feyerabend [Mun91a], chapter 5, pages 131–157. ISBN 94-010-5406-1, 94-011-3188-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-011-3188-9_5.

Hacking:1994:HNS

[Hac94a] Ian Hacking. How numerical sociology began by counting suicides: From medical pathology to social pathology. In Cohen [Coh94b], chapter 2, pages 101–133. ISBN 0-7923-2223-1, 94-017-3391-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 150. URL https:/

                                              /link.springer.com/chapter/10.
                                              1007/978-94-017-3391-5_2.

Hacking:1994:SST

[Hac94b] Ian Hacking. Styles of scientific thinking or reasoning: A new analytical tool for historians and philosophers of the sciences. In Trends in the Historiography of Science [GCN94], chapter 3, pages 31–48. ISBN 90-481-4264-4, 94-017-3596-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-017-
3596-4_3.

Hacking:2015:LTA

[Hac15a] Ian Hacking. Let’s not talk about objectivity. In Objectivity in Science: New Perspectives from Science and Technology Studies [PRT15], chapter 2, pages 19–33. ISBN 3-319-14348-4 (hardcover), 3-319-14349-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175 O25 2015.

URL https://link.springer.com/
chapter/10.1007/978-3-319-14349-
1_2.

Hacking:2015:PRN

[Hac15b] Ian Hacking. Probable reasoning and its novelties. In Arabatzis et al. [ARS15], chapter 12, pages 177–192. ISBN 3-319-14552-5 (paperback), 3-319-14553-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q125 .R45 2015. URL https://link.

springer.com/chapter/10.1007/
978-3-319-14553-2_12.

Hacohen:2017:JET

[Hac17] Malachi H. Hacohen. Jacob & Esau Today: The end of a two millennia paradigm? In Encouraging Openness: Essays for Joseph Agassi on the Occasion of His 90th Birthday [BAG17], chapter 14, pages 167–190. ISBN 3-319-57668-2 (hardcover), 3-319-57669-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN B833.E53

2017. URL https://link.springer.

                                              com/chapter/10.1007/978-3-319-
                                              57669-5_14.

Hagengruber:2016:MVW

[Hag16] Ruth Hagengruber. Measuring the value of women: A feminist analysis of economic categories and thought. In Meta-Philosophical Reflection on Feminist Philosophies of Science [AV16a], chapter 10, pages 171–183. ISBN 3-319-26346-3 (hardcover), 3-319-26348-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN HQ1154 .M48 2016; HQ1190 M48 2016. URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-3-319-26348-9_10.

Hahn:1991:PLT

[Hah91] Elke Hahn. The philosophy of living things: Schelling’s Naturphilosophie as a transition to the philosophy of identity. In Woodward and Cohen [WC91], pages 339–350. ISBN 94-010-5395-2, 94-011-3164-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-011-3164-3_31.

Hall:1971:CWU

[Hal71] Richard J. Hall. Can we use the history of science to decide between competing methodologies. In Buck and Cohen [BC71], pages 151–159. ISBN 90-277-0309-4, 94-010-3142-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-010-3142-4_10.

Haller:1982:PIE

[Hal82] Rudolf Haller. Poetic imagination and economy: Ernst Mach as theorist of science. In Scientific Philosophy Today: Essays in Honor of Mario Bunge [AC82], chapter 3, pages 71–84. ISBN 90-277-1262-X, 90-277-1263-8, 94-009-8462-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 67. URL https:

//link.springer.com/chapter/10.
1007/978-94-009-8462-2_3.

Halfmann:1984:DPS

[Hal84] Jost Halfmann. The dethroning of the philosophy of science: Ideological and technical functions of the metasciences. In Methodology, Metaphysics and the History of Science: In Memory of Benjamin Nelson [CW84b], chapter 5, pages 149–171. ISBN 94-009-6333-5, 94-009-6331-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-009-
6331-3_5.

Hall:1990:WGM

[Hal90a] A. Rupert Hall. Was Galileo a metaphysicist? In Levere and Shea [LS90], chapter 4, pages 105–121. ISBN 0-7923-0420-9, 94-009-1878-X. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 120. URL https:/

/link.springer.com/chapter/10.
1007/978-94-009-1878-8_4.

Hall:1990:ISN

[Hal90b] Marie Boas Hall. The inductive sciences in Nineteenth-Century England. In Levere and Shea [LS90], chapter 10, pages 225–247. ISBN 0-7923-0420-9, 94-009-1878-X. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 120. URL https://link.

springer.com/chapter/10.1007/
978-94-009-1878-8_10.

Haller:1991:FVC

[Hal91a] Rudolf Haller. The first Vienna Circle.

In Uebel [Ueb91b], chapter 7, pages 95–108. ISBN 0-7923-1276-7 (hardcover), 94-011-3182-1. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 133. URL https:/

                                              /link.springer.com/chapter/10.
                                              1007/978-94-011-3182-7_7.

Haller:1991:HSS

[Hal91b] Rudolf Haller. History and the system of science in Otto Neurath. In Uebel [Ueb91b], chapter 3, pages 33–40. ISBN 0-7923-1276-7 (hardcover), 94-011-3182-1. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 133. URL https://link.

                                              springer.com/chapter/10.1007/
                                              978-94-011-3182-7_3.

Haller:1991:NPG

[Hal91c] Rudolf Haller. The Neurath principle: Its grounds and consequences. In Uebel [Ueb91b], chapter 9, pages 117–129. ISBN 0-7923-1276-7 (hardcover), 94-011-3182-1. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 133. URL https://link.

                                              springer.com/chapter/10.1007/
                                              978-94-011-3182-7_9.

Haller:1991:ON

[Hal91d] Rudolf Haller. On Otto Neurath. In Uebel [Ueb91b], chapter 2, pages 25– 31. ISBN 0-7923-1276-7 (hardcover), 94-011-3182-1. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 133. URL https://link.

                                              springer.com/chapter/10.1007/
                                              978-94-011-3182-7_2.

Haller:1991:HAP

[Hal91e] Rudolf Haller. On the historiography of Austrian philosophy. In Uebel [Ueb91b], chapter 4, pages 41–50. ISBN 0-7923-1276-7 (hardcover), 94-011-3182-1. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 133. URL https://link.

springer.com/chapter/10.1007/
978-94-011-3182-7_4.

Haller:1991:TWE

[Hal91f] Rudolf Haller. Two ways of experiential justification. In Uebel [Ueb91b], chapter 15, pages 191–201. ISBN 0-7923-1276-7 (hardcover), 94-011-3182-1. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 133. URL https://link.

springer.com/chapter/10.1007/
978-94-011-3182-7_15.

Haller:1992:PIE

[Hal92] Rudolf Haller. Poetic imagination and economy: Ernst Mach as theorist of science. In Ernst Mach — A Deeper Look: Documents and New Perspectives [Bla92a], chapter 11, pages 215–228. ISBN 94-010-5230-1, 94-011-2771-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN C5.53;

QC6.9. URL https://link.springer.
com/chapter/10.1007/978-94-011-
2771-4_11.

Haller:1995:DCN

[Hal95a] Rudolf Haller. Dirt and crystal: Neurath on the language of science. In Gavroglu et al. [GSW95c], chapter 17, pages 287–299. ISBN 0-7923-2991-0 (set), 0-7923-2988-0, 94-017-2658-2. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 163.

URL https://link.springer.com/
chapter/10.1007/978-94-017-2658-
0_17.

Hallett:1995:HL

[Hal95b] Michael Hallett. Hilbert and logic. In Québec Studies in the Philosophy of Science: Part I: Logic, Mathematics, Physics and History of Science Essays in Honor of Hugues Leblanc [MC95], chapter 10, pages 135–187. ISBN 94-010-7204-3, 94-009-1575-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-009-1575-6_10.

Haller:1997:AO

[Hal97] Rudolf Haller. An autobiographical outline. In Austrian Philosophy Past and Present: Essays in Honor of Rudolf Haller [LM97], chapter 16, pages 237–249. ISBN 94-010-6412-1, 94-011-5720-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic).

LCCN ????  URL https://link.
springer.com/chapter/10.1007/
978-94-011-5720-9_16.

Haallsten:2007:WAE

[Hål07] Henrik Hållsten. What to ask of an explanation — theory. In Rethinking Explanation [PY07], chapter 2, pages 13–26. ISBN 1-4020-5580-3 (paperback), 1-4020-5581-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN BD237 .R48 2007; Q175.32.E97 R48 2007. URL https:

//link.springer.com/chapter/10.
1007/978-1-4020-5581-2_2.

Halimi:2016:SD

[Hal16] Brice Halimi. Sets and descent. In Objectivity, Realism, and Proof: FilMat Studies in the Philosophy of Mathematics [BS16], chapter 8, pages 123–142. ISBN 3-319-31642-7, 3-319-31644-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QA9

.O25 2016.  URL https://link.
springer.com/chapter/10.1007/
978-3-319-31644-4_8.

Hamm:1991:GMS

[Ham91] Ernst Hamm. Goethe’s morphology of stones: Between natural history and historical geology. In Woodward and Cohen [WC91], pages 329–338. ISBN 94-010-5395-2, 94-011-3164-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-011-3164-3_30.

Hamm:2007:SOC

[Ham07] Ernst P. Hamm. Steffens, Ørsted, and the chemical construction of the Earth. In Hans Christian Ørsted and The Romantic Legacy In Science: Ideas, Disciplines, Practices [BCK07], chapter 9, pages 159–175. ISBN 1-4020-2979-9 (hardcover), 1-4020-2987-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC515.O37 H36 2007. URL https:

//link.springer.com/chapter/10.
1007/978-1-4020-2987-5_9.

Hamou:2018:ECN

[Ham18] Philippe Hamou. Experimentum crucis:

Newton’s empiricism at the crossroads. In What Does it Mean to be an Empiricist?: Empiricisms in Eighteenth Century Sciences [BR18b], chapter 4, pages 47–69. ISBN 3-319-69858-3 (hardcover), 3-319-69860-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN B816 .W43 2018.

URL https://link.springer.com/
chapter/10.1007/978-3-319-69860-
1_4.

Hanson:1967:WDB

[Han67] Norwood Russell Hanson. What I don’t believe. In Cohen and Wartofsky [CW67], pages 467–489. ISBN 90-277-0013-3, 94-010-3508-3. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175 .B731 v.3. URL https://link.

                                              springer.com/chapter/10.1007/
                                              978-94-010-3508-8_28.

Hanly:1988:AR

[Han88] Charles Hanly. From animism to rationalism. In Philosophy, History and Social Action: Essays in Honor of Lewis Feuer with an autobiographic essay by Lewis Feuer [HOO88c], chapter 9, pages 221–234. ISBN 94-010-7793-2, 94-009-2873-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-009-2873-2_9.

Hannay:1991:FPW

[Han91] Alastair Hannay. Free of prejudice and wholly critical? In Beyond Reason: Essays on the Philosophy of Paul Feyerabend [Mun91a], chapter 4, pages 109–129. ISBN 94-010-5406-1, 94-011-3188-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

                                              ???? URL https://link.springer.
                                              com/chapter/10.1007/978-94-011-
                                              3188-9_4.

Hannay:1995:CP

[Han95] Alastair Hannay. Commitment and paradox. In Critical Rationalism, the Social Sciences and the Humanities: Essays for Joseph Agassi. Volume II [JL95b], chapter 11, pages 189–202. ISBN 0-7923-2961-9, 90-481-4430-2, 94-017-0441-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN B29 .C75 1995. URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-017-0441-0_11.

Hanzel:1999:CSL

[Han99a] Igor Hanzel. The Concept of Scientific Law in the Philosophy of Science and Epistemology: A Study of Theoretical Reason, volume 208 of Boston Studies in the Philosophy of Science. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 1999. ISBN 90-481-5275-5, 94-017-3265-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). xiv + 230 pp. LCCN ???? URL

https://link.springer.com/book/
10.1007/978-94-017-3265-9.

Hanzel:1999:C

[Han99b] Igor Hanzel. Conclusion. In The Concept of Scientific Law in the Philosophy of Science and Epistemology: A Study of Theoretical Reason [Han99a], chapter 12, pages 197–200. ISBN 90-481-5275-5, 94-017-3265-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-017-3265-9_12.

Hanzel:1999:EL

[Han99c] Igor Hanzel. Epistemology and logic.

In The Concept of Scientific Law in the Philosophy of Science and Epistemology: A Study of Theoretical Reason [Han99a], chapter 10, pages 157–174. ISBN 90-481-5275-5, 94-017-3265-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-017-3265-
9_10.

Hanzel:1999:EF

[Han99d] Igor Hanzel. The essential formality.

In The Concept of Scientific Law in the Philosophy of Science and Epistemology: A Study of Theoretical Reason [Han99a], chapter 7, pages 108–121. ISBN 90-481-5275-5, 94-017-3265-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-017-3265-
                                              9_7.

Hanzel:1999:HMC

[Han99e] Igor Hanzel. Hegel’s and Marx’s categories: A comparison. In The Concept of Scientific Law in the Philosophy of Science and Epistemology: A Study of Theoretical Reason [Han99a], chapter 11, pages 175–196. ISBN 90-481-5275-5, 94-017-3265-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-017-3265-9_11.

Hanzel:1999:ILEa

[Han99f] Igor Hanzel. The idealizational law of essence. In The Concept of Scientific Law in the Philosophy of Science and Epistemology: A Study of Theoretical Reason [Han99a], chapter 3, pages 33– 51. ISBN 90-481-5275-5, 94-017-3265-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-017-3265-
                                              9_3.

Hanzel:1999:ILEc

[Han99g] Igor Hanzel. The inherent law and essentialism. In The Concept of Scientific Law in the Philosophy of Science and Epistemology: A Study of Theoretical Reason [Han99a], chapter 8, pages 122–133. ISBN 90-481-5275-5, 94-017-3265-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

                                              ???? URL https://link.springer.
                                              com/chapter/10.1007/978-94-017-
                                              3265-9_8.

Hanzel:1999:ILEb

[Han99h] Igor Hanzel. The inherent law of essence. In The Concept of Scientific Law in the Philosophy of Science and Epistemology: A Study of Theoretical Reason [Han99a], chapter 5, pages 63– 84. ISBN 90-481-5275-5, 94-017-3265-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-017-3265-
9_5.

Hanzel:1999:LA

[Han99i] Igor Hanzel. The law of appearance.

In The Concept of Scientific Law in the Philosophy of Science and Epistemology: A Study of Theoretical Reason [Han99a], chapter 1, pages 3–13. ISBN 90-481-5275-5, 94-017-3265-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-017-3265-9_1.

Hanzel:1999:OF

[Han99j] Igor Hanzel. Ontological foundations.

In The Concept of Scientific Law in the Philosophy of Science and Epistemology: A Study of Theoretical Reason [Han99a], chapter 4, pages 52–59. ISBN 90-481-5275-5, 94-017-3265-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-017-3265-9_4.

Hanzel:1999:PM

[Han99k] Igor Hanzel. Popper’s moa. In The Concept of Scientific Law in the Philosophy of Science and Epistemology: A Study of Theoretical Reason [Han99a], chapter 2, pages 14–30. ISBN 90-481-5275-5, 94-017-3265-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-017-3265-9_2.

Hanzel:1999:SS

[Han99l] Igor Hanzel. Substance–subject. In The Concept of Scientific Law in the Philosophy of Science and Epistemology: A Study of Theoretical Reason [Han99a], chapter 9, pages 134–153. ISBN 90-481-5275-5, 94-017-3265-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-017-3265-
                                              9_9.

Hanzel:1999:TFM

[Han99m] Igor Hanzel. The three faces of measure. In The Concept of Scientific Law in the Philosophy of Science and Epistemology: A Study of Theoretical Reason [Han99a], chapter 6, pages 85–107. ISBN 90-481-5275-5, 94-017-3265-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-017-3265-
                                              9_6.

Hansson:2007:EAC

[Han07] Bengt Hansson. Explanations are about concepts and concept formation. In Rethinking Explanation [PY07], chapter 1, pages 3–11. ISBN 1-4020-5580-3 (paperback), 1-4020-5581-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN BD237 .R48 2007; Q175.32.E97 R48 2007. URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-1-4020-5581-2_1.

Handy:2014:PPC

[Han14] Carlos R. Handy. A perspective on physics in Cuba. In The History of Physics in Cuba [BRW14a], chapter 30, pages 407–411. ISBN 94-017-8040-4, 94-017-8041-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic).

LCCN QC5.53; QC6.9; QC9 C9H57 2014; Q124.6-127.2. URL https:/

/link.springer.com/chapter/10.
1007/978-94-017-8041-4_30.

Hanken:2015:HSE

[Han15] James Hanken. Is heterochrony still an effective paradigm for contemporary studies of evo-devo? In Conceptual Change in Biology: Scientific and Philosophical Perspectives on Evolution and Development [Lov15b], chapter 4, pages 97–110. ISBN 94-017-9411-1, 94-017-9412-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QH331 C66 2015; QH491; QH492 .C66

2014. URL https://link.springer.
com/chapter/10.1007/978-94-017-
9412-1_4.

Handelman:2017:PMH

[Han17] Sapir Handelman. The perfect mediator and his limitations: Toward a multifaceted approach to peacemaking in the Israeli–Palestinian struggle. In Encouraging Openness: Essays for Joseph Agassi on the Occasion of His 90th Birthday [BAG17], chapter 22, pages 263–286. ISBN 3-319-57668-2 (hardcover), 3-319-57669-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN B833.E53 2017.

URL https://link.springer.com/
chapter/10.1007/978-3-319-57669-
5_22.

Haq:1994:CES

[Haq94a] Syed Nomanul Haq. A critical edition of select text of the Kitāb Al-Ahjār. In Names, Natures and Things: The Alchemist Jābir ibn Hayyān and his Kitāb al-Ahjār (Book of Stones) [Haq94f], chapter 6, pages 117–162. ISBN 0-7923-2587-7, 94-011-1898-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 158

                                              QD25. URL https://link.springer.
                                              com/chapter/10.1007/978-94-011-
                                              1898-9_6.

Haq:1994:DCJ

[Haq94b] Syed Nomanul Haq. The doctrinal context of Jābir’s Kitāb Al-Ahjār: Substance, qualities and the science of balance. In Names, Natures and Things: The Alchemist Jābir ibn Hayyān and his Kitāb al-Ahjār (Book of Stones) [Haq94f], chapter 3, pages 49–80. ISBN 0-7923-2587-7, 94-011-1898-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 158 QD25. URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-011-1898-9_3.

Haq:1994:ENN

[Haq94c] Syed Nomanul Haq. Errata names, natures and things. In Names, Natures and Things: The Alchemist Jābir ibn Hayyān and his Kitāb al-Ahjār (Book of Stones) [Haq94f], chapter 1, pages 1–2. ISBN 0-7923-2587-7, 94-011-1898-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 158 QD25. URL https:/

                                              /link.springer.com/chapter/10.
                                              1007/978-94-011-1898-9_1.

Haq:1994:I

[Haq94d] Syed Nomanul Haq. Introduction. In Names, Natures and Things: The Alchemist Jābir ibn Hayyān and his Kitāb al-Ahjār (Book of Stones) [Haq94f], chapter 2, pages 3–47. ISBN 0-7923-2587-7, 94-011-1898-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 158

                                              QD25. URL https://link.springer.
                                              com/chapter/10.1007/978-94-011-
                                              1898-9_2.

Haq:1994:NNTa

[Haq94e] Syed Nomanul Haq. Names, natures and things: A prefatory note on the central theme of the Kitāb Al-Ahjār. In Names, Natures and Things: The Alchemist Jābir ibn Hayyān and his Kitāb al-Ahjār (Book of Stones) [Haq94f], chapter 4, pages 81–108. ISBN 0-7923-2587-7, 94-011-1898-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 158

QD25. URL https://link.springer.
com/chapter/10.1007/978-94-011-
1898-9_4.

Haq:1994:NNTb

[Haq94f] Syed Nomanul Haq. Names, Natures and Things: The Alchemist Jābir ibn Hayyān and his Kitāb al-Ahjār (Book of Stones), volume 158 of Boston Studies in the Philosophy of Science. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 1994. ISBN 0-7923-2587-7, 94-011-1898-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). xx + 284 pp. LCCN Q174 .B67 vol. 158 QD25. URL

https://link.springer.com/book/
10.1007/978-94-011-1898-9.

Haq:1994:PO

[Haq94g] Syed Nomanul Haq. Preliminary observations. In Names, Natures and Things: The Alchemist Jābir ibn Hayyān and his Kitāb al-Ahjār (Book of Stones) [Haq94f], chapter 5, pages 111–116. ISBN 0-7923-2587-7, 94-011-1898-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 158 QD25. URL https:/

/link.springer.com/chapter/10.
1007/978-94-011-1898-9_5.

Haq:1994:TCT

[Haq94h] Syed Nomanul Haq. Translation, commentary and textual notes. In Names, Natures and Things: The Alchemist Jābir ibn Hayyān and his Kitāb al-Ahjār (Book of Stones) [Haq94f], chapter 7, pages 163–248. ISBN 0-7923-2587-7, 94-011-1898-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 158

                                              QD25. URL https://link.springer.
                                              com/chapter/10.1007/978-94-011-
                                              1898-9_7.

Harris:1984:DSS

[Har84a] Errol Harris. The dialectical structure of scientific thinking. In Hegel and the Sciences [CW84a], chapter 12, pages 195–213. ISBN 94-009-6235-5, 94-009-6233-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

                                              ???? URL https://link.springer.
                                              com/chapter/10.1007/978-94-009-
                                              6233-0_12.

Hartman:1984:PSO

[Har84b] Hyman Hartman. Philosophy of science and the origin of life. In Methodology, Metaphysics and the History of Science: In Memory of Benjamin Nelson [CW84b], chapter 7, pages 183–213. ISBN 94-009-6333-5, 94-009-6331-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-009-6331-
                                              3_7.

Harre:1986:LE

[Har86] Rom Harré. Location and existence. In Cohen and Wartofsky [CW85], chapter 13, pages 243–257. ISBN 90-277-1971-3, 94-009-5345-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-009-5349-9_13.

Harper:1989:CNK

[Har89a]  William Harper.  Consilience and

natural kind reasoning. In An Intimate Relation: Studies in the History and Philosophy of Science Presented to Robert E. Butts on his 60th Birthday [BM89], chapter 7, pages 115–152. ISBN 94-010-7546-8, 94-009-2327-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-009-2327-
0_7.

Harris:1989:SI

[Har89b] Zellig Harris. From structure to information. In The Form of Information in Science: Analysis of an Immunology Sublanguage [HGR^{+}89], chapter 3, pages 64–84. ISBN 94-010-7777-0, 94-009-2837-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN B832.A-832.Z. URL https://link.

springer.com/chapter/10.1007/
978-94-009-2837-4_3.

Harris:1989:RTF

[Har89c] Zellig Harris. Reducing texts to formulas. In The Form of Information in Science: Analysis of an Immunology Sublanguage [HGR^{+}89], chapter 1, pages 1–24. ISBN 94-010-7777-0, 94-009-2837-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN B832.A-832.Z. URL https://link.

springer.com/chapter/10.1007/
978-94-009-2837-4_1.

Harris:1989:RFI

[Har89d] Zellig Harris. Result: Formulas of information. In The Form of Information in Science: Analysis of an Immunology Sublanguage [HGR^{+}89], chapter 2, pages 25–63. ISBN 94-010-7777-0, 94-009-2837-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN B832.A-832.Z. URL https://link.

springer.com/chapter/10.1007/
978-94-009-2837-4_2.

Harman:1995:TLG

[Har95]  Peter M. Harman.  Through the

looking-glass, and what Maxwell found there. In No Truth Except in the Details: Essays in Honor of Martin J. Klein [KS95], chapter 4, pages 79–93. ISBN 94-010-4097-4, 94-011-0217-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-011-0217-
9_4.

Hardy:1996:CBM

[Har96]  Lucien Hardy.  Contextuality in

Bohmian mechanics. In Bohmian Mechanics and Quantum Theory: an Appraisal [CFG96], chapter 4, pages 67–76. ISBN 90-481-4698-4, 94-015-8715-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC173.96-174.52. URL https://link.

springer.com/chapter/10.1007/
978-94-015-8715-0_4.

Hardy:1997:EPR

[Har97a]  Lucien Hardy.  Einstein–Podolsky–

Rosen reasoning in nonlocality theorems. In Cohen et al. [CHS97b], pages 105–111. ISBN 90-481-4809-X, 94-017-2732-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC173.96-174.52. URL https:

//link.springer.com/chapter/10.
1007/978-94-017-2732-7_8.

Harre:1997:RSS

[Har97b] Rom Harré. The redundancy of spacetime: Special relativity as a grammar and the strangeness of ‘c’. In Ginev and Cohen [GC97], chapter 7, pages 95–105. ISBN 0-7923-4444-8 (hardcover), 94-011-5788-X. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 192. URL https:/

/link.springer.com/chapter/10.
1007/978-94-011-5788-9_7.

Hardcastle:1999:ATN

[Har99] Gary L. Hardcastle. Accomplishing translation: The notion of evidence in the discipline of the history of science. In Fehér et al. [FKR99], pages 117–124. ISBN 90-481-5257-7, 94-015-9293-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-015-9293-
2_10.

Harrington:2001:CQC

[Har01] Anne Harrington. China, “Qi,” and the challenges of engaged scholarship. In Allen and MacLeod [AM01b], chapter 7, pages 105–122. ISBN 1-4020-0495-8 (hardcover), 94-017-2956-5. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 228; Q175.55.

URL https://link.springer.com/
chapter/10.1007/978-94-017-2956-
7_7.

Harre:2002:SWC

[Har02] Rom Harré. Science as the work of a community. In Hermeneutic Philosophy of Science, Van Gogh’s Eyes, and God: Essays in Honor of Patrick A. Heelan, S.J. [Bab02c], chapter 18, pages 219–229. ISBN 90-481-5926-1, 94-017-1767-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 225. URL https:/

/link.springer.com/chapter/10.
1007/978-94-017-1767-0_18.

Harre:2008:RCA

[Har08a] Rom Harré. Reasons for choosing among readings of equipollent theories. In Rethinking Scientific Change and Theory Comparison: Stabilities, Ruptures, Incommensurabilities? [SSHH08], chapter 16, pages 225–238. ISBN 1-4020-6274-5 (hardcover), 1-4020-6279-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175 .R397 2008. URL https://link.

springer.com/chapter/10.1007/
978-1-4020-6279-7_16.

Hartmann:2008:MHT

[Har08b] Stephan Hartmann. Modeling hightemperature superconductivity: Correspondence at bay? In Rethinking Scientific Change and Theory Comparison: Stabilities, Ruptures, Incommensurabilities? [SSHH08], chapter 8, pages 109–129. ISBN 1-4020-6274-5 (hardcover), 1-4020-6279-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175 .R397 2008.

URL https://link.springer.com/
chapter/10.1007/978-1-4020-6279-
7_8.

Harper:2010:TFK

[Har10] Donald Harper. The textual form of knowledge: Occult miscellanies in ancient and medieval Chinese manuscripts, Fourth Century B.C. to Tenth Century A.D. In Looking at it from Asia: the Processes that Shaped the Sources of History of Science [BE10b], chapter 2, pages 37–80. ISBN 90-481-3675-X (hardcover), 90-481-3676-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q127.A65 L66 2010. URL https:/

/link.springer.com/chapter/10.
1007/978-90-481-3676-6_2.

Harding:2015:OSB

[Har15a] Sandra Harding. Objectivity for sciences from below. In Objectivity in Science: New Perspectives from Science and Technology Studies [PRT15], chapter 3, pages 35–55. ISBN 3-319-14348-4 (hardcover), 3-319-14349-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175 O25 2015.

URL https://link.springer.com/
chapter/10.1007/978-3-319-14349-
1_3.

Harre:2015:MPC

[Har15b] Rom Harré. Mereological principles and chemical affordances. In Philosophy of Chemistry: Growth of a New Discipline [SM15b], chapter 8, pages 107–119. ISBN 94-017-9363-8 (hardcover), 94-017-9364-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QD6 .P455 2015; QD6 .S384 2014.

URL https://link.springer.com/
chapter/10.1007/978-94-017-9364-
3_8.

Harrell:2016:PFP

[Har16] Maralee Harrell. On the possibility of feminist philosophy of physics. In Meta-Philosophical Reflection on Feminist Philosophies of Science [AV16a], chapter 2, pages 15–34. ISBN 3-319-26346-3 (hardcover), 3-319-26348-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN HQ1154 .M48 2016; HQ1190 M48 2016. URL https:

//link.springer.com/chapter/10.
1007/978-3-319-26348-9_2.

Hasle:1997:LTL

[Has97] Per F. V. Hasle. Linguistic and tense logical considerations on the generality of a three-point structure of tenses. In Perspectives on Time [FSU97b], chapter 14, pages 323–344. ISBN 90-481-4774-3, 94-015-8875-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol.189.

URL https://link.springer.com/
chapter/10.1007/978-94-015-8875-
1_14.

Hashimoto:2011:AAM

[Has11] Takehiko Hashimoto. The adoption and adaptation of mechanical clocks in Japan. In Science between Europe and Asia: Historical Studies on the Transmission, Adoption and Adaptation of Knowledge [GR11b], chapter 9, pages 137–149. ISBN 90-481-9968-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q125 .S354 2010; Q175.52.A78 S35 2010. URL https:

//link.springer.com/chapter/10.
1007/978-90-481-9968-6_9.

Hattiangadi:1980:VCD

[Hat80] J. N. Hattiangadi. The vanishing context of discovery: Newton’s discovery of gravity. In Scientific Discovery, Logic, and Rationality [Nic80f], chapter 11, pages 257–265. ISBN 90-277-1070-8, 94-009-8986-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-009-
8986-3_11.

Hattiangadi:1983:MMR

[Hat83] J. N. Hattiangadi. A methodology without methodological rules. In Language, Logic and Method [CW83b], chapter 4, pages 103–151. ISBN 94-009-7704-2, 94-009-7702-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-009-7702-0_4.

Hattiangadi:1989:KS

[Hat89] J. N. Hattiangadi. Kuhn studies. In Freedom and Rationality: Essays in Honor of John Watkins From his Colleagues and Friends [DJ89], chapter 12, pages 191–205. ISBN 94-010-7571-9, 94-009-2380-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-009-
2380-5_12.

Hattiangadi:1995:FWW

[Hat95] Jagdish Hattiangadi. The first world war and 1991. In Critical Rationalism, the Social Sciences and the Humanities: Essays for Joseph Agassi. Volume II [JL95b], chapter 4, pages 59–78. ISBN 0-7923-2961-9, 90-481-4430-2, 94-017-0441-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN B29 .C75 1995. URL https:

//link.springer.com/chapter/10.
1007/978-94-017-0441-0_4.

Hattiangadi:2017:TPC

[Hat17] Jagdish Hattiangadi. Two philosophical contributions to modern physics. In Encouraging Openness: Essays for Joseph Agassi on the Occasion of His 90th Birthday [BAG17], chapter 9, pages 91–103. ISBN 3-319-57668-2 (hardcover), 3-319-57669-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN B833.E53 2017.

URL https://link.springer.com/
chapter/10.1007/978-3-319-57669-
5_9.

Hausman:2010:PCC

[Hau10] Daniel M. Hausman. Probabilistic causality and causal generalizations. In The Place of Probability in Science: In Honor of Ellery Eells (1953–2006) [Fet10a], chapter 2, pages 47–63. ISBN 90-481-3614-8 (hardcover), 90-481-3615-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175 .P563 2010. URL https://link.

                                              springer.com/chapter/10.1007/
                                              978-90-481-3615-5_2.

Havas:1969:CRT

[Hav69] Peter Havas. Causality requirements and the theory of relativity. In Cohen and Wartofsky [CW69a], pages 151–178. ISBN 94-010-3383-8, 94-010-3263-7, 94-010-3381-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-010-3381-7_2.  See

comments [Sta69].

Hawkins:1967:TI

[Haw67] David Hawkins. Taxonomy and information. In Cohen and Wartofsky [CW67], pages 41–55. ISBN 90-277-0013-3, 94-010-3508-3. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175 .B731 v.3. URL https://link.

                                              springer.com/chapter/10.1007/
                                              978-94-010-3508-8_2.

Hayward:2003:EMC

[Hay03] Rhodri Hayward. Emmanuel Mendes da Costa (1717–1791). In Simões et al. [SCD03b], chapter 5, pages 101–114. ISBN 1-4020-1259-4, 94-017-3584-0. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 v. 233.

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-017-3584-
                                              1_5.

Holton:1976:SPC

[HB76] Gerald James Holton and William A. Blanpied, editors. Science and Its Public: the Changing Relationship, volume 33 of Boston Studies in the Philosophy of Science. D. Reidel, Dordrecht, The Netherlands; Boston, MA, USA; Lancaster, UK; Tokyo, Japan, 1976. ISBN 90-277-0657-3 (hardcover), 90-277-0658-1 (paperback). xxv + 289 pp. LCCN Q174 .B67 vol. 33; Q175.5. Synthese Library (volume 96).

Hector:2018:TPP

[HCP18] Donald Hector, Carleton Christensen, and Jim Petrie. Toward a practical philosophy of engineering: Dealing with complex problems from the sustainability discourse. In Mitcham et al. [MLNZ18], chapter 11, pages 119–135. ISBN 3-319-62448-2, 3-319-62450-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN TA15

.P45 2018.  URL https://link.
springer.com/chapter/10.1007/
978-3-319-62450-1_11.

Herridge:1995:ECL

[HD95] D. F. Herridge and S. K. A. Danso.

Enhancing crop legume N_{2} fixation through selection and breeding. In No Truth Except in the Details: Essays in Honor of Martin J. Klein [KS95], chapter 3, pages 51–82. ISBN 94-010-4097-4, 94-011-0217-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-011-0055-7_3.

Heath:2004:DMS

[Hea04]  Thomas L. Heath.  Diophantus’

methods of solution. In Classics in the History of Greek Mathematics [Chr04a], chapter 17, pages 285–327. ISBN 1-4020-0081-2 (hardcover), 1-4020-2640-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QA22 .C596 2004. URL https:/

/link.springer.com/chapter/10.
1007/978-1-4020-2640-9_17.

Hecht:2001:LCP

[Hec01] Hartmut Hecht. Leibniz’ concept of possible worlds and the analysis of motion in Eighteenth-Century physics. In Between Leibniz, Newton, and Kant: Philosophy and Science in the Eighteenth Century [Lef01a], chapter 2, pages 27–45. ISBN 90-481-5774-9, 94-015-9729-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

                                              ???? URL https://link.springer.
                                              com/chapter/10.1007/978-94-015-
                                              9729-6_2.

Heelan:1974:QLC

[Hee74] Patrick A. Heelan. Quantum logic and classical logic: Their respective roles. In Logical and Epistemological Studies in Contemporary Physics [CW74a], chapter 11, pages 318–349. ISBN 90-277-0391-4 (hardcover), 90-277-0377-9 (paperback), 94-010-2656-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol.

                                              13; QC5.56. URL https://link.
                                              springer.com/chapter/10.1007/
                                              978-94-010-2656-7_11.

Heelan:1979:LGK

[Hee79] Patrick A. Heelan. The lattice of growth in knowledge. In The Structure and Development of Science [RA79], chapter 8, pages 205–211. ISBN 90-277-0995-5, 94-009-9459-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .S778 1979.

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-009-9459-
                                              1_8.

Heelan:1984:CPE

[Hee84] Patrick A. Heelan. Commentary on ‘particles or events?’. In Physical Sciences and History of Physics [CW84c], chapter 2, pages 29–34. ISBN 94-009-7180-X, 94-009-7178-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .P497 1984.

URL https://link.springer.com/
chapter/10.1007/978-94-009-7178-
3_2.

Heelan:1986:FCE

[Hee86a] Patrick A. Heelan. Fleck’s contribution to epistemology. In Cognition and Fact: Materials on Ludwik Fleck [CS86], chapter 15, pages 287–307. ISBN 94-010-8504-8, 94-009-4498-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-009-4498-5_15.

Heelan:1986:MP

[Hee86b] Patrick A. Heelan. Machine perception. In Philosophy and Technology II: Information Technology and Computers in Theory and Practice [MH86], chapter 10, pages 131–156. ISBN 94-010-8510-2, 94-009-4512-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-009-4512-8_10.

Heelan:1997:CHO

[Hee97] Patrick A. Heelan. Context, hermeneutics, and ontology in the experimental sciences. In Ginev and Cohen [GC97], chapter 8, pages 107–126. ISBN 0-7923-4444-8 (hardcover), 94-011-5788-X. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 192.

URL https://link.springer.com/
chapter/10.1007/978-94-011-5788-
9_8.

Heelan:1999:NPH

[Hee99] Patrick A. Heelan. Nietzsche’s perspectivalism: A hermeneutic philosophy of science. In Nietzsche, Epistemology, and Philosophy of Science: Nietzsche and the Sciences II [Bab99a], chapter 17, pages 203–220. ISBN 90-481-5234-8, 94-017-2428-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-017-2428-9_17.

Heelan:2002:A

[Hee02] Patrick A. Heelan. Afterword. In Hermeneutic Philosophy of Science, Van Gogh’s Eyes, and God: Essays in Honor of Patrick A. Heelan, S.J. [Bab02c], chapter 39, pages 445–459. ISBN 90-481-5926-1, 94-017-1767-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 225. URL https://link.

springer.com/chapter/10.1007/
978-94-017-1767-0_39.

Heffner:1987:CRV

[Hef87] John Heffner. Causal relations in visual perception. In Shimony and Nails [SN87], pages 193–214. ISBN 94-010-8168-9, 94-009-3735-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-009-3735-2_11.

Hefferlin:2006:PSM

[Hef06] Ray Hefferlin. The periodics systems of molecules. In Philosophy of Chemistry: Synthesis of a New Discipline [BSM06b], chapter 12, pages 221–243. ISBN 1-4020-2455-X, 1-4020-3256-0, 1-4020-3261-7. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 242 QD6. URL https:/

/link.springer.com/chapter/10.
1007/1-4020-3261-7_12.

Hegge:1987:TSL

[Heg87] Hjalmar Hegge. Theory of science in the light of Goethe’s science of nature. In Amrine et al. [AZW87], chapter 11, pages 195–218. ISBN 90-277-2265-X, 94-009-3761-X. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 97. US$79.00. URL https:

//link.springer.com/chapter/10.
1007/978-94-009-3761-1_11.

Hegel:1991:AWO

[Heg91] Ralf-Dietmar Hegel. August Weismann:

One of the first synthetic theorists of evolutionary biology. In Woodward and Cohen [WC91], pages 259–267. ISBN 94-010-5395-2, 94-011-3164-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-011-3164-3_23.

Heilbron:1988:EMC

[Hei88] John L. Heilbron. The earliest missionaries of the Copenhagen Spirit. In Science in Reflection: The Israel Colloquium: Studies in History, Philosophy, and Sociology of Science Volume 3 [UM88], chapter 16, pages 201–233. ISBN 90-277-2713-9, 94-009-2957-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-009-2957-
9_16.

Heims:1991:FLC

[Hei91] Steve J. Heims. Fritz London and the community of quantum physicists. In Woodward and Cohen [WC91], pages 177–190. ISBN 94-010-5395-2, 94-011-3164-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.

                                              com/chapter/10.1007/978-94-011-
                                              3164-3_16.

Heidelberger:1994:UNM

[Hei94a] Michael Heidelberger. The unity of nature and mind: Gustav Theodor Fechner’s non-reductive materialism. In Poggi and Bossi [PB94], chapter 14, pages 215–236. ISBN 0-7923-2336-X, 94-017-2921-2. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 152. URL https://link.

                                              springer.com/chapter/10.1007/
                                              978-94-017-2921-5_14.

Hein:1994:IBM

[Hei94b] Hilde Hein. Institutional blessing: The museum as canon-maker. In Artifacts, Representations and Social Practice: Essays for Marx Wartofsky [GC94], chapter 1, pages 1–19. ISBN 94-010-4390-6, 94-011-0902-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN B73 .A77 1994.

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-011-0902-
                                              4_1.

Hein:1997:ADS

[Hei97] Hilde Hein. The art of displaying science: Museum exhibitions. In The Elusive Synthesis: Aesthetics and Science [Tau97a], chapter 13, pages 267–288. ISBN 0-7923-4763-3, 0-7923-3904-5, 94-009-1786-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-009-1786-6_13.

Heidelberger:1998:HPS

[Hei98]  Michael Heidelberger.  From

Helmholtz’s philosophy of science to Hertz’s picture–theory. In Heinrich Hertz: Classical Physicist, Modern Philosopher [BHN98], chapter 2, pages 9–24. ISBN 90-481-4881-2, 94-015-8855-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-015-
8855-3_2.

Heinzmann:2000:CO

[Hei00] Gerhard Heinzmann. Convention and observability. In The Reality of the Unobservable: Observability, Unobservability and Their Impact on the Issue of Scientific Realism [AP00b], chapter 9, pages 137–144. ISBN 90-481-5458-8, 94-015-9391-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-015-9391-5_9.

Heilbron:2007:C

[Hei07a] John L. Heilbron. On connoisseurship. In Positioning the history of science [RG07], chapter 14, pages 73–76. ISBN 1-4020-5419-X (hardcover), 1-4020-5420-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q125 .P77 2007.

URL https://link.springer.com/
chapter/10.1007/1-4020-5420-3_14.

Heims:2007:CE

[Hei07b]  Steve Joshua Heims.  Concerning

energy. In Positioning the history of science [RG07], chapter 15, pages 77–85. ISBN 1-4020-5419-X (hardcover), 1-4020-5420-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q125 .P77 2007.

URL https://link.springer.com/
chapter/10.1007/1-4020-5420-3_15.

Heim:2008:CAR

[Hei08a] Susanne Heim. Calories — agricultural research, the food economy and war. In Plant Breeding and Agrarian Research in Kaiserwilhelm-Institutes 1933–1945: Calories, Caoutchouc, Careers [Hei08f], pages 15–95. ISBN 1-4020-6717-8 (hardcover), 1-4020-6718-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN SB123.25.G3 H44913 2008. URL https://link.

springer.com/chapter/10.1007/
978-1-4020-6718-1_2.

Heim:2008:CVW

[Hei08b] Susanne Heim. Caoutchouc — a vital war reserve. In Plant Breeding and Agrarian Research in Kaiserwilhelm-Institutes 1933–1945: Calories, Caoutchouc, Careers [Hei08f], pages 97–153. ISBN 1-4020-6717-8 (hardcover), 1-4020-6718-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN SB123.25.G3 H44913

2008. URL https://link.springer.
com/chapter/10.1007/978-1-4020-
6718-1_3.

Heim:2008:CHS

[Hei08c] Susanne Heim. Careers — Hans Stubbe and Klaus von Rosenstiel. In Plant Breeding and Agrarian Research in Kaiserwilhelm-Institutes 1933–1945: Calories, Caoutchouc, Careers [Hei08f], pages 155–192. ISBN 1-4020-6717-8 (hardcover), 1-4020-6718-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN SB123.25.G3 H44913 2008. URL https://link.

springer.com/chapter/10.1007/
978-1-4020-6718-1_4.

Heim:2008:C

[Hei08d] Susanne Heim. Conclusions. In Plant Breeding and Agrarian Research in Kaiserwilhelm-Institutes 1933–1945: Calories, Caoutchouc, Careers [Hei08f], pages 193–201. ISBN 1-4020-6717-8 (hardcover), 1-4020-6718-6 (e-book).

ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN SB123.25.G3 H44913 2008. URL https://link.

springer.com/chapter/10.1007/
978-1-4020-6718-1_5.

Heim:2008:IPA

[Hei08e] Susanne Heim. Introduction: Politics and agricultural research. In Plant Breeding and Agrarian Research in Kaiserwilhelm-Institutes 1933–1945: Calories, Caoutchouc, Careers [Hei08f], pages 1–13. ISBN 1-4020-6717-8 (hardcover), 1-4020-6718-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN SB123.25.G3 H44913

2008. URL https://link.springer.
com/chapter/10.1007/978-1-4020-
6718-1_1.

Heim:2008:PBA

[Hei08f]  Susanne Heim, editor.  Plant

Breeding and Agrarian Research in Kaiserwilhelm-Institutes 1933–1945: Calories, Caoutchouc, Careers, volume 260 of Boston Studies in the Philosophy of Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2008. ISBN 1-4020-6717-8 (hardcover), 1-4020-6718-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN SB123.25.G3 H44913 2008. URL

https://link.springer.com/book/
10.1007/978-1-4020-6718-1.

Heis:2013:ECK

[Hei11] Jeremy Heis. Ernst Cassirer, Kurt Lewin, and Hans Reichenbach. In The Berlin Group and the Philosophy of Logical Empiricism [MP13], chapter 4, pages 67–94. ISBN 94-007-5484-1 (hardcover), 94-007-5485-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175.32.A24 .M384 2011; Q175 .B47 2013. URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-007-5485-0_4.

Heilbron:2015:BB

[Hei15] J. L. Heilbron. Boscovich in Britain. In Arabatzis et al. [ARS15], chapter 8, pages 99–116. ISBN 3-319-14552-5 (paperback), 3-319-14553-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q125 .R45 2015.

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-3-319-14553-
                                              2_8.

Hein:2017:WBS

[Hei17] Hilde Hein. Worker bees and slave labor. In Encouraging Openness: Essays for Joseph Agassi on the Occasion of His 90th Birthday [BAG17], chapter 32, pages 399–416. ISBN 3-319-57668-2 (hardcover), 3-319-57669-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN B833.E53 2017.

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-3-319-57669-
                                              5_32.

Hellman:1974:GSD

[Hel74] C. Doris Hellman. Georg Samuel Dörffel. In For Dirk Struik: Scientific, Historical and Political Essays in Honor of Dirk J. Struik [CSW74], chapter 16, pages 201–203. ISBN 90-277-0379-5, 94-010-2115-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-010-2115-9_16.

Heller:1999:FBF

[Hel99] Ágnes Heller. Friction of bodies, friction of minds. In Fehér et al. [FKR99], pages 93–99. ISBN 90-481-5257-7, 94-015-9293-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-015-
9293-2_8.

Helm:2000:CT

[Hel00a] Georg Helm. Classical thermodynamics. In The Historical Development of Energetics [Hel00f], chapter 3, pages 125–160. ISBN 0-7923-5874-0, 94-011-4471-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 209. URL https:/

/link.springer.com/chapter/10.
1007/978-94-011-4471-1_3.

Helm:2000:EFM

[Hel00b] Georg Helm. The energetic foundation of mechanics. In The Historical Development of Energetics [Hel00f], chapter 6, pages 253–297. ISBN 0-7923-5874-0, 94-011-4471-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 209.

URL https://link.springer.com/
chapter/10.1007/978-94-011-4471-
1_6.

Helm:2000:ETC

[Hel00c] Georg Helm. The energetic treatment of chemistry. In The Historical Development of Energetics [Hel00f], chapter 5, pages 181–251. ISBN 0-7923-5874-0, 94-011-4471-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 209.

URL https://link.springer.com/
chapter/10.1007/978-94-011-4471-
1_5.

Helm:2000:EF

[Hel00d] Georg Helm. Energy factors. In The Historical Development of Energetics [Hel00f], chapter 7, pages 299–362. ISBN 0-7923-5874-0, 94-011-4471-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 209. URL https://link.

springer.com/chapter/10.1007/
978-94-011-4471-1_7.

Helm:2000:EFL

[Hel00e] Georg Helm. The establishment of the first law. In The Historical Development of Energetics [Hel00f], chapter 1, pages 65–105. ISBN 0-7923-5874-0, 94-011-4471-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 209. URL https:/

/link.springer.com/chapter/10.
1007/978-94-011-4471-1_1.

Helm:2000:HDE

[Hel00f] Georg Helm. The Historical Development of Energetics, volume 209 of Boston Studies in the Philosophy of Science. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 2000. ISBN 0-7923-5874-0, 94-011-4471-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). vii + 409 pp. LCCN Q174 .B67 vol. 209. URL

https://link.springer.com/book/
10.1007/978-94-011-4471-1.

Helm:2000:MAE

[Hel00g] Georg Helm. The mechanical approach to energetics and mechanical pictures. In The Historical Development of Energetics [Hel00f], chapter 8, pages 363–404. ISBN 0-7923-5874-0, 94-011-4471-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 209. URL https:/

/link.springer.com/chapter/10.
1007/978-94-011-4471-1_8.

Helm:2000:NIC

[Hel00h] Georg Helm. New initiatives, controversies and misplaced efforts. In The Historical Development of Energetics [Hel00f], chapter 4, pages 161–180. ISBN 0-7923-5874-0, 94-011-4471-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 209. URL https://link.

springer.com/chapter/10.1007/
978-94-011-4471-1_4.

Helm:2000:PSL

[Hel00i] Georg Helm. Preparation for the second law. In The Historical Development of Energetics [Hel00f], chapter 2, pages 107–123. ISBN 0-7923-5874-0, 94-011-4471-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 209. URL https:/

/link.springer.com/chapter/10.
1007/978-94-011-4471-1_2.

Helbing:2008:MNP

[Hel08]  Mario Otto Helbing.  Mechanics

and natural philosophy in Late 16th-Century Pisa: Cesalpino and Buonamici, humanist masters of the Faculty of Arts. In Mechanics and Natural Philosophy Before the Scientific Revolution [LR08], chapter 7, pages 185–193. ISBN 1-4020-5966-3 (hardcover), 1-4020-5967-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QA802 .M43 2008.

URL https://link.springer.com/
chapter/10.1007/978-1-4020-5967-
4_7.

Helling:2014:FKP

[Hel14] Ingeborg K. Helling. Felix Kaufmann in perspective: An introductory essay. In Felix Kaufmann’s Theory and Method in the Social Sciences [CH14], chapter 1, pages 1–101. ISBN 3-319-02844-8 (hardcover), 3-319-02845-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN H61 .K313 2014.

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-3-319-02845-
                                              3_1.

Hempel:1976:DEC

[Hem76] Carl G. Hempel. Dispositional explanation and the covering-law model: Response to Laird Addis. In Cohen et al. [CHMV76], pages 369–376. ISBN 90-277-0647-6, 90-277-0648-4 (paperback), 94-010-1449-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-010-1449-6_18.

Hempel:1983:VOS

[Hem83] Carl G. Hempel. Valuation and objectivity in science. In Cohen and Laudan [CL83], chapter 4, pages 73–100. ISBN 94-009-7055-2, 94-009-7057-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-009-7055-7_4.

Hemmendinger:1984:GPM

[Hem84] David Hemmendinger. Galileo and the phenomena: On making the evidence visible. In Physical Sciences and History of Physics [CW84c], chapter 7, pages 115–143. ISBN 94-009-7180-X, 94-009-7178-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .P497 1984. URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-009-7178-3_7.

Hempel:1986:VOS

[Hem86] Carl G. Hempel. Valuation and objectivity in science. In Cohen and Wartofsky [CW85], chapter 19, pages 277–304. ISBN 90-277-1971-3, 94-009-5345-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-009-
5345-1_19.

Hempel:1988:LDC

[Hem88] Carl G. Hempel. Limits of a deductive construal of the function of scientific theories. In Science in Reflection: The Israel Colloquium: Studies in History, Philosophy, and Sociology of Science Volume 3 [UM88], chapter 1, pages 1– 15. ISBN 90-277-2713-9, 94-009-2957-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-009-2957-
9_1.

Hentschel:1998:HHM

[Hen98] Klaus Hentschel. Heinrich Hertz’s mechanics: A model for Werner Heisenberg’s April 1925 paper on the anomalous Zeeman effect. In Heinrich Hertz: Classical Physicist, Modern Philosopher [BHN98], chapter 12, pages 183–223. ISBN 90-481-4881-2, 94-015-8855-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic).

LCCN ????  URL https://link.
springer.com/chapter/10.1007/
978-94-015-8855-3_12.

Hendry:2001:MRM

[Hen01] Robin Findlay Hendry. Mathematics, representation and molecular structure. In Klein [Kle01b], chapter 13, pages 221–236. ISBN 1-4020-0100-2, 94-015-9737-5. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 222; QD7. URL https:/

/link.springer.com/chapter/10.
1007/978-94-015-9737-1_13.

Hendry:2006:TDC

[Hen06] Robin Findlay Hendry. Is there downward causation in chemistry? In Philosophy of Chemistry: Synthesis of a New Discipline [BSM06b], chapter 9, pages 173–189. ISBN 1-4020-2455-X, 1-4020-3256-0, 1-4020-3261-7. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 242 QD6.

URL https://link.springer.com/
chapter/10.1007/1-4020-3261-7_9.

Herzberger:1974:SHI

[Her74] M. Herzberger. Some heretical ideas with respect to mathematics and physics. In For Dirk Struik: Scientific, Historical and Political Essays in Honor of Dirk J. Struik [CSW74], chapter 27, pages 345–348. ISBN 90-277-0379-5, 94-010-2115-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-010-2115-9_27.

Herzberger:1976:PP

[Her76] Hans G. Herzberger. Presuppositional policies. In Language in Focus: Foundations, Methods and Systems: Essays in Memory of Yehoshua Bar-Hillel [Kas76b], chapter 9, pages 139–164. ISBN 90-277-0644-1, 90-277-0645-X, 94-010-1876-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN P101-120. URL https://link.

springer.com/chapter/10.1007/
978-94-010-1876-0_9.

Herbst:2014:AOS

[Her14a] Marcel Herbst. Academic organization and scientific productivity. In The Institution of Science and the Science of Institutions: The Legacy of Joseph Ben-David [Her14b], chapter 2, pages 15–35. ISBN 94-007-7406-0 (hardcover), 94-007-7407-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 v.302; Q125 .I58

2014. URL https://link.springer.
com/chapter/10.1007/978-94-007-
7407-0_2.

Herbst:2014:ISS

[Her14b] Marcel Herbst. The Institution of Science and the Science of Institutions: The Legacy of Joseph Ben-David, volume 302 of Boston Studies in the Philosophy and History of Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2014. ISBN 94-007-7406-0 (hardcover), 94-007-7407-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). xii + 227 pp. LCCN Q174 .B67 v.302; Q125 .I58 2014. URL

https://link.springer.com/book/
10.1007/978-94-007-7407-0.

Herbst:2014:I

[Her14c] Marcel Herbst. Introduction. In The Institution of Science and the Science of Institutions: The Legacy of Joseph Ben-David [Her14b], chapter 1, pages 3–14. ISBN 94-007-7406-0 (hardcover), 94-007-7407-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 v.302; Q125 .I58

2014. URL https://link.springer.
com/chapter/10.1007/978-94-007-
7407-0_1.

Herbst:2014:LJB

[Her14d] Marcel Herbst. The legacy of Joseph Ben-David. In The Institution of Science and the Science of Institutions: The Legacy of Joseph Ben-David [Her14b], chapter 10, pages 187–211. ISBN 94-007-7406-0 (hardcover), 94-007-7407-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 v.302; Q125 .I58 2014.

URL https://link.springer.com/
chapter/10.1007/978-94-007-7407-
0_10.

Hestenes:1996:GV

[Hes96]  David Hestenes.  Grassmann’s vi-

sion. In Schubring [Sch96f], pages 243–254. ISBN 90-481-4758-1, 94-015-8753-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QA8.9-10.3. URL https://link.

                                              springer.com/chapter/10.1007/
                                              978-94-015-8753-2_20.

Hessenbruch:2007:MDK

[Hes07] Arne Hessenbruch. The making of a Danish Kantian: Science and the new civil society. In Hans Christian Ørsted and The Romantic Legacy In Science: Ideas, Disciplines, Practices [BCK07], chapter 3, pages 21–54. ISBN 1-4020-2979-9 (hardcover), 1-4020-2987-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC515.O37 H36 2007. URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-1-4020-2987-5_3.

Hessen:2009:SER

[Hes09] Boris Hessen. The social and economic roots of Newton’s Principia. In Freudenthal and McLaughlin [FM09c], chapter 2, pages 41–101. ISBN 1-4020-9603-8, 1-4020-9604-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174.6 .S63 2009.

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-1-4020-9604-
                                              4_2.

Hettema:2015:RDW

[Het15] Hinne Hettema. Reduction for a dappled world: Connecting chemical and physical theories. In Philosophy of Chemistry: Growth of a New Discipline [SM15b], chapter 2, pages 5–22. ISBN 94-017-9363-8 (hardcover), 94-017-9364-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QD6 .P455 2015; QD6 .S384 2014.

URL https://link.springer.com/
chapter/10.1007/978-94-017-9364-
3_2.

Heuser:1996:GPE

[Heu96]  Marie-Luise Heuser.  Geometrical

product — exponentiation — evolution. Justus Günther Grassmann and dynamist Naturphilosophie. In Schubring [Sch96f], pages 47–58. ISBN 90-481-4758-1, 94-015-8753-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QA8.9-10.3.

URL https://link.springer.com/
chapter/10.1007/978-94-015-8753-
2_6.

Heyd:1986:TFS

[Hey86] Michael Heyd. The third force in Seventeenth-Century thought: A comment. In Ullmann-Margalit [UM86c], chapter 4, pages 51–56. ISBN 90-277-2160-2, 90-277-2161-0 (paperback), 94-009-4566-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 95 Q175.

URL https://link.springer.com/
chapter/10.1007/978-94-009-4566-
1_4.

Hronszky:1988:SKS

[HFD88] Imre Hronszky, Márta Fehér, and Balázs Dajka, editors. Scientific knowledge socialized: selected proceedings of the 5th Joint International Conference on the History and Philosophy of Science Organized by the IUHPS, Veszprém, 1984, volume 108 of Boston Studies in the Philosophy of Science. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 1988. ISBN 90-277-2284-6. LCCN Q124.6 .I574 1984; Q174 .B67 vol. 108; Q175.4.

Harris:1989:FIS

[HGR^{+}89] Zellig Harris, Michael Gottfried, Thomas Ryckman, Paul Mattick, Jr., Anne Daladier, T. N. Harris, and S. Harris. The Form of Information in Science: Analysis of an Immunology Sublanguage, volume 104 of Boston Studies in the Philosophy of Science. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 1989. ISBN 94-010-7777-0, 94-009-2837-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). 616 (est.) pp. LCCN B832.A-832.Z.

                                              URL https://link.springer.com/
                                              book/10.1007/978-94-009-2837-4.

Harris:1989:CSA

[HH89] T. N. Harris and Susanna Harris. The cellular source of antibody: A review. In The Form of Information in Science: Analysis of an Immunology Sublanguage [HGR^{+}89], chapter 8, pages 192–215. ISBN 94-010-7777-0, 94-009-2837-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN B832.A-832.Z. URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-009-2837-4_8.

Hecht:1991:BSS

[HH91] Hartmut Hecht and Dieter Hoffmann. The Berlin ‘Society for Scientific Philosophy’ as organizational form of philosophizing in the medium of natural science. In Woodward and Cohen [WC91], pages 75–87. ISBN 94-010-5395-2, 94-011-3164-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-011-3164-3_7.

Hoyningen-Huene:1994:NBA

[HH94] Paul Hoyningen-Huene. Niels Bohr’s argument for the irreducibility of biology to physics. In Faye and Folse [FF94], chapter 10, pages 231–255. ISBN 0-7923-2378-5 (hardcover), 94-015-8106-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 153 1994; QC16 1994.

URL https://link.springer.com/
chapter/10.1007/978-94-015-8106-
6_10.

Herfel:1997:CDD

[HH97] William. E. Herfel and Clifford. A.

Hooker. Cognitive dynamics and the development of science. In Ginev and Cohen [GC97], chapter 9, pages 127– 172. ISBN 0-7923-4444-8 (hardcover), 94-011-5788-X. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 192. URL https://link.

springer.com/chapter/10.1007/
978-94-011-5788-9_9.

Hoyningen-Huene:2008:CBP

[HH08] Paul Hoyningen-Huene. Commentary on Bird’s paper. In Rethinking Scientific Change and Theory Comparison: Stabilities, Ruptures, Incommensurabilities? [SSHH08], chapter 3, pages 41–46. ISBN 1-4020-6274-5 (hardcover), 1-4020-6279-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175 .R397 2008. URL https:

//link.springer.com/chapter/10.
1007/978-1-4020-6279-7_3.

Hoyningen-Huene:2015:KDB

[HH15] Paul Hoyningen-Huene. Kuhn’s development before and after structure. In Kuhn’s Structure of Scientific Revolutions — 50 Years On [DB15b], chapter 13, pages 185–195. ISBN 3-319-13382-9, 3-319-13383-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175.3 .K84 2015.

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-3-319-13383-
                                              6_13.

Hoyningen-Huene:2001:IRM

[HHS01] Paul Hoyningen-Huene and Howard Sankey. Incommensurability and Related Matters, volume 216 of Boston Studies in the Philosophy of Science. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 2001. ISBN 90-481-5709-9, 94-015-9680-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). xxxiv + 337 pp. LCCN ????

                                              URL https://link.springer.com/
                                              book/10.1007/978-94-015-9680-0.

Hiebert:1970:GME

[Hie70] Erwin N. Hiebert. The genesis of Mach’s early views on atomism. In Ernst Mach: Physicist and Philosopher [CS70], chapter 6, pages 79–106. ISBN 90-481-8318-9, 90-277-0016-8, 94-017-1462-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

                                              ???? URL https://link.springer.
                                              com/chapter/10.1007/978-94-017-
                                              1462-4_6.

Hiebert:1994:DBS

[Hie94] Erwin N. Hiebert. On demarcations between science in context and the context of science. In Trends in the Historiography of Science [GCN94], chapter 6, pages 87–105. ISBN 90-481-4264-4, 94-017-3596-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-017-3596-4_6.

Hiebert:1995:EDR

[Hie95a] Erwin N. Hiebert. Electric discharge in rarefied gases: The dominion of experiment. Faraday. Plücker. Hittorf. In No Truth Except in the Details: Essays in Honor of Martin J. Klein [KS95], chapter 5, pages 95–134. ISBN 94-010-4097-4, 94-011-0217-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-011-0217-9_5.

Hiestand:1995:PGW

[Hie95b] Emily Hiestand. Poems from Green the Witch-Hazel Wood and from Alluvial. In Science, Mind and Art: Essays on science and the humanistic understanding in art, epistemology, religion and ethics In honor of Robert S. Cohen [GSW95b], chapter 4, pages 45–54. ISBN 0-7923-2990-2 (hardcover), 94-011-0469-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 165 BD418.3.

URL https://link.springer.com/
chapter/10.1007/978-94-011-0469-
2_4.

Hiebert:2007:RD

[Hie07] Erwin N. Hiebert. Reflections on a discipline. In Positioning the history of science [RG07], chapter 16, pages 87–94. ISBN 1-4020-5419-X (hardcover), 1-4020-5420-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q125 .P77 2007.

URL https://link.springer.com/
chapter/10.1007/1-4020-5420-3_16.

Hilpinen:1989:CCP

[Hil89] Risto Hilpinen. On the characterization of cognitive progress. In Imre Lakatos and Theories of Scientific Change [GGN89], chapter 6, pages 69–80. ISBN 94-010-7860-2, 94-009-3025-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-009-3025-4_6.

Hilts:1994:TSO

[Hil94] Victor L. Hilts. Towards the social organism: Herbert Spencer and William B. Carpenter on the analogical method. In Cohen [Coh94b], chapter 8, pages 275–303. ISBN 0-7923-2223-1, 94-017-3391-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 150. URL https:/

                                              /link.springer.com/chapter/10.
                                              1007/978-94-017-3391-5_8.

Hilpinen:1996:SFR

[Hil96] Risto Hilpinen. On some formulations of realism, or how many objects are there in the world? In Cohen et al. [CHR96], pages 1–10. ISBN 90-481-4493-0, 94-015-8638-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-015-8638-2_1.

Hilbert:2007:FPFa

[Hil07a] David Hilbert. The foundations of physics (first communication). In Renn et al. [RSS^{+}07b], pages 989–1001. ISBN 1-4020-3999-9 (hardcover), 1-4020-4000-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 no. 250 v. 4; QC173.6 .G469

                                              2007. URL https://link.springer.
                                              com/chapter/10.1007/978-1-4020-

4000-9_43. Originally published 20 November 2015.

Hilbert:2007:FPFb

[Hil07b] David Hilbert. The foundations of physics (first communication). In Renn et al. [RSS^{+}07b], pages 1003–1015. ISBN 1-4020-3999-9 (hardcover), 1-4020-4000-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 no. 250 v. 4; QC173.6 .G469

2007. URL https://link.springer.
com/chapter/10.1007/978-1-4020-

4000-9_44. Originally published as “Die Grundlagen der Physik (Erste Mitteilung)”, Nachrichten der Königlichen Gesellschaft der Wissenschaften zu Göttingen, Math.-phys. Klasse, (8) 395–407 (1916).

Hilbert:2007:FPS

[Hil07c] David Hilbert. The foundations of physics (second communication). In Renn et al. [RSS^{+}07b], pages 1017– 1038. ISBN 1-4020-3999-9 (hardcover), 1-4020-4000-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 no. 250 v. 4; QC173.6 .G469

2007. URL https://link.springer.
com/chapter/10.1007/978-1-4020-

4000-9_45. Originally published as “Die Grundlagen der Physik (Zweite Mitteilung)”, Nachrichten der Königlichen Gesellschaft der Wissenschaften zu Göttingen, Math.-phys. Klasse, 53–76 (1917).

Hintikka:1971:UPB

[Hin71] Jaakko Hintikka. Unknown probabilities, Bayesianism, and De Finetti’s Representation Theorem. In Buck and Cohen [BC71], pages 325–341. ISBN 90-277-0309-4, 94-010-3142-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-010-3142-4_22.

Hintikka:1976:CTT

[Hin76] Jaakko Hintikka. A counterexample to Tarski-type truth-definitions as applied to natural languages. In Language in Focus: Foundations, Methods and Systems: Essays in Memory of Yehoshua Bar-Hillel [Kas76b], chapter 6, pages 107–112. ISBN 90-277-0644-1, 90-277-0645-X, 94-010-1876-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN P101-120.

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-010-1876-
                                              0_6.

Hintikka:1989:PTK

[Hin89] Jaakko Hintikka. The paradox of transcendental knowledge. In An Intimate Relation: Studies in the History and Philosophy of Science Presented to Robert E. Butts on his 60th Birthday [BM89], chapter 12, pages 243–257. ISBN 94-010-7546-8, 94-009-2327-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-009-2327-
                                              0_12.

Hintikka:1994:AWP

[Hin94] Jaakko Hintikka. An anatomy of Wittgenstein’s picture theory. In Artifacts, Representations and Social Practice: Essays for Marx Wartofsky [GC94], chapter 14, pages 223–256. ISBN 94-010-4390-6, 94-011-0902-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN B73 .A77

                                              1994. URL https://link.springer.
                                              com/chapter/10.1007/978-94-011-
                                              0902-4_14.

Hintikka:1995:WEL

[Hin95] Jaakko Hintikka. What is elementary logic? Independence-friendly logic as the true core area of logic. In Gavroglu et al. [GSW95c], chapter 18, pages 301–326. ISBN 0-7923-2991-0 (set), 0-7923-2988-0, 94-017-2658-2. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 163.

                                              URL https://link.springer.com/

chapter/10.1007/978-94-017-2658-
0_18.

Hintikka:1997:IPW

[Hin97] Jaakko Hintikka. The idea of phenomenology in Wittgenstein and Husserl. In Austrian Philosophy Past and Present: Essays in Honor of Rudolf Haller [LM97], chapter 8, pages 101–123. ISBN 94-010-6412-1, 94-011-5720-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-011-
5720-9_8.

Hendricks:2006:IMP

[HJLP06] Vincent F. Hendricks, Klaus Frovin Jørgensen, Jesper Lützen, and Stig Andur Pedersen, editors. Interactions: Mathematics, Physics and Philosophy, 1860–1930, volume 251 of Boston Studies in the Philosophy of Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2006. ISBN 1-4020-5194-8 (hardcover), 1-4020-5195-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175 .B69 v.251 2006; Q175 .I51377 2006. URL http://www.

loc.gov/catdir/enhancements/
fy0825/2007425756-d.html;
http://www.loc.gov/catdir/
toc/fy0713/2007425756.html;
https://link.springer.com/book/
10.1007/978-1-4020-5195-1.

Hanen:1992:GAK

[HK92] Marsha Hanen and Jane Kelley.

Gender and archaeological knowledge. In Metaarchaeology: Reflections by Archaeologists and Philosophers [Emb92b], chapter 8, pages 195–225. ISBN 94-010-4806-1, 94-011-1826-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-011-1826-
                                              2_8.

Hasse:2017:BBR

[HL17] Hans Hasse and Johannes Lenhard. Boon and bane: On the role of adjustable parameters in simulation models. In Mathematics as a Tool: Tracing New Roles of Mathematics in the Sciences [LC17b], chapter 6, pages 93–115. ISBN 3-319-54468-3 (hardcover), 3-319-54469-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN GE45.M37; GE45.M38; QA8.4. URL https:/

                                              /link.springer.com/chapter/10.
                                              1007/978-3-319-54469-4_6.

Hankinson-Nelson:2016:FNF

[HN16] Lynn Hankinson-Nelson. Feminist and non-feminist philosophy of biology: Parallels, differences, and prospects for future engagements. In Meta-Philosophical Reflection on Feminist Philosophies of Science [AV16a], chapter 4, pages 55–74. ISBN 3-319-26346-3 (hardcover), 3-319-26348-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN HQ1154 .M48 2016; HQ1190 M48 2016. URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-3-319-26348-9_4.

Hochfeld:1979:CCI

[Hoc79]  Julian Hochfeld.  The concept of

class interest. In Polish Essays in the Methodology of the Social Sciences [Wia79b], chapter 2, pages 21–32. ISBN 90-277-0956-4, 94-009-9353-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .P655 1979.

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-009-9353-
                                              2_2.

Hodes:1974:NRH

[Hod74a] Jane Hodes. A note on Robert Hodes.

In For Dirk Struik: Scientific, Historical and Political Essays in Honor of Dirk J. Struik [CSW74], chapter 28, pages 349–352. ISBN 90-277-0379-5, 94-010-2115-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-010-
2115-9_28.

Hodes:1974:AMS

[Hod74b] Robert Hodes. Aims and methods of scientific research. In For Dirk Struik: Scientific, Historical and Political Essays in Honor of Dirk J. Struik [CSW74], chapter 29, pages 353–364. ISBN 90-277-0379-5, 94-010-2115-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-010-2115-
9_29.

Hodgson:1979:PLS

[Hod79] P. E. Hodgson. Presuppositions and limits of science. In The Structure and Development of Science [RA79], chapter 5, pages 133–147. ISBN 90-277-0995-5, 94-009-9459-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .S778 1979.

URL https://link.springer.com/
chapter/10.1007/978-94-009-9459-
1_5.

Hodge:1990:DSW

[Hod90] M. J. S. Hodge. Darwin studies at work: A re-examination of three decisive years (1835–37). In Levere and Shea [LS90], chapter 11, pages 249–274. ISBN 0-7923-0420-9, 94-009-1878-X. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 120.

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-009-1878-
                                              8_11.

Hodge:1992:BPI

[Hod92] M. J. S. Hodge. Biology and philosophy (including ideology): A study of Fisher and Wright. In Sarkar [Sar92a], chapter 9, pages 231–293. ISBN 0-7923-1777-7 (hardcover), 94-011-2856-1. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 142.

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-011-2856-
                                              8_9.

Hodge:1999:HPN

[Hod99] Joanna Hodge. Habermasian passion and the Nietzschean contagion. In Nietzsche, Theories of Knowledge, and Critical Theory: Nietzsche and the Sciences I [Bab99b], chapter 20, pages 273–279. ISBN 90-481-5233-X, 94-017-2430-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

                                              ???? URL https://link.springer.
                                              com/chapter/10.1007/978-94-017-
                                              2430-2_20.

Hofmann:1974:BUA

[Hof74] J. E. Hofmann. Über Ein Beispiel zur Unbestimmten Analytik und Seine Allgemeine Bedeutung. (German) [On an example of unspecified analytics and its general importance]. In For Dirk Struik: Scientific, Historical and Political Essays in Honor of Dirk J. Struik [CSW74], chapter 7, pages 75–89. ISBN 90-277-0379-5, 94-010-2115-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-010-2115-9_7.

Hoffding:1976:EEN

[Höf76] H. Höffding. Establishment and extension of the new world scheme: Giordano Bruno. In The Concepts of Space and Time: Their Structure and Their Development [Čap76], chapter 12, pages 57–63. ISBN 90-277-0375-2, 94-010-1727-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic).

LCCN BD632. URL https://link.
springer.com/chapter/10.1007/
978-94-010-1727-5_12.

Hoffmann:1992:EMP

[Hof92] Dieter Hoffmann. Ernst Mach in Prague. In Ernst Mach — A Deeper Look: Documents and New Perspectives [Bla92a], chapter 2, pages 29–55. ISBN 94-010-5230-1, 94-011-2771-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN C5.53; QC6.9.

URL https://link.springer.com/
chapter/10.1007/978-94-011-2771-
4_2.

Hoffmann:1998:HHB

[Hof98] Dieter Hoffmann. Heinrich Hertz and the Berlin School of Physics. In Heinrich Hertz: Classical Physicist, Modern Philosopher [BHN98], chapter 1, pages 1–8. ISBN 90-481-4881-2, 94-015-8855-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-015-8855-
3_1.

Hoffmann:2009:SFE

[Hof09] Christoph Hoffmann. The scent of filth:

Experiments, waste, and the set-up. In Going Amiss In Experimental Research [HSS09a], chapter 13, pages 239–251. ISBN 1-4020-8892-2 (hardcover), 1-4020-8893-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q172.5.E77 G647 2009. URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-1-4020-8893-3_13.

Hohlfeld:1983:CRS

[Hoh83] Rainer Hohlfeld. Cancer research. A study of praxis-related theoretical developments in chemistry, the biosciences and medicine. In Finalization in Science: The Social Orientation of Scientific Progress [Sch83b], chapter 5, pages 93–126. ISBN 94-009-7082-X, 94-009-7080-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

                                              ???? URL https://link.springer.
                                              com/chapter/10.1007/978-94-009-
                                              7080-9_5.

Holton:1970:MES

[Hol70] Gerald J. Holton. Mach, Einstein, and the search for reality. In Ernst Mach: Physicist and Philosopher [CS70], chapter 9, pages 165–199. ISBN 90-481-8318-9, 90-277-0016-8, 94-017-1462-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-017-1462-
                                              4_9.

Holz:1983:ISP

[Hol83] Hans Heinz Holz. The import of social, political, and anthropological considerations in an adequate philosophy of technology. In Philosophy and Technology [DR83], chapter 22, pages 319–322. ISBN 94-009-7126-5, 94-009-7124-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-009-7124-
                                              0_22.

Holton:1992:MME

[Hol92] Gerald Holton. More on Mach and Einstein. In Ernst Mach — A Deeper Look: Documents and New Perspectives [Bla92a], chapter 14, pages 263–276. ISBN 94-010-5230-1, 94-011-2771-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN C5.53;

QC6.9. URL https://link.springer.
com/chapter/10.1007/978-94-011-
2771-4_14.

Holmes:1995:CCC

[Hol95a] Frederic L. Holmes. Crystals and carriers: The chemical and physiological identification of hemoglobin. In No Truth Except in the Details: Essays in Honor of Martin J. Klein [KS95], chapter 8, pages 191–243. ISBN 94-010-4097-4, 94-011-0217-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-011-0217-9_8.

Holton:1995:EB

[Hol95b] Gerald Holton. Einstein and books. In No Truth Except in the Details: Essays in Honor of Martin J. Klein [KS95], chapter 11, pages 273–279. ISBN 94-010-4097-4, 94-011-0217-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-011-0217-9_11.

Holton:1995:MPH

[Hol95c] Gerald Holton. Michael Polanyi and the history of science. In Gavroglu et al. [GSW95c], chapter 11, pages 205–223. ISBN 0-7923-2991-0 (set), 0-7923-2988-0, 94-017-2658-2. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 163.

URL https://link.springer.com/
chapter/10.1007/978-94-017-2658-
0_11.

Holland:1996:QMU

[Hol96] Peter R. Holland. Is quantum mechanics universal? In Bohmian Mechanics and Quantum Theory: an Appraisal [CFG96], chapter 7, pages 99–110. ISBN 90-481-4698-4, 94-015-8715-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC173.96-174.52.

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-015-8715-
                                              0_7.

Holmes:1997:BEL

[Hol97] Frederic L. Holmes. Beautiful experiments in the life sciences. In The Elusive Synthesis: Aesthetics and Science [Tau97a], chapter 4, pages 83–101. ISBN 0-7923-4763-3, 0-7923-3904-5, 94-009-1786-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

                                              ???? URL https://link.springer.
                                              com/chapter/10.1007/978-94-009-
                                              1786-6_4.

Holton:2007:WES

[Hol07] Gerald Holton. The woman in Einstein’s shadow. In Positioning the history of science [RG07], chapter 17, pages 95–98. ISBN 1-4020-5419-X (hardcover), 1-4020-5420-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q125 .P77 2007.

                                              URL https://link.springer.com/
                                              chapter/10.1007/1-4020-5420-3_17.

Hon:1993:UNL

[Hon93] Giora Hon. The unnatural nature of the laws of nature: Symmetry and asymmetry. In Correspondence, Invariance and Heuristics: Essays in Honour of Heinz Post [FK93], chapter 9, pages 171–187. ISBN 90-481-4229-6, 94-017-1185-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-017-1185-
2_9.

Honner:1994:DD

[Hon94] John Honner. Description and deconstruction. In Faye and Folse [FF94], chapter 7, pages 141–153. ISBN 0-7923-2378-5 (hardcover), 94-015-8106-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 153 1994; QC16 1994. URL https:

//link.springer.com/chapter/10.
1007/978-94-015-8106-6_7.

Hong:1996:MPA

[Hon96a] Zhou Ji Hong. The mathematicized practical attitude. In Cohen et al. [CHR96], pages 419–426. ISBN 90-481-4493-0, 94-015-8638-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-015-8638-2_31.

Hongsen:1996:NWQ

[Hon96b] Wei Hongsen. Norbert Wiener at Qinghua University. In Chinese Studies in the History and Philosophy of Science and Technology [DC96], chapter 35, pages 447–451. ISBN 90-481-4546-5, 94-015-8717-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 180; U390 .N19 1996. URL https:

//link.springer.com/chapter/10.
1007/978-94-015-8717-4_35.

Hon:1998:HSP

[Hon98] Giora Hon. Hertz’s study of propagation vs. Rutherford’s study of structure: Two modes of experimentation and their theoretical underpinnings. In Heinrich Hertz: Classical Physicist, Modern Philosopher [BHN98], chapter 5, pages 59–72. ISBN 90-481-4881-2, 94-015-8855-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic).

                                              LCCN ????  URL https://link.
                                              springer.com/chapter/10.1007/
                                              978-94-015-8855-3_5.

Hon:2003:APE

[Hon03] Giora Hon. An attempt at a philosophy of experiment. In Descartes’s Mathematical Thought [Sas03d], chapter 18, pages 259–284. ISBN 90-481-6487-7, 94-017-1225-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic).

                                              LCCN ????  URL https://link.
                                              springer.com/chapter/10.1007/
                                              0-306-48123-5_18. See comments

[Cam03, Cam03].

Hon:2009:ELN

[Hon09a] Giora Hon. Error: The long neglect, the one-sided view, and a typology. In Going Amiss In Experimental Research [HSS09a], chapter 2, pages 11–26. ISBN 1-4020-8892-2 (hardcover), 1-4020-8893-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q172.5.E77 G647 2009. URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-1-4020-8893-3_2.

Hon:2009:LEF

[Hon09b] Giora Hon. Living extremely flat: The life of an automaton; John von Neumann’s conception of error of (in)animate systems. In Going Amiss In Experimental Research [HSS09a], chapter 5, pages 55–71. ISBN 1-4020-8892-2 (hardcover), 1-4020-8893-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q172.5.E77

                                              G647 2009.  URL https://link.
                                              springer.com/chapter/10.1007/
                                              978-1-4020-8893-3_5.

Hon:2016:BSP

[Hon16] Giora Hon. “Baseline” and “snapshot”: Philosophical reflections on an approach to historical case studies. In Sauer and Scholl [SS16b], chapter 3, pages 31–47. ISBN 3-319-30227-2, 3-319-30229-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q126.8 .P45 2016. URL https:

//link.springer.com/chapter/10.
1007/978-3-319-30229-4_3.

Hooker:1974:DNC

[Hoo74] C. A. Hooker. Defense of a nonconventionalist interpretation of classical mechanics. In Logical and Epistemological Studies in Contemporary Physics [CW74a], chapter 3, pages 123–191. ISBN 90-277-0391-4 (hardcover), 90-277-0377-9 (paperback), 94-010-2656-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 13; QC5.56.

URL https://link.springer.com/
chapter/10.1007/978-94-010-2656-
7_3.

Hook:1988:TGE

[Hoo88a] Sidney Hook. Toward greater equality.

In Philosophy, History and Social Action: Essays in Honor of Lewis Feuer with an autobiographic essay by Lewis Feuer [HOO88c], chapter 10, pages 235–243. ISBN 94-010-7793-2, 94-009-2873-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-009-
2873-2_10.

Hook:1988:LPW

[HOO88b] Sidney Hook, William L. O’Neill, and Roger O’Toole. List of published writings of Lewis S. Feuer. In Philosophy, History and Social Action:

Essays in Honor of Lewis Feuer with an autobiographic essay by Lewis Feuer [HOO88c], chapter 2, pages 87–101. ISBN 94-010-7793-2, 94-009-2873-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-009-2873-
2_2.

Hook:1988:PHS

[HOO88c] Sidney Hook, William L. O’Neill, and Roger O’Toole. Philosophy, History and Social Action: Essays in Honor of Lewis Feuer with an autobiographic essay by Lewis Feuer, volume 107 of Boston Studies in the Philosophy of Science. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 1988. ISBN 94-010-7793-2, 94-009-2873-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). 494 (est.) pp. LCCN ????

URL https://link.springer.com/
book/10.1007/978-94-009-2873-2.

Hooker:1991:BFA

[Hoo91] C. A. Hooker. Between formalism and anarchism: A reasonable middle way. In Beyond Reason: Essays on the Philosophy of Paul Feyerabend [Mun91a], chapter 3, pages 41–107. ISBN 94-010-5406-1, 94-011-3188-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-011-3188-
9_3.

Hooker:1994:BCE

[Hoo94] Clifford A. Hooker. Bohr and the crisis of empirical intelligibility: An essay on the depth of Bohr’s thought and our philosophical ignorance. In Faye and Folse [FF94], chapter 8, pages 155–199.

ISBN 0-7923-2378-5 (hardcover), 94-015-8106-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 153 1994; QC16 1994.

URL https://link.springer.com/
chapter/10.1007/978-94-015-8106-
6_8.

Hooykaas:1999:SSS

[Hoo99a] R. Hooykaas. And the sun stood still. In Fact, Faith and Fiction in the Development of Science: The Gifford Lectures Given in the University of St Andrews 1976 [Hoo99c], pages 147–181. ISBN 0-7923-5774-4, 94-015-9295-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 205.

URL https://link.springer.com/
chapter/10.1007/978-94-015-9295-
6_7.

Hooykaas:1999:CN

[Hoo99b] R. Hooykaas. Cleopatra’s nose. In Fact, Faith and Fiction in the Development of Science: The Gifford Lectures Given in the University of St Andrews 1976 [Hoo99c], pages 319–341. ISBN 0-7923-5774-4, 94-015-9295-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 205.

URL https://link.springer.com/
chapter/10.1007/978-94-015-9295-
6_11.

Hooykaas:1999:FFF

[Hoo99c] R. Hooykaas, editor. Fact, Faith and Fiction in the Development of Science: The Gifford Lectures Given in the University of St Andrews 1976, volume 205 of Boston Studies in the Philosophy of Science. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 1999. ISBN 0-7923-5774-4, 94-015-9295-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 205.

                                              URL https://link.springer.com/
                                              book/10.1007/978-94-015-9295-6.

Hooykaas:1999:HN

[Hoo99d] R. Hooykaas. Harmony in nature. In Fact, Faith and Fiction in the Development of Science: The Gifford Lectures Given in the University of St Andrews 1976 [Hoo99c], pages 27–77. ISBN 0-7923-5774-4, 94-015-9295-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 205.

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-015-9295-
                                              6_3.

Hooykaas:1999:I

[Hoo99e] R. Hooykaas. Introduction. In Fact, Faith and Fiction in the Development of Science: The Gifford Lectures Given in the University of St Andrews 1976 [Hoo99c], pages 5–16. ISBN 0-7923-5774-4, 94-015-9295-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 205.

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-015-9295-
                                              6_1.

Hooykaas:1999:NT

[Hoo99f] R. Hooykaas. On natural theology. In Fact, Faith and Fiction in the Development of Science: The Gifford Lectures Given in the University of St Andrews 1976 [Hoo99c], pages 17–25. ISBN 0-7923-5774-4, 94-015-9295-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 205.

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-015-9295-
                                              6_2.

Hooykaas:1999:PS

[Hoo99g] R. Hooykaas. The philosopher’s stone. In Fact, Faith and Fiction in the Development of Science: The Gifford Lectures Given in the University of St Andrews 1976 [Hoo99c], pages 79–98. ISBN 0-7923-5774-4, 94-015-9295-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 205. URL https://link.

springer.com/chapter/10.1007/
978-94-015-9295-6_4.

Hooykaas:1999:PMT

[Hoo99h] R. Hooykaas. Physical and mathematical theories. In Fact, Faith and Fiction in the Development of Science: The Gifford Lectures Given in the University of St Andrews 1976 [Hoo99c], pages 229–264. ISBN 0-7923-5774-4, 94-015-9295-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 205. URL https:/

/link.springer.com/chapter/10.
1007/978-94-015-9295-6_9.

Hooykaas:1999:TR

[Hoo99i] R. Hooykaas. The ‘thinking reed’. In Fact, Faith and Fiction in the Development of Science: The Gifford Lectures Given in the University of St Andrews 1976 [Hoo99c], pages 343–372. ISBN 0-7923-5774-4, 94-015-9295-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 205.

URL https://link.springer.com/
chapter/10.1007/978-94-015-9295-
6_12.

Hooykaas:1999:TH

[Hoo99j] R. Hooykaas. Thinking with the hands.

In Fact, Faith and Fiction in the Development of Science: The Gifford Lectures Given in the University of St Andrews 1976 [Hoo99c], pages 183–228. ISBN 0-7923-5774-4, 94-015-9295-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 205. URL https:/

                                              /link.springer.com/chapter/10.
                                              1007/978-94-015-9295-6_8.

Hooykaas:1999:TTE

[Hoo99k] R. Hooykaas. A tunnel through the Earth. In Fact, Faith and Fiction in the Development of Science: The Gifford Lectures Given in the University of St Andrews 1976 [Hoo99c], pages 117–145. ISBN 0-7923-5774-4, 94-015-9295-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 205. URL https:/

                                              /link.springer.com/chapter/10.
                                              1007/978-94-015-9295-6_6.

Hooykaas:1999:UFQ

[Hoo99l] R. Hooykaas. The undying fire qualitative theories of Stahl and Lavoisier. In Fact, Faith and Fiction in the Development of Science: The Gifford Lectures Given in the University of St Andrews 1976 [Hoo99c], pages 99–115. ISBN 0-7923-5774-4, 94-015-9295-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 205.

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-015-9295-
                                              6_5.

Hooykaas:1999:WNW

[Hoo99m] R. Hooykaas. Works of nature, works of art. In Fact, Faith and Fiction in the Development of Science: The Gifford Lectures Given in the University of St Andrews 1976 [Hoo99c], pages 265–318. ISBN 0-7923-5774-4, 94-015-9295-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 205. URL https:/

                                              /link.springer.com/chapter/10.
                                              1007/978-94-015-9295-6_10.

Hooper:2004:SCT

[Hoo04] Wallace Hooper. Seventeenth-century theories of the tides as a gauge of scientific change. In The Reception of the Galilean Science of Motion in Seventeenth-Century Europe [PT04], chapter 11, pages 199–242. ISBN 1-4020-2454-1, 1-4020-2455-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC133 .R43 2004.

URL https://link.springer.com/
chapter/10.1007/978-1-4020-2455-
9_11.

Hopson:1974:OAC

[Hop74] Ronald C. Hopson. The objects of acceptance: Competing scientific explanations. In Schaffner and Cohen [SC74a], pages 349–363. ISBN 90-277-0408-2, 90-277-0409-0, 94-010-2140-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175 .B7312.

URL https://link.springer.com/
chapter/10.1007/978-94-010-2140-
1_24.

Hoquet:2018:BNH

[Hoq18]  Thierry Hoquet.  Buffon’s natural

history: The catalogue of facts and the temptation of system. In What Does it Mean to be an Empiricist?: Empiricisms in Eighteenth Century Sciences [BR18b], chapter 8, pages 139–167. ISBN 3-319-69858-3 (hardcover), 3-319-69860-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN B816 .W43 2018. URL https://link.

springer.com/chapter/10.1007/
978-3-319-69860-1_8.

Horowitz:1988:LWF

[Hor88]  Irving Louis Horowitz.  Left-wing

fascism and right-wing Communism: The fission–fusion effect in American extremist ideologies. In Philosophy, History and Social Action: Essays in Honor of Lewis Feuer with an autobiographic essay by Lewis Feuer [HOO88c], chapter 11, pages 245–266. ISBN 94-010-7793-2, 94-009-2873-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-009-2873-
2_11.

Horz:1989:DSC

[Hör89] Herbert Hörz. Development of science as a change of types. In Imre Lakatos and Theories of Scientific Change [GGN89], chapter 3, pages 33–46. ISBN 94-010-7860-2, 94-009-3025-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-009-3025-4_3.

Horz:1991:DUU

[Hör91a] Herbert Hörz. Dialectical understanding of the unity of scientific knowledge. In Woodward and Cohen [WC91], pages 39–48. ISBN 94-010-5395-2, 94-011-3164-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-011-
3164-3_4.

Horz:1991:ECP

[Hör91b] Herbert Hörz. Ecology as a challenge to philosophy. In Beyond Reason: Essays on the Philosophy of Paul Feyerabend [Mun91a], chapter 20, pages 417–431. ISBN 94-010-5406-1, 94-011-3188-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-011-3188-
9_20.

Horng:1993:CMT

[Hor93] Wann-Sheng Horng. Chinese mathematics at the turn of the 19th Century: Jiao Xun, Wang Lai and Li Rui. In Philosophy and Conceptual History of Science in Taiwan [LF93], chapter 10, pages 167–208. ISBN 94-010-5103-8, 94-011-2500-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q127.T28 .P455 1993. URL https:

//link.springer.com/chapter/10.
1007/978-94-011-2500-0_10.

Horowitz:1994:SOR

[Hor94] Gregg M. Horowitz. “Suddenly one has the right eyes”: Illusion and iconoclasm in the early Gombrich. In Artifacts, Representations and Social Practice: Essays for Marx Wartofsky [GC94], chapter 2, pages 21–38. ISBN 94-010-4390-6, 94-011-0902-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN B73 .A77 1994.

URL https://link.springer.com/
chapter/10.1007/978-94-011-0902-
4_2.

Horz:1995:SHG

[Hör95]  Herbert Hörz.  Schopenhauer und

Helmholtz. In Science, Mind and Art: Essays on science and the humanistic understanding in art, epistemology, religion and ethics In honor of Robert S. Cohen [GSW95b], chapter 8, pages 99– 122. ISBN 0-7923-2990-2 (hardcover), 94-011-0469-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 165 BD418.3.

URL https://link.springer.com/
chapter/10.1007/978-94-011-0469-
2_8.

Horwich:2000:EM

[Hor00] Paul Horwich. On the existence of meanings. In Knowledge, Language and Logic: Questions for Quine [OK00], chapter 12, pages 151–162. ISBN 1-4020-0253-X, 94-011-3933-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-011-3933-5_12.

Horsten:2016:AIC

[Hor16] Leon Horsten. Absolute infinity in class theory and in theology. In Objectivity, Realism, and Proof: FilMat Studies in the Philosophy of Mathematics [BS16], chapter 7, pages 103–122. ISBN 3-319-31642-7, 3-319-31644-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QA9 .O25 2016.

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-3-319-31644-
                                              4_7.

Houng:1993:EMC

[Hou93] Yu-Houng Houng. Eliminative materialism and connectionism. In Philosophy and Conceptual History of Science in Taiwan [LF93], chapter 7, pages 91–119. ISBN 94-010-5103-8, 94-011-2500-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q127.T28 .P455 1993. URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-011-2500-0_7.

Howson:1976:DLP

[How76]  Colin Howson.  The development

of logical probability. In Essays in Memory of Imre Lakatos [CFW76], chapter 18, pages 277–298. ISBN 90-277-0655-7, 94-010-1451-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-010-1451-9_18.

Howson:1979:MNE

[How79] Colin Howson. Methodology in nonempirical disciplines. In The Structure and Development of Science [RA79], chapter 10, pages 257–266. ISBN 90-277-0995-5, 94-009-9459-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .S778 1979.

URL https://link.springer.com/
chapter/10.1007/978-94-009-9459-
1_10.

Howard:1994:WMC

[How94] Don Howard. What makes a classical concept classical? In Faye and Folse [FF94], chapter 9, pages 201–229. ISBN 0-7923-2378-5 (hardcover), 94-015-8106-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 153 1994; QC16 1994.

URL https://link.springer.com/
chapter/10.1007/978-94-015-8106-
6_9.

Howard:1997:STS

[How97] Don Howard. Space–time and separability: Problems of identity and individuation in fundamental physics. In Cohen et al. [CHS97b], pages 113–141. ISBN 90-481-4809-X, 94-017-2732-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC173.96-174.52. URL https://link.

springer.com/chapter/10.1007/
978-94-017-2732-7_9.

Howson:2003:BE

[How03] Colin Howson. Bayesian evidence.

In Descartes’s Mathematical Thought [Sas03d], chapter 21, pages 301–320. ISBN 90-481-6487-7, 94-017-1225-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/0-306-48123-5_21.

Howes:2015:OIV

[How15] Moira Howes. Objectivity, intellectual virtue, and community. In Objectivity in Science: New Perspectives from Science and Technology Studies [PRT15], chapter 9, pages 173–188. ISBN 3-319-14348-4 (hardcover), 3-319-14349-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175 O25 2015. URL https://link.

                                              springer.com/chapter/10.1007/
                                              978-3-319-14349-1_9.

Hoyle:1974:HCO

[Hoy74] Fred Hoyle. Highly condensed objects. In Seeger and Cohen [SC74b], pages 215–226. ISBN 90-277-0390-6, 90-277-0376-0 (paperback), 94-010-2126-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 11. URL https://link.

                                              springer.com/chapter/10.1007/
                                              978-94-010-2126-5_13.

Hinitikka:1974:AAFb

[HR74a] Jaakko Hinitikka and Unto Remes. Analysis as analysis of figures: Pappus’ terminology and his practice. In The Method of Analysis: Its Geometrical Origin and Its General Significance [HR74d], chapter 7, pages 70–83. ISBN 90-277-0543-7, 94-010-2296-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-010-2296-5_7.

Hinitikka:1974:AAFa

[HR74b] Jaakko Hinitikka and Unto Remes. Analysis as analysis of figures: The logic of the analytical method. In The Method of Analysis: Its Geometrical Origin and Its General Significance [HR74d], chapter 4, pages 31–40. ISBN 90-277-0543-7, 94-010-2296-8 (e-book).

ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-010-2296-5_4.

Hinitikka:1974:HSI

[HR74c] Jaakko Hinitikka and Unto Remes.

The historical significance of the idea of geometrical analysis. In The Method of Analysis: Its Geometrical Origin and Its General Significance [HR74d], chapter 1, pages 1–6. ISBN 90-277-0543-7, 94-010-2296-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-010-2296-5_1.

Hinitikka:1974:MAG

[HR74d] Jaakko Hinitikka and Unto Remes. The Method of Analysis: Its Geometrical Origin and Its General Significance, volume 25 of Boston Studies in the Philosophy of Science. D. Reidel, Dordrecht, The Netherlands; Boston, MA, USA; Lancaster, UK; Tokyo, Japan, 1974. ISBN 90-277-0543-7, 94-010-2296-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). xviii + 149 pp. LCCN ???? URL

https://link.springer.com/book/
10.1007/978-94-010-2296-5.

Hinitikka:1974:SMA

[HR74e] Jaakko Hinitikka and Unto Remes.

On the significance of the method of analysis in early modern science. In The Method of Analysis: Its Geometrical Origin and Its General Significance [HR74d], chapter 9, pages 105–117. ISBN 90-277-0543-7, 94-010-2296-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-010-2296-
5_9.

Hinitikka:1974:PTG

[HR74f] Jaakko Hinitikka and Unto Remes. Pappus and the tradition of geometrical analysis. In The Method of Analysis: Its Geometrical Origin and Its General Significance [HR74d], chapter 8, pages 84–104. ISBN 90-277-0543-7, 94-010-2296-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

                                              ???? URL https://link.springer.
                                              com/chapter/10.1007/978-94-010-
                                              2296-5_8.

Hinitikka:1974:PDA

[HR74g] Jaakko Hinitikka and Unto Remes. Pappus on the direction of analysis and synthesis. In The Method of Analysis: Its Geometrical Origin and Its General Significance [HR74d], chapter 2, pages 7–21. ISBN 90-277-0543-7, 94-010-2296-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

                                              ???? URL https://link.springer.
                                              com/chapter/10.1007/978-94-010-
                                              2296-5_2.

Hinitikka:1974:PR

[HR74h] Jaakko Hinitikka and Unto Remes. The problem of the ‘resolution’. In The Method of Analysis: Its Geometrical Origin and Its General Significance [HR74d], chapter 6, pages 49–69. ISBN 90-277-0543-7, 94-010-2296-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-010-2296-5_6.

Hinitikka:1974:RAC

[HR74i] Jaakko Hinitikka and Unto Remes. The role of auxiliary constructions. In The Method of Analysis: Its Geometrical Origin and Its General Significance [HR74d], chapter 5, pages 41–48. ISBN 90-277-0543-7, 94-010-2296-8 (e-book).

ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-010-2296-5_5.

Hinitikka:1974:WPS

[HR74j] Jaakko Hinitikka and Unto Remes.

What Pappus says and what he does: A comparison and an example. In The Method of Analysis: Its Geometrical Origin and Its General Significance [HR74d], chapter 3, pages 22–30. ISBN 90-277-0543-7, 94-010-2296-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-010-2296-5_3.

Hintikka:1976:AGA

[HR76] Jaakko Hintikka and Unto Remes. Ancient geometrical analysis and modern logic. In Essays in Memory of Imre Lakatos [CFW76], chapter 17, pages 253–276. ISBN 90-277-0655-7, 94-010-1451-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-010-
1451-9_17.

Hemmo:2016:AT

[HS16] Meir Hemmo and Orly Shenker. The arrow of time. In Cosmological and Psychological Time [DR16], chapter 9, pages 155–162. ISBN 3-319-22589-8 (hardcover), 3-319-22590-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN BD638 C67 2016.

URL https://link.springer.com/
chapter/10.1007/978-3-319-22590-
6_9.

Hon:2009:GAE

[HSS09a] Giora Hon, Jutta Schickore, and Friedrich Steinle. Going Amiss In Experimental Research, volume 267 of Boston Studies in the Philosophy of Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2009. ISBN 1-4020-8892-2 (hardcover), 1-4020-8893-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). x + 277 pp. LCCN Q172.5.E77 G647 2009. URL

https://link.springer.com/book/
10.1007/978-1-4020-8893-3.

Hon:2009:IMG

[HSS09b] Giora Hon, Jutta Schickore, and Friedrich Steinle. Introduction: Mapping “going amiss”. In Going Amiss In Experimental Research [HSS09a], chapter 1, pages 1–7. ISBN 1-4020-8892-2 (hardcover), 1-4020-8893-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q172.5.E77 G647

2009. URL https://link.springer.
com/chapter/10.1007/978-1-4020-
8893-3_1.

Hsu:1992:RWM

[Hsü92]  Elisabeth Hsü.  The reception of

Western medicine in China: Examples from Yunnan. In Science and Empires: Historical Studies about Scientific Development and European Expansion [PJM92], chapter 11, pages 89–101. ISBN 0-7923-1518-9 (hardcover), 94-010-5145-3, 94-011-2594-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 136.

URL https://link.springer.com/
chapter/10.1007/978-94-011-2594-
9_11.

Handy:2014:ACP

[HTG14] Carlos R. Handy and Carlos Trallero-Giner. Accomplishments in Cuban physics (up to 1995). In The History of Physics in Cuba [BRW14a], chapter 7, pages 235–245. ISBN 94-017-8040-4, 94-017-8041-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC5.53; QC6.9; QC9 C9H57 2014; Q124.6-127.2. URL https:/

/link.springer.com/chapter/10.
1007/978-94-017-8041-4_7.

Hubner:1976:DRI

[Hüb76] Kurt Hübner. Descartes’ rules of impact and their criticism. In Essays in Memory of Imre Lakatos [CFW76], chapter 19, pages 299–310. ISBN 90-277-0655-7, 94-010-1451-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-010-1451-9_19.

Hubner:1978:RWH

[Hüb78a] Kurt Hübner. Reply to Watkins’:

Hübner on falsification, “grue” and truth’. In Progress and Rationality in Science [RA78b], chapter 18, pages 393–396. ISBN 90-277-0922-X, 94-009-9866-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-009-
9866-7_18.

Hubner:1978:SCC

[Hüb78b] Kurt Hübner. Some critical comments on current Popperianism on the basis of a theory of system sets. In Progress and Rationality in Science [RA78b], chapter 12, pages 279–289. ISBN 90-277-0922-X, 94-009-9866-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-009-9866-7_12.

Hughes:1988:EII

[Hug88] Thomas P. Hughes. The era of independent inventors. In Science in Reflection: The Israel Colloquium: Studies in History, Philosophy, and Sociology of Science Volume 3 [UM88], chapter 13, pages 151–168. ISBN 90-277-2713-9, 94-009-2957-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-009-2957-9_13.

Hull:1976:CSP

[Hul76a] David L. Hull. Contemporary systematic philosophies. In Topics in the Philosophy of Biology [GM76], chapter 18, pages 396–440. ISBN 90-277-0596-8, 94-010-1829-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-010-1829-6_18.

Hull:1976:IAT

[Hul76b] David L. Hull. Informal aspects of theory reduction. In Cohen et al. [CHMV76], pages 653–670. ISBN 90-277-0647-6, 90-277-0648-4 (paperback), 94-010-1449-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic).

LCCN ????  URL https://link.
springer.com/chapter/10.1007/
978-94-010-1449-6_37.

Hultenschmidt:1996:HGC

[Hül96] Erika Hültenschmidt. Hermann Grassmann’s contribution to the construction of a German ,,Kulturnation“. Scientific school grammar between Latin tradition and French conceptions. In Schubring [Sch96f], pages 87–113. ISBN 90-481-4758-1, 94-015-8753-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QA8.9-10.3.

URL https://link.springer.com/
chapter/10.1007/978-94-015-8753-
2_9.

Humphreys:2011:CSE

[Hum11]  Paul Humphreys.  Computational

science and its effects. In Science in the Context of Application [CN11a], chapter 9, pages 131–142. ISBN 90-481-9050-9, 90-481-9051-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q180.55.M4 S385

2011. URL https://link.springer.
com/chapter/10.1007/978-90-481-
9051-5_9.

Huning:1983:THR

[Hun83] Alois Huning. Technology and human rights. In Philosophy and Technology [DR83], chapter 4, pages 49–57. ISBN 94-009-7126-5, 94-009-7124-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-009-7124-0_4.

Hunger:2006:HCM

[Hun06] Johannes Hunger. How classical models of explanation fail to cope with chemistry. In Philosophy of Chemistry: Synthesis of a New Discipline [BSM06b], chapter 7, pages 129–156. ISBN 1-4020-2455-X, 1-4020-3256-0, 1-4020-3261-7. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 242 QD6. URL https:/

/link.springer.com/chapter/10.
1007/1-4020-3261-7_7.

Hunter:2010:ETR

[Hun10] Matthew C. Hunter. Experiment, theory, representation: Robert Hooke’s material models. In Frigg and Hunter [FH10b], chapter 9, pages 193–219. ISBN 90-481-3850-7 (hardcover), 90-481-3851-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN BH301.M47 B49 2010; Q175 .B52 2010.

URL https://link.springer.com/
chapter/10.1007/978-90-481-3851-
7_9.

Huneman:2017:MMI

[Hun17] Philippe Huneman. Macroevolution and microevolution: Issues of time scale in evolutionary biology. In Time of Nature and the Nature of Time: Philosophical Perspectives of Time in Natural Sciences [BH17b], chapter 14, pages 315–357. ISBN 3-319-53723-7 (hardcover), 3-319-53725-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC173.59.S65 T56

2017. URL https://link.springer.
com/chapter/10.1007/978-3-319-
53725-2_14.

Huss:2017:POT

[Hus17] John Huss. Paleontology: Outrunning time. In Time of Nature and the Nature of Time: Philosophical Perspectives of Time in Natural Sciences [BH17b], chapter 10, pages 211–235. ISBN 3-319-53723-7 (hardcover), 3-319-53725-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC173.59.S65 T56 2017. URL https:

//link.springer.com/chapter/10.
1007/978-3-319-53725-2_10.

Hyder:2006:KHD

[Hyd06] David Hyder. Kant, Helmholtz and the determinacy of physical notes. In Hendricks et al. [HJLP06], pages 1–44. ISBN 1-4020-5194-8 (hardcover), 1-4020-5195-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175 .B69 v.251 2006; Q175 .I51377

2006. URL https://link.springer.
com/chapter/10.1007/978-1-4020-
5195-1_1.

Hyman:2010:AN

[Hym10] John Hyman. Art and neuroscience. In Frigg and Hunter [FH10b], chapter 11, pages 245–261. ISBN 90-481-3850-7 (hardcover), 90-481-3851-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN BH301.M47 B49 2010; Q175 .B52 2010. URL https:

//link.springer.com/chapter/10.
1007/978-90-481-3851-7_11.

Idel:2000:DSN

[Ide00] Moshe Idel. Deus Sive Natura — the metamorphosis of a dictum from Maimonides to Spinoza. In Maimonides and the Sciences [CL00], chapter 6, pages 87–110. ISBN 90-481-5348-4, 94-017-2128-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-017-
2128-8_6.

Intemann:2016:FVC

[IdMM16] Kristen Intemann and Inmaculada de Melo-Martín. Feminist values, commercial values, and the bias paradox in biomedical research. In Meta-Philosophical Reflection on Feminist Philosophies of Science [AV16a], chapter 5, pages 75–89. ISBN 3-319-26346-3 (hardcover), 3-319-26348-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN HQ1154 .M48 2016; HQ1190 M48 2016. URL https:

//link.springer.com/chapter/10.
1007/978-3-319-26348-9_5.

Ierodiakonou:1990:RSS

[Ier90] Katerina Ierodiakonou. Rediscovering some stoic arguments. In Greek Studies in the Philosophy and History of Science [Nic90], chapter 9, pages 137–148. ISBN 94-010-7403-8, 94-009-2015-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-009-2015-
6_9.

Irzik:2005:IMR

[IG05a] Gürol Irzik and Güven Güzeldere. An interview with Maria Reichenbach and David Kaplan. In Turkish Studies in the History and Philosophy of Science [IG05c], chapter 2, pages 7–24. ISBN 1-4020-3332-X (hardcover), 1-4020-3333-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 244.

URL https://link.springer.com/
chapter/10.1007/1-4020-3333-8_2.

Irzik:2005:IR

[IG05b] Gürol Irzik and Güven Güzeldere. Introductory remarks. In Turkish Studies in the History and Philosophy of Science [IG05c], chapter 1, pages 1–5. ISBN 1-4020-3332-X (hardcover), 1-4020-3333-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 244.

URL https://link.springer.com/
chapter/10.1007/1-4020-3333-8_1.

Irzik:2005:TSH

[IG05c] Gürol Irzik and Güven Güzeldere. Turkish Studies in the History and Philosophy of Science, volume 244 of Boston Studies in the Philosophy of Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2005. ISBN 1-4020-3332-X (hardcover), 1-4020-3333-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). xvii + 308 pp. LCCN Q174 .B67 vol. 244. URL

http://d-nb.info/1079906622/34;
 https://link.springer.com/book/
10.1007/1-4020-3333-8.

Ihde:1979:BRM

[Ihd79a] Don Ihde. Bach to rock, a musical odyssey. In Technics and Praxis [Ihd79i], chapter 8, pages 93–100. ISBN 90-277-0954-8, 94-009-9900-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-009-9900-8_8.

Ihde:1979:EIC

[Ihd79b] Don Ihde. The existential import of computer technology. In Technics and Praxis [Ihd79i], chapter 5, pages 53–65. ISBN 90-277-0954-8, 94-009-9900-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-009-9900-
8_5.

Ihde:1979:ETH

[Ihd79c] Don Ihde. The experience of technology: Human–machine relations. In Technics and Praxis [Ihd79i], chapter 1, pages 3–15. ISBN 90-277-0954-8, 94-009-9900-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-009-
9900-8_1.

Ihde:1979:HPT

[Ihd79d] Don Ihde. Heidegger’s philosophy of technology. In Technics and Praxis [Ihd79i], chapter 9, pages 103–129. ISBN 90-277-0954-8, 94-009-9900-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-009-9900-
8_9.

Ihde:1979:PIa

[Ihd79e] Don Ihde. A phenomenology of instrumentation. In Technics and Praxis [Ihd79i], chapter 2, pages 16–27. ISBN 90-277-0954-8, 94-009-9900-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-009-9900-8_2.

Ihde:1979:PIb

[Ihd79f] Don Ihde. A phenomenology of instrumentation. In Technics and Praxis [Ihd79i], chapter 3, pages 28–39. ISBN 90-277-0954-8, 94-009-9900-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-009-9900-8_3.

Ihde:1979:PIc

[Ihd79g] Don Ihde. A phenomenology of instrumentation. In Technics and Praxis [Ihd79i], chapter 4, pages 40–50. ISBN 90-277-0954-8, 94-009-9900-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-009-9900-8_4.

Ihde:1979:SCE

[Ihd79h] Don Ihde. The secular city and the existentialists. In Technics and Praxis [Ihd79i], chapter 11, pages 141–150. ISBN 90-277-0954-8, 94-009-9900-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-009-9900-
                                              8_11.

Ihde:1979:TP

[Ihd79i]  Don Ihde.  Technics and Praxis,

volume 24 of Boston Studies in the Philosophy of Science. D. Reidel, Dordrecht, The Netherlands; Boston, MA, USA; Lancaster, UK; Tokyo, Japan, 1979. ISBN 90-277-0954-8, 94-009-9900-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). xxviii + 163 pp. LCCN ???? URL

https://link.springer.com/book/
10.1007/978-94-009-9900-8.

Ihde:1979:THH

[Ihd79j] Don Ihde. Technology and the human:

Hans Jonas. In Technics and Praxis [Ihd79i], chapter 10, pages 130–140. ISBN 90-277-0954-8, 94-009-9900-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-009-9900-
8_10.

Ihde:1979:TTE

[Ihd79k] Don Ihde. Technology and the transformation of experience. In Technics and Praxis [Ihd79i], chapter 6, pages 66–81. ISBN 90-277-0954-8, 94-009-9900-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-009-9900-
8_6.

Ihde:1979:VO

[Ihd79l] Don Ihde. Vision and objectification. In Technics and Praxis [Ihd79i], chapter 7, pages 82–92. ISBN 90-277-0954-8, 94-009-9900-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-009-
9900-8_7.

Ihde:1983:HOP

[Ihd83] Don Ihde. The historical–ontological priority of technology over science. In Philosophy and Technology [DR83], chapter 16, pages 235–252. ISBN 94-009-7126-5, 94-009-7124-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-009-7124-0_16.

Ihde:1999:EH

[Ihd99a] Don Ihde. Expanding hermeneutics. In Fehér et al. [FKR99], pages 345–351. ISBN 90-481-5257-7, 94-015-9293-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-015-9293-
                                              2_28.

Ihde:1999:PRH

[Ihd99b] Don Ihde. Perceptual reasoning — hermeneutics and perception. In Fehér et al. [FKR99], pages 13–23. ISBN 90-481-5257-7, 94-015-9293-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-015-9293-2_2.

Ihsanoglu:1992:OES

[Ihs92]  Ekmeleddin Ihsanoğlu.  Ottomans

and European science. In Science and Empires: Historical Studies about Scientific Development and European Expansion [PJM92], chapter 6, pages 37–48. ISBN 0-7923-1518-9 (hardcover), 94-010-5145-3, 94-011-2594-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 136.

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-011-2594-
                                              9_6.

Ihsanoglu:2005:ISM

[İhs05] Ekmeleddin İhsanoğlu. Institutionalisation of science in the medreses of Pre-Ottoman and Ottoman Turkey. In Turkish Studies in the History and Philosophy of Science [IG05c], chapter 18, pages 265–283. ISBN 1-4020-3332-X (hardcover), 1-4020-3333-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 244.

URL https://link.springer.com/
chapter/10.1007/1-4020-3333-8_18.

Imai:2001:MBM

[Ima01] Michio Imai. Musil between Mach and Stumpf. In Blackmore et al. [BIT01c], pages 187–209. ISBN 90-481-5752-8, 94-015-9690-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-015-
9690-9_9.

Imocrante:2015:DMM

[Imo15] Marina Imocrante. Defending Maddy’s mathematical naturalism from Roland’s criticisms: The role of mathematical depth. In From Logic to Practice: Italian Studies in the Philosophy of Mathematics [LPV15], chapter 12, pages 223–239. ISBN 3-319-10433-0, 3-319-10434-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QA8.4 .F76 2015; Q174 .B67 v.308.

URL https://link.springer.com/
chapter/10.1007/978-3-319-10434-
8_12.

Imocrante:2016:EOA

[Imo16]  Marina Imocrante.  Epistemology,

ontology and application in Pincock’s account. In Objectivity, Realism, and Proof: FilMat Studies in the Philosophy of Mathematics [BS16], chapter 6, pages 81–99. ISBN 3-319-31642-7, 3-319-31644-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QA9 .O25 2016. URL https:

//link.springer.com/chapter/10.
1007/978-3-319-31644-4_6.

Inan:2005:DIR

[Ina05] Ilhan Inan. Discovery and inostensible De Re knowledge. In Turkish Studies in the History and Philosophy of Science [IG05c], chapter 11, pages 153–161. ISBN 1-4020-3332-X (hardcover), 1-4020-3333-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 244.

                                              URL https://link.springer.com/
                                              chapter/10.1007/1-4020-3333-8_11.

Ionicioiu:2015:QMK

[Ion15] Radu Ionicioiu. Quantum mechanics: Knocking at the gates of mathematical foundations. In Romanian Studies in Philosophy of Science [PST15], chapter 11, pages 167–179. ISBN 3-319-16654-9 (hardcover), 3-319-16655-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-3-319-16655-
                                              1_11.

Irurozqui:2001:DAP

[Iru01] Marta Irurozqui. “Desvío al Paraíso”: Citizenship and social Darwinism in Bolivia, 1880–1920. In The Reception of Darwinism in the Iberian World: Spain, Spanish America and Brazil [GPSR01], chapter 13, pages 205–227. ISBN 94-010-3885-6, 94-010-0602-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-010-0602-
                                              6_13.

Irzik:1995:PEW

[Irz95] Gürol Irzik. Popper’s epistemology and world three. In The Concept of Knowledge: The Ankara Seminar [KC95], chapter 7, pages 83–95. ISBN 90-481-4495-7, 94-017-3263-9 (e-book).

ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-017-3263-5_7.

Irzik:2003:ISE

[Irz03] Gürol Irzik. Improving “styles of experimentation” a comment on Ursula Klein. In Descartes’s Mathematical Thought [Sas03d], chapter 13, pages 203–207. ISBN 90-481-6487-7, 94-017-1225-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/0-306-48123-5_13.

See [Kle03].

Irzik:2015:IED

[Irz15] Gürol Irzik. The internal–external distinction sheds light on the history of the Twentieth-Century philosophy of science. In Arabatzis et al. [ARS15], chapter 14, pages 211–223. ISBN 3-319-14552-5 (paperback), 3-319-14553-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q125 .R45

2015. URL https://link.springer.
com/chapter/10.1007/978-3-319-
14553-2_14.

Ismael:2016:PTH

[Ism16] Jenann Ismael. From physical time to human time. In Cosmological and Psychological Time [DR16], chapter 6, pages 107–124. ISBN 3-319-22589-8 (hardcover), 3-319-22590-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN BD638 C67 2016.

URL https://link.springer.com/
chapter/10.1007/978-3-319-22590-
6_6.

Ismael:2017:PFL

[Ism17] Jenann Ismael. Passage, flow, and the logic of temporal perspectives. In Time of Nature and the Nature of Time: Philosophical Perspectives of Time in Natural Sciences [BH17b], chapter 2, pages 23–38. ISBN 3-319-53723-7 (hardcover), 3-319-53725-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC173.59.S65 T56

2017. URL https://link.springer.
com/chapter/10.1007/978-3-319-
53725-2_2.

Itagaki:1992:TBR

[Ita92] Ryoichi Itagaki. Three batches of reasons for Mach’s rejection of Einstein’s theory of relativity. In Ernst Mach — A Deeper Look: Documents and New Perspectives [Bla92a], chapter 15, pages 277–295. ISBN 94-010-5230-1, 94-011-2771-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN C5.53; QC6.9. URL https:

//link.springer.com/chapter/10.
1007/978-94-011-2771-4_15.

Ito:1998:WPG

[Ito98] Shuntaro Ito. Who are precursors of Galileo in his Pisan dynamics? — a criticism of Professor Moody’s paper. In Japanese Studies in the Philosophy of Science [NC98], chapter 11, pages 159–171. ISBN 94-010-6176-9, 94-011-5175-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-011-
5175-7_11.

Ito:2011:HEE

[Ito11] Akiko Ito. How electricity energizes the body: Electrotherapeutics and its analogy of life in the Japanese medical context. In Science between Europe and Asia: Historical Studies on the Transmission, Adoption and Adaptation of Knowledge [GR11b], chapter 16, pages 245–258. ISBN 90-481-9968-9 (e-book).

ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q125 .S354 2010; Q175.52.A78 S35 2010. URL https:

//link.springer.com/chapter/10.
1007/978-90-481-9968-6_16.

Izquierdo:1974:ORS

[Izq74] J. J. Izquierdo. Observational, rational and scientific medicine in Mexico. In For Dirk Struik: Scientific, Historical and Political Essays in Honor of Dirk J. Struik [CSW74], chapter 17, pages 205–211. ISBN 90-277-0379-5, 94-010-2115-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-010-
2115-9_17.

Jacquette:1997:HWK

[Jac97] Dale Jacquette. Haller on Wittgenstein and Kant. In Austrian Philosophy Past and Present: Essays in Honor of Rudolf Haller [LM97], chapter 3, pages 29–44. ISBN 94-010-6412-1, 94-011-5720-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-011-5720-
9_3.

Jacobsen:2007:PDR

[Jac07] Anja Skaar Jacobsen. Phrenology and Danish romanticism. In Hans Christian Ørsted and The Romantic Legacy In Science: Ideas, Disciplines, Practices [BCK07], chapter 4, pages 55–74. ISBN 1-4020-2979-9 (hardcover), 1-4020-2987-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC515.O37 H36 2007. URL https:

//link.springer.com/chapter/10.
1007/978-1-4020-2987-5_4.

Jaffe:2002:VVR

[Jaf02] Irma B. Jaffe. Virtue and virtual reality in John Trumbull’s Pantheon. In Hermeneutic Philosophy of Science, Van Gogh’s Eyes, and God: Essays in Honor of Patrick A. Heelan, S.J. [Bab02c], chapter 24, pages 287–298. ISBN 90-481-5926-1, 94-017-1767-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 225. URL https://link.

                                              springer.com/chapter/10.1007/
                                              978-94-017-1767-0_24.

Jahn:1991:IJF

[Jah91] Ilse Jahn. The influence of Jakob Friedrich Fries on Matthias Schleiden. In Woodward and Cohen [WC91], pages 357–365. ISBN 94-010-5395-2, 94-011-3164-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic).

                                              LCCN ????  URL https://link.
                                              springer.com/chapter/10.1007/
                                              978-94-011-3164-3_33.

Jahn:1994:ORB

[Jah94] Ilse Jahn. On the origin of romantic biology and its further development at the University of Jena between 1790 and 1850. In Poggi and Bossi [PB94], chapter 5, pages 75–89. ISBN 0-7923-2336-X, 94-017-2921-2. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 152. URL https:/

                                              /link.springer.com/chapter/10.
                                              1007/978-94-017-2921-5_5.

Jammer:1976:GEA

[Jam76] M. Jammer. Gradual emancipation from Aristotle: From Crescas to Gilbert. In The Concepts of Space and Time: Their Structure and Their Development [Čap76], chapter 13, pages 65–72. ISBN 90-277-0375-2, 94-010-1727-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic).

LCCN BD632. URL https://link.
springer.com/chapter/10.1007/
978-94-010-1727-5_13.

Jammer:1979:CPI

[Jam79] Max Jammer. A consideration of the philosophical implications of the new physics. In The Structure and Development of Science [RA79], chapter 3, pages 41–61. ISBN 90-277-0995-5, 94-009-9459-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .S778 1979. URL https:

//link.springer.com/chapter/10.
1007/978-94-009-9459-1_3.

Jami:1992:WMC

[Jam92] Catherine Jami. Western mathematics in China, Seventeenth Century and Nineteenth Century. In Science and Empires: Historical Studies about Scientific Development and European Expansion [PJM92], chapter 10, pages 79–88. ISBN 0-7923-1518-9 (hardcover), 94-010-5145-3, 94-011-2594-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 136.

URL https://link.springer.com/
chapter/10.1007/978-94-011-2594-
9_10.

Jami:2010:ISW

[Jam10] Catherine Jami. Imperial science written in Manchu in Early Qing China: Does it matter? In Looking at it from Asia: the Processes that Shaped the Sources of History of Science [BE10b], chapter 12, pages 371–391. ISBN 90-481-3675-X (hardcover), 90-481-3676-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q127.A65 L66

2010. URL https://link.springer.
com/chapter/10.1007/978-90-481-
3676-6_12.

Janis:1983:SC

[Jan83] Allen I. Janis. Simultaneity and conventionality. In Cohen and Laudan [CL83], chapter 5, pages 101–110. ISBN 94-009-7055-2, 94-009-7057-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-009-7055-7_5.

Janich:1984:CPT

[Jan84a] Peter Janich. Commentary on ‘protophysics of time and the principle of relativity’. In Physical Sciences and History of Physics [CW84c], chapter 10, pages 191–197. ISBN 94-009-7180-X, 94-009-7178-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .P497 1984. URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-009-7178-3_10.

Janik:1984:CJH

[Jan84b] Allan Janik. Comments on Jost Halfmann’s ‘Dethroning the Philosophy of Science: Ideological and Technical Functions of the Metasciences’. In Methodology, Metaphysics and the History of Science: In Memory of Benjamin Nelson [CW84b], chapter 6, pages 173–181. ISBN 94-009-6333-5, 94-009-6331-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

                                              ???? URL https://link.springer.
                                              com/chapter/10.1007/978-94-009-
                                              6331-3_6.

Janich:1985:C

[Jan85a]  Peter Janich.  Chronometry.  In

Protophysics of Time: Constructive Foundation and History of Time Measurement [Jan85e], chapter 3, pages 87–164. ISBN 94-010-8794-6, 94-009-5189-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-009-
5189-1_3.

Janich:1985:HC

[Jan85b] Peter Janich. On a history of chronometry. In Protophysics of Time: Constructive Foundation and History of Time Measurement [Jan85e], chapter 4, pages 165–213. ISBN 94-010-8794-6, 94-009-5189-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-009-
5189-1_4.

Janich:1985:MP

[Jan85c] Peter Janich. On the method of physics. In Protophysics of Time: Constructive Foundation and History of Time Measurement [Jan85e], chapter 2, pages 30–86. ISBN 94-010-8794-6, 94-009-5189-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-009-
5189-1_2.

Janich:1985:PCP

[Jan85d] Peter Janich. On the problem of chronometry in the present–day theory of science. In Protophysics of Time: Constructive Foundation and History of Time Measurement [Jan85e], chapter 1, pages 1–29. ISBN 94-010-8794-6, 94-009-5189-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-009-
5189-1_1.

Janich:1985:PTC

[Jan85e] Peter Janich. Protophysics of Time:

Constructive Foundation and History of Time Measurement, volume 30 of Boston Studies in the Philosophy of Science. D. Reidel, Dordrecht, The Netherlands; Boston, MA, USA; Lancaster, UK; Tokyo, Japan, 1985. ISBN 94-010-8794-6, 94-009-5189-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). xxvi + 241 pp. LCCN ????

URL https://link.springer.com/
book/10.1007/978-94-009-5189-1.

Janik:1989:BEC

[Jan89a]  Allan Janik.  Between Enlighten-

ment and counter-Enlightenment: The self-critical rationalism of G. C. Lichtenberg. In Style, Politics and the Future of Philosophy [Jan89l], chapter 13, pages 197–210. ISBN 94-010-7508-5, 94-009-2251-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-009-2251-8_13.

Janik:1989:CCH

[Jan89b] Allan Janik. Culture, controversy and the human studies. In Style, Politics and the Future of Philosophy [Jan89l], chapter 6, pages 109–115. ISBN 94-010-7508-5, 94-009-2251-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-009-2251-8_6.

Janik:1989:DTB

[Jan89c] Allan Janik. Discussing technology — breaking the ground. In Style, Politics and the Future of Philosophy [Jan89l], chapter 8, pages 133–145. ISBN 94-010-7508-5, 94-009-2251-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-009-2251-8_8.

Janik:1989:MC

[Jan89d] Allan Janik. Myth and certainty.

In Style, Politics and the Future of Philosophy [Jan89l], chapter 10, pages 159–171. ISBN 94-010-7508-5, 94-009-2251-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-009-
2251-8_10.

Janik:1989:NCW

[Jan89e] Allan Janik. Nyíri on the conservatism of Wittgenstein’s later philosophy. In Style, Politics and the Future of Philosophy [Jan89l], chapter 2, pages 40–58. ISBN 94-010-7508-5, 94-009-2251-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-009-2251-
8_2.

Janik:1989:ECC

[Jan89f] Allan Janik. On edification and cultural conversation: A critique of Rorty. In Style, Politics and the Future of Philosophy [Jan89l], chapter 4, pages 80–92. ISBN 94-010-7508-5, 94-009-2251-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-009-2251-
8_4.

Janik:1989:PC

[Jan89g] Allan Janik. The politics of conciliation. In Style, Politics and the Future of Philosophy [Jan89l], chapter 7, pages 116–132. ISBN 94-010-7508-5, 94-009-2251-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-009-
2251-8_7.

Janik:1989:PSL

[Jan89h] Allan Janik. Psychoanalysis: Science, literature or art? In Style, Politics and the Future of Philosophy [Jan89l], chapter 12, pages 190–196. ISBN 94-010-7508-5, 94-009-2251-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-009-2251-8_12.

Janik:1989:SDN

[Jan89i] Allan Janik. Self-deception, naturalism and certainty: Prolegomena to a critical hermeneutics. In Style, Politics and the Future of Philosophy [Jan89l], chapter 11, pages 172–189. ISBN 94-010-7508-5, 94-009-2251-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-009-2251-8_11.

Janik:1989:SCB

[Jan89j] Allan Janik. Socialization is creative because creativity is social. In Style, Politics and the Future of Philosophy [Jan89l], chapter 9, pages 146–158. ISBN 94-010-7508-5, 94-009-2251-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-009-2251-
                                              8_9.

Janik:1989:SIL

[Jan89k] Allan Janik. Style and idea in the later Heidegger: Rhetoric, politics and philosophy. In Style, Politics and the Future of Philosophy [Jan89l], chapter 1, pages 1–39. ISBN 94-010-7508-5, 94-009-2251-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic).

                                              LCCN ????  URL https://link.

springer.com/chapter/10.1007/
978-94-009-2251-8_1.

Janik:1989:SPF

[Jan89l] Allan Janik. Style, Politics and the Future of Philosophy, volume 114 of Boston Studies in the Philosophy of Science. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 1989. ISBN 94-010-7508-5, 94-009-2251-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). xvi + 280 pp. LCCN ????

URL https://link.springer.com/
book/10.1007/978-94-009-2251-8.

Janik:1989:TKW

[Jan89m] Allan Janik. Tacit knowledge, working life and scientific method. In Style, Politics and the Future of Philosophy [Jan89l], chapter 14, pages 211–224. ISBN 94-010-7508-5, 94-009-2251-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-009-2251-
8_14.

Janik:1989:TWM

[Jan89n] Allan Janik. Towards a Wittgensteinian metaphysics of the political. In Style, Politics and the Future of Philosophy [Jan89l], chapter 5, pages 93–108. ISBN 94-010-7508-5, 94-009-2251-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-009-2251-8_5.

Janik:1989:WMS

[Jan89o] Allan Janik. Wittgenstein, Marx and sociology. In Style, Politics and the Future of Philosophy [Jan89l], chapter 3, pages 59–79. ISBN 94-010-7508-5, 94-009-2251-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

                                              ???? URL https://link.springer.
                                              com/chapter/10.1007/978-94-009-
                                              2251-8_3.

Janich:1997:MC

[Jan97] Peter Janich. Methodical constructivism. In Ginev and Cohen [GC97], chapter 10, pages 173–190. ISBN 0-7923-4444-8 (hardcover), 94-011-5788-X. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 192. URL https://link.

                                              springer.com/chapter/10.1007/
                                              978-94-011-5788-9_10.

Janik:2002:WHH

[Jan02] Allan Janik. Wittgenstein, Hertz, and hermeneutics. In Hermeneutic Philosophy of Science, Van Gogh’s Eyes, and God: Essays in Honor of Patrick A. Heelan, S.J. [Bab02c], chapter 7, pages 79–95. ISBN 90-481-5926-1, 94-017-1767-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 225. URL https:/

                                              /link.springer.com/chapter/10.
                                              1007/978-94-017-1767-0_7.

Janis:2003:PSF

[Jan03a] Allen I. Janis. Physics and science fiction. In Renn et al. [RDS03], chapter 27, pages 545–554. ISBN 1-4020-1284-5, 1-4020-1285-3 (paperback). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 234. URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-010-0111-3_27.

Janssen:2003:TES

[Jan03b] Michel Janssen. The Trouton experiment, E = mc^{2}, and a slice of Minkowski space–time. In Renn et al. [RDS03], chapter 3, pages 27–54. ISBN 1-4020-1284-5, 1-4020-1285-3 (paperback). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 234. URL https://link.

springer.com/chapter/10.1007/
978-94-010-0111-3_3.

Janssen:2007:WDE

[Jan07] Michel Janssen. What did Einstein know and when did he know it? A Besso memo dated August 1913. In The Genesis of General Relativity: Volume 2. Einstein’s Zurich Notebook: Commentary and Essays [JRN^{+}07b], pages 785–837. ISBN 1-4020-3942-5. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 no. 250 v. 2; QC173.6 .G469 2007. URL https:

//link.springer.com/chapter/10.
1007/978-1-4020-4000-9_8.

Jarvie:1974:OA

[Jar74] I. C. Jarvie. On the objectivity of anthropology. In Seeger and Cohen [SC74b], pages 317–324. ISBN 90-277-0390-6, 90-277-0376-0 (paperback), 94-010-2126-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 11. URL https:

//link.springer.com/chapter/10.
1007/978-94-010-2126-5_17.

Jarvie:1976:TRS

[Jar76] I. C. Jarvie. Toulmin and the rationality of science. In Essays in Memory of Imre Lakatos [CFW76], chapter 20, pages 311–333. ISBN 90-277-0655-7, 94-010-1451-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-010-
1451-9_20.

Jarvie:1983:RSP

[Jar83] I. C. Jarvie. Realism and the supposed poverty of sociological theories. In Epistemology, Methodology, and the Social Sciences [CW83a], chapter 5, pages 107–121. ISBN 90-481-8376-6, 94-017-1458-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-017-
1458-7_5.

Jarvie:1986:AI

[Jar86a] I. C. Jarvie. Anthropologists and the irrational. In Thinking about Society: Theory and Practice [Jar86y], chapter 15, pages 233–256. ISBN 94-010-8894-2, 94-009-5424-7 (ebook). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN HM545.

URL https://link.springer.com/
chapter/10.1007/978-94-009-5424-
3_15.

Jarvie:1986:ASA

[Jar86b] I. C. Jarvie. Anthropology as science and the anthropology of science and of anthropology. In Thinking about Society: Theory and Practice [Jar86y], chapter 11, pages 162–182. ISBN 94-010-8894-2, 94-009-5424-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN HM545.

URL https://link.springer.com/
chapter/10.1007/978-94-009-5424-
3_11.

Jarvie:1986:ESA

[Jar86c] I. C. Jarvie. The emergence of social anthropology from philosophy. In Thinking about Society: Theory and Practice [Jar86y], chapter 6, pages 95–106. ISBN 94-010-8894-2, 94-009-5424-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic).

LCCN HM545. URL https://link.
springer.com/chapter/10.1007/
978-94-009-5424-3_6.

Jarvie:1986:EA

[Jar86d] I. C. Jarvie. Epistle to the anthropologists. In Thinking about Society: Theory and Practice [Jar86y], chapter 12, pages 183–196. ISBN 94-010-8894-2, 94-009-5424-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic).

LCCN HM545. URL https://link.
springer.com/chapter/10.1007/
978-94-009-5424-3_12.

Jarvie:1986:ESC

[Jar86e] I. C. Jarvie. Explorations in the social career of movies: Business and religion. In Thinking about Society: Theory and Practice [Jar86y], chapter 24, pages 368–389. ISBN 94-010-8894-2, 94-009-5424-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic).

LCCN HM545. URL https://link.
springer.com/chapter/10.1007/
978-94-009-5424-3_24.

Jarvie:1986:FM

[Jar86f] I. C. Jarvie. Freeman on Mead.

In Thinking about Society: Theory and Practice [Jar86y], chapter 16, pages 257–262. ISBN 94-010-8894-2, 94-009-5424-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic).

LCCN HM545. URL https://link.
springer.com/chapter/10.1007/
978-94-009-5424-3_16.

Jarvie:1986:TU

[Jar86g] I. C. Jarvie. Is technology unnatural?

In Thinking about Society: Theory and Practice [Jar86y], chapter 21, pages 321–327. ISBN 94-010-8894-2, 94-009-5424-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic).

LCCN HM545. URL https://link.
springer.com/chapter/10.1007/
978-94-009-5424-3_21.

Jarvie:1986:LFA

[Jar86h] I. C. Jarvie. Limits to functionalism and alternatives to it in anthropology. In Thinking about Society: Theory and Practice [Jar86y], chapter 8, pages 127–143. ISBN 94-010-8894-2, 94-009-5424-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic).

                                              LCCN HM545. URL https://link.
                                              springer.com/chapter/10.1007/
                                              978-94-009-5424-3_8.

Jarvie:1986:MCP

[Jar86i] I. C. Jarvie. Methodological and conceptual problems in the study of pornography and violence. In Thinking about Society: Theory and Practice [Jar86y], chapter 25, pages 390–475. ISBN 94-010-8894-2, 94-009-5424-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN HM545.

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-009-5424-
                                              3_25.

Jarvie:1986:NSSb

[Jar86j] I. C. Jarvie. Nationalism and the social sciences. In Thinking about Society: Theory and Practice [Jar86y], chapter 23, pages 352–367. ISBN 94-010-8894-2, 94-009-5424-7 (ebook). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN HM545.

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-009-5424-
                                              3_23.

Jarvie:1986:NSSa

[Jar86k] I. C. Jarvie. The notion of a social science. In Thinking about Society: Theory and Practice [Jar86y], chapter 1, pages 3–16. ISBN 94-010-8894-2, 94-009-5424-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic).

LCCN HM545. URL https://link.
springer.com/chapter/10.1007/
978-94-009-5424-3_1.

Jarvie:1986:OCA

[Jar86l]  I. C. Jarvie.  The objectivity of

criticism of the arts. In Thinking about Society: Theory and Practice [Jar86y], chapter 17, pages 265–281. ISBN 94-010-8894-2, 94-009-5424-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN HM545.

URL https://link.springer.com/
chapter/10.1007/978-94-009-5424-
3_17.

Jarvie:1986:LSI

[Jar86m] I. C. Jarvie. On the limits of symbolic interpretation in anthropology. In Thinking about Society: Theory and Practice [Jar86y], chapter 13, pages 197–211. ISBN 94-010-8894-2, 94-009-5424-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic).

LCCN HM545. URL https://link.
springer.com/chapter/10.1007/
978-94-009-5424-3_13.

Jarvie:1986:OA

[Jar86n] I. C. Jarvie. On the objectivity of anthropology. In Thinking about Society: Theory and Practice [Jar86y], chapter 9, pages 144–151. ISBN 94-010-8894-2, 94-009-5424-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic).

LCCN HM545. URL https://link.
springer.com/chapter/10.1007/
978-94-009-5424-3_9.

Jarvie:1986:TFS

[Jar86o] I. C. Jarvie. On theories of fieldwork and the scientific character of social anthropology. In Thinking about Society: Theory and Practice [Jar86y], chapter 7, pages 107–126.

ISBN 94-010-8894-2, 94-009-5424-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN HM545.

URL https://link.springer.com/
chapter/10.1007/978-94-009-5424-
3_7.

Jarvie:1986:PDB

[Jar86p] I. C. Jarvie. Popper on the difference between the natural and the social sciences. In Thinking about Society: Theory and Practice [Jar86y], chapter 5, pages 70–94. ISBN 94-010-8894-2, 94-009-5424-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic).

LCCN HM545. URL https://link.
springer.com/chapter/10.1007/
978-94-009-5424-3_5.

Jarvie:1986:PEI

[Jar86q] I. C. Jarvie. The problem of ethical integrity in participant observation. In Thinking about Society: Theory and Practice [Jar86y], chapter 10, pages 152–161. ISBN 94-010-8894-2, 94-009-5424-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic).

LCCN HM545. URL https://link.
springer.com/chapter/10.1007/
978-94-009-5424-3_10.

Jarvie:1986:PER

[Jar86r] I. C. Jarvie. The problem of the ethnographic real. In Thinking about Society: Theory and Practice [Jar86y], chapter 14, pages 212–232. ISBN 94-010-8894-2, 94-009-5424-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN HM545.

URL https://link.springer.com/
chapter/10.1007/978-94-009-5424-
3_14.

Jarvie:1986:RR

[Jar86s] I. C. Jarvie. Rationality and relativism. In Thinking about Society: Theory and Practice [Jar86y], chapter 4, pages 50– 69. ISBN 94-010-8894-2, 94-009-5424-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN HM545.

URL https://link.springer.com/
chapter/10.1007/978-94-009-5424-
3_4.

Jarvie:1986:RC

[Jar86t] I. C. Jarvie. The rationality of creativity. In Thinking about Society: Theory and Practice [Jar86y], chapter 18, pages 282–301. ISBN 94-010-8894-2, 94-009-5424-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic).

LCCN HM545. URL https://link.
springer.com/chapter/10.1007/
978-94-009-5424-3_18.

Jarvie:1986:RSP

[Jar86u] I. C. Jarvie. Realism and the supposed poverty of sociological theories. In Thinking about Society: Theory and Practice [Jar86y], chapter 3, pages 35– 49. ISBN 94-010-8894-2, 94-009-5424-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN HM545.

URL https://link.springer.com/
chapter/10.1007/978-94-009-5424-
3_3.

Jarvie:1986:SCT

[Jar86v] I. C. Jarvie. The social character of technological problems. In Thinking about Society: Theory and Practice [Jar86y], chapter 20, pages 314–320. ISBN 94-010-8894-2, 94-009-5424-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN HM545.

URL https://link.springer.com/
chapter/10.1007/978-94-009-5424-
3_20.

Jarvie:1986:SPS

[Jar86w] I. C. Jarvie. Social perception and social change. In Thinking about Society:

Theory and Practice [Jar86y], chapter 2, pages 17–34. ISBN 94-010-8894-2, 94-009-5424-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic).

LCCN HM545. URL https://link.
springer.com/chapter/10.1007/
978-94-009-5424-3_2.

Jarvie:1986:TSK

[Jar86x] I. C. Jarvie. Technology and the structure of knowledge. In Thinking about Society: Theory and Practice [Jar86y], chapter 19, pages 302–313. ISBN 94-010-8894-2, 94-009-5424-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN HM545.

URL https://link.springer.com/
chapter/10.1007/978-94-009-5424-
3_19.

Jarvie:1986:TAS

[Jar86y] I. C. Jarvie. Thinking about Society: Theory and Practice, volume 93 of Boston Studies in the Philosophy of Science. D. Reidel, Dordrecht, The Netherlands; Boston, MA, USA; Lancaster, UK; Tokyo, Japan, 1986. ISBN 94-010-8894-2, 94-009-5424-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). 544 (est.) pp.

LCCN HM545. URL https://link.
springer.com/book/10.1007/978-
94-009-5424-3.

Jarvie:1986:UTA

[Jar86z] I. C. Jarvie. Utopian thinking and the architect. In Thinking about Society: Theory and Practice [Jar86y], chapter 22, pages 328–351. ISBN 94-010-8894-2, 94-009-5424-7 (ebook). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN HM545.

URL https://link.springer.com/
chapter/10.1007/978-94-009-5424-
3_22.

Jarvie:1995:PSF

[Jar95] I. C. Jarvie. The place of the sciences and of the fine arts in the intellectual scheme of things. In Critical Rationalism, the Social Sciences and the Humanities: Essays for Joseph Agassi. Volume II [JL95b], chapter 12, pages 203–219. ISBN 0-7923-2961-9, 90-481-4430-2, 94-017-0441-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN B29 .C75 1995.

URL https://link.springer.com/
chapter/10.1007/978-94-017-0441-
0_12.

Jarvie:2009:PCR

[Jar09] Ian Jarvie. Popper’s continuing relevance. In Rethinking Popper [PC09], chapter 17, pages 217–235. ISBN 1-4020-9337-3, 1-4020-9338-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN B1649.P64 R48 2009; Q174 .B67 2009. URL https:

//link.springer.com/chapter/10.
1007/978-1-4020-9338-8_17.

Jarvie:2017:TGS

[Jar17] Ian C. Jarvie. Towards a general sociology of science. In Encouraging Openness: Essays for Joseph Agassi on the Occasion of His 90th Birthday [BAG17], chapter 30, pages 367–388. ISBN 3-319-57668-2 (hardcover), 3-319-57669-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN B833.E53 2017. URL https://link.

springer.com/chapter/10.1007/
978-3-319-57669-5_30.

Jaworski:1990:SCP

[Jaw90] Gary Dean Jaworski. Simmel’s contribution to Parsons’ action theory and its fate. In Georg Simmel and Contemporary Sociology [KPC90], chapter 8, pages 109–130. ISBN 94-010-6691-4 (hardcover), 94-009-0459-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN HM401-1281.

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-009-0459-
                                              0_8.

Jebeile:2017:IEM

[Jeb17] Julie Jebeile. Idealizations in empirical modeling. In Mathematics as a Tool: Tracing New Roles of Mathematics in the Sciences [LC17b], chapter 12, pages 213–232. ISBN 3-319-54468-3 (hardcover), 3-319-54469-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN GE45.M37; GE45.M38; QA8.4. URL https:/

                                              /link.springer.com/chapter/10.
                                              1007/978-3-319-54469-4_12.

Jedlowski:1990:SM

[Jed90] Paolo Jedlowski. Simmel on memory. In Georg Simmel and Contemporary Sociology [KPC90], chapter 9, pages 131–154. ISBN 94-010-6691-4 (hardcover), 94-009-0459-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN HM401-1281.

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-009-0459-
                                              0_9.

Jeffrey:1971:REP

[Jef71]  Richard C. Jeffrey.  Remarks on

explanatory power. In Buck and Cohen [BC71], pages 40–46. ISBN 90-277-0309-4, 94-010-3142-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-010-3142-4_3.

Jeronimo:2015:TEV

[Jer15] Helena Mateus Jerónimo. Technology and ecological values: Confronting normal waste as unavoidable matter in modern society. In New Perspectives on Technology, Values, and Ethics: Theoretical and Practical [Gon15a], chapter 10, pages 183–195. ISBN 3-319-21869-7 (hardcover), 3-319-21870-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-3-319-21870-
0_10.

Jhu:2007:CWA

[Jhu07] Dong-Won Kim Jhu. Challenges in writing about Twentieth Century East Asian physicists. In Positioning the history of science [RG07], chapter 7, pages 39–41. ISBN 1-4020-5419-X (hardcover), 1-4020-5420-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q125 .P77 2007.

URL https://link.springer.com/
chapter/10.1007/1-4020-5420-3_7.

Jiachang:1996:TPS

[Jia96] Luo Jiachang. Time: A philosophical survey. In Chinese Studies in the History and Philosophy of Science and Technology [DC96], chapter 8, pages 77–93. ISBN 90-481-4546-5, 94-015-8717-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 180; U390 .N19 1996.

URL https://link.springer.com/
chapter/10.1007/978-94-015-8717-
4_8.

Jie:1996:HVN

[Jie96] Song Jie. The historical value of the nine chapters on the mathematical art in society and the economy. In Chinese Studies in the History and Philosophy of Science and Technology [DC96], chapter 21, pages 261–266. ISBN 90-481-4546-5, 94-015-8717-5 (e-book).

ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 180; U390 .N19 1996. URL https:

//link.springer.com/chapter/10.
1007/978-94-015-8717-4_21.

Jimenez:1995:SLR

[Jim95] Carlos Alvarez Jimenez. Some logical remarks concerning the continuum problem. In Mexican Studies in the History and Philosophy of Science [RC95], chapter 11, pages 173–186. ISBN 94-010-6535-7 (hardcover), 94-009-0109-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-009-
0109-4_11.

Jingyan:1996:SSP

[Jin96] Lu Jingyan. Studies of the southpointing chariot: Survey of the past 80 years. In Chinese Studies in the History and Philosophy of Science and Technology [DC96], chapter 22, pages 267–278. ISBN 90-481-4546-5, 94-015-8717-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 180; U390 .N19 1996.

URL https://link.springer.com/
chapter/10.1007/978-94-015-8717-
4_22.

Jelved:2007:OSO

[JJ07] Karen Jelved and Andrew D. Jackson. The other side of ørsted: Civil obedience. In Hans Christian Ørsted and The Romantic Legacy In Science: Ideas, Disciplines, Practices [BCK07], chapter 2, pages 13–19. ISBN 1-4020-2979-9 (hardcover), 1-4020-2987-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC515.O37 H36

2007. URL https://link.springer.
com/chapter/10.1007/978-1-4020-
2987-5_2.

Jarvie:1995:CRM

[JL95a] I. C. Jarvie and Nathaniel Laor. Critical Rationalism, Metaphysics and Science: Essays for Joseph Agassi. Volume I, volume 161 of Boston Studies in the Philosophy of Science. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 1995. ISBN 94-010-4211-X, 94-011-0471-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). xxii + 266 pp. LCCN ???? URL

https://link.springer.com/book/
10.1007/978-94-011-0471-5.

Jarvie:1995:CRS

[JL95b] I. C. Jarvie and Nathaniel Laor. Critical Rationalism, the Social Sciences and the Humanities: Essays for Joseph Agassi. Volume II, volume 162 of Boston Studies in the Philosophy and History of Science. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 1995. ISBN 0-7923-2961-9, 90-481-4430-2, 94-017-0441-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). xvii + 278 pp. LCCN B29 .C75 1995. URL

https://link.springer.com/book/
10.1007/978-94-017-0441-0.

Janssen:2006:CRM

[JM06] Michel Janssen and Matthew Mecklenburg. From classical to relativistic mechanics: Electromagnetic models of the electron. In Hendricks et al. [HJLP06], pages 65–134. ISBN 1-4020-5194-8 (hardcover), 1-4020-5195-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175 .B69 v.251 2006; Q175 .I51377 2006.

URL https://link.springer.com/
chapter/10.1007/978-1-4020-5195-
1_3.

Jasinska:1979:MFM

[JN79] Aleksandra Jasińska and Leszek Nowak. The methodological foundations of Marx’s theory of class: A reconstruction. In Polish Essays in the Methodology of the Social Sciences [Wia79b], chapter 5, pages 75–104. ISBN 90-277-0956-4, 94-009-9353-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174

                                              .P655 1979. URL https://link.
                                              springer.com/chapter/10.1007/
                                              978-94-009-9353-2_5.

Janssen:2007:IVZ

[JNR^{+}07a] Michel Janssen, John Norton, Jürgen Renn, Sauer Tilmann, and John Stachel. Introduction to volumes 1 and 2: The Zurich Notebook and the genesis of General Relativity. In The Genesis of General Relativity: Volume 1. Einstein’s Zurich Notebook: Introduction and Source [JNR^{+}07b], pages 7–20. ISBN 1-4020-3999-9 (hardcover), 1-4020-4000-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 no. 250 v. 1; QC173.6 .G469

                                              2007. URL https://link.springer.
                                              com/chapter/10.1007/978-1-4020-
                                              4000-9_1.

Janssen:2007:GGRa

[JNR^{+}07b] Michel Janssen, John Norton, Jürgen Renn, Sauer Tilmann, John Stachel, and Lindy Divarci. The Genesis of General Relativity: Volume 1. Einstein’s Zurich Notebook: Introduction and Source, volume 250(1) of Boston Studies in the Philosophy of Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2007. ISBN 1-4020-3999-9 (hardcover), 1-4020-4000-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). 487 pp.

LCCN Q174 .B67 no. 250 v. 1; QC173.6 .G469 2007.

Johansson:1996:RWP

[Joh96a] Lars-Göran Johansson. Realism and wave–particle duality. In Cohen et al. [CHR96], pages 329–338. ISBN 90-481-4493-0, 94-015-8638-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-015-8638-2_24.

Johansson:1996:VFC

[Joh96b] Lars-Göran Johansson. Van Fraassen’s constructive empiricism — a critique. In Cohen et al. [CHR96], pages 339–342. ISBN 90-481-4493-0, 94-015-8638-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-015-8638-
2_25.

Johnson:2006:VT

[Joh06] Ann Johnson. Virtual tools. In Philosophy of Chemistry: Synthesis of a New Discipline [BSM06b], chapter 14, pages 273–291. ISBN 1-4020-2455-X, 1-4020-3256-0, 1-4020-3261-7. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 242 QD6. URL https:

//link.springer.com/chapter/10.
1007/1-4020-3261-7_14.

Johnson:2011:ENO

[Joh11] Ann Johnson. Everything new is old again: What place should applied science have in the history of science? In Science in the Context of Application [CN11a], chapter 26, pages 455–466. ISBN 90-481-9050-9, 90-481-9051-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q180.55.M4 S385 2011. URL https:

//link.springer.com/chapter/10.
1007/978-90-481-9051-5_26.

Johnson:2017:REC

[Joh17] Ann Johnson. Rational and empirical cultures of prediction. In Mathematics as a Tool: Tracing New Roles of Mathematics in the Sciences [LC17b], chapter 2, pages 23–35. ISBN 3-319-54468-3 (hardcover), 3-319-54469-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN GE45.M37; GE45.M38; QA8.4. URL https:/

/link.springer.com/chapter/10.
1007/978-3-319-54469-4_2.

Joly:2018:WEE

[Jol18] Bernard Joly. Was early Eighteenth-Century chemistry an empirical science? In What Does it Mean to be an Empiricist?: Empiricisms in Eighteenth Century Sciences [BR18b], chapter 11, pages 203–213. ISBN 3-319-69858-3 (hardcover), 3-319-69860-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN B816 .W43 2018.

URL https://link.springer.com/
chapter/10.1007/978-3-319-69860-
1_11.

Jonas:1986:PCP

[Jon86] Hans Jonas. Parallelism and complementarity: The psycho-physical problem in Spinoza and in the succession of Niels Bohr. In Spinoza and the Sciences [GN86], chapter 10, pages 237–247. ISBN 94-010-8511-0, 94-009-4514-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-009-
4514-2_10.

Jones:1991:TGE

[Jon91] Paul R. Jones. The training in Germany of English-speaking chemists in the Nineteenth Century and its profound influence in America and Britain. In Woodward and Cohen [WC91], pages 299–308. ISBN 94-010-5395-2, 94-011-3164-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-011-
3164-3_27.

Jost:2017:OOM

[Jos17] Jürgen Jost. Object oriented models vs. data analysis — is this the right alternative? In Mathematics as a Tool: Tracing New Roles of Mathematics in the Sciences [LC17b], chapter 14, pages 253–286. ISBN 3-319-54468-3 (hardcover), 3-319-54469-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN GE45.M37; GE45.M38; QA8.4. URL https:/

/link.springer.com/chapter/10.
1007/978-3-319-54469-4_14.

Janssen:2007:UKH

[JR07] Michel Janssen and Jürgen Renn. Untying the knot: How Einstein found his way back to field equations discarded in the Zurich Notebook. In The Genesis of General Relativity: Volume 2. Einstein’s Zurich Notebook: Commentary and Essays [JRN^{+}07b], pages 839–925. ISBN 1-4020-3942-5. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 no. 250 v. 2; QC173.6 .G469

2007. URL https://link.springer.
com/chapter/10.1007/978-1-4020-
4000-9_9.

Janssen:2007:CNG

[JRN^{+}07a] Michel Janssen, Jürgen Renn, John Norton, Sauer Tilmann, and John Stachel. A commentary on the notes on gravity in the Zurich Notebook. In The Genesis of General Relativity:

Volume 2. Einstein’s Zurich Notebook: Commentary and Essays [JRN^{+}07b], pages 489–714. ISBN 1-4020-3942-5. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 no. 250 v. 2; QC173.6 .G469 2007. URL https:

//link.springer.com/chapter/10.
1007/978-1-4020-4000-9_6.

Janssen:2007:GGRb

[JRN^{+}07b] Michel Janssen, Jürgen Renn, John Norton, Sauer Tilmann, and John Stachel. The Genesis of General Relativity: Volume 2. Einstein’s Zurich Notebook: Commentary and Essays, volume 250(2) of Boston Studies in the Philosophy of Science. Springer-Ver-lag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2007. ISBN 1-4020-3942-5. ISSN 0068-0346 (print), 2214-7942 (electronic). 489–938 pp. LCCN Q174 .B67 no. 250 v. 2; QC173.6 .G469 2007.

Jaeger:2003:CER

[JS03] Gregg Jaeger and Sahotra Sarkar. Coherence, entanglement, and reductionist explanation in quantum physics. In Renn et al. [RDS03], chapter 26, pages 523–542. ISBN 1-4020-1284-5, 1-4020-1285-3 (paperback). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 234. URL https:/

/link.springer.com/chapter/10.
1007/978-94-010-0111-3_26.

Jurkowitz:2008:IBM

[Jur08] Edward Jurkowitz. An instrumental Bayesianism meets the history of science. In Rethinking Scientific Change and Theory Comparison: Stabilities, Ruptures, Incommensurabilities? [SSHH08], chapter 9, pages 131–141. ISBN 1-4020-6274-5 (hardcover), 1-4020-6279-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175 .R397 2008. URL https://link.

springer.com/chapter/10.1007/
978-1-4020-6279-7_9.

Joas:2014:SCM

[JW14] Christian Joas and Georges Waysand.

Superconductivity — a challenge to modern physics. In History of Artificial Cold, Scientific, Technological and Cultural Issues [Gav14b], chapter 5, pages 83–92. ISBN 94-007-7198-3 (hardcover), 94-007-7199-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC278 H57 2014; TP494 .H577 2014. URL https:/

/link.springer.com/chapter/10.
1007/978-94-007-7199-4_5.

Jiachang:1996:RRR

[JX96] Luo Jiachang and Hu Xinhe. Relational realism on reform of the view of physical reality and its logical manifestation. In Cohen et al. [CHR96], pages 359–379. ISBN 90-481-4493-0, 94-015-8638-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-015-8638-
2_27.

Kaern:1990:IOS

[Kae90a] Michael Kaern. Introduction one:

Simmel as a puzzling figure. In Georg Simmel and Contemporary Sociology [KPC90], chapter 1, pages 1–8. ISBN 94-010-6691-4 (hardcover), 94-009-0459-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN HM401-1281. URL https://link.

springer.com/chapter/10.1007/
978-94-009-0459-0_1.

Kaern:1990:WHC

[Kae90b] Michael Kaern. The world as human construction. In Georg Simmel and Contemporary Sociology [KPC90], chapter 6, pages 75–98. ISBN 94-010-6691-4 (hardcover), 94-009-0459-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN HM401-1281.

URL https://link.springer.com/
chapter/10.1007/978-94-009-0459-
0_6.

Kahlow:1991:KTN

[Kah91] Andreas Kahlow. Knowledge transfer in the Nineteenth Century: Young, Navier, Roebling, and the Brooklyn Bridge. In Woodward and Cohen [WC91], pages 377–386. ISBN 94-010-5395-2, 94-011-3164-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-011-3164-3_35.

Kahle:2018:SS

[Kah18] Reinhard Kahle. Structure and structures. In Piazza and Pulcini [PP18], pages 109–120. ISBN 3-319-93341-8, 3-319-93342-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic).

LCCN QA8.4. URL https://link.
springer.com/chapter/10.1007/
978-3-319-93342-9_7.

Kaiser:1994:NBC

[Kai94] David Kaiser. Niels Bohr’s conceptual legacy in contemporary particle physics. In Faye and Folse [FF94], chapter 11, pages 257–268. ISBN 0-7923-2378-5 (hardcover), 94-015-8106-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 153 1994; QC16 1994. URL https:

//link.springer.com/chapter/10.
1007/978-94-015-8106-6_11.

Kaiser:1996:EVT

[Kai96] Matthias Kaiser. Empirical versus theoretical progress in science. In Cohen et al. [CHR96], pages 171–193. ISBN 90-481-4493-0, 94-015-8638-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-015-8638-2_12.

Kaiser:2007:MEI

[Kai07] David Kaiser. The mutual embrace:

Institutions and epistemology. In Positioning the history of science [RG07], chapter 18, pages 99–103. ISBN 1-4020-5419-X (hardcover), 1-4020-5420-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q125 .P77 2007. URL https://link.

springer.com/chapter/10.1007/1-
4020-5420-3_18.

Kaluszynska:1982:CAQ

[Kał82] Elżbieta Kałuszyńska. Characteristics of additive quantities. In Polish Essays in the Philosophy of the Natural Sciences [Kra82c], chapter 17, pages 197–207. ISBN 94-009-7705-0 (e-book), 90-277-1287-5. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-009-7705-
1_17.

Kalikow:1987:KLE

[Kal87] Theodora J. Kalikow. Konrad Lorenz as evolutionary epistemologist: The problem of intentionality. In Shimony and Nails [SN87], pages 119–144. ISBN 94-010-8168-9, 94-009-3735-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-009-3735-2_8.

Kalfas:1990:CCB

[Kal90] Vassilis Kalfas. Criteria concerning the birth of a new science: The case of Greek astronomy. In Greek Studies in the Philosophy and History of Science [Nic90], chapter 11, pages 171–185. ISBN 94-010-7403-8, 94-009-2015-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-009-2015-
6_11.

Kaloyerou:1996:OIB

[Kal96] P. N. Kaloyerou. An ontological interpretation of boson fields. In Bohmian Mechanics and Quantum Theory: an Appraisal [CFG96], chapter 11, pages 155–167. ISBN 90-481-4698-4, 94-015-8715-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC173.96-174.52. URL https://link.

springer.com/chapter/10.1007/
978-94-015-8715-0_11.

Kamminga:1993:TAC

[Kam93] Harmke Kamminga. Taking antecedent conditions seriously: A lesson in heuristics from biology. In Correspondence, Invariance and Heuristics: Essays in Honour of Heinz Post [FK93], chapter 3, pages 65–82. ISBN 90-481-4229-6, 94-017-1185-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-017-1185-2_3.

Kampis:1999:HL

[Kam99] George Kampis. The hermeneutics of life. In Fehér et al. [FKR99], pages 157–169. ISBN 90-481-5257-7, 94-015-9293-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-015-
9293-2_13.

Kamlah:2013:ERD

[Kam11]  Andreas Kamlah.  Everybody has

the right to do what he wants: Hans Reichenbach’s volitionism and its historical roots. In The Berlin Group and the Philosophy of Logical Empiricism [MP13], chapter 7, pages 151–175. ISBN 94-007-5484-1 (hardcover), 94-007-5485-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175.32.A24 .M384 2011; Q175 .B47

2013. URL https://link.springer.
com/chapter/10.1007/978-94-007-
5485-0_7.

Kangrga:1979:MMP

[Kan79] Milan Kangrga. The meaning of Marx’s philosophy. In Praxis: Yugoslav Essays in the Philosophy and Methodology of the Social Sciences [MP79], chapter 2, pages 45–61. ISBN 90-277-0968-8, 94-009-9355-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-009-
9355-6_2.

Kanitscheider:1984:ASP

[Kan84] B. Kanitscheider. Analytic and synthetic philosophy. In Rationality in Science and Politics [And84d], chapter 9, pages 159–190. ISBN 90-277-1953-5, 94-009-6254-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-009-6254-5_9.

Kaneko:1987:EIJ

[Kan87] Tsutomu Kaneko. Einstein’s impact on Japanese intellectuals. In Glick [Gli87a], chapter 10, pages 351–379. ISBN 90-277-2498-9, 94-009-3875-6. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 103; QC173.5. US$79.00 (US). URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-009-3875-5_10.

Kangnian:1996:NBC

[Kan96] Yan Kangnian. Niels Bohr in China. In Chinese Studies in the History and Philosophy of Science and Technology [DC96], chapter 33, pages 433–437. ISBN 90-481-4546-5, 94-015-8717-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 180; U390 .N19 1996.

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-015-8717-
                                              4_33.

Kanitscheider:2000:QMR

[Kan00]  Bernulf Kanitscheider.  Quantum

mechanics, realism and the ultimate observer. In The Reality of the Unobservable: Observability, Unobservability and Their Impact on the Issue of Scientific Realism [AP00b], chapter 23, pages 311–316. ISBN 90-481-5458-8, 94-015-9391-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

                                              ???? URL https://link.springer.
                                              com/chapter/10.1007/978-94-015-
                                              9391-5_23.

Kapitza:1980:AAF

[Kap80a] P. L. Kapitza. Alexandre Alexandrovich Friedmann. In Experiment, Theory, Practice: Articles and Addresses [Kap80j], chapter 36, pages 325–326. ISBN 90-277-1062-7, 94-009-8977-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-009-8977-
                                              1_36.

Kapitza:1980:AAS

[Kap80b] P. L. Kapitza. Applying the achievements of science and engineering. In Experiment, Theory, Practice: Articles and Addresses [Kap80j], chapter 18, pages 165–172. ISBN 90-277-1062-7, 94-009-8977-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-009-
8977-1_18.

Kapitza:1980:BFO

[Kap80c] P. L. Kapitza. Basic factors in the organization of science and how they are handled in the U.S.S.R. In Experiment, Theory, Practice: Articles and Addresses [Kap80j], chapter 20, pages 183–194. ISBN 90-277-1062-7, 94-009-8977-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-009-
8977-1_20.

Kapitza:1980:CJE

[Kap80d] P. L. Kapitza. The centenary of the journal of experimental and theoretical physics and the role of journals in the development of science. In Experiment, Theory, Practice: Articles and Addresses [Kap80j], chapter 19, pages 173–182. ISBN 90-277-1062-7, 94-009-8977-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-009-
8977-1_19.

Kapitza:1980:CSP

[Kap80e] P. L. Kapitza. Complex scientific problems. In Experiment, Theory, Practice: Articles and Addresses [Kap80j], chapter 15, pages 152–154. ISBN 90-277-1062-7, 94-009-8977-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/

                                              chapter/10.1007/978-94-009-8977-
                                              1_15.

Kapitza:1980:CEW

[Kap80f] P. L. Kapitza. The construction and early work of the Institute for Physical Problems. In Experiment, Theory, Practice: Articles and Addresses [Kap80j], chapter 10, pages 103–114. ISBN 90-277-1062-7, 94-009-8977-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-009-8977-
                                              1_10.

Kapitza:1980:ESW

[Kap80g] P. L. Kapitza. Effectiveness of scientific work. In Experiment, Theory, Practice: Articles and Addresses [Kap80j], chapter 17, pages 161–164. ISBN 90-277-1062-7, 94-009-8977-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-009-8977-1_17.

Kapitza:1980:EP

[Kap80h] P. L. Kapitza. Energy and physics. In Experiment, Theory, Practice: Articles and Addresses [Kap80j], chapter 8, pages 75–85. ISBN 90-277-1062-7, 94-009-8977-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

                                              ???? URL https://link.springer.
                                              com/chapter/10.1007/978-94-009-
                                              8977-1_8.

Kapitza:1980:ETPa

[Kap80i] P. L. Kapitza. Experiment, theory, practice. In Experiment, Theory, Practice: Articles and Addresses [Kap80j], chapter 16, pages 155–160. ISBN 90-277-1062-7, 94-009-8977-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-009-8977-1_16.

Kapitza:1980:ETPb

[Kap80j] P. L. Kapitza. Experiment, Theory, Practice: Articles and Addresses, volume 46 of Boston Studies in the Philosophy of Science. D. Reidel, Dordrecht, The Netherlands; Boston, MA, USA; Lancaster, UK; Tokyo, Japan, 1980. ISBN 90-277-1062-7, 94-009-8977-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). 464 (est.) pp. LCCN ???? URL

https://link.springer.com/book/
10.1007/978-94-009-8977-1.

Kapitza:1980:GP

[Kap80k] P. L. Kapitza. For the good of the people. In Experiment, Theory, Practice: Articles and Addresses [Kap80j], chapter 26, pages 224–225. ISBN 90-277-1062-7, 94-009-8977-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-009-8977-1_26.

Kapitza:1980:FS

[Kap80l] P. L. Kapitza. The future of science. In Experiment, Theory, Practice: Articles and Addresses [Kap80j], chapter 40, pages 345–364. ISBN 90-277-1062-7, 94-009-8977-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-009-
8977-1_40.

Kapitza:1980:GPE

[Kap80m] P. L. Kapitza. Global problems and energy. In Experiment, Theory, Practice: Articles and Addresses [Kap80j], chapter 42, pages 372–386. ISBN 90-277-1062-7, 94-009-8977-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-009-8977-1_42.

Kapitza:1980:GSP

[Kap80n] P. L. Kapitza. Global scientific problems of the immediate future. In Experiment, Theory, Practice: Articles and Addresses [Kap80j], chapter 41, pages 365–371. ISBN 90-277-1062-7, 94-009-8977-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

                                              ???? URL https://link.springer.
                                              com/chapter/10.1007/978-94-009-
                                              8977-1_41.

Kapitza:1980:HPE

[Kap80o] P. L. Kapitza. High power electronics. In Experiment, Theory, Practice: Articles and Addresses [Kap80j], chapter 6, pages 53–59. ISBN 90-277-1062-7, 94-009-8977-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

                                              ???? URL https://link.springer.
                                              com/chapter/10.1007/978-94-009-
                                              8977-1_6.

Kapitza:1980:HRP

[Kap80p] P. L. Kapitza. History of a Rutherford portrait, 1933–1934. In Experiment, Theory, Practice: Articles and Addresses [Kap80j], chapter 29, pages 244–250. ISBN 90-277-1062-7, 94-009-8977-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

                                              ???? URL https://link.springer.
                                              com/chapter/10.1007/978-94-009-
                                              8977-1_29.

Kapitza:1980:HAW

[Kap80q] P. L. Kapitza. How is atomic war to be prevented? In Experiment, Theory, Practice: Articles and Addresses [Kap80j], chapter 38, pages 337–341. ISBN 90-277-1062-7, 94-009-8977-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-009-8977-
1_38.

Kapitza:1980:MER

[Kap80r] P. L. Kapitza. In memory of Ernest Rutherford. In Experiment, Theory, Practice: Articles and Addresses [Kap80j], chapter 27, pages 229–230. ISBN 90-277-1062-7, 94-009-8977-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-009-8977-
1_27.

Kapitza:1980:MIP

[Kap80s] P. L. Kapitza. In memory of Ivan Petrovich Pavlov. In Experiment, Theory, Practice: Articles and Addresses [Kap80j], chapter 35, pages 323–324. ISBN 90-277-1062-7, 94-009-8977-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-009-8977-
1_35.

Kapitza:1980:LDL

[Kap80t] P. L. Kapitza. Lev Davydovitch Landau. In Experiment, Theory, Practice: Articles and Addresses [Kap80j], chapter 37, pages 327–333. ISBN 90-277-1062-7, 94-009-8977-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-009-8977-1_37.

Kapitza:1980:LWS

[Kap80u] P. L. Kapitza. Lomonosov and world science. In Experiment, Theory, Practice: Articles and Addresses [Kap80j], chapter 32, pages 279–299. ISBN 90-277-1062-7, 94-009-8977-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-009-8977-1_32.

Kapitza:1980:NML

[Kap80v] P. L. Kapitza. A new method for the liquefaction of helium. In Experiment, Theory, Practice: Articles and Addresses [Kap80j], chapter 2, pages 8–11. ISBN 90-277-1062-7, 94-009-8977-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-009-8977-
                                              1_2.

Kapitza:1980:LS

[Kap80w] P. L. Kapitza. On leadership in science. In Experiment, Theory, Practice: Articles and Addresses [Kap80j], chapter 14, pages 144–151. ISBN 90-277-1062-7, 94-009-8977-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-009-8977-1_14.

Kapitza:1980:SSR

[Kap80x] P. L. Kapitza. On some stages of research in the field of magnetism. In Experiment, Theory, Practice: Articles and Addresses [Kap80j], chapter 7, pages 60–74. ISBN 90-277-1062-7, 94-009-8977-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

                                              ???? URL https://link.springer.
                                              com/chapter/10.1007/978-94-009-
                                              8977-1_7.

Kapitza:1980:NBL

[Kap80y] P. L. Kapitza. On the nature of ball lightning. In Experiment, Theory, Practice: Articles and Addresses [Kap80j], chapter 5, pages 47–52. ISBN 90-277-1062-7, 94-009-8977-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-009-8977-1_5.

Kapitza:1980:ORI

[Kap80z] P. L. Kapitza. The organization of research at the Institute for Physical Problems. In Experiment, Theory, Practice: Articles and Addresses [Kap80j], chapter 11, pages 115–132. ISBN 90-277-1062-7, 94-009-8977-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-009-8977-
1_11.

Kapitza:1980:O

[Kap80-27] P. L. Kapitza. Oxygen. In Experiment, Theory, Practice: Articles and Addresses [Kap80j], chapter 4, pages 35–46. ISBN 90-277-1062-7, 94-009-8977-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-009-
8977-1_4.

Kapitza:1980:PIS

[Kap80-28] P. L. Kapitza. Philosophy and ideological struggle. In Experiment, Theory, Practice: Articles and Addresses [Kap80j], chapter 39, pages 342–344. ISBN 90-277-1062-7, 94-009-8977-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-009-8977-
1_39.

Kapitza:1980:PESb

[Kap80-29] P. L. Kapitza. Physical experimentation at school. In Experiment, Theory, Practice: Articles and Addresses [Kap80j], chapter 21, page 197. ISBN 90-277-1062-7, 94-009-8977-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-009-8977-1_21.

Kapitza:1980:PPF

[Kap80-30] P. L. Kapitza. The physicist and public figure Paul Langevin. In Experiment, Theory, Practice: Articles and Addresses [Kap80j], chapter 34, pages 315–322. ISBN 90-277-1062-7, 94-009-8977-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

                                              ???? URL https://link.springer.
                                              com/chapter/10.1007/978-94-009-
                                              8977-1_34.

Kapitza:1980:PSa

[Kap80-31] P. L. Kapitza. Planning in science. In Experiment, Theory, Practice: Articles and Addresses [Kap80j], chapter 13, pages 140–143. ISBN 90-277-1062-7, 94-009-8977-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

                                              ???? URL https://link.springer.
                                              com/chapter/10.1007/978-94-009-
                                              8977-1_13.

Kapitza:1980:PCT

[Kap80-32] P. L. Kapitza. Plasma and the controlled thermonuclear reaction. In Experiment, Theory, Practice: Articles and Addresses [Kap80j], chapter 9, pages 86–100. ISBN 90-277-1062-7, 94-009-8977-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

                                              ???? URL https://link.springer.
                                              com/chapter/10.1007/978-94-009-
                                              8977-1_9.

Kapitza:1980:PP

[Kap80-33] P. L. Kapitza. Problems in physics. In Experiment, Theory, Practice: Articles and Addresses [Kap80j], chapter 22, pages 198–203. ISBN 90-277-1062-7, 94-009-8977-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-009-
8977-1_22.

Kapitza:1980:PLH

[Kap80-34] P. L. Kapitza. Problems of liquid helium. In Experiment, Theory, Practice: Articles and Addresses [Kap80j], chapter 3, pages 12–34. ISBN 90-277-1062-7, 94-009-8977-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-009-8977-1_3.

Kapitza:1980:PESa

[Kap80-35] P. L. Kapitza. The production of and experiments in strong magnetic fields. In Experiment, Theory, Practice: Articles and Addresses [Kap80j], chapter 1, pages 3–7. ISBN 90-277-1062-7, 94-009-8977-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-009-
8977-1_1.

Kapitza:1980:PSb

[Kap80-36] P. L. Kapitza. Professor and student. In Experiment, Theory, Practice: Articles and Addresses [Kap80j], chapter 24, pages 215–219. ISBN 90-277-1062-7, 94-009-8977-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-009-
8977-1_24.

Kapitza:1980:RLR

[Kap80-37] P. L. Kapitza. Recollections of Lord Rutherford. In Experiment, Theory, Practice: Articles and Addresses [Kap80j], chapter 30, pages 251–270. ISBN 90-277-1062-7, 94-009-8977-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-009-8977-
                                              1_30.

Kapitza:1980:RAP

[Kap80-38] P. L. Kapitza. Remarks on the anniversary of the Physico–Technical Institute. In Experiment, Theory, Practice: Articles and Addresses [Kap80j], chapter 25, pages 220–223. ISBN 90-277-1062-7, 94-009-8977-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-009-8977-1_25.

Kapitza:1980:ROS

[Kap80-39] P. L. Kapitza. The role of an outstanding scientist in the development of science. In Experiment, Theory, Practice: Articles and Addresses [Kap80j], chapter 31, pages 271–276. ISBN 90-277-1062-7, 94-009-8977-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-009-8977-1_31.

Kapitza:1980:SAB

[Kap80-40] P. L. Kapitza. The scientific activity of Benjamin Franklin. In Experiment, Theory, Practice: Articles and Addresses [Kap80j], chapter 33, pages 300–314. ISBN 90-277-1062-7, 94-009-8977-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-009-
8977-1_33.

Kapitza:1980:SWR

[Kap80-41] P. L. Kapitza. The scientific work of Rutherford. In Experiment, Theory, Practice: Articles and Addresses [Kap80j], chapter 28, pages 231–243. ISBN 90-277-1062-7, 94-009-8977-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-009-8977-
1_28.

Kapitza:1980:SPC

[Kap80-42] P. L. Kapitza. Some principles of the creative upbringing and education of today’s youth. In Experiment, Theory, Practice: Articles and Addresses [Kap80j], chapter 23, pages 204–214. ISBN 90-277-1062-7, 94-009-8977-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-009-8977-
1_23.

Kapitza:1980:UST

[Kap80-43] P. L. Kapitza. The unity of science and technology. In Experiment, Theory, Practice: Articles and Addresses [Kap80j], chapter 12, pages 135–139. ISBN 90-277-1062-7, 94-009-8977-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-009-8977-
1_12.

Kapitza:1986:CSP

[Kap86a] P. L. Kapitza. Complex scientific problems. In Cohen and Wartofsky [CW85], chapter 14, pages 206–208.

ISBN 90-277-1971-3, 94-009-5345-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-009-5345-
1_14.

Kapitza:1986:ETP

[Kap86b] P. L. Kapitza. Experiment, theory, practice. In Cohen and Wartofsky [CW85], chapter 15, pages 209–214. ISBN 90-277-1971-3, 94-009-5345-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-009-5345-
1_15.

Karasmanis:1990:HMP

[Kar90] Vassilis Karasmanis. The hypotheses of mathematics in Plato’s Republic and his contribution to the axiomatization of geometry. In Greek Studies in the Philosophy and History of Science [Nic90], chapter 8, pages 121–135. ISBN 94-010-7403-8, 94-009-2015-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-009-2015-
6_8.

Karachalios:2010:CBE

[Kar10a] Andreas Karachalios. The controversy between Erich Hückel and Linus Pauling over the benzene problem. In Erich Hückel (1896–1980): From Physics to Quantum Chemistry [Kar10b], chapter 3, pages 115–139. ISBN 90-481-3559-1 (hardcover), 90-481-3560-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC174.12 .K37313 2010. URL https://link.

springer.com/chapter/10.1007/
978-90-481-3560-8_3.

Karachalios:2010:EHP

[Kar10b] Andreas Karachalios. Erich Hückel (1896–1980): From Physics to Quantum Chemistry, volume 283 of Boston Studies in the Philosophy of Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2010. ISBN 90-481-3559-1 (hardcover), 90-481-3560-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). xvii + 200 pp. LCCN QC174.12 .K37313 2010. URL

https://link.springer.com/book/
10.1007/978-90-481-3560-8.

Karachalios:2010:EHE

[Kar10c] Andreas Karachalios. Erich Hückel’s education and scientific awakening: The path to quantum chemistry. In Erich Hückel (1896–1980): From Physics to Quantum Chemistry [Kar10b], chapter 1, pages 1–48. ISBN 90-481-3559-1 (hardcover), 90-481-3560-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC174.12 .K37313

2010. URL https://link.springer.
com/chapter/10.1007/978-90-481-
3560-8_1.

Karachalios:2010:EHR

[Kar10d] Andreas Karachalios. Erich Hückel’s research agenda during the 1930s: Underpinning organic chemistry with quantum theory. In Erich Hückel (1896–1980): From Physics to Quantum Chemistry [Kar10b], chapter 2, pages 49–113. ISBN 90-481-3559-1 (hardcover), 90-481-3560-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC174.12 .K37313

2010. URL https://link.springer.
com/chapter/10.1007/978-90-481-
3560-8_2.

Karachalios:2010:HED

[Kar10e] Andreas Karachalios. Hückel’s efforts to disseminate his theory and its reception. In Erich Hückel (1896–1980): From Physics to Quantum Chemistry [Kar10b], chapter 5, pages 151–159. ISBN 90-481-3559-1 (hardcover), 90-481-3560-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC174.12 .K37313 2010. URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-90-481-3560-8_5.

Karachalios:2010:HPC

[Kar10f] Andreas Karachalios. Hückel’s professional career in National Socialist Germany. In Erich Hückel (1896–1980): From Physics to Quantum Chemistry [Kar10b], chapter 6, pages 161–169. ISBN 90-481-3559-1 (hardcover), 90-481-3560-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC174.12 .K37313 2010. URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-90-481-3560-8_6.

Karachalios:2010:LPB

[Kar10g] Andreas Karachalios. Linus Pauling’s breakthrough to the theory of aromatic compounds and Hückel’s reaction. In Erich Hückel (1896–1980): From Physics to Quantum Chemistry [Kar10b], chapter 4, pages 141–149. ISBN 90-481-3559-1 (hardcover), 90-481-3560-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC174.12 .K37313 2010. URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-90-481-3560-8_4.

Karachalios:2010:PY

[Kar10h] Andreas Karachalios. The postwar years. In Erich Hückel (1896–1980): From Physics to Quantum Chemistry [Kar10b], chapter 7, pages 171–173.

ISBN 90-481-3559-1 (hardcover), 90-481-3560-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC174.12 .K37313 2010. URL https:

//link.springer.com/chapter/10.
1007/978-90-481-3560-8_7.

Karachalios:2010:SCR

[Kar10i] Andreas Karachalios. Summary and concluding remarks. In Erich Hückel (1896–1980): From Physics to Quantum Chemistry [Kar10b], chapter 8, pages 175–177. ISBN 90-481-3559-1 (hardcover), 90-481-3560-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC174.12 .K37313

2010. URL https://link.springer.
com/chapter/10.1007/978-90-481-
3560-8_8.

Karstens:2017:LFD

[Kar17] Bart Karstens. “The lonely form dies”:

How epistemic virtues connect Roman Jakobson’s new science of language and his personality. In van Dongen and Paul [vDP17a], chapter 10, pages 149–171. ISBN 3-319-48892-9 (hardcover), 3-319-48893-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 v. 321; Q175.32.V57 E65

2017. URL https://link.springer.
com/chapter/10.1007/978-3-319-
48893-6_10.

Kasher:1976:CMR

[Kas76a] Asa Kasher. Conversational maxims and rationality. In Language in Focus: Foundations, Methods and Systems: Essays in Memory of Yehoshua Bar-Hillel [Kas76b], chapter 12, pages 197–216. ISBN 90-277-0644-1, 90-277-0645-X, 94-010-1876-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN P101-120. URL https://link.

springer.com/chapter/10.1007/
978-94-010-1876-0_12.

Kasher:1976:LFF

[Kas76b] Asa Kasher. Language in Focus: Foundations, Methods and Systems: Essays in Memory of Yehoshua Bar-Hillel, volume 43 of Boston Studies in the Philosophy of Science. D. Reidel, Dordrecht, The Netherlands; Boston, MA, USA; Lancaster, UK; Tokyo, Japan, 1976. ISBN 90-277-0644-1, 90-277-0645-X, 94-010-1876-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). xxviii + 685 pp. LCCN P101-120. URL

                                              https://link.springer.com/book/
                                              10.1007/978-94-010-1876-0.

Kasabova:2013:DB

[Kas11] Anita Kasabova. Dubislav and Bolzano. In The Berlin Group and the Philosophy of Logical Empiricism [MP13], chapter 10, pages 205–228. ISBN 94-007-5484-1 (hardcover), 94-007-5485-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175.32.A24 .M384 2011; Q175 .B47

                                              2013. URL https://link.springer.
                                              com/chapter/10.1007/978-94-007-
                                              5485-0_10.

Katz:1976:DBO

[Kat76] Jerrold J. Katz. The dilemma between orthodoxy and identity. In Language in Focus: Foundations, Methods and Systems: Essays in Memory of Yehoshua Bar-Hillel [Kas76b], chapter 10, pages 165–175. ISBN 90-277-0644-1, 90-277-0645-X, 94-010-1876-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN P101-120. URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-010-1876-0_10.

Katzir:2006:BPS

[Kat06a] Shaul Katzir. The Beginnings of Piezoelectricity: A Study in Mundane Physics, volume 246 of Boston Studies in the Philosophy of Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2006. ISBN 1-4020-4669-3 (hardcover), 1-4020-4670-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). xiii + 273 pp. LCCN QC595.5 .K37 2006.

URL https://link.springer.com/
book/10.1007/978-1-4020-4670-4.

Katzir:2006:EW

[Kat06b] Shaul Katzir. Empirical work in the 1890s. In The Beginnings of Piezoelectricity: A Study in Mundane Physics [Kat06a], chapter 6, pages 187–225. ISBN 1-4020-4669-3 (hardcover), 1-4020-4670-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC595.5 .K37 2006. URL https:/

/link.springer.com/chapter/10.
1007/978-1-4020-4670-4_6.

Katzir:2006:I

[Kat06c] Shaul Katzir. Introduction. In The Beginnings of Piezoelectricity: A Study in Mundane Physics [Kat06a], chapter 1, pages 1–13. ISBN 1-4020-4669-3 (hardcover), 1-4020-4670-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC595.5 .K37

2006. URL https://link.springer.
com/chapter/10.1007/978-1-4020-
4670-4_1.

Katzir:2006:DPE

[Kat06d] Shaul Katzir. the discovery of the piezoelectric effect. In The Beginnings of Piezoelectricity: A Study in Mundane Physics [Kat06a], chapter 2, pages 15–64. ISBN 1-4020-4669-3 (hardcover), 1-4020-4670-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC595.5 .K37

2006. URL https://link.springer.
com/chapter/10.1007/978-1-4020-
4670-4_2.

Katzir:2006:RDT

[Kat06e] Shaul Katzir. the road to the descriptive theory. In The Beginnings of Piezoelectricity: A Study in Mundane Physics [Kat06a], chapter 3, pages 65–97. ISBN 1-4020-4669-3 (hardcover), 1-4020-4670-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC595.5 .K37 2006.

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-1-4020-4670-
                                              4_3.

Katzir:2006:TEV

[Kat06f] Shaul Katzir. Theoretical elaboration of Voigt’s theory. In The Beginnings of Piezoelectricity: A Study in Mundane Physics [Kat06a], chapter 5, pages 149–185. ISBN 1-4020-4669-3 (hardcover), 1-4020-4670-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC595.5 .K37

                                              2006. URL https://link.springer.
                                              com/chapter/10.1007/978-1-4020-
                                              4670-4_5.

Katzir:2006:TMA

[Kat06g] Shaul Katzir. Theories and models about the causes of the piezoelectric phenomena. In The Beginnings of Piezoelectricity: A Study in Mundane Physics [Kat06a], chapter 4, pages 99–147. ISBN 1-4020-4669-3 (hardcover), 1-4020-4670-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC595.5 .K37 2006.

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-1-4020-4670-
                                              4_4.

Katsumori:2011:BDQ

[Kat11a] Makoto Katsumori. Bohr and the development of quantum theory: A brief review. In Niels Bohr’s Complementarity: Its Structure, History, and Intersections with Hermeneutics and Deconstruction [Kat11d], chapter 1, pages 1–10. ISBN 94-007-1747-4, 94-007-1748-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC174.17.C63 K38 2011. URL https:

//link.springer.com/chapter/10.
1007/978-94-007-1748-0_1.

Katsumori:2011:IDD

[Kat11b] Makoto Katsumori. Intersections with Derridean deconstruction. In Niels Bohr’s Complementarity: Its Structure, History, and Intersections with Hermeneutics and Deconstruction [Kat11d], chapter 6, pages 115–151. ISBN 94-007-1747-4, 94-007-1748-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC174.17.C63 K38 2011. URL https:

//link.springer.com/chapter/10.
1007/978-94-007-1748-0_6.

Katsumori:2011:IHP

[Kat11c] Makoto Katsumori. Intersections with hermeneutic philosophy. In Niels Bohr’s Complementarity: Its Structure, History, and Intersections with Hermeneutics and Deconstruction [Kat11d], chapter 5, pages 89–114. ISBN 94-007-1747-4, 94-007-1748-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC174.17.C63 K38

2011. URL https://link.springer.
com/chapter/10.1007/978-94-007-
1748-0_5.

Katsumori:2011:NBC

[Kat11d] Makoto Katsumori. Niels Bohr’s Complementarity: Its Structure, History, and Intersections with Hermeneutics and Deconstruction, volume 286 of Boston Studies in the Philosophy of Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2011. ISBN 94-007-1747-4, 94-007-1748-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). xiv + 176 pp. LCCN QC174.17.C63 K38 2011. URL

https://link.springer.com/book/
10.1007/978-94-007-1748-0.

Katsumori:2011:OBC

[Kat11e] Makoto Katsumori. An overview of Bohr’s complementarity. In Niels Bohr’s Complementarity: Its Structure, History, and Intersections with Hermeneutics and Deconstruction [Kat11d], chapter 2, pages 11–37. ISBN 94-007-1747-4, 94-007-1748-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC174.17.C63 K38

2011. URL https://link.springer.
com/chapter/10.1007/978-94-007-
1748-0_2.

Katsumori:2011:PHA

[Kat11f] Makoto Katsumori. A philosophical– historical analysis of complementarity. In Niels Bohr’s Complementarity: Its Structure, History, and Intersections with Hermeneutics and Deconstruction [Kat11d], chapter 4, pages 61–88. ISBN 94-007-1747-4, 94-007-1748-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC174.17.C63 K38

2011. URL https://link.springer.
com/chapter/10.1007/978-94-007-
1748-0_4.

Katsumori:2011:PIC

[Kat11g] Makoto Katsumori. Prior interpretations of complementarity. In Niels Bohr’s Complementarity: Its Structure, History, and Intersections with Hermeneutics and Deconstruction [Kat11d], chapter 3, pages 39–60. ISBN 94-007-1747-4, 94-007-1748-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC174.17.C63 K38

2011. URL https://link.springer.
com/chapter/10.1007/978-94-007-
1748-0_3.

Katz:2014:STH

[Kat14] Shaul Katz. The scion and its tree:

The Hebrew University of Jerusalem and its German epistemological and organizational origins. In The Institution of Science and the Science of Institutions: The Legacy of Joseph Ben-David [Her14b], chapter 7, pages 103–144. ISBN 94-007-7406-0 (hardcover), 94-007-7407-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 v.302; Q125 .I58 2014.

URL https://link.springer.com/
chapter/10.1007/978-94-007-7407-
0_7.

Kauffman:1971:APE

[Kau71] Stuart A. Kauffman. Articulation of parts explanation in biology and the rational search for them. In Buck and Cohen [BC71], pages 257–272. ISBN 90-277-0309-4, 94-010-3142-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-010-3142-4_18.

Kauffman:1974:EGC

[Kau74] Stuart Kauffman. Elsasser, generalized complementarity, and finite classes: A critique of his anti-reductionism. In Schaffner and Cohen [SC74a], pages 57– 65. ISBN 90-277-0408-2, 90-277-0409-0, 94-010-2140-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175 .B7312. URL https:/

/link.springer.com/chapter/10.
1007/978-94-010-2140-1_4.

Kauffman:1976:APE

[Kau76] Stuart A. Kauffman. Articulation of parts explanation in biology and the rational search for them. In Topics in the Philosophy of Biology [GM76], chapter 11, pages 245–263. ISBN 90-277-0596-8, 94-010-1829-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-010-1829-6_11.

Kauffman:1992:OOE

[Kau92] Stuart A. Kauffman. Origins of order in evolution: Self-organization and selection. In Understanding Origins: Contemporary Views on the Origin of Life, Mind and Society [VD92], chapter 8, pages 153–181. ISBN 90-481-4090-0, 94-015-8054-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-015-8054-0_8.

Kaufmann:1996:BDM

[Kau96] J. Nicolas Kaufmann. The belief–desire model of decision theory needs a third component: Prospective intentions. In Québec Studies in the Philosophy of Science: Part II: Biology, Psychology, Cognitive Science and Economics Essays in Honor of Hugues Leblanc [MC96], chapter 13, pages 215–227. ISBN 94-010-6537-3, 94-009-0113-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-009-0113-
1_13.

Kaufmann:2014:TMS

[Kau14] Felix Kaufmann. Theory and Method in the Social Sciences by Felix Kaufmann: An English translation. In Felix Kaufmann’s Theory and Method in the Social Sciences [CH14], chapter 2, pages 103–353. ISBN 3-319-02844-8 (hardcover), 3-319-02845-6 (e-book).

ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN H61 .K313 2014.

URL https://link.springer.com/
chapter/10.1007/978-3-319-02845-
3_2.

Kaye:1976:ICP

[Kay76] Kenneth Kaye. The IQ controversy and the philosophy of education. In Cohen et al. [CHMV76], pages 181–188. ISBN 90-277-0647-6, 90-277-0648-4 (paperback), 94-010-1449-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-010-1449-6_8.

Kay:2001:BRR

[Kay01] Lily E. Kay. Biopower: Reflections on the rise of molecular biology. In Allen and MacLeod [AM01b], chapter 6, pages 87–102. ISBN 1-4020-0495-8 (hardcover), 94-017-2956-5. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 228; Q175.55.

URL https://link.springer.com/
chapter/10.1007/978-94-017-2956-
7_6.

Kroes:1992:ITD

[KB92a] P. Kroes and M. Bakker. Introduction:

Technological development and science. In Technological Development and Science in the Industrial Age: New Perspectives on the Science–Technology Relationship [KB92b], chapter 1, pages 1–15. ISBN 90-481-4186-9, 94-015-8010-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-015-
8010-6_1.

Kroes:1992:TDS

[KB92b] Peter Kroes and Martijn Bakker. Technological Development and Science in the Industrial Age: New Perspectives on the Science–Technology Relationship, volume 144 of Boston Studies in the Philosophy of Science. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 1992. ISBN 90-481-4186-9, 94-015-8010-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). vii + 280 pp. LCCN ???? URL

https://link.springer.com/book/
10.1007/978-94-015-8010-6.

Kucuradi:1995:CKA

[KC95] Ioanna Kuçuradi and Robert S. Cohen. The Concept of Knowledge: The Ankara Seminar, volume 170 of Boston Studies in the Philosophy of Science. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 1995. ISBN 90-481-4495-7, 94-017-3263-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). lviii + 246 pp. LCCN ????

URL https://link.springer.com/
book/10.1007/978-94-017-3263-5.

Kearney:2002:PPG

[Kea02] Richard Kearney. Poetics of a possible God — faith or philosophy? In Hermeneutic Philosophy of Science, Van Gogh’s Eyes, and God: Essays in Honor of Patrick A. Heelan, S.J. [Bab02c], chapter 30, pages 351–360. ISBN 90-481-5926-1, 94-017-1767-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 225. URL https://link.

springer.com/chapter/10.1007/
978-94-017-1767-0_30.

Kekes:1991:RP

[Kek91] John Kekes. Reason and practice. In Beyond Reason: Essays on the Philosophy of Paul Feyerabend [Mun91a], chapter 6, pages 159–178. ISBN 94-010-5406-1, 94-011-3188-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-011-3188-9_6.

Kekes:2017:SPE

[Kek17] John Kekes. The secular problem of evil. In Encouraging Openness: Essays for Joseph Agassi on the Occasion of His 90th Birthday [BAG17], chapter 29, pages 349–363. ISBN 3-319-57668-2 (hardcover), 3-319-57669-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN B833.E53 2017.

URL https://link.springer.com/
chapter/10.1007/978-3-319-57669-
5_29.

Kelly:1994:DDO

[Kel94] Michael Kelly. Danto, Dutton, and our preunderstanding of tribal art and artifacts. In Artifacts, Representations and Social Practice: Essays for Marx Wartofsky [GC94], chapter 3, pages 39–52. ISBN 94-010-4390-6, 94-011-0902-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN B73 .A77 1994. URL https://link.

springer.com/chapter/10.1007/
978-94-011-0902-4_3.

Keller:2009:SOS

[Kel09] Evelyn Fox Keller. Self-organization, self-assembly, and the origin of life. In Mapping the Future of Biology: Evolving Concepts and Theories [BMP09b], chapter 9, pages 131–140. ISBN 1-4020-9635-6 (hardcover), 1-4020-9636-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QH331 .M295

2009. URL https://link.springer.
com/chapter/10.1007/978-1-4020-
9636-5_9.

Keller:2010:SCD

[Kel10] Agathe Keller. On Sanskrit commentaries dealing with mathematics (Fifth–Twelfth Century). In Looking at it from Asia: the Processes that Shaped the Sources of History of Science [BE10b], chapter 7, pages 211–244. ISBN 90-481-3675-X (hardcover), 90-481-3676-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q127.A65 L66 2010. URL https:/

                                              /link.springer.com/chapter/10.
                                              1007/978-90-481-3676-6_7.

Kent:1996:RCH

[Ken96] Adrian Kent. Remarks on consistent histories and Bohmian mechanics. In Bohmian Mechanics and Quantum Theory: an Appraisal [CFG96], chapter 24, pages 343–352. ISBN 90-481-4698-4, 94-015-8715-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC173.96-174.52.

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-015-8715-
                                              0_24.

Kennedy:2017:TGB

[Ken17] Juliette Kennedy. Turing, Gödel and the “bright abyss”. In Floyd and Bokulich [FB17], chapter 3, pages 63– 91. ISBN 3-319-53278-2, 3-319-53280-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QA29.T8 P45

                                              2017. URL https://link.springer.
                                              com/chapter/10.1007/978-3-319-
                                              53280-6_3.

Kerszberg:1988:HSH

[Ker88] Pierre Kerszberg. The hermeneutical status of the history of science: The views of Hélène Metzger. In Science in Reflection: The Israel Colloquium: Studies in History, Philosophy, and Sociology of Science Volume 3 [UM88], chapter 12, pages 145–150. ISBN 90-277-2713-9, 94-009-2957-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-009-2957-9_12.

Kramer-Friedrich:1986:IMI

[KF86] Sybille Kramer-Friedrich. Information measurement and information technology: A myth of the Twentieth Century. In Philosophy and Technology II: Information Technology and Computers in Theory and Practice [MH86], chapter 2, pages 17–28. ISBN 94-010-8510-2, 94-009-4512-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-009-
4512-8_2.

Kidd:2017:CHH

[Kid17] Ian James Kidd. Confidence, humility, and hubris in Victorian scientific naturalism. In van Dongen and Paul [vDP17a], chapter 2, pages 11–25. ISBN 3-319-48892-9 (hardcover), 3-319-48893-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 v. 321; Q175.32.V57 E65

2017. URL https://link.springer.
com/chapter/10.1007/978-3-319-
48893-6_2.

Kilinc:2003:YMC

[Kil03] Berna Kilinç. Yirmisekiz Mehmed Çelebi’s travelogue and the wonders that make a scientific centre. In Simões et al. [SCD03b], chapter 4, pages 77–100. ISBN 1-4020-1259-4, 94-017-3584-0. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174

.B67 v. 233. URL https://link.
springer.com/chapter/10.1007/
978-94-017-3584-1_4.

Kilinc:2005:OSS

[Kil05] Berna Kilinç. Ottoman science studies — a review. In Turkish Studies in the History and Philosophy of Science [IG05c], chapter 17, pages 251–263. ISBN 1-4020-3332-X (hardcover), 1-4020-3333-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 244.

                                              URL https://link.springer.com/
                                              chapter/10.1007/1-4020-3333-8_17.

Kim:1996:SI

[Kim96] Chin-Tai Kim. Science as ideology. In Cohen et al. [CHR96], pages 195–205. ISBN 90-481-4493-0, 94-015-8638-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-015-8638-
                                              2_13.

Kim:2011:CAH

[Kim11] Nanny Kim. Cultural attitudes and horse technologies: A view on chariots and stirrups from the eastern end of the Eurasian Continent. In Science between Europe and Asia: Historical Studies on the Transmission, Adoption and Adaptation of Knowledge [GR11b], chapter 5, pages 57–73. ISBN 90-481-9968-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q125 .S354 2010; Q175.52.A78 S35

                                              2010. URL https://link.springer.
                                              com/chapter/10.1007/978-90-481-
                                              9968-6_5.

Kinzel:2016:PHC

[Kin16] Katherina Kinzel. Pluralism in historiography: A case study of case studies. In Sauer and Scholl [SS16b], chapter 7, pages 123–149. ISBN 3-319-30227-2, 3-319-30229-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic).

LCCN Q126.8 .P45 2016. URL https:

//link.springer.com/chapter/10.
1007/978-3-319-30229-4_7.

Kirschenmann:1976:TFD

[Kir76] Peter Kirschenmann. Two forms of determinism. In Cohen et al. [CHMV76], pages 393–422. ISBN 90-277-0647-6, 90-277-0648-4 (paperback), 94-010-1449-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-010-
1449-6_20.

Kirschenmann:1982:STI

[Kir82] Peter P. Kirschenmann. Some thoughts on the ideal of exactness in science and philosophy. In Scientific Philosophy Today: Essays in Honor of Mario Bunge [AC82], chapter 4, pages 85–98. ISBN 90-277-1262-X, 90-277-1263-8, 94-009-8462-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 67. URL https:

//link.springer.com/chapter/10.
1007/978-94-009-8462-2_4.

Kirschner:2015:RFV

[Kir15] Marc W. Kirschner. The road to facilitated variation. In Conceptual Change in Biology: Scientific and Philosophical Perspectives on Evolution and Development [Lov15b], chapter 9, pages 199–217. ISBN 94-017-9411-1, 94-017-9412-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QH331 C66 2015; QH491; QH492 .C66

2014. URL https://link.springer.
com/chapter/10.1007/978-94-017-
9412-1_9.

Kiss:1999:MM

[Kis99] Olga Kiss. Meaningful mistakes. In Fehér et al. [FKR99], pages 125–133.

ISBN 90-481-5257-7, 94-015-9293-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-015-9293-
2_11.

Kisiel:2002:WHB

[Kis02] Theodore Kisiel. Was Heißt das — Die Bewandtnis?. (German) [What is it called? — the circumstance]. In Hermeneutic Philosophy of Science, Van Gogh’s Eyes, and God: Essays in Honor of Patrick A. Heelan, S.J. [Bab02c], chapter 10, pages 127–136. ISBN 90-481-5926-1, 94-017-1767-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 225. URL https://link.

springer.com/chapter/10.1007/
978-94-017-1767-0_10.

Kitchener:1974:BFS

[Kit74] Richard F. Kitchener. B. F. Skinner — the butcher, the baker, the behavior-shaper. In Schaffner and Cohen [SC74a], pages 87–98. ISBN 90-277-0408-2, 90-277-0409-0, 94-010-2140-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175 .B7312. URL https://link.

springer.com/chapter/10.1007/
978-94-010-2140-1_6.

Kitromilides:1990:ISM

[Kit90] Paschalis M. Kitromilides. The idea of science in the Modern Greek enlightenment. In Greek Studies in the Philosophy and History of Science [Nic90], chapter 12, pages 187–200. ISBN 94-010-7403-8, 94-009-2015-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-009-2015-
6_12.

Kivy:1994:DMR

[Kiv94] Peter Kivy. In defense of musical representation: Music, representation and the hybrid arts. In Artifacts, Representations and Social Practice: Essays for Marx Wartofsky [GC94], chapter 4, pages 53–67. ISBN 94-010-4390-6, 94-011-0902-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN B73 .A77 1994.

URL https://link.springer.com/
chapter/10.1007/978-94-011-0902-
4_4.

Kivistik:2001:RH

[Kiv01] Jaan Kivistik. On the raising of a hand.

In Estonian Studies in the History and Philosophy of Science [Vih01b], chapter 22, pages 295–304. ISBN 94-010-3863-5, 94-010-0672-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-010-0672-9_22.

Knolle:2014:PCC

[KJ14] Johannes Knolle and Christian Joas.

The physics of cold in the cold war — “on-line computing” between the ICBM program and superconductivity. In History of Artificial Cold, Scientific, Technological and Cultural Issues [Gav14b], chapter 7, pages 119–132. ISBN 94-007-7198-3 (hardcover), 94-007-7199-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC278 H57 2014; TP494 .H577 2014.

URL https://link.springer.com/
chapter/10.1007/978-94-007-7199-
4_7.

Koiv:2001:WT

[KK01] Madis Kõiv and Piret Kuusk. What is time? In Estonian Studies in the History and Philosophy of Science [Vih01b], chapter 17, pages 231–246. ISBN 94-010-3863-5, 94-010-0672-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-010-0672-
9_17.

Kiikeri:2007:BNI

[KK07] Mika Kiikeri and Tomi Kokkonen. Biological notions of innateness and explanation of language acquisition. In Rethinking Explanation [PY07], chapter 13, pages 177–192. ISBN 1-4020-5580-3 (paperback), 1-4020-5581-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN BD237 .R48 2007; Q175.32.E97 R48

2007. URL https://link.springer.
com/chapter/10.1007/978-1-4020-
5581-2_13.

Knuuttila:2016:CCL

[KL16] Tarja Knuuttila and Andrea Loettgers. Contrasting cases: The Lotka–Volterra model times three. In Sauer and Scholl [SS16b], chapter 8, pages 151–178. ISBN 3-319-30227-2, 3-319-30229-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q126.8

.P45 2016.  URL https://link.
springer.com/chapter/10.1007/
978-3-319-30229-4_8.

Knuuttila:2017:MSB

[KL17] Tarja Knuuttila and Andrea Loettgers. Mathematization in synthetic biology: Analogies, templates, and fictions. In Mathematics as a Tool: Tracing New Roles of Mathematics in the Sciences [LC17b], chapter 3, pages 37– 56. ISBN 3-319-54468-3 (hardcover), 3-319-54469-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN GE45.M37; GE45.M38; QA8.4.

URL https://link.springer.com/
chapter/10.1007/978-3-319-54469-
4_3.

Kleiner:1971:OTC

[Kle71] Scott A. Kleiner. Ontological and terminological commitment and the methodological commensurability of theories. In Buck and Cohen [BC71], pages 507–518. ISBN 90-277-0309-4, 94-010-3142-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-010-
3142-4_33.

Klein:1974:BMM

[Kle74] Martin J. Klein. Boltzmann, monocycles and mechanical explanation. In Seeger and Cohen [SC74b], pages 155– 175. ISBN 90-277-0390-6, 90-277-0376-0 (paperback), 94-010-2126-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 11.

URL https://link.springer.com/
chapter/10.1007/978-94-010-2126-
5_8.

Kleinhempel:1991:NCB

[Kle91] Friedrich Kleinhempel. ‘Nobody can become a real engineer who has not already become a whole person’. In Woodward and Cohen [WC91], pages 423–429. ISBN 94-010-5395-2, 94-011-3164-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-011-
3164-3_39.

Klein:2001:CPP

[Kle01a] Ursula Klein. The creative power of paper tools in early Nineteenth-Century chemistry. In Tools and modes of representation in the laboratory sciences [Kle01b], chapter 2, pages 13–34. ISBN 1-4020-0100-2, 94-015-9737-5. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 222;

QD7. URL https://link.springer.
com/chapter/10.1007/978-94-015-
9737-1_2.

Klein:2001:TMR

[Kle01b] Ursula Klein, editor. Tools and modes of representation in the laboratory sciences, volume 222 of Boston Studies in the Philosophy and History of Science. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 2001. ISBN 1-4020-0100-2, 94-015-9737-5. ISSN 0068-0346 (print), 2214-7942 (electronic). xv + 259 pp. LCCN Q174 .B67 vol.

222; QD7. URL http://www.loc.
gov/catdir/enhancements/fy0814/
2001050224-d.html;   http://www.
loc.gov/catdir/enhancements/
fy0814/2001050224-t.html;
https://link.springer.com/book/
10.1007/978-94-015-9737-1.

Klein:2003:SE

[Kle03] Ursula Klein. Styles of experimentation. In Descartes’s Mathematical Thought [Sas03d], chapter 10, pages 159–185. ISBN 90-481-6487-7, 94-017-1225-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/0-306-48123-5_10.

See comments [Bal03, Gil03, Irz03].

Klein:2009:TOM

[Kle09] Ursula Klein. In the thick of organic matter. In Going Amiss In Experimental Research [HSS09a], chapter 14, pages 253–272. ISBN 1-4020-8892-2 (hardcover), 1-4020-8893-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q172.5.E77 G647

2009. URL https://link.springer.
com/chapter/10.1007/978-1-4020-
8893-3_14.

Klein:2012:NAP

[Kle12] Rebekka A. Klein. The neurobiology of altruistic punishment: A moral assessment of its social utility. In Philosophy of Behavioral Biology [PR12b], chapter 14, pages 297–313. ISBN 94-007-1950-7, 94-007-1951-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QH331 .P4679 2012; Q175 .B657 v.282. URL https:

//link.springer.com/chapter/10.
1007/978-94-007-1951-4_14.

Klingenberg:1996:GMG

[Kli96] Wilhelm Klingenberg. Grassmannian manifolds in geometry. In Schubring [Sch96f], pages 281–283. ISBN 90-481-4758-1, 94-015-8753-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QA8.9-10.3.

URL https://link.springer.com/
chapter/10.1007/978-94-015-8753-
2_23.

Kloskowska:1979:CCA

[Klo79] Antonina Kloskowska. The conception of culture according to Karl Marx. In Polish Essays in the Methodology of the Social Sciences [Wia79b], chapter 3, pages 33–47. ISBN 90-277-0956-4, 94-009-9353-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .P655 1979. URL https:

//link.springer.com/chapter/10.
1007/978-94-009-9353-2_3.

Kremer-Marietti:1999:NCM

[KM99] Angèle Kremer-Marietti. Nietzsche’s critique of modern reason. In Nietzsche, Theories of Knowledge, and Critical Theory: Nietzsche and the Sciences I [Bab99b], chapter 7, pages 87–102. ISBN 90-481-5233-X, 94-017-2430-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-017-2430-
2_7.

Kroon:2001:RRF

[KN01] Fred Kroon and Robert Nola. Ramsification, reference fixing and incommensurability. In Incommensurability and Related Matters [HHS01], chapter 3, pages 91–121. ISBN 90-481-5709-9, 94-015-9680-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-015-
9680-0_3.

Knight:2007:SM

[Kni07] D. M. Knight. The spiritual in the material. In Hans Christian Ørsted and The Romantic Legacy In Science: Ideas, Disciplines, Practices [BCK07], chapter 19, pages 417–432. ISBN 1-4020-2979-9 (hardcover), 1-4020-2987-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC515.O37 H36 2007. URL https:

//link.springer.com/chapter/10.
1007/978-1-4020-2987-5_19.

Knoespel:1989:MTR

[Kno89] Kenneth J. Knoespel. The mythological transformations of Renaissance science: Physical allegory and the crisis of alchemical narrative. In Literature and Science as Modes of Expression [Amr89], chapter 5, pages 99–112. ISBN 94-010-7531-X, 94-009-2297-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-009-2297-6_5.

Knobloch:1994:ILM

[Kno94] Eberhard Knobloch. The infinite in Leibniz’s mathematics — the historiographical method of comprehension in context. In Trends in the Historiography of Science [GCN94], chapter 20, pages 265–278. ISBN 90-481-4264-4, 94-017-3596-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-017-
3596-4_20.

Knorr:2004:TFA

[Kno04a] Wilbur Knorr. Techniques of fractions in Ancient Egypt and Greece. In Classics in the History of Greek Mathematics [Chr04a], chapter 19, pages 337–365. ISBN 1-4020-0081-2 (hardcover), 1-4020-2640-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QA22 .C596 2004.

URL https://link.springer.com/
chapter/10.1007/978-1-4020-2640-
9_19.

Knorr:2004:CEP

[Kno04b] Wilbur R. Knorr. Construction as existence proof in ancient geometry. In Classics in the History of Greek Mathematics [Chr04a], chapter 6, pages 115–137. ISBN 1-4020-0081-2 (hardcover), 1-4020-2640-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QA22 .C596 2004.

URL https://link.springer.com/
chapter/10.1007/978-1-4020-2640-
9_6.

Knorr:2004:IMM

[Kno04c] Wilbur R. Knorr. The impact of modern mathematics on ancient mathematics. In Classics in the History of Greek Mathematics [Chr04a], chapter 13, pages 243–253. ISBN 1-4020-0081-2 (hardcover), 1-4020-2640-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QA22 .C596 2004.

URL https://link.springer.com/
chapter/10.1007/978-1-4020-2640-
9_13.

Knorr:2004:EHA

[Kno04d] Wilbur Richard Knorr. On the early history of axiomatics: The interaction of mathematics and philosophy in Greek antiquity. In Classics in the History of Greek Mathematics [Chr04a], chapter 4, pages 81–109. ISBN 1-4020-0081-2 (hardcover), 1-4020-2640-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QA22 .C596

2004. URL https://link.springer.
com/chapter/10.1007/978-1-4020-
2640-9_4.

Knobloch:2005:LUM

[Kno05] Eberhard Knobloch. Leibniz and the use of manuscripts: Text as process. In History of Science, History of Text [Che05a], chapter 2, pages 51–79. ISBN 1-281-33589-4, 1-4020-2320-0 (hardcover), 1-4020-2321-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 238

Q125. URL https://link.springer.
com/chapter/10.1007/1-4020-2321-
9_2.

Knowles:2016:SAP

[Kno16] Robert Knowles. Semantic assumptions in the philosophy of mathematics. In Objectivity, Realism, and Proof: FilMat Studies in the Philosophy of Mathematics [BS16], chapter 4, pages 43–65. ISBN 3-319-31642-7, 3-319-31644-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QA9 .O25 2016. URL https://link.

springer.com/chapter/10.1007/
978-3-319-31644-4_4.

Knudsen:1995:EEE

[Knu95] Ole Knudsen. Electromagnetic energy and the early history of the energy principle. In No Truth Except in the Details: Essays in Honor of Martin J. Klein [KS95], chapter 3, pages 55–78. ISBN 94-010-4097-4, 94-011-0217-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-011-0217-
9_3.

Knudsen:2007:OWC

[Knu07] Ole Knudsen. Ørsted’s work on the compressibility of liquids and gases, and his dynamic theory of matter. In Hans Christian Ørsted and The Romantic Legacy In Science: Ideas, Disciplines, Practices [BCK07], chapter 17, pages 387–398. ISBN 1-4020-2979-9 (hardcover), 1-4020-2987-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC515.O37 H36 2007.

URL https://link.springer.com/
chapter/10.1007/978-1-4020-2987-
5_17.

Kockelmans:1979:RLM

[Koc79] Joseph J. Kockelmans. Reflections on Lakatos’ methodology of scientific research programs. In The Structure and Development of Science [RA79], chapter 7, pages 187–203. ISBN 90-277-0995-5, 94-009-9459-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .S778 1979.

URL https://link.springer.com/
chapter/10.1007/978-94-009-9459-
1_7.

Koch:1989:H

[Koc89] Robert Koch. Hypotyposes. In Literature and Science as Modes of Expression [Amr89], chapter 4, pages 81–98. ISBN 94-010-7531-X, 94-009-2297-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-009-2297-6_4.

Koch:1991:BPC

[Koc91] Martin Koch. From Boltzmann to Planck: On continuity in scientific revolutions. In Woodward and Cohen [WC91], pages 141–150. ISBN 94-010-5395-2, 94-011-3164-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-011-3164-3_13.

Kockelmans:1997:HVE

[Koc97] Joseph J. Kockelmans. Hermeneutic vs. empiricist philosophy of science. In Ginev and Cohen [GC97], chapter 11, pages 191–215. ISBN 0-7923-4444-8 (hardcover), 94-011-5788-X. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 192.

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-011-5788-
                                              9_11.

Kockelmans:2002:INH

[Koc02] Joseph J. Kockelmans. On the interpretive nature of Hertz’s mechanics. In Hermeneutic Philosophy of Science, Van Gogh’s Eyes, and God: Essays in Honor of Patrick A. Heelan, S.J. [Bab02c], chapter 8, pages 97–115. ISBN 90-481-5926-1, 94-017-1767-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 225. URL https://link.

                                              springer.com/chapter/10.1007/
                                              978-94-017-1767-0_8.

Koc:2005:IGQ

[Koç05] Yalçin Koç. Implications of the geometry of quantum mechanical perfect correlation functions concerning “Bell’s Theorem without Inequalities”. In Turkish Studies in the History and Philosophy of Science [IG05c], chapter 5, pages 59–67. ISBN 1-4020-3332-X (hardcover), 1-4020-3333-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 244.

URL https://link.springer.com/
chapter/10.1007/1-4020-3333-8_5.

Kochan:2009:PC

[Koc09] Jeff Kochan. Popper’s communitarianism. In Rethinking Popper [PC09], chapter 22, pages 287–303. ISBN 1-4020-9337-3, 1-4020-9338-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN B1649.P64 R48 2009; Q174 .B67 2009. URL https:

//link.springer.com/chapter/10.
1007/978-1-4020-9338-8_22.

Koertge:1971:ITC

[Koe71] Noretta Koertge. Inter-theoretic criticism and the growth of science. In Buck and Cohen [BC71], pages 160–173. ISBN 90-277-0309-4, 94-010-3142-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-010-3142-
4_11.

Koertge:1974:PPS

[Koe74] N. Koertge. On Popper’s philosophy of social science. In Schaffner and Cohen [SC74a], pages 195–207. ISBN 90-277-0408-2, 90-277-0409-0, 94-010-2140-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175 .B7312.

URL https://link.springer.com/
chapter/10.1007/978-94-010-2140-
1_13.

Koertge:1976:RR

[Koe76] Noretta Koertge. Rational reconstructions. In Essays in Memory of Imre Lakatos [CFW76], chapter 22, pages 359–369. ISBN 90-277-0655-7, 94-010-1451-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

                                              ???? URL https://link.springer.
                                              com/chapter/10.1007/978-94-010-
                                              1451-9_22.

Koertge:1978:TNT

[Koe78] Noretta Koertge. Towards a new theory of scientific inquiry. In Progress and Rationality in Science [RA78b], chapter 11, pages 253–278. ISBN 90-277-0922-X, 94-009-9866-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-009-9866-7_11.

Koertge:1980:AMD

[Koe80] Noretta Koertge. Analysis as a method of discovery during the scientific revolution. In Scientific Discovery, Logic, and Rationality [Nic80f], chapter 4, pages 139–157. ISBN 90-277-1070-8, 94-009-8986-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

                                              ???? URL https://link.springer.
                                              com/chapter/10.1007/978-94-009-
                                              8986-3_4.

Koertge:1984:EPS

[Koe84] Noretta Koertge. Ethical problems in science communication. In Rationality in Science and Politics [And84d], chapter 10, pages 191–203. ISBN 90-277-1953-5, 94-009-6254-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-009-6254-5_10.

Koertge:1986:AMD

[Koe86] Noretta Koertge. Analysis as a method of discovery during the scientific revolution. In Cohen and Wartofsky [CW85], chapter 17, pages 238–256. ISBN 90-277-1971-3, 94-009-5345-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-009-5345-1_17.

Koertge:1989:SSS

[Koe89] Noretta Koertge. Stereotypes, statistics, and schemata. In Freedom and Rationality: Essays in Honor of John Watkins From his Colleagues and Friends [DJ89], chapter 7, pages 111–124. ISBN 94-010-7571-9, 94-009-2380-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-009-
2380-5_7.

Koertge:1991:ISF

[Koe91] Noretta Koertge. Ideology, science and a free society. In Beyond Reason: Essays on the Philosophy of Paul Feyerabend [Mun91a], chapter 10, pages 225–241. ISBN 94-010-5406-1, 94-011-3188-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-011-
3188-9_10.

Koertge:1993:IHR

[Koe93] Noretta Koertge. Ideology, heuristics and rationality in the context of discovery. In Correspondence, Invariance and Heuristics: Essays in Honour of Heinz Post [FK93], chapter 6, pages 125–136. ISBN 90-481-4229-6, 94-017-1185-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-017-1185-
                                              2_6.

Koertge:1995:TPS

[Koe95] Noretta Koertge. Towards a Popperian sociology of science: The problem of credit. In Critical Rationalism, Metaphysics and Science: Essays for Joseph Agassi. Volume I [JL95a], chapter 12, pages 193–210. ISBN 94-010-4211-X, 94-011-0471-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-011-0471-5_12.

Koertge:2009:MUP

[Koe09] Noretta Koertge. The moral underpinnings of Popper’s philosophy. In Rethinking Popper [PC09], chapter 24, pages 323–338. ISBN 1-4020-9337-3, 1-4020-9338-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN B1649.P64 R48 2009; Q174 .B67 2009.

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-1-4020-9338-
                                              8_24.

Kohn:1986:DPD

[Koh86] David Kohn. Darwin’s principle of divergence as internal dialogue. In Ullmann-Margalit [UM86a], chapter 11, pages 117–132. ISBN 90-277-2159-9 (paperback), 90-277-2158-0 (hardcover), 94-009-5496-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 94 Q175.

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-009-5496-
                                              0_11.

Kohnke:1990:FCS

[Kóh90] Klaus Christian Kóhnke. Four concepts of social science at Berlin University: Dilthey, Lazarus, Schmoller and Simmel. In Georg Simmel and Contemporary Sociology [KPC90], chapter 7, pages 99–107. ISBN 94-010-6691-4 (hardcover), 94-009-0459-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN HM401-1281.

URL https://link.springer.com/
chapter/10.1007/978-94-009-0459-
0_7.

Kohler:1991:MVC

[Köh91a] Eckehart Köhler. Metaphysics in the Vienna Circle. In Uebel [Ueb91b], chapter 10, pages 131–142. ISBN 0-7923-1276-7 (hardcover), 94-011-3182-1. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 133. URL https://link.

springer.com/chapter/10.1007/
978-94-011-3182-7_10.

Kohler:1991:NWL

[Köh91b] Eckehart Köhler. On Neurath’s writings on logic, ethics and physics. In Uebel [Ueb91b], chapter 8, pages 109– 115. ISBN 0-7923-1276-7 (hardcover), 94-011-3182-1. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 133. URL https://link.

springer.com/chapter/10.1007/
978-94-011-3182-7_8.

Kohak:1994:GR

[Koh94] Erazim Kohák. The good and the rational. In Artifacts, Representations and Social Practice: Essays for Marx Wartofsky [GC94], chapter 24, pages 397–417. ISBN 94-010-4390-6, 94-011-0902-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN B73 .A77 1994. URL https://link.

springer.com/chapter/10.1007/
978-94-011-0902-4_24.

Kohak:1995:KGE

[Koh95] Erazim Kohák. Knowing good and evil … (Genesis 3: 5B). In Science, Mind and Art: Essays on science and the humanistic understanding in art, epistemology, religion and ethics In honor of Robert S. Cohen [GSW95b], chapter 14, pages 243–254. ISBN 0-7923-2990-2 (hardcover), 94-011-0469-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 165 BD418.3. URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-011-0469-2_14.

Kohn:1997:ACD

[Koh97] David Kohn. The aesthetic construction of Darwin’s theory. In The Elusive Synthesis: Aesthetics and Science [Tau97a], chapter 2, pages 13–48. ISBN 0-7923-4763-3, 0-7923-3904-5, 94-009-1786-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

                                              ???? URL https://link.springer.
                                              com/chapter/10.1007/978-94-009-
                                              1786-6_2.

Kohak:1998:COT

[Koh98a] Erazim Kohák. Creation’s orphans: Toward a metaphysics of artifacts. In Philosophies of Nature: The Human Dimension: In Celebration of Erazim Kohák [CT98], chapter 23, pages 303–314. ISBN 90-481-4859-6, 94-017-2614-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic).

                                              LCCN BD581. URL https://link.
                                              springer.com/chapter/10.1007/
                                              978-94-017-2614-6_23.

Kohak:1998:EDE

[Koh98b] Erazim Kohák. The ecological dilemma: Ethical categories in a biocentric world. In Philosophies of Nature: The Human Dimension: In Celebration of Erazim Kohák [CT98], chapter 22, pages 293–301. ISBN 90-481-4859-6, 94-017-2614-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic).

LCCN BD581. URL https://link.
springer.com/chapter/10.1007/
978-94-017-2614-6_22.

Kohak:1998:HRN

[Koh98c] Erazim Kohák. Human rights and nature’s rightness. In Philosophies of Nature: The Human Dimension: In Celebration of Erazim Kohák [CT98], chapter 16, pages 221–230. ISBN 90-481-4859-6, 94-017-2614-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN BD581.

URL https://link.springer.com/
chapter/10.1007/978-94-017-2614-
6_16.

Kohak:1998:NPE

[Koh98d] Erazim Kohák. Nature as presence and experience. In Philosophies of Nature: The Human Dimension: In Celebration of Erazim Kohák [CT98], chapter 20, pages 273–284. ISBN 90-481-4859-6, 94-017-2614-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic).

LCCN BD581. URL https://link.
springer.com/chapter/10.1007/
978-94-017-2614-6_20.

Kohak:1998:PED

[Koh98e] Erazim Kohák. Phenomenology and ecology: Dependence and codependency. In Philosophies of Nature: The Human Dimension: In Celebration of Erazim Kohák [CT98], chapter 15, pages 209–219. ISBN 90-481-4859-6, 94-017-2614-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic).

LCCN BD581. URL https://link.
springer.com/chapter/10.1007/
978-94-017-2614-6_15.

Kohak:1998:SAA

[Koh98f] Erazim Kohák. Should auld acquaintance be forgot… ? In Philosophies of Nature: The Human Dimension: In Celebration of Erazim Kohák [CT98], chapter 18, pages 249–256. ISBN 90-481-4859-6, 94-017-2614-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN BD581.

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-017-2614-
                                              6_18.

Kohak:1998:TEE

[Koh98g] Erazim Kohák. Transcendental experience, everyday philosophy. In Philosophies of Nature: The Human Dimension: In Celebration of Erazim Kohák [CT98], chapter 17, pages 231–247. ISBN 90-481-4859-6, 94-017-2614-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic).

                                              LCCN BD581. URL https://link.
                                              springer.com/chapter/10.1007/
                                              978-94-017-2614-6_17.

Kohak:1998:TGR

[Koh98h] Erazim Kohák. The true and the good: Reflections on the primacy of practical reason. In Philosophies of Nature: The Human Dimension: In Celebration of Erazim Kohák [CT98], chapter 21, pages 285–291. ISBN 90-481-4859-6, 94-017-2614-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic).

                                              LCCN BD581. URL https://link.
                                              springer.com/chapter/10.1007/
                                              978-94-017-2614-6_21.

Kohak:1998:VEE

[Koh98i] Erazim Kohák. Varieties of ecological experience. In Philosophies of Nature: The Human Dimension: In Celebration of Erazim Kohák [CT98], chapter 19, pages 257–271. ISBN 90-481-4859-6, 94-017-2614-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic).

LCCN BD581. URL https://link.
springer.com/chapter/10.1007/
978-94-017-2614-6_19.

Kohak:1998:WNW

[Koh98j]  Erazim Kohák.  “Whose nature?

Which morality? A response. In Philosophies of Nature: The Human Dimension: In Celebration of Erazim Kohák [CT98], chapter 3, pages 35–37. ISBN 90-481-4859-6, 94-017-2614-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN BD581.

URL https://link.springer.com/
chapter/10.1007/978-94-017-2614-
6_3.

Kohn:1998:YYN

[Koh98k] Livia Kohn. Yin and yang: The natural dimension of evil. In Philosophies of Nature: The Human Dimension: In Celebration of Erazim Kohák [CT98], chapter 7, pages 91–105. ISBN 90-481-4859-6, 94-017-2614-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN BD581.

URL https://link.springer.com/
chapter/10.1007/978-94-017-2614-
6_7.

Kolman:1974:CSP

[Kol74] A. Kolman. The concept of ‘simplicity’ in the physico–mathematical sciences. In For Dirk Struik: Scientific, Historical and Political Essays in Honor of Dirk J. Struik [CSW74], chapter 30, pages 365–371. ISBN 90-277-0379-5, 94-010-2115-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-010-
2115-9_30.

Komlosi:1999:VFW

[Kom99] László I. Komlósi. Value-free worlds and value-imposing minds. Representational practices and the constructivist argument. In Fehér et al. [FKR99], pages 325–343. ISBN 90-481-5257-7, 94-015-9293-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

                                              ???? URL https://link.springer.
                                              com/chapter/10.1007/978-94-015-
                                              9293-2_27.

Konvitz:1988:NBA

[Kon88] Milton R. Konvitz. The nature of Bronson Alcott. In Philosophy, History and Social Action: Essays in Honor of Lewis Feuer with an autobiographic essay by Lewis Feuer [HOO88c], chapter 12, pages 267–274. ISBN 94-010-7793-2, 94-009-2873-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-009-2873-2_12.

Kopperschmidt:1999:NRP

[Kop99]  Josef Kopperschmidt.  Nietzsche’s

rhetorical philosophy as critique of impure reason. In Nietzsche, Theories of Knowledge, and Critical Theory: Nietzsche and the Sciences I [Bab99b], chapter 14, pages 199–207. ISBN 90-481-5233-X, 94-017-2430-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-017-2430-2_14.

Kordig:1971:OSC

[Kor71] Carl R. Kordig. Objectivity, scientific change, and self-reference. In Buck and Cohen [BC71], pages 519–523. ISBN 90-277-0309-4, 94-010-3142-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-010-3142-4_34.

Korner:1986:EAM

[Kör86] Stephan Körner. On the empirical application of mathematics and some of its philosophical aspects. In Ullmann-Margalit [UM86a], chapter 1, pages 1–11. ISBN 90-277-2159-9 (paperback), 90-277-2158-0 (hardcover), 94-009-5496-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 94 Q175. URL https:

//link.springer.com/chapter/10.
1007/978-94-009-5496-0_1.

Kosik:1976:DCS

[Kos76a] Karel Kosík. Dialectics of the Concrete:

A Study on Problems of Man and World, volume 52 of Boston Studies in the Philosophy of Science. D. Reidel, Dordrecht, The Netherlands; Boston, MA, USA; Lancaster, UK; Tokyo, Japan, 1976. ISBN 94-011-9766-0, 94-010-1520-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). viii + 158 pp. LCCN ???? URL

https://link.springer.com/book/
10.1007/978-94-010-1520-2.

Kosik:1976:DCT

[Kos76b] Karel Kosík. Dialectics of the concrete totality. In Dialectics of the Concrete: A Study on Problems of Man and World [Kos76a], chapter 1, pages 1–35. ISBN 94-011-9766-0, 94-010-1520-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-010-1520-
2_1.

Kosik:1976:EP

[Kos76c] Karel Kosík. Economics and philosophy. In Dialectics of the Concrete: A Study on Problems of Man and World [Kos76a], chapter 2, pages 36–92. ISBN 94-011-9766-0, 94-010-1520-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-010-1520-2_2.

Kosik:1976:PE

[Kos76d] Karel Kosík. Philosophy and economy. In Dialectics of the Concrete: A Study on Problems of Man and World [Kos76a], chapter 3, pages 93–132. ISBN 94-011-9766-0, 94-010-1520-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-010-1520-
                                              2_3.

Kosik:1976:PT

[Kos76e] Karel Kosík. Praxis and totality. In Dialectics of the Concrete: A Study on Problems of Man and World [Kos76a], chapter 4, pages 133–155. ISBN 94-011-9766-0, 94-010-1520-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-010-1520-2_4.

Kosok:1984:DHD

[Kos84] M. Kosok. The dynamics of Hegelian dialectics, and non-linearity in the sciences. In Hegel and the Sciences [CW84a], chapter 19, pages 311–347. ISBN 94-009-6235-5, 94-009-6233-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-009-6233-
                                              0_19.

Kostoula:1994:PMS

[Kos94] Anna Kostoula. The problem of method in the study of the influence a philosophy has on scientific practice. The case of thermoelectricity. In Trends in the Historiography of Science [GCN94], chapter 27, pages 379–384. ISBN 90-481-4264-4, 94-017-3596-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-017-3596-4_27.

Koutougos:1989:RPP

[Kou89] Aris Koutougos. Research programmes and paradigms as dialogue structures. In Imre Lakatos and Theories of Scientific Change [GGN89], chapter 26, pages 361–374. ISBN 94-010-7860-2, 94-009-3025-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-009-
3025-4_26.

Koutougos:1990:PWR

[Kou90]  Aris Koutougos.  The Plato–

Wittgenstein route to the pragmatics of falsification. In Greek Studies in the Philosophy and History of Science [Nic90], chapter 25, pages 385–396. ISBN 94-010-7403-8, 94-009-2015-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-009-2015-
6_25.

Kourany:2011:IES

[Kou11] Janet A. Kourany. Integrating the ethical into scientific rationality. In Science in the Context of Application [CN11a], chapter 22, pages 371–386. ISBN 90-481-9050-9, 90-481-9051-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q180.55.M4 S385

2011. URL https://link.springer.
com/chapter/10.1007/978-90-481-
9051-5_22.

Kovac:2006:PES

[Kov06] Jeffrey Kovac. Professional ethics in science. In Philosophy of Chemistry: Synthesis of a New Discipline [BSM06b], chapter 8, pages 157–169. ISBN 1-4020-2455-X, 1-4020-3256-0, 1-4020-3261-7. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 242 QD6. URL https:/

                                              /link.springer.com/chapter/10.
                                              1007/1-4020-3261-7_8.

Kowarsch:2016:EGF

[Kow16a] Martin Kowarsch. Elements of a guideline for future integrated economic assessments of the IPCC. In A Pragmatist Orientation for the Social Sciences in Climate Policy: How to Make Integrated Economic Assessments Serve Society [Kow16j], chapter 11, pages 275–299. ISBN 3-319-43279-6, 3-319-43281-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN GE170; GE220; HC79.E5; QC903 .K683 2016; QC981.8.C5 K69 2016.

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-3-319-43281-
                                              6_11.

Kowarsch:2016:ECE

[Kow16b] Martin Kowarsch. Ethics in climate economics: Balance or bias? In A Pragmatist Orientation for the Social Sciences in Climate Policy: How to Make Integrated Economic Assessments Serve Society [Kow16j], chapter 8, pages 197–227. ISBN 3-319-43279-6, 3-319-43281-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN GE170; GE220; HC79.E5; QC903 .K683 2016; QC981.8.C5 K69 2016.

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-3-319-43281-
                                              6_8.

Kowarsch:2016:EIW

[Kow16c] Martin Kowarsch. An evaluation of the IPCC WG III assessments. In A Pragmatist Orientation for the Social Sciences in Climate Policy: How to Make Integrated Economic Assessments Serve Society [Kow16j], chapter 10, pages 249–272. ISBN 3-319-43279-6, 3-319-43281-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN GE170; GE220; HC79.E5; QC903 .K683 2016; QC981.8.C5 K69 2016.

URL https://link.springer.com/
chapter/10.1007/978-3-319-43281-
6_10.

Kowarsch:2016:FVC

[Kow16d] Martin Kowarsch. Fact/value conflation and the danger of the traditional models. In A Pragmatist Orientation for the Social Sciences in Climate Policy: How to Make Integrated Economic Assessments Serve Society [Kow16j], chapter 5, pages 101–132. ISBN 3-319-43279-6, 3-319-43281-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN GE170; GE220; HC79.E5; QC903 .K683 2016; QC981.8.C5 K69 2016. URL https:

//link.springer.com/chapter/10.
1007/978-3-319-43281-6_5.

Kowarsch:2016:FPS

[Kow16e] Martin Kowarsch. Fundamental perils for scientific assessments. In A Pragmatist Orientation for the Social Sciences in Climate Policy: How to Make Integrated Economic Assessments Serve Society [Kow16j], chapter 3, pages 47– 77. ISBN 3-319-43279-6, 3-319-43281-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN GE170; GE220; HC79.E5; QC903 .K683 2016; QC981.8.C5 K69 2016. URL https:

//link.springer.com/chapter/10.
1007/978-3-319-43281-6_3.

Kowarsch:2016:I

[Kow16f] Martin Kowarsch. Introduction. In A Pragmatist Orientation for the Social Sciences in Climate Policy: How to Make Integrated Economic Assessments Serve Society [Kow16j], chapter 1, pages 1–15. ISBN 3-319-43279-6, 3-319-43281-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN GE170; GE220; HC79.E5; QC903 .K683 2016; QC981.8.C5 K69 2016.

URL https://link.springer.com/
chapter/10.1007/978-3-319-43281-
6_1.

Kowarsch:2016:NIE

[Kow16g] Martin Kowarsch. The need for an integrated economic assessment of climate policy options. In A Pragmatist Orientation for the Social Sciences in Climate Policy: How to Make Integrated Economic Assessments Serve Society [Kow16j], chapter 2, pages 19– 46. ISBN 3-319-43279-6, 3-319-43281-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN GE170; GE220; HC79.E5; QC903 .K683 2016; QC981.8.C5 K69 2016. URL https:

//link.springer.com/chapter/10.
1007/978-3-319-43281-6_2.

Kowarsch:2016:PII

[Kow16h] Martin Kowarsch. Potential implications of the IPCC reform: Deliberative learning and difficulties of in-depth policy assessment. In A Pragmatist Orientation for the Social Sciences in Climate Policy: How to Make Integrated Economic Assessments Serve Society [Kow16j], chapter 12, pages 301–326. ISBN 3-319-43279-6, 3-319-43281-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN GE170; GE220; HC79.E5; QC903 .K683 2016; QC981.8.C5 K69 2016.

URL https://link.springer.com/
chapter/10.1007/978-3-319-43281-
6_12.

Kowarsch:2016:POD

[Kow16i] Martin Kowarsch. Pragmatism: Objectivity despite fact/value entanglement. In A Pragmatist Orientation for the Social Sciences in Climate Policy: How to Make Integrated Economic Assessments Serve Society [Kow16j], chapter 6, pages 133–169. ISBN 3-319-43279-6, 3-319-43281-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN GE170; GE220; HC79.E5; QC903 .K683 2016; QC981.8.C5 K69 2016.

URL https://link.springer.com/
chapter/10.1007/978-3-319-43281-
6_6.

Kowarsch:2016:POS

[Kow16j] Martin Kowarsch. A Pragmatist Orientation for the Social Sciences in Climate Policy: How to Make Integrated Economic Assessments Serve Society, volume 323 of Boston Studies in the Philosophy and History of Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2016. ISBN 3-319-43279-6, 3-319-43281-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). xvii + 326 pp. LCCN GE170; GE220; HC79.E5; QC903 .K683 2016; QC981.8.C5 K69 2016.

URL https://link.springer.com/
book/10.1007/978-3-319-43281-6.

Kowarsch:2016:PAG

[Kow16k] Martin Kowarsch. Prevalent actionguiding models of scientific expertise in policy. In A Pragmatist Orientation for the Social Sciences in Climate Policy: How to Make Integrated Economic Assessments Serve Society [Kow16j], chapter 4, pages 81–100. ISBN 3-319-43279-6, 3-319-43281-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN GE170; GE220; HC79.E5; QC903 .K683 2016; QC981.8.C5 K69 2016. URL https:

//link.springer.com/chapter/10.
1007/978-3-319-43281-6_4.

Kowarsch:2016:TTE

[Kow16l] Martin Kowarsch. Trust them? The epistemic quality of climate economics. In A Pragmatist Orientation for the Social Sciences in Climate Policy: How to Make Integrated Economic Assessments Serve Society [Kow16j], chapter 9, pages 229–247. ISBN 3-319-43279-6, 3-319-43281-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN GE170; GE220; HC79.E5; QC903 .K683 2016; QC981.8.C5 K69 2016.

URL https://link.springer.com/
chapter/10.1007/978-3-319-43281-
6_9.

Kowarsch:2016:UEI

[Kow16m] Martin Kowarsch. Understanding and evaluating the IAM–Based economics. In A Pragmatist Orientation for the Social Sciences in Climate Policy: How to Make Integrated Economic Assessments Serve Society [Kow16j], chapter 7, pages 173–195. ISBN 3-319-43279-6, 3-319-43281-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN GE170; GE220; HC79.E5; QC903 .K683 2016; QC981.8.C5 K69 2016.

URL https://link.springer.com/
chapter/10.1007/978-3-319-43281-
6_7.

Kox:1995:ESH

[Kox95] A. J. Kox. Einstein, specific heats, and residual rays: The history of a retracted paper. In No Truth Except in the Details: Essays in Honor of Martin J. Klein [KS95], chapter 9, pages 245–257. ISBN 94-010-4097-4, 94-011-0217-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-011-0217-
9_9.

Koyre:1976:FWC

[Koy76a] A. Koyré. The finite world of Copernicus. In The Concepts of Space and Time: Their Structure and Their Development [Čap76], chapter 11, pages 51–55. ISBN 90-277-0375-2, 94-010-1727-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic).

LCCN BD632. URL https://link.
springer.com/chapter/10.1007/
978-94-010-1727-5_11.

Koyre:1976:ISF

[Koy76b] A. Koyré. The infinite space in the Fourteenth Century. In The Concepts of Space and Time: Their Structure and Their Development [Čap76], chapter 10, pages 47–50. ISBN 90-277-0375-2, 94-010-1727-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN BD632.

URL https://link.springer.com/
chapter/10.1007/978-94-010-1727-
5_10.

Knei-Paz:1986:RIH

[KP86]  Baruch Knei-Paz.  Restricted and

inclusive historical materialism: A comment. In Ullmann-Margalit [UM86c], chapter 6, pages 85–92. ISBN 90-277-2160-2, 90-277-2161-0 (paperback), 94-009-4566-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 95

Q175. URL https://link.springer.

                                              com/chapter/10.1007/978-94-009-
                                              4566-1_6.

Kaern:1990:GSC

[KPC90] Michael Kaern, Bernard S. Phillips, and Robert S. Cohen. Georg Simmel and Contemporary Sociology, volume 119 of Boston Studies in the Philosophy of Science. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 1990. ISBN 94-010-6691-4 (hardcover), 94-009-0459-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). vi + 388 pp. LCCN HM401-1281. URL

                                              https://link.springer.com/book/
                                              10.1007/978-94-009-0459-0.

Krajewski:1974:ISL

[Kra74] Władysław Krajewski. The idea of statistical law in Nineteenth Century science. In Methodological and Historical Essays in the Natural and Social Sciences [CW74b], chapter 20, pages 397–405. ISBN 90-277-0378-7, 94-010-2128-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 14. URL https:/

                                              /link.springer.com/chapter/10.
                                              1007/978-94-010-2128-9_20.

Krajewski:1982:FCC

[Kra82a] Władysław Krajewski. Four conceptions of causation. In Polish Essays in the Philosophy of the Natural Sciences [Kra82c], chapter 20, pages 223–235. ISBN 94-009-7705-0 (e-book), 90-277-1287-5. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-009-7705-
                                              1_20.

Krajewski:1982:HH

[Kra82b] Władysław Krajewski. On hypotheses and hypotheticism. In Scientific Philosophy Today: Essays in Honor of Mario Bunge [AC82], chapter 5, pages 99–109. ISBN 90-277-1262-X, 90-277-1263-8, 94-009-8462-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 67.

URL https://link.springer.com/
chapter/10.1007/978-94-009-8462-
2_5.

Krajewski:1982:PEP

[Kra82c] Władysław Krajewski. Polish Essays in the Philosophy of the Natural Sciences, volume 68 of Boston Studies in the Philosophy of Science. D. Reidel, Dordrecht, The Netherlands; Boston, MA, USA; Lancaster, UK; Tokyo, Japan, 1982. ISBN 94-009-7705-0 (e-book), 90-277-1287-5. ISSN 0068-0346 (print), 2214-7942 (electronic). xxviii + 491 pp. LCCN ???? URL

https://link.springer.com/book/
10.1007/978-94-009-7705-1.

Krafft:1991:NCP

[Kra91] Fritz Krafft. The new celestial physics of Johannes Kepler. In Unguru [Ung91b], pages 185–227. ISBN 94-010-5476-2, 94-011-3342-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC5.53; QC6.9.

URL https://link.springer.com/
chapter/10.1007/978-94-011-3342-
5_9.

Kramnick:1992:ECS

[Kra92a] Isaac Kramnick. Eighteenth-Century science and radical social theory: The case of Joseph Priestley’s scientific liberalism. In The Scientific Enterprise: The Bar-Hillel Colloquium: Studies in History, Philosophy, and Sociology of Science, Volume 4 [UM92], chapter 1, pages 1–31. ISBN 94-010-5190-9, 94-011-2688-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .S354 1992. URL https://link.

                                              springer.com/chapter/10.1007/
                                              978-94-011-2688-5_1.

Krankis:1992:HCI

[Kra92b] Eda Krankis. Hybrid careers and the interaction of science and technology. In Technological Development and Science in the Industrial Age: New Perspectives on the Science–Technology Relationship [KB92b], chapter 8, pages 177–204. ISBN 90-481-4186-9, 94-015-8010-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-015-8010-
                                              6_8.

Krajewski:1995:ATD

[Kra95] Władysław Krajewski. Approximative truth and depth as the main aims of science. In Misiek [Mis95c], pages 119–123. ISBN 94-010-4206-3, 94-011-0461-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

                                              ???? URL https://link.springer.
                                              com/chapter/10.1007/978-94-011-
                                              0461-6_8.

Krajewski:1997:MET

[Kra97] Władysław Krajewski. Must the Explanans be true? In Ginev and Cohen [GC97], chapter 12, pages 217–222. ISBN 0-7923-4444-8 (hardcover), 94-011-5788-X. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 192. URL https://link.

                                              springer.com/chapter/10.1007/
                                              978-94-011-5788-9_12.

Kraemer:2000:MUA

[Kra00] Joel I. Kraemer. Maimonides’ use of (Aristotelian) dialectic. In Maimonides and the Sciences [CL00], chapter 7, pages 111–130. ISBN 90-481-5348-4, 94-017-2128-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-017-
2128-8_7.

Kragh:2007:HSH

[Kra07] Helge Kragh. History, science, and history of science. In Positioning the history of science [RG07], chapter 19, pages 105–107. ISBN 1-4020-5419-X (hardcover), 1-4020-5420-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q125 .P77 2007.

URL https://link.springer.com/
chapter/10.1007/1-4020-5420-3_19.

Krause:2011:MNI

[Kra11] Décio Krause. The metaphysics of non-individuality. In Brazilian Studies in Philosophy and History of Science: An account of recent works [KV11], chapter 19, pages 257–267. ISBN 90-481-9421-0 (hardcover), 90-481-9422-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175.3 .B73

2011. URL https://link.springer.
com/chapter/10.1007/978-90-481-
9422-3_19.

Kragh:2015:RBS

[Kra15a] Helge Kragh. From Ørsted to Bohr:the sciences and the Danish university system, 1800–1920. In Sciences in the Universities of Europe, Nineteenth and Twentieth Centuries: Academic Landscapes [SDG15b], chapter 3, pages 31–47. ISBN 94-017-9635-1, 94-017-9636-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q183.4.E85 2015. URL https://link.

springer.com/chapter/10.1007/
978-94-017-9636-1_3.

Kragh:2015:SBB

[Kra15b] Helge Kragh. On scientific biography and biographies of scientists. In Arabatzis et al. [ARS15], chapter 18, pages 269–280. ISBN 3-319-14552-5 (paperback), 3-319-14553-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q125 .R45 2015.

URL https://link.springer.com/
chapter/10.1007/978-3-319-14553-
2_18.

Kresic:1979:PDC

[Kre79] Andrija Krešić. Political dictatorship: The conflict of politics and society. In Praxis: Yugoslav Essays in the Philosophy and Methodology of the Social Sciences [MP79], chapter 7, pages 121–137. ISBN 90-277-0968-8, 94-009-9355-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-009-
9355-6_7.

Kremer:1992:PPM

[Kre91] Richard L. Kremer. From psychophysics to phenomenalism: Mach and Hering on color vision. In The Invention of Physical Science: Intersections of Mathematics, Theology and Natural Philosophy Since the Seventeenth Century: Essays in Honor of Erwin N. Hiebert [NRS92], chapter 7, pages 147–173. ISBN 94-010-5097-X, 94-011-2488-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-011-
2488-1_7.

Krejci:1996:DRA

[Kre96] Rudolph Krejci. Dissolution of the realism/antirealism problem. In Cohen et al. [CHR96], pages 11–18. ISBN 90-481-4493-0, 94-015-8638-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-015-8638-2_2.

Krimbas:1990:EEU

[Kri90]  Costas B. Krimbas.  Evolutionary

epistemology on universals as innate classificatory devices. In Greek Studies in the Philosophy and History of Science [Nic90], chapter 14, pages 209–225. ISBN 94-010-7403-8, 94-009-2015-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-009-
2015-6_14.

Krishna:1992:CME

[Kri92] V. V. Krishna. The colonial “model”

and the emergence of national science in India: 1876–1920. In Science and Empires: Historical Studies about Scientific Development and European Expansion [PJM92], chapter 8, pages 57–72. ISBN 0-7923-1518-9 (hardcover), 94-010-5145-3, 94-011-2594-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 136.

URL https://link.springer.com/
chapter/10.1007/978-94-011-2594-
9_8.

Krips:1994:CBL

[Kri94] Henry Krips. A critique of Bohr’s local realism. In Faye and Folse [FF94], chapter 12, pages 269–277. ISBN 0-7923-2378-5 (hardcover), 94-015-8106-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 153 1994; QC16 1994. URL https:

//link.springer.com/chapter/10.
1007/978-94-015-8106-6_12.

Kristeller:2003:RZM

[Kri03] Paul O. Kristeller. Remarks on Zilsel’s ‘the methods of humanism’. In The Social Origins of Modern Science [Zil03], chapter 4, pages 65–70. ISBN 0-7923-6457-0, 94-011-4142-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 200; Q175.5. URL https://link.

springer.com/chapter/10.1007/
978-94-011-4142-0_4.  Edited 59

years after the author’s death by Diederick Raven, Wolfgang Krohn, and R. S. (Robert Sonné) Cohen.

Kroes:1989:PST

[Kro89] Peter Kroes. Philosophy of science and the technological dimension of science. In Imre Lakatos and Theories of Scientific Change [GGN89], chapter 27, pages 375–382. ISBN 94-010-7860-2, 94-009-3025-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-009-3025-4_27.

Krober:1991:HSG

[Krö91] Günter Kröber. History of science in the GDR: Institutions and programmatic positions. In Woodward and Cohen [WC91], pages 49–62. ISBN 94-010-5395-2, 94-011-3164-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-011-3164-3_5.

Kroes:1992:RDE

[Kro92] Peter Kroes. On the role of design in engineering theories; Pambour’s theory of the steam engine. In Technological Development and Science in the Industrial Age: New Perspectives on the Science–Technology Relationship [KB92b], chapter 4, pages 69–98. ISBN 90-481-4186-9, 94-015-8010-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-015-8010-6_4.

Kromer:2012:WSB

[Krö12] Ralf Krömer. Are we still Babylonians? The structure of the foundations of mathematics from a Wimsattian perspective. In Characterizing the Robustness of Science: After the Practice Turn in Philosophy of Science [STNW12], chapter 8, pages 189–206. ISBN 94-007-2758-5, 94-007-2759-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175 .C4595 2012. URL https://link.

springer.com/chapter/10.1007/
978-94-007-2759-5_8.

Kruger:1988:HPS

[Krü88] Lorenz Krüger. How philosophy and science came to differ. In Science in Reflection: The Israel Colloquium: Studies in History, Philosophy, and Sociology of Science Volume 3 [UM88], chapter 4, pages 37–50. ISBN 90-277-2713-9, 94-009-2957-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-009-2957-9_4.

Kruger:1989:RR

[Krü89] Lorenz Krüger. Reduction without reductionism? In An Intimate Relation: Studies in the History and Philosophy of Science Presented to Robert E. Butts on his 60th Birthday [BM89], chapter 18, pages 369–390. ISBN 94-010-7546-8, 94-009-2327-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-009-2327-0_18.

Kruger:1990:MTS

[Krü90] Lorenz Krüger. Method, theory, and statistics: The lesson of physics. In Statistics in Science: The Foundations of Statistical Methods in Biology, Physics and Economics [CC90], chapter 1, pages 1–13. ISBN 94-010-6765-1, 94-009-0619-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QA276-280. URL https:/

                                              /link.springer.com/chapter/10.
                                              1007/978-94-009-0619-8_1.

Kruger:1991:IBC

[Krü91] Hans-Peter Krüger. Introduction to the basic concepts of communicationoriented science studies. In Woodward and Cohen [WC91], pages 199–208. ISBN 94-010-5395-2, 94-011-3164-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-011-3164-
                                              3_18.

Krohn:1983:ACO

[KS83] Wolfgang Krohn and Wolf Schäfer. Agricultural chemistry. The origin and structure of a finalized science. In Finalization in Science: The Social Orientation of Scientific Progress [Sch83b], chapter 3, pages 17–52. ISBN 94-009-7082-X, 94-009-7080-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-009-7080-9_3.

Kox:1995:NTE

[KS95] A. J. Kox and Daniel M. Siegel. No Truth Except in the Details: Essays in Honor of Martin J. Klein, volume 167 of Boston Studies in the Philosophy of Science. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 1995. ISBN 94-010-4097-4, 94-011-0217-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). xxv + 382 + 13 + 1 pp. LCCN ???? URL

https://link.springer.com/book/
10.1007/978-94-011-0217-9.

Krampitz:1997:TTE

[KSW97] Karl-Heinz Krampitz, Uwe Scheffler, and Horst Wessel. Time, truth and existence. In Perspectives on Time [FSU97b], chapter 15, pages 345–365. ISBN 90-481-4774-3, 94-015-8875-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol.189. URL https://link.

springer.com/chapter/10.1007/
978-94-015-8875-1_15.

Kamenka:1976:PAC

[KT76] Eugene Kamenka and Alice Erh-Soon Tay. Participation, ‘authenticity’ and the contemporary vision of man, law and society. In Essays in Memory of Imre Lakatos [CFW76], chapter 21, pages 335–358. ISBN 90-277-0655-7, 94-010-1451-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-010-
1451-9_21.

Kilmister:1993:PCR

[KT93] Clive Kilmister and Barrie Tonkinson.

Pragmatic circles in relativistic time keeping. In Correspondence, Invariance and Heuristics: Essays in Honour of Heinz Post [FK93], chapter 11, pages 207–225. ISBN 90-481-4229-6, 94-017-1185-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-017-
1185-2_11.

Kurs:2001:PDB

[KT01] Ott Kurs and Erki Tammiksaar. In political draughts between science and the humanities. In Estonian Studies in the History and Philosophy of Science [Vih01b], chapter 4, pages 51–62. ISBN 94-010-3863-5, 94-010-0672-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-010-0672-9_4.

Kucuradi:1995:KO

[Kuç95] Ioanna Kuçuradi. Knowledge and its object. In The Concept of Knowledge: The Ankara Seminar [KC95], chapter 8, pages 97–102. ISBN 90-481-4495-7, 94-017-3263-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

                                              ???? URL https://link.springer.
                                              com/chapter/10.1007/978-94-017-
                                              3263-5_8.

Kuhn:1971:NL

[Kuh71] Thomas S. Kuhn. Notes on Lakatos. In Buck and Cohen [BC71], pages 137–146. ISBN 90-277-0309-4, 94-010-3142-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

                                              ???? URL https://link.springer.
                                              com/chapter/10.1007/978-94-010-
                                              3142-4_8.

Kuhn:1987:GRT

[Kuh87] Dorothea Kuhn. Goethe’s relationship to the theories of development of his time. In Amrine et al. [AZW87], chapter 1, pages 3–15. ISBN 90-277-2265-X, 94-009-3761-X. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 97. US$79.00. URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-009-3761-1_1.

Kuhn:2009:HGB

[Kuh09] Rick Kuhn. Henryk Grossman: A biographical sketch. In Freudenthal and McLaughlin [FM09c], chapter 6, pages 239–252. ISBN 1-4020-9603-8, 1-4020-9604-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174.6 .S63 2009. URL https:

//link.springer.com/chapter/10.
1007/978-1-4020-9604-4_6.

Kulu:2001:KHI

[Kul01] Hill Kulu. Knowledge, human interests and migration studies. In Estonian Studies in the History and Philosophy of Science [Vih01b], chapter 19, pages 265–270. ISBN 94-010-3863-5, 94-010-0672-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-010-
0672-9_19.

Kumar:1992:PSA

[Kum92] Deepak Kumar. Problems in science administration: A study of the scientific surveys in British India 1757–1900. In Science and Empires: Historical Studies about Scientific Development and European Expansion [PJM92], chapter 28, pages 269–280. ISBN 0-7923-1518-9 (hardcover), 94-010-5145-3, 94-011-2594-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 136.

URL https://link.springer.com/
chapter/10.1007/978-94-011-2594-
9_28.

Kumar:2011:AAS

[Kum11] Deepak Kumar. Adoption and adaption: A study of medical ideas and techniques in Colonial India. In Science between Europe and Asia: Historical Studies on the Transmission, Adoption and Adaptation of Knowledge [GR11b], chapter 15, pages 233–243. ISBN 90-481-9968-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q125 .S354 2010; Q175.52.A78 S35

2010. URL https://link.springer.
com/chapter/10.1007/978-90-481-
9968-6_15.

Kung:1995:TCK

[Kün95] Guido Küng. Two concepts of knowing. In The Concept of Knowledge: The Ankara Seminar [KC95], chapter 1, pages 3–10. ISBN 90-481-4495-7, 94-017-3263-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-017-
3263-5_1.

Kuorikoski:2007:EE

[Kuo07] Jaakko Kuorikoski. Explaining with equilibria 68. In Rethinking Explanation [PY07], chapter 11, pages 149–162. ISBN 1-4020-5580-3 (paperback), 1-4020-5581-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN BD237 .R48 2007; Q175.32.E97 R48

2007. URL https://link.springer.
com/chapter/10.1007/978-1-4020-
5581-2_11.

Kurtz:1988:MR

[Kur88] Paul Kurtz. Is Marxism a religion? In Philosophy, History and Social Action: Essays in Honor of Lewis Feuer with an autobiographic essay by Lewis Feuer [HOO88c], chapter 13, pages 275–295. ISBN 94-010-7793-2, 94-009-2873-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-009-2873-
2_13.

Kuroda:1998:OM

[Kur98a] Wataru Kuroda. Other minds. In Japanese Studies in the Philosophy of Science [NC98], chapter 2, pages 7–19. ISBN 94-010-6176-9, 94-011-5175-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-011-5175-
7_2.

Kurosaki:1998:IS

[Kur98b] Hiroshi Kurosaki. On inference in science. In Japanese Studies in the Philosophy of Science [NC98], chapter 9, pages 141–150. ISBN 94-010-6176-9, 94-011-5175-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-011-
5175-7_9.

Kuusk:2001:PRT

[Kuu01] Piret Kuusk. Physical reality, theoretical physics, and mathematics. In Estonian Studies in the History and Philosophy of Science [Vih01b], chapter 15, pages 203–214. ISBN 94-010-3863-5, 94-010-0672-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-010-0672-9_15.

Kuvacic:1979:MCI

[Kuv79] Ivan Kuvačić. Middle class ideology.

In Praxis: Yugoslav Essays in the Philosophy and Methodology of the Social Sciences [MP79], chapter 21, pages 333–356. ISBN 90-277-0968-8, 94-009-9355-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-009-
9355-6_21.

Kuznetsov:1987:A

[Kuz87a] B. G. Kuznetsov. Afterword. In Reason and Being [FC87], chapter 14, pages 395–426. ISBN 90-277-2181-5, 94-009-4590-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 17. URL https:/

                                              /link.springer.com/chapter/10.
                                              1007/978-94-009-4590-6_14.

Kuznetsov:1987:D

[Kuz87b] B. G. Kuznetsov. Descartes. In Reason and Being [FC87], chapter 6, pages 148–177. ISBN 90-277-2181-5, 94-009-4590-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 17. URL https:/

                                              /link.springer.com/chapter/10.
                                              1007/978-94-009-4590-6_6.

Kuznetsov:1987:DCS

[Kuz87c] B. G. Kuznetsov. Dynamism and the critique of stationary being. In Reason and Being [FC87], chapter 9, pages 239–275. ISBN 90-277-2181-5, 94-009-4590-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 17. URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-009-4590-6_9.

Kuznetsov:1987:EE

[Kuz87d] B. G. Kuznetsov. Einstein and Epicurus. In Reason and Being [FC87], chapter 4, pages 67–108. ISBN 90-277-2181-5, 94-009-4590-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 17.

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-009-4590-
                                              6_4.

Kuznetsov:1987:EA

[Kuz87e] B. G. Kuznetsov. Existence and actuality. In Reason and Being [FC87], chapter 11, pages 298–324. ISBN 90-277-2181-5, 94-009-4590-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 17.

URL https://link.springer.com/
chapter/10.1007/978-94-009-4590-
6_11.

Kuznetsov:1987:GCS

[Kuz87f] B. G. Kuznetsov. The genesis of classical science and the problem of nonidentity. In Reason and Being [FC87], chapter 8, pages 211–238. ISBN 90-277-2181-5, 94-009-4590-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 17.

URL https://link.springer.com/
chapter/10.1007/978-94-009-4590-
6_8.

Kuznetsov:1987:HEC

[Kuz87g] B. G. Kuznetsov. The Heraclitean– Eleatic clash. In Reason and Being [FC87], chapter 2, pages 31–46. ISBN 90-277-2181-5, 94-009-4590-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 17.

URL https://link.springer.com/
chapter/10.1007/978-94-009-4590-
6_2.

Kuznetsov:1987:HB

[Kuz87h] B. G. Kuznetsov. Heterogeneous being. In Reason and Being [FC87], chapter 10, pages 279–297. ISBN 90-277-2181-5, 94-009-4590-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 17.

URL https://link.springer.com/
chapter/10.1007/978-94-009-4590-
6_10.

Kuznetsov:1987:I

[Kuz87i] B. G. Kuznetsov. Introduction. In Reason and Being [FC87], chapter 1, pages 3–28. ISBN 90-277-2181-5, 94-009-4590-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 17. URL https:/

                                              /link.springer.com/chapter/10.
                                              1007/978-94-009-4590-6_1.

Kuznetsov:1987:NV

[Kuz87j] B. G. Kuznetsov. Nothing and the vacuum. In Reason and Being [FC87], chapter 13, pages 358–394. ISBN 90-277-2181-5, 94-009-4590-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 17.

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-009-4590-
                                              6_13.

Kuznetsov:1987:PB

[Kuz87k] B. G. Kuznetsov. Paradoxes of being. In Reason and Being [FC87], chapter 3, pages 47–66. ISBN 90-277-2181-5, 94-009-4590-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 17. URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-009-4590-6_3.

Kuznetsov:1987:RR

[Kuz87l] B. G. Kuznetsov. The rationalism of the Renaissance. In Reason and Being [FC87], chapter 5, pages 109–147. ISBN 90-277-2181-5, 94-009-4590-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 17.

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-009-4590-
                                              6_5.

Kuznetsov:1987:SE

[Kuz87m] B. G. Kuznetsov. Spinoza and Einstein. In Reason and Being [FC87], chapter 7, pages 178–210. ISBN 90-277-2181-5, 94-009-4590-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic).

LCCN Q174 .B67 vol. 17. URL https:

//link.springer.com/chapter/10.
1007/978-94-009-4590-6_7.

Kuznetsov:1987:URN

[Kuz87n] B. G. Kuznetsov. Understanding and reason in Nineteenth- and Twentieth-Century science. In Reason and Being [FC87], chapter 12, pages 325–357. ISBN 90-277-2181-5, 94-009-4590-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 17. URL https://link.

springer.com/chapter/10.1007/
978-94-009-4590-6_12.

Krause:2011:BSP

[KV11] Décio Krause and Antonio Videira.

Brazilian Studies in Philosophy and History of Science: An account of recent works, volume 290 of Boston Studies in the Philosophy of Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2011. ISBN 90-481-9421-0 (hardcover), 90-481-9422-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). xiii + 352 pp. LCCN Q175.3 .B73 2011. URL

https://link.springer.com/book/
10.1007/978-90-481-9422-3.

Kyburg:1974:R

[Kyb74] Henry E. Kyburg. Randomness. In Schaffner and Cohen [SC74a], pages 137–149. ISBN 90-277-0408-2, 90-277-0409-0, 94-010-2140-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175 .B7312.

URL https://link.springer.com/
chapter/10.1007/978-94-010-2140-
1_10.

Kvasz:2009:PPF

[KZ09] Ladislav Kvasz and Eugen Zeleňák. A problem for Popper’s fallibilism. In Rethinking Popper [PC09], chapter 6, pages 71–81. ISBN 1-4020-9337-3, 1-4020-9338-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN B1649.P64 R48 2009; Q174 .B67 2009.

URL https://link.springer.com/
chapter/10.1007/978-1-4020-9338-
8_6.

Laor:1986:CDT

[LA86] Nathaniel Laor and Joseph Agassi. The computer as a diagnostic tool in medicine. In Philosophy and Technology II: Information Technology and Computers in Theory and Practice [MH86], chapter 16, pages 227–238. ISBN 94-010-8510-2, 94-009-4512-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-009-4512-8_16.

Lackey:1994:TVH

[Lac94] Douglas P. Lackey. Two vignettes in the history of the mensuration of value. In Artifacts, Representations and Social Practice: Essays for Marx Wartofsky [GC94], chapter 5, pages 69–86. ISBN 94-010-4390-6, 94-011-0902-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN B73 .A77 1994.

URL https://link.springer.com/
chapter/10.1007/978-94-011-0902-
4_5.

Lacey:2001:IMS

[Lac01] Hugh Lacey. Incommensurability and “multicultural science”. In Incommensurability and Related Matters [HHS01], chapter 9, pages 225–239. ISBN 90-481-5709-9, 94-015-9680-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-015-9680-0_9.

Lagueux:1996:HCA

[Lag96] Maurice Lagueux. How could anyone be irrational? In Québec Studies in the Philosophy of Science: Part II: Biology, Psychology, Cognitive Science and Economics Essays in Honor of Hugues Leblanc [MC96], chapter 11, pages 177–192. ISBN 94-010-6537-3, 94-009-0113-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-009-
0113-1_11.

Lagrange:1997:DES

[Lag97a] J. L. Lagrange. Differential equations for the solution of all problems of dynamics. In Boissonnade and Vagliente [BV97], pages 223–236. ISBN 90-481-4779-4, 94-015-8903-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC5.53; QC6.9.

URL https://link.springer.com/
chapter/10.1007/978-94-015-8903-
1_12.

Lagrange:1997:ECE

[Lag97b] J. L. Lagrange. The equilibrium of compressible and elastic fluids. In Boissonnade and Vagliente [BV97], pages 164–167. ISBN 90-481-4779-4, 94-015-8903-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC5.53; QC6.9. URL https:

//link.springer.com/chapter/10.
1007/978-94-015-8903-1_8.

Lagrange:1997:EIF

[Lag97c] J. L. Lagrange. The equilibrium of incompressible fluids. In Boissonnade and Vagliente [BV97], pages 140–163. ISBN 90-481-4779-4, 94-015-8903-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC5.53; QC6.9.

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-015-8903-
                                              1_7.

Lagrange:1997:GFD

[Lag97d] J. L. Lagrange. A general formula of dynamics for the motion of a system of bodies moved by arbitrary forces. In Boissonnade and Vagliente [BV97], pages 184–190. ISBN 90-481-4779-4, 94-015-8903-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC5.53; QC6.9. URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-015-8903-1_10.

Lagrange:1997:GFS

[Lag97e] J. L. Lagrange. A general formula of statics and its application to the equilibrium of an arbitrary system of forces. In Boissonnade and Vagliente [BV97], pages 26–37. ISBN 90-481-4779-4, 94-015-8903-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC5.53; QC6.9. URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-015-8903-1_2.

Lagrange:1997:GMA

[Lag97f] J. L. Lagrange. A general method of approximation for the problems of dynamics based on the variation of arbitrary constants. In Boissonnade and Vagliente [BV97], pages 236–252. ISBN 90-481-4779-4, 94-015-8903-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC5.53; QC6.9.

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-015-8903-
                                              1_13.

Lagrange:1997:GPE

[Lag97g] J. L. Lagrange. The general properties of equilibrium of a system of bodies deduced from the preceding formula.

In Boissonnade and Vagliente [BV97], pages 37–59. ISBN 90-481-4779-4, 94-015-8903-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC5.53; QC6.9. URL https:

//link.springer.com/chapter/10.
1007/978-94-015-8903-1_3.

Lagrange:1997:GPM

[Lag97h] J. L. Lagrange. General properties of motion deduced from the preceding formula. In Boissonnade and Vagliente [BV97], pages 190–223. ISBN 90-481-4779-4, 94-015-8903-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC5.53; QC6.9.

URL https://link.springer.com/
chapter/10.1007/978-94-015-8903-
1_11.

Lagrange:1997:MGS

[Lag97i] J. L. Lagrange. A more general and simpler way to use the formula of equilibrium presented in section II. In Boissonnade and Vagliente [BV97], pages 60–82. ISBN 90-481-4779-4, 94-015-8903-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC5.53; QC6.9. URL https:

//link.springer.com/chapter/10.
1007/978-94-015-8903-1_4.

Lagrange:1997:MSF

[Lag97j] J. L. Lagrange. The motion of a system of free bodies treated as mass points and acted upon by forces of attraction. In Boissonnade and Vagliente [BV97], pages 311–442. ISBN 90-481-4779-4, 94-015-8903-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC5.53; QC6.9. URL https:

//link.springer.com/chapter/10.
1007/978-94-015-8903-1_15.

Lagrange:1997:MCE

[Lag97k] J. L. Lagrange. The motion of compressible and elastic fluids. In Boissonnade and Vagliente [BV97], pages 560–573. ISBN 90-481-4779-4, 94-015-8903-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC5.53; QC6.9. URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-015-8903-1_20.

Lagrange:1997:MCB

[Lag97l] J. L. Lagrange. The motion of constrained bodies which interact in an arbitrary fashion. In Boissonnade and Vagliente [BV97], pages 442–467. ISBN 90-481-4779-4, 94-015-8903-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC5.53; QC6.9.

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-015-8903-
                                              1_16.

Lagrange:1997:MIF

[Lag97m] J. L. Lagrange. The motion of incompressible fluids. In Boissonnade and Vagliente [BV97], pages 521–560. ISBN 90-481-4779-4, 94-015-8903-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC5.53; QC6.9.

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-015-8903-
                                              1_19.

Lagrange:1997:PHb

[Lag97n] J. L. Lagrange. The principles of hydrodynamics. In Boissonnade and Vagliente [BV97], pages 517–521. ISBN 90-481-4779-4, 94-015-8903-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC5.53; QC6.9.

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-015-8903-
                                              1_18.

Lagrange:1997:PHa

[Lag97o] J. L. Lagrange. The principles of hydrostatics. In Boissonnade and Vagliente [BV97], pages 136–140. ISBN 90-481-4779-4, 94-015-8903-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC5.53; QC6.9.

URL https://link.springer.com/
chapter/10.1007/978-94-015-8903-
1_6.

Lagrange:1997:RM

[Lag97p] J. L. Lagrange. Rotational motion.

In Boissonnade and Vagliente [BV97], pages 467–517. ISBN 90-481-4779-4, 94-015-8903-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC5.53; QC6.9. URL https:

//link.springer.com/chapter/10.
1007/978-94-015-8903-1_17.

Lagrange:1997:SVP

[Lag97q] J. L. Lagrange. The solution of various problems of statics. In Boissonnade and Vagliente [BV97], pages 82–136. ISBN 90-481-4779-4, 94-015-8903-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC5.53; QC6.9.

URL https://link.springer.com/
chapter/10.1007/978-94-015-8903-
1_5.

Lagrange:1997:VPD

[Lag97r] J. L. Lagrange. The various principles of dynamics. In Boissonnade and Vagliente [BV97], pages 169–184. ISBN 90-481-4779-4, 94-015-8903-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC5.53; QC6.9.

URL https://link.springer.com/
chapter/10.1007/978-94-015-8903-
1_9.

Lagrange:1997:VPS

[Lag97s] J. L. Lagrange. The various principles of statics. In Boissonnade and Vagliente [BV97], pages 11–26. ISBN 90-481-4779-4, 94-015-8903-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC5.53; QC6.9.

URL https://link.springer.com/
chapter/10.1007/978-94-015-8903-
1_1.

Lagrange:1997:VSO

[Lag97t] J. L. Lagrange. The very small oscillations of an arbitrary system of bodies. In Boissonnade and Vagliente [BV97], pages 253–305. ISBN 90-481-4779-4, 94-015-8903-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC5.53; QC6.9. URL https:

//link.springer.com/chapter/10.
1007/978-94-015-8903-1_14.

Laird:2008:NMV

[Lai08] Walter Roy Laird. Nature, mechanics, and voluntary movement in Giuseppe Moletti’s lectures on the pseudo-Aristotelian mechanica. In Mechanics and Natural Philosophy Before the Scientific Revolution [LR08], chapter 6, pages 173–183. ISBN 1-4020-5966-3 (hardcover), 1-4020-5967-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QA802 .M43 2008.

URL https://link.springer.com/
chapter/10.1007/978-1-4020-5967-
4_6.

Lakatos:1971:HSR

[Lak71a] Imre Lakatos. History of science and its rational reconstructions. In Buck and Cohen [BC71], pages 91–136. ISBN 90-277-0309-4, 94-010-3142-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-010-3142-4_7.

Lakatos:1971:RC

[Lak71b] Imre Lakatos. Replies to critics.

In Buck and Cohen [BC71], pages 174–182. ISBN 90-277-0309-4, 94-010-3142-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-010-
3142-4_12.

Lambek:1982:IHM

[Lam82] J. Lambek. The influence of Heraclitus on modern mathematics. In Scientific Philosophy Today: Essays in Honor of Mario Bunge [AC82], chapter 6, pages 111–121. ISBN 90-277-1262-X, 90-277-1263-8, 94-009-8462-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 67.

URL https://link.springer.com/
chapter/10.1007/978-94-009-8462-
2_6.

Lambek:1995:NIN

[Lam95] J. Lambek. On the nominalistic interpretation of natural languages. In Québec Studies in the Philosophy of Science: Part I: Logic, Mathematics, Physics and History of Science Essays in Honor of Hugues Leblanc [MC95], chapter 5, pages 69–78. ISBN 94-010-7204-3, 94-009-1575-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-009-1575-6_5.

Lamb:1997:CEI

[Lam97] David Lamb. Communication with extraterrestrial intelligence: SETI and scientific methodology. In Ginev and Cohen [GC97], chapter 13, pages 223– 251. ISBN 0-7923-4444-8 (hardcover), 94-011-5788-X. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 192. URL https://link.

                                              springer.com/chapter/10.1007/
                                              978-94-011-5788-9_13.

Lambert:2012:MSD

[Lam12] Jacques Lambert. Mechanism and surgery: Dionis’ Anatomy (1690). In Philosophy of Behavioral Biology [PR12b], chapter 11, pages 263–283. ISBN 94-007-1950-7, 94-007-1951-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QH331 .P4679 2012; Q175 .B657 v.282.

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-007-4345-
                                              8_11.

Lambert:2013:MSD

[Lam13] Jacques Lambert. Mechanism and surgery: Dionis’ Anatomy (1690). In Garber and Roux [GR13b], chapter 11, pages 263–283. CODEN ???? ISBN 94-007-4344-0 (hardcover), 94-007-4345-9 (e-book). ISSN ???? LCCN QC6 .M3616 2013; QC6 .M43 2013.

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-007-4345-
                                              8_11.

Lamasse:2014:RBF

[Lam14]  Stéphane Lamassé.  Relationships

between French “practical arithmetics” and teaching? In Bernard and Proust [BP14b], chapter 6, pages 125–153. ISBN 94-007-5121-4 (hardcover), 94-007-5122-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q181 .S35 2014; Q181 S387 2014.

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-007-5122-
                                              4_6.

Lanczos:1967:RPW

[Lan67] C. Lanczos. Rationalism and the physical world. In Cohen and Wartofsky [CW67], pages 181–198. ISBN 90-277-0013-3, 94-010-3508-3. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175 .B731 v.3. URL https://link.

springer.com/chapter/10.1007/
978-94-010-3508-8_10.

Landon:1990:SDG

[Lan90] Richard Landon. The Stillman Drake Galileo Collection. In Levere and Shea [LS90], chapter 14, pages 321–337. ISBN 0-7923-0420-9, 94-009-1878-X. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 120.

URL https://link.springer.com/
chapter/10.1007/978-94-009-1878-
8_14.

Langermann:1991:ARM

[Lan91] Y. Tzvi Langermann. The astronomy of Rabbi Moses Isserles. In Unguru [Ung91b], pages 83–98. ISBN 94-010-5476-2, 94-011-3342-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC5.53; QC6.9.

URL https://link.springer.com/
chapter/10.1007/978-94-011-3342-
5_5.

Lang:1994:ILF

[Lan94] Berel Lang. Irony, Ltd., and the future of art. In Artifacts, Representations and Social Practice: Essays for Marx Wartofsky [GC94], chapter 6, pages 87–103. ISBN 94-010-4390-6, 94-011-0902-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN B73 .A77 1994. URL https://link.

springer.com/chapter/10.1007/
978-94-011-0902-4_6.

Langermann:2000:MRA

[Lan00] Y. Tzvi Langermann. Maimonides’ repudiation of astrology. In Maimonides and the Sciences [CL00], chapter 8, pages 131–157. ISBN 90-481-5348-4, 94-017-2128-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-017-
2128-8_8.

Laor:1995:RRP

[Lao95] Nathaniel Laor. Religion and rational philosophy: Coming of age. In Critical Rationalism, Metaphysics and Science: Essays for Joseph Agassi. Volume I [JL95a], chapter 7, pages 91–129. ISBN 94-010-4211-X, 94-011-0471-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-011-0471-5_7.

Laor:2017:TPA

[Lao17] Nathaniel Laor. To transcend the Popper–Agassi impasse. In Encouraging Openness: Essays for Joseph Agassi on the Occasion of His 90th Birthday [BAG17], chapter 38, pages 493–510. ISBN 3-319-57668-2 (hardcover), 3-319-57669-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN B833.E53 2017. URL https://link.

springer.com/chapter/10.1007/
978-3-319-57669-5_38.

Lapierre:1995:GQI

[Lap95] Serge Lapierre. Generalized quantifiers and inferences. In Québec Studies in the Philosophy of Science: Part I: Logic, Mathematics, Physics and History of Science Essays in Honor of Hugues Leblanc [MC95], chapter 3, pages 41– 55. ISBN 94-010-7204-3, 94-009-1575-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-009-1575-
6_3.

Large:1999:HCD

[Lar99] Duncan Large. Hermes contra Dionysus: Michel Serres’s critique of Nietzsche. In Nietzsche, Epistemology, and Philosophy of Science: Nietzsche and the Sciences II [Bab99a], chapter 12, pages 151–159. ISBN 90-481-5234-8, 94-017-2428-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-017-
2428-9_12.

Larrere:2013:TPE

[Lar13]  Catherine Larrère.  Two philoso-

phies of the environmental crisis. In Bergandi [Ber13c], chapter 10, pages 141–149. ISBN 94-007-5066-8, 94-007-5067-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN GE42 .S77 2013; Q175 .B73 v. 296.

URL https://link.springer.com/
chapter/10.1007/978-94-007-5067-
8_10.

Lastowski:1982:IST

[Łas82] Krzysztof Łastowski. The idealizational status of theoretical biology. In Polish Essays in the Philosophy of the Natural Sciences [Kra82c], chapter 30, pages 353–364. ISBN 94-009-7705-0 (e-book), 90-277-1287-5. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-009-7705-
1_30.

Laszlo:2001:CFW

[Las01] Pierre Laszlo. Conventionalities in formula writing. In Klein [Kle01b], chapter 4, pages 47–60. ISBN 1-4020-0100-2, 94-015-9737-5. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 222; QD7. URL https:

//link.springer.com/chapter/10.
1007/978-94-015-9737-1_4.

Laszlo:2015:USW

[Las15] Pierre Laszlo. The University of Strasbourg and world wars. In Sciences in the Universities of Europe, Nineteenth and Twentieth Centuries: Academic Landscapes [SDG15b], chapter 6, pages 89–105. ISBN 94-017-9635-1, 94-017-9636-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q183.4.E85 2015. URL https://link.

springer.com/chapter/10.1007/
978-94-017-9636-1_6.

Latsis:1983:RSR

[Lat83] Spiro J. Latsis. The role and status of the rationality principle in the social sciences. In Epistemology, Methodology, and the Social Sciences [CW83a], chapter 6, pages 123–151. ISBN 90-481-8376-6, 94-017-1458-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-017-1458-7_6.

Lauer:1974:MCS

[Lau74] Quentin Lauer. The Marxist conception of science. In Methodological and Historical Essays in the Natural and Social Sciences [CW74b], chapter 19, pages 377–396. ISBN 90-277-0378-7, 94-010-2128-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 14. URL https:

//link.springer.com/chapter/10.
1007/978-94-010-2128-9_19.

Laudan:1980:WWL

[Lau80a] Larry Laudan. Why was the logic of discovery abandoned? In Scientific Discovery, Logic, and Rationality [Nic80f], chapter 6, pages 173–183. ISBN 90-277-1070-8, 94-009-8986-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-009-8986-
3_6.

Laudan:1980:MMW

[Lau80b] Rachel Laudan. The method of multiple working hypotheses and the development of plate tectonic theory. In Scientific Discovery: Case Studies [Nic80e], chapter 17, pages 331–343. ISBN 90-277-1093-7, 94-009-9015-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 60. URL https://link.

springer.com/chapter/10.1007/
978-94-009-9015-9_17.

Laudan:1983:DDP

[Lau83] Larry Laudan. The demise of the demarcation problem. In Cohen and Laudan [CL83], chapter 6, pages 111–127. ISBN 94-009-7055-2, 94-009-7057-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-009-
7055-7_6.

Laudan:1987:REP

[Lau87] Rachel Laudan. The rationality of entertainment and pursuit. In Pitt and Pera [PP87], chapter 9, pages 203–220. ISBN 90-277-2417-2, 94-009-3779-2. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol.

98. US$64.00. URL https://link.

springer.com/chapter/10.1007/
978-94-009-3779-6_9.

Laudan:1989:MAF

[Lau89] Larry Laudan. For method: Or, against Feyerabend. In An Intimate Relation: Studies in the History and Philosophy of Science Presented to Robert E. Butts on his 60th Birthday [BM89], chapter 15, pages 299–317. ISBN 94-010-7546-8, 94-009-2327-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-009-2327-0_15.

Laurier:1996:FNT

[Lau96] Daniel Laurier. Function, normality and temporality. In Québec Studies in the Philosophy of Science: Part II: Biology, Psychology, Cognitive Science and Economics Essays in Honor of Hugues Leblanc [MC96], chapter 3, pages 25–52. ISBN 94-010-6537-3, 94-009-0113-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-009-
0113-1_3.

Laugier:2009:SRL

[Lau09] Sandra Laugier. Science and realism: The legacy of Duhem and Meyerson in contemporary American philosophy of science. In Brenner and Gayon [BG09], chapter 4, pages 91–112. ISBN 1-4020-9367-5, 1-4020-9368-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .F746 2009.

URL https://link.springer.com/
chapter/10.1007/978-1-4020-9368-
5_4.

Lavine:1988:JCM

[Lav88] Thelma Z. Lavine. Judaism in the culture of modernism. In Philosophy, History and Social Action: Essays in Honor of Lewis Feuer with an autobiographic essay by Lewis Feuer [HOO88c], chapter 14, pages 297–311. ISBN 94-010-7793-2, 94-009-2873-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-009-2873-
2_14.

Lawvere:1996:GDC

[Law96] F. William Lawvere. Grassmann’s dialectics and category theory. In Schubring [Sch96f], pages 255–264. ISBN 90-481-4758-1, 94-015-8753-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QA8.9-10.3.

URL https://link.springer.com/
chapter/10.1007/978-94-015-8753-
2_21.

Layzer:1974:CE

[Lay74] David Layzer. Cosmic evolution. In Seeger and Cohen [SC74b], pages 203– 213. ISBN 90-277-0390-6, 90-277-0376-0 (paperback), 94-010-2126-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 11.

URL https://link.springer.com/
chapter/10.1007/978-94-010-2126-
5_12.

Layton:1992:EJS

[Lay92] Edwin T. Layton, Jr. Escape from the jail of shape; dimensionality and engineering science. In Technological Development and Science in the Industrial Age: New Perspectives on the Science–Technology Relationship [KB92b], chapter 3, pages 35–68. ISBN 90-481-4186-9, 94-015-8010-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-015-8010-6_3.

Lazerowitz:1977:LPF

[Laz77] Morris Lazerowitz. The Language of Philosophy: Freud and Wittgenstein, volume 55 of Boston Studies in the Philosophy of Science. D. Reidel, Dordrecht, The Netherlands; Boston, MA, USA; Lancaster, UK; Tokyo, Japan, 1977. ISBN 90-277-0826-6. xiv + 209 pp. LCCN B808.5 .L38 1977; Q174 .B67 vol. 55. Synthese Library (volume 117).

Levi-Civita:2007:NPG

[LC07] Tullio Levi-Civita. Notion of parallelism on a general manifold and consequent geometrical specification of the Riemannian curvature (excerpts). In Renn et al. [RSS^{+}07b], pages 1081–1088. ISBN 1-4020-3999-9 (hardcover), 1-4020-4000-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 no. 250 v. 4; QC173.6 .G469

                                              2007. URL https://link.springer.
                                              com/chapter/10.1007/978-1-4020-

4000-9_48. Originally published as “Nozione di parallelismo in una varietà qualunque e conseguente specificazione geometrica della curvatura riemanniana”, Circolo Mathematico di Palermo, Rendiconti 42, 173–204 (1916).

Lenhard:2017:IMT

[LC17a] Johannes Lenhard and Martin Carrier. Introduction: Mathematics as a tool. In Mathematics as a Tool: Tracing New Roles of Mathematics in the Sciences [LC17b], chapter 1, pages 1–19. ISBN 3-319-54468-3 (hardcover), 3-319-54469-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN GE45.M37; GE45.M38; QA8.4.

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-3-319-54469-
                                              4_1.

Lenhard:2017:MTT

[LC17b] Johannes Lenhard and Martin Carrier.

Mathematics as a Tool: Tracing New Roles of Mathematics in the Sciences, volume 327 of Boston Studies in the Philosophy and History of Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2017. ISBN 3-319-54468-3 (hardcover), 3-319-54469-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). x + 286 pp. LCCN GE45.M37; GE45.M38; QA8.4. URL

https://link.springer.com/book/
10.1007/978-3-319-54469-4.

LeBihan:2008:CCW

[Le 08] Soazig Le Bihan. Commentary on “Can We Consider Quantum Mechanics to Be a Description of Reality?”, by Herve Zwirn. In Rethinking Scientific Change and Theory Comparison: Stabilities, Ruptures, Incommensurabilities? [SSHH08], chapter 15, pages 219–222. ISBN 1-4020-6274-5 (hardcover), 1-4020-6279-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175 .R397 2008. URL https://

link.springer.com/chapter/10.
1007/978-1-4020-6279-7_15.  See

[Zwi08].

Lynch:1997:APA

[LE97] Michael Lynch and Samuel Y. Edgerton, Jr. Abstract painting and astronomical image processing. In The Elusive Synthesis: Aesthetics and Science [Tau97a], chapter 5, pages 103–124. ISBN 0-7923-4763-3, 0-7923-3904-5, 94-009-1786-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-009-1786-6_5.

Leavens:1996:TTP

[Lea96] C. Richard Leavens. The “tunnelingtime problem” for electrons. In Bohmian Mechanics and Quantum Theory: an Appraisal [CFG96], chapter 8, pages 111–129. ISBN 90-481-4698-4, 94-015-8715-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC173.96-174.52.

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-015-8715-
                                              0_8.

Leblanc:1982:FIL

[Leb82] Hugues Leblanc. Free intuitionistic logic: A formal sketch. In Scientific Philosophy Today: Essays in Honor of Mario Bunge [AC82], chapter 7, pages 123–138. ISBN 90-277-1262-X, 90-277-1263-8, 94-009-8462-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 67.

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-009-8462-
                                              2_7.

Leblanc:1995:AFL

[Leb95] Hugues Leblanc. On axiomatizing free logic — and inclusive logic in the bargain. In Québec Studies in the Philosophy of Science: Part I: Logic, Mathematics, Physics and History of Science Essays in Honor of Hugues Leblanc [MC95], chapter 1, pages 1–22. ISBN 94-010-7204-3, 94-009-1575-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-009-1575-
                                              6_1.

Lecrivain:1986:SCM

[Lec86] André Lecrivain. Spinoza and Cartesian mechanics. In Spinoza and the Sciences [GN86], chapter 2, pages 15–60. ISBN 94-010-8511-0, 94-009-4514-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-009-4514-2_2.

Lechner:1990:SDM

[Lec90] Frank J. Lechner. Social differentiation and modernity: On Simmel’s macrosociology. In Georg Simmel and Contemporary Sociology [KPC90], chapter 10, pages 155–179. ISBN 94-010-6691-4 (hardcover), 94-009-0459-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN HM401-1281.

URL https://link.springer.com/
chapter/10.1007/978-94-009-0459-
0_10.

Lecuyer:1994:PTN

[Lec94] Bernard-Pierre Lecuyer. Probabilistic thinking, the natural sciences and the social sciences: Changing configurations (1800–1850). In Cohen [Coh94b], chapter 3, pages 135–152. ISBN 0-7923-2223-1, 94-017-3391-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 150.

URL https://link.springer.com/
chapter/10.1007/978-94-017-3391-
5_3.

Leccabue:2014:MCC

[Lec14] Fabrizio Leccabue. My collaboration with Cuban physicists. In The History of Physics in Cuba [BRW14a], chapter 25, pages 381–386. ISBN 94-017-8040-4, 94-017-8041-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC5.53; QC6.9; QC9 C9H57 2014; Q124.6-127.2.

URL https://link.springer.com/
chapter/10.1007/978-94-017-8041-
4_25.

Leduc:2018:COT

[Led18] Christian Leduc. Conducting observations and tests: Lambert’s theory of empirical science. In What Does it Mean to be an Empiricist?: Empiricisms in Eighteenth Century Sciences [BR18b], chapter 12, pages 215–233. ISBN 3-319-69858-3 (hardcover), 3-319-69860-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN B816 .W43 2018. URL https://link.

                                              springer.com/chapter/10.1007/
                                              978-3-319-69860-1_12.

Leeds:1974:SOS

[Lee74]  Anthony Leeds.  ‘Subjective’ and

‘objective’ in social anthropological epistemology. In Seeger and Cohen [SC74b], pages 349–361. ISBN 90-277-0390-6, 90-277-0376-0 (paperback), 94-010-2126-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 11. URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-010-2126-5_20.

Leeds:1984:SAS

[Lee84] Anthony Leeds. Sociobiology, antisociobiology, epistemology, and human nature. In Methodology, Metaphysics and the History of Science: In Memory of Benjamin Nelson [CW84b], chapter 8, pages 215–234. ISBN 94-009-6333-5, 94-009-6331-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-009-6331-3_8.

Lee:1986:MBW

[Lee86] Emanoel C. G. Lee. Medicine and the Boer War: Social and political consequences. In Ullmann-Margalit [UM86c], chapter 9, pages 113–139. ISBN 90-277-2160-2, 90-277-2161-0 (paperback), 94-009-4566-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 95

Q175. URL https://link.springer.
com/chapter/10.1007/978-94-009-

4566-1_9. See comment [Dol86].

Lefevre:1996:TS

[Lef96] Wolfgang Lefèvre. Tools of science. In Abstraction and Representation: Essays on the Cultural Evolution of Thinking [Dam96b], chapter 12, pages 395–404. ISBN 0-7923-3816-2, 94-015-8624-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 175. URL https://link.

springer.com/chapter/10.1007/
978-94-015-8624-5_12.

Lefevre:2001:BLN

[Lef01a] Wolfgang Lefèvre. Between Leibniz, Newton, and Kant: Philosophy and Science in the Eighteenth Century, volume 220 of Boston Studies in the Philosophy and History of Science. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 2001. ISBN 90-481-5774-9, 94-015-9729-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). xvi + 288 pp. LCCN ????

URL https://link.springer.com/
book/10.1007/978-94-015-9729-6.

Lefevre:2001:NAS

[Lef01b] Wolfgang Lefèvre. Natural or artificial systems? In Between Leibniz, Newton, and Kant: Philosophy and Science in the Eighteenth Century [Lef01a], chapter 10, pages 191–209. ISBN 90-481-5774-9, 94-015-9729-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-015-9729-6_10.

Lefevre:2003:DMW

[Lef03] Wolfgang Lefèvre. Darwin, Marx, and warranted progress. In Renn et al. [RDS03], chapter 30, pages 593–613. ISBN 1-4020-1284-5, 1-4020-1285-3 (paperback). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 234. URL https://link.

                                              springer.com/chapter/10.1007/
                                              978-94-010-0111-3_30.

Legrenzi:2003:NP

[Leg03] Paolo Legrenzi. Naïve probability. In Descartes’s Mathematical Thought [Sas03d], chapter 2, pages 43–55. ISBN 90-481-6487-7, 94-017-1225-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

                                              URL https://link.springer.com/
                                              chapter/10.1007/0-306-48123-5_2.

Lehrer:2000:JCQ

[Leh00] Keith Lehrer. Justification, coherence and Quine. In Knowledge, Language and Logic: Questions for Quine [OK00], chapter 5, pages 57–61. ISBN 1-4020-0253-X, 94-011-3933-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-011-3933-5_5.

Leinfellner:1983:MPV

[Lei83] Werner Leinfellner. Marxian paradigms versus microeconomic structures. In Epistemology, Methodology, and the Social Sciences [CW83a], chapter 7, pages 153–201. ISBN 90-481-8376-6, 94-017-1458-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

                                              ???? URL https://link.springer.
                                              com/chapter/10.1007/978-94-017-
                                              1458-7_7.

LeitAo:2003:PBT

[Lei03] Henrique Leitão. A periphery between two centres? In Simões et al. [SCD03b], chapter 2, pages 19–46. ISBN 1-4020-1259-4, 94-017-3584-0. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 v. 233. URL https:/

/link.springer.com/chapter/10.
1007/978-94-017-3584-1_2.

Leijenhorst:2004:HGL

[Lei04] Cees Leijenhorst. Hobbes and the Galilean law of free fall. In The Reception of the Galilean Science of Motion in Seventeenth-Century Europe [PT04], chapter 9, pages 165–184. ISBN 1-4020-2454-1, 1-4020-2455-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC133 .R43 2004.

URL https://link.springer.com/
chapter/10.1007/978-1-4020-2455-
9_9.

Lektorsky:1995:KCO

[Lek95] Vladislav A. Lektorsky. Knowledge and cultural objects. In The Concept of Knowledge: The Ankara Seminar [KC95], chapter 15, pages 191–196. ISBN 90-481-4495-7, 94-017-3263-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-017-3263-
5_15.

Lekka:2015:BML

[Lek15a] Vasia Lekka. “Bodies that matter”:

Living in the Nineteenth Century. In The Neurological Emergence of Epilepsy: The National Hospital for the Paralysed and Epileptic (1870–1895) [Lek15e], chapter 2, pages 13–37. ISBN 3-319-06292-1, 3-319-06293-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol.305.

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-3-319-06293-
                                              8_2.

Lekka:2015:DEE

[Lek15b] Vasia Lekka. Discovering epilepsy and epileptics in Victorian London. In The Neurological Emergence of Epilepsy: The National Hospital for the Paralysed and Epileptic (1870–1895) [Lek15e], chapter 4, pages 71–112. ISBN 3-319-06292-1, 3-319-06293-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol.305.

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-3-319-06293-
                                              8_4.

Lekka:2015:EAN

[Lek15c] Vasia Lekka. Epilepsy in the age of neurology. In The Neurological Emergence of Epilepsy: The National Hospital for the Paralysed and Epileptic (1870–1895) [Lek15e], chapter 5, pages 113–164. ISBN 3-319-06292-1, 3-319-06293-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol.305. URL https:/

                                              /link.springer.com/chapter/10.
                                              1007/978-3-319-06293-8_5.

Lekka:2015:I

[Lek15d] Vasia Lekka. Introduction. In The Neurological Emergence of Epilepsy: The National Hospital for the Paralysed and Epileptic (1870–1895) [Lek15e], chapter 1, pages 1–11. ISBN 3-319-06292-1, 3-319-06293-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol.305.

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-3-319-06293-
                                              8_1.

Lekka:2015:NEE

[Lek15e] Vasia Lekka. The Neurological Emergence of Epilepsy: The National Hospital for the Paralysed and Epileptic (1870–1895), volume 305 of Boston Studies in the Philosophy and History of Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2015. ISBN 3-319-06292-1, 3-319-06293-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). xiii + 209 pp. LCCN Q174 .B67 vol.305.

URL https://link.springer.com/
book/10.1007/978-3-319-06293-8.

Lekka:2015:TTF

[Lek15f] Vasia Lekka. Towards the Twenty-First Century. In The Neurological Emergence of Epilepsy: The National Hospital for the Paralysed and Epileptic (1870–1895) [Lek15e], chapter 6, pages 165–185. ISBN 3-319-06292-1, 3-319-06293-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol.305. URL https:/

/link.springer.com/chapter/10.
1007/978-3-319-06293-8_6.

Lekka:2015:UAT

[Lek15g] Vasia Lekka. Unrolling the archives’ thread: Epilepsy and epileptics at the National Hospital for the Paralysed and Epileptic. In The Neurological Emergence of Epilepsy: The National Hospital for the Paralysed and Epileptic (1870–1895) [Lek15e], chapter 3, pages 39–70. ISBN 3-319-06292-1, 3-319-06293-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol.305. URL https:/

/link.springer.com/chapter/10.
1007/978-3-319-06293-8_3.

Lelas:2000:B

[Lel00a] Sr– dan Lelas. Biosynthesis. In Science and Modernity: Toward an Integral Theory of Science [Lel00m], chapter 5, pages 64–79. ISBN 1-4020-0247-5, 94-010-9036-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic).

                                              LCCN BD581. URL https://link.
                                              springer.com/chapter/10.1007/
                                              978-94-010-9036-0_5.

Lelas:2000:DK

[Lel00b] Sr– dan Lelas. Divine knowledge. In Science and Modernity: Toward an Integral Theory of Science [Lel00m], chapter 1, pages 3–16. ISBN 1-4020-0247-5, 94-010-9036-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN BD581. URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-010-9036-0_1.

Lelas:2000:E

[Lel00c] Sr– dan Lelas. Evolution. In Science and Modernity: Toward an Integral Theory of Science [Lel00m], chapter 6, pages 80–96. ISBN 1-4020-0247-5, 94-010-9036-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic).

                                              LCCN BD581. URL https://link.
                                              springer.com/chapter/10.1007/
                                              978-94-010-9036-0_6.

Lelas:2000:FSI

[Lel00d] Sr– dan Lelas. The first surrogate: Ideal language. In Science and Modernity: Toward an Integral Theory of Science [Lel00m], chapter 2, pages 17–31. ISBN 1-4020-0247-5, 94-010-9036-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN BD581.

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-010-9036-
                                              0_2.

Lelas:2000:H

[Lel00e] Sr– dan Lelas. Humans. In Science and Modernity: Toward an Integral Theory of Science [Lel00m], chapter 7, pages 99–114. ISBN 1-4020-0247-5, 94-010-9036-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic).

LCCN BD581. URL https://link.
springer.com/chapter/10.1007/
978-94-010-9036-0_7.

Lelas:2000:KN

[Lel00f] Sr– dan Lelas. Knowledge naturalised.

In Science and Modernity: Toward an Integral Theory of Science [Lel00m], chapter 4, pages 49–63. ISBN 1-4020-0247-5, 94-010-9036-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN BD581. URL https:

//link.springer.com/chapter/10.
1007/978-94-010-9036-0_4.

Lelas:2000:L

[Lel00g] Sr– dan Lelas. Linguosynthesis. In Science and Modernity: Toward an Integral Theory of Science [Lel00m], chapter 10, pages 149–170. ISBN 1-4020-0247-5, 94-010-9036-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic).

LCCN BD581. URL https://link.
springer.com/chapter/10.1007/
978-94-010-9036-0_10.

Lelas:2000:MSE

[Lel00h] Sr– dan Lelas. Modern science: Experiment. In Science and Modernity: Toward an Integral Theory of Science [Lel00m], chapter 12, pages 198–223. ISBN 1-4020-0247-5, 94-010-9036-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN BD581.

URL https://link.springer.com/
chapter/10.1007/978-94-010-9036-
0_12.

Lelas:2000:MSL

[Lel00i] Sr– dan Lelas. Modern science: Language. In Science and Modernity:

Toward an Integral Theory of Science [Lel00m], chapter 13, pages 224–247. ISBN 1-4020-0247-5, 94-010-9036-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN BD581.

URL https://link.springer.com/
chapter/10.1007/978-94-010-9036-
0_13.

Lelas:2000:MSS

[Lel00j] Sr– dan Lelas. Modern science: Sociosynthesis. In Science and Modernity: Toward an Integral Theory of Science [Lel00m], chapter 14, pages 248–267. ISBN 1-4020-0247-5, 94-010-9036-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN BD581.

URL https://link.springer.com/
chapter/10.1007/978-94-010-9036-
0_14.

Lelas:2000:N

[Lel00k] Sr– dan Lelas. Neurosynthesis. In Science and Modernity: Toward an Integral Theory of Science [Lel00m], chapter 8, pages 115–131. ISBN 1-4020-0247-5, 94-010-9036-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic).

LCCN BD581. URL https://link.
springer.com/chapter/10.1007/
978-94-010-9036-0_8.

Lelas:2000:SM

[Lel00l] Sr– dan Lelas. Science and modernity. In Science and Modernity: Toward an Integral Theory of Science [Lel00m], chapter 11, pages 173–197. ISBN 1-4020-0247-5, 94-010-9036-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN BD581.

URL https://link.springer.com/
chapter/10.1007/978-94-010-9036-
0_11.

Lelas:2000:SMT

[Lel00m] Sr– dan Lelas. Science and Modernity:

Toward an Integral Theory of Science, volume 214 of Boston Studies in the Philosophy of Science. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 2000. ISBN 1-4020-0247-5, 94-010-9036-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). 316 (est.) pp. LCCN BD581. URL

https://link.springer.com/book/
10.1007/978-94-010-9036-0.

Lelas:2000:SEM

[Lel00n] Sr– dan Lelas. Science and the end of modernity. In Science and Modernity: Toward an Integral Theory of Science [Lel00m], chapter 15, pages 271–279. ISBN 1-4020-0247-5, 94-010-9036-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN BD581.

URL https://link.springer.com/
chapter/10.1007/978-94-010-9036-
0_15.

Lelas:2000:SSO

[Lel00o] Sr– dan Lelas. The second surrogate:

Objective knowledge. In Science and Modernity: Toward an Integral Theory of Science [Lel00m], chapter 3, pages 32–46. ISBN 1-4020-0247-5, 94-010-9036-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic).

LCCN BD581. URL https://link.
springer.com/chapter/10.1007/
978-94-010-9036-0_3.

Lelas:2000:T

[Lel00p] Sr– dan Lelas. Technosynthesis. In Science and Modernity: Toward an Integral Theory of Science [Lel00m], chapter 9, pages 132–148. ISBN 1-4020-0247-5, 94-010-9036-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic).

                                              LCCN BD581. URL https://link.
                                              springer.com/chapter/10.1007/
                                              978-94-010-9036-0_9.

Lemay:1982:SLS

[Lem82] Fernand Lemay. Some lessons in the sun. In Scientific Philosophy Today: Essays in Honor of Mario Bunge [AC82], chapter 8, pages 139–150. ISBN 90-277-1262-X, 90-277-1263-8, 94-009-8462-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 67. URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-009-8462-2_8.

Lenneberg:1963:RLF

[Len63] Eric H. Lenneberg. The relationship of language to the formation of concepts. In Wartofsky [War63], pages 48–54. ISBN 94-010-3265-3, 94-010-3263-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174

                                              .B678 1970. URL https://link.
                                              springer.com/chapter/10.1007/
                                              978-94-010-3263-6_4.

Lenk:1982:IAC

[Len82] Hans Lenk. Interpretative action constructs. In Scientific Philosophy Today: Essays in Honor of Mario Bunge [AC82], chapter 9, pages 151–157. ISBN 90-277-1262-X, 90-277-1263-8, 94-009-8462-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 67. URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-009-8462-2_9.

Lenk:1983:NER

[Len83] Hans Lenk. Notes on extended responsibility and increased technological power. In Philosophy and Technology [DR83], chapter 14, pages 195–210. ISBN 94-009-7126-5, 94-009-7124-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-009-7124-
0_14.

Lenk:1986:SPN

[Len86] Hans Lenk. Socio-philosophical notes on the implications of computer revolution. In Philosophy and Technology II: Information Technology and Computers in Theory and Practice [MH86], chapter 17, pages 239–245. ISBN 94-010-8510-2, 94-009-4512-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-009-4512-8_17.

Lenoir:1987:ELF

[Len87] Timothy Lenoir. The eternal laws of form: Morphotypes and the conditions of existence in Goethe’s biological thought. In Amrine et al. [AZW87], chapter 2, pages 17–28. ISBN 90-277-2265-X, 94-009-3761-X. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 97. US$79.00.

URL https://link.springer.com/
chapter/10.1007/978-94-009-3761-
1_2.

Lenoir:1988:SIO

[Len88] Timothy Lenoir. Social interests and the organic physics of 1847. In Science in Reflection: The Israel Colloquium: Studies in History, Philosophy, and Sociology of Science Volume 3 [UM88], chapter 14, pages 169–191. ISBN 90-277-2713-9, 94-009-2957-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-009-2957-9_14.

Lennon:1989:PMA

[Len89] Thomas M. Lennon. Physical and metaphysical atomism: 1666–1682. In An Intimate Relation: Studies in the History and Philosophy of Science Presented to Robert E. Butts on his 60th Birthday [BM89], chapter 5, pages 81–95. ISBN 94-010-7546-8, 94-009-2327-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

                                              ???? URL https://link.springer.
                                              com/chapter/10.1007/978-94-009-
                                              2327-0_5.

Leplin:1980:RMT

[Lep80] Jarrett Leplin. The role of models in theory construction. In Scientific Discovery, Logic, and Rationality [Nic80f], chapter 12, pages 267–283. ISBN 90-277-1070-8, 94-009-8986-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-009-8986-3_12.

Lepage:1995:PPL

[Lep95] François Lepage. Partial propositional logic. In Québec Studies in the Philosophy of Science: Part I: Logic, Mathematics, Physics and History of Science Essays in Honor of Hugues Leblanc [MC95], chapter 2, pages 23– 39. ISBN 94-010-7204-3, 94-009-1575-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-009-1575-
                                              6_2.

Leroux:1995:HME

[Ler95] Jean Leroux. Helmholtz and modern empiricism. In Québec Studies in the Philosophy of Science: Part I: Logic, Mathematics, Physics and History of Science Essays in Honor of Hugues Leblanc [MC95], chapter 17, pages 287–296. ISBN 94-010-7204-3, 94-009-1575-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-009-
1575-6_17.

Lesne:2017:TVT

[Les17] Annick Lesne. Time variable and time scales in natural systems and their modeling. In Time of Nature and the Nature of Time: Philosophical Perspectives of Time in Natural Sciences [BH17b], chapter 4, pages 55–66. ISBN 3-319-53723-7 (hardcover), 3-319-53725-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC173.59.S65 T56 2017. URL https:

//link.springer.com/chapter/10.
1007/978-3-319-53725-2_4.

Levy:1971:CRT

[Lev71] Edwin Levy. Competing radical translations: Examples, limitations and implications. In Buck and Cohen [BC71], pages 590–605. ISBN 90-277-0309-4, 94-010-3142-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-010-3142-4_40.

Levi:1976:PB

[Lev76] Isaac Levi. A paradox for the birds.

In Essays in Memory of Imre Lakatos [CFW76], chapter 23, pages 371–378. ISBN 90-277-0655-7, 94-010-1451-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-010-1451-
9_23.

Levi:1983:RM

[Lev83a] Isaac Levi. Response to Margalit. In Language, Logic and Method [CW83b], chapter 9, pages 189–195. ISBN 94-009-7704-2, 94-009-7702-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-009-7702-0_9.

Levi:1983:RS

[Lev83b] Isaac Levi. Response to Scheffler. In Language, Logic and Method [CW83b], chapter 8, pages 185–188. ISBN 94-009-7704-2, 94-009-7702-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-009-7702-0_8.  See

[Sch83e].

Levi:1983:TFG

[Lev83c] Isaac Levi. Truth, fallibility and the growth of knowledge. In Language, Logic and Method [CW83b], chapter 5, pages 153–174. ISBN 94-009-7704-2, 94-009-7702-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

                                              ???? URL https://link.springer.
                                              com/chapter/10.1007/978-94-009-
                                              7702-0_5.

Levine:1983:PSJ

[Lev83d] Hillel Levine. Paradise not surrendered: Jewish reactions to Copernicus and the growth of modern science. In Epistemology, Methodology, and the Social Sciences [CW83a], chapter 8, pages 203–225. ISBN 90-481-8376-6, 94-017-1458-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

                                              ???? URL https://link.springer.
                                              com/chapter/10.1007/978-94-017-
                                              1458-7_8.

Levinson:1986:ITV

[Lev86]  Paul Levinson.  Information tech-

nologies as vehicles of evolution. In Philosophy and Technology II: Information Technology and Computers in Theory and Practice [MH86], chapter 3, pages 29–47. ISBN 94-010-8510-2, 94-009-4512-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-009-
4512-8_3.

Levine:1987:QP

[Lev87] Joseph Levine. Quine on psychology. In Shimony and Nails [SN87], pages 259–290. ISBN 94-010-8168-9, 94-009-3735-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic).

LCCN ????  URL https://link.
springer.com/chapter/10.1007/
978-94-009-3735-2_15. See comments

[Shi87b, Aga87a].

Levere:1990:LLI

[Lev90] Trevor H. Levere. Lavoisier: Language, instruments, and the chemical revolution. In Levere and Shea [LS90], chapter 9, pages 207–223. ISBN 0-7923-0420-9, 94-009-1878-X. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 120. URL https:/

/link.springer.com/chapter/10.
1007/978-94-009-1878-8_9.

Levi:1994:RC

[Lev94] Isaac Levi. Rationality and commitment. In Artifacts, Representations and Social Practice: Essays for Marx Wartofsky [GC94], chapter 15, pages 257–275. ISBN 94-010-4390-6, 94-011-0902-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN B73 .A77 1994. URL https://link.

springer.com/chapter/10.1007/
978-94-011-0902-4_15.

Levine:1995:JT

[Lev95] Hillel Levine. Is Judaism thisworldly? In Science, Mind and Art: Essays on science and the humanistic understanding in art, epistemology, religion and ethics In honor of Robert S. Cohen [GSW95b], chapter 15, pages 255– 277. ISBN 0-7923-2990-2 (hardcover), 94-011-0469-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 165 BD418.3.

URL https://link.springer.com/
chapter/10.1007/978-94-011-0469-
2_15.

Levere:2007:BER

[Lev07] Trevor H. Levere. Between Enlightenment and romanticism: The case of Dr. Thomas Beddoes. In Hans Christian Ørsted and The Romantic Legacy In Science: Ideas, Disciplines, Practices [BCK07], chapter 14, pages 259–272. ISBN 1-4020-2979-9 (hardcover), 1-4020-2987-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC515.O37 H36 2007. URL https:

//link.springer.com/chapter/10.
1007/978-1-4020-2987-5_14.

Levanon:2016:RAC

[Lev16] Tamar Levanon. Relation, action and the continuity of transition. In Cosmological and Psychological Time [DR16], chapter 7, pages 125–142. ISBN 3-319-22589-8 (hardcover), 3-319-22590-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN BD638 C67 2016. URL https://link.

springer.com/chapter/10.1007/
978-3-319-22590-6_7.

Lewontin:1976:ETG

[Lew76] R. C. Lewontin. Evolution and the theory of games. In Topics in the Philosophy of Biology [GM76], chapter 13, pages 286–311. ISBN 90-277-0596-8, 94-010-1829-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-010-
1829-6_13.

Lewis:1996:IGT

[Lew96] Albert C. Lewis. The influence of Grassmann’s theory of tides on the Ausdehnungslehre. In Schubring [Sch96f], pages 29–36. ISBN 90-481-4758-1, 94-015-8753-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QA8.9-10.3.

URL https://link.springer.com/
chapter/10.1007/978-94-015-8753-
2_4.

Ley:1984:SMH

[Ley84] Herman Ley. Some ‘moments’ of Hegel’s relation to the sciences. In Hegel and the Sciences [CW84a], chapter 14, pages 241–270. ISBN 94-009-6235-5, 94-009-6233-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-009-6233-0_14.

Lin:1993:PCH

[LF93] Cheng-Hung Lin and Daiwie Fu.

Philosophy and Conceptual History of Science in Taiwan, volume 141 of Boston Studies in the Philosophy of Science. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 1993. ISBN 94-010-5103-8, 94-011-2500-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). xiv + 276 pp.

LCCN Q127.T28 .P455 1993. URL

                                              https://link.springer.com/book/
                                              10.1007/978-94-011-2500-0.

Lopez-Guazo:2001:MES

[LG01] Laura Suárez Y. Lopez-Guazo. The Mexican Eugenics Society. In The Reception of Darwinism in the Iberian World: Spain, Spanish America and Brazil [GPSR01], chapter 9, pages 143–151. ISBN 94-010-3885-6, 94-010-0602-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

                                              ???? URL https://link.springer.
                                              com/chapter/10.1007/978-94-010-
                                              0602-6_9.

Li:1996:CR

[Li96] Li Li. On complementarity reality. In Cohen et al. [CHR96], pages 343–357. ISBN 90-481-4493-0, 94-015-8638-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-015-8638-
                                              2_26.

Li:2018:RBH

[Li18] Bocong Li. On relationships between the history and philosophy of engineering. In Mitcham et al. [MLNZ18], chapter 1, pages 3–12. ISBN 3-319-62448-2, 3-319-62450-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN TA15 .P45 2018.

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-3-319-62450-
                                              1_1.

Lianghui:1996:PSC

[Lia96a] Qiu Lianghui. A preliminary study of the characteristics of metallurgical technology in Ancient China. In Chinese Studies in the History and Philosophy of Science and Technology [DC96], chapter 18, pages 219–241. ISBN 90-481-4546-5, 94-015-8717-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 180; U390 .N19 1996. URL https:

//link.springer.com/chapter/10.
1007/978-94-015-8717-4_18.

Liangying:1996:ERS

[Lia96b] Xu Liangying. Essay on the role of science and democracy in society. In Chinese Studies in the History and Philosophy of Science and Technology [DC96], chapter 2, pages 5–11. ISBN 90-481-4546-5, 94-015-8717-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 180; U390 .N19 1996. URL https:

//link.springer.com/chapter/10.
1007/978-94-015-8717-4_2.

Liangying:1996:PDP

[Lia96c] Xu Liangying. The problem of demarcation of periods in the history of science. In Chinese Studies in the History and Philosophy of Science and Technology [DC96], chapter 15, pages 185–197. ISBN 90-481-4546-5, 94-015-8717-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 180; U390 .N19 1996.

URL https://link.springer.com/
chapter/10.1007/978-94-015-8717-
4_15.

Lidz:1989:OR

[Lid89] Charles W. Lidz. ‘Objectivity’ and rapport. In The Qualitative-Quantitative Distinction in the Social Sciences [GM89b], chapter 4, pages 43–56. ISBN 90-481-8460-6, 94-017-3444-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 112.

URL https://link.springer.com/
chapter/10.1007/978-94-017-3444-
8_4.

Lieb:1976:RPL

[Lie76] Hans-Heinrich Lieb. On relating pragmatics, linguistics, and non-semiotic disciplines. In Language in Focus: Foundations, Methods and Systems: Essays in Memory of Yehoshua Bar-Hillel [Kas76b], chapter 13, pages 217–249. ISBN 90-277-0644-1, 90-277-0645-X, 94-010-1876-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN P101-120. URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-010-1876-0_13.

Lih:1993:BQS

[Lih93a] Ko-Wei Lih. Bao Qi-Shou and his polyhedral Hun Yuan Tu. In Philosophy and Conceptual History of Science in Taiwan [LF93], chapter 11, pages 209–220. ISBN 94-010-5103-8, 94-011-2500-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q127.T28 .P455 1993. URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-011-2500-0_11.

Lih:1993:OGT

[Lih93b] Ko-Wei Lih. From one gnomon to two gnomons: A methodological study of the method of double differences. In Philosophy and Conceptual History of Science in Taiwan [LF93], chapter 9, pages 149–165. ISBN 94-010-5103-8, 94-011-2500-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q127.T28 .P455 1993. URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-011-2500-0_9.

Lijmbach:1999:HE

[Lij99] Susanne Lijmbach. A hermeneutical ethology? In Fehér et al. [FKR99], pages 199–205. ISBN 90-481-5257-7, 94-015-9293-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-015-
9293-2_16.

Limoges:1994:MED

[Lim94]  Camille Limoges.  Milne-Edwards,

Darwin, Durkheim and the division of labour: A case study in reciprocal conceptual exchanges between the social and the natural sciences. In Cohen [Coh94b], chapter 10, pages 317–343. ISBN 0-7923-2223-1, 94-017-3391-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 150. URL https://link.

springer.com/chapter/10.1007/
978-94-017-3391-5_10.

Lindholm:1982:RHS

[Lin82] Lynn M. Lindholm. Is realistic history of science possible? In Scientific Philosophy Today: Essays in Honor of Mario Bunge [AC82], chapter 10, pages 159–186. ISBN 90-277-1262-X, 90-277-1263-8, 94-009-8462-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 67.

URL https://link.springer.com/
chapter/10.1007/978-94-009-8462-
2_10.

Lindberg:1991:KIL

[Lin91] David C. Lindberg. Kepler and the incorporeality of light. In Unguru [Ung91b], pages 229–250. ISBN 94-010-5476-2, 94-011-3342-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC5.53; QC6.9.

URL https://link.springer.com/
chapter/10.1007/978-94-011-3342-
5_10.

Lin:1993:CLA

[Lin93a] Chao-Tien Lin. Confirmation logic and its applications. In Philosophy and Conceptual History of Science in Taiwan [LF93], chapter 4, pages 33–59. ISBN 94-010-5103-8, 94-011-2500-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q127.T28 .P455 1993. URL https:

//link.springer.com/chapter/10.
1007/978-94-011-2500-0_4.

Lin:1993:SPC

[Lin93b]  Chao-Tien Lin.  Solutions to the

paradoxes of confirmation, Goodman’s paradox, and two new theories of confirmation. In Philosophy and Conceptual History of Science in Taiwan [LF93], chapter 2, pages 15–20. ISBN 94-010-5103-8, 94-011-2500-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q127.T28 .P455

1993. URL https://link.springer.
com/chapter/10.1007/978-94-011-
2500-0_2.

Lin:1993:HIS

[Lin93c] Cheng-Hung Lin. Hempel on inductive shortcomings in Craigian method. In Philosophy and Conceptual History of Science in Taiwan [LF93], chapter 3, pages 21–31. ISBN 94-010-5103-8, 94-011-2500-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q127.T28 .P455 1993. URL https:

//link.springer.com/chapter/10.
1007/978-94-011-2500-0_3.

Lin:1993:PLA

[Lin93d] Cheng-Hung Lin. Popper’s logical analysis of basic statements. In Philosophy and Conceptual History of Science in Taiwan [LF93], chapter 5, pages 61–71. ISBN 94-010-5103-8, 94-011-2500-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q127.T28 .P455 1993. URL https:

//link.springer.com/chapter/10.
1007/978-94-011-2500-0_5.

Lipkin:1986:GTA

[Lip86] Harry J. Lipkin. On Gamow’s theory of alpha-decay: A comment. In Ullmann-Margalit [UM86a], chapter 15, pages 187–192. ISBN 90-277-2159-9 (paperback), 90-277-2158-0 (hardcover), 94-009-5496-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 94 Q175.

URL https://link.springer.com/
chapter/10.1007/978-94-009-5496-

0_15. See [Stu86].

Lipsey:2017:JAM

[Lip17] Richard G. Lipsey. Joseph Agassi, the M^{2}T seminar, and his influence on my work. In Encouraging Openness: Essays for Joseph Agassi on the Occasion of His 90th Birthday [BAG17], chapter 37, pages 467–489. ISBN 3-319-57668-2 (hardcover), 3-319-57669-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN B833.E53 2017.

URL https://link.springer.com/
chapter/10.1007/978-3-319-57669-
5_37.

Liu:1993:AS

[Liu93] Shih-Chao Liu. On the analytic and the synthetic. In Philosophy and Conceptual History of Science in Taiwan [LF93], chapter 1, pages 3–14. ISBN 94-010-5103-8, 94-011-2500-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q127.T28 .P455

1993. URL https://link.springer.
com/chapter/10.1007/978-94-011-
2500-0_1.

Livingston:1992:GOC

[Liv92] Paisley Livingston. Girard and the origin of culture. In Understanding Origins: Contemporary Views on the Origin of Life, Mind and Society [VD92], chapter 5, pages 91–110. ISBN 90-481-4090-0, 94-015-8054-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-015-8054-0_5.

Lizhi:1996:CPM

[Liz96] Fang Lizhi. The concept of ‘primordial motion’: Past and present. In Chinese Studies in the History and Philosophy of Science and Technology [DC96], chapter 9, pages 95–102. ISBN 90-481-4546-5, 94-015-8717-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 180; U390 .N19 1996. URL https:

//link.springer.com/chapter/10.
1007/978-94-015-8717-4_9.

Levy-Leblond:1992:WDP

[LL92] Jean-Marc Lévy-Leblond. Why does physics need mathematics? In The Scientific Enterprise: The Bar-Hillel Colloquium: Studies in History, Philosophy, and Sociology of Science, Volume 4 [UM92], chapter 9, pages 145–161. ISBN 94-010-5190-9, 94-011-2688-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174

.S354 1992. URL https://link.
springer.com/chapter/10.1007/
978-94-011-2688-5_9.

Lerch:2014:ECP

[LL14] Irving A. Lerch and Irving A. Lerch. Engaging Cuban physicists through the APS/CPS partnership. In The History of Physics in Cuba [BRW14a], chapter 29, pages 403–405. ISBN 94-017-8040-4, 94-017-8041-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC5.53; QC6.9; QC9 C9H57 2014; Q124.6-127.2.

URL https://link.springer.com/
chapter/10.1007/978-94-017-8041-
4_29.

Lloyd:1986:SVE

[Llo86] Genevieve Lloyd. Spinoza’s version of the eternity of the mind. In Spinoza and the Sciences [GN86], chapter 9, pages 211–233. ISBN 94-010-8511-0, 94-009-4514-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-009-
4514-2_9.

Lloyd:2004:MMM

[Llo04] G. E. R. Lloyd. The Meno and the mysteries of mathematics. In Classics in the History of Greek Mathematics [Chr04a], chapter 8, pages 169–183. ISBN 1-4020-0081-2 (hardcover), 1-4020-2640-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QA22 .C596 2004. URL https:/

/link.springer.com/chapter/10.
1007/978-1-4020-2640-9_8.

Llored:2015:IMC

[Llo15] Jean-Pierre Noël Llored. Investigating the meaning of the ceteris paribus clause in chemistry. In Philosophy of Chemistry: Growth of a New Discipline [SM15b], chapter 14, pages 219–233. ISBN 94-017-9363-8 (hardcover), 94-017-9364-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QD6 .P455 2015; QD6 .S384 2014.

URL https://link.springer.com/
chapter/10.1007/978-94-017-9364-
3_14.

Lehrer:1997:APP

[LM97] Keith Lehrer and Johann Christian Marek. Austrian Philosophy Past and Present: Essays in Honor of Rudolf Haller, volume 190 of Boston Studies in the Philosophy of Science. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 1997. ISBN 94-010-6412-1, 94-011-5720-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). xiii + 274 pp. LCCN ???? URL

https://link.springer.com/book/
10.1007/978-94-011-5720-9.

LaPalmeReyes:1995:CTA

[LMR95] Marie La Palme Reyes, John Macnamara, and Gonzalo E. Reyes. A category–theoretic approach to Aristotle’s term logic, with special reference to syllogisms. In Québec Studies in the Philosophy of Science: Part I: Logic, Mathematics, Physics and History of Science Essays in Honor of Hugues Leblanc [MC95], chapter 4, pages 57– 68. ISBN 94-010-7204-3, 94-009-1575-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-009-1575-
6_4.

Liu:2018:HPS

[LMZ18] Dachun Liu, Carl Mitcham, and Junhai Zhao. Humanities perspectives on science, technology, and engineering in China. In Mitcham et al. [MLNZ18], chapter 25, pages 301–316. ISBN 3-319-62448-2, 3-319-62450-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN TA15 .P45 2018.

URL https://link.springer.com/
chapter/10.1007/978-3-319-62450-
1_25.

Locke:1976:ISD

[Loc76a] J. Locke. On infinite space and its difference from matter. In The Concepts of Space and Time: Their Structure and Their Development [Čap76], chapter 19, pages 107–112. ISBN 90-277-0375-2, 94-010-1727-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN BD632.

URL https://link.springer.com/
chapter/10.1007/978-94-010-1727-
5_19.

Locke:1976:SD

[Loc76b] J. Locke. On succession and duration.

In The Concepts of Space and Time: Their Structure and Their Development [Čap76], chapter 37, pages 211–224. ISBN 90-277-0375-2, 94-010-1727-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN BD632.

URL https://link.springer.com/
chapter/10.1007/978-94-010-1727-
5_37.

Lochak:1997:RQR

[Loc97]  Georges Lochak.  “Les recettes

qui réussissent toujours”. (French) [“Recipes that always succeed”]. In Cohen et al. [CHS97b], pages 143–148. ISBN 90-481-4809-X, 94-017-2732-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC173.96-174.52. URL https://link.

springer.com/chapter/10.1007/
978-94-017-2732-7_10.

Lodynski:1982:SRD

[Łod82] Andrzej Łodyński. Some remarks in defense of the incommensurability thesis. In Polish Essays in the Philosophy of the Natural Sciences [Kra82c], chapter 8, pages 91–102. ISBN 94-009-7705-0 (e-book), 90-277-1287-5. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-009-7705-1_8.

Lohkivi:2001:HBC

[Lõh01] Endla Lõhkivi. Herman Boerhaave — Communis Europae Praeceptor. In Estonian Studies in the History and Philosophy of Science [Vih01b], chapter 10, pages 139–150. ISBN 94-010-3863-5, 94-010-0672-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-010-0672-9_10.

Lombardi:2015:OAC

[Lom15] Olimpia Lombardi. The ontological autonomy of the chemical world: Facing the criticisms. In Philosophy of Chemistry: Growth of a New Discipline [SM15b], chapter 3, pages 23–38. ISBN 94-017-9363-8 (hardcover), 94-017-9364-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QD6 .P455 2015; QD6 .S384 2014.

URL https://link.springer.com/
chapter/10.1007/978-94-017-9364-
3_3.

Longo:1993:RIM

[Lon93] Giuseppe Longo. Remarks on information and mind. In Bridging the Gap: Philosophy, Mathematics, and Physics: Lectures on the Foundations of Science [CCG93], chapter 7, pages 141–146. ISBN 94-010-5101-1, 94-011-2496-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-011-2496-
6_7.

Longino:2012:KWM

[Lon12] Helen Longino. Knowledge for what? Monist, pluralist, pragmatist approaches to the sciences of behavior. In Philosophy of Behavioral Biology [PR12b], chapter 2, pages 25–40. ISBN 94-007-1950-7, 94-007-1951-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QH331 .P4679 2012; Q175 .B657 v.282. URL https:

//link.springer.com/chapter/10.
1007/978-94-007-1951-4_2.

Lopes:1992:BMN

[Lop92] Maria Margaret Lopes. Brazilian museums of natural history and international exchanges in the transition to the 20th Century. In Science and Empires: Historical Studies about Scientific Development and European Expansion [PJM92], chapter 20, pages 193–200. ISBN 0-7923-1518-9 (hardcover), 94-010-5145-3, 94-011-2594-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 136.

URL https://link.springer.com/
chapter/10.1007/978-94-011-2594-
9_20.

Lorentz:2007:CG

[Lor07a] Hendrik A. Lorentz. Considerations on gravitation. In Renn et al. [RSS^{+}07a], pages 113–126. ISBN 1-4020-3999-9. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 no. 250 v. 3; QC173.6 .G469 2007. URL https:

//link.springer.com/chapter/10.
1007/978-1-4020-4000-9_13.

Lorentz:2007:ONQ

[Lor07b] Hendrik A. Lorentz. Old and new questions in physics (excerpt). In Renn et al. [RSS^{+}07a], pages 287–301. ISBN 1-4020-3999-9. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 no. 250 v. 3; QC173.6 .G469

2007. URL https://link.springer.
com/chapter/10.1007/978-1-4020-

4000-9_21. Originally published as “Alte unde neue Fragen der Physik”, Physikalische Zeitschrift 11, 1234–1257 (1910).

Love:2015:CCE

[Lov15a] Alan C. Love. Conceptual change and evolutionary developmental biology.

In Conceptual Change in Biology: Scientific and Philosophical Perspectives on Evolution and Development [Lov15b], chapter 1, pages 1–54. ISBN 94-017-9411-1, 94-017-9412-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QH331 C66 2015; QH491; QH492 .C66 2014. URL https:

//link.springer.com/chapter/10.
1007/978-94-017-9412-1_1.

Love:2015:CCB

[Lov15b] Alan C. Love. Conceptual Change in Biology: Scientific and Philosophical Perspectives on Evolution and Development, volume 307 of Boston Studies in the Philosophy and History of Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2015. ISBN 94-017-9411-1, 94-017-9412-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). xvii + 490 + 54 pp. LCCN QH331 C66 2015; QH491; QH492 .C66 2014. URL

https://link.springer.com/book/
10.1007/978-94-017-9412-1.

Lowy:1986:ESE

[Löw86] Ilana Löwy. The epistemology of the science of an epistemologist of the sciences: Ludwik Fleck’s professional outlook and its relationship to his philosophical works. In Cognition and Fact: Materials on Ludwik Fleck [CS86], chapter 22, pages 421–442. ISBN 94-010-8504-8, 94-009-4498-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-009-4498-
5_22.

Lowy:1991:ICS

[Löw91] Ilana Löwy. The immunological construction of the self. In Organism and the Origins of Self [Tau91i], chapter 3, pages 43–75. ISBN 0-7923-1185-X, 94-011-3406-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QR180-189.5. URL https:/

/link.springer.com/chapter/10.
1007/978-94-011-3406-4_3.

Lowy:2014:SRM

[Löw14] Ilana Löwy. The scientists’ role and medical innovations. In The Institution of Science and the Science of Institutions: The Legacy of Joseph Ben-David [Her14b], chapter 4, pages 47–61. ISBN 94-007-7406-0 (hardcover), 94-007-7407-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 v.302; Q125 .I58 2014.

URL https://link.springer.com/
chapter/10.1007/978-94-007-7407-
0_4.

Lumiste:2001:NPC

[LP01] Ülo Lumiste and Helmut Piirimäe.

Newton’s Principia in the curricula of the University of Tartu (Dorpat) in the early 1690s. In Estonian Studies in the History and Philosophy of Science [Vih01b], chapter 1, pages 3–18. ISBN 94-010-3863-5, 94-010-0672-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-010-0672-9_1.

Lolli:2015:LPI

[LPV15] Gabriele Lolli, Marco Panza, and Giorgio Venturi. From Logic to Practice: Italian Studies in the Philosophy of Mathematics, volume 308 of Boston Studies in the Philosophy and History of Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2015. ISBN 3-319-10433-0, 3-319-10434-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). xix + 336 + 27 + 10 pp. LCCN QA8.4 .F76 2015; Q174 .B67 v.308. URL

https://link.springer.com/book/
10.1007/978-3-319-10434-8.

Laird:2008:MNP

[LR08] Walter Roy Laird and Sophie Roux. Mechanics and Natural Philosophy Before the Scientific Revolution, volume 254 of Boston Studies in the Philosophy of Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2008. ISBN 1-4020-5966-3 (hardcover), 1-4020-5967-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). viii + 306 pp. LCCN QA802 .M43 2008. URL

https://link.springer.com/book/
10.1007/978-1-4020-5967-4.

Levere:1990:NES

[LS90] Trevor Harvey Levere and William R. Shea, editors. Nature, Experiment, and the Sciences: Essays on Galileo and the History of Science in Honour of Stillman Drake, volume 120 of Boston Studies in the Philosophy of Science. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 1990. ISBN 0-7923-0420-9, 94-009-1878-X. ISSN 0068-0346 (print), 2214-7942 (electronic). xiii + 349 pp. LCCN Q174 .B67 vol. 120.

URL https://link.springer.com/
book/10.1007/978-94-009-1878-8.

Lubanski:1982:IS

[Lub82a] Mieczysław Lubański. Information and signal. In Polish Essays in the Philosophy of the Natural Sciences [Kra82c], chapter 23, pages 265–273. ISBN 94-009-7705-0 (e-book), 90-277-1287-5. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-009-7705-1_23.

Lubomirski:1982:PM

[Lub82b] Andrzej Lubomirski. On the philosophy of mathematics. In Polish Essays in the Philosophy of the Natural Sciences [Kra82c], chapter 21, pages 239–248. ISBN 94-009-7705-0 (e-book), 90-277-1287-5. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-009-7705-
1_21.

Lubenow:2015:USC

[Lub15] William C. Lubenow. University societies and clubs in Nineteenth- and Twentieth-Century Britain and their role in the promotion of research. In Sciences in the Universities of Europe, Nineteenth and Twentieth Centuries: Academic Landscapes [SDG15b], chapter 12, pages 193–210. ISBN 94-017-9635-1, 94-017-9636-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q183.4.E85 2015.

URL https://link.springer.com/
chapter/10.1007/978-94-017-9636-
1_12.

Luckyj:1988:PKU

[Luc88] George S. N. Luckyj. Panteleimon Kulish: A Ukrainian romantic conservative. In Philosophy, History and Social Action: Essays in Honor of Lewis Feuer with an autobiographic essay by Lewis Feuer [HOO88c], chapter 15, pages 313–321. ISBN 94-010-7793-2, 94-009-2873-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-009-
2873-2_15.

Luegenbiehl:2018:CSP

[Lue18] Heinz C. Luegenbiehl. Chinese student perceptions of engineering ethics. In Mitcham et al. [MLNZ18], chapter 20, pages 237–245. ISBN 3-319-62448-2, 3-319-62450-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN TA15 .P45 2018. URL https:

//link.springer.com/chapter/10.
1007/978-3-319-62450-1_20.

Lugg:1989:HDI

[Lug89] Andrew Lugg. History, discovery and induction: Whewell on Kepler on the orbit of Mars. In An Intimate Relation: Studies in the History and Philosophy of Science Presented to Robert E. Butts on his 60th Birthday [BM89], chapter 14, pages 283–298. ISBN 94-010-7546-8, 94-009-2327-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-009-2327-0_14.

Lutzen:1998:HHG

[Lüt98] Jesper Lützen. Heinrich Hertz and the geometrisation of mechanics. In Heinrich Hertz: Classical Physicist, Modern Philosopher [BHN98], chapter 8, pages 103–121. ISBN 90-481-4881-2, 94-015-8855-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-015-
8855-3_8.

Lutzen:2006:MIH

[Lüt06] Jesper Lützen. A mechanical image: Heinrich Hertz’s Principles of Mechanics. In Hendricks et al. [HJLP06], pages 45–64. ISBN 1-4020-5194-8 (hardcover), 1-4020-5195-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175 .B69 v.251 2006; Q175 .I51377 2006. URL https:

//link.springer.com/chapter/10.
1007/978-1-4020-5195-1_2.

Lintsen:1992:MEN

[LvHV92] H. Lintsen, G. van Hooff, and G. Verbong. Mechanical engineering in The Netherlands in the Nineteenth Century: Technology without a professional community. In Technological Development and Science in the Industrial Age: New Perspectives on the Science–Technology Relationship [KB92b], chapter 7, pages 155–176. ISBN 90-481-4186-9, 94-015-8010-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-015-
8010-6_7.

Lenhard:2011:HEC

[LW11] Johannes Lenhard and Eric Winsberg. Holism and entrenchment in climate model validation. In Science in the Context of Application [CN11a], chapter 8, pages 115–130. ISBN 90-481-9050-9, 90-481-9051-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q180.55.M4 S385

2011. URL https://link.springer.
com/chapter/10.1007/978-90-481-
9051-5_8.

Lizhi:1996:CST

[LY96] Fang Lizhi and Zhou Youyuan. Concepts of space and time in Ancient China and in modern cosmology. In Chinese Studies in the History and Philosophy of Science and Technology [DC96], chapter 6, pages 55–60. ISBN 90-481-4546-5, 94-015-8717-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 180; U390 .N19 1996. URL https:

//link.springer.com/chapter/10.
1007/978-94-015-8717-4_6.

Ly:2008:COV

[Ly08] Igor Ly. Commentary on “From One Version to the Other: Intra-Theoretical Change”, by Anouk Barberousse. In Rethinking Scientific Change and Theory Comparison: Stabilities, Ruptures, Incommensurabilities? [SSHH08], chapter 7, pages 103–105. ISBN 1-4020-6274-5 (hardcover), 1-4020-6279-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175 .R397

2008. URL https://link.springer.
com/chapter/10.1007/978-1-4020-
6279-7_7.

Meyer-Abich:1983:WST

[MA83] Klaus Michael Meyer-Abich. What sort of technology permits the language of nature? Conditions for controlling nature-domination constitutionally. In Philosophy and Technology [DR83], chapter 15, pages 211–232. ISBN 94-009-7126-5, 94-009-7124-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-009-7124-0_15.

Meyer-Abich:1987:SKF

[MA87] Klaus Michael Meyer-Abich. Selfknowledge, freedom, and irony: The language of nature in Goethe. In Amrine et al. [AZW87], chapter 18, pages 351–371. ISBN 90-277-2265-X, 94-009-3761-X. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 97. US$79.00. URL https:

//link.springer.com/chapter/10.
1007/978-94-009-3761-1_18.

Muzio:2014:SCR

[MA14] Oscar Arés Muzio and Ernesto Altshuler. Superconductivity in Cuba: Reaching the frontline. In The History of Physics in Cuba [BRW14a], chapter 15, pages 301–306. ISBN 94-017-8040-4, 94-017-8041-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC5.53; QC6.9; QC9 C9H57 2014; Q124.6-127.2.

URL https://link.springer.com/
chapter/10.1007/978-94-017-8041-
4_15.

Mancosu:2015:PSG

[MA15] Paolo Mancosu and Andrew Arana.

Plane and solid geometry: A note on purity of methods. In From Logic to Practice: Italian Studies in the Philosophy of Mathematics [LPV15], chapter 2, pages 23–31. ISBN 3-319-10433-0, 3-319-10434-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QA8.4 .F76 2015; Q174 .B67 v.308.

URL https://link.springer.com/
chapter/10.1007/978-3-319-10434-
8_2.

Maas:2017:JRT

[Maa17] Ad Maas. Johan Rudolph Thorbecke’s revenge: Objectivity and the rise of the Dutch nation state. In van Dongen and Paul [vDP17a], chapter 11, pages 173–193. ISBN 3-319-48892-9 (hardcover), 3-319-48893-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 v. 321; Q175.32.V57 E65 2017. URL https:

//link.springer.com/chapter/10.
1007/978-3-319-48893-6_11.

Machamer:1971:O

[Mac71] Peter K. Machamer. Observation.

In Buck and Cohen [BC71], pages 187–201. ISBN 90-277-0309-4, 94-010-3142-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-010-
3142-4_13.

Mackinnon:1974:OCQ

[Mac74] Edward Mackinnon. Ontic commitments of quantum mechanics. In Logical and Epistemological Studies in Contemporary Physics [CW74a], chapter 9, pages 255–308. ISBN 90-277-0391-4 (hardcover), 90-277-0377-9 (paperback), 94-010-2656-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol.

13; QC5.56. URL https://link.
springer.com/chapter/10.1007/
978-94-010-2656-7_9.

MacKinnon:1980:DNQ

[Mac80] Edward MacKinnon. The discovery of a new quantum theory. In Scientific Discovery: Case Studies [Nic80e], chapter 14, pages 261–272. ISBN 90-277-1093-7, 94-009-9015-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 60.

URL https://link.springer.com/
chapter/10.1007/978-94-009-9015-
9_14.

Maccormac:1986:MMC

[Mac86a] Earl R. Maccormac. Men and machines: The computational metaphor. In Philosophy and Technology II: Information Technology and Computers in Theory and Practice [MH86], chapter 11, pages 157–170. ISBN 94-010-8510-2, 94-009-4512-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-009-
4512-8_11.

Macintyre:1986:PSP

[Mac86b] Alasdair Macintyre. Positivism, sociology, and practical reasoning: Notes on Durkheim’s suicide. In Cohen and Wartofsky [CW85], chapter 5, pages 87–104. ISBN 90-277-1971-3, 94-009-5345-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.

com/chapter/10.1007/978-94-009-
5349-9_5.

MacLachlan:1990:DAP

[Mac90] James MacLachlan. Drake against the philosophers. In Levere and Shea [LS90], chapter 5, pages 123–144. ISBN 0-7923-0420-9, 94-009-1878-X. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 120.

URL https://link.springer.com/
chapter/10.1007/978-94-009-1878-
8_5.

Mach:1992:PD

[Mac92] Ernst Mach. Philosophical documents.

In Ernst Mach — A Deeper Look: Documents and New Perspectives [Bla92a], chapter 5, pages 111–126. ISBN 94-010-5230-1, 94-011-2771-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN C5.53; QC6.9.

URL https://link.springer.com/
chapter/10.1007/978-94-011-2771-
4_5.

Machamer:1994:SSH

[Mac94a] Peter Machamer. Selection, system and historiography. In Trends in the Historiography of Science [GCN94], chapter 11, pages 149–160. ISBN 90-481-4264-4, 94-017-3596-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-017-3596-4_11.

MacIntyre:1994:TFR

[Mac94b] Alasdair MacIntyre. The theses on Feuerbach: A road not taken. In Artifacts, Representations and Social Practice: Essays for Marx Wartofsky [GC94], chapter 16, pages 277–290. ISBN 94-010-4390-6, 94-011-0902-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN B73 .A77

1994. URL https://link.springer.
com/chapter/10.1007/978-94-011-
0902-4_16.

Mackinnon:1994:BRD

[Mac94c] Edward Mackinnon. Bohr and the realism debates. In Faye and Folse [FF94], chapter 13, pages 279–302. ISBN 0-7923-2378-5 (hardcover), 94-015-8106-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 153 1994; QC16 1994.

URL https://link.springer.com/
chapter/10.1007/978-94-015-8106-
6_13.

Macrakis:1997:A

[Mac97a] Michael S. Macrakis. Appendices. In Scarcity’s Ways: The Origins of Capital: A Critical Essay on Thermodynamics, Statistical Mechanics and Economics [Mac97f], chapter 7, pages 183–213. ISBN 90-481-4919-3, 94-015-8861-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN HB1-846.8. URL https://

link.springer.com/chapter/10.
1007/978-94-015-8861-4_7.

Macrakis:1997:Bb

[Mac97b] Michael S. Macrakis. Bibliography. In Scarcity’s Ways: The Origins of Capital: A Critical Essay on Thermodynamics, Statistical Mechanics and Economics [Mac97f], chapter 8, pages 215–221. ISBN 90-481-4919-3, 94-015-8861-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN HB1-846.8. URL https://

link.springer.com/chapter/10.
1007/978-94-015-8861-4_8.

Macrakis:1997:Ba

[Mac97c] Michael S. Macrakis. Biology. In Scarcity’s Ways: The Origins of Capital: A Critical Essay on Thermodynamics, Statistical Mechanics and Economics [Mac97f], chapter 3, pages 61–91. ISBN 90-481-4919-3, 94-015-8861-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN HB1-846.8. URL https://

link.springer.com/chapter/10.
1007/978-94-015-8861-4_3.

Macrakis:1997:E

[Mac97d] Michael S. Macrakis. Economics.

In Scarcity’s Ways: The Origins of Capital: A Critical Essay on Thermodynamics, Statistical Mechanics and Economics [Mac97f], chapter 2, pages 17–60. ISBN 90-481-4919-3, 94-015-8861-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN HB1-846.8. URL https://

link.springer.com/chapter/10.
1007/978-94-015-8861-4_2.

Macrakis:1997:PI

[Mac97e] Michael S. Macrakis. Preface & introduction. In Scarcity’s Ways: The Origins of Capital: A Critical Essay on Thermodynamics, Statistical Mechanics and Economics [Mac97f], chapter 1, pages 1–16. ISBN 90-481-4919-3, 94-015-8861-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

HB1-846.8.  URL https://link.
springer.com/chapter/10.1007/
978-94-015-8861-4_1.

Macrakis:1997:SWO

[Mac97f] Michael S. Macrakis. Scarcity’s Ways:

The Origins of Capital: A Critical Essay on Thermodynamics, Statistical Mechanics and Economics, volume 176 of Boston Studies in the Philosophy of Science. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 1997. ISBN 90-481-4919-3, 94-015-8861-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). ix + 230 pp. LCCN HB1-846.8. URL

                                              https://link.springer.com/book/
                                              10.1007/978-94-015-8861-4.

Macrakis:1997:SC

[Mac97g] Michael S. Macrakis. Summary and conclusions. In Scarcity’s Ways: The Origins of Capital: A Critical Essay on Thermodynamics, Statistical Mechanics and Economics [Mac97f], chapter 6, pages 173–181. ISBN 90-481-4919-3, 94-015-8861-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN HB1-846.8. URL https://

                                              link.springer.com/chapter/10.
                                              1007/978-94-015-8861-4_6.

Macrakis:1997:TGTa

[Mac97h] Michael S. Macrakis. Thermodynamics ([Greek text]), statistical mechanics, and capital. In Scarcity’s Ways: The Origins of Capital: A Critical Essay on Thermodynamics, Statistical Mechanics and Economics [Mac97f], chapter 4, pages 93–123. ISBN 90-481-4919-3, 94-015-8861-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN HB1-846.8. URL https://

                                              link.springer.com/chapter/10.
                                              1007/978-94-015-8861-4_4.

Macrakis:1997:TGTb

[Mac97i] Michael S. Macrakis. Thermodynamics ([Greek text]), work, and capital. In Scarcity’s Ways: The Origins of Capital: A Critical Essay on Thermodynamics, Statistical Mechanics and Economics [Mac97f], chapter 5, pages 125–172. ISBN 90-481-4919-3, 94-015-8861-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN HB1-846.8. URL https://

link.springer.com/chapter/10.
1007/978-94-015-8861-4_5.

MacLeod:2000:SPW

[Mac00] Roy M. MacLeod, editor. Science and the Pacific War: science and survival in the Pacific, 1939–1945, volume 207 of Boston Studies in the Philosophy of Science. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 2000. ISBN 0-7923-5851-1. xiii + 320 pp. LCCN D810.S2

S35 2000. URL http://www.loc.
gov/catdir/enhancements/fy0829/
99016758-d.html;   http://www.
loc.gov/catdir/enhancements/
fy0829/99016758-t.html.

MacLeod:2001:SBS

[Mac01]  Roy MacLeod.  “Strictly for the

birds”: Science, the military and the Smithsonian’s Pacific Ocean Biological Survey Program, 1963–1970. In Allen and MacLeod [AM01b], chapter 19, pages 307–337. ISBN 1-4020-0495-8 (hardcover), 94-017-2956-5. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 228; Q175.55.

URL https://link.springer.com/
chapter/10.1007/978-94-017-2956-
7_19.

Mackinnon:2012:BMB

[Mac12a] Edward Mackinnon. Beyond a minimal basis. In Interpreting Physics: Language and the Classical/Quantum Divide [Mac12d], chapter 6, pages 151–186. ISBN 94-007-2368-7, 94-007-2369-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC174.13 .M335 2012. URL https:

//link.springer.com/chapter/10.
1007/978-94-007-2369-6_6.

Mackinnon:2012:CQC

[Mac12b] Edward Mackinnon. From categories to quantitative concepts. In Interpreting Physics: Language and the Classical/ Quantum Divide [Mac12d], chapter 2, pages 25–68. ISBN 94-007-2368-7, 94-007-2369-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC174.13 .M335 2012. URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-007-2369-6_2.

Mackinnon:2012:ICP

[Mac12c] Edward Mackinnon. The interpretation of classical physics. In Interpreting Physics: Language and the Classical/ Quantum Divide [Mac12d], chapter 4, pages 89–119. ISBN 94-007-2368-7, 94-007-2369-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC174.13 .M335 2012. URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-007-2369-6_4.

MacKinnon:2012:IPL

[Mac12d] Edward MacKinnon. Interpreting Physics: Language and the Classical/ Quantum Divide, volume 289 of Boston Studies in the Philosophy of Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2012. ISBN 94-007-2368-7, 94-007-2369-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). xiii + 268 pp. LCCN QC174.13 .M335 2012.

                                              URL https://link.springer.com/
                                              book/10.1007/978-94-007-2369-6.

Mackinnon:2012:IQM

[Mac12e] Edward Mackinnon. Interpreting quantum mechanics. In Interpreting Physics: Language and the Classical/Quantum Divide [Mac12d], chapter 7, pages 187–212. ISBN 94-007-2368-7, 94-007-2369-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC174.13 .M335 2012. URL https:

//link.springer.com/chapter/10.
1007/978-94-007-2369-6_7.

Mackinnon:2012:OQM

[Mac12f] Edward Mackinnon. Orthodox quantum mechanics. In Interpreting Physics: Language and the Classical/Quantum Divide [Mac12d], chapter 5, pages 129–149. ISBN 94-007-2368-7, 94-007-2369-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC174.13 .M335 2012. URL https:

//link.springer.com/chapter/10.
1007/978-94-007-2369-6_5.

Mackinnon:2012:PO

[Mac12g] Edward Mackinnon. A philosophical overview. In Interpreting Physics: Language and the Classical/Quantum Divide [Mac12d], chapter 1, pages 3–24. ISBN 94-007-2368-7, 94-007-2369-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC174.13 .M335 2012. URL https:

//link.springer.com/chapter/10.
1007/978-94-007-2369-6_1.

Mackinnon:2012:RR

[Mac12h] Edward Mackinnon. Realism and reductionism. In Interpreting Physics: Language and the Classical/Quantum Divide [Mac12d], chapter 8, pages 213–257. ISBN 94-007-2368-7, 94-007-2369-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC174.13 .M335 2012. URL https:

//link.springer.com/chapter/10.
1007/978-94-007-2369-6_8.

Mackinnon:2012:UCP

[Mac12i] Edward Mackinnon. The unification of classical physics. In Interpreting Physics: Language and the Classical/ Quantum Divide [Mac12d], chapter 3, pages 69–87. ISBN 94-007-2368-7, 94-007-2369-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC174.13 .M335 2012. URL https:

//link.springer.com/chapter/10.
1007/978-94-007-2369-6_3.

MacLeod:2017:SBL

[Mac17] Miles MacLeod. Systems biology in the light of uncertainty: The limits of computation. In Mathematics as a Tool: Tracing New Roles of Mathematics in the Sciences [LC17b], chapter 7, pages 117–136. ISBN 3-319-54468-3 (hardcover), 3-319-54469-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN GE45.M37; GE45.M38; QA8.4. URL https:/

/link.springer.com/chapter/10.
1007/978-3-319-54469-4_7.

Madden:1969:CNN

[Mad69] Edward H. Madden. Causality and the notion of necessity. In Cohen and Wartofsky [CW69b], pages 450–462. ISBN 94-010-3380-3, 94-010-3378-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-010-3378-
7_16.

Mahdi:1975:SPR

[Mah75] Muhsin Mahdi. Science, philosophy, and religion in Alfarabi’s enumeration of the sciences. In Murdoch and Sylla [MS75b], pages 113–147. ISBN 90-277-0587-9, 94-010-1781-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN B721 .I584 1975.

URL https://link.springer.com/
chapter/10.1007/978-94-010-1781-
7_5.

Mahoney:2011:WMC

[Mah11] Michael S. Mahoney. What makes computer science a science? In Science in the Context of Application [CN11a], chapter 23, pages 389–408. ISBN 90-481-9050-9, 90-481-9051-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q180.55.M4 S385

2011. URL https://link.springer.
com/chapter/10.1007/978-90-481-
9051-5_23.

Mahootian:2015:MCE

[Mah15] Farzad Mahootian. Metaphor in chemistry: An examination of chemical metaphor. In Philosophy of Chemistry: Growth of a New Discipline [SM15b], chapter 9, pages 121–139. ISBN 94-017-9363-8 (hardcover), 94-017-9364-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QD6 .P455 2015; QD6 .S384 2014. URL https:

//link.springer.com/chapter/10.
1007/978-94-017-9364-3_9.

Maienschein:2001:AHS

[Mai01] Jane Maienschein. Advocating the history of science. In Allen and MacLeod [AM01b], chapter 2, pages 23–36. ISBN 1-4020-0495-8 (hardcover), 94-017-2956-5. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 228; Q175.55. URL https:

//link.springer.com/chapter/10.
1007/978-94-017-2956-7_2.

Maienschein:2012:HPS

[Mai12] Jane Maienschein. History and philosophy of science at work: Making regenerative medicine research better. In Mauskopf and Schmaltz [MS12b], chapter 12, pages 201–220. ISBN 94-007-1744-X, 94-007-1745-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175 .I5134 2011.

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-007-1745-
                                              9_12.

Maiese:2016:RED

[Mai16] Michelle Maiese. The reason/emotion divide in contemporary philosophy of psychology. In Meta-Philosophical Reflection on Feminist Philosophies of Science [AV16a], chapter 7, pages 113–132. ISBN 3-319-26346-3 (hardcover), 3-319-26348-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN HQ1154 .M48 2016; HQ1190 M48 2016. URL https:/

                                              /link.springer.com/chapter/10.
                                              1007/978-3-319-26348-9_7.

Majer:1998:HHP

[Maj98] Ulrich Majer. Heinrich Hertz’s pictureconception of theories: Its elaboration by Hilbert, Weyl, and Ramsey. In Heinrich Hertz: Classical Physicist, Modern Philosopher [BHN98], chapter 13, pages 225–242. ISBN 90-481-4881-2, 94-015-8855-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

                                              ???? URL https://link.springer.
                                              com/chapter/10.1007/978-94-015-
                                              8855-3_13.

Majer:2006:HAA

[Maj06] Ulrich Majer. Hilbert’s axiomatic approach to the foundations of science — a failed research program? In Hendricks et al. [HJLP06], pages 155–184. ISBN 1-4020-5194-8 (hardcover), 1-4020-5195-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175 .B69 v.251 2006; Q175 .I51377

                                              2006. URL https://link.springer.
                                              com/chapter/10.1007/978-1-4020-
                                              5195-1_5.

Maki:1996:SRS

[Mäk96] Uskali Mäki. Scientific realism and some peculiarities of economics. In Cohen et al. [CHR96], pages 427–447. ISBN 90-481-4493-0, 94-015-8638-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-015-8638-
2_32.

Mali:1992:SPN

[Mal92] Joseph Mali. Science, politics, and the new science of politics: A comment. In The Scientific Enterprise: The Bar-Hillel Colloquium: Studies in History, Philosophy, and Sociology of Science, Volume 4 [UM92], chapter 2, pages 33–41. ISBN 94-010-5190-9, 94-011-2688-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .S354 1992. URL https:

//link.springer.com/chapter/10.
1007/978-94-011-2688-5_2.

Malament:2003:ROR

[Mal03] David B. Malament. On relative orbital rotation in relativity theory. In Renn et al. [RDS03], chapter 8, pages 175–190. ISBN 1-4020-1284-5, 1-4020-1285-3 (paperback). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 234. URL https:/

/link.springer.com/chapter/10.
1007/978-94-010-0111-3_8.

Malet:2012:BME

[Mal12] Antoni Malet. Between mathematics and experimental philosophy: Hydrostatics in Scotland about 1700. In Philosophy of Behavioral Biology [PR12b], chapter 7, pages 159–187. ISBN 94-007-1950-7, 94-007-1951-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QH331 .P4679 2012; Q175 .B657 v.282.

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-007-4345-
                                              8_7.

Malet:2013:BME

[Mal13] Antoni Malet. Between mathematics and experimental philosophy: Hydrostatics in Scotland about 1700. In Garber and Roux [GR13b], chapter 7, pages 159–187. CODEN ???? ISBN 94-007-4344-0 (hardcover), 94-007-4345-9 (e-book). ISSN ???? LCCN QC6 .M3616 2013; QC6 .M43 2013.

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-007-4345-
                                              8_7.

Martin:1969:UPO

[MAM69] Michael Martin, Judith B. Agassi, and Sidney W. Mintz. Understanding and participant observation in cultural and social anthropology. In Cohen and Wartofsky [CW69b], pages 303–349. ISBN 94-010-3380-3, 94-010-3378-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-010-3378-
                                              7_12.

Mamchur:1994:SFH

[Mam94] Elena A. Mamchur. Sociocultural factors and historiography of science. In Trends in the Historiography of Science [GCN94], chapter 15, pages 193–201. ISBN 90-481-4264-4, 94-017-3596-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-017-3596-
                                              4_15.

Manier:1971:FNF

[Man71] Edward Manier. Functionalism and the negative feedback model in biology.

In Buck and Cohen [BC71], pages 225–240. ISBN 90-277-0309-4, 94-010-3142-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-010-
3142-4_16.

Mangione:1981:WCL

[Man81] Corrado Mangione. What is contemporary logic talking about? In Italian Studies in the Philosophy of Science [Chi81b], chapter 7, pages 89–111. ISBN 90-277-1073-2, 94-009-8937-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-009-8937-
5_7.

Manuel:1986:MSU

[Man86] Frank Manuel. Metamorphoses of the scientist in utopia. In Ullmann-Margalit [UM86c], chapter 1, pages 1–16. ISBN 90-277-2160-2, 90-277-2161-0 (paperback), 94-009-4566-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 95 Q175.

URL https://link.springer.com/
chapter/10.1007/978-94-009-4566-
1_1.

Manicas:1989:EQ

[Man89] Peter T. Manicas. Explanation and quantification. In The Qualitative-Quantitative Distinction in the Social Sciences [GM89b], chapter 10, pages 179–205. ISBN 90-481-8460-6, 94-017-3444-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 112. URL https:/

/link.springer.com/chapter/10.
1007/978-94-017-3444-8_10.

Manning:2012:DHM

[Man12] Gideon Manning. Descartes’ healthy machines and the human exception. In Philosophy of Behavioral Biology [PR12b], chapter 10, pages 237–262. ISBN 94-007-1950-7, 94-007-1951-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QH331 .P4679 2012; Q175 .B657 v.282.

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-007-4345-
                                              8_10.

Manning:2013:DHM

[Man13] Gideon Manning. Descartes’ healthy machines and the human exception. In Garber and Roux [GR13b], chapter 10, pages 237–262. CODEN ???? ISBN 94-007-4344-0 (hardcover), 94-007-4345-9 (e-book). ISSN ???? LCCN QC6 .M3616 2013; QC6 .M43 2013.

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-007-4345-
                                              8_10.

Manafu:2015:NAE

[Man15a] Alexandru Manafu. A novel approach to emergence in chemistry. In Philosophy of Chemistry: Growth of a New Discipline [SM15b], chapter 4, pages 39–55. ISBN 94-017-9363-8 (hardcover), 94-017-9364-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QD6 .P455 2015; QD6 .S384 2014. URL https://link.

                                              springer.com/chapter/10.1007/
                                              978-94-017-9364-3_4.

Manafu:2015:PFE

[Man15b] Alexandru Manafu. The prospects for fusion emergence. In Romanian Studies in Philosophy of Science [PST15], chapter 14, pages 221–235. ISBN 3-319-16654-9 (hardcover), 3-319-16655-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-3-319-16655-
1_14.

Marcus:1963:MLM

[Mar63] Ruth Barcan Marcus. Modal logics I:

Modalities and intensional languages. In Wartofsky [War63], pages 77–96. ISBN 94-010-3265-3, 94-010-3263-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174

.B678 1970. URL https://link.
springer.com/chapter/10.1007/
978-94-010-3263-6_7.

Martin:1967:CAP

[Mar67] Michael Martin. Comments: Analytic premises and existential conclusions. In Cohen and Wartofsky [CW67], pages 376–386. ISBN 90-277-0013-3, 94-010-3508-3. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175

.B731 v.3.  URL https://link.
springer.com/chapter/10.1007/
978-94-010-3508-8_21.

Markovic:1969:PT

[Mar69]  Mihailo Marković.  The problem

of truth. In Cohen and Wartofsky [CW69a], pages 341–362. ISBN 94-010-3383-8, 94-010-3263-7, 94-010-3381-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-010-3381-
7_10.

Markovic:1974:SDT

[Mar74a] Mihailo Marković. Self-determination in theory and practice. In For Dirk Struik: Scientific, Historical and Political Essays in Honor of Dirk J. Struik [CSW74], chapter 41, pages 547–559. ISBN 90-277-0379-5, 94-010-2115-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-010-2115-
9_41.

Marlow:1974:INA

[Mar74b] A. R. Marlow. Implications of a new axiom set for quantum logic. In Logical and Epistemological Studies in Contemporary Physics [CW74a], chapter 12, pages 350–360. ISBN 90-277-0391-4 (hardcover), 90-277-0377-9 (paperback), 94-010-2656-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol.

13; QC5.56. URL https://link.
springer.com/chapter/10.1007/
978-94-010-2656-7_12.

Marchi:1976:MCE

[Mar76a] Peggy Marchi. Mathematics as a critical enterprise. In Essays in Memory of Imre Lakatos [CFW76], chapter 24, pages 379–393. ISBN 90-277-0655-7, 94-010-1451-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-010-
1451-9_24.

Margalit:1976:TCP

[Mar76b] Avishai Margalit. Talking with children, Piaget style. In Language in Focus: Foundations, Methods and Systems: Essays in Memory of Yehoshua Bar-Hillel [Kas76b], chapter 19, pages 457–471. ISBN 90-277-0644-1, 90-277-0645-X, 94-010-1876-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN P101-120. URL https:

//link.springer.com/chapter/10.
1007/978-94-010-1876-0_19.

Martin:1976:RPS

[Mar76c] Michael Martin. The relevance of philosophy of science for science education. In Cohen et al. [CHMV76], pages 293–300. ISBN 90-277-0647-6, 90-277-0648-4 (paperback), 94-010-1449-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-010-1449-
6_13.

Martin:1976:HSR

[Mar76d] R. M. Martin. On Harris’s systems of report and paraphrase. In Language in Focus: Foundations, Methods and Systems: Essays in Memory of Yehoshua Bar-Hillel [Kas76b], chapter 23, pages 541–568. ISBN 90-277-0644-1, 90-277-0645-X, 94-010-1876-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN P101-120. URL https:

//link.springer.com/chapter/10.
1007/978-94-010-1876-0_23.

Margolis:1978:AI

[Mar78a] Joseph Margolis. Action and ideology.

In Persons and Minds: The Prospects of Nonreductive Materialism [Mar78i], chapter 14, pages 243–262. ISBN 90-277-0863-0, 94-009-9801-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-009-9801-8_14.

Margolis:1978:GI

[Mar78b] Joseph Margolis. General introduction. In Persons and Minds: The Prospects of Nonreductive Materialism [Mar78i], chapter 1, pages 3–10. ISBN 90-277-0863-0, 94-009-9801-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-009-9801-8_1.

Margolis:1978:IT

[Mar78c] Joseph Margolis. The identity theory. In Persons and Minds: The Prospects of Nonreductive Materialism [Mar78i], chapter 4, pages 34–44. ISBN 90-277-0863-0, 94-009-9801-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-009-9801-8_4.

Margolis:1978:LAR

[Mar78d] Joseph Margolis. Language acquisition. I: Rationalists vs. empiricists. In Persons and Minds: The Prospects of Nonreductive Materialism [Mar78i], chapter 8, pages 97–117. ISBN 90-277-0863-0, 94-009-9801-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-009-9801-8_8.

Margolis:1978:LAI

[Mar78e] Joseph Margolis. Language acquisition. II: First and second languages and the theory of thought and perception. In Persons and Minds: The Prospects of Nonreductive Materialism [Mar78i], chapter 9, pages 118–145. ISBN 90-277-0863-0, 94-009-9801-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-009-9801-8_9.

Margolis:1978:MI

[Mar78f] Joseph Margolis. Materialism without identity. In Persons and Minds: The Prospects of Nonreductive Materialism [Mar78i], chapter 6, pages 60–77. ISBN 90-277-0863-0, 94-009-9801-5 (e-book).

ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-009-9801-8_6.

Margolis:1978:MSS

[Mar78g] Joseph Margolis. Mental states and sentience. In Persons and Minds: The Prospects of Nonreductive Materialism [Mar78i], chapter 11, pages 171–196. ISBN 90-277-0863-0, 94-009-9801-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-009-9801-
8_11.

Margolis:1978:NIC

[Mar78h] Joseph Margolis. The nature and identity of cultural entities. In Persons and Minds: The Prospects of Nonreductive Materialism [Mar78i], chapter 13, pages 225–242. ISBN 90-277-0863-0, 94-009-9801-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-009-9801-8_13.

Margolis:1978:PMP

[Mar78i] Joseph Margolis. Persons and Minds:

The Prospects of Nonreductive Materialism, volume 57 of Boston Studies in the Philosophy of Science. D. Reidel, Dordrecht, The Netherlands; Boston, MA, USA; Lancaster, UK; Tokyo, Japan, 1978. ISBN 90-277-0863-0, 94-009-9801-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). xiv + 310 pp. LCCN ???? URL

https://link.springer.com/book/
10.1007/978-94-009-9801-8.

Margolis:1978:PRP

[Mar78j] Joseph Margolis. Problems regarding persons. In Persons and Minds: The Prospects of Nonreductive Materialism [Mar78i], chapter 7, pages 81–96. ISBN 90-277-0863-0, 94-009-9801-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-009-9801-8_7.

Margolis:1978:PCB

[Mar78k] Joseph Margolis. Propositional content and the beliefs of animals. In Persons and Minds: The Prospects of Nonreductive Materialism [Mar78i], chapter 10, pages 146–170. ISBN 90-277-0863-0, 94-009-9801-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-009-9801-8_10.

Margolis:1978:PI

[Mar78l] Joseph Margolis. Psychophysical interaction. In Persons and Minds: The Prospects of Nonreductive Materialism [Mar78i], chapter 12, pages 199–224. ISBN 90-277-0863-0, 94-009-9801-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-009-9801-
                                              8_12.

Margolis:1978:RM

[Mar78m] Joseph Margolis. Radical materialism. In Persons and Minds: The Prospects of Nonreductive Materialism [Mar78i], chapter 5, pages 45–59. ISBN 90-277-0863-0, 94-009-9801-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-009-9801-8_5.

Margolis:1978:RMB

[Mar78n] Joseph Margolis. The relation of mind and body. In Persons and Minds: The Prospects of Nonreductive Materialism [Mar78i], chapter 3, pages 28–33. ISBN 90-277-0863-0, 94-009-9801-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-009-9801-8_3.

Margolis:1978:TPS

[Mar78o] Joseph Margolis. The theory of persons sketched. In Persons and Minds: The Prospects of Nonreductive Materialism [Mar78i], chapter 2, pages 11–25. ISBN 90-277-0863-0, 94-009-9801-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-009-9801-8_2.

Markovic:1979:DT

[Mar79a] Mihailo Marković. Dialectic today.

In Praxis: Yugoslav Essays in the Philosophy and Methodology of the Social Sciences [MP79], chapter 1, pages 3–43. ISBN 90-277-0968-8, 94-009-9355-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-009-
9355-6_1.

Martin:1979:X

[Mar79b] R. M. Martin. ‘And’. In Pragmatics, Truth, and Language [Mar79o], chapter 17, pages 224–236. ISBN 90-277-0993-9, 94-009-9457-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-009-9457-7_17.

Martin:1979:EAB

[Mar79c] R. M. Martin. Events and actions:

Brand and kim. In Pragmatics, Truth, and Language [Mar79o], chapter 11, pages 144–159. ISBN 90-277-0993-9, 94-009-9457-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

                                              ???? URL https://link.springer.
                                              com/chapter/10.1007/978-94-009-
                                              9457-7_11.

Martin:1979:ELF

[Mar79d] R. M. Martin. Existence and logical form. In Pragmatics, Truth, and Language [Mar79o], chapter 8, pages 95–109. ISBN 90-277-0993-9, 94-009-9457-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

                                              ???? URL https://link.springer.
                                              com/chapter/10.1007/978-94-009-
                                              9457-7_8.

Martin:1979:MN

[Mar79e] R. M. Martin. Mathematical nominalism. In Pragmatics, Truth, and Language [Mar79o], chapter 20, pages 271–282. ISBN 90-277-0993-9, 94-009-9457-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

                                              ???? URL https://link.springer.
                                              com/chapter/10.1007/978-94-009-
                                              9457-7_20.

Martin:1979:LLL

[Mar79f] R. M. Martin. Of logic, learning, and language. In Pragmatics, Truth, and Language [Mar79o], chapter 21, pages 283–295. ISBN 90-277-0993-9, 94-009-9457-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

                                              ???? URL https://link.springer.
                                              com/chapter/10.1007/978-94-009-
                                              9457-7_21.

Margin:1979:X

[Mar79g] R. M. Martin. Of ‘of’. In Pragmatics, Truth, and Language [Mar79o], chapter 10, pages 130–143. ISBN 90-277-0993-9, 94-009-9457-5 (e-book).

ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-009-9457-7_10.

Martin:1979:TNI

[Mar79h] R. M. Martin. Of time and the null individual. In Pragmatics, Truth, and Language [Mar79o], chapter 7, pages 82–94. ISBN 90-277-0993-9, 94-009-9457-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-009-
9457-7_7.

Martin:1979:DI

[Mar79i] R. M. Martin. On disquotation and intensionality. In Pragmatics, Truth, and Language [Mar79o], chapter 3, pages 30–41. ISBN 90-277-0993-9, 94-009-9457-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-009-
9457-7_3.

Martin:1979:PER

[Mar79j] R. M. Martin. On possibilia and essentiality: Ruth Marcus. In Pragmatics, Truth, and Language [Mar79o], chapter 14, pages 181–191. ISBN 90-277-0993-9, 94-009-9457-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-009-9457-7_14.

Martin:1979:SD

[Mar79k] R. M. Martin. On suppositio and denotation. In Pragmatics, Truth, and Language [Mar79o], chapter 6, pages 72–81. ISBN 90-277-0993-9, 94-009-9457-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

                                              ???? URL https://link.springer.
                                              com/chapter/10.1007/978-94-009-
                                              9457-7_6.

Martin:1979:LCT

[Mar79l] R. M. Martin. On the language of causal talk: Scriven and Suppes. In Pragmatics, Truth, and Language [Mar79o], chapter 15, pages 192–205. ISBN 90-277-0993-9, 94-009-9457-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-009-9457-
                                              7_15.

Martin:1979:TBM

[Mar79m] R. M. Martin. On truth, belief, and modes of description. In Pragmatics, Truth, and Language [Mar79o], chapter 4, pages 42–54. ISBN 90-277-0993-9, 94-009-9457-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-009-9457-7_4.

Martin:1979:PSR

[Mar79n] R. M. Martin. The pragmatics of self-reference. In Pragmatics, Truth, and Language [Mar79o], chapter 5, pages 55–71. ISBN 90-277-0993-9, 94-009-9457-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

                                              ???? URL https://link.springer.
                                              com/chapter/10.1007/978-94-009-
                                              9457-7_5.

Martin:1979:PTL

[Mar79o] R. M. Martin. Pragmatics, Truth, and Language, volume 38 of Boston Studies in the Philosophy of Science. D. Reidel, Dordrecht, The Netherlands; Boston, MA, USA; Lancaster, UK; Tokyo, Japan, 1979. ISBN 90-277-0993-9, 94-009-9457-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). 325 (est.) pp. LCCN ???? URL

https://link.springer.com/book/
10.1007/978-94-009-9457-7.

Martin:1979:RFS

[Mar79p] R. M. Martin. A reading of Frege on sense and designation. In Pragmatics, Truth, and Language [Mar79o], chapter 16, pages 206–223. ISBN 90-277-0993-9, 94-009-9457-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-009-9457-7_16.

Martin:1979:SHZ

[Mar79q] R. M. Martin. Some Hiżian heresies.

In Pragmatics, Truth, and Language [Mar79o], chapter 19, pages 257–270. ISBN 90-277-0993-9, 94-009-9457-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-009-9457-
7_19.

Martin:1979:SPT

[Mar79r] R. M. Martin. Some protolinguistic transformations. In Pragmatics, Truth, and Language [Mar79o], chapter 18, pages 237–256. ISBN 90-277-0993-9, 94-009-9457-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-009-
9457-7_18.

Martin:1979:SRC

[Mar79s] R. M. Martin. Some reminders concerning truth, satisfaction, and reference. In Pragmatics, Truth, and Language [Mar79o], chapter 2, pages 17–29. ISBN 90-277-0993-9, 94-009-9457-5 (e-book).

ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-009-9457-7_2.

Martin:1979:TAM

[Mar79t] R. M. Martin. Tense, aspect, and modality. In Pragmatics, Truth, and Language [Mar79o], chapter 9, pages 110–129. ISBN 90-277-0993-9, 94-009-9457-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

                                              ???? URL https://link.springer.
                                              com/chapter/10.1007/978-94-009-
                                              9457-7_9.

Martin:1979:TAK

[Mar79u] R. M. Martin. The truth about Kripke’s “truth”. In Pragmatics, Truth, and Language [Mar79o], chapter 13, pages 173–180. ISBN 90-277-0993-9, 94-009-9457-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

                                              ???? URL https://link.springer.
                                              com/chapter/10.1007/978-94-009-
                                              9457-7_13.

Martin:1979:TIS

[Mar79v] R. M. Martin. Truth and its illicit surrogates. In Pragmatics, Truth, and Language [Mar79o], chapter 1, pages 1– 16. ISBN 90-277-0993-9, 94-009-9457-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-009-9457-
                                              7_1.

Martin:1979:WMG

[Mar79w] R. M. Martin. Why I am not a Montague grammarian. In Pragmatics, Truth, and Language [Mar79o], chapter 12, pages 160–172. ISBN 90-277-0993-9, 94-009-9457-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-009-9457-7_12.

Marchi:1980:MAM

[Mar80] Peggy Marchi. The method of analysis in mathematics. In Scientific Discovery, Logic, and Rationality [Nic80f], chapter 5, pages 159–172. ISBN 90-277-1070-8, 94-009-8986-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-009-8986-3_5.

Margenau:1982:PDR

[Mar82] Henry Margenau. Physics and the doctrine of reductionism. In Scientific Philosophy Today: Essays in Honor of Mario Bunge [AC82], chapter 11, pages 187–199. ISBN 90-277-1262-X, 90-277-1263-8, 94-009-8462-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 67.

URL https://link.springer.com/
chapter/10.1007/978-94-009-8462-
2_11.

Margalit:1983:KPK

[Mar83a] Avishai Margalit. Knowledge in pursuit of knowledge — a few worries. In Language, Logic and Method [CW83b], chapter 7, pages 179–183. ISBN 94-009-7704-2, 94-009-7702-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-009-7702-0_7.

Margolis:1983:PTA

[Mar83b] Joseph Margolis. Pragmatism, transcendental arguments, and the technological. In Philosophy and Technology [DR83], chapter 20, pages 291–309. ISBN 94-009-7126-5, 94-009-7124-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-009-7124-
                                              0_20.

Markovic:1984:COR

[Mar84a] Mihailo Marković. Categories of objective reality. In Dialectical Theory of Meaning [Mar84d], chapter 3, pages 43–90. ISBN 94-009-6258-4, 94-009-6256-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

                                              ???? URL https://link.springer.
                                              com/chapter/10.1007/978-94-009-
                                              6256-9_3.

Markovic:1984:COC

[Mar84b] Mihailo Marković. Concepts and other categories of thought. In Dialectical Theory of Meaning [Mar84d], chapter 6, pages 131–167. ISBN 94-009-6258-4, 94-009-6256-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

                                              ???? URL https://link.springer.
                                              com/chapter/10.1007/978-94-009-
                                              6256-9_6.

Markovic:1984:CEC

[Mar84c] Mihailo Marković. Conditions of effective communication. In Dialectical Theory of Meaning [Mar84d], chapter 14, pages 372–395. ISBN 94-009-6258-4, 94-009-6256-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-009-6256-9_14.

Markovic:1984:DTM

[Mar84d] Mihailo Marković. Dialectical Theory of Meaning, volume 81 of Boston Studies in the Philosophy of Science. D. Reidel, Dordrecht, The Netherlands; Boston, MA, USA; Lancaster, UK; Tokyo, Japan, 1984. ISBN 94-009-6258-4, 94-009-6256-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). xxviii + 404 pp. LCCN ???? URL

https://link.springer.com/book/
10.1007/978-94-009-6256-9.

Markovic:1984:GDM

[Mar84e] Mihailo Marković. General definition of meaning. In Dialectical Theory of Meaning [Mar84d], chapter 13, pages 363–371. ISBN 94-009-6258-4, 94-009-6256-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-009-
6256-9_13.

Markovic:1984:GLP

[Mar84f] Mihailo Marković. General logical problems of constructing a theory of meaning. In Dialectical Theory of Meaning [Mar84d], chapter 2, pages 31–42. ISBN 94-009-6258-4, 94-009-6256-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-009-
6256-9_2.

Markovic:1984:GSM

[Mar84g] Mihailo Marković. The genesis of signs and meaning. In Dialectical Theory of Meaning [Mar84d], chapter 12, pages 331–362. ISBN 94-009-6258-4, 94-009-6256-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-009-
6256-9_12.

Markovic:1984:I

[Mar84h] Mihailo Marković. Introduction. In Dialectical Theory of Meaning [Mar84d], chapter 1, pages 1–27. ISBN 94-009-6258-4, 94-009-6256-8 (e-book).

ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-009-6256-9_1.

Markovic:1984:LM

[Mar84i] Mihailo Marković. Linguistic meaning. In Dialectical Theory of Meaning [Mar84d], chapter 10, pages 261–318. ISBN 94-009-6258-4, 94-009-6256-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-009-6256-
                                              9_10.

Markovic:1984:MCR

[Mar84j] Mihailo Marković. Meaning as a complex of relationships. In Dialectical Theory of Meaning [Mar84d], chapter 7, pages 171–178. ISBN 94-009-6258-4, 94-009-6256-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

                                              ???? URL https://link.springer.
                                              com/chapter/10.1007/978-94-009-
                                              6256-9_7.

Markovic:1984:MM

[Mar84k] Mihailo Marković. Mental meaning. In Dialectical Theory of Meaning [Mar84d], chapter 8, pages 179–187. ISBN 94-009-6258-4, 94-009-6256-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-009-6256-
                                              9_8.

Markovic:1984:OE

[Mar84l] Mihailo Marković. Objective experience. In Dialectical Theory of Meaning [Mar84d], chapter 5, pages 108–130. ISBN 94-009-6258-4, 94-009-6256-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-009-6256-
9_5.

Markovic:1984:OM

[Mar84m] Mihailo Marković. Objective meaning. In Dialectical Theory of Meaning [Mar84d], chapter 9, pages 188–260. ISBN 94-009-6258-4, 94-009-6256-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-009-6256-
9_9.

Markovic:1984:PM

[Mar84n] Mihailo Marković. Practical meaning. In Dialectical Theory of Meaning [Mar84d], chapter 11, pages 319–328. ISBN 94-009-6258-4, 94-009-6256-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-009-6256-
9_11.

Markovic:1984:S

[Mar84o] Mihailo Marković. Symbols. In Dialectical Theory of Meaning [Mar84d], chapter 4, pages 91–107. ISBN 94-009-6258-4, 94-009-6256-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-009-6256-9_4.

Martin:1984:HGP

[Mar84p] Michael Martin. How to be a good philosopher of science: A plea for empiricism in matters methodological. In Methodology, Metaphysics and the History of Science: In Memory of Benjamin Nelson [CW84b], chapter 2, pages 33–42. ISBN 94-009-6333-5, 94-009-6331-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-009-
6331-3_2.

Marcus:1986:SOP

[Mar86a] Ruth Barcan Marcus. Spinoza and the ontological proof. In Cohen and Wartofsky [CW85], chapter 8, pages 153–166. ISBN 90-277-1971-3, 94-009-5345-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-009-
5349-9_8.

Marcuse:1986:SP

[Mar86b] Herbert Marcuse. On science and phenomenology. In Cohen and Wartofsky [CW85], chapter 2, pages 19–30. ISBN 90-277-1971-3, 94-009-5345-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-009-5345-1_2.

Margalit:1986:PIC

[Mar86c] Avishai Margalit. The past of an illusion: A comment. In Ullmann-Margalit [UM86a], chapter 8, pages 89–94. ISBN 90-277-2159-9 (paperback), 90-277-2158-0 (hardcover), 94-009-5496-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 94 Q175. URL https:/

/link.springer.com/chapter/10.
1007/978-94-009-5496-0_8.

Margolis:1986:IAI

[Mar86d] Joseph Margolis. Information, artificial intelligence, and the praxical. In Philosophy and Technology II: Information Technology and Computers in Theory and Practice [MH86], chapter 12, pages 171–186. ISBN 94-010-8510-2, 94-009-4512-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-009-
4512-8_12.

Markiewicz:1986:LCI

[Mar86e] WŁadysŁaw Markiewicz. Lwów as cultural and intellectual background of the genesis of Fleck’s ideas. In Cognition and Fact: Materials on Ludwik Fleck [CS86], chapter 12, pages 223–229. ISBN 94-010-8504-8, 94-009-4498-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-009-4498-
5_12.

Markus:1986:LPC

[Mar86f] Gyorgy Markus. Language and Production: A Critique of the Paradigms, volume 96 of Boston Studies in the Philosophy of Science. D. Reidel, Dordrecht, The Netherlands; Boston, MA, USA; Lancaster, UK; Tokyo, Japan, 1986. ISBN 94-010-8532-3, 94-009-4574-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). xv + 190 pp. LCCN ???? URL

https://link.springer.com/book/
10.1007/978-94-009-4574-6.

Markus:1986:PLP

[Mar86g] Gyorgy Markus. On the paradigm of language: Positivism and hermeneutics as theories of objectivation. In Language and Production: A Critique of the Paradigms [Mar86f], chapter 1, pages 1–40. ISBN 94-010-8532-3, 94-009-4574-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-009-
4574-6_1.

Markus:1986:PPM

[Mar86h] Gyorgy Markus. On the paradigm of production: Marxian materialism and the problem of the constitution of the social world. In Language and Production: A Critique of the Paradigms [Mar86f], chapter 2, pages 41–125. ISBN 94-010-8532-3, 94-009-4574-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-009-4574-
                                              6_2.

Margolis:1987:GP

[Mar87] Joseph Margolis. Goethe and psychoanalysis. In Amrine et al. [AZW87], chapter 5, pages 83–100. ISBN 90-277-2265-X, 94-009-3761-X. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 97.

                                              US$79.00.  URL https://link.
                                              springer.com/chapter/10.1007/
                                              978-94-009-3761-1_5.

Margolis:1989:MDC

[Mar89] Joseph Margolis. Monistic and dualistic canons for the natural and human sciences. In The Qualitative-Quantitative Distinction in the Social Sciences [GM89b], chapter 9, pages 155–178. ISBN 90-481-8460-6, 94-017-3444-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 112. URL https://link.

                                              springer.com/chapter/10.1007/
                                              978-94-017-3444-8_9.

Marubini:1990:SLT

[Mar90] Ettore Marubini. Short and long term survival analysis in oncological research. In Statistics in Science: The Foundations of Statistical Methods in Biology, Physics and Economics [CC90], chapter 5, pages 73–87. ISBN 94-010-6765-1, 94-009-0619-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QA276-280.

URL https://link.springer.com/
chapter/10.1007/978-94-009-0619-
8_5.

Margolis:1991:SMF

[Mar91] Joseph Margolis. Scientific methods and Feyerabend’s advocacy of anarchism. In Beyond Reason: Essays on the Philosophy of Paul Feyerabend [Mun91a], chapter 24, pages 465–486. ISBN 94-010-5406-1, 94-011-3188-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-011-3188-
9_24.

Margalit:1992:HOR

[Mar92] Avishai Margalit. How to outsmart the rules: A comment. In The Scientific Enterprise: The Bar-Hillel Colloquium: Studies in History, Philosophy, and Sociology of Science, Volume 4 [UM92], chapter 8, pages 139–144. ISBN 94-010-5190-9, 94-011-2688-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .S354 1992.

URL https://link.springer.com/
chapter/10.1007/978-94-011-2688-
5_8.

Margolis:1994:DDP

[Mar94a] Joseph Margolis. Donald Davidson’s philosophical strategies. In Artifacts, Representations and Social Practice: Essays for Marx Wartofsky [GC94], chapter 17, pages 291–322. ISBN 94-010-4390-6, 94-011-0902-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN B73 .A77 1994.

URL https://link.springer.com/
chapter/10.1007/978-94-011-0902-
4_17.

Marino:1994:SKO

[Mar94b] Luigi Marino. Soemmerring, Kant and the organ of the soul. In Poggi and Bossi [PB94], chapter 8, pages 127–142. ISBN 0-7923-2336-X, 94-017-2921-2. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 152.

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-017-2921-
                                              5_8.

Markus:1994:EMB

[Mar94c] Gyorgy Markus. The end of a metaphor: The base and the superstructure. In Artifacts, Representations and Social Practice: Essays for Marx Wartofsky [GC94], chapter 25, pages 419–439. ISBN 94-010-4390-6, 94-011-0902-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN B73 .A77

                                              1994. URL https://link.springer.
                                              com/chapter/10.1007/978-94-011-
                                              0902-4_25.

Margolis:1995:MTK

[Mar95a] Joseph Margolis. The meaning of Thomas Kuhn’s “different worlds”. In Science, Mind and Art: Essays on science and the humanistic understanding in art, epistemology, religion and ethics In honor of Robert S. Cohen [GSW95b], chapter 9, pages 123–148. ISBN 0-7923-2990-2 (hardcover), 94-011-0469-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 165 BD418.3. URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-011-0469-2_9.

Marion:1995:KSH

[Mar95b] Mathieu Marion. Kronecker’s ‘safe haven of real mathematics’. In Québec Studies in the Philosophy of Science: Part I: Logic, Mathematics, Physics and History of Science Essays in Honor of Hugues Leblanc [MC95], chapter 11, pages 189–215. ISBN 94-010-7204-3, 94-009-1575-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-009-
1575-6_11.

Marquis:1995:ITB

[Mar95c] Jean-Pierre Marquis. If not-true and not being true are not identical, which one is false? In Québec Studies in the Philosophy of Science: Part I: Logic, Mathematics, Physics and History of Science Essays in Honor of Hugues Leblanc [MC95], chapter 6, pages 79– 94. ISBN 94-010-7204-3, 94-009-1575-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-009-1575-
6_6.

MartinezE:1995:AMC

[Mar95d] Rafael Martinez E. Among men, closest to God. In Mexican Studies in the History and Philosophy of Science [RC95], chapter 9, pages 129–149. ISBN 94-010-6535-7 (hardcover), 94-009-0109-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-009-
0109-4_9.

Marek:1997:HFP

[Mar97a] Johann Christian Marek. Haller on the first person. In Austrian Philosophy Past and Present: Essays in Honor of Rudolf Haller [LM97], chapter 6, pages 71–85. ISBN 94-010-6412-1, 94-011-5720-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-011-
5720-9_6.

Margolis:1997:OFL

[Mar97b] Joseph Margolis. Objectivity: False leads from T. S. Kuhn on the role of the aesthetic in the sciences. In The Elusive Synthesis: Aesthetics and Science [Tau97a], chapter 9, pages 189–202. ISBN 0-7923-4763-3, 0-7923-3904-5, 94-009-1786-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

                                              ???? URL https://link.springer.
                                              com/chapter/10.1007/978-94-009-
                                              1786-6_9.

Martinson:2001:FRP

[Mar01] Helle Martinson. Formation of R&D policy in a small country in a changing world. In Estonian Studies in the History and Philosophy of Science [Vih01b], chapter 5, pages 63–76. ISBN 94-010-3863-5, 94-010-0672-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-010-0672-9_5.

Margolis:2002:PHI

[Mar02] Joseph Margolis. Patrick Heelan’s interpretation of van Gogh’s “Bedroom at Arles”. In Hermeneutic Philosophy of Science, Van Gogh’s Eyes, and God: Essays in Honor of Patrick A. Heelan, S.J. [Bab02c], chapter 19, pages 233–238. ISBN 90-481-5926-1, 94-017-1767-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 225. URL https:/

                                              /link.springer.com/chapter/10.
                                              1007/978-94-017-1767-0_19.

Mariconda:2011:GMS

[Mar11a] Pablo Rubén Mariconda. Galileo and modern science. In Brazilian Studies in Philosophy and History of Science: An account of recent works [KV11], chapter 2, pages 57–69. ISBN 90-481-9421-0 (hardcover), 90-481-9422-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175.3 .B73 2011.

URL https://link.springer.com/
chapter/10.1007/978-90-481-9422-
3_2.

Martins:2011:RDT

[Mar11b] Lilian Al-Chueyr Pereira Martins. Regeneration as a difficulty for the theory of natural selection: Morgan’s changing attitudes, 1897–1932. In Brazilian Studies in Philosophy and History of Science: An account of recent works [KV11], chapter 7, pages 119–129. ISBN 90-481-9421-0 (hardcover), 90-481-9422-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175.3 .B73 2011. URL https:/

/link.springer.com/chapter/10.
1007/978-90-481-9422-3_7.

Martin:2012:CDM

[Mar12] Craig Martin. Causation in Descartes’ Les Météores and Late Renaissance Aristotelian meteorology. In Philosophy of Behavioral Biology [PR12b], chapter 9, pages 217–236. ISBN 94-007-1950-7, 94-007-1951-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QH331 .P4679 2012; Q175 .B657 v.282. URL https:

//link.springer.com/chapter/10.
1007/978-94-007-4345-8_9.

Martin:2013:CDM

[Mar13] Craig Martin. Causation in Descartes’ Les Météores and Late Renaissance Aristotelian meteorology. In Garber and Roux [GR13b], chapter 9, pages 217–236. CODEN ???? ISBN 94-007-4344-0 (hardcover), 94-007-4345-9 (e-book). ISSN ???? LCCN QC6 .M3616 2013; QC6 .M43 2013.

URL https://link.springer.com/

                                              chapter/10.1007/978-94-007-4345-
                                              8_9.

Marcum:2015:ENR

[Mar15] James A. Marcum. The evolving notion and role of Kuhn’s incommensurability thesis. In Kuhn’s Structure of Scientific Revolutions — 50 Years On [DB15b], chapter 9, pages 115–134. ISBN 3-319-13382-9, 3-319-13383-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175.3 .K84 2015.

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-3-319-13383-
                                              6_9.

Massey:1971:CPP

[Mas71] Gerald J. Massey. Is ‘congruence’ a peculiar predicate? In Buck and Cohen [BC71], pages 606–615. ISBN 90-277-0309-4, 94-010-3142-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-010-3142-4_41.

Massimi:2011:SRN

[Mas11] Michela Massimi. Structural realism: A neo-Kantian perspective. In Bokulich and Bokulich [BB11], chapter 1, pages 1–23. ISBN 90-481-9596-9, 90-481-9597-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175.32.S78 S35 2011. URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-90-481-9597-8_1.

Maschio:2015:DBS

[Mas15a] Samuele Maschio. On the distinction between sets and classes: A categorical perspective. In From Logic to Practice: Italian Studies in the Philosophy of Mathematics [LPV15], chapter 10, pages 185–199. ISBN 3-319-10433-0, 3-319-10434-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QA8.4 .F76 2015; Q174 .B67 v.308.

URL https://link.springer.com/
chapter/10.1007/978-3-319-10434-
8_10.

Massimi:2015:WLK

[Mas15b] Michela Massimi. Walking the line:

Kuhn between realism and relativism. In Kuhn’s Structure of Scientific Revolutions — 50 Years On [DB15b], chapter 10, pages 135–152. ISBN 3-319-13382-9, 3-319-13383-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175.3 .K84 2015.

URL https://link.springer.com/
chapter/10.1007/978-3-319-13383-
6_10.

Matheron:1986:SEA

[Mat86] Alexandre Matheron. Spinoza and Euclidean arithmetic: The example of the Fourth Proportional. In Spinoza and the Sciences [GN86], chapter 5, pages 125–150. ISBN 94-010-8511-0, 94-009-4514-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-009-
4514-2_5.

Mattick:1989:EAI

[Mat89] Paul Mattick, Jr. Extending the analysis: The informational environment of the science sentences. In The Form of Information in Science: Analysis of an Immunology Sublanguage [HGR^{+}89], chapter 6, pages 151–169. ISBN 94-010-7777-0, 94-009-2837-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN B832.A-832.Z.

URL https://link.springer.com/
chapter/10.1007/978-94-009-2837-
4_6.

Matthews:1995:MSE

[Mat95] Michael R. Matthews. Multicultural science education: The contribution of history and philosophy of science. In Science, Mind and Art: Essays on science and the humanistic understanding in art, epistemology, religion and ethics In honor of Robert S. Cohen [GSW95b], chapter 10, pages 149–168. ISBN 0-7923-2990-2 (hardcover), 94-011-0469-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 165 BD418.3. URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-011-0469-2_10.

Matjus:2001:EHP

[Mat01] Ülo Matjus. Edmund Husserl pursuing the paths of Descartes. In Estonian Studies in the History and Philosophy of Science [Vih01b], chapter 9, pages 123–138. ISBN 94-010-3863-5, 94-010-0672-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

                                              ???? URL https://link.springer.
                                              com/chapter/10.1007/978-94-010-
                                              0672-9_9.

Matthews:2017:FSO

[Mat17] Michael R. Matthews. Feng shui: An overlooked topic for critical rationalists. In Encouraging Openness: Essays for Joseph Agassi on the Occasion of His 90th Birthday [BAG17], chapter 6, pages 53–67. ISBN 3-319-57668-2 (hardcover), 3-319-57669-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN B833.E53 2017.

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-3-319-57669-
                                              5_6.

Maull:1980:CS

[Mau80] Nancy L. Maull. Comment on Schaffner. In Scientific Discovery: Case Studies [Nic80e], chapter 11, pages 207–209. ISBN 90-277-1093-7, 94-009-9015-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 60. URL https://link.

springer.com/chapter/10.1007/
978-94-009-9015-9_11. See [Sch80a]

and reply [Sch80b].

Maurer:1983:OMT

[Mau83] Reinhart Maurer. The origins of modern technology in millenarianism. In Philosophy and Technology [DR83], chapter 17, pages 253–265. ISBN 94-009-7126-5, 94-009-7124-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-009-7124-0_17.

Maull:1986:SCS

[Mau86] Nancy Maull. Spinoza in the century of science. In Spinoza and the Sciences [GN86], chapter 1, pages 3–13. ISBN 94-010-8511-0, 94-009-4514-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-009-4514-2_1.

Maudlin:1996:STQ

[Mau96] Tim Maudlin. Space–time in the quantum world. In Bohmian Mechanics and Quantum Theory: an Appraisal [CFG96], chapter 20, pages 285–307. ISBN 90-481-4698-4, 94-015-8715-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC173.96-174.52. URL https://link.

springer.com/chapter/10.1007/
978-94-015-8715-0_20.

Maxwell:1976:SCT

[Max76a] Grover Maxwell. Some current trends in philosophy of science: With special attention to confirmation, theoretical entities, and mind–body. In Cohen et al. [CHMV76], pages 565–584. ISBN 90-277-0647-6, 90-277-0648-4 (paperback), 94-010-1449-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-010-1449-6_32.

Maxwell:1976:ATO

[Max76b] J. C. Maxwell. Absolute time and the order of nature. In The Concepts of Space and Time: Their Structure and Their Development [Čap76], chapter 40, pages 231–232. ISBN 90-277-0375-2, 94-010-1727-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic).

LCCN BD632. URL https://link.
springer.com/chapter/10.1007/
978-94-010-1727-5_40.

Maxwell:1976:AS

[Max76c] J. C. Maxwell. On absolute space. In The Concepts of Space and Time: Their Structure and Their Development [Čap76], chapter 21, pages 121–123. ISBN 90-277-0375-2, 94-010-1727-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN BD632.

URL https://link.springer.com/
chapter/10.1007/978-94-010-1727-
5_21.

Maxwell:1991:FM

[Max91] Grover Maxwell. Feyerabend’s materialism. In Beyond Reason: Essays on the Philosophy of Paul Feyerabend [Mun91a], chapter 23, pages 453–463. ISBN 94-010-5406-1, 94-011-3188-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-011-3188-
9_23.

Max:1997:DT

[Max97] Ingolf Max. Dimensions of time.

In Perspectives on Time [FSU97b], chapter 16, pages 367–397. ISBN 90-481-4774-3, 94-015-8875-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol.189.

URL https://link.springer.com/
chapter/10.1007/978-94-015-8875-
1_16.

Mayr:1969:CTH

[May69] Ernst Mayr. Comments on ‘theories and hypotheses in biology’. In Cohen and Wartofsky [CW69a], pages 452–456. ISBN 94-010-3383-8, 94-010-3263-7, 94-010-3381-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-010-
3381-7_15.

Mayr:1974:TTN

[May74] Ernst Mayr. Teleological and teleonomic, a new analysis. In Methodological and Historical Essays in the Natural and Social Sciences [CW74b], chapter 6, pages 91–117. ISBN 90-277-0378-7, 94-010-2128-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 14.

URL https://link.springer.com/
chapter/10.1007/978-94-010-2128-
9_6.

Mayr:1976:SCD

[May76] Ernst Mayr. Species concepts and definitions. In Topics in the Philosophy of Biology [GM76], chapter 16, pages 353–371. ISBN 90-277-0596-8, 94-010-1829-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-010-
1829-6_16.

Mayr:1986:TTN

[May86] Ernst Mayr. Teleological and teleonomic, a new analysis. In Cohen and Wartofsky [CW85], chapter 10, pages 133–159. ISBN 90-277-1971-3, 94-009-5345-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

                                              ???? URL https://link.springer.
                                              com/chapter/10.1007/978-94-009-
                                              5345-1_10.

Mayer:1995:MHP

[May95] Arno J. Mayer. Memory and history: On the poverty of remembering and forgetting the Judeocide. In Science, Mind and Art: Essays on science and the humanistic understanding in art, epistemology, religion and ethics In honor of Robert S. Cohen [GSW95b], chapter 23, pages 385–400. ISBN 0-7923-2990-2 (hardcover), 94-011-0469-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 165 BD418.3. URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-011-0469-2_23.

Mays:2002:PHT

[May02] Wolfe Mays. Piaget and Husserl: On theory and praxis in science. In Hermeneutic Philosophy of Science, Van Gogh’s Eyes, and God: Essays in Honor of Patrick A. Heelan, S.J. [Bab02c], chapter 14, pages 177–186. ISBN 90-481-5926-1, 94-017-1767-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 225. URL https://link.

                                              springer.com/chapter/10.1007/
                                              978-94-017-1767-0_14.

Mehlberg:1980:AAT

[MC80a] Henry Mehlberg and Robert S. Cohen. Alternative approaches to time’s arrow. In Time, Causality, and the Quantum Theory: Studies in the Philosophy of Science. Volume 1: Essay on the Causal Theory of Time [Coh80], chapter 8, pages 147–160. ISBN 90-277-1074-0, 94-009-8935-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-009-
8935-1_8.

Mehlberg:1980:CNT

[MC80b] Henry Mehlberg and Robert S. Cohen.

The causal nature of time. In Time, Causality, and the Quantum Theory: Studies in the Philosophy of Science. Volume Two: Time in a Quantized Universe [MC80z], chapter 9, pages 243–249. ISBN 90-277-1076-7, 94-009-8988-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

BJ1-1725.  URL https://link.
springer.com/chapter/10.1007/
978-94-009-8988-7_9.

Mehlberg:1980:CAP

[MC80c] Henry Mehlberg and Robert S. Cohen.

Conceptual analysis of psychophysical parallelism. In Time, Causality, and the Quantum Theory: Studies in the Philosophy of Science. Volume 1: Essay on the Causal Theory of Time [Coh80], chapter 13, pages 269–285. ISBN 90-277-1074-0, 94-009-8935-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-009-8935-1_13.

Mehlberg:1980:C

[MC80d] Henry Mehlberg and Robert S. Cohen. Conclusion. In Time, Causality, and the Quantum Theory: Studies in the Philosophy of Science. Volume Two: Time in a Quantized Universe [MC80z], chapter 12, pages 259–261. ISBN 90-277-1076-7, 94-009-8988-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN BJ1-1725.

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-009-8988-
                                              7_12.

Mehlberg:1980:IVI

[MC80e] Henry Mehlberg and Robert S. Cohen. Introduction to volumes I and II. In Time, Causality, and the Quantum Theory: Studies in the Philosophy of Science. Volume 1: Essay on the Causal Theory of Time [Coh80], chapter 1, pages 1–36. ISBN 90-277-1074-0, 94-009-8935-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

                                              ???? URL https://link.springer.
                                              com/chapter/10.1007/978-94-009-
                                              8935-1_1.

Mehlberg:1980:IFK

[MC80f] Henry Mehlberg and Robert S. Cohen. The intuitive foundations of the knowledge of time. In Time, Causality, and the Quantum Theory: Studies in the Philosophy of Science. Volume 1: Essay on the Causal Theory of Time [Coh80], chapter 9, pages 163–186. ISBN 90-277-1074-0, 94-009-8935-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-009-8935-
                                              1_9.

Mehlberg:1980:KPI

[MC80g] Henry Mehlberg and Robert S. Cohen. Kant’s phenomenalist interpretation of the causal theory of time. In Time, Causality, and the Quantum Theory: Studies in the Philosophy of Science. Volume 1: Essay on the Causal Theory of Time [Coh80], chapter 3, pages 51– 69. ISBN 90-277-1074-0, 94-009-8935-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-009-8935-
1_3.

Mehlberg:1980:LNT

[MC80h] Henry Mehlberg and Robert S. Cohen.

Laws of nature and time’s arrow. In Time, Causality, and the Quantum Theory: Studies in the Philosophy of Science. Volume Two: Time in a Quantized Universe [MC80z], chapter 4, pages 152–202. CODEN ???? ISBN 90-277-1076-7, 94-009-8988-1 (e-book). ISSN ???? LCCN BJ1-1725.

URL https://link.springer.com/
chapter/10.1007/978-94-009-8988-
7_4.

Mehlberg:1980:LAC

[MC80i] Henry Mehlberg and Robert S. Cohen. Lechalas’ adaptation of the causal theory of time to the laws of pre-Einsteinian physics. In Time, Causality, and the Quantum Theory: Studies in the Philosophy of Science. Volume 1: Essay on the Causal Theory of Time [Coh80], chapter 4, pages 70– 90. ISBN 90-277-1074-0, 94-009-8935-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-009-8935-
1_4.

Mehlberg:1980:LBC

[MC80j] Henry Mehlberg and Robert S. Cohen.

Leibniz and the beginnings of the causal theory of time. In Time, Causality, and the Quantum Theory: Studies in the Philosophy of Science. Volume 1: Essay on the Causal Theory of Time [Coh80], chapter 2, pages 42–50. ISBN 90-277-1074-0, 94-009-8935-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-009-8935-1_2.

Mehlberg:1980:MT

[MC80k] Henry Mehlberg and Robert S. Cohen. The measurement of time. In Time, Causality, and the Quantum Theory: Studies in the Philosophy of Science. Volume Two: Time in a Quantized Universe [MC80z], chapter 6, pages 211–224. ISBN 90-277-1076-7, 94-009-8988-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

                                              BJ1-1725.  URL https://link.
                                              springer.com/chapter/10.1007/
                                              978-94-009-8988-7_6.

Mehlberg:1980:NPT

[MC80l] Henry Mehlberg and Robert S. Cohen. Non-physical time. In Time, Causality, and the Quantum Theory: Studies in the Philosophy of Science. Volume 1: Essay on the Causal Theory of Time [Coh80], chapter 11, pages 237–260. ISBN 90-277-1074-0, 94-009-8935-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-009-8935-
                                              1_11.

Mehlberg:1980:OST

[MC80m] Henry Mehlberg and Robert S. Cohen. The ontological status of time. In Time, Causality, and the Quantum Theory: Studies in the Philosophy of Science. Volume Two: Time in a Quantized Universe [MC80z], chapter 7, pages 225–234. ISBN 90-277-1076-7, 94-009-8988-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN BJ1-1725. URL https://link.

                                              springer.com/chapter/10.1007/
                                              978-94-009-8988-7_7.

Mehlberg:1980:PIQ

[MC80n] Henry Mehlberg and Robert S. Cohen.

Philosophical interpretations of quantum physics. In Time, Causality, and the Quantum Theory: Studies in the Philosophy of Science. Volume Two: Time in a Quantized Universe [MC80z], chapter 1, pages 3–74. CODEN ???? ISBN 90-277-1076-7, 94-009-8988-1 (e-book). ISSN ???? LCCN BJ1-1725.

URL https://link.springer.com/
chapter/10.1007/978-94-009-8988-
7_1.

Mehlberg:1980:PTa

[MC80o] Henry Mehlberg and Robert S. Cohen.

Physical time. In Time, Causality, and the Quantum Theory: Studies in the Philosophy of Science. Volume 1: Essay on the Causal Theory of Time [Coh80], chapter 10, pages 187–236. ISBN 90-277-1074-0, 94-009-8935-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-009-8935-1_10.

Mehlberg:1980:PES

[MC80p] Henry Mehlberg and Robert S. Cohen. The present empirical status of psychophysical parallelism. In Time, Causality, and the Quantum Theory: Studies in the Philosophy of Science. Volume 1: Essay on the Causal Theory of Time [Coh80], chapter 12, pages 261–268. ISBN 90-277-1074-0, 94-009-8935-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-009-
8935-1_12.

Mehlberg:1980:PCI

[MC80q] Henry Mehlberg and Robert S. Cohen.

The problem of causality in an indeterministic science. In Time, Causality, and the Quantum Theory: Studies in the Philosophy of Science. Volume Two: Time in a Quantized Universe [MC80z], chapter 2, pages 75–112. CODEN ???? ISBN 90-277-1076-7, 94-009-8988-1 (e-book). ISSN ???? LCCN BJ1-1725.

URL https://link.springer.com/
chapter/10.1007/978-94-009-8988-
7_2.

Mehlberg:1980:PTb

[MC80r] Henry Mehlberg and Robert S. Cohen. The psychology of time. In Time, Causality, and the Quantum Theory: Studies in the Philosophy of Science. Volume Two: Time in a Quantized Universe [MC80z], chapter 11, pages 255–258. ISBN 90-277-1076-7, 94-009-8988-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

BJ1-1725.  URL https://link.
springer.com/chapter/10.1007/
978-94-009-8988-7_11.

Mehlberg:1980:RT

[MC80s] Henry Mehlberg and Robert S. Cohen. The reality of time. In Time, Causality, and the Quantum Theory: Studies in the Philosophy of Science. Volume Two: Time in a Quantized Universe [MC80z], chapter 8, pages 235–242. ISBN 90-277-1076-7, 94-009-8988-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

BJ1-1725.  URL https://link.
springer.com/chapter/10.1007/
978-94-009-8988-7_8.

Mehlberg:1980:RPCa

[MC80t] Henry Mehlberg and Robert S. Cohen. The relativistic phase of the causal theory of time: The axiomatic systems of Robb and Carnap. In Time, Causality, and the Quantum Theory: Studies in the Philosophy of Science. Volume 1:

Essay on the Causal Theory of Time [Coh80], chapter 5, pages 91–104. ISBN 90-277-1074-0, 94-009-8935-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-009-8935-1_5.

Mehlberg:1980:RPCb

[MC80u] Henry Mehlberg and Robert S. Cohen. The relativistic phase of the causal theory of time: The work of Reichenbach. In Time, Causality, and the Quantum Theory: Studies in the Philosophy of Science. Volume 1: Essay on the Causal Theory of Time [Coh80], chapter 6, pages 105–133. ISBN 90-277-1074-0, 94-009-8935-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-009-8935-
1_6.

Mehlberg:1980:RA

[MC80v] Henry Mehlberg and Robert S. Cohen.

Relativity and the atom. In Time, Causality, and the Quantum Theory: Studies in the Philosophy of Science. Volume Two: Time in a Quantized Universe [MC80z], chapter 3, pages 113–151. CODEN ???? ISBN 90-277-1076-7, 94-009-8988-1 (e-book). ISSN ???? LCCN BJ1-1725. URL https:

//link.springer.com/chapter/10.
1007/978-94-009-8988-7_3.

Mehlberg:1980:RCE

[MC80w] Henry Mehlberg and Robert S. Cohen. Russell’s causal explanation of duration. In Time, Causality, and the Quantum Theory: Studies in the Philosophy of Science. Volume 1: Essay on the Causal Theory of Time [Coh80], chapter 7, pages 134–146. ISBN 90-277-1074-0, 94-009-8935-0 (e-book).

ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-009-8935-1_7.

Mehlberg:1980:ST

[MC80x] Henry Mehlberg and Robert S. Cohen. The symmetry of time. In Time, Causality, and the Quantum Theory: Studies in the Philosophy of Science. Volume Two: Time in a Quantized Universe [MC80z], chapter 10, pages 250–254. ISBN 90-277-1076-7, 94-009-8988-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

                                              BJ1-1725.  URL https://link.
                                              springer.com/chapter/10.1007/
                                              978-94-009-8988-7_10.

Mehlberg:1980:STB

[MC80y] Henry Mehlberg and Robert S. Cohen. The symmetry of time and the branch hypothesis. In Time, Causality, and the Quantum Theory: Studies in the Philosophy of Science. Volume Two: Time in a Quantized Universe [MC80z], chapter 5, pages 203–208. ISBN 90-277-1076-7, 94-009-8988-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN BJ1-1725. URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-009-8988-7_5.

Mehlberg:1980:TCQ

[MC80z] Henry Mehlberg and Robert S. Cohen. Time, Causality, and the Quantum Theory: Studies in the Philosophy of Science. Volume Two: Time in a Quantized Universe, volume 19-2 of Boston Studies in the Philosophy of Science. D. Reidel, Dordrecht, The Netherlands; Boston, MA, USA; Lancaster, UK; Tokyo, Japan, 1980. ISBN 90-277-1076-7, 94-009-8988-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). x + 294 pp. LCCN BJ1-1725.

URL https://link.springer.com/
book/10.1007/978-94-009-8988-7.

Marion:1995:QSP

[MC95] Mathieu Marion and Robert S. Cohen.

Québec Studies in the Philosophy of Science: Part I: Logic, Mathematics, Physics and History of Science Essays in Honor of Hugues Leblanc, volume 177 of Boston Studies in the Philosophy of Science. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 1995. ISBN 94-010-7204-3, 94-009-1575-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). xi + 331 pp. LCCN ????

URL https://link.springer.com/
book/10.1007/978-94-009-1575-6.

Marion:1996:QSP

[MC96] Mathieu Marion and Robert S. Cohen.

Québec Studies in the Philosophy of Science: Part II: Biology, Psychology, Cognitive Science and Economics Essays in Honor of Hugues Leblanc, volume 178 of Boston Studies in the Philosophy of Science. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 1996. ISBN 94-010-6537-3, 94-009-0113-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). xi + 308 pp. LCCN ???? URL

https://link.springer.com/book/
10.1007/978-94-009-0113-1.

Marx:2014:PCP

[MC14] Werner Marx and Manuel Cardona.

Physics in Cuba from the perspective of bibliometrics. In The History of Physics in Cuba [BRW14a], chapter 33, pages 423–437. ISBN 94-017-8040-4, 94-017-8041-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC5.53; QC6.9; QC9 C9H57 2014; Q124.6-127.2. URL https:/

                                              /link.springer.com/chapter/10.
                                              1007/978-94-017-8041-4_33.

McAllister:1997:SAP

[McA97]  James W. McAllister.  Scientists’

aesthetic preferences among theories: Conservative factors in revolutionary crises. In The Elusive Synthesis: Aesthetics and Science [Tau97a], chapter 8, pages 169–187. ISBN 0-7923-4763-3, 0-7923-3904-5, 94-009-1786-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-009-1786-6_8.

McBride:1994:MVB

[McB94] William McBride. The Marxian vision of a (better) possible future: End of a grand illusion? In Artifacts, Representations and Social Practice: Essays for Marx Wartofsky [GC94], chapter 26, pages 441–461. ISBN 94-010-4390-6, 94-011-0902-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN B73 .A77 1994. URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-011-0902-4_26.

McCarthy:1974:OCV

[McC74] Thomas McCarthy. The operation called Verstehen: Towards a redefinition of the problem. In Schaffner and Cohen [SC74a], pages 167–193. ISBN 90-277-0408-2, 90-277-0409-0, 94-010-2140-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175 .B7312. URL https:/

                                              /link.springer.com/chapter/10.
                                              1007/978-94-010-2140-1_12.

McCarthy:1994:CDS

[McC94] Thomas McCarthy. On the communicative dimension of social practice. In Artifacts, Representations and Social Practice: Essays for Marx Wartofsky [GC94], chapter 27, pages 463–482. ISBN 94-010-4390-6, 94-011-0902-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN B73 .A77

1994. URL https://link.springer.
com/chapter/10.1007/978-94-011-
0902-4_27.

McCall:1995:BIQ

[McC95a] Storrs McCall. A branched interpretation of quantum mechanics which differs from Everett’s. In Québec Studies in the Philosophy of Science: Part I: Logic, Mathematics, Physics and History of Science Essays in Honor of Hugues Leblanc [MC95], chapter 13, pages 229–241. ISBN 94-010-7204-3, 94-009-1575-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-009-
1575-6_13.

McCormmach:1995:LEH

[McC95b] Russell McCormmach. The last experiment of Henry Cavendish. In No Truth Except in the Details: Essays in Honor of Martin J. Klein [KS95], chapter 1, pages 1–30. ISBN 94-010-4097-4, 94-011-0217-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-011-
0217-9_1.

McClain:2002:PVB

[McC02] Ernest G. McClain. A priestly view of Bible arithmetic: Deity’s regulative aesthetic activity within Davidic musicology. In Hermeneutic Philosophy of Science, Van Gogh’s Eyes, and God: Essays in Honor of Patrick A. Heelan, S.J. [Bab02c], chapter 38, pages 429–443. ISBN 90-481-5926-1, 94-017-1767-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 225. URL https:/

                                              /link.springer.com/chapter/10.
                                              1007/978-94-017-1767-0_38.

Mcguire:1974:FPA

[Mcg74] J. E. Mcguire. Forces, powers, aethers, and fields. In Methodological and Historical Essays in the Natural and Social Sciences [CW74b], chapter 7, pages 119–159. ISBN 90-277-0378-7, 94-010-2128-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 14. URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-010-2128-9_7.

McGeachie:1987:OCI

[McG87] James McGeachie. Organicism, culture and ideology in Late Victorian Britain: The uses of complexity. In Burwick [Bur87b], chapter 8, pages 211–252. ISBN 90-277-2541-1, 94-009-3917-5. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol.

105. US$64.00(U.S.). URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-009-3917-2_8.

McGilvray:1996:MCO

[McG96] James A. McGilvray. Making colored objects. In Québec Studies in the Philosophy of Science: Part II: Biology, Psychology, Cognitive Science and Economics Essays in Honor of Hugues Leblanc [MC96], chapter 6, pages 81– 97. ISBN 94-010-6537-3, 94-009-0113-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-009-0113-
                                              1_6.

McKeon:1975:OSR

[McK75] Richard McKeon. The organization of sciences and the relations of cultures in the Twelfth and Thirteenth Centuries. In Murdoch and Sylla [MS75b], pages 151–192. ISBN 90-277-0587-9, 94-010-1781-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN B721 .I584 1975. URL https://link.

springer.com/chapter/10.1007/
978-94-010-1781-7_6.

McKenna:1992:SAG

[McK92] Andrew J. McKenna. Supplement to Apocalypse: Girard and Derrida. In Understanding Origins: Contemporary Views on the Origin of Life, Mind and Society [VD92], chapter 3, pages 45–76. ISBN 90-481-4090-0, 94-015-8054-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-015-8054-
0_3.

McLaughlin:1971:MFA

[McL71] Andrew McLaughlin. Method and factual agreement in science. In Buck and Cohen [BC71], pages 459–469. ISBN 90-277-0309-4, 94-010-3142-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-010-3142-4_29.

McMullin:1971:CNE

[McM71] Ernan McMullin. Capacities and natures: An exercise in ontology. In Buck and Cohen [BC71], pages 63–82. ISBN 90-277-0309-4, 94-010-3142-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-010-3142-4_5.

McMullin:1974:ES

[McM74a] Ernan McMullin. Empiricism at sea. In Methodological and Historical Essays in the Natural and Social Sciences [CW74b], chapter 2, pages 21–32. ISBN 90-277-0378-7, 94-010-2128-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 14.

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-010-2128-
                                              9_2.

McMullin:1974:LRC

[McM74b] Ernan McMullin. Logicality and rationality: A comment on Toulmin’s theory of science. In Seeger and Cohen [SC74b], pages 415–430. ISBN 90-277-0390-6, 90-277-0376-0 (paperback), 94-010-2126-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 11. URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-010-2126-5_24.

McMullin:1976:FTU

[McM76a] Ernan McMullin. The fertility of theory and the unit for appraisal in science. In Essays in Memory of Imre Lakatos [CFW76], chapter 25, pages 395–432. ISBN 90-277-0655-7, 94-010-1451-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-010-1451-
                                              9_25.

Mcmullin:1976:HPS

[Mcm76b] Ernan Mcmullin. History and philosophy of science: A marriage of convenience? In Cohen et al. [CHMV76], pages 585–601. ISBN 90-277-0647-6, 90-277-0648-4 (paperback), 94-010-1449-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-010-
1449-6_33.

McMullin:1978:PSR

[McM78] Ernán McMullin. Philosophy of science and its rational reconstructions. In Progress and Rationality in Science [RA78b], chapter 10, pages 221–252. ISBN 90-277-0922-X, 94-009-9866-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-009-9866-
7_10.

McMullin:1984:PSD

[McM84] Ernan McMullin. Is the progress of science dialectical? In Hegel and the Sciences [CW84a], chapter 13, pages 215–239. ISBN 94-009-6235-5, 94-009-6233-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-009-
6233-0_13.

McMullin:1986:ES

[McM86] Ernan McMullin. Empiricism at sea.

In Cohen and Wartofsky [CW85], chapter 9, pages 121–132. ISBN 90-277-1971-3, 94-009-5345-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-009-5345-1_9.

McReynolds:1980:CMH

[McR80] Paul McReynolds. The clock metaphor in the history of psychology. In Scientific Discovery: Case Studies [Nic80e], chapter 6, pages 97–112. ISBN 90-277-1093-7, 94-009-9015-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 60.

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-009-9015-
                                              9_6.

Medina:1996:PTA

[Med96] Manuel Medina. The philosophy of technology assessment. In Munévar [Mun96], chapter 10, pages 201–226. ISBN 94-010-6622-1, 94-009-0305-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-009-0305-
                                              0_10.

Megill:1974:AMU

[Meg74] Kenneth A. Megill. The appeal of Marxism in the United States. In For Dirk Struik: Scientific, Historical and Political Essays in Honor of Dirk J. Struik [CSW74], chapter 42, pages 561–573. ISBN 90-277-0379-5, 94-010-2115-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

                                              ???? URL https://link.springer.
                                              com/chapter/10.1007/978-94-010-
                                              2115-9_42.

Meixner:1997:OTF

[Mei97] Uwe Meixner. The objectivity of time–flux and the direction of time. In Perspectives on Time [FSU97b], chapter 4, pages 107–117. ISBN 90-481-4774-3, 94-015-8875-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol.189.

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-015-8875-
                                              1_4.

Meidan:2017:LPS

[Mei17a] Abraham Meidan. Lessons from the philosophy of science to data mining and vice versa. In Encouraging Openness: Essays for Joseph Agassi on the Occasion of His 90th Birthday [BAG17], chapter 10, pages 105–111. ISBN 3-319-57668-2 (hardcover), 3-319-57669-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN B833.E53 2017. URL https://link.

springer.com/chapter/10.1007/
978-3-319-57669-5_10.

Meital:2017:CIA

[Mei17b] Amir Meital. Can the Israeli academia replace mandatory conscription? In Encouraging Openness: Essays for Joseph Agassi on the Occasion of His 90th Birthday [BAG17], chapter 24, pages 291–300. ISBN 3-319-57668-2 (hardcover), 3-319-57669-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN B833.E53 2017.

URL https://link.springer.com/
chapter/10.1007/978-3-319-57669-
5_24.

Melcher:2003:IEM

[Mel03] Horst Melcher. Interpretations and equations of the Michelson experiment and its variations. In Renn et al. [RDS03], chapter 2, pages 13–25. ISBN 1-4020-1284-5, 1-4020-1285-3 (paperback). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 234. URL https://link.

springer.com/chapter/10.1007/
978-94-010-0111-3_2.

Menger:1970:MIM

[Men70] Karl Menger. Mathematical implications of Mach’s ideas: Positivistic geometry, the clarification of functional connections. In Ernst Mach: Physicist and Philosopher [CS70], chapter 7, pages 107–125. ISBN 90-481-8318-9, 90-277-0016-8, 94-017-1462-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-017-1462-4_7.

Mendelsohn:1974:SSS

[Men74] Everett Mendelsohn. Should science survive its success? In For Dirk Struik: Scientific, Historical and Political Essays in Honor of Dirk J. Struik [CSW74], chapter 31, pages 373–389. ISBN 90-277-0379-5, 94-010-2115-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-010-2115-
                                              9_31.

Mendelsohn:1976:PBV

[Men76] Everett I. Mendelsohn. Philosophical biology vs experimental biology: Spontaneous generation in the Seventeenth Century. In Topics in the Philosophy of Biology [GM76], chapter 2, pages 37– 65. ISBN 90-277-0596-8, 94-010-1829-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-010-1829-
                                              6_2.

Mendelsohn:1986:KPS

[Men86] Everett Mendelsohn. Knowledge and power in the sciences. In Ullmann-Margalit [UM86a], chapter 18, pages 225–240. ISBN 90-277-2159-9 (paperback), 90-277-2158-0 (hardcover), 94-009-5496-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 94 Q175.

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-009-5496-
                                              0_18.

Merzkirch:1970:MCD

[Mer70] Wolfgang F. Merzkirch. Mach’s contribution to the development of gas dynamics. In Ernst Mach: Physicist and Philosopher [CS70], chapter 3, pages 42–59. ISBN 90-481-8318-9, 90-277-0016-8, 94-017-1462-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-017-1462-4_3.

Merton:1976:AS

[Mer76] Robert K. Merton. The ambivalence of scientists. In Essays in Memory of Imre Lakatos [CFW76], chapter 26, pages 433–455. ISBN 90-277-0655-7, 94-010-1451-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-010-
1451-9_26.

Mermin:1997:HAV

[Mer97] N. David Mermin. How to ascertain the values of every member of a set of observables that cannot all have values. In Cohen et al. [CHS97b], pages 149–157. ISBN 90-481-4809-X, 94-017-2732-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC173.96-174.52. URL https:

//link.springer.com/chapter/10.
1007/978-94-017-2732-7_11.

Mestrovic:1989:SWI

[Meš89] Stjepan G. Meštrović. Schopenhauer’s will and idea in Durkheim’s methodology. In The Qualitative-Quantitative Distinction in the Social Sciences [GM89b], chapter 7, pages 99–118. ISBN 90-481-8460-6, 94-017-3444-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 112. URL https://link.

springer.com/chapter/10.1007/
978-94-017-3444-8_7.

Mestrovic:1990:SSR

[Mes90] Stjepan G. Mestrovic. Simmel’s sociology in relation to Schopenhauer’s philosophy. In Georg Simmel and Contemporary Sociology [KPC90], chapter 11, pages 181–197. ISBN 94-010-6691-4 (hardcover), 94-009-0459-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN HM401-1281.

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-009-0459-
                                              0_11.

Metaxopoulos:1989:CCL

[Met89] Emilio Metaxopoulos. A critical consideration of the Lakatosian concepts: “mature” and “immature” science. In Imre Lakatos and Theories of Scientific Change [GGN89], chapter 16, pages 203–214. ISBN 94-010-7860-2, 94-009-3025-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

                                              ???? URL https://link.springer.
                                              com/chapter/10.1007/978-94-009-
                                              3025-4_16.

Meyerson:1976:ETC

[Mey76] E. Meyerson. The elimination of time in classical science. In The Concepts of Space and Time: Their Structure and Their Development [Čap76], chapter 45, pages 255–264. ISBN 90-277-0375-2, 94-010-1727-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic).

                                              LCCN BD632. URL https://link.
                                              springer.com/chapter/10.1007/
                                              978-94-010-1727-5_45.

Meyerson:1985:AR

[Mey85a] Émile Meyerson. The appeal of relativism. In The Relativistic Deduction: Epistemological Implications of the Theory of Relativity [Mey85v], chapter 17, pages 120–131. ISBN 90-277-1699-4, 94-009-5211-2. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC173.585; Q174 .B67

vol. 83. URL https://link.springer.
com/chapter/10.1007/978-94-009-

5211-9_17. With a review by Albert Einstein.

Meyerson:1985:ATE

[Mey85b] Émile Meyerson. The aprioristic tendency and experience. In The Relativistic Deduction: Epistemological Implications of the Theory of Relativity [Mey85v], chapter 23, pages 179–197. ISBN 90-277-1699-4, 94-009-5211-2. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC173.585; Q174 .B67 vol. 83. URL https://link.

springer.com/chapter/10.1007/
978-94-009-5211-9_23.  With a

review by Albert Einstein.

Meyerson:1985:BP

[Mey85c] Émile Meyerson. Biological phenomena.

In The Relativistic Deduction: Epistemological Implications of the Theory of Relativity [Mey85v], chapter 10, pages 86–87. ISBN 90-277-1699-4, 94-009-5211-2. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC173.585; Q174 .B67 vol. 83. URL https:/

/link.springer.com/chapter/10.
1007/978-94-009-5211-9_10. With a

review by Albert Einstein.

Meyerson:1985:DR

[Mey85d] Émile Meyerson. The deducible and the real. In The Relativistic Deduction: Epistemological Implications of the Theory of Relativity [Mey85v], chapter 18, pages 132–142. ISBN 90-277-1699-4, 94-009-5211-2. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC173.585; Q174 .B67

vol. 83. URL https://link.springer.
com/chapter/10.1007/978-94-009-

5211-9_18. With a review by Albert Einstein.

Meyerson:1985:D

[Mey85e] Émile Meyerson. Diversity. In The Relativistic Deduction: Epistemological Implications of the Theory of Relativity [Mey85v], chapter 14, pages 104–110. ISBN 90-277-1699-4, 94-009-5211-2. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC173.585; Q174 .B67 vol. 83. URL https://link.

                                              springer.com/chapter/10.1007/
                                              978-94-009-5211-9_14.  With a

review by Albert Einstein.

Meyerson:1985:DSS

[Mey85f] Émile Meyerson. Dogmatism and skepticism in science. In The Relativistic Deduction: Epistemological Implications of the Theory of Relativity [Mey85v], chapter 25, pages 216–237. ISBN 90-277-1699-4, 94-009-5211-2. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC173.585; Q174 .B67 vol. 83. URL https://link.

                                              springer.com/chapter/10.1007/
                                              978-94-009-5211-9_25.  With a

review by Albert Einstein.

Meyerson:1985:EP

[Mey85g] Émile Meyerson. Electrical phenomena. In The Relativistic Deduction: Epistemological Implications of the Theory of Relativity [Mey85v], chapter 9, pages 80–85. ISBN 90-277-1699-4, 94-009-5211-2. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC173.585; Q174 .B67 vol. 83. URL https:/

                                              /link.springer.com/chapter/10.
                                              1007/978-94-009-5211-9_9. With a

review by Albert Einstein.

Meyerson:1985:EE

[Mey85h] Émile Meyerson. Essence and existence.

In The Relativistic Deduction: Epistemological Implications of the Theory of Relativity [Mey85v], chapter 13, pages 101–103. ISBN 90-277-1699-4, 94-009-5211-2. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC173.585; Q174 .B67 vol. 83. URL https:/

/link.springer.com/chapter/10.
1007/978-94-009-5211-9_13. With a

review by Albert Einstein.

Meyerson:1985:ER

[Mey85i] Émile Meyerson. The evolution of reason. In The Relativistic Deduction: Epistemological Implications of the Theory of Relativity [Mey85v], chapter 24, pages 198–215. ISBN 90-277-1699-4, 94-009-5211-2. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC173.585; Q174 .B67

vol. 83. URL https://link.springer.
com/chapter/10.1007/978-94-009-

5211-9_24. With a review by Albert Einstein.

Meyerson:1985:G

[Mey85j] Émile Meyerson. Gravitation. In The Relativistic Deduction: Epistemological Implications of the Theory of Relativity [Mey85v], chapter 7, pages 64–69. ISBN 90-277-1699-4, 94-009-5211-2. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC173.585; Q174 .B67 vol. 83. URL https://link.

springer.com/chapter/10.1007/
978-94-009-5211-9_7. With a review

by Albert Einstein.

Meyerson:1985:I

[Mey85k] Émile Meyerson. Interpretation. In The Relativistic Deduction: Epistemological Implications of the Theory of Relativity [Mey85v], chapter 15, pages 111–113. ISBN 90-277-1699-4, 94-009-5211-2. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC173.585; Q174 .B67 vol. 83. URL https://link.

springer.com/chapter/10.1007/
978-94-009-5211-9_15.  With a

review by Albert Einstein.

Meyerson:1985:M

[Mey85l] Émile Meyerson. Matter. In The Relativistic Deduction: Epistemological Implications of the Theory of Relativity [Mey85v], chapter 12, pages 95–100. ISBN 90-277-1699-4, 94-009-5211-2. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC173.585; Q174 .B67 vol. 83. URL https://link.

springer.com/chapter/10.1007/
978-94-009-5211-9_12.  With a

review by Albert Einstein.

Meyerson:1985:OF

[Mey85m] Émile Meyerson. The outlook for the future. In The Relativistic Deduction: Epistemological Implications of the Theory of Relativity [Mey85v], chapter 26, pages 238–251. ISBN 90-277-1699-4, 94-009-5211-2. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC173.585; Q174 .B67

vol. 83. URL https://link.springer.
com/chapter/10.1007/978-94-009-

5211-9_26. With a review by Albert Einstein.

Meyerson:1985:P

[Mey85n] Emile Meyerson. Preface. In The Relativistic Deduction: Epistemological Implications of the Theory of Relativity [Mey85v], chapter 1, pages 3–7. ISBN 90-277-1699-4, 94-009-5211-2. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC173.585; Q174 .B67 vol. 83. URL https://link.

springer.com/chapter/10.1007/
978-94-009-5211-9_1. With a review

by Albert Einstein.

Meyerson:1985:PI

[Mey85o] Émile Meyerson. The principle of inertia. In The Relativistic Deduction: Epistemological Implications of the Theory of Relativity [Mey85v], chapter 5, pages 34–44. ISBN 90-277-1699-4, 94-009-5211-2. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC173.585; Q174 .B67 vol. 83. URL https://link.

springer.com/chapter/10.1007/
978-94-009-5211-9_5. With a review

by Albert Einstein.

Meyerson:1985:PMT

[Mey85p] Émile Meyerson. Progress in making things rational. In The Relativistic Deduction: Epistemological Implications of the Theory of Relativity [Mey85v], chapter 22, pages 176–178. ISBN 90-277-1699-4, 94-009-5211-2. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC173.585; Q174 .B67 vol. 83. URL https://link.

springer.com/chapter/10.1007/
978-94-009-5211-9_22.  With a

review by Albert Einstein.

Meyerson:1985:Q

[Mey85q] Émile Meyerson. The quantitative. In The Relativistic Deduction: Epistemological Implications of the Theory of Relativity [Mey85v], chapter 2, pages 8– 17. ISBN 90-277-1699-4, 94-009-5211-2. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC173.585; Q174 .B67 vol. 83. URL https://link.

springer.com/chapter/10.1007/
978-94-009-5211-9_2. With a review

by Albert Einstein.

Meyerson:1985:REP

[Mey85r] Émile Meyerson. Rational explanation and the progress of mathematics. In The Relativistic Deduction: Epistemological Implications of the Theory of Relativity [Mey85v], chapter 21, pages 170–175. ISBN 90-277-1699-4, 94-009-5211-2. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC173.585; Q174 .B67 vol. 83. URL https://link.

                                              springer.com/chapter/10.1007/
                                              978-94-009-5211-9_21.  With a

review by Albert Einstein.

Meyerson:1985:R

[Mey85s] Émile Meyerson. Reality. In The Relativistic Deduction: Epistemological Implications of the Theory of Relativity [Mey85v], chapter 3, pages 18–26. ISBN 90-277-1699-4, 94-009-5211-2. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC173.585; Q174 .B67 vol. 83. URL https://link.

                                              springer.com/chapter/10.1007/
                                              978-94-009-5211-9_3. With a review

by Albert Einstein.

Meyerson:1985:RTA

[Mey85t] Émile Meyerson. Relativism, a theory about reality. In The Relativistic Deduction: Epistemological Implications of the Theory of Relativity [Mey85v], chapter 6, pages 45–63. ISBN 90-277-1699-4, 94-009-5211-2. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC173.585; Q174 .B67 vol. 83. URL https://link.

                                              springer.com/chapter/10.1007/
                                              978-94-009-5211-9_6. With a review

by Albert Einstein.

Meyerson:1985:RM

[Mey85u] Émile Meyerson. Relativism and mechanism. In The Relativistic Deduction: Epistemological Implications of the Theory of Relativity [Mey85v], chapter 20, pages 154–169. ISBN 90-277-1699-4, 94-009-5211-2. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC173.585; Q174 .B67

vol. 83. URL https://link.springer.
com/chapter/10.1007/978-94-009-

5211-9_20. With a review by Albert Einstein.

Meyerson:1985:RDE

[Mey85v] Émile Meyerson. The Relativistic Deduction: Epistemological Implications of the Theory of Relativity, volume 83 of Boston Studies in the Philosophy of Science. D. Reidel, Dordrecht, The Netherlands; Boston, MA, USA; Lancaster, UK; Tokyo, Japan, 1985. ISBN 90-277-1699-4, 94-009-5211-2. ISSN 0068-0346 (print), 2214-7942 (electronic). liii + 268 pp. LCCN QC173.585; Q174 .B67 vol. 83. URL

https://link.springer.com/book/
10.1007/978-94-009-5211-9. With a

review by Albert Einstein.

Meyerson:1985:RI

[Mey85w] Émile Meyerson. The relativistic imagination. In The Relativistic Deduction: Epistemological Implications of the Theory of Relativity [Mey85v], chapter 16, pages 114–119. ISBN 90-277-1699-4, 94-009-5211-2. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC173.585; Q174 .B67

vol. 83. URL https://link.springer.
com/chapter/10.1007/978-94-009-

5211-9_16. With a review by Albert Einstein.

Meyerson:1985:Sa

[Mey85x] Émile Meyerson. The spatial. In The Relativistic Deduction: Epistemological Implications of the Theory of Relativity [Mey85v], chapter 4, pages 27–33. ISBN 90-277-1699-4, 94-009-5211-2. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC173.585; Q174 .B67 vol. 83. URL https://link.

                                              springer.com/chapter/10.1007/
                                              978-94-009-5211-9_4. With a review

by Albert Einstein.

Meyerson:1985:Sb

[Mey85y] Émile Meyerson. The system. In The Relativistic Deduction: Epistemological Implications of the Theory of Relativity [Mey85v], chapter 19, pages 143–153. ISBN 90-277-1699-4, 94-009-5211-2. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC173.585; Q174 .B67 vol. 83. URL https://link.

                                              springer.com/chapter/10.1007/
                                              978-94-009-5211-9_19.  With a

review by Albert Einstein.

Meyerson:1985:T

[Mey85z] Émile Meyerson. Time. In The Relativistic Deduction: Epistemological Implications of the Theory of Relativity [Mey85v], chapter 8, pages 70–79. ISBN 90-277-1699-4, 94-009-5211-2. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC173.585; Q174 .B67 vol. 83. URL https://link.

                                              springer.com/chapter/10.1007/
                                              978-94-009-5211-9_8. With a review

by Albert Einstein.

Meyerson:1985:UE

[Mey85-27] Émile Meyerson. Universal explanation. In The Relativistic Deduction: Epistemological Implications of the Theory of Relativity [Mey85v], chapter 11, pages 88–94. ISBN 90-277-1699-4, 94-009-5211-2. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC173.585; Q174 .B67 vol. 83. URL https:/

                                              /link.springer.com/chapter/10.

1007/978-94-009-5211-9_11. With a

review by Albert Einstein.

Meyerson:1991:BP

[Mey91a] Émile Meyerson. Biological phenomena. In Explanation in the Sciences [Mey91d], chapter 7, pages 177–205. ISBN 94-010-5511-4, 94-011-3414-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-011-3414-
9_7.

Meyerson:1991:D

[Mey91b] Émile Meyerson. Deduction. In Explanation in the Sciences [Mey91d], chapter 3, pages 47–71. ISBN 94-010-5511-4, 94-011-3414-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-011-3414-9_3.

Meyerson:1991:EP

[Mey91c] Émile Meyerson. The epistemological paradox. In Explanation in the Sciences [Mey91d], chapter 17, pages 492–519. ISBN 94-010-5511-4, 94-011-3414-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-011-3414-
9_17.

Meyerson:1991:ES

[Mey91d] Émile Meyerson. Explanation in the Sciences, volume 128 of Boston Studies in the Philosophy of Science. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 1991. ISBN 94-010-5511-4, 94-011-3414-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). xxxvii + 623 pp. LCCN ???? URL

                                              https://link.springer.com/book/
                                              10.1007/978-94-011-3414-9.

Meyerson:1991:FSE

[Mey91e] Émile Meyerson. Forms of spatial explanation. In Explanation in the Sciences [Mey91d], chapter 8, pages 206–239. ISBN 94-010-5511-4, 94-011-3414-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

                                              ???? URL https://link.springer.
                                              com/chapter/10.1007/978-94-011-
                                              3414-9_8.

Meyerson:1991:HC

[Mey91f] Émile Meyerson. Hegel and Comte. In Explanation in the Sciences [Mey91d], chapter 13, pages 350–362. ISBN 94-010-5511-4, 94-011-3414-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-011-3414-9_13.

Meyerson:1991:HDK

[Mey91g] Émile Meyerson. Hegel, Descartes and Kant. In Explanation in the Sciences [Mey91d], chapter 14, pages 363–391. ISBN 94-010-5511-4, 94-011-3414-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-011-3414-
                                              9_14.

Meyerson:1991:HA

[Mey91h] Émile Meyerson. Hegel’s attempt. In Explanation in the Sciences [Mey91d], chapter 11, pages 263–310. ISBN 94-010-5511-4, 94-011-3414-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-011-3414-9_11.

Meyerson:1991:II

[Mey91i] Émile Meyerson. Identity and identification. In Explanation in the Sciences [Mey91d], chapter 5, pages 102–142. ISBN 94-010-5511-4, 94-011-3414-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-011-3414-
9_5.

Meyerson:1991:I

[Mey91j] Émile Meyerson. The irrational. In Explanation in the Sciences [Mey91d], chapter 6, pages 143–176. ISBN 94-010-5511-4, 94-011-3414-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-011-3414-9_6.

Meyerson:1991:OHR

[Mey91k] Émile Meyerson. The oneness of human reason. In Explanation in the Sciences [Mey91d], chapter 18, pages 520–543. ISBN 94-010-5511-4, 94-011-3414-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-011-3414-
9_18.

Meyerson:1991:PSE

[Mey91l] Émile Meyerson. The possibilities of scientific explanation. In Explanation in the Sciences [Mey91d], chapter 9, pages 240–246. ISBN 94-010-5511-4, 94-011-3414-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-011-
3414-9_9.

Meyerson:1991:RR

[Mey91m] Émile Meyerson. The rationality of the real. In Explanation in the Sciences [Mey91d], chapter 4, pages 72–101. ISBN 94-010-5511-4, 94-011-3414-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-011-3414-
9_4.

Meyerson:1991:RRR

[Mey91n] Émile Meyerson. The rationality of the real reconsidered. In Explanation in the Sciences [Mey91d], chapter 16, pages 437–491. ISBN 94-010-5511-4, 94-011-3414-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-011-
3414-9_16.

Meyerson:1991:SO

[Mey91o] Émile Meyerson. Schelling’s objections. In Explanation in the Sciences [Mey91d], chapter 12, pages 311–349. ISBN 94-010-5511-4, 94-011-3414-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-011-3414-
9_12.

Meyerson:1991:SPS

[Mey91p] Émile Meyerson. Science and philosophic systems. In Explanation in the Sciences [Mey91d], chapter 15, pages 395–436. ISBN 94-010-5511-4, 94-011-3414-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-011-
3414-9_15.

Meyerson:1991:SDC

[Mey91q] Émile Meyerson. Science demands the concept of thing. In Explanation in the Sciences [Mey91d], chapter 1, pages 9–31. ISBN 94-010-5511-4, 94-011-3414-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-011-
3414-9_1.

Meyerson:1991:SSE

[Mey91r] Émile Meyerson. Science seeks explanation. In Explanation in the Sciences [Mey91d], chapter 2, pages 32–44. ISBN 94-010-5511-4, 94-011-3414-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-011-3414-9_2.

Meyerson:1991:SP

[Mey91s] Émile Meyerson. State of potentiality. In Explanation in the Sciences [Mey91d], chapter 10, pages 247–260. ISBN 94-010-5511-4, 94-011-3414-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-011-3414-
9_10.

Meyer:1992:ME

[Mey92] Stefan Meyer. Mach and Einstein. In Ernst Mach — A Deeper Look: Documents and New Perspectives [Bla92a], chapter 7, pages 151–164. ISBN 94-010-5230-1, 94-011-2771-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN C5.53; QC6.9.

URL https://link.springer.com/
chapter/10.1007/978-94-011-2771-
4_7.

Meyer:2016:CP

[Mey16] Ulrich Meyer. Consciousness and the present. In Cosmological and Psychological Time [DR16], chapter 8, pages 143–153. ISBN 3-319-22589-8 (hardcover), 3-319-22590-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN BD638 C67 2016.

URL https://link.springer.com/
chapter/10.1007/978-3-319-22590-
6_8.

Magalhaes:2011:CVP

[MG11] João Carlos M. Magalhães and Cedric Gondro. A contemporary view of population genetics in evolution. In Brazilian Studies in Philosophy and History of Science: An account of recent works [KV11], chapter 21, pages 281–290. ISBN 90-481-9421-0 (hardcover), 90-481-9422-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175.3 .B73 2011.

URL https://link.springer.com/
chapter/10.1007/978-90-481-9422-
3_21.

Mitcham:1986:PTI

[MH86] Carl Mitcham and Alois Huning. Philosophy and Technology II: Information Technology and Computers in Theory and Practice, volume 90 of Boston Studies in the Philosophy of Science. D. Reidel, Dordrecht, The Netherlands; Boston, MA, USA; Lancaster, UK; Tokyo, Japan, 1986. ISBN 94-010-8510-2, 94-009-4512-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). 384 (est.) pp. LCCN ????

URL https://link.springer.com/
book/10.1007/978-94-009-4512-8.

Marin:2001:TDS

[MH01] Ricardo Campos Marín and Rafael Huertas. The theory of degeneration in Spain (1886–1920). In The Reception of Darwinism in the Iberian World: Spain, Spanish America and Brazil [GPSR01], chapter 11, pages 171–187. ISBN 94-010-3885-6, 94-010-0602-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-010-0602-
6_11.

Michalos:1971:CBV

[Mic71] Alex C. Michalos. Cost-benefit versus expected utility acceptance rules. In Buck and Cohen [BC71], pages 375–402. ISBN 90-277-0309-4, 94-010-3142-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-010-
3142-4_24.

Michalos:1974:RBM

[Mic74] Alex C. Michalos. Rationality between the maximizers and the satisficers. In Schaffner and Cohen [SC74a], pages 423–445. ISBN 90-277-0408-2, 90-277-0409-0, 94-010-2140-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175 .B7312.

URL https://link.springer.com/
chapter/10.1007/978-94-010-2140-
1_29.

Micunovic:1979:BPC

[Mić79] Dragoljub Mićunović. Bureaucracy and public communication. In Praxis: Yugoslav Essays in the Philosophy and Methodology of the Social Sciences [MP79], chapter 19, pages 303–316. ISBN 90-277-0968-8, 94-009-9355-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-009-9355-
6_19.

Michalos:1983:TAF

[Mic83] Alex C. Michalos. Technology assessment, facts and values. In Philosophy and Technology [DR83], chapter 5, pages 59–81. ISBN 94-009-7126-5, 94-009-7124-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

                                              ???? URL https://link.springer.
                                              com/chapter/10.1007/978-94-009-
                                              7124-0_5.

Michalski:1997:LTE

[Mic97] Krzysztof Michalski. Logic and Time: an Essay on Husserl’s Theory of Meaning, volume 185 of Boston Studies in the Philosophy of Science. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 1997. ISBN 0-7923-4082-5 (hardcover). viii + 162 pp. LCCN Q174 .B67 vol. 185 B3279.H94.

Michalski:1998:HNN

[Mic98] Krzysztof Michalski. Human nature and the nature of time: A Nietzschean metaphor and its consequences. In Philosophies of Nature: The Human Dimension: In Celebration of Erazim Kohák [CT98], chapter 8, pages 107–119. ISBN 90-481-4859-6, 94-017-2614-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic).

                                              LCCN BD581. URL https://link.
                                              springer.com/chapter/10.1007/
                                              978-94-017-2614-6_8.

Michelfelder:2018:CTH

[Mic18] Diane P. Michelfelder. Critical thinking and heuristics: What philosophy can learn from engineering about the back of the envelope. In Mitcham et al. [MLNZ18], chapter 2, pages 13–21. ISBN 3-319-62448-2, 3-319-62450-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN TA15

.P45 2018.  URL https://link.
springer.com/chapter/10.1007/
978-3-319-62450-1_2.

Mie:2007:FTM

[Mie07a] Gustav Mie. Foundations of a theory of matter (excerpts). In Renn et al. [RSS^{+}07b], pages 633–697. ISBN 1-4020-3999-9 (hardcover), 1-4020-4000-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 no. 250 v. 4; QC173.6 .G469

2007. URL https://link.springer.
com/chapter/10.1007/978-1-4020-

4000-9_36. Originally published as “Grundlagen einer Theorie der Materie”, Annalen der Physik, 37, 511–534 (1912), 40, 1–65 (1913).

Mie:2007:PRG

[Mie07b] Gustav Mie. The principle of the relativity of the gravitational potential. In Renn et al. [RSS^{+}07b], pages 729–743. ISBN 1-4020-3999-9 (hardcover), 1-4020-4000-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 no. 250 v. 4; QC173.6 .G469

2007. URL https://link.springer.
com/chapter/10.1007/978-1-4020-

4000-9_38. Originally published as “Das Prinzip von der Relativität der Gravitationspotentials” in Arbeiten aus den Gebieten der Physik, Mathematik, Chemie, Festschrift Julius Elster und Hans Geitel zum sechzigsten Geburtstag, Braunschweig, Friedrich Vieweg und Sohn, 251–268 (1915).

Mie:2007:RCE

[Mie07c] Gustav Mie. Remarks concerning Einstein’s theory of gravitation. In Renn et al. [RSS^{+}07b], pages 699–728. ISBN 1-4020-3999-9 (hardcover), 1-4020-4000-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 no. 250 v. 4; QC173.6 .G469

                                              2007. URL https://link.springer.
                                              com/chapter/10.1007/978-1-4020-

4000-9_37. Originally published as “Bemerkungen zu der Einsteinschen Gravitationstheorie”, Physikalische Zeitschrift 15(3), 115–122 (1914) and 15(3), 169–176 (1914).

Miller:1967:C

[Mil67] Mary B. Miller. Comments. In Cohen and Wartofsky [CW67], pages 264–271. ISBN 90-277-0013-3, 94-010-3508-3. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175 .B731 v.3.

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-010-3508-
                                              8_14.

Milic:1979:SOS

[Mil79] Vojin Milić. Some one-sided conceptions of social determinism. In Praxis: Yugoslav Essays in the Philosophy and Methodology of the Social Sciences [MP79], chapter 4, pages 81–88. ISBN 90-277-0968-8, 94-009-9355-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-009-9355-6_4.

Miller:1987:GCS

[Mil87] Douglas E. Miller. Goethe’s color studies in a new perspective: Die Farbenlehre in English. In Amrine et al. [AZW87], chapter 6, pages 101–112. ISBN 90-277-2265-X, 94-009-3761-X. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol.

97. US$79.00. URL https://link.

                                              springer.com/chapter/10.1007/
                                              978-94-009-3761-1_6.

Milosz:1988:I

[Mil88] Czeslaw Milosz. Idea. In Philosophy, History and Social Action: Essays in Honor of Lewis Feuer with an autobiographic essay by Lewis Feuer [HOO88c], chapter 16, pages 323–324. ISBN 94-010-7793-2, 94-009-2873-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-009-2873-
2_16.

Miller:1995:HLU

[Mil95] David Miller. How little uniformity need an inductive inference presuppose? In Critical Rationalism, Metaphysics and Science: Essays for Joseph Agassi. Volume I [JL95a], chapter 9, pages 143–148. ISBN 94-010-4211-X, 94-011-0471-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-011-
0471-5_9.

Miller:2009:EFE

[Mil09] David Miller. Out of error: Further essays on critical rationalism. In Rethinking Popper [PC09], chapter 31, pages 417–423. ISBN 1-4020-9337-3, 1-4020-9338-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN B1649.P64 R48 2009; Q174 .B67 2009.

URL https://link.springer.com/
chapter/10.1007/978-1-4020-9338-
8_31.

Milkov:2013:BGV

[Mil11a] Nikolay Milkov. The Berlin Group and the Vienna Circle: Affinities and divergences. In The Berlin Group and the Philosophy of Logical Empiricism [MP13], chapter 1, pages 3–32. ISBN 94-007-5484-1 (hardcover), 94-007-5485-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175.32.A24 .M384 2011; Q175 .B47

                                              2013. URL https://link.springer.
                                              com/chapter/10.1007/978-94-007-
                                              5485-0_1.

Milkov:2013:CHW

[Mil11b] Nikolay Milkov. Carl Hempel: Whose philosopher? In The Berlin Group and the Philosophy of Logical Empiricism [MP13], chapter 14, pages 293–309. ISBN 94-007-5484-1 (hardcover), 94-007-5485-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175.32.A24 .M384 2011; Q175 .B47

                                              2013. URL https://link.springer.
                                              com/chapter/10.1007/978-94-007-
                                              5485-0_14.

Miller:2012:HPS

[Mil12] David Marshall Miller. The history and philosophy of science history. In Mauskopf and Schmaltz [MS12b], chapter 3, pages 29–48. ISBN 94-007-1744-X, 94-007-1745-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175 .I5134 2011.

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-007-1745-
                                              9_3.

Milici:2015:GCA

[Mil15] Pietro Milici. A geometrical constructive approach to infinitesimal analysis: Epistemological potential and boundaries of tractional motion. In From Logic to Practice: Italian Studies in the Philosophy of Mathematics [LPV15], chapter 1, pages 3–21. ISBN 3-319-10433-0, 3-319-10434-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QA8.4 .F76 2015; Q174 .B67 v.308. URL https://link.

                                              springer.com/chapter/10.1007/
                                              978-3-319-10434-8_1.

Mink:1967:CST

[Min67] Louis O. Mink. Comment on Stephen Toulmin’s ‘conceptual revolutions in science’. In Cohen and Wartofsky [CW67], pages 348–355. ISBN 90-277-0013-3, 94-010-3508-3. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175 .B731 v.3. URL https://link.

springer.com/chapter/10.1007/
978-94-010-3508-8_19.

Minkowski:2007:MRP

[Min07] Hermann Minkowski. Mechanics and the relativity postulate. In Renn et al. [RSS^{+}07a], pages 275–285. ISBN 1-4020-3999-9. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 no. 250 v. 3; QC173.6

.G469 2007. URL https://link.
springer.com/chapter/10.1007/
978-1-4020-4000-9_20.  Originally

published as “Mechanik und Relativitätspostulat” appended to “Die Grundgleichungen für elektromagnetischen Vorgänge in bewegten Körpern” in Nachrichten der Königlichen Gesellschaft der Wissenschaften zu Göttingen, Mathematisch-Physikalische Klasse, 53–111 (1908).

Minkowski:2010:SSL

[Min10] Christopher Minkowski. Sanskrit scientific libraries and their uses: Examples and problems of the Early Modern Period. In Looking at it from Asia: the Processes that Shaped the Sources of History of Science [BE10b], chapter 3, pages 81–114. ISBN 90-481-3675-X (hardcover), 90-481-3676-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q127.A65 L66

2010. URL https://link.springer.
com/chapter/10.1007/978-90-481-
3676-6_3.

Mirabal:2014:MRP

[Mir14] Carlos A. Cabal Mirabal. Magnetic Resonance Project 35-26-7: A Cuban case of engineering physics and biophysics. In The History of Physics in Cuba [BRW14a], chapter 17, pages 315–322. ISBN 94-017-8040-4, 94-017-8041-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC5.53; QC6.9; QC9 C9H57 2014; Q124.6-127.2. URL https://link.

                                              springer.com/chapter/10.1007/
                                              978-94-017-8041-4_17.

Miroiu:2015:SMR

[Mir15] Adrian Miroiu. The simple majority rule in a three-valued logic framework. In Romanian Studies in Philosophy of Science [PST15], chapter 9, pages 131– 148. ISBN 3-319-16654-9 (hardcover), 3-319-16655-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic).

                                              LCCN ????  URL https://link.
                                              springer.com/chapter/10.1007/
                                              978-3-319-16655-1_9.

Misner:1974:STP

[Mis74] Charles W. Misner. Some topics for philosophical inquiry concerning the theories of mathematical geometrodynamics and of physical geometrodynamics. In Schaffner and Cohen [SC74a], pages 7–29. ISBN 90-277-0408-2, 90-277-0409-0, 94-010-2140-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175 .B7312. URL https://link.

                                              springer.com/chapter/10.1007/
                                              978-94-010-2140-1_2.

Mis:1982:MCD

[Miś82a] Andrzej Miś. Marxism and the controversy over the development of science. In Polish Essays in the Philosophy of the Natural Sciences [Kra82c], chapter 9, pages 103–111. ISBN 94-009-7705-0 (e-book), 90-277-1287-5. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-009-7705-1_9.

Misiek:1982:CM

[Mis82b] JóZef Misiek. On the concept of matter. In Polish Essays in the Philosophy of the Natural Sciences [Kra82c], chapter 18, pages 211–214. ISBN 94-009-7705-0 (e-book), 90-277-1287-5. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-009-7705-1_18.

Misiek:1995:AT

[Mis95a] Józef Misiek. Assessment of theories. In The Problem of Rationality in Science and its Philosophy: On Popper vs. Polanyi: The Polish Conferences 1988– 89 [Mis95c], pages 131–142. ISBN 94-010-4206-3, 94-011-0461-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-011-0461-6_10.

Misiek:1995:PR

[Mis95b] Józef Misiek. Personal rationality. In The Problem of Rationality in Science and its Philosophy: On Popper vs. Polanyi: The Polish Conferences 1988– 89 [Mis95c], pages 125–130. ISBN 94-010-4206-3, 94-011-0461-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-011-0461-6_9.

Misiek:1995:PRS

[Mis95c] Józef Misiek, editor. The Problem of Rationality in Science and its Philosophy: On Popper vs. Polanyi: The Polish Conferences 1988–89, volume 160 of Boston Studies in the Philosophy of Science. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 1995. ISBN 94-010-4206-3, 94-011-0461-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
book/10.1007/978-94-011-0461-6.

Miscevic:2000:QA

[Mis00] Nenad Miscevic. Quining the apriori. In Knowledge, Language and Logic: Questions for Quine [OK00], chapter 8, pages 95–107. ISBN 1-4020-0253-X, 94-011-3933-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-011-
3933-5_8.

Mitroff:1976:IPS

[Mit76a] Ian I. Mitroff. Integrating the philosophy and the social psychology of science or a plague on two houses divided. In Cohen et al. [CHMV76], pages 529–548. ISBN 90-277-0647-6, 90-277-0648-4 (paperback), 94-010-1449-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-010-1449-6_30.

Mittelstaedt:1976:CL

[Mit76b] Peter Mittelstaedt. The causal law. In Philosophical Problems of Modern Physics [Mit76g], chapter 6, pages 133–155. ISBN 90-277-0506-2, 94-010-9617-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC6 .M588 1976. URL https://link.

springer.com/chapter/10.1007/
978-94-010-9617-1_6.

Mittelstaedt:1976:CS

[Mit76c] Peter Mittelstaedt. The concept of substance. In Philosophical Problems of Modern Physics [Mit76g], chapter 5, pages 114–132. ISBN 90-277-0506-2, 94-010-9617-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC6 .M588 1976. URL https:

//link.springer.com/chapter/10.
1007/978-94-010-9617-1_5.

Mittelstaedt:1976:ERG

[Mit76d] Peter Mittelstaedt. Euclidean and Riemannian geometry. In Philosophical Problems of Modern Physics [Mit76g], chapter 3, pages 37–87. ISBN 90-277-0506-2, 94-010-9617-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC6 .M588 1976.

URL https://link.springer.com/
chapter/10.1007/978-94-010-9617-
1_3.

Mittelstaedt:1976:I

[Mit76e] Peter Mittelstaedt. Introduction. In Philosophical Problems of Modern Physics [Mit76g], chapter 1, pages 1–6. ISBN 90-277-0506-2, 94-010-9617-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC6

.M588 1976. URL https://link.
springer.com/chapter/10.1007/
978-94-010-9617-1_1.

Mittelstaedt:1976:LQL

[Mit76f] Peter Mittelstaedt. Logic and quantum logic. In Philosophical Problems of Modern Physics [Mit76g], chapter 7, pages 156–204. ISBN 90-277-0506-2, 94-010-9617-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC6 .M588 1976. URL https:

//link.springer.com/chapter/10.
1007/978-94-010-9617-1_7.

Mittelstaedt:1976:PPM

[Mit76g] Peter Mittelstaedt. Philosophical Problems of Modern Physics, volume 18 of Boston Studies in the Philosophy of Science. D. Reidel, Dordrecht, The Netherlands; Boston, MA, USA; Lancaster, UK; Tokyo, Japan, 1976. ISBN 90-277-0506-2, 94-010-9617-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). x + 211 + 1 pp. LCCN QC6 .M588 1976. URL

                                              https://link.springer.com/book/
                                              10.1007/978-94-010-9617-1.

Mittelstaedt:1976:QL

[Mit76h] Peter Mittelstaedt. Quantum logic. In Cohen et al. [CHMV76], pages 501–514. ISBN 90-277-0647-6, 90-277-0648-4 (paperback), 94-010-1449-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-010-1449-6_28.

Mittelstaedt:1976:QMM

[Mit76i] Peter Mittelstaedt. The quantummechanical measurement process. In Philosophical Problems of Modern Physics [Mit76g], chapter 4, pages 88–113. ISBN 90-277-0506-2, 94-010-9617-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC6 .M588 1976. URL https://link.

                                              springer.com/chapter/10.1007/
                                              978-94-010-9617-1_4.

Mittelstaedt:1976:ST

[Mit76j] Peter Mittelstaedt. Space and time. In Philosophical Problems of Modern Physics [Mit76g], chapter 2, pages 7– 36. ISBN 90-277-0506-2, 94-010-9617-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC6 .M588

                                              1976. URL https://link.springer.

com/chapter/10.1007/978-94-010-
9617-1_2.

Mitcham:1983:RPO

[Mit83] Carl Mitcham. The religious and political origins of modern technology. In Philosophy and Technology [DR83], chapter 18, pages 267–273. ISBN 94-009-7126-5, 94-009-7124-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-009-7124-0_18.

Mitcham:1986:ITP

[Mit86a] Carl Mitcham. Information technology and the problem of incontinence. In Philosophy and Technology II: Information Technology and Computers in Theory and Practice [MH86], chapter 18, pages 247–255. ISBN 94-010-8510-2, 94-009-4512-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-009-4512-8_18.

Mitcham:1986:IIT

[Mit86b] Carl Mitcham. Introduction: Information technology and computers as themes in the philosophy of technology. In Philosophy and Technology II: Information Technology and Computers in Theory and Practice [MH86], chapter 1, pages 1–14. ISBN 94-010-8510-2, 94-009-4512-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-009-
4512-8_1.

Mittelstrass:1989:WPW

[Mit89] Jürgen Mittelstrass. World pictures:

The world of the history and philosophy of science. In An Intimate Relation:

Studies in the History and Philosophy of Science Presented to Robert E. Butts on his 60th Birthday [BM89], chapter 16, pages 319–341. ISBN 94-010-7546-8, 94-009-2327-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-009-2327-0_16.

Mittelstaedt:1997:QMP

[Mit97] Peter Mittelstaedt. Is quantum mechanics a probabilistic theory? In Cohen et al. [CHS97b], pages 159–175. ISBN 90-481-4809-X, 94-017-2732-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC173.96-174.52.

URL https://link.springer.com/
chapter/10.1007/978-94-017-2732-
7_12.

Mittelstaedt:2000:ISI

[Mit00] Peter Mittelstaedt. Individualistic and statistical interpretation of quantum mechanics. In The Reality of the Unobservable: Observability, Unobservability and Their Impact on the Issue of Scientific Realism [AP00b], chapter 24, pages 317–327. ISBN 90-481-5458-8, 94-015-9391-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-015-9391-5_24.

Mittelstrass:2004:EMW

[Mit04] Jürgen Mittelstrass. Die Entdeckung der Möglichkeit von Wissenschaft. (German) [The discovery of possibility in science]. In Classics in the History of Greek Mathematics [Chr04a], chapter 2, pages 19–44. ISBN 1-4020-0081-2 (hardcover), 1-4020-2640-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QA22 .C596 2004.

URL https://link.springer.com/
chapter/10.1007/978-1-4020-2640-
9_2.

Mitcham:2015:RTE

[Mit15] Carl Mitcham. Rationality in technology and in ethics. In New Perspectives on Technology, Values, and Ethics: Theoretical and Practical [Gon15a], chapter 4, pages 63–87. ISBN 3-319-21869-7 (hardcover), 3-319-21870-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-3-319-21870-
0_4.

Muller-Lauter:1999:JWB

[ML99] Wolfgang Müller-Lauter. On judging in a world of becoming: A reflection on the ‘great change’ in Nietzsche’s philosophy. In Nietzsche, Theories of Knowledge, and Critical Theory: Nietzsche and the Sciences I [Bab99b], chapter 12, pages 165–185. ISBN 90-481-5233-X, 94-017-2430-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-017-2430-2_12.

Mitcham:2018:PEE

[MLNZ18] Carl Mitcham, Bocong Li, Byron Newberry, and Baichun Zhang, editors. Philosophy of Engineering, East and West, volume 330 of Boston Studies in the Philosophy and History of Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2018. ISBN 3-319-62448-2, 3-319-62450-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). xl + 333 pp. LCCN TA15 .P45 2018. URL

https://link.springer.com/book/
10.1007/978-3-319-62450-1.

Magnasco:1992:WIB

[MM91] Valerio Magnasco and Roy McWeeny. Weak interactions between molecules and their physical interpretation. In The Invention of Physical Science: Intersections of Mathematics, Theology and Natural Philosophy Since the Seventeenth Century: Essays in Honor of Erwin N. Hiebert [NRS92], chapter 5, pages 133–169. ISBN 94-010-5097-X, 94-011-2488-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-3-642-58183-0_5.

Miller:1994:TIC

[MM94] Alfred E. Miller and Maria G. Miller. Translators’ introduction and commentary. In Kant’s Theory of Natural Science [Pla94b], chapter 1, pages 1–162. ISBN 94-010-4492-9, 94-011-1126-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

                                              ???? URL https://link.springer.
                                              com/chapter/10.1007/978-94-011-
                                              1126-3_1.

Mickiewicz-Olczyk:1982:GHE

[MO82] Elżbieta Mickiewicz-Olczyk. Genetic and historical explanation in biology. In Polish Essays in the Philosophy of the Natural Sciences [Kra82c], chapter 29, pages 335–351. ISBN 94-009-7705-0 (ebook), 90-277-1287-5. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

                                              ???? URL https://link.springer.
                                              com/chapter/10.1007/978-94-009-
                                              7705-1_29.

Mocek:1991:TFB

[Moc91] Reinhard Mocek. Two faces of biologism: Some reflections on a difficult period in the history of biology in Germany. In Woodward and Cohen [WC91], pages 279–291. ISBN 94-010-5395-2, 94-011-3164-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-011-3164-3_25.

Mody:2015:WDS

[Mod15] Cyrus C. M. Mody. What do scientists and engineers do all day? On the structure of scientific normalcy. In Kuhn’s Structure of Scientific Revolutions — 50 Years On [DB15b], chapter 7, pages 91–104. ISBN 3-319-13382-9, 3-319-13383-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175.3 .K84 2015. URL https:

//link.springer.com/chapter/10.
1007/978-3-319-13383-6_7.

Molder:2001:ILE

[Möl01] Bruno Mölder. Inter-level explanation and the category-mistake. In Estonian Studies in the History and Philosophy of Science [Vih01b], chapter 21, pages 283–294. ISBN 94-010-3863-5, 94-010-0672-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-010-
0672-9_21.

Molinini:2018:POC

[Mol18] Daniele Molinini. Parsimony, ontological commitment and the import of mathematics. In Piazza and Pulcini [PP18], pages 189–207. ISBN 3-319-93341-8, 3-319-93342-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QA8.4. URL https:

//link.springer.com/chapter/10.
1007/978-3-319-93342-9_11.

Monk:1980:PRS

[Mon80] Robert Monk. Productive reasoning and the structure of scientific research.

In Scientific Discovery, Logic, and Rationality [Nic80f], chapter 15, pages 337–354. ISBN 90-277-1070-8, 94-009-8986-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-009-
8986-3_15.

Montminy:1996:VMV

[Mon96] Martin Montminy. Verificationism and the molecular view of language. In Québec Studies in the Philosophy of Science: Part II: Biology, Psychology, Cognitive Science and Economics Essays in Honor of Hugues Leblanc [MC96], chapter 17, pages 279–293. ISBN 94-010-6537-3, 94-009-0113-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-009-0113-
1_17.

Montserrat:2001:EMA

[Mon01] Marcelo Montserrat. The evolutionist mentality in Argentina: An ideology of progress. In The Reception of Darwinism in the Iberian World: Spain, Spanish America and Brazil [GPSR01], chapter 1, pages 1–27. ISBN 94-010-3885-6, 94-010-0602-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-010-0602-6_1.

Mongin:2009:DTE

[Mon09] Philippe Mongin. Duhemian themes in expected utility theory. In Brenner and Gayon [BG09], chapter 13, pages 303–357. ISBN 1-4020-9367-5, 1-4020-9368-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .F746 2009. URL https:

//link.springer.com/chapter/10.
1007/978-1-4020-9368-5_13.

Moore:1967:CW

[Moo67] Asher Moore. The center of the world.

In Cohen and Wartofsky [CW67], pages 356–375. ISBN 90-277-0013-3, 94-010-3508-3. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175

.B731 v.3.  URL https://link.
springer.com/chapter/10.1007/
978-94-010-3508-8_20.

Moore:1990:SRS

[Moo90] O. K. Moore. Simmel on the ratio of subjective values to objective cultural possibilities. In Georg Simmel and Contemporary Sociology [KPC90], chapter 12, pages 199–224. ISBN 94-010-6691-4 (hardcover), 94-009-0459-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN HM401-1281.

URL https://link.springer.com/
chapter/10.1007/978-94-009-0459-
0_12.

Moon:2018:EPH

[Moo18] Suzanne Moon. Engineering and the postcolonial: Historical perspectives and ethical practices. In Mitcham et al. [MLNZ18], chapter 16, pages 187–200. ISBN 3-319-62448-2, 3-319-62450-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN TA15 .P45 2018. URL https://link.

springer.com/chapter/10.1007/
978-3-319-62450-1_16.

Morgenbesser:1963:PCA

[Mor63]  Sidney Morgenbesser.  Perception:

Cause and achievement. In Wartofsky [War63], pages 206–212. ISBN 94-010-3265-3, 94-010-3263-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B678 1970.

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-010-3263-
                                              6_17.

Morrison:1971:TKT

[Mor71] Paul G. Morrison. Two kinds of theory in the social sciences. In Buck and Cohen [BC71], pages 565–572. ISBN 90-277-0309-4, 94-010-3142-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-010-3142-4_38.

Morrison:1974:OC

[Mor74] Philip Morrison. Open or closed? In Seeger and Cohen [SC74b], pages 189– 202. ISBN 90-277-0390-6, 90-277-0376-0 (paperback), 94-010-2126-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 11.

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-010-2126-
                                              5_11.

More:1976:DBE

[Mor76] H. More. On the difference between extension and matter. In The Concepts of Space and Time: Their Structure and Their Development [Čap76], chapter 15, pages 85–87. ISBN 90-277-0375-2, 94-010-1727-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN BD632.

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-010-1727-
                                              5_15.

Moran:1980:WIH

[Mor80] Bruce T. Moran. Wilhelm IV of Hesse– Kassel: Informal communication and the aristocratic context of discovery. In Scientific Discovery: Case Studies [Nic80e], chapter 5, pages 67–96. ISBN 90-277-1093-7, 94-009-9015-4 (e-book).

ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 60.

URL https://link.springer.com/
chapter/10.1007/978-94-009-9015-
9_5.

Moravcsik:1983:NLF

[Mor83] Julius M. Moravcsik. Natural languages and formal languages: A tenable dualism. In Language, Logic and Method [CW83b], chapter 12, pages 225–239. ISBN 94-009-7704-2, 94-009-7702-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-009-7702-
0_12.

Morrison:1989:HCC

[Mor89] Margaret Morrison. Hypotheses and certainty in cartesian science. In An Intimate Relation: Studies in the History and Philosophy of Science Presented to Robert E. Butts on his 60th Birthday [BM89], chapter 3, pages 43–64. ISBN 94-010-7546-8, 94-009-2327-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-009-2327-0_3.

Morange:2009:ADM

[Mor09] Michel Morange. Articulating different modes of explanation: The present boundary in biological research. In Mapping the Future of Biology: Evolving Concepts and Theories [BMP09b], chapter 2, pages 15–26. ISBN 1-4020-9635-6 (hardcover), 1-4020-9636-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QH331 .M295

2009. URL https://link.springer.
com/chapter/10.1007/978-1-4020-
9636-5_2.

Morrison:2011:BPA

[Mor11] Margaret Morrison. Between the pure and applied: The search for the elusive middle ground. In Science in the Context of Application [CN11a], chapter 3, pages 31–45. ISBN 90-481-9050-9, 90-481-9051-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q180.55.M4 S385 2011. URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-90-481-9051-5_3.

Moriconi:2018:SRT

[Mor18] Enrico Moriconi. Some remarks on true undecidable sentences. In Piazza and Pulcini [PP18], pages 3–15. ISBN 3-319-93341-8, 3-319-93342-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QA8.4.

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-3-319-93342-
                                              9_1.

Moser:1989:EF

[Mos89] Walter Moser. Experiment and fiction. In Literature and Science as Modes of Expression [Amr89], chapter 3, pages 61–80. ISBN 94-010-7531-X, 94-009-2297-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

                                              ???? URL https://link.springer.
                                              com/chapter/10.1007/978-94-009-
                                              2297-6_3.

Mosterin:1996:PC

[Mos96] Jesús Mosterín. Philosophy and cosmology. In Munévar [Mun96], chapter 3, pages 57–88. ISBN 94-010-6622-1, 94-009-0305-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

                                              ???? URL https://link.springer.
                                              com/chapter/10.1007/978-94-009-
                                              0305-0_3.

Mosterin:2000:OCS

[Mos00] Jesús Mosterín. Observation, construction and speculation in cosmology. In The Reality of the Unobservable: Observability, Unobservability and Their Impact on the Issue of Scientific Realism [AP00b], chapter 16, pages 219–229. ISBN 90-481-5458-8, 94-015-9391-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-015-
9391-5_16.

Motycka:1982:EMS

[Mot82] Alina Motycka. The epistemological and methodological sense of the concept of rationality. In Polish Essays in the Philosophy of the Natural Sciences [Kra82c], chapter 2, pages 19–30. ISBN 94-009-7705-0 (e-book), 90-277-1287-5. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-009-7705-1_2.

Mota:2015:MFG

[Mot15] Teresa Salomé Mota. From the museum to the field: Geology teaching in the Faculty of Sciences of the University of Lisbon. In Sciences in the Universities of Europe, Nineteenth and Twentieth Centuries: Academic Landscapes [SDG15b], chapter 20, pages 345–360. ISBN 94-017-9635-1, 94-017-9636-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q183.4.E85

2015. URL https://link.springer.
com/chapter/10.1007/978-94-017-
9636-1_20.

Moulines:1982:SP

[Mou82] Carlos-Ulises Moulines. A study in protophysics. In Scientific Philosophy Today: Essays in Honor of Mario Bunge [AC82], chapter 13, pages 207–224. ISBN 90-277-1262-X, 90-277-1263-8, 94-009-8462-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 67.

URL https://link.springer.com/
chapter/10.1007/978-94-009-8462-
2_13.

Moulin:1986:FS

[Mou86] Anne-Marie Moulin. Fleck’s style. In Cognition and Fact: Materials on Ludwik Fleck [CS86], chapter 21, pages 407–419. ISBN 94-010-8504-8, 94-009-4498-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-009-
4498-5_21.

Moulines:1989:ERP

[Mou89] C. Ulises Moulines. The emergence of a research programme in classical thermodynamics. In Imre Lakatos and Theories of Scientific Change [GGN89], chapter 9, pages 111–121. ISBN 94-010-7860-2, 94-009-3025-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-009-3025-4_9.

Moulin:1992:PSN

[Mou92] Anne Marie Moulin. Patriarchal science: The network of the Overseas Pasteur Institutes. In Science and Empires: Historical Studies about Scientific Development and European Expansion [PJM92], chapter 31, pages 307–322. ISBN 0-7923-1518-9 (hardcover), 94-010-5145-3, 94-011-2594-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 136. URL https://link.

springer.com/chapter/10.1007/
978-94-011-2594-9_31.

Markovic:1979:PYE

[MP79] Mihailo Marković and Gajo Petrović.

Praxis: Yugoslav Essays in the Philosophy and Methodology of the Social Sciences, volume 36 of Boston Studies in the Philosophy of Science. D. Reidel, Dordrecht, The Netherlands; Boston, MA, USA; Lancaster, UK; Tokyo, Japan, 1979. ISBN 90-277-0968-8, 94-009-9355-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). xxxviii + 416 pp. LCCN ???? URL

https://link.springer.com/book/
10.1007/978-94-009-9355-6.

Milkov:2013:BGP

[MP13] Nikolay Milkov and Volker Peckhaus.

The Berlin Group and the Philosophy of Logical Empiricism, volume 273 of Boston Studies in the Philosophy and History of Science. Springer-Ver-lag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2013. ISBN 94-007-5484-1 (hardcover), 94-007-5485-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). x + 332 pp. LCCN Q175.32.A24 .M384 2011; Q175 .B47 2013. URL

https://link.springer.com/book/
10.1007/978-94-007-5485-0.

May:2016:DNS

[MP16] Robert C. May and Marco Panza.

Discussion note on: “Semantic Nominalism: How I Learned to Stop Worrying and Love Universals” by G. Aldo Antonelli. In Objectivity, Realism, and Proof: FilMat Studies in the Philosophy of Mathematics [BS16], chapter 3, pages 33–41. ISBN 3-319-31642-7, 3-319-31644-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QA9 .O25 2016.

URL https://link.springer.com/

                                              chapter/10.1007/978-3-319-31644-

4_3. See [Ant16].

Maggiora:1992:EES

[MPC91] Gerald M. Maggiora, James D. Petke, and Ralph E. Christoffersen. Electronic excited states of biomolecular systems: Ab initio FSGO-based quantum mechanical methods with applications to photosynthetic and related systems. In The Invention of Physical Science: Intersections of Mathematics, Theology and Natural Philosophy Since the Seventeenth Century: Essays in Honor of Erwin N. Hiebert [NRS92], chapter 3, pages 65–102. ISBN 94-010-5097-X, 94-011-2488-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

                                              ???? URL https://link.springer.
                                              com/chapter/10.1007/978-3-642-
                                              58183-0_3.

Marcus:1963:D

[MQK^{+}63] Ruth Barcan Marcus, Willard Van Orman Quine, Saul Kripke, John McCarthy, and Dagfinn Føllesdal. Discussion. In Wartofsky [War63], pages 105–116. ISBN 94-010-3265-3, 94-010-3263-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B678 1970. URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-010-3263-6_9.

Murdoch:1975:I

[MS75a] John E. Murdoch and Edith Dudley Sylla. Introduction. In Murdoch and Sylla [MS75b], pages 1–30. ISBN 90-277-0587-9, 94-010-1781-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN B721 .I584 1975.

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-010-1781-
                                              7_1.

Murdoch:1975:CCM

[MS75b] John Emery Murdoch and Edith Dudley Sylla, editors. The Cultural Context of Medieval Learning: Proceedings of the First International Colloquium on Philosophy, Science, and Theology in the Middle Ages — September 1973, volume 26 of Boston Studies in the Philosophy of Science. D. Reidel, Dordrecht, The Netherlands; Boston, MA, USA; Lancaster, UK; Tokyo, Japan, 1975. ISBN 90-277-0587-9, 94-010-1781-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN B721 .I584 1975. URL

https://link.springer.com/book/
10.1007/978-94-010-1781-7.

Margalit:1983:RLR

[MS83] Avishai Margalit and Israel Scheffler.

Rejoinder to Levi’s reply. In Language, Logic and Method [CW83b], chapter 10, pages 197–198. ISBN 94-009-7704-2, 94-009-7702-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-009-
7702-0_10.

Marcovich:2011:CII

[MS11] Anne Marcovich and Terry Shinn. The cognitive, instrumental and institutional origins of nanoscale research: The place of biology. In Science in the Context of Application [CN11a], chapter 14, pages 221–242. ISBN 90-481-9050-9, 90-481-9051-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q180.55.M4 S385

2011. URL https://link.springer.
com/chapter/10.1007/978-90-481-
9051-5_14.

Mauskopf:2012:I

[MS12a] Seymour Mauskopf and Tad Schmaltz.

Introduction. In Mauskopf and Schmaltz [MS12b], chapter 1, pages 1– 10. ISBN 94-007-1744-X, 94-007-1745-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175 .I5134

2011. URL https://link.springer.
com/chapter/10.1007/978-94-007-
1745-9_1.

Mauskopf:2012:IHP

[MS12b] Seymour H. Mauskopf and Tad M. Schmaltz, editors. Integrating History and Philosophy of Science: Problems and Prospects, volume 263 of Boston Studies in the Philosophy of Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2012. ISBN 94-007-1744-X, 94-007-1745-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). x + 249 pp. LCCN Q175 .I5134 2011. URL

https://link.springer.com/book/
10.1007/978-94-007-1745-9.

Mundici:2017:TM

[MS17] Daniele Mundici and Wilfried Sieg. Turing, the mathematician. In Floyd and Bokulich [FB17], chapter 2, pages 39–62. ISBN 3-319-53278-2, 3-319-53280-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QA29.T8 P45 2017. URL https:/

/link.springer.com/chapter/10.
1007/978-3-319-53280-6_2.

Muelder:1995:GRC

[Mue95] Walter George Muelder. A gloss on Robert S. Cohen’s ambiguities of science. In Science, Mind and Art: Essays on science and the humanistic understanding in art, epistemology, religion and ethics In honor of Robert S. Cohen [GSW95b], chapter 16, pages 279–294. ISBN 0-7923-2990-2 (hardcover), 94-011-0469-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 165 BD418.3. URL https://link.

springer.com/chapter/10.1007/
978-94-011-0469-2_16.

Mugnai:1981:LSR

[Mug81] Massimo Mugnai. Leibniz on the structure of relations. In Italian Studies in the Philosophy of Science [Chi81b], chapter 21, pages 389–409. ISBN 90-277-1073-2, 94-009-8937-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-009-8937-5_21.

Mulcrone:1974:RTT

[Mul74] Thomas Mulcrone. Remarks on twoby-two matric semigroups. In For Dirk Struik: Scientific, Historical and Political Essays in Honor of Dirk J. Struik [CSW74], chapter 8, pages 91–99. ISBN 90-277-0379-5, 94-010-2115-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-010-2115-9_8.

Muller:1991:NTP

[Mül91a] Karl H. Müller. Neurath’s theory of pictorial–statistical representation. In Uebel [Ueb91b], chapter 18, pages 223– 251. ISBN 0-7923-1276-7 (hardcover), 94-011-3182-1. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 133. URL https://link.

springer.com/chapter/10.1007/
978-94-011-3182-7_18.

Muller:1991:IGE

[Mül91b] Martin Müller. On the interdisciplinary genesis of experimental methods in Nineteenth-Century German psychology. In Woodward and Cohen [WC91], pages 129–140. ISBN 94-010-5395-2, 94-011-3164-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

                                              ???? URL https://link.springer.
                                              com/chapter/10.1007/978-94-011-
                                              3164-3_12.

Muller:1994:WLO

[Mül94] Gerhard H. Müller. Wechselwirkung in the life and other sciences: A word, new claims and a concept around 1800 … and much later. In Poggi and Bossi [PB94], chapter 1, pages 1–14. ISBN 0-7923-2336-X, 94-017-2921-2. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 152.

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-017-2921-
                                              5_1.

Mulligan:1998:RHH

[Mul98] Joseph F. Mulligan. The reception of Heinrich Hertz’s Principles of Mechanics by his contemporaries. In Heinrich Hertz: Classical Physicist, Modern Philosopher [BHN98], chapter 11, pages 173–181. ISBN 90-481-4881-2, 94-015-8855-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic).

                                              LCCN ????  URL https://link.
                                              springer.com/chapter/10.1007/
                                              978-94-015-8855-3_11.

Munson:1976:BA

[Mun76] Ronald Munson. Biological adaptation. In Topics in the Philosophy of Biology [GM76], chapter 15, pages 330–350. ISBN 90-277-0596-8, 94-010-1829-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-010-1829-
                                              6_15.

Munz:1989:WEB

[Mun89] Peter Munz. Watkins’s evolutionism between Hume and Kant. In Freedom and Rationality: Essays in Honor of John Watkins From his Colleagues and Friends [DJ89], chapter 14, pages 225–246. ISBN 94-010-7571-9, 94-009-2380-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-009-
2380-5_14.

Munevar:1991:BRE

[Mun91a] Gonzalo Munévar. Beyond Reason:

Essays on the Philosophy of Paul Feyerabend, volume 132 of Boston Studies in the Philosophy of Science. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 1991. ISBN 94-010-5406-1, 94-011-3188-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). xxi + 538 pp. LCCN ????

URL https://link.springer.com/
book/10.1007/978-94-011-3188-9.

Munevar:1991:SFF

[Mun91b] Gonzalo Munévar. Science in Feyerabend’s free society. In Beyond Reason: Essays on the Philosophy of Paul Feyerabend [Mun91a], chapter 7, pages 179–198. ISBN 94-010-5406-1, 94-011-3188-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-011-
3188-9_7.

Munevar:1996:SSP

[Mun96] Gonzalo Munévar, editor. Spanish Studies in the Philosophy of Science, volume 186 of Boston Studies in the Philosophy of Science. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 1996. ISBN 94-010-6622-1, 94-009-0305-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). xviii + 363 pp. LCCN ???? URL

                                              https://link.springer.com/book/
                                              10.1007/978-94-009-0305-0.

Munch:2014:FGD

[Mün14]  Richard Münch.  Faded grandeur:

Disciplinary differentiation, interdisciplinarity and renewal in the German academic system. In The Institution of Science and the Science of Institutions: The Legacy of Joseph Ben-David [Her14b], chapter 6, pages 83–101. ISBN 94-007-7406-0 (hardcover), 94-007-7407-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 v.302; Q125 .I58 2014.

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-007-7407-
                                              0_6.

Muntean:2015:SPD

[Mun15] Ioan Muntean. Structural pluralism and S-dualities: A project in string realism. In Romanian Studies in Philosophy of Science [PST15], chapter 13, pages 199–219. ISBN 3-319-16654-9 (hardcover), 3-319-16655-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-3-319-16655-1_13.

Murdoch:1975:SIF

[Mur75] John E. Murdoch. From social into intellectual factors: An aspect of the unitary character of late medieval learning. In Murdoch and Sylla [MS75b], pages 271–348. ISBN 90-277-0587-9, 94-010-1781-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN B721 .I584 1975. URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-010-1781-7_9.

Murdoch:1994:BED

[Mur94] Dugald Murdoch. The Bohr–Einstein dispute. In Faye and Folse [FF94], chapter 14, pages 303–324. ISBN 0-7923-2378-5 (hardcover), 94-015-8106-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 153 1994; QC16 1994. URL https:

//link.springer.com/chapter/10.
1007/978-94-015-8106-6_14.

Murphey:2012:DQP

[Mur12a] Murray Murphey. The Development of Quine’s Philosophy, volume 291 of Boston Studies in the Philosophy and History of Science. Springer-Ver-lag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2012. ISBN 94-007-2423-3, 94-007-2424-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). xxv + 283 pp. LCCN B945.Q54 M87 2012. URL

https://link.springer.com/book/
10.1007/978-94-007-2424-2.

Murphey:2012:C

[Mur12b] Murray G. Murphey. Conclusion.

In The Development of Quine’s Philosophy [Mur12a], chapter 5, pages 227–245. ISBN 94-007-2423-3, 94-007-2424-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN B945.Q54 M87 2012. URL https:/

/link.springer.com/chapter/10.
1007/978-94-007-2424-2_5.

Murphey:2012:E

[Mur12c] Murray G. Murphey. Erratum. In The Development of Quine’s Philosophy [Mur12a], chapter 6, pages E1–E2. ISBN 94-007-2423-3, 94-007-2424-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN B945.Q54 M87 2012. URL https:/

                                              /link.springer.com/chapter/10.
                                              1007/978-94-007-2424-2_6.

Murphey:2012:WOR

[Mur12d] Murray G. Murphey. From word and object to roots of reference. In The Development of Quine’s Philosophy [Mur12a], chapter 3, pages 105–176. ISBN 94-007-2423-3, 94-007-2424-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN B945.Q54 M87 2012. URL https:/

                                              /link.springer.com/chapter/10.
                                              1007/978-94-007-2424-2_3.

Murphey:2012:LR

[Mur12e] Murray G. Murphey. Last revisions. In The Development of Quine’s Philosophy [Mur12a], chapter 4, pages 177–226. ISBN 94-007-2423-3, 94-007-2424-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN B945.Q54 M87 2012. URL https:/

                                              /link.springer.com/chapter/10.
                                              1007/978-94-007-2424-2_4.

Murphey:2012:QLP

[Mur12f] Murray G. Murphey. Quine: Logic and philosophy. In The Development of Quine’s Philosophy [Mur12a], chapter 1, pages 1–45. ISBN 94-007-2423-3, 94-007-2424-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN B945.Q54 M87 2012.

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-007-2424-
                                              2_1.

Murphey:2012:SO

[Mur12g] Murray G. Murphey. Semantics and ontology. In The Development of Quine’s Philosophy [Mur12a], chapter 2, pages 47–104. ISBN 94-007-2423-3, 94-007-2424-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN B945.Q54 M87 2012. URL https:/

/link.springer.com/chapter/10.
1007/978-94-007-2424-2_2.

Musgrave:1976:MM

[Mus76] Alan Musgrave. Method or madness?

In Essays in Memory of Imre Lakatos [CFW76], chapter 27, pages 457–491. ISBN 90-277-0655-7, 94-010-1451-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-010-1451-
9_27.

Musgrave:1978:ESF

[Mus78] Alan Musgrave. Evidential support, falsification, heuristics, and anarchism. In Progress and Rationality in Science [RA78b], chapter 8, pages 181–201. ISBN 90-277-0922-X, 94-009-9866-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-009-9866-
7_8.

Musgrave:1989:DH

[Mus89a] Alan Musgrave. Deductive heuristics.

In Imre Lakatos and Theories of Scientific Change [GGN89], chapter 2, pages 15–32. ISBN 94-010-7860-2, 94-009-3025-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-009-
3025-4_2.

Musgrave:1989:SSS

[Mus89b] Alan Musgrave. Saving science from scepticism. In Freedom and Rationality: Essays in Honor of John Watkins From his Colleagues and Friends [DJ89], chapter 17, pages 297–323. ISBN 94-010-7571-9, 94-009-2380-5 (e-book).

ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-009-2380-5_17.

Musgrave:1991:MAI

[Mus91] Alan Musgrave. The myth of astronomical instrumentalism. In Beyond Reason: Essays on the Philosophy of Paul Feyerabend [Mun91a], chapter 11, pages 243–280. ISBN 94-010-5406-1, 94-011-3188-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

                                              ???? URL https://link.springer.
                                              com/chapter/10.1007/978-94-011-
                                              3188-9_11.

Musgrave:1995:RI

[Mus95] Alan Musgrave. Realism and idealisation. In Misiek [Mis95c], pages 143–166. ISBN 94-010-4206-3, 94-011-0461-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

                                              ???? URL https://link.springer.
                                              com/chapter/10.1007/978-94-011-
                                              0461-6_11.

Musgrave:1996:RTO

[Mus96] Alan Musgrave. Realism, truth and objectivity. In Cohen et al. [CHR96], pages 19–44. ISBN 90-481-4493-0, 94-015-8638-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

                                              ???? URL https://link.springer.
                                              com/chapter/10.1007/978-94-015-
                                              8638-2_3.

Musgrave:2009:EPB

[Mus09] Alan Musgrave. Experience and perceptual belief. In Rethinking Popper [PC09], chapter 1, pages 5–19. ISBN 1-4020-9337-3, 1-4020-9338-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN B1649.P64 R48 2009; Q174 .B67 2009. URL https:

//link.springer.com/chapter/10.
1007/978-1-4020-9338-8_1.

Muursepp:2001:SMC

[Müü01] Peeter Müürsepp. Science and magic:

Causality. In Estonian Studies in the History and Philosophy of Science [Vih01b], chapter 12, pages 165–178. ISBN 94-010-3863-5, 94-010-0672-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-010-0672-
9_12.

Maturana:1980:ACR

[MV80a] Humberto R. Maturana and Francisco J. Varela. Autopoiesis and Cognition: The Realization of the Living, volume 42 of Boston Studies in the Philosophy of Science. D. Reidel, Dordrecht, The Netherlands; Boston, MA, USA; Lancaster, UK; Tokyo, Japan, 1980. ISBN 90-277-1016-3, 94-009-8947-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). xxx + 146 pp. LCCN ???? URL

https://link.springer.com/book/
10.1007/978-94-009-8947-4.

Maturana:1980:CFG

[MV80b] Humberto R. Maturana and Francisco J. Varela. Cognitive function in general. In Autopoiesis and Cognition: The Realization of the Living [MV80a], chapter 3, pages 8–14. ISBN 90-277-1016-3, 94-009-8947-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-009-8947-4_3.

Maturana:1980:CFP

[MV80c] Humberto R. Maturana and Francisco J. Varela. Cognitive function in particular. In Autopoiesis and Cognition: The Realization of the Living [MV80a], chapter 4, pages 15–40. ISBN 90-277-1016-3, 94-009-8947-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-009-8947-4_4.

Maturana:1980:C

[MV80d] Humberto R. Maturana and Francisco J. Varela. Conclusions. In Autopoiesis and Cognition: The Realization of the Living [MV80a], chapter 6, pages 48–56. ISBN 90-277-1016-3, 94-009-8947-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-009-8947-4_6.

Maturana:1980:DT

[MV80e] Humberto R. Maturana and Francisco J. Varela. Dispensability of teleonomy. In Autopoiesis and Cognition: The Realization of the Living [MV80a], chapter 10, pages 85–87. ISBN 90-277-1016-3, 94-009-8947-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-009-8947-
4_10.

Maturana:1980:DA

[MV80f] Humberto R. Maturana and Francisco J. Varela. Diversity of autopoiesis. In Autopoiesis and Cognition: The Realization of the Living [MV80a], chapter 12, pages 96–111. ISBN 90-277-1016-3, 94-009-8947-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-009-8947-4_12.

Maturana:1980:EA

[MV80g] Humberto R. Maturana and Francisco J. Varela. Embodiments of autopoiesis. In Autopoiesis and Cognition: The Realization of the Living [MV80a], chapter 11, pages 88–95. ISBN 90-277-1016-3, 94-009-8947-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-009-8947-
4_11.

Maturana:1980:Ia

[MV80h] Humberto R. Maturana and Francisco J. Varela. Introduction. In Autopoiesis and Cognition: The Realization of the Living [MV80a], chapter 1, pages 5–6. ISBN 90-277-1016-3, 94-009-8947-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-009-8947-4_1.

Maturana:1980:Ib

[MV80i] Humberto R. Maturana and Francisco J. Varela. Introduction. In Autopoiesis and Cognition: The Realization of the Living [MV80a], chapter 8, pages 73–76. ISBN 90-277-1016-3, 94-009-8947-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-009-8947-4_8.

Maturana:1980:MLO

[MV80j] Humberto R. Maturana and Francisco J. Varela. On machines, living and otherwise. In Autopoiesis and Cognition: The Realization of the Living [MV80a], chapter 9, pages 77–84. ISBN 90-277-1016-3, 94-009-8947-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-009-8947-4_9.

Maturana:1980:PA

[MV80k] Humberto R. Maturana and Francisco J. Varela. Presence of autopoiesis. In Autopoiesis and Cognition: The Realization of the Living [MV80a], chapter 13, pages 112–123. ISBN 90-277-1016-3, 94-009-8947-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-009-8947-4_13.

Maturana:1980:P

[MV80l] Humberto R. Maturana and Francisco J. Varela. The problem. In Autopoiesis and Cognition: The Realization of the Living [MV80a], chapter 2, page 7. ISBN 90-277-1016-3, 94-009-8947-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic).

                                              LCCN ????  URL https://link.
                                              springer.com/chapter/10.1007/
                                              978-94-009-8947-4_2.

Maturana:1980:PNC

[MV80m] Humberto R. Maturana and Francisco J. Varela. Problems in the neurophysiology of cognition. In Autopoiesis and Cognition: The Realization of the Living [MV80a], chapter 5, pages 41– 47. ISBN 90-277-1016-3, 94-009-8947-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-009-8947-
                                              4_5.

Martin:1976:RT

[MW76] Robert L. Martin and Peter W. Woodruff. On representing ‘true-in-L’ in L. In Language in Focus: Foundations, Methods and Systems: Essays in Memory of Yehoshua Bar-Hillel [Kas76b], chapter 7, pages 113–117. ISBN 90-277-0644-1, 90-277-0645-X, 94-010-1876-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN P101-120. URL https://link.

springer.com/chapter/10.1007/
978-94-010-1876-0_7.

Myllyntaus:2003:DS

[Myl03] Timo Myllyntaus. Discovering Switzerland. In Simões et al. [SCD03b], chapter 12, pages 299–328. ISBN 1-4020-1259-4, 94-017-3584-0. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 v. 233.

URL https://link.springer.com/
chapter/10.1007/978-94-017-3584-
1_12.

Myrvold:1997:DPE

[Myr97] Wayne C. Myrvold. The decision problem for entanglement. In Cohen et al. [CHS97b], pages 177–190. ISBN 90-481-4809-X, 94-017-2732-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC173.96-174.52.

URL https://link.springer.com/
chapter/10.1007/978-94-017-2732-
7_13.

Mitcham:2018:EPE

[MZ18] Carl Mitcham and Kang Zhang.

Engineering policy: Exploratory reflections. In Mitcham et al. [MLNZ18], chapter 21, pages 247–259. ISBN 3-319-62448-2, 3-319-62450-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN TA15 .P45 2018.

URL https://link.springer.com/
chapter/10.1007/978-3-319-62450-
1_21.

Nielsen:2007:IKP

[NA07] Keld Nielsen and Hanne Andersen. The influence of Kant’s philosophy on the young H. C. Ørsted. In Hans Christian Ørsted and The Romantic Legacy In Science: Ideas, Disciplines, Practices [BCK07], chapter 6, pages 97–114. ISBN 1-4020-2979-9 (hardcover), 1-4020-2987-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC515.O37 H36 2007. URL https:

//link.springer.com/chapter/10.
1007/978-1-4020-2987-5_6.

Nadeau:1996:EI

[Nad96a] Robert Nadeau. Economics and intentionality. In Québec Studies in the Philosophy of Science: Part II: Biology, Psychology, Cognitive Science and Economics Essays in Honor of Hugues Leblanc [MC96], chapter 10, pages 159–176. ISBN 94-010-6537-3, 94-009-0113-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-009-
0113-1_10.

Nadenik:1996:RGI

[Nád96b] Zbyněk Nádeník. Reception of Grassmann’s ideas in Bohemia. In Schubring [Sch96f], pages 147–153. ISBN 90-481-4758-1, 94-015-8753-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QA8.9-10.3.

URL https://link.springer.com/
chapter/10.1007/978-94-015-8753-
2_12.

Naess:1991:PFG

[Nae91] Arne Naess. Paul Feyerabend — a green hero? In Beyond Reason: Essays on the Philosophy of Paul Feyerabend [Mun91a], chapter 19, pages 403–416. ISBN 94-010-5406-1, 94-011-3188-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/

chapter/10.1007/978-94-011-3188-
9_19.

Nagasaka:1971:EPR

[Nag71] Gen-Ichiro Nagasaka. The Einstein– Podolsky–Rosen Paradox reexamined. In Buck and Cohen [BC71], pages 437–445. ISBN 90-277-0309-4, 94-010-3142-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-010-
3142-4_27.

Nagai:1984:PCF

[Nag84a] Hiroshi Nagai. A philosophical conception of finality in biology. In Rationality in Science and Politics [And84d], chapter 11, pages 205–219. ISBN 90-277-1953-5, 94-009-6254-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-009-6254-5_11.

Nagasaka:1984:QTM

[Nag84b] Gen-Ichiro Nagasaka. Quantum theory of measurement: A non-quantum mechanical approach. In Physical Sciences and History of Physics [CW84c], chapter 8, pages 145–158. ISBN 94-009-7180-X, 94-009-7178-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .P497 1984.

URL https://link.springer.com/
chapter/10.1007/978-94-009-7178-
3_8.

Nagai:1998:PMC

[Nag98] Hiroshi Nagai. Philosophical meanings of the concept of evolution. In Japanese Studies in the Philosophy of Science [NC98], chapter 12, pages 173–188. ISBN 94-010-6176-9, 94-011-5175-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-011-5175-
7_12.

Nails:1986:ABS

[Nai86] Debra Nails. Annotated bibliography of Spinoza and the sciences. In Spinoza and the Sciences [GN86], chapter 13, pages 305–314. ISBN 94-010-8511-0, 94-009-4514-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-009-
4514-2_13.

Nails:1995:SPU

[Nai95] Debra Nails. Socrates and Plato: Understanding the world and changing it. In Science, Mind and Art: Essays on science and the humanistic understanding in art, epistemology, religion and ethics In honor of Robert S. Cohen [GSW95b], chapter 17, pages 295– 313. ISBN 0-7923-2990-2 (hardcover), 94-011-0469-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 165 BD418.3.

URL https://link.springer.com/
chapter/10.1007/978-94-011-0469-
2_17.

Naibo:2015:CG

[Nai15] Alberto Naibo. Constructibility and geometry. In From Logic to Practice: Italian Studies in the Philosophy of Mathematics [LPV15], chapter 8, pages 123–161. ISBN 3-319-10433-0, 3-319-10434-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QA8.4 .F76 2015; Q174 .B67 v.308.

URL https://link.springer.com/
chapter/10.1007/978-3-319-10434-
8_8.

Nakayama:1974:HSS

[Nak74] Shigeru Nakayama. History of science:

A subject for the frustrated. In For Dirk Struik: Scientific, Historical and Political Essays in Honor of Dirk J. Struik [CSW74], chapter 18, pages 213–224. ISBN 90-277-0379-5, 94-010-2115-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-010-
2115-9_18.

Nakayama:1995:COT

[Nak95]  Shigeru Nakayama.  Cosmological

outlooks and technological transfers: A comparative view from eastern periphery. In Gavroglu et al. [GSW95c], chapter 12, pages 225–233. ISBN 0-7923-2991-0 (set), 0-7923-2988-0, 94-017-2658-2. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 163. URL https://link.

springer.com/chapter/10.1007/
978-94-017-2658-0_12.

Nakamura:1998:IE

[Nak98] Hidekichi Nakamura. On the individuation of events. In Japanese Studies in the Philosophy of Science [NC98], chapter 3, pages 21–32. ISBN 94-010-6176-9, 94-011-5175-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-011-5175-7_3.

Napinen:2001:PRB

[Näp01] Leo Näpinen. The problem of the relationship between human and physical realities in Ilya Prigogine’s paradigm of self-organisation. In Estonian Studies in the History and Philosophy of Science [Vih01b], chapter 11, pages 151–164. ISBN 94-010-3863-5, 94-010-0672-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-010-0672-
9_11.

Naraniecki:2009:LOS

[Nar09] Alexander J. Naraniecki. Logic and the open society: Revising the place of Tarski’s theory of truth within Popper’s political philosophy. In Rethinking Popper [PC09], chapter 20, pages 257–271. ISBN 1-4020-9337-3, 1-4020-9338-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN B1649.P64 R48 2009; Q174 .B67 2009.

URL https://link.springer.com/
chapter/10.1007/978-1-4020-9338-
8_20.

Navarro:2015:PCE

[Nav15] Jaume Navarro. A peripheral centre. Early quantum physics at Cambridge. In Sciences in the Universities of Europe, Nineteenth and Twentieth Centuries: Academic Landscapes [SDG15b], chapter 19, pages 327–344. ISBN 94-017-9635-1, 94-017-9636-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q183.4.E85

2015. URL https://link.springer.
com/chapter/10.1007/978-94-017-
9636-1_19.

Navarro-Brotons:2012:MFS

[NB12] Victor Navarro-Brotons. Matter and form in Sixteenth-Century Spain: Some case studies. In Philosophy of Behavioral Biology [PR12b], chapter 4, pages 99–116. ISBN 94-007-1950-7, 94-007-1951-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QH331 .P4679 2012; Q175 .B657 v.282.

URL https://link.springer.com/
chapter/10.1007/978-94-007-4345-
8_4.

Navarro-Brotons:2013:MFS

[NB13] Victor Navarro-Brotons. Matter and form in Sixteenth-Century Spain: Some case studies. In Garber and Roux [GR13b], chapter 4, pages 99–116. CODEN ???? ISBN 94-007-4344-0 (hardcover), 94-007-4345-9 (e-book). ISSN ???? LCCN QC6 .M3616 2013; QC6 .M43 2013. URL https://link.

springer.com/chapter/10.1007/
978-94-007-4345-8_4.

Nagasaka:1998:JSP

[NC98] Francis Gen-Ichiro Nagasaka and Robert S. Cohen. Japanese Studies in the Philosophy of Science, volume 45 of Boston Studies in the Philosophy of Science. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 1998. ISBN 94-010-6176-9, 94-011-5175-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). xviii + 202 pp. LCCN ????

URL https://link.springer.com/
book/10.1007/978-94-011-5175-7.

Neale:2000:ME

[Nea00] Stephen Neale. On a milestone of empiricism. In Knowledge, Language and Logic: Questions for Quine [OK00], chapter 18, pages 237–346. ISBN 1-4020-0253-X, 94-011-3933-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-011-3933-5_18.

Nedelmann:1990:GSA

[Ned90a] Birgitta Nedelmann. Georg Simmel as an analyst of autonomous dynamics: The merry-go-round of fashion. In Georg Simmel and Contemporary Sociology [KPC90], chapter 14, pages 243–257. ISBN 94-010-6691-4 (hardcover), 94-009-0459-2 (e-book).

ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN HM401-1281.

URL https://link.springer.com/
chapter/10.1007/978-94-009-0459-
0_14.

Nedelmann:1990:CEG

[Ned90b] Birgitta Nedelmann. On the concept of “Erleben” in Georg Simme’s sociology. In Georg Simmel and Contemporary Sociology [KPC90], chapter 13, pages 225–241. ISBN 94-010-6691-4 (hardcover), 94-009-0459-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN HM401-1281.

URL https://link.springer.com/
chapter/10.1007/978-94-009-0459-
0_13.

Nederbragt:2012:MDR

[Ned12]  Hubertus Nederbragt.  Multiple

derivability and the reliability and stabilization of theories. In Characterizing the Robustness of Science: After the Practice Turn in Philosophy of Science [STNW12], chapter 5, pages 121–145. ISBN 94-007-2758-5, 94-007-2759-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175 .C4595 2012. URL https://link.

springer.com/chapter/10.1007/
978-94-007-2759-5_5.

Needham:1997:FTP

[Nee97] Paul Needham. Fleeting things and permanent stuff: A Priorean project in real time. In Perspectives on Time [FSU97b], chapter 5, pages 119–141. ISBN 90-481-4774-3, 94-015-8875-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol.189. URL https://link.

springer.com/chapter/10.1007/
978-94-015-8875-1_5.

Needham:2006:ATC

[Nee06]  Paul Needham.  Aristole’s theory

of chemical reaction and chemical substances. In Philosophy of Chemistry: Synthesis of a New Discipline [BSM06b], chapter 3, pages 43–67. ISBN 1-4020-2455-X, 1-4020-3256-0, 1-4020-3261-7. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 242 QD6. URL https:/

/link.springer.com/chapter/10.
1007/1-4020-3261-7_3.

Needham:2011:CCR

[Nee11a] Paul Needham. The Carnot Cycle and reversible changes. In Commentary on the Principles of Thermodynamics by Pierre Duhem [Nee11c], chapter 6, pages 85–97. ISBN 94-007-0310-4, 94-007-0311-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC311 .D79 2011. URL https:

//link.springer.com/chapter/10.
1007/978-94-007-0311-7_6.

Needham:2011:CTA

[Nee11b] Paul Needham. Carnot’s theorem and absolute temperature. In Commentary on the Principles of Thermodynamics by Pierre Duhem [Nee11c], chapter 7, pages 99–119. ISBN 94-007-0310-4, 94-007-0311-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC311 .D79 2011. URL https:

//link.springer.com/chapter/10.
1007/978-94-007-0311-7_7.

Needham:2011:CPT

[Nee11c] Paul Needham. Commentary on the Principles of Thermodynamics by Pierre Duhem, volume 277 of Boston Studies in the Philosophy of Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2011. ISBN 94-007-0310-4, 94-007-0311-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). xv + 215 pp. LCCN QC311 .D79 2011. URL

https://link.springer.com/book/
10.1007/978-94-007-0311-7.

Needham:2011:C

[Nee11d] Paul Needham. Connections. In Commentary on the Principles of Thermodynamics by Pierre Duhem [Nee11c], chapter 11, pages 171–190. ISBN 94-007-0310-4, 94-007-0311-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC311

.D79 2011.  URL https://link.
springer.com/chapter/10.1007/
978-94-007-0311-7_11.

Needham:2011:ECT

[Nee11e] Paul Needham. Entropy and Clausius’ Theorem. In Commentary on the Principles of Thermodynamics by Pierre Duhem [Nee11c], chapter 8, pages 121–140. ISBN 94-007-0310-4, 94-007-0311-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC311 .D79 2011. URL https:

//link.springer.com/chapter/10.
1007/978-94-007-0311-7_8.

Needham:2011:ESL

[Nee11f] Paul Needham. An examination of the second law of thermodynamics. In Commentary on the Principles of Thermodynamics by Pierre Duhem [Nee11c], chapter 1, pages 3–22. ISBN 94-007-0310-4, 94-007-0311-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC311 .D79 2011.

URL https://link.springer.com/
chapter/10.1007/978-94-007-0311-
7_1.

Needham:2011:HBP

[Nee11g] Paul Needham. Historical background and principal applications of Gibbs’ theory. In Commentary on the Principles of Thermodynamics by Pierre Duhem [Nee11c], chapter 2, pages 23– 34. ISBN 94-007-0310-4, 94-007-0311-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC311 .D79

2011. URL https://link.springer.
com/chapter/10.1007/978-94-007-
0311-7_2.

Needham:2011:PD

[Nee11h] Paul Needham. Preliminary definitions.

In Commentary on the Principles of Thermodynamics by Pierre Duhem [Nee11c], chapter 3, pages 37–49. ISBN 94-007-0310-4, 94-007-0311-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC311 .D79 2011.

URL https://link.springer.com/
chapter/10.1007/978-94-007-0311-
7_3.

Needham:2011:PCE

[Nee11i] Paul Needham. The principle of the conservation of energy. In Commentary on the Principles of Thermodynamics by Pierre Duhem [Nee11c], chapter 4, pages 51–64. ISBN 94-007-0310-4, 94-007-0311-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC311 .D79 2011. URL https:

//link.springer.com/chapter/10.
1007/978-94-007-0311-7_4.

Needham:2011:PSE

[Nee11j] Paul Needham. Properties of a system at equilibrium *. In Commentary on the Principles of Thermodynamics by Pierre Duhem [Nee11c], chapter 9, pages 143–155. ISBN 94-007-0310-4, 94-007-0311-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic).

LCCN QC311 .D79 2011. URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-007-0311-7_9.

Needham:2011:PSM

[Nee11k] Paul Needham. Properties of a system in motion. In Commentary on the Principles of Thermodynamics by Pierre Duhem [Nee11c], chapter 10, pages 157–170. ISBN 94-007-0310-4, 94-007-0311-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC311 .D79 2011. URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-007-0311-7_10.

Needham:2011:SDE

[Nee11l] Paul Needham. Stability and displacement of equilibrium. In Commentary on the Principles of Thermodynamics by Pierre Duhem [Nee11c], chapter 12, pages 191–209. ISBN 94-007-0310-4, 94-007-0311-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC311 .D79 2011. URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-007-0311-7_12.

Needham:2011:WQH

[Nee11m] Paul Needham. Work and quantity of heat. In Commentary on the Principles of Thermodynamics by Pierre Duhem [Nee11c], chapter 5, pages 65–82. ISBN 94-007-0310-4, 94-007-0311-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC311 .D79 2011.

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-007-0311-
                                              7_5.

Needham:2015:OSM

[Nee15] Paul Needham. One substance or more? In Philosophy of Chemistry: Growth of a New Discipline [SM15b], chapter 7, pages 91–105. ISBN 94-017-9363-8 (hardcover), 94-017-9364-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QD6 .P455 2015; QD6 .S384 2014. URL https://link.

springer.com/chapter/10.1007/
978-94-017-9364-3_7.

Negri:1981:CSR

[Neg81] Maurizio Negri. Constructive sequent reduction in Gentzen’s first consistency proof for arithmetic. In Italian Studies in the Philosophy of Science [Chi81b], chapter 10, pages 153–168. ISBN 90-277-1073-2, 94-009-8937-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-009-8937-5_10.

Neira:2015:VRR

[Nei15] Paula Neira. Values regarding results of the information and communication technologies: Internal values. In New Perspectives on Technology, Values, and Ethics: Theoretical and Practical [Gon15a], chapter 3, pages 47–60. ISBN 3-319-21869-7 (hardcover), 3-319-21870-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-3-319-
21870-0_3.

Nelson:1967:EMR

[Nel67] Benjamin Nelson. The early modern revolution in science and philosophy. In Cohen and Wartofsky [CW67], pages 1– 40. ISBN 90-277-0013-3, 94-010-3508-3. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175 .B731 v.3.

URL https://link.springer.com/
chapter/10.1007/978-94-010-3508-
8_1.

Nemeth:1991:ONU

[Nem91a] Elisabeth Nemeth. Otto Neurath’s utopias — the will to hope. In Uebel [Ueb91b], chapter 22, pages 285–292. ISBN 0-7923-1276-7 (hardcover), 94-011-3182-1. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 133. URL https://link.

                                              springer.com/chapter/10.1007/
                                              978-94-011-3182-7_22.

Nemeth:1991:UPE

[Nem91b] Elisabeth Nemeth. The unity of planned economy and the unity of science. In Uebel [Ueb91b], chapter 21, pages 275–283. ISBN 0-7923-1276-7 (hardcover), 94-011-3182-1. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 133. URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-011-3182-7_21.

Nenci:2018:JCE

[Nen18] Elio Nenci. A journey to the center of the Earth: Cosmology and the centrobaric theory from antiquity to the Renaissance. In Feldhay et al. [FRSV18a], chapter 6, pages 139– 178. ISBN 3-319-90343-8 (hardcover), 3-319-90345-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic).

                                              LCCN QA802. URL https://link.
                                              springer.com/chapter/10.1007/
                                              978-3-319-90345-3_6.

Nersessian:1989:SDC

[Ner89] Nancy J. Nersessian. Scientific discovery and commensurability of meaning. In Imre Lakatos and Theories of Scientific Change [GGN89], chapter 23, pages 323–334. ISBN 94-010-7860-2, 94-009-3025-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

                                              ???? URL https://link.springer.

com/chapter/10.1007/978-94-009-
3025-4_23.

Nersessian:2001:CFC

[Ner01] Nancy J. Nersessian. Concept formation and commensurability. In Incommensurability and Related Matters [HHS01], chapter 11, pages 275–301. ISBN 90-481-5709-9, 94-015-9680-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-015-9680-0_11.

Netz:2004:I

[Net04] Reviel Netz. Introduction. In Classics in the History of Greek Mathematics [Chr04a], chapter 5, pages 113–114. ISBN 1-4020-0081-2 (hardcover), 1-4020-2640-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QA22 .C596 2004. URL https:/

/link.springer.com/chapter/10.
1007/978-1-4020-2640-9_5.

Netz:2005:LTG

[Net05] Reviel Netz. The limits of text in Greek mathematics. In History of Science, History of Text [Che05a], chapter 8, pages 161–176. ISBN 1-281-33589-4, 1-4020-2320-0 (hardcover), 1-4020-2321-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 238 Q125.

URL https://link.springer.com/
chapter/10.1007/1-4020-2321-9_8.

Neumann:1976:DES

[Neu76a] C. Neumann. On the definition of the equality of successive intervals of time. In The Concepts of Space and Time: Their Structure and Their Development [Čap76], chapter 41, pages 233–234. ISBN 90-277-0375-2, 94-010-1727-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN BD632.

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-010-1727-
                                              5_41.

Neumann:1976:NAF

[Neu76b] C. Neumann. On the necessity of the absolute frame of reference. In The Concepts of Space and Time: Their Structure and Their Development [Čap76], chapter 22, pages 125–127. ISBN 90-277-0375-2, 94-010-1727-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN BD632.

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-010-1727-
                                              5_22.

Neubauer:1989:NC

[Neu89] John Neubauer. Nature as construct. In Literature and Science as Modes of Expression [Amr89], chapter 7, pages 129–140. ISBN 94-010-7531-X, 94-009-2297-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

                                              ???? URL https://link.springer.
                                              com/chapter/10.1007/978-94-009-
                                              2297-6_7.

Neurath:1991:STO

[Neu91] Paul Neurath. Sociological thought with Otto Neurath. In Uebel [Ueb91b], chapter 17, pages 209–222. ISBN 0-7923-1276-7 (hardcover), 94-011-3182-1. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 133. URL https://link.

                                              springer.com/chapter/10.1007/
                                              978-94-011-3182-7_17.

Neville:1998:CN

[Nev98] Robert Cummings Neville. The contingency of nature. In Philosophies of Nature: The Human Dimension: In Celebration of Erazim Kohák [CT98], chapter 9, pages 121–137. ISBN 90-481-4859-6, 94-017-2614-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN BD581.

URL https://link.springer.com/
chapter/10.1007/978-94-017-2614-
6_9.

Neville:2002:PPC

[Nev02]  Robert Cummings Neville.  Phe-

nomenology and pragmatism: A continuing dialogue with Patrick A. Heelan. In Hermeneutic Philosophy of Science, Van Gogh’s Eyes, and God: Essays in Honor of Patrick A. Heelan, S.J. [Bab02c], chapter 28, pages 323–334. ISBN 90-481-5926-1, 94-017-1767-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 225. URL https:/

/link.springer.com/chapter/10.
1007/978-94-017-1767-0_28.

Newton:1976:ASA

[New76a] I. Newton. On absolute space and absolute motion. In The Concepts of Space and Time: Their Structure and Their Development [Čap76], chapter 18, pages 97–105. ISBN 90-277-0375-2, 94-010-1727-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic).

LCCN BD632. URL https://link.
springer.com/chapter/10.1007/
978-94-010-1727-5_18.

Newton:1976:T

[New76b] I. Newton. On time. In The Concepts of Space and Time: Their Structure and Their Development [Čap76], chapter 36, pages 209–210. ISBN 90-277-0375-2, 94-010-1727-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic).

LCCN BD632. URL https://link.
springer.com/chapter/10.1007/
978-94-010-1727-5_36.

Newman:2015:DEP

[New15] Stuart A. Newman. Development and evolution: The physics connection. In Conceptual Change in Biology: Scientific and Philosophical Perspectives on Evolution and Development [Lov15b], chapter 19, pages 421–440. ISBN 94-017-9411-1, 94-017-9412-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QH331 C66 2015; QH491; QH492 .C66 2014. URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-017-9412-1_19.

Nieto-Galan:2001:AAD

[NG01a] Agustí Nieto-Galan. Artisans and artists in dyeing and printing. In Colouring textiles: a history of natural dyestuffs in industrial Europe [NG01c], chapter 5, pages 153–180. ISBN 0-7923-7022-8, 94-017-1081-3. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN TP919 .N54 2001. URL https://

                                              link.springer.com/chapter/10.
                                              1007/978-94-017-1081-7_5.

Nieto-Galan:2001:CSE

[NG01b] Agustí Nieto-Galan. Circulating skills in a European network: The “Republic of Chemist–Dyers”. In Colouring textiles: a history of natural dyestuffs in industrial Europe [NG01c], chapter 4, pages 123–151. ISBN 0-7923-7022-8, 94-017-1081-3. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN TP919

                                              .N54 2001.  URL https://link.
                                              springer.com/chapter/10.1007/
                                              978-94-017-1081-7_4.

Nieto-Galan:2001:CTH

[NG01c] Agustí Nieto-Galan. Colouring textiles: a history of natural dyestuffs in industrial Europe, volume 217 of Boston Studies in the Philosophy of Science. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 2001. ISBN 0-7923-7022-8, 94-017-1081-3. ISSN 0068-0346 (print), 2214-7942 (electronic). xxv + 246 pp. LCCN TP919

.N54 2001. URL http://www.loc.
gov/catdir/enhancements/fy0813/
2001033541-d.html;   http://www.
loc.gov/catdir/enhancements/
fy0813/2001033541-t.html;
https://link.springer.com/book/
10.1007/978-94-017-1081-7.

Nieto-Galan:2001:NDK

[NG01d] Agustí Nieto-Galan. Natural dyestuffs and the kingdoms of nature. In Colouring textiles: a history of natural dyestuffs in industrial Europe [NG01c], chapter 1, pages 1–41. ISBN 0-7923-7022-8, 94-017-1081-3. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN TP919 .N54 2001. URL https://

link.springer.com/chapter/10.
1007/978-94-017-1081-7_1.

Nieto-Galan:2001:SND

[NG01e] Agustí Nieto-Galan. The “science”

of natural dyestuffs in the laboratory. In Colouring textiles: a history of natural dyestuffs in industrial Europe [NG01c], chapter 3, pages 85–121. ISBN 0-7923-7022-8, 94-017-1081-3. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN TP919 .N54 2001.

URL https://link.springer.com/
chapter/10.1007/978-94-017-1081-
7_3.

Nieto-Galan:2001:SDP

[NG01f] Agustí Nieto-Galan. Sites of dyeing and printing textiles: From the workshop to the factory system. In Colouring textiles: a history of natural dyestuffs in industrial Europe [NG01c], chapter 2, pages 43–84. ISBN 0-7923-7022-8, 94-017-1081-3. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN TP919

                                              .N54 2001.  URL https://link.
                                              springer.com/chapter/10.1007/
                                              978-94-017-1081-7_2.

Nieto-Galan:2001:TAL

[NG01g] Agustí Nieto-Galan. Towards the “artificial”: A long-standing technological change. In Colouring textiles: a history of natural dyestuffs in industrial Europe [NG01c], chapter 6, pages 181–204. ISBN 0-7923-7022-8, 94-017-1081-3. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN TP919 .N54 2001.

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-017-1081-
                                              7_6.

Nieto-Galan:2003:UBC

[NG03] Agustí Nieto-Galan. Under the banner of Catalan industry. In Simões et al. [SCD03b], chapter 8, pages 189–212. ISBN 1-4020-1259-4, 94-017-3584-0. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 v. 233.

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-017-3584-
                                              1_8.

Nickelsen:2009:CBH

[NG09] Kärin Nickelsen and Gerd Graßhoff. Concepts from the bench: Hans Krebs, Kurt Henseleit and the urea cycle. In Going Amiss In Experimental Research [HSS09a], chapter 7, pages 91–117. ISBN 1-4020-8892-2 (hardcover), 1-4020-8893-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q172.5.E77 G647 2009. URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-1-4020-8893-3_7.

Nieto-Galan:2015:CPR

[NG15a]  Agustí Nieto-Galan.  Centers and

peripheries revisited: STEP and the mainstream historiography of science. In Arabatzis et al. [ARS15], chapter 6, pages 69–84. ISBN 3-319-14552-5 (paperback), 3-319-14553-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q125 .R45 2015.

URL https://link.springer.com/
chapter/10.1007/978-3-319-14553-
2_6.

Nieto-Galan:2015:RRM

[NG15b]  Agustí Nieto-Galan.  Reform and

repression: Manuel Lora-Tamayo and the Spanish university in the 1960s. In Sciences in the Universities of Europe, Nineteenth and Twentieth Centuries: Academic Landscapes [SDG15b], chapter 10, pages 159–174. ISBN 94-017-9635-1, 94-017-9636-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q183.4.E85 2015.

URL https://link.springer.com/
chapter/10.1007/978-94-017-9636-
1_10.

Ng:2017:CRA

[Ng17] Margaret Ng. Critical rationalism applied: A footnote to the siblinghood of humanity. In Encouraging Openness: Essays for Joseph Agassi on the Occasion of His 90th Birthday [BAG17], chapter 20, pages 243–249. ISBN 3-319-57668-2 (hardcover), 3-319-57669-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN B833.E53 2017. URL https://link.

springer.com/chapter/10.1007/
978-3-319-57669-5_20.

Nianzu:1996:DMP

[Nia96] Dai Nianzu. The development of modern physics in China: The 50th anniversary of the founding of the Chinese physical society. In Chinese Studies in the History and Philosophy of Science and Technology [DC96], chapter 17, pages 207–218. ISBN 90-481-4546-5, 94-015-8717-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 180; U390 .N19 1996. URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-015-8717-4_17.

Nickles:1976:TGP

[Nic76] Thomas Nickles. Theory generalization problem reduction and the unity of science. In Cohen et al. [CHMV76], pages 33–75. ISBN 90-277-0647-6, 90-277-0648-4 (paperback), 94-010-1449-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-010-1449-
                                              6_3.

Nickles:1980:CSC

[Nic80a] Thomas Nickles. Can scientific constraints be violated rationally? In Scientific Discovery, Logic, and Rationality [Nic80f], chapter 13, pages 285–315. ISBN 90-277-1070-8, 94-009-8986-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

                                              ???? URL https://link.springer.
                                              com/chapter/10.1007/978-94-009-
                                              8986-3_13.

Nickles:1980:DWP

[Nic80b] Thomas Nickles. Discussion of Wartofsky’s paper. In Scientific Discovery: Case Studies [Nic80e], chapter 2, pages 17–20. ISBN 90-277-1093-7, 94-009-9015-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 60. URL https:/

                                              /link.springer.com/chapter/10.
                                              1007/978-94-009-9015-9_2.

Nickles:1980:IES

[Nic80c] Thomas Nickles. Introductory essay: Scientific discovery and the future of philosophy of science. In Scientific Discovery, Logic, and Rationality [Nic80f], chapter 1, pages 1–59. ISBN 90-277-1070-8, 94-009-8986-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-009-8986-3_1.

Nickles:1980:REH

[Nic80d] Thomas Nickles. The rational explanation of historical discoveries. In Scientific Discovery: Case Studies [Nic80e], chapter 3, pages 21–49. ISBN 90-277-1093-7, 94-009-9015-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 60.

URL https://link.springer.com/
chapter/10.1007/978-94-009-9015-
9_3.

Nickles:1980:SDC

[Nic80e] Thomas Nickles. Scientific Discovery: Case Studies, volume 60 of Boston Studies in the Philosophy of Science. D. Reidel, Dordrecht, The Netherlands; Boston, MA, USA; Lancaster, UK; Tokyo, Japan, 1980. ISBN 90-277-1093-7, 94-009-9015-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). xxvi + 382 + 1 pp. LCCN Q174 .B67 vol. 60. URL

https://link.springer.com/book/
10.1007/978-94-009-9015-9.

Nickles:1980:SDL

[Nic80f] Thomas Nickles. Scientific Discovery, Logic, and Rationality, volume 56 of Boston Studies in the Philosophy of Science. D. Reidel, Dordrecht, The Netherlands; Boston, MA, USA; Lancaster, UK; Tokyo, Japan, 1980. ISBN 90-277-1070-8, 94-009-8986-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). xii + 388 pp. LCCN ????

                                              URL https://link.springer.com/
                                              book/10.1007/978-94-009-8986-3.

Nickles:1987:MHR

[Nic87] Thomas Nickles. Methodology, heuristics, and rationality. In Pitt and Pera [PP87], chapter 5, pages 103–132. ISBN 90-277-2417-2, 94-009-3779-2. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol.

98. US$64.00. URL https://link.

                                              springer.com/chapter/10.1007/
                                              978-94-009-3779-6_5.

Nicholas:1989:RSB

[Nic89a] John M. Nicholas. Realism for shopkeepers; behaviouralist notes on constructive empiricism. In An Intimate Relation: Studies in the History and Philosophy of Science Presented to Robert E. Butts on his 60th Birthday [BM89], chapter 23, pages 459–476. ISBN 94-010-7546-8, 94-009-2327-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-009-2327-
                                              0_23.

Nicolacopoulos:1989:TLG

[Nic89b] Pantelis D. Nicolacopoulos. Through the looking glass: Philosophy, research programmes and the scientific community. In Imre Lakatos and Theories of Scientific Change [GGN89], chapter 15, pages 189–202. ISBN 94-010-7860-2, 94-009-3025-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

                                              ???? URL https://link.springer.
                                              com/chapter/10.1007/978-94-009-
                                              3025-4_15.

Nicolacopoulos:1990:GSP

[Nic90] Pantelis Nicolacopoulos. Greek Studies in the Philosophy and History of Science, volume 121 of Boston Studies in the Philosophy of Science. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 1990. ISBN 94-010-7403-8, 94-009-2015-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). xviii + 439 pp. LCCN ???? URL

https://link.springer.com/book/
10.1007/978-94-009-2015-6.

Nickles:2008:DSC

[Nic08] Thomas Nickles. Disruptive scientific change. In Rethinking Scientific Change and Theory Comparison: Stabilities, Ruptures, Incommensurabilities? [SSHH08], chapter 24, pages 351–379. ISBN 1-4020-6274-5 (hardcover), 1-4020-6279-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175 .R397 2008.

URL https://link.springer.com/
chapter/10.1007/978-1-4020-6279-
7_24.

Nickles:2012:DRD

[Nic12] Thomas Nickles. Dynamic robustness and design in nature and artifact. In Characterizing the Robustness of Science: After the Practice Turn in Philosophy of Science [STNW12], chapter 14, pages 329–360. ISBN 94-007-2758-5, 94-007-2759-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175 .C4595 2012.

URL https://link.springer.com/
chapter/10.1007/978-94-007-2759-
5_14.

Nicolai:2016:MST

[Nic16] Carlo Nicolai. More on systems of truth and predicative comprehension. In Objectivity, Realism, and Proof: FilMat Studies in the Philosophy of Mathematics [BS16], chapter 14, pages 265–290. ISBN 3-319-31642-7, 3-319-31644-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QA9 .O25 2016. URL https://link.

                                              springer.com/chapter/10.1007/
                                              978-3-319-31644-4_14.

Nicoglou:2017:TD

[Nic17] Antonine Nicoglou. The timing of development. In Time of Nature and the Nature of Time: Philosophical Perspectives of Time in Natural Sciences [BH17b], chapter 15, pages 359–390. ISBN 3-319-53723-7 (hardcover), 3-319-53725-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC173.59.S65 T56 2017. URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-3-319-53725-2_15.

Nicolai:2018:EPA

[Nic18] Carlo Nicolai. On expressive power over arithmetic. In Piazza and Pulcini [PP18], pages 39–55. ISBN 3-319-93341-8, 3-319-93342-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QA8.4. URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-3-319-93342-9_3.

Niebyl:1974:SCS

[Nie74] Karl H. Niebyl. Scientific concepts and social structure in Ancient Greece. In Seeger and Cohen [SC74b], pages 365– 381. ISBN 90-277-0390-6, 90-277-0376-0 (paperback), 94-010-2126-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 11.

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-010-2126-
                                              5_21.

Niedzwiedzki:1982:RIP

[Nie82] Wiktor Niedźwiedzki. Realism and instrumentalism: On a priori conditions of science. In Polish Essays in the Philosophy of the Natural Sciences [Kra82c], chapter 4, pages 43–50. ISBN 94-009-7705-0 (e-book), 90-277-1287-5. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-009-7705-1_4.

Niiniluoto:1989:CVS

[Nii89]  Ilkka Niiniluoto.  Corroboration,

verisimilitude, and the success of science. In Imre Lakatos and Theories of Scientific Change [GGN89], chapter 18, pages 229–243. ISBN 94-010-7860-2, 94-009-3025-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-009-
3025-4_18.

Niiniluoto:1996:QAI

[Nii96] Ilkka Niiniluoto. Queries about internal realism. In Cohen et al. [CHR96], pages 45–54. ISBN 90-481-4493-0, 94-015-8638-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-015-
8638-2_4.

Niiniluoto:2003:PLC

[Nii03]  Ilkka Niiniluoto.  Probability and

logic comments on Colin Howson. In Descartes’s Mathematical Thought [Sas03d], chapter 24, pages 333–337. ISBN 90-481-6487-7, 94-017-1225-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/0-306-48123-5_24.

Niiniluoto:2017:VWH

[Nii17] Ilkka Niiniluoto. Verisimilitude: Why and how. In Encouraging Openness: Essays for Joseph Agassi on the Occasion of His 90th Birthday [BAG17], chapter 7, pages 71–79. ISBN 3-319-57668-2 (hardcover), 3-319-57669-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN B833.E53 2017.

URL https://link.springer.com/
chapter/10.1007/978-3-319-57669-
5_7.

Nijhout:2015:DPP

[Nij15] H. Frederik Nijhout. A developmentalphysiological perspective on the development and evolution of phenotypic plasticity. In Conceptual Change in Biology: Scientific and Philosophical Perspectives on Evolution and Development [Lov15b], chapter 7, pages 147–173. ISBN 94-017-9411-1, 94-017-9412-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QH331 C66 2015; QH491; QH492 .C66

2014. URL https://link.springer.
com/chapter/10.1007/978-94-017-
9412-1_7.

Nikolic:1974:RBE

[Nik74] Milenko Nikolić. The relation between Eudoxus’ theory of proportions and Dedekind’s theory of cuts. In For Dirk Struik: Scientific, Historical and Political Essays in Honor of Dirk J. Struik [CSW74], chapter 19, pages 225–243. ISBN 90-277-0379-5, 94-010-2115-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-010-2115-
9_19.

Niklas:2015:AAD

[Nik15] Karl J. Niklas. Adaptive aspects of development: A 30-year perspective on the relevance of biomechanical and allometric analyses. In Conceptual Change in Biology: Scientific and Philosophical Perspectives on Evolution and Development [Lov15b], chapter 2, pages 57–76. ISBN 94-017-9411-1, 94-017-9412-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QH331 C66 2015; QH491; QH492 .C66

2014. URL https://link.springer.
com/chapter/10.1007/978-94-017-
9412-1_2.

Nickles:2002:JBE

[NMN02] Thomas Nickles and Gaye McCollum-Nickles. James on bootstraps, evolution, and life. In Hermeneutic Philosophy of Science, Van Gogh’s Eyes, and God: Essays in Honor of Patrick A. Heelan, S.J. [Bab02c], chapter 31, pages 361–376. ISBN 90-481-5926-1, 94-017-1767-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 225. URL https:/

/link.springer.com/chapter/10.
1007/978-94-017-1767-0_31.

Nunes-Neto:2011:FEB

[NNEH11] Nei Freitas Nunes-Neto and Charbel Niño El-Hani. Functional explanations in biology, ecology, and earth system science: Contributions from philosophy of biology. In Brazilian Studies in Philosophy and History of Science: An account of recent works [KV11], chapter 13, pages 185–199. ISBN 90-481-9421-0 (hardcover), 90-481-9422-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175.3 .B73 2011. URL https:/

/link.springer.com/chapter/10.
1007/978-90-481-9422-3_13.

Nelson:1974:SCE

[NNW74] Benjamin Nelson, Joseph Needham, and Max Weber. Sciences and civilizations, ‘East’ and ‘West’ Joseph. In Seeger and Cohen [SC74b], pages 445– 492. ISBN 90-277-0390-6, 90-277-0376-0 (paperback), 94-010-2126-0 (e-book).

ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 11.

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-010-2126-
                                              5_26.

Noe:1992:MRV

[Noé92] Keiichi Noé. Mach’s relativism versus apriorism and the mechanistic world view. In Ernst Mach — A Deeper Look: Documents and New Perspectives [Bla92a], chapter 12, pages 229–241. ISBN 94-010-5230-1, 94-011-2771-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN C5.53;

                                              QC6.9. URL https://link.springer.
                                              com/chapter/10.1007/978-94-011-
                                              2771-4_12.

Nolte:1991:NCP

[Nol91] Peter Nolte. A new correspondence of the philosopher F. W. J. Schelling. In Woodward and Cohen [WC91], pages 351–356. ISBN 94-010-5395-2, 94-011-3164-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

                                              ???? URL https://link.springer.
                                              com/chapter/10.1007/978-94-011-
                                              3164-3_32.

Nola:1999:NNS

[Nol99] Robert Nola. Nietzsche’s naturalism: Science and belief. In Nietzsche, Epistemology, and Philosophy of Science: Nietzsche and the Sciences II [Bab99a], chapter 7, pages 91–100. ISBN 90-481-5234-8, 94-017-2428-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-017-2428-9_7.

Nola:2003:ADT

[Nol03a] Robert Nola. An archaeological dig through Foucault’s texts. In Rescuing Reason: A Critique of Anti-Rationalist Views of Science and Knowledge [Nol03k], chapter 9, pages 365–416. ISBN 1-4020-1043-5, 94-010-0289-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 230; Q175.32.R45. URL https:

//link.springer.com/chapter/10.
1007/978-94-010-0289-9_9.

Nola:2003:CTS

[Nol03b] Robert Nola. The critical tradition and some of its discontents. In Rescuing Reason: A Critique of Anti-Rationalist Views of Science and Knowledge [Nol03k], chapter 2, pages 19–73. ISBN 1-4020-1043-5, 94-010-0289-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 230; Q175.32.R45. URL https://link.

springer.com/chapter/10.1007/
978-94-010-0289-9_2.

Nola:2003:ECS

[Nol03c] Robert Nola. The Edinburgh Connection I: The strong programme and the social causes of scientific belief. In Rescuing Reason: A Critique of Anti-Rationalist Views of Science and Knowledge [Nol03k], chapter 6, pages 205–260. ISBN 1-4020-1043-5, 94-010-0289-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 230; Q175.32.R45.

URL https://link.springer.com/
chapter/10.1007/978-94-010-0289-
9_6.

Nola:2003:ECI

[Nol03d] Robert Nola. The Edinburgh Connection II: Strong and wrong. In Rescuing Reason: A Critique of Anti-Rationalist Views of Science and Knowledge [Nol03k], chapter 7, pages 261–295. ISBN 1-4020-1043-5, 94-010-0289-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 230; Q175.32.R45.

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-010-0289-
                                              9_7.

Nola:2003:E

[Nol03e] Robert Nola. Epilogue. In Rescuing Reason: A Critique of Anti-Rationalist Views of Science and Knowledge [Nol03k], chapter 12, pages 539–542. ISBN 1-4020-1043-5, 94-010-0289-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 230; Q175.32.R45. URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-010-0289-9_12.

Nola:2003:GPK

[Nol03f] Robert Nola. Genealogy, power and knowledge. In Rescuing Reason: A Critique of Anti-Rationalist Views of Science and Knowledge [Nol03k], chapter 10, pages 417–460. ISBN 1-4020-1043-5, 94-010-0289-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 230; Q175.32.R45. URL https://link.

                                              springer.com/chapter/10.1007/
                                              978-94-010-0289-9_10.

Nola:2003:I

[Nol03g] Robert Nola. Introduction. In Rescuing Reason: A Critique of Anti-Rationalist Views of Science and Knowledge [Nol03k], chapter 1, pages 1–11. ISBN 1-4020-1043-5, 94-010-0289-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 230; Q175.32.R45. URL https://link.

                                              springer.com/chapter/10.1007/
                                              978-94-010-0289-9_1.

Nola:2003:NNR

[Nol03h] Robert Nola. Naturalism and norms of reason and method. In Rescuing Reason: A Critique of Anti-Rationalist Views of Science and Knowledge [Nol03k], chapter 4, pages 117–172. ISBN 1-4020-1043-5, 94-010-0289-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 230; Q175.32.R45. URL https:

//link.springer.com/chapter/10.
1007/978-94-010-0289-9_4.

Nola:2003:NGB

[Nol03i] Robert Nola. Nietzsche’s genealogy of belief and morality. In Rescuing Reason: A Critique of Anti-Rationalist Views of Science and Knowledge [Nol03k], chapter 11, pages 465–537. ISBN 1-4020-1043-5, 94-010-0289-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 230; Q175.32.R45. URL https://link.

springer.com/chapter/10.1007/
978-94-010-0289-9_11.

Nola:2003:PK

[Nol03j] Robert Nola. The problem of knowledge. In Rescuing Reason: A Critique of Anti-Rationalist Views of Science and Knowledge [Nol03k], chapter 3, pages 75–116. ISBN 1-4020-1043-5, 94-010-0289-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 230; Q175.32.R45.

URL https://link.springer.com/
chapter/10.1007/978-94-010-0289-
9_3.

Nola:2003:RRC

[Nol03k] Robert Nola. Rescuing Reason: A Critique of Anti-Rationalist Views of Science and Knowledge, volume 230 of Boston Studies in the Philosophy of Science. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 2003. ISBN 1-4020-1043-5, 94-010-0289-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). xi + 559 pp. LCCN Q174 .B67 vol. 230; Q175.32.R45. URL

                                              https://link.springer.com/book/
                                              10.1007/978-94-010-0289-9.

Nola:2003:SGC

[Nol03l] Robert Nola. Some German connections: Marx and Mannheim. In Rescuing Reason: A Critique of Anti-Rationalist Views of Science and Knowledge [Nol03k], chapter 5, pages 179–203. ISBN 1-4020-1043-5, 94-010-0289-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 230; Q175.32.R45.

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-010-0289-
                                              9_5.

Nola:2003:WCS

[Nol03m] Robert Nola. The Wittgenstein connection: The social and the rational. In Rescuing Reason: A Critique of Anti-Rationalist Views of Science and Knowledge [Nol03k], chapter 8, pages 297–360. ISBN 1-4020-1043-5, 94-010-0289-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 230; Q175.32.R45.

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-010-0289-
                                              9_8.

Nola:2008:OMI

[Nol08] Robert Nola. The optimistic metainduction and ontological continuity: the case of the electron. In Rethinking Scientific Change and Theory Comparison: Stabilities, Ruptures, Incommensurabilities? [SSHH08], chapter 12, pages 159–202. ISBN 1-4020-6274-5 (hardcover), 1-4020-6279-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175 .R397

                                              2008. URL https://link.springer.

com/chapter/10.1007/978-1-4020-
6279-7_12.

North:1981:SA

[Nor81] John D. North. Science and analogy. In Grmek et al. [GCC81b], pages 115–140. ISBN 90-277-1123-2, 94-010-1284-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-010-1284-
3_5.

North:1991:OTM

[Nor91] John D. North. One truth or more?

Demarcation in the universe of discourse. In Unguru [Ung91b], pages 253–293. ISBN 94-010-5476-2, 94-011-3342-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC5.53; QC6.9. URL https://link.

springer.com/chapter/10.1007/
978-94-011-3342-5_11.

Nordmann:1998:ECD

[Nor98a] Alfred Nordmann. “Everything could be different”: The principles of mechanics and the limits of physics. In Heinrich Hertz: Classical Physicist, Modern Philosopher [BHN98], chapter 10, pages 155–171. ISBN 90-481-4881-2, 94-015-8855-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-015-8855-3_10.

Nordmann:1998:HHB

[Nor98b] Alfred Nordmann. Heinrich Hertz — a bibliography. In Heinrich Hertz: Classical Physicist, Modern Philosopher [BHN98], chapter 16, pages 281–305. ISBN 90-481-4881-2, 94-015-8855-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-015-8855-
                                              3_16.

Norton:2003:SF

[Nor03] John D. Norton. The N-Stein family. In Renn et al. [RDS03], chapter 4, pages 55–67. ISBN 1-4020-1284-5, 1-4020-1285-3 (paperback). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 234. URL https:/

                                              /link.springer.com/chapter/10.
                                              1007/978-94-010-0111-3_4.

Nordmann:2006:MM

[Nor06] Alfred Nordmann. From metaphysics to metachemistry. In Philosophy of Chemistry: Synthesis of a New Discipline [BSM06b], chapter 19, pages 347–362. ISBN 1-4020-2455-X, 1-4020-3256-0, 1-4020-3261-7. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 242 QD6. URL https:/

                                              /link.springer.com/chapter/10.
                                              1007/1-4020-3261-7_19.

Nordstrom:2007:IGM

[Nor07a]  Gunnar Nordström.  Inertial and

gravitational mass in relativistic mechanics. In Renn et al. [RSS^{+}07a], pages 499–521. ISBN 1-4020-3999-9. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 no. 250 v. 3; QC173.6 .G469 2007. URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-1-4020-4000-9_29.

Nordstrom:2007:TGS

[Nor07b] Gunnar Nordström. On the theory of gravitation from the standpoint of the Principle of Relativity. In Renn et al. [RSS^{+}07a], pages 523–542. ISBN 1-4020-3999-9. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 no. 250 v. 3; QC173.6 .G469 2007.

URL https://link.springer.com/
chapter/10.1007/978-1-4020-4000-
9_30.

Nordstrom:2007:PRG

[Nor07c] Gunnar Nordström. The Principle of Relativity and gravitation. In Renn et al. [RSS^{+}07a], pages 489–497. ISBN 1-4020-3999-9. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 no. 250 v. 3; QC173.6 .G469 2007.

URL https://link.springer.com/
chapter/10.1007/978-1-4020-4000-
9_28.

Norton:2007:ENE

[Nor07d] John D. Norton. Einstein, Nordström, and the early demise of scalar, Lorentz covariant theories of gravitation. In Renn et al. [RSS^{+}07a], pages 413–487. ISBN 1-4020-3999-9. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 no. 250 v. 3; QC173.6 .G469

2007. URL https://link.springer.
com/chapter/10.1007/978-1-4020-
4000-9_27.

Norton:2007:WWE

[Nor07e] John D. Norton. What was Einstein’s “fateful prejudice”? In The Genesis of General Relativity: Volume 2. Einstein’s Zurich Notebook: Commentary and Essays [JRN^{+}07b], pages 715–783. ISBN 1-4020-3942-5. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 no. 250 v. 2; QC173.6 .G469

2007. URL https://link.springer.
com/chapter/10.1007/978-1-4020-
4000-9_7.

Nordmann:2011:SCT

[Nor11] Alfred Nordmann. Science in the context of technology. In Science in the Context of Application [CN11a], chapter 27, pages 467–482. ISBN 90-481-9050-9, 90-481-9051-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q180.55.M4 S385

2011. URL https://link.springer.
com/chapter/10.1007/978-90-481-
9051-5_27.

Northcott:2012:GTC

[Nor12] Robert Northcott. Genetic traits and causal explanation. In Philosophy of Behavioral Biology [PR12b], chapter 4, pages 65–82. ISBN 94-007-1950-7, 94-007-1951-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QH331 .P4679 2012; Q175 .B657 v.282.

URL https://link.springer.com/
chapter/10.1007/978-94-007-1951-
4_4.

Norton:2013:FVA

[Nor13] Bryan G. Norton. Facts, values, and analogies: A Darwinian approach to environmental choice. In Bergandi [Ber13c], chapter 5, pages 63–82. ISBN 94-007-5066-8, 94-007-5067-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN GE42 .S77 2013; Q175 .B73 v. 296. URL https:/

/link.springer.com/chapter/10.
1007/978-94-007-5067-8_5.

Noutsos:1990:HTN

[Nou90] Panayiotis Noutsos. The history of the theory of natural sciences: A paradigm. In Greek Studies in the Philosophy and History of Science [Nic90], chapter 13, pages 201–208. ISBN 94-010-7403-8, 94-009-2015-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-009-
2015-6_13.

Novakovic:1974:TOT

[Nov74] Stanisa Novakovic. Is the transition from an old theory to a new one of a sudden and unexpected character? In Methodological and Historical Essays in the Natural and Social Sciences [CW74b], chapter 9, pages 173–196. ISBN 90-277-0378-7, 94-010-2128-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 14. URL https://link.

springer.com/chapter/10.1007/
978-94-010-2128-9_9.

Nowak:1979:PEK

[Now79a] Leszek Nowak. The problem of explanation in Karl Marx’s Capital. In Polish Essays in the Methodology of the Social Sciences [Wia79b], chapter 4, pages 49–73. ISBN 90-277-0956-4, 94-009-9353-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .P655 1979. URL https:

//link.springer.com/chapter/10.
1007/978-94-009-9353-2_4.

Nowak:1979:MDC

[Now79b] Stefan Nowak. Methodological dilemmas of contemporary sociology. In Polish Essays in the Methodology of the Social Sciences [Wia79b], chapter 8, pages 157–171. ISBN 90-277-0956-4, 94-009-9353-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .P655 1979. URL https:

//link.springer.com/chapter/10.
1007/978-94-009-9353-2_8.

Nowakowa:1982:DCE

[Now82] Izabella Nowakowa. Dialectical correspondence and essential truth. In Polish Essays in the Philosophy of the Natural Sciences [Kra82c], chapter 12, pages 135–146. ISBN 94-009-7705-0 (e-book), 90-277-1287-5. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-009-7705-
1_12.

Nowotny:2001:HTS

[Now01] Helga Nowotny. How to tell a story in the sciences: Settings and lessons. In Allen and MacLeod [AM01b], chapter 8, pages 123–135. ISBN 1-4020-0495-8 (hardcover), 94-017-2956-5. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 228; Q175.55.

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-017-2956-
                                              7_8.

Napoletani:2017:FOB

[NPS17] Domenico Napoletani, Marco Panza, and Daniele C. Struppa. Forcing optimality and Brandt’s principle. In Mathematics as a Tool: Tracing New Roles of Mathematics in the Sciences [LC17b], chapter 13, pages 233–251. ISBN 3-319-54468-3 (hardcover), 3-319-54469-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN GE45.M37; GE45.M38; QA8.4.

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-3-319-54469-
                                              4_13.

Nye:1992:IPS

[NRS92] Mary Jo Nye, Joan L. Richards, and Roger H. Stuewer. The Invention of Physical Science: Intersections of Mathematics, Theology and Natural Philosophy Since the Seventeenth Century: Essays in Honor of Erwin N. Hiebert, volume 139 of Boston Studies in the Philosophy of Science. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 1992. ISBN 94-010-5097-X, 94-011-2488-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). xxxiv + 278 pp. LCCN ???? URL

                                              https://link.springer.com/book/
                                              10.1007/978-94-011-2488-1.

Nadel-Turonski:1982:PPM

[NT82] Tadeusz Nadel-Turoński. Principles of physics as metalaws. In Polish Essays in the Philosophy of the Natural Sciences [Kra82c], chapter 24, pages 277–285. ISBN 94-009-7705-0 (e-book), 90-277-1287-5. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-009-7705-
1_24.

Nye:1992:PCC

[Nye91] Mary Jo Nye. Physics and chemistry:

Commensurate or incommensurate sciences? In The Invention of Physical Science: Intersections of Mathematics, Theology and Natural Philosophy Since the Seventeenth Century: Essays in Honor of Erwin N. Hiebert [NRS92], chapter 9, pages 205–224. ISBN 94-010-5097-X, 94-011-2488-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-011-2488-1_9.

Nye:2001:PTM

[Nye01] Mary Jo Nye. Paper tools and molecular architecture in the chemistry of Linus Pauling. In Klein [Kle01b], chapter 8, pages 117–132. ISBN 1-4020-0100-2, 94-015-9737-5. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 222; QD7. URL https:

//link.springer.com/chapter/10.
1007/978-94-015-9737-1_8.

Nye:2007:PLH

[Nye07] Mary Jo Nye. Parallel lives and the history of science. In Positioning the history of science [RG07], chapter 20, pages 109–112. ISBN 1-4020-5419-X (hardcover), 1-4020-5420-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q125 .P77 2007.

                                              URL https://link.springer.com/
                                              chapter/10.1007/1-4020-5420-3_20.

Nye:2015:BHS

[Nye15] Mary Jo Nye. Biography and the history of science. In Arabatzis et al. [ARS15], chapter 19, pages 281–296. ISBN 3-319-14552-5 (paperback), 3-319-14553-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q125 .R45 2015. URL https://link.

                                              springer.com/chapter/10.1007/
                                              978-3-319-14553-2_19.

Nyiri:1997:HWA

[Nyí97] J. C. Nyíri. Haller on Wittgenstein on art. In Austrian Philosophy Past and Present: Essays in Honor of Rudolf Haller [LM97], chapter 2, pages 21–28. ISBN 94-010-6412-1, 94-011-5720-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-011-5720-
                                              9_2.

Onnes:1991:FELc

[OB91a] H. Kamerlingh Onnes and J. D. A. Boks. Further experiments with liquid helium. In Gavroglu and Goudaroulis [GG91], chapter 9, pages 195–200. ISBN 94-010-7433-X (hardcover), 94-009-2079-2. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 124. URL https://link.

                                              springer.com/chapter/10.1007/
                                              978-94-009-2079-8_9.

Onnes:1991:FELe

[OB91b] H. Kamerlingh Onnes and J. D. A. Boks. Further experiments with liquid helium. In Gavroglu and Goudaroulis [GG91], chapter 11, pages 221–226. ISBN 94-010-7433-X (hardcover), 94-009-2079-2. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 124. URL https://link.

springer.com/chapter/10.1007/
978-94-009-2079-8_11.

Oberman:1975:RRC

[Obe75] Heiko A. Oberman. Reformation and revolution: Copernicus’s discovery in an era of change. In Murdoch and Sylla [MS75b], pages 397–435. ISBN 90-277-0587-9, 94-010-1781-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN B721 .I584 1975.

URL https://link.springer.com/
chapter/10.1007/978-94-010-1781-
7_11.

Oberheim:2008:CNP

[Obe08]  Eric Oberheim.  Commentary on

“Nonsense and Paradigm Change”, by Aristides Baltas. In Rethinking Scientific Change and Theory Comparison: Stabilities, Ruptures, Incommensurabilities? [SSHH08], chapter 5, pages 71–83. ISBN 1-4020-6274-5 (hardcover), 1-4020-6279-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175 .R397 2008. URL https:

//link.springer.com/chapter/10.
1007/978-1-4020-6279-7_5.  See

[Bal08].

Obregon:1992:TNS

[Obr92] Diana Obregón. Trade and the natural sciences in the United States of Colombia. In Science and Empires: Historical Studies about Scientific Development and European Expansion [PJM92], chapter 16, pages 147–152. ISBN 0-7923-1518-9 (hardcover), 94-010-5145-3, 94-011-2594-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 136.

URL https://link.springer.com/
chapter/10.1007/978-94-011-2594-
9_16.

OConnor:1999:NES

[O’C99] Brian O’Connor. Nietzsche and Enlightenment science: A dialectical reading. In Nietzsche, Theories of Knowledge, and Critical Theory: Nietzsche and the Sciences I [Bab99b], chapter 17, pages 243–251. ISBN 90-481-5233-X, 94-017-2430-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-017-2430-2_17.

OConnor:2002:HAS

[O’C02] Tony O’Connor. Human agency and the social sciences: From contextual phenomenology to genealogy. In Hermeneutic Philosophy of Science, Van Gogh’s Eyes, and God: Essays in Honor of Patrick A. Heelan, S.J. [Bab02c], chapter 15, pages 187–194. ISBN 90-481-5926-1, 94-017-1767-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 225. URL https://link.

                                              springer.com/chapter/10.1007/
                                              978-94-017-1767-0_15.

Ochoa:1996:FIP

[Och96] Juan Carlos García-Bermejo Ochoa. Four instrumental proposals. In Munévar [Mun96], chapter 12, pages 243–257. ISBN 94-010-6622-1, 94-009-0305-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

                                              ???? URL https://link.springer.
                                              com/chapter/10.1007/978-94-009-
                                              0305-0_12.

ODonovan:2002:GRS

[O’D02] Leo J. O’Donovan. Getting at the rapture of seeing: Ellsworth Kelly and visual experience. In Hermeneutic Philosophy of Science, Van Gogh’s Eyes, and God: Essays in Honor of Patrick A. Heelan, S.J. [Bab02c], chapter 25, pages 299–304. ISBN 90-481-5926-1, 94-017-1767-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 225.

URL https://link.springer.com/
chapter/10.1007/978-94-017-1767-
0_25.

Oddie:1989:PIM

[Odd89a] Graham Oddie. Partial interpretation, meaning variance, and incommensurability. In Imre Lakatos and Theories of Scientific Change [GGN89], chapter 22, pages 305–322. ISBN 94-010-7860-2, 94-009-3025-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-009-
3025-4_22.

Oddie:1989:UT

[Odd89b] Graham Oddie. The unity of theories.

In Freedom and Rationality: Essays in Honor of John Watkins From his Colleagues and Friends [DJ89], chapter 19, pages 343–368. ISBN 94-010-7571-9, 94-009-2380-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-009-2380-5_19.

Oesper:1991:TSW

[Oes91] Peter Oesper. Things are seldom what they seem: The story of non-phosphorylating glycolysis. In Woodward and Cohen [WC91], pages 321–327. ISBN 94-010-5395-2, 94-011-3164-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-011-
3164-3_29.

OHear:2009:PIT

[O’H09]  Anthony O’Hear.  Popperian in-

dividualism today. In Rethinking Popper [PC09], chapter 16, pages 205–215. ISBN 1-4020-9337-3, 1-4020-9338-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN B1649.P64 R48 2009; Q174 .B67 2009.

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-1-4020-9338-
                                              8_16.

Ohmori:1998:DLS

[Ohm98] Shozo Ohmori. Double look: Science superposed on a perceptual world. In Japanese Studies in the Philosophy of Science [NC98], chapter 5, pages 67–86. ISBN 94-010-6176-9, 94-011-5175-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-011-5175-
                                              7_5.

Ohrstrom:1997:PIR

[Øhr97] Peter Øhrstrøm. A. N. Prior’s ideas on the relation between semantics and axiomatics for temporal logic. In Perspectives on Time [FSU97b], chapter 19, pages 443–457. ISBN 90-481-4774-3, 94-015-8875-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol.189.

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-015-8875-
                                              1_19.

Oide:1998:SSI

[Oid98] Akira Oide. Structure of statistical inference. In Japanese Studies in the Philosophy of Science [NC98], chapter 8, pages 115–139. ISBN 94-010-6176-9, 94-011-5175-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-011-5175-7_8.

Ones:2014:CUD

[OJ14] Isarelis Pérez Ones and Jorge Núñez Jover. Can universities develop advanced technology and solve social problems? In The History of Physics in Cuba [BRW14a], chapter 10, pages 269–276. ISBN 94-017-8040-4, 94-017-8041-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC5.53; QC6.9; QC9 C9H57 2014; Q124.6-127.2. URL https://link.

springer.com/chapter/10.1007/
978-94-017-8041-4_10.

Orenstein:2000:KLL

[OK00] Alex Orenstein and Petr Kotatko.

Knowledge, Language and Logic: Questions for Quine, volume 210 of Boston Studies in the Philosophy of Science. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 2000. ISBN 1-4020-0253-X, 94-011-3933-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). 444 (est.) pp. LCCN ????

URL https://link.springer.com/
book/10.1007/978-94-011-3933-5.

Okruhlik:1989:FAP

[Okr89] Kathleen Okruhlik. The foundation of all philosophy: Newton’s third rule. In An Intimate Relation: Studies in the History and Philosophy of Science Presented to Robert E. Butts on his 60th Birthday [BM89], chapter 6, pages 97–113. ISBN 94-010-7546-8, 94-009-2327-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-009-
2327-0_6.

Oliveri:2016:TVT

[Oli16] Gianluigi Oliveri. True v or not true v, that is the question. In Objectivity, Realism, and Proof: FilMat Studies in the Philosophy of Mathematics [BS16], chapter 9, pages 143–164. ISBN 3-319-31642-7, 3-319-31644-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QA9 .O25 2016.

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-3-319-31644-
                                              4_9.

Olszewski:1982:NTR

[Ols82] Eugeniusz Olszewski. The notion of technological research and its place among other informational activities. In Polish Essays in the Philosophy of the Natural Sciences [Kra82c], chapter 34, pages 407–412. ISBN 94-009-7705-0 (ebook), 90-277-1287-5. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

                                              ???? URL https://link.springer.
                                              com/chapter/10.1007/978-94-009-
                                              7705-1_34.

Olson:1987:NGE

[Ols87] Richard Olson. On the nature of God’s existence, wisdom and power: The interplay between organic and mechanistic imagery in Anglican natural theology — 1640–1740. In Burwick [Bur87b], chapter 1, pages 1–48. ISBN 90-277-2541-1, 94-009-3917-5. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 105. US$64.00(U.S.). URL https://link.

                                              springer.com/chapter/10.1007/
                                              978-94-009-3917-2_1.

Olson:1998:TRN

[Ols98] Alan M. Olson. Theological reflections on the nature of nature: Revolution, reformation, restoration. In Philosophies of Nature: The Human Dimension: In Celebration of Erazim Kohák [CT98], chapter 10, pages 139–149. ISBN 90-481-4859-6, 94-017-2614-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN BD581.

URL https://link.springer.com/
chapter/10.1007/978-94-017-2614-
6_10.

Olson:2005:KA

[Ols05] David R. Olson. Knowledge and its artifacts. In History of Science, History of Text [Che05a], chapter 11, pages 231–245. ISBN 1-281-33589-4, 1-4020-2320-0 (hardcover), 1-4020-2321-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 238 Q125. URL https:/

/link.springer.com/chapter/10.
1007/1-4020-2321-9_11.

Omodeo:2018:SPI

[Omo18a] Pietro D. Omodeo. The social position and intellectual identity of the Renaissance mathematician–physicist Giovanni Battista Benedetti: A case study in the socio-political history of mechanics. In Feldhay et al. [FRSV18a], chapter 7, pages 181–213. ISBN 3-319-90343-8 (hardcover), 3-319-90345-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QA802.

URL https://link.springer.com/
chapter/10.1007/978-3-319-90345-
3_7.

Omodeo:2018:SPC

[Omo18b] Pietro D. Omodeo. Socio-political coordinates of early–modern mechanics: A preliminary discussion. In Feldhay et al. [FRSV18a], chapter 3, pages 55– 78. ISBN 3-319-90343-8 (hardcover), 3-319-90345-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic).

LCCN QA802. URL https://link.

                                              springer.com/chapter/10.1007/
                                              978-3-319-90345-3_3.

Onnes:1991:EES

[Onn91a] H. Kamerlingh Onnes. Expression of the equation of state of gases and liquids by means of series.”. In Gavroglu and Goudaroulis [GG91], chapter 6, pages 146–163. ISBN 94-010-7433-X (hardcover), 94-009-2079-2. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 124. URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-009-2079-8_6.

Onnes:1991:FELd

[Onn91b] H. Kamerlingh Onnes. “Further experiments with liquid helium.”. In Gavroglu and Goudaroulis [GG91], chapter 10, pages 201–220. ISBN 94-010-7433-X (hardcover), 94-009-2079-2. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 124.

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-009-2079-
                                              8_10.

Onnes:1991:FELf

[Onn91c] H. Kamerlingh Onnes. Further experiments with liquid helium. B. On the change in the resistance of pure metals at very low temperatures, etc. III. The resistance of platinum at helium temperatures. In Gavroglu and Goudaroulis [GG91], chapter 14, pages 252–260. ISBN 94-010-7433-X (hardcover), 94-009-2079-2. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 124. URL https:/

                                              /link.springer.com/chapter/10.
                                              1007/978-94-009-2079-8_14.

Onnes:1991:FELg

[Onn91d] H. Kamerlingh Onnes. Further experiments with liquid helium. C. On the change of electric resistance of pure metals at very low temperatures etc. IV. The resistance of pure mercury at helium temperatures. In Gavroglu and Goudaroulis [GG91], chapter 15, pages 261–263. ISBN 94-010-7433-X (hardcover), 94-009-2079-2. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 124. URL https:/

/link.springer.com/chapter/10.
1007/978-94-009-2079-8_15.

Onnes:1991:FELh

[Onn91e] H. Kamerlingh Onnes. Further experiments with liquid helium. D. On the change of the electrical resistance of pure metals at very low temperatures, etc. V. The disappearance of the resistance of mercury. In Gavroglu and Goudaroulis [GG91], chapter 16, pages 264–266. ISBN 94-010-7433-X (hardcover), 94-009-2079-2. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 124. URL https:/

/link.springer.com/chapter/10.
1007/978-94-009-2079-8_16.

Onnes:1991:FELb

[Onn91f] H. Kamerlingh Onnes. “Further experiments with liquid helium. E. A helium-cryostat. Remarks on the preceding communications.”. In Gavroglu and Goudaroulis [GG91], chapter 8, pages 188–194. ISBN 94-010-7433-X (hardcover), 94-009-2079-2. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 124. URL https:

//link.springer.com/chapter/10.
1007/978-94-009-2079-8_8.

Onnes:1991:FELi

[Onn91g] H. Kamerlingh Onnes. Further experiments with liquid helium. G. On the electrical resistance of pure metals, etc. VI. On the sudden change in the rate at which the resistance of mercury disappears. In Gavroglu and Goudaroulis [GG91], chapter 17, pages 267–272. ISBN 94-010-7433-X (hardcover), 94-009-2079-2. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 124. URL https:/

                                              /link.springer.com/chapter/10.
                                              1007/978-94-009-2079-8_17.

Onnes:1991:FELk

[Onn91h] H. Kamerlingh Onnes. Further experiments with liquid helium. H. On the electrical resistance etc. (continued). VIII. The sudden disappearance of the ordinary resistance of tin, and the super-conductive state of lead. In Gavroglu and Goudaroulis [GG91], chapter 19, pages 315–332. ISBN 94-010-7433-X (hardcover), 94-009-2079-2. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 124.

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-009-2079-
                                              8_19.

Onnes:1991:FELj

[Onn91i] H. Kamerlingh Onnes. Further experiments with liquid helium. H. On the electrical resistance of pure metals etc. VII. The potential difference necessary for the electric current through mercury below 4.19^{∘} K. In Gavroglu and Goudaroulis [GG91], chapter 18, pages 273–314. ISBN 94-010-7433-X (hardcover), 94-009-2079-2. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 124. URL https:/

                                              /link.springer.com/chapter/10.
                                              1007/978-94-009-2079-8_18.

Onnes:1991:FELl

[Onn91j] H. Kamerlingh Onnes. Further experiments with liquid helium. I. The Hall-effect, and the magnetic change in resistance at low temperatures. IX. The appearance of galvanic resistance in supra-conductors which are brought into a magnetic field, at a threshold value of the field. In Gavroglu and Goudaroulis [GG91], chapter 20, pages 333–339. ISBN 94-010-7433-X (hardcover), 94-009-2079-2. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 124. URL https:/

/link.springer.com/chapter/10.
1007/978-94-009-2079-8_20.

Onnes:1991:FELm

[Onn91k] H. Kamerlingh Onnes. Further experiments with liquid helium. J. The imitation of an Ampère molecular current or of a permanent magnet by means of a supra-conductor. In Gavroglu and Goudaroulis [GG91], chapter 21, pages 340–355. ISBN 94-010-7433-X (hardcover), 94-009-2079-2. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 124.

URL https://link.springer.com/
chapter/10.1007/978-94-009-2079-
8_21.

Onnes:1991:FELn

[Onn91l] H. Kamerlingh Onnes. Further experiments with liquid helium. L. The persistence of currents without electromotive force in supra-conducting circuits. In Gavroglu and Goudaroulis [GG91], chapter 22, pages 356–362. ISBN 94-010-7433-X (hardcover), 94-009-2079-2. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 124. URL https://link.

springer.com/chapter/10.1007/
978-94-009-2079-8_22.

Onnes:1991:FELa

[Onn91m] H. Kamerlingh Onnes. Further experiments with liquid helium. P. On the lowest temperature yet obtained. In Gavroglu and Goudaroulis [GG91], chapter 3, pages 59–88. ISBN 94-010-7433-X (hardcover), 94-009-2079-2. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 124.

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-009-2079-
                                              8_3.

Onnes:1991:IAM

[Onn91n] H. Kamerlingh Onnes. The importance of accurate measurements at very low temperatures. In Gavroglu and Goudaroulis [GG91], chapter 2, pages 31–58. ISBN 94-010-7433-X (hardcover), 94-009-2079-2. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 124. URL https:/

                                              /link.springer.com/chapter/10.
                                              1007/978-94-009-2079-8_2.

Onnes:1991:CLL

[Onn91o] H. Kamerlingh Onnes. On the Cryogenic Laboratory at Leiden and on the production of very low temperatures. In Gavroglu and Goudaroulis [GG91], chapter 1, pages 3–30. ISBN 94-010-7433-X (hardcover), 94-009-2079-2. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 124.

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-009-2079-
                                              8_1.

Onnes:1991:RLH

[Onn91p] H. Kamerlingh Onnes. Remarks on the liquefaction of hydrogen, on thermodynamical similarity and on the use of vacuum vessels. In Gavroglu and Goudaroulis [GG91], chapter 5, pages 125–145. ISBN 94-010-7433-X (hardcover), 94-009-2079-2. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 124. URL https:/

                                              /link.springer.com/chapter/10.
                                              1007/978-94-009-2079-8_5.

Onnes:1991:LH

[Onn91q] H. Kamerlingh Onnes. “The liquefaction of helium.”. In Gavroglu and Goudaroulis [GG91], chapter 7, pages 164–187. ISBN 94-010-7433-X (hardcover), 94-009-2079-2. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 124. URL https:/

/link.springer.com/chapter/10.
1007/978-94-009-2079-8_7.

Onnes:1991:TGE

[Onn91r] H. Kamerlingh Onnes. Théorie générale de l’état fluide. In Gavroglu and Goudaroulis [GG91], chapter 4, pages 89–124. ISBN 94-010-7433-X (hardcover), 94-009-2079-2. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 124. URL https:/

/link.springer.com/chapter/10.
1007/978-94-009-2079-8_4.

Olczyk:1982:DAC

[OP82] Sławoj Olczyk and Maciej Przanowski.

Difficulties with absolutism: The case of Von Weizsäcker’s philosophy. In Polish Essays in the Philosophy of the Natural Sciences [Kra82c], chapter 35, pages 413–430. ISBN 94-009-7705-0 (ebook), 90-277-1287-5. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-009-
7705-1_35.

Onnes:1991:RMI

[OP91a] H. Kamerlingh Onnes and Albert Perrier. Researches on magnetism. III. On para- and dia-magnetism at very low temperatures. In Gavroglu and Goudaroulis [GG91], chapter 26, pages 464–471. ISBN 94-010-7433-X (hardcover), 94-009-2079-2. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 124. URL https:/

/link.springer.com/chapter/10.
1007/978-94-009-2079-8_26.

Onnes:1991:RML

[OP91b] H. Kamerlingh Onnes and Albert Perrier. “Researches on the magnetization of liquid and solid oxygen”. In Gavroglu and Goudaroulis [GG91], chapter 25, pages 422–463. ISBN 94-010-7433-X (hardcover), 94-009-2079-2. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 124.

URL https://link.springer.com/
chapter/10.1007/978-94-009-2079-
8_25.

Otte:1997:ASM

[OP97] Michael Otte and Marco Panza, editors. Analysis and Synthesis in Mathematics: History and Philosophy, volume 196 of Boston Studies in the Philosophy of Science. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 1997. ISBN 0-7923-4570-3. xiii + 440 pp. LCCN Q174 .B67 vol. 196 QA300.

Ordonez:1996:SND

[Ord96]  Javier Ordóñez.  The story of a

non-discovery: Helmholtz and the conservation of energy. In Munévar [Mun96], chapter 1, pages 1–18. ISBN 94-010-6622-1, 94-009-0305-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-009-0305-0_1.

Orenstein:2000:PBQ

[Ore00] Alex Orenstein. Plato’s beard, Quine’s stubble and Ockham’s razor. In Knowledge, Language and Logic: Questions for Quine [OK00], chapter 15, pages 195–212. ISBN 1-4020-0253-X, 94-011-3933-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-011-
3933-5_15.

Orlean:1992:OM

[Orl92] André Orléan. The origin of money. In Understanding Origins: Contemporary Views on the Origin of Life, Mind and Society [VD92], chapter 6, pages 113–143. ISBN 90-481-4090-0, 94-015-8054-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-015-
8054-0_6.

Oruka:1995:CFP

[Oru95] H. Odera Oruka. Cultural fundamentals in philosophy obstacles in philosophical dialogues. In The Concept of Knowledge: The Ankara Seminar [KC95], chapter 13, pages 167–181. ISBN 90-481-4495-7, 94-017-3263-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-017-3263-5_13.

Orvell:1989:LAT

[Orv89] Miles Orvell. Literature and the authority of technology. In Literature and Science as Modes of Expression [Amr89], chapter 10, pages 169–175. ISBN 94-010-7531-X, 94-009-2297-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-009-2297-
6_10.

Odling-Smee:2009:NCE

[OS09] John Odling-Smee. Niche construction in evolution, ecosystems and developmental biology. In Mapping the Future of Biology: Evolving Concepts and Theories [BMP09b], chapter 6, pages 69–92. ISBN 1-4020-9635-6 (hardcover), 1-4020-9636-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QH331 .M295 2009.

URL https://link.springer.com/
chapter/10.1007/978-1-4020-9636-
5_6.

Osborne:1992:SZA

[Osb92] Michael A. Osborne. The Société Zoologique d’Acclimatation and the New French Empire: Science and political economy. In Science and Empires: Historical Studies about Scientific Development and European Expansion [PJM92], chapter 30, pages 299–306. ISBN 0-7923-1518-9 (hardcover), 94-010-5145-3, 94-011-2594-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 136.

URL https://link.springer.com/
chapter/10.1007/978-94-011-2594-
9_30.

Ostien:1976:OCT

[Ost76] Philip A. Ostien. Observationality and the comparability of theories. In Cohen et al. [CHMV76], pages 271–289. ISBN 90-277-0647-6, 90-277-0648-4 (paperback), 94-010-1449-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-010-1449-6_12.

Ostwald:1992:MC

[Ost92] Wilhelm Ostwald. Mach correspondence. In Ernst Mach — A Deeper Look: Documents and New Perspectives [Bla92a], chapter 4, pages 71–110. ISBN 94-010-5230-1, 94-011-2771-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN C5.53; QC6.9.

URL https://link.springer.com/
chapter/10.1007/978-94-011-2771-
4_4.

OToole:1988:OWP

[O’T88] Roger O’Toole. Organizational weapons and political sects. In Philosophy, History and Social Action: Essays in Honor of Lewis Feuer with an autobiographic essay by Lewis Feuer [HOO88c], chapter 17, pages 325–348. ISBN 94-010-7793-2, 94-009-2873-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-009-2873-
2_17.

Otte:1994:HTS

[Ott94] Michael Otte. Historiographical trends in the social history of mathematics and science. In Trends in the Historiography of Science [GCN94], chapter 22, pages 295–315. ISBN 90-481-4264-4, 94-017-3596-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-017-
3596-4_22.

Owen:1999:SVA

[Owe99] David Owen. Science, value, and the ascetic ideal. In Nietzsche, Epistemology, and Philosophy of Science: Nietzsche and the Sciences II [Bab99a], chapter 14, pages 169–177. ISBN 90-481-5234-8, 94-017-2428-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-017-2428-9_14.

Oyama:1992:PRO

[Oya92] Susan Oyama. Is phylogeny recapitulating ontogeny? In Understanding Origins: Contemporary Views on the Origin of Life, Mind and Society [VD92], chapter 12, pages 227–232. ISBN 90-481-4090-0, 94-015-8054-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-015-8054-
0_12.

Oyama:2009:CPM

[Oya09] Susan Oyama. Compromising positions: The minding of matter. In Mapping the Future of Biology: Evolving Concepts and Theories [BMP09b], chapter 3, pages 27–46. ISBN 1-4020-9635-6 (hardcover), 1-4020-9636-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QH331 .M295 2009. URL https:/

/link.springer.com/chapter/10.
1007/978-1-4020-9636-5_3.

Padovani:2013:GTT

[Pad11] Flavia Padovani. Genidentity and topology of time: Kurt Lewin and Hans Reichenbach. In The Berlin Group and the Philosophy of Logical Empiricism [MP13], chapter 5, pages 97–122. ISBN 94-007-5484-1 (hardcover), 94-007-5485-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175.32.A24 .M384 2011; Q175 .B47

2013. URL https://link.springer.
com/chapter/10.1007/978-94-007-
5485-0_5.

Page:1996:ATC

[Pag96] Don N. Page. Attaching theories of consciousness to Bohmian quantum mechanics. In Bohmian Mechanics and Quantum Theory: an Appraisal [CFG96], chapter 14, pages 197–210. ISBN 90-481-4698-4, 94-015-8715-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC173.96-174.52. URL https://link.

springer.com/chapter/10.1007/
978-94-015-8715-0_14.

Pagin:2000:PI

[Pag00] Peter Pagin. Publicness and indeterminacy. In Knowledge, Language and Logic: Questions for Quine [OK00], chapter 13, pages 163–180. ISBN 1-4020-0253-X, 94-011-3933-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-011-3933-5_13.

Pais:1995:PW

[Pai95] Abraham Pais. The power of the word.

In No Truth Except in the Details: Essays in Honor of Martin J. Klein [KS95], chapter 14, pages 319–332. ISBN 94-010-4097-4, 94-011-0217-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-011-0217-
9_14.

Palter:1976:PSH

[Pal76] Robert Palter. Philosophy of science, history of science, and science education. In Cohen et al. [CHMV76], pages 313–321. ISBN 90-277-0647-6, 90-277-0648-4 (paperback), 94-010-1449-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-010-1449-
6_15.

Palm:2001:SFC

[Pal01]  Viktor Palm.  Some fundamental

criteria of the scientific method and the internal structure of science. In Estonian Studies in the History and Philosophy of Science [Vih01b], chapter 7, pages 91–110. ISBN 94-010-3863-5, 94-010-0672-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-010-0672-9_7.

Pallo:2003:ANC

[Pal03] Gábor Palló. Accommodation to a new centre. In Simões et al. [SCD03b], chapter 13, pages 329–344. ISBN 1-4020-1259-4, 94-017-3584-0. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 v. 233.

URL https://link.springer.com/
chapter/10.1007/978-94-017-3584-
1_13.

Palmerino:2004:GTF

[Pal04a] Carla Rita Palmerino. Galileo’s theories of free fall and projectile motion as interpreted by Pierre Gassendi. In The Reception of the Galilean Science of Motion in Seventeenth-Century Europe [PT04], chapter 8, pages 137–164. ISBN 1-4020-2454-1, 1-4020-2455-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC133 .R43 2004.

URL https://link.springer.com/
chapter/10.1007/978-1-4020-2455-
9_8.

Palmerino:2004:I

[Pal04b] Carla Rita Palmerino. Introduction. In The Reception of the Galilean Science of Motion in Seventeenth-Century Europe [PT04], chapter 1, pages 1–9. ISBN 1-4020-2454-1, 1-4020-2455-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC133 .R43

2004. URL https://link.springer.
com/chapter/10.1007/978-1-4020-
2455-9_1.

Palmerino:2012:CST

[Pal12] Carla Rita Palmerino. The composition of space, time and matter according to Isaac Newton and John Keill. In Philosophy of Behavioral Biology [PR12b], chapter 5, pages 117–142. ISBN 94-007-1950-7, 94-007-1951-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QH331 .P4679 2012; Q175 .B657 v.282. URL https:

//link.springer.com/chapter/10.
1007/978-94-007-4345-8_5.

Palmerino:2013:CST

[Pal13] Carla Rita Palmerino. The composition of space, time and matter according to Isaac Newton and John Keill. In Garber and Roux [GR13b], chapter 5, pages 117–142. CODEN ???? ISBN 94-007-4344-0 (hardcover), 94-007-4345-9 (e-book). ISSN ???? LCCN QC6 .M3616 2013; QC6 .M43 2013.

URL https://link.springer.com/
chapter/10.1007/978-94-007-4345-
8_5.

Pallo:2015:UMC

[Pal15] Gabor Palló. University models in changing political contexts. In Sciences in the Universities of Europe, Nineteenth and Twentieth Centuries: Academic Landscapes [SDG15b], chapter 8, pages 127–144. ISBN 94-017-9635-1, 94-017-9636-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q183.4.E85 2015.

URL https://link.springer.com/
chapter/10.1007/978-94-017-9636-
1_8.

Pandit:1983:CAM

[Pan83a] G. L. Pandit. Consequences and alternative methodologies. In The Structure and Growth of Scientific Knowledge: A Study in the Methodology of Epistemic Appraisal [Pan83f], chapter 4, pages 111–143. ISBN 90-481-8375-8, 94-015-7680-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 73. URL https:/

/link.springer.com/chapter/10.
1007/978-94-015-7680-2_4.

Pandit:1983:ESL

[Pan83b] G. L. Pandit. Epistemic structuralism: The limit to radical alternatives to traditional epistemology. In The Structure and Growth of Scientific Knowledge: A Study in the Methodology of Epistemic Appraisal [Pan83f], chapter 2, pages 16–39. ISBN 90-481-8375-8, 94-015-7680-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 73. URL https:/

/link.springer.com/chapter/10.
1007/978-94-015-7680-2_2.

Pandit:1983:IMI

[Pan83c] G. L. Pandit. Introduction: Methodology, ideology, and scientific revolutions. In The Structure and Growth of Scientific Knowledge: A Study in the Methodology of Epistemic Appraisal [Pan83f], chapter 1, pages 1–15. ISBN 90-481-8375-8, 94-015-7680-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 73.

URL https://link.springer.com/
chapter/10.1007/978-94-015-7680-
2_1.

Pandit:1983:NMV

[Pan83d] G. L. Pandit. The nature of methodological variance: From commensurable canons to incommensurable strategies. In The Structure and Growth of Scientific Knowledge: A Study in the Methodology of Epistemic Appraisal [Pan83f], chapter 5, pages 144–183. ISBN 90-481-8375-8, 94-015-7680-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 73. URL https://link.

springer.com/chapter/10.1007/
978-94-015-7680-2_5.

Pandit:1983:PSG

[Pan83e] G. L. Pandit. Problems of structure and growth: Towards an interactive model of the growth of scientific knowledge. In The Structure and Growth of Scientific Knowledge: A Study in the Methodology of Epistemic Appraisal [Pan83f], chapter 3, pages 40–110. ISBN 90-481-8375-8, 94-015-7680-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 73.

URL https://link.springer.com/
chapter/10.1007/978-94-015-7680-
2_3.

Pandit:1983:SGS

[Pan83f] G. L. Pandit. The Structure and Growth of Scientific Knowledge: A Study in the Methodology of Epistemic Appraisal, volume 73 of Boston Studies in the Philosophy of Science. D. Reidel, Dordrecht, The Netherlands; Boston, MA, USA; Lancaster, UK; Tokyo, Japan, 1983. ISBN 90-481-8375-8, 94-015-7680-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). xxii + 231 pp. LCCN Q174 .B67 vol. 73. URL

https://link.springer.com/book/
10.1007/978-94-015-7680-2.

Panova:1984:SLS

[Pan84] Elena Panova. Substance and its logical significance. In Methodology, Metaphysics and the History of Science: In Memory of Benjamin Nelson [CW84b], chapter 9, pages 235–245. ISBN 94-009-6333-5, 94-009-6331-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-009-6331-3_9.

Pandit:1991:E

[Pan91a] G. L. Pandit. Epilogue. In Methodological Variance: Essays in Epistemological Ontology and the Methodology of Science [Pan91f], chapter 10, pages 327–362. ISBN 94-010-5400-2, 94-011-3174-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

                                              ???? URL https://link.springer.
                                              com/chapter/10.1007/978-94-011-
                                              3174-2_10.

Pandit:1991:FMI

[Pan91b] G. L. Pandit. Falsifiability and methodological invariance in science. In Methodological Variance: Essays in Epistemological Ontology and the Methodology of Science [Pan91f], chapter 7, pages 198–213. ISBN 94-010-5400-2, 94-011-3174-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-011-3174-2_7.

Pandit:1991:HKH

[Pan91c] G. L. Pandit. Human knowledge and human interaction. In Methodological Variance: Essays in Epistemological Ontology and the Methodology of Science [Pan91f], chapter 2, pages 18–54. ISBN 94-010-5400-2, 94-011-3174-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-011-3174-
                                              2_2.

Pandit:1991:ITN

[Pan91d] G. L. Pandit. Indeterminacy of translation: A non–Quinean function of content-indeterminacy. In Methodological Variance: Essays in Epistemological Ontology and the Methodology of Science [Pan91f], chapter 3, pages 55–80.

ISBN 94-010-5400-2, 94-011-3174-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-011-3174-
2_3.

Pandit:1991:MES

[Pan91e] G. L. Pandit. Methodological essentialism in science and in philosophy. In Methodological Variance: Essays in Epistemological Ontology and the Methodology of Science [Pan91f], chapter 5, pages 99–150. ISBN 94-010-5400-2, 94-011-3174-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-011-3174-2_5.

Pandit:1991:MVE

[Pan91f] G. L. Pandit. Methodological Variance:

Essays in Epistemological Ontology and the Methodology of Science, volume 131 of Boston Studies in the Philosophy of Science. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 1991. ISBN 94-010-5400-2, 94-011-3174-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). xxiv + 431 pp. LCCN ????

URL https://link.springer.com/
book/10.1007/978-94-011-3174-2.

Pandit:1991:MTP

[Pan91g] G. L. Pandit. The methodology of theory–problem interactive systems. In Methodological Variance: Essays in Epistemological Ontology and the Methodology of Science [Pan91f], chapter 8, pages 214–264. ISBN 94-010-5400-2, 94-011-3174-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-011-3174-2_8.

Pandit:1991:VIS

[Pan91h] G. L. Pandit. Of variance and invariance in science: Empirical science as an enterprise comprising NFCPS systems. In Methodological Variance: Essays in Epistemological Ontology and the Methodology of Science [Pan91f], chapter 6, pages 151–197. ISBN 94-010-5400-2, 94-011-3174-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-011-3174-2_6.

Pandit:1991:IEC

[Pan91i] G. L. Pandit. On the impossibility of any enterprise concerning self-knowledge within traditional epistemology. In Methodological Variance: Essays in Epistemological Ontology and the Methodology of Science [Pan91f], chapter 4, pages 81–96. ISBN 94-010-5400-2, 94-011-3174-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-011-3174-2_4.

Pandit:1991:OOS

[Pan91j] G. L. Pandit. On the objects of our subjective knowledge. In Methodological Variance: Essays in Epistemological Ontology and the Methodology of Science [Pan91f], chapter 1, pages 3–17. ISBN 94-010-5400-2, 94-011-3174-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-011-3174-
                                              2_1.

Pandit:1991:RPS

[Pan91k] G. L. Pandit. The resolving power of a scientific theory as a basis of its epistemic appraisal. In Methodological Variance: Essays in Epistemological Ontology and the Methodology of Science [Pan91f], chapter 9, pages 265–326. ISBN 94-010-5400-2, 94-011-3174-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-011-
3174-2_9.

Pancaldi:1994:TNM

[Pan94] Giuliano Pancaldi. The technology of nature: Marx’s thoughts on Darwin. In Cohen [Coh94b], chapter 7, pages 257–274. ISBN 0-7923-2223-1, 94-017-3391-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 150. URL https:/

/link.springer.com/chapter/10.
1007/978-94-017-3391-5_7.

Panaccio:1996:BSO

[Pan96]  Claude Panaccio.  Belief-sentences:

Outline of a nominalistic approach. In Québec Studies in the Philosophy of Science: Part II: Biology, Psychology, Cognitive Science and Economics Essays in Honor of Hugues Leblanc [MC96], chapter 16, pages 265–277. ISBN 94-010-6537-3, 94-009-0113-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-009-0113-
1_16.

Pantsar:2016:MSC

[Pan16] Markus Pantsar. The modal status of contextually a priori arithmetical truths. In Objectivity, Realism, and Proof: FilMat Studies in the Philosophy of Mathematics [BS16], chapter 5, pages 67–79. ISBN 3-319-31642-7, 3-319-31644-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic).

LCCN QA9 .O25 2016. URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-3-319-31644-4_5.

Paolucci:1984:HCM

[Pao84] Henry Paolucci. Hegel and the celestial mechanics of Newton and Einstein. In Hegel and the Sciences [CW84a], chapter 5, pages 55–85. ISBN 94-009-6235-5, 94-009-6233-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-009-6233-0_5.

Papapetrou:1974:GRS

[Pap74a] A. Papapetrou. General relativity — some puzzling questions. In Logical and Epistemological Studies in Contemporary Physics [CW74a], chapter 14, pages 376–387. ISBN 90-277-0391-4 (hardcover), 90-277-0377-9 (paperback), 94-010-2656-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol.

                                              13; QC5.56. URL https://link.
                                              springer.com/chapter/10.1007/
                                              978-94-010-2656-7_14.

Pappenheim:1974:RVQ

[Pap74b] Fritz Pappenheim. Relative values and the quest for socio-political standards. In For Dirk Struik: Scientific, Historical and Political Essays in Honor of Dirk J. Struik [CSW74], chapter 43, pages 575–580. ISBN 90-277-0379-5, 94-010-2115-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

                                              ???? URL https://link.springer.
                                              com/chapter/10.1007/978-94-010-
                                              2115-9_43.

Papineau:1989:PBG

[Pap89] David Papineau. Has Popper been a good thing? In Imre Lakatos and Theories of Scientific Change [GGN89], chapter 31, pages 431–440. ISBN 94-010-7860-2, 94-009-3025-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-009-3025-4_31.

Papagounos:1990:TT

[Pap90] G. Papagounos. On theoretical terms.

In Greek Studies in the Philosophy and History of Science [Nic90], chapter 22, pages 353–360. ISBN 94-010-7403-8, 94-009-2015-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-009-
2015-6_22.

Papathanassiou:1994:AGD

[Pap94] Maria K. Papathanassiou. Archaeoastronomy in Greece: Data, problems and perspectives. In Trends in the Historiography of Science [GCN94], chapter 32, pages 433–443. ISBN 90-481-4264-4, 94-017-3596-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-017-3596-4_32.

Papineau:2003:CGG

[Pap03] David Papineau. Comments on Gerd Gigerenzer. In Descartes’s Mathematical Thought [Sas03d], chapter 8, pages 141–151. ISBN 90-481-6487-7, 94-017-1225-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/0-306-48123-
5_8.

Papanelopoulou:2014:FTF

[Pap14] Faidra Papanelopoulou. ‘Fresher than fresh’. Consumer attitudes towards the development of the cold chain in Post-2WW Greece. In History of Artificial Cold, Scientific, Technological and Cultural Issues [Gav14b], chapter 14, pages 267–280. ISBN 94-007-7198-3 (hardcover), 94-007-7199-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC278 H57 2014; TP494 .H577 2014. URL https:/

/link.springer.com/chapter/10.
1007/978-94-007-7199-4_14.

Papanelopoulou:2015:LPC

[Pap15] Faidra Papanelopoulou. Louis Paul Cailletet, the liquefaction of oxygen and the emergence of an ‘in-between discipline’: Low-temperature research. In Arabatzis et al. [ARS15], chapter 2, pages 9–22. ISBN 3-319-14552-5 (paperback), 3-319-14553-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q125 .R45 2015.

URL https://link.springer.com/
chapter/10.1007/978-3-319-14553-
2_2.

Parikh:1983:PVP

[Par83] Rohit Parikh. The problem of vague predicates. In Language, Logic and Method [CW83b], chapter 13, pages 241–261. ISBN 94-009-7704-2, 94-009-7702-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-009-
7702-0_13.

Parsons:2000:IIO

[Par00] Terence Parsons. Indeterminacy of identity of objects: An exercise in metaphysical aesthetics. In Knowledge, Language and Logic: Questions for Quine [OK00], chapter 16, pages 213–224. ISBN 1-4020-0253-X, 94-011-3933-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-011-
3933-5_16.

Parve:2001:VNP

[Par01] Valdar Parve. Value-neutral paternalism. In Estonian Studies in the History and Philosophy of Science [Vih01b], chapter 20, pages 271–282. ISBN 94-010-3863-5, 94-010-0672-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-010-0672-9_20.

Parker:2009:DRR

[Par09a] Wendy Parker. Distinguishing real results from instrumental artifacts: The case of the missing rain. In Going Amiss In Experimental Research [HSS09a], chapter 10, pages 161–177. ISBN 1-4020-8892-2 (hardcover), 1-4020-8893-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q172.5.E77 G647 2009. URL https:

//link.springer.com/chapter/10.
1007/978-1-4020-8893-3_10.

Parrochia:2009:FPT

[Par09b] Daniel Parrochia. French philosophy of technology. In Brenner and Gayon [BG09], chapter 2, pages 51–70. ISBN 1-4020-9367-5, 1-4020-9368-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .F746 2009.

URL https://link.springer.com/
chapter/10.1007/978-1-4020-9368-
5_2.

Parusnikova:2009:RNP

[Par09c] Zuzana Parusniková. Ratio Negativa — the Popperian challenge. In Rethinking Popper [PC09], chapter 3, pages 31–45. ISBN 1-4020-9337-3, 1-4020-9338-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN B1649.P64 R48 2009; Q174 .B67 2009.

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-1-4020-9338-
                                              8_3.

Paraoanu:2015:QV

[Par15a] Gheorghe S. Paraoanu. The quantum vacuum. In Romanian Studies in Philosophy of Science [PST15], chapter 12, pages 181–197. ISBN 3-319-16654-9 (hardcover), 3-319-16655-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-3-319-16655-1_12.

Parvu:2015:TSP

[Pǎr15b] Ilie Pǎrvu. The tradition of scientific philosophy in Romania. In Romanian Studies in Philosophy of Science [PST15], chapter 1, pages 3–24. ISBN 3-319-16654-9 (hardcover), 3-319-16655-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

                                              ???? URL https://link.springer.
                                              com/chapter/10.1007/978-3-319-
                                              16655-1_1.

Parikh:2017:TCT

[Par17a] Rohit Parikh. Is there a Church–Turing Thesis for social algorithms? In Floyd and Bokulich [FB17], chapter 15, pages 339–357. ISBN 3-319-53278-2, 3-319-53280-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QA29.T8 P45 2017. URL https:/

                                              /link.springer.com/chapter/10.
                                              1007/978-3-319-53280-6_15.

Parusnikova:2017:HD

[Par17b] Zuzana Parusniková. Horror dogmatis. In Encouraging Openness: Essays for Joseph Agassi on the Occasion of His 90th Birthday [BAG17], chapter 4, pages 27–39. ISBN 3-319-57668-2 (hardcover), 3-319-57669-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN B833.E53 2017.

URL https://link.springer.com/
chapter/10.1007/978-3-319-57669-
5_4.

Pascal:1976:RM

[Pas76] B. Pascal. The relativity of magnitude.

In The Concepts of Space and Time: Their Structure and Their Development [Čap76], chapter 16, page 89. ISBN 90-277-0375-2, 94-010-1727-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN BD632.

URL https://link.springer.com/
chapter/10.1007/978-94-010-1727-
5_16.

Past:2001:EPC

[Pas01] Vello Past. The emergence of physical chemistry. In Estonian Studies in the History and Philosophy of Science [Vih01b], chapter 3, pages 35–50. ISBN 94-010-3863-5, 94-010-0672-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-010-0672-9_3.

Paslaru:2015:CME

[Pâs15] Viorel Pâslaru. Causal and mechanistic explanations, and a lesson from ecology. In Romanian Studies in Philosophy of Science [PST15], chapter 17, pages 269–289. ISBN 3-319-16654-9 (hardcover), 3-319-16655-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-3-319-16655-1_17.

Paterson:1976:VVA

[Pat76a] A. M. Paterson. Velikovsky versus academic lag (the problem of hypothesis). In Cohen et al. [CHMV76], pages 487–498. ISBN 90-277-0647-6, 90-277-0648-4 (paperback), 94-010-1449-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-010-1449-
                                              6_27.

Pattee:1976:PTB

[Pat76b] H. H. Pattee. Physical theories of biological co-ordination. In Topics in the Philosophy of Biology [GM76], chapter 7, pages 153–173. ISBN 90-277-0596-8, 94-010-1829-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-010-1829-6_7.

Patryas:1982:PAE

[Pat82] Wojciech Patryas. The pluralistic approach to empirical testing and the special forms of experiment. In Polish Essays in the Philosophy of the Natural Sciences [Kra82c], chapter 11, pages 127–133. ISBN 94-009-7705-0 (e-book), 90-277-1287-5. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-009-7705-
                                              1_11.

Paty:1986:ES

[Pat86] Michel Paty. Einstein and Spinoza. In Spinoza and the Sciences [GN86], chapter 12, pages 267–302. ISBN 94-010-8511-0, 94-009-4514-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-009-4514-2_12.

Paty:1987:SRR

[Pat87] Michel Paty. The scientific reception of relativity in France. In Glick [Gli87a], chapter 4, pages 113–167. ISBN 90-277-2498-9, 94-009-3875-6. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 103; QC173.5. US$79.00

(US). URL https://link.springer.
com/chapter/10.1007/978-94-009-
3875-5_4.

Patellis:1990:FTM

[Pat90] Ioli Patellis. Frege: Theory of meaning or philosophy of science? In Greek Studies in the Philosophy and History of Science [Nic90], chapter 24, pages 373–383. ISBN 94-010-7403-8, 94-009-2015-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-009-
2015-6_24.

Paty:1992:DPM

[Pat92] Michel Paty. Les débuts de la physique mathématique et théorique au brésil et l’influence de la tradition française. (French) [The beginnings of mathematical and theoretical physics in Brazil and the influence of French tradition]. In Science and Empires: Historical Studies about Scientific Development and European Expansion [PJM92], chapter 19, pages 173–191. ISBN 0-7923-1518-9 (hardcover), 94-010-5145-3, 94-011-2594-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 136.

URL https://link.springer.com/
chapter/10.1007/978-94-011-2594-
9_19.

Patiniotis:2003:STG

[Pat03a] Manolis Patiniotis. Scientific travels of the Greek scholars in the Eighteenth Century. In Simões et al. [SCD03b], chapter 3, pages 47–75. ISBN 1-4020-1259-4, 94-017-3584-0. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 v. 233. URL https:/

                                              /link.springer.com/chapter/10.
                                              1007/978-94-017-3584-1_3.

Paty:2003:CQS

[Pat03b] Michel Paty. The concept of quantum state: New views on old phenomena. In Renn et al. [RDS03], chapter 23, pages 451–478. ISBN 1-4020-1284-5, 1-4020-1285-3 (paperback). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 234. URL https:/

                                              /link.springer.com/chapter/10.
                                              1007/978-94-010-0111-3_23.

Patiniotis:2011:EAN

[Pat11a] Manolis Patiniotis. Eclecticism and appropriation of the new scientific methods by the Greek–speaking scholars in the Ottoman Empire. In Science between Europe and Asia: Historical Studies on the Transmission, Adoption and Adaptation of Knowledge [GR11b], chapter 13, pages 193–206. ISBN 90-481-9968-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q125 .S354 2010; Q175.52.A78 S35

                                              2010. URL https://link.springer.
                                              com/chapter/10.1007/978-90-481-
                                              9968-6_13.

Paty:2011:I

[Pat11b] Michel Paty. Introduction. In Brazilian Studies in Philosophy and History of Science: An account of recent works [KV11], chapter 1, pages 1–56. ISBN 90-481-9421-0 (hardcover), 90-481-9422-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175.3 .B73 2011. URL https:/

                                              /link.springer.com/chapter/10.
                                              1007/978-90-481-9422-3_1.

Patiniotis:2015:NHE

[Pat15] Manolis Patiniotis. Neo-Hellenic Enlightenment: In search of a European identity. In Arabatzis et al. [ARS15], chapter 9, pages 117–130. ISBN 3-319-14552-5 (paperback), 3-319-14553-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q125 .R45

2015. URL https://link.springer.
com/chapter/10.1007/978-3-319-
14553-2_9.

Pauri:1997:PWR

[Pau97] Massimo Pauri. The physical worldview and the reality of becoming. In Perspectives on Time [FSU97b], chapter 12, pages 267–297. ISBN 90-481-4774-3, 94-015-8875-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol.189. URL https:

//link.springer.com/chapter/10.
1007/978-94-015-8875-1_12.

Pauri:2000:LKQ

[Pau00] Massimo Pauri. Leibniz, Kant and the quantum. In The Reality of the Unobservable: Observability, Unobservability and Their Impact on the Issue of Scientific Realism [AP00b], chapter 20, pages 257–282. ISBN 90-481-5458-8, 94-015-9391-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-015-9391-5_20.

Paul:2017:WWW

[Pau17] Herman Paul. Weber, Wöhler, and Waitz: Virtue language in Late Nineteenth-Century physics, chemistry, and history. In van Dongen and Paul [vDP17a], chapter 7, pages 91–107. ISBN 3-319-48892-9 (hardcover), 3-319-48893-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 v. 321; Q175.32.V57 E65

2017. URL https://link.springer.
com/chapter/10.1007/978-3-319-
48893-6_7.

Poggi:1994:RSS

[PB94] Stefano Poggi and Maurizio Bossi, editors. Romanticism in science: science in Europe, 1790–1840, volume 152 of Boston Studies in the Philosophy of Science. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 1994. ISBN 0-7923-2336-X, 94-017-2921-2. ISSN 0068-0346 (print), 2214-7942 (electronic). xv + 245 pp. LCCN Q174

                                              .B67 vol. 152. URL http://www.loc.
                                              gov/catdir/enhancements/fy0821/
                                              93001728-d.html;   http://www.
                                              loc.gov/catdir/enhancements/
                                              fy0821/93001728-t.html;   https:
                                              //link.springer.com/book/10.
                                              1007/978-94-017-2921-5.

Paul:2007:DDC

[PB07] Diane B. Paul and John Beatty. Discarding dichotomies, creating community: Sam Schweber and Darwin studies. In Positioning the history of science [RG07], chapter 21, pages 113–118. ISBN 1-4020-5419-X (hardcover), 1-4020-5420-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q125 .P77 2007.

                                              URL https://link.springer.com/
                                              chapter/10.1007/1-4020-5420-3_21.

Putnam:1969:EK

[PC69] Ruth Anna Putnam and John Compton. On empirical knowledge. In Cohen and Wartofsky [CW69b], pages 392–418. ISBN 94-010-3380-3, 94-010-3378-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

                                              ???? URL https://link.springer.
                                              com/chapter/10.1007/978-94-010-
                                              3378-7_14.

Parusnikova:2009:RP

[PC09] Zuzana Parusniková and Robert S.

Cohen. Rethinking Popper, volume 272 of Boston Studies in the Philosophy of Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2009. ISBN 1-4020-9337-3, 1-4020-9338-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). xii + 431 pp. LCCN B1649.P64 R48 2009; Q174 .B67 2009.

URL https://link.springer.com/
book/10.1007/978-1-4020-9338-8.

Pears:1986:PR

[Pea86] David Pears. Practical reasoning. In Ullmann-Margalit [UM86c], chapter 7, pages 93–106. ISBN 90-277-2160-2, 90-277-2161-0 (paperback), 94-009-4566-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 95 Q175. URL https:/

/link.springer.com/chapter/10.
1007/978-94-009-4566-1_7.

Pechenkin:1994:TDV

[Pec94] Alexander A. Pechenkin. The twodimensional view of the history of chemistry. In Trends in the Historiography of Science [GCN94], chapter 26, pages 369–377. ISBN 90-481-4264-4, 94-017-3596-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-017-
3596-4_26.

Peckhaus:1996:IHG

[Pec96] Volker Peckhaus. The influence of Hermann Günther Grassmann and Robert Grassmann on Ernst Schröder’s algebra of logic. In Schubring [Sch96f], pages 217–227. ISBN 90-481-4758-1, 94-015-8753-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QA8.9-10.3. URL https:/

/link.springer.com/chapter/10.
1007/978-94-015-8753-2_18.

Peckhaus:2013:TMK

[Pec11] Volker Peckhaus. The third man: Kurt Grelling and the Berlin Group. In The Berlin Group and the Philosophy of Logical Empiricism [MP13], chapter 11, pages 231–244. ISBN 94-007-5484-1 (hardcover), 94-007-5485-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175.32.A24 .M384 2011; Q175 .B47 2013. URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-007-5485-0_11.

Peijnenburg:2003:CDC

[Pei03] Jeanne Peijnenburg. On the concept of discovery comments on Gerd Gigerenzer. In Descartes’s Mathematical Thought [Sas03d], chapter 9, pages 153–158. ISBN 90-481-6487-7, 94-017-1225-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

                                              ???? URL https://link.springer.
                                              com/chapter/10.1007/0-306-48123-
                                              5_9.

Pelham:1995:RRS

[Pel95] Judy Pelham. A reconstruction of Russell’s substitution theory. In Québec Studies in the Philosophy of Science: Part I: Logic, Mathematics, Physics and History of Science Essays in Honor of Hugues Leblanc [MC95], chapter 9, pages 123–133. ISBN 94-010-7204-3, 94-009-1575-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

                                              ???? URL https://link.springer.
                                              com/chapter/10.1007/978-94-009-
                                              1575-6_9.

Pelayo:2001:RES

[Pel01] Francisco Pelayo. Repercussions of evolutionism in the Spanish Natural History Society. In The Reception of Darwinism in the Iberian World: Spain, Spanish America and Brazil [GPSR01], chapter 6, pages 95–110. ISBN 94-010-3885-6, 94-010-0602-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-010-0602-
6_6.

Pensky:1999:TIH

[Pen99] Max Pensky. Truth and interest: On Habermas’s postscript to Nietzsche’s theory of knowledge. In Nietzsche, Theories of Knowledge, and Critical Theory: Nietzsche and the Sciences I [Bab99b], chapter 19, pages 265–272. ISBN 90-481-5233-X, 94-017-2430-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-017-2430-
2_19.

Pera:1981:IMS

[Per81]  Marcello Pera.  Inductive method

and scientific discovery. In Grmek et al. [GCC81b], pages 141–165. ISBN 90-277-1123-2, 94-010-1284-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-010-1284-3_6.

Pera:1984:JSP

[Per84] Marcello Pera. The justification of scientific progress. In Rationality in Science and Politics [And84d], chapter 12, pages 221–243. ISBN 90-277-1953-5, 94-009-6254-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-009-6254-5_12.

Perovich:1986:GSM

[Per86] Anthony N. Perovich, Jr. Genius, scientific method, and the stability of synthetic a priori principles. In Cohen and Wartofsky [CW85], chapter 17, pages 327–339. ISBN 90-277-1971-3, 94-009-5345-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-009-
5349-9_17.

Pera:1987:RDG

[Per87] Marcello Pera. The rationality of discovery: Galvani’s animal electricity. In Pitt and Pera [PP87], chapter 8, pages 177–201. ISBN 90-277-2417-2, 94-009-3779-2. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 98. US$64.00. URL https:

//link.springer.com/chapter/10.
1007/978-94-009-3779-6_8.

Pera:1989:MSD

[Per89] Marcello Pera. Methodological sophisticationism: A degenerating project. In Imre Lakatos and Theories of Scientific Change [GGN89], chapter 14, pages 169–187. ISBN 94-010-7860-2, 94-009-3025-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-009-
3025-4_14.

Perovich:1991:IVO

[Per91] Anthony N. Perovich, Jr. Incommensurability, its varieties and its ontological consequences. In Beyond Reason: Essays on the Philosophy of Paul Feyerabend [Mun91a], chapter 14, pages 313–328. ISBN 94-010-5406-1, 94-011-3188-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-011-
3188-9_14.

Perminov:1994:CSR

[Per94] Vassili Y. Perminov. The conception of the scientific research programs and the real history of mathematics. In Trends in the Historiography of Science [GCN94], chapter 23, pages 317–325. ISBN 90-481-4264-4, 94-017-3596-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-017-3596-
4_23.

Pereyra:1995:FEH

[Per95] Carlos Pereyra. Functional explanations in history. In Mexican Studies in the History and Philosophy of Science [RC95], chapter 5, pages 59–70. ISBN 94-010-6535-7 (hardcover), 94-009-0109-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-009-
0109-4_5.

Peres:1997:BIP

[Per97] Asher Peres. Bell inequalities with postselection. In Cohen et al. [CHS97b], pages 191–196. ISBN 90-481-4809-X, 94-017-2732-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC173.96-174.52. URL https:

//link.springer.com/chapter/10.
1007/978-94-017-2732-7_14.

Persson:2007:IE

[Per07] Johannes Persson. IBE and EBI.

In Rethinking Explanation [PY07], chapter 10, pages 137–147. ISBN 1-4020-5580-3 (paperback), 1-4020-5581-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN BD237 .R48 2007; Q175.32.E97 R48

2007. URL https://link.springer.

                                              com/chapter/10.1007/978-1-4020-
                                              5581-2_10.

Peregrin:2018:IM

[Per18] Jaroslav Peregrin. Intensionality in mathematics. In Piazza and Pulcini [PP18], pages 57–70. ISBN 3-319-93341-8, 3-319-93342-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QA8.4. URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-3-319-93342-9_4.

Pester:1991:TGP

[Pes91] Reinhardt Pester. The transformation of German philosophy in the context of scientific research in the Nineteenth Century. In Woodward and Cohen [WC91], pages 235–246. ISBN 94-010-5395-2, 94-011-3164-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-011-3164-3_21.

Pestre:2007:PPI

[Pes07] Dominique Pestre. Public participation and industrial technoscience today: The difficult question of accountability. In Positioning the history of science [RG07], chapter 22, pages 119–133. ISBN 1-4020-5419-X (hardcover), 1-4020-5420-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q125 .P77 2007.

                                              URL https://link.springer.com/
                                              chapter/10.1007/1-4020-5420-3_22.

Pessoa:2011:CSS

[Pes11]  Osvaldo Pessoa, Jr.  The causal

strength of scientific advances. In Brazilian Studies in Philosophy and History of Science: An account of recent works [KV11], chapter 16, pages 223–231. ISBN 90-481-9421-0 (hardcover), 90-481-9422-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175.3 .B73 2011.

URL https://link.springer.com/
chapter/10.1007/978-90-481-9422-
3_16.

Petersen:1969:PSC

[Pet69a] Aage Petersen. On the philosophical significance of the correspondence argument. In Cohen and Wartofsky [CW69a], pages 242–252. ISBN 94-010-3383-8, 94-010-3263-7, 94-010-3381-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-010-3381-
7_6.

Petrovic:1969:DMP

[Pet69b] Gajo Petrović. Dialectical materialism and the philosophy of praxis. In Cohen and Wartofsky [CW69b], pages 261–276. ISBN 94-010-3380-3, 94-010-3378-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-010-
3378-7_10.

Petrie:1976:MMM

[Pet76] Hugh G. Petrie. Metaphorical models of mastery: or, how to learn to do the problems at the end of the chapter of the physics textbook. In Cohen et al. [CHMV76], pages 301–312. ISBN 90-277-0647-6, 90-277-0648-4 (paperback), 94-010-1449-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-010-1449-6_14.

Petrovic:1979:PCR

[Pet79]  Gajo Petrović.  The philosophical

concept of revolution. In Praxis: Yugoslav Essays in the Philosophy and Methodology of the Social Sciences [MP79], chapter 9, pages 151–164. ISBN 90-277-0968-8, 94-009-9355-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-009-9355-
6_9.

Peterfreund:1987:OBE

[Pet87] Stuart Peterfreund. Organicism and the birth of energy. In Burwick [Bur87b], chapter 5, pages 113–152. ISBN 90-277-2541-1, 94-009-3917-5. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 105. US$64.00(U.S.).

URL https://link.springer.com/
chapter/10.1007/978-94-009-3917-
2_5.

Pettit:1989:DTP

[Pet89] Philip Pettit. Decision theory, political theory and the hats hypothesis. In Freedom and Rationality: Essays in Honor of John Watkins From his Colleagues and Friends [DJ89], chapter 2, pages 23–34. ISBN 94-010-7571-9, 94-009-2380-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-009-
2380-5_2.

Petitjean:1992:AMF

[Pet92a] Patrick Petitjean. Autour de la mission française pour la création de l’Université de São Paulo (1934). (French) [Around the French Mission for the creation of the University of Sao Paulo (1934)]. In Science and Empires: Historical Studies about Scientific Development and European Expansion [PJM92], chapter 34, pages 339–362. ISBN 0-7923-1518-9 (hardcover), 94-010-5145-3, 94-011-2594-5 (e-book).

ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 136.

URL https://link.springer.com/
chapter/10.1007/978-94-011-2594-
9_34.

Petitjean:1992:SET

[Pet92b] Patrick Petitjean. Sciences et empires:

un thème prometteur, des enjeux cruciaux. (French) [Science and empires: A promising theme, crucial issues]. In Science and Empires: Historical Studies about Scientific Development and European Expansion [PJM92], chapter 1, pages 3–12. ISBN 0-7923-1518-9 (hardcover), 94-010-5145-3, 94-011-2594-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 136. URL https://link.

springer.com/chapter/10.1007/
978-94-011-2594-9_1.

Petrov:2003:BHK

[Pet03a] Assen Petrov. Are bifurcations of human knowledge possible? In Bulgarian Studies in the Philosophy of Science [Gin03a], chapter 4, pages 43–49. ISBN 1-4020-1496-1, 94-017-0961-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-017-0961-3_4.

Petrov:2003:DCN

[Pet03b] Sava Petrov. The danger of catching nature in contradiction. In Bulgarian Studies in the Philosophy of Science [Gin03a], chapter 1, pages 3–15. ISBN 1-4020-1496-1, 94-017-0961-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-017-0961-3_1.

Pfarr:1984:PTP

[Pfa84] Joachim Pfarr. Protophysics of time and the principle of relativity. In Physical Sciences and History of Physics [CW84c], chapter 9, pages 159–189. ISBN 94-009-7180-X, 94-009-7178-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .P497

                                              1984. URL https://link.springer.
                                              com/chapter/10.1007/978-94-009-
                                              7178-3_9.

Pfalzgraf:1996:AGG

[Pfa96] Jochen Pfalzgraf. An application of Grassmann geometry to a problem in robotics. In Schubring [Sch96f], pages 337–343. ISBN 90-481-4758-1, 94-015-8753-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QA8.9-10.3. URL https:/

                                              /link.springer.com/chapter/10.
                                              1007/978-94-015-8753-2_27.

Phillips:1990:ITS

[Phi90a] Bernard S. Phillips. Introduction two: Simmel as a puzzling figure for contemporary sociology. In Georg Simmel and Contemporary Sociology [KPC90], chapter 2, pages 9–12. ISBN 94-010-6691-4 (hardcover), 94-009-0459-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN HM401-1281.

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-009-0459-
                                              0_2.

Phillips:1990:SIF

[Phi90b] Bernard S. Phillips. Simmel, individuality, and fundamental change. In Georg Simmel and Contemporary Sociology [KPC90], chapter 15, pages 259–281. ISBN 94-010-6691-4 (hardcover), 94-009-0459-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN HM401-1281. URL https://link.

springer.com/chapter/10.1007/
978-94-009-0459-0_15.

Phili:1994:JLF

[Phi94] Christine Phili. John Landen: First attempt for the algebrization of infinitesimal calculus. In Trends in the Historiography of Science [GCN94], chapter 21, pages 279–293. ISBN 90-481-4264-4, 94-017-3596-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-017-3596-4_21.

Peoples:1995:BNF

[PHL95] M. B. Peoples, D. F. Herridge, and J. K. Ladha. Biological nitrogen fixation: An efficient source of nitrogen for sustainable agricultural production? In No Truth Except in the Details: Essays in Honor of Martin J. Klein [KS95], chapter 1, pages 3–28. ISBN 94-010-4097-4, 94-011-0217-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-011-0055-7_1.

Piatek:1982:CIH

[Pia82] Zdzisława Piatek. Chomsky’s inconsistencies in his critique of evolutionary conceptions of language. In Polish Essays in the Philosophy of the Natural Sciences [Kra82c], chapter 31, pages 365–373. ISBN 94-009-7705-0 (e-book), 90-277-1287-5. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-009-7705-
1_31.

Piazza:1990:SGH

[Pia90] Alberto Piazza. Statistics in genetics: Human migrations detected by multivariate techniques. In Statistics in Science: The Foundations of Statistical Methods in Biology, Physics and Economics [CC90], chapter 7, pages 103–118. ISBN 94-010-6765-1, 94-009-0619-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QA276-280. URL https:/

/link.springer.com/chapter/10.
1007/978-94-009-0619-8_7.

Piatek:1995:EST

[Pia95] Zdzisława Piatek. Is evolutionism a scientific theory? In Misiek [Mis95c], pages 167–174. ISBN 94-010-4206-3, 94-011-0461-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-011-
0461-6_12.

Pickering:2012:RSD

[Pic12] Andrew Pickering. The robustness of science and the dance of agency. In Characterizing the Robustness of Science: After the Practice Turn in Philosophy of Science [STNW12], chapter 13, pages 317–327. ISBN 94-007-2758-5, 94-007-2759-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175 .C4595 2012.

URL https://link.springer.com/
chapter/10.1007/978-94-007-2759-
5_13.

Pietroski:1995:OTE

[Pie95] Paul M. Pietroski. Other things equal, the chances improve. In Québec Studies in the Philosophy of Science: Part I: Logic, Mathematics, Physics and History of Science Essays in Honor of Hugues Leblanc [MC95], chapter 15, pages 259–273. ISBN 94-010-7204-3, 94-009-1575-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-009-
1575-6_15.

Pietsch:2012:UDH

[Pie12] Wolfgang Pietsch. The underdetermination debate: How lack of history leads to bad philosophy. In Mauskopf and Schmaltz [MS12b], chapter 7, pages 83–106. ISBN 94-007-1744-X, 94-007-1745-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175 .I5134 2011. URL https:

//link.springer.com/chapter/10.
1007/978-94-007-1745-9_7.

Pietsch:2016:TMR

[Pie16] Wolfgang Pietsch. Two modes of reasoning with case studies. In Sauer and Scholl [SS16b], chapter 4, pages 49–67. ISBN 3-319-30227-2, 3-319-30229-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q126.8 .P45 2016. URL https:/

/link.springer.com/chapter/10.
1007/978-3-319-30229-4_4.

Piller:1997:HP

[Pil97]  Christian Piller.  Humeanism and

prudence. In Austrian Philosophy Past and Present: Essays in Honor of Rudolf Haller [LM97], chapter 13, pages 189–202. ISBN 94-010-6412-1, 94-011-5720-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-011-
5720-9_13.

Pingshu:1996:CYC

[Pin96] Gao Pingshu. Cai Yuanpei’s contributions to China’s science. In Chinese Studies in the History and Philosophy of Science and Technology [DC96], chapter 30, pages 395–417. ISBN 90-481-4546-5, 94-015-8717-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 180; U390 .N19 1996. URL https:

//link.springer.com/chapter/10.
1007/978-94-015-8717-4_30.

Pinar:2001:DB

[Pin01] Susana Pinar. Darwinism and botany. In The Reception of Darwinism in the Iberian World: Spain, Spanish America and Brazil [GPSR01], chapter 7, pages 111–126. ISBN 94-010-3885-6, 94-010-0602-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-010-
0602-6_7.

Pingyi:2010:STC

[Pin10] Chu Pingyi. Scientific texts in contest, 1600–1800. In Looking at it from Asia: the Processes that Shaped the Sources of History of Science [BE10b], chapter 5, pages 141–166. ISBN 90-481-3675-X (hardcover), 90-481-3676-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q127.A65 L66

2010. URL https://link.springer.
com/chapter/10.1007/978-90-481-
3676-6_5.

Pincock:2011:MSR

[Pin11] Christopher Pincock. Mathematical structural realism. In Bokulich and Bokulich [BB11], chapter 4, pages 67–79. ISBN 90-481-9596-9, 90-481-9597-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175.32.S78 S35 2011. URL https:

//link.springer.com/chapter/10.
1007/978-90-481-9597-8_4.

Pistone:2015:RFL

[Pis15] Paolo Pistone. Rule-following and the limits of formalization: Wittgenstein’s considerations through the lens of logic. In From Logic to Practice: Italian Studies in the Philosophy of Mathematics [LPV15], chapter 6, pages 91–110. ISBN 3-319-10433-0, 3-319-10434-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QA8.4 .F76 2015; Q174 .B67 v.308.

URL https://link.springer.com/
chapter/10.1007/978-3-319-10434-
8_6.

Pitt:1987:GRC

[Pit87] Joseph C. Pitt. Galileo and rationality:

The case of the tides. In Pitt and Pera [PP87], chapter 6, pages 135–153. ISBN 90-277-2417-2, 94-009-3779-2. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol.

98. US$64.00. URL https://link.

springer.com/chapter/10.1007/
978-94-009-3779-6_6.

Pitt:1989:APS

[Pit89] Joseph C. Pitt. Apologia Pro Simplicio:

Galileo and the limits of knowledge. In An Intimate Relation: Studies in the History and Philosophy of Science Presented to Robert E. Butts on his 60th Birthday [BM89], chapter 1, pages 1–22. ISBN 94-010-7546-8, 94-009-2327-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-009-
2327-0_1.

Pitowsky:1992:WDP

[Pit92] Itamar Pitowsky. Why does physics need mathematics? A comment. In The Scientific Enterprise: The Bar-Hillel Colloquium: Studies in History, Philosophy, and Sociology of Science, Volume 4 [UM92], chapter 10, pages 163–167. ISBN 94-010-5190-9, 94-011-2688-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .S354 1992. URL https://link.

                                              springer.com/chapter/10.1007/
                                              978-94-011-2688-5_10.

Pizzi:1981:SEI

[Piz81] Claudio Pizzi. ‘Since’, ‘even if’, ‘as if’. In Italian Studies in the Philosophy of Science [Chi81b], chapter 6, pages 73– 87. ISBN 90-277-1073-2, 94-009-8937-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-009-8937-
                                              5_6.

Petitjean:1992:SEH

[PJM92] Patrick Petitjean, Catherine Jami, and Anne Marie Moulin. Science and Empires: Historical Studies about Scientific Development and European Expansion, volume 136 of Boston Studies in the Philosophy of Science. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 1992. ISBN 0-7923-1518-9 (hardcover), 94-010-5145-3, 94-011-2594-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). xiii + 411 pp. LCCN Q174 .B67 vol. 136.

                                              URL https://link.springer.com/
                                              book/10.1007/978-94-011-2594-9.

Planck:1992:MPP

[Pla92]  Max Planck.  The Mach–Planck

polemics [1908–1909]. In Ernst Mach — A Deeper Look: Documents and New Perspectives [Bla92a], chapter 6, pages 127–150. ISBN 94-010-5230-1, 94-011-2771-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN C5.53; QC6.9. URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-011-2771-4_6.

Plaass:1994:KTNa

[Pla94a] Peter Plaass. Kant’s theory of natural science. In Kant’s Theory of Natural Science [Pla94b], chapter 3, pages 189–332. ISBN 94-010-4492-9, 94-011-1126-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-011-
1126-3_3.

Plaass:1994:KTNb

[Pla94b] Peter Plaass. Kant’s Theory of Natural Science, volume 159 of Boston Studies in the Philosophy of Science. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 1994. ISBN 94-010-4492-9, 94-011-1126-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). xvi + 367 pp. LCCN ???? URL

https://link.springer.com/book/
10.1007/978-94-011-1126-3.

Plazowski:1995:RB

[Pła95] Jan Werszowiec Płazowski. Rationality and beyond. In Misiek [Mis95c], pages 175–193. ISBN 94-010-4206-3, 94-011-0461-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-011-
0461-6_13.

Peoples:1995:ELF

[PLH95] M. B. Peoples, J. K. Ladha, and D. F. Herridge. Enhancing legume N_{2} fixation through plant and soil management. In No Truth Except in the Details: Essays in Honor of Martin J. Klein [KS95], chapter 4, pages 83–101. ISBN 94-010-4097-4, 94-011-0217-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-011-0055-
                                              7_4.

Plotinus:1976:CRT

[Plo76] Plotinus. Criticism of the relational theories of time. In The Concepts of Space and Time: Their Structure and Their Development [Čap76], chapter 29, pages 173–178. ISBN 90-277-0375-2, 94-010-1727-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic).

                                              LCCN BD632. URL https://link.
                                              springer.com/chapter/10.1007/
                                              978-94-010-1727-5_29.

Pietruska-Madej:1982:CAC

[PM82] Elżbieta Pietruska-Madej. Continuity and anticumulative changes in the growth of science. In Polish Essays in the Philosophy of the Natural Sciences [Kra82c], chapter 7, pages 75–90. ISBN 94-009-7705-0 (e-book), 90-277-1287-5. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-009-7705-1_7.

Perez:2014:PUO

[PM14] Luis M. Méndez Pérez and Carlos A. Cabal Mirabal. Physics at the university of oriente. In The History of Physics in Cuba [BRW14a], chapter 8, pages 247–260. ISBN 94-017-8040-4, 94-017-8041-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC5.53; QC6.9; QC9 C9H57 2014; Q124.6-127.2. URL https:/

                                              /link.springer.com/chapter/10.
                                              1007/978-94-017-8041-4_8.

Perrier:1991:MIX

[PO91] Albert Perrier and H. Kamerlingh Onnes. Magnetic investigations. XIII. The susceptibility of liquid mixtures of oxygen and nitrogen, and the influence of the mutual distance of the molecules upon paramagnetism. In Gavroglu and Goudaroulis [GG91], chapter 27, pages 472–489. ISBN 94-010-7433-X (hardcover), 94-009-2079-2. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 124. URL https:/

/link.springer.com/chapter/10.
1007/978-94-009-2079-8_27.

Poellner:1999:CFN

[Poe99] Peter Poellner. Causation and force in Nietzsche. In Nietzsche, Epistemology, and Philosophy of Science: Nietzsche and the Sciences II [Bab99a], chapter 22, pages 287–297. ISBN 90-481-5234-8, 94-017-2428-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-017-2428-9_22.

Poggi:1994:NBR

[Pog94] Stefano Poggi. Neurology and biology in the Romantic Age in Germany: Carus, Burdach, Gall, von Baer. In Poggi and Bossi [PB94], chapter 9, pages 143–160. ISBN 0-7923-2336-X, 94-017-2921-2. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 152. URL https://link.

springer.com/chapter/10.1007/
978-94-017-2921-5_9.

Poggiolesi:2016:COM

[Pog16]  Francesca Poggiolesi.  A critical

overview of the most recent logics of grounding. In Objectivity, Realism, and Proof: FilMat Studies in the Philosophy of Mathematics [BS16], chapter 15, pages 291–309. ISBN 3-319-31642-7, 3-319-31644-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QA9 .O25 2016. URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-3-319-31644-4_15.

Poincare:2007:DEE

[Poi07] Henri Poincaré. On the dynamics of the electron (excerpts). In Renn et al. [RSS^{+}07a], pages 253–271. ISBN 1-4020-3999-9. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 no. 250 v. 3; QC173.6 .G469 2007.

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-1-4020-4000-

9_19. Originally published as “Sur la dynamique de l’électron” in Rendiconti del Circolo Matematico di Palermo 21, 129–175 (1906), dated Paris July 1905.

Pole:1971:SSI

[Pol71] Nelson Pole. ‘Self-supporting’ inductive arguments. In Buck and Cohen [BC71], pages 496–503. ISBN 90-277-0309-4, 94-010-3142-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

                                              ???? URL https://link.springer.
                                              com/chapter/10.1007/978-94-010-
                                              3142-4_32.

Polanyi:1974:GS

[Pol74a] Michael Polanyi. Genius in science. In Methodological and Historical Essays in the Natural and Social Sciences [CW74b], chapter 4, pages 57–71. ISBN 90-277-0378-7, 94-010-2128-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 14.

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-010-2128-
                                              9_4.

Polikarov:1974:DCM

[Pol74b] A. Polikarov. The divergent–convergent method — a heuristic approach to problem-solving. In Methodological and Historical Essays in the Natural and Social Sciences [CW74b], chapter 11, pages 211–233. ISBN 90-277-0378-7, 94-010-2128-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 14. URL https:

//link.springer.com/chapter/10.
1007/978-94-010-2128-9_11.

Polanyi:1976:LIS

[Pol76] Michael Polanyi. Life’s irreducible structure. In Topics in the Philosophy of Biology [GM76], chapter 5, pages 128–142. ISBN 90-277-0596-8, 94-010-1829-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-010-
1829-6_5.

Polanco:1992:WSH

[Pol92] Xavier Polanco. World-science: How is the history of world-science to be written? In Science and Empires: Historical Studies about Scientific Development and European Expansion [PJM92], chapter 24, pages 225–242. ISBN 0-7923-1518-9 (hardcover), 94-010-5145-3, 94-011-2594-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 136.

URL https://link.springer.com/
chapter/10.1007/978-94-011-2594-
9_24.

Polikarov:1995:SQC

[Pol95] Azarya Polikarov. Some questions concerning limitations of the range of validity of Kuhn’s model of the history of science. In Gavroglu et al. [GSW95c], chapter 13, pages 235–240. ISBN 0-7923-2991-0 (set), 0-7923-2988-0, 94-017-2658-2. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 163. URL https:/

/link.springer.com/chapter/10.
1007/978-94-017-2658-0_13.

Polikarov:2003:PSP

[Pol03] Azarya Polikarov. The proliferation and synthesis of physical theories. In Bulgarian Studies in the Philosophy of Science [Gin03a], chapter 5, pages 53– 68. ISBN 1-4020-1496-1, 94-017-0961-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-017-0961-
                                              3_5.

Pollock:2016:SCT

[Pol16] Samantha Pollock. The significance of a categoricity theorem for formal theories and informal beliefs. In Objectivity, Realism, and Proof: FilMat Studies in the Philosophy of Mathematics [BS16], chapter 17, pages 319–344. ISBN 3-319-31642-7, 3-319-31644-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QA9 .O25 2016.

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-3-319-31644-
                                              4_17.

Popkin:1974:BCS

[Pop74] Richard Popkin. Bible criticism and social science. In Methodological and Historical Essays in the Natural and Social Sciences [CW74b], chapter 17, pages 339–360. ISBN 90-277-0378-7, 94-010-2128-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 14. URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-010-2128-9_17.

Popper:1984:AIO

[Pop84] Karl R. Popper. Against induction: One of many arguments. In Rationality in Science and Politics [And84d], chapter 13, pages 245–247. ISBN 90-277-1953-5, 94-009-6254-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-009-6254-5_13.

Popkin:1986:TFS

[Pop86a] Richard Popkin. The third force in Seventeenth-Century thought: Skepticism, science and millenarianism. In Ullmann-Margalit [UM86c], chapter 3, pages 21–50. ISBN 90-277-2160-2, 90-277-2161-0 (paperback), 94-009-4566-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 95 Q175. URL https:/

/link.springer.com/chapter/10.
1007/978-94-009-4566-1_3.

Popkin:1986:SNL

[Pop86b] Richard H. Popkin. Some new light on the roots of Spinoza’s science of Bible study. In Spinoza and the Sciences [GN86], chapter 7, pages 171–188. ISBN 94-010-8511-0, 94-009-4514-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-009-4514-
2_7.

Popper:1987:MF

[Pop87] Karl Popper. The myth of the framework. In Pitt and Pera [PP87], chapter 2, pages 35–62. ISBN 90-277-2417-2, 94-009-3779-2. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 98. US$64.00. URL https:

//link.springer.com/chapter/10.
1007/978-94-009-3779-6_2.

Popkin:1988:MEH

[Pop88]  Richard Popkin.  Millenarianism

in England, Holland and America: Jewish–Christian relations in Amsterdam, London and Newport, Rhode Island. In Philosophy, History and Social Action: Essays in Honor of Lewis Feuer with an autobiographic essay by Lewis Feuer [HOO88c], chapter 18, pages 349–371. ISBN 94-010-7793-2, 94-009-2873-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-009-
2873-2_18.

Popkin:1992:NOF

[Pop92] Richard H. Popkin. Newton and the origins of fundamentalism. In The Scientific Enterprise: The Bar-Hillel Colloquium: Studies in History, Philosophy, and Sociology of Science, Volume 4 [UM92], chapter 15, pages 241–259. ISBN 94-010-5190-9, 94-011-2688-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174

.S354 1992. URL https://link.
springer.com/chapter/10.1007/
978-94-011-2688-5_15.

Portmann:1987:GCM

[Por87] Adolf Portmann. Goethe and the concept of metamorphosis. In Amrine et al. [AZW87], chapter 8, pages 133–145. ISBN 90-277-2265-X, 94-009-3761-X. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol.

97. US$79.00. URL https://link.

springer.com/chapter/10.1007/
978-94-009-3761-1_8.

Porter:1994:QMS

[Por94] Theodore M. Porter. From Quetelet to Maxwell: Social statistics and the origins of statistical physics. In Cohen [Coh94b], chapter 11, pages 345–362. ISBN 0-7923-2223-1, 94-017-3391-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 150. URL https://link.

springer.com/chapter/10.1007/
978-94-017-3391-5_11.

Post:1974:CFR

[Pos74] E. J. Post. Comments on ‘the formal representation of physical quantities’. In Logical and Epistemological Studies in Contemporary Physics [CW74a], chapter 6, pages 210–213. ISBN 90-277-0391-4 (hardcover), 90-277-0377-9 (paperback), 94-010-2656-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol.

13; QC5.56. URL https://link.
springer.com/chapter/10.1007/
978-94-010-2656-7_6.

Posner:1976:DME

[Pos76a] Roland Posner. Discourse as a means to enlightenment. In Language in Focus: Foundations, Methods and Systems: Essays in Memory of Yehoshua Bar-Hillel [Kas76b], chapter 27, pages 641–660. ISBN 90-277-0644-1, 90-277-0645-X, 94-010-1876-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN P101-120. URL https://link.

springer.com/chapter/10.1007/
978-94-010-1876-0_27.

Post:1976:NPC

[Pos76b] H. R. Post. Novel predictions as a criterion of merit. In Essays in Memory of Imre Lakatos [CFW76], chapter 28, pages 493–495. ISBN 90-277-0655-7, 94-010-1451-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-010-
1451-9_28.

Post:1978:OVS

[Pos78] Heinz Post. Objectivism vs sociologism.

In Progress and Rationality in Science [RA78b], chapter 14, pages 311–318. ISBN 90-277-0922-X, 94-009-9866-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-009-9866-
7_14.

Post:1993:CIH

[Pos93] H. R. Post. Correspondence, invariance and heuristics: In praise of conservative induction. In Correspondence, Invariance and Heuristics: Essays in Honour of Heinz Post [FK93], chapter 1, pages 1–43. ISBN 90-481-4229-6, 94-017-1185-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-017-
1185-2_1.

Post:1995:AVR

[Pos95a] Evert Jan Post. Absolute versus relative indeterminism. In Quantum Reprogramming: Ensembles and Single Systems: a Two-Tier Approach to Quantum Mechanics [Pos95p], chapter 17, pages 264–282. ISBN 0-7923-3565-1 (hardcover), 94-015-8410-9. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 181; Q174; Q174.12; QC174.12 P675 1995. URL https:

//link.springer.com/chapter/10.
1007/978-94-015-8410-4_17.

Post:1995:ATC

[Pos95b] Evert Jan Post. Arrowed time and cyclic time. In Quantum Reprogramming: Ensembles and Single Systems: a Two-Tier Approach to Quantum Mechanics [Pos95p], chapter 12, pages 189–197. ISBN 0-7923-3565-1 (hardcover), 94-015-8410-9. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 181; Q174; Q174.12; QC174.12 P675 1995. URL https:

//link.springer.com/chapter/10.
1007/978-94-015-8410-4_12.

Post:1995:CNA

[Pos95c] Evert Jan Post. Classical nonclassical asymptotics. In Quantum Reprogramming: Ensembles and Single Systems: a Two-Tier Approach to Quantum Mechanics [Pos95p], chapter 11, pages 169–187. ISBN 0-7923-3565-1 (hardcover), 94-015-8410-9. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 181; Q174; Q174.12; QC174.12 P675 1995. URL https:

//link.springer.com/chapter/10.
1007/978-94-015-8410-4_11.

Post:1995:CQM

[Pos95d] Evert Jan Post. Compatibility of quantum mechanics and relativity. In Quantum Reprogramming: Ensembles and Single Systems: a Two-Tier Approach to Quantum Mechanics [Pos95p], chapter 15, pages 240–254. ISBN 0-7923-3565-1 (hardcover), 94-015-8410-9. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 181; Q174; Q174.12; QC174.12 P675 1995. URL https:

//link.springer.com/chapter/10.
1007/978-94-015-8410-4_15.

Post:1995:CVC

[Pos95e] Evert Jan Post. Copenhagen versus Copenhagen. In Quantum Reprogramming: Ensembles and Single Systems: a Two-Tier Approach to Quantum Mechanics [Pos95p], chapter 4, pages 60–73. ISBN 0-7923-3565-1 (hardcover), 94-015-8410-9. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 181; Q174; Q174.12; QC174.12 P675 1995. URL https:

//link.springer.com/chapter/10.
1007/978-94-015-8410-4_4.

Post:1995:DMM

[Pos95f] Evert Jan Post. The Diffeo-4 mandate of Michelson–Sagnac. In Quantum Reprogramming: Ensembles and Single Systems: a Two-Tier Approach to Quantum Mechanics [Pos95p], chapter 18, pages 283–296. ISBN 0-7923-3565-1 (hardcover), 94-015-8410-9. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 181; Q174; Q174.12; QC174.12 P675 1995.

URL https://link.springer.com/
chapter/10.1007/978-94-015-8410-
4_18.

Post:1995:EEF

[Pos95g] Evert Jan Post. Epilogue for extrapolating a favor of fortune. In Quantum Reprogramming: Ensembles and Single Systems: a Two-Tier Approach to Quantum Mechanics [Pos95p], chapter 19, pages 297–307. ISBN 0-7923-3565-1 (hardcover), 94-015-8410-9. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 181; Q174; Q174.12; QC174.12 P675 1995.

URL https://link.springer.com/
chapter/10.1007/978-94-015-8410-
4_19.

Post:1995:FPI

[Pos95h] Evert Jan Post. Fitting period integrals to physics. In Quantum Reprogramming: Ensembles and Single Systems: a Two-Tier Approach to Quantum Mechanics [Pos95p], chapter 8, pages 126–143. ISBN 0-7923-3565-1 (hardcover), 94-015-8410-9. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 181; Q174; Q174.12; QC174.12 P675 1995. URL https:

//link.springer.com/chapter/10.
1007/978-94-015-8410-4_8.

Post:1995:ICB

[Pos95i]  Evert Jan Post.  Implications of

cooperative behavior. In Quantum Reprogramming: Ensembles and Single Systems: a Two-Tier Approach to Quantum Mechanics [Pos95p], chapter 9, pages 144–150. ISBN 0-7923-3565-1 (hardcover), 94-015-8410-9. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 181; Q174; Q174.12; QC174.12 P675 1995.

URL https://link.springer.com/
chapter/10.1007/978-94-015-8410-
4_9.

Post:1995:IR

[Pos95j] Evert Jan Post. Introductory remarks. In Quantum Reprogramming: Ensembles and Single Systems: a Two-Tier Approach to Quantum Mechanics [Pos95p], chapter 1, pages 1–14. ISBN 0-7923-3565-1 (hardcover), 94-015-8410-9. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 181; Q174; Q174.12; QC174.12 P675 1995. URL https:

//link.springer.com/chapter/10.
1007/978-94-015-8410-4_1.

Post:1995:LCA

[Pos95k] Evert Jan Post. Larmor and cyclotron aspects of flux quanta. In Quantum Reprogramming: Ensembles and Single Systems: a Two-Tier Approach to Quantum Mechanics [Pos95p], chapter 7, pages 119–125. ISBN 0-7923-3565-1 (hardcover), 94-015-8410-9. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 181; Q174; Q174.12; QC174.12 P675 1995.

URL https://link.springer.com/
chapter/10.1007/978-94-015-8410-
4_7.

Post:1995:ORF

[Pos95l] Evert Jan Post. Optimizing reduction to familiar concepts. In Quantum Reprogramming: Ensembles and Single Systems: a Two-Tier Approach to Quantum Mechanics [Pos95p], chapter 14, pages 224–238. ISBN 0-7923-3565-1 (hardcover), 94-015-8410-9. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 181; Q174; Q174.12; QC174.12 P675 1995.

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-015-8410-
                                              4_14.

Post:1995:PIU

[Pos95m] Evert Jan Post. Period integrals: A universal tool of physics. In Quantum Reprogramming: Ensembles and Single Systems: a Two-Tier Approach to Quantum Mechanics [Pos95p], chapter 6, pages 85–118. ISBN 0-7923-3565-1 (hardcover), 94-015-8410-9. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 181; Q174; Q174.12; QC174.12 P675 1995.

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-015-8410-
                                              4_6.

Post:1995:PR

[Pos95n] Evert Jan Post. The psychology of the 1925 revolution. In Quantum Reprogramming: Ensembles and Single Systems: a Two-Tier Approach to Quantum Mechanics [Pos95p], chapter 2, pages 16–34. ISBN 0-7923-3565-1 (hardcover), 94-015-8410-9. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 181; Q174; Q174.12; QC174.12 P675 1995.

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-015-8410-
                                              4_2.

Post:1995:QC

[Pos95o] Evert Jan Post. Quantum cohomology. In Quantum Reprogramming: Ensembles and Single Systems: a Two-Tier Approach to Quantum Mechanics [Pos95p], chapter 13, pages 198–223. ISBN 0-7923-3565-1 (hardcover), 94-015-8410-9. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 181; Q174; Q174.12; QC174.12 P675 1995. URL https:

//link.springer.com/chapter/10.
1007/978-94-015-8410-4_13.

Post:1995:QRE

[Pos95p] Evert Jan Post. Quantum Reprogramming: Ensembles and Single Systems: a Two-Tier Approach to Quantum Mechanics, volume 181 of Boston Studies in the Philosophy of Science. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 1995. ISBN 0-7923-3565-1 (hardcover), 94-015-8410-9. ISSN 0068-0346 (print), 2214-7942 (electronic). xiv + 320 pp. LCCN Q174 .B67 vol. 181; Q174; Q174.12; QC174.12 P675

1995. URL http://books.google.
com/books?id=lK3aAAAAMAAJ;
http://catalog.hathitrust.org/
api/volumes/oclc/32469387.html;
 http://catdir.loc.gov/catdir/
enhancements/fy0823/95017463d.
html;   http://catdir.loc.gov/
catdir/enhancements/fy0823/
95017463t.html;   https://link.
springer.com/book/10.1007/978-
94-015-8410-4.

Post:1995:QUG

[Pos95q] Evert Jan Post. Quantum understanding in global perspective. In Quantum Reprogramming: Ensembles and Single Systems: a Two-Tier Approach to Quantum Mechanics [Pos95p], chapter 16, pages 255–263. ISBN 0-7923-3565-1 (hardcover), 94-015-8410-9. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 181; Q174; Q174.12; QC174.12 P675 1995.

URL https://link.springer.com/

                                              chapter/10.1007/978-94-015-8410-
                                              4_16.

Post:1995:RC

[Pos95r] Evert Jan Post. Reassessing Copenhagen. In Quantum Reprogramming: Ensembles and Single Systems: a Two-Tier Approach to Quantum Mechanics [Pos95p], chapter 3, pages 35–59. ISBN 0-7923-3565-1 (hardcover), 94-015-8410-9. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 181; Q174; Q174.12; QC174.12 P675 1995. URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-015-8410-4_3.

Post:1995:TFS

[Pos95s] Evert Jan Post. A tale of fine structure coincidences. In Quantum Reprogramming: Ensembles and Single Systems: a Two-Tier Approach to Quantum Mechanics [Pos95p], chapter 10, pages 151–168. ISBN 0-7923-3565-1 (hardcover), 94-015-8410-9. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 181; Q174; Q174.12; QC174.12 P675 1995. URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-015-8410-4_10.

Post:1995:NPE

[Pos95t] Evert Jan Post. Von Neumann, Popper– EPR, Bohm, Bell, Aspect. In Quantum Reprogramming: Ensembles and Single Systems: a Two-Tier Approach to Quantum Mechanics [Pos95p], chapter 5, pages 74–83. ISBN 0-7923-3565-1 (hardcover), 94-015-8410-9. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 181; Q174; Q174.12; QC174.12 P675 1995.

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-015-8410-
                                              4_5.

Poser:2018:OTA

[Pos18] Hans Poser. Ontology of technical artifacts: A proposal. In Mitcham et al. [MLNZ18], chapter 3, pages 23– 36. ISBN 3-319-62448-2, 3-319-62450-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN TA15 .P45

2018. URL https://link.springer.
com/chapter/10.1007/978-3-319-
62450-1_3.

Potrc:1997:HBE

[Pot97] Matjaz Potrc. Haller and Brentano’s empiricism. In Austrian Philosophy Past and Present: Essays in Honor of Rudolf Haller [LM97], chapter 5, pages 55–69. ISBN 94-010-6412-1, 94-011-5720-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-011-
5720-9_5.

Piattelli-Palmarini:1981:ECP

[PP81]  M. Piattelli-Palmarini.  Equilibria,

crystals, programs, energetic models, and organizational models. In Italian Studies in the Philosophy of Science [Chi81b], chapter 19, pages 341–359. ISBN 90-277-1073-2, 94-009-8937-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-009-8937-
5_19.

Pitt:1987:RCS

[PP87] Joseph C. Pitt and Marcello Pera, editors. Rational Changes in Science: Essays on Scientific Reasoning, volume 98 of Boston Studies in the Philosophy of Science. D. Reidel, Dordrecht, The Netherlands; Boston, MA, USA; Lancaster, UK; Tokyo, Japan, 1987. ISBN 90-277-2417-2, 94-009-3779-2. ISSN 0068-0346 (print), 2214-7942 (electronic). xiii + 224 pp. LCCN Q174 .B67 vol. 98. US$64.00.

URL https://link.springer.com/
book/10.1007/978-94-009-3779-6.

Plebanski:2003:VTT

[PP03] Jerzy F. Plebański and Maciej Przanowski. On vacuum twisting type-N again. In Renn et al. [RDS03], chapter 16, pages 361–373. ISBN 1-4020-1284-5, 1-4020-1285-3 (paperback). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 234. URL https://link.

springer.com/chapter/10.1007/
978-94-010-0111-3_16.

Piazza:2015:DAT

[PP15a] Mario Piazza and Gabriele Pulcini. A deflationary account of the truth of the Gödel sentence G. In From Logic to Practice: Italian Studies in the Philosophy of Mathematics [LPV15], chapter 5, pages 71–90. ISBN 3-319-10433-0, 3-319-10434-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QA8.4 .F76 2015; Q174 .B67 v.308. URL https://link.

springer.com/chapter/10.1007/
978-3-319-10434-8_5. See erratum

[PP15b].

Piazza:2015:ECD

[PP15b] Mario Piazza and Gabriele Pulcini. Erratum: Chapter 5: A Deflationary Account of the Truth of the Gödel Sentence G. In From Logic to Practice: Italian Studies in the Philosophy of Mathematics [LPV15], chapter 16, page E1. ISBN 3-319-10433-0, 3-319-10434-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QA8.4 .F76 2015; Q174 .B67 v.308.

URL https://link.springer.com/

chapter/10.1007/978-3-319-10434-

8_16. See [PP15a].

Piazza:2016:WSA

[PP16] Mario Piazza and Gabriele Pulcini.

What’s so special about the Gödel sentence G? In Objectivity, Realism, and Proof: FilMat Studies in the Philosophy of Mathematics [BS16], chapter 13, pages 245–263. ISBN 3-319-31642-7, 3-319-31644-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QA9 .O25 2016. URL https:

//link.springer.com/chapter/10.
1007/978-3-319-31644-4_13.

Piazza:2018:TEE

[PP18] Mario Piazza and Gabriele Pulcini, editors. Truth, Existence and Explanation: FilMat 2016 Studies in the Philosophy of Mathematics, volume 334 of Boston Studies in the Philosophy and History of Science. Springer-Ver-lag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2018. ISBN 3-319-93341-8, 3-319-93342-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QA8.4.

URL https://link.springer.com/
book/10.1007/978-3-319-93342-9.

Peishan:1996:QCG

[PQQSe96] Li Peishan, Meng Qinzhe, Huang Qinghe, and Huang Shu-e. The Qingdao Conference of 1956 on genetics: The historical background and fundamental experiences. In Chinese Studies in the History and Philosophy of Science and Technology [DC96], chapter 5, pages 41–54. ISBN 90-481-4546-5, 94-015-8717-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 180; U390 .N19 1996.

URL https://link.springer.com/
chapter/10.1007/978-94-015-8717-
4_5.

Popescu:1997:APD

[PR97] Sandu Popescu and Daniel Rohrlich. Action and passion at a distance. In Cohen et al. [CHS97b], pages 197–206. ISBN 90-481-4809-X, 94-017-2732-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC173.96-174.52. URL https://link.

                                              springer.com/chapter/10.1007/
                                              978-94-017-2732-7_15.

Plaisance:2012:PBBa

[PR12a] Kathryn S. Plaisance and Thomas A. C. Reydon. The philosophy of behavioral biology. In Philosophy of Behavioral Biology [PR12b], chapter 1, pages 3–24. ISBN 94-007-1950-7, 94-007-1951-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QH331 .P4679 2012; Q175 .B657 v.282.

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-007-1951-
                                              4_1.

Plaisance:2012:PBBb

[PR12b] Kathryn S. Plaisance and Thomas A. C. Reydon. Philosophy of Behavioral Biology, volume 282 of Boston Studies in the Philosophy of Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2012. ISBN 94-007-1950-7, 94-007-1951-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). viii + 373 pp. LCCN QH331 .P4679 2012; Q175 .B657 v.282.

                                              URL https://link.springer.com/
                                              book/10.1007/978-94-007-1951-4.

Parikh:2017:JTB

[PR17] Rohit Parikh and Adriana Renero. Justified true belief: Plato, Gettier, and Turing. In Floyd and Bokulich [FB17], chapter 4, pages 93–102. ISBN 3-319-53278-2, 3-319-53280-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QA29.T8 P45 2017.

URL https://link.springer.com/
chapter/10.1007/978-3-319-53280-
6_4.

Pratt:1976:BC

[Pra76] Vernon Pratt. Biological classification.

In Topics in the Philosophy of Biology [GM76], chapter 17, pages 372–395. ISBN 90-277-0596-8, 94-010-1829-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-010-1829-
6_17.

Prawitz:1993:RHP

[Pra93] Dag Prawitz. Remarks on Hilbert’s program for the foundation of mathematics. In Bridging the Gap: Philosophy, Mathematics, and Physics: Lectures on the Foundations of Science [CCG93], chapter 4, pages 87–98. ISBN 94-010-5101-1, 94-011-2496-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-011-2496-6_4.

Provine:1992:RFS

[Pro92] William B. Provine. The R. A. Fisher– Sewall Wright controversy. In Sarkar [Sar92a], chapter 8, pages 201–229. ISBN 0-7923-1777-7 (hardcover), 94-011-2856-1. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 142. URL https://link.

springer.com/chapter/10.1007/
978-94-011-2856-8_8.

Proust:1994:TCE

[Pro94] Joëlle Proust. Time and conscious experience. In Artifacts, Representations and Social Practice: Essays for Marx Wartofsky [GC94], chapter 18, pages 323–341. ISBN 94-010-4390-6, 94-011-0902-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN B73 .A77 1994. URL https:

//link.springer.com/chapter/10.
1007/978-94-011-0902-4_18.

Proust:1995:FMI

[Pro95] Joëlle Proust. Functionalism and multirealizability on interaction between structure and function. In Science, Mind and Art: Essays on science and the humanistic understanding in art, epistemology, religion and ethics In honor of Robert S. Cohen [GSW95b], chapter 11, pages 169–185. ISBN 0-7923-2990-2 (hardcover), 94-011-0469-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 165 BD418.3. URL https:

//link.springer.com/chapter/10.
1007/978-94-011-0469-2_11.

Prosperi:2000:CEQ

[Pro00] Giovanni Maria Prosperi. Common experience and quantum theory. In The Reality of the Unobservable: Observability, Unobservability and Their Impact on the Issue of Scientific Realism [AP00b], chapter 26, pages 343–351. ISBN 90-481-5458-8, 94-015-9391-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-015-9391-
5_26.

Proctor:2001:HEA

[Pro01] Robert N. Proctor. Human experimental abuse, in and out of context. In Allen and MacLeod [AM01b], chapter 15, pages 235–254. ISBN 1-4020-0495-8 (hardcover), 94-017-2956-5. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 228; Q175.55.

URL https://link.springer.com/
chapter/10.1007/978-94-017-2956-
7_15.

Proust:2009:WMF

[Pro09] Joëlle Proust. What is a mental function? In Brenner and Gayon [BG09], chapter 11, pages 227–253. ISBN 1-4020-9367-5, 1-4020-9368-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .F746 2009.

URL https://link.springer.com/
chapter/10.1007/978-1-4020-9368-
5_11.

Proust:2010:MML

[Pro10]  Christine Proust.  Mesopotamian

metrological lists and tables: Forgotten sources. In Looking at it from Asia: the Processes that Shaped the Sources of History of Science [BE10b], chapter 8, pages 245–276. ISBN 90-481-3675-X (hardcover), 90-481-3676-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q127.A65 L66

2010. URL https://link.springer.
com/chapter/10.1007/978-90-481-
3676-6_8.

Proust:2014:DMA

[Pro14] Christine Proust. Does a master always write for his students? Some evidence from Old Babylonian scribal schools. In Bernard and Proust [BP14b], chapter 4, pages 69–94. ISBN 94-007-5121-4 (hardcover), 94-007-5122-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q181 .S35 2014; Q181 S387 2014. URL https://link.

springer.com/chapter/10.1007/
978-94-007-5122-4_4.

Proietti:2017:TLT

[Pro17a] Carlo Proietti. Time of logics and time of physics. In Time of Nature and the Nature of Time: Philosophical Perspectives of Time in Natural Sciences [BH17b], chapter 3, pages 39–53. ISBN 3-319-53723-7 (hardcover), 3-319-53725-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC173.59.S65 T56 2017. URL https:

//link.springer.com/chapter/10.
1007/978-3-319-53725-2_3.

Proudfoot:2017:TFW

[Pro17b] Diane Proudfoot. Turing and free will: A new take on an old debate. In Floyd and Bokulich [FB17], chapter 13, pages 305–321. ISBN 3-319-53278-2, 3-319-53280-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QA29.T8 P45 2017. URL https:/

/link.springer.com/chapter/10.
1007/978-3-319-53280-6_13.

Padovani:2015:OSN

[PRT15] Flavia Padovani, Alan Richardson, and Jonathan Y. Tsou. Objectivity in Science: New Perspectives from Science and Technology Studies, volume 310 of Boston Studies in the Philosophy and History of Science. Springer-Ver-lag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2015. ISBN 3-319-14348-4 (hardcover), 3-319-14349-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). v + 226 pp. LCCN Q175 O25 2015. URL

https://link.springer.com/book/
10.1007/978-3-319-14349-1.

Przelecki:1995:DPV

[Prz95] Marian Przełecki. In defence of the positivist view of science. In Misiek [Mis95c], pages 195–204. ISBN 94-010-4206-3, 94-011-0461-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-011-0461-6_14.

Park:2001:PWD

[PS01a] Buhm Soon Park and DeWitt Stetten, Jr. A principle written in diagrams: The Aufbau Principle for molecules and its visual representations, 1927– 1932. In Klein [Kle01b], chapter 11, pages 179–198. ISBN 1-4020-0100-2, 94-015-9737-5. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 222; QD7. URL https:/

/link.springer.com/chapter/10.
1007/978-94-015-9737-1_11.

Puig-Samper:2001:DSP

[PS01b] Miguel Ángel Puig-Samper. Darwinism in Spanish physical anthropology. In The Reception of Darwinism in the Iberian World: Spain, Spanish America and Brazil [GPSR01], chapter 8, pages 127–141. ISBN 94-010-3885-6, 94-010-0602-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-010-
0602-6_8.

Panza:2015:IPI

[PS15] Marco Panza and Andrea Sereni.

On the indispensable premises of the indispensability argument. In From Logic to Practice: Italian Studies in the Philosophy of Mathematics [LPV15], chapter 13, pages 241–276. ISBN 3-319-10433-0, 3-319-10434-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QA8.4 .F76 2015; Q174 .B67 v.308. URL https://link.

springer.com/chapter/10.1007/
978-3-319-10434-8_13.

Psillos:2007:CEM

[Psi07] Stathis Psillos. Causal explanation and manipulation. In Rethinking Explanation [PY07], chapter 6, pages 93–107.

ISBN 1-4020-5580-3 (paperback), 1-4020-5581-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN BD237 .R48 2007; Q175.32.E97 R48

2007. URL https://link.springer.
com/chapter/10.1007/978-1-4020-
5581-2_6.

Parvu:2015:RSP

[PST15] Ilie Pârvu, Gabriel Sandu, and Iulian D. Toader. Romanian Studies in Philosophy of Science, volume 313 of Boston Studies in the Philosophy and History of Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2015. ISBN 3-319-16654-9 (hardcover), 3-319-16655-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). viii + 301 pp. LCCN ????

URL https://link.springer.com/
book/10.1007/978-3-319-16655-1.

Podolsky:1999:NCH

[PT99] Scott H. Podolsky and Alfred I. Tauber. Nietzsche’s conception of health: The idealization of struggle. In Nietzsche, Epistemology, and Philosophy of Science: Nietzsche and the Sciences II [Bab99a], chapter 23, pages 299–311. ISBN 90-481-5234-8, 94-017-2428-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-017-2428-
9_23.

Palmerino:2004:RGS

[PT04] Carla Rita Palmerino and J. M. M. H. Thijssen. The Reception of the Galilean Science of Motion in Seventeenth-Century Europe, volume 239 of Boston Studies in the Philosophy of Science. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 2004. ISBN 1-4020-2454-1, 1-4020-2455-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). ix + 287 pp. LCCN QC133 .R43 2004.

URL https://link.springer.com/
book/10.1007/978-1-4020-2455-9.

Pulte:2001:ONO

[Pul01] Helmut Pulte. Order of nature and orders of science. In Between Leibniz, Newton, and Kant: Philosophy and Science in the Eighteenth Century [Lef01a], chapter 4, pages 61–92. ISBN 90-481-5774-9, 94-015-9729-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-015-9729-6_4.

Pulte:2006:SBH

[Pul06] Helmut Pulte. The space between Helmholtz and Einstein: Moritz Schlick on spatial intuition and the foundations of geometry. In Hendricks et al. [HJLP06], pages 185–206. ISBN 1-4020-5194-8 (hardcover), 1-4020-5195-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175 .B69 v.251 2006; Q175 .I51377 2006.

URL https://link.springer.com/
chapter/10.1007/978-1-4020-5195-
1_6.

Pulte:2013:JFF

[Pul11] Helmut Pulte. J. F. Fries’ philosophy of science, the New Friesian School and the Berlin Group: On divergent scientific philosophies, difficult relations and missed opportunities. In The Berlin Group and the Philosophy of Logical Empiricism [MP13], chapter 3, pages 43–66. ISBN 94-007-5484-1 (hardcover), 94-007-5485-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175.32.A24 .M384 2011; Q175 .B47 2013. URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-007-5485-0_3.

Puri:1993:TCS

[Pur93] Avinash K. Puri. Tales from the classroom: The see-saw. In Correspondence, Invariance and Heuristics: Essays in Honour of Heinz Post [FK93], chapter 8, pages 159–169. ISBN 90-481-4229-6, 94-017-1185-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-017-1185-2_8.

Putnam:1967:IHE

[Put67a] Hilary Putnam. The ‘innateness hypothesis’ and explanatory models in linguistics. In Cohen and Wartofsky [CW67], pages 91–101. ISBN 90-277-0013-3, 94-010-3508-3. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175 .B731 v.3. URL https://link.

                                              springer.com/chapter/10.1007/
                                              978-94-010-3508-8_5.

Putnam:1967:C

[Put67b] Ruth Anna Putnam. Comments. In Cohen and Wartofsky [CW67], pages 173–180. ISBN 90-277-0013-3, 94-010-3508-3. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175

                                              .B731 v.3.  URL https://link.
                                              springer.com/chapter/10.1007/
                                              978-94-010-3508-8_9.

Putnam:1969:LE

[Put69] Hilary Putnam. Is logic empirical? In Cohen and Wartofsky [CW69a], pages 216–241. ISBN 94-010-3383-8, 94-010-3263-7, 94-010-3381-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-010-3381-7_5.

Putnam:1976:PLP

[Put76] Hilary Putnam. Philosophy of language and philosophy of science. In Cohen et al. [CHMV76], pages 603–610. ISBN 90-277-0647-6, 90-277-0648-4 (paperback), 94-010-1449-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-010-1449-6_34.

Putnam:1986:IHE

[Put86a] Hilary Putnam. The ‘innateness hypothesis’ and explanatory models in linguistics. In Cohen and Wartofsky [CW85], chapter 4, pages 41–51. ISBN 90-277-1971-3, 94-009-5345-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-009-5345-1_4.

Putnam:1986:LE

[Put86b] Hilary Putnam. Is logic empirical?

In Cohen and Wartofsky [CW85], chapter 7, pages 75–100. ISBN 90-277-1971-3, 94-009-5345-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-009-5345-1_7.

Putnam:1986:MOM

[Put86c] Hilary Putnam. Meaning and our mental life. In Ullmann-Margalit [UM86a], chapter 3, pages 17–32. ISBN 90-277-2159-9 (paperback), 90-277-2158-0 (hardcover), 94-009-5496-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 94

Q175. URL https://link.springer.
com/chapter/10.1007/978-94-009-
5496-0_3.

Pentzopoulou-Valalas:1990:ECE

[PV90] Theresa Pentzopoulou-Valalas. Experience and causal explanation in medical empiricism. In Greek Studies in the Philosophy and History of Science [Nic90], chapter 6, pages 91–107. ISBN 94-010-7403-8, 94-009-2015-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-009-2015-6_6.

Price:2017:DQS

[PW17] Huw Price and Ken Wharton. Dispelling the quantum spooks: A clue that Einstein missed? In Time of Nature and the Nature of Time: Philosophical Perspectives of Time in Natural Sciences [BH17b], chapter 7, pages 123–137. ISBN 3-319-53723-7 (hardcover), 3-319-53725-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC173.59.S65 T56

2017. URL https://link.springer.
com/chapter/10.1007/978-3-319-
53725-2_7.

Persson:2007:RE

[PY07] Johannes Persson and Petri Ylikoski. Rethinking Explanation, volume 252 of Boston Studies in the Philosophy of Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2007. ISBN 1-4020-5580-3 (paperback), 1-4020-5581-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). xiv + 207 pp. LCCN BD237 .R48 2007; Q175.32.E97 R48 2007. URL

https://link.springer.com/book/
10.1007/978-1-4020-5581-2.

Pyenson:1987:RRG

[Pye87] Lewis Pyenson. The relativity revolution in Germany. In Glick [Gli87a], chapter 3, pages 59–111. ISBN 90-277-2498-9, 94-009-3875-6. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 103; QC173.5. US$79.00

(US). URL https://link.springer.
com/chapter/10.1007/978-94-009-
3875-5_3.

Pyenson:1992:TSE

[Pye92] Lewis Pyenson. Typologie des stratégies d’expansion en sciences exactes. (French) [Typology of expansion strategies in exact sciences]. In Science and Empires: Historical Studies about Scientific Development and European Expansion [PJM92], chapter 22, pages 211–217. ISBN 0-7923-1518-9 (hardcover), 94-010-5145-3, 94-011-2594-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 136. URL https://link.

springer.com/chapter/10.1007/
978-94-011-2594-9_22.

Qiongbing:1996:NSF

[Qio96] Wu Qiongbing. Is natural science free from morality? In Cohen et al. [CHR96], pages 245–264. ISBN 90-481-4493-0, 94-015-8638-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-015-8638-2_17.

Quay:1984:TSP

[Qua84] Paul M. Quay. Temporality and the structure of physics as human endeavor. In Physical Sciences and History of Physics [CW84c], chapter 11, pages 199–230. ISBN 94-009-7180-X, 94-009-7178-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .P497 1984. URL https://link.

springer.com/chapter/10.1007/
978-94-009-7178-3_11.

Quesada:1995:KD

[Que95] Francisco Miró Quesada. Knowledge and destiny. In The Concept of Knowledge: The Ankara Seminar [KC95], chapter 17, pages 219–229. ISBN 90-481-4495-7, 94-017-3263-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-017-3263-5_17.

Queralto:1997:SIN

[Que97] Ramón Queraltó. Since indeterminacy: The new picture of the physical world at the end of modernity. In Ginev and Cohen [GC97], chapter 14, pages 253–262. ISBN 0-7923-4444-8 (hardcover), 94-011-5788-X. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 192. URL https:/

                                              /link.springer.com/chapter/10.
                                              1007/978-94-011-5788-9_14.

Queralto:2000:SRO

[Que00] Ramón Queraltó. Scientific realism, objectivity, and ‘technological realism’. In The Reality of the Unobservable: Observability, Unobservability and Their Impact on the Issue of Scientific Realism [AP00b], chapter 10, pages 145–151. ISBN 90-481-5458-8, 94-015-9391-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

                                              ???? URL https://link.springer.
                                              com/chapter/10.1007/978-94-015-
                                              9391-5_10.

Quine:1963:C

[Qui63] Willard Van Orman Quine. Comments. In Wartofsky [War63], pages 97–104. ISBN 94-010-3265-3, 94-010-3263-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174

                                              .B678 1970. URL https://link.

springer.com/chapter/10.1007/
978-94-010-3263-6_8.

Quinn:1974:WDR

[Qui74] Philip L. Quinn. What Duhem really meant. In Methodological and Historical Essays in the Natural and Social Sciences [CW74b], chapter 3, pages 33–56. ISBN 90-277-0378-7, 94-010-2128-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 14. URL https://link.

springer.com/chapter/10.1007/
978-94-010-2128-9_3.

Quine:1976:WPO

[Qui76]  Willard Quine.  Whither physical

objects? In Essays in Memory of Imre Lakatos [CFW76], chapter 29, pages 497–504. ISBN 90-277-0655-7, 94-010-1451-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-010-
1451-9_29.

Quintanilla:1982:MFC

[Qui82] Miguel A. Quintanilla. Materialist foundations of critical rationalism. In Scientific Philosophy Today: Essays in Honor of Mario Bunge [AC82], chapter 14, pages 225–237. ISBN 90-277-1262-X, 90-277-1263-8, 94-009-8462-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 67. URL https:/

/link.springer.com/chapter/10.
1007/978-94-009-8462-2_14.

Quinn:1983:GDM

[Qui83] Philip L. Quinn. Grünbaum on determinism and the moral life. In Cohen and Laudan [CL83], chapter 7, pages 129–151. ISBN 94-009-7055-2, 94-009-7057-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-009-
7055-7_7.

Quintanilla:1996:IT

[Qui96] Miguel A. Quintanilla. The incompleteness of technics. In Munévar [Mun96], chapter 4, pages 89–102. ISBN 94-010-6622-1, 94-009-0305-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-009-0305-0_4.

Quine:2000:YIE

[Qui00a] W. V. Quine. I, you, and it: An epistemological triangle. In Knowledge, Language and Logic: Questions for Quine [OK00], chapter 1, pages 1–6. ISBN 1-4020-0253-X, 94-011-3933-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-011-3933-
5_1.

Quine:2000:QR

[Qui00b] W. V. Quine. Quine’s responses. In Knowledge, Language and Logic: Questions for Quine [OK00], chapter 21, pages 407–430. ISBN 1-4020-0253-X, 94-011-3933-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic).

LCCN ????  URL https://link.
springer.com/chapter/10.1007/
978-94-011-3933-5_21.

Radnitzky:1978:OCS

[RA78a] Gerard Radnitzky and Gunnar Andersson. Objective criteria of scientific progress? Inductivism, falsificationism, and relativism. In Progress and Rationality in Science [RA78b], chapter 1, pages 3–19. ISBN 90-277-0922-X, 94-009-9866-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-009-
9866-7_1.

Radnitzky:1978:PRS

[RA78b] Gerard Radnitzky and Gunnar Andersson. Progress and Rationality in Science, volume 58 of Boston Studies in the Philosophy of Science. D. Reidel, Dordrecht, The Netherlands; Boston, MA, USA; Lancaster, UK; Tokyo, Japan, 1978. ISBN 90-277-0922-X, 94-009-9866-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). 426 (est.) pp. LCCN ???? URL

https://link.springer.com/book/
10.1007/978-94-009-9866-7.

Radnitzky:1979:SDS

[RA79] Gerard Radnitzky and Gunnar Andersson. The Structure and Development of Science, volume 59 of Boston Studies in the Philosophy of Science. D. Reidel, Dordrecht, The Netherlands; Boston, MA, USA; Lancaster, UK; Tokyo, Japan, 1979. ISBN 90-277-0995-5, 94-009-9459-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). ix + 293 pp. LCCN Q174 .S778 1979. URL

https://link.springer.com/book/
10.1007/978-94-009-9459-1.

Ruiz:2001:DHC

[RA01] Rosaura Ruiz and Francisco J. Ayala.

Darwinism: Its hard core. In The Reception of Darwinism in the Iberian World: Spain, Spanish America and Brazil [GPSR01], chapter 15, pages 239–261. ISBN 94-010-3885-6, 94-010-0602-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-010-
0602-6_15.

Reif-Acherman:2014:DIM

[RA14]  Simón Reif-Acherman.  Domestic

ice-making machines 1830–1930. In History of Artificial Cold, Scientific, Technological and Cultural Issues [Gav14b], chapter 8, pages 135–170. ISBN 94-007-7198-3 (hardcover), 94-007-7199-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC278 H57 2014; TP494 .H577 2014.

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-007-7199-
                                              4_8.

Radnitzky:1976:PPS

[Rad76] Gerard Radnitzky. Popperian philosophy of science as an antidote against relativism. In Essays in Memory of Imre Lakatos [CFW76], chapter 30, pages 505–546. ISBN 90-277-0655-7, 94-010-1451-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

                                              ???? URL https://link.springer.
                                              com/chapter/10.1007/978-94-010-
                                              1451-9_30.

Radnitzky:1979:JTV

[Rad79] Gerard Radnitzky. Justifying a theory vs. giving good reasons for preferring a theory. In The Structure and Development of Science [RA79], chapter 9, pages 213–256. ISBN 90-277-0995-5, 94-009-9459-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .S778 1979. URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-009-9459-1_9.

Radnitzky:1981:PRR

[Rad81]  Gerard Radnitzky.  Progress and

rationality in research (science from the viewpoint of Popperian methodology). In Grmek et al. [GCC81b], pages 43–102. ISBN 90-277-1123-2, 94-010-1284-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-010-
1284-3_3.

Radnitzky:1982:APC

[Rad82] Gerard Radnitzky. Analytic philosophy as the confrontation between Wittgensteinians and Popper. In Scientific Philosophy Today: Essays in Honor of Mario Bunge [AC82], chapter 15, pages 239–286. ISBN 90-277-1262-X, 90-277-1263-8, 94-009-8462-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 67.

URL https://link.springer.com/
chapter/10.1007/978-94-009-8462-
2_15.

Radnitzky:1989:FLE

[Rad89a] Gerard Radnitzky. Falsificationism looked at from an “economic” point of view. In Imre Lakatos and Theories of Scientific Change [GGN89], chapter 28, pages 383–395. ISBN 94-010-7860-2, 94-009-3025-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-009-
3025-4_28.

Radnitzky:1989:SPM

[Rad89b] Gerard Radnitzky. Science as a particular mode of thinking and the ‘taming of the state’. In Freedom and Rationality: Essays in Honor of John Watkins From his Colleagues and Friends [DJ89], chapter 10, pages 163–181. ISBN 94-010-7571-9, 94-009-2380-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-009-2380-5_10.

Radev:2003:TWH

[Rad03] Slavian Radev. Three words: Hypertext and argumentation readings of Tractatus Logico-Philosophicus. In Bulgarian Studies in the Philosophy of Science [Gin03a], chapter 11, pages 163–166. ISBN 1-4020-1496-1, 94-017-0961-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

                                              ???? URL https://link.springer.
                                              com/chapter/10.1007/978-94-017-
                                              0961-3_11.

Radder:2012:ENS

[Rad12a] Hans Radder. Experimentation in the natural sciences. In The Material Realization of Science: From Habermas to Experimentation and Referential Realism [Rad12d], chapter 3, pages 53– 71. ISBN 94-007-4106-5 (hardcover), 94-007-4107-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175 .R33 2012. URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-007-4107-2_3.

Radder:2012:HOT

[Rad12b] Hans Radder. Habermas on objectivity and truth: Analysis and critique. In The Material Realization of Science: From Habermas to Experimentation and Referential Realism [Rad12d], chapter 2, pages 31–50. ISBN 94-007-4106-5 (hardcover), 94-007-4107-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175 .R33

                                              2012. URL https://link.springer.
                                              com/chapter/10.1007/978-94-007-
                                              4107-2_2.

Radder:2012:HPN

[Rad12c] Hans Radder. Habermas’s philosophy of the natural sciences. In The Material Realization of Science: From Habermas to Experimentation and Referential Realism [Rad12d], chapter 1, pages 3–29. ISBN 94-007-4106-5 (hardcover), 94-007-4107-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175 .R33 2012. URL https:

//link.springer.com/chapter/10.
1007/978-94-007-4107-2_1.

Radder:2012:MRS

[Rad12d] Hans Radder. The Material Realization of Science: From Habermas to Experimentation and Referential Realism, volume 294 of Boston Studies in the Philosophy of Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2012. ISBN 94-007-4106-5 (hardcover), 94-007-4107-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). xix + 205 pp. LCCN Q175

.R33 2012. URL http://www.loc.
gov/catdir/enhancements/fy1313/
2012935860-d.html;   http://www.
loc.gov/catdir/enhancements/
fy1313/2012935860-t.html;
https://link.springer.com/book/
10.1007/978-94-007-4107-2.

Radder:2012:SAT

[Rad12e] Hans Radder. Specification and application: Two case studies from the history and philosophy of quantum mechanics. In The Material Realization of Science: From Habermas to Experimentation and Referential Realism [Rad12d], chapter 5, pages 113–154. ISBN 94-007-4106-5 (hardcover), 94-007-4107-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175 .R33 2012. URL https://link.

springer.com/chapter/10.1007/
978-94-007-4107-2_5.

Radder:2012:VRR

[Rad12f] Hans Radder. Verifiability and reference, relativism and realism. In The Material Realization of Science: From Habermas to Experimentation and Referential Realism [Rad12d], chapter 4, pages 73–111. ISBN 94-007-4106-5 (hardcover), 94-007-4107-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175 .R33 2012.

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-007-4107-
                                              2_4.

Raff:2015:EDE

[Raf15] Rudolf A. Raff. Evo-devo and the evolution of marine larvae: From the modern world to the dawn of the Metazoa. In Conceptual Change in Biology: Scientific and Philosophical Perspectives on Evolution and Development [Lov15b], chapter 11, pages 243–258. ISBN 94-017-9411-1, 94-017-9412-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QH331 C66 2015; QH491; QH492 .C66 2014.

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-017-9412-
                                              1_11.

Raina:2010:FJM

[Rai10] Dhruv Raina. The French Jesuit manuscripts on Indian astronomy: The narratology and mystery surrounding a Late Seventeenth–Early Eighteenth Century project. In Looking at it from Asia: the Processes that Shaped the Sources of History of Science [BE10b], chapter 4, pages 115–140. ISBN 90-481-3675-X (hardcover), 90-481-3676-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q127.A65 L66

                                              2010. URL https://link.springer.
                                              com/chapter/10.1007/978-90-481-
                                              3676-6_4.

Raina:2011:TBW

[Rai11] Dhruv Raina. Travelling both ways: The adaptation of disciplines, scientific textbooks and institutions. In Science between Europe and Asia: Historical Studies on the Transmission, Adoption and Adaptation of Knowledge [GR11b], chapter 11, pages 165–176. ISBN 90-481-9968-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q125 .S354 2010; Q175.52.A78 S35

2010. URL https://link.springer.
com/chapter/10.1007/978-90-481-
9968-6_11.

Ramberg:2001:PTF

[Ram01] Peter J. Ramberg. Paper tools and fictional worlds: Prediction, synthesis and auxiliary hypotheses in chemistry. In Klein [Kle01b], chapter 5, pages 61– 78. ISBN 1-4020-0100-2, 94-015-9737-5. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 222;

QD7. URL https://link.springer.
com/chapter/10.1007/978-94-015-
9737-1_5.

Rantala:1989:CR

[Ran89] V. Rantala. Counterfactual reduction.

In Imre Lakatos and Theories of Scientific Change [GGN89], chapter 25, pages 347–360. ISBN 94-010-7860-2, 94-009-3025-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-009-
3025-4_25.

Rapp:1981:APT

[Rap81a] Friedrich Rapp. Analytical Philosophy of Technology, volume 63 of Boston Studies in the Philosophy of Science. D. Reidel, Dordrecht, The Netherlands; Boston, MA, USA; Lancaster, UK; Tokyo, Japan, 1981. ISBN 90-277-1222-0, 94-009-8412-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). xiv + 202 pp. LCCN ????

                                              URL https://link.springer.com/
                                              book/10.1007/978-94-009-8412-7.

Rapp:1981:DVC

[Rap81b] Friedrich Rapp. Differing versions of the concept of ‘technology’. In Analytical Philosophy of Technology [Rap81a], chapter 2, pages 23–36. ISBN 90-277-1222-0, 94-009-8412-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-009-8412-7_2.

Rapp:1981:MA

[Rap81c] Friedrich Rapp. Methodological analysis. In Analytical Philosophy of Technology [Rap81a], chapter 3, pages 37–65. ISBN 90-277-1222-0, 94-009-8412-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

                                              ???? URL https://link.springer.
                                              com/chapter/10.1007/978-94-009-
                                              8412-7_3.

Rapp:1981:RMT

[Rap81d] Friedrich Rapp. The road to modern technology. In Analytical Philosophy of Technology [Rap81a], chapter 4, pages 66–117. ISBN 90-277-1222-0, 94-009-8412-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

                                              ???? URL https://link.springer.
                                              com/chapter/10.1007/978-94-009-
                                              8412-7_4.

Rapp:1981:SPT

[Rap81e] Friedrich Rapp. Stages of the philosophy of technology. In Analytical Philosophy of Technology [Rap81a], chapter 1, pages 1–22. ISBN 90-277-1222-0, 94-009-8412-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-009-8412-7_1.

Rapp:1981:TW

[Rap81f] Friedrich Rapp. The technological world. In Analytical Philosophy of Technology [Rap81a], chapter 5, pages 118–186. ISBN 90-277-1222-0, 94-009-8412-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-009-
8412-7_5.

Rapp:1982:DR

[Rap82] Friedrich Rapp. Distrust of reason. In Scientific Philosophy Today: Essays in Honor of Mario Bunge [AC82], chapter 16, pages 287–298. ISBN 90-277-1262-X, 90-277-1263-8, 94-009-8462-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 67. URL https:/

/link.springer.com/chapter/10.
1007/978-94-009-8462-2_16.

Rapp:1983:PTA

[Rap83] Friedrich Rapp. The prospects for technology assessment. In Philosophy and Technology [DR83], chapter 10, pages 141–150. ISBN 94-009-7126-5, 94-009-7124-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-009-
7124-0_10.

Rapp:1986:TLI

[Rap86] Friedrich Rapp. The theory–ladenness of information. In Philosophy and Technology II: Information Technology and Computers in Theory and Practice [MH86], chapter 4, pages 49–62. ISBN 94-010-8510-2, 94-009-4512-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-009-4512-8_4.

Rapp:1997:DS

[Rap97] Friedrich Rapp. The dynamics of science. In Ginev and Cohen [GC97], chapter 15, pages 263–274. ISBN 0-7923-4444-8 (hardcover), 94-011-5788-X. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 192. URL https://link.

                                              springer.com/chapter/10.1007/
                                              978-94-011-5788-9_15.

Rashed:1974:ALA

[Ras74] Roshdi Rashed. Algébre et linguistique: l’analyse combinatoire dans la science arabe. (French) [Algebra and linguistics: Combinatorial analysis in Arabic science]. In Seeger and Cohen [SC74b], pages 383–399. ISBN 90-277-0390-6, 90-277-0376-0 (paperback), 94-010-2126-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 11. URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-010-2126-5_22.

Rashed:1975:RAX

[Ras75] Roshdi Rashed. Recommencements de l’algèbre aux XIe et XIIe siècles. (French) [repetitions of algebra in the Eleventh and Twelfth Centuries]. In Murdoch and Sylla [MS75b], pages 33– 60. ISBN 90-277-0587-9, 94-010-1781-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN B721 .I584

                                              1975. URL https://link.springer.
                                              com/chapter/10.1007/978-94-010-
                                              1781-7_2.

Rashed:1992:SCS

[Ras92]  Roshdi Rashed.  Science classique

et science moderne a l’époque de l’expansion de la science européenne. (French) [Classical science and modern science at the time of the expansion of European science]. In Science and Empires: Historical Studies about Scientific Development and European Expansion [PJM92], chapter 4, pages 19–30. ISBN 0-7923-1518-9 (hardcover), 94-010-5145-3, 94-011-2594-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 136.

URL https://link.springer.com/
chapter/10.1007/978-94-011-2594-
9_4.

Rashed:1994:ASA

[Ras94a] Roshdi Rashed. Analysis and synthesis according to Ibn Al-Haytham. In Artifacts, Representations and Social Practice: Essays for Marx Wartofsky [GC94], chapter 8, pages 121–140. ISBN 94-010-4390-6, 94-011-0902-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN B73 .A77 1994.

URL https://link.springer.com/
chapter/10.1007/978-94-011-0902-
4_8.

Rashed:1994:BA

[Ras94b] Roshdi Rashed. The beginnings of algebra. In The Development of Arabic Mathematics: Between Arithmetic and Algebra [Ras94c], chapter 2, pages 8–84. ISBN 90-481-4338-1, 94-017-3274-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QA27.A67 R37 1994. URL https:

//link.springer.com/chapter/10.
1007/978-94-017-3274-1_2.

Rashed:1994:DAM

[Ras94c] Roshdi Rashed. The Development of Arabic Mathematics: Between Arithmetic and Algebra, volume 156 of Boston Studies in the Philosophy of Science. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 1994. ISBN 90-481-4338-1, 94-017-3274-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). x + 372 pp. LCCN QA27.A67 R37 1994. URL

                                              https://link.springer.com/book/
                                              10.1007/978-94-017-3274-1.

Rashed:1994:I

[Ras94d] Roshdi Rashed. Introduction. In The Development of Arabic Mathematics: Between Arithmetic and Algebra [Ras94c], chapter 1, pages 1–7. ISBN 90-481-4338-1, 94-017-3274-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QA27.A67 R37

                                              1994. URL https://link.springer.
                                              com/chapter/10.1007/978-94-017-
                                              3274-1_1.

Rashed:1994:NTC

[Ras94e] Roshdi Rashed. Number theory and combinatorial analysis. In The Development of Arabic Mathematics: Between Arithmetic and Algebra [Ras94c], chapter 5, pages 205–331. ISBN 90-481-4338-1, 94-017-3274-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QA27.A67 R37

                                              1994. URL https://link.springer.
                                              com/chapter/10.1007/978-94-017-
                                              3274-1_5.

Rashed:1994:NA

[Ras94f] Roshdi Rashed. Numerical analysis. In The Development of Arabic Mathematics: Between Arithmetic and Algebra [Ras94c], chapter 3, pages 85–146. ISBN 90-481-4338-1, 94-017-3274-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QA27.A67 R37 1994. URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-017-3274-1_3.

Rashed:1994:NE

[Ras94g] Roshdi Rashed. Numerical equations. In The Development of Arabic Mathematics: Between Arithmetic and Algebra [Ras94c], chapter 4, pages 147–204. ISBN 90-481-4338-1, 94-017-3274-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QA27.A67 R37 1994. URL https:

//link.springer.com/chapter/10.
1007/978-94-017-3274-1_4.

Rashed:1995:CSB

[Ras95] Roshdi Rashed. Conic sections and burning mirrors: An example of the application of ancient and classical mathematics. In Gavroglu et al. [GSW95c], chapter 20, pages 357–376. ISBN 0-7923-2991-0 (set), 0-7923-2988-0, 94-017-2658-2. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 163. URL https:/

/link.springer.com/chapter/10.
1007/978-94-017-2658-0_20.

Rashed:2000:ASM

[Ras00] Roshdi Rashed. Al-Sijzī and Maimonides: A mathematical and philosophical commentary on Proposition II–14 in Apollonius’ conic sections. In Maimonides and the Sciences [CL00], chapter 9, pages 159–172. ISBN 90-481-5348-4, 94-017-2128-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-017-2128-8_9.

Ratliff:1970:MCA

[Rat70] Floyd Ratliff. On Mach’s contributions to the analysis of sensations. In Ernst Mach: Physicist and Philosopher [CS70], chapter 2, pages 23–41. ISBN 90-481-8318-9, 90-277-0016-8, 94-017-1462-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-017-
1462-4_2.

Ratner:1988:JDP

[Rat88] Sidney Ratner. John Dewey’s philosophy of war and peace. In Philosophy, History and Social Action: Essays in Honor of Lewis Feuer with an autobiographic essay by Lewis Feuer [HOO88c], chapter 19, pages 373–390. ISBN 94-010-7793-2, 94-009-2873-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-009-2873-
                                              2_19.

Ratcliff:2018:FEM

[Rat18] Marc Ratcliff. Fictitious empiricism, material experiments. Conditions for thinking the Enlightenment “issue of empiricism”. In What Does it Mean to be an Empiricist?: Empiricisms in Eighteenth Century Sciences [BR18b], chapter 3, pages 31–43. ISBN 3-319-69858-3 (hardcover), 3-319-69860-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN B816

                                              .W43 2018.  URL https://link.
                                              springer.com/chapter/10.1007/
                                              978-3-319-69860-1_3.

Rauh:1991:NKE

[Rau91] Hans-Christoph Rauh. Neo-Kantianism and epistemology: On the formation of a philosophical discipline in Nineteenth-Century Germany. In Woodward and Cohen [WC91], pages 223–234. ISBN 94-010-5395-2, 94-011-3164-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-011-3164-3_20.

Ravetz:1991:ICP

[Rav91] Jerome R. Ravetz. Ideological commitments in the philosophy of science. In Beyond Reason: Essays on the Philosophy of Paul Feyerabend [Mun91a], chapter 16, pages 355–377. ISBN 94-010-5406-1, 94-011-3188-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-011-3188-9_16.

Ravesteijn:2018:DCW

[Rav18] Wim Ravesteijn. The Dao of Chinese water management and development: Challenges and perspectives. In Mitcham et al. [MLNZ18], chapter 22, pages 261–269. ISBN 3-319-62448-2, 3-319-62450-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN TA15 .P45 2018. URL https:

//link.springer.com/chapter/10.
1007/978-3-319-62450-1_22.

Rayski:1982:CPP

[Ray82] Jerzy Rayski. Controversial problems of the probabilistic interpretation of quantum phenomena. In Polish Essays in the Philosophy of the Natural Sciences [Kra82c], chapter 26, pages 299–308. ISBN 94-009-7705-0 (e-book), 90-277-1287-5. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-009-7705-
1_26.

Ray:2000:RML

[Ray00] Greg Ray. De Re modality: Lessons from Quine. In Knowledge, Language and Logic: Questions for Quine [OK00], chapter 19, pages 347–365. ISBN 1-4020-0253-X, 94-011-3933-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-011-3933-
5_19.

Raz:2016:GTN

[Räz16] Tim Räz. Gone till November: A disagreement in Einstein scholarship. In Sauer and Scholl [SS16b], chapter 9, pages 179–200. ISBN 3-319-30227-2, 3-319-30229-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q126.8 .P45 2016. URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-3-319-30229-4_9.

Root-Bernstein:1991:SNP

[RB91] Robert S. Root-Bernstein. Self, nonself and the paradoxes of autoimmunity. In Organism and the Origins of Self [Tau91i], chapter 8, pages 159–209. ISBN 0-7923-1185-X, 94-011-3406-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QR180-189.5.

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-011-3406-
                                              4_8.

Root-Bernstein:1997:SAS

[RB97] Robert S. Root-Bernstein. The sciences and arts share a common creative aesthetic. In The Elusive Synthesis: Aesthetics and Science [Tau97a], chapter 3, pages 49–82. ISBN 0-7923-4763-3, 0-7923-3904-5, 94-009-1786-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-009-1786-6_3.

Rowbottom:2009:WAP

[RB09] Darrell P. Rowbottom and Otávio Bueno. Why advocate pancritical rationalism? In Rethinking Popper [PC09], chapter 7, pages 81–89. ISBN 1-4020-9337-3, 1-4020-9338-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN B1649.P64 R48 2009; Q174 .B67 2009. URL https:

//link.springer.com/chapter/10.
1007/978-1-4020-9338-8_7.

Ramirez:1995:MSH

[RC95] Santiago Ramirez and Robert S. Cohen.

Mexican Studies in the History and Philosophy of Science, volume 172 of Boston Studies in the Philosophy of Science. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 1995. ISBN 94-010-6535-7 (hardcover), 94-009-0109-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). 292 (est.) pp. LCCN ???? URL

https://link.springer.com/book/
10.1007/978-94-009-0109-4.

Renn:2003:RFR

[RDS03] Jürgen Renn, Lindy Divarci, and Petra Schröter, editors. Revisiting the Foundations of Relativistic Physics: Festschrift in Honor of John Stachel, volume 234 of Boston Studies in the Philosophy of Science. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 2003. ISBN 1-4020-1284-5, 1-4020-1285-3 (paperback). ISSN 0068-0346 (print), 2214-7942 (electronic). xxii + 649 pp. LCCN Q174 .B67 vol.

234. URL http://www.gbv.de/dms/
bowker/toc/9781402012846.pdf;
http://www.loc.gov/catdir/
enhancements/fy0813/2003046757-
d.html;   http://www.loc.gov/
catdir/enhancements/fy0813/
2003046757-t.html;   https://
link.springer.com/book/10.1007/
978-94-010-0111-3.

Renn:2018:MMC

[RDS^{+}18] Jürgen Renn, Peter Damerow, Matthias Schemmel, Christoph Lehner, and Matteo Valleriani. Mental models as cognitive instruments in the transformation of knowledge. In Feldhay et al. [FRSV18a], chapter 1, pages 3–28. ISBN 3-319-90343-8 (hardcover), 3-319-90345-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic).

LCCN QA802. URL https://link.
springer.com/chapter/10.1007/
978-3-319-90345-3_1.

Recanati:2000:OA

[Rec00] François Recanati. Opacity and the attitudes. In Knowledge, Language and Logic: Questions for Quine [OK00], chapter 20, pages 367–406. ISBN 1-4020-0253-X, 94-011-3933-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-011-3933-5_20.

Reck:2013:HCC

[Rec11] Erich H. Reck. Hempel, Carnap, and the covering law model. In The Berlin Group and the Philosophy of Logical Empiricism [MP13], chapter 15, pages 311–324. ISBN 94-007-5484-1 (hardcover), 94-007-5485-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175.32.A24 .M384 2011; Q175 .B47 2013. URL https:

//link.springer.com/chapter/10.
1007/978-94-007-5485-0_15.

Redhead:1993:EPS

[Red93] Michael Redhead. Is the end of physics in sight? In Correspondence, Invariance and Heuristics: Essays in Honour of Heinz Post [FK93], chapter 16, pages 327–341. ISBN 90-481-4229-6, 94-017-1185-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-017-
1185-2_16.

Redei:2003:TAT

[Réd03] Miklós Rédei. Thinking about thought experiments in physics: Comment on “Experiments and Thought Experiments in Natural Science”. In Descartes’s Mathematical Thought [Sas03d], chapter 16, pages 237–241. ISBN 90-481-6487-7, 94-017-1225-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/0-306-48123-5_16.

Reed:1987:WIM

[Ree87a] Edward S. Reed. Why ideas are not in the mind: An introduction to ecological epistemology. In Shimony and Nails [SN87], pages 215–229. ISBN 94-010-8168-9, 94-009-3735-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-009-3735-2_12.  See

comment [Shi87c].

Reeves:1987:EPE

[Ree87b] Barbara J. Reeves. Einstein politicized:

The early reception of relativity in Italy. In Glick [Gli87a], chapter 6, pages 189–229. ISBN 90-277-2498-9, 94-009-3875-6. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 103; QC173.5. US$79.00

(US). URL https://link.springer.
com/chapter/10.1007/978-94-009-
3875-5_6.

Rees:2014:IRT

[Ree14] Jonathan Rees. The invention of refrigerated transport and the development of the international dressed meat trade. In History of Artificial Cold, Scientific, Technological and Cultural Issues [Gav14b], chapter 13, pages 251–265.

ISBN 94-007-7198-3 (hardcover), 94-007-7199-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC278 H57 2014; TP494 .H577 2014.

URL https://link.springer.com/
chapter/10.1007/978-94-007-7199-
4_13.

Regan:2013:ARE

[Reg13] Tom Regan. Animal rights and environmental ethics. In Bergandi [Ber13c], chapter 8, pages 117–126. ISBN 94-007-5066-8, 94-007-5067-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN GE42 .S77 2013; Q175 .B73 v. 296. URL https:/

/link.springer.com/chapter/10.
1007/978-94-007-5067-8_8.

Rehberg:1999:NTC

[Reh99] Andrea Rehberg. Nietzsche’s transvaluation of causality. In Nietzsche, Epistemology, and Philosophy of Science: Nietzsche and the Sciences II [Bab99a], chapter 21, pages 279–286. ISBN 90-481-5234-8, 94-017-2428-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-017-2428-
9_21.

Reich:1996:EVC

[Rei96] Karin Reich. The emergence of vector calculus in physics: The early decades. In Schubring [Sch96f], pages 197–210. ISBN 90-481-4758-1, 94-015-8753-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QA8.9-10.3. URL https://link.

springer.com/chapter/10.1007/
978-94-015-8753-2_16.

Reinhardt:2011:EMM

[Rei11] Carsten Reinhardt. Expertise in methods, methods of expertise. In Science in the Context of Application [CN11a], chapter 10, pages 143–159. ISBN 90-481-9050-9, 90-481-9051-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q180.55.M4 S385

2011. URL https://link.springer.
com/chapter/10.1007/978-90-481-
9051-5_10.

Renn:1990:GTE

[Ren90] Jürgen Renn. Galileo’s theorem of equivalence: The missing keystone of his theory of motion. In Levere and Shea [LS90], chapter 3, pages 77–103. ISBN 0-7923-0420-9, 94-009-1878-X. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 120.

URL https://link.springer.com/
chapter/10.1007/978-94-009-1878-
8_3.

Renn:1995:HEI

[Ren95] Jürgen Renn. Historical epistemology and interdisciplinarity. In Gavroglu et al. [GSW95c], chapter 14, pages 241–251. ISBN 0-7923-2991-0 (set), 0-7923-2988-0, 94-017-2658-2. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 163.

URL https://link.springer.com/
chapter/10.1007/978-94-017-2658-
0_14.

Renzong:1996:HKW

[Ren96] Qiu Renzong. How to know what rises up is the moon? — on the concept of realism and the irrelevancy of quantum mechanics to the debate on realism vs. antirealism. In Cohen et al. [CHR96], pages 55–73. ISBN 90-481-4493-0, 94-015-8638-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-015-
8638-2_5.

Renn:2007:CPD

[Ren07a] Jürgen Renn. Classical physics in disarray: The emergence of the riddle of gravitation. In The Genesis of General Relativity: Volume 1. Einstein’s Zurich Notebook: Introduction and Source [JNR^{+}07b], pages 21–80. ISBN 1-4020-3999-9 (hardcover), 1-4020-4000-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 no. 250 v. 1; QC173.6 .G469 2007.

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-1-4020-4000-
                                              9_2.

Renn:2007:P

[Ren07b] Jürgen Renn. Preface. In The Genesis of General Relativity: Volume 1. Einstein’s Zurich Notebook: Introduction and Source [JNR^{+}07b], pages 1– 6. ISBN 1-4020-3999-9 (hardcover), 1-4020-4000-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 no. 250 v. 1; QC173.6 .G469 2007.

Renn:2007:SAS

[Ren07c] Jürgen Renn. The summit almost scaled: Max Abraham as a pioneer of a relativistic theory of gravitation. In Renn et al. [RSS^{+}07a], pages 305–330. ISBN 1-4020-3999-9. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 no. 250 v. 3; QC173.6 .G469

                                              2007. URL https://link.springer.
                                              com/chapter/10.1007/978-1-4020-
                                              4000-9_22.

Renn:2007:TWG

[Ren07d] Jürgen Renn. The third way to General Relativity: Einstein and Mach in context. In Renn et al. [RSS^{+}07a], pages 21–75. ISBN 1-4020-3999-9. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 no. 250 v. 3; QC173.6 .G469 2007. URL https:

//link.springer.com/chapter/10.
1007/978-1-4020-4000-9_11.

Renn:2015:HSG

[Ren15] Jürgen Renn. The history of science and the globalization of knowledge. In Arabatzis et al. [ARS15], chapter 16, pages 241–252. ISBN 3-319-14552-5 (paperback), 3-319-14553-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q125 .R45 2015.

URL https://link.springer.com/
chapter/10.1007/978-3-319-14553-
2_16.

Rescher:1979:SIR

[Res79] Nicholas Rescher. Some issues regarding the completeness of science and the limits of scientific knowledge. In The Structure and Development of Science [RA79], chapter 2, pages 19–40. ISBN 90-277-0995-5, 94-009-9459-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .S778 1979.

URL https://link.springer.com/
chapter/10.1007/978-94-009-9459-
1_2.

Rescher:1983:UFS

[Res83] Nicholas Rescher. The unpredictability of future science. In Cohen and Laudan [CL83], chapter 8, pages 153–168. ISBN 94-009-7055-2, 94-009-7057-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-009-7055-7_8.

Rescher:1989:PG

[Res89] Nicholas Rescher. The philosophers of gambling. In An Intimate Relation: Studies in the History and Philosophy of Science Presented to Robert E. Butts on his 60th Birthday [BM89], chapter 10, pages 203–220. ISBN 94-010-7546-8, 94-009-2327-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-009-2327-0_10.

Rescher:1997:LLR

[Res97] Nicholas Rescher. The law of logarithmic returns and its implications. In Ginev and Cohen [GC97], chapter 16, pages 275–287. ISBN 0-7923-4444-8 (hardcover), 94-011-5788-X. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 192.

URL https://link.springer.com/
chapter/10.1007/978-94-011-5788-
9_16.

Rescher:2013:BGU

[Res11] Nicholas Rescher. The Berlin Group and the USA: A narrative of personal interactions. In The Berlin Group and the Philosophy of Logical Empiricism [MP13], chapter 2, pages 33–39. ISBN 94-007-5484-1 (hardcover), 94-007-5485-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175.32.A24 .M384 2011; Q175 .B47

2013. URL https://link.springer.
com/chapter/10.1007/978-94-007-
5485-0_2.

Rescher:2017:ESM

[Res17a] Nicholas Rescher. Evaluative selection in multimerit situations. In Encouraging Openness: Essays for Joseph Agassi on the Occasion of His 90th Birthday [BAG17], chapter 2, pages 13–17. ISBN 3-319-57668-2 (hardcover), 3-319-57669-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN B833.E53 2017. URL https://link.

springer.com/chapter/10.1007/
978-3-319-57669-5_2.

Rescorla:2017:OTB

[Res17b] Michael Rescorla. From Ockham to Turing — and back again. In Floyd and Bokulich [FB17], chapter 12, pages 279–304. ISBN 3-319-53278-2, 3-319-53280-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QA29.T8 P45 2017. URL https:/

/link.springer.com/chapter/10.
1007/978-3-319-53280-6_12.

Reti:1974:LVB

[Ret74] Ladislao Reti. Leonardo Da Vinci and the beginnings of factories with a central source of power. In Seeger and Cohen [SC74b], pages 97–115. ISBN 90-277-0390-6, 90-277-0376-0 (paperback), 94-010-2126-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 11.

URL https://link.springer.com/
chapter/10.1007/978-94-010-2126-
5_5.

Rey:2018:EWJ

[Rey18] Anne-Lise Rey. The experiments of Willem Jacob ’s Gravesande: A validation of Leibnizian dynamics against Newton? In What Does it Mean to be an Empiricist?: Empiricisms in Eighteenth Century Sciences [BR18b], chapter 5, pages 71–85. ISBN 3-319-69858-3 (hardcover), 3-319-69860-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN B816

.W43 2018.  URL https://link.
springer.com/chapter/10.1007/
978-3-319-69860-1_5.

Ryckman:1989:SFI

[RG89] Thomas Ryckman and Michael Gottfried. Sublanguage formulas as information units. In The Form of Information in Science: Analysis of an Immunology Sublanguage [HGR^{+}89], chapter 4, pages 85–111. ISBN 94-010-7777-0, 94-009-2837-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN B832.A-832.Z.

URL https://link.springer.com/
chapter/10.1007/978-94-009-2837-
4_4.

Renn:2007:PHS

[RG07] Jürgen Renn and Kostas Gavroglu. Positioning the history of science, volume 248 of Boston Studies in the Philosophy of Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2007. ISBN 1-4020-5419-X (hardcover), 1-4020-5420-3 (ebook). ISSN 0068-0346 (print), 2214-7942 (electronic). vii + 188 pp. LCCN Q125 .P77 2007.

Raina:1992:TCS

[RH92] Dhruv Raina and S. Irfan Habib. Technical content and social context: Locating technical institutes. The first two decades in the history of the Kala Bhavan, Baroda (1890–1910). In Science and Empires: Historical Studies about Scientific Development and European Expansion [PJM92], chapter 14, pages 121–136. ISBN 0-7923-1518-9 (hardcover), 94-010-5145-3, 94-011-2594-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 136. URL https:/

/link.springer.com/chapter/10.
1007/978-94-011-2594-9_14.

Reid:2014:VCC

[RH14] Rosalind Reid and Brian Hayes. Viva la ciencia: Cuba’s creative scientists aim to make knowledge their country’s sugar substitute. In The History of Physics in Cuba [BRW14a], chapter 32, pages 417–420. ISBN 94-017-8040-4, 94-017-8041-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic).

LCCN QC5.53; QC6.9; QC9 C9H57 2014; Q124.6-127.2. URL https:/

/link.springer.com/chapter/10.
1007/978-94-017-8041-4_32.

Rheinberger:2005:WWR

[Rhe05] Hans-Jörg Rheinberger. Writing works:

A reaction to Michael Cahn’s paper. In History of Science, History of Text [Che05a], chapter 4, pages 95–103. ISBN 1-281-33589-4, 1-4020-2320-0 (hardcover), 1-4020-2321-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 238

Q125. URL https://link.springer.
com/chapter/10.1007/1-4020-2321-
9_4.

Rheinberger:2009:ER

[Rhe09] Hans-Jörg Rheinberger. Experimental reorientations. In Going Amiss In Experimental Research [HSS09a], chapter 6, pages 75–90. ISBN 1-4020-8892-2 (hardcover), 1-4020-8893-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q172.5.E77 G647

2009. URL https://link.springer.
com/chapter/10.1007/978-1-4020-
8893-3_6.

Rheinberger:2011:ROR

[Rhe11] Hans-Jörg Rheinberger. Recent orientations and reorientations in the life sciences. In Science in the Context of Application [CN11a], chapter 11, pages 161–168. ISBN 90-481-9050-9, 90-481-9051-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q180.55.M4 S385 2011. URL https:

//link.springer.com/chapter/10.
1007/978-90-481-9051-5_11.

Rheinberger:2015:CEU

[Rhe15] Hans-Jörg Rheinberger. Curating the European university. In Arabatzis et al.

[ARS15], chapter 24, pages 357–365. ISBN 3-319-14552-5 (paperback), 3-319-14553-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q125 .R45 2015. URL https://link.

springer.com/chapter/10.1007/
978-3-319-14553-2_24.

Richards:1992:GTM

[Ric91] Joan L. Richards. God, truth, and mathematics in Nineteenth Century England. In The Invention of Physical Science: Intersections of Mathematics, Theology and Natural Philosophy Since the Seventeenth Century: Essays in Honor of Erwin N. Hiebert [NRS92], chapter 3, pages 51–78. ISBN 94-010-5097-X, 94-011-2488-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-011-2488-1_3.

Richardson:2002:PPT

[Ric02] William J. Richardson. Psychoanalytic praxis and the truth of pain. In Hermeneutic Philosophy of Science, Van Gogh’s Eyes, and God: Essays in Honor of Patrick A. Heelan, S.J. [Bab02c], chapter 29, pages 339–350. ISBN 90-481-5926-1, 94-017-1767-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 225. URL https://link.

springer.com/chapter/10.1007/
978-94-017-1767-0_29.

Richards:2007:CT

[Ric07] Joan Richards. The character of truth. In Positioning the history of science [RG07], chapter 23, pages 135–137. ISBN 1-4020-5419-X (hardcover), 1-4020-5420-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q125 .P77 2007.

URL https://link.springer.com/
chapter/10.1007/1-4020-5420-3_23.

Rickles:2012:QGM

[Ric12] Dean Rickles. Quantum gravity meets &HPS. In Mauskopf and Schmaltz [MS12b], chapter 11, pages 163–199. ISBN 94-007-1744-X, 94-007-1745-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175 .I5134

2011. URL https://link.springer.
com/chapter/10.1007/978-94-007-
1745-9_11.

Richardson:2015:TMU

[Ric15] Alan Richardson. From troubled marriage to uneasy colocation: Thomas Kuhn, epistemological revolutions, romantic narratives, and history and philosophy of science. In Kuhn’s Structure of Scientific Revolutions — 50 Years On [DB15b], chapter 4, pages 39–50. ISBN 3-319-13382-9, 3-319-13383-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175.3 .K84 2015. URL https:

//link.springer.com/chapter/10.
1007/978-3-319-13383-6_4.

Richmond:2017:MCT

[Ric17] Sheldon Richmond. The mystique of computer technology and the waning of critical enquiry. In Encouraging Openness: Essays for Joseph Agassi on the Occasion of His 90th Birthday [BAG17], chapter 11, pages 113–127. ISBN 3-319-57668-2 (hardcover), 3-319-57669-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN B833.E53 2017. URL https://link.

springer.com/chapter/10.1007/
978-3-319-57669-5_11.

Riepe:1974:DSS

[Rie74] Dale Riepe. Dirk Struik and the sociology of science. In For Dirk Struik:

Scientific, Historical and Political Essays in Honor of Dirk J. Struik [CSW74], chapter 44, pages 581–591. ISBN 90-277-0379-5, 94-010-2115-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-010-2115-
9_44.

Riedel:1999:STP

[Rie99] Manfred Riedel. Scientific theory or practical doctrine? In Nietzsche, Theories of Knowledge, and Critical Theory: Nietzsche and the Sciences I [Bab99b], chapter 13, pages 187–197. ISBN 90-481-5233-X, 94-017-2430-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-017-2430-
2_13.

Rimini:1997:HAV

[Rim97] Alberto Rimini. The history approach viewed by an external observer. In Cohen et al. [CHS97b], pages 217–229. ISBN 90-481-4809-X, 94-017-2732-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC173.96-174.52. URL https://link.

springer.com/chapter/10.1007/
978-94-017-2732-7_17.

Rip:1992:STD

[Rip92] Arie Rip. Science and technology as dancing partners. In Technological Development and Science in the Industrial Age: New Perspectives on the Science–Technology Relationship [KB92b], chapter 10, pages 231–270. ISBN 90-481-4186-9, 94-015-8010-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/

chapter/10.1007/978-94-015-8010-
6_10.

Rip:2011:PSS

[Rip11] Arie Rip. Protected spaces of science:

Their emergence and further evolution in a changing world. In Science in the Context of Application [CN11a], chapter 13, pages 197–220. ISBN 90-481-9050-9, 90-481-9051-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q180.55.M4 S385

2011. URL https://link.springer.
com/chapter/10.1007/978-90-481-
9051-5_13.

Ritter:2001:EHA

[Rit01] Christopher Ritter. An early history of Alexander Crum Brown’s graphical formulas. In Klein [Kle01b], chapter 3, pages 35–46. ISBN 1-4020-0100-2, 94-015-9737-5. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 222; QD7. URL https:/

/link.springer.com/chapter/10.
1007/978-94-015-9737-1_3.

Ritter:2005:RST

[Rit05] Jim Ritter. Reading Strasbourg 368:

A thrice-told tale. In History of Science, History of Text [Che05a], chapter 9, pages 177–200. ISBN 1-281-33589-4, 1-4020-2320-0 (hardcover), 1-4020-2321-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 238 Q125.

URL https://link.springer.com/
chapter/10.1007/1-4020-2321-9_9.

Rivadulla:1996:BIS

[Riv96] Andrés Rivadulla. Bayesian induction and statistical inference. In Munévar [Mun96], chapter 5, pages 103–121. ISBN 94-010-6622-1, 94-009-0305-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-009-0305-
                                              0_5.

Roper:1995:BFH

[RL95] Margaret M. Roper and J. K. Ladha. Biological n 2 fixation by heterotrophic and phototrophic bacteria in association with straw. In No Truth Except in the Details: Essays in Honor of Martin J. Klein [KS95], chapter 10, pages 211–224. ISBN 94-010-4097-4, 94-011-0217-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

                                              ???? URL https://link.springer.
                                              com/chapter/10.1007/978-94-011-
                                              0055-7_10.

Redhead:1997:REA

[RL97] Michael Redhead and Patrick La Rivière. The relativistic EPR argument. In Cohen et al. [CHS97b], pages 207–215. ISBN 90-481-4809-X, 94-017-2732-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC173.96-174.52. URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-017-2732-7_16.

Renn:2018:BLA

[RM18] Jürgen Renn and Peter McLaughlin. The balance, the lever and the Aristotelian origins of mechanics. In Feldhay et al. [FRSV18a], chapter 5, pages 111– 137. ISBN 3-319-90343-8 (hardcover), 3-319-90345-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic).

                                              LCCN QA802. URL https://link.
                                              springer.com/chapter/10.1007/
                                              978-3-319-90345-3_5.

Robinson:2003:GNH

[Rob03] David C. Robinson. Geometry, null hypersurfaces and new variables. In Renn et al. [RDS03], chapter 15, pages 349–360. ISBN 1-4020-1284-5, 1-4020-1285-3 (paperback). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 234. URL https:/

/link.springer.com/chapter/10.
1007/978-94-010-0111-3_15.

Rockmore:1999:HNC

[Roc99] Tom Rockmore. Habermas, Nietzsche, and cognitive perspective. In Nietzsche, Theories of Knowledge, and Critical Theory: Nietzsche and the Sciences I [Bab99b], chapter 21, pages 281–288. ISBN 90-481-5233-X, 94-017-2430-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-017-2430-
2_21.

Rocke:2001:CAE

[Roc01] Alan J. Rocke. Chemical atomism and the evolution of chemical theory in the Nineteenth Century. In Klein [Kle01b], chapter 1, pages 1–11. ISBN 1-4020-0100-2, 94-015-9737-5. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 222; QD7. URL https:

//link.springer.com/chapter/10.
1007/978-94-015-9737-1_1.

RocheldeCamargo:2011:EPP

[Roc11] Kenneth Rochel de Camargo, Jr.

Echoes from the past: The persisting shadow of classical determinism in contemporary health sciences. In Brazilian Studies in Philosophy and History of Science: An account of recent works [KV11], chapter 18, pages 245–256. ISBN 90-481-9421-0 (hardcover), 90-481-9422-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175.3 .B73 2011. URL https:/

/link.springer.com/chapter/10.
1007/978-90-481-9422-3_18.

Rodriguez:2015:RTI

[Rod15] Hannot Rodríguez. Risk and trust in institutions that regulate strategic technological innovations: Challenges for a socially legitimate risk analysis. In New Perspectives on Technology, Values, and Ethics: Theoretical and Practical [Gon15a], chapter 8, pages 147–166. ISBN 3-319-21869-7 (hardcover), 3-319-21870-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

                                              ???? URL https://link.springer.
                                              com/chapter/10.1007/978-3-319-
                                              21870-0_8.

Rodrig:2017:RSH

[Rod17] Miriam Farhi Rodrig. En route to the siblinghood of humanity? In Encouraging Openness: Essays for Joseph Agassi on the Occasion of His 90th Birthday [BAG17], chapter 21, pages 251–262. ISBN 3-319-57668-2 (hardcover), 3-319-57669-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN B833.E53 2017.

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-3-319-57669-
                                              5_21.

Roe:2001:RNE

[Roe01] Shirley A. Roe. Radical nature in the encyclopédie. In Allen and MacLeod [AM01b], chapter 3, pages 37–59. ISBN 1-4020-0495-8 (hardcover), 94-017-2956-5. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 228; Q175.55. URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-017-2956-7_3.

Rogers:1971:MCT

[Rog71] Ben Rogers. Material conditions on tests of statistical hypotheses. In Buck and Cohen [BC71], pages 403–412. ISBN 90-277-0309-4, 94-010-3142-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-010-3142-
4_25.

Roger:1981:TSD

[Rog81] Jacques Roger. Two scientific discoveries: Their genesis and destiny. In Grmek et al. [GCC81b], pages 229–237. ISBN 90-277-1123-2, 94-010-1284-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-010-1284-
3_11.

Rolin:2016:VSS

[Rol16] Kristina Rolin. Values in the social sciences: The case of feminist research. In Meta-Philosophical Reflection on Feminist Philosophies of Science [AV16a], chapter 8, pages 133–150. ISBN 3-319-26346-3 (hardcover), 3-319-26348-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN HQ1154 .M48 2016; HQ1190 M48 2016.

URL https://link.springer.com/
chapter/10.1007/978-3-319-26348-
9_8.

Roman:1969:SP

[Rom69] P. Roman. Symmetry in physics.

In Cohen and Wartofsky [CW69a], pages 363–369. ISBN 94-010-3383-8, 94-010-3263-7, 94-010-3381-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-010-3381-7_11.

Rongzhao:1996:LHT

[Ron96] Mei Rongzhao. Liu Hui’s theories of mathematics. In Chinese Studies in the History and Philosophy of Science and Technology [DC96], chapter 19, pages 243–254. ISBN 90-481-4546-5, 94-015-8717-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 180; U390 .N19

1996. URL https://link.springer.
com/chapter/10.1007/978-94-015-
8717-4_19.

Ropohl:1983:CTD

[Rop83] Günter Ropohl. A critique of technological determinism. In Philosophy and Technology [DR83], chapter 6, pages 83–96. ISBN 94-009-7126-5, 94-009-7124-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-009-
7124-0_6.

Ropohl:1986:IDM

[Rop86] Günter Ropohl. Information does not make sense or: The relevance gap in information technology and its social dangers. In Philosophy and Technology II: Information Technology and Computers in Theory and Practice [MH86], chapter 5, pages 63–74. ISBN 94-010-8510-2, 94-009-4512-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-009-4512-8_5.

Ropolyi:1999:AST

[Rop99] László Ropolyi. Against the selfish theory. In Fehér et al. [FKR99], pages 307–314. ISBN 90-481-5257-7, 94-015-9293-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-015-
9293-2_25.

Roque:2011:QAD

[Roq11] Tatiana Roque. The qualitative analysis of differential equations and the development of dynamical systems theory. In Brazilian Studies in Philosophy and History of Science: An account of recent works [KV11], chapter 24, pages 313–324. ISBN 90-481-9421-0 (hardcover), 90-481-9422-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175.3 .B73 2011.

URL https://link.springer.com/
chapter/10.1007/978-90-481-9422-
3_24.

Rorty:1986:SHA

[Ror86] Richard Rorty. Should Hume be answered or bypassed? In Cohen and Wartofsky [CW85], chapter 18, pages 341–352. ISBN 90-277-1971-3, 94-009-5345-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-009-
5349-9_18.

Rosen:1974:RES

[Ros74a] Edward Rosen. Rheticus as editor of Sacrobosco. In For Dirk Struik: Scientific, Historical and Political Essays in Honor of Dirk J. Struik [CSW74], chapter 20, pages 245–248. ISBN 90-277-0379-5, 94-010-2115-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-010-2115-9_20.

Rosenbaum:1974:ES

[Ros74b] R. A. Rosenbaum. Is Euclid on the skids? In For Dirk Struik: Scientific, Historical and Political Essays in Honor of Dirk J. Struik [CSW74], chapter 21, pages 249–260. ISBN 90-277-0379-5, 94-010-2115-5 (e-book).

ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-010-2115-9_21.

Rose:1976:DAO

[Ros76] Lynn E. Rose. The domination of astronomy over other disciplines. In Cohen et al. [CHMV76], pages 469–476. ISBN 90-277-0647-6, 90-277-0648-4 (paperback), 94-010-1449-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-010-1449-6_25.

Rossi:1981:FPH

[Ros81] Paolo Rossi. Francesco Patrizi: Heavenly spheres and flocks of cranes. In Italian Studies in the Philosophy of Science [Chi81b], chapter 20, pages 363–388. ISBN 90-277-1073-2, 94-009-8937-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

                                              ???? URL https://link.springer.
                                              com/chapter/10.1007/978-94-009-
                                              8937-5_20.

Rosenkrantz:1986:KBW

[Ros86] Barbara Gutmann Rosenkrantz. Koch’s bacillus: Was there a technological fix? In Ullmann-Margalit [UM86c], chapter 11, pages 147–160. ISBN 90-277-2160-2, 90-277-2161-0 (paperback), 94-009-4566-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 95

                                              Q175. URL https://link.springer.
                                              com/chapter/10.1007/978-94-009-

4566-1_11. See comment [AA86].

Roseberg:1991:HEM

[Rös91] Ulrich Röseberg. Historical explanations in modern physics? The lesson of quantum mechanics. In Woodward and Cohen [WC91], pages 163–175. ISBN 94-010-5395-2, 94-011-3164-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-011-3164-3_15.

Roseberg:1994:HHC

[Rös94] Ulrich Röseberg. Hidden historicity:

The challenge of Bohr’s philosophical thought. In Faye and Folse [FF94], chapter 15, pages 325–343. ISBN 0-7923-2378-5 (hardcover), 94-015-8106-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 153 1994; QC16 1994. URL https:

//link.springer.com/chapter/10.
1007/978-94-015-8106-6_15.

Roseberg:1995:DTJ

[Rös95]  Ulrich Röseberg.  Did they just

misunderstood each other? Logical empiricists and Bohr’s complementarity argument. In Gavroglu et al. [GSW95c], chapter 6, pages 105–123. ISBN 0-7923-2991-0 (set), 0-7923-2988-0, 94-017-2658-2. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 163. URL https:/

/link.springer.com/chapter/10.
1007/978-94-017-2658-0_6.

Rosen:1998:RHN

[Ros98] Stanley Rosen. Remarks on human nature in Plato. In Philosophies of Nature: The Human Dimension: In Celebration of Erazim Kohák [CT98], chapter 11, pages 151–162. ISBN 90-481-4859-6, 94-017-2614-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic).

LCCN BD581. URL https://link.
springer.com/chapter/10.1007/
978-94-017-2614-6_11.

Rossler:1999:I

[Ros99] Otto E. Rossler. Interfaciology. In Fehér et al. [FKR99], pages 85–92. ISBN 90-481-5257-7, 94-015-9293-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-015-9293-
                                              2_7.

Rosende:2009:PRS

[Ros09] Diego L. Rosende. Popper on refutability: Some philosophical and historical questions. In Rethinking Popper [PC09], chapter 11, pages 135–154. ISBN 1-4020-9337-3, 1-4020-9338-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN B1649.P64 R48 2009; Q174 .B67 2009. URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-1-4020-9338-8_11.

Rotenstreich:1986:OSC

[Rot86a]  Nathan Rotenstreich.  Origin and

spontaneity: A comment. In Ullmann-Margalit [UM86a], chapter 6, pages 65–73. ISBN 90-277-2159-9 (paperback), 90-277-2158-0 (hardcover), 94-009-5496-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 94 Q175.

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-009-5496-
                                              0_6.

Rotenstreich:1986:PIT

[Rot86b] Nathan Rotenstreich. The proto-ideas and their aftermath. In Cognition and Fact: Materials on Ludwik Fleck [CS86], chapter 9, pages 161–178. ISBN 94-010-8504-8, 94-009-4498-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-009-4498-5_9.

Rotenstreich:1987:CSD

[Rot87a] Nathan Rotenstreich. Conditioning situations and decisions. In Time and Meaning in History [Rot87j], chapter 6, pages 120–134. ISBN 94-010-8212-X, 94-009-3845-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic).

LCCN ????  URL https://link.
springer.com/chapter/10.1007/
978-94-009-3845-8_6.

Rotenstreich:1987:CI

[Rot87b] Nathan Rotenstreich. Contexts and individuals. In Time and Meaning in History [Rot87j], chapter 5, pages 94–119. ISBN 94-010-8212-X, 94-009-3845-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-009-
3845-8_5.

Rotenstreich:1987:EV

[Rot87c] Nathan Rotenstreich. Evaluation and values. In Time and Meaning in History [Rot87j], chapter 7, pages 135–158. ISBN 94-010-8212-X, 94-009-3845-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-009-3845-
8_7.

Rotenstreich:1987:IAE

[Rot87d] Nathan Rotenstreich. Interaction, actions and events. In Time and Meaning in History [Rot87j], chapter 4, pages 79–93. ISBN 94-010-8212-X, 94-009-3845-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-009-
3845-8_4.

Rotenstreich:1987:PD

[Rot87e] Nathan Rotenstreich. Progress and direction. In Time and Meaning in History [Rot87j], chapter 3, pages 53– 78. ISBN 94-010-8212-X, 94-009-3845-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-009-3845-
                                              8_3.

Rotenstreich:1987:RT

[Rot87f] Nathan Rotenstreich. The rhythm of time. In Time and Meaning in History [Rot87j], chapter 8, pages 159–174. ISBN 94-010-8212-X, 94-009-3845-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-009-3845-
                                              8_8.

Rotenstreich:1987:SI

[Rot87g] Nathan Rotenstreich. The settings and ideologies. In Time and Meaning in History [Rot87j], chapter 9, pages 175–205. ISBN 94-010-8212-X, 94-009-3845-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

                                              ???? URL https://link.springer.
                                              com/chapter/10.1007/978-94-009-
                                              3845-8_9.

Rotenstreich:1987:SH

[Rot87h] Nathan Rotenstreich. The status of history. In Time and Meaning in History [Rot87j], chapter 1, pages 1–24. ISBN 94-010-8212-X, 94-009-3845-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-009-3845-8_1.

Rotenstreich:1987:SP

[Rot87i] Nathan Rotenstreich. The subject and process. In Time and Meaning in History [Rot87j], chapter 2, pages 25–52. ISBN 94-010-8212-X, 94-009-3845-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-009-3845-
8_2.

Rotenstreich:1987:TMH

[Rot87j]  Nathan Rotenstreich.  Time and

Meaning in History, volume 101 of Boston Studies in the Philosophy of Science. D. Reidel, Dordrecht, The Netherlands; Boston, MA, USA; Lancaster, UK; Tokyo, Japan, 1987. ISBN 94-010-8212-X, 94-009-3845-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). 236 (est.) pp. LCCN ????

URL https://link.springer.com/
book/10.1007/978-94-009-3845-8.

Rottschaefer:1987:WSN

[Rot87k] William A. Rottschaefer. Wilfrid Sellars on the nature of thought. In Shimony and Nails [SN87], pages 145–161. ISBN 94-010-8168-9, 94-009-3735-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-009-3735-2_9.

Rousseau:1989:DN

[Rou89] G. S. Rousseau. Discourses of the nerve.

In Literature and Science as Modes of Expression [Amr89], chapter 2, pages 29–60. ISBN 94-010-7531-X, 94-009-2297-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-009-
2297-6_2.

Rouner:1995:SRQ

[Rou95] Leroy S. Rouner. Science, religion, and the quest for truth. In Science, Mind and Art: Essays on science and the humanistic understanding in art, epistemology, religion and ethics In honor of Robert S. Cohen [GSW95b], chapter 18, pages 315–325. ISBN 0-7923-2990-2 (hardcover), 94-011-0469-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 165 BD418.3. URL https:

//link.springer.com/chapter/10.
1007/978-94-011-0469-2_18.

Roux:2004:CM

[Rou04] Sophie Roux. Cartesian mechanics. In The Reception of the Galilean Science of Motion in Seventeenth-Century Europe [PT04], chapter 3, pages 25–66. ISBN 1-4020-2454-1, 1-4020-2455-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC133 .R43

2004. URL https://link.springer.
com/chapter/10.1007/978-1-4020-
2455-9_3.

Roux:2012:EDF

[Rou12] Sophie Roux. An empire divided: French natural philosophy (1670–1690). In Philosophy of Behavioral Biology [PR12b], chapter 3, pages 55–95. ISBN 94-007-1950-7, 94-007-1951-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QH331 .P4679 2012; Q175 .B657 v.282. URL https:

//link.springer.com/chapter/10.
1007/978-94-007-4345-8_3.

Roux:2013:EDF

[Rou13] Sophie Roux. An empire divided: French natural philosophy (1670— 1690). In Garber and Roux [GR13b], chapter 3, pages 55–95. CODEN ???? ISBN 94-007-4344-0 (hardcover), 94-007-4345-9 (e-book). ISSN ???? LCCN QC6 .M3616 2013; QC6 .M43

2013. URL https://link.springer.
com/chapter/10.1007/978-94-007-
4345-8_3.

Roush:2015:RSR

[Rou15] Sherrilyn Roush. The rationality of science in relation to its history. In Kuhn’s Structure of Scientific Revolutions — 50 Years On [DB15b], chapter 6, pages 71–90. ISBN 3-319-13382-9, 3-319-13383-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175.3 .K84 2015. URL https:/

/link.springer.com/chapter/10.
1007/978-3-319-13383-6_6.

Roubach:2016:HPT

[Rou16] Michael Roubach. Heidegger’s primordial temporality and other notions of time. In Cosmological and Psychological Time [DR16], chapter 10, pages 165–175. ISBN 3-319-22589-8 (hardcover), 3-319-22590-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN BD638 C67 2016.

URL https://link.springer.com/
chapter/10.1007/978-3-319-22590-
6_10.

Rowe:1996:RGW

[Row96] David E. Rowe. On the reception of Grassmann’s work in Germany during the 1870’s. In Schubring [Sch96f], pages 131–145. ISBN 90-481-4758-1, 94-015-8753-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QA8.9-10.3. URL https:/

/link.springer.com/chapter/10.
1007/978-94-015-8753-2_11.

Rowe:2006:EAE

[Row06] David E. Rowe. Einstein’s allies and enemies: Debating relativity in Germany, 1916–1920. In Hendricks et al.

[HJLP06], pages 231–280. ISBN 1-4020-5194-8 (hardcover), 1-4020-5195-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175 .B69 v.251 2006; Q175 .I51377 2006.

URL https://link.springer.com/
chapter/10.1007/978-1-4020-5195-
1_8.

Rowlinson:2014:JDH

[Row14] J. S. Rowlinson. James Dewar and his route to the liquefaction of hydrogen. In History of Artificial Cold, Scientific, Technological and Cultural Issues [Gav14b], chapter 3, pages 53–64. ISBN 94-007-7198-3 (hardcover), 94-007-7199-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC278 H57 2014; TP494 .H577 2014.

URL https://link.springer.com/
chapter/10.1007/978-94-007-7199-
4_3.

Rozeboom:1971:NDC

[Roz71] William W. Rozeboom. New dimensions of confirmation theory II: The structure of uncertainty. In Buck and Cohen [BC71], pages 342–374. ISBN 90-277-0309-4, 94-010-3142-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-010-3142-4_23.

Roca-Rosell:2015:AIU

[RR15] Antoni Roca-Rosell. The Autonomous Industrial University of Barcelona (1933–1934?) and the frustrated expectations of democracy in pre-war Spain. In Sciences in the Universities of Europe, Nineteenth and Twentieth Centuries: Academic Landscapes [SDG15b], chapter 9, pages 145–158. ISBN 94-017-9635-1, 94-017-9636-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q183.4.E85

2015. URL https://link.springer.
com/chapter/10.1007/978-94-017-
9636-1_9.

Ruddick:1969:CC

[RS69] William Ruddick and M. M. Schuster.

Causal connection. In Cohen and Wartofsky [CW69b], pages 419–449. ISBN 94-010-3380-3, 94-010-3378-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-010-3378-
7_15.

Renn:2003:ES

[RS03] Jürgen Renn and Tilman Sauer.

Eclipses of the stars. In Renn et al. [RDS03], chapter 5, pages 69–92. ISBN 1-4020-1284-5, 1-4020-1285-3 (paperback). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 234. URL https://link.

springer.com/chapter/10.1007/
978-94-010-0111-3_5.

Rothbart:2006:VIT

[RS06] Daniel Rothbart and John Schreifels.

Visualizing instrumental techniques of surface chemistry. In Philosophy of Chemistry: Synthesis of a New Discipline [BSM06b], chapter 16, pages 309–324. ISBN 1-4020-2455-X, 1-4020-3256-0, 1-4020-3261-7. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 242 QD6. URL https:

//link.springer.com/chapter/10.
1007/1-4020-3261-7_16.

Renn:2007:PCP

[RS07a] Jürgen Renn and Tilman Sauer. Pathways out of classical physics: Einstein’s double strategy in his search for the gravitational field equation. In The Genesis of General Relativity: Volume 1. Einstein’s Zurich Notebook:

Introduction and Source [JNR^{+}07b], pages 113–312. ISBN 1-4020-3999-9 (hardcover), 1-4020-4000-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 no. 250 v. 1; QC173.6 .G469 2007. URL https:

//link.springer.com/chapter/10.
1007/978-1-4020-4000-9_4.

Renn:2007:GTC

[RS07b] Jürgen Renn and Matthias Schemmel. Gravitation in the twilight of classical physics: An introduction. In Renn et al. [RSS^{+}07a], pages 1–18. ISBN 1-4020-3999-9. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 no. 250 v. 3; QC173.6 .G469 2007.

URL https://link.springer.com/
chapter/10.1007/978-1-4020-4000-
9_10.

Renn:2007:HFP

[RS07c] Jürgen Renn and John Stachel. Hilbert’s foundation of physics: From a theory of everything to a constituent of General Relativity. In Renn et al. [RSS^{+}07b], pages 857–973. ISBN 1-4020-3999-9 (hardcover), 1-4020-4000-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 no. 250 v. 4; QC173.6 .G469 2007.

URL https://link.springer.com/
chapter/10.1007/978-1-4020-4000-
9_41.

Renn:2007:GGRa

[RSS^{+}07a] Jürgen Renn, Matthias Schemmel, Christopher Smeenk, Christopher Martin, and Lindy Divarci, editors. The Genesis of General Relativity. Volume 3. Gravitation in the Twilight of Classical Physics: Between Mechanics, Field Theory, and Astronomy, volume 250(3) of Boston Studies in the Philosophy of Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2007. ISBN 1-4020-3999-9. ISSN 0068-0346 (print), 2214-7942 (electronic). 619 pp. LCCN Q174 .B67 no. 250 v. 3; QC173.6 .G469 2007.

Renn:2007:GGRb

[RSS^{+}07b] Jürgen Renn, Matthias Schemmel, Christopher Smeenk, Christopher Martin, and Lindy Divarci, editors. The Genesis of General Relativity. Volume 4. Gravitation in the Twilight of Classical Physics: the Promise of Mathematics, volume 250(4) of Boston Studies in the Philosophy of Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2007. ISBN 1-4020-3999-9 (hardcover), 1-4020-4000-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). 621–1152 pp. LCCN Q174 .B67 no. 250 v. 4; QC173.6

.G469 2007. URL http://d-nb.info/
985765933/34;   http://www.loc.
gov/catdir/enhancements/fy0824/
2007440325-d.html;   http://www.
loc.gov/catdir/enhancements/
fy0824/2007440325-t.html;
https://link.springer.com/book/
10.1007/978-1-4020-4000-9.

Ruse:1989:NDF

[RT89] Michael Ruse and Paul Thompson.

Neo-Darwinism: Form and content. In An Intimate Relation: Studies in the History and Philosophy of Science Presented to Robert E. Butts on his 60th Birthday [BM89], chapter 25, pages 495–512. ISBN 94-010-7546-8, 94-009-2327-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-009-
2327-0_25.

Reinhardt:2001:APT

[RT01] Carsten Reinhardt and Anthony S.

Travis. Aspects of paper tools in the industrial–academic context: Constitutions and structures of aniline dyes, 1860–1880. In Klein [Kle01b], chapter 6, pages 79–94. ISBN 1-4020-0100-2, 94-015-9737-5. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 222; QD7. URL https:/

                                              /link.springer.com/chapter/10.
                                              1007/978-94-015-9737-1_6.

Rubinstein:1983:FET

[Rub83] Benjamin B. Rubinstein. Freud’s early theories of hysteria. In Cohen and Laudan [CL83], chapter 9, pages 169–190. ISBN 94-009-7055-2, 94-009-7057-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-009-7055-
                                              7_9.

Rubin:1988:FVC

[Rub88] Vera Cooper Rubin. To L. F. from V. C. R., 1984. In Philosophy, History and Social Action: Essays in Honor of Lewis Feuer with an autobiographic essay by Lewis Feuer [HOO88c], chapter 20, pages 391–393. ISBN 94-010-7793-2, 94-009-2873-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-009-2873-2_20.

Ruben:1989:OE

[Rub89] David-Hillel Ruben. The ontology of explanation. In Freedom and Rationality: Essays in Honor of John Watkins From his Colleagues and Friends [DJ89], chapter 5, pages 67–85. ISBN 94-010-7571-9, 94-009-2380-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-009-2380-5_5.

Rudich:1967:A

[Rud67] Norman Rudich. An answer. In Cohen and Wartofsky [CW67], pages 425–432. ISBN 90-277-0013-3, 94-010-3508-3. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175 .B731 v.3.

URL https://link.springer.com/
chapter/10.1007/978-94-010-3508-
8_24.

Rudwick:1986:GCS

[Rud86] Martin J. S. Rudwick. The group construction of scientific knowledge: Gentlemen-specialists and the Devonian controversy. In Ullmann-Margalit [UM86a], chapter 16, pages 193–217. ISBN 90-277-2159-9 (paperback), 90-277-2158-0 (hardcover), 94-009-5496-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 94 Q175. URL https:/

/link.springer.com/chapter/10.
1007/978-94-009-5496-0_16.

Ruf:1969:WWA

[Ruf69] Henry Ruf. Wolniewicz on Wittgenstein and Aristotle. In Cohen and Wartofsky [CW69b], pages 218–225. ISBN 94-010-3380-3, 94-010-3378-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-010-3378-7_7.

Ruiz:1995:PM

[Rui95a] Santiago Ramirez Ruiz. Psychoanalysis and Marxism. In Mexican Studies in the History and Philosophy of Science [RC95], chapter 4, pages 49–57. ISBN 94-010-6535-7 (hardcover), 94-009-0109-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-009-
0109-4_4.

RuizdeEsparza:1995:CSR

[Rui95b] Jose Ruiz de Esparza. Contributions of Santiago Ramírez to the history of the Colegio De Minería. In Mexican Studies in the History and Philosophy of Science [RC95], chapter 2, pages 23–28. ISBN 94-010-6535-7 (hardcover), 94-009-0109-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

                                              ???? URL https://link.springer.
                                              com/chapter/10.1007/978-94-009-
                                              0109-4_2.

Ruse:1976:RG

[Rus76a] Michael Ruse. Reduction in genetics. In Cohen et al. [CHMV76], pages 633–651. ISBN 90-277-0647-6, 90-277-0648-4 (paperback), 94-010-1449-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-010-1449-6_36.

Russell:1976:EDN

[Rus76b] B. Russell. Early defense of Newton’s absolute space. In The Concepts of Space and Time: Their Structure and Their Development [Čap76], chapter 23, pages 129–134. ISBN 90-277-0375-2, 94-010-1727-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic).

                                              LCCN BD632. URL https://link.
                                              springer.com/chapter/10.1007/
                                              978-94-010-1727-5_23.

Russell:1976:CTM

[Rus76c] B. Russell. On change, time and motion. In The Concepts of Space and Time: Their Structure and Their Development [Čap76], chapter 44, pages 251–254. ISBN 90-277-0375-2, 94-010-1727-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic).

                                              LCCN BD632. URL https://link.

springer.com/chapter/10.1007/
978-94-010-1727-5_44.

Russell:1976:ZP

[Rus76d] B. Russell. On Zeno’s paradoxes. In The Concepts of Space and Time: Their Structure and Their Development [Čap76], chapter 42, pages 235–243. ISBN 90-277-0375-2, 94-010-1727-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN BD632.

URL https://link.springer.com/
chapter/10.1007/978-94-010-1727-
5_42.

Rus:1979:IRM

[Rus79] Veljko Rus. Institutionalization of the revolutionary movement. In Praxis: Yugoslav Essays in the Philosophy and Methodology of the Social Sciences [MP79], chapter 17, pages 273–287. ISBN 90-277-0968-8, 94-009-9355-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-009-9355-
6_17.

Ruse:1980:OPC

[Rus80] Michael Ruse. Ought philosophers consider scientific discovery? A Darwinian case-study. In Scientific Discovery: Case Studies [Nic80e], chapter 8, pages 131–149. ISBN 90-277-1093-7, 94-009-9015-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 60. URL https:/

/link.springer.com/chapter/10.
1007/978-94-009-9015-9_8.

Ruse:1982:TR

[Rus82] Michael Ruse. Teleology redux. In Scientific Philosophy Today: Essays in Honor of Mario Bunge [AC82], chapter 17, pages 299–309. ISBN 90-277-1262-X, 90-277-1263-8, 94-009-8462-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 67. URL https:/

/link.springer.com/chapter/10.
1007/978-94-009-8462-2_17.

Ruse:2009:SOC

[Rus09] Michael Ruse. Self-organization and complexity in evolutionary theory, or, in this life the bread always falls jammy side down. In Mapping the Future of Biology: Evolving Concepts and Theories [BMP09b], chapter 10, pages 141–154. ISBN 1-4020-9635-6 (hardcover), 1-4020-9636-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QH331 .M295 2009.

URL https://link.springer.com/
chapter/10.1007/978-1-4020-9636-
5_10.

Ruse:2013:EVC

[Rus13] Michael Ruse. Evolution versus creation: A sibling rivalry? In Bergandi [Ber13c], chapter 2, pages 29–38. ISBN 94-007-5066-8, 94-007-5067-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN GE42 .S77 2013; Q175 .B73 v. 296. URL https:/

/link.springer.com/chapter/10.
1007/978-94-007-5067-8_2.

Rutte:1991:EPV

[Rut91a] Heiner Rutte. Ethics and the problem of value in the Vienna Circle. In Uebel [Ueb91b], chapter 11, pages 143–157. ISBN 0-7923-1276-7 (hardcover), 94-011-3182-1. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 133. URL https://link.

springer.com/chapter/10.1007/
978-94-011-3182-7_11.

Rutte:1991:NCS

[Rut91b] Heiner Rutte. Neurath contra Schlick.

On the discussion of truth in the Vienna Circle. In Uebel [Ueb91b], chapter 13, pages 169–174. ISBN 0-7923-1276-7 (hardcover), 94-011-3182-1. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 133. URL https:

//link.springer.com/chapter/10.
1007/978-94-011-3182-7_13.

Rutte:1991:NEH

[Rut91c] Heiner Rutte. On Neurath’s empiricism and his critique of empiricism. In Uebel [Ueb91b], chapter 14, pages 175–190. ISBN 0-7923-1276-7 (hardcover), 94-011-3182-1. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 133. URL https://link.

springer.com/chapter/10.1007/
978-94-011-3182-7_14.

Rutte:1991:PON

[Rut91d] Heiner Rutte. The philosopher Otto Neurath. In Uebel [Ueb91b], chapter 6, pages 81–94. ISBN 0-7923-1276-7 (hardcover), 94-011-3182-1. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 133. URL https:

//link.springer.com/chapter/10.
1007/978-94-011-3182-7_6.

Ruthenberg:2015:RRR

[Rut15] Klaus Ruthenberg. Radicals, reactions, realism. In Philosophy of Chemistry: Growth of a New Discipline [SM15b], chapter 12, pages 183–199. ISBN 94-017-9363-8 (hardcover), 94-017-9364-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QD6 .P455 2015; QD6 .S384 2014. URL https:

//link.springer.com/chapter/10.
1007/978-94-017-9364-3_12.

Romein-Verschoor:1974:WB

[RV74] Annie Romein-Verschoor. What is Burgerlijk? In For Dirk Struik: Scientific, Historical and Political Essays in Honor of Dirk J. Struik [CSW74], chapter 45, pages 593–604. ISBN 90-277-0379-5, 94-010-2115-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-010-2115-9_45.

Ryan:1994:BFS

[Rya94] Cheyney Ryan. The bread of faithful speech. In Artifacts, Representations and Social Practice: Essays for Marx Wartofsky [GC94], chapter 28, pages 483–500. ISBN 94-010-4390-6, 94-011-0902-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN B73 .A77 1994. URL https://link.

springer.com/chapter/10.1007/
978-94-011-0902-4_28.

Schejter:1982:MPB

[SA82] Abel Schejter and Joseph Agassi. Molecular phylogenetics: Biological parsimony and methodological extravagance. In Scientific Philosophy Today: Essays in Honor of Mario Bunge [AC82], chapter 19, pages 333–356. ISBN 90-277-1262-X, 90-277-1263-8, 94-009-8462-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 67. URL https:

//link.springer.com/chapter/10.
1007/978-94-009-8462-2_19.

Stotz:2012:CSR

[SA12] Karola Stotz and Colin Allen. From cell–surface receptors to higher learning: A whole world of experience. In Philosophy of Behavioral Biology [PR12b], chapter 5, pages 85–123. ISBN 94-007-1950-7, 94-007-1951-5 (e-book).

ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QH331 .P4679 2012; Q175 .B657 v.282. URL https:

//link.springer.com/chapter/10.
1007/978-94-007-1951-4_5.

Saarloos:2017:VCV

[Saa17] Léjon Saarloos. Virtues of courage and virtues of restraint: Tyndall, Tait and the use of the imagination in Late Victorian science. In van Dongen and Paul [vDP17a], chapter 8, pages 109–128. ISBN 3-319-48892-9 (hardcover), 3-319-48893-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 v. 321; Q175.32.V57 E65 2017. URL https:

//link.springer.com/chapter/10.
1007/978-3-319-48893-6_8.

Sabelli:1984:MDS

[Sab84] H. C. Sabelli. Mathematical dialectics, scientific logic and the psychoanalysis of thinking. In Hegel and the Sciences [CW84a], chapter 20, pages 349–359. ISBN 94-009-6235-5, 94-009-6233-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-009-6233-
0_20.

Sabloff:1992:ICC

[Sab92] Jeremy A. Sabloff. Interpreting the collapse of classic Maya civilization: A case study of changing archaeological perspectives. In Metaarchaeology: Reflections by Archaeologists and Philosophers [Emb92b], chapter 4, pages 99–119. ISBN 94-010-4806-1, 94-011-1826-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-011-
1826-2_4.

Sachs:1967:EMG

[Sac67] Mendel Sachs. On the elementarity of measurement in general relativity: Toward a general theory. In Cohen and Wartofsky [CW67], pages 56–80. ISBN 90-277-0013-3, 94-010-3508-3. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175 .B731 v.3.

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-010-3508-
                                              8_3.

Sachsse:1983:CWA

[Sac83] Hans Sachsse. Comment: What is alternative technology? A reply to Professor Stanley Carpenter. In Philosophy and Technology [DR83], chapter 9, pages 137–139. ISBN 94-009-7126-5, 94-009-7124-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

                                              ???? URL https://link.springer.
                                              com/chapter/10.1007/978-94-009-
                                              7124-0_9.

Sadovsky:1989:MSK

[Sad89] Vadim Sadovsky. Models of scientific knowledge. In An Intimate Relation: Studies in the History and Philosophy of Science Presented to Robert E. Butts on his 60th Birthday [BM89], chapter 19, pages 391–404. ISBN 94-010-7546-8, 94-009-2327-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-009-2327-0_19.

Sagal:1987:NEH

[Sag87] Paul T. Sagal. Naturalistic epistemology and the harakiri of philosophy. In Shimony and Nails [SN87], pages 321–332. ISBN 94-010-8168-9, 94-009-3735-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic).

LCCN ????  URL https://link.
springer.com/chapter/10.1007/
978-94-009-3735-2_19. See comments

[Shi87d, Aga87b].

Saito:2004:BIE

[Sai04a] Ken Saito. Book II of Euclid’s Elements in the light of the theory of conic sections. In Classics in the History of Greek Mathematics [Chr04a], chapter 7, pages 139–168. ISBN 1-4020-0081-2 (hardcover), 1-4020-2640-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QA22 .C596 2004.

URL https://link.springer.com/
chapter/10.1007/978-1-4020-2640-
9_7.

Saito:2004:I

[Sai04b] Ken Saito. Introduction. In Classics in the History of Greek Mathematics [Chr04a], chapter 9, pages 187–190. ISBN 1-4020-0081-2 (hardcover), 1-4020-2640-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QA22 .C596 2004. URL https:/

/link.springer.com/chapter/10.
1007/978-1-4020-2640-9_9.

Sakamoto:1998:MPC

[Sak98] Hyakudai Sakamoto. Mind, privacy and causality. In Japanese Studies in the Philosophy of Science [NC98], chapter 4, pages 33–66. ISBN 94-010-6176-9, 94-011-5175-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-011-5175-7_4.

Sakul:2011:GOO

[Şak11] Kahraman Şakul. General observations on the Ottoman military industry, 1774–1839: Problems of organization and standardization. In Science between Europe and Asia: Historical Studies on the Transmission, Adoption and Adaptation of Knowledge [GR11b], chapter 4, pages 41–55. ISBN 90-481-9968-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q125 .S354 2010; Q175.52.A78 S35

2010. URL https://link.springer.
com/chapter/10.1007/978-90-481-
9968-6_4.

Salmon:1971:ERC

[Sal71] Wesley C. Salmon. Explanation and relevance: Comments on James G. Greeno’s ‘Theoretical Entities in Statistical Explanation’. In Buck and Cohen [BC71], pages 27–39. ISBN 90-277-0309-4, 94-010-3142-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-010-3142-4_2.

Salmon:1974:FPC

[Sal74] Wesley C. Salmon. Fundamental problems in the concept of randomness. In Schaffner and Cohen [SC74a], page 101. ISBN 90-277-0408-2, 90-277-0409-0, 94-010-2140-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175 .B7312. URL https:/

/link.springer.com/chapter/10.
1007/978-94-010-2140-1_7.

Salamun:1984:PIC

[Sal84] Kurt Salamun. The problem of ideology and critical rationalism. In Rationality in Science and Politics [And84d], chapter 14, pages 249–260. ISBN 90-277-1953-5, 94-009-6254-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-009-6254-5_14.

Saldana:1992:SPA

[Sal92a] Juan-José Saldaña. Science et pouvoir au XIXe siècle. La France et le Mexique en perspective. (French) [Science and power in the Nineteenth Century. France and Mexico in perspective]. In Science and Empires: Historical Studies about Scientific Development and European Expansion [PJM92], chapter 17, pages 153–164. ISBN 0-7923-1518-9 (hardcover), 94-010-5145-3, 94-011-2594-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 136.

URL https://link.springer.com/
chapter/10.1007/978-94-011-2594-
9_17.

Salmon:1992:PMP

[Sal92b] Merrilee H. Salmon. Philosophical models for postprocessual archaeology. In Metaarchaeology: Reflections by Archaeologists and Philosophers [Emb92b], chapter 9, pages 227–241. ISBN 94-010-4806-1, 94-011-1826-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-011-1826-
2_9.

Salmon:1992:EAU

[Sal92c] Wesley C. Salmon. Explanation in archaeology: An update. In Metaarchaeology: Reflections by Archaeologists and Philosophers [Emb92b], chapter 10, pages 243–253. ISBN 94-010-4806-1, 94-011-1826-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-011-
1826-2_10.

Salomon:2001:DSP

[Sal01] Jean-Jacques Salomon. The dark side of progress. In Allen and MacLeod [AM01b], chapter 10, pages 153–166. ISBN 1-4020-0495-8 (hardcover), 94-017-2956-5. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 228; Q175.55. URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-017-2956-7_10.

Salverda:2018:CPM

[Sal18] Josephine Salverda. Can proofs by mathematical induction be explanatory? In Piazza and Pulcini [PP18], pages 163–188. ISBN 3-319-93341-8, 3-319-93342-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic).

                                              LCCN QA8.4. URL https://link.
                                              springer.com/chapter/10.1007/
                                              978-3-319-93342-9_10.

Sambursky:1976:SDE

[Sam76a] S. Sambursky. The stoic doctrine of eternal recurrence. In The Concepts of Space and Time: Their Structure and Their Development [Čap76], chapter 28, pages 167–171. ISBN 90-277-0375-2, 94-010-1727-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic).

                                              LCCN BD632. URL https://link.
                                              springer.com/chapter/10.1007/
                                              978-94-010-1727-5_28.

Sambursky:1976:SIS

[Sam76b] S. Sambursky. The stoic idea of space. In The Concepts of Space and Time: Their Structure and Their Development [Čap76], chapter 5, pages 31–32. ISBN 90-277-0375-2, 94-010-1727-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN BD632.

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-010-1727-
                                              5_5.

Sambursky:1976:SVT

[Sam76c] S. Sambursky. The stoic views of time.

In The Concepts of Space and Time: Their Structure and Their Development [Čap76], chapter 27, pages 159–166. ISBN 90-277-0375-2, 94-010-1727-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN BD632.

URL https://link.springer.com/
chapter/10.1007/978-94-010-1727-
5_27.

Sambursky:1986:OR

[Sam86]  Shmuel Sambursky.  Opening re-

marks. In Ullmann-Margalit [UM86c], chapter 14, pages 181–182. ISBN 90-277-2160-2, 90-277-2161-0 (paperback), 94-009-4566-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 95 Q175.

URL https://link.springer.com/
chapter/10.1007/978-94-009-4566-
1_14.

Sambursky:1992:MSA

[Sam92] Shmuel Sambursky. Man as spectator and actor in the drama of existence. In The Scientific Enterprise: The Bar-Hillel Colloquium: Studies in History, Philosophy, and Sociology of Science, Volume 4 [UM92], chapter 16, pages 263–267. ISBN 94-010-5190-9, 94-011-2688-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .S354 1992. URL https:

//link.springer.com/chapter/10.
1007/978-94-011-2688-5_16.

Samols:1996:RFQ

[Sam96] Trevor M. Samols. A realistic formulation of quantum field theory. In Bohmian Mechanics and Quantum Theory: an Appraisal [CFG96], chapter 13, pages 191–196. ISBN 90-481-4698-4, 94-015-8715-9 (e-book).

ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC173.96-174.52.

URL https://link.springer.com/
chapter/10.1007/978-94-015-8715-
0_13.

Sani:1981:NCT

[San81] Andrea Sani. Necessary and contingent truths in Leibniz. In Italian Studies in the Philosophy of Science [Chi81b], chapter 22, pages 411–422. ISBN 90-277-1073-2, 94-009-8937-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-009-8937-5_22.

Santas:1990:KBP

[San90] Gerasimos Santas. Knowledge and belief in Plato’s Republic. In Greek Studies in the Philosophy and History of Science [Nic90], chapter 3, pages 45– 59. ISBN 94-010-7403-8, 94-009-2015-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-009-2015-
6_3.

Sangwan:1992:NHC

[San92a] Satpal Sangwan. Natural history in colonial context: Profit or pursuit? British botanical enterprise in India 1778–1820. In Science and Empires: Historical Studies about Scientific Development and European Expansion [PJM92], chapter 29, pages 281–298. ISBN 0-7923-1518-9 (hardcover), 94-010-5145-3, 94-011-2594-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 136.

URL https://link.springer.com/
chapter/10.1007/978-94-011-2594-
9_29.

Santone:1992:EMI

[San92b] Michael A. Santone, Jr. Ernst Mach’s influence on four Japanese thinkers. In Ernst Mach — A Deeper Look: Documents and New Perspectives [Bla92a], chapter 17, pages 333–361. ISBN 94-010-5230-1, 94-011-2771-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN C5.53; QC6.9.

URL https://link.springer.com/
chapter/10.1007/978-94-011-2771-
4_17.

Sandkuhler:1997:HRK

[San97a]  Hans Jörg Sandkühler.  The hu-

man right to know and the reality of knowledge. In Ginev and Cohen [GC97], chapter 17, pages 289–303. ISBN 0-7923-4444-8 (hardcover), 94-011-5788-X. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 192. URL https://link.

springer.com/chapter/10.1007/
978-94-011-5788-9_17.

Sankey:1997:KOR

[San97b] Howard Sankey. Kuhn’s ontological relativism. In Ginev and Cohen [GC97], chapter 18, pages 305–320. ISBN 0-7923-4444-8 (hardcover), 94-011-5788-X. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 192. URL https://link.

springer.com/chapter/10.1007/
978-94-011-5788-9_18.

Sankey:2008:SRE

[San08] Howard Sankey. Some reflections on experimental incommensurability. In Rethinking Scientific Change and Theory Comparison: Stabilities, Ruptures, Incommensurabilities? [SSHH08], chapter 23, pages 341–347. ISBN 1-4020-6274-5 (hardcover), 1-4020-6279-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175 .R397

2008. URL https://link.springer.
com/chapter/10.1007/978-1-4020-
6279-7_23.

Santos:2014:ESP

[San14] Elena Vigil Santos. Experimental semiconductor physics: The will to contribute to the country’s economic development. In The History of Physics in Cuba [BRW14a], chapter 13, pages 289–293. ISBN 94-017-8040-4, 94-017-8041-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC5.53; QC6.9; QC9 C9H57 2014; Q124.6-127.2. URL https://link.

springer.com/chapter/10.1007/
978-94-017-8041-4_13.

Sandu:2015:CTP

[San15] Gabriel Sandu. On a combination of truth and probability: Probabilistic independence-friendly logic. In Romanian Studies in Philosophy of Science [PST15], chapter 7, pages 105–124. ISBN 3-319-16654-9 (hardcover), 3-319-16655-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-3-319-
16655-1_7.

SanMauro:2018:CTT

[San18] Luca San Mauro. Church–Turing thesis, in practice. In Piazza and Pulcini [PP18], pages 225–248. ISBN 3-319-93341-8, 3-319-93342-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QA8.4. URL https:

//link.springer.com/chapter/10.
1007/978-3-319-93342-9_13.

Sarkar:1991:LCD

[Sar91]  Sahotra Sarkar.  Lamarck contre

Darwin, reduction versus statistics:

Conceptual issues in the controversy over directed mutagenesis in bacteria. In Organism and the Origins of Self [Tau91i], chapter 11, pages 235–271. ISBN 0-7923-1185-X, 94-011-3406-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QR180-189.5.

URL https://link.springer.com/
chapter/10.1007/978-94-011-3406-
4_11.

Sarkar:1992:FEG

[Sar92a] Sahotra Sarkar, editor. The founders of evolutionary genetics: a centenary reappraisal, volume 142 of Boston Studies in the Philosophy of Science. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 1992. ISBN 0-7923-1777-7 (hardcover), 94-011-2856-1. ISSN 0068-0346 (print), 2214-7942 (electronic). 300 pp. LCCN Q174

.B67 vol. 142. URL http://www.loc.
gov/catdir/enhancements/fy0823/
92012823-d.html;   http://www.
loc.gov/catdir/enhancements/
fy0823/92012823-t.html;   https:
//link.springer.com/book/10.
1007/978-94-011-2856-8.

Sarkar:1992:FTE

[Sar92b] Sahotra Sarkar. The founders of theoretical evolutionary genetics: Editor’s introduction. In The founders of evolutionary genetics: a centenary reappraisal [Sar92a], chapter 1, pages 1–22. ISBN 0-7923-1777-7 (hardcover), 94-011-2856-1. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 142. URL https://link.

springer.com/chapter/10.1007/
978-94-011-2856-8_1.

Sarkar:1992:HBC

[Sar92c] Sahotra Sarkar. Haldane as biochemist:

The Cambridge decade, 1923–1932. In The founders of evolutionary genetics: a centenary reappraisal [Sar92a], chapter 4, pages 53–81. ISBN 0-7923-1777-7 (hardcover), 94-011-2856-1. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 142.

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-011-2856-
                                              8_4.

Sarkar:1996:PHM

[Sar96] Sahotra Sarkar, editor. The Philosophy and History of Molecular Biology: New Perspectives, volume 183 of Boston Studies in the Philosophy of Science. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 1996. ISBN 0-7923-3947-9 (hardcover). LCCN Q174 .B67 vol. 183 QH506.

Sarkar:1997:FFM

[Sar97] Sahotra Sarkar. Form and function in the molecularization of biology. In The Elusive Synthesis: Aesthetics and Science [Tau97a], chapter 7, pages 153–168. ISBN 0-7923-4763-3, 0-7923-3904-5, 94-009-1786-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-009-1786-6_7.

Sarnowsky:2008:CIH

[Sar08] Jürgen Sarnowsky. Concepts of impetus and the history of mechanics. In Mechanics and Natural Philosophy Before the Scientific Revolution [LR08], chapter 4, pages 121–145. ISBN 1-4020-5966-3 (hardcover), 1-4020-5967-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QA802 .M43

                                              2008. URL https://link.springer.
                                              com/chapter/10.1007/978-1-4020-
                                              5967-4_4.

Sarewitz:2011:PES

[Sar11] Daniel Sarewitz. Political effectiveness in science and technology. In Science in the Context of Application [CN11a], chapter 18, pages 301–315. ISBN 90-481-9050-9, 90-481-9051-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q180.55.M4 S385

2011. URL https://link.springer.
com/chapter/10.1007/978-90-481-
9051-5_18.

Simoes:2015:Ib

[SAR15] Ana Simões, Theodore Arabatzis, and Jürgen Renn. Introduction. In Arabatzis et al. [ARS15], chapter 1, pages 1–5. ISBN 3-319-14552-5 (paperback), 3-319-14553-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q125 .R45 2015. URL https:

//link.springer.com/chapter/10.
1007/978-3-319-14553-2_1.

Sasaki:1992:SJE

[Sas92]  Chikara Sasaki.  Science and the

Japanese Empire 1868–1945: An overview. In Science and Empires: Historical Studies about Scientific Development and European Expansion [PJM92], chapter 25, pages 243–246. ISBN 0-7923-1518-9 (hardcover), 94-010-5145-3, 94-011-2594-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 136.

URL https://link.springer.com/
chapter/10.1007/978-94-011-2594-
9_25.

Sassower:1995:PPV

[Sas95] Raphael Sassower. Pedagogy as psychology: A view from within. In Critical Rationalism, the Social Sciences and the Humanities: Essays for Joseph Agassi. Volume II [JL95b], chapter 5, pages 79–99. ISBN 0-7923-2961-9, 90-481-4430-2, 94-017-0441-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN B29 .C75 1995.

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-017-0441-
                                              0_5.

Sasaki:2003:G

[Sas03a] Chikara Sasaki. The Géométrie of 1637. In Descartes’s Mathematical Thought [Sas03d], chapter 6, pages 205–280. ISBN 90-481-6487-7, 94-017-1225-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-017-1225-
                                              5_6.

Sasaki:2003:C

[Sas03b] Chikara Sasaki. Conclusion. In Descartes’s Mathematical Thought [Sas03d], chapter 11, pages 419–438. ISBN 90-481-6487-7, 94-017-1225-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-017-1225-
                                              5_11.

Sasaki:2003:DJM

[Sas03c] Chikara Sasaki. Descartes and Jesuit mathematical education. In Descartes’s Mathematical Thought [Sas03d], chapter 2, pages 13–44. ISBN 90-481-6487-7, 94-017-1225-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic).

                                              LCCN ????  URL https://link.
                                              springer.com/chapter/10.1007/
                                              978-94-017-1225-5_2.

Sasaki:2003:DMT

[Sas03d] Chikara Sasaki. Descartes’s Mathematical Thought, volume 237 of Boston Studies in the Philosophy of Science. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 2003. ISBN 90-481-6487-7, 94-017-1225-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). xiv + 496 pp. LCCN ????

URL https://link.springer.com/
book/10.1007/978-94-017-1225-5.

Sasaki:2003:FAR

[Sas03e] Chikara Sasaki. The first attempt at reforming mathematics. In Descartes’s Mathematical Thought [Sas03d], chapter 4, pages 95–157. ISBN 90-481-6487-7, 94-017-1225-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-017-1225-5_4.

Sasaki:2003:IC

[Sas03f] Chikara Sasaki. Interim consideration.

In Descartes’s Mathematical Thought [Sas03d], chapter 7, pages 281–285. ISBN 90-481-6487-7, 94-017-1225-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-017-1225-
5_7.

Sasaki:2003:I

[Sas03g] Chikara Sasaki. Introduction. In Descartes’s Mathematical Thought [Sas03d], chapter 1, pages 1–9. ISBN 90-481-6487-7, 94-017-1225-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-017-1225-5_1.

Sasaki:2003:MBR

[Sas03h] Chikara Sasaki. The mathematical background of the Regulae ad Directionem Ingenii. In Descartes’s Mathematical Thought [Sas03d], chapter 5, pages 159–203. ISBN 90-481-6487-7, 94-017-1225-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-017-1225-5_5.

Sasaki:2003:MTC

[Sas03i] Chikara Sasaki. The mathematical thought of Christoph Clavius. In Descartes’s Mathematical Thought [Sas03d], chapter 3, pages 45–93. ISBN 90-481-6487-7, 94-017-1225-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-017-1225-5_3.

Sasaki:2003:MUSb

[Sas03j] Chikara Sasaki. ‘Mathesis universalis’ in the Seventeenth Century. In Descartes’s Mathematical Thought [Sas03d], chapter 10, pages 359–418. ISBN 90-481-6487-7, 94-017-1225-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-017-1225-
                                              5_10.

Sasaki:2003:MUSa

[Sas03k] Chikara Sasaki. ‘Mathesis universalis’ in the Sixteenth Century. In Descartes’s Mathematical Thought [Sas03d], chapter 9, pages 333–358. ISBN 90-481-6487-7, 94-017-1225-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-017-1225-
                                              5_9.

Sasaki:2003:UMA

[Sas03l] Chikara Sasaki. ‘Universal mathematics’ in Aristotle. In Descartes’s Mathematical Thought [Sas03d], chapter 8, pages 289–331. ISBN 90-481-6487-7, 94-017-1225-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-017-1225-5_8.

Sassower:2017:CC

[Sas17] Raphael Sassower. Capitalism at a crossroad. In Encouraging Openness: Essays for Joseph Agassi on the Occasion of His 90th Birthday [BAG17], chapter 35, pages 439–453. ISBN 3-319-57668-2 (hardcover), 3-319-57669-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN B833.E53

2017. URL https://link.springer.
com/chapter/10.1007/978-3-319-
57669-5_35.

Saunders:1993:WPC

[Sau93] Simon Saunders. To what physics corresponds. In Correspondence, Invariance and Heuristics: Essays in Honour of Heinz Post [FK93], chapter 15, pages 295–325. ISBN 90-481-4229-6, 94-017-1185-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-017-
1185-2_15.

Saunders:1998:HP

[Sau98] Simon Saunders. Hertz’s principles.

In Heinrich Hertz: Classical Physicist, Modern Philosopher [BHN98], chapter 9, pages 123–154. ISBN 90-481-4881-2, 94-015-8855-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-015-8855-3_9.

Saunders:2002:GP

[Sau02] Barbara Saunders. Grammar(s) of perception. In Hermeneutic Philosophy of Science, Van Gogh’s Eyes, and God: Essays in Honor of Patrick A. Heelan, S.J. [Bab02c], chapter 26, pages 305–313. ISBN 90-481-5926-1, 94-017-1767-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 225. URL https:/

                                              /link.springer.com/chapter/10.
                                              1007/978-94-017-1767-0_26.

Saunders:2003:IGC

[Sau03] Simon Saunders. Indiscernibles, general covariance, and other symmetries: The case for non-reductive relationalism. In Renn et al. [RDS03], chapter 7, pages 151–173. ISBN 1-4020-1284-5, 1-4020-1285-3 (paperback). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 234. URL https:/

                                              /link.springer.com/chapter/10.
                                              1007/978-94-010-0111-3_7.

Sauer:2007:EEH

[Sau07] Tilman Sauer. Einstein equations and Hilbert action: What is missing on page 8 of the proofs for Hilbert’s first communication on the foundations of physics? In Renn et al. [RSS^{+}07b], pages 975–988. ISBN 1-4020-3999-9 (hardcover), 1-4020-4000-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 no. 250 v. 4; QC173.6 .G469 2007. URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-1-4020-4000-9_42.

Sauer:2016:MPS

[Sau16] Tilman Sauer. Multiple perspectives on the Stern–Gerlach experiment. In Sauer and Scholl [SS16b], chapter 12, pages 251–263. ISBN 3-319-30227-2, 3-319-30229-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q126.8 .P45 2016. URL https:

//link.springer.com/chapter/10.
1007/978-3-319-30229-4_12.

Savan:1986:SST

[Sav86] David Savan. Spinoza: Scientist and theorist of scientific method. In Spinoza and the Sciences [GN86], chapter 4, pages 95–123. ISBN 94-010-8511-0, 94-009-4514-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-009-
4514-2_4.

Savoia:2015:SHC

[Sav15] Paolo Savoia. Seeing and hearing:

Charcot, Freud and the objectivity of hysteria. In Objectivity in Science: New Perspectives from Science and Technology Studies [PRT15], chapter 7, pages 123–144. ISBN 3-319-14348-4 (hardcover), 3-319-14349-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175 O25 2015.

URL https://link.springer.com/
chapter/10.1007/978-3-319-14349-
1_7.

Sawada:1998:MHJ

[Saw98] Nobushige Sawada. The mind as human jobs. In Japanese Studies in the Philosophy of Science [NC98], chapter 1, pages 1–6. ISBN 94-010-6176-9, 94-011-5175-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-011-
5175-7_1.

Sawyer:2017:DAB

[Saw17] Charles Sawyer. Drive around the block.

In Encouraging Openness: Essays for Joseph Agassi on the Occasion of His 90th Birthday [BAG17], chapter 42, pages 547–561. ISBN 3-319-57668-2 (hardcover), 3-319-57669-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN B833.E53 2017.

URL https://link.springer.com/
chapter/10.1007/978-3-319-57669-
5_42.

Sayan:2005:IAS

[Say05]  Erdinç Sayan.  Idealizations and

approximations in science, and the Bayesian theory of confirmation. In Turkish Studies in the History and Philosophy of Science [IG05c], chapter 8, pages 103–112. ISBN 1-4020-3332-X (hardcover), 1-4020-3333-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 244.

URL https://link.springer.com/
chapter/10.1007/1-4020-3333-8_8.

Schweber:2015:QCQ

[SB15] Sam Schweber and Gal BenPorat. Quantum chemistry and the quantum revolution. In Arabatzis et al. [ARS15], chapter 5, pages 41–66. ISBN 3-319-14552-5 (paperback), 3-319-14553-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q125

.R45 2015.  URL https://link.
springer.com/chapter/10.1007/
978-3-319-14553-2_5.

Schaffner:1974:PPB

[SC74a] Kenneth F. Schaffner and Robert S. Cohen, editors. PSA 1972: proceedings of the 1972 Biennial Meeting Philosophy of Science Association (Olds Plaza Hotel, East Lansing, Michigan, October 27–29, 1972), volume 20 of Boston Studies in the Philosophy of Science. D. Reidel, Dordrecht, The Netherlands; Boston, MA, USA; Lancaster, UK; Tokyo, Japan, 1974. ISBN 90-277-0408-2, 90-277-0409-0, 94-010-2140-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175 .B7312.

URL https://link.springer.com/
book/10.1007/978-94-010-2140-1.

Seeger:1974:PFS

[SC74b] Raymond J. Seeger and Robert S. Cohen, editors. Philosophical Foundations of Science: Proceedings of Section L, 1969, American Association for the Advancement of Science, volume 11 of Boston Studies in the Philosophy of Science. D. Reidel, Dordrecht, The Netherlands; Boston, MA, USA; Lancaster, UK; Tokyo, Japan, 1974. ISBN 90-277-0390-6, 90-277-0376-0 (paperback), 94-010-2126-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 11.

URL https://link.springer.com/
book/10.1007/978-94-010-2126-5.

Sotolongo-Costa:2014:PCS

[SC14] Oscar Sotolongo-Costa. The physics of complex systems in Cuba. In The History of Physics in Cuba [BRW14a], chapter 16, pages 307–313. ISBN 94-017-8040-4, 94-017-8041-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC5.53; QC6.9; QC9 C9H57 2014; Q124.6-127.2.

URL https://link.springer.com/
chapter/10.1007/978-94-017-8041-
4_16.

Scaff:1990:GST

[Sca90a] Lawrence A. Scaff. Georg Simmel’s theory of culture. In Georg Simmel and Contemporary Sociology [KPC90], chapter 16, pages 283–296. ISBN 94-010-6691-4 (hardcover), 94-009-0459-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN HM401-1281.

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-009-0459-
                                              0_16.

Scaltsas:1990:SAA

[Sca90b] Theodore Scaltsas. Soul as attunement: An analogy or a model? In Greek Studies in the Philosophy and History of Science [Nic90], chapter 7, pages 109–119. ISBN 94-010-7403-8, 94-009-2015-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

                                              ???? URL https://link.springer.
                                              com/chapter/10.1007/978-94-009-
                                              2015-6_7.

Scazzieri:2003:EHS

[Sca03]  Roberto Scazzieri.  Experiments,

heuristics and social diversity: A comment on Reinhard Selten. In Descartes’s Mathematical Thought [Sas03d], chapter 6, pages 85–98. ISBN 90-481-6487-7, 94-017-1225-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

                                              URL https://link.springer.com/
                                              chapter/10.1007/0-306-48123-5_6.

Simoes:2003:TL

[SCD03a] Ana Simões, Ana Carneiro, and Maria Paula Diogo. Travels of learning. In Travels of learning: a geography of science in Europe [SCD03b], chapter 1, pages 1–18. ISBN 1-4020-1259-4, 94-017-3584-0. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174

                                              .B67 v. 233. URL https://link.
                                              springer.com/chapter/10.1007/
                                              978-94-017-3584-1_1.

Simoes:2003:TLG

[SCD03b] Ana Simões, Ana Carneiro, and Maria Paula Diogo, editors. Travels of learning: a geography of science in Europe, volume 233 of Boston Studies in the Philosophy of Science. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 2003. ISBN 1-4020-1259-4, 94-017-3584-0. ISSN 0068-0346 (print), 2214-7942 (electronic). xiv + 353 pp. LCCN Q174 .B67 v. 233. URL

https://link.springer.com/book/
10.1007/978-94-017-3584-1.

Scerri:1993:CRC

[Sce93] Eric R. Scerri. Correspondence and reduction in chemistry. In Correspondence, Invariance and Heuristics: Essays in Honour of Heinz Post [FK93], chapter 2, pages 45–64. ISBN 90-481-4229-6, 94-017-1185-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-017-1185-2_2.

Scerri:2001:PTU

[Sce01] Eric Scerri. The periodic table: The ultimate paper tool in chemistry. In Klein [Kle01b], chapter 10, pages 163–177. ISBN 1-4020-0100-2, 94-015-9737-5. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 222;

QD7. URL https://link.springer.
com/chapter/10.1007/978-94-015-
9737-1_10.

Scerri:2006:NDP

[Sce06] Eric R. Scerri. Normative and descriptive philosophy of science and the role of chemistry. In Philosophy of Chemistry: Synthesis of a New Discipline [BSM06b], chapter 6, pages 119–128. ISBN 1-4020-2455-X, 1-4020-3256-0, 1-4020-3261-7. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 242 QD6. URL https:/

/link.springer.com/chapter/10.
1007/1-4020-3261-7_6.

Schiller:1963:DIQ

[Sch63] Ralph Schiller. Deterministic interpretations of the quantum theory. In Wartofsky [War63], pages 144–155. ISBN 94-010-3265-3, 94-010-3263-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174

                                              .B678 1970. URL https://link.
                                              springer.com/chapter/10.1007/
                                              978-94-010-3263-6_11.

Schlesinger:1967:NK

[Sch67] G. Schlesinger. Natural kinds. In Cohen and Wartofsky [CW67], pages 108–122. ISBN 90-277-0013-3, 94-010-3508-3. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175 .B731 v.3.

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-010-3508-
                                              8_7.

Schaffner:1974:UST

[Sch74] Kenneth F. Schaffner. The unity of science and theory construction in molecular biology. In Seeger and Cohen [SC74b], pages 497–533. ISBN 90-277-0390-6, 90-277-0376-0 (paperback), 94-010-2126-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 11. URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-010-2126-5_27.

Schmitt:1975:PSS

[Sch75] Charles B. Schmitt. Philosophy and science in Sixteenth-Century universities: Some preliminary comments. In Murdoch and Sylla [MS75b], pages 485–537. ISBN 90-277-0587-9, 94-010-1781-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN B721 .I584 1975. URL https://link.

                                              springer.com/chapter/10.1007/
                                              978-94-010-1781-7_13.

Schaffner:1976:RBP

[Sch76a] Kenneth F. Schaffner. Reductionism in biology: Prospects and problems. In Cohen et al. [CHMV76], pages 613–632. ISBN 90-277-0647-6, 90-277-0648-4 (paperback), 94-010-1449-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-010-1449-6_35.

Schaffner:1976:WCM

[Sch76b] Kenneth F. Schaffner. The Watson– Crick model and reductionism. In Topics in the Philosophy of Biology [GM76], chapter 4, pages 101–127. ISBN 90-277-0596-8, 94-010-1829-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-010-1829-
6_4.

Scheibe:1976:CPC

[Sch76c] Erhard Scheibe. Conditions of progress and the comparability of theories. In Essays in Memory of Imre Lakatos [CFW76], chapter 31, pages 547–568. ISBN 90-277-0655-7, 94-010-1451-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-010-1451-
9_31.

Schnelle:1976:BAT

[Sch76d] Helmut Schnelle. Basic aspects of the theory of grammatical form. In Language in Focus: Foundations, Methods and Systems: Essays in Memory of Yehoshua Bar-Hillel [Kas76b], chapter 17, pages 377–404. ISBN 90-277-0644-1, 90-277-0645-X, 94-010-1876-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

                                              P101-120.  URL https://link.
                                              springer.com/chapter/10.1007/
                                              978-94-010-1876-0_17.

Schopenhauer:1976:NAT

[Sch76e] A. Schopenhauer. On the necessary attributes of time and space. In The Concepts of Space and Time: Their Structure and Their Development [Čap76], chapter 39, pages 227–230. ISBN 90-277-0375-2, 94-010-1727-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN BD632.

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-010-1727-
                                              5_39.

Schaff:1979:SIC

[Sch79] Adam Schaff. Structuralism as an intellectual current. In Polish Essays in the Methodology of the Social Sciences [Wia79b], chapter 6, pages 105–132. ISBN 90-277-0956-4, 94-009-9353-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174

                                              .P655 1979. URL https://link.
                                              springer.com/chapter/10.1007/
                                              978-94-009-9353-2_6.

Schaffner:1980:DBS

[Sch80a] Kenneth F. Schaffner. Discovery in the biomedical sciences: Logic or irrational intuition? In Scientific Discovery: Case Studies [Nic80e], chapter 10, pages 171–205. ISBN 90-277-1093-7, 94-009-9015-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 60. URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-009-9015-9_10.  See

comment [Mau80] and reply [Sch80b].

Schaffner:1980:RM

[Sch80b]  Kenneth F. Schaffner.  Reply to

Maull. In Scientific Discovery: Case Studies [Nic80e], chapter 12, pages 211–212. ISBN 90-277-1093-7, 94-009-9015-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 60. URL https:/

/link.springer.com/chapter/10.
1007/978-94-009-9015-9_12.  See

[Sch80a, Mau80].

Scheibe:1982:IC

[Sch82] Erhard Scheibe. Invariance and covariance. In Scientific Philosophy Today: Essays in Honor of Mario Bunge [AC82], chapter 18, pages 311–331. ISBN 90-277-1262-X, 90-277-1263-8, 94-009-8462-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 67. URL https:

//link.springer.com/chapter/10.
1007/978-94-009-8462-2_18.

Schafer:1983:FDR

[Sch83a] Wolf Schäfer. The finalization debate:

A reply to our critics. In Finalization in Science: The Social Orientation of Scientific Progress [Sch83b], chapter 14, pages 275–299. ISBN 94-009-7082-X, 94-009-7080-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic).

LCCN ????  URL https://link.
springer.com/chapter/10.1007/
978-94-009-7080-9_14.

Schafer:1983:FSS

[Sch83b] Wolf Schäfer. Finalization in Science:

The Social Orientation of Scientific Progress, volume 77 of Boston Studies in the Philosophy of Science. D. Reidel, Dordrecht, The Netherlands; Boston, MA, USA; Lancaster, UK; Tokyo, Japan, 1983. ISBN 94-009-7082-X, 94-009-7080-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). xviii + 318 pp. LCCN ???? URL

https://link.springer.com/book/
10.1007/978-94-009-7080-9.

Schafer:1983:NF

[Sch83c] Wolf Schäfer. Normative finalization. In Finalization in Science: The Social Orientation of Scientific Progress [Sch83b], chapter 9, pages 207–231. ISBN 94-009-7082-X, 94-009-7080-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-009-7080-
                                              9_9.

Schaffner:1983:CTV

[Sch83d] Kenneth F. Schaffner. Clinical trials: The validation of theory and therapy. In Cohen and Laudan [CL83], chapter 10, pages 191–208. ISBN 94-009-7055-2, 94-009-7057-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-009-7055-7_10.

Scheffler:1983:FFD

[Sch83e] Israel Scheffler. Fallible is as fallible does. In Language, Logic and Method [CW83b], chapter 6, pages 175–178. ISBN 94-009-7704-2, 94-009-7702-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-009-7702-

0_6. See response [Lev83b].

Schirmacher:1983:PET

[Sch83f]  Wolfgang Schirmacher.  From the

phenomenon to the event of technology. In Philosophy and Technology [DR83], chapter 19, pages 275–289. ISBN 94-009-7126-5, 94-009-7124-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-009-7124-0_19.

Schweber:1984:CTS

[Sch84] S. S. Schweber. Commentary on ‘temporality and the structure of physics as human endeavor’. In Physical Sciences and History of Physics [CW84c], chapter 12, pages 231–238. ISBN 94-009-7180-X, 94-009-7178-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .P497 1984.

URL https://link.springer.com/
chapter/10.1007/978-94-009-7178-
3_12.

Schirmacher:1986:PEP

[Sch86a] Wolfgang Schirmacher. Privacy as an ethical problem in the computer society. In Philosophy and Technology II: Information Technology and Computers in Theory and Practice [MH86], chapter 19, pages 257–268. ISBN 94-010-8510-2, 94-009-4512-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-009-4512-8_19.

Schnelle:1986:LFI

[Sch86b] Thomas Schnelle. Ludwik Fleck and the influence of the philosophy of Lwów. In Cognition and Fact: Materials on Ludwik Fleck [CS86], chapter 13, pages 231–265. ISBN 94-010-8504-8, 94-009-4498-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-009-
4498-5_13.

Schnelle:1986:MPS

[Sch86c] Thomas Schnelle. Microbiology and philosophy of science, Lwów and the German Holocaust: Stations of a life — Ludwik Fleck 1896–1961. In Cognition and Fact: Materials on Ludwik Fleck [CS86], chapter 1, pages 3–36. ISBN 94-010-8504-8, 94-009-4498-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-009-4498-5_1.

Schweber:1986:DPD

[Sch86d] Silvan S. Schweber. On Darwin’s principle of divergence: A comment. In Ullmann-Margalit [UM86a], chapter 12, pages 133–136. ISBN 90-277-2159-9 (paperback), 90-277-2158-0 (hardcover), 94-009-5496-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 94 Q175.

URL https://link.springer.com/
chapter/10.1007/978-94-009-5496-
0_12.

Schweber:1986:DCC

[Sch86e] Silvan S. Schweber. On the Devonian controversy: A comment. In Ullmann-Margalit [UM86a], chapter 17, pages 219–223. ISBN 90-277-2159-9 (paperback), 90-277-2158-0 (hardcover), 94-009-5496-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 94 Q175.

URL https://link.springer.com/
chapter/10.1007/978-94-009-5496-
0_17.

Schiffrin:1988:CTR

[Sch88] Harold Z. Schiffrin. China today: Retreat from Mao and return to Marx? In Philosophy, History and Social Action: Essays in Honor of Lewis Feuer with an autobiographic essay by Lewis Feuer [HOO88c], chapter 21, pages 395–404. ISBN 94-010-7793-2, 94-009-2873-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-009-2873-
2_21.

Schleichert:1989:RBC

[Sch89] Hubert Schleichert. The relationship between consciousness and language. In An Intimate Relation: Studies in the History and Philosophy of Science Presented to Robert E. Butts on his 60th Birthday [BM89], chapter 22, pages 441–458. ISBN 94-010-7546-8, 94-009-2327-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic).

LCCN ????  URL https://link.
springer.com/chapter/10.1007/
978-94-009-2327-0_22.

Scheiner:1992:ISH

[Sch91a] Steve Scheiner. Ab initio studies of hydrogen bonding. In The Invention of Physical Science: Intersections of Mathematics, Theology and Natural Philosophy Since the Seventeenth Century: Essays in Honor of Erwin N. Hiebert [NRS92], chapter 6, pages 171–227. ISBN 94-010-5097-X, 94-011-2488-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-3-642-
58183-0_6.

Schnadelbach:1991:AF

[Sch91b] Herbert Schnädelbach. Against Feyerabend. In Beyond Reason: Essays on the Philosophy of Paul Feyerabend [Mun91a], chapter 21, pages 433–448. ISBN 94-010-5406-1, 94-011-3188-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-011-3188-
9_21.

Schabas:1994:PEM

[Sch94a] Margaret Schabas. From political economy to market mechanics: The Jevonian moment in the history of economics. In Cohen [Coh94b], chapter 6, pages 235–255. ISBN 0-7923-2223-1, 94-017-3391-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 150. URL https:

//link.springer.com/chapter/10.
1007/978-94-017-3391-5_6.

Schweber:1994:DAI

[Sch94b] S. S. Schweber. Darwin and the agronomists: An influence of political economy on scientific thought. In Cohen [Coh94b], chapter 9, pages 305–316. ISBN 0-7923-2223-1, 94-017-3391-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 150. URL https:/

/link.springer.com/chapter/10.
1007/978-94-017-3391-5_9.

Scharfstein:1995:ODF

[Sch95a] Ben-Ami Scharfstein. On our difficulties in finding the right words. In Critical Rationalism, Metaphysics and Science: Essays for Joseph Agassi. Volume I [JL95a], chapter 2, pages 15–27. ISBN 94-010-4211-X, 94-011-0471-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-011-0471-5_2.

Schulmann:1995:PCE

[Sch95b] Robert Schulmann. From periphery to center: Einstein’s path from Bern to Berlin (1902–1914). In No Truth Except in the Details: Essays in Honor of Martin J. Klein [KS95], chapter 10, pages 259–271. ISBN 94-010-4097-4, 94-011-0217-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-011-
0217-9_10.

Schweber:1995:PCC

[Sch95c] Sylvan S. Schweber. Physics, community and the crisis in physical theory. In Gavroglu et al. [GSW95c], chapter 7, pages 125–152. ISBN 0-7923-2991-0 (set), 0-7923-2988-0, 94-017-2658-2. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 163.

URL https://link.springer.com/
chapter/10.1007/978-94-017-2658-
0_7.

Schagrin:1996:REM

[Sch96a] Morton L. Schagrin. Rumford’s experiments on the materiality of light. In Cohen et al. [CHR96], pages 395–403. ISBN 90-481-4493-0, 94-015-8638-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-015-8638-
2_29.

Schlagel:1996:RRA

[Sch96b] Richard H. Schlagel. Resolving the realist–antirealist dilemma. In Cohen et al. [CHR96], pages 405–417. ISBN 90-481-4493-0, 94-015-8638-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-015-8638-2_30.

Schlote:1996:HGG

[Sch96c] Karl-Heinz Schlote. Hermann Günther Grassmann and the theory of hypercomplex number systems. In Schubring [Sch96f], pages 165–173. ISBN 90-481-4758-1, 94-015-8753-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QA8.9-10.3.

URL https://link.springer.com/
chapter/10.1007/978-94-015-8753-
2_14.

Scholz:1996:IJG

[Sch96d]  Erhard Scholz.  The influence of

Justus Grassmann’s crystallographic works on Hermann Grassmann. In Schubring [Sch96f], pages 37–45. ISBN 90-481-4758-1, 94-015-8753-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QA8.9-10.3.

URL https://link.springer.com/
chapter/10.1007/978-94-015-8753-
2_5.

Schubring:1996:CBH

[Sch96e] Gert Schubring. The cooperation between Hermann and Robert Grassmann on the foundations of mathematics. In Hermann Günther Graßmann (1809– 1877): Visionary Mathematician, Scientist and Neohumanist Scholar: Papers from a Sesquicentennial Conference [Sch96f], pages 59–70. ISBN 90-481-4758-1, 94-015-8753-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QA8.9-10.3.

URL https://link.springer.com/
chapter/10.1007/978-94-015-8753-
2_7.

Schubring:1996:HGG

[Sch96f] Gert Schubring, editor. Hermann Günther Graßmann (1809–1877): Visionary Mathematician, Scientist and Neohumanist Scholar: Papers from a Sesquicentennial Conference, volume 187 of Boston Studies in the Philosophy of Science. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 1996. ISBN 90-481-4758-1, 94-015-8753-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QA8.9-10.3. URL

https://link.springer.com/book/
10.1007/978-94-015-8753-2.

Schubring:1996:RFG

[Sch96g] Gert Schubring. Remarks on the fate of Grassmann’s Nachlaß. In Hermann Günther Graßmann (1809–1877): Visionary Mathematician, Scientist and Neohumanist Scholar: Papers from a Sesquicentennial Conference [Sch96f], pages 19–26. ISBN 90-481-4758-1, 94-015-8753-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QA8.9-10.3. URL https:/

/link.springer.com/chapter/10.
1007/978-94-015-8753-2_3.

Schwartze:1996:HGL

[Sch96h] Heinz Schwartze. On Hermann Grassmann’s life and his work as a mathematics teacher. In Schubring [Sch96f], pages 7–18. ISBN 90-481-4758-1, 94-015-8753-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QA8.9-10.3. URL https://link.

springer.com/chapter/10.1007/
978-94-015-8753-2_2.

Scheibe:1997:PRT

[Sch97a] Erhard Scheibe. The problem of reduction in the theory of special relativity. In Ginev and Cohen [GC97], chapter 19, pages 321–342. ISBN 0-7923-4444-8 (hardcover), 94-011-5788-X. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 192.

URL https://link.springer.com/
chapter/10.1007/978-94-011-5788-
9_19.

Schramm:1997:IK

[Sch97b] Alfred Schramm. Inductive knowledge. In Austrian Philosophy Past and Present: Essays in Honor of Rudolf Haller [LM97], chapter 15, pages 221–235. ISBN 94-010-6412-1, 94-011-5720-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

                                              ???? URL https://link.springer.
                                              com/chapter/10.1007/978-94-011-
                                              5720-9_15.

Schiemann:1998:LWI

[Sch98] Gregor Schiemann. The loss of world in the image: Origin and development of the concept of image in the thought of Hermann von Helmholtz and Heinrich Hertz. In Heinrich Hertz: Classical Physicist, Modern Philosopher [BHN98], chapter 3, pages 25–38. ISBN 90-481-4881-2, 94-015-8855-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-015-8855-3_3.

Schacht:1999:NTK

[Sch99a] Richard Schacht. Nietzsche: Truth and knowledge. In Nietzsche, Epistemology, and Philosophy of Science: Nietzsche and the Sciences II [Bab99a], chapter 2, pages 25–38. ISBN 90-481-5234-8, 94-017-2428-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

                                              ???? URL https://link.springer.
                                              com/chapter/10.1007/978-94-017-
                                              2428-9_2.

Schmid:1999:NMC

[Sch99b] Holger Schmid. The Nietzschean metacritique of knowledge. In Nietzsche, Theories of Knowledge, and Critical Theory: Nietzsche and the Sciences I [Bab99b], chapter 11, pages 153–163. ISBN 90-481-5233-X, 94-017-2430-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-017-2430-
                                              2_11.

Schulz:1999:WHB

[Sch99c] Reinhard Schulz. Was Heisst Biologie Verstehen? Vorüberlegungen zu einer Hermeneutischen Biologie. (German) [What does biology mean? Preliminary considerations for a hermeneutic biology]. In Fehér et al. [FKR99], pages 171–182. ISBN 90-481-5257-7, 94-015-9293-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-015-
9293-2_14.

Schwendtner:1999:PZH

[Sch99d] Tibor Schwendtner. Parallelen zwischen Heideggers und Kuhns Wissenschaftsauffassung. (German) [Parallels between Heidegger and Kuhn’s theory of science]. In Fehér et al. [FKR99], pages 35–57. ISBN 90-481-5257-7, 94-015-9293-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-015-
9293-2_4.

Scheibe:2000:OSR

[Sch00] Erhard Scheibe. The origin of scientific realism: Boltzmann, Planck, Einstein. In The Reality of the Unobservable: Observability, Unobservability and Their Impact on the Issue of Scientific Realism [AP00b], chapter 2, pages 31– 44. ISBN 90-481-5458-8, 94-015-9391-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-015-9391-
5_2.

Schuller:2001:SCA

[Sch01] Volkmar Schüller. Samuel Clarke’s annotations in Jacques Rohault’s Traité de Physique, and how they contributed to popularising Newton’s physics. In Between Leibniz, Newton, and Kant: Philosophy and Science in the Eighteenth Century [Lef01a], chapter 5, pages 95–110. ISBN 90-481-5774-9, 94-015-9729-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

                                              ???? URL https://link.springer.
                                              com/chapter/10.1007/978-94-015-
                                              9729-6_5.

Scharff:2002:CPH

[Sch02a] Robert C. Scharff. Comte and the possibility of a hermeneutics of science. In Hermeneutic Philosophy of Science, Van Gogh’s Eyes, and God: Essays in Honor of Patrick A. Heelan, S.J. [Bab02c], chapter 9, pages 117–126. ISBN 90-481-5926-1, 94-017-1767-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 225. URL https://link.

                                              springer.com/chapter/10.1007/
                                              978-94-017-1767-0_9.

Schulkin:2002:CNS

[Sch02b] Jay Schulkin. Cognitive neuroscience of social sensibility. In Hermeneutic Philosophy of Science, Van Gogh’s Eyes, and God: Essays in Honor of Patrick A. Heelan, S.J. [Bab02c], chapter 27, pages 315–322. ISBN 90-481-5926-1, 94-017-1767-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 225.

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-017-1767-
                                              0_27.

Schweber:2003:AEF

[Sch03] Silvan S. Schweber. Albert Einstein and the founding of Brandeis University. In Renn et al. [RDS03], chapter 31, pages 615–640. ISBN 1-4020-1284-5, 1-4020-1285-3 (paperback). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 234. URL https:/

                                              /link.springer.com/chapter/10.
                                              1007/978-94-010-0111-3_31.

Scholz:2006:CCM

[Sch06a] Erhard Scholz. The changing concept of matter in H. Weyl’s thought, 1918– 1930. In Hendricks et al. [HJLP06], pages 281–306. ISBN 1-4020-5194-8 (hardcover), 1-4020-5195-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175 .B69 v.251 2006; Q175 .I51377 2006. URL https:

//link.springer.com/chapter/10.
1007/978-1-4020-5195-1_9.

Schummer:2006:PC

[Sch06b] Joachim Schummer. The philosophy of chemistry. In Philosophy of Chemistry: Synthesis of a New Discipline [BSM06b], chapter 2, pages 19–39. ISBN 1-4020-2455-X, 1-4020-3256-0, 1-4020-3261-7. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 242 QD6. URL https:/

/link.springer.com/chapter/10.
1007/1-4020-3261-7_2.

Schemmel:2007:CBC

[Sch07a] Matthias Schemmel. The continuity between classical and relativistic cosmology in the work of Karl Schwarzschild. In Renn et al. [RSS^{+}07a], pages 155–18. ISBN 1-4020-3999-9. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 no. 250 v. 3; QC173.6 .G469

2007. URL https://link.springer.
com/chapter/10.1007/978-1-4020-
4000-9_16.

Schwarzschild:2007:TRU

[Sch07b] Karl Schwarzschild. Things at rest in the universe. In Renn et al. [RSS^{+}07a], pages 183–190. ISBN 1-4020-3999-9. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 no. 250 v. 3; QC173.6 .G469 2007. URL https:

//link.springer.com/chapter/
10.1007/978-1-4020-4000-9_17.

Originally published as “Was in der Welt ruht” in Die Zeit, Vienna 11, No. 142, 181–183, 19 June 1897.

Schweber:2007:P

[Sch07c] Sam Schweber. Postscript. In Positioning the history of science [RG07], chapter 32, pages 185–188. ISBN 1-4020-5419-X (hardcover), 1-4020-5420-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q125 .P77 2007. URL https://link.

springer.com/chapter/10.1007/1-
4020-5420-3_32.

Schweder:2007:SNU

[Sch07d] Rebecca Schweder. Some notes on unificationism and probabilistic explanation. In Rethinking Explanation [PY07], chapter 8, pages 119–128. ISBN 1-4020-5580-3 (paperback), 1-4020-5581-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN BD237 .R48 2007; Q175.32.E97 R48

2007. URL https://link.springer.
com/chapter/10.1007/978-1-4020-
5581-2_8.

Schemmel:2008:EGTa

[Sch08a]  Matthias Schemmel.  The English

Galileo: Thomas Harriot’s work on motion as an example of preclassical mechanics. Volume 1: Interpretation, volume 268(1) of Boston Studies in the Philosophy of Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2008. ISBN 1-4020-5498-X (set: cased), 1-4020-5499-8 (set: e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). xx + 762 (two volumes) pp. LCCN QC133 .S36 2008; Q174 .B67 v.268. URL

https://link.springer.com/book/
10.1007/978-1-4020-5499-0.

Schemmel:2008:EGTb

[Sch08b]  Matthias Schemmel.  The English

Galileo: Thomas Harriot’s work on motion as an example of preclassical mechanics. Volume 2: Sources, volume 268(2) of Boston Studies in the Philosophy of Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2008. ISBN 1-4020-5498-X (set: cased), 1-4020-5499-8 (set: e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). xx + 762 (two volumes) pp. LCCN QC133 .S36 2008; Q174 .B67 v.268. URL

https://link.springer.com/book/
10.1007/978-1-4020-5499-0.

Schiefsky:2008:TPH

[Sch08c] Mark J. Schiefsky. Theory and practice in Heron’s Mechanics. In Mechanics and Natural Philosophy Before the Scientific Revolution [LR08], chapter 1, pages 15–49. ISBN 1-4020-5966-3 (hardcover), 1-4020-5967-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QA802 .M43 2008.

URL https://link.springer.com/
chapter/10.1007/978-1-4020-5967-
4_1.

Schmuhl:2008:KWI

[Sch08d] Hans-Walter Schmuhl, editor. The Kaiser Wilhelm Institute for Anthropology, Human Heredity, and Eugenics, 1927–1945: Crossing Boundaries, volume 259 of Boston Studies in the Philosophy of Science. Springer-Ver-lag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2008. ISBN 1-4020-6599-X (hardcover), 1-4020-6600-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN HQ755.5.G3 S37 2008. URL

https://link.springer.com/book/
10.1007/978-1-4020-6600-9.

Schickore:2009:EHC

[Sch09] Jutta Schickore. Error as historiographical challenge: The infamous globule hypothesis. In Going Amiss In Experimental Research [HSS09a], chapter 3, pages 27–45. ISBN 1-4020-8892-2 (hardcover), 1-4020-8893-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q172.5.E77 G647

                                              2009. URL https://link.springer.
                                              com/chapter/10.1007/978-1-4020-
                                              8893-3_3.

Schuld:2010:LSC

[Sch10] Dawna Schuld. Lost in space: Consciousness and experiment in the work of Irwin and Turrell. In Frigg and Hunter [FH10b], chapter 10, pages 221–244. ISBN 90-481-3850-7 (hardcover), 90-481-3851-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN BH301.M47 B49 2010; Q175 .B52 2010. URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-90-481-3851-7_10.

Schiemann:2011:EMC

[Sch11] Gregor Schiemann. An epoch-making change in the development of science? A critique of the “epochal-break-thesis”. In Science in the Context of Application [CN11a], chapter 25, pages 431–453. ISBN 90-481-9050-9, 90-481-9051-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q180.55.M4 S385 2011. URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-90-481-9051-5_25.

Schickore:2012:SMA

[Sch12a] Jutta Schickore. Scientists’ methods accounts: S. Weir Mitchell’s research on the venom of poisonous snakes. In Mauskopf and Schmaltz [MS12b], chapter 10, pages 141–161. ISBN 94-007-1744-X, 94-007-1745-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175 .I5134 2011.

URL https://link.springer.com/
chapter/10.1007/978-94-007-1745-
9_10.

Schulz:2012:HEP

[Sch12b] Armin W. Schulz. Heuristic evolutionary psychology. In Philosophy of Behavioral Biology [PR12b], chapter 10, pages 217–234. ISBN 94-007-1950-7, 94-007-1951-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QH331 .P4679 2012; Q175 .B657

v.282. URL https://link.springer.
com/chapter/10.1007/978-94-007-
1951-4_10.

Schummer:2015:MPC

[Sch15] Joachim Schummer. The methodological pluralism of chemistry and its philosophical implications. In Philosophy of Chemistry: Growth of a New Discipline [SM15b], chapter 5, pages 57–72. ISBN 94-017-9363-8 (hardcover), 94-017-9364-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QD6 .P455 2015; QD6 .S384 2014. URL https://link.

springer.com/chapter/10.1007/
978-94-017-9364-3_5.

Schliesser:2018:FME

[Sch18] Eric Schliesser. Four methods of empirical inquiry in the aftermath of Newton’s challenge. In What Does it Mean to be an Empiricist?: Empiricisms in Eighteenth Century Sciences [BR18b], chapter 2, pages 15–30. ISBN 3-319-69858-3 (hardcover), 3-319-69860-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN B816 .W43 2018. URL https://link.

springer.com/chapter/10.1007/
978-3-319-69860-1_2.

Scott:1980:PCD

[Sco80] William T. Scott. The personal character of the discovery of mechanisms in cloud physics. In Scientific Discovery: Case Studies [Nic80e], chapter 15, pages 273–289. ISBN 90-277-1093-7, 94-009-9015-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 60. URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-009-9015-9_15.

Scott:1995:PNR

[Sco95] William T. Scott. On Polanyi’s notion of rationality. In Misiek [Mis95c], pages 205–214. ISBN 94-010-4206-3, 94-011-0461-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

                                              ???? URL https://link.springer.
                                              com/chapter/10.1007/978-94-011-
                                              0461-6_15.

Scriba:1974:JPE

[Scr74a] Christoph J. Scriba. John Pell’s English edition of J. H. Rahn’s Teutsche Algebra. In For Dirk Struik: Scientific, Historical and Political Essays in Honor of Dirk J. Struik [CSW74], chapter 22, pages 261–274. ISBN 90-277-0379-5, 94-010-2115-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-010-2115-9_22.

Scriven:1974:ERV

[Scr74b] Michael Scriven. The exact role of value judgments in science. In Schaffner and Cohen [SC74a], pages 219–247. ISBN 90-277-0408-2, 90-277-0409-0, 94-010-2140-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175 .B7312. URL https://link.

                                              springer.com/chapter/10.1007/
                                              978-94-010-2140-1_15.

Scully:1998:NGE

[Scu98] Stephen Scully. The nature of the Gods in Early Greek poetic thought. In Philosophies of Nature: The Human Dimension: In Celebration of Erazim Kohák [CT98], chapter 12, pages 163–176. ISBN 90-481-4859-6, 94-017-2614-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic).

LCCN BD581. URL https://link.
springer.com/chapter/10.1007/
978-94-017-2614-6_12.

Simoes:2015:Ia

[SDG15a] Ana Simões, Maria Paula Diogo, and Kostas Gavroglu. Introduction. In Sciences in the Universities of Europe, Nineteenth and Twentieth Centuries: Academic Landscapes [SDG15b], chapter 1, pages 1–7. ISBN 94-017-9635-1, 94-017-9636-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q183.4.E85 2015. URL https:

//link.springer.com/chapter/10.
1007/978-94-017-9636-1_1.

Simoes:2015:SUE

[SDG15b] Ana Simões, Maria Paula Diogo, and Kostas Gavroglu. Sciences in the Universities of Europe, Nineteenth and Twentieth Centuries: Academic Landscapes, volume 309 of Boston Studies in the Philosophy and History of Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2015. ISBN 94-017-9635-1, 94-017-9636-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). x + 393 pp. LCCN Q183.4.E85 2015. URL

https://link.springer.com/book/
10.1007/978-94-017-9636-1.

Sarlemijn:1992:PRA

[SdV92] Andries Sarlemijn and Marc de Vries. The piecemeal rationality of application-oriented research. In Technological Development and Science in the Industrial Age: New Perspectives on the Science–Technology Relationship [KB92b], chapter 5, pages 99–131. ISBN 90-481-4186-9, 94-015-8010-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-015-8010-
6_5.

Seeger:1970:MCA

[See70] Raymond J. Seeger. On Mach’s curiosity about shockwaves. In Ernst Mach: Physicist and Philosopher [CS70], chapter 4, pages 60–68. ISBN 90-481-8318-9, 90-277-0016-8, 94-017-1462-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-017-1462-
4_4.

Seeger:1974:PIL

[See74] Raymond J. Seeger. On the physical insights of Leonardo Da Vinci. In Seeger and Cohen [SC74b], pages 41– 53. ISBN 90-277-0390-6, 90-277-0376-0 (paperback), 94-010-2126-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 11.

URL https://link.springer.com/
chapter/10.1007/978-94-010-2126-
5_3.

Seebohm:2002:HT

[See02] Thomas M. Seebohm. The hermeneutics of texts. In Hermeneutic Philosophy of Science, Van Gogh’s Eyes, and God: Essays in Honor of Patrick A. Heelan, S.J. [Bab02c], chapter 11, pages 137–152. ISBN 90-481-5926-1, 94-017-1767-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 225. URL https:/

/link.springer.com/chapter/10.
1007/978-94-017-1767-0_11.

Segre:1991:STC

[Seg91] Michael Segre. Science at the Tuscan Court, 1642–1667. In Unguru [Ung91b], pages 295–308. ISBN 94-010-5476-2, 94-011-3342-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC5.53; QC6.9. URL https:

//link.springer.com/chapter/10.
1007/978-94-011-3342-5_12.

Segre:1995:PLF

[Seg95] Michael Segre. Peano, logicism, and formalism. In Critical Rationalism, Metaphysics and Science: Essays for Joseph Agassi. Volume I [JL95a], chapter 8, pages 133–142. ISBN 94-010-4211-X, 94-011-0471-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-011-0471-5_8.

Segal:2000:FAI

[Seg00] Gabriel Segal. Four arguments for the indeterminacy of translation. In Knowledge, Language and Logic: Questions for Quine [OK00], chapter 10, pages 131–139. ISBN 1-4020-0253-X, 94-011-3933-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-011-
3933-5_10.

Segala:2003:BAE

[Seg03] Marco Segala. Babbage, the Analytical Engine and the Turin Academy of Sciences. In Simões et al. [SCD03b], chapter 10, pages 233–247. ISBN 1-4020-1259-4, 94-017-3584-0. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 v. 233.

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-017-3584-
                                              1_10.

Segre:2009:APD

[Seg09] Michael Segre. Applying Popperian didactics. In Rethinking Popper [PC09], chapter 29, pages 389–395. ISBN 1-4020-9337-3, 1-4020-9338-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN B1649.P64 R48 2009; Q174 .B67 2009. URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-1-4020-9338-8_29.

Segal:2015:VHT

[Seg15] Judy Z. Segal. The view from here and there: Objectivity and the rhetoric of breast cancer. In Objectivity in Science: New Perspectives from Science and Technology Studies [PRT15], chapter 11, pages 211–226. ISBN 3-319-14348-4 (hardcover), 3-319-14349-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175 O25

                                              2015. URL https://link.springer.
                                              com/chapter/10.1007/978-3-319-
                                              14349-1_11.

Segre:2017:CRT

[Seg17] Michael Segre. Critical rationalism as therapy. In Encouraging Openness: Essays for Joseph Agassi on the Occasion of His 90th Birthday [BAG17], chapter 40, pages 529–541. ISBN 3-319-57668-2 (hardcover), 3-319-57669-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN B833.E53

                                              2017. URL https://link.springer.
                                              com/chapter/10.1007/978-3-319-
                                              57669-5_40.

Seidenfeld:1992:RFD

[Sei92] T. Seidenfeld. R. A. Fisher on the design of experiments and statistical estimation. In Sarkar [Sar92a], chapter 2, pages 23–36. ISBN 0-7923-1777-7 (hardcover), 94-011-2856-1. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 142. URL https:

//link.springer.com/chapter/10.
1007/978-94-011-2856-8_2.

Seibt:1997:ETS

[Sei97] Johanna Seibt. Existence in time: From substance to process. In Perspectives on Time [FSU97b], chapter 6, pages 143–182. ISBN 90-481-4774-3, 94-015-8875-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol.189. URL https:/

/link.springer.com/chapter/10.
1007/978-94-015-8875-1_6.

Selsam:1974:JEF

[Sel74] Howard Selsam. Jonathan Edwards on the freedom of the will. In For Dirk Struik: Scientific, Historical and Political Essays in Honor of Dirk J. Struik [CSW74], chapter 32, pages 391–402. ISBN 90-277-0379-5, 94-010-2115-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-010-2115-
9_32.

Selten:2003:EFE

[Sel03]  Reinhard Selten.  Emergence and

future of experimental economics. In Descartes’s Mathematical Thought [Sas03d], chapter 4, pages 63–70. ISBN 90-481-6487-7, 94-017-1225-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/0-306-48123-5_4.

Sengor:2005:RID

[Şen05] A. M. C. Şengör. Repeated independent discovery and ‘objective evidence’ in science: An example from geology. In Turkish Studies in the History and Philosophy of Science [IG05c], chapter 9, pages 113–135. ISBN 1-4020-3332-X (hardcover), 1-4020-3333-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 244.

URL https://link.springer.com/
chapter/10.1007/1-4020-3333-8_9.

Sepper:1987:GAN

[Sep87] Dennis L. Sepper. Goethe against Newton: Towards saving the phenomenon. In Amrine et al. [AZW87], chapter 10, pages 175–193. ISBN 90-277-2265-X, 94-009-3761-X. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 97. US$79.00. URL https:

//link.springer.com/chapter/10.
1007/978-94-009-3761-1_10.

Servi:1981:SCI

[Ser81] Gisèle Fischer Servi. Semantics for a class of intuitionistic modal calculi. In Italian Studies in the Philosophy of Science [Chi81b], chapter 5, pages 59– 72. ISBN 90-277-1073-2, 94-009-8937-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-009-8937-
5_5.

Sesiano:2004:I

[Ses04] Jacques Sesiano. Introduction. In Classics in the History of Greek Mathematics [Chr04a], chapter 14, pages 257–263. ISBN 1-4020-0081-2 (hardcover), 1-4020-2640-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QA22 .C596 2004.

URL https://link.springer.com/
chapter/10.1007/978-1-4020-2640-
9_14.

Settle:1976:BTA

[Set76] Tom Settle. The bicentenary of technology assessment. In Cohen et al. [CHMV76], pages 437–447. ISBN 90-277-0647-6, 90-277-0648-4 (paperback), 94-010-1449-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic).

LCCN ????  URL https://link.
springer.com/chapter/10.1007/
978-94-010-1449-6_22.

Settle:1982:LMS

[Set82] Tom Settle. Letter to Mario: The self and its mind. In Scientific Philosophy Today: Essays in Honor of Mario Bunge [AC82], chapter 20, pages 357–379. ISBN 90-277-1262-X, 90-277-1263-8, 94-009-8462-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 67.

URL https://link.springer.com/
chapter/10.1007/978-94-009-8462-
2_20.

Settle:1995:YCS

[Set95] Tom Settle. You can’t have science as your religion! In Critical Rationalism, Metaphysics and Science: Essays for Joseph Agassi. Volume I [JL95a], chapter 6, pages 59–90. ISBN 94-010-4211-X, 94-011-0471-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-011-0471-5_6.

Seymour:1996:TTE

[Sey96] Michel Seymour. “Three thought experiments revisited”. In Québec Studies in the Philosophy of Science: Part II: Biology, Psychology, Cognitive Science and Economics Essays in Honor of Hugues Leblanc [MC96], chapter 8, pages 113–138. ISBN 94-010-6537-3, 94-009-0113-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic).

LCCN ????  URL https://link.
springer.com/chapter/10.1007/
978-94-009-0113-1_8.

Shrader-Frechette:1983:TAP

[SF83] Kristin Shrader-Frechette. Technology assessment and the problem of quantification. In Philosophy and Technology [DR83], chapter 11, pages 151–164. ISBN 94-009-7126-5, 94-009-7124-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-009-7124-
0_11.

Shrader-Frechette:1994:UDG

[SF94] Kristin Shrader-Frechette. Unsafe at any depth: Geological methods, subjective judgments, and nuclear waste disposal. In Artifacts, Representations and Social Practice: Essays for Marx Wartofsky [GC94], chapter 29, pages 501–524. ISBN 94-010-4390-6, 94-011-0902-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN B73 .A77 1994. URL https://link.

springer.com/chapter/10.1007/
978-94-011-0902-4_29.

Schroeder-Gudehus:1992:SEP

[SG92] Brigitte Schroeder-Gudehus. Sciences exactes et politique extérieure. (French) [Exact sciences and foreign policy]. In Science and Empires: Historical Studies about Scientific Development and European Expansion [PJM92], chapter 23, pages 219–223. ISBN 0-7923-1518-9 (hardcover), 94-010-5145-3, 94-011-2594-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 136.

URL https://link.springer.com/

chapter/10.1007/978-94-011-2594-
9_23.

Shavit:2011:TOD

[SG11] Ayelet Shavit and James Griesemer.

Transforming objects into data: How minute technicalities of recording “species location” entrench a basic challenge for biodiversity. In Science in the Context of Application [CN11a], chapter 12, pages 169–193. ISBN 90-481-9050-9, 90-481-9051-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q180.55.M4 S385

2011. URL https://link.springer.
com/chapter/10.1007/978-90-481-
9051-5_12.

Shaik:1992:CCD

[SH91] Sason S. Shaik and Philippe C. Hiberty.

Curve crossing diagrams as general models for chemical reactivity and structure. In The Invention of Physical Science: Intersections of Mathematics, Theology and Natural Philosophy Since the Seventeenth Century: Essays in Honor of Erwin N. Hiebert [NRS92], chapter 8, pages 269–322. ISBN 94-010-5097-X, 94-011-2488-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-3-642-58183-0_8.

Shapere:1974:DRP

[Sha74a] Dudley Shapere. Discovery, rationality, and progress in science: A perspective in the philosophy of science. In Schaffner and Cohen [SC74a], pages 407–419. ISBN 90-277-0408-2, 90-277-0409-0, 94-010-2140-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175 .B7312.

URL https://link.springer.com/
chapter/10.1007/978-94-010-2140-
1_28.

Shapere:1974:NSF

[Sha74b] Dudley Shapere. Natural science and the future of metaphysics. In Methodological and Historical Essays in the Natural and Social Sciences [CW74b], chapter 8, pages 161–171. ISBN 90-277-0378-7, 94-010-2128-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 14.

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-010-2128-
                                              9_8.

Shapley:1974:ALP

[Sha74c] John Shapley. An aspect of Leonardo’s painting. In Seeger and Cohen [SC74b], pages 3–16. ISBN 90-277-0390-6, 90-277-0376-0 (paperback), 94-010-2126-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 11. URL https://link.

                                              springer.com/chapter/10.1007/
                                              978-94-010-2126-5_1.

Shapere:1980:CSC

[Sha80] Dudley Shapere. The character of scientific change. In Scientific Discovery, Logic, and Rationality [Nic80f], chapter 2, pages 61–116. ISBN 90-277-1070-8, 94-009-8986-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-009-8986-3_2.

Shapere:1984:AGL

[Sha84a] Dudley Shapere. Alteration of goals and language in the development of science. In Reason and the Search for Knowledge: Investigations in the Philosophy of Science [Sha84l], chapter 15, pages 325–341. ISBN 90-277-1641-2, 94-010-9731-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-010-
9731-4_15.

Shapere:1984:CSC

[Sha84b] Dudley Shapere. The character of scientific change. In Reason and the Search for Knowledge: Investigations in the Philosophy of Science [Sha84l], chapter 11, pages 205–260. ISBN 90-277-1641-2, 94-010-9731-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-010-9731-4_11.

Shapere:1984:COS

[Sha84c] Dudley Shapere. The concept of observation in science and philosophy. In Reason and the Search for Knowledge: Investigations in the Philosophy of Science [Sha84l], chapter 16, pages 342–351. ISBN 90-277-1641-2, 94-010-9731-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-010-
9731-4_16.

Shapere:1984:ISA

[Sha84d] Dudley Shapere. Interpretations of science in America. In Reason and the Search for Knowledge: Investigations in the Philosophy of Science [Sha84l], chapter 8, pages 156–166. ISBN 90-277-1641-2, 94-010-9731-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-010-9731-4_8.

Shapere:1984:MIM

[Sha84e] Dudley Shapere. Mathematical ideals and metaphysical concepts. In Reason and the Search for Knowledge: Investigations in the Philosophy of Science [Sha84l], chapter 2, pages 26–36. ISBN 90-277-1641-2, 94-010-9731-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-010-9731-4_2.

Shapere:1984:MSC

[Sha84f] Dudley Shapere. Meaning and scientific change. In Reason and the Search for Knowledge: Investigations in the Philosophy of Science [Sha84l], chapter 5, pages 58–101. ISBN 90-277-1641-2, 94-010-9731-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-010-
9731-4_5.

Shapere:1984:MSP

[Sha84g] Dudley Shapere. Modern science and the philosophical tradition. In Reason and the Search for Knowledge: Investigations in the Philosophy of Science [Sha84l], chapter 19, pages 408–417. ISBN 90-277-1641-2, 94-010-9731-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-010-9731-
4_19.

Shapere:1984:NTPa

[Sha84h] Dudley Shapere. Notes toward a postpositivistic interpretation of science, Part I. In Reason and the Search for Knowledge: Investigations in the Philosophy of Science [Sha84l], chapter 6, pages 102–119. ISBN 90-277-1641-2, 94-010-9731-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-010-
9731-4_6.

Shapere:1984:NTPb

[Sha84i] Dudley Shapere. Notes toward a post-positivistic interpretation of science, part II. In Reason and the Search for Knowledge: Investigations in the Philosophy of Science [Sha84l], chapter 17, pages 352–382. ISBN 90-277-1641-2, 94-010-9731-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-010-9731-4_17.

Shapere:1984:PC

[Sha84j] Dudley Shapere. The paradigm concept. In Reason and the Search for Knowledge: Investigations in the Philosophy of Science [Sha84l], chapter 4, pages 49–57. ISBN 90-277-1641-2, 94-010-9731-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-010-
9731-4_4.

Shapere:1984:PAL

[Sha84k] Dudley Shapere. Philosophy and the analysis of language. In Reason and the Search for Knowledge: Investigations in the Philosophy of Science [Sha84l], chapter 1, pages 3–25. ISBN 90-277-1641-2, 94-010-9731-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-010-9731-4_1.

Shapere:1984:RSK

[Sha84l] Dudley Shapere. Reason and the Search for Knowledge: Investigations in the Philosophy of Science, volume 78 of Boston Studies in the Philosophy of Science. D. Reidel, Dordrecht, The Netherlands; Boston, MA, USA; Lancaster, UK; Tokyo, Japan, 1984. ISBN 90-277-1641-2, 94-010-9731-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). 477 (est.) pp. LCCN ????

                                              URL https://link.springer.com/
                                              book/10.1007/978-94-010-9731-4.

Shapere:1984:RRQ

[Sha84m] Dudley Shapere. Reason, reference, and the quest for knowledge. In Reason and the Search for Knowledge: Investigations in the Philosophy of Science [Sha84l], chapter 18, pages 383–407. ISBN 90-277-1641-2, 94-010-9731-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-010-9731-
                                              4_18.

Shapere:1984:RCD

[Sha84n] Dudley Shapere. Remarks on the concepts of domain and field. In Reason and the Search for Knowledge: Investigations in the Philosophy of Science [Sha84l], chapter 14, pages 320–324. ISBN 90-277-1641-2, 94-010-9731-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-010-9731-
                                              4_14.

Shapere:1984:STT

[Sha84o] Dudley Shapere. Scientific theories and their domains. In Reason and the Search for Knowledge: Investigations in the Philosophy of Science [Sha84l], chapter 13, pages 273–319. ISBN 90-277-1641-2, 94-010-9731-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-010-9731-4_13.

Shapere:1984:SLS

[Sha84p] Dudley Shapere. The scope and limits of scientific change. In Reason and the Search for Knowledge: Investigations in the Philosophy of Science [Sha84l], chapter 12, pages 261–272. ISBN 90-277-1641-2, 94-010-9731-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-010-9731-4_12.

Shapere:1984:STL

[Sha84q] Dudley Shapere. Space, time, and language. In Reason and the Search for Knowledge: Investigations in the Philosophy of Science [Sha84l], chapter 7, pages 123–155. ISBN 90-277-1641-2, 94-010-9731-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-010-
9731-4_7.

Shapere:1984:SSR

[Sha84r] Dudley Shapere. The structure of scientific revolutions. In Reason and the Search for Knowledge: Investigations in the Philosophy of Science [Sha84l], chapter 3, pages 37–48. ISBN 90-277-1641-2, 94-010-9731-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-010-9731-4_3.

Shapere:1984:UMC

[Sha84s] Dudley Shapere. Unity and method in contemporary science. In Reason and the Search for Knowledge: Investigations in the Philosophy of Science [Sha84l], chapter 9, pages 167–181. ISBN 90-277-1641-2, 94-010-9731-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-010-9731-
                                              4_9.

Shapere:1984:WCT

[Sha84t] Dudley Shapere. What can the theory of knowledge learn from the history of knowledge? In Reason and the Search for Knowledge: Investigations in the Philosophy of Science [Sha84l], chapter 10, pages 182–202. ISBN 90-277-1641-2, 94-010-9731-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-010-9731-4_10.

Shaham:1986:ELN

[Sha86a] Jacob Shaham. Eddington and the large numbers. In Ullmann-Margalit [UM86c], chapter 16, pages 195–207. ISBN 90-277-2160-2, 90-277-2161-0 (paperback), 94-009-4566-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 95

                                              Q175. URL https://link.springer.
                                              com/chapter/10.1007/978-94-009-
                                              4566-1_16.

Shapin:1986:HSS

[Sha86b] Steven Shapin. History of science and its sociological reconstructions. In Cognition and Fact: Materials on Ludwik Fleck [CS86], chapter 18, pages 325–386. ISBN 94-010-8504-8, 94-009-4498-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

                                              ???? URL https://link.springer.
                                              com/chapter/10.1007/978-94-009-
                                              4498-5_18.

Sharlin:1988:LWB

[Sha88] Harold Issadore Sharlin. Life and work: A biography of Lord Kelvin reconsidered. In Philosophy, History and Social Action: Essays in Honor of Lewis Feuer with an autobiographic essay by Lewis Feuer [HOO88c], chapter 22, pages 405–417. ISBN 94-010-7793-2, 94-009-2873-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-009-
2873-2_22.

Shad:1990:GSN

[Sha90] Jeffrey A. Shad, Jr. The groundwork of Simmel’s new “storey” beneath historical materialism. In Georg Simmel and Contemporary Sociology [KPC90], chapter 17, pages 297–317. ISBN 94-010-6691-4 (hardcover), 94-009-0459-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN HM401-1281.

URL https://link.springer.com/
chapter/10.1007/978-94-009-0459-
0_17.

Shah:1993:MBA

[Sha93] K. Tahir Shah. Minds and brains, algorithms and machines. In Bridging the Gap: Philosophy, Mathematics, and Physics: Lectures on the Foundations of Science [CCG93], chapter 6, pages 125–139. ISBN 94-010-5101-1, 94-011-2496-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-011-
2496-6_6.

Shapiro:1995:RR

[Sha95] Theodore Shapiro. Remembering and reality. In Science, Mind and Art: Essays on science and the humanistic understanding in art, epistemology, religion and ethics In honor of Robert S. Cohen [GSW95b], chapter 24, pages 401–409. ISBN 0-7923-2990-2 (hardcover), 94-011-0469-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 165 BD418.3. URL https://link.

                                              springer.com/chapter/10.1007/
                                              978-94-011-0469-2_24.

Shanahan:1996:RAE

[Sha96]  Timothy Shanahan.  Realism and

antirealism in evolutionary biology. In Cohen et al. [CHR96], pages 449–466. ISBN 90-481-4493-0, 94-015-8638-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-015-8638-
                                              2_33.

Shapere:2000:TE

[Sha00]  Dudley Shapere.  Testability and

empiricism. In The Reality of the Unobservable: Observability, Unobservability and Their Impact on the Issue of Scientific Realism [AP00b], chapter 11, pages 153–164. ISBN 90-481-5458-8, 94-015-9391-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-015-9391-5_11.

Shapere:2001:RRC

[Sha01] Dudley Shapere. Reasons, radical change and incommensurability in science. In Incommensurability and Related Matters [HHS01], chapter 7, pages 181–206. ISBN 90-481-5709-9, 94-015-9680-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

                                              ???? URL https://link.springer.
                                              com/chapter/10.1007/978-94-015-
                                              9680-0_7.

Shapin:2015:KSM

[Sha15] Steven Shapin. Kuhn’s structure: A moment in modern naturalism. In Kuhn’s Structure of Scientific Revolutions — 50 Years On [DB15b], chapter 2, pages 11–21. ISBN 3-319-13382-9, 3-319-13383-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175.3 .K84 2015.

URL https://link.springer.com/
chapter/10.1007/978-3-319-13383-
6_2.

Shapiro:2016:MPP

[Sha16]  Stewart Shapiro.  Mathematics in

philosophy, philosophy in mathematics: Three case studies. In Objectivity, Realism, and Proof: FilMat Studies in the Philosophy of Mathematics [BS16], chapter 1, pages 1–9. ISBN 3-319-31642-7, 3-319-31644-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QA9 .O25 2016.

URL https://link.springer.com/
chapter/10.1007/978-3-319-31644-
4_1.

Shehaby:1975:ISL

[She75] Nabil Shehaby. The influence of stoic logic on Al-Jas.s.ās.’s legal theory. In Murdoch and Sylla [MS75b], pages 61– 85. ISBN 90-277-0587-9, 94-010-1781-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN B721 .I584

1975. URL https://link.springer.
com/chapter/10.1007/978-94-010-
1781-7_3.

Shea:1982:YHQ

[She82] William Shea. The young Hegel’s quest for a philosophy of science, or pitting Kepler against Newton. In Scientific Philosophy Today: Essays in Honor of Mario Bunge [AC82], chapter 21, pages 381–397. ISBN 90-277-1262-X, 90-277-1263-8, 94-009-8462-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 67.

URL https://link.springer.com/
chapter/10.1007/978-94-009-8462-
2_21.

Shea:1987:QSR

[She87] William R. Shea. The quest for scientific rationality: Some historical considerations. In Pitt and Pera [PP87], chapter 7, pages 155–176. ISBN 90-277-2417-2, 94-009-3779-2. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol.

98. US$64.00. URL https://link.

                                              springer.com/chapter/10.1007/
                                              978-94-009-3779-6_7.

Shea:1989:CCC

[She89] William R. Shea. Cartesian clarity and Cartesian motion. In An Intimate Relation: Studies in the History and Philosophy of Science Presented to Robert E. Butts on his 60th Birthday [BM89], chapter 2, pages 23–42. ISBN 94-010-7546-8, 94-009-2327-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-009-2327-0_2.

Shea:1990:GGA

[She90] William R. Shea. Galileo Galilei: An astronomer at work. In Levere and Shea [LS90], chapter 2, pages 51–76. ISBN 0-7923-0420-9, 94-009-1878-X. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 120.

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-009-1878-
                                              8_2.

Shea:1995:TRM

[She95a] William R. Shea. Technology and the rise of the mechanical philosophy. In Québec Studies in the Philosophy of Science: Part I: Logic, Mathematics, Physics and History of Science Essays in Honor of Hugues Leblanc [MC95], chapter 18, pages 297–307. ISBN 94-010-7204-3, 94-009-1575-6 (e-book).

ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-009-1575-6_18.

Shearmur:1995:PMM

[She95b]  Jeremy Shearmur.  Philosophical

method, modified essentialism and the open society. In Critical Rationalism, the Social Sciences and the Humanities: Essays for Joseph Agassi. Volume II [JL95b], chapter 2, pages 19–39. ISBN 0-7923-2961-9, 90-481-4430-2, 94-017-0441-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN B29 .C75 1995. URL https:

//link.springer.com/chapter/10.
1007/978-94-017-0441-0_2.

Shea:2004:RDE

[She04] William R. Shea. The “rational”

Descartes and the “empirical” Galileo. In The Reception of the Galilean Science of Motion in Seventeenth-Century Europe [PT04], chapter 4, pages 67–82. ISBN 1-4020-2454-1, 1-4020-2455-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC133 .R43

2004. URL https://link.springer.
com/chapter/10.1007/978-1-4020-
2455-9_4.

Shearmur:2009:CRE

[She09]  Jeremy Shearmur.  Critical ratio-

nalism and ethics. In Rethinking Popper [PC09], chapter 25, pages 339–356. ISBN 1-4020-9337-3, 1-4020-9338-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN B1649.P64 R48 2009; Q174 .B67 2009.

URL https://link.springer.com/
chapter/10.1007/978-1-4020-9338-
8_25.

Shimony:1963:C

[Shi63] Abner Shimony. Comments. In Wartofsky [War63], pages 174–177. ISBN 94-010-3265-3, 94-010-3263-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B678 1970.

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-010-3263-
                                              6_13.

Shimony:1976:CTE

[Shi76] Abner Shimony. Comments on two epistemological theses of Thomas Kuhn. In Essays in Memory of Imre Lakatos [CFW76], chapter 32, pages 569–588. ISBN 90-277-0655-7, 94-010-1451-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-010-1451-
                                              9_32.

Shimony:1983:RPB

[Shi83] Abner Shimony. Reflections on the philosophy of Bohr, Heisenberg, and Schrödinger. In Cohen and Laudan [CL83], chapter 11, pages 209–221. ISBN 94-009-7055-2, 94-009-7057-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-009-7055-
                                              7_11.

Shimony:1986:RPB

[Shi86] Abner Shimony. Reflections on the philosophy of Bohr, Heisenberg, and Schrödinger. In Cohen and Wartofsky [CW85], chapter 20, pages 305–317. ISBN 90-277-1971-3, 94-009-5345-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-009-5345-
                                              1_20.

Shimony:1987:CA

[Shi87a] Abner Shimony. Comment on Agassi.

In Shimony and Nails [SN87], pages 352–355. ISBN 94-010-8168-9, 94-009-3735-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-009-

3735-2_23. See [Aga87c].

Shimony:1987:CL

[Shi87b] Abner Shimony. Comment on Levine.

In Shimony and Nails [SN87], pages 291–294. ISBN 94-010-8168-9, 94-009-3735-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-009-

3735-2_16. See [Lev87].

Shimony:1987:CR

[Shi87c] Abner Shimony. Comment on Reed.

In Shimony and Nails [SN87], pages 230–234. ISBN 94-010-8168-9, 94-009-3735-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-009-

3735-2_13. See [Ree87a].

Shimony:1987:CS

[Shi87d] Abner Shimony. Comment on Sagal.

In Shimony and Nails [SN87], pages 333–336. ISBN 94-010-8168-9, 94-009-3735-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-009-

3735-2_20. See [Sag87].

Shimony:1987:CC

[Shi87e] Abner Shimony. Comment on Čapek.

In Shimony and Nails [SN87], pages 112–114. ISBN 94-010-8168-9, 94-009-3735-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-009-

3735-2_6. See [Čap87a] and reply [Čap87b].

Shimony:1987:CWb

[Shi87f] Abner Shimony. Comment on Wartofsky. In Shimony and Nails [SN87], pages 375–377. ISBN 94-010-8168-9, 94-009-3735-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic).

LCCN ????  URL https://link.
springer.com/chapter/10.1007/
978-94-009-3735-2_25. See [War87].

Shimony:1987:CWa

[Shi87g] Abner Shimony. Comment on Webb. In Shimony and Nails [SN87], pages 71– 72. ISBN 94-010-8168-9, 94-009-3735-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-009-3735-

2_3. See [Web87].

Shimony:1987:IE

[Shi87h] Abner Shimony. Integral epistemology. In Shimony and Nails [SN87], pages 299–318. ISBN 94-010-8168-9, 94-009-3735-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-009-
3735-2_18.

Shimony:1987:I

[Shi87i] Abner Shimony. Introduction. In Shimony and Nails [SN87], pages 1–13. ISBN 94-010-8168-9, 94-009-3735-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-009-3735-
2_1.

Shimony:1990:TNS

[Shi90] Abner Shimony. The theory of natural selection as a null theory. In Statistics in Science: The Foundations of Statistical Methods in Biology, Physics and Economics [CC90], chapter 2, pages 15–26. ISBN 94-010-6765-1, 94-009-0619-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QA276-280. URL https:/

/link.springer.com/chapter/10.
1007/978-94-009-0619-8_2.

Shimony:1994:TPP

[Shi94] Abner Shimony. Ten philosophical poems. In Artifacts, Representations and Social Practice: Essays for Marx Wartofsky [GC94], chapter 19, pages 343–348. ISBN 94-010-4390-6, 94-011-0902-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN B73 .A77 1994. URL https://link.

springer.com/chapter/10.1007/
978-94-011-0902-4_19.

Shimony:1995:C

[Shi95] Abner Shimony. The confrontation.

In Critical Rationalism, Metaphysics and Science: Essays for Joseph Agassi. Volume I [JL95a], chapter 3, pages 29– 32. ISBN 94-010-4211-X, 94-011-0471-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-011-0471-
5_3.

Shiying:1996:RIQ

[Shi96] Mao Shiying. A realistic interpretation of quantum mechanics. In Cohen et al. [CHR96], pages 381–393. ISBN 90-481-4493-0, 94-015-8638-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-015-8638-2_28.

Shimony:1998:RBP

[Shi98]  Abner Shimony.  The relationship

between physics and philosophy. In Philosophies of Nature: The Human Dimension: In Celebration of Erazim Kohák [CT98], chapter 13, pages 177–184. ISBN 90-481-4859-6, 94-017-2614-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic).

                                              LCCN BD581. URL https://link.
                                              springer.com/chapter/10.1007/
                                              978-94-017-2614-6_13.

Shinn:2007:WNS

[Shi07] Terry Shinn. What’s new in science? In Positioning the history of science [RG07], chapter 25, pages 143–148. ISBN 1-4020-5419-X (hardcover), 1-4020-5420-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q125 .P77 2007. URL https://link.

                                              springer.com/chapter/10.1007/1-
                                              4020-5420-3_25.

Shramko:1997:TN

[Shr97] Yaroslav V. Shramko. Time and negation. In Perspectives on Time [FSU97b], chapter 17, pages 399–416. ISBN 90-481-4774-3, 94-015-8875-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol.189. URL https://link.

                                              springer.com/chapter/10.1007/
                                              978-94-015-8875-1_17.

Siegel:1963:OAH

[Sie63] Armand Siegel. Operational aspects of hidden-variable quantum theories with a postscript on the impact of recent scientific trends on art. In Wartofsky [War63], pages 156–173. ISBN 94-010-3265-3, 94-010-3263-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B678 1970.

URL https://link.springer.com/
chapter/10.1007/978-94-010-3263-
6_12.

Siemens:1971:LET

[Sie71] Warren D. Siemens. A logical empiricist theory of scientific change? In Buck and Cohen [BC71], pages 524–535. ISBN 90-277-0309-4, 94-010-3142-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-010-3142-
4_35.

Siemianowski:1982:TKC

[Sie82] Andrzej Siemianowski. On two kinds of conventionalism with respect to empirical sciences. In Polish Essays in the Philosophy of the Natural Sciences [Kra82c], chapter 3, pages 31–42. ISBN 94-009-7705-0 (e-book), 90-277-1287-5. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-009-7705-1_3.

Siegel:1984:TMC

[Sie84] Armand Siegel. Tracking down the misplaced concreton in the neurosciences. In Methodology, Metaphysics and the History of Science: In Memory of Benjamin Nelson [CW84b], chapter 10, pages 247–262. ISBN 94-009-6333-5, 94-009-6331-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-009-
6331-3_10.

Siebrand:1986:SRM

[Sie86] Heine Siebrand. Spinoza and the rise of modern science in The Netherlands. In Spinoza and the Sciences [GN86], chapter 3, pages 61–91. ISBN 94-010-8511-0, 94-009-4514-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-009-4514-2_3.

Siegel:1995:TCM

[Sie95] Daniel M. Siegel. Text and context in Maxwell’s electromagnetic theory. In No Truth Except in the Details: Essays in Honor of Martin J. Klein [KS95], chapter 12, pages 281–297. ISBN 94-010-4097-4, 94-011-0217-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-011-0217-9_12.

Siegel:2001:IRR

[Sie01] Harvey Siegel. Incommensurability, rationality and relativism: In science, culture and science education. In Incommensurability and Related Matters [HHS01], chapter 8, pages 207–224. ISBN 90-481-5709-9, 94-015-9680-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-015-9680-
                                              0_8.

Sigurdsson:1992:EPP

[Sig91] Skuli Sigurdsson. Equivalence, pragmatic Platonism, and discovery of the calculus. In The Invention of Physical Science: Intersections of Mathematics, Theology and Natural Philosophy Since the Seventeenth Century: Essays in Honor of Erwin N. Hiebert [NRS92], chapter 5, pages 97–116. ISBN 94-010-5097-X, 94-011-2488-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-011-2488-1_5.

Sigurdsson:1994:UGA

[Sig94] Skuli Sigurdsson. Unification, geometry and ambivalence: Hilbert, Weyl and the Göttingen community. In Trends in the Historiography of Science [GCN94], chapter 25, pages 355–367. ISBN 90-481-4264-4, 94-017-3596-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-017-3596-4_25.

Sigurdsson:2007:R

[Sig07] Skúli Sigurdsson. On the road. In Positioning the history of science [RG07], chapter 26, pages 149–157. ISBN 1-4020-5419-X (hardcover), 1-4020-5420-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q125 .P77 2007. URL https://link.

springer.com/chapter/10.1007/1-
4020-5420-3_26.

Silverman:1989:TCS

[Sil89] David Silverman. Telling convincing stories: A plea for cautious positivism in case-studies. In The Qualitative-Quantitative Distinction in the Social Sciences [GM89b], chapter 5, pages 57– 77. ISBN 90-481-8460-6, 94-017-3444-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 112. URL https://link.

springer.com/chapter/10.1007/
978-94-017-3444-8_5.

Silber:1995:BC

[Sil95] John Silber. On Bob Cohen. In Science, Mind and Art: Essays on science and the humanistic understanding in art, epistemology, religion and ethics In honor of Robert S. Cohen [GSW95b], chapter 20, pages 345–347. ISBN 0-7923-2990-2 (hardcover), 94-011-0469-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 165 BD418.3. URL https:

//link.springer.com/chapter/10.
1007/978-94-011-0469-2_20.

Silva:2011:JAN

[Sil11] Cibelle Celestino Silva. Jean Antoine Nollet’s contributions to the institutionalization of physics during the 18th Century. In Brazilian Studies in Philosophy and History of Science: An account of recent works [KV11], chapter 8, pages 131–140. ISBN 90-481-9421-0 (hardcover), 90-481-9422-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175.3 .B73

2011. URL https://link.springer.
com/chapter/10.1007/978-90-481-
9422-3_8.

Simon:1977:AVD

[Sim77a] Herbert A. Simon. Aggregation of variables in dynamic systems. In Models of Discovery: and Other Topics in the Methods of Science [Sim77m], chapter 12, pages 183–213. ISBN 90-277-0970-X, 94-010-9521-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-010-9521-1_12.

Simon:1977:APT

[Sim77b] Herbert A. Simon. The axiomatization of physical theories. In Models of Discovery: and Other Topics in the Methods of Science [Sim77m], chapter 24, pages 388–402. ISBN 90-277-0970-X, 94-010-9521-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-010-9521-1_24.

Simon:1977:ANM

[Sim77c] Herbert A. Simon. The axioms of Newtonian mechanics. In Models of Discovery: and Other Topics in the Methods of Science [Sim77m], chapter 20, pages 349–369. ISBN 90-277-0970-X, 94-010-9521-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-010-9521-1_20.

Simon:1977:COI

[Sim77d] Herbert A. Simon. Causal ordering and identifiability. In Models of Discovery: and Other Topics in the Methods of Science [Sim77m], chapter 5, pages 53– 80. ISBN 90-277-0970-X, 94-010-9521-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-010-9521-
1_5.

Simon:1977:CC

[Sim77e] Herbert A. Simon. Cause and counterfactual. In Models of Discovery: and Other Topics in the Methods of Science [Sim77m], chapter 8, pages 107–131. ISBN 90-277-0970-X, 94-010-9521-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-010-9521-
1_8.

Simon:1977:DTP

[Sim77f] Herbert A. Simon. Definable terms and primitives in axiom systems. In Models of Discovery: and Other Topics in the Methods of Science [Sim77m], chapter 22, pages 376–386. ISBN 90-277-0970-X, 94-010-9521-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-010-9521-1_22.

Simon:1977:DAC

[Sim77g] Herbert A. Simon. Discussion: The axiomatization of classical mechanics. In Models of Discovery: and Other Topics in the Methods of Science [Sim77m], chapter 21, pages 370–375. ISBN 90-277-0970-X, 94-010-9521-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-010-9521-1_21.

Simon:1977:DMP

[Sim77h] Herbert A. Simon. Discussion: The Meno Paradox. In Models of Discovery: and Other Topics in the Methods of Science [Sim77m], chapter 19, pages 338–341. ISBN 90-277-0970-X, 94-010-9521-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-010-
9521-1_19.

Simon:1977:DSD

[Sim77i] Herbert A. Simon. Does scientific discovery have a logic? In Models of Discovery: and Other Topics in the Methods of Science [Sim77m], chapter 18, pages 326–337. ISBN 90-277-0970-X, 94-010-9521-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-010-9521-1_18.

Simon:1977:IST

[Sim77j] Herbert A. Simon. Identifiability and the status of theoretical terms. In Models of Discovery: and Other Topics in the Methods of Science [Sim77m], chapter 26, pages 422–440. ISBN 90-277-0970-X, 94-010-9521-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-010-9521-1_26.

Simon:1977:LHD

[Sim77k] Herbert A. Simon. The logic of heuristic decision making. In Models of Discovery: and Other Topics in the Methods of Science [Sim77m], chapter 10, pages 154–175. ISBN 90-277-0970-X, 94-010-9521-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-010-
9521-1_10.

Simon:1977:LRD

[Sim77l] Herbert A. Simon. The logic of rational decision. In Models of Discovery: and Other Topics in the Methods of Science [Sim77m], chapter 9, pages 137–153. ISBN 90-277-0970-X, 94-010-9521-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-010-9521-
1_9.

Simon:1977:MDO

[Sim77m] Herbert A. Simon. Models of Discovery:

and Other Topics in the Methods of Science, volume 54 of Boston Studies in the Philosophy of Science. D. Reidel, Dordrecht, The Netherlands; Boston, MA, USA; Lancaster, UK; Tokyo, Japan, 1977. ISBN 90-277-0970-X, 94-010-9521-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). 475 (est.) pp. LCCN ???? URL

https://link.springer.com/book/
10.1007/978-94-010-9521-1.

Simon:1977:NAE

[Sim77n] Herbert A. Simon. A note on almosteverywhere definability. In Models of Discovery: and Other Topics in the Methods of Science [Sim77m], chapter 23, page 387. ISBN 90-277-0970-X, 94-010-9521-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic).

                                              LCCN ????  URL https://link.
                                              springer.com/chapter/10.1007/
                                              978-94-010-9521-1_23.

Simon:1977:JPT

[Sim77o] Herbert A. Simon. On judging the plausibility of theories. In Models of Discovery: and Other Topics in the Methods of Science [Sim77m], chapter 4, pages 25–45. ISBN 90-277-0970-X, 94-010-9521-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-010-9521-1_4.

Simon:1977:DCR

[Sim77p] Herbert A. Simon. On the definition of the causal relation. In Models of Discovery: and Other Topics in the Methods of Science [Sim77m], chapter 6, pages 81–92. ISBN 90-277-0970-X, 94-010-9521-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-010-9521-1_6.

Simon:1977:OCS

[Sim77q] Herbert A. Simon. The organization of complex systems. In Models of Discovery: and Other Topics in the Methods of Science [Sim77m], chapter 14, pages 245–261. ISBN 90-277-0970-X, 94-010-9521-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-010-
9521-1_14.

Simon:1977:PHC

[Sim77r] Herbert A. Simon. Prediction and hindsight as confirmatory evidence. In Models of Discovery: and Other Topics in the Methods of Science [Sim77m], chapter 3, pages 20–24. ISBN 90-277-0970-X, 94-010-9521-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-010-9521-1_3.

Simon:1977:RET

[Sim77s] Herbert A. Simon. Ramsey eliminability and the testability of scientific theories. In Models of Discovery: and Other Topics in the Methods of Science [Sim77m], chapter 25, pages 403–421. ISBN 90-277-0970-X, 94-010-9521-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-010-9521-
1_25.

Simon:1977:SDP

[Sim77t] Herbert A. Simon. Scientific discovery and the psychology of problem solving. In Models of Discovery: and Other Topics in the Methods of Science [Sim77m], chapter 16, pages 286–303. ISBN 90-277-0970-X, 94-010-9521-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-010-9521-
1_16.

Simon:1977:SCC

[Sim77u] Herbert A. Simon. Spurious correlation: A causal interpretation. In Models of Discovery: and Other Topics in the Methods of Science [Sim77m], chapter 7, pages 93–106. ISBN 90-277-0970-X, 94-010-9521-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-010-9521-1_7.

Simon:1977:STB

[Sim77v] Herbert A. Simon. Statistical tests as a basis for ‘yes-no’ choices. In Models of Discovery: and Other Topics in the Methods of Science [Sim77m], chapter 2, pages 14–19. ISBN 90-277-0970-X, 94-010-9521-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-010-9521-1_2.

Simon:1977:SIS

[Sim77w] Herbert A. Simon. The structure of ill-structured problems. In Models of Discovery: and Other Topics in the Methods of Science [Sim77m], chapter 17, pages 304–325. ISBN 90-277-0970-X, 94-010-9521-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-010-9521-1_17.

Simon:1977:STH

[Sim77x] Herbert A. Simon. Symmetric tests of the hypothesis that the mean of one normal population exceeds that of another. In Models of Discovery: and Other Topics in the Methods of Science [Sim77m], chapter 1, pages 7–13. ISBN 90-277-0970-X, 94-010-9521-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-010-9521-1_1.

Simon:1977:TAD

[Sim77y] Herbert A. Simon. Theory of automata:

Discussion. In Models of Discovery: and Other Topics in the Methods of Science [Sim77m], chapter 11, page 182. ISBN 90-277-0970-X, 94-010-9521-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-010-9521-1_11.

Simon:1977:TPS

[Sim77z] Herbert A. Simon. The theory of problem solving. In Models of Discovery: and Other Topics in the Methods of Science [Sim77m], chapter 13, pages 214–244. ISBN 90-277-0970-X, 94-010-9521-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-010-
9521-1_13.

Simon:1977:TC

[Sim77-27] Herbert A. Simon. Thinking by computers. In Models of Discovery: and Other Topics in the Methods of Science [Sim77m], chapter 15, pages 268–285. ISBN 90-277-0970-X, 94-010-9521-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-010-9521-
1_15.

Simpson:1994:CDR

[Sim94] Lorenzo C. Simpson. Community and difference: Reflections in the wake of Rodney King. In Artifacts, Representations and Social Practice: Essays for Marx Wartofsky [GC94], chapter 30, pages 525–542. ISBN 94-010-4390-6, 94-011-0902-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic).

LCCN B73 .A77 1994. URL https:

//link.springer.com/chapter/10.
1007/978-94-011-0902-4_30.

Simon:1999:GTN

[Sim99] Josef Simon. Grammar and truth: On Nietzsche’s relationship to the speculative sentential grammar of the metaphysical tradition. In Nietzsche, Theories of Knowledge, and Critical Theory: Nietzsche and the Sciences I [Bab99b], chapter 10, pages 129–151. ISBN 90-481-5233-X, 94-017-2430-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-017-2430-
2_10.

Simanke:2011:FPM

[Sim11a] Richard Theisen Simanke. Freudian psychoanalysis as a model for overcoming the duality between natural and human sciences. In Brazilian Studies in Philosophy and History of Science: An account of recent works [KV11], chapter 15, pages 211–221. ISBN 90-481-9421-0 (hardcover), 90-481-9422-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175.3 .B73

2011. URL https://link.springer.
com/chapter/10.1007/978-90-481-
9422-3_15.

Simon:2011:PAM

[Sim11b] Samuel Simon. The problem of adequacy of mathematics to physics: The relativity theory case. In Brazilian Studies in Philosophy and History of Science: An account of recent works [KV11], chapter 25, pages 325–340. ISBN 90-481-9421-0 (hardcover), 90-481-9422-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175.3 .B73 2011. URL https:/

/link.springer.com/chapter/10.
1007/978-90-481-9422-3_25.

Simoes:2015:DUC

[Sim15] Ana Simões. Different undertakings, common practices: Some directions for the history of science. In Arabatzis et al. [ARS15], chapter 20, pages 297– 312. ISBN 3-319-14552-5 (paperback), 3-319-14553-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q125 .R45 2015. URL https:

//link.springer.com/chapter/10.
1007/978-3-319-14553-2_20.

Sivin:1984:RNT

[Siv84] N. Sivin. Reflections on ‘nature on trial’. In Methodology, Metaphysics and the History of Science: In Memory of Benjamin Nelson [CW84b], chapter 14, pages 323–330. ISBN 94-009-6333-5, 94-009-6331-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-009-
6331-3_14.

Skagestad:1983:PPP

[Ska83] Peter Skagestad. Peirce and Pearson:

Pragmatism vs. instrumentalism. In Language, Logic and Method [CW83b], chapter 14, pages 263–282. ISBN 94-009-7704-2, 94-009-7702-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-009-7702-0_14.

Skagestad:1987:PCT

[Ska87] Peter Skagestad. Peirce’s conception of truth: A framework for naturalistic epistemology? In Shimony and Nails [SN87], pages 73–90. ISBN 94-010-8168-9, 94-009-3735-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-009-3735-2_4.

Skarda:1992:PCC

[Ska92] Christine A. Skarda. Perception, connectionism, and cognitive science. In Understanding Origins: Contemporary Views on the Origin of Life, Mind and Society [VD92], chapter 14, pages 265–271. ISBN 90-481-4090-0, 94-015-8054-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

                                              ???? URL https://link.springer.
                                              com/chapter/10.1007/978-94-015-
                                              8054-0_14.

Sklar:1974:EPU

[Skl74] Lawrence Sklar. The evolution of the problem of the unity of science. In Seeger and Cohen [SC74b], pages 535– 545. ISBN 90-277-0390-6, 90-277-0376-0 (paperback), 94-010-2126-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 11.

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-010-2126-
                                              5_28.

Sklar:1976:TSM

[Skl76] Lawrence Sklar. Thermodynamics, statistical mechanics and the complexity of reductions. In Cohen et al. [CHMV76], pages 15–32. ISBN 90-277-0647-6, 90-277-0648-4 (paperback), 94-010-1449-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-010-1449-
                                              6_2.

Sklar:2006:WDS

[Skl06] Lawrence Sklar. Why does the standard measure work in statistical mechanics? In Hendricks et al. [HJLP06], pages 307–320. ISBN 1-4020-5194-8 (hardcover), 1-4020-5195-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175 .B69 v.251 2006; Q175 .I51377 2006. URL https:

//link.springer.com/chapter/10.
1007/978-1-4020-5195-1_10.

Skolimowski:1976:ER

[Sko76a] Henryk Skolimowski. Evolutionary rationality. In Cohen et al. [CHMV76], pages 191–213. ISBN 90-277-0647-6, 90-277-0648-4 (paperback), 94-010-1449-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-010-
1449-6_9.

Skolimowski:1976:TAC

[Sko76b] Henryk Skolimowski. Technology assessment as a critique of a civilization. In Cohen et al. [CHMV76], pages 459–465. ISBN 90-277-0647-6, 90-277-0648-4 (paperback), 94-010-1449-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-010-1449-
6_24.

Skvorecky:1988:CLF

[Škv88] Josef Škvorecký. The case of Lewis S. Feuer, crime writer. In Philosophy, History and Social Action: Essays in Honor of Lewis Feuer with an autobiographic essay by Lewis Feuer [HOO88c], chapter 23, pages 419–423. ISBN 94-010-7793-2, 94-009-2873-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-009-2873-
2_23.

Skyrms:1983:ZPM

[Sky83] Brian Skyrms. Zeno’s paradox of measure. In Cohen and Laudan [CL83], chapter 12, pages 223–254. ISBN 94-009-7055-2, 94-009-7057-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-009-7055-7_12.

Szumilewicz-Lachman:1982:OMA

[SL82]  Irena Szumilewicz-Lachman.  Once

more about empirical support. In Polish Essays in the Philosophy of the Natural Sciences [Kra82c], chapter 5, pages 51–62. ISBN 94-009-7705-0 (ebook), 90-277-1287-5. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

                                              ???? URL https://link.springer.
                                              com/chapter/10.1007/978-94-009-
                                              7705-1_5.

Szumilewicz-Lachman:1984:PVR

[SL84] Irena Szumilewicz-Lachman. Poincaré versus Le Roy on incommensurability. In Rationality in Science and Politics [And84d], chapter 15, pages 261–275. ISBN 90-277-1953-5, 94-009-6254-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-009-6254-
                                              5_15.

Slade:1989:CTL

[Sla89] Joseph W. Slade. Conceptualizing technology in literary terms: Some American examples. In Literature and Science as Modes of Expression [Amr89], chapter 9, pages 153–168. ISBN 94-010-7531-X, 94-009-2297-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

                                              URL https://link.springer.com/

chapter/10.1007/978-94-009-2297-
6_9.

Sleinis:1999:BNL

[Sle99] E. E. Sleinis. Between Nietzsche and Leibniz: Perspectivism and irrationalism. In Nietzsche, Theories of Knowledge, and Critical Theory: Nietzsche and the Sciences I [Bab99b], chapter 5, pages 67–76. ISBN 90-481-5233-X, 94-017-2430-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-017-2430-2_5.

Slobodkin:1976:SE

[Slo76] Lawrence B. Slobodkin. The strategy of evolution. In Topics in the Philosophy of Biology [GM76], chapter 12, pages 267–285. ISBN 90-277-0596-8, 94-010-1829-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-010-
1829-6_12.

Slobodkin:1983:PES

[Slo83] Lawrence Slobodkin. The peculiar evolutionary strategy of man. In Epistemology, Methodology, and the Social Sciences [CW83a], chapter 9, pages 227–248. ISBN 90-481-8376-6, 94-017-1458-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-017-
1458-7_9.

Spencer:1974:UAH

[SM74] Domina Eberle Spencer and Parry Moon. A unified approach to hypernumbers. In For Dirk Struik: Scientific, Historical and Political Essays in Honor of Dirk J. Struik [CSW74], chapter 9, pages 101–119. ISBN 90-277-0379-5, 94-010-2115-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-010-2115-
9_9.

Sfendoni-Mentzou:1989:PPO

[SM89] D. Sfendoni-Mentzou. Popper’s propensities: An ontological interpretation of probability. In Imre Lakatos and Theories of Scientific Change [GGN89], chapter 32, pages 441–455. ISBN 94-010-7860-2, 94-009-3025-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-009-3025-4_32.

Sfendoni-Mentzou:1990:MCC

[SM90] Demetra Sfendoni-Mentzou. Models of change: A common ground for Ancient Greek philosophy and modern science. In Greek Studies in the Philosophy and History of Science [Nic90], chapter 10, pages 149–169. ISBN 94-010-7403-8, 94-009-2015-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-009-
2015-6_10.

Sagan:1991:EUS

[SM91] Dorion Sagan and Lynn Margulis. Epilogue: The uncut self. In Organism and the Origins of Self [Tau91i], chapter 18, pages 361–374. ISBN 0-7923-1185-X, 94-011-3406-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QR180-189.5. URL https:/

/link.springer.com/chapter/10.
1007/978-94-011-3406-4_18.

Sfendoni-Mentzou:1996:RTP

[SM96] Demetra Sfendoni-Mentzou. The reality of thirdness — a potential–pragmatic account of laws of nature. In Cohen et al. [CHR96], pages 75–95. ISBN 90-481-4493-0, 94-015-8638-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-015-8638-2_6.

Smeenk:2007:MTM

[SM07] Christopher Smeenk and Christopher Martin. Mie’s theories of matter and gravitation. In Renn et al. [RSS^{+}07b], pages 623–632. ISBN 1-4020-3999-9 (hardcover), 1-4020-4000-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 no. 250 v. 4; QC173.6 .G469 2007. URL https:

//link.springer.com/chapter/10.
1007/978-1-4020-4000-9_35.

Scerri:2015:I

[SM15a] Eric Scerri and Lee McIntyre. Introduction. In Philosophy of Chemistry: Growth of a New Discipline [SM15b], chapter 1, pages 1–4. ISBN 94-017-9363-8 (hardcover), 94-017-9364-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QD6 .P455 2015; QD6 .S384 2014. URL https:

//link.springer.com/chapter/10.
1007/978-94-017-9364-3_1.

Scerri:2015:PCG

[SM15b] Eric Scerri and Lee McIntyre. Philosophy of Chemistry: Growth of a New Discipline, volume 306 of Boston Studies in the Philosophy and History of Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2015. ISBN 94-017-9363-8 (hardcover), 94-017-9364-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). xii + 233 pp. LCCN QD6 .P455 2015; QD6 .S384 2014. URL

                                              https://link.springer.com/book/
                                              10.1007/978-94-017-9364-3.

Smart:1989:MO

[Sma89a] J. J. C. Smart. Methodology and ontology. In Imre Lakatos and Theories of Scientific Change [GGN89], chapter 4, pages 47–57. ISBN 94-010-7860-2, 94-009-3025-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

                                              ???? URL https://link.springer.
                                              com/chapter/10.1007/978-94-009-
                                              3025-4_4.

Smart:1989:NU

[Sma89b] J. J. C. Smart. Negative utilitarianism. In Freedom and Rationality: Essays in Honor of John Watkins From his Colleagues and Friends [DJ89], chapter 3, pages 35–46. ISBN 94-010-7571-9, 94-009-2380-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

                                              ???? URL https://link.springer.
                                              com/chapter/10.1007/978-94-009-
                                              2380-5_3.

Smart:1992:WFR

[Sma92] J. J. C. Smart. Wittgenstein, following a rule, and scientific psychology. In The Scientific Enterprise: The Bar-Hillel Colloquium: Studies in History, Philosophy, and Sociology of Science, Volume 4 [UM92], chapter 7, pages 123–137. ISBN 94-010-5190-9, 94-011-2688-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174

                                              .S354 1992. URL https://link.
                                              springer.com/chapter/10.1007/
                                              978-94-011-2688-5_7.

Small:1999:WS

[Sma99] Robin Small. We sensualists. In Nietzsche, Epistemology, and Philosophy of Science: Nietzsche and the Sciences II [Bab99a], chapter 6, pages 73–89. ISBN 90-481-5234-8, 94-017-2428-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-017-2428-9_6.

Smith:1982:SC

[Smi82] John Maynard Smith. Symbolism and chance. In Scientific Philosophy Today: Essays in Honor of Mario Bunge [AC82], chapter 12, pages 201–206. ISBN 90-277-1262-X, 90-277-1263-8, 94-009-8462-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 67. URL https:

//link.springer.com/chapter/10.
1007/978-94-009-8462-2_12.

Smith:1989:QSQ

[Smi89] Charles W. Smith. The qualitative significance of quantitative representation. In The Qualitative-Quantitative Distinction in the Social Sciences [GM89b], chapter 3, pages 29–42. ISBN 90-481-8460-6, 94-017-3444-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 112.

URL https://link.springer.com/
chapter/10.1007/978-94-017-3444-
8_3.

Smith:1990:ADP

[Smi90] A. Mark Smith. Alhazen’s debt to Ptolemy’s optics. In Levere and Shea [LS90], chapter 6, pages 147–164. ISBN 0-7923-0420-9, 94-009-1878-X. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 120.

URL https://link.springer.com/
chapter/10.1007/978-94-009-1878-
8_6.

Smith:1992:JBH

[Smi92] John Maynard Smith. J. B. S. Haldane. In Sarkar [Sar92a], chapter 3, pages 37–51. ISBN 0-7923-1777-7 (hardcover), 94-011-2856-1. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 142. URL https:/

                                              /link.springer.com/chapter/10.
                                              1007/978-94-011-2856-8_3.

Smith:1994:GAW

[Smi94] Gary Smith. A genealogy of ‘aura’: Walter Benjamin’s idea of beauty. In Artifacts, Representations and Social Practice: Essays for Marx Wartofsky [GC94], chapter 7, pages 105–119. ISBN 94-010-4390-6, 94-011-0902-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN B73 .A77 1994.

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-011-0902-
                                              4_7.

Smith:1995:CR

[Smi95a] Charles W. Smith. A closet realist. In Science, Mind and Art: Essays on science and the humanistic understanding in art, epistemology, religion and ethics In honor of Robert S. Cohen [GSW95b], chapter 12, pages 187–201. ISBN 0-7923-2990-2 (hardcover), 94-011-0469-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 165 BD418.3. URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-011-0469-2_12.

Smith:1995:RMW

[Smi95b] Ruth L. Smith. Relational morality: Which relations, which morals? In Science, Mind and Art: Essays on science and the humanistic understanding in art, epistemology, religion and ethics In honor of Robert S. Cohen [GSW95b], chapter 19, pages 327–343. ISBN 0-7923-2990-2 (hardcover), 94-011-0469-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 165 BD418.3. URL https:

//link.springer.com/chapter/10.
1007/978-94-011-0469-2_19.

Smith:1997:NHT

[Smi97] Barry Smith. The Neurath–Haller thesis: Austria and the rise of scientific philosophy. In Austrian Philosophy Past and Present: Essays in Honor of Rudolf Haller [LM97], chapter 1, pages 1–20. ISBN 94-010-6412-1, 94-011-5720-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-011-
5720-9_1.

Smolin:2003:TSE

[Smo03] Lee Smolin. Time, structure and evolution in cosmology. In Renn et al. [RDS03], chapter 10, pages 221–274. ISBN 1-4020-1284-5, 1-4020-1285-3 (paperback). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 234. URL https://link.

springer.com/chapter/10.1007/
978-94-010-0111-3_10.

Shimony:1987:NES

[SN87] Abner Shimony and Debra Nails, editors. Naturalistic Epistemology: A Symposium of Two Decades, volume 100 of Boston Studies in the Philosophy of Science. D. Reidel, Dordrecht, The Netherlands; Boston, MA, USA; Lancaster, UK; Tokyo, Japan, 1987. ISBN 94-010-8168-9, 94-009-3735-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
book/10.1007/978-94-009-3735-2.

Schwarz:2011:PET

[SN11] Astrid Schwarz and Alfred Nordmann. The political economy of technoscience. In Science in the Context of Application [CN11a], chapter 19, pages 317–336. ISBN 90-481-9050-9, 90-481-9051-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q180.55.M4 S385 2011. URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-90-481-9051-5_19.

Sneed:1989:MMG

[Sne89] Joseph D. Sneed. Machine models for the growth of knowledge: Theory nets in Prolog. In Imre Lakatos and Theories of Scientific Change [GGN89], chapter 19, pages 245–268. ISBN 94-010-7860-2, 94-009-3025-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-009-3025-4_19.

Snelders:1994:RDS

[Sne94] H. A. M. Snelders. Romanticism and Dutch scientists. In Poggi and Bossi [PB94], chapter 11, pages 175–188. ISBN 0-7923-2336-X, 94-017-2921-2. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 152.

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-017-2921-
                                              5_11.

Sober:1991:OIU

[Sob91] Elliott Sober. Organisms, individuals, and units of selection. In Organism and the Origins of Self [Tau91i], chapter 13, pages 275–296. ISBN 0-7923-1185-X, 94-011-3406-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QR180-189.5. URL https:/

                                              /link.springer.com/chapter/10.
                                              1007/978-94-011-3406-4_13.

Sober:2010:ETR

[Sob10] Elliott Sober. Evolutionary theory and the reality of macro-probabilities. In The Place of Probability in Science: In Honor of Ellery Eells (1953–2006) [Fet10a], chapter 6, pages 133–161. ISBN 90-481-3614-8 (hardcover), 90-481-3615-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175 .P563 2010. URL https://link.

springer.com/chapter/10.1007/
978-90-481-3615-5_6.

Soderqvist:2007:PVS

[Söd07]  Thomas Söderqvist.  Plutarchian

versus Socratic scientific biography. In Positioning the history of science [RG07], chapter 27, pages 159–162. ISBN 1-4020-5419-X (hardcover), 1-4020-5420-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q125 .P77 2007. URL https://link.

springer.com/chapter/10.1007/1-
4020-5420-3_27.

Solomonoff:1963:C

[Sol63] Ray Solomonoff. Comments. In Wartofsky [War63], pages 42–45. ISBN 94-010-3265-3, 94-010-3263-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B678 1970.

URL https://link.springer.com/
chapter/10.1007/978-94-010-3263-
6_2.

Soll:1984:CKI

[Sol84] Ivan Soll. Comments on Kosok’s interpretation of Hegel’s logic. In Hegel and the Sciences [CW84a], chapter 21, pages 361–364. ISBN 94-009-6235-5, 94-009-6233-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-009-
6233-0_21.

Soler:2008:IEP

[Sol08a] Léna Soler. The incommensurability of experimental practices: An incommensurability of what? An incommensurability of a third type? In Rethinking Scientific Change and Theory Comparison: Stabilities, Ruptures, Incommensurabilities? [SSHH08], chapter 22, pages 299–339. ISBN 1-4020-6274-5 (hardcover), 1-4020-6279-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175 .R397

2008. URL https://link.springer.
com/chapter/10.1007/978-1-4020-
6279-7_22.

Soler:2008:I

[Sol08b] Léna Soler. Introduction. In Rethinking Scientific Change and Theory Comparison: Stabilities, Ruptures, Incommensurabilities? [SSHH08], chapter 1, pages 1–17. ISBN 1-4020-6274-5 (hardcover), 1-4020-6279-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175

.R397 2008. URL https://link.
springer.com/chapter/10.1007/
978-1-4020-6279-7_1.

Soler:2012:ISS

[Sol12a] Léna Soler. Introduction: The solidity of scientific achievements: Structure of the problem, difficulties, philosophical implications. In Characterizing the Robustness of Science: After the Practice Turn in Philosophy of Science [STNW12], chapter 1, pages 1–60. ISBN 94-007-2758-5, 94-007-2759-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175 .C4595 2012. URL https://link.

springer.com/chapter/10.1007/
978-94-007-2759-5_1.

Soler:2012:RRR

[Sol12b] Léna Soler. Robustness of results and robustness of derivations: The internal architecture of a solid experimental proof. In Characterizing the Robustness of Science: After the Practice Turn in Philosophy of Science [STNW12], chapter 10, pages 227–266. ISBN 94-007-2758-5, 94-007-2759-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175 .C4595 2012.

URL https://link.springer.com/
chapter/10.1007/978-94-007-2759-
5_10.

Sommers:1974:LEL

[Som74] Fred Sommers. The logical and the extra-logical. In Methodological and Historical Essays in the Natural and Social Sciences [CW74b], chapter 12, pages 235–252. ISBN 90-277-0378-7, 94-010-2128-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 14. URL https:

//link.springer.com/chapter/10.
1007/978-94-010-2128-9_12.

Sommers:1976:LPD

[Som76a] Fred Sommers. Leibniz’s program for the development of logic. In Essays in Memory of Imre Lakatos [CFW76], chapter 33, pages 589–615. ISBN 90-277-0655-7, 94-010-1451-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-010-1451-9_33.

Sommers:1976:PLS

[Som76b] Fred Sommers. On predication and logical syntax. In Language in Focus: Foundations, Methods and Systems: Essays in Memory of Yehoshua Bar-Hillel [Kas76b], chapter 3, pages 41–53. ISBN 90-277-0644-1, 90-277-0645-X, 94-010-1876-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN P101-120. URL https://link.

                                              springer.com/chapter/10.1007/
                                              978-94-010-1876-0_3.

Somenzi:1981:SDV

[Som81] Vittorio Somenzi. Scientific discovery from the viewpoint of evolutionary epistemology. In Grmek et al. [GCC81b], pages 167–177. ISBN 90-277-1123-2, 94-010-1284-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

                                              ???? URL https://link.springer.
                                              com/chapter/10.1007/978-94-010-
                                              1284-3_7.

Sorensen:2017:SMT

[Sør17]  Henrik Kragh Sørensen.  Shaping

mathematics as a tool: The search for a mathematical model for quasi-crystals. In Mathematics as a Tool: Tracing New Roles of Mathematics in the Sciences [LC17b], chapter 5, pages 69– 90. ISBN 3-319-54468-3 (hardcover), 3-319-54469-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN GE45.M37; GE45.M38; QA8.4.

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-3-319-54469-
                                              4_5.

Sosa:1974:WLC

[Sos74] Ernest Sosa. What is a logical constant? In Methodological and Historical Essays in the Natural and Social Sciences [CW74b], chapter 13, pages 253–256. ISBN 90-277-0378-7, 94-010-2128-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 14. URL https://link.

                                              springer.com/chapter/10.1007/
                                              978-94-010-2128-9_13.

Sosa:1995:BB

[Sos95] Ernest Sosa. Back to basics. In The Concept of Knowledge: The Ankara Seminar [KC95], chapter 3, pages 21– 29. ISBN 90-481-4495-7, 94-017-3263-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-017-3263-
5_3.

Sotirov:2003:LLS

[Sot03] Vladimir Sotirov. Leibniz’s logical systems: A reconstruction. In Bulgarian Studies in the Philosophy of Science [Gin03a], chapter 7, pages 85–93. ISBN 1-4020-1496-1, 94-017-0961-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-017-0961-3_7.

Soyhun:2005:IDB

[Soy05] A. Karanfil Soyhun. Implications of the distinction between semantics and pragmatics for philosophy of science. In Turkish Studies in the History and Philosophy of Science [IG05c], chapter 12, pages 163–176. ISBN 1-4020-3332-X (hardcover), 1-4020-3333-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 244.

URL https://link.springer.com/
chapter/10.1007/1-4020-3333-8_12.

Skyrms:2010:LN

[SP10] Brian Skyrms and Robin Pemantle.

Learning to network. In The Place of Probability in Science: In Honor of Ellery Eells (1953–2006) [Fet10a], chapter 12, pages 277–287. ISBN 90-481-3614-8 (hardcover), 90-481-3615-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175 .P563

2010. URL https://link.springer.
com/chapter/10.1007/978-90-481-
3615-5_12.

Spassov:2003:ITR

[Spa03] Spas Spassov. The information technology revolution: A new techno–economic paradigm. In Bulgarian Studies in the Philosophy of Science [Gin03a], chapter 3, pages 31–41. ISBN 1-4020-1496-1, 94-017-0961-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic).

LCCN ????  URL https://link.
springer.com/chapter/10.1007/
978-94-017-0961-3_3.

Silva:2018:HSM

[SPB18] Édison Renato Silva, Domício Proença, Jr., and Roberto Bartholo. Herbert Simon meets Billy Vaughn Koen and Joan van Aken: From sciences of the artificial to engineering heuristics and design propositions. In Mitcham et al. [MLNZ18], chapter 13, pages 149–155. ISBN 3-319-62448-2, 3-319-62450-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN TA15

.P45 2018.  URL https://link.
springer.com/chapter/10.1007/
978-3-319-62450-1_13.

Spiekermann:1999:NCT

[Spi99] Klaus Spiekermann. Nietzsche and critical theory. In Nietzsche, Theories of Knowledge, and Critical Theory: Nietzsche and the Sciences I [Bab99b], chapter 16, pages 225–242. ISBN 90-481-5233-X, 94-017-2430-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-017-2430-2_16.

Spohn:2009:CCC

[Spo09] Wolfgang Spohn. Causation, Coherence, and Concepts: A Collection of Essays, volume 256 of Boston Studies in the Philosophy of Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2009. ISBN 1-4020-5473-4 (hardcover), 1-4020-5474-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). xvi + 386 pp. LCCN BD541; BD591 .S66 2009; Q175.3.

URL https://link.springer.com/
book/10.1007/978-1-4020-5474-7.

Spohn:2010:CN

[Spo10] Wolfgang Spohn. Chance and necessity.

In The Place of Probability in Science: In Honor of Ellery Eells (1953–2006) [Fet10a], chapter 5, pages 101–131. ISBN 90-481-3614-8 (hardcover), 90-481-3615-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175 .P563 2010. URL https://link.

springer.com/chapter/10.1007/
978-90-481-3615-5_5.

Sprung:1991:BBS

[Spr91a] Helga Sprung. Bourgeois Berlin salons:

Meeting places for culture and the sciences. In Woodward and Cohen [WC91], pages 401–414. ISBN 94-010-5395-2, 94-011-3164-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-011-3164-3_37.

Sprung:1991:BPT

[Spr91b] Lothar Sprung. The Berlin psychological tradition: Between experiment and quasi-experimental design, 1850–1990. In Woodward and Cohen [WC91], pages 107–116. ISBN 94-010-5395-2, 94-011-3164-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic).

LCCN ????  URL https://link.
springer.com/chapter/10.1007/
978-94-011-3164-3_10.

Squires:1996:LBM

[Squ96] Euan J. Squires. Local Bohmian mechanics. In Bohmian Mechanics and Quantum Theory: an Appraisal [CFG96], chapter 9, pages 131–140. ISBN 90-481-4698-4, 94-015-8715-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC173.96-174.52. URL https://link.

                                              springer.com/chapter/10.1007/
                                              978-94-015-8715-0_9.

Sanchez-Ron:1987:RSR

[SR87] José M. Sánchez-Ron. The reception of special relativity in Great Britain. In Glick [Gli87a], chapter 2, pages 27–58. ISBN 90-277-2498-9, 94-009-3875-6. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 103; QC173.5. US$79.00 (US). URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-009-3875-5_2.

Sanchez-Ron:2007:SPH

[SR07]  José M. Sánchez-Ron.  Schweber,

physicist, historian and moral example. In Positioning the history of science [RG07], chapter 24, pages 139–142. ISBN 1-4020-5419-X (hardcover), 1-4020-5420-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q125 .P77 2007.

                                              URL https://link.springer.com/
                                              chapter/10.1007/1-4020-5420-3_24.

Sanchez-Ron:2015:NIL

[SR15]  José M. Sánchez-Ron.  The non-

introduction of low-temperature physics in Spain: Julio Palacios and Heike Kamerlingh Onnes. In Arabatzis et al. [ARS15], chapter 10, pages 131– 157. ISBN 3-319-14552-5 (paperback), 3-319-14553-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic).

LCCN Q125 .R45 2015. URL https:

//link.springer.com/chapter/10.
1007/978-3-319-14553-2_10.

Scholl:2016:TMI

[SR16] Raphael Scholl and Tim Räz. Towards a methodology for integrated history and philosophy of science. In Sauer and Scholl [SS16b], chapter 5, pages 69– 91. ISBN 3-319-30227-2, 3-319-30229-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q126.8 .P45

2016. URL https://link.springer.
com/chapter/10.1007/978-3-319-
30229-4_5.

Sredniawa:1987:RTR

[Sre87] Bronisław Sredniawa. The reception of the theory of relativity in Poland. In Glick [Gli87a], chapter 9, pages 327–350. ISBN 90-277-2498-9, 94-009-3875-6. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 103; QC173.5. US$79.00

(US). URL https://link.springer.
com/chapter/10.1007/978-94-009-
3875-5_9.

Shokichi:1992:IPD

[SRI^{+}92] Iyanaga Shokichi, Roshdi Rashed, Ekmeleddin Ihsanoğlu, V. V. Krishna, Nakayama Shigeru, and Anne Marie Moulin. Integration problems: Discussion. In Science and Empires: Historical Studies about Scientific Development and European Expansion [PJM92], chapter 9, pages 73–78. ISBN 0-7923-1518-9 (hardcover), 94-010-5145-3, 94-011-2594-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 136.

URL https://link.springer.com/
chapter/10.1007/978-94-011-2594-
9_9.

Siegel:1974:AVR

[SS74] Armand Siegel and Mildred Siegel. The accelerator and the virgin: The rise & fall of two cults. In For Dirk Struik: Scientific, Historical and Political Essays in Honor of Dirk J. Struik [CSW74], chapter 33, pages 403–415. ISBN 90-277-0379-5, 94-010-2115-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-010-2115-
                                              9_33.

Siegmund-Schultze:1991:MIF

[SS91] Reinhard Siegmund-Schultze. Mathematics and ideology in Fascist Germany. In Woodward and Cohen [WC91], pages 89–95. ISBN 94-010-5395-2, 94-011-3164-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

                                              ???? URL https://link.springer.
                                              com/chapter/10.1007/978-94-011-
                                              3164-3_8.

Shimony:2003:QNL

[SS03] Abner Shimony and Howard Stein. On quantum non-locality, special relativity, and counterfactual reasoning. In Renn et al. [RDS03], chapter 25, pages 499–521. ISBN 1-4020-1284-5, 1-4020-1285-3 (paperback). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 234. URL https:/

                                              /link.springer.com/chapter/10.
                                              1007/978-94-010-0111-3_25.

Sauer:2016:I

[SS16a] Tilman Sauer and Raphael Scholl. Introduction. In The Philosophy of Historical Case Studies [SS16b], chapter 1, pages 1–8. ISBN 3-319-30227-2, 3-319-30229-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic).

LCCN Q126.8 .P45 2016. URL https:

//link.springer.com/chapter/10.
1007/978-3-319-30229-4_1.

Sauer:2016:PHC

[SS16b] Tilman Sauer and Raphael Scholl, editors. The Philosophy of Historical Case Studies, volume 319 of Boston Studies in the Philosophy and History of Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2016. ISBN 3-319-30227-2, 3-319-30229-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). viii + 296 pp. LCCN Q126.8 .P45 2016.

URL https://link.springer.com/
book/10.1007/978-3-319-30229-4.

Soler:2008:RSC

[SSHH08] Léna Soler, Howard Sankey, and Paul Hoyningen-Huene. Rethinking Scientific Change and Theory Comparison: Stabilities, Ruptures, Incommensurabilities?, volume 255 of Boston Studies in the Philosophy of Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2008. ISBN 1-4020-6274-5 (hardcover), 1-4020-6279-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). x + 384 pp. LCCN Q175 .R397 2008. URL

https://link.springer.com/book/
10.1007/978-1-4020-6279-7.

Skolimowski:1969:KLR

[ST69] Henryk Skolimowski and Stephen Toulmin. Knowledge, language, and rationality. In Cohen and Wartofsky [CW69b], pages 174–207. ISBN 94-010-3380-3, 94-010-3378-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-010-3378-7_5.

Stachel:1969:CCR

[Sta69] John Stachel. Comments on ‘Causality Requirements and the Theory of Relativity’. In Cohen and Wartofsky [CW69a], pages 179–198. ISBN 94-010-3383-8, 94-010-3263-7, 94-010-3381-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-010-3381-

7_3. See [Hav69].

Stachel:1974:COC

[Sta74a] John Stachel. Comments on ‘ontic commitments of quantum mechanics’. In Logical and Epistemological Studies in Contemporary Physics [CW74a], chapter 10, pages 309–317. ISBN 90-277-0391-4 (hardcover), 90-277-0377-9 (paperback), 94-010-2656-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol.

                                              13; QC5.56. URL https://link.
                                              springer.com/chapter/10.1007/
                                              978-94-010-2656-7_10.

Stachel:1974:CFR

[Sta74b] John Stachel. Comments on ‘the formal representation of physical quantities’. In Logical and Epistemological Studies in Contemporary Physics [CW74a], chapter 7, pages 214–223. ISBN 90-277-0391-4 (hardcover), 90-277-0377-9 (paperback), 94-010-2656-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol.

                                              13; QC5.56. URL https://link.
                                              springer.com/chapter/10.1007/
                                              978-94-010-2656-7_7.

Stachel:1974:ISC

[Sta74c] John Stachel. Introduction to the symposium on cosmology. In Seeger and Cohen [SC74b], pages 179–180. ISBN 90-277-0390-6, 90-277-0376-0 (paperback), 94-010-2126-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 11.

URL https://link.springer.com/
chapter/10.1007/978-94-010-2126-
5_9.

Stachel:1974:NCS

[Sta74d] John Stachel. A note on the concept of scientific practice. In For Dirk Struik: Scientific, Historical and Political Essays in Honor of Dirk J. Struik [CSW74], chapter 34, pages 417–433. ISBN 90-277-0379-5, 94-010-2115-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-010-2115-
9_34.

Stachel:1974:RFG

[Sta74e] John Stachel. The rise and fall of geometrodynamics. In Schaffner and Cohen [SC74a], pages 31–54. ISBN 90-277-0408-2, 90-277-0409-0, 94-010-2140-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175 .B7312. URL https://link.

springer.com/chapter/10.1007/
978-94-010-2140-1_3.

Stachel:1976:LQL

[Sta76a] John Stachel. The’ logic of’ quantum logic. In Cohen et al. [CHMV76], pages 515–526. ISBN 90-277-0647-6, 90-277-0648-4 (paperback), 94-010-1449-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-010-1449-
6_29.

Stalnaker:1976:IC

[Sta76b] Robert C. Stalnaker. Indicative conditionals. In Language in Focus: Foundations, Methods and Systems: Essays in Memory of Yehoshua Bar-Hillel [Kas76b], chapter 11, pages 179–196. ISBN 90-277-0644-1, 90-277-0645-X, 94-010-1876-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN P101-120. URL https://link.

springer.com/chapter/10.1007/
978-94-010-1876-0_11.

Staniszkis:1979:ASP

[Sta79] Jadwiga Staniszkis. Adaptational superstructure — the problem of negative self-regulation. In Polish Essays in the Methodology of the Social Sciences [Wia79b], chapter 12, pages 225–248. ISBN 90-277-0956-4, 94-009-9353-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174

.P655 1979. URL https://link.
springer.com/chapter/10.1007/
978-94-009-9353-2_12.

Stasiak:1982:PI

[Sta82] Makary Krzysztof Stasiak. Practical idealization. In Polish Essays in the Philosophy of the Natural Sciences [Kra82c], chapter 14, pages 153–165. ISBN 94-009-7705-0 (e-book), 90-277-1287-5. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-009-7705-
1_14.

Stachel:1983:CSL

[Sta83a] John Stachel. Comments on ‘Some Logical Problems Suggested by Empirical Theories’ by Professor Dalla Chiara. In Language, Logic and Method [CW83b], chapter 3, pages 91–102. ISBN 94-009-7704-2, 94-009-7702-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-009-7702-
0_3.

Stachel:1983:SRM

[Sta83b] John Stachel. Special relativity from measuring rods. In Cohen and Laudan [CL83], chapter 13, pages 255–272. ISBN 94-009-7055-2, 94-009-7057-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-009-7055-
7_13.

Stachel:1986:EE

[Sta86a] John Stachel. Eddington and Einstein. In Ullmann-Margalit [UM86c], chapter 18, pages 225–250. ISBN 90-277-2160-2, 90-277-2161-0 (paperback), 94-009-4566-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 95 Q175.

URL https://link.springer.com/
chapter/10.1007/978-94-009-4566-
1_18.

Stachel:1986:NCS

[Sta86b] John Stachel. A note on the concept of scientific practice. In Cohen and Wartofsky [CW85], chapter 11, pages 160–176. ISBN 90-277-1971-3, 94-009-5345-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-009-
5345-1_11.

Stadler:1991:ASB

[Sta91a] Friedrich Stadler. Aspects of the social background and position of the Vienna Circle at the University of Vienna. In Uebel [Ueb91b], chapter 5, pages 51– 77. ISBN 0-7923-1276-7 (hardcover), 94-011-3182-1. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 133. URL https://link.

springer.com/chapter/10.1007/
978-94-011-3182-7_5.

Stadler:1991:ONM

[Sta91b] Friedrich Stadler. Otto Neurath — Moritz Schlick: On the philosophical and political antagonisms in the Vienna Circle. In Uebel [Ueb91b], chapter 12, pages 159–168. ISBN 0-7923-1276-7 (hardcover), 94-011-3182-1. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 133. URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-011-3182-7_12.

Stadler:1991:ONE

[Sta91c]  Friedrich Stadler.  Otto Neurath:

Encyclopedist, adult educationalist and school reformer. In Uebel [Ueb91b], chapter 19, pages 255–264. ISBN 0-7923-1276-7 (hardcover), 94-011-3182-1. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 133. URL https://link.

                                              springer.com/chapter/10.1007/
                                              978-94-011-3182-7_19.

Stadler:1992:VEM

[Sta92a] Friedrich Stadler. The “Verein Ernst Mach” — what was it really? In Ernst Mach — A Deeper Look: Documents and New Perspectives [Bla92a], chapter 18, pages 363–377. ISBN 94-010-5230-1, 94-011-2771-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN C5.53; QC6.9.

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-011-2771-
                                              4_18.

Staudenmaier:1992:STW

[Sta92b] John M. Staudenmaier. Science and technology: Who gets a say? In Technological Development and Science in the Industrial Age: New Perspectives on the Science–Technology Relationship [KB92b], chapter 9, pages 205–230. ISBN 90-481-4186-9, 94-015-8010-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-015-8010-
6_9.

Stachel:1994:CCS

[Sta94a] John Stachel. Changes in the concepts of space and time brought about by relativity. In Artifacts, Representations and Social Practice: Essays for Marx Wartofsky [GC94], chapter 9, pages 141–162. ISBN 94-010-4390-6, 94-011-0902-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN B73 .A77 1994. URL https://link.

springer.com/chapter/10.1007/
978-94-011-0902-4_9.

Stachel:1994:SDH

[Sta94b] John Stachel. Scientific discoveries as historical artifacts. In Trends in the Historiography of Science [GCN94], chapter 10, pages 139–148. ISBN 90-481-4264-4, 94-017-3596-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-017-3596-4_10.

Stapp:1994:QTP

[Sta94c] Henry P. Stapp. Quantum theory and the place of mind in nature. In Faye and Folse [FF94], chapter 16, pages 345–352. ISBN 0-7923-2378-5 (hardcover), 94-015-8106-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 153 1994; QC16 1994. URL https:

//link.springer.com/chapter/10.
1007/978-94-015-8106-6_16.

Stachel:1997:FPQ

[Sta97] John Stachel. Feynman paths and quantum entanglement: Is there any more to the mystery. In Cohen et al. [CHS97b], pages 245–256. ISBN 90-481-4809-X, 94-017-2732-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC173.96-174.52.

URL https://link.springer.com/
chapter/10.1007/978-94-017-2732-
7_19.

Stachel:2007:FTA

[Sta07a] John Stachel. The first two acts. In The Genesis of General Relativity: Volume 1. Einstein’s Zurich Notebook: Introduction and Source [JNR^{+}07b], pages 81–111. ISBN 1-4020-3999-9 (hardcover), 1-4020-4000-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 no. 250 v. 1; QC173.6 .G469 2007. URL https:

//link.springer.com/chapter/10.
1007/978-1-4020-4000-9_3.

Stachel:2007:PD

[Sta07b] John Stachel. Problems not disciplines. In Positioning the history of science [RG07], chapter 28, pages 163–167. ISBN 1-4020-5419-X (hardcover), 1-4020-5420-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q125 .P77 2007.

URL https://link.springer.com/
chapter/10.1007/1-4020-5420-3_28.

Stachel:2007:SNG

[Sta07c] John Stachel. The story of Newstein or: Is gravity just another pretty force? In Renn et al. [RSS^{+}07b], pages 1041–1078. ISBN 1-4020-3999-9 (hardcover), 1-4020-4000-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 no. 250 v. 4; QC173.6 .G469 2007. URL https:

//link.springer.com/chapter/10.
1007/978-1-4020-4000-9_46.

Stachel:2009:GC

[Sta09] John J. Stachel. Going Critical: Volume One: The Challenge of Practice, volume 201 of Boston Studies in the Philosophy of Science. Springer-Ver-lag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2009. ISBN 1-4020-1308-6 (hardcover). 407 (est.) pp. LCCN ????

Stachel:2015:EH

[Sta15] John Stachel. Einstein and Hilbert. In Arabatzis et al. [ARS15], chapter 4, pages 33–39. ISBN 3-319-14552-5 (paperback), 3-319-14553-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q125 .R45 2015.

URL https://link.springer.com/
chapter/10.1007/978-3-319-14553-
2_4.

Stanley:2017:RSV

[Sta17] Matthew Stanley. Religious and scientific virtues: Maxwell, Eddington, and persistence. In van Dongen and Paul [vDP17a], chapter 4, pages 49–61. ISBN 3-319-48892-9 (hardcover), 3-319-48893-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 v. 321; Q175.32.V57 E65

2017. URL https://link.springer.
com/chapter/10.1007/978-3-319-
48893-6_4.

Stegmuller:1979:CAD

[Ste79] Wolfgang Stegmüller. A combined approach to the dynamics of theories. In The Structure and Development of Science [RA79], chapter 6, pages 151–186. ISBN 90-277-0995-5, 94-009-9459-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .S778 1979. URL https://link.

springer.com/chapter/10.1007/
978-94-009-9459-1_6.

Stent:1988:CLS

[Ste88] Gunther Stent. Cognitive limits of science. In Science in Reflection: The Israel Colloquium: Studies in History, Philosophy, and Sociology of Science Volume 3 [UM88], chapter 3, pages 23– 36. ISBN 90-277-2713-9, 94-009-2957-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-009-2957-
                                              9_3.

Stepan:1992:PAE

[Ste92] Nancy Leys Stepan. The Pan American experiment in eugenics. In Science and Empires: Historical Studies about Scientific Development and European Expansion [PJM92], chapter 21, pages 201–208. ISBN 0-7923-1518-9 (hardcover), 94-010-5145-3, 94-011-2594-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 136. URL https://link.

                                              springer.com/chapter/10.1007/
                                              978-94-011-2594-9_21.

Stepin:1994:SEF

[Ste94] Vyacheslav S. Stepin. Social environment, foundations of science, and the possible histories of science. In Trends in the Historiography of Science [GCN94], chapter 9, pages 129–137. ISBN 90-481-4264-4, 94-017-3596-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-017-3596-
                                              4_9.

Stein:1997:MEI

[Ste97] Howard Stein. Maximal extension of an impossibility theorem concerning quantum measurement. In Cohen et al. [CHS97b], pages 231–243. ISBN 90-481-4809-X, 94-017-2732-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC173.96-174.52.

URL https://link.springer.com/
chapter/10.1007/978-94-017-2732-
7_18.

Steinhart:1999:WPP

[Ste99] Eric Steinhart. The will to power and parallel distributed processing. In Nietzsche, Epistemology, and Philosophy of Science: Nietzsche and the Sciences II [Bab99a], chapter 24, pages 313–322. ISBN 90-481-5234-8, 94-017-2428-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-017-2428-
9_24.

Stern:2000:MLS

[Ste00] Josef Stern. Maimonides on language and the science of language. In Maimonides and the Sciences [CL00], chapter 10, pages 173–226. ISBN 90-481-5348-4, 94-017-2128-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-017-2128-8_10.

Stefanov:2003:KCS

[Ste03] Anguel S. Stefanov. On Kant’s conception of space and time. In Bulgarian Studies in the Philosophy of Science [Gin03a], chapter 12, pages 169–185. ISBN 1-4020-1496-1, 94-017-0961-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-017-0961-
3_12.

Stenwall:2007:AK

[Ste07] Robin Stenwall. Aspect kinds. In Rethinking Explanation [PY07], chapter 14, pages 193–203. ISBN 1-4020-5580-3 (paperback), 1-4020-5581-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN BD237 .R48 2007; Q175.32.E97 R48 2007.

URL https://link.springer.com/
chapter/10.1007/978-1-4020-5581-
2_14.

Steinle:2009:HEM

[Ste09a] Friedrich Steinle. How experiments make concepts fail: Faraday and magnetic curves. In Going Amiss In Experimental Research [HSS09a], chapter 8, pages 119–135. ISBN 1-4020-8892-2 (hardcover), 1-4020-8893-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q172.5.E77

G647 2009.  URL https://link.
springer.com/chapter/10.1007/
978-1-4020-8893-3_8.

Stelzer:2009:PCJ

[Ste09b] Harald Stelzer. Popper and communitarianism: Justification and criticism of moral standards. In Rethinking Popper [PC09], chapter 21, pages 273–285. ISBN 1-4020-9337-3, 1-4020-9338-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN B1649.P64 R48 2009; Q174 .B67 2009.

URL https://link.springer.com/
chapter/10.1007/978-1-4020-9338-
8_21.

Sterelny:2009:NPN

[Ste09c] Kim Sterelny. Novelty, plasticity and niche construction: The influence of phenotypic variation on evolution. In Mapping the Future of Biology: Evolving Concepts and Theories [BMP09b], chapter 7, pages 93–110. ISBN 1-4020-9635-6 (hardcover), 1-4020-9636-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QH331 .M295

2009. URL https://link.springer.
com/chapter/10.1007/978-1-4020-
9636-5_7.

Stegenga:2012:RCD

[Ste12] Jacob Stegenga. Rerum concordia discors: Robustness and discordant multimodal evidence. In Characterizing the Robustness of Science: After the Practice Turn in Philosophy of Science [STNW12], chapter 9, pages 207–226. ISBN 94-007-2758-5, 94-007-2759-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175 .C4595 2012. URL https://link.

springer.com/chapter/10.1007/
978-94-007-2759-5_9.

Stearns:2015:CPP

[Ste15a] Stephen C. Stearns. The concept of phenotypic plasticity and the evolution of phenotypic plasticity in life history traits. In Conceptual Change in Biology: Scientific and Philosophical Perspectives on Evolution and Development [Lov15b], chapter 6, pages 131–146. ISBN 94-017-9411-1, 94-017-9412-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QH331 C66 2015; QH491; QH492 .C66

2014. URL https://link.springer.
com/chapter/10.1007/978-94-017-
9412-1_6.

Stefanov:2015:REA

[Şte15b] Gheorghe Ştefanov. On rule embedding artifacts. In Romanian Studies in Philosophy of Science [PST15], chapter 5, pages 71–85. ISBN 3-319-16654-9 (hardcover), 3-319-16655-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-3-319-16655-1_5.

Steigerwald:2015:ECG

[Ste15c] Joan Steigerwald. The ethos of critique in German idealism. In Objectivity in Science: New Perspectives from Science and Technology Studies [PRT15], chapter 5, pages 79–99. ISBN 3-319-14348-4 (hardcover), 3-319-14349-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175 O25 2015.

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-3-319-14349-
                                              1_5.

Sterrett:2017:TIH

[Ste17] Susan G. Sterrett. Turing on the integration of human and machine intelligence. In Floyd and Bokulich [FB17], chapter 14, pages 323–338. ISBN 3-319-53278-2, 3-319-53280-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QA29.T8 P45

                                              2017. URL https://link.springer.
                                              com/chapter/10.1007/978-3-319-
                                              53280-6_14.

Stern:2018:PNA

[Ste18] Johannes Stern. Penrose’s new argument and paradox. In Piazza and Pulcini [PP18], pages 17–38. ISBN 3-319-93341-8, 3-319-93342-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QA8.4.

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-3-319-93342-
                                              9_2.

Stites:1974:LVS

[Sti74] Raymond S. Stites. Leonardo Da Vinci and the sublimatory process. In Seeger and Cohen [SC74b], pages 17–40. ISBN 90-277-0390-6, 90-277-0376-0 (paperback), 94-010-2126-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 11.

URL https://link.springer.com/
chapter/10.1007/978-94-010-2126-
5_2.

Stichweh:1994:UTR

[Sti94] Rudolf Stichweh. The unity of teaching and research. In Poggi and Bossi [PB94], chapter 12, pages 189–202. ISBN 0-7923-2336-X, 94-017-2921-2. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 152.

URL https://link.springer.com/
chapter/10.1007/978-94-017-2921-
5_12.

Soler:2012:CRS

[STNW12] Léna Soler, Emiliano Trizio, Thomas Nickles, and William Wimsatt. Characterizing the Robustness of Science: After the Practice Turn in Philosophy of Science, volume 292 of Boston Studies in the Philosophy of Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2012. ISBN 94-007-2758-5, 94-007-2759-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). x + 372 pp. LCCN Q175 .C4595 2012. URL

https://link.springer.com/book/
10.1007/978-94-007-2759-5.

Stock:1975:EPW

[Sto75] Brian Stock. Experience, praxis, work, and planning in Bernard of Clairvaux: Observations on the Sermones in Cantica. In Murdoch and Sylla [MS75b], pages 219–268. ISBN 90-277-0587-9, 94-010-1781-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN B721 .I584 1975. URL https:

//link.springer.com/chapter/10.
1007/978-94-010-1781-7_8.

Stojanovic:1979:THM

[Sto79] Svetozar Stojanovic. A tension in historical materialism. In Praxis:

Yugoslav Essays in the Philosophy and Methodology of the Social Sciences [MP79], chapter 3, pages 63–79. ISBN 90-277-0968-8, 94-009-9355-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-009-9355-6_3.

Stone:1992:CEM

[Sto91] Anthony J. Stone. Classical electrostatics in molecular interactions. In The Invention of Physical Science: Intersections of Mathematics, Theology and Natural Philosophy Since the Seventeenth Century: Essays in Honor of Erwin N. Hiebert [NRS92], chapter 4, pages 103–131. ISBN 94-010-5097-X, 94-011-2488-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-3-642-58183-0_4.

Stoeckler:1997:SEL

[Sto97] Manfred Stoeckler. Symmetries and explanations: The lessons of elementary particle physics. In Ginev and Cohen [GC97], chapter 20, pages 343–355. ISBN 0-7923-4444-8 (hardcover), 94-011-5788-X. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 192. URL https://link.

springer.com/chapter/10.1007/
978-94-011-5788-9_20.

Stoltzner:1999:WED

[Stö99] Michael Stöltzner. To what extent does formal teleology still make sense? In Fehér et al. [FKR99], pages 227–246. ISBN 90-481-5257-7, 94-015-9293-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-015-9293-
2_18.

Stoutland:2000:ISQ

[Sto00] Frederick Stoutland. Individual and social in Quine’s philosophy of language. In Knowledge, Language and Logic: Questions for Quine [OK00], chapter 14, pages 181–194. ISBN 1-4020-0253-X, 94-011-3933-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-011-3933-5_14.

Stoltzner:2001:ON

[Stö01] Michael Stöltzner. Otto Neurath 1913– 1915. In Blackmore et al. [BIT01c], pages 105–122. ISBN 90-481-5752-8, 94-015-9690-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-015-
9690-9_5.

Stoltzner:2003:DTE

[Stö03] Michael Stöltzner. The dynamics of thought experiments: A comment on David Atkinson. In Descartes’s Mathematical Thought [Sas03d], chapter 17, pages 243–258. ISBN 90-481-6487-7, 94-017-1225-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic).

LCCN ????  URL https://link.
springer.com/chapter/10.1007/0-
306-48123-5_17.

Stoltzner:2013:DRA

[Stö11] Michael Stöltzner. Did Reichenbach anticipate quantum mechanical indeterminism? In The Berlin Group and the Philosophy of Logical Empiricism [MP13], chapter 6, pages 123–150. ISBN 94-007-5484-1 (hardcover), 94-007-5485-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175.32.A24 .M384 2011; Q175 .B47

                                              2013. URL https://link.springer.
                                              com/chapter/10.1007/978-94-007-
                                              5485-0_6.

Struik:1974:SRC

[Str74] Ruth Rebekka Struik. Some remarks on the concept of limit. In For Dirk Struik: Scientific, Historical and Political Essays in Honor of Dirk J. Struik [CSW74], chapter 10, pages 121–129. ISBN 90-277-0379-5, 94-010-2115-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-010-2115-
                                              9_10.

Strong:1976:EDJ

[Str76] John V. Strong. The erkenntnistheoretikers dilemma: J. B. Stallo’s attack on atomism in his concepts and theories of modern physics (1881). In Cohen et al. [CHMV76], pages 105–123. ISBN 90-277-0647-6, 90-277-0648-4 (paperback), 94-010-1449-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-010-1449-6_5.

Strawinski:1982:FDC

[Str82] Witold Strawiński. A formal definition of the concept of simplicity. In Polish Essays in the Philosophy of the Natural Sciences [Kra82c], chapter 16, pages 187–195. ISBN 94-009-7705-0 (e-book), 90-277-1287-5. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-009-7705-
                                              1_16.

Stroker:1983:PTP

[Str83] Elisabeth Ströker. Philosophy of technology: Problems of a philosophical discipline. In Philosophy and Technology [DR83], chapter 23, pages 323–336. ISBN 94-009-7126-5, 94-009-7124-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-009-7124-
0_23.

Stroker:1984:DPC

[Str84]  Elisabeth Ströker.  Does Popper’s

conventionalism contradict his critical rationalism? Objections against Popper in German philosophy and some metacritical remarks. In Methodology, Metaphysics and the History of Science: In Memory of Benjamin Nelson [CW84b], chapter 11, pages 263–282. ISBN 94-009-6333-5, 94-009-6331-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-009-6331-
3_11.

Strombach:1986:IEO

[Str86]  Werner Strombach.  “Information”

in epistemological and ontological perspective. In Philosophy and Technology II: Information Technology and Computers in Theory and Practice [MH86], chapter 6, pages 75–81. ISBN 94-010-8510-2, 94-009-4512-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-009-4512-8_6.

Strickland:1994:RTR

[Str94] Stuart Walker Strickland. Reopening the texts of romantic science: The language of experience in J. W. Ritter’s Beweis. In Trends in the Historiography of Science [GCN94], chapter 28, pages 385–396. ISBN 90-481-4264-4, 94-017-3596-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

                                              ???? URL https://link.springer.
                                              com/chapter/10.1007/978-94-017-
                                              3596-4_28.

Struik:1995:SSP

[Str95] Dirk J. Struik. Some sociological problems in the history of mathematics. In Gavroglu et al. [GSW95c], chapter 21, pages 377–383. ISBN 0-7923-2991-0 (set), 0-7923-2988-0, 94-017-2658-2. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 163.

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-017-2658-
                                              0_21.

Strong:1999:PHN

[Str99] Tracy B. Strong. A postscript on Habermas, Nietzsche, and politics. In Nietzsche, Theories of Knowledge, and Critical Theory: Nietzsche and the Sciences I [Bab99b], chapter 24, pages 307–313. ISBN 90-481-5233-X, 94-017-2430-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

                                              ???? URL https://link.springer.
                                              com/chapter/10.1007/978-94-017-
                                              2430-2_24.

Strathmann:2015:DFR

[Str15] Richard R. Strathmann. Do functional requirements for embryos and larvae have a place in evo-devo? In Conceptual Change in Biology: Scientific and Philosophical Perspectives on Evolution and Development [Lov15b], chapter 3, pages 77–95. ISBN 94-017-9411-1, 94-017-9412-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QH331 C66 2015; QH491; QH492 .C66

                                              2014. URL https://link.springer.
                                              com/chapter/10.1007/978-94-017-
                                              9412-1_3.

Strollo:2018:MSD

[Str18] Andrea Strollo. Making sense of deflationism from a formal perspective: Conservativity and relative interpretability. In Piazza and Pulcini [PP18], pages 89–105. ISBN 3-319-93341-8, 3-319-93342-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QA8.4. URL https:

//link.springer.com/chapter/10.
1007/978-3-319-93342-9_6.

Stuewer:1976:CRP

[Stu76] Roger H. Stuewer. On Compton’s research program. In Essays in Memory of Imre Lakatos [CFW76], chapter 34, pages 617–633. ISBN 90-277-0655-7, 94-010-1451-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-010-
1451-9_34.

Stuewer:1986:GTA

[Stu86] Roger H. Stuewer. Gamow’s theory of alpha-decay. In Ullmann-Margalit [UM86a], chapter 14, pages 147–186. ISBN 90-277-2159-9 (paperback), 90-277-2158-0 (hardcover), 94-009-5496-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 94 Q175. URL https:/

/link.springer.com/chapter/10.
1007/978-94-009-5496-0_14.  See

comment [Lip86].

Stuewer:1995:SSC

[Stu95] Roger H. Stuewer. The Seventh Solvay Conference: Nuclear physics at the crossroads. In No Truth Except in the Details: Essays in Honor of Martin J. Klein [KS95], chapter 15, pages 333–362. ISBN 94-010-4097-4, 94-011-0217-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

                                              ???? URL https://link.springer.
                                              com/chapter/10.1007/978-94-011-
                                              0217-9_15.

Stubenberg:1997:AVA

[Stu97] Leopold Stubenberg. Austria vs. Australia: Two versions of the identity theory. In Austrian Philosophy Past and Present: Essays in Honor of Rudolf Haller [LM97], chapter 9, pages 125–146. ISBN 94-010-6412-1, 94-011-5720-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

                                              ???? URL https://link.springer.
                                              com/chapter/10.1007/978-94-011-
                                              5720-9_9.

Stuewer:2007:PH

[Stu07] Roger H. Stuewer. Physicist-historians. In Positioning the history of science [RG07], chapter 29, pages 169–172. ISBN 1-4020-5419-X (hardcover), 1-4020-5420-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q125 .P77 2007. URL https://link.

                                              springer.com/chapter/10.1007/1-
                                              4020-5420-3_29.

Suarez:2008:HNN

[Suá08] Mauricio Suárez. Harré needs no realism. In Rethinking Scientific Change and Theory Comparison: Stabilities, Ruptures, Incommensurabilities? [SSHH08], chapter 17, pages 239–243. ISBN 1-4020-6274-5 (hardcover), 1-4020-6279-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175 .R397 2008. URL https://link.

                                              springer.com/chapter/10.1007/
                                              978-1-4020-6279-7_17.

Suchting:1974:RL

[Suc74] W. A. Suchting. Regularity and law. In Methodological and Historical Essays in the Natural and Social Sciences [CW74b], chapter 5, pages 73–90. ISBN 90-277-0378-7, 94-010-2128-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 14.

URL https://link.springer.com/
chapter/10.1007/978-94-010-2128-
9_5.

Such:1982:TDF

[Suc82] Jan Such. Are there definitively falsifying procedures in science? In Polish Essays in the Philosophy of the Natural Sciences [Kra82c], chapter 10, pages 113–126. ISBN 94-009-7705-0 (e-book), 90-277-1287-5. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-009-7705-
1_10.

Sun:2018:DMS

[Sun18] Lie Sun. Decision making in the 120MN Shanghai hydraulic forging press project: Walking a tightrope between politics and technology. In Mitcham et al. [MLNZ18], chapter 23, pages 271–280. ISBN 3-319-62448-2, 3-319-62450-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN TA15 .P45 2018. URL https:

//link.springer.com/chapter/10.
1007/978-3-319-62450-1_23.

Suppes:1974:EIR

[Sup74] Patrick Suppes. The essential but implicit role of modal concepts in science. In Schaffner and Cohen [SC74a], pages 305–314. ISBN 90-277-0408-2, 90-277-0409-0, 94-010-2140-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175 .B7312.

URL https://link.springer.com/
chapter/10.1007/978-94-010-2140-

1_19. See comments [Bre74a].

Supek:1979:SCI

[Sup79] Rudi Supek. Some contradictions and insufficiencies of Yugoslav selfmanaging socialism. In Praxis: Yugoslav Essays in the Philosophy and Methodology of the Social Sciences [MP79], chapter 16, pages 249–271. ISBN 90-277-0968-8, 94-009-9355-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-009-9355-
                                              6_16.

Suppe:1991:OOF

[Sup91] Frederick Suppe. The observational origins of Feyerabend’s anarchistic epistemology. In Beyond Reason: Essays on the Philosophy of Paul Feyerabend [Mun91a], chapter 13, pages 297–311. ISBN 94-010-5406-1, 94-011-3188-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-011-3188-
                                              9_13.

Suppes:2003:TEB

[Sup03] Patrick Suppes. From theory to experiment and back again. In Descartes’s Mathematical Thought [Sas03d], chapter 1, pages 1–41. ISBN 90-481-6487-7, 94-017-1225-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic).

                                              LCCN ????  URL https://link.
                                              springer.com/chapter/10.1007/0-
                                              306-48123-5_1.

Susnjic:1979:IL

[Šuš79] Djuro Šušnjić. Ideas and life. In Praxis: Yugoslav Essays in the Philosophy and Methodology of the Social Sciences [MP79], chapter 12, pages 199–208. ISBN 90-277-0968-8, 94-009-9355-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-009-9355-
6_12.

SanMauro:2015:NM

[SV15] Luca San Mauro and Giorgio Venturi.

Naturalness in mathematics. In From Logic to Practice: Italian Studies in the Philosophy of Mathematics [LPV15], chapter 14, pages 277–313. ISBN 3-319-10433-0, 3-319-10434-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QA8.4 .F76 2015; Q174 .B67 v.308. URL https://link.

springer.com/chapter/10.1007/
978-3-319-10434-8_14.

Schiaffonati:2018:WDB

[SV18] Viola Schiaffonati and Mario Verdicchio. What do bridges and software tell us about the philosophy of engineering? In Mitcham et al. [MLNZ18], chapter 12, pages 137–148. ISBN 3-319-62448-2, 3-319-62450-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN TA15 .P45 2018.

URL https://link.springer.com/
chapter/10.1007/978-3-319-62450-
1_12.

Sanginga:1995:MBF

[SVD95] N. Sanginga, B. Vanlauwe, and S. K. A.

Danso. Management of biological N_{2} fixation in alley cropping systems: Estimation and contribution to N balance. In No Truth Except in the Details: Essays in Honor of Martin J. Klein [KS95], chapter 6, pages 119–141. ISBN 94-010-4097-4, 94-011-0217-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-011-0055-7_6.

Svobodny:2015:UCE

[Svo15] Petr Svobodný. Universities in Central Europe: Changing perspectives in the troubled Twentieth Century. In Sciences in the Universities of Europe, Nineteenth and Twentieth Centuries: Academic Landscapes [SDG15b], chapter 7, pages 107–123. ISBN 94-017-9635-1, 94-017-9636-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q183.4.E85 2015.

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-017-9636-
                                              1_7.

Swann:2009:PSI

[Swa09] Joanna Swann. Popperian selectionism and its implications for education, or ‘what to do about the myth of learning by instruction from without?’. In Rethinking Popper [PC09], chapter 28, pages 379–388. ISBN 1-4020-9337-3, 1-4020-9338-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN B1649.P64 R48 2009; Q174 .B67 2009.

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-1-4020-9338-
                                              8_28.

Swartz:2017:ED

[Swa17] Ronald Swartz. Education for democracy. In Encouraging Openness: Essays for Joseph Agassi on the Occasion of His 90th Birthday [BAG17], chapter 25, pages 301–304. ISBN 3-319-57668-2 (hardcover), 3-319-57669-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN B833.E53 2017.

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-3-319-57669-
                                              5_25.

Swerdlow:1990:RCP

[Swe90] Noel M. Swerdlow. Regiomontanus on the critical problems of astronomy.

In Levere and Shea [LS90], chapter 7, pages 165–195. ISBN 0-7923-0420-9, 94-009-1878-X. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 120. URL https://link.

springer.com/chapter/10.1007/
978-94-009-1878-8_7.

Swerdlow:1994:BND

[Swe94] Noel M. Swerdlow. Blackstone’s “Newtonian” dissent. In Cohen [Coh94b], chapter 5, pages 205–234. ISBN 0-7923-2223-1, 94-017-3391-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 150.

URL https://link.springer.com/
chapter/10.1007/978-94-017-3391-
5_5.

Swimmer:1996:CGN

[Swi96] Alvin Swimmer. The completion of Grassmann’s Natur-Wissenschaftliche Methode. In Schubring [Sch96f], pages 265–280. ISBN 90-481-4758-1, 94-015-8753-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QA8.9-10.3. URL https://link.

springer.com/chapter/10.1007/
978-94-015-8753-2_22.

Swindal:1999:NCT

[Swi99] James Swindal. Nietzsche, critical theory and a theory of knowledge. In Nietzsche, Theories of Knowledge, and Critical Theory: Nietzsche and the Sciences I [Bab99b], chapter 18, pages 253–264. ISBN 90-481-5233-X, 94-017-2430-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-017-
2430-2_18.

Sylla:1975:AHS

[Syl75] Edith Dudley Sylla. Autonomous and handmaiden science: St. Thomas Aquinas and William of Ockham on the physics of the Eucharist. In Murdoch and Sylla [MS75b], pages 349–396. ISBN 90-277-0587-9, 94-010-1781-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN B721

.I584 1975.  URL https://link.
springer.com/chapter/10.1007/
978-94-010-1781-7_10.

Sylvan:1989:QQD

[Syl89] David J. Sylvan. The qualitativequantitative distinction in political science. In The Qualitative-Quantitative Distinction in the Social Sciences [GM89b], chapter 6, pages 79–97. ISBN 90-481-8460-6, 94-017-3444-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 112.

URL https://link.springer.com/
chapter/10.1007/978-94-017-3444-
8_6.

Sylla:1991:OCM

[Syl91] Edith D. Sylla. The Oxford calculators and mathematical physics: John Dumbleton’s Summa Logicae et Philosophiae Naturalis, parts II and III. In Unguru [Ung91b], pages 129–161. ISBN 94-010-5476-2, 94-011-3342-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC5.53; QC6.9.

URL https://link.springer.com/
chapter/10.1007/978-94-011-3342-
5_7.

Sylvan:1997:IRP

[Syl97a] Richard Sylvan. Issues in regional philosophy. Austrian philosophy? And its austral image? In Austrian Philosophy Past and Present: Essays in Honor of Rudolf Haller [LM97], chapter 10, pages 147–166. ISBN 94-010-6412-1, 94-011-5720-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-011-
5720-9_10.

Sylvan:1997:MSC

[Syl97b] Richard Sylvan. Metaphysics: Destroyed or in-de(con)structible. In Austrian Philosophy Past and Present: Essays in Honor of Rudolf Haller [LM97], chapter 11, pages 167–176. ISBN 94-010-6412-1, 94-011-5720-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-011-5720-
9_11.

Sylla:2008:OFT

[Syl08] Edith Dudley Sylla. The origin and fate of Thomas Bradwardine’s De Proportionibus Velocitatum in Motibus in relation to the history of mathematics. In Mechanics and Natural Philosophy Before the Scientific Revolution [LR08], chapter 3, pages 67–119. ISBN 1-4020-5966-3 (hardcover), 1-4020-5967-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QA802 .M43

2008. URL https://link.springer.
com/chapter/10.1007/978-1-4020-
5967-4_3.

Synowiecki:1982:PCO

[Syn82] Adam Synowiecki. The problem of the chemical organization of matter in the light of a closed development model. In Polish Essays in the Philosophy of the Natural Sciences [Kra82c], chapter 32, pages 377–390. ISBN 94-009-7705-0 (ebook), 90-277-1287-5. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-009-
7705-1_32.

Szacki:1979:CHS

[Sza79a] Jerzy Szacki. On so-called historicism in the social sciences. In Polish Essays in the Methodology of the Social Sciences [Wia79b], chapter 10, pages 195–209. ISBN 90-277-0956-4, 94-009-9353-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174

                                              .P655 1979. URL https://link.
                                              springer.com/chapter/10.1007/
                                              978-94-009-9353-2_10.

Szaniawski:1979:SMR

[Sza79b] Klemens Szaniawski. Sociology and models of rational behavior. In Polish Essays in the Methodology of the Social Sciences [Wia79b], chapter 11, pages 211–224. ISBN 90-277-0956-4, 94-009-9353-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .P655 1979. URL https://link.

                                              springer.com/chapter/10.1007/
                                              978-94-009-9353-2_11.

Szabo:1984:HEH

[Sza84] Arpad K. Szabó. How to explore the history of ancient mathematics? In Methodology, Metaphysics and the History of Science: In Memory of Benjamin Nelson [CW84b], chapter 12, pages 283–294. ISBN 94-009-6333-5, 94-009-6331-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

                                              ???? URL https://link.springer.
                                              com/chapter/10.1007/978-94-009-
                                              6331-3_12.

Szabo:2003:TEB

[Sza03] László E. Szabó. From theory to experiments and back again … and back again …: Comments on Patrick Suppes. In Descartes’s Mathematical Thought [Sas03d], chapter 3, pages 57– 62. ISBN 90-481-6487-7, 94-017-1225-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/0-306-48123-5_3.

Szabo:2004:WMD

[Sza04] Á. Szabó. Wie ist die Mathematik zu Einer Deduktiven Wissenschaft Geworden?. (German) [How did math become a deductive science?]. In Classics in the History of Greek Mathematics [Chr04a], chapter 3, pages 45–80. ISBN 1-4020-0081-2 (hardcover), 1-4020-2640-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QA22 .C596 2004. URL https:/

/link.springer.com/chapter/10.
1007/978-1-4020-2640-9_3.

Sztompka:1979:MFS

[Szt79a] Piotr Sztompka. Marxism, functionalism and systems-approach. In Polish Essays in the Methodology of the Social Sciences [Wia79b], chapter 7, pages 133–156. ISBN 90-277-0956-4, 94-009-9353-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .P655 1979. URL https://link.

springer.com/chapter/10.1007/
978-94-009-9353-2_7.

Sztompka:1979:STC

[Szt79b] Piotr Sztompka. Strategy of theory– construction in sociology. In Polish Essays in the Methodology of the Social Sciences [Wia79b], chapter 9, pages 173–194. ISBN 90-277-0956-4, 94-009-9353-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .P655 1979. URL https://link.

springer.com/chapter/10.1007/
978-94-009-9353-2_9.

Szubka:2000:QDP

[Szu00] Tadeusz Szubka. Quine and Davidson on perceptual knowledge. In Knowledge, Language and Logic: Questions for Quine [OK00], chapter 2, pages 7– 19. ISBN 1-4020-0253-X, 94-011-3933-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-011-3933-
5_2.

Szymanski:1982:TIL

[Szy82] Jerzy Szymański. Testing idealizational laws. In Polish Essays in the Philosophy of the Natural Sciences [Kra82c], chapter 13, pages 147–151. ISBN 94-009-7705-0 (e-book), 90-277-1287-5. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-009-7705-1_13.

Tabakov:2003:LBT

[Tab03] Martin Tabakov. The logic between two centuries. In Bulgarian Studies in the Philosophy of Science [Gin03a], chapter 8, pages 95–125. ISBN 1-4020-1496-1, 94-017-0961-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-017-0961-3_8.

Tachibana:2017:HCW

[Tac17] Kiichi Tachibana. How can we attain both democracy and constitutionalism? In Encouraging Openness: Essays for Joseph Agassi on the Occasion of His 90th Birthday [BAG17], chapter 26, pages 305–318. ISBN 3-319-57668-2 (hardcover), 3-319-57669-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN B833.E53 2017.

URL https://link.springer.com/
chapter/10.1007/978-3-319-57669-
5_26.

Tadic:1979:BRO

[Tad79] Ljubomir Tadić. Bureaucracy — reified organization. In Praxis: Yugoslav Essays in the Philosophy and Methodology of the Social Sciences [MP79], chapter 18, pages 289–301. ISBN 90-277-0968-8, 94-009-9355-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-009-9355-6_18.

Tamas:1986:LC

[Tam86] György Tamás. The Logic of Categories, volume 85 of Boston Studies in the Philosophy of Science. D. Reidel, Dordrecht, The Netherlands; Boston, MA, USA; Lancaster, UK; Tokyo, Japan, 1986. ISBN 90-277-1742-7. xiv + 524 pp. LCCN BC172 .T3513 1986; Q174 .B67 vol. 85.

Tammaru:2001:SR

[Tam01] Jüri Tammaru. Symmetry and rationality. In Estonian Studies in the History and Philosophy of Science [Vih01b], chapter 13, pages 179–184. ISBN 94-010-3863-5, 94-010-0672-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-010-0672-9_13.

Taminiaux:2002:MPR

[Tam02] Jacques Taminiaux. Merleau-Ponty’s reading of Heidegger. In Hermeneutic Philosophy of Science, Van Gogh’s Eyes, and God: Essays in Honor of Patrick A. Heelan, S.J. [Bab02c], chapter 21, pages 251–263. ISBN 90-481-5926-1, 94-017-1767-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 225.

URL https://link.springer.com/

chapter/10.1007/978-94-017-1767-
0_21.

Tamargo:2014:CUR

[Tam14] Maria C. Tamargo. Cuban/US research interactions since 1995. In The History of Physics in Cuba [BRW14a], chapter 31, pages 413–416. ISBN 94-017-8040-4, 94-017-8041-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC5.53; QC6.9; QC9 C9H57 2014; Q124.6-127.2.

URL https://link.springer.com/
chapter/10.1007/978-94-017-8041-
4_31.

Tanaka:1992:MEK

[Tan92] Setsuko Tanaka. Mach, Einstein, and Kuwaki. In Ernst Mach — A Deeper Look: Documents and New Perspectives [Bla92a], chapter 16, pages 297–332. ISBN 94-010-5230-1, 94-011-2771-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN C5.53;

QC6.9. URL https://link.springer.
com/chapter/10.1007/978-94-011-
2771-4_16.

Tankler:2001:UBT

[Tan01] Hain Tankler. A university between two cultures. In Estonian Studies in the History and Philosophy of Science [Vih01b], chapter 2, pages 19–34. ISBN 94-010-3863-5, 94-010-0672-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-010-0672-9_2.

Tapia:1992:TES

[Tap91] O. Tapia. Theoretical evaluation of solvent effects. In The Invention of Physical Science: Intersections of Mathematics, Theology and Natural Philosophy Since the Seventeenth Century: Essays in Honor of Erwin N.

Hiebert [NRS92], chapter 11, pages 435–458. ISBN 94-010-5097-X, 94-011-2488-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-3-642-
58183-0_11.

Tassy:2017:MPT

[Tas17] Pascal Tassy. The making of paleontological time. In Time of Nature and the Nature of Time: Philosophical Perspectives of Time in Natural Sciences [BH17b], chapter 12, pages 253–271. ISBN 3-319-53723-7 (hardcover), 3-319-53725-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC173.59.S65 T56 2017. URL https:

//link.springer.com/chapter/10.
1007/978-3-319-53725-2_12.

Tauber:1991:ECS

[Tau91a] Alfred I. Tauber. Editorial comments to Sober by Sarkar. In Organism and the Origins of Self [Tau91i], chapter 12, pages 273–274. ISBN 0-7923-1185-X, 94-011-3406-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QR180-189.5. URL https:/

/link.springer.com/chapter/10.
1007/978-94-011-3406-4_12.

Tauber:1991:ECF

[Tau91b] Alfred I. Tauber. Editor’s comments to Foster and Sarkar. In Organism and the Origins of Self [Tau91i], chapter 9, pages 211–212. ISBN 0-7923-1185-X, 94-011-3406-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QR180-189.5. URL https:/

/link.springer.com/chapter/10.
1007/978-94-011-3406-4_9.

Tauber:1991:ECG

[Tau91c] Alfred I. Tauber. Editor’s comments to Gilbert. In Organism and the Origins of Self [Tau91i], chapter 16, pages 337–340. ISBN 0-7923-1185-X, 94-011-3406-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QR180-189.5. URL https:/

                                              /link.springer.com/chapter/10.
                                              1007/978-94-011-3406-4_16.  See

[Gil91].

Tauber:1991:ECL

[Tau91d] Alfred I. Tauber. Editor’s comments to löwy. In Organism and the Origins of Self [Tau91i], chapter 2, page 41. ISBN 0-7923-1185-X, 94-011-3406-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QR180-189.5.

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-011-3406-
                                              4_2.

Tauber:1991:ECR

[Tau91e] Alfred I. Tauber. Editor’s comments to Root–Bernstein. In Organism and the Origins of Self [Tau91i], chapter 7, pages 157–158. ISBN 0-7923-1185-X, 94-011-3406-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QR180-189.5. URL https:/

                                              /link.springer.com/chapter/10.
                                              1007/978-94-011-3406-4_7.

Tauber:1991:ECV

[Tau91f] Alfred I. Tauber. Editor’s comments to Varela, Chernyak and Tauber. In Organism and the Origins of Self [Tau91i], chapter 4, pages 77–78. ISBN 0-7923-1185-X, 94-011-3406-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QR180-189.5.

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-011-3406-
                                              4_4.

Tauber:1991:ECW

[Tau91g] Alfred I. Tauber. Editor’s comments to Williamson. In Organism and the Origins of Self [Tau91i], chapter 14, pages 297–298. ISBN 0-7923-1185-X, 94-011-3406-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QR180-189.5. URL https:/

/link.springer.com/chapter/10.
1007/978-94-011-3406-4_14.  See

[Wil91].

Tauber:1991:ISC

[Tau91h] Alfred I. Tauber. Introduction: Speculations concerning the origins of the self. In Organism and the Origins of Self [Tau91i], chapter 1, pages 1–39. ISBN 0-7923-1185-X, 94-011-3406-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QR180-189.5.

URL https://link.springer.com/
chapter/10.1007/978-94-011-3406-
4_1.

Tauber:1991:OOS

[Tau91i] Alfred I. Tauber. Organism and the Origins of Self, volume 129 of Boston Studies in the Philosophy of Science. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 1991. ISBN 0-7923-1185-X, 94-011-3406-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). xix + 384 pp. LCCN QR180-189.5.

URL https://link.springer.com/
book/10.1007/978-94-011-3406-4.

Tauber:1995:TV

[Tau95] Alfred I. Tauber. On the transvaluation of values. In Science, Mind and Art: Essays on science and the humanistic understanding in art, epistemology, religion and ethics In honor of Robert S. Cohen [GSW95b], chapter 21, pages 349–367. ISBN 0-7923-2990-2 (hardcover), 94-011-0469-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol.

165 BD418.3. URL https://link.

                                              springer.com/chapter/10.1007/
                                              978-94-011-0469-2_21.

Tauber:1997:ESA

[Tau97a] Alfred I. Tauber. The Elusive Synthesis: Aesthetics and Science, volume 182 of Boston Studies in the Philosophy of Science. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 1997. ISBN 0-7923-4763-3, 0-7923-3904-5, 94-009-1786-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). 328 (est.) pp. LCCN ???? URL

                                              https://link.springer.com/book/
                                              10.1007/978-94-009-1786-6.

Tauber:1997:DDH

[Tau97b] Alfred I. Tauber. From Descartes’ dream to Husserl’s nightmare. In The Elusive Synthesis: Aesthetics and Science [Tau97a], chapter 14, pages 289–312. ISBN 0-7923-4763-3, 0-7923-3904-5, 94-009-1786-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-009-1786-6_14.

Tauber:1997:I

[Tau97c] Alfred I. Tauber. Introduction. In The Elusive Synthesis: Aesthetics and Science [Tau97a], chapter 1, pages 1–12. ISBN 0-7923-4763-3, 0-7923-3904-5, 94-009-1786-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

                                              ???? URL https://link.springer.
                                              com/chapter/10.1007/978-94-009-
                                              1786-6_1.

Tauber:1998:ECS

[Tau98] Alfred I. Tauber. Ecology and the claims for a science-based ethics. In Philosophies of Nature: The Human Dimension: In Celebration of Erazim Kohák [CT98], chapter 14, pages 185–206. ISBN 90-481-4859-6, 94-017-2614-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic).

LCCN BD581. URL https://link.
springer.com/chapter/10.1007/
978-94-017-2614-6_14.

Taureck:1999:HCN

[Tau99] Bernhard H. F. Taureck. Habermas’s critique of Nietzsche’s critique of reason. In Nietzsche, Theories of Knowledge, and Critical Theory: Nietzsche and the Sciences I [Bab99b], chapter 22, pages 289–294. ISBN 90-481-5233-X, 94-017-2430-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-017-
2430-2_22.

Taylor:1992:EPR

[Tay92]  Charles Taylor.  Explanation and

practical reason. In The Scientific Enterprise: The Bar-Hillel Colloquium: Studies in History, Philosophy, and Sociology of Science, Volume 4 [UM92], chapter 12, pages 179–201. ISBN 94-010-5190-9, 94-011-2688-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .S354 1992.

URL https://link.springer.com/
chapter/10.1007/978-94-011-2688-
5_12.

Tang:1996:ARC

[TB96] Paul C. L. Tang and Ralph W. Brown III. Anti-realism and the complementarity model of mind–brain. In Cohen et al. [CHR96], pages 467–479. ISBN 90-481-4493-0, 94-015-8638-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-015-8638-2_34.

Thyssen:2015:MRE

[TB15] Pieter Thyssen and Koen Binnemans. Mendeleev and the rare-earth crisis. In Philosophy of Chemistry: Growth of a New Discipline [SM15b], chapter 11, pages 155–182. ISBN 94-017-9363-8 (hardcover), 94-017-9364-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QD6 .P455 2015; QD6 .S384 2014. URL https://link.

                                              springer.com/chapter/10.1007/
                                              978-94-017-9364-3_11.

Tomasi:1992:EEP

[TBC91] Jacopo Tomasi, Rosanna Bonaccorsi, and Roberto Cammi. The extramolecular electrostatic potential. An indicator of the chemical reactivity. In The Invention of Physical Science: Intersections of Mathematics, Theology and Natural Philosophy Since the Seventeenth Century: Essays in Honor of Erwin N. Hiebert [NRS92], chapter 7, pages 229–268. ISBN 94-010-5097-X, 94-011-2488-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic).

                                              LCCN ????  URL https://link.
                                              springer.com/chapter/10.1007/
                                              978-3-642-58183-0_7.

Tegtmeier:1997:DTP

[Teg97] Erwin Tegtmeier. Direction of time: A problem of ontology, not of physics. In Perspectives on Time [FSU97b], chapter 7, pages 183–191. ISBN 90-481-4774-3, 94-015-8875-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol.189.

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-015-8875-
                                              1_7.

Telesio:1976:ITM

[Tel76] B. Telesio. Independence of time from motion. In The Concepts of Space and Time: Their Structure and Their Development [Čap76], chapter 32, pages 187–188. ISBN 90-277-0375-2, 94-010-1727-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic).

LCCN BD632. URL https://link.
springer.com/chapter/10.1007/
978-94-010-1727-5_32.

Teller:2008:CII

[Tel08] Paul Teller. Of course idealizations are incommensurable! In Rethinking Scientific Change and Theory Comparison: Stabilities, Ruptures, Incommensurabilities? [SSHH08], chapter 18, pages 247–264. ISBN 1-4020-6274-5 (hardcover), 1-4020-6279-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175 .R397 2008.

URL https://link.springer.com/
chapter/10.1007/978-1-4020-6279-
7_18.

Tempczyk:1982:SLP

[Tem82] Michał Tempczyk. Structural laws in physics. In Polish Essays in the Philosophy of the Natural Sciences [Kra82c], chapter 25, pages 287–297. ISBN 94-009-7705-0 (e-book), 90-277-1287-5. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-009-7705-1_25.

Tempczyk:1995:LDC

[Tem95]  Michał Tempczyk.  Limitations of

determinism in classical mechanics. In Misiek [Mis95c], pages 215–219. ISBN 94-010-4206-3, 94-011-0461-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-011-0461-6_16.

Tillers:1988:PIL

[TG88] Peter Tillers and Eric D. Green, editors. Probability and Inference in the Law of Evidence: the Uses and Limits of Bayesianism, volume 109 of Boston Studies in the Philosophy of Science. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 1988. ISBN 90-277-2689-2. LCCN Q174 .B67 vol. 109; K2261 .P76 1987.

terHark:2009:PTS

[tH09] Michel ter Hark. Popper’s theory of the searchlight: A historical assessment of its significance. In Rethinking Popper [PC09], chapter 14, pages 175–184. ISBN 1-4020-9337-3, 1-4020-9338-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN B1649.P64 R48 2009; Q174 .B67 2009.

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-1-4020-9338-
                                              8_14.

Thao:1984:IGO

[Tha84a] Trân Duc Thao. The indicative gesture as the original form of consciousness. In Investigations into the Origin of Language and Consciousness [Tha84b], chapter 1, pages 3–29. ISBN 94-009-6238-X, 94-009-6236-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .T736 1984.

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-009-6236-
                                              1_1.

Thao:1984:IOL

[Tha84b] Trân Duc Thao. Investigations into the Origin of Language and Consciousness, volume 44 of Boston Studies in the Philosophy of Science. D. Reidel, Dordrecht, The Netherlands; Boston, MA, USA; Lancaster, UK; Tokyo, Japan, 1984. ISBN 94-009-6238-X, 94-009-6236-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). 244 + 2 pp. LCCN Q174 .T736 1984. URL

https://link.springer.com/book/
10.1007/978-94-009-6236-1.

Thao:1984:MPO

[Tha84c] Trân Duc Thao. Marxism and psychoanalysis — the origins of the Oedipal crisis. In Investigations into the Origin of Language and Consciousness [Tha84b], chapter 3, pages 145–197. ISBN 94-009-6238-X, 94-009-6236-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .T736

1984. URL https://link.springer.
com/chapter/10.1007/978-94-009-
6236-1_3.

Thao:1984:SL

[Tha84d] Trân Duc Thao. Syncretic language.

In Investigations into the Origin of Language and Consciousness [Tha84b], chapter 2, pages 33–141. ISBN 94-009-6238-X, 94-009-6236-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .T736 1984.

URL https://link.springer.com/
chapter/10.1007/978-94-009-6236-
1_2.

Thao:1986:DAB

[Tha86a] Trân Duc Thao. The dialectic of animal behavior as the becoming of sense certainty. In Phenomenology and Dialectical Materialism [Coh86d], chapter 6, pages 143–178. ISBN 94-010-8795-4, 94-009-5191-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-009-5191-4_6.

Thao:1986:DHS

[Tha86b] Trân Duc Thao. The dialectic of human societies as the becoming of reason. In Phenomenology and Dialectical Materialism [Coh86d], chapter 7, pages 179–218. ISBN 94-010-8795-4, 94-009-5191-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

                                              ???? URL https://link.springer.
                                              com/chapter/10.1007/978-94-009-
                                              5191-4_7.

Thao:1986:IPT

[Tha86c] Trân Duc Thao. Introduction to part two. In Phenomenology and Dialectical Materialism [Coh86d], chapter 5, pages 133–142. ISBN 94-010-8795-4, 94-009-5191-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

                                              ???? URL https://link.springer.
                                              com/chapter/10.1007/978-94-009-
                                              5191-4_5.

Thao:1986:IE

[Tha86d] Trân Duc Thao. The intuition of essences. In Phenomenology and Dialectical Materialism [Coh86d], chapter 1, pages 3–11. ISBN 94-010-8795-4, 94-009-5191-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic).

                                              LCCN ????  URL https://link.
                                              springer.com/chapter/10.1007/
                                              978-94-009-5191-4_1.

Thao:1986:PR

[Tha86e] Trân Duc Thao. The problems of reason. In Phenomenology and Dialectical Materialism [Coh86d], chapter 3, pages 69–120. ISBN 94-010-8795-4, 94-009-5191-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

                                              ???? URL https://link.springer.
                                              com/chapter/10.1007/978-94-009-
                                              5191-4_3.

Thao:1986:RP

[Tha86f] Trân Duc Thao. The result of phenomenology. In Phenomenology and Dialectical Materialism [Coh86d], chapter 4, pages 121–130. ISBN 94-010-8795-4, 94-009-5191-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-009-5191-4_4.

Thao:1986:TCC

[Tha86g] Trân Duc Thao. The thematization of concrete consciousness. In Phenomenology and Dialectical Materialism [Coh86d], chapter 2, pages 13–68. ISBN 94-010-8795-4, 94-009-5191-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-009-5191-
4_2.

Thiel:2013:DCM

[Thi11] Christian Thiel. Dubislav and classical monadic quantificational logic. In The Berlin Group and the Philosophy of Logical Empiricism [MP13], chapter 8, pages 179–189. ISBN 94-007-5484-1 (hardcover), 94-007-5485-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175.32.A24 .M384 2011; Q175 .B47 2013. URL https:

//link.springer.com/chapter/10.
1007/978-94-007-5485-0_8.

Thomson:1967:C

[Tho67] Judith Jarvis Thomson. Comments.

In Cohen and Wartofsky [CW67], pages 324–330. ISBN 90-277-0013-3, 94-010-3508-3. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175

.B731 v.3.  URL https://link.
springer.com/chapter/10.1007/
978-94-010-3508-8_17.

Thomason:1976:NQT

[Tho76] Richmond H. Thomason. Necessity, quotation, and truth: An indexical theory. In Language in Focus: Foundations, Methods and Systems: Essays in Memory of Yehoshua Bar-Hillel [Kas76b], chapter 8, pages 119–138. ISBN 90-277-0644-1, 90-277-0645-X, 94-010-1876-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN P101-120. URL https://link.

                                              springer.com/chapter/10.1007/
                                              978-94-010-1876-0_8.

Thomas:1995:RLP

[Tho95] R. J. Thomas. Role of legumes in providing n for sustainable tropical pasture systems. In No Truth Except in the Details: Essays in Honor of Martin J. Klein [KS95], chapter 5, pages 103–118. ISBN 94-010-4097-4, 94-011-0217-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-011-0055-
                                              7_5.

Thompson:1996:AIA

[Tho96] Evan Thompson. Artificial intelligence, artificial life, and the symbol–matter problem. In Québec Studies in the Philosophy of Science: Part II: Biology, Psychology, Cognitive Science and Economics Essays in Honor of Hugues Leblanc [MC96], chapter 5, pages 63– 80. ISBN 94-010-6537-3, 94-009-0113-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-009-0113-
                                              1_5.

Thomson:2001:MTM

[Tho01] Ann Thomson. Materialistic theories of mind and brain. In Between Leibniz, Newton, and Kant: Philosophy and Science in the Eighteenth Century [Lef01a], chapter 8, pages 149–173. ISBN 90-481-5774-9, 94-015-9729-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-015-9729-
6_8.

Thompson:2010:SLM

[Tho10] C. Michele Thompson. Sinification as limitation: Minh Ma.ng’s prohibition on use of Nôm and the resulting marginalization of Nôm medical texts. In Looking at it from Asia: the Processes that Shaped the Sources of History of Science [BE10b], chapter 13, pages 393–412. ISBN 90-481-3675-X (hardcover), 90-481-3676-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q127.A65 L66

2010. URL https://link.springer.
com/chapter/10.1007/978-90-481-
3676-6_13.

Thoms:2014:IFF

[Tho14] Ulrike Thoms. The introduction of frozen foods in West Germany and its integration into the daily diet. In History of Artificial Cold, Scientific, Technological and Cultural Issues [Gav14b], chapter 11, pages 201–229. ISBN 94-007-7198-3 (hardcover), 94-007-7199-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC278 H57 2014; TP494 .H577 2014.

URL https://link.springer.com/
chapter/10.1007/978-94-007-7199-
4_11.

Tisza:1963:LSP

[Tis63] Laszlo Tisza. The logical structure of physics. In Wartofsky [War63], pages 55–71. ISBN 94-010-3265-3, 94-010-3263-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B678 1970. URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-010-3263-6_5.

Tisza:1967:C

[Tis67] Laszlo Tisza. Comments. In Cohen and Wartofsky [CW67], pages 228–235. ISBN 90-277-0013-3, 94-010-3508-3. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175 .B731 v.3.

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-010-3508-
                                              8_12.

Tisza:1974:CCH

[Tis74] Laszlo Tisza. Comments on C. A. Hooker: Systematic realism. In Logical and Epistemological Studies in Contemporary Physics [CW74a], chapter 4, pages 192–195. ISBN 90-277-0391-4 (hardcover), 90-277-0377-9 (paperback), 94-010-2656-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol.

                                              13; QC5.56. URL https://link.
                                              springer.com/chapter/10.1007/
                                              978-94-010-2656-7_4.

Tisza:2003:CWL

[Tis03] Laszlo Tisza. Can we learn from history? Do we want to? In Renn et al. [RDS03], chapter 28, pages 555–568. ISBN 1-4020-1284-5, 1-4020-1285-3 (paperback). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 234. URL https://link.

                                              springer.com/chapter/10.1007/
                                              978-94-010-0111-3_28.

Thomson-Jones:2011:SAS

[TJ11] Martin Thomson-Jones. Structuralism about scientific representation. In Bokulich and Bokulich [BB11], chapter 7, pages 119–141. ISBN 90-481-9596-9, 90-481-9597-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175.32.S78 S35

2011. URL https://link.springer.
com/chapter/10.1007/978-90-481-
9597-8_7.

Tuyn:1991:FEL

[TO91] W. Tuyn and H. Kamerlingh Onnes.

Further experiments with liquid helium, AA. The disturbance of supraconductivity by magnetic fields and currents. The hypothesis of Silsbee. In Gavroglu and Goudaroulis [GG91], chapter 23, pages 363–387. ISBN 94-010-7433-X (hardcover), 94-009-2079-2. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 124.

URL https://link.springer.com/
chapter/10.1007/978-94-009-2079-
8_23.

Toader:2015:AHI

[Toa15] Iulian D. Toader. Against harmony:

Infinite idealizations and causal explanation. In Romanian Studies in Philosophy of Science [PST15], chapter 18, pages 291–301. ISBN 3-319-16654-9 (hardcover), 3-319-16655-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-3-319-16655-
1_18.

Tobies:1996:RHG

[Tob96] Renate Tobies. The reception of H.

Grassmann’s mathematical achievements by A. Clebsch and his school. In Schubring [Sch96f], pages 117–130. ISBN 90-481-4758-1, 94-015-8753-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QA8.9-10.3.

URL https://link.springer.com/

                                              chapter/10.1007/978-94-015-8753-
                                              2_10.

Toellner:1981:LPA

[Toe81] Richard Toellner. Logical and psychological aspects of the discovery of the circulation of the blood. In Grmek et al. [GCC81b], pages 239–259. ISBN 90-277-1123-2, 94-010-1284-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-010-1284-3_12.

Tokumaru:2009:PAP

[Tok09] Natsuka Tokumaru. Popper’s analysis of the problems of induction and demarcation and Mises’ justification of the theoretical social sciences. In Rethinking Popper [PC09], chapter 13, pages 161–174. ISBN 1-4020-9337-3, 1-4020-9338-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN B1649.P64 R48 2009; Q174 .B67 2009.

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-1-4020-9338-
                                              8_13.

Tondl:1973:CSC

[Ton73a] Ladislav Tondl. The concept of ‘science’ as cognitive activity. In Scientific Procedures: A Contribution Concerning the Methodological Problems of Scientific Concepts and Scientific Explanation [Ton73f], chapter 1, pages 1–13. ISBN 90-277-0147-4 (hardcover), 90-277-0323-X (paperback), 94-010-2916-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 10; Q175 .T663s 1973.

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-010-2916-
                                              2_1.

Tondl:1973:CLS

[Ton73b] Ladislav Tondl. Concepts of the language of science. In Scientific Procedures: A Contribution Concerning the Methodological Problems of Scientific Concepts and Scientific Explanation [Ton73f], chapter 4, pages 46–125. ISBN 90-277-0147-4 (hardcover), 90-277-0323-X (paperback), 94-010-2916-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 10; Q175 .T663s 1973. URL https:

//link.springer.com/chapter/10.
1007/978-94-010-2916-2_4.

Tondl:1973:EBA

[Ton73c] Ladislav Tondl. The empirical basis and the analysis of the ‘universe of discourse’. In Scientific Procedures: A Contribution Concerning the Methodological Problems of Scientific Concepts and Scientific Explanation [Ton73f], chapter 3, pages 23–45. ISBN 90-277-0147-4 (hardcover), 90-277-0323-X (paperback), 94-010-2916-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 10; Q175 .T663s 1973. URL https:

//link.springer.com/chapter/10.
1007/978-94-010-2916-2_3.

Tondl:1973:AMS

[Ton73d] Ladislav Tondl. On the approach to models of scientific procedures. In Scientific Procedures: A Contribution Concerning the Methodological Problems of Scientific Concepts and Scientific Explanation [Ton73f], chapter 2, pages 14–22. ISBN 90-277-0147-4 (hardcover), 90-277-0323-X (paperback), 94-010-2916-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 10; Q175 .T663s 1973. URL https://link.

                                              springer.com/chapter/10.1007/
                                              978-94-010-2916-2_2.

Tondl:1973:SE

[Ton73e] Ladislav Tondl. Scientific explanation. In Scientific Procedures: A Contribution Concerning the Methodological Problems of Scientific Concepts and Scientific Explanation [Ton73f], chapter 5, pages 126–264. ISBN 90-277-0147-4 (hardcover), 90-277-0323-X (paperback), 94-010-2916-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 10; Q175 .T663s 1973. URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-010-2916-2_5.

Tondl:1973:SPC

[Ton73f] Ladislav Tondl. Scientific Procedures: A Contribution Concerning the Methodological Problems of Scientific Concepts and Scientific Explanation, volume 10 of Boston Studies in the Philosophy of Science. D. Reidel, Dordrecht, The Netherlands; Boston, MA, USA; Lancaster, UK; Tokyo, Japan, 1973. ISBN 90-277-0147-4 (hardcover), 90-277-0323-X (paperback), 94-010-2916-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). xiii + 268 pp. LCCN Q174 .B67 vol. 10; Q175 .T663s 1973.

                                              URL https://link.springer.com/
                                              book/10.1007/978-94-010-2916-2.

Tondl:1981:CSP

[Ton81a] Ladislav Tondl. The concept of semantics and prerequisites for the investigation of semantic problems. In Problems of Semantics: A Contribution to the Analysis of the Language Science [Ton81f], chapter 2, pages 10–30. ISBN 90-277-0316-7, 94-009-8364-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-009-8364-9_2.

Tondl:1981:CS

[Ton81b] Ladislav Tondl. The criterion of sense.

In Problems of Semantics: A Contribution to the Analysis of the Language Science [Ton81f], chapter 6, pages 166–228. ISBN 90-277-0316-7, 94-009-8364-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-009-
8364-9_6.

Tondl:1981:OER

[Ton81c] Ladislav Tondl. An outline of the evaluation of the results of scientific activity in terms of semantic information. In Problems of Semantics: A Contribution to the Analysis of the Language Science [Ton81f], chapter 9, pages 274–303. ISBN 90-277-0316-7, 94-009-8364-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-009-8364-
9_9.

Tondl:1981:OSE

[Ton81d] Ladislav Tondl. An outline of the semantic evaluation of graphic communication. In Problems of Semantics: A Contribution to the Analysis of the Language Science [Ton81f], chapter 12, pages 338–353. ISBN 90-277-0316-7, 94-009-8364-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-009-
8364-9_12.

Tondl:1981:PIS

[Ton81e] Ladislav Tondl. The problem of informational synonymity. In Problems of Semantics: A Contribution to the Analysis of the Language Science [Ton81f], chapter 11, pages 325–337. ISBN 90-277-0316-7, 94-009-8364-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-009-8364-9_11.

Tondl:1981:PSC

[Ton81f] Ladislav Tondl. Problems of Semantics: A Contribution to the Analysis of the Language Science, volume 66 of Boston Studies in the Philosophy of Science. D. Reidel, Dordrecht, The Netherlands; Boston, MA, USA; Lancaster, UK; Tokyo, Japan, 1981. ISBN 90-277-0316-7, 94-009-8364-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). xiv + 407 pp. LCCN ????

URL https://link.springer.com/
book/10.1007/978-94-009-8364-9.

Tondl:1981:SC

[Ton81g] Ladislav Tondl. Semantic concepts. In Problems of Semantics: A Contribution to the Analysis of the Language Science [Ton81f], chapter 3, pages 31–60. ISBN 90-277-0316-7, 94-009-8364-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-009-8364-9_3.

Tondl:1981:SPS

[Ton81h] Ladislav Tondl. The semantic problem — sources and themes. In Problems of Semantics: A Contribution to the Analysis of the Language Science [Ton81f], chapter 1, pages 1–9. ISBN 90-277-0316-7, 94-009-8364-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-009-8364-9_1.

Tondl:1981:SSP

[Ton81i] Ladislav Tondl. Semantics and some problems of ontology. In Problems of Semantics: A Contribution to the Analysis of the Language Science [Ton81f], chapter 8, pages 247–273. ISBN 90-277-0316-7, 94-009-8364-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-009-8364-
9_8.

Tondl:1981:SLC

[Ton81j] Ladislav Tondl. The semantics of logical concepts. In Problems of Semantics: A Contribution to the Analysis of the Language Science [Ton81f], chapter 4, pages 61–102. ISBN 90-277-0316-7, 94-009-8364-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-009-
8364-9_4.

Tondl:1981:SPA

[Ton81k] Ladislav Tondl. The semantics of preference attitudes. In Problems of Semantics: A Contribution to the Analysis of the Language Science [Ton81f], chapter 10, pages 304–324. ISBN 90-277-0316-7, 94-009-8364-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-009-8364-9_10.

Tondl:1981:SD

[Ton81l] Ladislav Tondl. Sense and denotation.

In Problems of Semantics: A Contribution to the Analysis of the Language Science [Ton81f], chapter 5, pages 103–165. ISBN 90-277-0316-7, 94-009-8364-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

                                              ???? URL https://link.springer.
                                              com/chapter/10.1007/978-94-009-
                                              8364-9_5.

Tondl:1981:V

[Ton81m] Ladislav Tondl. Vagueness. In Problems of Semantics: A Contribution to the Analysis of the Language Science [Ton81f], chapter 7, pages 229–246. ISBN 90-277-0316-7, 94-009-8364-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-009-8364-
                                              9_7.

Tondl:1997:CS

[Ton97] Ladislav Tondl. Cognition as a system. In Ginev and Cohen [GC97], chapter 21, pages 357–375. ISBN 0-7923-4444-8 (hardcover), 94-011-5788-X. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 192.

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-011-5788-
                                              9_21.

Toomer:2004:LGM

[Too04] G. J. Toomer. Lost Greek mathematical works in Arabic translation. In Classics in the History of Greek Mathematics [Chr04a], chapter 16, pages 275–284. ISBN 1-4020-0081-2 (hardcover), 1-4020-2640-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QA22 .C596 2004.

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-1-4020-2640-
                                              9_16.

Toon:2010:MMB

[Too10] Adam Toon. Models as make-believe. In Frigg and Hunter [FH10b], chapter 5, pages 71–96. ISBN 90-481-3850-7 (hardcover), 90-481-3851-5 (e-book).

ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN BH301.M47 B49 2010; Q175 .B52 2010. URL https:

//link.springer.com/chapter/10.
1007/978-90-481-3851-7_5.

Topitsch:1984:EFC

[Top84] Ernst Topitsch. On early forms of critical rationalism. In Rationality in Science and Politics [And84d], chapter 16, pages 277–287. ISBN 90-277-1953-5, 94-009-6254-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-009-6254-5_16.

Tornebohm:1976:ISP

[Tör76] Håkan Törnebohm. Inquiring systems and paradigms. In Essays in Memory of Imre Lakatos [CFW76], chapter 35, pages 635–654. ISBN 90-277-0655-7, 94-010-1451-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-010-
1451-9_35.

Torretti:1982:TKM

[Tor82] Roberto Torretti. Three kinds of mathematical fictionalism. In Scientific Philosophy Today: Essays in Honor of Mario Bunge [AC82], chapter 22, pages 399–414. ISBN 90-277-1262-X, 90-277-1263-8, 94-009-8462-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 67.

URL https://link.springer.com/
chapter/10.1007/978-94-009-8462-
2_22.

Torretti:1983:CSS

[Tor83] Roberto Torretti. Causality and spacetime structure in relativity. In Cohen and Laudan [CL83], chapter 14, pages 273–293. ISBN 94-009-7055-2, 94-009-7057-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-009-
7055-7_14.

Torres:1995:PPH

[Tor95] Carlos Torres. The philosophy and the programme of Hilbert. In Mexican Studies in the History and Philosophy of Science [RC95], chapter 10, pages 151–172. ISBN 94-010-6535-7 (hardcover), 94-009-0109-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-009-0109-4_10.

Toronyai:1999:KSR

[Tor99] Gábor Toronyai. Kuhn, Scheler and the revolutionary genesis of modern science. In Fehér et al. [FKR99], pages 59–74. ISBN 90-481-5257-7, 94-015-9293-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-015-9293-
2_5.

Torretti:2000:SRS

[Tor00] Roberto Torretti. ‘Scientific realism’ and scientific practice. In The Reality of the Unobservable: Observability, Unobservability and Their Impact on the Issue of Scientific Realism [AP00b], chapter 6, pages 113–122. ISBN 90-481-5458-8, 94-015-9391-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-015-9391-5_6.

Tosi:1981:IPF

[Tos81] Piero Tosi. Intuitive proofs and firstorder derivations: Some notes on the metamathematics of first-order number theory. In Italian Studies in the Philosophy of Science [Chi81b], chapter 9, pages 137–152. ISBN 90-277-1073-2, 94-009-8937-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-009-8937-5_9.

Toulmin:1967:CRS

[Tou67] Stephen Toulmin. Conceptual revolutions in science. In Cohen and Wartofsky [CW67], pages 331–347. ISBN 90-277-0013-3, 94-010-3508-3. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175 .B731 v.3.

URL https://link.springer.com/
chapter/10.1007/978-94-010-3508-
8_18.

Toulmin:1971:LSC

[Tou71] Stephen Toulmin. From logical systems to conceptual populations. In Buck and Cohen [BC71], pages 552–564. ISBN 90-277-0309-4, 94-010-3142-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-010-3142-4_37.

Toulmin:1974:RSD

[Tou74a] Stephen Toulmin. Rationality and scientific discovery. In Schaffner and Cohen [SC74a], pages 387–406. ISBN 90-277-0408-2, 90-277-0409-0, 94-010-2140-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175 .B7312. URL https://link.

springer.com/chapter/10.1007/
978-94-010-2140-1_27.

Toulmin:1974:SSH

[Tou74b] Stephen Toulmin. Scientific strategies and historical change. In Seeger and Cohen [SC74b], pages 401–414. ISBN 90-277-0390-6, 90-277-0376-0 (paperback), 94-010-2126-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 11.

URL https://link.springer.com/
chapter/10.1007/978-94-010-2126-
5_23.

Toulmin:1976:HPT

[Tou76] Stephen Toulmin. History, praxis and the ‘third world’. In Essays in Memory of Imre Lakatos [CFW76], chapter 36, pages 655–675. ISBN 90-277-0655-7, 94-010-1451-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-010-1451-9_36.

Toulmin:1986:CRS

[Tou86a] Stephen Toulmin. Conceptual revolutions in science. In Cohen and Wartofsky [CW85], chapter 6, pages 58–74. ISBN 90-277-1971-3, 94-009-5345-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-009-
5345-1_6.

Toulmin:1986:LFH

[Tou86b] Stephen Toulmin. Ludwik Fleck and the historical interpretation of science. In Cognition and Fact: Materials on Ludwik Fleck [CS86], chapter 14, pages 267–285. ISBN 94-010-8504-8, 94-009-4498-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-009-
4498-5_14.

Toulmin:2002:HNS

[Tou02] Stephen Toulmin. The hermeneutics of the natural sciences. In Hermeneutic Philosophy of Science, Van Gogh’s Eyes, and God: Essays in Honor of Patrick A. Heelan, S.J. [Bab02c], chapter 2, pages 25–29. ISBN 90-481-5926-1, 94-017-1767-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 225.

URL https://link.springer.com/
chapter/10.1007/978-94-017-1767-
0_2.

Townsend:1971:FPT

[Tow71] Burke Townsend. Feyerabend’s pragmatic theory of observation and the comparability of alternative theories. In Buck and Cohen [BC71], pages 202–211. ISBN 90-277-0309-4, 94-010-3142-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-010-
3142-4_14.

Tzavaras:1990:DFT

[TP90] A. Tzavaras and G. Papagounos. The development of Freudian theory: The role of the ‘centre’ and the ‘excentric’ in theory production and diffusion. In Greek Studies in the Philosophy and History of Science [Nic90], chapter 15, pages 227–237. ISBN 94-010-7403-8, 94-009-2015-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-009-
2015-6_15.

Tranchini:2015:PIR

[Tra15] Luca Tranchini. Paradox and inconsistency: Revising Tennant’s distinction through Schroeder–Heister’s assumption rules. In From Logic to Practice: Italian Studies in the Philosophy of Mathematics [LPV15], chapter 7, pages 111–121. ISBN 3-319-10433-0, 3-319-10434-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QA8.4 .F76 2015; Q174 .B67 v.308.

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-3-319-10434-
                                              8_7.

Trizio:2008:SRV

[Tri08] Emiliano Trizio. Scientific revolutions: the view from inside and the view from outside. In Rethinking Scientific Change and Theory Comparison: Stabilities, Ruptures, Incommensurabilities? [SSHH08], chapter 25, pages 381–384. ISBN 1-4020-6274-5 (hardcover), 1-4020-6279-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175 .R397 2008.

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-1-4020-6279-
                                              7_25.

Trizio:2012:ARC

[Tri12] Emiliano Trizio. Achieving robustness to confirm controversial hypotheses: A case study in cell biology. In Characterizing the Robustness of Science: After the Practice Turn in Philosophy of Science [STNW12], chapter 4, pages 105–120. ISBN 94-007-2758-5, 94-007-2759-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175 .C4595 2012. URL https:/

                                              /link.springer.com/chapter/10.
                                              1007/978-94-007-2759-5_4.

Tsou:2015:IOS

[TRP15] Jonathan Y. Tsou, Alan Richardson, and Flavia Padovani. Introduction: Objectivity in science. In Objectivity in Science: New Perspectives from Science and Technology Studies [PRT15], chapter 1, pages 1–15. ISBN 3-319-14348-4 (hardcover), 3-319-14349-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175 O25 2015.

URL https://link.springer.com/
chapter/10.1007/978-3-319-14349-
1_1.

Truitt:1974:IEM

[Tru74] Willis H. Truitt. Ideology, expression, and mediation. In For Dirk Struik: Scientific, Historical and Political Essays in Honor of Dirk J. Struik [CSW74], chapter 35, pages 435–463. ISBN 90-277-0379-5, 94-010-2115-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-010-2115-9_35.

Truesdell:1982:DEE

[Tru82] C. Truesdell. The disastrous effects of experiment upon the early development of thermodynamics. In Scientific Philosophy Today: Essays in Honor of Mario Bunge [AC82], chapter 23, pages 415–423. ISBN 90-277-1262-X, 90-277-1263-8, 94-009-8462-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 67.

URL https://link.springer.com/
chapter/10.1007/978-94-009-8462-
2_23.

Truitt:1994:PUD

[Tru94] Willis H. Truitt. Partisanship, universalism, and the dialectics of moral consciousness. In Artifacts, Representations and Social Practice: Essays for Marx Wartofsky [GC94], chapter 31, pages 543–550. ISBN 94-010-4390-6, 94-011-0902-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN B73 .A77 1994. URL https:

//link.springer.com/chapter/10.
1007/978-94-011-0902-4_31.

Truitt:1995:NPN

[Tru95] Willis H. Truitt. Neo-pragmatism and the new aesthetic. In Science, Mind and Art: Essays on science and the humanistic understanding in art, epistemology, religion and ethics In honor of Robert S. Cohen [GSW95b], chapter 2, pages 19–27. ISBN 0-7923-2990-2 (hardcover), 94-011-0469-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 165 BD418.3. URL https:

//link.springer.com/chapter/10.
1007/978-94-011-0469-2_2.

Tsinorema:1990:WRR

[Tsi90] Stavroula F. Tsinorema. Wittgenstein, rationality and relativism. In Greek Studies in the Philosophy and History of Science [Nic90], chapter 26, pages 397–414. ISBN 94-010-7403-8, 94-009-2015-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-009-
2015-6_26.

Tsou:2015:RCK

[Tso15] Jonathan Y. Tsou. Reconsidering the Carnap–Kuhn connection. In Kuhn’s Structure of Scientific Revolutions — 50 Years On [DB15b], chapter 5, pages 51–69. ISBN 3-319-13382-9, 3-319-13383-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175.3 .K84 2015. URL https:

//link.springer.com/chapter/10.
1007/978-3-319-13383-6_5.

Tuana:2016:CCT

[Tua16] Nancy Tuana. Climate change through the lens of feminist philosophy. In Meta-Philosophical Reflection on Feminist Philosophies of Science [AV16a], chapter 3, pages 35–53. ISBN 3-319-26346-3 (hardcover), 3-319-26348-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN HQ1154 .M48 2016; HQ1190 M48 2016. URL https:

//link.springer.com/chapter/10.
1007/978-3-319-26348-9_3.

Tucker:1976:ASM

[Tuc76] John Tucker. Against some methods.

In Essays in Memory of Imre Lakatos [CFW76], chapter 37, pages 677–689. ISBN 90-277-0655-7, 94-010-1451-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-010-1451-
9_37.

Tulczyjew:2003:NHC

[Tul03] Wlodzimierz M. Tulczyjew. A note on holonomic constraints. In Renn et al. [RDS03], chapter 20, pages 403–419. ISBN 1-4020-1284-5, 1-4020-1285-3 (paperback). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 234. URL https://link.

springer.com/chapter/10.1007/
978-94-010-0111-3_20.

Tulodziecki:2016:ZGD

[Tul16] Dana Tulodziecki. From zymes to germs: Discarding the realist/antirealist framework. In Sauer and Scholl [SS16b], chapter 13, pages 265–283. ISBN 3-319-30227-2, 3-319-30229-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q126.8

.P45 2016.  URL https://link.
springer.com/chapter/10.1007/
978-3-319-30229-4_13.

Tuomela:1976:CDE

[Tuo76] Raimo Tuomela. Causes and deductive explanation. In Cohen et al. [CHMV76], pages 325–360. ISBN 90-277-0647-6, 90-277-0648-4 (paperback), 94-010-1449-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-010-
1449-6_16.

Tuomela:1982:ICF

[Tuo82] Raimo Tuomela. Individualism and concept formation in the social sciences. In Scientific Philosophy Today: Essays in Honor of Mario Bunge [AC82], chapter 24, pages 425–438. ISBN 90-277-1262-X, 90-277-1263-8, 94-009-8462-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 67. URL https:

//link.springer.com/chapter/10.
1007/978-94-009-8462-2_24.

Tuomela:1987:SPP

[Tuo87] Raimo Tuomela. Science, protoscience, and pseudoscience. In Pitt and Pera [PP87], chapter 4, pages 83–101. ISBN 90-277-2417-2, 94-009-3779-2. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol.

98. US$64.00. URL https://link.

springer.com/chapter/10.1007/
978-94-009-3779-6_4.

Turner:1986:BEC

[Tur86a]  Stephen P. Turner.  Beyond the

enlightenment: Comte and the new problem of social science. In The Search for a Methodology of Social Science: Durkheim, Weber, and the Nineteenth-Century Problem of Cause, Probability, and Action [Tur86l], chapter 2, pages 6–28. ISBN 90-481-8417-7, 94-017-3461-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-017-
3461-5_2.

Turner:1986:CFC

[Tur86b] Stephen P. Turner. Collective forces, causation, and probability. In The Search for a Methodology of Social Science: Durkheim, Weber, and the Nineteenth-Century Problem of Cause, Probability, and Action [Tur86l], chapter 7, pages 124–143. ISBN 90-481-8417-7, 94-017-3461-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-017-3461-5_7.

Turner:1986:DI

[Tur86c] Stephen P. Turner. Durkheim’s individual. In The Search for a Methodology of Social Science: Durkheim, Weber, and the Nineteenth-Century Problem of Cause, Probability, and Action [Tur86l], chapter 8, pages 144–160. ISBN 90-481-8417-7, 94-017-3461-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-017-3461-
5_8.

Turner:1986:EA

[Tur86d] Stephen P. Turner. The end of the ascent. In The Search for a Methodology of Social Science: Durkheim, Weber, and the Nineteenth-Century Problem of Cause, Probability, and Action [Tur86l], chapter 12, pages 219–227. ISBN 90-481-8417-7, 94-017-3461-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-017-3461-
5_12.

Turner:1986:I

[Tur86e] Stephen P. Turner. The interregnum.

In The Search for a Methodology of Social Science: Durkheim, Weber, and the Nineteenth-Century Problem of Cause, Probability, and Action [Tur86l], chapter 5, pages 92–104. ISBN 90-481-8417-7, 94-017-3461-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-017-3461-5_5.

Turner:1986:LSE

[Tur86f] Stephen P. Turner. Large-scale explanations: Aggregation and interpretation. In The Search for a Methodology of Social Science: Durkheim, Weber, and the Nineteenth-Century Problem of Cause, Probability, and Action [Tur86l], chapter 11, pages 198–218. ISBN 90-481-8417-7, 94-017-3461-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-017-3461-
                                              5_11.

Turner:1986:MAC

[Tur86g] Stephen P. Turner. Mill and ‘the ascent to causes’. In The Search for a Methodology of Social Science: Durkheim, Weber, and the Nineteenth-Century Problem of Cause, Probability, and Action [Tur86l], chapter 3, pages 29– 59. ISBN 90-481-8417-7, 94-017-3461-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-017-3461-
                                              5_3.

Turner:1986:OPA

[Tur86h] Stephen P. Turner. Objective possibility and adequate cause. In The Search for a Methodology of Social Science: Durkheim, Weber, and the Nineteenth-Century Problem of Cause, Probability, and Action [Tur86l], chapter 9, pages 163–179. ISBN 90-481-8417-7, 94-017-3461-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-017-
3461-5_9.

Turner:1986:QRT

[Tur86i] Stephen P. Turner. Quetelet: Rates and their explanation. In The Search for a Methodology of Social Science: Durkheim, Weber, and the Nineteenth-Century Problem of Cause, Probability, and Action [Tur86l], chapter 4, pages 60–91. ISBN 90-481-8417-7, 94-017-3461-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-017-
3461-5_4.

Turner:1986:RA

[Tur86j] Stephen P. Turner. Rationality and action. In The Search for a Methodology of Social Science: Durkheim, Weber, and the Nineteenth-Century Problem of Cause, Probability, and Action [Tur86l], chapter 10, pages 180–197. ISBN 90-481-8417-7, 94-017-3461-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-017-3461-
5_10.

Turner:1986:RTA

[Tur86k] Stephen P. Turner. Realism, teleology, and action. In The Search for a Methodology of Social Science: Durkheim, Weber, and the Nineteenth-Century Problem of Cause, Probability, and Action [Tur86l], chapter 6, pages 107–123. ISBN 90-481-8417-7, 94-017-3461-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-017-
3461-5_6.

Turner:1986:SMS

[Tur86l] Stephen P. Turner. The Search for a Methodology of Social Science: Durkheim, Weber, and the Nineteenth-Century Problem of Cause, Probability, and Action, volume 92 of Boston Studies in the Philosophy of Science. D. Reidel, Dordrecht, The Netherlands; Boston, MA, USA; Lancaster, UK; Tokyo, Japan, 1986. ISBN 90-481-8417-7, 94-017-3461-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). xv + 255 pp. LCCN ???? URL

                                              https://link.springer.com/book/
                                              10.1007/978-94-017-3461-5.

Turner:1986:TG

[Tur86m] Stephen P. Turner. Two generations. In The Search for a Methodology of Social Science: Durkheim, Weber, and the Nineteenth-Century Problem of Cause, Probability, and Action [Tur86l], chapter 1, pages 3–5. ISBN 90-481-8417-7, 94-017-3461-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-017-3461-5_1.

Turner:1996:OCW

[Tur96] R. Steven Turner. The origins of colorimetry: What did Helmholtz and Maxwell learn from Grassmann? In Schubring [Sch96f], pages 71–86. ISBN 90-481-4758-1, 94-015-8753-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QA8.9-10.3.

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-015-8753-
                                              2_8.

Turkheimer:2012:GWA

[Tur12] Eric Turkheimer. Genome wide association studies of behavior are social science. In Philosophy of Behavioral Biology [PR12b], chapter 3, pages 43– 64. ISBN 94-007-1950-7, 94-007-1951-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QH331 .P4679 2012; Q175 .B657 v.282. URL https:

//link.springer.com/chapter/10.
1007/978-94-007-1951-4_3.

Turetzky:2016:PST

[Tur16] Philip Turetzky. The passive syntheses of time. In Cosmological and Psychological Time [DR16], chapter 11, pages 177–202. ISBN 3-319-22589-8 (hardcover), 3-319-22590-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN BD638 C67 2016.

URL https://link.springer.com/
chapter/10.1007/978-3-319-22590-
6_11.

Torres:1995:SP

[TV95] Carlos Torres and Jaime Oscar Falcon Vege. To show and to prove. In Mexican Studies in the History and Philosophy of Science [RC95], chapter 16, pages 249–264. ISBN 94-010-6535-7 (hardcover), 94-009-0109-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-009-0109-4_16.

Tversky:1986:CIJ

[Tve86] Amos Tversky. Cognitive illusions in judgment and choice. In Ullmann-Margalit [UM86a], chapter 7, pages 75–87. ISBN 90-277-2159-9 (paperback), 90-277-2158-0 (hardcover), 94-009-5496-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 94 Q175.

URL https://link.springer.com/
chapter/10.1007/978-94-009-5496-
0_7.

Twersky:2000:AME

[Twe00] Isadore Twersky. Aspects of Maimonides’ epistemology: Halakah and science. In Maimonides and the Sciences [CL00], chapter 11, pages 227–243. ISBN 90-481-5348-4, 94-017-2128-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

                                              ???? URL https://link.springer.
                                              com/chapter/10.1007/978-94-017-
                                              2128-8_11.

Tymieniecka:1995:KCS

[Tym95] Anna-Teresa Tymieniecka. Knowledge and cognition in the self-individualizing progress of life. In The Concept of Knowledge: The Ankara Seminar [KC95], chapter 16, pages 197–217. ISBN 90-481-4495-7, 94-017-3263-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-017-3263-
                                              5_16.

Tympas:2015:HCT

[Tym15] Aristotle Tympas. On the hazardousness of the concept ‘technology’: Notes on a conversation between the history of science and the history of technology. In Arabatzis et al. [ARS15], chapter 22, pages 329–342. ISBN 3-319-14552-5 (paperback), 3-319-14553-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q125 .R45 2015.

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-3-319-14553-
                                              2_22.

Uebel:1991:ONN

[Ueb91a] Thomas E. Uebel. Otto Neurath and the Neurath reception: Puzzle and promise. In Uebel [Ueb91b], chapter 1, pages 3–22. ISBN 0-7923-1276-7 (hardcover), 94-011-3182-1. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 133. URL https:/

/link.springer.com/chapter/10.
1007/978-94-011-3182-7_1.

Uebel:1991:RFV

[Ueb91b] Thomas E. (Thomas Ernst) Uebel, editor. Rediscovering the forgotten Vienna Circle: Austrian studies on Otto Neurath and the Vienna Circle, volume 133 of Boston Studies in the Philosophy of Science. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 1991. ISBN 0-7923-1276-7 (hardcover), 94-011-3182-1. ISSN 0068-0346 (print), 2214-7942 (electronic). xii + 326 pp. LCCN Q174 .B67 vol. 133. URL

https://link.springer.com/book/
10.1007/978-94-011-3182-7.

Uebel:1995:PWV

[Ueb95] Thomas E. Uebel. Physicalism in Wittgenstein and the Vienna Circle. In Gavroglu et al. [GSW95c], chapter 19, pages 327–356. ISBN 0-7923-2991-0 (set), 0-7923-2988-0, 94-017-2658-2. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 163.

URL https://link.springer.com/
chapter/10.1007/978-94-017-2658-
0_19.

Uebel:1997:DTN

[Ueb97] Thomas E. Uebel. From the Duhem thesis to the Neurath principle. In Austrian Philosophy Past and Present: Essays in Honor of Rudolf Haller [LM97], chapter 7, pages 87–100. ISBN 94-010-6412-1, 94-011-5720-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-011-5720-9_7.

Ugaglia:2018:EVC

[Uga18] Monica Ugaglia. Existence vs. conceivability in Aristotle: Are straight lines infinitely extendible? In Piazza and Pulcini [PP18], pages 249–272. ISBN 3-319-93341-8, 3-319-93342-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QA8.4.

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-3-319-93342-
                                              9_14.

Ullmann-Margalit:1986:KSI

[UM86a] Edna Ullmann-Margalit, editor. The Kaleidoscope of Science: The Israel Colloquium: Studies in History, Philosophy, and Sociology of Science. Volume 1, volume 94 of Boston Studies in the Philosophy of Science. D. Reidel, Dordrecht, The Netherlands; Boston, MA, USA; Lancaster, UK; Tokyo, Japan, 1986. ISBN 90-277-2159-9 (paperback), 90-277-2158-0 (hardcover), 94-009-5496-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). xii + 250 pp. LCCN Q174 .B67 vol. 94 Q175.

                                              URL https://link.springer.com/
                                              book/10.1007/978-94-009-5496-0.

Ullmann-Margalit:1986:PRB

[UM86b] Edna Ullmann-Margalit. Practical reasoning — the bottom line: A comment. In The Prism of Science: The Israel Colloquium: Studies in History, Philosophy, and Sociology of Science. Volume 2 [UM86c], chapter 8, pages 107–111. ISBN 90-277-2160-2, 90-277-2161-0 (paperback), 94-009-4566-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 95

                                              Q175. URL https://link.springer.
                                              com/chapter/10.1007/978-94-009-
                                              4566-1_8.

Ullmann-Margalit:1986:PSI

[UM86c] Edna Ullmann-Margalit, editor. The Prism of Science: The Israel Colloquium: Studies in History, Philosophy, and Sociology of Science. Volume 2, volume 95(2) of Boston Studies in the Philosophy of Science. D. Reidel, Dordrecht, The Netherlands; Boston, MA, USA; Lancaster, UK; Tokyo, Japan, 1986. ISBN 90-277-2160-2, 90-277-2161-0 (paperback), 94-009-4566-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). ix + 250 pp. LCCN Q174 .B67 vol. 95 Q175.

URL  http://www.springerlink.
com/content/978-94-009-4566-1;
https://link.springer.com/book/
10.1007/978-94-009-4566-1.

Ullmann-Margalit:1988:SRI

[UM88] Edna Ullmann-Margalit. Science in Reflection: The Israel Colloquium: Studies in History, Philosophy, and Sociology of Science Volume 3, volume 110 of Boston Studies in the Philosophy of Science. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 1988. ISBN 90-277-2713-9, 94-009-2957-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). 239 (est.) pp. LCCN ????

URL https://link.springer.com/
book/10.1007/978-94-009-2957-9.

Ullmann-Margalit:1992:SEB

[UM92] Edna Ullmann-Margalit. The Scientific Enterprise: The Bar-Hillel Colloquium: Studies in History, Philosophy, and Sociology of Science, Volume 4, volume 146 of Boston Studies in the Philosophy of Science. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 1992. ISBN 94-010-5190-9, 94-011-2688-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). ix + 302 pp. LCCN Q174 .S354 1992. URL

                                              https://link.springer.com/book/
                                              10.1007/978-94-011-2688-5.

Unguru:1991:EMO

[Ung91a] Sabetai Unguru. Experiment in medieval optics. In Physics, Cosmology and Astronomy, 1300–1700: Tension and Accommodation [Ung91b], pages 163–181. ISBN 94-010-5476-2, 94-011-3342-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC5.53; QC6.9. URL https://link.

                                              springer.com/chapter/10.1007/
                                              978-94-011-3342-5_8.

Unguru:1991:PCA

[Ung91b] Sabetai Unguru, editor. Physics, Cosmology and Astronomy, 1300–1700: Tension and Accommodation, volume 126 of Boston Studies in the Philosophy of Science. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 1991. ISBN 94-010-5476-2, 94-011-3342-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC5.53; QC6.9.

                                              URL https://link.springer.com/
                                              book/10.1007/978-94-011-3342-5.

Unguru:1994:MAA

[Ung94]  Sabetai Unguru.  Is mathematics

ahistorical? An attempt to an answer motivated by Greek mathematics. In Trends in the Historiography of Science [GCN94], chapter 16, pages 203–219. ISBN 90-481-4264-4, 94-017-3596-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-017-3596-
                                              4_16.

Unguru:2004:HAM

[Ung04a] Sabetai Unguru. History of ancient mathematics: Some reflections on the state of the art. In Classics in the History of Greek Mathematics [Chr04a], chapter 25, pages 451–461. ISBN 1-4020-0081-2 (hardcover), 1-4020-2640-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QA22 .C596

2004. URL https://link.springer.
com/chapter/10.1007/978-1-4020-
2640-9_25.

Unguru:2004:I

[Ung04b] Sabetai Unguru. Introduction. In Classics in the History of Greek Mathematics [Chr04a], chapter 21, pages 383–384. ISBN 1-4020-0081-2 (hardcover), 1-4020-2640-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QA22 .C596 2004.

URL https://link.springer.com/
chapter/10.1007/978-1-4020-2640-
9_21.

Unguru:2004:NRH

[Ung04c] Sabetai Unguru. On the need to rewrite the history of Greek mathematics. In Classics in the History of Greek Mathematics [Chr04a], chapter 22, pages 385–431. ISBN 1-4020-0081-2 (hardcover), 1-4020-2640-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QA22 .C596 2004.

URL https://link.springer.com/
chapter/10.1007/978-1-4020-2640-
9_22.

Ungureanu:2015:MTO

[Ung15] Manuela L. Ungureanu. A momentous triangle: Ontology, methodology and phenomenology in the philosophy of language. In Romanian Studies in Philosophy of Science [PST15], chapter 4, pages 55–70. ISBN 3-319-16654-9 (hardcover), 3-319-16655-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-3-319-16655-1_4.

Unt:2001:CIQ

[Unt01] Anto Unt. The Copenhagen interpretation of quantum mechanics and common sense. In Estonian Studies in the History and Philosophy of Science [Vih01b], chapter 18, pages 247–262. ISBN 94-010-3863-5, 94-010-0672-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-010-0672-
                                              9_18.

Urbach:1978:OPR

[Urb78] Peter Urbach. The objective promise of a research programme. In Progress and Rationality in Science [RA78b], chapter 5, pages 99–113. ISBN 90-277-0922-X, 94-009-9866-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-009-9866-7_5.

Urbach:1989:BAM

[Urb89] Peter Urbach. The Bayesian alternative to the methodology of scientific research programmes. In Imre Lakatos and Theories of Scientific Change [GGN89], chapter 29, pages 399–412. ISBN 94-010-7860-2, 94-009-3025-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-009-3025-
                                              4_29.

Urchs:1997:TE

[Urc97] Max Urchs. Tense and existence. In Perspectives on Time [FSU97b], chapter 8, pages 193–200. ISBN 90-481-4774-3, 94-015-8875-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol.189.

URL https://link.springer.com/
chapter/10.1007/978-94-015-8875-
1_8.

Ursua:1996:BOC

[Urs96] Nicanor Ursúa. The biologization of our culture. In Munévar [Mun96], chapter 11, pages 227–242. ISBN 94-010-6622-1, 94-009-0305-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-009-0305-0_11.

Uus:2001:GMM

[Uus01] Undo Uus. The glory and misery of modern science. In Estonian Studies in the History and Philosophy of Science [Vih01b], chapter 8, pages 111–122. ISBN 94-010-3863-5, 94-010-0672-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-010-0672-
9_8.

Uyemov:1997:SAP

[Uye97] A. I. Uyemov. System approach to the problem of the classification of sciences and scientific researches. In Ginev and Cohen [GC97], chapter 22, pages 377–390. ISBN 0-7923-4444-8 (hardcover), 94-011-5788-X. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 192. URL https:/

/link.springer.com/chapter/10.
1007/978-94-011-5788-9_22.

Vajda:1999:WGN

[Vaj99] Mihály Vajda. Worum Geht es in den Naturwissenschaften?. (German)

[What is science about?]. In Fehér et al. [FKR99], pages 75–84. ISBN 90-481-5257-7, 94-015-9293-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-015-9293-2_6.

Valentini:1996:PWT

[Val96] Antony Valentini. Pilot-wave theory of fields, gravitation and cosmology. In Bohmian Mechanics and Quantum Theory: an Appraisal [CFG96], chapter 3, pages 45–66. ISBN 90-481-4698-4, 94-015-8715-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC173.96-174.52. URL https:

//link.springer.com/chapter/10.
1007/978-94-015-8715-0_3.

Valadier:1999:SNR

[Val99] Paul Valadier. Science as new religion. In Nietzsche, Epistemology, and Philosophy of Science: Nietzsche and the Sciences II [Bab99a], chapter 19, pages 241–252. ISBN 90-481-5234-8, 94-017-2428-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-017-
2428-9_19.

Valleriani:2010:AEA

[Val10a] Matteo Valleriani. Artist–engineers’ apprenticeship and Galileo. In Galileo Engineer [Val10b], chapter 1, pages 3–20. ISBN 90-481-8644-7 (hardcover), 90-481-8645-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic).

LCCN ????  URL https://link.
springer.com/chapter/10.1007/
978-90-481-8645-7_1.

Valleriani:2010:GE

[Val10b] Matteo Valleriani. Galileo Engineer, volume 269 of Boston Studies in the Philosophy of Science. Springer-Ver-lag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2010. ISBN 90-481-8644-7 (hardcover), 90-481-8645-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). xxii + 320 pp. LCCN ???? URL

https://link.springer.com/book/
10.1007/978-90-481-8645-7.

Valleriani:2010:GPC

[Val10c] Matteo Valleriani. Galileo’s private course on fortifications. In Galileo Engineer [Val10b], chapter 3, pages 71–113. ISBN 90-481-8644-7 (hardcover), 90-481-8645-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-90-481-
8645-7_3.

Valleriani:2010:IM

[Val10d] Matteo Valleriani. Instruments and machines. In Galileo Engineer [Val10b], chapter 2, pages 21–70. ISBN 90-481-8644-7 (hardcover), 90-481-8645-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-90-481-8645-
7_2.

Valleriani:2010:KVA

[Val10e] Matteo Valleriani. The knowledge of the Venetian Arsenal. In Galileo Engineer [Val10b], chapter 4, pages 117–153. ISBN 90-481-8644-7 (hardcover), 90-481-8645-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-90-481-
8645-7_4.

Valleriani:2010:PTN

[Val10f] Matteo Valleriani. Pneumatics, the thermoscope and the new atomistic conception of heat. In Galileo Engineer [Val10b], chapter 5, pages 155–190. ISBN 90-481-8644-7 (hardcover), 90-481-8645-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-90-481-
8645-7_5.

Valleriani:2010:WGE

[Val10g] Matteo Valleriani. Was Galileo an engineer? In Galileo Engineer [Val10b], chapter 6, pages 193–211. ISBN 90-481-8644-7 (hardcover), 90-481-8645-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-90-481-8645-
7_6.

Vamvacas:2009:ACC

[Vam09a] Constantine J. Vamvacas. Anaxagoras of Clazomenae (ca. 500—428 B.C.). In The Founders of Western Thought — The Presocratics: A diachronic parallelism between Presocratic Thought and Philosophy and the Natural Sciences [Vam09g], chapter 12, pages 197–208. ISBN 1-4020-9790-5 (hardcover), 1-4020-9791-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175 .V33413 2009. URL https:/

/link.springer.com/chapter/10.
1007/978-1-4020-9791-1_12.

Vamvacas:2009:AMCa

[Vam09b] Constantine J. Vamvacas. Anaximander of Miletus (ca. 610—546 B.C.). In The Founders of Western Thought — The Presocratics: A diachronic parallelism between Presocratic Thought and Philosophy and the Natural Sciences [Vam09g], chapter 5, pages 35–43. ISBN 1-4020-9790-5 (hardcover), 1-4020-9791-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175 .V33413 2009. URL https:/

/link.springer.com/chapter/10.
1007/978-1-4020-9791-1_5.

Vamvacas:2009:AMCb

[Vam09c] Constantine J. Vamvacas. Anaximenes of Miletus (ca. 585—525 B.C.). In The Founders of Western Thought — The Presocratics: A diachronic parallelism between Presocratic Thought and Philosophy and the Natural Sciences [Vam09g], chapter 6, pages 45–51. ISBN 1-4020-9790-5 (hardcover), 1-4020-9791-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175 .V33413 2009. URL https:/

/link.springer.com/chapter/10.
1007/978-1-4020-9791-1_6.

Vamvacas:2009:DAC

[Vam09d] Constantine J. Vamvacas. Democritus of Abdera (ca. 460—360 B.C.). In The Founders of Western Thought — The Presocratics: A diachronic parallelism between Presocratic Thought and Philosophy and the Natural Sciences [Vam09g], chapter 13, pages 209–248. ISBN 1-4020-9790-5 (hardcover), 1-4020-9791-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175 .V33413 2009. URL https:/

/link.springer.com/chapter/10.
1007/978-1-4020-9791-1_13.

Vamvacas:2009:EAC

[Vam09e] Constantine J. Vamvacas. Empedocles of Acragas (ca. 494—434 B.C.). In The Founders of Western Thought — The Presocratics: A diachronic parallelism between Presocratic Thought and Philosophy and the Natural Sciences [Vam09g], chapter 11, pages 167–196. ISBN 1-4020-9790-5 (hardcover), 1-4020-9791-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175 .V33413 2009. URL https:/

/link.springer.com/chapter/10.
1007/978-1-4020-9791-1_11.

Vamvacas:2009:E

[Vam09f] Constantine J. Vamvacas. Epilogue. In The Founders of Western Thought — The Presocratics: A diachronic parallelism between Presocratic Thought and Philosophy and the Natural Sciences [Vam09g], chapter 14, pages 249–254. ISBN 1-4020-9790-5 (hardcover), 1-4020-9791-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175 .V33413 2009. URL https:/

/link.springer.com/chapter/10.
1007/978-1-4020-9791-1_14.

Vamvacas:2009:FWT

[Vam09g] Constantine J. Vamvacas. The Founders of Western Thought — The Presocratics: A diachronic parallelism between Presocratic Thought and Philosophy and the Natural Sciences, volume 257 of Boston Studies in the Philosophy of Science. Springer-Ver-lag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2009. ISBN 1-4020-9790-5 (hardcover), 1-4020-9791-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). xvi + 293 pp. LCCN Q175 .V33413 2009.

URL https://link.springer.com/
book/10.1007/978-1-4020-9791-1.

Vamvacas:2009:HEC

[Vam09h] Constantine J. Vamvacas. Heraclitus of Ephesus (ca. 540—480 B.C.). In The Founders of Western Thought — The Presocratics: A diachronic parallelism between Presocratic Thought and Philosophy and the Natural Sciences [Vam09g], chapter 9, pages 101–134. ISBN 1-4020-9790-5 (hardcover), 1-4020-9791-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175 .V33413 2009. URL https:/

/link.springer.com/chapter/10.
1007/978-1-4020-9791-1_9.

Vamvacas:2009:I

[Vam09i] Constantine J. Vamvacas. Introduction.

In The Founders of Western Thought — The Presocratics: A diachronic parallelism between Presocratic Thought and Philosophy and the Natural Sciences [Vam09g], chapter 1, pages 1–2. ISBN 1-4020-9790-5 (hardcover), 1-4020-9791-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175 .V33413 2009. URL https:/

/link.springer.com/chapter/10.
1007/978-1-4020-9791-1_1.

Vamvacas:2009:IP

[Vam09j] Constantine J. Vamvacas. Introduction to the presocratics. In The Founders of Western Thought — The Presocratics: A diachronic parallelism between Presocratic Thought and Philosophy and the Natural Sciences [Vam09g], chapter 3, pages 19–27. ISBN 1-4020-9790-5 (hardcover), 1-4020-9791-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175 .V33413

2009. URL https://link.springer.
com/chapter/10.1007/978-1-4020-
9791-1_3.

Vamvacas:2009:J

[Vam09k] Constantine J. Vamvacas. The juncture.

In The Founders of Western Thought — The Presocratics: A diachronic parallelism between Presocratic Thought and Philosophy and the Natural Sciences [Vam09g], chapter 2, pages 3–18. ISBN 1-4020-9790-5 (hardcover), 1-4020-9791-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175 .V33413 2009. URL https:/

/link.springer.com/chapter/10.
1007/978-1-4020-9791-1_2.

Vamvacas:2009:PEC

[Vam09l] Constantine J. Vamvacas. Parmenides of Elea (ca. 515—450 B.C.). In The Founders of Western Thought — The Presocratics: A diachronic parallelism between Presocratic Thought and Philosophy and the Natural Sciences [Vam09g], chapter 10, pages 135–165. ISBN 1-4020-9790-5 (hardcover), 1-4020-9791-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175 .V33413 2009. URL https:/

                                              /link.springer.com/chapter/10.
                                              1007/978-1-4020-9791-1_10.

Vamvacas:2009:PSC

[Vam09m] Constantine J. Vamvacas. Pythagoras of Samos (ca. 570—496 B.C.). In The Founders of Western Thought — The Presocratics: A diachronic parallelism between Presocratic Thought and Philosophy and the Natural Sciences [Vam09g], chapter 7, pages 53–84. ISBN 1-4020-9790-5 (hardcover), 1-4020-9791-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175 .V33413 2009. URL https:/

                                              /link.springer.com/chapter/10.
                                              1007/978-1-4020-9791-1_7.

Vamvacas:2009:TMC

[Vam09n] Constantine J. Vamvacas. Thales of Miletus (ca. 625—546 B.C.). In The Founders of Western Thought — The Presocratics: A diachronic parallelism between Presocratic Thought and Philosophy and the Natural Sciences [Vam09g], chapter 4, pages 29–33. ISBN 1-4020-9790-5 (hardcover), 1-4020-9791-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175 .V33413 2009. URL https:/

                                              /link.springer.com/chapter/10.
                                              1007/978-1-4020-9791-1_4.

Vamvacas:2009:XCC

[Vam09o] Constantine J. Vamvacas. Xenophanes of Colophon (ca. 570—470, B.C.). In The Founders of Western Thought — The Presocratics: A diachronic parallelism between Presocratic Thought and Philosophy and the Natural Sciences [Vam09g], chapter 8, pages 85–99. ISBN 1-4020-9790-5 (hardcover), 1-4020-9791-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175 .V33413 2009. URL https:/

/link.springer.com/chapter/10.
1007/978-1-4020-9791-1_8.

VanHeijenoort:1967:LCL

[Van67] Jean Van Heijenoort. Logic as calculus and logic as language. In Cohen and Wartofsky [CW67], pages 440–446. ISBN 90-277-0013-3, 94-010-3508-3. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175 .B731 v.3.

URL https://link.springer.com/
chapter/10.1007/978-94-010-3508-
8_26.

VanFraassen:1974:BSM

[Van74a] Bas C. Van Fraassen. Bressan and Suppes on modality. In Schaffner and Cohen [SC74a], pages 323–330. ISBN 90-277-0408-2, 90-277-0409-0, 94-010-2140-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175 .B7312. URL https:/

/link.springer.com/chapter/10.
1007/978-94-010-2140-1_21.

VanFraassen:1974:FRP

[Van74b] Bas C. Van Fraassen. The formal representation of physical quantities. In Logical and Epistemological Studies in Contemporary Physics [CW74a], chapter 5, pages 196–209. ISBN 90-277-0391-4 (hardcover), 90-277-0377-9 (paperback), 94-010-2656-4 (e-book).

ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol.

13; QC5.56. URL https://link.
springer.com/chapter/10.1007/
978-94-010-2656-7_5.

VanFraassen:1974:LQL

[Van74c] Bas C. Van Fraassen. The labyrinth of quantum logics. In Logical and Epistemological Studies in Contemporary Physics [CW74a], chapter 8, pages 224–254. ISBN 90-277-0391-4 (hardcover), 90-277-0377-9 (paperback), 94-010-2656-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 13; QC5.56.

URL https://link.springer.com/
chapter/10.1007/978-94-010-2656-
7_8.

VanSpronsen:1974:CSH

[Van74d] Johannes Willem Van Spronsen. Could the specific heat of the elements have contributed to the discovery of the periodic system? In For Dirk Struik: Scientific, Historical and Political Essays in Honor of Dirk J. Struik [CSW74], chapter 23, pages 275–292. ISBN 90-277-0379-5, 94-010-2115-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-010-2115-
9_23.

VanFraassen:1983:CFJ

[Van83] Bas C. Van Fraassen. Calibration: A frequency justification for personal probability. In Cohen and Laudan [CL83], chapter 15, pages 295–319. ISBN 94-009-7055-2, 94-009-7057-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-009-7055-
7_15.

VanZandt:1986:RES

[Van86] Joe D. Van Zandt. Res Extensa and the space–time continuum. In Spinoza and the Sciences [GN86], chapter 11, pages 249–266. ISBN 94-010-8511-0, 94-009-4514-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-009-
4514-2_11.

vandeVate:1991:NST

[van91] Dwight van de Vate, Jr. A new slant on the tower experiment. In Beyond Reason: Essays on the Philosophy of Paul Feyerabend [Mun91a], chapter 22, pages 449–452. ISBN 94-010-5406-1, 94-011-3188-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-011-
3188-9_22.

VanFraassen:1993:SPK

[Van93] Bas Van Fraassen. Symmetries of probability kinematics. In Bridging the Gap: Philosophy, Mathematics, and Physics: Lectures on the Foundations of Science [CCG93], chapter 14, pages 285–314. ISBN 94-010-5101-1, 94-011-2496-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-011-
2496-6_14.

Vanderveken:1995:NFL

[Van95] Daniel Vanderveken. A new formulation of the logic of propositions. In Québec Studies in the Philosophy of Science: Part I: Logic, Mathematics, Physics and History of Science Essays in Honor of Hugues Leblanc [MC95], chapter 7, pages 95–105. ISBN 94-010-7204-3, 94-009-1575-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

                                              ???? URL https://link.springer.
                                              com/chapter/10.1007/978-94-009-
                                              1575-6_7.

VanTongeren:1999:NSS

[Van99] Paul J. M. Van Tongeren. Nietzsche’s symptomatology of skepticism. In Nietzsche, Epistemology, and Philosophy of Science: Nietzsche and the Sciences II [Bab99a], chapter 5, pages 61–71. ISBN 90-481-5234-8, 94-017-2428-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-017-2428-9_5.

VanBrakel:2006:KLP

[Van06] J. Van Brakel. Kant’s legacy for the philosophy of chemistry. In Philosophy of Chemistry: Synthesis of a New Discipline [BSM06b], chapter 4, pages 69–91. ISBN 1-4020-2455-X, 1-4020-3256-0, 1-4020-3261-7. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 242 QD6. URL https:

                                              //link.springer.com/chapter/10.
                                              1007/1-4020-3261-7_4.

Vanpaemel:2008:MMP

[Van08] Geert Vanpaemel. Mechanics and mechanical philosophy in some Jesuit mathematical textbooks of the Early 17th Century. In Mechanics and Natural Philosophy Before the Scientific Revolution [LR08], chapter 11, pages 259–274. ISBN 1-4020-5966-3 (hardcover), 1-4020-5967-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QA802 .M43 2008.

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-1-4020-5967-
                                              4_11.

Vanpaemel:2015:GML

[Van15] Geert Vanpaemel. The German model of laboratory science and the European periphery (1860–1914). In Sciences in the Universities of Europe, Nineteenth and Twentieth Centuries: Academic Landscapes [SDG15b], chapter 13, pages 211–225. ISBN 94-017-9635-1, 94-017-9636-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q183.4.E85 2015. URL https:

//link.springer.com/chapter/10.
1007/978-94-017-9636-1_13.

Varela:1991:OMS

[Var91] Francisco J. Varela. Organism: A meshwork of selfless selves. In Organism and the Origins of Self [Tau91i], chapter 5, pages 79–107. ISBN 0-7923-1185-X, 94-011-3406-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QR180-189.5. URL https:/

/link.springer.com/chapter/10.
1007/978-94-011-3406-4_5.

Varela:1992:WPM

[Var92] Francisco J. Varela. Whence perceptual meaning? A cartography of current ideas. In Understanding Origins: Contemporary Views on the Origin of Life, Mind and Society [VD92], chapter 13, pages 235–263. ISBN 90-481-4090-0, 94-015-8054-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-015-8054-0_13.

Vasantha:1992:OOC

[Vas92] A. Vasantha. The “Oriental–Occidental Controversy” of 1839 and its impact on Indian science. In Science and Empires: Historical Studies about Scientific Development and European Expansion [PJM92], chapter 7, pages 49–56. ISBN 0-7923-1518-9 (hardcover), 94-010-5145-3, 94-011-2594-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 136.

URL https://link.springer.com/
chapter/10.1007/978-94-011-2594-
9_7.

Vazquez:1996:SP

[Váz96] Juan Vázquez. Scientific progress. In Munévar [Mun96], chapter 16, pages 333–343. ISBN 94-010-6622-1, 94-009-0305-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-009-
0305-0_16.

Varela:1992:UOC

[VD92] Francisco J. Varela and Jean-Pierre Dupuy. Understanding Origins: Contemporary Views on the Origin of Life, Mind and Society, volume 130 of Boston Studies in the Philosophy and History of Science. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 1992. ISBN 90-481-4090-0, 94-015-8054-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). viii + 329 pp. LCCN ????

URL https://link.springer.com/
book/10.1007/978-94-015-8054-0.

vanDelft:2014:CLH

[vD14] Dirk van Delft. The cryogenic laboratory of Heike Kamerlingh Onnes: An early case of big science. In History of Artificial Cold, Scientific, Technological and Cultural Issues [Gav14b], chapter 4, pages 65–81. ISBN 94-007-7198-3 (hardcover), 94-007-7199-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC278 H57 2014; TP494 .H577 2014. URL https:/

/link.springer.com/chapter/10.
1007/978-94-007-7199-4_4.

vanDelft:2015:PLL

[vD15] Dirk van Delft. The Physics Laboratory of Leiden University. In Sciences in the Universities of Europe, Nineteenth and Twentieth Centuries: Academic Landscapes [SDG15b], chapter 18, pages 311–326. ISBN 94-017-9635-1, 94-017-9636-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q183.4.E85 2015. URL https:

//link.springer.com/chapter/10.
1007/978-94-017-9636-1_18.

vanDongen:2017:EVV

[vD17] Jeroen van Dongen. The epistemic virtues of the virtuous theorist: On Albert Einstein and his autobiography. In van Dongen and Paul [vDP17a], chapter 5, pages 63–77. ISBN 3-319-48892-9 (hardcover), 3-319-48893-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 v. 321; Q175.32.V57 E65 2017.

URL https://link.springer.com/
chapter/10.1007/978-3-319-48893-
6_5.

vanderAvoird:1992:IFP

[vdA91] Ad van der Avoird. Intermolecular forces and the properties of molecular solids. In The Invention of Physical Science: Intersections of Mathematics, Theology and Natural Philosophy Since the Seventeenth Century: Essays in Honor of Erwin N. Hiebert [NRS92], chapter 10, pages 391–433. ISBN 94-010-5097-X, 94-011-2488-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-3-642-58183-0_10.

Videira:2011:CCD

[VdOM11] Antonio Videira and André L.

de O. Mendonça. Contextualizing the contexts of discovery and justification: How to do science studies in Brazil. In Brazilian Studies in Philosophy and History of Science: An account of recent works [KV11], chapter 17, pages 233–243. ISBN 90-481-9421-0 (hardcover), 90-481-9422-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175.3 .B73 2011.

URL https://link.springer.com/
chapter/10.1007/978-90-481-9422-
3_17.

vandePoel:2015:VET

[vdP15] Ibo van de Poel. Values in engineering and technology. In New Perspectives on Technology, Values, and Ethics: Theoretical and Practical [Gon15a], chapter 2, pages 29–46. ISBN 3-319-21869-7 (hardcover), 3-319-21870-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-3-319-21870-
0_2.

vanDongen:2017:EVS

[vDP17a] Jeroen van Dongen and Herman Paul, editors. Epistemic Virtues in the Sciences and the Humanities, volume 321 of Boston Studies in the Philosophy and History of Science. Springer International Publishing, Cham, Switzerland, 2017. ISBN 3-319-48892-9 (hardcover), 3-319-48893-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). vi + 198 pp. LCCN Q174 .B67 v. 321; Q175.32.V57 E65 2017. URL

https://link.springer.com/book/
10.1007/978-3-319-48893-6.

vanDongen:2017:IEV

[vDP17b] Jeroen van Dongen and Herman Paul. Introduction: Epistemic virtues in the sciences and the humanities. In Epistemic Virtues in the Sciences and the Humanities [vDP17a], chapter 1, pages 1–10. ISBN 3-319-48892-9 (hardcover), 3-319-48893-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 v. 321; Q175.32.V57 E65 2017. URL https:

//link.springer.com/chapter/10.
1007/978-3-319-48893-6_1.

vanderWaerden:2004:DSP

[vdW04] B. L. van der Waerden. Defence of a “shocking” point of view. In Classics in the History of Greek Mathematics [Chr04a], chapter 23, pages 433–445. ISBN 1-4020-0081-2 (hardcover), 1-4020-2640-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QA22 .C596 2004. URL https:/

/link.springer.com/chapter/10.
1007/978-1-4020-2640-9_23.

vanEss:1975:BIT

[vE75] Josef van Ess. The beginnings of Islamic theology. In Murdoch and Sylla [MS75b], pages 87–111. ISBN 90-277-0587-9, 94-010-1781-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN B721 .I584 1975.

URL https://link.springer.com/
chapter/10.1007/978-94-010-1781-
7_4.

Veltfort:2014:BSR

[Vel14] Theodore Veltfort. The beginning of semiconductor research in Cuba. In The History of Physics in Cuba [BRW14a], chapter 21, pages 353–356. ISBN 94-017-8040-4, 94-017-8041-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC5.53; QC6.9; QC9 C9H57 2014; Q124.6-127.2.

URL https://link.springer.com/
chapter/10.1007/978-94-017-8041-
4_21.

Vemulapalli:2006:PCC

[Vem06] G. K. Vemulapalli. Physics in the crucible of chemistry. In Philosophy of Chemistry: Synthesis of a New Discipline [BSM06b], chapter 10, pages 191–204. ISBN 1-4020-2455-X, 1-4020-3256-0, 1-4020-3261-7. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 242 QD6. URL https:

                                              //link.springer.com/chapter/10.
                                              1007/1-4020-3261-7_10.

Venturi:2016:FMR

[Ven16] Giorgio Venturi. Forcing, multiverse and realism. In Objectivity, Realism, and Proof: FilMat Studies in the Philosophy of Mathematics [BS16], chapter 12, pages 211–241. ISBN 3-319-31642-7, 3-319-31644-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QA9 .O25 2016.

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-3-319-31644-
                                              4_12.

Verdugo:2009:PTU

[Ver09] Carlos Verdugo. Popper’s thesis of the unity of scientific method: Method versus techniques. In Rethinking Popper [PC09], chapter 12, pages 155–160. ISBN 1-4020-9337-3, 1-4020-9338-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN B1649.P64 R48 2009; Q174 .B67 2009.

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-1-4020-9338-
                                              8_12.

vonFritz:2004:DIH

[vF04] Kurt von Fritz. The discovery of incommensurability by Hippasus of Metapontum. In Classics in the History of Greek Mathematics [Chr04a], chapter 11, pages 211–231. ISBN 1-4020-0081-2 (hardcover), 1-4020-2640-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QA22 .C596

2004. URL https://link.springer.
com/chapter/10.1007/978-1-4020-
2640-9_11.

Vizgin:1987:RTR

[VG87] V. P. Vizgin and G. E. Gorelik. The reception of the theory of relativity in Russia and the USSR. In Glick [Gli87a], chapter 8, pages 265–326. ISBN 90-277-2498-9, 94-009-3875-6. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 103; QC173.5. US$79.00 (US). URL https:

//link.springer.com/chapter/10.
1007/978-94-009-3875-5_8.

vonHelmholtz:1977:EWP

[vH77a] Hermann von Helmholtz, editor. Epistemological Writings: The Paul Hertz/ Moritz Schlick Centenary Edition of 1921 with Notes and Commentary by the Editors, volume 37 of Boston Studies in the Philosophy of Science. D. Reidel, Dordrecht, The Netherlands; Boston, MA, USA; Lancaster, UK; Tokyo, Japan, 1977. ISBN 90-277-0582-8, 94-010-1115-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL

https://link.springer.com/book/
10.1007/978-94-010-1115-0.

vonHelmholtz:1977:FP

[vH77b] Hermann von Helmholtz. The facts in perception. In Epistemological Writings: The Paul Hertz/Moritz Schlick Centenary Edition of 1921 with Notes and Commentary by the Editors [vH77a], pages 115–185. ISBN 90-277-0582-8, 94-010-1115-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-010-1115-0_4.

vonHelmholtz:1977:NME

[vH77c] Hermann von Helmholtz. Numbering and measuring from an epistemological viewpoint. In Epistemological Writings: The Paul Hertz/Moritz Schlick Centenary Edition of 1921 with Notes and Commentary by the Editors [vH77a], pages 72–114. ISBN 90-277-0582-8, 94-010-1115-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

                                              ???? URL https://link.springer.
                                              com/chapter/10.1007/978-94-010-
                                              1115-0_3.

vonHelmholtz:1977:FUG

[vH77d] Hermann von Helmholtz. On the facts underlying geometry. In Epistemological Writings: The Paul Hertz/Moritz Schlick Centenary Edition of 1921 with Notes and Commentary by the Editors [vH77a], pages 39–71. ISBN 90-277-0582-8, 94-010-1115-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-010-1115-0_2.

vonHelmholtz:1977:OSA

[vH77e] Hermann von Helmholtz. On the origin and significance of the axioms of geometry. In Epistemological Writings: The Paul Hertz/Moritz Schlick Centenary Edition of 1921 with Notes and Commentary by the Editors [vH77a], pages 1–38. ISBN 90-277-0582-8, 94-010-1115-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

                                              ???? URL https://link.springer.
                                              com/chapter/10.1007/978-94-010-
                                              1115-0_1.

vonHayek:2001:HBM

[vHBM01] Friedrich von Hayek, Gustav Bergmann, and Josef Mayerhöfer. Von Hayek, Bergmann, and Mayerhöfer. In Blackmore et al. [BIT01c], pages 123–138. ISBN 90-481-5752-8, 94-015-9690-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-015-9690-
9_6.

Vickers:1992:CRO

[Vic92] Brian Vickers. Critical reactions to the occult sciences during the Renaissance. In The Scientific Enterprise: The Bar-Hillel Colloquium: Studies in History, Philosophy, and Sociology of Science, Volume 4 [UM92], chapter 3, pages 43–92. ISBN 94-010-5190-9, 94-011-2688-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .S354 1992. URL https:

//link.springer.com/chapter/10.
1007/978-94-011-2688-5_3.

Vihalemm:2001:CIS

[Vih01a] Rein Vihalemm. Chemistry as an interesting subject for the philosophy of science. In Estonian Studies in the History and Philosophy of Science [Vih01b], chapter 14, pages 185–200. ISBN 94-010-3863-5, 94-010-0672-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-010-0672-
9_14.

Vihalemm:2001:ESH

[Vih01b] Rein Vihalemm. Estonian Studies in the History and Philosophy of Science, volume 219 of Boston Studies in the Philosophy of Science. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 2001. ISBN 94-010-3863-5, 94-010-0672-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). xiv + 306 pp. LCCN ???? URL

                                              https://link.springer.com/book/
                                              10.1007/978-94-010-0672-9.

Villemaire:2002:BNP

[Vil02a] Diane Davis Villemaire. Burtt’s new philosophy of mind. In E. A. Burtt, Historian and Philosopher: A Study of the Author of The Metaphysical Foundations of Modern Physical Science [Vil02g], chapter 8, pages 95–129. ISBN 90-481-5937-7, 94-017-1331-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-017-1331-3_8.

Villemaire:2002:BND

[Vil02b] Diane Davis Villemaire. Burtt’s Newton and the debate over the rationality of early modern science. In E. A. Burtt, Historian and Philosopher: A Study of the Author of The Metaphysical Foundations of Modern Physical Science [Vil02g], chapter 10, pages 155–186. ISBN 90-481-5937-7, 94-017-1331-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-017-1331-
                                              3_10.

Villemaire:2002:BPA

[Vil02c] Diane Davis Villemaire. Burtt’s philosophy after 1935. In E. A. Burtt, Historian and Philosopher: A Study of the Author of The Metaphysical Foundations of Modern Physical Science [Vil02g], chapter 12, pages 197–213. ISBN 90-481-5937-7, 94-017-1331-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-017-1331-
                                              3_12.

Villemaire:2002:CSP

[Vil02d] Diane Davis Villemaire. The Chicago School of Pragmatism. In E. A. Burtt, Historian and Philosopher: A Study of the Author of The Metaphysical Foundations of Modern Physical Science [Vil02g], chapter 6, pages 67–81. ISBN 90-481-5937-7, 94-017-1331-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-017-1331-3_6.

Villemaire:2002:CUY

[Vil02e] Diane Davis Villemaire. Columbia University in the 1920s: The young radical philosophers. In E. A. Burtt, Historian and Philosopher: A Study of the Author of The Metaphysical Foundations of Modern Physical Science [Vil02g], chapter 2, pages 7–14. ISBN 90-481-5937-7, 94-017-1331-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-017-1331-3_2.

Villemaire:2002:C

[Vil02f] Diane Davis Villemaire. Conclusion.

In E. A. Burtt, Historian and Philosopher: A Study of the Author of The Metaphysical Foundations of Modern Physical Science [Vil02g], chapter 14, pages 233–245. ISBN 90-481-5937-7, 94-017-1331-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-017-
1331-3_14.

Villemaire:2002:BHP

[Vil02g] Diane Davis Villemaire. E. A. Burtt, Historian and Philosopher: A Study of the Author of The Metaphysical Foundations of Modern Physical Science, volume 226 of Boston Studies in the Philosophy of Science. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 2002. ISBN 90-481-5937-7, 94-017-1331-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). x + 299 pp. LCCN ???? URL

https://link.springer.com/book/
10.1007/978-94-017-1331-3.

Villemaire:2002:GHM

[Vil02h] Diane Davis Villemaire. FROM the genesis of hypothesis (1935) TO the metaphysics of empirical theology (1987?). In E. A. Burtt, Historian and Philosopher: A Study of the Author of The Metaphysical Foundations of Modern Physical Science [Vil02g], chapter 9, pages 131–153. ISBN 90-481-5937-7, 94-017-1331-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-017-1331-3_9.

Villemaire:2002:IJH

[Vil02i] Diane Davis Villemaire. The idealists: James Harvey Robinson and the genetic method of history, John Dewey and reflective thinking. In E. A. Burtt, Historian and Philosopher: A Study of the Author of The Metaphysical Foundations of Modern Physical Science [Vil02g], chapter 3, pages 15–29. ISBN 90-481-5937-7, 94-017-1331-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-017-1331-3_3.

Villemaire:2002:I

[Vil02j] Diane Davis Villemaire. Introduction. In E. A. Burtt, Historian and Philosopher: A Study of the Author of The Metaphysical Foundations of Modern Physical Science [Vil02g], chapter 1, pages 1–6. ISBN 90-481-5937-7, 94-017-1331-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-017-
1331-3_1.

Villemaire:2002:MFM

[Vil02k] Diane Davis Villemaire. The metaphysical foundations of modern physical science. In E. A. Burtt, Historian and Philosopher: A Study of the Author of The Metaphysical Foundations of Modern Physical Science [Vil02g], chapter 5, pages 47–65. ISBN 90-481-5937-7, 94-017-1331-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-017-
1331-3_5.

Villemaire:2002:RH

[Vil02l] Diane Davis Villemaire. Religious humanism. In E. A. Burtt, Historian and Philosopher: A Study of the Author of The Metaphysical Foundations of Modern Physical Science [Vil02g], chapter 7, pages 83–93. ISBN 90-481-5937-7, 94-017-1331-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-017-
1331-3_7.

Villemaire:2002:WHM

[Vil02m] Diane Davis Villemaire. What happened to the metaphysical foundations and Burtt’s interest in the history of science? In E. A. Burtt, Historian and Philosopher: A Study of the Author of The Metaphysical Foundations of Modern Physical Science [Vil02g], chapter 11, pages 187–196. ISBN 90-481-5937-7, 94-017-1331-6 (e-book).

ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-017-1331-3_11.

Vilain:2004:CHG

[Vil04] Christiane Vilain. Christiaan Huygens’ Galilean mechanics. In The Reception of the Galilean Science of Motion in Seventeenth-Century Europe [PT04], chapter 10, pages 185–198. ISBN 1-4020-2454-1, 1-4020-2455-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC133 .R43 2004.

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-1-4020-2455-
                                              9_10.

Vilain:2008:CRM

[Vil08] Christiane Vilain. Circular and rectilinear motion in the mechanica and in the 16th century. In Mechanics and Natural Philosophy Before the Scientific Revolution [LR08], chapter 5, pages 149–172. ISBN 1-4020-5966-3 (hardcover), 1-4020-5967-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QA802 .M43 2008.

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-1-4020-5967-
                                              4_5.

Vincenti:1992:EKT

[Vin92]  Walter G. Vincenti.  Engineering

knowledge, type of design, and level of hierarchy: Further thoughts about what engineers know . . . . . . . In Technological Development and Science in the Industrial Age: New Perspectives on the Science–Technology Relationship [KB92b], chapter 2, pages 17–34. ISBN 90-481-4186-9, 94-015-8010-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-015-8010-6_2.

Visser:1992:MUD

[Vis92] Henk Visser. Mach, Utrecht, and Dutch philosophy. In Ernst Mach — A Deeper Look: Documents and New Perspectives [Bla92a], chapter 20, pages 403–430. ISBN 94-010-5230-1, 94-011-2771-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN C5.53; QC6.9. URL https:

//link.springer.com/chapter/10.
1007/978-94-011-2771-4_20.

Visser:2001:WMS

[Vis01] Henk Visser. Wittgenstein’s Machist sources. In Blackmore et al. [BIT01c], pages 139–158. ISBN 90-481-5752-8, 94-015-9690-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-015-
9690-9_7.

Vishveshwara:2003:RRD

[Vis03] C. V. Vishveshwara. Rigidly rotating disk revisited. In Renn et al. [RDS03], chapter 13, pages 305–316. ISBN 1-4020-1284-5, 1-4020-1285-3 (paperback). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 234. URL https://link.

springer.com/chapter/10.1007/
978-94-010-0111-3_13.

VanDenDaele:1983:SCL

[VK83] Wolfgang Van Den Daele and Wolfgang Krohn. Science in a crisis of legitimation. In Finalization in Science: The Social Orientation of Scientific Progress [Sch83b], chapter 11, pages 237–250. ISBN 94-009-7082-X, 94-009-7080-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-009-7080-
9_11.

vanLeeuwen:2016:AMT

[vL16a] Joyce van Leeuwen. The Aristotelian Mechanics: Text and Diagrams, volume 316 of Boston Studies in the Philosophy and History of Science. Springer-Ver-lag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2016. ISBN 3-319-25923-7 (hardcover), 3-319-25925-3 (electronic). ISSN 0068-0346 (print), 2214-7942 (electronic). ix + 253 pp. LCCN QA805 .L44 2016. URL

                                              https://link.springer.com/book/
                                              10.1007/978-3-319-25925-3.

vanLeeuwen:2016:C

[vL16b] Joyce van Leeuwen. Conclusion. In The Aristotelian Mechanics: Text and Diagrams [vL16a], chapter 7, pages 189–191. ISBN 3-319-25923-7 (hardcover), 3-319-25925-3 (electronic). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QA805 .L44 2016.

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-3-319-25925-
                                              3_7.

vanLeeuwen:2016:DR

[vL16c] Joyce van Leeuwen. Diagrammatic reasoning. In The Aristotelian Mechanics: Text and Diagrams [vL16a], chapter 5, pages 139–157. ISBN 3-319-25923-7 (hardcover), 3-319-25925-3 (electronic). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QA805 .L44 2016.

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-3-319-25925-
                                              3_5.

vanLeeuwen:2016:EMT

[vL16d] Joyce van Leeuwen. Early modern transformations of mechanics. In The Aristotelian Mechanics: Text and Diagrams [vL16a], chapter 6, pages 159–188. ISBN 3-319-25923-7 (hardcover), 3-319-25925-3 (electronic).

ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QA805 .L44 2016.

URL https://link.springer.com/
chapter/10.1007/978-3-319-25925-
3_6.

vanLeeuwen:2016:ED

[vL16e] Joyce van Leeuwen. Editing diagrams.

In The Aristotelian Mechanics: Text and Diagrams [vL16a], chapter 4, pages 73–137. ISBN 3-319-25923-7 (hardcover), 3-319-25925-3 (electronic). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QA805 .L44 2016.

URL https://link.springer.com/
chapter/10.1007/978-3-319-25925-
3_4.

vanLeeuwen:2016:GCT

[vL16f] Joyce van Leeuwen. General characteristics of the transmission. In The Aristotelian Mechanics: Text and Diagrams [vL16a], chapter 2, pages 25– 35. ISBN 3-319-25923-7 (hardcover), 3-319-25925-3 (electronic). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QA805 .L44 2016. URL https:

//link.springer.com/chapter/10.
1007/978-3-319-25925-3_2.

vanLeeuwen:2016:I

[vL16g] Joyce van Leeuwen. Introduction. In The Aristotelian Mechanics: Text and Diagrams [vL16a], chapter 1, pages 1–24. ISBN 3-319-25923-7 (hardcover), 3-319-25925-3 (electronic). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QA805 .L44 2016. URL https:

//link.springer.com/chapter/10.
1007/978-3-319-25925-3_1.

vanLeeuwen:2016:TT

[vL16h] Joyce van Leeuwen. The textual transmission. In The Aristotelian Mechanics: Text and Diagrams [vL16a], chapter 3, pages 37–72. ISBN 3-319-25923-7 (hardcover), 3-319-25925-3 (electronic). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QA805 .L44 2016.

URL https://link.springer.com/
chapter/10.1007/978-3-319-25925-
3_3.

Vlastos:1990:ROP

[Vla90] Gregory Vlastos. The role of observation in Plato’s conception of astronomy. In Greek Studies in the Philosophy and History of Science [Nic90], chapter 1, pages 1–27. ISBN 94-010-7403-8, 94-009-2015-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-009-
2015-6_1.

Vlahakis:1994:PME

[Vla94] Yiorgos N. Vlahakis. Problems and methodology of exploring the scientific thought during the Greek Enlightenment (1750–1821). In Trends in the Historiography of Science [GCN94], chapter 29, pages 397–404. ISBN 90-481-4264-4, 94-017-3596-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-017-3596-4_29.

vonMolnar:1989:WEH

[vM89] Géza von Molnár. “What ever happened to ethics?”. In Literature and Science as Modes of Expression [Amr89], chapter 6, pages 113–127. ISBN 94-010-7531-X, 94-009-2297-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-009-2297-
6_6.

Verpooten:2012:SEU

[VN12] Jan Verpooten and Mark Nelissen.

Sensory exploitation: Underestimated in the evolution of art as once in sexual selection theory? In Philosophy of Behavioral Biology [PR12b], chapter 9, pages 189–216. ISBN 94-007-1950-7, 94-007-1951-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QH331 .P4679 2012; Q175 .B657

v.282. URL https://link.springer.
com/chapter/10.1007/978-94-007-
1951-4_9.

Vodichev:2015:URC

[Vod15] Evgeny Vodichev. Universities in Russia: Current reforms through the prism of Soviet heritage and international practice. In Sciences in the Universities of Europe, Nineteenth and Twentieth Centuries: Academic Landscapes [SDG15b], chapter 11, pages 175–190. ISBN 94-017-9635-1, 94-017-9636-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q183.4.E85

2015. URL https://link.springer.
com/chapter/10.1007/978-94-017-
9636-1_11.

Vogt:1991:SGS

[Vog91] Annette Vogt. Soviet–German scientific relations before World War II: Fruitful cooperation in different social orders. In Woodward and Cohen [WC91], pages 387–399. ISBN 94-010-5395-2, 94-011-3164-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-011-
3164-3_36.

Vogel:2004:BQG

[Vog04] Kurt Vogel. Zur Berechnung der Quadratischen Gleichungen bei den Babyloniern. (German) [On calculating the quadratic equations of the Babylonians]. In Classics in the History of Greek Mathematics [Chr04a], chapter 15, pages 265–273. ISBN 1-4020-0081-2 (hardcover), 1-4020-2640-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QA22 .C596

2004. URL https://link.springer.
com/chapter/10.1007/978-1-4020-
2640-9_15.

Voizard:1996:ICW

[Voi96] Alain Voizard. “If cows had wings, we’d carry big umbrellas.” an almost number-free note on Newcomb’s problem. In Québec Studies in the Philosophy of Science: Part II: Biology, Psychology, Cognitive Science and Economics Essays in Honor of Hugues Leblanc [MC96], chapter 12, pages 193–213. ISBN 94-010-6537-3, 94-009-0113-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-009-
0113-1_12.

Vollmer:2006:SMR

[Vol06] S. H. Vollmer. Space in molecular representation; or how pictures represent objects. In Philosophy of Chemistry: Synthesis of a New Discipline [BSM06b], chapter 15, pages 293–308. ISBN 1-4020-2455-X, 1-4020-3256-0, 1-4020-3261-7. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 242 QD6. URL https:/

/link.springer.com/chapter/10.
1007/1-4020-3261-7_15.

Volkov:2014:DDM

[Vol14] Alexei Volkov. Didactical dimensions of mathematical problems: Weighted distribution in a Vietnamese mathematical treatise. In Bernard and Proust [BP14b], chapter 10, pages 247–271. ISBN 94-007-5121-4 (hardcover), 94-007-5122-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q181 .S35 2014; Q181 S387 2014.

URL https://link.springer.com/
chapter/10.1007/978-94-007-5122-
4_10.

VonKutschera:1976:EIC

[Von76]  Franz Von Kutschera.  Epistemic

interpretation of conditionals. In Language in Focus: Foundations, Methods and Systems: Essays in Memory of Yehoshua Bar-Hillel [Kas76b], chapter 21, pages 487–501. ISBN 90-277-0644-1, 90-277-0645-X, 94-010-1876-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN P101-120.

URL https://link.springer.com/
chapter/10.1007/978-94-010-1876-
0_21.

VonEngelhardt:1984:CSS

[Von84a] Dietrich Von Engelhardt. The chemical system of substances, forces and processes in Hegel’s philosophy of nature and the science of his time. In Hegel and the Sciences [CW84a], chapter 4, pages 41–54. ISBN 94-009-6235-5, 94-009-6233-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-009-
6233-0_4.

VonEngelhardt:1984:HPU

[Von84b] Dietrich Von Engelhardt. Hegel’s philosophical understanding of illness. In Hegel and the Sciences [CW84a], chapter 9, pages 123–141. ISBN 94-009-6235-5, 94-009-6233-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-009-6233-0_9.

VonBrentano:1991:LAL

[Von91] Margherita Von Brentano. Letter to an anti-liberal liberal. In Beyond Reason: Essays on the Philosophy of Paul Feyerabend [Mun91a], chapter 8, pages 199–212. ISBN 94-010-5406-1, 94-011-3188-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

                                              ???? URL https://link.springer.
                                              com/chapter/10.1007/978-94-011-
                                              3188-9_8.

VonWeizsacker:1999:NPM

[Von99] Carl Friedrich Von Weizsäcker. Nietzsche: Perceptions of modernity. In Nietzsche, Epistemology, and Philosophy of Science: Nietzsche and the Sciences II [Bab99a], chapter 18, pages 221–240. ISBN 90-481-5234-8, 94-017-2428-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

                                              ???? URL https://link.springer.
                                              com/chapter/10.1007/978-94-017-
                                              2428-9_18.

Votsis:2011:SRC

[Vot11] Ioannis Votsis. Structural realism: Continuity and its limits. In Bokulich and Bokulich [BB11], chapter 6, pages 105–117. ISBN 90-481-9596-9, 90-481-9597-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175.32.S78 S35 2011. URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-90-481-9597-8_6.

Vranicki:1979:TFI

[Vra79] Predrag Vranicki. Theoretical foundations for the idea of self-management. In Praxis: Yugoslav Essays in the Philosophy and Methodology of the Social Sciences [MP79], chapter 15, pages 229–247. ISBN 90-277-0968-8, 94-009-9355-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-009-
9355-6_15.

Vrcan:1979:SEI

[Vrc79] Srdjan Vrcan. Social equality and inequality in the bourgeois world and in socialism. In Praxis: Yugoslav Essays in the Philosophy and Methodology of the Social Sciences [MP79], chapter 20, pages 317–332. ISBN 90-277-0968-8, 94-009-9355-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-009-
9355-6_20.

Vromen:1990:GSC

[Vro90]  Suzanne Vromen.  Georg Simmel

and the cultural dilemma of women. In Georg Simmel and Contemporary Sociology [KPC90], chapter 18, pages 319–340. ISBN 94-010-6691-4 (hardcover), 94-009-0459-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN HM401-1281.

URL https://link.springer.com/
chapter/10.1007/978-94-009-0459-
0_18.

vanSanten:1992:OIC

[vSB91] R. A. van Santen and E. J. Baerends.

Orbital interactions and chemical reactivity of metal particles and metal surfaces. In The Invention of Physical Science: Intersections of Mathematics, Theology and Natural Philosophy Since the Seventeenth Century: Essays in Honor of Erwin N. Hiebert [NRS92], chapter 9, pages 323–389. ISBN 94-010-5097-X, 94-011-2488-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-3-642-58183-0_9.

vonWeizsacker:1969:UP

[vW69] Carl-Friedrich von Weizsäcker. The unity of physics. In Cohen and Wartofsky [CW69a], pages 460–473. ISBN 94-010-3383-8, 94-010-3263-7, 94-010-3381-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

                                              ???? URL https://link.springer.
                                              com/chapter/10.1007/978-94-010-
                                              3381-7_17.

Weizsacker:1984:UN

[vW84] C. F. v. Weizsäcker. The unity of nature. In Physical Sciences and History of Physics [CW84c], chapter 13, pages 239–254. ISBN 94-009-7180-X, 94-009-7178-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .P497 1984. URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-009-7178-3_13.

vonWeizsacker:1987:GMS

[vW87] Carl Friedrich von Weizsäcker. Goethe and modern science. In Amrine et al. [AZW87], chapter 7, pages 115–132. ISBN 90-277-2265-X, 94-009-3761-X. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol.

97. US$79.00. URL https://link.

                                              springer.com/chapter/10.1007/
                                              978-94-009-3761-1_7.

vonWeizsacker:1994:KTN

[vW94] Carl Friedrich von Weizsäcker. Kant’s theory of natural science according to P. Plaass. In Kant’s Theory of Natural Science [Pla94b], chapter 2, pages 165–187. ISBN 94-010-4492-9, 94-011-1126-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

                                              ???? URL https://link.springer.
                                              com/chapter/10.1007/978-94-011-
                                              1126-3_2.

vanZantwijk:2013:DDW

[vZ11] Temilo van Zantwijk. “Demonstrations”, not “deductions”: Walter Dubislav on transcendental arguments. In The Berlin Group and the Philosophy of Logical Empiricism [MP13], chapter 9, pages 191–204. ISBN 94-007-5484-1 (hardcover), 94-007-5485-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175.32.A24 .M384 2011; Q175 .B47

2013. URL https://link.springer.
com/chapter/10.1007/978-94-007-
5485-0_9.

Wagner:2015:ROE

[Wag15] Günter P. Wagner. Reinventing the organism: Evolvability and homology in post-Dahlem evolutionary biology. In Conceptual Change in Biology: Scientific and Philosophical Perspectives on Evolution and Development [Lov15b], chapter 15, pages 327–342. ISBN 94-017-9411-1, 94-017-9412-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QH331 C66 2015; QH491; QH492 .C66 2014.

URL https://link.springer.com/
chapter/10.1007/978-94-017-9412-
1_15.

Waisman:2014:ALM

[Wai14] Dina Waisman. Andrea Levialdi in memoriam. In The History of Physics in Cuba [BRW14a], chapter 22, pages 357–359. ISBN 94-017-8040-4, 94-017-8041-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC5.53; QC6.9; QC9 C9H57 2014; Q124.6-127.2. URL https://link.

springer.com/chapter/10.1007/
978-94-017-8041-4_22.

Wake:2015:HMT

[Wak15a] David B. Wake. Homoplasy, a moving target. In Conceptual Change in Biology: Scientific and Philosophical Perspectives on Evolution and Development [Lov15b], chapter 5, pages 111–127. ISBN 94-017-9411-1, 94-017-9412-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QH331 C66 2015; QH491; QH492 .C66

2014. URL https://link.springer.
com/chapter/10.1007/978-94-017-
9412-1_5.

Wake:2015:HIE

[Wak15b] Marvalee H. Wake. Hierarchies and integration in evolution and development. In Conceptual Change in Biology: Scientific and Philosophical Perspectives on Evolution and Development [Lov15b], chapter 18, pages 405–420. ISBN 94-017-9411-1, 94-017-9412-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QH331 C66 2015; QH491; QH492 .C66 2014.

URL https://link.springer.com/
chapter/10.1007/978-94-017-9412-
1_18.

Wallace:1967:ERP

[Wal67] William A. Wallace. Elementarity and reality in particle physics. In Cohen and Wartofsky [CW67], pages 236–263. ISBN 90-277-0013-3, 94-010-3508-3. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175 .B731 v.3.

URL https://link.springer.com/
chapter/10.1007/978-94-010-3508-
8_13.

Wallace:1976:GRE

[Wal76] William A. Wallace. Galileo and reasoning ex suppositione: the methodology of the two new sciences. In Cohen et al. [CHMV76], pages 79–104. ISBN 90-277-0647-6, 90-277-0648-4 (paperback), 94-010-1449-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-010-1449-6_4.

Wallace:1981:AMG

[Wal81a] William A. Wallace. Anneliese Maier:

Galileo and theories of impetus. In Prelude to Galileo: Essays on Medieval and Sixteenth-Century Sources of Galileo’s Thought [Wal81q], chapter 15, pages 320–340. ISBN 90-277-1215-8, 90-277-1216-6 (paperback), 94-009-8404-9. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 62.

URL https://link.springer.com/
chapter/10.1007/978-94-009-8404-
2_15.

Wallace:1981:CSC

[Wal81b] William A. Wallace. The calculatores in the Sixteenth Century. In Prelude to Galileo: Essays on Medieval and Sixteenth-Century Sources of Galileo’s Thought [Wal81q], chapter 5, pages 78–90. ISBN 90-277-1215-8, 90-277-1216-6 (paperback), 94-009-8404-9. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 62.

URL https://link.springer.com/
chapter/10.1007/978-94-009-8404-
2_5.

Wallace:1981:CFC

[Wal81c] William A. Wallace. Causes and forces at the Collegio Romano. In Prelude to Galileo: Essays on Medieval and Sixteenth-Century Sources of Galileo’s Thought [Wal81q], chapter 7, pages 110–126. ISBN 90-277-1215-8, 90-277-1216-6 (paperback), 94-009-8404-9. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 62.

URL https://link.springer.com/
chapter/10.1007/978-94-009-8404-
2_7.

Wallace:1981:CMS

[Wal81d] William A. Wallace. The concept of motion in the Sixteenth Century. In Prelude to Galileo: Essays on Medieval and Sixteenth-Century Sources of Galileo’s Thought [Wal81q], chapter 4, pages 64–77. ISBN 90-277-1215-8, 90-277-1216-6 (paperback), 94-009-8404-9. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 62.

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-009-8404-
                                              2_4.

Wallace:1981:DMS

[Wal81e] William A. Wallace. The development of mechanics to the Sixteenth Century. In Prelude to Galileo: Essays on Medieval and Sixteenth-Century Sources of Galileo’s Thought [Wal81q], chapter 3, pages 51–63. ISBN 90-277-1215-8, 90-277-1216-6 (paperback), 94-009-8404-9. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 62.

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-009-8404-
                                              2_3.

Wallace:1981:EDS

[Wal81f] William A. Wallace. The enigma of Domingo De Soto. In Prelude to Galileo: Essays on Medieval and Sixteenth-Century Sources of Galileo’s Thought [Wal81q], chapter 6, pages 91–109. ISBN 90-277-1215-8, 90-277-1216-6 (paperback), 94-009-8404-9. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 62.

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-009-8404-
                                              2_6.

Wallace:1981:EMG

[Wal81g] William A. Wallace. Ernest Moody: Galileo and nominalism. In Prelude to Galileo: Essays on Medieval and Sixteenth-Century Sources of Galileo’s Thought [Wal81q], chapter 16, pages 341–348. ISBN 90-277-1215-8, 90-277-1216-6 (paperback), 94-009-8404-9. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 62.

URL https://link.springer.com/
chapter/10.1007/978-94-009-8404-
2_16.

Wallace:1981:GAM

[Wal81h] William A. Wallace. Galileo and Albertus Magnus. In Prelude to Galileo: Essays on Medieval and Sixteenth-Century Sources of Galileo’s Thought [Wal81q], chapter 12, pages 264–285. ISBN 90-277-1215-8, 90-277-1216-6 (paperback), 94-009-8404-9. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 62.

URL https://link.springer.com/
chapter/10.1007/978-94-009-8404-
2_12.

Wallace:1981:GRE

[Wal81i] William A. Wallace. Galileo and reasoning Ex Suppositione. In Prelude to Galileo: Essays on Medieval and Sixteenth-Century Sources of Galileo’s Thought [Wal81q], chapter 8, pages 129–159. ISBN 90-277-1215-8, 90-277-1216-6 (paperback), 94-009-8404-9. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 62.

URL https://link.springer.com/
chapter/10.1007/978-94-009-8404-
2_8.

Wallace:1981:GCN

[Wal81j] William A. Wallace. Galileo and the causality of nature. In Prelude to Galileo: Essays on Medieval and Sixteenth-Century Sources of Galileo’s Thought [Wal81q], chapter 13, pages 286–299. ISBN 90-277-1215-8, 90-277-1216-6 (paperback), 94-009-8404-9. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 62.

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-009-8404-
                                              2_13.

Wallace:1981:GDP

[Wal81k] William A. Wallace. Galileo and the Doctores Parisienses. In Prelude to Galileo: Essays on Medieval and Sixteenth-Century Sources of Galileo’s Thought [Wal81q], chapter 10, pages 192–252. ISBN 90-277-1215-8, 90-277-1216-6 (paperback), 94-009-8404-9. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 62.

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-009-8404-
                                              2_10.

Wallace:1981:GS

[Wal81l] William A. Wallace. Galileo and the Scotists. In Prelude to Galileo: Essays on Medieval and Sixteenth-Century Sources of Galileo’s Thought [Wal81q], chapter 11, pages 253–263. ISBN 90-277-1215-8, 90-277-1216-6 (paperback), 94-009-8404-9. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 62. URL https://link.

                                              springer.com/chapter/10.1007/
                                              978-94-009-8404-2_11.

Wallace:1981:GT

[Wal81m] William A. Wallace. Galileo and the Thomists. In Prelude to Galileo: Essays on Medieval and Sixteenth-Century Sources of Galileo’s Thought [Wal81q], chapter 9, pages 160–191. ISBN 90-277-1215-8, 90-277-1216-6 (paperback), 94-009-8404-9. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 62. URL https://link.

springer.com/chapter/10.1007/
978-94-009-8404-2_9.

Wallace:1981:MAM

[Wal81n] William A. Wallace. The medieval accomplishment in mechanics and optics. In Prelude to Galileo: Essays on Medieval and Sixteenth-Century Sources of Galileo’s Thought [Wal81q], chapter 2, pages 29–48. ISBN 90-277-1215-8, 90-277-1216-6 (paperback), 94-009-8404-9. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 62. URL https://link.

springer.com/chapter/10.1007/
978-94-009-8404-2_2.

Wallace:1981:PSM

[Wal81o] William A. Wallace. The philosophical setting of medieval science. In Prelude to Galileo: Essays on Medieval and Sixteenth-Century Sources of Galileo’s Thought [Wal81q], chapter 1, pages 3–28. ISBN 90-277-1215-8, 90-277-1216-6 (paperback), 94-009-8404-9. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 62.

URL https://link.springer.com/
chapter/10.1007/978-94-009-8404-
2_1.

Wallace:1981:PDG

[Wal81p] William A. Wallace. Pierre Duhem:

Galileo and the science of motion. In Prelude to Galileo: Essays on Medieval and Sixteenth-Century Sources of Galileo’s Thought [Wal81q], chapter 14, pages 303–319. ISBN 90-277-1215-8, 90-277-1216-6 (paperback), 94-009-8404-9. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 62.

URL https://link.springer.com/
chapter/10.1007/978-94-009-8404-
2_14.

Wallace:1981:PGE

[Wal81q] William A. Wallace. Prelude to Galileo: Essays on Medieval and Sixteenth-Century Sources of Galileo’s Thought, volume 62 of Boston Studies in the Philosophy of Science. D. Reidel, Dordrecht, The Netherlands; Boston, MA, USA; Lancaster, UK; Tokyo, Japan, 1981. ISBN 90-277-1215-8, 90-277-1216-6 (paperback), 94-009-8404-9. ISSN 0068-0346 (print), 2214-7942 (electronic). xvi + 369 pp. LCCN Q174 .B67 vol. 62. URL

                                              https://link.springer.com/book/
                                              10.1007/978-94-009-8404-2.

Walter:1984:NTC

[Wal84] E. V. Walter. Nature on trial: The case of the rooster that laid an egg. In Methodology, Metaphysics and the History of Science: In Memory of Benjamin Nelson [CW84b], chapter 13, pages 295–321. ISBN 94-009-6333-5, 94-009-6331-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

                                              ???? URL https://link.springer.
                                              com/chapter/10.1007/978-94-009-
                                              6331-3_13.

Wallace:1990:DSG

[Wal90]  William A. Wallace.  The dating

and significance of Galileo’s Pisan manuscripts. In Levere and Shea [LS90], chapter 1, pages 3–50. ISBN 0-7923-0420-9, 94-009-1878-X. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 120.

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-009-1878-
                                              8_1.

Wallace:1992:DRM

[Wal92a] William A. Wallace. Demonstration and its requirements in MS 27. In Galileo’s logic of discovery and proof: the background, content, and use of his appropriated treatises on Aristotle’s Posterior analytics [Wal92c], chapter 4, pages 134–190. ISBN 0-7923-1577-4, 0-7923-1579-0 (set), 94-015-8040-5. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 137.

URL https://link.springer.com/
chapter/10.1007/978-94-015-8040-
3_4.

Wallace:1992:E

[Wal92b] William A. Wallace. Epilogue. In Galileo’s logic of discovery and proof: the background, content, and use of his appropriated treatises on Aristotle’s Posterior analytics [Wal92c], chapter 7, pages 299–303. ISBN 0-7923-1577-4, 0-7923-1579-0 (set), 94-015-8040-5. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 137.

URL https://link.springer.com/
chapter/10.1007/978-94-015-8040-
3_7.

Wallace:1992:GLD

[Wal92c] William A. Wallace. Galileo’s logic of discovery and proof: the background, content, and use of his appropriated treatises on Aristotle’s Posterior analytics, volume 137 of Boston Studies in the Philosophy of Science. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 1992. ISBN 0-7923-1577-4, 0-7923-1579-0 (set), 94-015-8040-5. ISSN 0068-0346 (print), 2214-7942 (electronic). xxiii + 323 pp. LCCN Q174 .B67 vol. 137.

URL https://link.springer.com/
book/10.1007/978-94-015-8040-3.

Wallace:1992:GLM

[Wal92d] William A. Wallace. Galileo’s logical methodology. In Galileo’s logic of discovery and proof: the background, content, and use of his appropriated treatises on Aristotle’s Posterior analytics [Wal92c], chapter 1, pages 1–29. ISBN 0-7923-1577-4, 0-7923-1579-0 (set), 94-015-8040-5. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 137. URL https:/

                                              /link.springer.com/chapter/10.
                                              1007/978-94-015-8040-3_1.

Wallace:1992:GLT

[Wal92e] William A. Wallace. Galileo’s Logical Treatises: A Translation, with Notes and Commentary, of His Appropriated Latin Questions on Aristotle’s Posterior Analytics, volume 138 of Boston Studies in the Philosophy of Science. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 1992. ISBN 90-481-4116-8, 94-015-8036-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). xix + 249 pp. LCCN ????

                                              URL https://link.springer.com/
                                              book/10.1007/978-94-015-8036-6.

Wallace:1992:GNS

[Wal92f] William A. Wallace. Galileo’s new sciences of mechanics and local motion. In Galileo’s logic of discovery and proof: the background, content, and use of his appropriated treatises on Aristotle’s Posterior analytics [Wal92c], chapter 6, pages 238–298. ISBN 0-7923-1577-4, 0-7923-1579-0 (set), 94-015-8040-5. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 137.

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-015-8040-
                                              3_6.

Wallace:1992:GSN

[Wal92g] William A. Wallace. Galileo’s search for a new science of the heavens. In Galileo’s logic of discovery and proof: the background, content, and use of his appropriated treatises on Aristotle’s Posterior analytics [Wal92c], chapter 5, pages 193–237. ISBN 0-7923-1577-4, 0-7923-1579-0 (set), 94-015-8040-5. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 137.

URL https://link.springer.com/
chapter/10.1007/978-94-015-8040-
3_5.

Wallace:1992:I

[Wal92h] William A. Wallace. Introduction. In Galileo’s Logical Treatises: A Translation, with Notes and Commentary, of His Appropriated Latin Questions on Aristotle’s Posterior Analytics [Wal92e], chapter 1, pages 1–83. ISBN 90-481-4116-8, 94-015-8036-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-015-8036-6_1.

Wallace:1992:SOU

[Wal92i] William A. Wallace. Science and opinion as understood in MS 27. In Galileo’s logic of discovery and proof: the background, content, and use of his appropriated treatises on Aristotle’s Posterior analytics [Wal92c], chapter 3, pages 84–133. ISBN 0-7923-1577-4, 0-7923-1579-0 (set), 94-015-8040-5. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 137.

URL https://link.springer.com/
chapter/10.1007/978-94-015-8040-
3_3.

Wallace:1992:ULI

[Wal92j] William A. Wallace. The understanding of logic implicit in MS 27. In Galileo’s logic of discovery and proof: the background, content, and use of his appropriated treatises on Aristotle’s Posterior analytics [Wal92c], chapter 2, pages 33–83. ISBN 0-7923-1577-4, 0-7923-1579-0 (set), 94-015-8040-5.

ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 137.

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-015-8040-
                                              3_2.

Wallner:1995:ACR

[Wal95] Mathias Wallner. American creativity research in a bipolar world: A look at one chapter in world history and history of science. In Gavroglu et al. [GSW95c], chapter 15, pages 253–272. ISBN 0-7923-2991-0 (set), 0-7923-2988-0, 94-017-2658-2. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 163. URL https:/

                                              /link.springer.com/chapter/10.
                                              1007/978-94-017-2658-0_15.

Wallin:2007:EE

[Wal07a] Annika Wallin. Explanation and environment. In Rethinking Explanation [PY07], chapter 12, pages 163–175. ISBN 1-4020-5580-3 (paperback), 1-4020-5581-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN BD237 .R48 2007; Q175.32.E97 R48

                                              2007. URL https://link.springer.
                                              com/chapter/10.1007/978-1-4020-
                                              5581-2_12.

Walter:2007:BVF

[Wal07b] Scott Walter. Breaking in the 4-vectors: The four-dimensional movement in gravitation, 1905–1910. In Renn et al. [RSS^{+}07a], pages 193–252. ISBN 1-4020-3999-9. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 no. 250 v. 3; QC173.6 .G469 2007.

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-1-4020-4000-
                                              9_18.

Wang:2017:BSP

[Wan17] Jessica Wang. “Broken symmetry”: Physics, aesthetics, and moral virtue in Nuclear Age America. In van Dongen and Paul [vDP17a], chapter 3, pages 27–47. ISBN 3-319-48892-9 (hardcover), 3-319-48893-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 v. 321; Q175.32.V57 E65 2017. URL https:

//link.springer.com/chapter/10.
1007/978-3-319-48893-6_3.

Wang:2018:CAT

[Wan18a] Bin Wang. Conflicts and adaptations in technology transfer to Modern China: The Jiaoji Railway Case. In Mitcham et al. [MLNZ18], chapter 17, pages 201–213. ISBN 3-319-62448-2, 3-319-62450-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN TA15 .P45 2018. URL https://link.

springer.com/chapter/10.1007/
978-3-319-62450-1_17.

Wang:2018:TEP

[Wan18b] Dazhou Wang. Toward an experimental philosophy of engineering. In Mitcham et al. [MLNZ18], chapter 4, pages 37– 50. ISBN 3-319-62448-2, 3-319-62450-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN TA15 .P45

2018. URL https://link.springer.
com/chapter/10.1007/978-3-319-
62450-1_4.

Wang:2018:ACA

[Wan18c] Nan Wang. Ancient Chinese attitudes toward technics: Chinese philosophy of technology prior to the 1800s. In Mitcham et al. [MLNZ18], chapter 6, pages 61–73. ISBN 3-319-62448-2, 3-319-62450-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN TA15 .P45 2018. URL https:

//link.springer.com/chapter/10.
1007/978-3-319-62450-1_6.

Wartofsky:1963:PBC

[War63] Marx W. Wartofsky, editor. Proceedings of the Boston Colloquium for the Philosophy of Science 1961/1962, volume 1 of Boston Studies in the Philosophy of Science. D. Reidel, Dordrecht, The Netherlands; Boston, MA, USA; Lancaster, UK; Tokyo, Japan, 1963. ISBN 94-010-3265-3, 94-010-3263-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B678 1970.

                                              URL https://link.springer.com/
                                              book/10.1007/978-94-010-3263-6.

Wartofsky:1967:MHS

[War67a] Marx Wartofsky. Metaphysics as heuristic for science. In Cohen and Wartofsky [CW67], pages 123–172. ISBN 90-277-0013-3, 94-010-3508-3. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175 .B731 v.3.

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-010-3508-
                                              8_8.

Wartofsky:1967:CIV

[War67b] Marx W. Wartofsky. Comments: Illustration vs. experimental test. In Cohen and Wartofsky [CW67], pages 433–439. ISBN 90-277-0013-3, 94-010-3508-3. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175 .B731 v.3.

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-010-3508-
                                              8_25.

Wartofsky:1974:SR

[War74] Marx W. Wartofsky. Is science rational? In For Dirk Struik: Scientific, Historical and Political Essays in Honor of Dirk J. Struik [CSW74], chapter 36, pages 465–475. ISBN 90-277-0379-5, 94-010-2115-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-010-2115-9_36.

Wartofsky:1976:RBP

[War76] Marx W. Wartofsky. The relation between philosophy of science and history of science. In Essays in Memory of Imre Lakatos [CFW76], chapter 39, pages 717–737. ISBN 90-277-0655-7, 94-010-1451-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-010-
1451-9_39.

Wartofsky:1979:APS

[War79a] Marx W. Wartofsky. Action and passion: Spinoza’s construction of a scientific psychology. In Models: Representation and the Scientific Understanding [War79k], chapter 13, pages 231–254. ISBN 90-277-0947-5, 94-009-9357-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 48; Q175.3.

URL https://link.springer.com/
chapter/10.1007/978-94-009-9357-
0_13.

Wartofsky:1979:ATC

[War79b] Marx W. Wartofsky. Art and technology: Conflicting models of education? The uses of a cultural myth. In Models: Representation and the Scientific Understanding [War79k], chapter 17, pages 338–356. ISBN 90-277-0947-5, 94-009-9357-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 48; Q175.3.

URL https://link.springer.com/
chapter/10.1007/978-94-009-9357-
0_17.

Wartofsky:1979:AHP

[War79c] Marx W. Wartofsky. Art as humanizing praxis. In Models: Representation and the Scientific Understanding [War79k], chapter 18, pages 357–369. ISBN 90-277-0947-5, 94-009-9357-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol.

                                              48; Q175.3. URL https://link.
                                              springer.com/chapter/10.1007/
                                              978-94-009-9357-0_18.

Wartofsky:1979:DDM

[War79d] Marx W. Wartofsky. Diderot and the development of materialist monism. In Models: Representation and the Scientific Understanding [War79k], chapter 16, pages 297–337. ISBN 90-277-0947-5, 94-009-9357-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol.

                                              48; Q175.3. URL https://link.
                                              springer.com/chapter/10.1007/
                                              978-94-009-9357-0_16.

Wartofsky:1979:PLG

[War79e] Marx W. Wartofsky. From praxis to logos: Genetic epistemology and physics. In Models: Representation and the Scientific Understanding [War79k], chapter 9, pages 154–174. ISBN 90-277-0947-5, 94-009-9357-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol.

                                              48; Q175.3. URL https://link.
                                              springer.com/chapter/10.1007/
                                              978-94-009-9357-0_9.

Wartofsky:1979:HCI

[War79f] Marx W. Wartofsky. Hume’s concept of identity and the principium individuations. In Models: Representation and the Scientific Understanding [War79k], chapter 15, pages 277–296. ISBN 90-277-0947-5, 94-009-9357-9 (e-book).

ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol.

48; Q175.3. URL https://link.
springer.com/chapter/10.1007/
978-94-009-9357-0_15.

Wartofsky:1979:MAI

[War79g] Marx W. Wartofsky. Matter, action and interaction. In Models: Representation and the Scientific Understanding [War79k], chapter 5, pages 90–103. ISBN 90-277-0947-5, 94-009-9357-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 48; Q175.3. URL https:

//link.springer.com/chapter/10.
1007/978-94-009-9357-0_5.

Wartofsky:1979:MHS

[War79h] Marx W. Wartofsky. Metaphysics as heuristic for science. In Models: Representation and the Scientific Understanding [War79k], chapter 4, pages 40–89. ISBN 90-277-0947-5, 94-009-9357-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 48; Q175.3.

URL https://link.springer.com/
chapter/10.1007/978-94-009-9357-
0_4.

Wartofsky:1979:MMP

[War79i] Marx W. Wartofsky. The model muddle: Proposals for an immodest realism. In Models: Representation and the Scientific Understanding [War79k], chapter 1, pages 1–11. ISBN 90-277-0947-5, 94-009-9357-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol.

48; Q175.3. URL https://link.
springer.com/chapter/10.1007/
978-94-009-9357-0_1.

Wartofsky:1979:MMV

[War79j] Marx W. Wartofsky. Models, metaphysics and the vagaries of empiricism. In Models: Representation and the Scientific Understanding [War79k], chapter 3, pages 24–39. ISBN 90-277-0947-5, 94-009-9357-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol.

                                              48; Q175.3. URL https://link.
                                              springer.com/chapter/10.1007/
                                              978-94-009-9357-0_3.

Wartofsky:1979:MRS

[War79k] Marx W. Wartofsky. Models: Representation and the Scientific Understanding, volume 48 of Boston Studies in the Philosophy of Science. D. Reidel, Dordrecht, The Netherlands; Boston, MA, USA; Lancaster, UK; Tokyo, Japan, 1979. ISBN 90-277-0947-5, 94-009-9357-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). xxvi + 398 pp. LCCN Q174 .B67 vol. 48; Q175.3. URL

                                              https://link.springer.com/book/
                                              10.1007/978-94-009-9357-0.

Wartofsky:1979:NNI

[War79l] Marx W. Wartofsky. Nature, number and individuals: Motive and method in Spinoza’s philosophy. In Models: Representation and the Scientific Understanding [War79k], chapter 14, pages 255–276. ISBN 90-277-0947-5, 94-009-9357-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 48; Q175.3.

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-009-9357-
                                              0_14.

Wartofsky:1979:PRF

[War79m] Marx W. Wartofsky. Perception, representation, and the forms of action: Towards an historical epistemology.

In Models: Representation and the Scientific Understanding [War79k], chapter 11, pages 188–210. ISBN 90-277-0947-5, 94-009-9357-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol.

48; Q175.3. URL https://link.
springer.com/chapter/10.1007/
978-94-009-9357-0_11.

Wartofsky:1979:PRU

[War79n] Marx W. Wartofsky. Pictures, representation, and the understanding. In Models: Representation and the Scientific Understanding [War79k], chapter 10, pages 175–187. ISBN 90-277-0947-5, 94-009-9357-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol.

48; Q175.3. URL https://link.
springer.com/chapter/10.1007/
978-94-009-9357-0_10.

Wartofsky:1979:REO

[War79o] Marx W. Wartofsky. Reduction, explanation and ontology. In Models: Representation and the Scientific Understanding [War79k], chapter 2, pages 12–23. ISBN 90-277-0947-5, 94-009-9357-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 48; Q175.3.

URL https://link.springer.com/
chapter/10.1007/978-94-009-9357-
0_2.

Wartofsky:1979:RBP

[War79p] Marx W. Wartofsky. The relation between philosophy of science and history of science. In Models: Representation and the Scientific Understanding [War79k], chapter 7, pages 119–139. ISBN 90-277-0947-5, 94-009-9357-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 48; Q175.3. URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-009-9357-0_7.

Wartofsky:1979:RRV

[War79q] Marx W. Wartofsky. Rules and representation: The virtues of constancy and fidelity put in perspective. In Models: Representation and the Scientific Understanding [War79k], chapter 12, pages 211–230. ISBN 90-277-0947-5, 94-009-9357-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 48; Q175.3.

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-009-9357-
                                              0_12.

Wartofsky:1979:TTM

[War79r] Marx W. Wartofsky. Telos and technique: Models as modes of action. In Models: Representation and the Scientific Understanding [War79k], chapter 8, pages 140–153. ISBN 90-277-0947-5, 94-009-9357-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol.

                                              48; Q175.3. URL https://link.
                                              springer.com/chapter/10.1007/
                                              978-94-009-9357-0_8.

Wartofsky:1979:TCM

[War79s] Marx W. Wartofsky. Towards a critical materialism. In Models: Representation and the Scientific Understanding [War79k], chapter 6, pages 104–118. ISBN 90-277-0947-5, 94-009-9357-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 48; Q175.3. URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-009-9357-0_6.

Wartofsky:1980:SJC

[War80] Marx W. Wartofsky. Scientific judgment: Creativity and discovery in scientific thought. In Scientific Discovery: Case Studies [Nic80e], chapter 1, pages 1–16. ISBN 90-277-1093-7, 94-009-9015-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 60. URL https:/

/link.springer.com/chapter/10.
1007/978-94-009-9015-9_1.

Wartofsky:1986:PRF

[War86]  Marx W. Wartofsky.  Perception,

representation, and the forms of action: Towards an historical epistemology. In Cohen and Wartofsky [CW85], chapter 16, pages 215–237. ISBN 90-277-1971-3, 94-009-5345-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-009-5345-1_16.

Wartofsky:1987:EH

[War87] Marx W. Wartofsky. Epistemology historicized. In Shimony and Nails [SN87], pages 357–374. ISBN 94-010-8168-9, 94-009-3735-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-009-3735-2_24.  See

comment [Shi87f].

Wartofsky:1991:HGR

[War91] Marx W. Wartofsky. How to be a good realist. In Beyond Reason: Essays on the Philosophy of Paul Feyerabend [Mun91a], chapter 2, pages 25–40. ISBN 94-010-5406-1, 94-011-3188-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-011-3188-
9_2.

Warnke:1992:WI

[War92] Martin Warnke. Works of the imagination. In The Scientific Enterprise: The Bar-Hillel Colloquium: Studies in History, Philosophy, and Sociology of Science, Volume 4 [UM92], chapter 5, pages 101–116. ISBN 94-010-5190-9, 94-011-2688-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .S354 1992. URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-011-2688-5_5.

Wartofsky:1996:TSC

[War96] Marx W. Wartofsky. Three stages of constitution: Historical changes in the ontological status of the scientific object. In Cohen et al. [CHR96], pages 207–218. ISBN 90-481-4493-0, 94-015-8638-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

                                              ???? URL https://link.springer.
                                              com/chapter/10.1007/978-94-015-
                                              8638-2_14.

Ward:2012:EPP

[War12] Chuck Ward. Evolutionary psychology and the problem of neural plasticity. In Philosophy of Behavioral Biology [PR12b], chapter 11, pages 235–254. ISBN 94-007-1950-7, 94-007-1951-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QH331 .P4679 2012; Q175 .B657 v.282.

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-007-1951-
                                              4_11.

Wasserstein:1991:HDC

[Was91] Abraham Wasserstein. Hunches that did not come off: Some problems in Greek science. In Unguru [Ung91b], pages 19–44. ISBN 94-010-5476-2, 94-011-3342-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic).

LCCN QC5.53; QC6.9. URL https:

//link.springer.com/chapter/10.
1007/978-94-011-3342-5_2.

Waschkies:2004:I

[Was04] Hans-Joachim Waschkies. Introduction. In Classics in the History of Greek Mathematics [Chr04a], chapter 1, pages 3–18. ISBN 1-4020-0081-2 (hardcover), 1-4020-2640-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QA22 .C596 2004. URL https:

//link.springer.com/chapter/10.
1007/978-1-4020-2640-9_1.

Watanabe:1963:MMB

[Wat63] Satosi Watanabe. A model of mind– body relation in terms of modular logic. In Wartofsky [War63], pages 1–41. ISBN 94-010-3265-3, 94-010-3263-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174

.B678 1970. URL https://link.
springer.com/chapter/10.1007/
978-94-010-3263-6_1.

Watkins:1976:HCT

[Wat76] J. W. N. Watkins. The human condition: Two criticisms of Hobbes. In Essays in Memory of Imre Lakatos [CFW76], chapter 38, pages 691–716. ISBN 90-277-0655-7, 94-010-1451-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-010-1451-
9_38.

Watkins:1978:CPC

[Wat78a] John Watkins. Corroboration and the problem of content-comparison. In Progress and Rationality in Science [RA78b], chapter 16, pages 339–378. ISBN 90-277-0922-X, 94-009-9866-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-009-9866-
                                              7_16.

Watkins:1978:PAS

[Wat78b] John Watkins. The Popperian approach to scientific knowledge. In Progress and Rationality in Science [RA78b], chapter 2, pages 23–43. ISBN 90-277-0922-X, 94-009-9866-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-009-9866-7_2.

Watanabe:1983:TP

[Wat83] Satosi Watanabe. Theory of propensity. In Language, Logic and Method [CW83b], chapter 15, pages 283–308. ISBN 94-009-7704-2, 94-009-7702-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-009-7702-
                                              0_15.

Watkins:1987:NVS

[Wat87] John Watkins. A new view of scientific rationality. In Pitt and Pera [PP87], chapter 3, pages 63–82. ISBN 90-277-2417-2, 94-009-3779-2. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 98.

                                              US$64.00.  URL https://link.
                                              springer.com/chapter/10.1007/
                                              978-94-009-3779-6_3.

Watkins:1989:MSR

[Wat89] John Watkins. The methodology of scientific research programmes: A retrospect. In Imre Lakatos and Theories of Scientific Change [GGN89], chapter 1, pages 3–13. ISBN 94-010-7860-2, 94-009-3025-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-009-
3025-4_1.

Watson:1992:EAR

[Wat92a] Patty Jo Watson. Explanation in archaeology: Reactions and rebuttals. In Metaarchaeology: Reflections by Archaeologists and Philosophers [Emb92b], chapter 5, pages 121–140. ISBN 94-010-4806-1, 94-011-1826-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-011-1826-
2_5.

Watson:1992:PAS

[Wat92b] Richard A. Watson. The place of archaeology in science. In Metaarchaeology: Reflections by Archaeologists and Philosophers [Emb92b], chapter 11, pages 255–267. ISBN 94-010-4806-1, 94-011-1826-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-011-
1826-2_11.

Watkins:1995:EHF

[Wat95a] John Watkins. Epiphenomenalism and human freedom. In Critical Rationalism, Metaphysics and Science: Essays for Joseph Agassi. Volume I [JL95a], chapter 4, pages 33–39. ISBN 94-010-4211-X, 94-011-0471-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-011-0471-5_4.

Watkins:1995:ST

[Wat95b] John Watkins. Scientific theoryhood. In Misiek [Mis95c], pages 221–229. ISBN 94-010-4206-3, 94-011-0461-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-011-0461-6_17.

Watkins:1996:SRV

[Wat96] John Watkins. Scientific realism versus common-sense realism? In Cohen et al. [CHR96], pages 219–226. ISBN 90-481-4493-0, 94-015-8638-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-015-8638-2_15.

Watanabe:1998:CTI

[Wat98]  Satosi Watanabe.  Causality and

temporal irreversibility. In Japanese Studies in the Philosophy of Science [NC98], chapter 7, pages 99–114. ISBN 94-010-6176-9, 94-011-5175-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-011-5175-7_7.

Watkins:2001:RFE

[Wat01a] Elizabeth S. Watkins. Radioactive fallout and emerging environmentalism: Cold war fears and public health concerns, 1954–1963. In Allen and MacLeod [AM01b], chapter 18, pages 291–306. ISBN 1-4020-0495-8 (hardcover), 94-017-2956-5. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 228; Q175.55.

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-017-2956-
                                              7_18.

Watkins:2001:KEF

[Wat01b] Eric Watkins. Kant on extension and force: Critical appropriations of Leibniz and Newton. In Between Leibniz, Newton, and Kant: Philosophy and Science in the Eighteenth Century [Lef01a], chapter 6, pages 111–127. ISBN 90-481-5774-9, 94-015-9729-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-015-9729-
6_6.

Wawrzynczak:1982:OSM

[Waw82] Ryszard Wawrzyńczak. An outline of a simulation model of science as a part of the model of action. In Polish Essays in the Philosophy of the Natural Sciences [Kra82c], chapter 33, pages 391–405. ISBN 94-009-7705-0 (e-book), 90-277-1287-5. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-009-7705-
1_33.

Woodward:1991:WVSb

[WC91] William R. Woodward and Robert S.

Cohen, editors. World Views and Scientific Discipline Formation: Science Studies in the German Democratic Republic: Papers from a German– American Summer Institute, 1988, volume 134 of Boston Studies in the Philosophy of Science. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 1991. ISBN 94-010-5395-2, 94-011-3164-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
book/10.1007/978-94-011-3164-3.

Webb:1983:GTC

[Web83] Judson C. Webb. Gödel’s theorems and Church’s thesis: A prologue to mechanism. In Language, Logic and Method [CW83b], chapter 16, pages 309–353. ISBN 94-009-7704-2, 94-009-7702-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

                                              ???? URL https://link.springer.
                                              com/chapter/10.1007/978-94-009-
                                              7702-0_16.

Webb:1987:IKG

[Web87] Judson Webb. Immanuel Kant and the greater glory of geometry. In Shimony and Nails [SN87], pages 17–70. ISBN 94-010-8168-9, 94-009-3735-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-009-3735-2_2.  See

[Shi87g].

Weber:1993:INE

[Web93] Tullio Weber. Indeterminism, nonseparability and the Einstein–Podolsky– Rosen Paradox. In Bridging the Gap: Philosophy, Mathematics, and Physics: Lectures on the Foundations of Science [CCG93], chapter 10, pages 199–209. ISBN 94-010-5101-1, 94-011-2496-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-011-2496-
                                              6_10.

Weber:2008:CSC

[Web08] Marcel Weber. Commentary on “Is Science Cumulative? A Physicist Viewpoint”, by Bernard d’Espagnat. In Rethinking Scientific Change and Theory Comparison: Stabilities, Ruptures, Incommensurabilities? [SSHH08], chapter 11, pages 153–156. ISBN 1-4020-6274-5 (hardcover), 1-4020-6279-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175 .R397 2008. URL https://

                                              link.springer.com/chapter/10.

1007/978-1-4020-6279-7_11.  See

[d’E08].

Weber:2011:BBO

[Web11] Jutta Weber. Black-boxing organisms, exploiting the unpredictable: Control paradigms in human–machine translations. In Science in the Context of Application [CN11a], chapter 24, pages 409–429. ISBN 90-481-9050-9, 90-481-9051-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q180.55.M4 S385 2011. URL https:

//link.springer.com/chapter/10.
1007/978-90-481-9051-5_24.

Weber:2012:BTI

[Web12]  Marcel Weber.  Behavioral traits,

the intentional stance, and biological functions: What neuroscience explains. In Philosophy of Behavioral Biology [PR12b], chapter 15, pages 317–327. ISBN 94-007-1950-7, 94-007-1951-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QH331 .P4679 2012; Q175 .B657 v.282.

URL https://link.springer.com/
chapter/10.1007/978-94-007-1951-
4_15.

Wedin:1986:TAN

[Wed86] Michael V. Wedin. Tracking Aristotle’s Noûs. In Cohen and Wartofsky [CW85], chapter 9, pages 167–197. ISBN 90-277-1971-3, 94-009-5345-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-009-5349-9_9.

Weiler:1976:PV

[Wei76] Gershon Weiler. Points of view. In Language in Focus: Foundations, Methods and Systems: Essays in Memory of Yehoshua Bar-Hillel [Kas76b], chapter 28, pages 661–674. ISBN 90-277-0644-1, 90-277-0645-X, 94-010-1876-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

P101-120.  URL https://link.
springer.com/chapter/10.1007/
978-94-010-1876-0_28.

Weingartner:1982:NTI

[Wei82] Paul Weingartner. A new theory of intension. In Scientific Philosophy Today: Essays in Honor of Mario Bunge [AC82], chapter 25, pages 439–464. ISBN 90-277-1262-X, 90-277-1263-8, 94-009-8462-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 67. URL https:

//link.springer.com/chapter/10.
1007/978-94-009-8462-2_25.

Weiss:1984:TVF

[Wei84] Donald D. Weiss. Toward the vindication of Friedrich Engels. In Methodology, Metaphysics and the History of Science: In Memory of Benjamin Nelson [CW84b], chapter 15, pages 331–358. ISBN 94-009-6333-5, 94-009-6331-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-009-
6331-3_15.

Weiler:1995:RMP

[Wei95] G. Weiler. Reason and myth in politics. In Critical Rationalism, the Social Sciences and the Humanities: Essays for Joseph Agassi. Volume II [JL95b], chapter 3, pages 41–57. ISBN 0-7923-2961-9, 90-481-4430-2, 94-017-0441-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN B29 .C75 1995. URL https:

//link.springer.com/chapter/10.
1007/978-94-017-0441-0_3.

Wei:1996:MC

[Wei96] Zhang Wei. Millikan and China. In Chinese Studies in the History and Philosophy of Science and Technology [DC96], chapter 34, pages 439–446. ISBN 90-481-4546-5, 94-015-8717-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 180; U390 .N19 1996.

URL https://link.springer.com/
chapter/10.1007/978-94-015-8717-
4_34.

Weingart:2001:LDS

[Wei01a] Peter Weingart. The loss of distance: Science in transition. In Allen and MacLeod [AM01b], chapter 11, pages 167–184. ISBN 1-4020-0495-8 (hardcover), 94-017-2956-5. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 228; Q175.55.

URL https://link.springer.com/
chapter/10.1007/978-94-017-2956-
7_11.

Weininger:2001:AAR

[Wei01b] Stephen J. Weininger. Affinity, additivity and the reification of the bond. In Klein [Kle01b], chapter 14, pages 237–251. ISBN 1-4020-0100-2, 94-015-9737-5. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 222; QD7. URL https:/

/link.springer.com/chapter/10.
1007/978-94-015-9737-1_14.

Weil:2004:WBE

[Wei04] André Weil. Who betrayed Euclid?

In Classics in the History of Greek Mathematics [Chr04a], chapter 24, pages 447–449. ISBN 1-4020-0081-2 (hardcover), 1-4020-2640-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QA22 .C596 2004.

URL https://link.springer.com/
chapter/10.1007/978-1-4020-2640-
9_24.

Weisberg:2006:WH

[Wei06] Michael Weisberg. Water is not h 2 o. In Philosophy of Chemistry: Synthesis of a New Discipline [BSM06b], chapter 18, pages 337–345. ISBN 1-4020-2455-X, 1-4020-3256-0, 1-4020-3261-7. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 242 QD6.

URL https://link.springer.com/
chapter/10.1007/1-4020-3261-7_18.

Weininger:2007:LSB

[Wei07] Stephen J. Weininger. Letting the scientists back in. In Positioning the history of science [RG07], chapter 30, pages 173–176. ISBN 1-4020-5419-X (hardcover), 1-4020-5420-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q125 .P77 2007.

URL https://link.springer.com/
chapter/10.1007/1-4020-5420-3_30.

Weirich:2010:PDR

[Wei10] Paul Weirich. Probabilities in decision rules. In The Place of Probability in Science: In Honor of Ellery Eells (1953–2006) [Fet10a], chapter 13, pages 289–319. ISBN 90-481-3614-8 (hardcover), 90-481-3615-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175 .P563 2010.

URL https://link.springer.com/
chapter/10.1007/978-90-481-3615-
5_13.

Weingart:2011:SPM

[Wei11] Peter Weingart. Science, the public and the media — views from everywhere. In Science in the Context of Application [CN11a], chapter 20, pages 337–348. ISBN 90-481-9050-9, 90-481-9051-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q180.55.M4 S385 2011. URL https:

//link.springer.com/chapter/10.
1007/978-90-481-9051-5_20.

Weisz:2014:OTC

[Wei14] George Weisz. The ongoing tension:

Clinical practice and clinical research. In The Institution of Science and the Science of Institutions: The Legacy of Joseph Ben-David [Her14b], chapter 5, pages 63–79. ISBN 94-007-7406-0 (hardcover), 94-007-7407-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 v.302; Q125 .I58 2014. URL https://link.

springer.com/chapter/10.1007/
978-94-007-7407-0_5.

Welshon:1999:POR

[Wel99] Robert C. Welshon. Perspectivist ontology and de re knowledge. In Nietzsche, Epistemology, and Philosophy of Science: Nietzsche and the Sciences II [Bab99a], chapter 3, pages 39–46. ISBN 90-481-5234-8, 94-017-2428-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-017-2428-9_3.

Wells:2015:MMM

[Wel15] Robert Wells. “Those that have most money must have least learning”: Undergraduate education at the University of Oxford in the Eighteenth and early Nineteenth Centuries. In Sciences in the Universities of Europe, Nineteenth and Twentieth Centuries: Academic Landscapes [SDG15b], chapter 2, pages 11–29. ISBN 94-017-9635-1, 94-017-9636-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q183.4.E85 2015. URL https://link.

springer.com/chapter/10.1007/
978-94-017-9636-1_2.

Wenzhao:1996:BAC

[Wen96] Lin Wenzhao. A brief account of Chinese studies of the history of science and technology in China. In Chinese Studies in the History and Philosophy of Science and Technology [DC96], chapter 16, pages 199–205. ISBN 90-481-4546-5, 94-015-8717-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 180; U390 .N19 1996. URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-015-8717-4_16.

Wendt:2014:SCB

[Wen14] Helge Wendt. Scientific cooperation between the German Academy of Sciences in Berlin (DAW) and Cuba in the 1960s and 1970s. In The History of Physics in Cuba [BRW14a], chapter 26, pages 387–393. ISBN 94-017-8040-4, 94-017-8041-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC5.53; QC6.9; QC9 C9H57 2014; Q124.6-127.2. URL https:/

                                              /link.springer.com/chapter/10.
                                              1007/978-94-017-8041-4_26.

Wernicke:1969:SCA

[Wer69] Carl Wernicke. The symptom complex of aphasia. In Cohen and Wartofsky [CW69b], pages 34–97. ISBN 94-010-3380-3, 94-010-3378-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-010-3378-7_2.

Werskey:2001:SCS

[Wer01] Gary Werskey. “The social context of science”: SOC SCI 119 as a way of life and learning. In Allen and MacLeod [AM01b], chapter 13, pages 203–214. ISBN 1-4020-0495-8 (hardcover), 94-017-2956-5. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 228; Q175.55. URL https:

//link.springer.com/chapter/10.
1007/978-94-017-2956-7_13.

Wessels:1976:LMP

[Wes76] Linda Wessels. Laws and meaning postulates. In Cohen et al. [CHMV76], pages 215–234. ISBN 90-277-0647-6, 90-277-0648-4 (paperback), 94-010-1449-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-010-
1449-6_10.

Westphal:1987:W

[Wes87] Jonathan Westphal. Whiteness. In Amrine et al. [AZW87], chapter 16, pages 319–339. ISBN 90-277-2265-X, 94-009-3761-X. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 97. US$79.00. URL https:

//link.springer.com/chapter/10.
1007/978-94-009-3761-1_16.

Wessely:1990:SIL

[Wes90a] Anna Wessely. Simmel’s influence on Lukács’s conception of the sociology of art. In Georg Simmel and Contemporary Sociology [KPC90], chapter 20, pages 357–373. ISBN 94-010-6691-4 (hardcover), 94-009-0459-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN HM401-1281.

URL https://link.springer.com/
chapter/10.1007/978-94-009-0459-
0_20.

Wessely:1990:SM

[Wes90b] Anna Wessely. Simmel’s metaphysics.

In Georg Simmel and Contemporary Sociology [KPC90], chapter 21, pages 375–383. ISBN 94-010-6691-4 (hardcover), 94-009-0459-2 (e-book).

ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN HM401-1281.

URL https://link.springer.com/
chapter/10.1007/978-94-009-0459-
0_21.

Wessel:1991:RET

[Wes91] Karl-Friedrich Wessel. Relating evolutionary theory to the natural sciences. In Woodward and Cohen [WC91], pages 29–37. ISBN 94-010-5395-2, 94-011-3164-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-011-
3164-3_3.

Westfall:1992:INT

[Wes92] Richard S. Westfall. Isaac Newton: Theologian. In The Scientific Enterprise: The Bar-Hillel Colloquium: Studies in History, Philosophy, and Sociology of Science, Volume 4 [UM92], chapter 14, pages 223–239. ISBN 94-010-5190-9, 94-011-2688-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .S354 1992.

URL https://link.springer.com/
chapter/10.1007/978-94-011-2688-
5_14.

Westfall:1994:CSC

[Wes94] Richard S. Westfall. Charting the scientific community. In Trends in the Historiography of Science [GCN94], chapter 1, pages 1–14. ISBN 90-481-4264-4, 94-017-3596-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-017-3596-4_1.

Wessely:1995:RPR

[Wes95] Anna Wessely. The reader’s progress: Remarks on Arnold Hauser’s philosophy of art history. In Science, Mind and Art: Essays on science and the humanistic understanding in art, epistemology, religion and ethics In honor of Robert S. Cohen [GSW95b], chapter 3, pages 29–43. ISBN 0-7923-2990-2 (hardcover), 94-011-0469-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 165 BD418.3. URL https://link.

springer.com/chapter/10.1007/
978-94-011-0469-2_3.

Wettersten:1982:PMB

[Wet82] John Wettersten. The place of Mario Bunge. In Scientific Philosophy Today: Essays in Honor of Mario Bunge [AC82], chapter 26, pages 465–486. ISBN 90-277-1262-X, 90-277-1263-8, 94-009-8462-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 67. URL https:

//link.springer.com/chapter/10.
1007/978-94-009-8462-2_26.

Wetzels:1987:ASO

[Wet87] Walter D. Wetzels. Art and science:

Organicism and Goethe’s classical aesthetics. In Burwick [Bur87b], chapter 3, pages 71–85. ISBN 90-277-2541-1, 94-009-3917-5. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 105. US$64.00(U.S.).

URL https://link.springer.com/
chapter/10.1007/978-94-009-3917-
2_3.

Wettersten:1995:PRE

[Wet95] John Wettersten. Preliminary report on efforts of psychologism to gain influence in proper epistemological, methodological and psychological societies. In Critical Rationalism, the Social Sciences and the Humanities: Essays for Joseph Agassi. Volume II [JL95b], chapter 8, pages 129–152. ISBN 0-7923-2961-9, 90-481-4430-2, 94-017-0441-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN B29 .C75 1995. URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-017-0441-0_8.

Wettersten:2009:PSR

[Wet09] John Wettersten. Popper and Sen on rationality and economics: Two (independent) wrong turns can be remedied with the same program. In Rethinking Popper [PC09], chapter 27, pages 369–378. ISBN 1-4020-9337-3, 1-4020-9338-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN B1649.P64 R48 2009; Q174 .B67 2009.

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-1-4020-9338-
                                              8_27.

Wettersten:2017:LOM

[Wet17] John R. Wettersten. The legends of one methodology of science used throughout its history and its independence from the institutions in which science has been conducted. In Encouraging Openness: Essays for Joseph Agassi on the Occasion of His 90th Birthday [BAG17], chapter 18, pages 221–233. ISBN 3-319-57668-2 (hardcover), 3-319-57669-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN B833.E53 2017. URL https://link.

                                              springer.com/chapter/10.1007/
                                              978-3-319-57669-5_18.

Weyl:2007:PIG

[Wey07] Hermann Weyl. Purely infinitesimal geometry (excerpt). In Renn et al. [RSS^{+}07b], pages 1089–1105. ISBN 1-4020-3999-9 (hardcover), 1-4020-4000-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 no. 250 v. 4; QC173.6 .G469 2007.

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-1-4020-4000-

9_49. Originally published as “Reine Infinitesimalgeometrie”, Mathematisch Zeitschrift 2, 384–411 (1918).

Wilholt:2011:CST

[WG11] Torsten Wilholt and Hans Glimell.

Conditions of science: The three-way tension of freedom, accountability and utility. In Science in the Context of Application [CN11a], chapter 21, pages 351–370. ISBN 90-481-9050-9, 90-481-9051-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q180.55.M4 S385 2011. URL https:

//link.springer.com/chapter/10.
1007/978-90-481-9051-5_21.

Wheeler:1974:MAC

[Whe74]  John Archibald Wheeler.  From

Mendeléev’s atom to the collapsing star. In Seeger and Cohen [SC74b], pages 257–301. ISBN 90-277-0390-6, 90-277-0376-0 (paperback), 94-010-2126-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 11. URL https:/

/link.springer.com/chapter/10.
1007/978-94-010-2126-5_15.

Whitfield:1988:ASG

[Whi88] Tephen J. Whitfield. After strange gods: Radical Jews in modern America. In Philosophy, History and Social Action: Essays in Honor of Lewis Feuer with an autobiographic essay by Lewis Feuer [HOO88c], chapter 24, pages 425–460. ISBN 94-010-7793-2, 94-009-2873-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-009-
2873-2_24.

Whitlock:1999:RJB

[Whi99] Greg Whitlock. Roger J. Boscovich and Friedrich Nietzsche: A re-examination.

In Nietzsche, Epistemology, and Philosophy of Science: Nietzsche and the Sciences II [Bab99a], chapter 16, pages 187–201. ISBN 90-481-5234-8, 94-017-2428-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-017-
2428-9_16.

Whittemore:2001:MCS

[Whi01] Gilbert F. Whittemore. The multidimensional chess of science and society: A postwar debate over plutonium exposure. In Allen and MacLeod [AM01b], chapter 17, pages 277–289. ISBN 1-4020-0495-8 (hardcover), 94-017-2956-5. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 228; Q175.55. URL https:

//link.springer.com/chapter/10.
1007/978-94-017-2956-7_17.

Wiatr:1979:MST

[Wia79a] Jerzy J. Wiatr. The Marxist social theory and the challenges of our time. In Polish Essays in the Methodology of the Social Sciences [Wia79b], chapter 1, pages 1–19. ISBN 90-277-0956-4, 94-009-9353-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .P655 1979. URL https://link.

springer.com/chapter/10.1007/
978-94-009-9353-2_1.

Wiatr:1979:PEM

[Wia79b] Jerzy J. Wiatr. Polish Essays in the Methodology of the Social Sciences, volume 29 of Boston Studies in the Philosophy of Science. D. Reidel, Dordrecht, The Netherlands; Boston, MA, USA; Lancaster, UK; Tokyo, Japan, 1979. ISBN 90-277-0956-4, 94-009-9353-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). xiii + 271 pp. LCCN Q174 .P655 1979. URL

https://link.springer.com/book/
10.1007/978-94-009-9353-2.

Wicklein:1991:MAD

[Wic91] Gerald Wicklein. The Middle Ages:

Darkness in the sciences? In Woodward and Cohen [WC91], pages 191–198. ISBN 94-010-5395-2, 94-011-3164-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-011-3164-
3_17.

Wiechert:1992:MRV

[Wie92] Emil Wiechert. Mach’s relativity vs.

Einstein’s relativity. In Ernst Mach — A Deeper Look: Documents and New Perspectives [Bla92a], chapter 8, pages 165–182. ISBN 94-010-5230-1, 94-011-2771-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN C5.53; QC6.9. URL https:

//link.springer.com/chapter/10.
1007/978-94-011-2771-4_8.

Wiesel:1995:RYP

[Wie95] Elie Wiesel. Rabbi Yehuda-the-Prince.

In Science, Mind and Art: Essays on science and the humanistic understanding in art, epistemology, religion and ethics In honor of Robert S. Cohen [GSW95b], chapter 22, pages 369– 384. ISBN 0-7923-2990-2 (hardcover), 94-011-0469-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 165 BD418.3.

URL https://link.springer.com/
chapter/10.1007/978-94-011-0469-
2_22.

Wieber:2012:MMD

[Wie12] Frédéric Wieber. Multiple means of determination and multiple constraints of construction: Robustness and strategies for modeling macromolecular objects. In Characterizing the Robustness of Science: After the Practice Turn in Philosophy of Science [STNW12], chapter 11, pages 267–288. ISBN 94-007-2758-5, 94-007-2759-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175 .C4595 2012.

URL https://link.springer.com/
chapter/10.1007/978-94-007-2759-
5_11.

Wigner:1974:PEL

[Wig74] Eugene P. Wigner. Physics and the explanation of life. In Seeger and Cohen [SC74b], pages 119–132. ISBN 90-277-0390-6, 90-277-0376-0 (paperback), 94-010-2126-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 11. URL https:

//link.springer.com/chapter/10.
1007/978-94-010-2126-5_6.

Williamson:1991:SC

[Wil91]  Donald I. Williamson.  Sequential

chimeras. In Organism and the Origins of Self [Tau91i], chapter 15, pages 299–336. ISBN 0-7923-1185-X, 94-011-3406-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QR180-189.5. URL https://link.

springer.com/chapter/10.1007/
978-94-011-3406-4_15. See comments

[Tau91g].

Wilson:2000:WDN

[Wil00] Mark Wilson. Where did the notion that forces are unobservable come from? In The Reality of the Unobservable: Observability, Unobservability and Their Impact on the Issue of Scientific Realism [AP00b], chapter 17, pages 231–239. ISBN 90-481-5458-8, 94-015-9391-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-015-
9391-5_17.

Wilson:2001:ESP

[Wil01] David B. Wilson. Enlightenment Scotland’s philosophico–chemical physics. In Between Leibniz, Newton, and Kant: Philosophy and Science in the Eighteenth Century [Lef01a], chapter 7, pages 129–145. ISBN 90-481-5774-9, 94-015-9729-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-015-
9729-6_7.

Wilson:2007:WNG

[Wil07] Andrew D. Wilson. The way from nature to God. In Hans Christian Ørsted and The Romantic Legacy In Science: Ideas, Disciplines, Practices [BCK07], chapter 1, pages 1–11. ISBN 1-4020-2979-9 (hardcover), 1-4020-2987-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QC515.O37 H36 2007. URL https:

//link.springer.com/chapter/10.
1007/978-1-4020-2987-5_1.

Wilson:2012:EFS

[Wil12] David Sloan Wilson. Evolving the future: Sketching a science of intentional change. In Philosophy of Behavioral Biology [PR12b], chapter 7, pages 149–165. ISBN 94-007-1950-7, 94-007-1951-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QH331 .P4679 2012; Q175 .B657 v.282.

URL https://link.springer.com/
chapter/10.1007/978-94-007-1951-
4_7.

Wilson:2018:PSE

[Wil18] Catherine Wilson. Philosophical and scientific empiricism and rationalism in the Seventeenth and Eighteenth Centuries. In What Does it Mean to be an Empiricist?: Empiricisms in Eighteenth Century Sciences [BR18b], chapter 7, pages 123–138. ISBN 3-319-69858-3 (hardcover), 3-319-69860-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN B816 .W43 2018.

URL https://link.springer.com/
chapter/10.1007/978-3-319-69860-
1_7.

Wimsatt:1971:SPC

[Wim71] William C. Wimsatt. Some problems with the concept of ‘feedback’. In Buck and Cohen [BC71], pages 241–256. ISBN 90-277-0309-4, 94-010-3142-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-010-3142-
4_17.

Wimsatt:1974:CO

[Wim74] William C. Wimsatt. Complexity and organization. In Schaffner and Cohen [SC74a], pages 67–86. ISBN 90-277-0408-2, 90-277-0409-0, 94-010-2140-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175 .B7312.

URL https://link.springer.com/
chapter/10.1007/978-94-010-2140-
1_5.

Wimsatt:1976:CO

[Wim76] William C. Wimsatt. Complexity and organization. In Topics in the Philosophy of Biology [GM76], chapter 8, pages 174–193. ISBN 90-277-0596-8, 94-010-1829-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-010-
1829-6_8.

Wimsatt:1980:RRS

[Wim80] William C. Wimsatt. Reductionistic research strategies and their biases in the units of selection controversy. In Scientific Discovery: Case Studies [Nic80e], chapter 13, pages 213–259. ISBN 90-277-1093-7, 94-009-9015-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 60. URL https://link.

springer.com/chapter/10.1007/
978-94-009-9015-9_13.

Wimsatt:1986:FA

[Wim86] William C. Wimsatt. Forms of aggregativity. In Cohen and Wartofsky [CW85], chapter 14, pages 259–291. ISBN 90-277-1971-3, 94-009-5345-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-009-5349-
9_14.

Wimsatt:1992:GGC

[Wim92] William C. Wimsatt. Golden generalities and co-opted anomalies: Haldane vs. Muller and the Drosophila Group on the theory and practice of linkage mapping. In Sarkar [Sar92a], chapter 6, pages 107–166. ISBN 0-7923-1777-7 (hardcover), 94-011-2856-1. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 142. URL https:

//link.springer.com/chapter/10.
1007/978-94-011-2856-8_6.

Wimsatt:2012:RMI

[Wim12a] William C. Wimsatt. Robustness:

Material, and inferential, in the natural and human sciences. In Characterizing the Robustness of Science: After the Practice Turn in Philosophy of Science [STNW12], chapter 3, pages 89–104. ISBN 94-007-2758-5, 94-007-2759-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175 .C4595 2012. URL https://link.

                                              springer.com/chapter/10.1007/
                                              978-94-007-2759-5_3.

Wimsatt:2012:RRO

[Wim12b] William C. Wimsatt. Robustness, reliability, and overdetermination (1981). In Characterizing the Robustness of Science: After the Practice Turn in Philosophy of Science [STNW12], chapter 2, pages 61–87. ISBN 94-007-2758-5, 94-007-2759-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175 .C4595 2012. URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-007-2759-5_2.

Wimsatt:2015:ETT

[Wim15] William C. Wimsatt. Entrenchment as a theoretical tool in evolutionary developmental biology. In Conceptual Change in Biology: Scientific and Philosophical Perspectives on Evolution and Development [Lov15b], chapter 17, pages 365–402. ISBN 94-017-9411-1, 94-017-9412-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QH331 C66 2015; QH491; QH492 .C66

                                              2014. URL https://link.springer.
                                              com/chapter/10.1007/978-94-017-
                                              9412-1_17.

Winnie:1971:TA

[Win71] John A. Winnie. Theoretical analyticity. In Buck and Cohen [BC71], pages 289–305. ISBN 90-277-0309-4, 94-010-3142-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

                                              ???? URL https://link.springer.
                                              com/chapter/10.1007/978-94-010-
                                              3142-4_20.

Winner:1983:TPT

[Win83a] Langdon Winner. Technē and politeia:

The technical constitution of society. In Philosophy and Technology [DR83], chapter 7, pages 97–111. ISBN 94-009-7126-5, 94-009-7124-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-009-7124-0_7.

Winner:1983:TFL

[Win83b] Langdon Winner. Technologies as forms of life. In Epistemology, Methodology, and the Social Sciences [CW83a], chapter 10, pages 249–263. ISBN 90-481-8376-6, 94-017-1458-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-017-1458-7_10.

Winner:1986:MIR

[Win86] Langdon Winner. Myth information:

Romantic politics in the computer revolution. In Philosophy and Technology II: Information Technology and Computers in Theory and Practice [MH86], chapter 20, pages 269–289. ISBN 94-010-8510-2, 94-009-4512-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-009-4512-
8_20.

Winther:2015:EDT

[Win15] Rasmus Grønfeldt Winther. Evo-devo as a trading zone. In Conceptual Change in Biology: Scientific and Philosophical Perspectives on Evolution and Development [Lov15b], chapter 21, pages 459–482. ISBN 94-017-9411-1, 94-017-9412-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QH331 C66 2015; QH491; QH492 .C66

                                              2014. URL https://link.springer.
                                              com/chapter/10.1007/978-94-017-
                                              9412-1_21.

Winchester:2017:JAE

[Win17a] Ian S. Winchester. Joseph Agassi’s educational thoughts in interchange (1970–1987). In Encouraging Openness: Essays for Joseph Agassi on the Occasion of His 90th Birthday [BAG17], chapter 23, pages 287–290. ISBN 3-319-57668-2 (hardcover), 3-319-57669-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN B833.E53 2017. URL https://link.

                                              springer.com/chapter/10.1007/
                                              978-3-319-57669-5_23.

Winston:2017:CMI

[Win17b] Patrick Henry Winston. On computing machinery and intelligence. In Floyd and Bokulich [FB17], chapter 11, pages 265–278. ISBN 3-319-53278-2, 3-319-53280-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QA29.T8 P45 2017. URL https:/

                                              /link.springer.com/chapter/10.
                                              1007/978-3-319-53280-6_11.

Wiredu:1995:KTF

[Wir95] Kwasi Wiredu. Knowledge, truth and fallibility. In The Concept of Knowledge: The Ankara Seminar [KC95], chapter 11, pages 127–148. ISBN 90-481-4495-7, 94-017-3263-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-017-3263-5_11.

Wisdom:1969:MRM

[Wis69] J. O. Wisdom. On methods of refutation in metaphysics. In Cohen and Wartofsky [CW69b], pages 523–537.

ISBN 94-010-3380-3, 94-010-3378-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-010-3378-
7_18.

Wisdom:1971:OBB

[Wis71] J. O. Wisdom. Observations as the building blocks of science in 20thcentury scientific thought. In Buck and Cohen [BC71], pages 212–222. ISBN 90-277-0309-4, 94-010-3142-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-010-3142-4_15.

Wise:2007:SH

[Wis07] M. Norton Wise. Science as history.

In Positioning the history of science [RG07], chapter 31, pages 177–183. ISBN 1-4020-5419-X (hardcover), 1-4020-5420-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q125 .P77 2007. URL https://link.

springer.com/chapter/10.1007/1-
4020-5420-3_31.

Wise:2011:TPS

[Wis11] M. Norton Wise. Thoughts on politicization of science through commercialization. In Science in the Context of Application [CN11a], chapter 17, pages 283–299. ISBN 90-481-9050-9, 90-481-9051-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q180.55.M4 S385 2011. URL https:

//link.springer.com/chapter/10.
1007/978-90-481-9051-5_17.

Wittich:1986:LFU

[Wit86] Dieter Wittich. On Ludwik Fleck’s use of social categories in knowledge. In Cognition and Fact: Materials on Ludwik Fleck [CS86], chapter 17, pages 317–323. ISBN 94-010-8504-8, 94-009-4498-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-009-
4498-5_17.

Wang:2018:FAE

[WLC18] Guoyu Wang, Lei Li, and Xu Cao. Feasibility and acceptability in engineering. In Mitcham et al. [MLNZ18], chapter 5, pages 51–59. ISBN 3-319-62448-2, 3-319-62450-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN TA15 .P45 2018. URL https:

//link.springer.com/chapter/10.
1007/978-3-319-62450-1_5.

Weiss:1991:RMV

[WO91] Pierre Weiss and H. Kamerlingh Onnes. Researches on magnetization at very low temperatures. In Gavroglu and Goudaroulis [GG91], chapter 24, pages 388–421. ISBN 94-010-7433-X (hardcover), 94-009-2079-2. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 124. URL https:/

/link.springer.com/chapter/10.
1007/978-94-009-2079-8_24.

Wegener:1997:NTM

[WØ97] Mogens Wegener and Peter Øhrstrøm. A new tempo–modal logic for emerging truth. In Perspectives on Time [FSU97b], chapter 18, pages 417–441. ISBN 90-481-4774-3, 94-015-8875-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol.189. URL https://link.

springer.com/chapter/10.1007/
978-94-015-8875-1_18.

Wolniewicz:1969:PBW

[Wol69] Bogusław Wolniewicz. A parallelism between Wittgensteinian and Aristotelian ontologies. In Cohen and Wartofsky [CW69b], pages 208–217. ISBN 94-010-3380-3, 94-010-3378-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-010-3378-
7_6.

Wolff:1976:Xa

[Wol76a] Kurt H. Wolff. 1951. In Surrender and Catch: Experience and Inquiry Today [Wol76-32], chapter 26, pages 219–220. ISBN 90-277-0765-0, 94-010-1526-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-010-1526-
4_26.

Wolf:1976:XX

[Wol76b] Kurt H. Wolff. 1961–62. In Surrender and Catch: Experience and Inquiry Today [Wol76-32], chapter 27, pages 221–245. ISBN 90-277-0765-0, 94-010-1526-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-010-
1526-4_27.

Wolff:1976:Xb

[Wol76c] Kurt H. Wolff. 1962–63. In Surrender and Catch: Experience and Inquiry Today [Wol76-32], chapter 29, pages 247–258. ISBN 90-277-0765-0, 94-010-1526-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-010-
1526-4_29.

Wolff:1976:Ck

[Wol76d] Kurt H. Wolff. 1964 — 65, concluded.

In Surrender and Catch: Experience and Inquiry Today [Wol76-32], chapter 32, pages 290–306. ISBN 90-277-0765-0, 94-010-1526-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-010-1526-4_32.

Wolff:1976:Xc

[Wol76e] Kurt H. Wolff. 1964–65. In Surrender and Catch: Experience and Inquiry Today [Wol76-32], chapter 30, pages 259–288. ISBN 90-277-0765-0, 94-010-1526-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

                                              ???? URL https://link.springer.
                                              com/chapter/10.1007/978-94-010-
                                              1526-4_30.

Wolff:1976:Xd

[Wol76f] Kurt H. Wolff. 1965–66. In Surrender and Catch: Experience and Inquiry Today [Wol76-32], chapter 34, pages 308–335. ISBN 90-277-0765-0, 94-010-1526-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

                                              ???? URL https://link.springer.
                                              com/chapter/10.1007/978-94-010-
                                              1526-4_34.

Wolff:1976:Xe

[Wol76g] Kurt H. Wolff. 1971. In Surrender and Catch: Experience and Inquiry Today [Wol76-32], chapter 35, pages 336–345. ISBN 90-277-0765-0, 94-010-1526-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-010-1526-
                                              4_35.

Wolff:1976:Cm

[Wol76h] Kurt H. Wolff. 1971, concluded. In Surrender and Catch: Experience and Inquiry Today [Wol76-32], chapter 38, pages 350–391. ISBN 90-277-0765-0, 94-010-1526-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-010-
1526-4_38.

Wolff:1976:L

[Wol76i] Kurt H. Wolff. And a letter. In Surrender and Catch: Experience and Inquiry Today [Wol76-32], chapter 36, pages 346–347. ISBN 90-277-0765-0, 94-010-1526-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-010-
1526-4_36.

Wolff:1976:TE

[Wol76j] Kurt H. Wolff. And there is an end. In Surrender and Catch: Experience and Inquiry Today [Wol76-32], chapter 39, pages 392–393. ISBN 90-277-0765-0, 94-010-1526-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-010-
1526-4_39.

Wolff:1976:BHT

[Wol76k] Kurt H. Wolff. Beginning: In Hegel and today. In Surrender and Catch: Experience and Inquiry Today [Wol76-32], chapter 18, pages 115–138. ISBN 90-277-0765-0, 94-010-1526-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-010-1526-4_18.

Wolff:1976:Cd

[Wol76l] Kurt H. Wolff. Chapter 12. In Surrender and Catch: Experience and Inquiry Today [Wol76-32], chapter 12, page 70. ISBN 90-277-0765-0, 94-010-1526-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-010-1526-
                                              4_12.

Wolff:1976:Ce

[Wol76m] Kurt H. Wolff. Chapter 14. In Surrender and Catch: Experience and Inquiry Today [Wol76-32], chapter 14, pages 94–95. ISBN 90-277-0765-0, 94-010-1526-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

                                              ???? URL https://link.springer.
                                              com/chapter/10.1007/978-94-010-
                                              1526-4_14.

Wolff:1976:Cf

[Wol76n] Kurt H. Wolff. Chapter 16. In Surrender and Catch: Experience and Inquiry Today [Wol76-32], chapter 16, page 109. ISBN 90-277-0765-0, 94-010-1526-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-010-1526-
                                              4_16.

Wolff:1976:Cg

[Wol76o] Kurt H. Wolff. Chapter 19. In Surrender and Catch: Experience and Inquiry Today [Wol76-32], chapter 19, pages 139–143. ISBN 90-277-0765-0, 94-010-1526-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

                                              ???? URL https://link.springer.
                                              com/chapter/10.1007/978-94-010-
                                              1526-4_19.

Wolff:1976:Ch

[Wol76p] Kurt H. Wolff. Chapter 21. In Surrender and Catch: Experience and Inquiry Today [Wol76-32], chapter 21, page 152. ISBN 90-277-0765-0, 94-010-1526-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

                                              URL https://link.springer.com/

chapter/10.1007/978-94-010-1526-
4_21.

Wolff:1976:Ci

[Wol76q] Kurt H. Wolff. Chapter 23. In Surrender and Catch: Experience and Inquiry Today [Wol76-32], chapter 23, pages 177–178. ISBN 90-277-0765-0, 94-010-1526-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-010-
1526-4_23.

Wolff:1976:Cj

[Wol76r] Kurt H. Wolff. Chapter 28. In Surrender and Catch: Experience and Inquiry Today [Wol76-32], chapter 28, page 246. ISBN 90-277-0765-0, 94-010-1526-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-010-1526-
4_28.

Wolff:1976:Cl

[Wol76s] Kurt H. Wolff. Chapter 33. In Surrender and Catch: Experience and Inquiry Today [Wol76-32], chapter 33, page 307. ISBN 90-277-0765-0, 94-010-1526-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-010-1526-
4_33.

Wolff:1976:Ca

[Wol76t] Kurt H. Wolff. Chapter 5. In Surrender and Catch: Experience and Inquiry Today [Wol76-32], chapter 5, pages 28– 29. ISBN 90-277-0765-0, 94-010-1526-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-010-1526-
4_5.

Wolff:1976:Cb

[Wol76u] Kurt H. Wolff. Chapter 7. In Surrender and Catch: Experience and Inquiry Today [Wol76-32], chapter 7, page 36. ISBN 90-277-0765-0, 94-010-1526-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-010-1526-
                                              4_7.

Wolff:1976:Cc

[Wol76v] Kurt H. Wolff. Chapter 9. In Surrender and Catch: Experience and Inquiry Today [Wol76-32], chapter 9, pages 43– 44. ISBN 90-277-0765-0, 94-010-1526-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-010-1526-
                                              4_9.

Wolff:1976:DP

[Wol76w] Kurt H. Wolff. “Dear phantasy”. In Surrender and Catch: Experience and Inquiry Today [Wol76-32], chapter 37, pages 348–349. ISBN 90-277-0765-0, 94-010-1526-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

                                              ???? URL https://link.springer.
                                              com/chapter/10.1007/978-94-010-
                                              1526-4_37.

Wolff:1976:CRO

[Wol76x] Kurt H. Wolff. On the cunning of reason in our time. In Surrender and Catch: Experience and Inquiry Today [Wol76-32], chapter 20, pages 144–151. ISBN 90-277-0765-0, 94-010-1526-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-010-1526-
                                              4_20.

Wolff:1976:PN

[Wol76y] Kurt H. Wolff. Prefatory note. In Surrender and Catch: Experience and Inquiry Today [Wol76-32], chapter 17, pages 113–114. ISBN 90-277-0765-0, 94-010-1526-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-010-
1526-4_17.

Wolff:1976:PPS

[Wol76z] Kurt H. Wolff. Publishing papers by students and identifying their authors. In Surrender and Catch: Experience and Inquiry Today [Wol76-32], chapter 31, page 289. ISBN 90-277-0765-0, 94-010-1526-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic).

LCCN ????  URL https://link.
springer.com/chapter/10.1007/
978-94-010-1526-4_31.

Wolff:1976:RT

[Wol76-27] Kurt H. Wolff. Recovery: Trying with others. In Surrender and Catch: Experience and Inquiry Today [Wol76-32], chapter 25, pages 217–218. ISBN 90-277-0765-0, 94-010-1526-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-010-1526-4_25.

Wolff:1976:SPS

[Wol76-28] Kurt H. Wolff. Sociology, phenomenology, and surrender — and — catch. In Surrender and Catch: Experience and Inquiry Today [Wol76-32], chapter 22, pages 153–176. ISBN 90-277-0765-0, 94-010-1526-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-010-
1526-4_22.

Wolff:1976:SAE

[Wol76-29] Kurt H. Wolff. Surrender and aesthetic experience. In Surrender and Catch: Experience and Inquiry Today [Wol76-32], chapter 15, pages 96–108. ISBN 90-277-0765-0, 94-010-1526-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-010-1526-
                                              4_15.

Wolff:1976:SCa

[Wol76-30] Kurt H. Wolff. Surrender and catch. In Surrender and Catch: Experience and Inquiry Today [Wol76-32], chapter 2, pages 9–17. ISBN 90-277-0765-0, 94-010-1526-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

                                              ???? URL https://link.springer.
                                              com/chapter/10.1007/978-94-010-
                                              1526-4_2.

Wolff:1976:SCb

[Wol76-31] Kurt H. Wolff. ‘Surrender’ and ‘catch’. In Surrender and Catch: Experience and Inquiry Today [Wol76-32], chapter 4, pages 20–27. ISBN 90-277-0765-0, 94-010-1526-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic).

                                              LCCN ????  URL https://link.
                                              springer.com/chapter/10.1007/
                                              978-94-010-1526-4_4.

Wolff:1976:SCE

[Wol76-32] Kurt H. Wolff. Surrender and Catch: Experience and Inquiry Today, volume 51 of Boston Studies in the Philosophy of Science. D. Reidel, Dordrecht, The Netherlands; Boston, MA, USA; Lancaster, UK; Tokyo, Japan, 1976. ISBN 90-277-0765-0, 94-010-1526-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). 422 (est.) pp. LCCN ???? URL

https://link.springer.com/book/
10.1007/978-94-010-1526-4.

Wolff:1976:SCS

[Wol76-33] Kurt H. Wolff. Surrender and community study: The study of Loma. In Surrender and Catch: Experience and Inquiry Today [Wol76-32], chapter 13, pages 71–93. ISBN 90-277-0765-0, 94-010-1526-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-010-
1526-4_13.

Wolff:1976:SRb

[Wol76-34] Kurt H. Wolff. Surrender and rebellion.

In Surrender and Catch: Experience and Inquiry Today [Wol76-32], chapter 11, pages 49–69. ISBN 90-277-0765-0, 94-010-1526-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-010-1526-4_11.

Wolff:1976:SRa

[Wol76-35] Kurt H. Wolff. Surrender and religion.

In Surrender and Catch: Experience and Inquiry Today [Wol76-32], chapter 8, pages 37–42. ISBN 90-277-0765-0, 94-010-1526-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic).

LCCN ????  URL https://link.
springer.com/chapter/10.1007/
978-94-010-1526-4_8.

Wolff:1976:SB

[Wol76-36] Kurt H. Wolff. Surrender and the body. In Surrender and Catch: Experience and Inquiry Today [Wol76-32], chapter 24, pages 179–213. ISBN 90-277-0765-0, 94-010-1526-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-010-1526-4_24.

Wolff:1976:SRO

[Wol76-37] Kurt H. Wolff. Surrender as a response to our crisis. In Surrender and Catch: Experience and Inquiry Today [Wol76-32], chapter 6, pages 30–35. ISBN 90-277-0765-0, 94-010-1526-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-010-1526-
                                              4_6.

Wolff:1976:TB

[Wol76-38] Kurt H. Wolff. There is a beginning. In Surrender and Catch: Experience and Inquiry Today [Wol76-32], chapter 1, pages 3–6. ISBN 90-277-0765-0, 94-010-1526-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

                                              ???? URL https://link.springer.
                                              com/chapter/10.1007/978-94-010-
                                              1526-4_1.

Wolff:1976:B

[Wol76-39] Kurt H. Wolff. (this book). In Surrender and Catch: Experience and Inquiry Today [Wol76-32], chapter 10, page 45. ISBN 90-277-0765-0, 94-010-1526-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-010-1526-
                                              4_10.

Wolff:1976:TLL

[Wol76-40] Kurt H. Wolff. Too literal, not literal enough. In Surrender and Catch: Experience and Inquiry Today [Wol76-32], chapter 3, pages 18–19. ISBN 90-277-0765-0, 94-010-1526-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-010-1526-4_3.

Wolniewicz:1986:LFP

[Wol86] Bogusław Wolniewicz. Ludwik Fleck and Polish philosophy. In Cognition and Fact: Materials on Ludwik Fleck [CS86], chapter 11, pages 217–221. ISBN 94-010-8504-8, 94-009-4498-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-009-4498-
5_11.

Wolters:1989:MMB

[Wol89] Gereon Wolters. Mesmer in a mountain bar: Anthropological difference, Butts, and Mesmerism. In An Intimate Relation: Studies in the History and Philosophy of Science Presented to Robert E. Butts on his 60th Birthday [BM89], chapter 13, pages 259–282. ISBN 94-010-7546-8, 94-009-2327-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-009-2327-
0_13.

Wolff:1995:ALL

[Wol95a] Kurt H. Wolff. “Authenticity in Loma and of ‘Loma’ ”. In Transformation in the Writing: A Case of Surrenderand-Catch [Wol95l], chapter 3, pages 29–43. ISBN 90-481-4478-7, 94-015-8412-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN B829.5.A-829.5.Z. URL https://

link.springer.com/chapter/10.
1007/978-94-015-8412-8_3.

Wolff:1995:FNS

[Wol95b] Kurt H. Wolff. First notes on surrender and autobiography. In Transformation in the Writing: A Case of Surrenderand-Catch [Wol95l], chapter 2, pages 16–28. ISBN 90-481-4478-7, 94-015-8412-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN B829.5.A-829.5.Z. URL https://

link.springer.com/chapter/10.
1007/978-94-015-8412-8_2.

Wolff:1995:NS

[Wol95c] Kurt H. Wolff. From nothing to sociology. In Transformation in the Writing: A Case of Surrender-and-Catch [Wol95l], chapter 11, pages 169–193. ISBN 90-481-4478-7, 94-015-8412-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN B829.5.A-829.5.Z. URL https://

link.springer.com/chapter/10.
1007/978-94-015-8412-8_11.

Wolff:1995:NEF

[Wol95d] Kurt H. Wolff. “Nineteen-Eighty-Four” and surrender-and-catch. In Transformation in the Writing: A Case of Surrender-and-Catch [Wol95l], chapter 12, pages 194–203. ISBN 90-481-4478-7, 94-015-8412-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN B829.5.A-829.5.Z.

URL https://link.springer.com/
chapter/10.1007/978-94-015-8412-
8_12.

Wolff:1995:SKS

[Wol95e] Kurt H. Wolff. The sociology of knowledge and surrender-and-catch. In Transformation in the Writing: A Case of Surrender-and-Catch [Wol95l], chapter 5, pages 55–67. ISBN 90-481-4478-7, 94-015-8412-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN B829.5.A-829.5.Z.

URL https://link.springer.com/
chapter/10.1007/978-94-015-8412-
8_5.

Wolff:1995:SCC

[Wol95f] Kurt H. Wolff. Surrender-and-catch and critical theory. In Transformation in the Writing: A Case of Surrenderand-Catch [Wol95l], chapter 10, pages 149–166. ISBN 90-481-4478-7, 94-015-8412-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN B829.5.A-829.5.Z. URL https://

link.springer.com/chapter/10.
1007/978-94-015-8412-8_10.

Wolff:1995:SCH

[Wol95g] Kurt H. Wolff. Surrender-and-catch and hermeneutics. In Transformation in the Writing: A Case of Surrenderand-Catch [Wol95l], chapter 9, pages 131–148. ISBN 90-481-4478-7, 94-015-8412-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN B829.5.A-829.5.Z. URL https://

link.springer.com/chapter/10.
1007/978-94-015-8412-8_9.

Wolff:1995:SCPb

[Wol95h] Kurt H. Wolff. Surrender-and-catch and phenomenology. In Transformation in the Writing: A Case of Surrenderand-Catch [Wol95l], chapter 8, pages 112–130. ISBN 90-481-4478-7, 94-015-8412-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN B829.5.A-829.5.Z. URL https://

link.springer.com/chapter/10.
1007/978-94-015-8412-8_8.

Wolff:1995:SCS

[Wol95i] Kurt H. Wolff. Surrender-and-catch and sociology. In Transformation in the Writing: A Case of Surrenderand-Catch [Wol95l], chapter 6, pages 68–91. ISBN 90-481-4478-7, 94-015-8412-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN B829.5.A-829.5.Z. URL https://

link.springer.com/chapter/10.
1007/978-94-015-8412-8_6.

Wolff:1995:SCPa

[Wol95j] Kurt H. Wolff. Surrender-and-catch, poetry, sociology, morality. In Transformation in the Writing: A Case of Surrender-and-Catch [Wol95l], chapter 7, pages 92–111. ISBN 90-481-4478-7, 94-015-8412-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN B829.5.A-829.5.Z.

URL https://link.springer.com/
chapter/10.1007/978-94-015-8412-
8_7.

Wolff:1995:PHE

[Wol95k] Kurt H. Wolff. “The personal history of an Émigré”. In Transformation in the Writing: A Case of Surrenderand-Catch [Wol95l], chapter 1, pages 3–15. ISBN 90-481-4478-7, 94-015-8412-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN B829.5.A-829.5.Z. URL https://

link.springer.com/chapter/10.
1007/978-94-015-8412-8_1.

Wolff:1995:TWC

[Wol95l] Kurt H. Wolff. Transformation in the Writing: A Case of Surrender-and-Catch, volume 166 of Boston Studies in the Philosophy of Science. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 1995. ISBN 90-481-4478-7, 94-015-8412-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). xxiii + 223 pp. LCCN B829.5.A-829.5.Z. URL

https://link.springer.com/book/
10.1007/978-94-015-8412-8.

Wolff:1995:TSS

[Wol95m] Kurt H. Wolff. “Trying sociology,” sociology of knowledge, surrender-and-catch. In Transformation in the Writing: A Case of Surrender-and-Catch [Wol95l], chapter 4, pages 44–52. ISBN 90-481-4478-7, 94-015-8412-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN B829.5.A-829.5.Z.

URL https://link.springer.com/
chapter/10.1007/978-94-015-8412-
8_4.

Wolenski:1997:HWK

[Wol97]  Jan Woleński.  Haller on Wiener

Kreis. In Austrian Philosophy Past and Present: Essays in Honor of Rudolf Haller [LM97], chapter 4, pages 45–54. ISBN 94-010-6412-1, 94-011-5720-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-011-5720-
9_4.

Wolters:2003:HEC

[Wol03]  Gereon Wolters.  O happy error:

a comment on Giora Hon. In Descartes’s Mathematical Thought [Sas03d], chapter 20, pages 295–299. ISBN 90-481-6487-7, 94-017-1225-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/0-306-48123-5_20.

See [Hon03].

Wolenski:2017:SGT

[Wol17] Jan Woleński. The story of the German translation of Tarski’s book on truth. In Encouraging Openness: Essays for Joseph Agassi on the Occasion of His 90th Birthday [BAG17], chapter 8, pages 81–90. ISBN 3-319-57668-2 (hardcover), 3-319-57669-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN B833.E53 2017.

URL https://link.springer.com/

                                              chapter/10.1007/978-3-319-57669-
                                              5_8.

Wolfe:2018:LME

[Wol18] Charles T. Wolfe. From Locke to materialism: Empiricism, the brain and the stirrings of ontology. In What Does it Mean to be an Empiricist?: Empiricisms in Eighteenth Century Sciences [BR18b], chapter 13, pages 235–263. ISBN 3-319-69858-3 (hardcover), 3-319-69860-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN B816 .W43 2018. URL https://link.

                                              springer.com/chapter/10.1007/
                                              978-3-319-69860-1_13.

Woodward:1991:WVSa

[Woo91] William R. Woodward. World views and scientific discipline formation: How East German science studies contributed to the fall of the cultural wall. In Woodward and Cohen [WC91], pages 1–15. ISBN 94-010-5395-2, 94-011-3164-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

                                              ???? URL https://link.springer.
                                              com/chapter/10.1007/978-94-011-
                                              3164-3_1.

Woodruff:2000:IOH

[Woo00] Peter Woodruff. Indefinite objects of higher order. In Knowledge, Language and Logic: Questions for Quine [OK00], chapter 17, pages 225–235. ISBN 1-4020-0253-X, 94-011-3933-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-011-3933-
                                              5_17.

Woodbridge:2002:RWS

[Woo02] F. J. E. Woodbridge. The realists who sharpened the edges of Burtt’s ideas.

In E. A. Burtt, Historian and Philosopher: A Study of the Author of The Metaphysical Foundations of Modern Physical Science [Vil02g], chapter 4, pages 31–45. ISBN 90-481-5937-7, 94-017-1331-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-017-
1331-3_4.

Woodyard:2006:NPS

[Woo06] Jack R. Woodyard. A new paradigm for Schrödinger and Kohn. In Philosophy of Chemistry: Synthesis of a New Discipline [BSM06b], chapter 13, pages 245–269. ISBN 1-4020-2455-X, 1-4020-3256-0, 1-4020-3261-7. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 242 QD6. URL https:

//link.springer.com/chapter/10.
1007/1-4020-3261-7_13.

Worrall:1976:ILP

[Wor76] John Worrall. Imre Lakatos (1922– 1974): Philosopher of mathematics and philosopher of science. In Essays in Memory of Imre Lakatos [CFW76], chapter 1, pages 1–8. ISBN 90-277-0655-7, 94-010-1451-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-010-1451-9_1.

Worrall:1978:RPE

[Wor78a] John Worrall. Research programmes, empirical support, and the Duhem problem: Replies to criticism. In Progress and Rationality in Science [RA78b], chapter 15, pages 321–338. ISBN 90-277-0922-X, 94-009-9866-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/

                                              chapter/10.1007/978-94-009-9866-
                                              7_15.

Worrall:1978:WWM

[Wor78b] John Worrall. The ways in which the methodology of scientific research programmes improves on Popper’s methodology. In Progress and Rationality in Science [RA78b], chapter 3, pages 45–70. ISBN 90-277-0922-X, 94-009-9866-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

                                              ???? URL https://link.springer.
                                              com/chapter/10.1007/978-94-009-
                                              9866-7_3.

Worrall:1989:WBP

[Wor89] John Worrall. Why both Popper and Watkins fail to solve the problem of induction. In Freedom and Rationality: Essays in Honor of John Watkins From his Colleagues and Friends [DJ89], chapter 16, pages 257–296. ISBN 94-010-7571-9, 94-009-2380-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-009-2380-5_16.

Worrall:1991:FF

[Wor91] John Worrall. Feyerabend and the facts. In Beyond Reason: Essays on the Philosophy of Paul Feyerabend [Mun91a], chapter 15, pages 329–353. ISBN 94-010-5406-1, 94-011-3188-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-011-3188-
                                              9_15.

Woodward:1994:RNT

[WP94] William R. Woodward and Reinhardt Pester. From romantic Naturphilosophie to a theory of scientific method for the medical disciplines. In Poggi and Bossi [PB94], chapter 10, pages 161–173. ISBN 0-7923-2336-X, 94-017-2921-2. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 152. URL https://link.

springer.com/chapter/10.1007/
978-94-017-2921-5_10.

Wallner:1996:CSE

[WP96] F. Wallner and M. F. Peschl. Cognitive science — an experiment in constructive realism; constructive realism — an experiment in cognitive science. In Cohen et al. [CHR96], pages 103–116. ISBN 90-481-4493-0, 94-015-8638-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-015-8638-
2_8.

Wilson:1995:NTS

[WPJ95] Kate J. Wilson, Mark B. Peoples, and Richard A. Jefferson. New techniques for studying competition by Rhizobia and for assessing nitrogen fixation in the field. In No Truth Except in the Details: Essays in Honor of Martin J. Klein [KS95], chapter 12, pages 241–253. ISBN 94-010-4097-4, 94-011-0217-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-011-
0055-7_12.

Wray:2015:KSE

[Wra15] K. Brad Wray. Kuhn’s social epistemology and the sociology of science. In Kuhn’s Structure of Scientific Revolutions — 50 Years On [DB15b], chapter 12, pages 167–183. ISBN 3-319-13382-9, 3-319-13383-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175.3 .K84 2015.

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-3-319-13383-
                                              6_12.

Wright:1976:F

[Wri76] Larry Wright. Functions. In Topics in the Philosophy of Biology [GM76], chapter 10, pages 213–242. ISBN 90-277-0596-8, 94-010-1829-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-010-1829-6_10.

Wrightsman:1980:LSB

[Wri80] Bruce Wrightsman. The legitimation of scientific belief: Theory justification by Copernicus. In Scientific Discovery: Case Studies [Nic80e], chapter 4, pages 51–66. ISBN 90-277-1093-7, 94-009-9015-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 60. URL https:/

                                              /link.springer.com/chapter/10.
                                              1007/978-94-009-9015-9_4.

Wright:1996:MRE

[Wri96] John Wright. Metaphysical realism and the explanation of the success of science. In Cohen et al. [CHR96], pages 227–243. ISBN 90-481-4493-0, 94-015-8638-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

                                              ???? URL https://link.springer.
                                              com/chapter/10.1007/978-94-015-
                                              8638-2_16.

Wani:1995:SAS

[WRL95] S. P. Wani, O. P. Rupela, and K. K. Lee. Sustainable agriculture in the semi-arid tropics through biological nitrogen fixation in grain legumes. In No Truth Except in the Details: Essays in Honor of Martin J. Klein [KS95], chapter 2, pages 29–49. ISBN 94-010-4097-4, 94-011-0217-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-011-0055-7_2.

Wrong:1988:CAR

[Wro88] Dennis H. Wrong. The concept of alienation revisited. In Philosophy, History and Social Action: Essays in Honor of Lewis Feuer with an autobiographic essay by Lewis Feuer [HOO88c], chapter 25, pages 461–473. ISBN 94-010-7793-2, 94-009-2873-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-009-2873-
2_25.

Wulun:1996:DST

[Wul96] Jin Wulun. Double structure of tao reality. In Cohen et al. [CHR96], pages 317–328. ISBN 90-481-4493-0, 94-015-8638-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-015-
8638-2_23.

Wunderlich:1976:TIT

[Wun76] Dieter Wunderlich. Towards an integrated theory of grammatical and pragmatical meaning. In Language in Focus: Foundations, Methods and Systems: Essays in Memory of Yehoshua Bar-Hillel [Kas76b], chapter 14, pages 251–277. ISBN 90-277-0644-1, 90-277-0645-X, 94-010-1876-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN P101-120. URL https:

//link.springer.com/chapter/10.
1007/978-94-010-1876-0_14.

Wunderlich:2001:KSP

[Wun01] Falk Wunderlich. Kant’s second paralogism in context. In Between Leibniz, Newton, and Kant: Philosophy and Science in the Eighteenth Century [Lef01a], chapter 9, pages 175–188. ISBN 90-481-5774-9, 94-015-9729-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-015-9729-
                                              6_9.

Wussing:1991:HMA

[Wuß91]  Hans Wußing.  Historiography of

mathematics: Aims, methods, tasks. In Woodward and Cohen [WC91], pages 63–73. ISBN 94-010-5395-2, 94-011-3164-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

                                              ???? URL https://link.springer.
                                              com/chapter/10.1007/978-94-011-
                                              3164-3_6.

Wuthrich:2016:HUC

[Wüt16] Adrian Wüthrich. Heisenberg’s Umdeutung: A case for a (quantum-)dialogue between history and philosophy of science. In Sauer and Scholl [SS16b], chapter 14, pages 285–296. ISBN 3-319-30227-2, 3-319-30229-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q126.8 .P45 2016.

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-3-319-30229-
                                              4_14.

Weber:2007:AEP

[WvB07] Erik Weber and Jeroen van Bouwel. Assessing the explanatory power of causal explanations. In Rethinking Explanation [PY07], chapter 7, pages 109–118. ISBN 1-4020-5580-3 (paperback), 1-4020-5581-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN BD237 .R48 2007; Q175.32.E97 R48 2007. URL https:

//link.springer.com/chapter/10.
1007/978-1-4020-5581-2_7.

Weinstein:1990:DCG

[WW90] Deena Weinstein and Michael A. Weinstein. Dimensions of conflict: Georg Simmel on modern life. In Georg Simmel and Contemporary Sociology [KPC90], chapter 19, pages 341–355. ISBN 94-010-6691-4 (hardcover), 94-009-0459-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN HM401-1281. URL https://link.

springer.com/chapter/10.1007/
978-94-009-0459-0_19.

Wesley:2003:TAD

[WW03] Daniel H. Wesley and John A. Wheeler.

Towards an action-at-a-distance concept of spacetime. In Renn et al. [RDS03], chapter 21, pages 421–436. ISBN 1-4020-1284-5, 1-4020-1285-3 (paperback). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 234. URL https://link.

springer.com/chapter/10.1007/
978-94-010-0111-3_21.

Wylie:1992:HDR

[Wyl92] Alison Wylie. On “heavily decomposing red herrings”: Scientific method in archaeology and the ladening of evidence with theory. In Metaarchaeology: Reflections by Archaeologists and Philosophers [Emb92b], chapter 12, pages 269–288. ISBN 94-010-4806-1, 94-011-1826-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-011-
1826-2_12.

Wylie:2015:PPC

[Wyl15] Alison Wylie. A plurality of pluralisms: Collaborative practice in archaeology. In Objectivity in Science: New Perspectives from Science and Technology Studies [PRT15], chapter 10, pages 189–210. ISBN 3-319-14348-4 (hardcover), 3-319-14349-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175 O25 2015.

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-3-319-14349-
                                              1_10.

Walther:1976:SL

[WZ76] Eric Walther and Eddy M. Zemach. Substance logic. In Language in Focus: Foundations, Methods and Systems: Essays in Memory of Yehoshua Bar-Hillel [Kas76b], chapter 4, pages 55–74. ISBN 90-277-0644-1, 90-277-0645-X, 94-010-1876-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN P101-120. URL https://link.

                                              springer.com/chapter/10.1007/
                                              978-94-010-1876-0_4.

Xiucheng:1996:CSM

[Xiu96] Le Xiucheng. A comparison of the structures of the mathematical systems of China and the West: Several revelations from information theory. In Chinese Studies in the History and Philosophy of Science and Technology [DC96], chapter 20, pages 255–260. ISBN 90-481-4546-5, 94-015-8717-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 180; U390 .N19 1996.

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-015-8717-
                                              4_20.

Yalan:1996:EAM

[Yal96] Chen Yalan. The effect of the autocratic monarchy of the Qing Dynasty on science and technology. In Chinese Studies in the History and Philosophy of Science and Technology [DC96], chapter 26, pages 315–326. ISBN 90-481-4546-5, 94-015-8717-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 180; U390 .N19 1996. URL https:

//link.springer.com/chapter/10.
1007/978-94-015-8717-4_26.

Yalcin:2005:QRR

[Yal05] Ümıt D. Yalçin. Quine’s robust relativism. In Turkish Studies in the History and Philosophy of Science [IG05c], chapter 6, pages 71–85. ISBN 1-4020-3332-X (hardcover), 1-4020-3333-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 244. URL https:/

/link.springer.com/chapter/10.
1007/1-4020-3333-8_6.

Yang:1993:DGR

[Yan93] Tsui-Hua Yang. The development of geology in Republican China, 1912– 1937. In Philosophy and Conceptual History of Science in Taiwan [LF93], chapter 12, pages 221–244. ISBN 94-010-5103-8, 94-011-2500-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q127.T28 .P455

1993. URL https://link.springer.
com/chapter/10.1007/978-94-011-
2500-0_12.

Yanase:1998:CMN

[Yan98] M. M. Yanase. Comment on the Machida–Namiki–Araki theory. In Japanese Studies in the Philosophy of Science [NC98], chapter 10, pages 151–157. ISBN 94-010-6176-9, 94-011-5175-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

                                              ???? URL https://link.springer.
                                              com/chapter/10.1007/978-94-011-
                                              5175-7_10.

Yavetz:2017:TCS

[Yav17] Ido Yavetz. Trigonometry, construction by straightedge and compass, and the applied mathematics of the Almagest. In Mathematics as a Tool: Tracing New Roles of Mathematics in the Sciences [LC17b], chapter 4, pages 57– 67. ISBN 3-319-54468-3 (hardcover), 3-319-54469-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN GE45.M37; GE45.M38; QA8.4.

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-3-319-54469-
                                              4_4.

Yuan:2018:RPT

[YCMY18] Deyu Yuan, Dongming Cao, Carl Mitcham, and Wenjuan Yin. Roots of the philosophy of technology in China. In Mitcham et al. [MLNZ18], chapter 24, pages 283–299. ISBN 3-319-62448-2, 3-319-62450-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN TA15 .P45 2018.

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-3-319-62450-
                                              1_24.

Yehezkely:2017:CNO

[Yeh17] Chen Yehezkely. Civil nationalism: The only buffer against fascism and fundamentalism. In Encouraging Openness: Essays for Joseph Agassi on the Occasion of His 90th Birthday [BAG17], chapter 27, pages 319–332. ISBN 3-319-57668-2 (hardcover), 3-319-57669-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN B833.E53 2017. URL https://link.

                                              springer.com/chapter/10.1007/
                                              978-3-319-57669-5_27.

Yilmaz:1974:PPS

[Yil74]  Hüseyin Yilmaz.  Perception and

philosophy of science. In Logical and Epistemological Studies in Contemporary Physics [CW74a], chapter 1, pages 1–91. ISBN 90-277-0391-4 (hardcover), 90-277-0377-9 (paperback), 94-010-2656-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 13; QC5.56.

URL https://link.springer.com/
chapter/10.1007/978-94-010-2656-
7_1.

Yilmaz:2005:GTR

[Yil05] Hüseyın Yilmaz. General theory of relativity and the 5th test. In Turkish Studies in the History and Philosophy of Science [IG05c], chapter 4, pages 39–58. ISBN 1-4020-3332-X (hardcover), 1-4020-3333-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 244.

URL https://link.springer.com/
chapter/10.1007/1-4020-3333-8_4.

Yinon:2016:COD

[Yin16] Dror Yinon. Change’s order: On Deleuze’s notion of time. In Cosmological and Psychological Time [DR16], chapter 12, pages 203–218. ISBN 3-319-22589-8 (hardcover), 3-319-22590-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN BD638 C67

2016. URL https://link.springer.
com/chapter/10.1007/978-3-319-
22590-6_12.

Yin:2018:EPE

[Yin18] Ruiyu Yin. From engineering to the philosophy of engineering: Philosophical reflections of an engineer. In Mitcham et al. [MLNZ18], chapter 7, pages 75–81. ISBN 3-319-62448-2, 3-319-62450-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN TA15 .P45 2018. URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-3-319-62450-1_7.

Ylikoski:2007:ICE

[Yli07] Petri Ylikoski. The idea of contrastive explanandum. In Rethinking Explanation [PY07], chapter 3, pages 27–42. ISBN 1-4020-5580-3 (paperback), 1-4020-5581-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN BD237 .R48 2007; Q175.32.E97 R48

                                              2007. URL https://link.springer.
                                              com/chapter/10.1007/978-1-4020-
                                              5581-2_3.

Yosida:1998:SLT

[Yos98] Natuhiko Yosida. Scientific laws as tools for taxonomy. In Japanese Studies in the Philosophy of Science [NC98], chapter 6, pages 87–98. ISBN 94-010-6176-9, 94-011-5175-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-011-5175-7_6.

Yourgrau:1969:VPU

[You69] Wolfgang Yourgrau. Verification or proof — an undecided issue? In Cohen and Wartofsky [CW69a], pages 370–399. ISBN 94-010-3383-8, 94-010-3263-7, 94-010-3381-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-010-3381-7_12.

Youschkevitch:1974:NFC

[You74] A. P. Youschkevitch. La notion de fonction chez Condorcet. (French) [The concept of function in Condorcet]. In For Dirk Struik: Scientific, Historical and Political Essays in Honor of Dirk J. Struik [CSW74], chapter 11, pages 131–139. ISBN 90-277-0379-5, 94-010-2115-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-010-
2115-9_11.

Yourgrau:1976:CLI

[You76] Wolfgang Yourgrau. Cosmology and logic — an intractable issue? In Essays in Memory of Imre Lakatos [CFW76], chapter 40, pages 739–753. ISBN 90-277-0655-7, 94-010-1451-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-010-1451-9_40.

Young:1995:CSR

[You95] Alfred F. Young. Common sense and the rights of man in America. In Science, Mind and Art: Essays on science and the humanistic understanding in art, epistemology, religion and ethics In honor of Robert S. Cohen [GSW95b], chapter 25, pages 411–439. ISBN 0-7923-2990-2 (hardcover), 94-011-0469-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 165 BD418.3. URL https:

//link.springer.com/chapter/10.
1007/978-94-011-0469-2_25.

Yturbe:1995:HSI

[Ytu95] Corina Yturbe. The history of science:

Internal or external? In Mexican Studies in the History and Philosophy of Science [RC95], chapter 6, pages 71–87. ISBN 94-010-6535-7 (hardcover), 94-009-0109-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-009-
0109-4_6.

Yusheng:1996:HHF

[Yus96a] Wang Yusheng. Hua Hengfang: Forerunner and disseminator of modern science in China. In Chinese Studies in the History and Philosophy of Science and Technology [DC96], chapter 29, pages 369–393. ISBN 90-481-4546-5, 94-015-8717-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 180; U390 .N19

                                              1996. URL https://link.springer.
                                              com/chapter/10.1007/978-94-015-
                                              8717-4_29.

Yusheng:1996:LSF

[Yus96b] Wang Yusheng. Li Shanlan: Forerunner of modern science in China. In Chinese Studies in the History and Philosophy of Science and Technology [DC96], chapter 28, pages 345–368. ISBN 90-481-4546-5, 94-015-8717-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 180; U390 .N19 1996. URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-015-8717-4_28.

Yuzhi:1996:HDC

[Yuz96] Gong Yuzhi. Historical development of the Chinese Communist Party’s scientific policy (prior to the founding of the People’s Republic). In Chinese Studies in the History and Philosophy of Science and Technology [DC96], chapter 3, pages 13–25. ISBN 90-481-4546-5, 94-015-8717-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 180; U390 .N19 1996. URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-015-8717-4_3.

Zabierowski:1995:OPI

[Zab95] Mirosław Zabierowski. On the objectivity of the Popperian interpretation of quantum mechanics. In Misiek [Mis95c], pages 231–237. ISBN 94-010-4206-3, 94-011-0461-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-011-0461-6_18.

Zaddach:1996:RPB

[Zad96] Arno Zaddach. Regressive products and Bourbaki. In Schubring [Sch96f], pages 285–295. ISBN 90-481-4758-1, 94-015-8753-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN QA8.9-10.3. URL https:/

/link.springer.com/chapter/10.
1007/978-94-015-8753-2_24.

Zahar:1978:CEC

[Zah78] Elie Zahar. ‘Crucial’ experiments: A case study. In Progress and Rationality in Science [RA78b], chapter 4, pages 71–97. ISBN 90-277-0922-X, 94-009-9866-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-009-
9866-7_4.

Zahar:1989:JWE

[Zah89] Elie Zahar. John Watkins on the empirical basis and the corroboration of scientific theories. In Freedom and Rationality: Essays in Honor of John Watkins From his Colleagues and Friends [DJ89], chapter 18, pages 325–341. ISBN 94-010-7571-9, 94-009-2380-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-009-
2380-5_18.

Zajonc:1987:FTA

[Zaj87] Arthur G. Zajonc. Facts as theory: Aspects of Goethe’s philosophy of science.

In Amrine et al. [AZW87], chapter 12, pages 219–245. ISBN 90-277-2265-X, 94-009-3761-X. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 97. US$79.00. URL https:

//link.springer.com/chapter/10.
1007/978-94-009-3761-1_12.

Zalc:1986:SCF

[Zal86] Bernard Zalc. Some comments on Fleck’s interpretation of the Bordet– Wassermann reaction in view of present biochemical knowledge. In Cognition and Fact: Materials on Ludwik Fleck [CS86], chapter 20, pages 399–406. ISBN 94-010-8504-8, 94-009-4498-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-009-4498-
5_20.

Zambelli:1992:AAC

[Zam92a] Paola Zambelli. Albert’s ‘Auctoritas’: Contemporaries and collaborators. In The Speculum astronomiae and its enigma: astrology, theology, and science in Albertus Magnus and his contemporaries [Zam92h], chapter 5, pages 45–50. ISBN 0-7923-1380-1 (hardcover), 94-017-3467-4. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 135. URL https:/

/link.springer.com/chapter/10.
1007/978-94-017-3467-7_5.

Zambelli:1992:AB

[Zam92b] Paola Zambelli. Albert’s biblionomia. In The Speculum astronomiae and its enigma: astrology, theology, and science in Albertus Magnus and his contemporaries [Zam92h], chapter 11, pages 105–111. ISBN 0-7923-1380-1 (hardcover), 94-017-3467-4. ISSN 0068-0346 (print), 2214-7942 (electronic).

LCCN Q174 .B67 vol. 135. URL https:

//link.springer.com/chapter/10.
1007/978-94-017-3467-7_11.

Zambelli:1992:DDS

[Zam92c] Paola Zambelli. Are “deaf and dumb”

stars and their movers at the origins of modern science? Another historiographical case-study. In The Speculum astronomiae and its enigma: astrology, theology, and science in Albertus Magnus and his contemporaries [Zam92h], chapter 8, pages 75–81. ISBN 0-7923-1380-1 (hardcover), 94-017-3467-4. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 135.

URL https://link.springer.com/
chapter/10.1007/978-94-017-3467-
7_8.

Zambelli:1992:AAU

[Zam92d] Paola Zambelli. Astrology in Albert’s undisputed works. In The Speculum astronomiae and its enigma: astrology, theology, and science in Albertus Magnus and his contemporaries [Zam92h], chapter 7, pages 61–74. ISBN 0-7923-1380-1 (hardcover), 94-017-3467-4. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 135.

URL https://link.springer.com/
chapter/10.1007/978-94-017-3467-
7_7.

Zambelli:1992:AED

[Zam92e] Paola Zambelli. Astrology in the Early Dominican School and Gerard of Feltre. In The Speculum astronomiae and its enigma: astrology, theology, and science in Albertus Magnus and his contemporaries [Zam92h], chapter 6, pages 51–59. ISBN 0-7923-1380-1 (hardcover), 94-017-3467-4. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 135. URL https:/

/link.springer.com/chapter/10.
1007/978-94-017-3467-7_6.

Zambelli:1992:C

[Zam92f] Paola Zambelli. Conclusion. In The Speculum astronomiae and its enigma: astrology, theology, and science in Albertus Magnus and his contemporaries [Zam92h], chapter 13, pages 121–125. ISBN 0-7923-1380-1 (hardcover), 94-017-3467-4. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 135. URL https://link.

springer.com/chapter/10.1007/
978-94-017-3467-7_13.

Zambelli:1992:DPM

[Zam92g] Paola Zambelli. Divine Providence and the meaning of “interrogationes”. In The Speculum astronomiae and its enigma: astrology, theology, and science in Albertus Magnus and his contemporaries [Zam92h], chapter 10, pages 95–101. ISBN 0-7923-1380-1 (hardcover), 94-017-3467-4. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 135. URL https:

//link.springer.com/chapter/10.
1007/978-94-017-3467-7_10.

Zambelli:1992:SAE

[Zam92h] Paola Zambelli. The Speculum astronomiae and its enigma: astrology, theology, and science in Albertus Magnus and his contemporaries, volume 135 of Boston Studies in the Philosophy of Science. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 1992. ISBN 0-7923-1380-1 (hardcover), 94-017-3467-4. ISSN 0068-0346 (print), 2214-7942 (electronic). xvi + 352 pp. LCCN Q174

.B67 vol. 135. URL http://www.loc.
gov/catdir/enhancements/fy0826/
91024566-d.html;   http://www.

loc.gov/catdir/enhancements/
fy0826/91024566-t.html;   https:
//link.springer.com/book/10.
1007/978-94-017-3467-7.

Zambelli:1992:FCD

[Zam92i] Paola Zambelli. Further condemnations, debates and “consultationes”. In The Speculum astronomiae and its enigma: astrology, theology, and science in Albertus Magnus and his contemporaries [Zam92h], chapter 2, pages 11–23. ISBN 0-7923-1380-1 (hardcover), 94-017-3467-4. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 135. URL https:/

/link.springer.com/chapter/10.
1007/978-94-017-3467-7_2.

Zambelli:1992:LTS

[Zam92j] Paola Zambelli. The literary tradition of the speculum and its role as a reference book. In The Speculum astronomiae and its enigma: astrology, theology, and science in Albertus Magnus and his contemporaries [Zam92h], chapter 12, pages 113–120. ISBN 0-7923-1380-1 (hardcover), 94-017-3467-4. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 135. URL https:

//link.springer.com/chapter/10.
1007/978-94-017-3467-7_12.

Zambelli:1992:MSA

[Zam92k] Paola Zambelli. Mandonnet, the speculum astronomiae and the condemnation of 1277. In The Speculum astronomiae and its enigma: astrology, theology, and science in Albertus Magnus and his contemporaries [Zam92h], chapter 1, pages 3–9. ISBN 0-7923-1380-1 (hardcover), 94-017-3467-4. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 135. URL https:/

/link.springer.com/chapter/10.
1007/978-94-017-3467-7_1.

Zambelli:1992:MHA

[Zam92l] Paola Zambelli. Mandonnet’s hypothesis: Acquiescence and doubts. In The Speculum astronomiae and its enigma: astrology, theology, and science in Albertus Magnus and his contemporaries [Zam92h], chapter 3, pages 25–31. ISBN 0-7923-1380-1 (hardcover), 94-017-3467-4. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 135. URL https://link.

                                              springer.com/chapter/10.1007/
                                              978-94-017-3467-7_3.

Zambelli:1992:HGA

[Zam92m] Paola Zambelli. Not the heavens, but God alone is endowed with life and the stars are simply his instruments. In The Speculum astronomiae and its enigma: astrology, theology, and science in Albertus Magnus and his contemporaries [Zam92h], chapter 9, pages 83–94. ISBN 0-7923-1380-1 (hardcover), 94-017-3467-4. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 135. URL https:/

                                              /link.springer.com/chapter/10.
                                              1007/978-94-017-3467-7_9.

Zambelli:1992:TCH

[Zam92n] Paola Zambelli. Thorndike’s consistency. His researches on the speculum astronomiae from 1923 to 1955. In The Speculum astronomiae and its enigma: astrology, theology, and science in Albertus Magnus and his contemporaries [Zam92h], chapter 4, pages 33–42. ISBN 0-7923-1380-1 (hardcover), 94-017-3467-4. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 135. URL https://link.

                                              springer.com/chapter/10.1007/
                                              978-94-017-3467-7_4.

Zanstra:1974:PMO

[Zan74] Herman Zanstra. On the philosophical meaning of observational errors. In For Dirk Struik: Scientific, Historical and Political Essays in Honor of Dirk J. Struik [CSW74], chapter 37, pages 477–490. ISBN 90-277-0379-5, 94-010-2115-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-010-
2115-9_37.

Zilberman:1988:BM

[ZC88a] David B. Zilberman and Robert S.

Cohen. The birth of ‘meaning’. In The Birth of Meaning in Hindu Thought [Coh88b], chapter 2, pages 71–140. ISBN 94-010-7141-1, 94-009-1431-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN B131 .Z55 1988; Q174 .B67 vol. 102 B131.

URL https://link.springer.com/
chapter/10.1007/978-94-009-1431-
5_2.

Zilberman:1988:CSW

[ZC88b] David B. Zilberman and Robert S.

Cohen. The canonical self in the world of knowledge. In The Birth of Meaning in Hindu Thought [Coh88b], chapter 4, pages 180–217. ISBN 94-010-7141-1, 94-009-1431-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN B131 .Z55 1988; Q174 .B67 vol. 102

B131. URL https://link.springer.
com/chapter/10.1007/978-94-009-
1431-5_4.

Zilberman:1988:DKN

[ZC88c] David B. Zilberman and Robert S.

Cohen. Dialectics in Kant and in the Nyāya-Sūtra. In The Birth of Meaning in Hindu Thought [Coh88b], chapter 3, pages 141–179. ISBN 94-010-7141-1, 94-009-1431-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN B131 .Z55 1988; Q174 .B67 vol. 102

B131. URL https://link.springer.
com/chapter/10.1007/978-94-009-
1431-5_3.

Zilberman:1988:FHV

[ZC88d] David B. Zilberman and Robert S. Cohen. The family of Hindu ‘visions’ as cultural entities. In The Birth of Meaning in Hindu Thought [Coh88b], chapter 9, pages 330–349. ISBN 94-010-7141-1, 94-009-1431-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN B131 .Z55 1988; Q174 .B67 vol. 102 B131. URL https:

//link.springer.com/chapter/10.
1007/978-94-009-1431-5_9.

Zilberman:1988:HST

[ZC88e] David B. Zilberman and Robert S. Cohen. Hindu systems of thought as epistemic disciplines. In The Birth of Meaning in Hindu Thought [Coh88b], chapter 1, pages 1–70. ISBN 94-010-7141-1, 94-009-1431-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN B131 .Z55 1988; Q174 .B67 vol. 102 B131. URL https:

//link.springer.com/chapter/10.
1007/978-94-009-1431-5_1.

Zilberman:1988:HVB

[ZC88f] David B. Zilberman and Robert S. Cohen. Hindu values and Buddhism. In The Birth of Meaning in Hindu Thought [Coh88b], chapter 7, pages 263–298. ISBN 94-010-7141-1, 94-009-1431-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN B131 .Z55 1988; Q174 .B67 vol. 102

B131. URL https://link.springer.
com/chapter/10.1007/978-94-009-
1431-5_7.

Zilberman:1988:BS

[ZC88g] David B. Zilberman and Robert S.

Cohen. Is the Bodhisattva a skeptic? In The Birth of Meaning in Hindu Thought [Coh88b], chapter 6, pages 247–262. ISBN 94-010-7141-1, 94-009-1431-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN B131 .Z55 1988; Q174 .B67 vol. 102

B131. URL https://link.springer.
com/chapter/10.1007/978-94-009-
1431-5_6.

Zilberman:1988:RAV

[ZC88h] David B. Zilberman and Robert S.

Cohen. Revelation in advaita vedānta as an experiment in the semantic destruction of language. In The Birth of Meaning in Hindu Thought [Coh88b], chapter 5, pages 218–246. ISBN 94-010-7141-1, 94-009-1431-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN B131 .Z55 1988; Q174 .B67 vol. 102 B131. URL https:

//link.springer.com/chapter/10.
1007/978-94-009-1431-5_5.

Zilberman:1988:UCT

[ZC88i] David B. Zilberman and Robert S.

Cohen. Understanding cultural traditions through types of thinking. In The Birth of Meaning in Hindu Thought [Coh88b], chapter 8, pages 299–329. ISBN 94-010-7141-1, 94-009-1431-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN B131 .Z55 1988; Q174 .B67 vol. 102 B131.

URL https://link.springer.com/
chapter/10.1007/978-94-009-1431-
5_8.

Zhenghai:1996:WDZ

[ZC96] Song Zhenghai and Chen Chuankang.

Why did Zheng He’s sea voyage fail to lead the Chinese to make the ‘great geographic discovery’? In Chinese Studies in the History and Philosophy of Science and Technology [DC96], chapter 25, pages 303–314. ISBN 90-481-4546-5, 94-015-8717-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 180; U390 .N19 1996. URL https:

//link.springer.com/chapter/10.
1007/978-94-015-8717-4_25.

Zeleny:1974:KMM

[Zel74] Jindřich Zelený. Kant, Marx and the modern rationality. In Methodological and Historical Essays in the Natural and Social Sciences [CW74b], chapter 18, pages 361–375. ISBN 90-277-0378-7, 94-010-2128-7 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 14.

URL https://link.springer.com/
chapter/10.1007/978-94-010-2128-
9_18.

Zeleny:1995:ADT

[Zel95] Jindřich Zelený. Analytical and/or dialectical thinking. In The Concept of Knowledge: The Ankara Seminar [KC95], chapter 14, pages 183–189. ISBN 90-481-4495-7, 94-017-3263-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-017-3263-
5_14.

Zemach:1986:MOM

[Zem86]  Eddy M. Zemach.  Meaning and

our mental life: A comment. In Ullmann-Margalit [UM86a], chapter 4, pages 33–37. ISBN 90-277-2159-9 (paperback), 90-277-2158-0 (hardcover), 94-009-5496-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 94 Q175.

URL https://link.springer.com/
chapter/10.1007/978-94-009-5496-
0_4.

Zemach:1988:ESR

[Zem88] Eddy M. Zemach. Epistemic and semantic reflections on scientific realism. In Science in Reflection: The Israel Colloquium: Studies in History, Philosophy, and Sociology of Science Volume 3 [UM88], chapter 8, pages 97–101. ISBN 90-277-2713-9, 94-009-2957-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-009-2957-
9_8.

Zeman:2009:KPH

[Zem09] Vladimír Zeman. Karl Popper and Hans Albert — the broad scope of critical rationalism. In Rethinking Popper [PC09], chapter 8, pages 91–100. ISBN 1-4020-9337-3, 1-4020-9338-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN B1649.P64 R48 2009; Q174 .B67 2009. URL https:

//link.springer.com/chapter/10.
1007/978-1-4020-9338-8_8.

Zenzen:1999:SHM

[Zen99] Michael J. Zenzen. Science, hermeneutics, and metaphorical thought. In Fehér et al. [FKR99], pages 281–291. ISBN 90-481-5257-7, 94-015-9293-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-015-9293-
2_22.

Zenneck:2007:G

[Zen07] Jonathan Zenneck. Gravitation. In Renn et al. [RSS^{+}07a], pages 77–112. ISBN 1-4020-3999-9. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 no. 250 v. 3; QC173.6 .G469

2007. URL https://link.springer.
com/chapter/10.1007/978-1-4020-
4000-9_12.

Zhengkun:1996:TFS

[Zhe96] Yin Zhengkun. Truth and fiction in scientific theory. In Cohen et al. [CHR96], pages 265–277. ISBN 90-481-4493-0, 94-015-8638-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-015-8638-2_18.

Zilsel:2003:SOM

[Zil03] Edgar Zilsel. The Social Origins of Modern Science, volume 200 of Boston Studies in the Philosophy of Science. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 2003. ISBN 0-7923-6457-0, 94-011-4142-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). lix + 267 pp. LCCN Q174 .B67 vol.

200; Q175.5. URL http://www.loc.
gov/catdir/enhancements/fy0813/
00042657-d.html;   http://www.
loc.gov/catdir/enhancements/
fy0813/00042657-t.html;   https:
//link.springer.com/book/10.
1007/978-94-011-4142-0.  Edited

59 years after the author’s death by Diederick Raven, Wolfgang Krohn, and R. S. (Robert Sonné) Cohen.

Zimmerli:1983:FVR

[Zim83]  Walther Ch. Zimmerli.  Forecast,

value, and the recent phenomenon of non-acceptance: The limits of a philosophy of technology assessment. In Philosophy and Technology [DR83], chapter 12, pages 165–184. ISBN 94-009-7126-5, 94-009-7124-9 (e-book).

ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-009-7124-0_12.

Zimmerli:1986:WBD

[Zim86] Walther Ch. Zimmerli. Who is to blame for data pollution? On individual moral responsibility with information technology. In Philosophy and Technology II: Information Technology and Computers in Theory and Practice [MH86], chapter 21, pages 291–305. ISBN 94-010-8510-2, 94-009-4512-4 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-009-4512-8_21.

Zimmermann:1992:IPI

[Zim92] Francis Zimmermann. Integration problems: Introductory report. In Science and Empires: Historical Studies about Scientific Development and European Expansion [PJM92], chapter 5, pages 33–35. ISBN 0-7923-1518-9 (hardcover), 94-010-5145-3, 94-011-2594-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 136.

URL https://link.springer.com/
chapter/10.1007/978-94-011-2594-
9_5.

Zimmerli:1999:NCT

[Zim99] Walther Ch. Zimmerli. Nietzsche’s critique of truth and science: A comprehensive approach. In Nietzsche, Epistemology, and Philosophy of Science: Nietzsche and the Sciences II [Bab99a], chapter 20, pages 253–277. ISBN 90-481-5234-8, 94-017-2428-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-017-2428-
9_20.

Ziman:2002:NMI

[Zim02] John Ziman. No man is an island. In Hermeneutic Philosophy of Science, Van Gogh’s Eyes, and God: Essays in Honor of Patrick A. Heelan, S.J. [Bab02c], chapter 17, pages 203–217. ISBN 90-481-5926-1, 94-017-1767-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 225. URL https://link.

                                              springer.com/chapter/10.1007/
                                              978-94-017-1767-0_17.

Zinovev:1973:Ca

[Zin73a] A. A. Zinov’ev. Classes. In Foundations of the Logical Theory of Scientific Knowledge (Complex Logic) [Zin73g], chapter 14, pages 166–184. ISBN 90-277-0324-8, 94-010-2501-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-010-2501-0_14.

Zinovev:1973:Cb

[Zin73b] A. A. Zinov’ev. Conclusion. In Foundations of the Logical Theory of Scientific Knowledge (Complex Logic) [Zin73g], chapter 22, page 263. ISBN 90-277-0324-8, 94-010-2501-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-010-2501-0_22.

Zinovev:1973:CS

[Zin73c] A. A. Zinov’ev. Conditional sentences. In Foundations of the Logical Theory of Scientific Knowledge (Complex Logic) [Zin73g], chapter 12, pages 157–161. ISBN 90-277-0324-8, 94-010-2501-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

                                              URL https://link.springer.com/

chapter/10.1007/978-94-010-2501-
0_12.

Zinovev:1973:EAO

[Zin73d] A. A. Zinov’ev. Empirical and abstract objects. In Foundations of the Logical Theory of Scientific Knowledge (Complex Logic) [Zin73g], chapter 9, pages 109–122. ISBN 90-277-0324-8, 94-010-2501-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-010-
2501-0_9.

Zinovev:1973:EL

[Zin73e] A. A. Zinov’ev. Existential logic. In Foundations of the Logical Theory of Scientific Knowledge (Complex Logic) [Zin73g], chapter 15, pages 185–186. ISBN 90-277-0324-8, 94-010-2501-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-010-2501-
0_15.

Zinovev:1973:FGT

[Zin73f] A. A. Zinov’ev. Formalization of the general theory of logical entailment. In Foundations of the Logical Theory of Scientific Knowledge (Complex Logic) [Zin73g], chapter 7, pages 75–89. ISBN 90-277-0324-8, 94-010-2501-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-010-2501-0_7.

Zinovev:1973:FLT

[Zin73g] A. A. Zinov’ev. Foundations of the Logical Theory of Scientific Knowledge (Complex Logic), volume 9 of Boston Studies in the Philosophy of Science. D. Reidel, Dordrecht, The Netherlands; Boston, MA, USA; Lancaster, UK; Tokyo, Japan, 1973. ISBN 90-277-0324-8, 94-010-2501-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). xxii + 304 pp. LCCN ????

URL https://link.springer.com/
book/10.1007/978-94-010-2501-0.

Zinovev:1973:GTL

[Zin73h] A. A. Zinov’ev. The general theory of logical entailment. In Foundations of the Logical Theory of Scientific Knowledge (Complex Logic) [Zin73g], chapter 6, pages 64–74. ISBN 90-277-0324-8, 94-010-2501-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-010-2501-0_6.

Zinovev:1973:LO

[Zin73i] A. A. Zinov’ev. Logic and ontology. In Foundations of the Logical Theory of Scientific Knowledge (Complex Logic) [Zin73g], chapter 20, pages 247–256. ISBN 90-277-0324-8, 94-010-2501-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-010-2501-
0_20.

Zinovev:1973:LTS

[Zin73j] A. A. Zinov’ev. The logical theory of scientific knowledge. In Foundations of the Logical Theory of Scientific Knowledge (Complex Logic) [Zin73g], chapter 1, pages 1–7. ISBN 90-277-0324-8, 94-010-2501-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-010-2501-0_1.

Zinovev:1973:MS

[Zin73k] A. A. Zinov’ev. Modal sentences. In Foundations of the Logical Theory of Scientific Knowledge (Complex Logic) [Zin73g], chapter 16, pages 187–209. ISBN 90-277-0324-8, 94-010-2501-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-010-2501-
0_16.

Zinovev:1973:PE

[Zin73l] A. A. Zinov’ev. Physical entailment. In Foundations of the Logical Theory of Scientific Knowledge (Complex Logic) [Zin73g], chapter 18, pages 226–238. ISBN 90-277-0324-8, 94-010-2501-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-010-2501-
0_18.

Zinovev:1973:R

[Zin73m] A. A. Zinov’ev. Relations. In Foundations of the Logical Theory of Scientific Knowledge (Complex Logic) [Zin73g], chapter 17, pages 210–225. ISBN 90-277-0324-8, 94-010-2501-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-010-2501-0_17.

Zinovev:1973:Sb

[Zin73n] A. A. Zinov’ev. Sentences. In Foundations of the Logical Theory of Scientific Knowledge (Complex Logic) [Zin73g], chapter 4, pages 32–51. ISBN 90-277-0324-8, 94-010-2501-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-010-2501-0_4.

Zinovev:1973:SQ

[Zin73o] A. A. Zinov’ev. Sentences with quantifiers. In Foundations of the Logical Theory of Scientific Knowledge (Complex Logic) [Zin73g], chapter 10, pages 123–145. ISBN 90-277-0324-8, 94-010-2501-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

                                              ???? URL https://link.springer.
                                              com/chapter/10.1007/978-94-010-
                                              2501-0_10.

Zinovev:1973:SL

[Zin73p] A. A. Zinov’ev. Sentential logic. In Foundations of the Logical Theory of Scientific Knowledge (Complex Logic) [Zin73g], chapter 5, pages 52–63. ISBN 90-277-0324-8, 94-010-2501-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-010-2501-0_5.

Zinovev:1973:Sa

[Zin73q] A. A. Zinov’ev. Signs. In Foundations of the Logical Theory of Scientific Knowledge (Complex Logic) [Zin73g], chapter 2, pages 8–20. ISBN 90-277-0324-8, 94-010-2501-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-010-2501-0_2.

Zinovev:1973:SPS

[Zin73r] A. A. Zinov’ev. Subject–predicate structures. In Foundations of the Logical Theory of Scientific Knowledge (Complex Logic) [Zin73g], chapter 8, pages 90–108. ISBN 90-277-0324-8, 94-010-2501-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

                                              ???? URL https://link.springer.
                                              com/chapter/10.1007/978-94-010-
                                              2501-0_8.

Zinovev:1973:Ta

[Zin73s] A. A. Zinov’ev. Terms. In Foundations of the Logical Theory of Scientific Knowledge (Complex Logic) [Zin73g], chapter 3, pages 21–31. ISBN 90-277-0324-8, 94-010-2501-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-010-2501-0_3.

Zinovev:1973:Tb

[Zin73t] A. A. Zinov’ev. Theories. In Foundations of the Logical Theory of Scientific Knowledge (Complex Logic) [Zin73g], chapter 19, pages 239–246. ISBN 90-277-0324-8, 94-010-2501-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-010-2501-0_19.

Zinovev:1973:TQ

[Zin73u] A. A. Zinov’ev. Theory of quantifiers.

In Foundations of the Logical Theory of Scientific Knowledge (Complex Logic) [Zin73g], chapter 11, pages 146–156. ISBN 90-277-0324-8, 94-010-2501-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-010-2501-
0_11.

Zinovev:1973:TT

[Zin73v] A. A. Zinov’ev. Theory of terms. In Foundations of the Logical Theory of Scientific Knowledge (Complex Logic) [Zin73g], chapter 13, pages 162–165. ISBN 90-277-0324-8, 94-010-2501-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.com/
chapter/10.1007/978-94-010-2501-
0_13.

Zinovev:1973:UL

[Zin73w] A. A. Zinov’ev. The universality of logic. In Foundations of the Logical Theory of Scientific Knowledge (Complex Logic) [Zin73g], chapter 21, pages 257–262. ISBN 90-277-0324-8, 94-010-2501-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

                                              ???? URL https://link.springer.
                                              com/chapter/10.1007/978-94-010-
                                              2501-0_21.

Zinovev:1983:NTT

[Zin83a] A. A. Zinov’ev. The non-traditional theory of quantifiers. In Language, Logic and Method [CW83b], chapter 17, pages 355–408. ISBN 94-009-7704-2, 94-009-7702-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic).

                                              LCCN ????  URL https://link.
                                              springer.com/chapter/10.1007/
                                              978-94-009-7702-0_17.

Zinoviev:1983:LP

[Zin83b] Aleksandr Zinoviev. Logical Physics, volume 74 of Boston Studies in the Philosophy of Science. D. Reidel, Dordrecht, The Netherlands; Boston, MA, USA; Lancaster, UK; Tokyo, Japan, 1983. ISBN 90-277-0734-0 (hardcover), 90-277-0743-0. xvii + 282 pp. LCCN BC57 .Z56 1983; Q174 .B67 vol. 74. Translated from the Russian Logicheskaia fizika, by O. A. Germogenova. Edited by Robert S. Cohen.

Zivotic:1979:BTT

[Živ79] Miladin Životić. Between two types of modern culture. In Praxis: Yugoslav Essays in the Philosophy and Methodology of the Social Sciences [MP79], chapter 11, pages 187–197. ISBN 90-277-0968-8, 94-009-9355-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-009-9355-6_11.

Ziche:2013:POO

[ZM11] Paul Ziche and Thomas Müller. Paul Oppenheim on order — the career of a logico-philosophical concept. In The Berlin Group and the Philosophy of Logical Empiricism [MP13], chapter 13, pages 265–291. ISBN 94-007-5484-1 (hardcover), 94-007-5485-X (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175.32.A24 .M384 2011; Q175 .B47 2013. URL https:

//link.springer.com/chapter/10.
1007/978-94-007-5485-0_13.

Zolo:1989:C

[Zol89a] Danilo Zolo. Conclusion. In Reflexive Epistemology: The Philosophical Legacy of Otto Neurath [Zol89g], chapter 8, pages 167–184. ISBN 94-010-7588-3, 94-009-2415-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-009-
2415-4_8.

Zolo:1989:ENM

[Zol89b] Danilo Zolo. Enlightenment, neo– Marxism, conventionalism: Towards a critique of Cartesian rationalism. In Reflexive Epistemology: The Philosophical Legacy of Otto Neurath [Zol89g], chapter 2, pages 15–26. ISBN 94-010-7588-3, 94-009-2415-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-009-2415-4_2.

Zolo:1989:EPP

[Zol89c] Danilo Zolo. Evaluation, prescription, and political decision. In Reflexive Epistemology: The Philosophical Legacy of Otto Neurath [Zol89g], chapter 7, pages 145–166. ISBN 94-010-7588-3, 94-009-2415-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

                                              ???? URL https://link.springer.
                                              com/chapter/10.1007/978-94-009-
                                              2415-4_7.

Zolo:1989:LRP

[Zol89d] Danilo Zolo. Linguistic reflexivity and ‘pseudorationalism’. In Reflexive Epistemology: The Philosophical Legacy of Otto Neurath [Zol89g], chapter 3, pages 27–60. ISBN 94-010-7588-3, 94-009-2415-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

                                              ???? URL https://link.springer.
                                              com/chapter/10.1007/978-94-009-
                                              2415-4_3.

Zolo:1989:NVP

[Zol89e] Danilo Zolo. Neurath versus Popper. In Reflexive Epistemology: The Philosophical Legacy of Otto Neurath [Zol89g], chapter 4, pages 61–82. ISBN 94-010-7588-3, 94-009-2415-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-009-2415-4_4.

Zolo:1989:PA

[Zol89f] Danilo Zolo. The problem of assessment. In Reflexive Epistemology: The Philosophical Legacy of Otto Neurath [Zol89g], chapter 1, pages 1–13. ISBN 94-010-7588-3, 94-009-2415-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-009-2415-4_1.

Zolo:1989:REP

[Zol89g] Danilo Zolo. Reflexive Epistemology: The Philosophical Legacy of Otto Neurath, volume 118 of Boston Studies in the Philosophy of Science. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 1989. ISBN 94-010-7588-3, 94-009-2415-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). xx + 206 pp. LCCN ???? URL

https://link.springer.com/book/
10.1007/978-94-009-2415-4.

Zolo:1989:SWE

[Zol89h] Danilo Zolo. Strengths and weaknesses of an empirical sociology. In Reflexive Epistemology: The Philosophical Legacy of Otto Neurath [Zol89g], chapter 6, pages 107–144. ISBN 94-010-7588-3, 94-009-2415-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-009-
2415-4_6.

Zolo:1989:USH

[Zol89i] Danilo Zolo. The unity of science as a historico–sociological goal: From the primacy of physics to the epistemological priority of sociology. In Reflexive Epistemology: The Philosophical Legacy of Otto Neurath [Zol89g], chapter 5, pages 83–106. ISBN 94-010-7588-3, 94-009-2415-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-009-
2415-4_5.

Zilsel:2003:CPN

[ZRKC03a] Edgar Zilsel, Diederick Raven, Wolfgang Krohn, and Robert S. Cohen. Concerning ‘phenomenology and natural science’. In The Social Origins of Modern Science [Zil03], chapter 12, pages 214–215. ISBN 0-7923-6457-0, 94-011-4142-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic).

LCCN Q174 .B67 vol. 200; Q175.5.

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-011-4142-

0_12. Edited 59 years after the author’s death by Diederick Raven, Wolfgang Krohn, and R. S. (Robert Sonné) Cohen.

Zilsel:2003:CM

[ZRKC03b] Edgar Zilsel, Diederick Raven, Wolfgang Krohn, and Robert S. Cohen. Copernicus and mechanics. In The Social Origins of Modern Science [Zil03], chapter 7, pages 123–127. ISBN 0-7923-6457-0, 94-011-4142-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 200; Q175.5. URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-011-4142-0_7. Edited

59 years after the author’s death by Diederick Raven, Wolfgang Krohn, and R. S. (Robert Sonné) Cohen.

Zilsel:2003:GCP

[ZRKC03c] Edgar Zilsel, Diederick Raven, Wolfgang Krohn, and Robert S. Cohen. The genesis of the concept of physical law. In The Social Origins of Modern Science [Zil03], chapter 6, pages 96–122. ISBN 0-7923-6457-0, 94-011-4142-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 200; Q175.5. URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-011-4142-0_6. Edited

59 years after the author’s death by Diederick Raven, Wolfgang Krohn, and R. S. (Robert Sonné) Cohen.

Zilsel:2003:GCS

[ZRKC03d] Edgar Zilsel, Diederick Raven, Wolfgang Krohn, and Robert S. Cohen. The genesis of the concept of scientific progress and scientific cooperation. In The Social Origins of Modern Science [Zil03], chapter 8, pages 128–168. ISBN 0-7923-6457-0, 94-011-4142-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 200; Q175.5. URL https:

//link.springer.com/chapter/10.
1007/978-94-011-4142-0_8. Edited

59 years after the author’s death by Diederick Raven, Wolfgang Krohn, and R. S. (Robert Sonné) Cohen.

Zilsel:2003:HBE

[ZRKC03e] Edgar Zilsel, Diederick Raven, Wolfgang Krohn, and Robert S. Cohen. History and biological evolution. In The Social Origins of Modern Science [Zil03], chapter 13, pages 216–220. ISBN 0-7923-6457-0, 94-011-4142-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 200; Q175.5. URL https:

//link.springer.com/chapter/10.
1007/978-94-011-4142-0_13. Edited

59 years after the author’s death by Diederick Raven, Wolfgang Krohn, and R. S. (Robert Sonné) Cohen.

Zilsel:2003:MH

[ZRKC03f] Edgar Zilsel, Diederick Raven, Wolfgang Krohn, and Robert S. Cohen. The methods of humanism. In The Social Origins of Modern Science [Zil03], chapter 3, pages 22–64. ISBN 0-7923-6457-0, 94-011-4142-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 200; Q175.5. URL https://link.

springer.com/chapter/10.1007/
978-94-011-4142-0_3.  Edited 59

years after the author’s death by Diederick Raven, Wolfgang Krohn, and R. S. (Robert Sonné) Cohen.

Zilsel:2003:OWG

[ZRKC03g] Edgar Zilsel, Diederick Raven, Wolfgang Krohn, and Robert S. Cohen. The origins of William Gilbert’s scientific method. In The Social Origins of Modern Science [Zil03], chapter 5, pages 71–95. ISBN 0-7923-6457-0, 94-011-4142-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 200; Q175.5.

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-011-4142-

0_5. Edited 59 years after the author’s death by Diederick Raven, Wolfgang Krohn, and R. S. (Robert Sonné) Cohen.

Zilsel:2003:PNS

[ZRKC03h] Edgar Zilsel, Diederick Raven, Wolfgang Krohn, and Robert S. Cohen. Phenomenology and natural science. In The Social Origins of Modern Science [Zil03], chapter 11, pages 209–213. ISBN 0-7923-6457-0, 94-011-4142-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 200; Q175.5. URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-011-4142-0_11. Edited

59 years after the author’s death by Diederick Raven, Wolfgang Krohn, and R. S. (Robert Sonné) Cohen.

Zilsel:2003:PPH

[ZRKC03i] Edgar Zilsel, Diederick Raven, Wolfgang Krohn, and Robert S. Cohen. Physics and the problem of historico– sociological laws. In The Social Origins of Modern Science [Zil03], chapter 10, pages 200–208. ISBN 0-7923-6457-0, 94-011-4142-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 200; Q175.5.

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-011-4142-

0_10. Edited 59 years after the author’s death by Diederick Raven, Wolfgang Krohn, and R. S. (Robert Sonné) Cohen.

Zilsel:2003:PE

[ZRKC03j] Edgar Zilsel, Diederick Raven, Wolfgang Krohn, and Robert S. Cohen. Problems of empiricism. In The Social Origins of Modern Science [Zil03], chapter 9, pages 171–199. ISBN 0-7923-6457-0, 94-011-4142-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 200; Q175.5. URL https://link.

springer.com/chapter/10.1007/
978-94-011-4142-0_9.  Edited 59

years after the author’s death by Diederick Raven, Wolfgang Krohn, and R. S. (Robert Sonné) Cohen.

Zilsel:2003:SHS

[ZRKC03k] Edgar Zilsel, Diederick Raven, Wolfgang Krohn, and Robert S. Cohen. Science and the humanistic studies. In The Social Origins of Modern Science [Zil03], chapter 14, pages 221–223. ISBN 0-7923-6457-0, 94-011-4142-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 200; Q175.5. URL https:

//link.springer.com/chapter/10.
1007/978-94-011-4142-0_14. Edited

59 years after the author’s death by Diederick Raven, Wolfgang Krohn, and R. S. (Robert Sonné) Cohen.

Zilsel:2003:SRSa

[ZRKC03l] Edgar Zilsel, Diederick Raven, Wolfgang Krohn, and Robert S. Cohen. The social roots of science. In The Social Origins of Modern Science [Zil03], chapter 1, pages 3–6. ISBN 0-7923-6457-0, 94-011-4142-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 200; Q175.5. URL https://link.

                                              springer.com/chapter/10.1007/
                                              978-94-011-4142-0_1.  Edited 59

years after the author’s death by Diederick Raven, Wolfgang Krohn, and R. S. (Robert Sonné) Cohen.

Zilsel:2003:SRSb

[ZRKC03m] Edgar Zilsel, Diederick Raven, Wolfgang Krohn, and Robert S. Cohen. The sociological roots of science. In The Social Origins of Modern Science [Zil03], chapter 2, pages 7–21. ISBN 0-7923-6457-0, 94-011-4142-8 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 200; Q175.5. URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-011-4142-0_2. Edited

59 years after the author’s death by Diederick Raven, Wolfgang Krohn, and R. S. (Robert Sonné) Cohen.

Zucker:1969:SCW

[Zuc69] Francis J. Zucker. Supplementary comments to Weizsäcker’s paper. In Cohen and Wartofsky [CW69a], pages 474–482. ISBN 94-010-3383-8, 94-010-3263-7, 94-010-3381-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-010-3381-7_18.

Zumeng:1996:CZH

[Zum96] Jin Zumeng. A critique of ‘Zhang Heng’s theory of a spherical Earth’. In Chinese Studies in the History and Philosophy of Science and Technology [DC96], chapter 32, pages 427–432. ISBN 90-481-4546-5, 94-015-8717-5 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 180; U390 .N19 1996.

URL https://link.springer.com/
chapter/10.1007/978-94-015-8717-
4_32.

Zuoxiu:1996:EPR

[Zuo96] He Zuoxiu. On the Einstein, Podolsky and Rosen Paradox and the relevant philosophical problems. In Cohen et al. [CHR96], pages 299–305. ISBN 90-481-4493-0, 94-015-8638-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

//link.springer.com/chapter/10.
1007/978-94-015-8638-2_21.

Zvan:1979:EHR

[Žva79] Antun Žvan. Ecstasy and hangover of a revolution. In Praxis: Yugoslav Essays in the Philosophy and Methodology of the Social Sciences [MP79], chapter 22, pages 357–369. ISBN 90-277-0968-8, 94-009-9355-2 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN

???? URL https://link.springer.
com/chapter/10.1007/978-94-009-
9355-6_22.

Zwirn:2008:CWC

[Zwi08] Hervé Zwirn. Can we consider quantum mechanics to be a description of reality? In Rethinking Scientific Change and Theory Comparison: Stabilities, Ruptures, Incommensurabilities? [SSHH08], chapter 14, pages 209–217. ISBN 1-4020-6274-5 (hardcover), 1-4020-6279-6 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q175 .R397 2008. URL https://

link.springer.com/chapter/10.
1007/978-1-4020-6279-7_14.  See

[Le 08].

Zwirn:2009:FPE

[Zwi09] Hervé Zwirn. Foundations of physics:

The empirical blindness. In Brenner and Gayon [BG09], chapter 6, pages 141–163. ISBN 1-4020-9367-5, 1-4020-9368-3 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .F746 2009. URL https://link.

                                              springer.com/chapter/10.1007/
                                              978-1-4020-9368-5_6.

Zycinski:1995:HRS

[Życ95a] Józef Życiński. How to de-ruse sociobiological theory of knowledge? In Misiek [Mis95c], pages 239–252. ISBN 94-010-4206-3, 94-011-0461-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-011-0461-6_19.

Zycinski:1995:TKR

[Życ95b] Józef Życiński. Tacit knowing and the rationality of science. In Misiek [Mis95c], pages 253–263. ISBN 94-010-4206-3, 94-011-0461-1 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ???? URL https:

                                              //link.springer.com/chapter/10.
                                              1007/978-94-011-0461-6_20.

Zytkow:1982:ICS

[Żyt82] J. M. Żytkow. An interpretation of a concept in science by a set of operational procedures. In Polish Essays in the Philosophy of the Natural Sciences [Kra82c], chapter 15, pages 169–185. ISBN 94-009-7705-0 (e-book), 90-277-1287-5. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

                                              URL https://link.springer.com/
                                              chapter/10.1007/978-94-009-7705-
                                              1_15.